A Complete Bibliography of IEEE/ACM Transactions on Networking

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

19 June 2024 Version 1.84

Title word cross-reference

+ [CDRV11]. 1 [BB16, LWL17, WNV13]. 1 + 1 [BCO17]. 1 + N [Kam10]. 10/7 + ϵ [SZ07]. 2 [AMG^{+}17, CPGZ15, JYT^{+}15, KKL03, LWL17, NBV17, WYL24, ZGLC20]. 2 − 1/N [HYZH16]. 3 [CQLW22, CXW^{+}18, FWN^{+}22, HR14, JYT^{+}15, KG05, LS93b, LZL^{+}14, LWK^{+}16, LJL^{+}16, LDY^{+}16, WJYL16, WWL24b, YJZW15, ZCF^{+}24]. 4 [DM15, YJ15]. 5 [AMCD19, AdVS20, MCC^{+}19, SCPB19, SKA^{+}18]. 60 [NKNK17]. = [CDRV11]. [w, f] [NWP09]. ^{+} [BLC21, ZY21]. ^{2} [GAA08]. ^{3} [LLM^{+}24]. α [ABC^{+}16, KDYV12]. c [XZL^{+}21]. d [HMM^{+}20, LQ13]. f [JPH08]. F^{2} [CZX^{+}17]. K [HS16, HSM^{+}20, ADT22, BCE^{+}19, GCWC17, KWS^{+}11, LLX^{+}17, LLG^{+}17, OGLK14, SXQ^{+}23, TXW^{+}21, VTBK21, YZL^{+}19, YBX^{+}10, ZH08a, ZWL^{+}16, ZSZN21, ZL16]. k <= 6 [YBX^{+}10]. L^{2} [CHML15]. log_{2} N [ZGS10]. m [LWK^{+}16]. μ [DGLM16]. N [CN08, OdG96, SL95, LZZ^{+}22a]. N × N [NMC07]. O(1) [Guo04]. O(log W) [LS07]. p [EM09, Kam10, MJ13, SJ12, WYH10]. q [Zha17]. R × W [AF99]. θ [XK06b].

-based [WBP^{+}11]. -bit [BB16]. -by-1 [LZZ^{+}22a]. -cast [JPH08]. -Composite [Zha17]. -Connected [LWK^{+}16]. -Connectivity [ADT22, YBX^{+}10, ZH08a]. -Cover [ZWL^{+}16]. -Coverage [ZWL^{+}16, XK06b]. -cycle [WYH10]. -cycles [EM09, Kam10, MJ13]. -D [JYT^{+}15, CXW^{+}18, JYT^{+}15, LZL^{+}14, NBV17, WJYL16, YJZW15]. -D/ [JYT^{+}15].

-dense [OGLK14]. -Dimensional [HMM^{+}20, LQ13]. -distributors [NWP09]. -diverse [SYR05]. -Dominating [LWK^{+}16]. -GHz [NKNK17]. -hop [WNV13, ZL16, HS16]. -hub [CN08]. -Learning [BLC21]. -NET [DGLM16]. -optimal [KDYV12]. -Persistent [HSM^{+}20]. -priority [LS93b]. -route [KKL03]. -Sink [GCWC17, KWS^{+}11]. -Source [HR14]. -structures [SJ12]. -Terminal [HR14, ZSZN21]. -to- [LWL17].

.onion [MRMR17].

/0 [DKC^{+}15].

1 [CM16, HBH93, JID^{+}07, NBTD07]. 100 [BLC12]. 100-Gb [BLC12]. 100-Gb/s [BLC12]. 10GBase [PYL^{+}17]. 10GBase-T [PYL^{+}17]. 10GbE [FC17]. 18-Year [ACZP21].

2 [KKM^{+}97]. 2K [TMGB19].

3-Year-Old [DLY^{+}22]. 3G [LXW^{+}17]. 3G/4G [LXW^{+}17].

40 [HM06]. 48 [BKH^{+}93]. 4G [LGC16, LXW^{+}17, YZBR14].

50 [PCB^{+}98]. 50-Gb [PCB^{+}98]. 50-Gb/s [PCB^{+}98]. 5G [AZP^{+}23, AMSB^{+}24, BLM^{+}17, CFP^{+}21, FMSM24, HLZY23, HLHL22, PPTP21, YYL23, ZLZ21a, ZLL^{+}23a].

60 [GHK18, SMM11, ZWGC17, ZWZM18]. 60-GHz [GHK18, SMM11, ZWZM18]. 6G [MLS^{+}23]. 6LB [DPT^{+}18]. 6Scan [HCW^{+}23].

802.11 [BOGS^{+}16, BK17, HKV^{+}13, HDM13, JS12, LLY^{+}16, PL17, TS08]. 802.11-based [LLM11b].

802.11-scheduled [JP09]. 802.11-Type [BK17]. 802.11a [QCS07]. 802.11a/h [QCS07]. 802.11ac [LCLC18]. 802.11ax [LR22]. 802.11e [BCGM07, TB10, RKA08]. 802.11ec [MGK14]. 802.11n [APB^{+}13, PLL13]. 802.12 [Kim98]. 802.16 [CAL09].

92 [HBS96].

[HLHL22].

A-MAC [VA07]. A2 [Kri14]. AAL [Kam96]. ABC [ST13]. ABD [TKZ94]. ABE [LPM23]. Abnormally [SKG12]. ABOI [YWHW24]. ABR [BFMF01, GM00, KJF^{+}00, KR99, SDW00, ZSSK02]. absolute [VRK09, WXBZ04]. Abstract [YZZ^{+}21, CDO97]. Abstraction [CWHW18, GXW^{+}19, MSTL17, MKG^{+}17, TML^{+}18, YLH17, NLB15, RCGS09, RM08]. abstractions [RD11a]. Abstracts [Tow06a]. Accelerated [XCW^{+}20a, XCW^{+}20b, WZL^{+}13]. Accelerating [BBK12, KAT^{+}22, LXX^{+}24, LZZ^{+}22b, RRS23, SCHG22, ZWZM18]. Acceleration [WZL^{+}23c, ZXW^{+}19]. accepted [CTG00]. Access [AD18, BBF18, CBdV^{+}17, CLGSS17, CG21, CH93, CGYZ17, CP18, CCG20, CEFS21, DRMP18, Dai22, DF20, DPTP24, EF17, EE18, GHW22, GSM16, IGHT17, KPK^{+}16, LWL17, LPD^{+}18, LK16b, LYL^{+}22b, LJSB22, MGK20, NST^{+}16, PPTP21, QZL^{+}16, SA21, SX16, URZ^{+}14, WZLM22, XZL20, XHZ^{+}19, YHCL21, YSY16, ZGHH19, ZW22, AD14, ALMR14, BCP00, BCL12, BB06, BS97, BD97, BP96, CZ12, Cha10, CL09a, CLD10, CG04, CFZ97, CPR99, CEFS99, CR98, DLT16, DHSS14, FTZ^{+}13, GS13, GRB09, HA16, HSM^{+}13, IW08, IZC00, JCJ95, JS09, JL12b, KS10, KH15, KYY^{+}12, KT07, KAZ01, KS12, LC97, LBB08, LA95c, LK13, LKZ^{+}04, LE06, MHRR12, MLS12, MH97, MW06, MAS09, PT96, PV10, PPV12, PWK^{+}13, RB02, SMGP15, SYDM09, SCN12, SC09, SL14, SK12b, SMM11, SKUB12, SS03, SL07b, SAS^{+}16c, Tha04]. access [TS08, TH97, VA06, VA07, WBEGS05, WZL^{+}13, XHN04, YKZ^{+}13, YJ15, YHE04, YM05, ZSK12, ZLSK15, dAF04]. Access-Based [DPTP24]. Access-Point [LWL17]. accessed [CDI^{+}04]. accessibility [ABA^{+}16]. Accessible [AZP^{+}23]. accessing [LO02a]. account [SL15c]. account-aided [SL15c]. Accountability [HRLY21]. Accountable [XHZ^{+}19]. Accounting [BSSU18]. Accumulation [KS19, XHK^{+}05]. Accumulation-based [XHK^{+}05]. Accumulative [GVGV17]. Accuracy [DKN21, HCFC20, LL18, PJDS18, TCTP20, XLW^{+}18, AD96, BM09]. Accuracy-Aware [TCTP20]. Accurate [CXK^{+}23, DYW^{+}16, FBRL18, GDC^{+}17, HZL^{+}23, LXW^{+}20, LYY^{+}22, LP24, LCZH17, LCL^{+}20, SL16a, SXQ^{+}23, WMCW22, WHLL23, XCQ^{+}23, XLW^{+}17a, XWW^{+}18, XPW^{+}18, XWW^{+}19, XCW^{+}20a, XCW^{+}20b, XOW^{+}23, YYT23, YZL^{+}19, ZCF^{+}24, ZDB^{+}17, GS97, HQY^{+}16, KS09a, KZ97, SL15a, TZZ^{+}14, XXBC14]. Accurately [MRM17]. ACF [RAA^{+}24]. achievable [JP09, KN05, SGR13]. Achieve [LL17a, CCG00, Kok10, XCR11, XCR15]. Achieved [YM16]. Achieves [CLS^{+}18, HMNK13]. Achieving [AZ03, BFF07, BM08, CNG^{+}16, CKLS22, EW08, GXL^{+}21, HLX^{+}15, HL15, JGS^{+}15, JZC11, KLO97, KLR^{+}20, LCZH17, SGD05, Van17, WGZC21, XHC^{+}18, ZRP^{+}22, ZYS^{+}23, ZBdV23, ZZHZ13, JGLS14, LLL06, MS03, NTS12, SS03, XME15, ZS05]. ACK [CQW^{+}18, DXX^{+}23]. Acknowledgment [LZX^{+}21, SR02]. acknowledgments [KWH11]. ACORN [APB^{+}13]. Acoustic [DPG^{+}24, GSL^{+}24, JHLW24, LYC^{+}19, QLF23, ZZL^{+}21, ZPCS11]. Acquiring [ALY^{+}20]. Across [ALY^{+}20, JWL^{+}18, LS17, WHC^{+}22, ZND^{+}16, BCMR04, EST93, SW04, ZNZT16]. Action [HVT18, TES19, LHK^{+}12]. Action-Based [TES19]. Activating [ALYX22]. activation [AABD13, KKJ06, KBS11]. Active [DXX^{+}23, SYW^{+}22, BCP00, EVF06, cFSKS02, HXLZ11, HGG06, KS04, LBS05a, LAJS07, SBDR08, ZAS12]. Activity [FHQ^{+}17, NZW24, CAO11, Tre11]. Activity-Oriented [NZW24]. Actor [XYT^{+}21, GAA08]. Actuator [SSY19, RKNS10]. Acyclic [HR14, LCP^{+}20, SPLM17, CER12, RGKS10]. Ad [BVBV17, ÇTD22, CDW19, Gan20, GDC^{+}17, HL99, MYMY17, PP17, QJZ^{+}16, RZS14, WCC14, AHK08, AS07a, AS07b, BCGC15, BCB99, BNJR12, BNJ16, CE09, CZF^{+}16, CFM13, CDM13, CW10, CMGL11, DLL^{+}11, DBT05, EFK07, GMP08, GGL09b, GGL09a, GGH11, GT02, GMYP16, HHL06, HS06a, JS11, KK07, KDHK15, KWZ08, LH07, LPKF10, LMP08, LZF09, Li09, LLLT10, LPF12, LLNC09, LSMS06, LR09, LCL^{+}12b, LNL^{+}16, LKZ^{+}04, MQ05, NL07, PS05, RM08, RSR10, RKNS10, SLP07, SPH04, SRR08, SMS07, SSHK11, SS10, SL12, SS07, UN11, WCY04, WTS^{+}13, YD07, YLL10, ZSFZ11, ZW10, vRWZ09]. Ad-Hoc [ÇTD22, LSMS06, SS07]. ADAM [AKS^{+}13]. Adaptation [CGZL20, EAH^{+}18, GA24, GLL^{+}18, KW17, Nee19, QWL21, SUS20, TL16, WZH^{+}24, XWJ22, CK10a, FSM14, GM03, KVR02, LRG10, PA12, PD16b, PLL13, SMGP15, fTL06, WZR08, WH11]. Adapting [MGCK15, LyT98, VG04]. Adaptive [BOY00, BSG^{+}18, BTD^{+}17, BNJR12, CZC^{+}22, CYTH22, CqLL98, CQLW22, CZTX23, CYX^{+}17, DPG^{+}24, DZL^{+}20, DHS^{+}23, FBQ^{+}23, FK99, Fuk20, GXW^{+}19, GLM^{+}16, HK24, HHL^{+}19, HGZ^{+}23, IYYI18, JSZ14, JK21, KL08, KVF^{+}12, LCM04, LCK^{+}18, LLY01, LK16b, LWW^{+}19b, LXL^{+}22b, LHL^{+}23a, LZX^{+}24, MGS^{+}21, Mil98, MA98, NZTD02, PILR05, PRH17, QYC^{+}24, RUH^{+}18, RR93, RAA^{+}24, SXQ^{+}23, SK21, SGJ17, TCTP20, TL16, TWTD17, TJD23, WJ17, WCC14, WLL^{+}16a, XCC^{+}17, YXC^{+}18, YJL^{+}19, YL97, YWW^{+}24, ZCZ^{+}21, ZCPG^{+}23, ZCF^{+}24, ZLZ21a, ZWJ^{+}22, ZZZ^{+}24, ZXW^{+}19, ZHWH21, AK00, Ada98, AJDH01, AAM05, AB05, AKS^{+}13, ABJ^{+}13, BCP13, BCMR04, BL94, CN10a, CDFG06, CLA07, CGK10, DM14, DRR98, EL11, EF08, FGM^{+}13, HCL09, KMT05, KMR95, KT06, KAZ01, KS04, LL09, LSV01, LZZR12, LCL^{+}12b, LS03b, ML06, NL99, PYL99, RD11b, SKY10, SPH04]. adaptive [SM05, ST13, SV11, VAS00, VA07, WS06, WD05, YSZL15, YHE04, YZBR14, ZLC12]. adaptive-rate [LS03b]. Adaptively [ZRD^{+}23, GL93]. Adding [BVL^{+}19]. Additions [VCVC17]. additive [GR12, RS07, VR13, XSZ^{+}07]. Address [HWHW18, KGH^{+}20, NKL^{+}23, CGW^{+}12, EFK07, GIL^{+}15, KIR06, MPL09, RW07, SMP^{+}14]. address-light [KIR06]. addressable [LMT16, SG96]. Addresses [KRRR17, SYW^{+}22, KLPS06]. Addressing [LZL^{+}20, SVG16, AQJRS16, CDM13, LK95]. Adjacent [BTH11]. Adjustable [DLC^{+}18a]. Adjusting [AMS22b, EF17, SAS^{+}16b]. Adjustment [LBGL20]. Admission [BCE^{+}19, DSYN24, GJCB18, ASCG08, AZR97, BLCT97, CCL99, CNP13, DM96, EF08, EM93, FKT98, FMT03, GKPS06, GT99, GT03, JDSZ97, LLD96, LAN97, LWF96, MH02, ML12, NKS08, PDSK04, QK01, RV01, SR01, WD05, WWL02, XSC01]. ADMs [SZ07]. Adoption [NMD^{+}17, JWSH15, SJGH10]. advance [CV12, CFS09, CFS11, LW13, TCPV13]. advanced [IK07]. Advantageous [ZML^{+}19]. Adversarial [HLP^{+}16, HWLL21, TSS21, XOW^{+}23, YXH^{+}21, ZCPP22, LJA14]. adversary [ZAS12]. Advertisement [LCH22, LZL12]. advertisements [KLMW11]. Advertising [YCGH17]. Advising [YWRK19]. AEGIS [ZWTC16, LTS10]. Aeolus [HZB^{+}22]. Aerial [CJ18, WCL^{+}22]. affected [BCP13]. Affects [VBHT17]. Affinity [YN20, SKT96]. affinity-based [SKT96]. After [BCLS17, ZAW^{+}22, SZM08]. Against [CLX^{+}22, HS19, JZW^{+}18, LMSS24, LZZ^{+}22a, OPGT16, SS21, YXL18b, ZCPP22, ANSX13, AC09, BKLS08, FTV^{+}10, GJVZ06, KRLL11, LWL^{+}11, LLY^{+}12, OF11, WZR08, WJS07, YLLY05, YKGF08, YGKX10]. Age [BU21, BSS19, GKCR21, HDF19, KSUB^{+}18, KSM19, LLL^{+}22b, MAE20, MKAE20, NSY23, OS21, TKM20a, TW23, WD22, ZKEN23, YWLL09]. age-based [YWLL09]. Age-Dependent [BU21, NSY23]. Age-of-Information [TW23]. Agent [PWL^{+}24, QYC^{+}24, WLY^{+}24]. agents [HBS96, La16, LMG04]. AggreFlow [GXL^{+}21]. Aggregate [CSD22, DJM97, LMS04b, QK01, SG13, TMH11, XG05]. aggregate-level [LMS04b]. Aggregated [TXL^{+}18, KL03, LRJ08]. aggregates [JS06, RBGK03, SS05]. Aggregation [BSSU18, CKA16, CGC^{+}18, FZX^{+}23, GHK^{+}23, JSXN18, LNM^{+}09, PJDS18, RPPA22, SVL^{+}16, WLD^{+}24, YZY^{+}20, YYB^{+}22, ZYH^{+}21, AS01, Cob02, FK03, HCL09, HY08, JS14, LNC04, OÇ10, PT10, TX08, TMP07, WMYR16, XLR13, XLWT12, YAA09]. Aggregation-Based [YYB^{+}22]. Aggregator [FBRL18]. Aggressive [ZWH^{+}17, EW08]. Agile [TL16, ZLZ^{+}21b, LCG^{+}14]. agility [VVP^{+}13]. Aging [JYC^{+}16, KLC^{+}18]. Agnostic [BCC^{+}17, DRW^{+}22, SXEZ21]. Agreement [JWZ23, LYZ^{+}23b, WCH^{+}24, XFCW18, LLY06]. agreements [LGGZ10, SYR05]. ahead [GSW99]. AI [ZLY23]. AI-Based [ZLY23]. Aided [CGYZ17, HCL18, LW22, YN19, ZGLC20, SL15c, SLL15]. AIDPS [DPG^{+}24]. Air [BBR19, SRMB^{+}23, XWL^{+}18, ZWGC17, KRH^{+}08]. airborne [MHRR12]. AirNN [SRMB^{+}23]. AirSync [BRM^{+}13]. airtime [CSN06]. Akamai [SCKB09]. alarming [BGK^{+}16]. ALBA [VHNPM96]. alerts [VG08]. Algebra [CBSK07, Sob02]. Algebra-based [CBSK07]. Algebraic [DMC06, KM03, Sob05]. Algorithm [ADT22, ER20, BBHH^{+}18, CLS^{+}18, CWH^{+}16, CLV17, CQLW22, CWZY21, CMP^{+}14, EAH^{+}18, FMH^{+}21a, Fuk20, JLW^{+}24, JD19, KA20, KSRW22, KLE16, LMODF18, LCSS17, LCS^{+}18, LW20, MLB21, NTD17, NLNL16, QLF23, RBPS21, SAMB18, SG17a, SZMD17, SKA^{+}18, SJSB22, SBLS19, SRB^{+}20, WLX^{+}17, WDL^{+}23, WWL24b, XXW^{+}23, YZL^{+}19, YN20, YN18, ZWZC23, ZJWY17, AA93, AEB02, ASCG08, AAV09, AOM04, BTC01, BS08, BSS11b, CHCH00, CLK01, CLW95, CAL09, CK09, DRR98, EAB01, EAB02, GW94, GLAM97, GVC97, GL10, HL05, HLW13, IPG97, JDSZ97, JMS08, Jia98, JW10, JYT^{+}15, JLS09, KJF^{+}00, Kar03, KD00, KG05, Kri14, KLNS93, KS04, LCY96, LLLS07, LGC16, LS06b, Lev95, LAN97, LHB^{+}05, LCW^{+}15, LDGL13, LW13, LL99, MBA06, MOY00, McK99, MMC05, Mil98, Mne08, NST01, NM06, PZGLA98, PHL15, Pil01, RMM99, RS00, RW96]. algorithm [ST09, SSHK11, SNS12, SZN00, SC10, WC08, XGF^{+}14, YSL^{+}14, YSTL11, ZA95, ZFC13]. Algorithmic [ABBH^{+}16, BSP21, CKS17, LFC18, vRWZ09, BCN02, KWZ08, Tha01]. algorithmically [YRRR12]. Algorithms [AGBS23, AP17, BCER20, BBO^{+}05, BBR19, BRK^{+}22, CCK16, CKA16, CJV16, CGC^{+}17, DRMP18, DMMS14, DWCZ17, DCZG19, Gan20, GCWC17, GFW^{+}18, GJWZ16, GHW22, GHW14, GSM16, HH17, IKS17, KSM19, KSSD24, KHYA20, KRSY02, LXX^{+}17, LT16, LTP10, MLJ^{+}22, MGS^{+}21, MKS17, MJ17, NGL22, PBT^{+}20, PG18, RpLP^{+}17, RR19a, SS17, SJZ^{+}24, SZ22, SG05, SPM^{+}17, SBTH19, SGJ17, TAH17, TW23, VLM16, WCW19, XL99, XL23, XLX^{+}21, YK23, YLL21, YLWH20, ZYL^{+}17, ZLL24, ZQL^{+}23, AA99, AS08, AZ11, AC06, ARS16, BBG11, BCP00, BO07a, BB94, BV96, BCN02, BZ97, BRS10, BSYS12, Bor05, BLB10, BGPS06, BLS07, CN10a, CRB09, CBSK07, CLSC15, CRB12, CKV11, CL08, CGGS97, CFS09, CK10b, CBLVW06, CYL16, ES96, EOSM10, EVF06, cFSKS02, GS13, GV97, GO99, GLS09, HIM07, HL15, JAS10, JW11]. algorithms [JGMB03, KWCR10, KA98, LCM04, LL09, LDFK12, LMR07, LH05, LWCY12, LNA07, LR09, LÜ14, Low03, MBL10, MSA^{+}16, MPL09, MR02, Mil95, ME96, Mod99, MMS01, MJ15, MLC07, NSS96, NST00, NTS12, PM96, PPSV13, QZZ^{+}13, RRG10, RL93, RVR93, RB09a, SK12a, SFAS05, Sob02, STC12, SIYL09, SV98a, SV98b, SV98c, SS05, SR14, TCS13, VK04, VAGT13, VL10, VAS00, Voi07, WPL06, WJLH06, XY10a, XL05, XCX^{+}06, XZTT08, Yua02, ZXTT08, ZCW15, ZL16, ZS05]. alias [KHLC13]. aliases [GS09]. AlignTrack [CW23]. All-Channel [BMBK21]. All-Optical [WJ17, ZWZ^{+}24, SAS96, ARK09, BTH11, CV12, CL05, MBLN93, MA98, PG95, Pan99, RSM09, RS95a, SMG05a, SS04a, TWHR11, THBR14, WQC06, WS05, XL99]. All-Terrain [CXK^{+}23]. all-to-all [LS06c, PEA09, ZQ99]. Alleviating [WLL^{+}16b]. allocating [XL99]. Allocation [AMCD19, AMSB^{+}24, BDR22, CWAO21, CYH^{+}18, CBHS20, CL19, CGAC20, DEP17, DHHD18, DTN^{+}21, DRQ^{+}16, DJB^{+}22, ENT^{+}24, FMSM24, FHMS18, FMPS20, HKLM17, HSO19, JTL^{+}17, JTL^{+}18, JD19, KK16b, KRS^{+}17, KKH^{+}22, LFZ^{+}22, LAV16, MZK^{+}17, ML23, MKOY24, NLB19, PL17, PPTP21, RTLC17, SJ21, SC18a, WDL^{+}23, WTJR22, XLL^{+}20, YYB^{+}22, YK23, ZLG^{+}20, ZSS^{+}20, ZWJ^{+}22, ZLZ^{+}23, ZRH18, ZGLC20, AS08, AK15, ACKZ14, BLV10, BF01, BGK97, BI00, BS08, BLEM^{+}12, CDFG06, CRL96, CSSJ14, CJJ09, CLK01, CLA07, CL13, CAL09, CLL^{+}14, CF98, CR14, CG15a, DS04, DGK05, ES07, FGK10, FP14, FMT03, GSKR99, GM00, GLLJ16, HZC07, HSS08, HG14, JS11, JWSLC13, JJSS04, JBR16, KKEE13, KS10, KK00, KMHS09, LOP97, LM95, LMS05a, LMS06, LNB01, LMG04, LCH95, LSZW13, LLE15a, LTS05, LPCVC13, MBL10, MCS99]. allocation [MPF^{+}15, NDGL06, NM09, NMR03, PLD16, STKL01, ST09, SSAK12, Smi02, SK97, TNRP11, VGP14, Wan04, WSW12, Wil96, WM95, XL11a, YMR00, YJ15, YZBR14, YJH05, ZB95, ZS05]. allocations [Low00, SSZ03]. Almost [LSDT19]. Aloha [BBF18, HYLS21, CL16b, LH95, LYL^{+}22b, WZL^{+}13, IZC00, LZC09, MMR09, MMP17]. alone [GV06]. Along [CCK16]. Alpaca [KRRR17]. alphabet [CFS06]. Alternate [Zap04, RM02]. Alternates [RS21]. Alternating [CH20]. Alternative [LPM23, OdG97, SSG18, WF93b, CT96, MM13, SD00]. Alternatives [DNCK20]. Always [RGKS10, ZMLL21]. Amazon [CGYZ17]. Ambient [GSH^{+}22]. AMI [GKB^{+}16]. Amidst [GA24]. Among [CLS^{+}19, DSYN24, DCN^{+}19, LGZ^{+}23, SJZ^{+}24, WLY^{+}23, LZXF14]. amortization [MSWL06]. amount [SSZ05]. amplifier [RIM98]. Amplify [IK09, BJ15]. Amplify-and-forward [IK09, BJ15]. Analog [WM16, WLD^{+}24, ZZ17]. Analyses [FAF^{+}17, TT17, NPQ06, PYL99]. Analysis [AB23, ALR^{+}24, ATE21, AAA18, AAAR19, AVPG14, BL15, BFG^{+}14, BRK^{+}22, CG21, CAZG20, CSC94, CLL^{+}18, CB11, CMR17, CLZ^{+}20, CG04, CKZC19, CRS99, DM03, DSA^{+}14, DKC^{+}15, DKM^{+}17, DTN^{+}21, ER23, FLH^{+}17, HH18, HYLS21, HRM22, IM03, JR21, JSW^{+}20, JE18, KV98, KP21, KS09b, LR22, LTZ08, LY10, LCP^{+}20, LZX^{+}21, LLCJ22, LCH20a, LFY^{+}19, LXC05, LZC^{+}17, LFL^{+}23, MLS^{+}23, Mar04, MS17, MBL19, ML22a, NSY20, NSP^{+}16, NMH99, OKAS23, PJMM22, PB93, RMPG16, RS04, RW93, RSB01, RZVZ06, RLZ10, RA95, RSLL23, RW95, Rum93, SARM24, SQ16, SS17, SR18, SLD^{+}23, SWH19, TL22, TYP^{+}15, VBC^{+}17, VHT21, VLM16, WSXL16, WVZ17, WCCM18, WLS^{+}18, WWW20b, WD22, WWYY18, WCY00, YXAZ^{+}18, ZMW^{+}22, ZFW14, ZSH^{+}16, ZFW^{+}17a, ZLW^{+}20, AZLB16, AS07a, AS07b, ALMR14, AOM04, BBM93, BO00]. analysis [BLPS10, BC01a, Bar95, BCL12, BBMELH08, BT93, BLB10, BH06, BD96, BL94, CFPP96, CJ14, CH04, CC95, CRL96, CLM99, CMM95, CZFF98, CK10b, DT15, DLH^{+}14, ENW96, FTV^{+}10, Fan05, FGK10, FHT^{+}10, GMP13, GYB^{+}04, GSKR99, GP94, GXWW11, GMWD13, GMD15, GS10b, GS11, HS03, HGE04, HSE97, Hon94, HBH93, ITSO01, IW08, IK09, ILS97, JS12, JRL15, KVR98, KHG^{+}14, KqL99, KS01a, KK03a, KH15, Kop96, KqL98, Kum98, KAMG07, KSM05, KT08, LS93a, LBRA05, LSV01, LBB08, LS93c, LA95b, LYS93, qLH93b, qLH93a, LD95, qLlH97, LK05, LZ09, LW13, LM96, LRL07, LRL08, LJ09, LT94a, LR03, LCW05, LMP96, LT94b, MBA06, MMR09, MCR10, MH02, Mar96, MBC^{+}94, McM95, MDMM09, MP94, Nee09, NL07, NR13, PG94b, PWHL16, PJ13, PS98, QCS07, QY12, RP06, RKA08, RG98]. analysis [RRB06, RW96, SLC^{+}07, SL94, She95, SM00, SMM11, SMSM06, SMP^{+}14, STC12, SV98b, Świ96, TMMS01, Tia05, TdWC^{+}94, VR13, VC14, VA09, WL07, WSKV08, WHW^{+}11, WVG12, WJZ^{+}12, WDCL15, WTSW97, WY95, WS08, WLR10, XHN04, Xin07, XWG14, YRRR12, YBG^{+}12, ZS03, ZS04, ZNN^{+}10, ZKL11, ZHLL06, ZFC15, DKL01]. Analytic [SKV03, AdE07, ES03, KL09, PMW10, Pax94]. Analytical [BK17, BP96, CL19, KGdV^{+}21, MSRG18, PPV17, SS16, TCPV13, ZML^{+}19, ZMLL21, AA04, CDPLCA16, Fan05, KEAAH08, LS97c, LMS04b, LC94b, ZY07a]. Analytics [FDM^{+}17, KSSD24, LX21, RWL^{+}22, ZWJ^{+}22]. Analyzing [BCR^{+}12, CKR^{+}09, JWSH18, LYL^{+}22b, PT94, SW04, YHCL21, ZLB17, CS98, ZLC12]. anchor [HA96]. AnchorHash [MVB^{+}21]. angles [PLR15]. anisotropic [LL10]. Anomalies [VBC^{+}17, WWYY18, XBM^{+}23, KR08]. Anomalous [VSR11, LBP^{+}16]. Anomaly [BDWS12, MWW^{+}21, NDN^{+}18, PBGMFM22, VRR24, XLW^{+}17a, XLW^{+}18, ZZX^{+}21b, MG16, PS09, TMH11, XY09a]. Anonymity [CS17, JV17, KHH^{+}18, MV16, TWS^{+}22]. Anonymizability [FZQ^{+}22]. Anonymization [CGL16, JLSB16, RW07]. Anonymizing [FZW^{+}20]. Anonymous [CCF17, LXL^{+}17b, LMP96, ZCZ^{+}20, ZFW^{+}17b, MYYR13, VT12]. Answering [TBV^{+}13]. Antenna [TAH17, PLR15, STKL01]. Antennas [CLV17, ZP18, KAZ01, LTS10, LZF09, SS07, ZJS^{+}12]. Anti [LW17, PBKG11]. Anti-Inference [LW17]. anti-jamming [PBKG11]. Anticipating [XXN^{+}19]. Any [TG96, GO02, YASS15]. anycast [KLSS10, KLS11a, LMP08]. anycasting [ZAFB00]. anypath [DFGV11, LDFK12]. AoI [CGC^{+}24, LZC22, LYKT21, TW22, WLY^{+}24]. AP [GB18, KLC15, LWC^{+}14, PJDS18, WQY^{+}17]. AP-Atoms [PJDS18]. App [TES19]. Appearance [SCW^{+}21]. Application [DPT^{+}18, HK24, JR22, LAV16, Le 18, LBZ^{+}20, MCM^{+}23, NRB22, PCW^{+}16, SKZ03, WPL06, WLD^{+}16, WDL^{+}23, WLLZ16, ZAFB00, ZCZC17, BL15, BLCT97, BLS07, DM03, DW11, FJL^{+}97, GP96b, KL95, KLT15, LWL^{+}11, MH02, RPGE04, RSU^{+}09, RW95, dSeSGM95, Tre11, WEK97, XY09b, YW07, ZNK^{+}13]. Application-Aware [DPT^{+}18, YW07]. Application-Awareness [NRB22]. Application-layer [ZAFB00, BL15, BLS07, LWL^{+}11, XY09b]. Application-level [WLLZ16, RPGE04]. Application-Oblivious [LBZ^{+}20]. Application-oriented [WPL06, GP96b]. application-specific [WEK97]. Application-Tailored [MCM^{+}23]. Applications [AWH^{+}22, BBHH^{+}18, CVHM22, CBZ16, CJL^{+}19, DSM^{+}17, DGLM16, FMH^{+}21b, FKCA18, GXW^{+}19, HCW^{+}16, KL12, LTDM17, LYSZ16, LSSC22, LDY^{+}16, QCMY16, SS16, TXW^{+}19, VHT21, WLS^{+}18, WJH^{+}21, XL23, XLX^{+}21, XXW^{+}23, YXZ19, ZCW^{+}22, AAM05, ACC^{+}94, AS02, BRISCSP11, BMS14a, BH06, CBSK07, CJH^{+}11, CZZY12, CPS^{+}12, CH15, CDS02, DFMR15, FHT^{+}10, GCZ98, HS06a, HLSG04, HL05, Jia98, JYT^{+}15, KCCM16, LL95, LZ06, MR96, NSW11, PGV16, RL07, RHMF16, SZG^{+}13, SMLN^{+}03, TLS^{+}12, WXBZ04, WS06, WMS09, Wu94, WWL^{+}15, YL97, YKR11, ZT12, ZPCS11]. applied [BBM93, HBH93]. Applying [SP94, YYFC24]. Approach [ACLX17, AY20, AWM^{+}20, ABMT23, AM19, BES22, BB16, BAB20, BBR^{+}22, BFG^{+}14, BDR22, CZP18, CNM20, CLQ^{+}19, CLZ^{+}20, DLW^{+}17, DNS23, DCN^{+}19, DMT^{+}19, DZL^{+}20, DF20, DLC^{+}18b, DZL^{+}18, DME23, DJB^{+}22, EMAL17, FLBR^{+}19, GYSPR14, GSKN18, HSS^{+}21, HZ20, LZL^{+}21, Li24, LPS19, LGS^{+}23, LL17b, LN19, LZZ^{+}22b, LYL^{+}22b, LDL^{+}22, MGLH18, MFR^{+}20, MMP17, QDD^{+}17, RRS22, SNZ^{+}23, SS16, SdVK16, SYZP19, SSG18, SLSC20, SX16, TY18, WYL23, WHLL23, WN16, WTJR22, WBM^{+}18, XWW^{+}18, XOYL20, XWYL23, YDLT18, ZYH^{+}21, ZGZ22, ZHCC24, ZLN^{+}17, vDJJ^{+}22, AP93b, AS94, AK01, AA96, AF99, AdE07, BLV10, BGSSW13, BO07b, BCN02, BYH^{+}15, BLEM^{+}12, CSMW02, CGM04, CM03, CZFF98, CS98, Coh94, CK07, CN09, DLT16, DMC06, DJM97, ES03, Fan05, GLAMM11, GG94, GSA15, GT03, GLLJ16, HD07, HKV^{+}13, HBU95, HL15]. approach [JLM15, KL13, KKS^{+}08, KLS03, KM03, KR99, KWZ08, KL09, LM13, LCH^{+}06, LEYS11, LHZ^{+}16, LTY06, LS06c, LFV10, LyT98, LS06e, LMT10, LSXS16, LV93, MLLY06, MRM99, MQ05, MLT12, MSBZ10, NL16, NZTD02, OS05, PM09, PG93, PG94a, PA12, RSM09, RVS^{+}02, RSS09, SLP07, SHZ16, SK10b, SK12b, SBP03, SM05, SKCW10, SBDR08, SPB16, SA01b, TT09, TK12, TWHR11, VNS02, VT12, WMYR16, XXBC14, XSHS12, YXF^{+}13, YMO97, YMKC08, YWZZ16, ZM09, ZQ00, ZWDS00, ZRLD05, ZRP00, ZY16, ZWO^{+}96]. approaches [DXT^{+}12, EM09, JK15, LT02, LESZ98, MLT11]. Approaching [EPS21, JW11, JW23, OY13]. Appropriately [ABS^{+}16]. Approximate [CSD22, DKN21, Hon94, RAA^{+}24, SNC23, Świ96, WYL^{+}22, AAG14, BBM93, CKR93, LBRA05, SZG^{+}13, SSZ03]. Approximating [LTS05, LWK^{+}16, PBV17, RCGT06, WLS97, ZWL^{+}16, CD96]. Approximation [AP17, BRS10, BLS07, CWH^{+}16, GZL^{+}17, GFW^{+}18, KWCR10, Kar10, LXX^{+}17, NTD17, PPSV13, SK12a, SZMD17, SGJ17, WLK^{+}17, XNHM22, XLX^{+}21, XXW^{+}23, YLWH20, ZY21, ZTH^{+}23, ZLL24, DXT^{+}12, JLRS16, LB04, SZ07, XZTT08]. Approximations [ER23, RS19, SBGJ18, MHXT10, MM94, RV01, SBD11]. Apps [MKG^{+}17]. Apr [ZND^{+}16]. AQM [DXX^{+}23, EW08, LBS05b, SCR08]. Arbitrarily [ADT22, XCZL20]. Arbitrary [GLS21, MZZ^{+}23, VPC17, XCC^{+}17, XZC^{+}19, BLEM^{+}12, HH10b, MKS16, MR98, MOY00, MFB99, OY95, PEA09, RLA06, RS97b, TNF97]. ARC [AA04]. ARCH [KZDM07]. ARCH-based [KZDM07]. architectural [ZWO^{+}96]. Architecture [ANTR17, CCC17, CWM^{+}17, CSR^{+}20, DRW^{+}22, ENT^{+}24, JPS^{+}17, KA20, LSL^{+}18, MAPZ18, MRJ20, MKG^{+}17, NRB22, RD11a, WLC^{+}20, YDY^{+}24, YXL^{+}19, BKH^{+}93, BCL10, BSS11b, BS00, CT01, CSS^{+}14, CEFS99, CS99b, CS00, CL08, DDPP00, DEF^{+}96, HA97, HW99, HXLZ11, IM03, Kim94, LSLL14, LK10, LCG^{+}14, LXX^{+}14, MD04, Mar96, MSH95, OKM94, Pad95, SP94, SLG^{+}16, SH07, SSZ03, fTL06, WZLX12, WJLH06, Wu94, YCB07, YWA08, ZAFB00]. Architectures [AAR18, EMAL17, EL24, LXLC20, PKVI17, SGH^{+}19, AMKY99, CLA07, CFS09, CT96, GLH95, RS04, RVR93, RG98, RSB01, WF93b]. Area [BFG^{+}14, ÇTD22, CCX^{+}23, DQYG23, LHY^{+}23, MFT^{+}20, SRI^{+}18, WZLM22, ZRP^{+}22, AIN^{+}15, BSNI06, BCC07, DEF^{+}96, ES96, FCAB00, GT00, HL98b, HL05, HK96, Jia98, KV96, KKM^{+}97, LM01, Med95, MBRM96, Pax94, PF95, RVS^{+}02, YNDM09, ZWDS00]. Areas [BPVRSP16, CCW^{+}17, DLLL16, VG04]. Ark [LFL^{+}23]. Armed [AM19, FM23, HVT18, GKJ12]. Army [NLRS21]. ARQ [CFG08, CGK10, KEY99, LZ09, SEK15, Spi97]. ARQ/FEC [KEY99]. Array [CDKZ21, KAZ01, TYJ16, WZLX12]. arrayed [NPQ06]. Arrival [DYJ^{+}23, ODT09]. arrivals [CFG08, LBS11, vDP93]. Art [GNK^{+}21]. ARTEMIS [SKG^{+}18]. Artificial [ZGY^{+}16]. AS-aware [AYM14]. AS-level [GIL^{+}15, OPW^{+}10, SFFF03]. Ascent [SNC23]. ASHs [WEK97]. ASN.1 [TNF97]. Aspects [LFC18, VCM04]. aspiration [JKJ13]. aspiration-based [JKJ13]. Assessing [GCM^{+}16, MTK03, NZCM11, XB07, DXT^{+}12, PS09, SNXT13]. Assessment [ZLHM22, CJ07, DT15, LJC05, WK13]. assigned [AJ06]. Assigning [BPVRSP16]. Assignment [AdSD16, AAF^{+}16, BSRdA16, DGW^{+}17, FM20, GYLH17, HDF19, LFZ^{+}22, MS95, TAH17, WLX^{+}17, WZZC17, ZGLC20, AZ09, AAV09, BPPP12, BB94, BB95, CV12, CM05b, CMV10, CL05, HRCW08, HBU95, KT07, LHL15, LMS06, LS01, LHM02, LR09, MK98, NBTD07, OB03, PT96, RS95a, RPF^{+}14, SMG05a, SMG06, SSHK11, SKCW10, wTjCjC97, WQC06, XWWC16, ZOM03, ZA95, ZQ00, ZY07b, ZT12, ZM04]. assignments [Hu93, Tha01]. Assisted [FLH^{+}17, GHK^{+}23, HNP23, SRCT23, TZL^{+}24, WZW^{+}20, WSX^{+}23, XWJ22, XXZ^{+}23, YWW^{+}23, ZCF^{+}24, ZML^{+}19, ZWC^{+}24, AJF11, BJY11, CY14, GZT03, HPR06, PPV04, RPGE04, RHC^{+}12, WLCW16, WLY^{+}24]. Associated [YYB^{+}22]. Association [AP17, GCMP20, HBSX20, LWC^{+}14, LPS19, SSNS17, AKSS12, AWFT15, BHL07, BDWS12, KDYV12, RD11b, SKS16]. assurance [BB06]. assured [WMYR16]. Assuring [YDW18]. Asymmetric [HPP^{+}23, HKS16, PKVI17, WHZJ20, XZL^{+}21, LCW^{+}15, Ram96, RM08]. Asymmetry [LHL^{+}23a, KS09a]. Asymmetry-Aware [LHL^{+}23a, KS09a]. Asymptotic [LZF09, LZC^{+}17, SMSM06, TL06, YK23, ZH08a, ZFW14, AEJV13, BCGC15, JGLS14, JGS^{+}15, KS01a, LLW^{+}14, PL02, SWL06, WL07, ZH08b]. Asymptotically [FM20, GLS21, LSS07, PLS07, SX10, TJHL21, CSSJ14]. Asymptotics [MHL19, JMMT12, SD15a]. Asynchronous [BESW08, CLWZ17, CWL^{+}21, CLW19, GR20b, Kri14, MSP^{+}07, MMP17, NLNL16, SARM24, WN17, WCM^{+}21, AK01, BJY11, BJ15, CK11, JC13, KLS11a, OSW97, Tur09]. Asynchronously [MAPZ18]. Atlantic [MHRR12]. ATM [PK01, AS94, AKS96, AJDH01, AMKY99, AL98, BBM93, BGVC00, BLCT97, BM97, BIS00, BI00, BL94, BS00, CT95, CFPP96, CU95a, CC95, CRL96, CqLL98, CHCH00, CC96, CPSWL96, CDM93, DM95, DK98, DJM97, FC99, GP96a, GCZ96, GCZ98, GH93, GM00, GP94, HW99, HLG94, HK96, IMG98, JK96, KV98, KKM^{+}97, KJF^{+}00, KR00, KMS^{+}01, KWC93, Kim94, KL95, KqL99, KEY99, KS98, LMR99, LS93c, LM95, LA95b, LLD96, LMSKZ99, LS97c, LMS99, LV93, MR98, MSB97, Med95, MMR96, MR96, MG95, MK96, MK98, NML98, NMH99, OWMM97, Pad95, PYL99, PB93, PG94b, PS98, RRK96, RLKT98, RB95, Ros96, RL94, SMT98, Ses97, SV99, SCY98, SS98, SBP03, SG94, SSD93, SC95, SK97, SDW00, SZT01, THP94, TdWC^{+}94, TG97, WF93a, WLL01, WM95, XM99, ZVN99]. ATM [ZSSK02, ZF96, ZKO93]. ATM-based [RLKT98]. ATM-oriented [ZVN99]. Atomic [TLS^{+}12, YLYL17, GHR14, LO99, YL16]. Atoms [PJDS18]. attachments [LT94a]. Attack [CH21, FWN^{+}22, GWYS19, GCZY18, HLZ^{+}21, LZZ^{+}22a, LJHB18, XNHM22, YFM^{+}22, YLK^{+}17, ZXW^{+}19, KSA12, KSV07, Kon06, LMR07, LLY^{+}12, SKCW10, WS05]. attack-aware [SKCW10]. attack-resistant [LMR07]. Attackers [HWJZ21]. Attacks [ABBH^{+}16, ABBF19, ACDP17, CLX^{+}22, CSSG23, DEP17, DAFZ^{+}18, FLS^{+}22, JZW^{+}18, KSC^{+}23, LSSC22, LZC^{+}24, LHL^{+}23b, OPGT16, QYZX22, QLQ^{+}22, RSLL23, SCN^{+}22, SS21, SVG16, WCCM18, WWW^{+}18, WXC^{+}24, XCL^{+}22, XS21, YXH^{+}21, ZLX^{+}21, ZCPP22, ZSLZ21, AHK08, AAS14, AC09, CLSS09, DT15, FTV^{+}10, FAB12, KVF^{+}12, KK06a, OF11, RSU^{+}09, TEML09, WZR08, WNV13, WXW15, XY09b, YRRR12, YLLY05, YKGF08, YGKX10]. Attained [HV22]. Attaining [CS17]. attains [MAN15]. Attention [LAL^{+}24]. attenuation [XK06a]. Attitude [WZLM22]. Attitude-Aware [WZLM22]. Attribute [ER20, CDW19, KRRR17, LZL^{+}21].

Attribute-Based [CDW19]. Attribute-Encoded [KRRR17]. Attributes [WZH^{+}18]. Attribution [NDN^{+}18]. Auction [CZD^{+}22, GCD23, LLM^{+}24, NBV17, NS21, SZW^{+}16, TPW^{+}18, WHC^{+}19, ZYH^{+}21, ZFLC18, CL13, HGW^{+}16, IGHT15, KS10, WHTC15, ZWTC16]. Auction-Based [LLM^{+}24, CL13]. Auctions [DRQ^{+}16, ZHW^{+}17, AAG14, DRJ^{+}14, MT06]. Audio [FML23]. Audio-Visual [FML23]. Auditing [LMD16, YCC21b]. Augmentation [AHP21]. Augmented [LXW^{+}19, RRS^{+}14]. Augmenting [KAA^{+}18]. Authentication [CCF17, KLY^{+}23, LWL^{+}23, LZC^{+}24, LCZ^{+}23, LYC^{+}19, LYZ^{+}23b, WCQ^{+}20, WCH^{+}24, XFCW18, XZL20, XTHL21, BAL10, BGH^{+}95, FHH10, LLY06, OF11]. Auto [FDM^{+}17, TSN^{+}21, APB^{+}13]. auto-configuration [APB^{+}13]. Auto-Scaling [FDM^{+}17]. Auto-Tuning [TSN^{+}21]. autoconfiguration [CDM13]. autocorrelation [HH98]. Automata [LT16, vDJJ^{+}22, LRC15, PM96]. automata-based [PM96]. Automata-Theoretic [vDJJ^{+}22]. Automated [HK94, HLP^{+}16, LFF^{+}19, ZZL^{+}22, GXWW11, YWZZ16]. Automatic [BVL^{+}19, FLM^{+}22, ZKVM14, CGW^{+}12, QY12]. Autonomous [LGCG^{+}21, PLM19, SC17, YWW^{+}23, ZCF^{+}24, DEH^{+}07, Gao01, SKG12]. autonomy [FJB07]. AUV [QLF23]. Availability [ISS22, LSC^{+}21, MBL19, NBV17, QDD^{+}17, XDZ^{+}23, ZZZ^{+}07, ZYS^{+}23, ABA^{+}16, BSP07, Con11, DCGN03, DFMR15, GS10a, Gro99, JHR05, MDL^{+}13, LL20]. Availability-aware [ZZZ^{+}07]. Available [LTCS22, CDS02, JD03, LRL07, LRL08, SKKA01]. Average [DSYN24, CFS06]. averaging [Kri14, MSP^{+}07]. Avoid [CXL^{+}24]. Avoidance [HS19, LYS^{+}18, BB95, FJ93, MGK14, MNR03, PM96, TYP^{+}15, YSL^{+}14]. Avoided [CWAO21]. Avoiding [FB07, SDV06, VKO17]. AVQ [KS04]. Aware [ADT22, ABS^{+}16, AMS22a, BDR22, CNM20, CLQ^{+}19, CWZ^{+}17, CAD^{+}17, CYX^{+}17, DPT^{+}18, DF20, DLC^{+}18b, DKSC18, EFA19, FBFB17, FLBR^{+}19, GLL^{+}18, HHL18, HW22, HWC22, JWZ^{+}21, JSXN18, KAT^{+}22, KCM16, KTvdSK18, LS22, LXL^{+}22a, LCH20b, LCH22, LBGL20, LHZ^{+}19, LHL^{+}23a, LYKT21, LKMK20, MHR^{+}20, MKOY24, Nee16b, PHC20, RMDJ16, RZE^{+}21, SMEH20, SRS21, SG17a, SYG^{+}22, SKA^{+}18, SGVO18, TCTP20, TZX^{+}22, WXN^{+}17, WT17, WRT^{+}21, WZLM22, WHYC23, WZH^{+}24, WN16, WCW^{+}17, XXCC17, XPL^{+}17, XL23, YO17, YLH17, YPA19, YCZ^{+}23, YDCF^{+}22, YZGC23, YXZ19, ZGL^{+}19, ZHGF19, ZLZ21a, ZMD^{+}20, ZLW^{+}17, ZHWH21, AD14, AYM14, BJY11, Bor05, BLB10, CLC12, CKV11, DYH13, DV09, DLT^{+}15, FHSZ13, GLZC12, HLL13, HDM13, KT11, KS09a, LSZW13, LSS07, LG13b, LMW16, MLS12, MW05, PZS^{+}16, PLL13, RSM09, RD11a, SM14, SRB10, SJ10, SKCW10, SPB16, SZL^{+}14, TNRP11, TWL05]. aware [WH11, WA11, XTMM11, YSZL15, YCV15, YW07, YWZZ16, ZZZ^{+}07, dOSAU04, YFB02]. Awareness [CES22, NRB22, WLL^{+}16b, ZV16, ZLX^{+}23]. AWG [BHN11, GYLH17, YLH15, YDLM20]. AWG-Based [GYLH17, YDLM20, BHN11, YLH15]. AWGR [YWHW24]. AWGR-Based [YWHW24]. AWS [ISS22]. axiomatic [HSE97]. axiomatized [BYH^{+}15]. axioms [STC12].

babies [KHW12]. Back

[ABJ^{+}13, BT23, FWN^{+}22, MMT16, Van17, BSS11b, JJS13a, LEY14, MBF^{+}02, MS15, OWMM97, YSTL11, YSRL11]. Back-Off [BT23, Van17]. Back-Pressure [MMT16, BSS11b, JJS13a, LEY14, OWMM97, YSTL11, YSRL11]. Back-pressure-based [ABJ^{+}13]. Backbone [CLQ^{+}19, LWK^{+}18, SZMD17, ZZT^{+}17, BBG^{+}10, BDWS12, HM04, JID^{+}07, MIB^{+}08, RDO^{+}07, VWT^{+}14, WKA^{+}13]. Backbones [ZLL24, KLOS09, MTK03, NBTD07]. Background [CDK^{+}17, WH11]. Backhaul [BLM^{+}17, GHW22, LL17a, SSNS17]. Backhaul-Limited [LL17a]. Backlog [Nee16b, ZL16]. Backoff [BBF18, SD15b, HSM^{+}13, Kon06, KSM05]. Backpressure [AWKN16, HZCL16, KZH^{+}20, RpLP^{+}17, YN18, CYL16, HMNK13, LSLL14, SM16, SPB16]. backpressured [KGL03]. Backscatter [CGZL20, CZZ^{+}21, FHQ^{+}17, GLL^{+}18, GSH^{+}22, JWH^{+}24, JHM^{+}19, JHJL21, JHM^{+}21, LCW^{+}24, TZL^{+}24, YWZG23, ZZW^{+}24]. Backscatters [HWJZ21]. Backup [ACA16, BCO17, DPSA21, HSO19, HSS^{+}21, KRS^{+}17, WHYC23, BL04, GPM03, JLM15, LTP10, RC08, SZM08]. backup-bandwidth [SZM08]. Backup-Sharing [ACA16]. BACs [CXZ^{+}22]. Bad [La17, WXJ^{+}17, JAW11]. Balance [CKLS22]. Balanced [LJL^{+}16, CLY06, GGFS02, HD07, HY10, JMS08, YCL09]. Balancer [BWK^{+}22, JIN^{+}12]. Balancers [HKV23]. Balancing [CWGT14, CZ12, CCZL23, DPT^{+}18, FMH^{+}21a, GXL^{+}21, HV22, KAT^{+}22, KPK^{+}16, LYS^{+}18, LY22, LHL^{+}23a, PGMR18, PJDS18, SRK22, SG17a, SRCDL19, VJV14, VKO20, WXN^{+}17, YDCF^{+}22, YN20, ZDCW18, AWFT15, BD07, BHL07, HA16, KDYV12, LLW^{+}15, MOR13, MSS16, SMG05b, SK10b, Smi08, WL07, WSW12, YCV15]. ball [NST01]. ball-and-string [NST01]. Ballot [HBH93]. Band [GPBL24, JJJ^{+}23, TZPZ23, XLZ^{+}19, ZLWM18, CR98, MG97a, SKK07, Wan04]. Bandit [AM19, FM23, HVT18, SdVS22, TSN^{+}21, WN16, XWYL23, XLL^{+}20, YPA21]. Bandit-Learning [TSN^{+}21]. Bandits [KJG18, LAV16, GKJ12]. Bandwidth [BKLS08, CBHS20, CGAC20, DRCM^{+}17, FAWW22, HK96, KK00, KK03b, LNG^{+}21, LA95a, LNB01, LGHL17, LTCS22, MR02, SLH^{+}06, YLH17, YK23, ZWJ^{+}22, ZCM14, AA93, AS09, AS08, AC09, BBG11, BB94, BK00, BI00, CDFG06, CL04, CLS07, CLSS09, CAL09, Coh94, DZH03, DJM97, EM93, GS10a, GLLJ16, HBB09, HTC04, JD03, JVJ05, JJSS04, KKL03, KL03, KLS03, KZDM07, LM97, LRJ08, LOP97, LBL07, LZW^{+}15, LZL11, LFV10, LS06e, LW13, LRL07, LRL08, Low00, LFL14, LNC04, LYL07, MPF^{+}15, MJ13, PLD16, PGV16, LZKT99, RB09b, SLP07, SRR08, SCY98, SSM06, Smi02, SK06, SZM08, SL08, SK97, SSZ03, SC10, WL08, Wil96, WXW15, YMR00, ZB95, ZEV07a, ZS05]. Bandwidth- [SLH^{+}06]. Bandwidth-allocation [LNB01]. bandwidth-based [CLSS09]. Bandwidth-delay [KK03b, LM97]. bandwidth-efficient [GS10a, SLP07]. bandwidth-flooding [AC09]. bandwidth-guaranteed [KL03, LRJ08]. bandwidth-intensive [PGV16]. bandwidths [BW98, KWC93]. bang [ST04]. banyan [AMKY99, GP94, JSuRKH03, Kop96, PYL99, PG94b, RCGT06, WY95]. bar [Geo08]. Bargained [BO16]. Bargaining [CZP18, SFS^{+}22, BS09, MRHWS14, SAM10]. Barrier [NDS19, VMCB22, ZWYD18, GZCX16]. Barycentric [PWLC23]. Base [BSSU18, AKSS12, LMS06, PT96, SH12, SKS16]. base-station [LMS06]. Based [AAT^{+}23, AEG^{+}17, AAAR19, AWH^{+}22, AMSB^{+}24, BCO17, BPW23, BSG^{+}18, BCD19, CG21, CP17, CWAO21, CCK16, CM16, CLQ^{+}19, CHS^{+}20, CLX^{+}22, CLY^{+}17, CWZY21, CES22, CMY^{+}17, CMY^{+}18, CDW19, DTM^{+}17, DCN^{+}19, DPTP24, DZL^{+}18, DKSC18, DJB^{+}22, DHS^{+}23, EE18, FGRQ18, FLG^{+}20, FAWW22, FMCS20, FBFB17, FLZ^{+}23, FAS^{+}23, GSW^{+}23, GKB^{+}16, GYLH17, GYSZ19, GND17, GWYS19, HKS16, HTW^{+}22, HLZY23, HWZ^{+}23, HFF^{+}24, HNP23, HZHZ18, HHL^{+}19, HWHW18, HMM^{+}20, HYLS21, HLHL22, HGZ^{+}23, JCR21, JHL22, JSZ14, JLW^{+}24, JE18, KSSK18, KHYA20, KLR^{+}20, KLE16, LR22, LPJ^{+}17, LCK^{+}18, LLZ^{+}17, LMODF18, LCP^{+}20, LWT^{+}21, LZL^{+}21, LTZ^{+}22, LGDC23, LLT^{+}16, LWAL17, LXL^{+}19, LDL^{+}22, LYC^{+}19, LYZ^{+}23b, LFXY23, LLM^{+}24, Ma16b, MLS^{+}23, MBL19, MZZ^{+}23, MRM17, MKOY24, NLB19, NLT^{+}18, OL16, PLM19, PPS^{+}22, QYZX22, QLY23, RAPP22, SQ16, SCPB19, SYL^{+}17, TES19, TZX^{+}21, TH21, TXW^{+}21, TZL23, WSX^{+}21]. Based [WWC^{+}18, WZH^{+}18, WLS^{+}18, WUZ^{+}19, WLS23, WFZ^{+}23, WYL24, WCH^{+}24, WXH^{+}20, WCZZ17, WMT^{+}22, XCZ^{+}17, XNHM22, XLT^{+}22, XYQ^{+}17, XYT^{+}21, XWW^{+}23, XHZ^{+}19, XCV^{+}20, YSC18, YDY^{+}24, YYB^{+}22, YDLM20, YLF^{+}21, YGL^{+}19, YLA^{+}18, YWRK19, ZYL^{+}17, ZGL^{+}19, ZGYB20, ZY21, ZCPP22, ZWJ^{+}22, ZWR^{+}23, ZLL^{+}23a, ZLY23, ZHCC24, ZWC^{+}24, ZWS^{+}17, ZSZ^{+}17, ZKEN23, ZQL^{+}23, ZFW^{+}17b, ZZLM23, ZCM14, ZLM^{+}23, AIN^{+}15, AP93b, ACR12, AA96, AN05, AHL96, AK15, AWKN16, AAS14, AS02, AdE07, AGGT16, ABJ^{+}13, ALMR14, ARS16, BM09, BLC12, BV10, BS97, BLCT97, BTC01, BHN11, BRS10, BCGM07, BSS^{+}11a, BESW08, CLP12, CJW11, CSLH13, CW16, CqLL98, CU95b, CBSK07, CLS07, CJV16, CH15, CTG00, CEFS99, CS98, CSN06, CLSS09, CLA07, CL09b, CL13, COS95, CWW^{+}15, DM03, DC13, DM15, DHSS14, ES07, ES03, FCA^{+}06, FJ07, FGM^{+}13, FNQ00, FML09, FCT03, GDW^{+}16, GMZR13]. based [GGPS96, GGM11, GMD15, GT99, GT03, Gro99, GZCF06, GS09, GCS06a, HHL06, HTAZ16, HM06, HM04, HCL09, HY10, HK11, IKDD15, IBM95, JDSZ97, JJS13a, JHR05, JYT^{+}15, JGMB03, JVJ05, JJ08, JKJ13, JJL15, Kam10, KKS^{+}08, KLC15, KG10, KWE^{+}10, KG05, KWH11, KT06, KAZ01, KZDM07, KqL98, LA02, LBS05a, LBB08, LS93c, LL95, LZSS10, LML11, LMP08, LYRL07, LM01, LHB^{+}05, LLM11b, LCL^{+}13b, LM15, LHZ^{+}16, LH03, LHC05, LS06e, LÜ14, LLY^{+}12, Liu10, LCL^{+}12b, LCG^{+}14, LDHT02, LR03, LQCC16, MVRZ09, MR09, MQ05, MBG^{+}02, MNR03, Med95, MLT11, MWC16, ML12, MKT96, MRD08, MW00, MK98, MJ14, NKS08, NST01, NM06, Nee13, NPQ06, NABZ12, NTS12, NBT98, OAN15, OMA^{+}10, OJRCC02, PM09, PM96, PS05, PJ13, QK01, RRK07, RLKT98, RVS^{+}02, RX07, RSR11, RLZ10, RV01, SKT96, SRS03]. based [SL05, SM16, ST09, SNS12, SCY98, SM00, SSAK12, SR94, SSD93, SAM12, SML04, SL08, SIYL09, SK97, SCKB09, SAS^{+}16c, TR98, TS08, TMP07, TYLH09, TYP^{+}15, VL05, VA06, WZR08, WYL09, WLC^{+}10, WLL^{+}11, WBP^{+}11, WZLX12, WSW12, WKWV16, WWTK11, WM95, XHK^{+}05, XSC03, XLZC14, YKZ^{+}13, YWLL09, YLH15, YDS06a, YSTL11, YMKC08, YNDM09, YM05, ZCD97, ZWDS00, ZNN^{+}10, ZYL^{+}14, ZTS11, ZHLL06, ZY16, CW19, GSL^{+}24, YWHW24]. basestation [STKL01]. Basic [Kar03, SK13, LL99]. batches [vDP93]. Batching [SC22b]. Battle [KKS19]. Bayesian [ABMT23, WJK^{+}12]. BCCC [LGY16]. BCMIX [ZLM^{+}23]. Be [OLZ17, YM16, SHJ10]. Beacon [DLY^{+}22]. Beaconing [GYSPR14]. beaconless [RKNS10]. Beam [GHK18, SKE19, ZP18, ZW22]. beamformer [ASKR16]. Beamforming [CZW^{+}21, CDKZ21, DJK22, KKH^{+}22, MAPZ18, MBN^{+}21, AKS^{+}13, ZJS^{+}12]. Beating [ZGY^{+}16]. Before [DJCA21, CTG00]. Behavior [HDQ^{+}16, HCL^{+}17, LXW^{+}17, WQY^{+}17, XWH^{+}16, XWG14, ZWH^{+}17, ZSZ^{+}17, ANSX13, BOGS^{+}16, BBLV06b, BPS99, CS98, DM95, EJ14, GSD09, HBS96, IW08, JWSH15, KEW06, KS13, LYWL08, LT95, LBP^{+}16, Pax97, SSD93, SENB09, TL06, VL05, XZB08, YR01, BBLV06a]. Behavioral [PJMM22]. Behaviors [JLW^{+}24, JSuRKH03, TWL06, XY09a]. Behind [VKO17, SCKB09]. Being [CKLS22]. belief [KL12, SS04b]. benchmark [TT07]. benchmarked [AIN^{+}15]. Benchmarking [MLJ^{+}22]. Benefit [NTD17, AADS05]. Benefits [DKS19, ZNK^{+}13, AMS^{+}08, AMSS08, BSF16, JWSH15, KBV^{+}13, LLCL11]. Benes [PS93]. Benign [LMSS24]. BER [ALJ99, Wan04]. BER-scheduling [Wan04]. Best [LPM23, CF98, KL07, PWK^{+}13, SL08, YD04]. best-effort [CF98, KL07, PWK^{+}13, SL08]. Better [GHBSWV17, PLR^{+}19, ZMLL21, MA12]. Between [CL19, CKLS22, KLLT18, LLZ^{+}23b, LCY^{+}19, ZCW^{+}22, BMB^{+}11, DGK05, DEH^{+}07, GT10, HFC^{+}13, LMS04a, MCL^{+}11, PMH95, QTWW16, SHHA09, SYP01, XL05]. Beware [HHD22, WH11]. Beyond [FMSM24, JHJL21, OS21, PWK^{+}13, QWL21, SMC02, TMGB19, YLL10, ZLL^{+}23a, BLC12, RTK^{+}16]. BFAST [DLW^{+}17]. BGP [BFF07, BGMB^{+}20, EKD12, FR07, GCH^{+}15, LBP^{+}16, SKG^{+}18, SCHG22, SVG16, VVP^{+}13]. Bi [CLS^{+}21, DME23]. Bi-Directional [CLS^{+}21]. Bi-Virus [DME23]. Biased [KE21]. Bidding [ZDCW18]. Bidirectional [WG16, GLNP01, LMS00, LH95, Lie97, RM08]. Bifurcation [CMR17]. Big [ST04, YWRK19]. Big-bang [ST04]. Bilateral [AJF11, LDL^{+}22]. Bilevel [CGAC20]. billiards [BH05]. Billion [LSDT19, NTD17]. Billion-Scale [LSDT19, NTD17]. Bin [PG21, RTLC17]. Binary [LK95, XLD^{+}24, LL09, LHL15, LMT16, ST13]. BIND [KZ97]. binding [MR96]. bio [FLMM10]. bio-inspired [FLMM10]. biological [HS06a]. Biology [DGLM16]. BioNetwork [CAP15]. bipartite [XWG14]. Birkhoff [CLY06, VIT21, YZLH17]. Birth [LAV16]. Birth-and-Death [LAV16]. Bisection [KHYA20]. bistatic [GZCX16]. Bit [HS18, KIL24, MLT12, XLT^{+}22, ZZ17, BSH^{+}11, BB16, CR99, HJL^{+}12, HH10b, LCL^{+}13b, RT99, SJ95, SA01b, XSSK08]. bit-rate [SA01b]. Bit-Selecting [XLT^{+}22]. BitCoding [HS18]. Bitcoin [JW23, SCN^{+}22, SARM24]. Bitmap [EVF06, LQCC16]. Bitmaps [WMCW22]. bitonic [LC94a]. Bitrate [JYL^{+}19, LOFH21, MGS^{+}21, SUS20]. Bitrates [WZH^{+}24]. bits [HJL^{+}12]. BitTorrent [CKC^{+}13, FLC09, LDH^{+}12, PWMC12]. BitTorrent-like [FLC09, LDH^{+}12, PWMC12]. Bivariate [REM17]. Blackout [NEH^{+}22]. Blessing [ZDCW18]. Blind [LCL17, WCH^{+}24, XOYL20, YN20, HKB14]. Block [CZM^{+}24, HZB^{+}22, LZX^{+}24, SKE19, WXX^{+}24, LYC11, XL98]. Block-Oriented [CZM^{+}24]. Block-Wise [LZX^{+}24, WXX^{+}24]. Blockages [DF20]. Blockchain [AWH^{+}22, AMSB^{+}24, HFF^{+}24, LFC^{+}22, PWL^{+}24, RSLL23, VRR24, WFZ^{+}23, WWL^{+}24a, ZY21, ZLM^{+}23]. Blockchain-Based [AWH^{+}22, HFF^{+}24, ZY21, ZLM^{+}23]. Blockchain-Enabled [AMSB^{+}24]. Blockchains [RV21]. Blocker [LXL^{+}17a].

Blocking [GCW21, JSuRKH03, LNC98, OL16, PLS^{+}21, SS04a, Xin07, vDP93, BIS00, CTG00, CLW95, CCKK16, CKR93, FT07, KL09, LL09, LWL04, LXC05, LC94b, TCPV13, WF93a, YLH15, Zeg95, ZRP00]. bloom [FDG^{+}10, Mit02, ALY^{+}20, AAS14, DKT06, EF17, HKLS12, KLC^{+}18, LYW^{+}18, QHZC18, QCMY16, RR19b, RKK14, RK15, ZZ17]. Bloom-filter-based [AAS14]. BLUE [cFSKS02]. Bluetooth [AHX19, CZZ^{+}21]. Body [GSM16, MFT^{+}20, TSS14]. BOLA [SUS20]. Bolt [WZL^{+}23b]. Bonding [BMBK21, SC22b]. bonds [ACOR99]. book [GSW99]. book-ahead [GSW99]. Boolean [BHA^{+}20, DMDM17, LWQ^{+}18, WXC16]. Boosting [JYL^{+}19, LYKT21]. Bootstrapping [HRLY21]. border [AVS04]. Born [QZC^{+}22]. borrowing [JR96]. BOSE [WXX^{+}24]. Both [CGC^{+}24, LGDC18, WXM21, WGL22, LLE16, YD04]. botnet [DKC^{+}15]. Botnets [WCCM18]. bots [GXWW11]. Bottleneck [HWF^{+}20, WXH^{+}20, YLL21, JK05]. Bottleneck-Based [WXH^{+}20]. Bound [FAWW22, LGDC23, SG18, ZMW^{+}22, ABA^{+}16, FP95, KWS^{+}11, KCB03, wTjCjC97, TG97, YYZ06]. Bound-Based [LGDC23]. boundaries [CGMS13, LE12a]. Boundary [LLJ^{+}14, BNS11, HGE04, LM01, LZL^{+}14]. boundary-point [HGE04]. Bounded [CGC^{+}17, CZK^{+}21, HKV23, LZC09, LL20, MGZ^{+}23, TSS14, CE09, CZC^{+}13, CFS11, HL05, JR14, Jia98, JKJ13, LWF96, Pil01, SS09, ZSK12]. bounded-hop [SS09]. Bounded-mean-delay [LZC09]. Bounding [FT07, KDHK15]. Bounds [AK96, BHA^{+}20, CLW16, HH17, KGH^{+}20, LLW^{+}14, SSM03, TML22, ZGZC20, AJV06, AGLM10, BBC^{+}02, CBL06a, LNS11, LPF12, Liu10, RTK^{+}16, SKK07, SS05, TCS13, WKZL96, XL05, YS93, ZSCJ14]. Box [LGS^{+}23]. BPSO [HWZ^{+}23].

BPSO-vote-HMM [HWZ^{+}23]. BRA [RM08]. braiding [SKHL12]. Branch [WWMZ20]. Breach [TXL^{+}18, ZWYD18]. Breach-Free [ZWYD18]. Break [DJCA21]. breaking [TGRR07]. breathing [WKWV16]. BRICK [HXLZ11]. bridge [ZTS94]. Bridging [ZCW^{+}22, HFC^{+}13]. bring [YASS15]. Bringing [WCWZ17]. broad [CR98, MG97a]. broad-band [CR98, MG97a]. Broadband [AD18, AK00, CGK10, CG15a, DM96, LA95a, LZZR12, MSH95, OKM94, Ord99, SYDM09, YMR00]. Broadcast [BP19, CLWZ17, CGR^{+}18, DKSC18, KSUB^{+}18, KS19, LTDM17, QLF23, SPLM17, SPM^{+}17, TZL23, YZY^{+}18, ASW00, AF99, AGGT16, BCB99, BK06, BC01b, CH11, CCA96, FSM14, GMP08, GGK99, HH10a, HL96a, LPP11, Mod99, PM96, PEA09, PS94, RS97b, SKR^{+}09, SS09, SMSM06, SV11, SKUB12, SZT01, VCM04, YS15]. broadcast-and-select [BC01b, Mod99, PM96, PS94]. broadcast-video [HL96a]. Broadcasting [Hou15, HH10b, PP17, SSA08, YZY^{+}21, CFM13, FWL08, TCS04, XAST12, ZG14]. Broadcasts [DKSC18, NST^{+}16, ME96]. Broker [SCPB19]. brokerage [SZG09]. Brownian [LSMS06]. browsing [XY09a]. Bruijn [SR94]. BS [HHA17]. BSP [HTJ^{+}21]. bucket [GQ16]. Buckets [EFFK18]. Budget [AIL23, LZY^{+}22, SJ21, ZLM16, HZL16, LSM^{+}14]. Budget-feasible [ZLM16]. Budget-Limited [LZY^{+}22]. budgeting [BM00]. Budgets [CMW^{+}20]. Buffer [BDR22, CLC12, GSW^{+}23, HWC22, LLM11b, OL16, BT93, CH97, CGK94, EW08, FJ07, Geo08, HM06, JVJ05, JJSS04, KS01a, LBS11, cLqL97, Low00, MV09, PV04, PDT09, LZKT99, RRK96, SV99, SEMO09, SC95, THP94, WM95, ZY07a]. Buffer-Aware [BDR22, CLC12]. Buffer-Based [GSW^{+}23]. buffer/bandwidth [LZKT99]. Bufferbloat [YLL21]. Buffered [BHC^{+}21, Geo08, SRCDL19, YXAZ^{+}18, CC95, HSG^{+}08, LC94b, OWMM97, SPC10, TT09, Tur09]. buffered-type [OWMM97]. Buffering [SLD14, VB94]. buffers [AGL16, BBG^{+}10, IKM08, Kim94, LMS12, LMSKZ99, VS97, VSR11, XME15, ZKO93]. Bugs [BZS23]. Build [QSW24]. Building [HZB^{+}22, KT08, PSST21, YW07, GS09]. Built [QW23]. Built-In [QW23]. Bulk [YCW^{+}19, ZDB^{+}17, ZLZL16, BKTN03, LSS^{+}13]. bundling [MDL^{+}13]. Burst [LT95, SR18, SRC^{+}20, YCZ^{+}23, BV10, HH10a, LQXX07, RLZ10, Zal09]. Burst-Aware [YCZ^{+}23]. burst-based [RLZ10]. burst-switched [BV10]. Burstable [JSW^{+}20]. Burstiness [JHM^{+}21, YXCH21, KA95, Val01]. bursty [JK96, JPS04, Nee09, WM96, YZ10]. BUS [SZT01, BB96, LH95, NSS96, SS93, SS94a]. Buses [KP21]. buy [KKP15]. Bypassing [PLT14, TJHL21]. byte [BKH^{+}93, CB99]. byte-interleaving [BKH^{+}93]. Byzantine [YKGK13, ZYY^{+}21]. Byzantine-Resilient [ZYY^{+}21]. Byzantine-resistant [YKGK13].

C [SG94]. CA [BK17, JZC11, JP13, Kon06, LK16a, NTS12, SKK07, Van17, VBHT17]. CA-based [HK11]. Cable [WWMZ20, WWMZ22]. CAC [CGMS13, ZTS11]. Cache [ADR18, BRK^{+}22, DJS^{+}17, DCN^{+}19, DMT^{+}19, LI21, LL17a, MMIY20, NCM18, PLD16, WTK^{+}17, XNHM22, YXC^{+}18, BD96, FCAB00, GMWD13, GMD15, KRS00, NSCR06, PP02, RW04, RV00, PLS^{+}21, WSX^{+}21]. cache-friendly [RW04]. cachecast [SPGM13]. Caches [CNM20, CG21, TJHL21, XWYL23, CDPLCA16]. Caching [ATEY22, ATE22, ATE23, ACLX17, AAAR19, ADR18, AN20, BSG^{+}18, BRK^{+}22, CYH^{+}18, CEFS21, CES22, DJS^{+}17, DD24, ER23, GR20b, IYYI18, KYM22, KLKP16, LMSR19, LDH^{+}12, LAJ20, LPWP22, MYH21, MLS^{+}23, MJ17, MHL19, NLRS21, NGL22, PD16a, PLS^{+}21, PDI20, QTE20, RTNS21, RT17, SSN^{+}23, SPLP20, SNC23, SLS^{+}23, SZWW22, TEE16, TE16, TJHL21, VWNT17, WBWV16, WWC^{+}18, WQL^{+}21, ZLL^{+}23b, AS14, AD14, BK06, HS08, JSBM02, MAN15, PMAN16, PD16b, RSB01]. calculations [KS01a, SS98]. Calculus [LYL^{+}22b, TL22, CBL06a, LBL07, MSB97, SKZ03, ZM09]. Call [LHL^{+}23b, ASCG08, AL98, BLCT97, BIS00, DM96, FCL97, HKT95, IPG97, KL09, LLD96, LAN97, MR96, PDSK04, Pil01, RV01, Rum93, RS95b, Smi95, VG04, WWL02]. call-in [RS95b]. calls [CCY^{+}14, CTG00, GSW99]. Camera [RWL^{+}22]. Campaigning [KK16a, KSK17]. campaigns [DZNT14]. Can [AQK^{+}19, HLH^{+}18, RS05, YM16, CPS13, LLY^{+}13, SHJ10, SSFM08, XCR11, XCR15]. Cancellation [CHS^{+}20, LPR17, LHW19, BSS14, GNP^{+}13, YASS15]. Candidate [YY20, WYH10]. Cap [WMX17]. capabilities [SAS16a, SYP01, SSM06, WNV13]. Capability [LLZ^{+}17, MHS^{+}17, RRK96]. Capability-Based [LLZ^{+}17]. capable [TEML09]. Capacitated [VLDM17, KNP05]. Capacity [AGLM10, ACKZ14, AHP21, BBLV06a, BMY^{+}17, CVV17, CCL11, DWCZ17, DHHD18, DZH19, GGL09b, GGL09a, HCL^{+}17, HW12, HR14, KAK19, KV09, LM95, LPF12, LL17a, MS08, SJ21, SV06, TZL^{+}24, XME15, ZFW14, ZRP^{+}22, ZZLW16, AJV06, ALMR14, AJ06, BBLV06b, BB96, CZF^{+}16, CJ97, CDS02, DSTM12, DTM15, DFZ06, DRM04, GHW14, GT02, HBB09, HKL06, HBU95, HM04, IMG98, JVY06, JLS09, KD10, Kuc14, LK16a, LPKF10, LCH95, Li09, LLLT10, LPW14, LSMS06, LTS05, LE06, MM94, MK98, PD16b, PDT09, QY04, RP13, RDO^{+}07, RK06, SKKA01, SLS10, SMS07, SR01, Smi02, UN11, WHW^{+}11, XK06a, XM99, ZH08b, ZLW16a, dFV02]. capacity-delay [CZF^{+}16]. capacity-estimation [DRM04]. capacity-varying [SR01]. CAPS [HHL^{+}19]. capture [CT04b]. Capturing [HPV09, CZM14, GSK08]. Car [CXZ^{+}22]. Cardinality [CN19, GLLL17, HOZL16, RAA^{+}24, WMCW22, XCZ^{+}17, XZC^{+}19, XCZL20, ZL14]. cards [LMP96, PZS^{+}16]. Carrier [LPD^{+}18, SAC^{+}18, YZZ^{+}21, BSH^{+}11, KNSV13, MVRZ09, SCN12, ZS13]. Carry [PK01, SMT98]. Carry-over [PK01, SMT98]. carrying [FRC98, LSC99a]. cascades [La16]. Cascading [LQZ^{+}24, XS21]. Case [DJK23, HKV23, JHL22, TML22, ZHCL17, AS07a, BGVC00, BM93, BS15, CPGZ15, DYH13, ESG11, GSKR99, JK05, Kim98, Lee96, LH10, PG93, PG94a, RIM98, RVB12, SM08, SMM11, SPR08b, SPR08a, Val01, WLS97]. Cast [ZSL^{+}21, JPH08]. Catch [AQK^{+}19]. Categorized [LLG^{+}17]. Category [LLL^{+}17, LCX^{+}19, ZCZ^{+}20]. causality [KS13]. Cause [FLM^{+}22, WWYY18, YBG^{+}12]. Caused [TRVG20, DSA^{+}14]. Causes [MRMR17, AST11, CB97, MG95]. Cayley [PC19]. CBFQ [BTC01]. CBID [HDQ^{+}16]. CBR [ITSO01, Lee96, LyT98, PS98]. CCDN [ZLW^{+}16b]. cDeepArch [YXL^{+}19]. CDF [JJL15]. CDF-based [JJL15]. CDMA [ALJ99, CT04b, CS99b, FT07, GKB^{+}16, Hu93, KMT05, KCB03, KG05, LMS06, fTL06, Wan04, YD07]. CDMA-Based [GKB^{+}16]. CDN [AAAR19, LYS^{+}18, SCKB09, TWWG19]. CDN-Based [AAAR19]. CDNs [CDL^{+}19, LLZ^{+}23a]. CE [RAA^{+}24]. CE-ACF [RAA^{+}24]. CEDAR [QSS^{+}15]. Cedos [MKG^{+}17]. Cell [AP17, BWG^{+}20, BAB20, CZX18, GKS05, GZJ^{+}18, HRM22, KLP16, LA95b, LCK^{+}18, MAPZ18, NLRS21, PK01, RTNS21, Ros96, SFM^{+}18, YZL^{+}18, BLCT97, BHN11, CHCH00, CG15b, FCL97, KDYV12, Kuc14, KAMG07, LMSKZ99, LLY^{+}12, MBG^{+}02, RrBG94, RKA08, SMT98, TG97, WF93a, WKWV16, YWK07, ZF96, DMMS14]. cell-based [MBG^{+}02]. cell-breathing [WKWV16]. cell-counting-based [LLY^{+}12]. Cell-Free [BWG^{+}20]. cell-scheduling [CHCH00]. cell-switching [RrBG94]. cells [ASKR16, GH93, MS95, SAS^{+}16c]. Cellular [AAT^{+}23, AEG^{+}17, AMG^{+}17, GHRH18, KSAK18, KPK^{+}16, KIL24, LKS^{+}16, LCK^{+}18, MRJ20, NV21, SFM^{+}18, TSN^{+}21, WXC^{+}24, WLL^{+}16b, XLW^{+}17b, YLWH20, ZLL^{+}23a, ZJWY17, AZR97, AS96, CSC94, DM15, DRJ^{+}14, GH04, HRCW08, JR96, KAEAS14, KMZR12, LPKF10, LS06b, LSC99a, LSC99b, LC04a, LCZC13, LG13b, MBL10, MGCK15, MSA^{+}16, McM95, MAS09, PMH95, RP13, SEK15, SJL^{+}13, SJL^{+}16, SKS16, TPC09, TEML09, XSC01, XSC03]. censorship [DSA^{+}14]. Center [AB21, AGCFV18, BWES22, CZP18, CZX^{+}17, CWM^{+}17, CLM^{+}18, CHFH20, CXW^{+}18, GXL^{+}21, HHL^{+}19, HZC^{+}19, HLL^{+}21, LLCJ22, LWC^{+}23, LHZ^{+}19, LHL^{+}23a, LFXY23, MBI^{+}17, QFH^{+}18, SS17, SRC^{+}20, SLD^{+}23, TJL^{+}19, WXN^{+}17, WLX^{+}17, WN17, XLAC16, ZWGC17, ZCB^{+}17, ZDZ^{+}24, ZLW^{+}16b, ZFW^{+}17b, ZHWH21, CKL16, CGW^{+}12, CSS^{+}14, CYG^{+}14, JRL15, LGW^{+}11, LLW^{+}12, LZW^{+}15, WFGZ13]. Centers [APC21, BCC^{+}17, DXX^{+}23, HTW^{+}19, HCW^{+}16, LGY16, MGZ^{+}23, WJ17, YLH17, BMB^{+}11, LZXF14, LWAT13, PMH95]. Central [SRCDL19, CS98]. central-limit-theorem-based [CS98]. Centrality [ML18]. Centralized [AS08, CGC^{+}17, DC13, HRM22, LWC^{+}23, ZZ17, ZXZ^{+}19, BLL07, HKV^{+}13, LNB00, SD15a]. Centric [ANTR17, DSM^{+}17, GTU19, LSCT17, MYMY17, PD16a, PGMR18, SS16, SGH^{+}19, WBWV16, XHZ^{+}19, XHY^{+}22, ZHZ^{+}24, ZLW^{+}16b, AK09, AGL16, CT04b, LM13, RJJ^{+}11, YLLY05]. Certificate [LLL^{+}22a]. Certificateless [CJS^{+}20]. Certificates [WQL^{+}21]. Certified [LYZ^{+}23b]. Chain [DJK22, DJK23, EMAL17, GR20a, HJG18, KLE16, LW20, QZL^{+}16, REM17, SARM24, GMWD13, ZS04, SJWH^{+}17]. Chaining [BSM21, LYL21]. Chains [FBM^{+}21, JWL^{+}18, KZH^{+}20, KLLT18, WHYC23, ZLZ^{+}21b]. ChainSGD [ZGZ22]. ChainSGD-Reduce [ZGZ22]. ChainSketch [HZL^{+}23]. Challenge [CQW^{+}18]. challenges [SRR08]. challenging [ML12]. change [CG04, SR01]. changers [KS01b]. changes [CCY^{+}14, CF94, CTVD14, SNC^{+}07, TSGR08]. changing [AC06, SP94]. Channel [BCP00, BMBK21, BSP21, CE19, CLW16, CSL21, CBZ16, CHS^{+}20, CJ18, DZ18, EE18, FLS^{+}22, GLL^{+}18, GWYS19, GSM16, GHZ20a, GHZ^{+}20b, HXZ23, JHJL21, KIW^{+}17, KW17, LSC99a, LHL^{+}21, LCLC18, MLS12, TKM20a, TMH97, WZLM22, WLL^{+}16b, XCL^{+}22, ZYL^{+}14, ZK19, ZSS^{+}20, ZY21, ZWC^{+}24, ZGLC20, AK15, AGGT16, AAV09, BGK97, Bor05, CL09a, CLM^{+}16, CK07, CFS09, FTZ^{+}13, GV93, HSM^{+}13, HL98b, IZC00, JR96, KKV16, KT07, Kuc14, LSC99b, LLLT10, LyT98, LR09, MRM99, MHSC95, NAA^{+}16, PT96, RW93, TS08, TCS04, WXW15]. channel-assignment [LR09]. Channel-Aware [GLL^{+}18, MLS12, Bor05]. Channel-hopping-based [ZYL^{+}14]. Channels [GV17, GLY17, HH18, KLP16, NST^{+}16, QTE20, SAMB18, TG23, XZL^{+}21, YSY16, YLY^{+}16, ZCZ^{+}20, AZLB16, AZ06a, BLEM^{+}12, CAK12, ÇM15, Coh94, CG15a, ESP05, GK16, Hou14, JLRS16, KVR98, KL07, KHTK00, KN05, LCQL14, NMR03, OES16, SL12, SKUB12, SV06, TMH97, YS15]. Chaos [ZGY^{+}16]. Characteristics [CNDK18, EE18, KE21, QTE20, SBLS19, CKR^{+}09, LH95, TWL04]. Characterization [DD24, LL98, MIB^{+}08, WCCM18, AW97, cFCcFW05, LLY01, LBX11, RRK07, SJL^{+}13, SH14, VAM^{+}06, WTXT11]. Characterizing [BMS14b, CFS^{+}10, FK07, ISS22, KN05, SJL^{+}16, SRS08, WW16]. Charger [DWL^{+}18, DLY^{+}21, LXX^{+}17]. Chargers [JLS^{+}17]. Charging [CLHY22, CGC^{+}24, DLC^{+}17, DMLC18, DLC^{+}18a, DWL^{+}18, JLS^{+}17, LXX^{+}17, LYD^{+}21, MLX18, SLD^{+}22, XSH^{+}15]. CharmSeeker [ZZL^{+}22]. Chase [CZM^{+}24, CLWZ17, CWL^{+}21]. chat [GXWW11]. cheap [SK12b]. Cheat [BLL07]. Cheat-proof [BLL07]. cheating [LWL^{+}11]. Check [RAA^{+}24]. Checker [TSS21]. checksums [SGPH98]. Cheetah [BWK^{+}22]. Chemistry [MSTL17]. Chemistry-Inspired [MSTL17]. Chen [FM06]. Chen-Stein [FM06]. Chinese [Su15]. chip [AIN^{+}15]. Chipless [JWH^{+}24]. Chips [DGLM16]. Chirp [LTCS22, XZL^{+}21]. Chirp-Train [LTCS22]. Choices [NGRF19, KM08]. CHOKe [EJ14, TWL04]. chord [FLMM10, SMLN^{+}03]. Chunk [RBPS21, Liu10, ZCL11]. chunk-based [Liu10]. chunk-scheduling [ZCL11]. Chunk-Sharing [RBPS21]. Chunked [TY18]. Chunking [HSKY23, LK16b]. Churn [BSSU18, XXCC17, BQ08, EKD12]. Circuit [MFT^{+}20, TZX^{+}21, VIT21, CJ14, CHA95, Coh94, LT02, RZZ06, VS97, VL99, WCY00, Zal09]. circuit-switched [LT02, RZZ06, WCY00]. Circular [VTBK21]. cislunar [WBP^{+}11]. CISTs [PC19]. Cities [DLC^{+}18b, WHC^{+}22]. Class [DKSC18, LTCS22, MGK20, YK23, ALMR14, CLA07, JM00, KG16, LMS05a, LMS05b, Med95, SG94, VR13]. class-based [CLA07]. ClassBench [MLJ^{+}22, TT07]. ClassBench-ng [MLJ^{+}22]. Classes [CLCL23, KK16b]. Classification [BSM21, BSP21, DBL^{+}19, FLH^{+}17, HS18, KAHKB17, LQZ^{+}24, LYH^{+}23, MLJ^{+}22, NLT^{+}18, RRS22, VBC^{+}17, VLZL16, WHLL23, XOYL20, XLD^{+}24, XLT^{+}22, YDLT18, YWZ^{+}23, ZXW^{+}21, ZLY23, ZWZC23, BV05a, CSLH13, CW16, CKKK09, GXWW11, JID^{+}07, KNR^{+}16, LCL^{+}13b, LLJ^{+}14, LMT16, LS07, LQCC16, MLT11, NABZ12, SMP^{+}14, ST13, SSZ05, TT07, Tre11, WLCC07, XLZC14, ZCX^{+}15]. Classifier [WQY^{+}17, FMMR10]. Classifiers [DNCK20, DKN21, LMT10, LS09, MLT12, WXC16]. classless [GCS06a]. CleanG [MRJ20]. Client [LHW19, PJDS18, WZD24, AYM14, AWFT15, BCL^{+}09, GCZ98, MR96, ZT12]. client-level [MR96]. client-membership [MR96]. client/server [GCZ98]. Clients [KHAWC17, LHW19, BK06, FHSZ13, JS06]. Clique [LLAS19]. CloakLoRa [HXZ23]. Clock [HPP^{+}23, HKS16, WHZJ20, GTS^{+}09, GCS06a, KL95, LSW15, MMH^{+}15, Mil98, RVB12, SA01b, XL95]. Clocks [ML22a, KMH12, Mil95, VRK09]. clone [LG13a]. Clos [GYLH17, HL00, LNC98, LC96, OJRCC02, Smi08, SRCDL19, WXN^{+}17, YLH15]. Clos-Network [SRCDL19, OJRCC02]. Close [LOFH21]. Closed [GLMM04, NGK19, fTL06, NKL^{+}23]. Closed-Loop [NGK19, fTL06]. Closer [AQK^{+}19, LOFH21]. closets [Bej09]. closures [Ber00]. Clothing [HHD22]. Cloud [AQK^{+}19, CLTM22, CPKL17, CNM20, CPS17, CJLF16, CDL^{+}19, DKS19, DJB^{+}22, ECL^{+}20, FLM^{+}22, FLTM18, FLBR^{+}19, FSSC18, GGZC19, GSW^{+}23, HKLM17, JTL^{+}17, JTL^{+}18, LLWB16, LSSC22, LLT^{+}16, LS17, LSC^{+}21, LW22, MCdG23, PCW^{+}16, PG18, QLY23, RTLC17, SRS21, SC18a, SC18b, SZW^{+}16, WFC18, WLS^{+}18, WLTJ19, XWYL23, XRL^{+}22, YCC21b, YWH21, YWZ^{+}23, ZLRC20, ZLZ^{+}21b, ZWR^{+}23, ZLWH17, ZLW18, ZLW^{+}17, ZFLC18, ZCM14, DGG^{+}02, HTAZ16, LK14, SAS^{+}16c, Szy16, WLCW16, WRS^{+}15, WXW15]. Cloud-Aided [LW22]. cloud-assisted [WLCW16]. Cloud-Based [WLS^{+}18, HTAZ16, SAS^{+}16c]. Cloud-Ready [ZLW^{+}17]. Cloudlets [CSR^{+}17]. CloudNet [WRS^{+}15]. Clouds [ALR^{+}24, FMH^{+}21b, HTL^{+}19, JSW^{+}20, LLZ^{+}23a, LK16b, SSR^{+}20, TZX^{+}22, WZX^{+}22, WFZ^{+}23, WWYY18, ZHW^{+}17, ZGS^{+}24, ZLX^{+}23, ZWX^{+}24, DBDJ14, JLX^{+}16, LWLL16, MS14, WWL^{+}15]. CLP [RRK96]. clue [ABBHP01]. Cluster [CL16a, SCS^{+}22, YSTL11, ZFW^{+}17a, LAN97, LNA07, LÜ14]. Cluster-based [YSTL11, LÜ14]. Clustered [AD18, EKSV16, SK10b]. Clustering [FAS^{+}23, GZL^{+}17, LSCT17, SL17, TSN^{+}21, BMM^{+}09, BLB10, CAO11, GMZR13, YD07, YF05]. Clusters [BPW23, DRW^{+}22, FC17, PG21, WLS23, WWL24b, JIN^{+}12]. CN [SCN12]. Co [AQK^{+}19, KSSD24, LS22, QLSW19, Kuc14]. co-channel [Kuc14]. Co-Evolution [QLSW19]. Co-Residency [AQK^{+}19]. Co-Scheduler [LS22]. Co-Scheduling [KSSD24]. Coalescing [CM16, FC17]. coalition [SSAK12]. coalition-based [SSAK12]. coalitional [SSA11]. coax [CLG^{+}00a, LS97b]. coaxial [CR98]. Cocktail [ZLL^{+}23b]. CocoSketch [MZZ^{+}23]. COD [CT96]. Code [BB95, CMY^{+}17, MWC16, CDO97, CSLH13, CWW^{+}15, Hos98, Hu93, KCA97, OF11]. Code-based [MWC16, CWW^{+}15]. Coded [ACLX17, AAA18, BPA21, BP19, CLCL23, GR20b, KK21, LYL^{+}22b, LPWP22, EGKM16, SLS^{+}23, SGVO18, XLAC16, YPA21, YHCL21, ACKZ14, FDG^{+}10, FSM14, GH93, KWH11, LRM^{+}06, MAN15, NLB15, PMAN16, RGG11, SM14, SV11, THMK12]. CodedReduce [RPPA22]. Codes [SKE19, SWH19, TY18, WSL^{+}24, AD11, DPR06, ESG11, Far95, Fel95, McA94, Sho06, SV15, WCAB15, YS15, YSJL14]. Coding [APC21, ABS^{+}16, BTP^{+}17, BK06, CYTH22, CE19, CJS^{+}20, CCCC17, CMY^{+}18, EFA19, EBJM18, GLA19, HHL^{+}19, HZLZ22, KSAK18, KW17, LWL17, LK16b, PP17, PBT^{+}20, QDD^{+}17, RV21, RRS^{+}14, RKPP16, SQ16, VPC17, WGvdS17, WMT^{+}22, YKB^{+}23, ZSH^{+}16, CFS06, CLC12, CZZY12, CGK10, CBL06b, DMC06, DYH13, DFZ06, FWL08, GV93, Hou15, HK11, Kam10, KRLL11, KRH^{+}08, KWS10, KBV^{+}13, KM03, KWH11, LE13, LSB06, LZZR12, LK14, LP07, MRHWS14, OF11, OWKS16, PRR06, PCL15, QY12, RGKR10, RJCE06, SM14, SRB10, WM16, WJK06, XY10a, XL11b, YYZ06, YSZL15, YASS15, YZBR14, YMKC08, ZNK^{+}13]. coding-aware [SM14, SRB10]. Coding-Based [HHL^{+}19, Kam10]. CoEdge [ZCZ^{+}21]. Coexistence [CLGSS17, GSPV^{+}18, GKCR21, LHL^{+}21, MSRG18, BSS^{+}11a, LMSKZ99]. coexisting [KCTI08, ZS13]. Coflow [JHL22, LS22, LFXY23, TJL^{+}19, TZX^{+}21, WZH^{+}18]. Coflow-Aware [LS22]. Coflows [SG18]. Cognitive [BAB20, BDR22, BMY^{+}17, CLW16, CCLL17, DAFZ^{+}18, DZL^{+}18, GJCB18, LSL^{+}18, RZS14, SPQZ20, SCH23, AK14, AK15, CAO11, CZM14, FEC13, GSA15, GMYP16, HW12, KKEE13, KS10, KNK^{+}14, LZES14, LWT^{+}15, SKY10, STC12, TW10, WSW12, YKZ^{+}13, YGC10, ZYL^{+}14]. COIN [XQG^{+}22]. Cold [DQW^{+}23, LX24]. Cold-Start [DQW^{+}23]. Collaboration [WYL24]. Collaborative [AD18, GND17, HSL20, IGHT17, KJG18, WQL^{+}21, WJH^{+}21, WFZ^{+}23, XWH^{+}16, YWH21, ZGHH19, ZLWM18, FAB12, GGM11, LLY06, VA06]. collapse [AVS04]. Collected [Kar06]. Collection [ALYX22, CCG20, LXL^{+}19, LCY^{+}19, XYQ^{+}17, XXZ^{+}22b, GIKK11, JC13, LFZS11, XLR13, YCV15, YZP^{+}14].

Collective [RDR17, ZJ12]. Collision [CW23, ODJ23, XXCC17, CT04b, HDM13, JL12b, MGK14, SCN12]. Collision-Aware [XXCC17, HDM13]. Collisions [LCL^{+}20, SSH^{+}23, TWL22, JW11]. Colluding [HWJZ21]. collusion [LMP96, ZW10]. collusion-resistant [ZW10]. colocated [KS06]. colored [JRY09]. Coloring [CL17, LCK^{+}18, NSW11, CHM^{+}05]. Combat [WSC^{+}23]. Combating [HWLL21, TWL22, YLL21, FTV^{+}10, YMKC08]. Combination [KSC^{+}23]. Combinatorial [ÇY07, GKJ12, YPA21, YOY97, HKLS12, HS03, ZWTC16]. Combined [AABD13, BSM21, SKS16, XLT^{+}22, YASS15]. Combining [AdSD16, AN20, VCVC17, YSRL11]. Come [OLZ17, ODC^{+}16]. Comments [CBAT06, Far95, GLG04, HL05, Kar03, LRJ08, LYL07, OdG96, PK01, ZCW15]. Commercial [ALY^{+}20, LZL^{+}20, LGGZ10]. Commitment [SZ22]. Commoditized [RFGL17]. Commodity [BCC^{+}17, CZZ^{+}21, HCFC20, HCW^{+}16, JYL^{+}19, LPD^{+}18, LYL^{+}22a, XFCW18, YWZG23]. common [BM09, RW93]. commons [KAS16]. Communication [ACC^{+}14, AD18, ALPK21, AM19, BSP21, CDHM17, CWLH20, CLS^{+}19, CZK^{+}21, DTM^{+}17, DGW^{+}17, GLS21, GHZ20a, GHZ^{+}20b, JYL^{+}19, JHM^{+}19, JHLW24, KYM22, KIW^{+}17, KP21, LWW^{+}19a, LDD21, LGZ^{+}23, LCY^{+}19, LZC22, MWW^{+}21, PLT^{+}20, QLF23, RVR93, RZE^{+}21, SKE19, SLC22, SPR^{+}20, SBTH19, VBC^{+}17, WWL^{+}24a, WCM^{+}21, WCW^{+}17, XZL^{+}21, XTHL21, YPA19, YEMJ24, ZWJ^{+}20, ZFW^{+}17b, AA96, AKK13, ABJ^{+}13, BMB^{+}11, BCP00, BSNI06, BBL95, CS00, CBLVW06, DT93, GS97, GPM03, GL10, GF95, HJL^{+}12, HLHD^{+}04, HN10, JK05, KS95, KPP93, Kri14, LM13, LBHO07, LTB04, LO96, LH14, LNC93, LYL07, MKS16, MSP^{+}07, MDMM09, MP08, MP93, MW98, NOF14, ORS93a, RLA06, RS12, SZG^{+}13, SS04b, VGP14, YS93, YKGK13, ZYL^{+}14, ZPCS11]. Communication-Aware [YPA19]. Communication-Efficient [WCM^{+}21]. Communications [AN20, CXK^{+}23, CLS^{+}21, CDW19, GV17, LYZ^{+}23a, LFF^{+}19, NV21, SKA^{+}18, VBHT17, WCWZ17, WCK^{+}20, XLZ^{+}19, Ban99, CPGZ15, CJJ09, FHH10, FUDA03, FMT03, HL98a, HA96, HTC04, JCJ95, JR96, LZ09, LyT98, MHS95, MTK03, RPV13, SKE16, SL07b, WBP^{+}11, WGL00, WZL^{+}13, ZJ12]. communities [DPMK11]. Community [CLL^{+}18, CWZY21, DMDM17, FZW^{+}20, GZY23, PSST21, ZCZC17, DPBT11, MPF^{+}15]. Compact [AGCFV18, Hos98, KRRR17, QCMY16, XCZ^{+}17, YXL^{+}19, MWQ^{+}10, YLCP11]. CompactDFA [BBHK14]. Comparative [AT03, Kum98, CFPP96, CJ14, RrBG94, WS08]. compare [LS97c]. Comparison [LVB96, BPSK97, BO03, Far95, JGKT07, LNB00, LESZ98, McA94, MV14, RPGE04, RS95b, TAB^{+}15, ZCD97]. compatibility [QL16b]. compatible [GSRS^{+}15]. Compensated [YZLH17]. Compensation [DLR^{+}18, HK94]. compete [NJW16]. Competition [CBHS20, CQLW22, GHR14, KAS16, KS20, Ma16a, Ma16b, VNM22, GS16, LMW16]. Competitive [BBMELH08, BFG^{+}14, GV17, ORS93a, SZ22, SLF21, VN20, BCN02, CFS11]. Competitiveness [RTLC17]. Compiling [LZS^{+}22]. Complementary [SC18a, RS12]. Complete [AMS22b, FHMS18, WM95]. Completely [RR19a, RS21, SSWK13]. completeness [CBLVW06, OPW^{+}10]. Completion [CLY^{+}17, DZL^{+}20, HHL^{+}19, LFXY23, SG18, SV15, TXW^{+}21, WZL^{+}23a, XWW^{+}18, XWW^{+}19, XCW^{+}20a, XCW^{+}20b, XOW^{+}23, ZLN^{+}17, NAA^{+}16, Rum93]. Complex [CWZY21, HK94, Ili00, LRC15, SSV13]. Complexity [ABBH^{+}16, AZ09, DRMP18, DJS^{+}17, LFC18, LW13, SG17b, TAH99, VLM16, BSYS12, BSS11b, BMS14b, CN08, FMSM^{+}11, Guo04, GLS09, HLW13, JGLS14, JGS^{+}15, KR00, KV05, LSB06, LMS04a, LLS10, MP08, Val07, XL05, XCX^{+}06, ZCW15]. compliance [SBDR10]. Compliant [KLC^{+}18, LDRS18, BLPS10, RVS09]. Component [SWL^{+}18, WLL^{+}11]. Component-based [WLL^{+}11]. Composite [GLC^{+}16, Zha17]. Compositional [LN19]. Compound [PWWP18, RMPG16]. compounding [LMS04b]. Comprehensive [LJZ^{+}23, PCW^{+}16, SARM24, LBB08, SZM08, ZQ00]. Compressed [LLT^{+}16, Mit02, XLR13, ZLWM18, BLC12, BBK12, LMR99, LCY96, LyT98, ZG14]. Compressing [RTK^{+}16, DLT16, MLT12]. Compression [RT17, ZCPG^{+}23, BSF16, TSR14, THDD05]. compression-transmission [TSR14]. Compressive [CZC^{+}22, LLL^{+}16, WLW^{+}17, YCC21b, RZWQ12, ZL15]. compressors [CCL09]. Compromised [ZWY^{+}18]. Computable [ER23]. Computation [ALR^{+}24, CJLF16, CZX18, GZJ^{+}18, JD20, KYM22, LFC18, LCY^{+}19, PLT^{+}20, SRMB^{+}23, VLM16, VLDM17, YZL^{+}18, ZMW^{+}22, BL04, CSS08, FC99, Ili00, Nai97, NST00, RRG10, RGKS10, SGR13, Sob02, WB11]. Computation-Intensive [ALR^{+}24]. Computational [CK10b, GS97, LYW^{+}18, RRS22, WM96, ZLZL16, CN08, XL05]. Computations [ZWJ^{+}20, GLA93]. Compute [DRW^{+}22, MHR^{+}20, QSW24, CLW95]. computer [CSEZ93, GEHM02, Lev95, Mil95, SC95, WLS97]. Computing [BFS21, BWES22, CPKL17, CJLF16, CZX18, CYH^{+}18, CZD^{+}22, CKR93, CVM^{+}15, DEH^{+}07, DJB^{+}22, GO02, GZJ^{+}18, GXS^{+}21, HNP23, JD19, JD20, JD22, KYM22, KAK19, KR20, LLWB16, LYMA^{+}17, LXX^{+}24, LW22, NDGL06, NLB19, PCW^{+}16, QLY23, RMDJ16, SLC22, SZW^{+}16, SJWH^{+}17, TJD23, WUZ^{+}19, WRT^{+}21, WYL23, WXX^{+}24, WYL24, YDY^{+}24, YZL^{+}18, YPA19, YPA21, ZLRC20, ZSL^{+}21, ZWR^{+}23, ZTH^{+}23, ZLL24, ZLWH17, BBO^{+}05, JL12a, KL09, XGF^{+}14, ZRP00]. CompVM [SC18a]. concatenation [OSZ^{+}06]. Concave [CSN^{+}23, RS07]. concentration [CM93, MGR02]. concentrator [LT94a]. Concept [GA24, LAN97]. Conception [DLY^{+}22]. Concepts [VK04, CSMW02]. Concerned [WZL^{+}23a]. Concerns [LZL^{+}20]. Concise [PT12]. Concurrency [TZL^{+}24]. Concurrent [CLWZ17, CLS^{+}19, DJK23, GH04, IAS06, OJRCC02, RCOC03, SE21, WCK^{+}20, XWL^{+}18, ZWH^{+}17, LK10, NM09]. condition [FP97]. Conditional [CH20]. Conditions [CHS^{+}20, KV05, OPGT16, CGMS13, KCTI08, LZC09, MDL07, ML12, RLA06, SCKB09]. cone [LHB^{+}05, RB09a]. cone-based [LHB^{+}05]. conference [TWL05]. conferences [RVR93]. conferencing [CPS^{+}12, LZL11, ZLS96]. confidential [OÇ10, SKE16]. Confidentiality [SEK15]. Confidentiality-preserving [SEK15]. Config [BHC^{+}21]. configurable [BWH^{+}07, WWT05]. Configuration [APSG14, GPBL24, LTN^{+}19, SCHG22, SMC^{+}20, TTM22, ZXW^{+}20a, ZWJ^{+}22, ZZL^{+}22, ZJWY17, APB^{+}13, CGW^{+}12, CAH08, GQ16, KIR08, RBGK03, SS93, SS94a, TD03, YKKY08, ZBA16]. configurations [KSG11, KHČ^{+}09]. Configure [ZDZ^{+}24]. Configuring [PC19]. Confinement [NS16]. Conflict [DZ20, LS05b, NV21, PM96, PEA09, SHHA09, ZZW^{+}15]. Conflict-Free [NV21, PEA09]. Conflicting [WWW20b]. conformance [MP93, MP94]. Congested [Kop96, ZMWX18, BM93, WWTK11]. congested-queue [Kop96]. Congestible [Ma16b]. Congestion [AY20, AMS22a, BES22, CDHM17, CXL^{+}24, CL16a, CJ97, CDK^{+}17, DTM^{+}17, DS04, GKPS06, GR20a, JLW^{+}24, KIL24, LPJ^{+}17, Li24, LYS^{+}18, LYZ^{+}17, PWDL05, PT00, PLM^{+}16, PHC20, QAZ12, RS12, RAPP22, SG17a, SJZ^{+}24, SXEZ21, TWH24, WXN^{+}17, WXH^{+}20, WLL^{+}16b, WBM^{+}18, XCV^{+}20, YLH17, YSC18, YLF^{+}21, ZBC^{+}22, ZV16, ZCW^{+}22, ZMLR23, ZZW^{+}23, ZKEN23, ZQL^{+}23, AMP01, AVS04, AB05, AdE07, BO07a, BM93, BNJ16, BV05b, BYH^{+}15, BESW08, CGM04, CCV03, CBD02, CFM^{+}09, ES07, FJ93, FF99, GP96a, GLG04, GMSK09, HSH^{+}06, HPV09, HLW13, ILS97, JRL15, JGMB03, JVJ05, JBDF07, JJ08, JT01, KMR95, KK05, KKS^{+}08, KG99, KS03, KK06b, LMS00, LAJS07, LPIH11, LS99, LS06d, LSXS16, LR03, LSM^{+}14, MNR03, MOY00, MKT96, MW00, PM09, PILR05, RCS14, RJJ^{+}11, RX07, RKT02a, RS95b, ST05, SL05, SSM03, SWL06]. congestion [SLD14, TKN06, TWLC07, TWLC10, THP94, TLS^{+}12, TC06, Tia05, Voi07, WFGZ13, XHK^{+}05, XFS06, YSL^{+}14, YOY97, YS07, YDS06b, ZKL07]. Congestion-Aware [PHC20, SG17a, WXN^{+}17, YLH17]. congestion-based [JVJ05, JJ08]. congestion-controlled [GMSK09]. Congestion-dependent [PT00, RS12]. congestion-driven [MOY00]. congestion-free [ILS97, YOY97]. Congestion-Resilient [YSC18]. CoNICE [JR22]. connect [FJ07]. Connected [BTP^{+}17, DNS23, FSGH17, FWK17, Fuk20, GZL^{+}17, GZDG06, JR22, LWK^{+}16, RS21, SCC^{+}17, SLH^{+}19, WLK^{+}17, CB11, CCF04, HS06b, RYS12, SPR08b, SPR08a, ZG08, ZLW16a]. Connection [BWK^{+}22, BIS00, CGS93, HYLS21, SR01, CCL99, CZFF98, GS10a, LWL04, MH02, QY04, RS08, RLKT98, XHN04, YJ15]. Connection-Based [HYLS21]. connection-oriented [CZFF98, GS10a, LWL04]. connectionless [CPSWL96, KMS^{+}01, OKM94]. Connections [CMY^{+}17, CMY^{+}18, LKS^{+}16, RUH^{+}18, YWRK19, ZWH^{+}17, Ban99, CDFG06, CL04, ESG11, FP14, KKL03, KS12, LLY09, MMS01, Pax94, ZQ99]. Connectivity [ADT22, BB16, DYJ20, FFX^{+}17, FWK17, HNP23, JYT^{+}15, PSST21, RZS14, SQS20, WWMZ22, WYZ^{+}24, ZFW^{+}17a, ZM18, ZAW^{+}22, AG16, DBT05, HLP11, KLT15, LZF09, SKG12, SQ12, WLWL13, XK06b, YBX^{+}10, ZH08a]. Connectivity-based [JYT^{+}15]. connectors [Zeg95]. connects [DMK05]. Conquer [ZZW^{+}23, CJV16]. conscious [MPFK02]. Consensus [DBW^{+}20, JR22, SARM24, SCS^{+}22, WWL^{+}24a]. conservation [BYH^{+}15]. conserve [KD10]. conserving [CPR99, GK16, TG96]. Consider [SC18b]. consideration [YYZ06]. Considerations [GR20a]. Considering [SAC^{+}18, ZLL24, BH06, CH15, LZXF14]. Consistency [BWK^{+}22, KAT^{+}22, LLZ^{+}23b, GMD15]. Consistency-Aware [KAT^{+}22]. Consistent [AAT^{+}23, CB99, MVB^{+}21, MSM16, HAGL16, LDK12, WL07]. Consistently [ZP18]. Constant [BU21, HCFC20, WLK^{+}17, YXL^{+}18a, YNZ^{+}17, BSS09]. Constant-Rate [HCFC20]. Constrained [CE19, CWH^{+}16, CZX18, DTM^{+}17, DMLC18, DLY^{+}21, DRCM^{+}17, FFZ^{+}18, GJWZ16, KHYA20, LYZ^{+}23a, MHXT10, TXW^{+}19, WN16, ZBdV23, CKS16, CM05a, CCLT02, CSS08, Hou15, HH10b, KWCR10, KKP15, KLS11a, KK03b, LE12b, LCW^{+}15, LH10, MCLG07, PZGLA98, RMM99, RS00, SCR08, SG05, XZTT08]. Constraint [CMY^{+}17, GZCF06, LYLW22, SLH^{+}19, TTM22, YXL18b, ZSLZ21, DBL13, HMM11, JL12b, Kuc14, KLT15, NMC07]. Constraint-based [GZCF06]. Constraints [AIL23, BU21, CLTM22, CBV^{+}18, CGR^{+}18, CDL^{+}19, DPM^{+}18, KSM19, LWL17, LLS^{+}23, LZC20, PLT^{+}20, SJ21, SLC22, SCHG22, SLF21, SK21, TKM20b, TJD23, VN22, WDL^{+}23, Bej04, CTH10, GS10b, JF04, LS03b, MSS16, PPSV13, WQC06, WLLZ16, XSZ^{+}07, ZM09, ZOM03]. Construct [WLK^{+}17]. Constructing [LLZ^{+}23a, LHC05, WMFS10]. Construction [CMY^{+}17, Dat17, EF17, FAWW22, HK24, HZ20, LWK^{+}18, TWN^{+}20, TMGB19, YNZ^{+}17, ZYL^{+}17, ZML^{+}19, CL08, hCgKsYwT96, DLT^{+}15, RMM99, SK03, ST08, TAB^{+}15, WKA^{+}13, ZXTT08]. Constructions [CCL06, CCL09, NPQ06, SS10]. Constructive [DLZL17, RPP^{+}19, WHM^{+}13]. consumers [XYLL14]. consuming [SSZ05]. Consumption [GYSPR14, LS16, YHH^{+}21, CK09, CMFA14, SGSB^{+}15]. contact [WMS09, ZLSK15]. Contactless [CCX^{+}23]. Contacts [HCL^{+}17]. Contagion [HAG19]. Container [ZLW18]. Containerized [SMEH20]. containers [LZXF14]. containment [WNV13]. Content [AS14, ATE21, ATEY22, ATE22, ATE23, ASKL18, AGL16, AAG^{+}16, ADR18, AN20, BSG^{+}18, DRCM^{+}17, DJS^{+}17, DCN^{+}19, GTU19, GSM^{+}17, HSL20, KS20, KLKP16, LMSR19, LYS^{+}18, LZC^{+}17, MYMY17, MHH20, MDL^{+}13, MJ17, PD16a, PLS^{+}21, SS16, SZWW22, TEE16, VWNT17, VNM22, XHY^{+}22, ZLW^{+}16b, ZLL^{+}23b, ACR12, AJF11, BCMR04, CKS16, CKR^{+}09, CG04, CY14, CKC^{+}13, LMT16, MCL^{+}11, MOR13, MJ14, RB02, SG96, SD15a, SJ10, SYJ09, TM13, WS08]. content-based [MJ14]. Content-Caching [KLKP16]. Content-Centric [GTU19, MYMY17, PD16a, SS16, ZLW^{+}16b, AGL16].

Contention [CSN06, KLVL19, ZZ17, ASSK13, DM03, SG96, YWK07, YD07, YDS10, YCL09]. Contention-based [CSN06, DM03]. Contention-Free [ZZ17]. Contest [SH23]. Context [DKSC18, KTvdSK18, LZY^{+}22, LG13b, SMEH20, SKA^{+}18, WZ16, LMW16]. Context-Aware [DKSC18, KTvdSK18, SKA^{+}18, LG13b]. Contexts [RMDJ16]. Contextual [YPA21]. Contextual-Combinatorial [YPA21]. Contiguous [SC22b]. Contiguous-Resource [SC22b]. Continuous [CK11, CMY^{+}18, GLM^{+}16, JZW^{+}18, WLL^{+}24, And04, AS02, GZT03, qLH93a, NABZ12, TX08, VNS02]. Continuum [MCdG23]. Contract [AN20, MGLH18, SL14]. contracts [RS12]. contributory [MSWL06]. Control [AAT^{+}23, ACDP17, AWH^{+}22, AHEK24, BT23, BD97, CLTM22, CCE^{+}17, CDHM17, CS17, CLX^{+}22, CLZ^{+}23, CL16a, CKZC19, CDK^{+}17, DTM^{+}17, Dai22, DSYN24, EML12, FLTM18, GJCB18, GKB^{+}16, GSM16, GCMP20, GDL^{+}22, HS19, HCW^{+}16, HBSX20, IKS17, JLW^{+}24, KA20, KLK^{+}20, KIL24, KES13, KLP16, LAV16, LPJ^{+}17, LLZ^{+}23b, LZX^{+}21, LXL^{+}22a, Li24, LJHB18, LAJ20, LLM23, LGCG^{+}21, LYZ^{+}17, MGK20, MMG22, PLM^{+}16, QZL^{+}16, QYZX22, RAPP22, SJZ^{+}24, SM18, SX16, SXEZ21, TWH24, URZ^{+}14, WLTJ19, WN17, WXH^{+}20, WBM^{+}18, XGQ^{+}19, XCL^{+}22, XHZ^{+}19, YLF^{+}21, YN18, ZBC^{+}22, ZV16, ZGL^{+}19, ZLX^{+}21, ZSL^{+}21, ZCW^{+}22, ZZW^{+}23, ZZLW16, ZRH18, ZQL^{+}23, AK01, ACOR99, AA04, AMSS08, AMP01, AAM05, ASCG08, AB05, AABD13, AADS05, AZR97, AL98, AOM04, BBG11, BCP00, BCL12, BHL07, BM93, BLCT97, BFMF01, BLT02, BS08, BCGM07, BSP07, BYH^{+}15, BESW08, CFP^{+}09, CGM04]. control [CDFG06, CBD02, CLM99, CH93, CFM^{+}09, CLD10, CYG^{+}14, CLK01, CSN06, CCKK16, CWW^{+}15, DLT16, DM14, DS04, DK98, DM96, EF08, EM93, ES07, EOSM10, FKT98, FF99, FMT03, GP96a, GKPS06, GHK02, GNP^{+}13, GP96b, GT99, GT03, GMY13, HP01, HIM07, HSH^{+}06, HRCW08, HDM13, HLW13, JR14, JDSZ97, JCJ95, JGMB03, JT01, KMR95, KK16a, Kar03, KK05, KWS10, KR99, KA95, KG05, KEY99, KqL98, KS03, KK06b, LA02, LCM04, LMR99, LMS12, LMS05b, LPIH11, LS06b, LA95c, LCH95, LHB^{+}05, LH05, LM15, LWF96, LyT98, LS06d, LT95, LJNK12, LSXS16, LR03, LL99, LKZ^{+}04, LRG10, MGK14, MOR13, MPS01, MH02, ML12, MLS12, MKT96, MW98, MW00, NM09, NKS08, NML08, Nee09, NS98, PWDL05, PM09, PDSK04, PG93, PG94a, PV10, PSA96, PPV12, PFC96, Pil01, QAZ12, QCS07]. control [QK01, QS04, QS05, RKZG10, RS97a, RJJ^{+}11, RLA06, RSS09, RX07, RV01, RS95b, RYS12, SMGP15, SEK15, SKE16, ST05, SL05, SKKA01, SWL06, SL07a, SBP03, SHN16, SMM11, SKS16, SR01, STC12, SDW00, SL07b, TKN06, TPC09, Tan16, TWL06, TWLC07, TWLC10, TAJ^{+}10, THP94, Tia05, TdWC^{+}94, TLP^{+}16, Voi07, VL05, VA06, VA07, VA09, WBEGS05, WPL06, WKWV16, WCH95, WD05, WLL01, WWL02, WFGZ13, XY10b, XHK^{+}05, XSC01, XSC03, XFS06, XC08, YWK07, YKZ^{+}13, YJ15, YHE04, YS07, YJH05, YM05, ZSSK02, ZS03, ZKL07, ZLW16a, dAF04, AMS^{+}08]. control-plane [TLP^{+}16]. Control-Theoretic [WBM^{+}18, EML12, KR99, LyT98]. controllability [JPS04, JS06]. Controllable [LLM23, LNLM24]. Controlled [CL07, DKN21, TR17, AQJRS16, BBM93, BKTN03, GMSK09, Hon94, KV98, KVR98, LAPS08, LL95, LKC11, LK13, ML06, XSC01, YL97]. Controller [DQYG23, GDL^{+}22, GDWX23, GDJX24, HBSX20, JM17, TW22, WLX^{+}17, WXH^{+}18, ZML^{+}19, ZMLL21, BL94, CC96, CCL99, HP00, KR99, LLD96, PILR05]. Controller-Assisted [ZML^{+}19]. controllers [RCS14, SSM03, SLD14, YDS06b]. Controlling [WD22]. controls [Smi95]. conventional [CFPP96]. Convergence [CMP^{+}14, DTN^{+}21, DME23, FSGH17, JJJ^{+}23, KSRW22, KHAWC17, ML18, Nee16b, Nee19, Nee22, ZWL^{+}22, FB07, Kar03, LABJ01, qLlH97, LLE15a, LR03, LL99, MMH^{+}15, YMO97]. Convergence-Preserving [JJJ^{+}23]. Convergent [LLX19a, SLJJ16, BS08]. Converse [Nee22]. conversion [CL05, DMK05, Hos98, KA98, NPQ06, QY04, RM02, RS98, RZVZ06, SAS96]. converter [SAS99, ZY07b]. Converters [ZDZ^{+}24, CM05b, NPY07, SJGH10, XL99]. convertible [ZZZ^{+}07]. Converting [KA20]. Converting-Space-Converting [KA20]. Convex [PDI20, SLF21, VL16, Ber00, CGMS13, LMS05b]. Convolutional [ZXW^{+}21]. cooperate [KKEE13]. cooperate-to-join [KKEE13]. Cooperation [DZL^{+}18, KNK^{+}14, MQ05, SR14, XL23, WFH12]. Cooperative [CFP^{+}21, CGYZ16, CSR^{+}17, EFA19, LKS^{+}16, LZC^{+}24, LNL^{+}16, LSL17, SKY10, SJWH^{+}17, SAM10, SSAK12, WSZL20, XWWC16, YXY^{+}18, ZCZ^{+}21, ZS13, AK14, AVPG14, CFG08, CBL13, CPGZ15, CW10, EH11, GMY13, GMYP16, GLLJ16, HS06b, IK09, KEW06, LZES14, MCL^{+}11, MEWP13, SSHK11, SYJ09, SMSM06, WQZ^{+}13]. Cooperatively [LLZ^{+}23a]. Coordinate [BCD19, CLY^{+}17, CGMS13, KBS11, LZSS10, LHC05, TYLH09]. coordinate-convex [CGMS13]. coordinate-free [KBS11]. Coordinated [AM19, LK02, MAPZ18, PD16a, RTNS21, WLL^{+}16b, CRB12, LK05, LPCVC13, YJ15, YHE04]. Coordinates [JPM^{+}19, DJ14, SBNRS14]. Coordination [CWZ^{+}17, DMMS14, KLP16, LCK^{+}18, SFM^{+}18, TW22, XLZ^{+}19, CHH06, GR01, MGK12, MDL07, RD11a]. Copa [JLW^{+}24]. copy [MHSC95, Ses97, SM00, SPR08b, SPR08a, ZKO93]. Core [CHO^{+}19, CWAO21, LNG^{+}21, LLCJ22, NEH^{+}22, YYB^{+}22, CHM^{+}05, EKD12, LBS11, LC04b, ZBA16, SSZ03]. Core-Stateless [LNG^{+}21, SSZ03]. correcting [BDS07]. Correction [BBLV06a, CWLH20, AD11, BMB^{+}11, Kri14, SCY08]. Corrections [AMS^{+}08, DKN97, JTL^{+}18, LCS^{+}18, SM19, XCR15, ZND^{+}16, ZCW15]. Correctness [Sob17]. Correlated [CKA16, HDF19, LCH20b, NDN^{+}18, ZFW^{+}17a, AT03, CMGL11, CBL06b, CBLVW06, Nee16a, PG94b, TSR14, VR13]. correlated/unbalanced [PG94b]. Correlation [KWH^{+}17, LCH23, CAO11, qLH93b, qLH93a, VA06, WA11, ZZHZ13]. correlation-based [VA06]. Correlations [La17, WZL^{+}23a]. CoSchd [WLL^{+}16b]. CoShare [WTJR22]. Cost [AdSD16, ALYX22, BWS10, CNM20, CCW^{+}17, CZL^{+}19, CLH^{+}24, CKS17, hCgKsYwT96, CR14, CDL^{+}19, DPM^{+}18, DZNT14, ECL^{+}20, GRS00, JWW^{+}23, LZL^{+}18, LZW^{+}21, LTZ^{+}22, LGS^{+}23, LS17, LSC^{+}21, LYW^{+}18, MCES19, PBT^{+}20, RG98, RS21, SZ22, SLH^{+}19, SBTH19, SRB^{+}20, WTXT11, WLTJ19, WZL^{+}23b, WZH^{+}24, WLW^{+}17, XLAC16, XYQ^{+}17, XXZ^{+}22b, YCW^{+}19, ZND^{+}16, ZWL^{+}22, ZRH18, ZCM14, AADS05, CMN12, CK00, CDM93, DFGV11, FEC13, HSE97, JLX^{+}16, KK93, LGW^{+}11, LPP11, Lin97, LRM^{+}06, PZGLA98, RV01, SHZ16, SML04, XY10a, XK06a, YYZ06, ZQ99, ZKL11, ZNZT16, ZWYY10]. Cost-Aware [CNM20, WZH^{+}24]. cost-benefit [AADS05]. Cost-effective [BWS10, CR14, DZNT14, GRS00, LGW^{+}11, SHZ16, ZQ99]. Cost-Efficient [ECL^{+}20, LZW^{+}21, WZL^{+}23b, WLW^{+}17, YCW^{+}19]. Cost-Minimized [LTZ^{+}22].

Cost-minimizing [hCgKsYwT96]. cost-performance [SML04]. cost/performance [CDM93]. Costs [KAA^{+}18, LCY^{+}19, PGMR18, SLF21, YEMJ24, ZHW^{+}17, ZLL24, CSG14, FK07, HA96, Ili00, LZ13]. COTS [LGZ^{+}23, OLZ17, WXJ^{+}17, YLL^{+}17]. could [PES^{+}12]. council [RSZ04]. Count [RAA^{+}24, ECN09, WJS07]. Counter [CCC17, EFFK18, KE21, LXL^{+}22b, NS16, SXQ^{+}23, TWL06, HXLZ11, KK06a, LCL12a, LT94a, RSU^{+}09, WZLX12, CCC17]. Counter-Intuitive [KE21, TWL06]. counter-rotating [LT94a]. counterfeits [GSN^{+}16]. Countering [XS21]. Counterintuitive [QTE20]. Countermeasure [AHX19, QLQ^{+}22, CHL16, KVF^{+}12]. Countermeasures [LHL^{+}23b, MRMR17, XCL^{+}22]. counterpart [XCX^{+}06]. Counters [LXL^{+}22b, RR19b, SXQ^{+}23]. Counting [GLC^{+}16, LI21, LXL^{+}22b, RR19b, EVF06, FDG^{+}10, HLZ^{+}14, LLY^{+}12, RKK14, ZCY16]. CountMax [YXY^{+}18]. country [DSA^{+}14]. country-wide [DSA^{+}14]. Counts [FBRL18, WLD^{+}16]. Coupled [CAK12, FSGH17, NLNL16, WN17, BMS14a]. couplers [GT00]. Coupon [MV08]. covariance [DL04]. Cover [ZWL^{+}16, GZDG06]. Coverage [ADT22, ÇTD22, GCWC17, GFW^{+}18, GSW^{+}23, LHY^{+}23, LJSB22, LKMK20, NK20, PBV17, QSW24, SK10b, WY06, ZWL^{+}16, ZWYD18, ZSLZ21, GZCX16, KBS11, KBS12, MP94, TXL^{+}12, XK06b, YKR11, YBX^{+}10, ZG08]. Coverage-Aware [ADT22, LKMK20]. Coverage-time [SK10b]. Covering [DNS23]. Covert [HXZ23, JHLW24, WXW15]. Cow [WTK^{+}17]. CPHR [WBWV16]. CQBT [RW96]. CQF [YCZ^{+}23]. CR [YCL09]. CRC [SGPH98]. Create [NST^{+}16]. creation [SLL^{+}11]. Credentials [WCH^{+}24]. Credit [BTC01, AS02, LYS93, LMP96]. Credit-based [BTC01]. Criteria [AWM^{+}20, SF23, RPF^{+}14, WC08]. criterion [AOM04, LK05, SD15b]. Critic [XYT^{+}21]. Critic-Based [XYT^{+}21]. Critical [BC01a, BCLS17, CLK^{+}24, DQYG23, FM06, SGL^{+}22, YS21, YXZ17, DZNT14, GGL09b, LKC^{+}13, SNXT13, TT09, TKI^{+}15, ZH08a, ZTS11, ZPCS11]. critical-load-based [ZTS11]. CRMA [SS94b]. CRNs [QDD^{+}17]. Cross [BAB20, BDR22, CBL13, CWLH20, CLS^{+}19, CH11, CGK10, DQW^{+}23, GHZ20a, GHZ^{+}20b, HLZ^{+}21, HTW^{+}22, HK11, JYL^{+}19, KIW^{+}17, KT06, LML11, LZS^{+}22, PNRMC13, RGG11, WLLD05, WVG12, WS05, XZL^{+}21, ZBC^{+}22, CK10a, CDFG06, CL03, CBL15, CCF04, DMK05, EOSM10, FJ07, Geo08, LSLL14, LS06d, PDE08, SLP07, SHHA09, SH07, SPB16, VA09, XE13]. cross-bar [Geo08]. cross-connect [FJ07]. cross-connects [DMK05]. Cross-Datacenter [ZBC^{+}22]. Cross-domain [CBL13, CBL15]. Cross-Language [LZS^{+}22]. Cross-Layer [BAB20, BDR22, CH11, CGK10, HK11, KT06, LML11, PNRMC13, RGG11, WLLD05, WVG12, CK10a, CDFG06, EOSM10, LSLL14, LS06d, PDE08, SLP07, SHHA09, SH07, SPB16, XE13]. cross-path [CL03]. Cross-Slice [DQW^{+}23]. Cross-talk [WS05]. Cross-Technology [CWLH20, CLS^{+}19, GHZ20a, GHZ^{+}20b, HLZ^{+}21, JYL^{+}19, KIW^{+}17, XZL^{+}21]. Cross-Tier [HTW^{+}22]. crossbar [HSG^{+}08, Kok10, NMC07, RCGT06, Tur09]. Crossing [CE09, CFP^{+}21]. crosspoint [SPC10]. Crosstalk [BHN11, CWAO21, CTH10, JSuRKH03]. Crosstalk-Avoided [CWAO21]. crosstalk-free [JSuRKH03]. Crosstalk-preventing [BHN11]. Crowd [JSXN18, JHS^{+}19, LXW^{+}20, LL17b, LLX^{+}19b, NL16]. Crowd-Sourcing [LL17b, NL16]. Crowdcast [WZW^{+}20]. crowded [SJL^{+}16]. Crowdfunding [SCH23]. crowds [CZCC14, RS05]. Crowdsensing [AWH^{+}22, FMK^{+}18, GS19, HHL18, LZY^{+}22, LZY20, WLY^{+}24, WLL^{+}16a, WWX^{+}19, XGW^{+}20, ZCPG^{+}23, ZYH^{+}21, HZL16, YXFT16, LYW^{+}21, WZL^{+}23a]. Crowdsourced [HZ20, JZ18, PWL^{+}22, TPW^{+}18, WWW^{+}18]. Crowdsourcing [CBV^{+}18, KTvdSK18, MGLH18, SYG^{+}22, SH23, XZC^{+}20, YXFT16, ZLM16]. CRT [CLP12]. CRT-based [CLP12]. Crying [KHW12]. Cryptography [vRDHSP17]. CSI [JM17, TWL^{+}21, ZWS^{+}17]. CSI-Based [ZWS^{+}17]. CSIsnoop [ZK19]. CSMA [ASSK13, BT23, BK17, CCL11, GSK08, GS13, HL15, HK11, JZC11, JW10, JW11, JP13, KL12, KNSV13, KLC15, Kon06, KLE16, LK16a, MAE19, MAE20, NTS12, QZZ^{+}13, SKK07, SCN12, SK21, SGJ17, SBGJ18, Van17, VBHT17, YLY^{+}16, ZZLM23]. CSMA-based [KLC15]. CSMA-CA [JP13]. CSMA-like [HL15]. CSMA/CA [BK17, HK11, JZC11, Kon06, LK16a, NTS12, SKK07, Van17, VBHT17]. CSMA/CA-based [HK11]. CSMA/CN [SCN12]. CSMA/PJ [ZZLM23]. CTC [HGZJ21]. CubicRing [ZLG^{+}17]. Cuckoo [RAA^{+}24, SXQ^{+}23]. cueing [BS02]. Cumulants [SJSB22]. cumulative [KWH11]. Curing [HAG19]. current [BB06, SGD05]. Curse [NGRF19, ZDCW18]. Curve [vRDHSP17, CAH08, LMS04b, SZN00]. CurveLight [YYT23]. curves [Wil96]. Customer [HDQ^{+}16, ZSZ^{+}17, SSA11]. Customers [AGMY21]. Customization [CBDCP19]. Customized [LPD^{+}18]. Cut [QFH^{+}18, Tas99, GL10]. Cut-Through [QFH^{+}18, Tas99]. cutoff [CSC94]. Cutting [CDL^{+}19, LLJ^{+}14]. cuttings [ST13]. Cyber [DLR^{+}18, GSKN18, LLX19a, PPS^{+}22, SHZ16]. Cyber-Physical [DLR^{+}18, GSKN18, PPS^{+}22, SHZ16]. Cybersecurity [MLB21]. cyberspace [CWSB05]. Cycle [BY06, CLWZ17, CWL^{+}21, CGC^{+}18, CNG^{+}16, KSSK18, CHML15, HLL13, KWCR10, SG96, WYH10]. Cycle-logical [BY06]. Cycled [CGC^{+}17, HLX^{+}15, LHC^{+}16, ODC^{+}16]. Cycles [SCHG22, Sob17, ARK09, CJ07, EM09, GR12, GR14, Kam10, MJ13]. Cyclic [GPBL24, LM96, LW11]. cycling [GTS^{+}09]. Cyclopathic [BY06]. CYRF [SL05].

D [HBH93, AMG^{+}17, CPGZ15, CQLW22, CXW^{+}18, FWN^{+}22, JYT^{+}15, KZ97, LZL^{+}14, LJL^{+}16, LDY^{+}16, NBV17, WJYL16, WXJ^{+}17, WYL24, WWL24b, YJZW15, ZCF^{+}24, ZGLC20]. D-BIND [KZ97]. D-Watch [WXJ^{+}17]. D/ [JYT^{+}15]. D2D [HAB^{+}22, KKH^{+}22, LBP^{+}17]. D2D-Enabled [HAB^{+}22]. DA&FD [ZGL^{+}19]. DA&FD-Deadline-Aware [ZGL^{+}19]. DAC [CGW^{+}12]. DACM [AJ06]. damage [KSA12]. Dampens [ZSS^{+}20]. Dampers [ML22a]. Dark [XYA^{+}21, YPA21]. DART [EFK07, XCV^{+}20]. Data [AB21, AGCFV18, AEG^{+}17, APC21, BCC^{+}17, BSRdA16, BWES22, BZS23, CZP18, CWH^{+}16, CZX^{+}17, CWM^{+}17, CGC^{+}18, CLM^{+}18, CLS^{+}21, CKZC19, CHFH20, CL19, CCG20, CDK^{+}17, CXW^{+}18, DSL^{+}18, DRW^{+}22, DLC^{+}18b, DLZL17, DXX^{+}23, EPS21, FMH^{+}21b, GWQ^{+}23, GYSZ19, GS19, GXL^{+}21, GXS^{+}21, HTW^{+}19, HK24, HFF^{+}24, HCW^{+}16, HHL^{+}19, HZC^{+}19, HLL^{+}21, HK14, JR21, JLSB16, JSXN18, JHS^{+}19, JWSH18, KYM22, LLZ^{+}23b, LGY16, LCL16, LLZ^{+}17, LCL^{+}18, LXW^{+}20, LDW^{+}20, LZS^{+}22, LS22, LLCJ22, LPS19, LCH20b, LCH23, LXX^{+}24, LWC^{+}23, LCY^{+}19, LHZ^{+}19, LAJ20, LX21, LSC^{+}21, LHL^{+}23a, LX24, LSL^{+}18, LSSK17, LSL17, LSL^{+}21, LFXY23, MGZ^{+}23, MS17, MBI^{+}17, NLS19, NHLB21, PBSS23, PWL^{+}22, PKK18, PJDS18, PRH17, QZL^{+}16, QFH^{+}18, RRS23, RLZ^{+}18, RDR17, SS17, SRC^{+}20, SLD^{+}23, SC22a, SYG^{+}22, SBC^{+}17, TJL^{+}19, TML^{+}18, TXW^{+}21, TXL^{+}18, VHT21, VPC17, VWNT17, VNM22, WJ17, WXN^{+}17, WLX^{+}17, WMX17, WZL^{+}23a]. Data [WLY^{+}23, WFZ^{+}23, WSX^{+}23, WN17, WWX^{+}19, XLAC16, XCZL20, XCQ^{+}23, XWW^{+}18, XPW^{+}18, XWW^{+}19, XCW^{+}20a, XCW^{+}20b, XQG^{+}22, XTW^{+}22, XXZ^{+}22b, YLH17, YLF^{+}21, YWRK19, ZCPG^{+}23, ZHC16, ZWGC17, ZCB^{+}17, ZCZC17, ZHZ^{+}18, ZZH19, ZHGF19, ZYH^{+}21, ZHCC24, ZDB^{+}17, ZQ23, ZLZL16, ZDZ^{+}24, ZBZ^{+}19, ZLW^{+}16b, ZFW^{+}17b, ZHWH21, AC16, AK09, AF99, AJDH01, AZ11, BMB^{+}11, BV96, BK00, BKTN03, BK06, Bor05, CKL16, CDI^{+}04, CT04b, CGW^{+}12, CZM14, CSS^{+}14, CYG^{+}14, CS15, CLL^{+}14, CM05c, CBL06b, CBLVW06, FML09, GIKK11, GIL^{+}15, HLX^{+}15, HRCW08, HY08, IGHT15, JCJ95, JC13, JRL15, KqL99, KR08, KWS^{+}11, LM13, LSS^{+}13, LLW^{+}09, LLY^{+}13, LGS09, LGW^{+}11, LLW^{+}12, LZZR12, LZXF14, LZW^{+}15, LS97b, LWAT13, LÜ14, LFZS11, LNL^{+}16, MEWP13, MG95, NCK15, ODT09, OSZ^{+}06, OÇ10, RP13, RVV^{+}15, SMH95, SLC^{+}07, SK13, SX10, SGPH98, TXL^{+}12, TX08, TRKN12, TAH99, VL97]. data [VCM04, WZY^{+}16, WCH95, WMFS10, WFGZ13, XLR13, XC08, YCV15, YAA09, YG10, ZM09]. data-center [LGW^{+}11, WFGZ13]. data-centric [AK09, CT04b, LM13]. Data-Driven [JHS^{+}19, LPS19, PJDS18]. data-gathering [LÜ14]. Data-Intensive [FMH^{+}21b, KYM22]. data-latency-bound [KWS^{+}11]. data-offloading [IGHT15]. Data-Parallel [LS22]. Data-Rate [DRW^{+}22]. Database [BPL20, HL98a, HA97, MD04]. Database-Driven [BPL20]. Datacenter [AHP21, DPSA21, FMH^{+}21a, HZB^{+}22, LZL^{+}18, LS22, RDZ^{+}19, SG17a, SLWW19, XCV^{+}20, YCW^{+}19, ZBC^{+}22, ZWH^{+}17, ZZW^{+}23, ZZZ^{+}14]. Datacenters [BHC^{+}21, CLL^{+}19, FSSC18, JWL^{+}18, LPJ^{+}17, LGHL17, LW20, MHR^{+}20, SG18, SC17, SC18b, ZDCW18, ZLW^{+}20, GLLJ16, SSWK13]. datagram [AC98, EAB01, WCH95]. Datapath [ZZZ^{+}24]. Dataplanes [TCTP20]. Datasets [DLL^{+}20]. Datum [RLZ^{+}18]. Day [ABBF19, FAF^{+}17, LSM^{+}14, WLC^{+}10]. DC [BHC^{+}21, ZGYB20, ZYS^{+}23]. DCF [LLY^{+}16, SD15b, ZTS11]. DCN [CYX^{+}17]. DCNs [ZGL^{+}19]. DDCA [CXL^{+}24]. DDoS [CLX^{+}22, CLH^{+}24, FAB12, LZW^{+}21, LJHB18, RSU^{+}09, WCCM18, XY09b]. DDoS-Attack [LJHB18]. DDoS-resilient [RSU^{+}09]. DDoS-shield [RSU^{+}09]. De-Anonymizability [FZQ^{+}22]. De-Anonymization [JLSB16, CGL16]. De-Anonymizing [FZW^{+}20]. De-Compositional [LN19]. De-randomizing [BV05b]. Deadline [CE19, LWL17, SK21, YSZL15, ZGL^{+}19, ATB^{+}10, AS02, FP97, LLLS07, LE12b, WLLZ16]. Deadline-aware [YSZL15]. Deadline-Constrained [CE19, LE12b]. deadline-credit-based [AS02]. deadline-driven [ATB^{+}10]. deadline-ordered [FP97]. Deadlines [GLS21, LFXY23, RL18, ZCB^{+}17, ZLWH17, HR95, MKS16, ZB95]. Deadlock [HZC^{+}19, IZC00]. deadlocks [KGL03, MG95]. Death [LAV16, TT17]. Debugging [ZBZ^{+}19]. Decentralized [CN10a, CCZL23, CVV17, CL17, CXL^{+}24, DD24, DPR06, DBL13, FXQ^{+}21, HK14, JD19, KLKP16, MAN15, MGK20, SQS20, SK21, WYL23, ZZLW16, AVPG14, LCM04, LYRL07, LDGL13, ST09, YKZ^{+}13]. Decentralizing [MVCS16]. Decision [CCK16, KE21, LZL^{+}21, SC17, WUZ^{+}19, XZC^{+}20, XLD^{+}24, AS94, ACR12, RV01].

Decision-Making [KE21, XZC^{+}20]. decision-supporting [ACR12]. decisions [ZZG^{+}16]. Declarative [LCL^{+}12b]. DECO [KYM22]. decodable [SV15]. Decoding [CW23, JHM^{+}19, OLZ17, XZG20]. Decomposition [ALR^{+}24, APSG14, JK15, KSNR20, SCS^{+}22, VIT21, ES05, GT03, LWL04, SAM10, TK12, YDS06a, ZRP00]. decoupled [RYS12]. Decoupling [GHBSWV17, GHK18, LNG^{+}21]. Decreases [ZHCL17]. Decreasing [LTCS22]. Decreasing- [LTCS22]. Dedicated [YWRK19, LW13]. Deduplication [EGKM16, XHY^{+}22]. Deep [AY20, BPW23, CCZL23, FGR^{+}17, HLZ^{+}21, HFF^{+}24, HNP23, HGZ^{+}23, LLZ^{+}19, LYH^{+}23, SNZ^{+}23, SGL^{+}22, TWY^{+}20, WZW^{+}20, WLY^{+}24, WHC^{+}22, XOW^{+}23, YXL^{+}19, ZCPP22, ZGZ22, ZGS^{+}24, ZQW^{+}23, ARS16, BAC12]. DeepCast [WZW^{+}20]. DeepCC [ZCW^{+}22]. Default [ZXC^{+}18]. Defects [RHX^{+}20]. Defending [FWN^{+}22, LWL^{+}11, YLLY05, YKGF08]. Defense [CLX^{+}22, CLH^{+}24, HLZ^{+}21, LZW^{+}21, LLX19a, SS21, WJS07, YFM^{+}22, ZCPP22, AC09, CLSS09, YGKX10]. Defenses [KSC^{+}23, WXC^{+}24, YLK^{+}17, ZLX^{+}21]. Deferral [VBHT17]. Deficit [TL22, KWJY16, LMS04a, SNS12, SV96]. deficit-based [SNS12]. Defined [AAR18, ACDP17, BTK^{+}17, CLK^{+}24, CPKL17, CYH^{+}18, CKZC19, CSR^{+}17, DQYG23, FXHY21, FLMS18, GJD18, GSM^{+}17, GDL^{+}22, GDWX23, GDJX24, HNW17, HSL20, KLKT16, LZZ^{+}22a, LXZ^{+}21, MGZ^{+}23, MSM16, NJK^{+}19, SM17, SM19, SBC^{+}17, SWH19, SGL^{+}22, TML^{+}18, TTM22, TTCT19, WMP^{+}18, WDR^{+}20, WBY^{+}17, WGZC21, XHC^{+}18, YXC^{+}18, YXCH21, YZGC23, YLK^{+}17, YXY^{+}18, ZXW^{+}20a, ZZH19, ZLX^{+}21, ZZX^{+}21b, HA16, LNL^{+}16]. defining [CWSB05]. definitions [TG97]. Deflection [YZLH17, BBFG95, BP96, CFC01, Lie97, PYL99, VL99]. Deflection-Compensated [YZLH17]. Defragmentation [BCO17, ZYZ16]. Degenerate [LSMS06]. Degeneration [GZY23]. Degradation [AEG^{+}17, DAA19, LD95]. degradations [VC12]. Degraded [VWT^{+}14]. DeGrading [CH21]. Degree [KK16b, La17, TMGB19, OR11, ZSCJ14]. Déjà [SPGM13]. Delay [ATE21, BLC21, BBF18, BBC^{+}02, BM22, CFG08, CGC^{+}17, CLQ^{+}19, CDK^{+}17, DTM^{+}17, Dat17, DPTP24, DV09, EE18, FZ16, FFZ^{+}18, FM22, FqL98, GDC^{+}17, GLA19, GS10b, GS11, HLZY23, HYLS21, ITSO01, JK96, JV17, JJS13a, JLW^{+}24, KIL24, KLE16, LNG^{+}21, LSS^{+}13, LK16b, LWAL17, LYD^{+}21, LPM23, Liu10, LZC20, LDZC20, MYMY17, MMT16, MNR03, McM95, MKG^{+}17, Nee09, Nee13, PYL^{+}17, PBSS23, PJM^{+}19, REM17, RTNS21, SBD11, SG17b, SMS07, SH14, TWN^{+}20, TWH24, TML22, TW22, WHW^{+}11, WLD^{+}16, WJ17, WLTJ19, WDL^{+}23, XL95, XPL^{+}17, XXZ^{+}22b, XE13, YSC16, YSC18, YLF^{+}21, YLY^{+}16, ZS03, ZKL07, ZHCL17, ZCZC17, AB05, AWKN16, AD11, AABD13, ALMR14, BBG11, BO00, BS15, BLS07, BBM^{+}10, BSS^{+}11a, BSS11b, BWS10, CZF^{+}16, CS99a, ÇM15, CLC^{+}01, CU95a, CCL09, CFM^{+}09, CS14, CMGL11, CK09, CYL16, DSR02, DL04, EMPS06]. delay [FP95, FSM14, GS13, GIKK11, GCS06b, HPV09, Hou15, HL05, HMM11, HMNK13, HLW13, HL15, HKT95, JR14, JGLS14, JGS^{+}15, Jia98, JS14, KR00, KLSS10, KLS11a, KCB03, KK03b, KCCM16, KS98, LM97, LS97a, LL98, LDK13, LLY01, LM01, LLE16, LK14, LWF96, LZC09, LHC05, LSMS06, LJNK12, LWR15, LDHT02, LLS09, LNC04, MJ15, MH97, NMC07, Nee08, NTS12, ORS93b, PZGLA98, PPSV13, Pil01, RMM99, RS00, RZZ06, SSM03, SAKS13, Smi08, SV15, SS05, TS08, TG97, UN11, WMS09, WVG12, WDCL15, WH97, WKZL96, XL05, YW11, YCV15, ZS04, ZNN^{+}10, ZW14, ZM04]. Delay-Aware [CLQ^{+}19, YCV15]. delay-bandwidth [LNC04]. Delay-Based [JLW^{+}24, LWAL17, YLF^{+}21, JJS13a, MNR03, Nee13, BSS^{+}11a]. delay-boundary [LM01]. Delay-Bounded [CGC^{+}17, HL05, Jia98, Pil01]. delay-capacity [LSMS06]. Delay-Constrained [DTM^{+}17, FFZ^{+}18, Hou15, PZGLA98, RMM99, RS00]. delay-endurable [YW11]. delay-friendliness [BBM^{+}10]. Delay-Guaranteed [KIL24, XE13]. Delay-independent [ZKL07]. Delay-optimal [SBD11]. Delay-Oriented [HLZY23]. delay-power [BBG11]. Delay-Sensitive [XXZ^{+}22b, KLS11a, LL98]. delay-throughput [CMGL11]. Delay-Tolerant [MKG^{+}17, WDL^{+}23, LSS^{+}13, AD11, AABD13, BWS10, CS14, SAKS13, UN11, WMS09]. Delayed [CL19, GLA19, JM17, LABJ01, MS17, SSG18]. Delays [ER23, HPP^{+}23, TSS14, VPC17, BR06, BLC11, CAH08, JT01, LKC^{+}13, RLA06, SBP03, Tia05, YDS06b]. deliver [LLY^{+}13]. Delivering [CS99a, GZT03]. Delivery [ASKL18, BSG^{+}18, CKZC19, DZH19, GLA19, GSKN18, HSL20, KCM16, MBN^{+}21, SCN^{+}24, XHY^{+}22, ZLX^{+}23, BCMR04, CF98, DLH^{+}14, LQ13, MOR13, RKNS10, SD15a, TYLH09, ZWDS00]. delivery-guaranteed [TYLH09]. deluge [TRKN12]. Delving [WCCM18]. Demand [AMS22a, AJ06, CZP18, CN16, CM18, GXW^{+}19, HH18, KLK^{+}20, LCU^{+}20, NST^{+}16, SAKMB21, SMEH20, SJ10, SC18b, TE16, ZZLW16, AF99, BK06, DYX12, LZW^{+}15, MEVSS03, MW05, PWMC12, PL02, TM13, ZEV07a, ZEV07b]. Demand-Aware [AMS22a, SJ10]. Demands [PG21, TWWG19, ZKEN23, AC06, CAQ07, FGL^{+}01, MG97a, YNDM09, ZBA16]. demultiplexer [BKH^{+}93]. demultiplexer/descrambler [BKH^{+}93]. Demystifying [LL13]. Denial [AAS14, XS21, AHK08, KK06a, YLLY05]. Denial-of-service [AAS14, YLLY05]. Dense [BPST18, CLZ^{+}20, GB18, LL17a, NSY20, NSY23, SRBBG17, SFM^{+}18, GMP13, OGLK14]. Densely [GZJ^{+}18]. Densified [MKS17]. Density [LMP08, AGLM10, ZW14]. Density-based [LMP08]. departure [CLC^{+}01]. departures [LBS11]. Dependability [MBL19, WLS97]. Dependability-Based [MBL19]. Dependable [FMCS20, GWQ^{+}23, GPM03, MMS01]. dependence [GB99, HL96b, RVA00]. dependencies [HSPH09]. Dependent [ALR^{+}24, AGMY21, BU21, BSP21, CXL18, JZW^{+}18, NSY23, TTM22, CLW95, CKR93, CNP13, ENW96, LB04, PT00, RS12, SD00, THBR14]. Deployability [LXLC20]. Deployable [WYL^{+}22, ZSL^{+}17]. Deployed [DYW^{+}16, GZJ^{+}18, TWY^{+}20, WY06]. Deploying [BDHR10, KLLT18, MGS^{+}21]. Deployment [BBR19, CCK16, CQLW22, CLP^{+}17, DYJ^{+}23, DLLL16, FBM^{+}21, LXZ^{+}21, MCES19, NKL^{+}23, RGY^{+}22, SMEH20, XLH^{+}17, XXZ^{+}22b, ZLZ21a, ZWR^{+}23, ZGLC20, CFD06, HPR06, LC97, SHZ16, SLO^{+}14, TBV^{+}13, YBX^{+}10, YBX^{+}12, ZSK12]. deployments [Kuc14]. Depot [JLS^{+}17]. derived [Pax94]. Deriving [FGL^{+}01]. Descent [SSN^{+}23]. descrambler [BKH^{+}93]. Describing [LBFE09]. Description [MVCS16]. descriptor [DK98]. descriptors [RB95]. Design [AMI^{+}07, AdSD16, AKS96, AHX19, ACCF12, AMS22a, AOM04, ACA16, BCL10, BI00, BM22, BLB10, CPS17, CC95, CWH^{+}16, CLV17, CC96, CKZC19, DMT^{+}19, ENT^{+}24, FML09, GYB^{+}04, GV17, GJVZ06, HLS^{+}14b, HCW^{+}16, ILS97, JCJ95, JIN^{+}12, JWH^{+}24, JE18, KNP05, Kim94, KH15, KS01b, KLKP16, LLD96, LLS^{+}23, LZY20, LAL^{+}24, MLB21, NBV17, OPGT16, PCW^{+}16, SK10a, SK11, SPQZ20, SS17, SDM20, SZG^{+}13, SH23, SG94, SBTH19, TWWG19, VKPI17, WWMZ20, WWMZ22, WY95, WXW11, XZC^{+}20, ZZL^{+}24, ZSH^{+}16, ZL16, ZWS^{+}17, ZSZ^{+}17, AIN^{+}15, AM16, APSKPMGM12, BFM^{+}96, BO07b, BJY11, BPK^{+}10, BL94, ÇY07, CLM99, CLD10, CJV16, CDM93, DJ16, ES96, FCA^{+}06, FLC09, FCT03, GMP13, GW94, Geo08, GS98, Gro99, GBL12, HD07, JLM15, KR99, KH07, LA95b, LLY^{+}16, LY94, LYC11, LZXF14, LLE15b, LLE16, LW13, LÜ14, MOZ05]. design [MGR02, Med95, MMR96, NL16, NOF14, OR11, PDE08, PWHL16, RP06, Ros05, RW96, SGSB^{+}15, SL14, SHZ16, SK12b, SPB16, SV98c, SD15b, SSR^{+}11, Tia05, TMP07, TAB^{+}15, VLMN09, WC08, WXR13, WYH10, YFB02, YOY97, ZLLY03]. designer [LO99]. Designing [BCER20, BQ08, IKM08, LP07, SX16, MPL09, MBRM96]. Designs [KR20, ZKEN23, KS13, PPV12, RGG11, TdWC^{+}94, ZQ99]. Desirable [LCL^{+}20]. Destination [FFX^{+}17, FWK17, JWZ^{+}21, AQJRS16, CLS07, LTY06, ZVN99]. destination-controlled [AQJRS16]. destinations [SAKS13]. Destructive [RPP^{+}19, BB96]. detect [KKS^{+}08]. detectable [LHC^{+}16]. DetectDUI} [CXZ^{+}22]. Detecting [AEG^{+}17, DPMK11, FHQ^{+}17, HWJZ21, LLW^{+}09, RHMF16, RKT02a, TLP^{+}16, YRRR12, ZWY^{+}18, KR08, LBP^{+}16, ZZH^{+}10]. Detection [BCER20, CLL^{+}18, CXZ^{+}22, CWZY21, CDH^{+}10, DPG^{+}24, DYJ^{+}23, GA24, GZY23, GLY17, HZL^{+}23, HSKY23, LXW^{+}20, LSSC22, Li24, LCL^{+}20, LYL^{+}22a, MSRG18, MWW^{+}21, NDN^{+}18, OL16, PBGMFM22, RDZ^{+}19, SL16b, TXHL23, VRR24, XLW^{+}17a, XLW^{+}18, YWW^{+}23, ZCZC17, ZCZF20, ZZX^{+}21b, ZCPP22, ZMLR23, ZWS^{+}17, ARK09, ACCF12, BAC12, BBHHR10, CRB09, CH11, DLL^{+}11, FCA^{+}06, Far95, Fel95, FJ93, FAB12, KLZ12, KSV07, LG13a, LZL^{+}14, LS05b, LCQL14, McA94, OÇ10, PS09, RLP06, RCOC03, SG94, TMH11, Tre11, WS93, XY09a, ZY16, ZGTG05]. Determination [FWK17, BSH^{+}11]. Determining [FFX^{+}17, RMDJ16]. Deterministic [Kim98, Le 18, LWP^{+}19, PP17, TC06, WKZL96, BCB99, CZFF98, GK16, KS98, RK06]. detour [LXY^{+}14]. detours [DFGV11]. Deviation [FM20, PPV12, VR13]. deviation-proof [PPV12]. deviations [PS09]. Device [ACC^{+}14, AN20, CCW^{+}17, CN19, GCD23, HW22, KSAK18, KCM16, LPD^{+}18, SYZP19, WXJ^{+}17, ZLZ^{+}23, HQW^{+}16]. Device-Customized [LPD^{+}18]. Device-Free [CCW^{+}17, WXJ^{+}17, HQW^{+}16]. Device-to-Device [AN20, GCD23, KSAK18, KCM16, SYZP19]. Devices [CZZ^{+}21, CW19, CLS^{+}19, DZL^{+}20, JYL^{+}19, JWZ23, LGZ^{+}23, LSL17, MQL^{+}22, SM17, SM19, WXJ^{+}17, WLY^{+}23, XFCW18, XCL^{+}18, YN19, ZCZ^{+}21, ZWY^{+}18, ZLN^{+}17, BBHHR10, SGSB^{+}15, ZS13]. Devolution [HBSX20]. DFAs [FDG^{+}11]. DFL [ZZZ^{+}14]. DGoS [CH21]. DHCP [WWW^{+}20a]. DHT [SENB09]. DHTs [YKGK13]. DHTTP [RW04]. Diagnosability [CH20]. Diagnosis [FLM^{+}22, LC94a, QJCR20, LLL10, ZCB09]. Diamond [CXW^{+}18]. Difference [FBRL18, ZMW^{+}22, CAO11, DLT16, VRK09]. differences [CM05c]. Different [LXL^{+}22b, SLH^{+}19, BK06, SSM06]. Differential [DJB^{+}22, FDG^{+}11, SG13, HMM11, MGG^{+}05]. Differentiated [FT06, SJ12, CZ06, DTM15, DSR02, FK99, JJ08, LLY01, LC04b, LLY09, PILR05, WXBZ04, YR01, ZZZ^{+}07]. differentiated-services [FK99, PILR05].

Differentiation [LPM23, SSNS17, ZMWX18, CCV03, CHM^{+}05, CLS07, CAL09, DSR02, LH14, MLLY06]. Difficulties [FP01]. DiffServ [dOSAU04]. DiffServ-aware [dOSAU04]. diffusing [GLA93]. Diffusion [AC16, BJK20, GCW21, JKJ13, LYLW22, OJSY16, IGE^{+}03, SA05]. Digital [HGZJ21, PSST21, WL99]. digraphs [LZ13]. dilated [AMKY99]. Dilemmas [XZC^{+}20]. Dimension [XCL^{+}18, KBS12]. Dimensional [HMM^{+}20, NK20, TPW^{+}18, TSN^{+}21, YCC^{+}21a, AS07a, AS07b, CLL^{+}14, LS94, LQ13, LPF12, LWT^{+}15, LS05b, WLCC07]. Dimensioning [BK00, GL93, NS03, XWYL23, DBDJ14, KT11, LBRA05, LY94, MV09, MG97b]. Dimming [XLP^{+}23]. DiR [FT06]. Direct [LCP^{+}20, CKV11, DG01, LC97]. Directed [CLL^{+}18, HR14, IGE^{+}03, MYH21, SPLM17, AS01, CYG^{+}14, CER12]. Directional [CLV17, CLS^{+}21, DWL^{+}18, LYD^{+}21, NKNK17, TAH17, ZWZM18, LZF09, SMM11]. directory [Bar95, GRB09]. Disaggregation [CFC^{+}24]. Disaster [HS19, WCL^{+}22, WSX^{+}23, XWX^{+}24]. Disasters [TRVG20, NZCM11]. discard [Rum93]. discarding [Kam96, KqL99]. Discharging [CCKK16]. discipline [FP95, Mil98]. disciplines [FP97, LMS04b, She95]. Disclosure [HHD22, FSH^{+}13]. discontinuity [MMH^{+}15]. Discount [HLZ^{+}14]. discover [SA04]. discovered [SQ12]. discovering [HSFK09]. Discovery [AAR18, CBZ16, CLV17, DMDM17, KSC^{+}23, LZY20, PWL^{+}22, SKE19, WML^{+}18, ZWZM18, Bej09, BGJ^{+}04, CK11, EDBN12, GB10, LL13, MWC16, NSW11, SNXT13, VAGT13]. Discrete [NDS19, TXW^{+}21, HS03, qLH93b, LMS99, XC08]. discrete-time [HS03, LMS99]. DISCS [CLY^{+}17]. Disjoint [KLVL19, YRB^{+}18, GR16, JRY09, TKI^{+}15, XCX^{+}06, XGF^{+}14]. Disk [LWK^{+}16, SZMD17, VTBK21, WLK^{+}17]. Dispatchers [VKO20]. Dispatching [GVM23, HTL^{+}19, OJRCC02]. DisPath [ABK15]. Dispersed [YPA19]. dispersion [CFS11, DRM04, LZ06]. Disrupting [XCL^{+}22]. Disruption [HK14, GLZC12, ZNK^{+}13]. Disruption-Tolerant [HK14, GLZC12, ZNK^{+}13]. Dissatisfaction [FS17]. Dissecting [WXC^{+}24]. disseminating [SB07]. Dissemination [DLZL17, JCR21, KK16b, PJM^{+}19, WLY^{+}23, ZDB^{+}17, ZYY^{+}21, CHLS07, FGM^{+}13, HLX^{+}15, KG10, STQ13, SX10, VGKG10]. Distance [FX17, LJL^{+}16, QL16a, WZZC17, FJJ^{+}01, LWL^{+}11, LH03, LDGL13, ST08]. distance-based [LH03]. Distance-Sensitive [LJL^{+}16]. Distanceless [DLLL16]. distances [LCW05, ST04]. Distillation [LZX^{+}24, XLD^{+}24]. Distinct [RAA^{+}24, LS93b]. distinction [QTWW16]. distinguishing [UZ93]. Distortion [BU21, FHSZ13, CC06, PSK^{+}15]. Distortion-aware [FHSZ13]. distortion-resistant [PSK^{+}15]. Distributed [ATE22, ATE23, ADT22, AAA18, ALPK21, BBG11, BPW23, BV96, BFS21, BGK97, BGK^{+}16, BL04, BZM08, BSS09, CLTM22, CT01, CMP16, CKA16, CGYZ17, CGC^{+}17, CLV17, CGC^{+}18, CLY^{+}17, CJZS14, CL16b, DRMP18, DLL^{+}20, DGNK21, EE18, EOSM10, FMSM24, FZX^{+}23, FX17, FMK^{+}18, Gan20, GGZC19, GCD23, GHW22, GYSPR14, GVM23, GTC^{+}24, GSM16, GMYP16, GCMP20, HZC07, HRCW08, HKLM17, HK24, HWM^{+}24, HJL^{+}20, Hu93, HGZ^{+}23, IKS17, JC13, JWL^{+}18, JTL^{+}17, JTL^{+}18, JLRS16, KK07, KDYV12, KR05, KNE^{+}17, KR20, LMD16, LLY06, LMR07, LHZ^{+}16, LYMA^{+}17, LDW^{+}20, LR09, LCSS17, LCS^{+}18, LPWP22, LDL^{+}22, LY22, MG97a, MCdG23, MCMdlO23, MNZ23, MBN^{+}21, NM09, NSY20, NSY23, NLS19, Nee16a, PD16a, PCW23, QW23, QZZ^{+}13, QLY23, QTE20, RV21, RS97a, RSZ04, RPPA22, RLZ^{+}18, RSR10, EGKM16, SGH^{+}19, SLC22, SC17, SLD^{+}23, SSY19, SE21]. Distributed [SCL^{+}23, SLO^{+}14, SVL^{+}16, TZP^{+}10, TZL23, WSW12, WFC18, WWC^{+}18, WLS^{+}18, WCW19, WLC^{+}20, WQL^{+}21, WRT^{+}21, WLS23, WFZ^{+}23, WN17, WSZL20, WSL^{+}24, XY10a, XSC01, XCC^{+}17, XZC^{+}19, XWH^{+}16, XGQ^{+}19, YSC18, YWK07, YJZW15, YZL^{+}18, YNZ^{+}17, YZY^{+}21, YSY16, ZCH^{+}24, ZLG^{+}17, ZML^{+}19, ZMLL21, ZSL^{+}21, ZQ23, ZKEN23, ZYY^{+}21, AK01, AS08, BRM^{+}13, BM09, BGSSW13, CLC^{+}01, CS14, CHLS07, DC13, DPR06, EAB01, EDM16, FLMM10, GM00, GMS16, GL10, GLS09, GBC^{+}95, HG14, HL05, Jia98, JW10, JW11, JLX^{+}16, KV96, KBS11, Kri14, Kuc14, LNB00, LWKD03, LHB^{+}05, LLS10, LSXS16, LPCVC13, LXC05, MDL07, MOR13, MRM99, MD04, MBC^{+}94, MPL09, MSP^{+}07, MLS12, MV14, OAN15, PDE08, Pil01, QSS^{+}15, RJCE06, RGKS10, RS00, RSB01, SAS16a, ST05, SG13, SKR^{+}09, SNS12, WL08, WTS^{+}13, WWL^{+}15, XY10b, XC08, XLZC14, XME15, YLLY05, YAA09, ZGG05]. distributed [ZKL07, ZT12, ZSCJ14, ZCW15, ZLW16a, ZHLL06, vDP93]. Distributed-Caching [ATE23]. Distributing [SCN^{+}24]. Distribution [ATE21, AGBS23, AHP21, BP19, CZC^{+}22, HHA17, LH07, MFR^{+}20, MJ17, TW23, ACR12, AJF11, BGH^{+}95, ÇY07, FHT^{+}10, FC99, KLC15, LL95, LY94, LMW16, MP08, SLP07, SJ10, SYJ09, TG97, VAS00, WVG12]. Distribution-Oblivious [TW23]. distributions [CT95, DLT05, FCL97, LDHT02, LGD^{+}10]. Distributively [QSW24]. distributors [NWP09]. Disturbance [YZGC23]. Disturbance-Aware [YZGC23]. Diverse [LML10, CS99a, CS99b, hCgKsYwT96, LGGZ10, ZKH10, SYR05]. Diversity [BTP^{+}17, CSL21, PBT^{+}20, QWL21, ZMD^{+}20, ZMMG22, AK14, BNJ16, FGK10, HSH^{+}06, IK09, SKRK12, TW10]. Divide [XCV^{+}20, ZZW^{+}23, CJV16]. Divide-and-Conquer [ZZW^{+}23]. division [CJW11, FT06, SYP01, Tha04]. DMFSGD [LDGL13]. DML [WGL22]. DNN [ZCZ^{+}21]. DNNs [MCdG23]. DNS [FHQ^{+}17, GYJ^{+}16, JSBM02, KSG11, LJZ^{+}23, MRMR17, PJMM22, YRRR12]. DNSSEC [vRDHSP17]. Do [TH21, HLS14a, SSFM08, TMH97]. Does [YASS15]. DoF [CHS^{+}20]. DoF-Based [CHS^{+}20]. Domain [MBL19, YZY^{+}20, ZWCL17, CE09, CBL13, CBL15, Jia06, cLqL97, LJC05, MJ01, RVS^{+}02, YRRR12, YCB07]. domain-based [RVS^{+}02]. domain-flux [YRRR12]. dominant [ES03, WWTK11]. Dominating [Fuk20, LWK^{+}16, SCC^{+}17, SLH^{+}19, WLK^{+}17]. Donation [TH21]. Donation-Based [TH21]. Doppler [DLR^{+}18]. DORE [AMG^{+}17]. DoS-limiting [YWA08]. Double [CZD^{+}22, DRQ^{+}16, GCD23, NS21, SZG09, WHC^{+}19, WYZ^{+}24, CKS16, CSC04, IGHT15, LT94a, PT94]. Double-Auction [NS21, IGHT15]. double-link [CSC04]. double-loop [PT94]. Double-Structure [WYZ^{+}24]. Down [AMS22b]. Downclocking [XZG21]. Downgraded [FLS^{+}22]. Downlink [HLZY23, KW17, LPKF10, LMS05a, LWL17, OES16, SC22a, SdVS22, BSYS12, CK10b, LMS06, OY13, RP13, WKWV16]. downlinks [Nee08]. download [CE08]. Downloading [WN16]. downstream [LT95]. DozyAP [HLS^{+}14b]. DPDK [FBQ^{+}23]. DPI [ABBH^{+}16]. DQDB [CMM95, HL98b, Sha97]. Drafting [SCKB09]. DRAM [WZLX12]. DRAM-based [WZLX12]. Drift [GA24, JE18, LYLW22, LNLM24]. Drift-Based [JE18]. Drift-Diffusion [LYLW22]. Drift-Plus-Penalty [LNLM24]. drifts [KMH12]. Drink [CXZ^{+}22]. Driven [BPL20, DKM^{+}17, GXW^{+}19, JHS^{+}19, LPS19, LJHB18, PJDS18, WCZZ17, XYT^{+}21, XRL^{+}22, ZZX^{+}21a, ZZW^{+}23, ZJL^{+}18, ATB^{+}10, BOY00, BPK^{+}10, CC06, GLAMM11, LGS09, MR09, MGK12, MOY00, PV04, PBKG11, RSS09, RHQZ13, SK12b, VNS02, WZL^{+}13]. Driving [CXZ^{+}22]. Drone [CJ18, MMG22, TZL^{+}24]. Drone-Assisted [TZL^{+}24]. Drones [DNS23, GCMP20, ZCF^{+}24]. Drop [RMPG16, HGG06, TRKN12]. Drop-Tail [RMPG16]. dropping [CSC94, CLA07, KCB03]. drops [CCKK16]. DRS [FDM^{+}17]. DSA [STKL01]. DSASync [KS12]. DTN [ER20, BCL10, CS15, PS15, WBP^{+}11]. DTN-FLOW [CS15]. DTN-meteo [PS15]. DTNs [BLV10, CS15, YSC16]. DTRAB [FTV^{+}10]. DTRACK [CTVD14]. Dual [BS19, EPS21, HNP23, KHYA20, PPTP21, QYZX22, RC08, SRCT23, SCR08, KRKH10, LGW^{+}11, NSS96, SS93, Voi07, PC19, SS94a]. Dual-CISTs [PC19]. Dual-link [RC08, KRKH10]. Dual-resource [SCR08]. Dual-Stacked [BS19]. Dual-Stage [SRCT23]. Dual-Terminal [QYZX22]. duality [Low03]. duct [ZOM03]. duct-layer [ZOM03]. due [Lee96]. Duplex [CDGZ20, CDKZ21, DZ18, LHW19, MZK^{+}17, MMP17, OBS17, WVZ17, YXAZ^{+}18, ZG14]. Duplicate [HSM^{+}21, LHC^{+}16]. Duration [AAA18, SAKMB21, ZGL^{+}19, MS14]. Duration-Based [ZGL^{+}19]. durations [LH07]. During [FGR^{+}17, GDL^{+}22, FB07, Rum93, RS95b, SJL^{+}16, SDV06, THP94]. Duty [CLWZ17, CWL^{+}21, CGC^{+}17, CGC^{+}18, CNG^{+}16, HLL13, BGK^{+}16, CHML15, GTS^{+}09, HLX^{+}15, KWCR10, LHC^{+}16, ODC^{+}16]. Duty-Cycle [CGC^{+}18, KWCR10]. Duty-cycle-aware [HLL13]. Duty-Cycled [CGC^{+}17, HLX^{+}15, LHC^{+}16, ODC^{+}16]. duty-cycling [GTS^{+}09]. DVB [RLZ10]. DVSR [GYB^{+}04]. DX [LPJ^{+}17]. DyMo [BRY^{+}19]. Dynamic [ALR^{+}24, ATEY22, ATE22, ATE23, BJK20, BTD^{+}17, BRY^{+}19, BPL20, BLEM^{+}12, CCG00, CFC^{+}24, CE19, CSS06, CXL^{+}24, CZ06, CWZ^{+}17, CTG00, CH98, CL05, DRQ^{+}16, EL24, FMH^{+}21a, FBM^{+}21, FLTM18, FSM14, FLG^{+}23, GHK^{+}23, GKT93, GLG04, HC02, HTW^{+}19, HS14, HS16, HGM^{+}17, HVT18, HA97, HWM^{+}24, HCW^{+}23, IKS17, JVJ05, JD22, KP23, KKJ06, KLK^{+}20, KL03, KZH^{+}20, LAV16, LMG04, LS99, LGDC19, LC04b, LQZ^{+}24, LWAT13, LSCT17, LLL^{+}16, LYW^{+}21, LSHZ16, LSL^{+}18, LGCG^{+}21, MWW^{+}21, NHLB21, PKK18, RV21, RTLC17, RB02, RKPP16, SMG05a, SKE16, STKL01, SGH^{+}19, SZW^{+}16, TWWG19, TSR14, TWTD17, VKO17, VGP14, WLX^{+}17, WLTJ19, WUZ^{+}19, WLW^{+}20, WD22, WYL23, WZD24, Wil96, XXCC17, XWYL23, YDLT18, YZY^{+}21, YZZ^{+}21, YEMJ24, ZKEN23, ZQL^{+}23, ZLL^{+}23b, ZLM^{+}23, AKA10, AC98, CAQ07, CZ12, CKL16, CDI^{+}04, CJ14, CCLT02]. dynamic [Con11, CDS02, CYL16, DC13, DT93, DRJ^{+}14, EFK07, GM03, GSKR99, HKLS12, HLG94, IS00, JJ08, KD10, KEAAH08, KZDM07, KS12, LT02, LLY06, LYWL08, LKL00, LCL^{+}13b, LPP11, MSWL06, MR98, MG97b, MJ13, MR96, MW06, NST00, NST01, NM06, NXTY10, PWK^{+}13, RMM99, RRG10, RD11b, SMG06, SC09, SLG^{+}16, Sob05, STQ13, SNC^{+}07, SC10, fTL06, WRS^{+}15, WWL02, WXW11, WLZ11, Xin07, YG10, ZKL11, ZHC16, LRJ08]. Dynamical [DME23]. Dynamically [KLC^{+}18, VG04, Med95]. Dynamics [JK05, LCL17, LLX19a, MHB^{+}21, MSTL17, RZS14, VBHT17, VNM22, EML12, HLP11, JD03, JKJ13, JBR16, LS05a, LYS11, Pax99, SJGH10, SLD14, TAJ^{+}10].

early [FJ93, KKM^{+}97, ZGTG05]. Earn [TH21]. Ears [CW19]. Earthquakes [ZLB17]. EASE [GV06]. Easy [CWHW18, LMSS24, ABK15, WBEGS05]. Easy-pass [WBEGS05]. eavesdropping [YSJL14]. eBA [LGHL17]. eBPF [PDV^{+}24]. eBPFlow [PDV^{+}24]. ECC [LYZ^{+}23b]. Echo [JWZ23, TdWC^{+}94]. ECN [KS03, SR18]. Economic [CW12, FS17, LXLC20, MLM15, SC09]. Economics [DKS19, HSS^{+}21, LSSK17, SS06, MCL^{+}10, WL10]. Ecosystem [LZL^{+}20, DD11, MLM15]. EDAL [YCV15]. EDCA [TB10]. EDF [FKT98]. Edge [ATEY22, ATE22, ATE23, AAT^{+}23, AZP^{+}23, AN20, BBR^{+}22, CTG^{+}20, CPS17, CJLF16, CZX18, CZD^{+}22, CGC^{+}24, DJB^{+}22, FMH^{+}21b, GHK^{+}23, GZJ^{+}18, GXS^{+}21, HTL^{+}19, HNP23, JD20, JD22, KSRW22, KSSD24, KKH^{+}22, LLCJ22, LXX^{+}24, LW22, MCdG23, PCW23, QWL21, SCL^{+}23, TGD^{+}20, TJD23, WUZ^{+}19, WZW^{+}20, WJH^{+}21, WYL23, WHYC23, WXX^{+}24, WLD^{+}24, WYL24, XHZ^{+}19, YZL^{+}18, YPA21, YWW^{+}23, YWH21, ZCZ^{+}21, ZCF^{+}24, ZLRC20, ZLG^{+}20, ZWJ^{+}22, ZWR^{+}23, ZTH^{+}23, ZLL^{+}23b, CHM^{+}05, CC06, FCA^{+}06, GR16, MFB99, NKS08, WBEGS05]. Edge-Assisted [WZW^{+}20, ZCF^{+}24]. Edge-Based [AAT^{+}23, XHZ^{+}19, ZWJ^{+}22, FCA^{+}06]. Edge-Cloud [YWH21, ZLRC20, ZWR^{+}23]. Edge-Clouds [HTL^{+}19]. Edge-Core [LLCJ22]. edge-independent [GR16]. edge-redundant [MFB99]. EdgeDuet [YWW^{+}23]. Editorial [Amm02, Amm03, Tow06b, Zeg03b, Zeg03a, Zeg04, Zeg05a, Zeg05b]. Effect [LWR^{+}16, MHH20, VNM22, ZSS^{+}20, CT04b, LZ06, SBP03]. Effective [BW98, CZL^{+}19, EM93, FZ16, KWC93, QLQ^{+}22, BWS10, CR14, DZNT14, GNP^{+}13, GRS00, LPIH11, LBL07, LGW^{+}11, SHZ16, SL08, ZQ99, ZQ00]. effective-bandwidth-based [SL08]. effectiveness [CN08, JSBM02, KYY^{+}12, SKT96]. Effects [KA98, La17, SS16, VC14, BB96, CJ14, ECN09, KV98, KVR98, Kop96, LAJS07, LTZ08, MK10, PL02, Rum93]. efficacy [KGGZ11, YMKC08]. Efficiency [BBZ^{+}18, GXL^{+}21, HCW^{+}23, JSZ14, KHAWC17, LHL^{+}21, LHY^{+}23, PYL^{+}17, SRBBG17, TES19, WLC16, XCZL20, ZLW^{+}20, ZW22, ZBdV23, BTC05, DHSS14, HLX^{+}15, JR14, JP13, JWSLC13, LNS11, LMS04b, MRHWS14, PFC96, PT10, SS94a, SL07a, SL12, SS03, VHvdH01]. Efficient [ACR12, APC21, BCS^{+}19, BBD^{+}18, BFS21, BCN02, Bej04, BSNI06, BCE^{+}19, BEK^{+}22, BPVRSP16, BAB20, BZM^{+}22, BKTN03, BFK^{+}18, BBHH^{+}18, CSLH13, CCLL17, CBV^{+}18, CM16, CM05b, CZZY12, CZM14, CJLF16, CXK^{+}23, CNG^{+}16, CCA96, CLL^{+}14, CG15a, CCG20, DLW^{+}17, DJCA21, ECL^{+}20, EF17, EDBN12, FRC98, FKT98, FC17, FWL08, FM20, GW94, GQ16, GCWC17, GGPS96, GCZ98, GLY17, GP98, GZJ^{+}18, GSH^{+}22, HAGL16, HTW^{+}19, HTW^{+}22, HGM^{+}17, HSM^{+}20, HZL^{+}23, IGHT17, JCR21, JD17, JYC^{+}16, JLZJ19, KW19, KNE^{+}17, KWH11, KLR^{+}20, LWKD03, LCL13a, LMODF18, LWW^{+}19a, LZW^{+}21, LDD21, LCX^{+}16, LGHL17, LCX^{+}19, LXL^{+}19, LXW^{+}19, LL20, LYL^{+}22a, LORS06, LGW^{+}17, LSL^{+}21, LFL^{+}23, MAE19, MCC^{+}19, MCdG23, MPF^{+}15, ME96, MMS01, Nai97, NSS96, NXTY10, NSCR06, PYL^{+}17, PKVI17, PCW23, PMH95, PPTP21, PP02, QFH^{+}18, QLY23, SK03]. Efficient [SL16a, SV96, SKHL12, SYW^{+}22, SPR08b, SPR08a, SV98a, SXEZ21, SGJ17, SBGJ18, VAGT13, VCVC17, WSX^{+}21, WF93a, WL08, WSXL16, WLX^{+}17, WZL^{+}23b, WCM^{+}21, WWL24b, WCAB15, WTJR22, WLW^{+}17, WSL^{+}24, XQG^{+}22, XLWT12, XHZ^{+}19, YCW^{+}19, YBQZ18, YGL^{+}19, YYC^{+}21, ZHGF19, ZCZ^{+}20, ZLRC20, ZLG^{+}20, ZWR^{+}23, ZHCC24, ZRH18, ZMMG22, ZPCS11, ZLWH17, ZKEN23, ZFW^{+}17b, AB09, AS02, BCL10, BO07b, Bej09, BK06, BIS00, BBL95, BSP07, CDO97, CDM13, CRV13, CHCH00, CLM^{+}16, CFC01, CK10b, DT93, DM96, EH11, GTS^{+}09, GS10a, GKT97, GV06, GPM03, GBL12, GZDG06, HLS^{+}14b, HKLM07, Hos98, IKDD15, KVF^{+}12, LLLS07, LSW15, LLW^{+}12, LWCY12, LSZW13, LXY^{+}14, LS97b, LR09, LCZC13, LSXS16, LQCC16, Pad95, PPPW05, QS04, RW04, RSS09, RSSZ13, RKNS10, SLP07, SFAS05, SA01a, SLL15, SLH^{+}06, SYJ09, TKN06, UBPE02, VL97, VG04]. efficient [WMS09, XLR13, XLZC14, YCV15, ZM09, ZBA16, ZL14, ZZHZ13]. Efficiently [CDI^{+}04, TXW^{+}21, KL09]. Effort [JWH^{+}24, JHS^{+}19, LPM23, CF98, KL07, PWK^{+}13, SL08, YD04]. Efforts [TWL^{+}21]. egress [TGRR07]. eICIC [DMMS14, ZJWY17]. Elasecutor [LX21]. Elastic [AAF^{+}16, BCO17, CNM20, CWAO21, CGAC20, FSSC18, LX21, SRCT23, WZZC17, WSL^{+}24, XWW^{+}23, YJL^{+}19, YYB^{+}22, ZYZ16, ZWX^{+}24, ZHT^{+}19, ZLW^{+}17, AS14, AK01, BK00, FT07, JS11, LA02, Low00, NDGL06, YWK07, DKL01]. election [RSZ04]. Electrical [LS22]. Electrical/Optical [LS22]. Electronic [ALYX22, ZHCC24, TZZ^{+}14]. elements [LL95]. Elephant [XCZ^{+}17, YZL^{+}19]. elephants [MGG^{+}05, MK10]. Elevate [CWHW18]. Elicitation [JHS^{+}19]. Eliminate [AAR18]. Eliminating [CLZ^{+}23, JWH^{+}24, SPGM13]. Elimination [FGRQ18, GZY23, TML22, XDZ^{+}23, HKCL13, LCW^{+}15]. Elliptic [vRDHSP17]. Elmo [SSR^{+}20]. Email [HZCB17]. eMBB [AdVS20]. embedded [HW99]. embedded-processor [HW99]. Embedding [AM16, BFK^{+}18, GJWZ16, LGS^{+}23, PHC20, QL16a, RS19, VLM16, YLH17, BO03, CRB12, EDM16, JK15, LZSS10, QM99, ST04, ST08, SZL^{+}14]. Embeddings [RS20]. Embracing [WXJ^{+}17]. Emergency [JR22]. Emerging [LLZ^{+}23a, KR05]. EMI [WWL24b]. EMIT [BCS^{+}19]. Empathy [DDP^{+}19]. Empirical [AB23, WWW^{+}20a, CBAT06, PFTK00, PS09, WK13]. Empirically [Pax94]. Employing [ZBXH13, IZC00, QY12]. Empowered [LW22, PWL^{+}24]. Empowering [XLD^{+}24]. Emulation [CWLH20, HGZJ21, NDS19, SZT01]. en-route [YG10]. Enable [AMG^{+}17, RGY^{+}22, AB07]. Enabled [AMSB^{+}24, CWL^{+}21, DLZL17, HAB^{+}22, HHA17, LXZ^{+}21, QZL^{+}16, SNZ^{+}23, YZP^{+}14, MMG22, RZE^{+}21]. Enables [XWY^{+}18]. Enabling [CBDCP19, CLL^{+}19, CZW^{+}21, DLLL16, GSM^{+}17, HKC^{+}20, KIL24, Kuc14, LZW^{+}21, LW17, SACH21, SLD^{+}23, SYW^{+}22, SGL^{+}22, TZCB23, WJYL16, WPZM16, WHZJ20, WWX^{+}19, XLZ^{+}19, ZZW^{+}23, ZZW^{+}24, AB09, BRM^{+}13, PPPW05, SLC^{+}07]. Enclaves [KHH^{+}18]. Encoded [HS18, KRRR17, HH10b]. encoder [LS03b]. Encoding [BBHH^{+}18, CCLL17, HNW17, HCW^{+}23, CSLH13, FDG^{+}11, LSB06, TNF97]. encodings [RKH^{+}16]. encounter [AWKN16, GV06]. encounter-based [AWKN16]. Encrypted [ADR18, FGRQ18, LYH^{+}23, WHLL23, YWZ^{+}23, FTV^{+}10]. Encryption [HKC^{+}20, LZ23, ASW00]. End [AEG^{+}17, BO00, BM22, BVBV17, CLTM22, CCV03, CZD^{+}22, CZK^{+}21, DCGN03, FZ16, JD03, JT01, KLOS11, KS03, LR03, MHS^{+}17, MLC07, Pax97, Pax99, SK21, SS05, WJ17, CZFF98, CBL06a, DL04, FK99, FF99, HGE04, IAS06, Kam96, KS12, KK06b, LT02, LK02, LE12b, MHL^{+}14, MW00, MK10, MK98, NXTY10, Ord99, RKT02a, SZKT98, SKKA01, TWL06, WVG12, XYLL14, YL98, ZWDS00, ZCB09, ZL16, ZM04]. end-consumers [XYLL14]. end-of-packet [Kam96]. end-point [KK06b, MK10]. End-to-End [AEG^{+}17, BM22, BVBV17, CLTM22, CZD^{+}22, CZK^{+}21, FZ16, MHS^{+}17, SK21, WJ17, BO00, CCV03, DCGN03, JD03, JT01, KLOS11, KS03, LR03, MLC07, Pax97, Pax99, SS05, CZFF98, CBL06a, DL04, FK99, FF99, HGE04, IAS06, KS12, LT02, LK02, LE12b, MHL^{+}14, MW00, MK98, NXTY10, Ord99, RKT02a, SZKT98, SKKA01, TWL06, WVG12, ZWDS00, ZCB09, ZL16, ZM04]. Endangered [DPSA21]. Endpoint [LJHB18, GKPS06]. Endpoint-Driven [LJHB18]. endpoints [TRKN10]. endurable [YW11]. Energy [ACC^{+}14, BAB20, CM16, CZX18, CPR99, DSM^{+}17, DHSS14, EH11, FC17, FFZ^{+}18, FM20, GCWC17, GV17, GYSPR14, GHZ20a, HTW^{+}19, IKDD15, JYC^{+}16, JWW^{+}23, KS19, LS16, LDD21, LHY^{+}23, MAE19, MCC^{+}19, MSRG18, MBN^{+}21, Nee16b, PYL^{+}17, PHC20, RMDJ16, RZE^{+}21, SLP07, SCC^{+}17, SZL^{+}14, SBGJ18, TPC09, TT17, UBPE02, WSZL20, WCW^{+}17, YHH^{+}21, YN19, ZBA16, ZHGF19, ZLW^{+}20, AIN^{+}15, BD07, BTC05, BCL10, CLP12, CFM13, CSSJ14, CMN12, CK09, FMT03, HLL13, HLX^{+}15, HA16, HH10b, HN13, KWCR10, KE16, KD10, KLS11a, KCCM16, LWCY12, LSZW13, LXY^{+}14, LHZ^{+}16, LSS07, LLS10, LFZS11, LCQL14, MCLG07, RPF^{+}14, SGSB^{+}15, SS09, SL12, SHN16, SK13, TSR14, UN11, VGP14, WMS09, XLR13, XSHS12, XSH^{+}15, YCV15, ZM09, ZH08b]. Energy-Aware [Nee16b, PHC20, RMDJ16, RZE^{+}21, SZL^{+}14, LSS07]. Energy-conserving [CPR99]. Energy-Constrained [CZX18, HH10b, KLS11a, MCLG07]. Energy-Efficient [BAB20, FM20, HTW^{+}19, JYC^{+}16, LDD21, MCC^{+}19, ZHGF19, EH11, IKDD15, UBPE02, ZBA16, BCL10, LWCY12, LSZW13, LXY^{+}14, WMS09, XLR13, YCV15, ZM09]. Energy-Harvesting [YN19, HN13, KE16, SK13, TSR14, VGP14]. energy-renewal [XSHS12]. Energy-robustness [TPC09]. energy-time [LCQL14]. Enforcement [ABS^{+}16, BVL^{+}19, LLZ^{+}17, LWW^{+}19b, NHLB21, WSXL16, XXZ^{+}22a, LS97a]. Enforcing [SJZ^{+}24, SBNRS14]. Engine [DLW^{+}17, PES^{+}12, Kai93]. Engineering [CKS17, CLP^{+}17, CGAC20, LRG10, OOM^{+}18, SAC^{+}18, SACH21, TZCB23, TWY^{+}20, TXW^{+}19, WDR^{+}20, YZGC23, ZYS^{+}23, ZHZ^{+}24, ZXZ^{+}19, CN09, DJ12, HL96b, LCM04, MW05, MLC07, SHHA09, SAM10, SGD05, XCR11, XCR15, dOSAU04]. Engineering-Centric [ZHZ^{+}24]. Engines [ABBH^{+}16, BBCD14, BN05]. enhance [BJ15, FGM^{+}13, KVR02]. Enhanced [BLM^{+}17, DMMS14, EE18, FLH^{+}17, GM00, LWP^{+}19, MCC^{+}19, MR96, YCZ^{+}23]. Enhancement [AZP^{+}23, ZLW^{+}20, AWKN16, KT06, ML06]. enhancements [ZRK06]. Enhancing [AB23, ABA^{+}16, CPKL17, CLA07, CYL16, FDG^{+}10, KSSK18, LCLC18, LL18, LZX^{+}24, MCM^{+}23, PD16b, TZL^{+}24, XSM22, XLP^{+}23, YD04, ZMH17, ZLZ^{+}23, ZJL^{+}19, ZXTT08, ZT12]. Enough [BHC^{+}21, HLH^{+}18, XSSK08]. enqueueing [HLG94]. Ensemble [ZLL^{+}23b]. ensure [SNS12]. Ensuring [CMP^{+}14, Smi95, WLY^{+}24, ZLSK15]. Entanglement [ZZL^{+}24, ZWZ^{+}24]. Enterprise [ALPK21, SSK^{+}17, SX16, AYS^{+}13, CFP^{+}09, CG04, SSR^{+}11]. Entities [ZLG^{+}20]. Entries [TXW^{+}21, XTW^{+}22]. Entropy [HCL09, CKKK09, RTK^{+}16]. Entry [RPV13]. Enumerating [VTBK21]. Enumeration [TRVG20, WYH10]. envelope [LK14]. envelopes [FKT98, QK01]. Environment [CL16a, CWZ^{+}17, HZ20, MAE20, XLW^{+}17b, AEJV13, CS99a, LC96, LS97c, RD11a]. environmental [LFZS11]. Environments [JR22, LTN^{+}19, ODJ23, RMDJ16, TSS21, YWW^{+}24, AK00, CK10a, JL12b, LTB04, LPP11, QYZS06, SCY15, STQ13]. Epidemic [CP17, DME23, KG10, SSV13, VGKG10]. Epidemic-based [KG10]. epidemic-style [VGKG10]. Epidemics [HAG19, EKSV16, KK16a]. EPONs [SC10]. Equal [GCZY18, HJL^{+}12, CKS17]. Equal-Cost-MultiPath [CKS17]. equalization [YTL12]. equation [DW11, RX07, VL05]. equation-based [RX07, VL05]. equations [MGG^{+}05]. equilibria [IW08]. Equilibrium [JSW^{+}20, Low00, RKPP16, TWLC07, TWLC10, ALW09, MS08, SRP^{+}11]. equivalence [CDS02]. Equivalent [SYP01, DJM97, YDS06a]. Era [MLJ^{+}22]. Erasure [ACLX17, AAA18, SGVO18, XLAC16, AGGT16, BDS07, DPR06, NSS96, YS15]. Erasure-Coded [ACLX17, AAA18, SGVO18, XLAC16]. ERICA [KJF^{+}00]. Erlang [MM94]. Erratum [SK11]. Error [CWLH20, LL20, PSA96, VNS02, VA09, WMO^{+}23, ZMW^{+}22, ZGYB20, BBM93, BMB^{+}11, BBHHR10, CLM99, CZZY12, Far95, Fel95, GP96b, KEY99, LNB00, LESZ98, McA94, RW93, SG94, SCY08]. error-controlled [BBM93]. Error-Tolerant [ZGYB20]. errors [HJL^{+}12]. ESM [LLW^{+}12]. ESN [MMG22]. essential [CZC^{+}13]. Establishment [ZWZ^{+}24, CGS93, EST93, HMvdLM07, LXC05, MRM99, RS08, TWL05]. estimated [OMA^{+}10]. estimates [LWR15, ZVN99]. Estimating [CZC^{+}22, DSL^{+}18, DLT05, GTS^{+}09, GMWD13, MG16, SNC^{+}07, XZC^{+}19, XCL^{+}19, XCZL20, XCQ^{+}23, ZRLD05, CZZY12, LZ13, ZDR04]. Estimation [BCE^{+}19, BLCT97, CN19, DZL^{+}20, EFFK18, GLLL17, HPP^{+}23, HOZL16, HSM^{+}21, JHJL21, LCZH17, LLL^{+}17, LXL^{+}17a, LL20, LTCS22, MVCS16, OS21, RAA^{+}24, SRCT23, SXQ^{+}23, TW22, WMCW22, WMO^{+}23, XCC^{+}17, XXCC17, XCZ^{+}17, YEMJ24, ZCZ^{+}20, ZLN^{+}17, CDS02, DMS06, DRM04, DJM97, ES03, FJJ^{+}01, GSN^{+}16, GCS06b, HKLM07, JHR05, LDK12, LPIH11, LAN97, LWCY12, LTY06, LFV10, LRL07, LRL08, LWR^{+}16, LFL14, MR98, ODT09, PV04, RVA00, SL15a, ST08, TC06, WF93a, WYL09, WTXT11, ZKH10, ZL14]. Estimator [HSM^{+}20, Val01, VG05, YLCP11]. Ethernet [BSH^{+}11, Bej09, CM16, CSR^{+}20, ECN09, EBJM18, FC17, GB10, JRL15, LTWW94, NM06, PYL^{+}17, QL16b, QGCL11, WTSW97]. Euclidean [LZSS10, ST04]. EV [TZZ^{+}14]. EV-Loc [TZZ^{+}14]. evacuation [GPLT15, Tas99]. evaluate [LMS99]. Evaluating [CLH^{+}24, DM95, SRS01, Zeg95, LNA07]. Evaluation [AB23, AMKY99, CRB09, CM16, GBG^{+}16, KGdV^{+}21, LAL^{+}24, MKOY24, SDM20, AC06, ASSK13, BIV01, BLPS10, BP96, BD96, CK10a, CAK12, CHA95, CBSK07, CZCC14, DM14, EF08, FSH^{+}13, GS97, HLS^{+}14b, JCJ95, LLY^{+}16, LLY01, LC04a, LLS07, LS03b, LNR94, MW98, PP93a, RLKT98, RLZ10, TYJ16, WM96, YFB02, YMKC08, ZR09]. Event [AA05, EPB14, GSW^{+}23, NDS19, QYZX22, WZL^{+}13]. event-driven [WZL^{+}13]. Event-to-sink [AA05]. Event-Triggered [QYZX22]. Events [DDP^{+}19, SDSY19, SLD^{+}22, JBDF07, SJL^{+}16, Ste08]. Every [FBRL18, WLD^{+}16]. Eviction [SSG18, PP02]. evidence [CB97]. Evolution [CQLW22, MLM15, OGLK14, QLSW19, QZC^{+}22, Cha10, CG04, DD11, EKD12, GCM^{+}16, WL10]. Evolutionary [DJB^{+}22, QYC^{+}24, ACP05]. Evolvability [LXLC20]. Evolving [KKS19, LFY^{+}19]. Exact [BS15, ER23, KHYA20, LSDT19, LWF96, LÜ14, Val07, YZZ^{+}21, HXLZ11, VK04]. example [CSEZ93]. Examples [CWZ^{+}23, CSMW02]. excess [DSTM12, DTM15, HGG06]. Exchange [VPC17, YDLM20, FHH10, IBM95, Lie97, OdG97]. exchanges [AJF11]. Exchanging [BCO17]. Exclusion [BPL20, RC08]. Exclusive [SAKMB21]. Execution [HTJ^{+}21, QW23, GZDG06, WF93b]. Executor [LX21]. existence [TWLC07]. Existing [MBI^{+}17, Far95, McA94]. exit [LMSKZ99, MSWL06]. Expandable [LGY16, TYL94]. Expansion [BBHH^{+}18]. Expected [CCLL17, BQ08]. Expedited [SSG18, BBC^{+}02, Jia06]. Expeditus [WXN^{+}17]. Experience [LOFH21, PGMR18, XYT^{+}21, FGL^{+}01, Kar06, TBV^{+}13]. Experience-Driven [XYT^{+}21]. Experiences [HKV^{+}13, BFM^{+}96]. Experimental [AMG^{+}17, BMBK21, CZGKB24, DRMP18, ENW96, GBG^{+}16, LLS07, LLS^{+}23, MKOY24, PP93a, BKH^{+}93, CK10a, CAK12, FSH^{+}13, HJL^{+}12, KS13, LGD^{+}10, TAB^{+}15, TYP^{+}15]. experimentation [BCL10, Mar96]. Experiments [DNS23, SPQZ20, CRB09, DYH13]. Explicit [CF98, HCW^{+}16, KVR02, KAA^{+}18, SDW00, Van17, CRL96, CLK01, CBLVW06, DRR98, GM00, KK05, KR99, LMR99, LAJS07, LP07, SBP03, SL08]. explicit-rate [LMR99]. exploit [HSH^{+}06, SKRK12]. Exploiting [AK14, AZP^{+}23, BJ15, CKS16, CPGZ15, CGYZ16, DSTM12, DTM15, EBJM18, HZCB17, JR22, KWH^{+}17, KNR^{+}16, LJJ^{+}19, MSA^{+}16, NST^{+}16, OKAS23, TXL^{+}12, WHM^{+}13, WHZJ20, WZL^{+}23a, ZLG^{+}17, ZMD^{+}20, ZQW^{+}23, PD07]. Exploits [CQW^{+}18, SBLS19]. Exploration [NG16, NMD^{+}17, SXEZ21, WLW^{+}17, AIN^{+}15, OZPZ09]. Exploring [AG16, DQYG23, LE12a, LLL^{+}22a, LSSC22, SCN^{+}22, SCC^{+}17, VFBD11, WXR13]. Explosion [YXL^{+}18a, PLT14]. Exponential [BBF18, HKV23, LBS05b, TSS14, Van19, CE09, CFM13, KSM05, YS93]. Exponential-RED [LBS05b]. Exponentially [ZHCL17]. exposed [VJV14]. express [MG97a]. Expression [LT16, LN19, MPN^{+}14, XZC^{+}19, BAC12, FDG^{+}11, PLT14]. expressive [KNR^{+}16]. expressiveness [FJB07]. Extend [CH15]. extended [AKS96, HS03, LTWW94, SKT96]. Extending [WSC08]. Extensible [TML^{+}18, BWH^{+}07]. extension [DW11, MBC^{+}94, PFC96]. Extensions [NK20]. External [LHW^{+}20, ML23]. Externalities [LCDW21, ST09]. externalities-based [ST09]. Externality [ZYH^{+}21]. extra [SYP01]. extra-stage [SYP01]. Extracting [DDP^{+}19, DJ14]. Extraction [ABBF19, LDY^{+}16, BDWS12]. Extremely [BHC^{+}21]. eyeball [MCL^{+}11]. Eyeballs [BS19].

Fabric [GWQ^{+}23]. fabrics

[AMI^{+}07, CTH10, WYHL09]. Face [CN16, LLNC09]. Facebook [RHMF16]. FaceChange [CS17]. facility [KNP05, LGD^{+}10, VL97]. Factor [LCZ^{+}23, SC18b, WLK^{+}17, WW16, AdE07]. Factorization [FLBR^{+}19, XLW^{+}17a, LDGL13]. Fading [GV17, HH18, TG23, YYC^{+}21, YYFC24, AK00, AZLB16, ESP05, Hou14, JLRS16, OES16, RGG11, Tan16, ZKH10, ZAS12]. Failure [ABMT23, BHA^{+}20, CZX^{+}17, KLKT16, KLR^{+}20, LLCJ22, LSC^{+}21, LAL^{+}24, OL16, SACH21, WHYC23, XDZ^{+}23, YY20, ZXZ^{+}19, ARK09, ARK11, BTH11, GS98, LYRL07, LJ09, MJ13, MLC07, PF95, RC08, Ste08, TWHR11, THRW12, THBR14, XGF^{+}14]. Failure-Aware [WHYC23]. failure-independent [MJ13]. Failures [BS19, BCLS17, EGR^{+}16, FS17, GDL^{+}22, GDWX23, GDJX24, LGDC18, MHS^{+}17, SCS^{+}22, TRVG20, VTBK21, XGQ^{+}19, YXL18b, ZAW^{+}22, vDJJ^{+}22, AEG^{+}13, BKLS08, BFF07, CSC04, JRY09, JLM15, KRLL11, KRKH10, LML10, LLM11a, MIB^{+}08, NAA^{+}16, NLY^{+}07, WQGW09]. Fair [CLGSS17, CM03, CL15, CGAC20, DM96, ES07, FHMS18, GB18, GLLJ16, HLHL22, IGHT17, KAEAS14, LBS99, ML23, MW00, PL17, PCW23, ST05, AS08, BZ97, BTC01, BI00, BSS^{+}11a, CGEN98, DS04, GYB^{+}04, GGC93, GVC97, HG14, JS11, KV96, LLE15a, LM96, LFZS11, LCZC13, MSA^{+}16, MV14, NDGL06, PLR15, PCL15, RSSZ13, SV96, SV98a, SV98c, SZN00, SSZ03, TKN06, Val07, WCAB15, YXF^{+}13, YLLY05]. Fair-efficient [DM96]. Fairness [BHL07, JSZ14, LWC^{+}14, NML08, SRBBG17, SJZ^{+}24, WTK^{+}17, YWW^{+}24, AVS04, ALW09, AWFT15, BB06, BS97, BS09, CY14, CGGS97, FP14, JZC11, JJL15, JWSLC13, KK93, KH15, LCS12, LMS04a, LPW14, Mar03, MOY00, MV16, PWDL05, RL07, RKT02b, SNS12, Smi95, SS03, WPL06, ZS05]. fairness-efficiency [JWSLC13]. FairTorrent [SNS12]. False [CES22, LXW^{+}20, OÇ10]. family [BGH^{+}95, LYY^{+}22]. FANET [WWL^{+}24a]. farms [RPF^{+}14]. FASA [WZL^{+}13]. Fast [AAT^{+}23, AGBS23, And04, BES22, BN05, BFS21, BPST18, CWL^{+}21, CL17, CLM^{+}18, CSA^{+}21, CCF04, Con11, DBL^{+}19, DLZL17, EGR^{+}16, Fel95, GSM^{+}17, GLM^{+}16, GLC^{+}16, GSN^{+}16, GGK99, GXS^{+}21, HKLM07, HKLS12, HZG^{+}18, HLH^{+}18, JLZJ19, KRKH10, LBRA05, LLWB16, LYY^{+}22, LP24, LK14, LWC^{+}23, LT16, LXL^{+}17b, LCY^{+}19, LCL^{+}20, LL20, MBL10, MPN^{+}14, NLY^{+}07, RPPA22, SL15a, SL16b, SBTH19, TCS13, TRVG20, WMCW22, WHLL23, WQZ^{+}13, WGZC21, XLW^{+}17a, XFCW18, XCW^{+}20a, XCW^{+}20b, XTW^{+}22, XCV^{+}20, YDY^{+}24, YXL18b, YDLT18, YBQZ18, YWH21, ZZH19, ZGYB20, ZXW^{+}21, ZL13b, ZWZC23, AA93, AB07, ABK15, BKLS08, CM93, CSS08, CL08, CG15b, FHH10, FDG^{+}11, GIKK11, GR16, HLZ^{+}14, KLS09a, KHČ^{+}09, LTY06, LXX^{+}14, MPL09, WL08, WY95, WXW11, WJLH06]. Faster [AB21, ZXTT08, PP93b]. FastND [ZWZM18]. Fat [QFH^{+}18, YNDM09]. Fat-Tree [QFH^{+}18, YNDM09]. Fault [Ban99, CWM^{+}17, KSSK18, LWK^{+}18, QJCR20, RDZ^{+}19, SZMD17, WS93, WLK^{+}17, WLC^{+}20, WYZ^{+}24, ZZT^{+}17, AA96, BDHR10, HIM07, HK94, KS95, LCW05, MP94, Pad95, PT94, RCOC03, SS09, SS04b, WKA^{+}13, WMYR16, ZZZ^{+}14]. Fault-Tolerance [WYZ^{+}24, AA96]. Fault-Tolerant [CWM^{+}17, LWK^{+}18, SZMD17, WLK^{+}17, WLC^{+}20, ZZT^{+}17, HIM07, Pad95, SS09, WKA^{+}13, WMYR16]. Faults [LMSS24, WBY^{+}17, BR06, LC94a]. FAVE [LL20]. fBm [JBDF07]. FDDI [RW95, WLS97]. FDoF [LCLC18]. FDQ [KV96]. feasibility [BSS14, BE06, CGMS13, CZZY12, JJ08, KGGZ11, RCGS09, SSWK13, SPGM13]. Feasible [SGVO18, FUDA03, ZLM16]. featherlight [YW11]. Feature [GCW21, SL17, WLS23, XLP^{+}23, FTV^{+}10, LS93a, ZWO^{+}96]. feature-rich [LS93a]. Features [DMDM17]. FEBA [CAL09]. FEC [AJDH01, CGK10, FKCA18, KL07, KEY99, YMKC08]. FEC/ARQ [CGK10]. FECs [CZTX23]. Federated [AMSB^{+}24, DTN^{+}21, DQW^{+}23, GHK^{+}23, GA24, HAB^{+}22, JJJ^{+}23, KSRW22, LXX^{+}24, LZZ^{+}22b, LYZ^{+}23a, LZX^{+}24, NT24, WXX^{+}24, WZD24, YKB^{+}23, ZLZ^{+}23]. Federation [CTG^{+}20, LWLL16]. Federations [DKS19]. feed [BS15, RVB12]. feed-forward [BS15, RVB12]. Feedback [BM93, BCGM07, DAFZ^{+}18, GLA19, GBG^{+}16, HY10, KIL24, OL16, SdVS22, AGGT16, BFMF01, CG15a, HP00, JJL15, KK05, KqL98, LMR99, LGS09, NB99, OY13, QAZ12, QS05, RR93, SSM03, SBP03, XAST12, ZLS96, ZS03]. Feedback-Based [OL16, BCGM07, HY10]. feedback-driven [LGS09]. feedback-synchronization [ZS03]. FeICIC [LCS^{+}18, LCSS17]. Femtocell [LBGL20, RPV13, WKWV16]. Femtocells [KPK^{+}16, AYS^{+}13]. FERN [LAL^{+}24]. Festive [JSZ14]. Few [SACH21]. Fi [BMBK21, BTD^{+}17, CZGKB24, HLS^{+}14b, JYC^{+}16, MGLH18, MSRG18, SPR^{+}20, WCWZ17, XLZ^{+}19, XS21, YCGH17, ZZW^{+}24]. FIB [KNE^{+}17, YXL^{+}18a]. Fiber [BLM^{+}17, CHO^{+}19, Dat17, TWN^{+}20, CR98, CLG^{+}00a, LS97b, NZCM11]. fiber-coax [CLG^{+}00a]. fiber/coax [LS97b]. fibers [SML04]. fidelity [LDK13, XLR13]. Field [BVBV17, LBP^{+}17, NSY20, WD22, WYL23, WYL24, BCL10, HTAZ16, SSV13, SH14]. FIFO [BS15, CCL06, LC03, SG96, VS97]. FIFO-multiplexing [BS15]. Fighting [ZGY^{+}16]. File [DD24, HGZ^{+}23, HSKY23, SLS^{+}23, WN16, CE08, FLC09, LBS11, NAA^{+}16, PLD16, SRS08]. file-sharing [PLD16, SRS08]. files [SKR^{+}09]. Filesystem [ECL^{+}20]. Filling [HHSS16]. Filter [EF17, FLG^{+}23, KLC^{+}18, MCZ^{+}22, QHZC18, RR19b, ZZ17, AAS14, CAO11, RKK14, RK15, WLCC07, WXW11, LFL^{+}23, GBL12]. filterbank [PWK^{+}13]. filtered [LCH95]. Filtering [FLH^{+}17, RFGL17, BL15, CDRV11, KMH12, SAM12, TAB^{+}15, WJS07, YG10]. Filterless [AAF^{+}16, ENT^{+}24, LZ23]. Filters [ALY^{+}20, LYW^{+}18, QCMY16, RAA^{+}24, DKT06, FDG^{+}10, HKLS12, LRC15, Mit02, RSR11]. Find [BZS23]. FINDERS [YW11]. Finding [BZS23, CMW^{+}20, CM05c, DLL^{+}20, Fuk20, LLZ^{+}19, SK12b, TKI^{+}15, WXC16, XSZ^{+}07, YZL^{+}19, GLAM97, XCX^{+}06]. Fine [BKLM06, CCW^{+}17, CS17, FTZ^{+}13, LWT^{+}21, LSL^{+}21, PKK18, WCM^{+}21, XWY^{+}18, KHG^{+}14, KLSV12, FMK^{+}18]. Fine-Grained [CCW^{+}17, CS17, LWT^{+}21, LSL^{+}21, PKK18, WCM^{+}21, XWY^{+}18, BKLM06, FTZ^{+}13, KHG^{+}14, KLSV12]. FineComb [LGKV14]. Finger [LCZ^{+}23]. Fingerprint [LLZ^{+}19, WCH^{+}24]. Fingerprint-Based [WCH^{+}24]. Fingerprinting [MQL^{+}22, SNLL16, SL17, TWL^{+}21]. finishing [HK96]. Finite [AHEK24, SC17, SLJJ16, WWW20b, YN18, AZ06a, CSC94, KS01a, LMS12, LRC15, LC94b, Nai97, SK13, XME15]. finite-buffered [LC94b]. Finite-Markov [SC17]. finite-state [Nai97]. Finite-Time [SLJJ16]. FIPP [MJ13, MJ13]. FireCol [FAB12]. firewall [CBL13]. Fireworks [LLZ^{+}19]. First [QSW24, YHH^{+}21, CAO11, FqL98, GCM^{+}16, KWJY16, LBX11, Mne08]. first-difference [CAO11]. first-order [FqL98]. Fit [YLCP11]. FitLoc [CCW^{+}17]. fitting [SC10]. Five [TGD^{+}20]. FiWi [ALMR14, BLM^{+}17]. Fixed [LB04, NLNL16, RKA08, RM02, URZ^{+}14, KIR08, KAMG07, LSM^{+}14, RrBG94]. Fixed-alternate [RM02]. fixed-budget [LSM^{+}14]. Fixed-Point [NLNL16, KAMG07]. flash [CZCC14, RS05]. FlashLinQ [WTS^{+}13]. FlashTrie [BLC12]. FlEC [MCM^{+}23]. Fleet [KSSD24]. FlexDATE [YZGC23]. flexibility [CSS^{+}14]. Flexible [CBV^{+}18, CZTX23, GDWX23, KW19, LZW^{+}21, MJ17, SE21, Smi02, XXZ^{+}22a, YZGC23, ATB^{+}10, CCL09, DYH13, LC97, LJNK12, SQZ09, SAS^{+}16c]. FlexiS [Li24]. Flight [GCMP20, MMG22, RFGL17, MHRR12]. flights [LKC^{+}13, TG09]. FlipTracer [JHM^{+}19]. Floating [GHK^{+}23, WZLM22].

Flooding [CLWZ17, CWL^{+}21, CNG^{+}16, OKAS23, AC09, CL07, CHLS07, FAB12, WHM^{+}13, ZZHZ13]. Flow [BCC^{+}17, BT23, CLK^{+}24, CCC17, CG97, DHS^{+}23, EPD94, GTC^{+}24, GDL^{+}22, HHL^{+}19, HZL^{+}23, KW17, LYY^{+}22, LL20, LAJ20, LFL^{+}23, NS16, SL16a, SSG18, SM18, WXH^{+}18, XYQ^{+}17, YCZ^{+}23, YLK^{+}17, YXZ17, YXH^{+}21, ZGL^{+}19, ZXC^{+}18, ZHWH21, ACOR99, AADS05, AdE07, BBM93, BFMF01, CM12, CqLL98, CS15, CLK01, CCKK16, Cob02, DLT05, FRC98, FK03, GSK08, GHK02, GS98, HKLM07, HCL09, HLZ^{+}14, HLW13, JJS13b, Kar03, KL13, KLS03, KA95, LDK12, LDK13, qLP97, LCL12a, LM15, LYS11, LL99, MFL^{+}04, MW98, MK98, NM09, Nee09, PG93, PG94a, PFC96, QS04, QS05, SDW00, TAJ^{+}10, TMP07, WPL06, WLL13, WSMJ04, YF05, ZSSK02, ZS03, CS15]. Flow-Aware [ZHWH21]. flow-based [CqLL98]. flow-level [LDK12, LYS11]. flow-switched [FRC98]. Flowlet [FMH^{+}21a]. FlowMate [YF05]. Flows [BBD^{+}18, CMY^{+}18, DGNK21, DWCZ17, HCFC20, HWF^{+}20, XCZ^{+}17, YZL^{+}19, ZGL^{+}19, ZGYB20, ZSL^{+}21, BH05, CAK12, CZFF98, CGEN98, CNP13, DW11, DS04, DGK05, EVF06, FCA^{+}06, GLMM04, Guo04, GMSK09, HZC07, HKB14, KKL03, LNB01, LEYS11, Lia06, NDGL06, NJW16, NCK15, RVV^{+}15, RKT02a, SM14, TL06, WL99]. fluctuation [CH15]. Fluctuations [ZHT^{+}19, LD95]. Fluid [VHT21, BBM93, EMPS06, LDH^{+}12, RCGT06, TGT01]. FluidNet [SAS^{+}16c]. fluids [KWC93]. flux [YRRR12]. Fly [ZBZ^{+}19]. Flying [GCMP20]. FM [CXK^{+}23]. FMTCP [CWW^{+}15]. Fog [HAB^{+}22, JD19, KAK19, LGCG^{+}21, NLB19, WRT^{+}21, ZCH^{+}24, AWM^{+}20]. FoGMatch [AWM^{+}20]. Fogs [SMEH20]. folklore [SMC02]. Forecasting [PCW^{+}16, KZDM07, PS15]. Forensic [NSP^{+}16]. Forensics [CXL18, CZM14]. Forests [HS14, WMFS10]. Forge [BMB19]. Forking [BPA21]. forks [SMH95]. Form [PPTP21]. Formal [CWZ^{+}23, SR02, KLNS93, LM13, LCH^{+}06, WJZ^{+}12]. Formation [ÇTD22]. forms [SG13]. Formulation [CAD^{+}17, BM00, CMN12, CSEZ93, KS01b, MHXT10]. formulations [WYH10]. Forward [AD11, HLH^{+}18, WCK^{+}20, WCM^{+}21, BJ15, BS15, CD96, IK09, RS12, RVB12, SCY08, Tas96]. Forward-Secure [WCM^{+}21]. ForwardDiffsig [BAL10]. forwarder [SHHP00]. Forwarding [BSSU18, CLS^{+}21, CNM^{+}17, DLW^{+}17, GPBL24, JFM^{+}22, KK21, LNG^{+}21, LLY^{+}22, PRH17, SRCDL19, WSX^{+}21, WWC^{+}18, WBY^{+}17, XBM^{+}23, YBQZ18, ZCZC17, ZZX^{+}21b, AAS14, AAV09, BM09, BN05, BBC^{+}02, CLP12, CHML15, CB11, EST93, Jia06, LHC^{+}16, LS10, LCB^{+}10, RTK^{+}16, SMG05b, SAKS13, XCR11, XCR15]. Foundation [CLV17, LRL07, LRL08, SXLL08]. foundations [NR98]. Fountain [BP19, AD11, CWW^{+}15, CWL^{+}21, DLZL17]. Fountain-Coded [BP19]. Fountain-Enabled [CWL^{+}21, DLZL17]. FPGA [XLT^{+}22]. FPGA-Based [XLT^{+}22]. fractal [TG09]. fraction [Lee96]. Fractional [SYZP19]. fragment [LNM^{+}09]. fragmentation [NAA^{+}16]. fragmented [SMC02]. Frame [WG16, CFG08, DK98, SGSB^{+}15]. frames [JMS08, WM16]. Framework [AGM^{+}17, AMG^{+}17, BMB19, BWK^{+}22, CLL^{+}18, CGYZ16, CLX^{+}22, CL19, CDW19, Dai22, FMK^{+}18, FLG^{+}23, HKC^{+}20, JWZ^{+}21, JWH^{+}24, KW19, LYMA^{+}17, LPS19, LYH^{+}23, NLS19, NLB19, NLT^{+}18, PBGMFM22, RPPA22, RAPP22, RSLL23, SAMB18, SM17, SM19, SZW^{+}16, SE21, TMGB19, VKPI17, VRR24, WT17, WFY^{+}18, XWH^{+}16, XYT^{+}21, XHZ^{+}19, YCC^{+}21a, YLS^{+}17, YHCL21, ZJL^{+}18, ZQW^{+}23, ZQL^{+}23, AW04, APB^{+}13, BB06, CLS07, CYG^{+}14, CL13, CAH08, DM96, DJM97, FJL^{+}97, FLMM10, FNQ00, GS10a, GV97, GT99, GLSB08, HA16, HS03, HSFK09, JWSLC13, KS10, KH07, LK02, LZ13, LNA07, LWT^{+}15, LCZC13, LMS04b, LMW16, MMR96, PSK^{+}15, PILR05, RL07, RS08, RHC^{+}12, RRR02, RL94, SPH04, SRS03, SRP^{+}11, SC09, SLG^{+}16, SQZ09, SS07, Tha01, WZR08, YMR00, YJ15, YKKY08, ZLC12, ZWTC16]. Frameworks [LYY^{+}22, ZLW18]. framing [FJL^{+}97, MMC05]. Free [BBD^{+}18, BWG^{+}20, BFK^{+}18, CCW^{+}17, CLW19, CCZZ17, CGL16, FLMS18, KIW^{+}17, NV21, QZX^{+}17, QLSW19, RpLP^{+}17, RS21, SBGJ18, WXJ^{+}17, WXM21, WGZC21, YFM^{+}22, ZZ17, ZWGC17, ZWYD18, ZGZC20, GLAM97, GLA93, GBC^{+}95, HQW^{+}16, ILS97, JSuRKH03, KBS11, LL10, MJ14, PEA09, THBR14, VS97, YOY97]. Freelance [CVV17]. Frequency [BCE^{+}19, DPM^{+}18, KAHKB17, LSHZ16, SXQ^{+}23, KL95, cLqL97, qLP97, wTjCjC97, TYP^{+}15, XL11a]. frequency-based [TYP^{+}15]. frequency-domain [cLqL97]. Fresh [ATEY22]. Freshness [JE18, PBSS23, TKM20b, TTM23]. Friendliness [JLW^{+}24, TKXP20, BBM^{+}10]. Friendly [MRR^{+}14, MBN^{+}21, JGMB03, RW04]. friends [HLS14a]. Friendship [ZQW^{+}23]. FSA [RSR11]. FSA-based [RSR11]. FSR [WJZ^{+}12]. FTrack [XZG20]. Fu [WCQ^{+}20]. Full [ABK15, CDGZ20, CDKZ21, DZ18, LHW19, MZK^{+}17, MMP17, OBS17, WVZ17, YXAZ^{+}18, BRM^{+}13, SRS03, YBX^{+}10, ZG14]. Full-Duplex [CDGZ20, CDKZ21, DZ18, LHW19, MZK^{+}17, MMP17, OBS17, WVZ17, YXAZ^{+}18]. full-length [SRS03]. fully [PYL99, SN15]. FUN [ZSH^{+}16]. Function [EMAL17, FBM^{+}21, KLR^{+}20, LYL21, QW23, WHYC23, YZY^{+}20, ZLZ^{+}21b, CHH06, HH98, KLT15, LZ13, MDL07, OWMM97, UN11]. Functional [ACLX17]. functionality [TEML09]. Functions [CFC^{+}24, CWHW18, FM22, KLLT18, LDS^{+}24, NGRF19, PDV^{+}24, VLM16, WZL^{+}23c, BS08, FqL98, KS03, qLH93b, qLH93a, SGR13]. Fundamental [BHA^{+}20, CVV17, JJL15, JK21, KEW06, LZL^{+}20, LW17, WVZ17, SH12, SD15b, WKZL96, XL05]. fundamentals [WPL06]. Fusion [GND17, LWR15, MVCS16, SCW^{+}21, LWR^{+}16, TXL^{+}12]. Fusion-Based [GND17]. FUSO [CLM^{+}18]. Future [LXLC20, MRJ20, SMD20]. fuzzy [BLCT97, CFPP96, CC96, CCL99, HP00, RrBG94]. fuzzy-logic [HP00].

G [CM16, RW95, AMCD19, AdVS20, DM15, KG05, MCC^{+}19, SCPB19, SKA^{+}18, YBG^{+}12, YJ15]. G-RCA [YBG^{+}12]. G.826 [SS96]. Gain [ATE21, KS19, KA98, TW10, fTL06, YASS15]. Gains [CDKZ21, MZK^{+}17, WVZ17, SJ95, SPGM13]. Game [AWM^{+}20, DZ20, DJB^{+}22, GKCR21, LCK^{+}18, LBP^{+}17, LCSS17, LCS^{+}18, NSY20, RRS^{+}14, WYL23, XZC^{+}20, BGSSW13, CSMW02, CLD10, CL16b, DJ12, DM96, FK13, GS16, GLLJ16, IW08, Kon06, KG05, LWT^{+}15, MLLY06, MW06, NOF14, RSS09, SRP^{+}11, She95, VT12, XC08, YMR00, YXF^{+}13]. game-theoretic [BGSSW13, CL16b, DJ12, Kon06, NOF14, RSS09, She95, VT12, YXF^{+}13]. game-theoretical [LWT^{+}15]. Games [CBDCP19, DKS19, HHSS16, MYH21, PPTP21, ZCdV^{+}18, ZBdV23, AKSS12, ACKZ14, CFS^{+}10, cFCcFW05, GMS16, HTAZ16, Lia06, MRHWS14, SSA11, TLS^{+}12]. Gaming [LLT^{+}16, BLL07]. gamma [FNQ00, SRS03]. gamma-based [FNQ00, SRS03]. Gap [CSL21, WWW^{+}20a, ZCW^{+}22, HFC^{+}13, ZSK12]. Gaps [YN18]. gated [SC10]. gateway [KLNS93, TL06]. gateways [FJ93, GQ16]. Gathering [LSL^{+}21, CBL06b, CBLVW06, FML09, LÜ14, SP94, WMFS10, ZHC16]. Gaussian [ACZP21, LLLT10, SL12, SKUB12]. GB [YN20, HM06]. GB-PANDAS [YN20]. Gb/s [BLC12, HM06, PCB^{+}98]. Gb/s-based [HM06]. GBAR [FNQ00, Hey97]. GEM [GMP13]. GEMNET [IBM95]. Gen2 [LYDA19]. GenePrint [HQY^{+}16]. General [CHS^{+}20, CZD^{+}22, CMY^{+}17, CMY^{+}18, DWCZ17, JWZ^{+}21, LFL^{+}23, RSLL23, SJWH^{+}17, TKM20b, WJYL16, XWH^{+}16, YLY^{+}16, BS08, CT95, EM93, FCL97, FqL98, GS10a, GGH11, GS10b, GBC^{+}95, HS03, HW12, HGW^{+}16, LS06c, PWDL05, SKZ03, SV98b, Tha04, YJZW15, ZBA16, FST^{+}09]. general-purpose [GBC^{+}95]. Generalized [Ali06, BMvU03, GV97, HC07, JYC^{+}16, LWCY12, LM96, LJNK12, MBF^{+}02, PBT^{+}20, SSV13, SM18, WSC^{+}23, XLX^{+}21, AS07a, AS07b, IBM95, JMMT12, JAS10, JC13, Kar10, MMR09, NJW16, PG93, PG94a, SCP99, Ste08, Zeg95]. generate [FUDA03]. generated [CKR^{+}09, YRRR12]. Generating [BMB19, CDO97, ZAS12]. Generation [CZP18, DRW^{+}22, SQS20, AMI^{+}07, ALMR14, DDPP00, DHSS14, KLNS93, MD04, MP93, MP94, Ram96, Ses97, THDD05, UZ93, VV09, VA07, WLC^{+}10, ZKVM14]. Generic [AGCFV18, AGM^{+}17, KBS12, LYY^{+}22, WWL24b, WMT^{+}22, ZLZ^{+}21b, ZDB^{+}17, CGW^{+}12, CK07, HQY^{+}16, MP08, YBG^{+}12, ZZZM03]. genetic [ES96, WC08]. GeneWave [XFCW18]. Genus [WJYL16]. Geo [JWL^{+}18, RLZ^{+}18, WFY^{+}18, JLX^{+}16, WWL^{+}15]. Geo-Distributed [JWL^{+}18, RLZ^{+}18, JLX^{+}16, WWL^{+}15]. Geo-Social [WFY^{+}18]. GeoCAM [LWT^{+}21]. geocasting [LLNC09]. Geodesics [JPM^{+}19]. Geographic [CLQ^{+}19, LQ13, KWZ08, MHRR12, TK12, GMP13]. Geographical [ZDCW18, AEG^{+}13, LLW^{+}15].

Geolocation [LWT^{+}21, GZCF06].

Geometric [LCK^{+}18, MCZ^{+}22, BCGC15, NT00, SBDR08, TYJ16, WLL13]. Geometric-Min [MCZ^{+}22]. geometrically [vDP93]. Geometry [MMP17]. georouting [RKNS10]. Ghost [WWW^{+}18]. GHz [GHK18, NKNK17, SMM11, ZWGC17, ZWZM18]. Giants [DQW^{+}23]. Gigabit [CM16, ALMR14]. gigabit-class [ALMR14]. GIST [FST^{+}09]. Glass [LHW^{+}20]. Global [CQW^{+}18, Cha10, NST^{+}16, PJDS18, QFH^{+}18, RLA06, TZL23, YDS06b, FJJ^{+}01, GR01, GYJ^{+}16, LGC16, MD04, SMS07]. Globally [LLX19a, SLWW19, AB05, BS08]. GLP [WFY^{+}18]. go [VS97, ZLSK15]. goal [RSS09, WC08]. goal-driven [RSS09]. Good [BO16, La17]. Google [GCM^{+}16, XYLL14]. Googling [TRKN10]. GOP [FNQ00]. Gossip [HHL06, LWQ^{+}18, BGPS06, DMC06]. Gossip-based [HHL06]. Gossiping [LLAS19, SLJJ16]. GPF [HLHL22]. GPRS [DM03]. GPS [PDSK04, Val07, YTJQ05]. GPU [HLHL22, XCW^{+}20a, XCW^{+}20b]. GPU-Accelerated [XCW^{+}20b]. GPU-Based [HLHL22]. GPUs [ARS16, VKPI17]. Graceful [CMP^{+}14, RZC11, CVM^{+}15, SDV06]. GRADES [SSN^{+}23]. Gradient [CYTH22, FZX^{+}23, RPPA22, SSN^{+}23, TAH99]. gradients [CJH^{+}11]. grading [CFS09]. Gradual [PIST19]. Gradually [OMA^{+}10]. Grained [CCW^{+}17, CS17, LWT^{+}21, LSL^{+}21, PKK18, WCM^{+}21, XWY^{+}18, BKLM06, FTZ^{+}13, KHG^{+}14, KLSV12]. Grant [CLW19]. Grant-Free [CLW19]. granularities [SSM06]. Granularity [GYSZ19, LHL^{+}23a, LCW^{+}24, QHZC18, AD96]. Graph [BMB19, CL17, DNS23, JCR21, JPM^{+}19, LCP^{+}20, LZB^{+}23, LWK^{+}16, LJJ^{+}19, LAL^{+}24, LCW05, OKAS23, PSST21, SSY19, SCS^{+}22, TMGB19, WLK^{+}17, WYL24, YZHZ21, ZCH^{+}24, ZYL^{+}17, ZWJ^{+}20, BCR^{+}12, GDW^{+}16, GSA15, MSS16, ST08, ZCD97, ZZZM03]. Graph-Based [JCR21, LCP^{+}20, ZYL^{+}17, ZCD97]. Graph-theoretic [LCW05, GSA15]. graphical [LJ09]. Graphics [LLT^{+}16, VLZL16]. graphlet [HFC^{+}13]. Graphs [BMB19, BFK^{+}18, DAFZ^{+}18, Fuk20, MYH21, SZMD17, WWL24b, WW16, ZSZN21, AS01, CER12, JVY06, MFB99, SR94, TLS^{+}12, WGL00, XWG14, ZZW^{+}15]. grating [NPQ06]. gray [CSLH13]. gray-code-based [CSLH13]. greed [She95]. Greedy [CSD22, FBFB17, QL16a, TK12, WJYL16, WW16, BCR^{+}12, JGS^{+}15, JLRS16, LNS11, SKUB12, JLS09]. Green [BBCD14, LZ13]. Greener [ACC^{+}14]. Greening [LLW^{+}15]. Greenput [CLS^{+}18]. Grid [CLQ^{+}19, HHA17, Tod94, FZX^{+}23]. Grid-Based [CLQ^{+}19]. GridFTP [NRB22]. grids [DBDJ14]. grids/clouds [DBDJ14]. Groomed [SS17]. Grooming [AdSD16, BBMELH08, CRD08, GRS00, RS04, SK10a, SK12a, Xin07, ZQ00, ZZZM03, SK11]. Group [CH20, CGYZ16, GCX^{+}17, LX97, QJZ^{+}16, WFY^{+}18, ZXC^{+}18, AGKK03, BOY00, BO03, LNC93, MW98, ODT09, SYR05, SL07b, WGL00, ZLLY03]. Group-Level [WFY^{+}18]. GroupCast [EFA19]. Grouping [LCX^{+}16]. Groups [GBG^{+}16, HWF^{+}20, VTBK21, XCL^{+}19, ACR12, BKTN03, CBD02, LLY06, NB99, WQZ^{+}13]. groupware [BSSS01]. growing [SP94]. growth [DTM15, NS03, PPK15]. Guarantee [LLL^{+}22b, LGHL17, ZY21, ZWX^{+}24, BBC^{+}02, CLK01, HR95, Jia06, KLC15, LC03, WZLX12, WWL02, XL95]. Guaranteed [BWK^{+}22, FM20, KIL24, KLS09a, LZZ^{+}22b, MKOY24, TD03, ZAW^{+}22, Ban99, BKLS08, BDHR10, CLY06, GV97, HSG^{+}08, HTC04, JF04, KKL03, KKLS05, KK00, KL03, LQ13, LRJ08, LV00, LYL07, RKNS10, SS05, Szy16, TYLH09, WYHL09, XE13]. guaranteed-rate [SS05, Szy16].

Guaranteeing [LZW^{+}15, ZCB^{+}17, KCB03, RRR02, SCP99, ZB95]. Guarantees [BM22, CLW19, CKZC19, DZH19, Gan20, IYYI18, MPMC^{+}22, SCN^{+}24, ZBdV23, AL98, CL03, CLC^{+}01, CCLT02, CRV13, CS99b, Cob02, EDM16, cFKSS99, GP98, KBS11, KA03, KKSS12, Kim98, KZ97, KLS03, KS98, LLLS07, LLE15b, Ord99, Smi08, TX08, Tur09, WFS09, XL11b, YL98]. Guard [GPBL24]. Guided [HLZ^{+}21]. guidelines [BPK^{+}10]. Gyroscope [LGZ^{+}23].

h [XXW^{+}23, HDM13, QCS07]. h-Hop [XXW^{+}23]. H-RCA [HDM13]. Half [CDGZ20, LHW19]. Half- [CDGZ20]. Half-Duplex [LHW19]. Halfin [LY22]. Hamming [QHZC18]. handlers [WEK97]. handling [CU95a, NLY^{+}07, VNS02]. handoff [BCN02, LSC99a]. handoffs [AS96, WLL01]. handover [NCT14]. Happy [BS19]. Hard [BZS23, DHHD18, GLS21, LWL17, CAP15, JGKT07, MKS16]. hard-state [JGKT07]. Hard-to-Find [BZS23]. Hardness [RS20, CD96, DXT^{+}12]. Hardware [AN05, FS17, FLH^{+}17, MSTL17, MKOY24, NLB15, PDV^{+}24, PKVI17, RHX^{+}20, DYH13, KR00, KM10, LXX^{+}14]. hardware-aware [DYH13]. Hardware-based [AN05]. Hardware/Software [PDV^{+}24]. HARMLESS [CSR^{+}20]. Harmonize [ZZLM23]. harmonizing [ZS13]. Harnessing [GHZ^{+}20b, LZY^{+}22, RHX^{+}20]. harsh [AK00]. Harvest [SCC^{+}17]. Harvesting [CWH^{+}16, GV17, TT17, WSZL20, YN19, HN13, KE16, LHZ^{+}16, LFZS11, SK13, TSR14, VGP14]. Hash [LYDA19, LCL^{+}20, MVB^{+}21, WBWV16, BLC12, XLZC14, ZGG05]. hash-based [BLC12]. Hash-Routing [WBWV16]. Hashed [VL97]. Hashing [YBQZ18, CKKK09, KM08, KM10, MPL09, WL07]. haul [LWR15, LWR^{+}16]. having [DM03].

HAWAII [RVS^{+}02]. Hazard [RL23]. HDS [GXS^{+}21]. headaches [CCKK16]. Header [CAS^{+}20, FLH^{+}17, KR08, THDD05]. headers [CV96]. healing [FCT03, MK98, SF95, Wu94, XM99]. Health [ZHCC24, JL12a]. Healthcare [HFF^{+}24]. heap [IK07]. Heartbeat [RUH^{+}18]. Heat [BJK20]. Heat-Diffusion [BJK20]. heavily [Świ96]. Heavy [BCER20, BEK^{+}22, HS19, HH18, HZL^{+}23, JE18, LWAL17, MMT14, MMT16, XCQ^{+}23, BMvU03, JMMT12, LLE16, LGD^{+}10, NAA^{+}16, NJW16, WZY^{+}16]. Heavy-Hitter [BCER20]. Heavy-Tailed [LWAL17, MMT14, MMT16, BMvU03, JMMT12, LGD^{+}10, NAA^{+}16, NJW16]. Heavy-Traffic [HH18, JE18, LLE16, WZY^{+}16]. HeavyKeeper [YZL^{+}19]. Heavyweight [MGS^{+}21]. helper [OWKS16]. Hermes [ZLG^{+}20]. Hershel [SNLL16]. heterogeneity [LZXF14]. Heterogeneous [BPW23, BTD^{+}17, CCLL17, CLW19, CDGZ20, CBHS20, CLS^{+}19, DSYN24, DJS^{+}17, FKCA18, FMK^{+}18, HTL^{+}19, HWM^{+}24, HSL20, KLP16, LLCJ22, LLS^{+}23, LFF^{+}19, MYMY17, ML23, PKVI17, SLC22, SJZ^{+}24, VKO20, WHC^{+}19, WXX^{+}24, XLL21, XWX^{+}24, YLH17, YWW^{+}24, ZCZ^{+}21, ZWYD18, ZTH^{+}23, ZJWY17, ZZT^{+}17, BBM93, BGJ^{+}04, CS99b, GGL09b, GGL09a, GHK02, GCZ98, Hou14, KK16a, KT08, LH05, LEYS11, LPW14, LZW^{+}15, MJ01, MDL07, MH02, NML08, PD07, PS15, LZKT99, RS04, RCS14, STL04, Tan16, TWLC07, TWLC10, Tia05, TL06, TWL05, YCV15, YDS06b, ZWTC16, ZZZM03, ZM04, vDP93]. HetNets [LCS^{+}18, BLM^{+}17, DMMS14, HTW^{+}22, KHAWC17, LCSS17, SSNS17]. Heuristic [SBTH19, Yua02, BLS07, CFM13, LÜ14, RL94, ZA95]. heuristics [SB07]. Hidden [MQL^{+}22, QJCR20, BB95, JS12, RCFC15, VJV14, XY09a]. hide [WL16]. hide-and-seek [WL16]. Hierarchical [BZ97, GMD15, GXS^{+}21, KAK19, KTvdSK18, OOM^{+}18, Ros05, SF95, SL07b, YWH21, ZWGC17, CH04, CRD08, CH97, FC99, HA97, LNA07, RPGE04, RSB01, SL15c, SZN00, VL97, VAM^{+}06, WFH12, ZR09]. hierarchies [SMV93]. Hierarchy [CT04b, XL98]. High [ABBF19, AS09, BHC^{+}21, BSM21, BWK^{+}22, BBR^{+}22, BTK^{+}17, CWM^{+}17, CZL^{+}19, CGZL20, DLW^{+}17, EBJM18, GSW^{+}23, GB18, Gro99, HM06, HCFC20, HCW^{+}23, HSM^{+}20, HSM^{+}21, KLE16, LDK13, LXW^{+}17, LOFH21, LSC^{+}21, LCW^{+}24, MCMdlO23, MZZ^{+}23, PLS^{+}21, PJDS18, PG21, RW07, RAA^{+}24, SRBBG17, SD15a, SBLS19, TXHL23, TWH24, TZL^{+}24, VIT21, WJYL16, WLC^{+}20, WMO^{+}23, WNV13, XLW^{+}18, XLZC14, XHC^{+}18, YDY^{+}24, ZP18, ZZX^{+}21a, ZRP^{+}22, ZYS^{+}23, ZLY23, ZZZ^{+}24, ZWZC23, AA93, AACD^{+}96, ACP05, BS97, BK00, BQ08, CS15, CCL99, CS98, CGS93, CGEN98, CR98, CBL06b, CT96, EM93, EVF06, FqL98, GYB^{+}04, GLH95, GGH11, GP96b, GGK99, HKT95, IK07, ILS97, JR14, KV96, KL13, KHW12, LS93a, LM97, cLqL97, LH95, LKC11, LH13, LYS93, LCH95, LLS07, LNM^{+}09, LS06e, LBS05b, LT94b, LXX^{+}14, PWDL05, PLT14, RDO^{+}07, SFAS05, SLC^{+}07]. high [Smi02, SS03, SSZ03, SXLL08, WEK97, WTSW97, WXW15, XLR13, YLCP11, ZTS94]. High-Accuracy [HCFC20, PJDS18]. High-bandwidth [AS09, AA93, LS06e, WXW15]. High-Bitrate [LOFH21]. high-capacity [RDO^{+}07, Smi02]. High-Concurrency [TZL^{+}24]. High-Coverage [GSW^{+}23]. High-Data [LSC^{+}21]. High-Efficiency [HCW^{+}23]. High-fidelity [LDK13, XLR13]. High-Granularity [LCW^{+}24]. High-Order [KLE16]. High-Performance [CWM^{+}17, MCMdlO23, MZZ^{+}23, PLS^{+}21, SBLS19, TXHL23, WLC^{+}20, ZZX^{+}21a, ZZZ^{+}24, ZWZC23, SD15a, WNV13, ACP05, GYB^{+}04, WEK97]. high-reliability [GGH11]. high-resolution [CBL06b]. High-Speed [BHC^{+}21, BSM21, BWK^{+}22, DLW^{+}17, EBJM18, HSM^{+}20, HSM^{+}21, RAA^{+}24, VIT21, WMO^{+}23, HM06, RW07, AACD^{+}96, BK00, CCL99, CS98, CGS93, CGEN98, EVF06, FqL98, GP96b, GGK99, IK07, ILS97, KV96, KL13, LS93a, cLqL97, LH95, LYS93, LCH95, LLS07, LNM^{+}09, LBS05b, LT94b, LXX^{+}14, PLT14, SFAS05, SLC^{+}07, SS03, SSZ03, SXLL08, YLCP11]. High-Throughput [CGZL20, PG21, TWH24, XLZC14, CS15, KHW12]. high-variability [WTSW97]. High-Volume [ABBF19]. Higher [GZY23]. Higher-Order [GZY23]. Highly [NKNK17, WLK^{+}17, ZWH^{+}17, CDI^{+}04, KLS09b, KLOS09, SMM11]. Highly-Directional [NKNK17]. HighwayNoC [EPS21]. Hijacking [FLS^{+}22, SKG^{+}18, ZZH^{+}10]. Hindering [LYKT21]. histogram [SSD93]. histogram-based [SSD93]. histories [GV06]. history [WZL^{+}13]. hit [GMWD13, TR98]. Hitless [ZWZ20]. Hitter [BCER20]. Hitters [BEK^{+}22, XCQ^{+}23]. Hitting [VTBK21]. HLB [YDCF^{+}22]. HMM [HWZ^{+}23]. Hoc [BVBV17, ÇTD22, CDW19, Gan20, GDC^{+}17, MYMY17, PP17, QJZ^{+}16, RZS14, WCC14, AHK08, AS07a, AS07b, BCGC15, BCB99, BNJR12, BNJ16, CE09, CZF^{+}16, CFM13, CDM13, CW10, CMGL11, DLL^{+}11, DBT05, EFK07, GMP08, GGL09b, GGL09a, GGH11, GT02, GMYP16, HL99, HHL06, HS06a, JS11, KK07, KDHK15, KWZ08, LH07, LPKF10, LMP08, LZF09, Li09, LLLT10, LPF12, LLNC09, LSMS06, LR09, LCL^{+}12b, LNL^{+}16, LKZ^{+}04, MQ05, NL07, PS05, RM08, RSR10, RKNS10, SLP07, SPH04, SRR08, SMS07, SSHK11, SS10, SL12, SS07, UN11, WCY04, WTS^{+}13, YD07, YLL10, ZSFZ11, ZW10, vRWZ09]. HOL [CCKK16]. holding [FCL97]. holes [LL10]. Holistic [BWES22, KH07]. Homogeneous [LWK^{+}18, ZWL^{+}16, KG16]. Homomorphic [CJS^{+}20]. Hop [BP19, BVBV17, GJCB18, GZL^{+}17, GVGV17, GEHM02, GLS21, HS16, KS19, KGH^{+}20, LAJ20, LYKT21, OBS17, SK21, SPM^{+}17, TWH24, XXW^{+}23, YXAZ^{+}18, YS07, BB96, BESW08, CF94, CFD06, DV09, GSK08, GS10b, HIM07, HBU95, JMI95, JS09, KN05, KS09b, LHB^{+}05, LRL07, LRL08, LJNK12, MKT96, NL07, NSCR06, PEA09, RA95, SKE16, SS09, Sob02, SV11, TMH97, WJS07, WNV13, XCR11, XCR15, ZL16, CZK^{+}21]. Hop-by-Hop [KGH^{+}20, YS07, CFD06, MKT96, Sob02, XCR11, XCR15]. hop-count [WJS07]. hop-limit [HBU95]. Hopless [LDZ^{+}17]. Hopping [CLW16, CSL21, SL15b, ZYL^{+}14]. hops [GO02]. Horizon [AHEK24, WWW20b]. Hose [YLH17, CL08, CL09b, KLOS11, KLS11b, KRSY02]. hose-model [CL08, CL09b]. hoses [DGG^{+}02]. host [FJJ^{+}01, HFC^{+}13, LZSS10, SC95]. host-based [LZSS10]. host-level [HFC^{+}13]. Hosting [ECL^{+}20]. hosts [GZCF06, SZ08]. hot [TSGR08]. hot-potato [TSGR08]. Hotspots [BMBK21, CJ18]. hour [Med95]. houses [KSG11]. HSM [HYK^{+}23]. HSOP [CHW^{+}20]. HTTP [BL15, BBK12, CL04, HOT97, JSZ14, TL16, ZWH^{+}17]. HTTP-Based [JSZ14]. HTTP-like [CL04]. Hu [WCQ^{+}20]. Hu-Fu [WCQ^{+}20]. hub [CN08, Kim98, LS03a]. Huffman [FDG^{+}10]. Huffman-coded [FDG^{+}10]. Human [TWL^{+}21, ZHZ^{+}18, LHK^{+}12, RSH^{+}11]. Humans [GXWW11]. Hungry [DSM^{+}17]. hurts [AGL16]. Hurwitz [AOM04]. Hybrid [AHP21, BAB20, CDGZ20, CZW^{+}21, CHW^{+}20, CSR^{+}20, ECL^{+}20, FLH^{+}17, GXS^{+}21, GSL^{+}24, HWZ^{+}23, HVT18, HAB^{+}22, HCL^{+}17, HRM22, KPK^{+}16, LZS^{+}22, LS22, LGS^{+}23, LFY^{+}19, PCW23, SYDM09, TTCT19, VVC^{+}17, XLH^{+}17, XHC^{+}18, YY20, ZGHH19, ZDZ^{+}24, BD97, CqLL98, CR98, CGK10, CLG^{+}00a, HA16, KEY99, LPKF10, LBHO07, LGC16, LS97b, LNL^{+}16, LXX^{+}14, Mil98, RWA^{+}08, SPH04, SEK15, SM08, SYR05, TCPV13, ZA11, ZR09, ZRK06]. HyCloud [ECL^{+}20]. HyPaFilter [FLH^{+}17]. Hyper [WCC14, WXW15]. hyper-space [WXW15]. Hyperbolic [BFK^{+}18, ST08, PPK15]. hypercube [VB94]. Hyperfractal [PJM^{+}19]. HyperTester [ZZX^{+}21a]. hypotheses [HDM10]. Hysteresis [AB23].

I-Seismograph [ZLB17]. I/O [qLP97]. IaaS [GLLJ16, ZHW^{+}17]. iBGP [VCD15]. iBUS [ASKR16]. ICE [EFFK18]. iChat [XYLL14]. ICN [ANTR17, LOFH21, SGS20, SS21]. ICN/WDM [ANTR17]. ICTCP [WFGZ13]. ID [CDPLCA16]. Ideal [EPS21, ML22a, DRJ^{+}14]. Identifiability [BHA^{+}20, MHL^{+}14]. Identification [CW19, HDQ^{+}16, HWF^{+}20, HZHZ18, LZB^{+}23, MCZ^{+}22, SL16b, SLSC20, WLD^{+}16, WQY^{+}17, CPR99, HQY^{+}16, KS95, KL13, LL09, LHL15, SL15b, WWTK11, YSL^{+}14, YWZZ16, ZL15, WMP^{+}18]. Identifying [AYL21, CCZZ17, DSL^{+}18, DSM^{+}17, SVG16, SSH^{+}23, WJK^{+}12, WWL24b, GR12, HLS14a, LCL13a]. Idle [WFC18]. IDMaps [FJJ^{+}01]. IDs [CCLL17]. IEEE [BJ15, BB06, BK17, CCG00, CSN06, CAL09, CLL^{+}14, CLG^{+}00a, HJL^{+}12, HKV^{+}13, HDM10, JS12, Kim98, KAMG07, LR22, QCS07, RKA08, SD15b, TS08, TB10, Tow06a, TYP^{+}15, WH11, ZTS11]. IEEE802.11 [NL07]. if [AQK^{+}19]. IGP [NBTD07]. IGPs [VVP^{+}12]. iHorology [VMCB22]. II [DTM15, PG94b]. IIoT [HWZ^{+}23]. ILP [BD96, TMP07, WYH10]. Image [RBS02]. Imbalance [SPLP20]. Imitation [YCC^{+}21a]. immediate [TCPV13]. immortal [XSHS12]. immune [CF94, XGF^{+}14]. Impact [AHK08, CBD02, CMGL11, CMP^{+}14, CDK^{+}17, DQYG23, DBT05, JWSH18, LBS11, LSSC22, MSRG18, QJZ^{+}16, SSNS17, TSGR08, WGL22, YYL23, vRDHSP17, ANSX13, BMS14b, CM12, CJH^{+}11, CDRV11, GS13, KV05, Lab97, LS06d, MGR02, RKT02b, STM^{+}12, SRS01, SNSW12, SS96, XFS06]. impacts [KS13]. Impairment [ZLW^{+}17, CKV11, KT11, RSM09]. Impairment- [ZLW^{+}17]. impairment-aware [CKV11, KT11, RSM09]. Impatient [AGMY21]. Imperfect [XWW^{+}23, KNSV13, LS06d]. Implantable [SDM20]. Implementation [AHX19, LZX^{+}21, ML18, SLD^{+}23, VKPI17, YZZ^{+}21, ZWS^{+}17, ZSZ^{+}17, AP93b, AKS96, ASSK13, BKH^{+}93, BFM^{+}96, BD96, CK10a, Fel95, GYB^{+}04, JIN^{+}12, LLY^{+}16, LO96, LY10, PP93a, PWHL16, RP06, SZG^{+}13, TYL94, WJZ^{+}12, WXW11]. Implementations [HLP^{+}16, SXEZ21, BG98, GP98]. Implementing [TNML93, Kar06, VL97]. implication [SGSB^{+}15, ZH08b]. Implications [ACZP21, FJB07, AW97, HL96b, LDH^{+}12, LMS04b, WDCL15]. Importance [HSO19, PV04, DT93]. Important [SC18b]. Impromptu [CCK16]. Improve [FC17, ODJ23, RZS14, BCL^{+}09, BV05b, DSTM12, TXL^{+}12]. Improved [BT93, CGGS97, CCCC17, DTM^{+}17, EFFK18, GDJX24, HV22, KSSD24, LNS11, Mil95, PCV08, SG18, SS98, XCZL20, BP96, FSM14]. Improvement [JLW^{+}24, SR18, CFM13, HL05, WLCC07]. improvements [VC14]. Improving [ANTR17, CLP12, JSZ14, LL17b, ML18, MRJ20, RPP^{+}19, RHX^{+}20, TKM20a, VVP^{+}13, ZGG05, BPSK97, cFKSS99, SBDR08]. IMS [LHL^{+}23b]. In-Band [JJJ^{+}23, TZPZ23]. In-Car [CXZ^{+}22]. in-flight [MHRR12]. In-Memory [ZLG^{+}17]. In-Network [BWES22, FZX^{+}23, KAT^{+}22, LJHB18, NLS19, PLS^{+}21, PLM^{+}16, RAPP22, VLM16, WBWV16, WBM^{+}18, XLD^{+}24, JS14, SGR13]. In-Order [GLA19, SRCDL19]. In-service [ZF96]. in/out [RKH^{+}16]. inaccurate [GO99, KK03b]. Inapproximability [RS20]. Inbound [NKL^{+}23]. incapable [PGV16]. Incast [ZHWH21, WFGZ13]. Incentive [JSXN18, LFC^{+}22, LSM^{+}14, MLLY06, PWL^{+}24, SYG^{+}22, SJWH^{+}17, YXFT16, SL14, ZLC12, ZLM16]. Incentives [WGvdS17, CKC^{+}13, SYJ09]. Incentivized [KSK17, DRJ^{+}14]. Incentivizing [LBP^{+}17, MGLH18, LMW16]. Incipient [Li24]. Incomplete [FMK^{+}18, LCH23, SFS^{+}22, XTW^{+}22, GB10, KO13]. incompleteness [GIL^{+}15]. Incorporating [SH23]. Incorrect [MKAE20]. increase [TR98]. increased [CJ97, PFC96]. increases [GT02, WLWL13]. Increasing [AP93a, KCA97]. increment [RKK14]. Incremental [CLP^{+}17, CFD06, HPR06, LXZ^{+}21, XLH^{+}17, BN05, RVV^{+}15]. increments [VR13]. independence [KNR^{+}16]. Independent [CER12, SPH04, TML^{+}18, GR12, GR14, GR16, IAS06, JLRS16, MJ13, ZKL07]. Independently [XXW^{+}23]. index [LBFE09]. Indexes [SNC23]. Indexing [GGZC19, WN16, XQG^{+}22]. Indicator [CES22, Kam96]. Indicator-Based [CES22]. Indirect [CKV11, YXCH21]. indirection [SAZ^{+}04]. Individual [LMSR19, XG05, GKJ12, LWLL16]. Indoor [GND17, LJJ^{+}19, WLW^{+}17, ZJL^{+}18, ZSL^{+}17, STKL01]. induced [LD95]. Inducing [YD07]. Industrial [CZL^{+}19, SSY19]. Industry [QZL^{+}16]. inelastic [AS14, HZC07, JS11]. Inertial [XCL^{+}18]. infection [La16]. Inference [BCD19, FAWW22, HWLL21, LW17, MVCS16, QJCR20, SCN^{+}24, SJSB22, YFM^{+}22, ZCZ^{+}21, ZQW^{+}23, BMM^{+}09, GDC^{+}16, LDHT02, NXTY10, WJK^{+}12]. Inferring [LGDC23, LHW^{+}20, MHL^{+}14, MQL^{+}22, TXW^{+}21, ZK19, AdE07, Gao01, KS13, LCB^{+}10, SCKB09]. infinity [ECN09]. inflated [GJVZ06]. Influence [GBMV21, LSDT19, TWTD17, ZND^{+}16, ZZS^{+}16, ZNZT16]. Influences [SYG^{+}22]. influential [HLS14a]. Information [ANTR17, BCC^{+}17, BU21, BSS19, BSSU18, CXL18, CKA16, CWZY21, CMW^{+}20, CL19, FHMS18, GZY23, GHZ^{+}20b, HDF19, Hua17, JCR21, JE18, KSUB^{+}18, KSM19, KK16b, KE21, LCK^{+}18, LLL^{+}22b, LCH23, LJL^{+}16, LCX^{+}19, LXL^{+}19, MAE20, MKAE20, MRD08, OS21, OBS17, PJM^{+}19, RCR^{+}18, SM05, SFS^{+}22, TKM20a, TKM20b, TTM23, TW23, WBWV16, WWW20b, WD22, XHZ^{+}19, XHY^{+}22, YN19, ZK19, ZQW^{+}23, ZY16, ABA^{+}16, BYH^{+}15, CCE^{+}06a, CCE^{+}06b, CLC^{+}01, CSS06, CHLS07, GB10, GLG04, GK16, GO99, HKL06, JJS13b, KK16a, KO13, KBS12, KL13, KKP15, KG10, KL03, KK03b, LS99, Lia06, LJC05, LP07, PMH95, PJ13, SZG09, SP94, SK06, STQ13, SB07, SSA08, SXLL08, SN15, Tow06a, VGKG10, XW11, YW11, YJZW15, YZP^{+}14, ZRLD05, LRJ08]. Information-Agnostic [BCC^{+}17]. Information-Based [LCK^{+}18]. information-bound [ABA^{+}16]. Information-Centric [ANTR17, WBWV16]. information-theoretic [SXLL08, ZRLD05]. information-theoretical [KL13]. Information-theory [MRD08]. Informed [BCMR04, BK06]. Infrastructure [HSS^{+}21, LAJ20, LSL17, MJ14, NK20, NSC^{+}22, DBDJ14, NZCM11, RPZ^{+}09, SD15a, SAZ^{+}04]. Infrastructure-free [MJ14]. infrastructureless [GMS16]. Infrastructures [KLK^{+}20, CW12, LAPS08]. Ingress [WGZC21]. Inhomogeneous [CCMW19, AGLM10]. Input [CWZ^{+}23, HYZH16, AC16, AZ03, Bar95, BMvU03, GKS05, GSD09, JK96, KKLS05, KK03a, LS94, LS06a, LLLS07, LMNM01, qLH93b, qLH93a, LCH95, MBG^{+}02, MBG^{+}03, McK99, MSS02, MS03, Mne08, Nai97, NMH99, OWMM97, PB93, PDT09, TGT01, TT09]. input-output [MSS02]. Input-Queued [HYZH16, AZ03, GKS05, GSD09, KKLS05, KK03a, LLLS07, LMNM01, MBG^{+}02, MBG^{+}03, McK99, MS03]. input/output [LS06a, Nai97, OWMM97]. input/output-queued [LS06a]. Inputs [KE21, HH98, YTJQ05]. Insecurity [LHL^{+}23b]. Insensitive [RPF^{+}14]. insider [CHL16]. Insights [PWMC12, KAMG07]. Inspecting [WBY^{+}17]. Inspection [FGR^{+}17, ZGS^{+}24, ARS16, BAC12, FMMR10]. Inspired [LZY20, MSTL17, SRB^{+}20, FLMM10]. Instabilities [CJL^{+}19, MFL^{+}04, RAL04]. instability [AST11, LMJ98, LMSKZ99, SDV06]. Installation [SSG18]. Instance [EMAL17, ZFLC18]. Instances [JSW^{+}20, ZWL^{+}22, LS14]. instantaneous [GMWD13, GSW99, SCY98]. instantaneous-request [GSW99]. instantly [SV15]. INT [TZPZ23]. Integer [CMY^{+}17]. Integrated [CTG^{+}20, GJWZ16, GHW22, HLSG04, MFT^{+}20, SX16, WC08, YZL^{+}18, AK01, ASKR16, BLT02, GLAMM11, GVC97, JDSZ97, KIR06, MRD08, MLC07, PG93, PG94a, RR93]. Integrating [SRCT23, AP93a, TZZ^{+}14]. Integration [OSW97, OÇ10, SL08, Bej04]. Integrity [LLX^{+}17, YCC21b, CL12, GEHM02]. integrity-preserving [CL12]. Intelligence [AZP^{+}23, HH17]. Intelligent [AWM^{+}20, DLC^{+}18b, PWL^{+}24, SRMB^{+}23, WSX^{+}23, YXH^{+}21, CHL16, CDH^{+}10, NS98]. Intensities [LJJ^{+}19]. Intensity [YS21]. Intensive [ALR^{+}24, FMH^{+}21b, KYM22, PGV16].

Inter

[CWAO21, DPSA21, DMMS14, GSKN18, KLP16, LCK^{+}18, LWL17, LZL^{+}18, LHW19, SFM^{+}18, WZL^{+}23a, YCW^{+}19, ZCB^{+}17, ZYS^{+}23, ZWCL17, CS15, CZ06, LJC05, PLD16, WLL01, YCB07]. Inter-Cell [KLP16, LCK^{+}18, SFM^{+}18, DMMS14]. Inter-Client [LHW19]. Inter-Core [CWAO21]. Inter-Data [WZL^{+}23a, ZCB^{+}17]. Inter-Datacenter [DPSA21, LZL^{+}18, YCW^{+}19]. Inter-DC [ZYS^{+}23]. Inter-Delivery [GSKN18]. Inter-Domain [ZWCL17, LJC05, YCB07]. inter-ISP [PLD16]. inter-landmark [CS15]. Inter-Mode [CWAO21]. Inter-Session [LWL17]. inter-SLA [CZ06]. inter-switch [WLL01]. interacting [GLMM04]. Interaction [BH05, RCS14]. Interactions [LFY^{+}19, MQL^{+}22, SNZ^{+}23, TLP^{+}16, ZWO^{+}96]. Interactive [MWW^{+}21, WLS^{+}18, NABZ12, ZT12]. Interactivity [TES19, ZT12]. interconnected [PMH95]. Interconnecting [LS14]. Interconnection [RGY^{+}22, CHA95, CTH10, LGW^{+}11, ZSK12]. Interconnections [GNK^{+}21, MYC^{+}19, BB96]. Interconnects [YWHW24, HD07]. Interdependent [La16, La17, ZM18]. interdomain [GSW02, LGGZ10, SAM10, TGRR07, WQGW09, WJZ^{+}12, ZZG^{+}16]. Interest [SGS20, GLAMM11]. interest-driven [GLAMM11]. Interesting [LGDC23]. Interface [XYA^{+}21, ZWC^{+}24]. Interface-Based [ZWC^{+}24]. Interfaces [KP21]. Interference [BMY^{+}17, CMP16, CHS^{+}20, DMMS14, DLZL17, HS16, KWH^{+}17, KLK^{+}20, KLP16, LCK^{+}18, LHW19, LBGL20, QCS07, RPP^{+}19, RL23, SFM^{+}18, SPR^{+}20, SMM11, TKM20b, YNZ^{+}17, YYFC24, ZLZ21a, AK00, AYS^{+}13, BCP13, BE08, BB95, BB96, BRS10, BSS14, BS08, DM15, GNP^{+}13, GS10b, JC13, KDHK15, LPCVC13, RK06, RD11b, RSSZ13, SAS16a, SH14, TYP^{+}15, WHM^{+}13, WK13, YASS15, YC12, ZL13a, ZL16, vRWZ09]. interference-affected [BCP13]. Interference-Aware [ZLZ21a]. interference-limited [BE08]. Interference-Managed [KLK^{+}20]. Interference-Resilient [SPR^{+}20]. interferences [DBT05]. Interferers [BVBV17]. Interfering [ODJ23]. interlayer [WCAB15]. Interleaved [Le 18, Kar10]. interleaving [BKH^{+}93]. intermeeting [CE09]. Intermittent [FBQ^{+}23]. Intermittently [JR22, CB11, RYS12, SPR08b, SPR08a]. Intermittently-Connected [JR22]. internal [LDHT02, WYHL09]. Internet [FST^{+}09, ASKL18, AQJRS16, ACZP21, AVS04, ALWD05, AB05, AC09, AW97, AFT11, BCS^{+}19, BBG^{+}10, BS02, CSMW02, CM12, CQLW22, CHW^{+}20, CWSB05, CTVD14, DSA^{+}14, DD11, EDBN12, EPB14, FHT^{+}10, FK99, FF99, FP01, FAF^{+}17, FJJ^{+}01, FWN^{+}22, Gao01, GR01, GXWW11, GIL^{+}15, GZCF06, GS09, GS04, HSH^{+}06, HSFK09, HFKC12, HRLY21, HM04, HCW^{+}23, IGHT17, JWSH18, JT01, KS20, KHLC13, KG99, LA02, LMJ98, LABJ01, LCM04, LSS^{+}13, LMS05b, LL13, LPIH11, LWW^{+}19a, LCP^{+}20, LXLC20, LHC05, LDD21, LSM^{+}14, LSK20, LBP^{+}16, MCL^{+}10, MCL^{+}11, MLM15, Ma16a, MT06, MHH20, MTK03, MHRR12, MYC^{+}19, NR13, NG16, OZPZ09, OPW^{+}10, OGLK14, PLR^{+}19, PJMM22, Pax97, Pax99, PPS^{+}22, QYZS06, QZC^{+}22, QML^{+}24, RBS02, RB02, RZWQ12, SDM20, SA04, SP94, SRP^{+}11, STM^{+}12, SJ10, ST08, SSW10, SKG12, SFFF03, SLO^{+}14, Sob02]. Internet [SVL^{+}16, SLD14, SMLN^{+}03, SAZ^{+}04, SXLL08, Szy16, TG09, TRKN10, TGD^{+}20, TH96, VMCB22, VC12, VC14, VWNT17, WL10, WCCM18, XHN04, XLW^{+}17a, XWW^{+}18, XPW^{+}18, XZB08, XWG14, YFB02, YDS06b, YXZ19, YYC^{+}21, ZCD97, ZNN^{+}10, ZLB17, ZGHH19, ZSK12, ZLSK15, ZGTG05]. Internet-Based [PPS^{+}22]. Internet-like [QYZS06]. Internet-scale [KHLC13]. Internet-style [AB05]. Internet-Telephony [CHW^{+}20]. Internet-Wide [HCW^{+}23, LL13, STM^{+}12]. Internets [EST93]. Internetwork [RT99]. interoperability [CLG00b, HLSG04]. Interparticipant [ZLS96]. Interpath [KLVL19]. Interplanetary [ER20]. interpolation [LDK13]. Interrelation [LYKT21]. Intersection [CFP^{+}21, DMDM17]. intersession [KWS10, MRHWS14]. interval [NM06]. interval-based [NM06]. Intra [GSM16, WG16, WZL^{+}23a, ZWH^{+}17, RGKR10]. Intra- [WZL^{+}23a]. Intra-Body [GSM16]. Intra-Datacenter [ZWH^{+}17]. Intra-Frame [WG16]. intra-session [RGKR10]. intrabody [SMGP15]. intralayer [LE13]. intriguing [LMSKZ99]. Intrinsic [CCMW19, qLP97, RCW15]. Introduction [CCE^{+}06a, CCE^{+}06b]. Intrusion [DPG^{+}24, ZCPP22, KLZ12]. Intrusive [CW19]. intserv [LS03b]. Intuitive [KE21, TWL06]. inverse [RRG10]. inversion [CLW95]. Invertible [TXHL23]. Inverting [HV06]. Investigating [LGD^{+}10]. Investigation [YDY^{+}24]. Investments [MLB21, JAW11]. Invisible [LLL^{+}22a]. Invoking [ABS^{+}16]. IoT [AWM^{+}20, CWZ^{+}17, CLS^{+}19, JYL^{+}19, JWZ23, LGZ^{+}23, MQL^{+}22, PWWP18, QYC^{+}24, WHZJ20, WQL^{+}21, WFZ^{+}23, WXC^{+}24, XSM22, XXZ^{+}22b, YHH^{+}21, YWH21, YYL23, YWZG23, YEMJ24, ZCH^{+}24]. IoT-Fog [AWM^{+}20]. IP [AM16, AN05, AMP01, AEB02, AAM05, AAB05, ABK15, AJ06, BLC12, BR06, BGJ^{+}04, CSG14, CJ14, CqLL98, CRS18, CL09b, CMP^{+}14, EAB02, EGR^{+}16, FGL^{+}01, Goo08, GR16, GS09, HL03, HHD22, HWHW18, JID^{+}07, KMS^{+}01, KP96, KRKH10, KLOS09, KLPS06, KHČ^{+}09, KGGZ11, LM97, LMS00, LSV99, LZ06, LXY^{+}14, LWT^{+}21, LTY06, LXX^{+}14, MIB^{+}08, MGG^{+}05, MPL09, NTR18, NML98, NABZ12, PP93a, PCB^{+}98, RRK07, RW07, RTK^{+}16, RS07, RS21, SK03, SFAS05, SWKA01, SAC^{+}18, SPS^{+}02, SXLL08, TAG08, TSGR08, WLLD05, WBEGS05, WJS07, YBG^{+}12, YXL^{+}18a, YY20, ZZH^{+}10, ZBA16, ZHLL06, ZLTX17]. IP-Based [LWT^{+}21, CL09b]. IP-Over-WDM [ZLTX17]. IP/ATM [CqLL98]. IP/SDN [YY20]. IPACT [SC10]. iPath [GDC^{+}16]. IPC [PP93a, WF93b]. IPID [FLS^{+}22]. IPO [AJ06]. IPs [HHD22]. IPsec [YW07]. IPTV [KLMW11]. IPv4 [KHLC13, PT12, SMP^{+}14]. IPv6 [HRLY21, HCW^{+}23, NG16, PT12, SYW^{+}22, SBLS19, TWY^{+}20]. IrDA [BH06]. IRS [PWL^{+}24]. IS-IS [SGD05]. ISCOD [BK06]. ISDN [LA95a, OKM94]. iSLIP [McK99]. isochronous [HL98b]. isolated [XGF^{+}14]. isolated-failure-immune [XGF^{+}14]. isolation [YWLL09]. ISP [CMN12, CAD^{+}17, CLP^{+}17, DJ16, MCL^{+}10, MRR^{+}14, PLD16, PIST19, STM^{+}12, SMWA04, TGD^{+}20, VNM22]. ISP-Friendly [MRR^{+}14]. ISPs [LYSZ16, SS06]. iSPY [ZZH^{+}10]. issue [CCE^{+}06a, CCE^{+}06b, Tow06a]. issues [AMI^{+}07, CRL96, CW12, GP96a, KGPL13, NBK02]. Item [LP24]. Items [DSL^{+}18, DLL^{+}20, LX24]. iterated [LGC16]. iteration [Mne08]. Iterative [HYZH16, SFS^{+}22, XXBC14, YMO97, Mne08, NM09, PZGLA98, RW95, WJK^{+}12]. ITP [RBS02]. IXPs [GNK^{+}21].

jammers [CHL16]. Jamming [CSSG23, DEP17, DHK16, HLZ^{+}21, LZZ^{+}22a, TNRP11, ZXW^{+}19, ZZLM23, CH11, PBKG11, YSJL14]. Jamming-aware [TNRP11]. Jamming-Resistant [DHK16]. JET [MSWL06]. JITRA [FMSM24]. Jitter [LM01, MPS01, MSB97, BBC^{+}02, FqL98, LS97a, PS98, SA01b]. Jitter-based [LM01]. Job [BPW23, DHHD18, FM20, GHBSWV17, HTL^{+}19, HTJ^{+}21, HKV23, LS22, Van19, WLS23, XLL21, RPF^{+}14]. Job-Assignment [FM20]. Jobs [FLG^{+}20, PG18, PG21, ZTH^{+}23, ZLWH17, MS14]. Join [WFC18, KKEE13, MSWL06, WL07]. join-exit-tree [MSWL06]. Join-the-Idle-Queue [WFC18]. Joint [AdVS20, AHP21, CKL16, CCE^{+}17, CYH^{+}18, CLZ^{+}20, CDKZ21, CG15b, DBDJ14, FXHY21, FFZ^{+}18, GV93, GLA19, GSM16, GCMP20, HDF19, JR14, JLS^{+}17, JD22, KYM22, KT07, LMS06, LSXS16, LH10, PT96, RTNS21, SJ21, TEE16, TJL^{+}19, TMGB19, URZ^{+}14, WDR^{+}20, WLD^{+}24, XLAC16, XCC^{+}17, XYL^{+}17, YN18, YLWH20, YWW^{+}24, ZXW^{+}20a, ZDCW18, ZCZ^{+}20, ZXC^{+}18, ZMMG22, ZWX^{+}24, ZGLC20, CSSJ14, DT15, HSH^{+}06, LR09, LRG10, NM06, PA12, YZBR14, ZS05]. joint-ONU [NM06]. Jointly [GMY13, HHA17, CFM^{+}09]. journey [CH15]. June [Tow06a]. Just [FMSM24]. Just-In-Time [FMSM24].

KAD [SENB09]. Kafe [HLH^{+}18]. Kalman [KMH12]. Keep [LOFH21]. Kelly [MMIY20, XXN^{+}19]. Kernel [ZZZ^{+}24]. Kernels [HLH^{+}18]. Key [ASW00, AGBS23, JWZ23, LXL^{+}17b, LYZ^{+}23b, MFR^{+}20, MZZ^{+}23, WQL^{+}21, WCH^{+}24, WSL^{+}24, XFCW18, YM16, ZLG^{+}17, Zha17, ZLHM22, BGH^{+}95, ÇY07, FHH10, HMvdLM07, LLY06, MSWL06, MP08, SLP07, SIYL09, STL04, TWL05, WGL00, WQZ^{+}13, ZAS12, FLZ^{+}23]. Key-Value [ZLG^{+}17]. Keying [GSH^{+}22]. Keystroke [YFM^{+}22]. KISS [FMMR10]. Knapsack [ZSLZ21]. Knife [NLRS21]. Knowledge [CN16, LZX^{+}24, SLS^{+}23, SCW^{+}21, TWWG19, WZH^{+}18, XLD^{+}24]. KPI [MPMC^{+}22]. Kraken [FSSC18].

KryptoKnight [BGH^{+}95].

L7 [GBL12, LBZ^{+}20]. L7-filter [GBL12]. LAA [GSPV^{+}18, MSRG18]. Label [SSFM08, CO94, COS95]. label-based [COS95]. lack [Sha97]. Lagrangean [SYDM09]. Lagrangian [KHYA20]. Lagrangians [AIL23]. LAN [CS00, CPSWL96, FTZ^{+}13, OY95, OWMM97, RIM98, SZ08, SZT01, WTSW97]. LAN/MAN [RIM98]. landmark [CS15]. Landmarks [LWT^{+}21]. Lanes [GSM^{+}17]. Language [LZS^{+}22, SBM^{+}18, AP93b]. language-based [AP93b]. LANs [AKS^{+}13, BHL07, Bej09, CSN06, CHH06, HSM^{+}13, HKV^{+}13, KS12, LJSB22, QCS07, SA01a, YWK07, ZBXH13]. Large [ADT22, AAG^{+}16, BRY^{+}19, CZW^{+}21, CXL^{+}24, CWZY21, DGC^{+}20, DLLL16, GCD23, GLM^{+}16, GLY17, GLLL17, GBG^{+}16, HV22, HAB^{+}22, HOZL16, HGZ^{+}23, JD17, LXL^{+}17b, LXW^{+}19, MLB21, MHL19, PJMM22, QLY23, RWL^{+}22, RL23, SJL^{+}13, SLD^{+}23, SBTH19, SXLL08, Van19, VKO20, VR13, WWW^{+}20a, XXCC17, XCZL20, XTW^{+}22, XLW^{+}17b, YHH^{+}21, YKKY08, YGL^{+}19, ZFW14, AKA10, AF99, AVPG14, Bej09, BS00, CZF^{+}16, CKR^{+}09, CL03, CL04, CL07, CC95, CCL11, CLM^{+}16, CKR93, DZNT14, DLH^{+}14, GSN^{+}16, Goo08, HMvdLM07, JC13, JYT^{+}15, KS09b, LYWL08, LTB04, LZL12, LCL13a, LS05a, LGD^{+}10, LS10, LCQL14, MWQ^{+}10, MA12, MGG^{+}05, MV14, MG95, MH97, NSW11, NB99, PYL99, PS05, PLS07, PJ13, SW04, SLS10, SQZ09, TK12, WDCL15, XY09a, XW11, XK06b, YBG^{+}12, ZSFZ11, ZW14, ZL13b, ZL14, ZKO93]. Large-Scale [AAG^{+}16, BRY^{+}19, CZW^{+}21, CXL^{+}24, CWZY21, DGC^{+}20, GCD23, GLM^{+}16, GLY17, GLLL17, HAB^{+}22, HOZL16, HGZ^{+}23, LXL^{+}17b, LXW^{+}19, PJMM22, QLY23, RWL^{+}22, SLD^{+}23, Van19, VKO20, WWW^{+}20a, XXCC17, YHH^{+}21, YGL^{+}19, ZFW14, SJL^{+}13, SXLL08, YKKY08, AKA10, AF99, BS00, CZF^{+}16, CKR^{+}09, CL03, CC95, CCL11, CLM^{+}16, DZNT14, DLH^{+}14, GSN^{+}16, Goo08, HMvdLM07, JC13, JYT^{+}15, LYWL08, LTB04, LZL12, LGD^{+}10, LCQL14, MA12, PYL99, PS05, PLS07, PJ13, SQZ09, TK12, WDCL15, XY09a, XW11, ZSFZ11, ZW14, ZL13b, ZL14, ZKO93]. Largest [TXW^{+}21, KWJY16]. largest-deficit-first [KWJY16]. LaScaDa [CHFH20]. laser [ZA11]. Last [DPSA21, PPV17]. Last-Mile [PPV17]. Last-Minute [DPSA21]. LASTor [AYM14]. Latencies [FBRL18, RS97b]. Latency [ACLX17, BS19, BLM^{+}17, CLTM22, CZGKB24, CGC^{+}18, CZL^{+}19, CKZC19, CZK^{+}21, CDL^{+}19, DZL^{+}20, FKCA18, FBFB17, GKB^{+}16, HTL^{+}19, HGB^{+}19, ISS22, LLZ^{+}23a, LPJ^{+}17, LPWP22, SL16a, SV98b, TKXP20, TMMS01, TJD23, WFC18, WHYC23, XLAC16, XYL^{+}17, YTL12, ZRD^{+}23, ZLN^{+}17, AYM14, CKS16, CM03, CB11, CJH^{+}11, CMFA14, GMP08, IM08, KWS^{+}11, KLSV12, LDK12, LDK13, LGKV14, LMS04a, MMC05, OdG96, QSS^{+}15, RSR10, SRR08, SL95, SS93, SKV03, Szy16, ZGG05]. Latency-Based [LPJ^{+}17]. Latency-Bounded [CZK^{+}21]. latency-constrained [CKS16]. Latency-Optimal [FBFB17]. Latency-rate [SV98b]. Latency-Sensitive [FKCA18, HTL^{+}19]. Latent [DMDM17, SDSY19]. lateral [SCY15]. LATS [NL99]. Lattice [BBLV06b, BBLV06a]. Law [TSS14, CE09, MOR13]. laws [AK09, SBNRS14, SFFF03, YGC10]. Layer [ALY^{+}20, BAB20, BDR22, CWLH20, DJCA21, GGZC19, GTU19, HOZL16, HZHZ18, KSNR20, LFC18, LZC^{+}24, LTN^{+}19, MLS^{+}23, XZL20, XTHL21, YZZ^{+}21, YYL23, AK00, AKS96, AZLB16, AC09, AAV09, BL15, BLS07, CK10a, CRB09, CDFG06, CR99, CHL16, CH11, CCF04, CGK10, EOSM10, HQY^{+}16, HK11, JZC11, KT06, LSLL14, LML11, LWL^{+}11, LS06d, LJ09, PDE08, PNRMC13, QL16b, RGG11, RSU^{+}09, SLP07, SAS16a, SHHA09, SH07, SPB16, SS07, VA09, WLLD05, WVG12, XY09b, XE13, ZOM03, ZAFB00, ZL15]. layer-2 [QL16b]. layer-2.5 [AAV09]. Layer-Based [MLS^{+}23]. Layered [YJH05, BKLM06, KK12, LLM11a, WCAB15]. layering [CW16, RKT02b]. layers [AP93a, PDE08]. layout [DJ14, GCZ96]. Lazy [AIL23, CHLS07, LCL16, CHML15]. LB [CKLS22]. LBDP [LZL^{+}14]. LBP [EAH^{+}18]. LBS [JZW^{+}18]. LBSs [JWZ^{+}21]. LC [GJWZ16]. LC-VNE [GJWZ16]. LDAP [WSKV08]. LDPC [TY18]. Leakage [MRMR17, GK16]. Lean [WZL^{+}23c]. Learn [BRK^{+}22]. Learners [LLS^{+}23]. Learning [AY20, AIL23, AMSB^{+}24, BLC21, BPW23, BBR^{+}22, BBZ^{+}18, BRK^{+}22, CE19, CCZL23, CXK^{+}23, CN19, DAFZ^{+}18, DHK16, DM15, DTN^{+}21, DQW^{+}23, EL24, FAWW22, FMK^{+}18, GHK^{+}23, GA24, HLZ^{+}21, HTW^{+}22, HLZY23, HFF^{+}24, HNP23, HAB^{+}22, HWM^{+}24, HCL18, HGZ^{+}23, JHL22, JJJ^{+}23, KTvdSK18, KJG18, KAA^{+}18, LYLW22, LXX^{+}24, LYH^{+}23, LL17b, LZZ^{+}22b, LYZ^{+}23a, LZX^{+}24, MCMdlO23, MNZ23, MGS^{+}21, MMG22, NT24, PWL^{+}24, PM96, PCW23, RRS23, RPPA22, SNZ^{+}23, SRCT23, SCN^{+}24, SDSY19, SCPB19, SZWW22, SKA^{+}18, SdVS22, SGL^{+}22, TSN^{+}21, TWY^{+}20, TJD23, WL16, WZW^{+}20, WLC^{+}20, WRT^{+}21, WLS23, WHYC23, WXX^{+}24, WZD24, WLY^{+}24, WLD^{+}24, WCZZ17, WHC^{+}22, XYA^{+}21, XOYL20, XXZ^{+}23, XWYL23, XYT^{+}21, XRL^{+}22, YCC^{+}21a, YZGC23, YN19, YKB^{+}23, ZCPP22, ZGZ22, ZWL^{+}22, ZLZ^{+}23, ZQW^{+}23, ZDZ^{+}24, ZQL^{+}23, ZLL^{+}23b, GV06, HZL16, JKJ13, MSBZ10, NABZ12, XC08, YDS10, FM22].

Learning-Aided [HCL18, YN19]. Learning-Based [AMSB^{+}24, BPW23, FAWW22, HNP23, SCPB19, WCZZ17, ZCPP22, ZQL^{+}23, DM15]. Learning-Driven [XRL^{+}22]. Learning-Guided [HLZ^{+}21]. Learning-NUM [FM22]. lease [AAS10]. Leases [SAKMB21]. Leasing [SAMB18]. Least [ZND^{+}16, ZNZT16, DFGV11, LHK^{+}12]. least-action [LHK^{+}12]. least-cost [DFGV11]. Ledger [LCP^{+}20]. Ledgers [FMSM24]. LEDs [WG16]. Left [VKO17]. Legacy [CSR^{+}20, GSRS^{+}15]. legacy-compatible [GSRS^{+}15]. legitimate [HFKC12]. Length [GR20a, HKV23, CT95, CH98, ES07, HC02, JMMT12, JMI95, Le 02, MP93, NTS12, SRS03, UZ93, WLC^{+}10]. length-based [WLC^{+}10]. Lengths [AMS22a, YN18]. LEO [EAB01, EAB02, LLZ^{+}23a, TKN06, WCH95]. less [BQ08]. less-structured [BQ08]. lessons [KKM^{+}97]. Level [CWHW18, DZL^{+}18, FGRQ18, HGZJ21, HS18, NTR18, VMCB22, WFY^{+}18, ZCZ^{+}20, AL98, AdE07, BCL12, BSF16, Bor05, CLM99, FJL^{+}97, GIL^{+}15, HFC^{+}13, KL95, LDK12, LYC11, LYS11, LMS04b, LCB^{+}10, MR96, OPW^{+}10, RPGE04, RD11a, SYR05, SFFF03, Tas96, TZP^{+}10, TNML93, WLC^{+}10, WTSW97, WLLZ16, YC12]. Leverage [DZL^{+}20]. Leveraging [KD10, LAL^{+}24, OBS17, PDV^{+}24, PBGMFM22, RS19, SAS16a, TES19, YPA21]. Levy [RSH^{+}11, LKC^{+}13, TG09]. Levy-walk [RSH^{+}11]. Lexicographically [GGFS02]. LGCC [TWH24]. License [SAKMB21]. Licensing [SA21]. Licklider [WBP^{+}11]. life [VFBD11]. Lifetime [CAD^{+}17, DLY^{+}22, KBS11, PBV17, ZWL^{+}16, ZG08, CT04a, HSS08, HY08, IKDD15, KLSS10, LYRL07, LJW^{+}07, LH10, TX08, WSC08, WMFS10, YCV15, ZCJ^{+}13]. Lifetime-Aware [CAD^{+}17]. lifetime-balancing [YCV15]. lifetime-based [LYRL07]. lifetimes [FM06, WYL09, YCL15]. LIFO [HMNK13]. LIFO-backpressure [HMNK13]. Light [BWG^{+}20, GBG^{+}16, PPV04, YYT23, ZHCL17, BGH^{+}95, BMvU03, FJL^{+}97, KIR06, LJC05, NJW16, SSM06, WBEGS05]. light-path [LJC05]. Light-Tailed [ZHCL17, BMvU03, NJW16]. light-trees [SSM06]. Light-Weight [GBG^{+}16, PPV04, BGH^{+}95, FJL^{+}97, WBEGS05]. Lightpath [BLRC05, CHO^{+}19, LLM14, LXC05, XGF^{+}14]. Lightwave [SR94, BSSLB95, GW94, IBM95, JMI95, Lab97, PS93, TMH97]. Lightweight [CCF17, CMP^{+}14, QLQ^{+}22, YXY^{+}18, ZGS^{+}24, CS14, LTY06]. like [CBD02, CL04, FLC09, HL15, LDH^{+}12, PWMC12, QYZS06, SWL06]. Likelihood [BB16]. Limit [CQW^{+}18, CW23, CCG00, CS98, DM95, HBU95, XW11]. Limitations [RX07, SSNS17, SCH23, ZAS12]. Limited [LZY^{+}22, LL17a, TRVG20, AGL16, BE08, CSS06, HZL16, NPQ06, NPY07, OY13, QY04, RS98, RZVZ06, TS09]. limited-range [NPY07]. limiting [CK09, YWA08]. Limits [CVV17, JK21, NJM^{+}19, RDZ^{+}19, BBLV06a, BBLV06b, BBL95, GGM11, HL03, JJL15, KEW06, LLW^{+}14, SK13, WKZL96]. Line [CCK16, LJSB22, VBHT17, XLW^{+}18, BSH^{+}11, BCN02, cFCcFW05, FCT03, MK98, PZS^{+}16, QM99, SMG06, VWT^{+}14, VLMN09, YKKY08, YF05, ZY07b]. Line-of-Sight [LJSB22]. Linear [CMY^{+}17, CMY^{+}18, Dat17, DME23, EBJM18, LL17a, NCM18, PP17, PWLC23, SKE19, WMT^{+}22, YNZ^{+}17, YLY^{+}16, Ada98, BSSLB95, BM00, CCL09, FKT98, GKJ12, KS01b, LLS09, OWKS16, PS93, SLH^{+}06, VJV14, XK06a]. linear-memory [LLS09]. linearity [qLP97]. linearly [GR12, GR14]. linecards [IKM08]. Lines [Dat17, TWN^{+}20, CCL09]. LineSwitch [ACDP17]. Link [CZGKB24, CMP^{+}14, DGW^{+}17, EGR^{+}16, FJ95, GJWZ16, KLLT18, LCH95, LGDC18, LGDC19, Lin93, LCZH17, RpLP^{+}17, RHX^{+}20, SCS^{+}22, TRVG20, VTBK21, XCR11, XCR15, YRB^{+}18, YXL18b, YY20, YYC^{+}21, YYFC24, ZMD^{+}20, vDJJ^{+}22, ARK09, AT03, BTH11, BCP00, BR06, BKLS08, BRS10, BFF07, BSS09, CLM99, CJH^{+}11, CSC04, CJZS14, CRB12, CL09b, DT15, DV09, FB07, GDW^{+}16, GR12, JK15, JHR05, KRLL11, KS09a, KRKH10, Kum98, LLM11a, LWL^{+}11, MHL^{+}14, NLY^{+}07, NBTD07, PDSK04, QZZ^{+}13, RCGS09, RC08, RW93, RS07, SRS01, SYR05, SKUB12, Ste08, SZN00, Tas96, UBPE02, VVP^{+}12, WYL09, WCH95, WK13, XL98, YCL15, ZWYY10, ZZHZ13, WMP^{+}18]. Link-Disjoint [YRB^{+}18]. link-level [Tas96]. Link-Reversal [RpLP^{+}17]. Link-sharing [FJ95, SZN00, XL98]. Link-State [CMP^{+}14, XCR11, XCR15, FB07, VVP^{+}12]. link-weighted [LWL^{+}11]. Links [CM16, DZ18, FC17, XCL^{+}22, Zha17, ZZLM23, AAM05, BPSK97, EVF06, GMLP10, HSFK09, Hou15, ML06, Ram96, RLZ10, SNXT13, VC12, WWTK11, ZL13a, ZW14]. Lip [LYC^{+}19]. LIRU [ZWCL17]. List [CG21]. List-Based [CG21]. lists [DLT16]. LiteFlow [ZZZ^{+}24]. LiteNap [XZG21]. little [PES^{+}12]. Live [CJW11, CBZ16, CZM^{+}24, ER23, MRR^{+}14, RWL^{+}22, SQ16, TH21, WZH^{+}24, CZCC14, SLL15, VAM^{+}06, WXR13, WLCW16, WRS^{+}15, WLR10, WLZ11]. Lived [RUH^{+}18, CDFG06, GLMM04]. livelocks [KGL03]. LiveRender [LLT^{+}16]. LLR [VHNPM96]. LMMC [YJH05]. LMS [AC16, PPV04]. Load [ATE23, BWK^{+}22, BPST18, CWGT14, CCZL23, DPT^{+}18, GCZY18, GXL^{+}21, HV22, HKV23, JCR21, KAT^{+}22, KPK^{+}16, LK16b, LJL^{+}16, LYS^{+}18, LY22, LHL^{+}23a, PJDS18, SRK22, SPLP20, SG17a, SMG05b, SRCDL19, VKO20, WL07, WXN^{+}17, WLL^{+}16b, YDCF^{+}22, YN20, ZDCW18, AWFT15, BHL07, CLY06, HA16, HY10, JMS08, JIN^{+}12, KL08, KDYV12, LLW^{+}15, MOR13, MSS16, NL99, Smi08, Wil96, YCL09, ZTS11]. load-adaptive [NL99]. Load-Aware [YDCF^{+}22]. Load-Balanced [LJL^{+}16, HY10, JMS08, YCL09]. Load-Balancer [BWK^{+}22]. Load-Balancing [CWGT14, SRCDL19, WL07]. Load-Optimal [BPST18]. Load-Splitting [ATE23]. loaded [Świ96]. Loads [CBdV^{+}17, LVB96]. LOC [ZJL^{+}18, CDPLCA16, TZZ^{+}14]. Loc/ID [CDPLCA16]. Local [BES22, BPST18, CWZY21, GHW22, HA96, LKS^{+}16, LESZ98, MOY00, QGCL11, WW16, YZY^{+}18, AZ06b, BM97, BCR^{+}12, BCC07, ES96, GT00, JCJ95, JMI95, KO13, Kum98, LGC16, NLY^{+}07, PJ13, SAS16a, SKR^{+}09, SSA08, THRW12]. local-area [ES96]. Local-Neighborhood [CWZY21]. Locality [BSSU18, QHZC18, XPL^{+}17, CG04, DLT^{+}15, WZY^{+}16]. Locality-Aware [XPL^{+}17, DLT^{+}15]. Locality-Sensitive [QHZC18]. Localizability [PWLC23, YS21, ZMW^{+}22, YLL10]. Localization [ABMT23, BB16, BZM^{+}22, CCW^{+}17, CXK^{+}23, GND17, HMM^{+}20, KLKT16, LL18, PWLC23, RDZ^{+}19, SYL^{+}17, SWL^{+}18, TWL^{+}21, WXJ^{+}17, XCS^{+}18, XWY^{+}18, ZZX^{+}21b, ZXH^{+}13, ZJL^{+}18, ARK11, BTH11, CZC^{+}13, GGM11, KO13, LL10, STM^{+}12, SDW14, SCY15, SS04b, TWHR11, THRW12, TZZ^{+}14, WLL^{+}11, WS05, XXBC14, ZZZ^{+}14]. Localized [LH05, XWW^{+}19, ZYL^{+}17, LZL^{+}14, NZTD02]. Localizing [AEG^{+}17, MHS^{+}17, ZZL^{+}21]. Locally [FSGH17, KLS09b, BMS14a, SAS^{+}16b]. Locating [GV06, SCS^{+}22]. Location [GJWZ16, GCX^{+}17, GXS^{+}21, JZW^{+}18, WPZM16, WFY^{+}18, ACR12, AHL96, BSNI06, CH15, GS16, HL98a, HA97, KBS12, KRS00, LSZW13, Lin97, MRD08, PS05, RLP06, SIYL09, VG04]. location-aware [LSZW13]. location-based [ACR12, CH15, PS05, SIYL09]. Location-Constrained [GJWZ16]. locking [JR96]. Log [ACZP21, SBD11, SKR^{+}09]. Log-Normal [ACZP21]. Logarithmic [NMC07, Val07]. Logic [ABS^{+}16, HP00]. Logical [CN16, WGL22, ZLTX17, BY06, KS01b, LQCC16]. Logs [SDSY19]. Long [CDFG06, DSYN24, HCL^{+}17, RUH^{+}18, SENB09, WDL^{+}23, ZZLM23, AAM05, ENW96, GLMM04, GB99, HL96b, LWR15, LWR^{+}16, RVA00, VLMN09, VL05]. long-haul [LWR15, LWR^{+}16]. long-line [VLMN09]. Long-Lived [RUH^{+}18, CDFG06, GLMM04]. Long-Range [HCL^{+}17, ENW96, GB99, HL96b, RVA00]. long-run [VL05]. Long-Term [DSYN24, WDL^{+}23]. Longer [QCMY16]. Longest [DKT06, HWHW18, RT17, BBHK14, DKN96, DKN97, LBX11, PT12]. longest-matching [DKN96, DKN97]. longest-queue-first [LBX11]. Longitudinal [ACZP21, BS19, FAF^{+}17, LXW^{+}17, LJZ^{+}23]. Look [AQK^{+}19, YHH^{+}21]. lookahead [BAC12]. Looking [LHW^{+}20]. Lookup [HWHW18, LLY^{+}22, LP24, QCMY16, SBLS19, WLL^{+}16a, YXL^{+}18a, YBQZ18, AN05, BLC12, MPL09, PT12, SK03, SFAS05, SMLN^{+}03, ZGG05, ZHLL06]. Lookups [GYSZ19, LSV99, LXX^{+}14]. Loop [BBD^{+}18, FLMS18, GLA93, NGK19, RpLP^{+}17, RS21, WGZC21, ZGZC20, GLAM97, MBF^{+}02, PT94, fTL06]. loop-back [MBF^{+}02]. Loop-Free [BBD^{+}18, FLMS18, RpLP^{+}17, WGZC21, ZGZC20, GLA93, GLAM97]. Loopback [CSC04]. loops [FB07]. LoRa [CW23, GSH^{+}22, HXZ23, XZG20, XZG21].

Lord [HSFK09]. Loss [AEG^{+}17, CLM^{+}18, FLM^{+}22, KS01a, LPM23, MH02, NJM^{+}19, QJCR20, WLD^{+}16, BLCT97, BSS^{+}11a, CN10a, CH04, CU95a, CTG00, CLW95, CKR93, DLPT06, GS98, HC02, HAGL16, KK00, LM97, LMS00, LA95b, LGKV14, LMSKZ99, LB04, LWR15, MEVSS03, MG97b, MMR96, NR13, NBT98, PL02, SL94, SS98, SBDR08, VS97, VSR11, Wil96, XFS06, XK06a, XG05, ZF96, vDP93]. loss- [BSS^{+}11a]. loss-free [VS97]. loss-load [Wil96]. Losses [LTDM17, NSP^{+}16, AAB05, AT03, BV05b, CCV03, KS03, YMKC08]. Lossless [VVP^{+}12, ZMLR23, ZWCL17, KGL03, LCY96]. Lossy [CBL06b, GLA19, RT17, AAM05, JS14, KL07, Kum98, ML06]. LOTOS [MBC^{+}94]. Low [BES22, BSSS21, BLM^{+}17, BSYS12, CCW^{+}17, CGC^{+}18, CGR^{+}18, CZL^{+}19, CLH^{+}24, CNG^{+}16, DRMP18, DRW^{+}22, GSW^{+}23, GLA19, GLS09, HGB^{+}19, JGLS14, JLZJ19, KLC^{+}18, KK06a, KLE16, LLZ^{+}23a, LYSZ16, LOFH21, LLS10, LCZH17, LSC^{+}21, LPWP22, LS10, LYW^{+}18, SRI^{+}18, SRR08, SS09, SBTH19, TSN^{+}21, TWH24, TKXP20, TWL^{+}21, WCWZ17, WFC18, WHZJ20, WZLM22, XYL^{+}17, YSC16, ZCW15, ZRP^{+}22, ZDB^{+}17, ZMD^{+}20, ZMMG22, AYM14, BM09, CHML15, CPS13, HLW13, HL15, JGS^{+}15, KR00, KMH12, KK06b, LQ13, LH13, LMS04a, qLP97, LPP11, LBS05b, NTS12, PLS07, QSS^{+}15, RSR10, Szy16, YDS10]. Low- [LOFH21, LBS05b]. low-accuracy [BM09]. Low-Complexity [DRMP18, BSYS12, GLS09, JGLS14, LLS10, ZCW15, HLW13]. Low-Cost [CCW^{+}17, CLH^{+}24, LSC^{+}21, SBTH19, LPP11]. Low-Delay [YSC16]. Low-Dimensional [TSN^{+}21]. Low-Duty-Cycle [CNG^{+}16, CHML15]. Low-energy [SS09]. Low-Latency [BLM^{+}17, CGC^{+}18, HGB^{+}19, LLZ^{+}23a, XYL^{+}17, AYM14, QSS^{+}15, RSR10].

Low-Overhead [GSW^{+}23]. Low-Power [BSSS21, DRW^{+}22, KLC^{+}18, SRI^{+}18, WHZJ20, ZDB^{+}17, ZMD^{+}20, ZMMG22, LS10, PLS07]. low-precision [KMH12]. low-priority [KK06b]. Low-rate [KK06a]. Lower [CLW16, ZGZC20, AGLM10, wTjCjC97]. Lowering [VMCB22]. LP [KK06b]. LPWANs [TWL22]. LRD [YTJQ05]. LRU [QTE20]. LSQ [VKO20]. LSRP [AZ06b]. LTE [LCS^{+}18, AY20, BRY^{+}19, BLM^{+}17, CLGSS17, DMMS14, DM15, KLP16, LCSS17, LPCVC13, LYZ^{+}23b, MSRG18, PLR15, PL17, WT17]. LTE-A [LCS^{+}18, BLM^{+}17, LCSS17]. LTE-LAA [MSRG18]. LTE-Multicast [BRY^{+}19]. LTE/802.11 [PL17]. LTE/WiFi [CLGSS17]. LTP [WBP^{+}11]. Luminaries [LJJ^{+}19]. Lyapunov [AN20, WN16].

M [CM16, RW95]. M/G/1 [CM16]. M/G/1/N [RW95]. M2M [WZL^{+}13]. MAC [AK00, AGM^{+}17, BCS^{+}19, BJY11, BCGM07, CRB09, CHL16, CSS06, CLG^{+}00a, GKB^{+}16, GSL^{+}24, HDM10, JZC11, KIR06, LKC11, NSY20, NSY23, ODC^{+}16, PLM19, RWA^{+}08, RSSZ13, SRBBG17, SA01a, SS07, TS08, VA07, Wan04, YD07, YDS10, YZZ^{+}21, ZB95, ZT03, ZKEN23, ZZLM23]. MAC-layer [CHL16, JZC11]. Machine [BPW23, CN19, CYX^{+}17, HTW^{+}19, LYLW22, MCMdlO23, SRCT23, SCN^{+}24, SKA^{+}18, WLC^{+}20, WLS23, XOYL20, XLL^{+}20, ZWL^{+}22, ZQL^{+}23, LWLL16, MSBZ10, NABZ12, SJL^{+}13]. Machine-Learning-Assisted [SRCT23]. machine-learning-based [NABZ12]. machine-to-machine [SJL^{+}13]. Machines [HKLM17, Nai97, WRS^{+}15]. macro [CK10b]. Made [LMSS24, ABK15]. Maelstrom [BMB^{+}11]. Maintaining [GDL^{+}22, JRY09, FK99]. Maintenance [ADT22, DYJ20, SQS20, AA93, AACD^{+}96, FEC13, SLL^{+}11]. Make [DJCA21, CPS13].

Make-Before-Break [DJCA21]. Making [ABBH^{+}16, AC06, BEK^{+}22, CF94, KE21, LLY^{+}16, LZL^{+}21, NSC^{+}22, She95, XZC^{+}20, XSHS12]. Malicious [AQK^{+}19, FHQ^{+}17, RHMF16, SKG12, SAM12]. malware [EKSV16, KSA12]. MAN [RIM98]. Managed [KLK^{+}20, NRB22]. Management [ACC^{+}14, CHO^{+}19, CMR17, DZ20, DXX^{+}23, GZJ^{+}18, HTW^{+}19, HNP23, HWC22, HGZ^{+}23, JD22, RMPG16, SC17, SCL^{+}23, WBM^{+}18, YXC^{+}18, YXCH21, ZWR^{+}23, ZLL^{+}23a, ZLHM22, ASW00, AYS^{+}13, ACP05, AGKK03, AJ06, BCP13, BLPS10, BRISCSP11, CqLL98, CHH06, CH97, CL16b, DC13, DM15, DGG^{+}02, DJM97, cFSKS02, FJ95, GP96a, GMYP16, HL99, HL98a, HM06, HKV^{+}13, HBS96, HA97, IPG97, IS00, IK07, KJF^{+}00, KS04, LBS05a, LBB08, LAJS07, LKL00, LGS09, LH03, LJC05, LSM^{+}14, Low03, MSWL06, MPFK02, MS08, MRD08, MW05, RrBG94, RRK96, SM14, SV99, SL15c, SCY98, SIYL09, STL04, VG04, WL08, WQZ^{+}13, YBG^{+}12]. manager [CU95a, LYS93]. Managing [DRCM^{+}17, PD07, RLZ^{+}18, SBM^{+}18, dFV02, KS12, YC12]. MANETs [CPS13, DPMK11, GLAMM11, JHR05, LJNK12, LZC^{+}17, PDE08, SL15c, WGvdS17]. Manhattan [LK95]. Manner [FXQ^{+}21]. MANO [MCC^{+}19]. Many [SK11, HLHD^{+}04, SK10a, SK12a, XSZ^{+}07]. Many-to-Many [SK11, SK10a, SK12a]. Manycast [PGV16]. manycasting [BV10]. Map [HZ20, CS14]. Mapping [BBR19, GJWZ16, JPM^{+}19, MYC^{+}19, WWW^{+}18, WZX^{+}22, YXCH21, CRB12, DK98, FJ07, JK15, PPK15]. Mappings [GHRH18, CDPLCA16, TR98]. MapReduce [FLG^{+}20, FC17, VN22, WZY^{+}16]. maps [DJ14, GS09, MG16]. MapTask [WZY^{+}16]. Market [LSK20, NLB19, RLZ^{+}18, SA21, VNM22, ZLWH17, GS16, KAS16, MQ05, SL14, XB07].

Market-Based [NLB19, MQ05].

Marketing [NTD17, DZNT14]. Markets [Ma16b, MHB^{+}21, NS21, SAKMB21, TH21, XRL^{+}22, ZMWX18, AAS10, HGW^{+}16, IGHT15, RPV13]. Marking [SR18, CHM^{+}05, CFM^{+}09, EW08, FK99, Goo08, TC06, YDS06a]. Markov [AS94, GMWD13, KWC93, KLE16, REM17, RCFC15, RV01, SRS03, SC17, WUZ^{+}19, XY09a, ZS04]. Markov-chain [ZS04]. Markov-Chain-Based [KLE16]. Markov-modulated [SRS03]. Markovian [EM93, ODT09, OES16]. marks [KS03]. MARS [ER20]. MASK [FLZ^{+}23]. mass [RS95b]. Massive [BSRdA16, BCLS17, CEC^{+}19, GCMP20, LZY^{+}22, OBS17, RWL^{+}22, XOYL20, ZSS^{+}20, ZAW^{+}22]. Massive-MIMO [GCMP20]. match [BBHK14, CW16]. Matching [Hua17, LS06a, LT16, LN19, LDL^{+}22, MPN^{+}14, Mne08, MHL19, RT17, WZL^{+}23b, YDW18, BBK12, BBHK14, BESW08, DKT06, DKN96, DKN97, FDG^{+}11, LH13, qLlH97, LK10, LS03b, PLT14, PT12, TT09]. matching-based [BESW08]. matchings [BE06]. mathematical [ZLC12]. Matrices [TR17, OMA^{+}10, RZWQ12, SNC^{+}07, ZRLD05]. Matrix [CLY^{+}17, FLBR^{+}19, Lia06, SYZP19, TMGB19, WDR^{+}20, XCW^{+}20a, ZHZ^{+}24, ZLN^{+}17, LDGL13]. Matter [DDP^{+}19]. matters [MSS^{+}12]. Max [KAA^{+}18, LCS12, MMT14, MS15, VL16, AS08, GL10, JMMT12, LPW14, LJA14, Mar03, MRHWS14, NDGL06, NJW16, RL07, YXF^{+}13, YLLY05, ÇTD22, CLK01]. Max-min [LCS12, AS08, GL10, LPW14, Mar03, NDGL06, RL07, YXF^{+}13, YLLY05, CLK01]. MAX-Tree [ÇTD22]. Max-Weight [KAA^{+}18, MMT14, VL16, JMMT12, LJA14, NJW16]. Maximal [ÇTD22, CSD22, VTBK21, WWL24b, WW16, BCR^{+}12, BESW08, CLSC15, JLS09, LNS11, MP94, Nee09, RB09a]. Maximizable [GS03]. Maximization [CSD22, CGYZ16, FM22, GCX^{+}17, JSW^{+}20, KSNR20, KSRW22, KTvdSK18, LSDT19, LXX^{+}17, LZC20, LNLM24, MLX18, NCM18, NZW24, RR19a, SYZP19, SSM20, SGJ17, TWTD17, WWC^{+}18, XLH^{+}17, XSZ^{+}22, XXW^{+}23, XWX^{+}24, YLF^{+}21, ZND^{+}16, ZHW^{+}17, BMS14a, BZM08, CPS^{+}12, EML12, JW10, LLCL11, LWLL16, LCZC13, Nee13, PPSV13, RRK96, SN15, TCS13, ZNZT16, ZG08]. Maximize [DSYN24, LHL^{+}21, LH10]. Maximized [ZFLC18]. Maximizing [BMY^{+}17, CGR^{+}18, CZTX23, CN10b, DPSA21, KK16b, KLT15, LLM14, LQXX07, LZES14, LHY^{+}23, LJW^{+}07, LJSB22, NTD17, OJSY16, RL18, WZH^{+}24, ZCJ^{+}13, ZSZN21, CSS06, HY08, HN10, IKDD15, KLSS10, VGP14]. Maximum [BB16, BSP07, CT04a, CLS^{+}18, CSL21, HKV23, KSA12, LWAL17, LZC20, SGR13, VLDM17, ZWL^{+}16, ZSLZ21, CKKK09, CK09, GR14, JLRS16, KKL03, LMMN07, LLL06, Lia06, MBG^{+}03, NTS12, OR11, WMFS10]. maximum-degree [OR11]. maximum-lifetime [WMFS10]. MaxWeight [LLM23]. MComIoV [LDD21]. MCR [FBFB17]. MCS [LLM^{+}24]. MDFE [MVCS16]. MDLdroid [ZGZ22]. Me [AQK^{+}19, XXZ^{+}22a]. Mean [HTAZ16, LBP^{+}17, NSY20, WD22, WYL23, WYL24, CTG00, HH98, LLE16, LZC09, SSV13]. Mean-field [HTAZ16, SSV13]. Means [FZQ^{+}22, BMM^{+}09]. Measure [WYZ^{+}24, MOZ05]. measured [DL04, KZDM07]. Measurement [BPK^{+}10, CCK16, CCC17, CLZ^{+}23, CJL^{+}19, DLH^{+}14, DHS^{+}23, EFFK18, GTC^{+}24, GMSK09, HHD22, HSM^{+}20, JID^{+}07, LXW^{+}17, LYY^{+}22, LJZ^{+}23, MGK12, MZZ^{+}23, NKS08, NS16, QK01, RRK07, SL16a, WSKV08, WLD^{+}16, WLS^{+}18, WDR^{+}20, XWW^{+}19, XTW^{+}22, XOW^{+}23, YHH^{+}21, YXY^{+}18, ZNN^{+}10, ZZS^{+}16, ZLW^{+}20, ZLW^{+}19, AKS96, BMVB09, BLCT97, ES03, GXWW11, GT99, GT03, JD03, JDSZ97, KS09a, KYY^{+}12, qLlH97, LCL12a, LHC05, NCT14, PBKG11, RW07, RKT02a, SJL^{+}13, SNSW12, SBDR08, SQZ09, WZR08, WDCL15, XYLL14, YCM11]. measurement-analytic [ES03]. Measurement-Based [CCK16, NKS08, QK01, RRK07, ZNN^{+}10, BLCT97, GT99, GT03, JDSZ97, WZR08]. Measurement-driven [BPK^{+}10, MGK12, PBKG11]. Measurements [BBEF^{+}21, MHS^{+}17, MRMR17, QJCR20, RFGL17, WZL^{+}23c, XPW^{+}18, YJL^{+}19, ZMW^{+}22, AdE07, GCS06b, KHG^{+}14, KLSV12, LDK13, LTY06, MHL^{+}14, MSA^{+}16, NR13, NXTY10, SCY98, WJK^{+}12, ZKH10]. measures [AK96, ANSX13, PS09, TJ95, WLS97]. Measuring [AFT11, FZQ^{+}22, GMLP10, HBB09, LLZ^{+}23b, NKL^{+}23, SMWA04, ZL13a, ZLB17, LGKV14]. MeasuRouting [RHC^{+}12]. MEC [DPTP24, PLT^{+}20, ZWR^{+}23]. Mechanical [YLL^{+}17]. Mechanism [AB21, AWH^{+}22, CZD^{+}22, EL24, GBG^{+}16, JSXN18, LZX^{+}21, LCH20a, LCH20b, LFC^{+}22, PK01, QW23, SC18a, SS21, XQG^{+}22, XXN^{+}19, ZRH18, ZLWH17, BLPS10, BCB99, CLSS09, CO94, FY07, HGW^{+}16, IGHT15, NL16, SMT98, SA04, SK12b, SMP^{+}14, WKWV16, XL11a, ZWTC16]. Mechanisms [GCD23, MCM^{+}23, NS21, SYG^{+}22, TPW^{+}18, VHT21, XRL^{+}22, ZCPP22, BPSK97, ÇY07, CFPP96, CY14, CLA07, FHH10, GKPS06, HGE04, LSM^{+}14, TYP^{+}15, WZR08, WHTC15, YXFT16, ZLM16]. media [AS02, BS02, CG04, KAZ01, LA95c, MEVSS03, PWMC12, PSA96, RVR93, SKR^{+}09, SZG^{+}13, VNS02, VAM^{+}06, WLCW16, WWL^{+}15, YJH05, ZEV07b]. Medical [SDM20]. Medium [MGK20, PV10, SMGP15, URZ^{+}14, YHE04, YSY16, ZW22, BBL95, CLD10, IZC00, JCJ95, KH15, MLS12, PPV12, SMM11, SS03, VA06, VA07]. Meet [CVHM22]. Meeting [LFXY23]. Meets [ALR^{+}24, CCZL23, FKCA18, HZCL16, KSAK18]. mega [LZXF14]. Membership [QCMY16, RAA^{+}24, AGKK03, HKLS12, KEAAH08, MR96]. Memento [BEK^{+}22]. Memory [DLW^{+}17, DD24, HWC22, KGH^{+}20, LSSC22, XCZL20, YBQZ18, ZLG^{+}17, AS09, CH97, CH98, Geo08, HKLM07, LH13, LMT16, LLS09, MAN15, PV10, SFAS05, SSZ05, XLZC14, YLCP11]. Memory-Efficient [DLW^{+}17, YBQZ18, XLZC14]. Memory-Rate [DD24, MAN15]. merchant [AMI^{+}07]. Merlin [SBM^{+}18]. Mesh [CZL^{+}19, FLBR^{+}19, AK14, AK15, ATB^{+}10, AAV09, AST11, BTH11, BLB10, BL04, BLRC05, BZM08, CYK09, CSC04, CCF04, CAL09, CK09, Con11, DPBT11, DSTM12, DYX12, EFK07, EM09, FCT03, GM03, GMSK09, HTC04, HMM11, IMG98, Kam10, KS09a, KS11, KN05, KMHS09, KHW12, LBRA05, LCS12, LWKD03, LCG^{+}14, LYL07, LLY09, LRG10, MVRZ09, MR09, MPF^{+}15, MBF^{+}02, MHRR12, ME96, MJ15, PNRMC13, PA12, PCV08, PCL15, RGKR10, RDO^{+}07, RCGS09, RJJ^{+}11, SJ12, SYR05, SMM11, SSM06, SZM08, TWHR11, THBR14, TH97, Wu94, XTMM11, Xin07, ZOM03, ZZZ^{+}07, ZKL11, ZZZM03]. mesh-based [MR09]. Mesh-Structured [FLBR^{+}19]. mesh-survivable [IMG98]. Message [LDD21, RKNS10, SDSY19, ZLX^{+}23, ZYY^{+}21, FGM^{+}13, HR95, PHL15, Rum93, ZB95]. Message-efficient [RKNS10]. message-passing [PHL15]. Messages [AAR18, BC01b, FJ94, JMI95, MGK14]. Messaging [WCM^{+}21, WEK97]. Meta [DQW^{+}23, SdVS22]. Meta-Scheduling [SdVS22]. Metadata [HGZ^{+}23, QW23]. METANET [OY95]. MetaRing [OSW97].

Metaverse [NZW24]. meteo [PS15]. metering [WMYR16]. Method [FAS^{+}23, HKS16, HWZ^{+}23, LL20, LKMK20, QYC^{+}24, SYL^{+}17, WSC^{+}23, ZWCL17, ZWY^{+}18, FM06, HGE04, JL98, KVR02, NR98, PPPW05, TAH99, ZDR04]. Methodology [KLKP16, CLM99, DRM04, FGL^{+}01, GIL^{+}15, JD03, SV98c, TB10]. Methods [CG21, LTCS22, LNR94, BL04, DT15, JF04, TMH11, WM96]. Metric [MKAE20, QHZC18, WMP^{+}18]. Metrics [LGDC23, LDZC20, GR12, GS03, JHR05, MHL^{+}14, PNRMC13, PA12, RS07]. metro [QGCL11, JLZJ19]. Metronome [FBQ^{+}23]. metropolitan [HL98b, KV96]. metropolitan-area [KV96]. mice [MGG^{+}05, MK10]. Micro [SMEH20, SR18, SRC^{+}20, CK10b]. Micro-Burst [SR18, SRC^{+}20]. Micro-Services [SMEH20]. MicroCast [LKS^{+}16]. Microfluidic [DGLM16]. Micromovement [XWL^{+}18]. MICRON [RS08]. microscopic [LGKV14]. Microsecond [VMCB22]. Microsecond-Level [VMCB22]. MIDAR [KHLC13]. Middle [PMN19]. Middlebox [FGRQ18, HKC^{+}20, WLW^{+}20, XXZ^{+}22a, FGR^{+}17]. Middlebox-Based [FGRQ18]. Middleboxes [HSO19, KRS^{+}17, LDS^{+}24, TML^{+}18, YDW18]. Middleware [BTK^{+}17, SHZ16]. MiFi [BB06, MGZ^{+}23]. Migratability [YXZ17]. Migrating [CSR^{+}20, NG16, YL98]. Migration [BFG^{+}14, CYX^{+}17, EMAL17, WLCW16, WUZ^{+}19, ZLZ21a, SLO^{+}14, WRS^{+}15]. migrations [RZC11, VVP^{+}12]. Mile [PPV17]. Military [HK14]. Milking [WTK^{+}17]. Millimeter [SKE19, YXAZ^{+}18, YLWH20, ZW22, ZWZM18, ZXW^{+}20b, AWFT15]. Millimeter-Wave [YXAZ^{+}18, YLWH20, ZW22, ZWZM18, ZXW^{+}20b, AWFT15]. milliseconds [BFF07]. MIMO [BRM^{+}13, BJY11, BSS14, CAS^{+}20, CZW^{+}21, CW10, GB18, GNP^{+}13, GHK18, GCMP20, LSCT17, OBS17, PLL13, QZZ^{+}13, SYZP19, XWJ22, XOYL20, ZP18, ZK19, ZSS^{+}20]. MIMO-assisted [BJY11]. MIMO-aware [PLL13]. Min [MCZ^{+}22, AS08, CCLT02, GL10, LCS12, LPW14, Mar03, MRHWS14, NDGL06, RL07, YXF^{+}13, YLLY05, CLK01]. min-max [GL10, MRHWS14, RL07]. Mind [WTK^{+}17]. Minimal [AMS22a, CMP^{+}14, CDK^{+}17, GPLT15, CVM^{+}15, Ili00, MP93]. minimal-length [MP93]. Minimization [AY20, CGC^{+}24, DPTP24, GR20a, HS14, HS16, JWW^{+}23, RTNS21, TW22, TW23, AAZZ12, BO07a, LLS10, SV15, ZL16]. Minimize [ACLX17, PLD16, dOSAU04]. Minimized [LTZ^{+}22]. Minimizing [CMN12, CE08, GMP08, HDF19, KSUB^{+}18, KLSS10, LS16, LYD^{+}21, LGS^{+}23, LZC22, SG18, WXH^{+}18, WYHL09, XYQ^{+}17, XXZ^{+}22b, ZWL^{+}16, ZWJ^{+}20, hCgKsYwT96, CK09, LMT10, SZ07, VL10, ZWO^{+}96]. Minimum [AdSD16, DPM^{+}18, FSH^{+}13, GCWC17, KWS^{+}11, KKH^{+}22, LWK^{+}16, LS17, LRM^{+}06, MJ15, OdG96, OR11, ORS93b, PBT^{+}20, SL95, SZ22, SZMD17, SLH^{+}19, SRB^{+}20, WCY04, WHYC23, YRB^{+}18, ZWYY10, BLS07, CFM13, CLK01, cFKSS99, FEC13, Geo08, HLL13, KWCR10, NPY07, PZGLA98, TK12, UN11, Wan04, XY10a, XGF^{+}14, YYZ06, ZH08b]. Minimum-Cost [LS17, SRB^{+}20, LRM^{+}06, ZWYY10, PZGLA98]. Minimum-delay [MJ15, BLS07]. minimum-energy [HLL13]. Minimum-latency [OdG96, SL95]. minimum-maintenance-cost [FEC13]. Minimum-power [WCY04, Wan04]. Minimum-Weight [YRB^{+}18]. Mining [JW23, ZSZ^{+}17, LLW^{+}09]. Minorization [SYZP19]. Minute [DPSA21, SKG^{+}18]. Misalignment [SC18b]. misbehavior [CRB09]. misconfigurations [LLW^{+}09]. miser [BRS06]. Misreporting [ZSS^{+}20]. Missing [LCQL14, SL16b, XWW^{+}19, ZCZF20, HSFK09, LCL13a, ZL15].

Missing-Tag [ZCZF20, LCQL14]. mission [EML12]. Missions [KSSD24]. Mitigate [ZHWH21]. Mitigating [HLL^{+}21, KKV16, KG99, SRC^{+}20, TEML09, ECN09, WZR08]. Mitigation [AS19, BPA20, BPA21, LJHB18, ZLW^{+}19, AYS^{+}13, CH11, LPCVC13]. Mix [JV17, SD00]. mix-dependent [SD00]. Mixed [ZGL^{+}19, ZSL^{+}21, BSH^{+}11, VWT^{+}14, VSR11]. Mixed-Cast [ZSL^{+}21]. mixed-line-rate [BSH^{+}11, VWT^{+}14]. Mixes [OPGT16]. Mixing [LYZ^{+}23a, DMK05, RVR93]. Mixnet [ZLM^{+}23]. MLSR [AEB02]. mm [DF20]. mm-Wave [DF20]. mmWave [DJK22, DJK23, GHW22, LJSB22, SKA^{+}18]. MNCM [TT09]. Mobile [ADT22, AZP^{+}23, AP17, CBDCP19, CPKL17, CPS17, CJLF16, CBZ16, CS17, CZX18, CLHY22, CW19, CJ18, CSR^{+}17, DYJ20, FFZ^{+}18, GCWC17, GFW^{+}18, GCX^{+}17, GS19, GZJ^{+}18, GXS^{+}21, HHL18, HHA17, HMM^{+}20, IGHT17, JLS^{+}17, JSXN18, JHS^{+}19, JWSH18, JD20, KTvdSK18, LLY^{+}13, LXW^{+}20, LXX^{+}17, LGZ^{+}23, LYS^{+}18, LYW^{+}21, LW22, LSL17, LLX^{+}19b, MCdG23, MS17, MBN^{+}21, MKG^{+}17, PP17, QLY23, SCW^{+}21, SCL^{+}23, SGS20, TEE16, TE16, TPW^{+}18, WPZM16, WWW^{+}18, WUZ^{+}19, WLY^{+}24, WYL24, WLL^{+}24, WCZZ17, WML^{+}18, WWX^{+}19, WHC^{+}22, XGW^{+}20, XFCW18, XCL^{+}18, XXZ^{+}23, XLW^{+}17b, YCC^{+}21a, YZL^{+}18, YXL^{+}19, YWW^{+}23, YWW^{+}24, ZGHH19, ZLRC20, ZYH^{+}21, ZGZ22, ZWR^{+}23, ZRD^{+}23, ZQW^{+}23, ACR12, AWKN16, AKSS12, ACCF12, CE09, CZF^{+}16, CPGZ15, CDH^{+}10, CFZ97, CMGL11, FHH10, Fan05, GGL09b, GGL09a, GH04, GV06, HL98a, HLS14a, HAGL16, HH10a, HSPH09, HH10b, IGHT15, KLZ12, KSA12, KD10, KG10, LH07]. mobile [LKC^{+}13, LSC99a, LC04a, LCL^{+}12b, LKZ^{+}04, MD04, McM95, MWC16, MEWP13, NL99, NCT14, PD16b, PMH95, RM08, SMS07, SHN16, SK06, SPR08b, SPR08a, TRKN12, TLP^{+}16, UN11, WSC08, WWL02]. Mobile-Edge [CJLF16]. Mobile-Edge-Cloud [MCdG23]. mobiles [KAEAS14]. Mobility [BPVRSP16, CCZL23, GT02, JYC^{+}16, QTWW16, TXL^{+}18, WLWL13, YYL23, ZFW^{+}17a, ZHZ^{+}18, ZLL^{+}23a, AW04, AGL16, AS07a, AS07b, BCB99, BLDF09, BLB10, CMGL11, CPS13, HL99, HSPH09, IPG97, LBB08, LKL00, LH03, LSMS06, LH10, MYYR13, MHS95, MSA^{+}16, PS15, RVS^{+}02, RSH^{+}11, VG04, WA11]. mobility-aware [BLB10, WA11]. mobility-transparent [BCB99]. MobiScatter [TZL^{+}24]. MobiSpace [LW11]. MobiT [YSC18]. Möbius [SJSB22]. Modal [RZE^{+}21]. Mode [CWAO21, EL24, RBPS21, AKS96, MBG^{+}02, XWG14]. Mode-Suppression [RBPS21]. Model [BMY^{+}17, CTG^{+}20, CH20, CMP16, CM16, GHBSWV17, GVM23, GCZY18, GCW21, HS16, HAG19, HSO19, HWZ^{+}23, HWM^{+}24, HH17, KGdV^{+}21, LXX^{+}24, MLB21, OOM^{+}18, PBGMFM22, RHQZ13, SXEZ21, SGJ17, VHT21, WLD^{+}24, WWTK11, YLH17, YYFC24, ZGYB20, AIN^{+}15, Ada98, AS07a, AAB05, AAZZ12, ASSK13, BBM93, BPPP12, BBFG95, CAK12, CT95, CHA95, CBAT06, CJZS14, CL08, CL09b, CDPLCA16, EMPS06, FJ07, FNQ00, FK03, HS06a, HAGL16, Hey97, HLP11, IK09, JC13, KZ97, KLOS11, KLS11b, KRSY02, KV09, LV06, LDH^{+}12, LWL04, LLLT10, LNC93, Low03, LC94b, MGG^{+}05, NST01, NCT14, PFTK00, PMW10, QZZ^{+}13, RCFC15, SSV13, SWL06, SSD93, SV98b, TYJ16, TCPV13, XY09a, YWLL09, YMKC08, ZY07a, ZCL11, ZFC15, ZZZM03, ZY16]. Model-Agnostic [SXEZ21]. Model-Based [ZGYB20, WWTK11, AIN^{+}15, YMKC08]. Model-driven [RHQZ13]. Modeling [AGM^{+}17, BK17, BBCD14, CR99, CBAT06, CCY^{+}14, FCL97, Fan05, FFX^{+}17, GSK08, GYSPR14, HOT97, HL03, HSPH09, HMvdLM07, JSW^{+}20, KL07, LBHO07, LMSR19, LRC15, LC04a, LZX^{+}21, LTN^{+}19, LFY^{+}19, MJ01, MCLG07, MDL07, MS17, MSRG18, NGK19, PFTK00, PPV17, PWWP18, SRS03, SJGH10, SLD^{+}23, TS08, WLL13, WLS^{+}18, WWW20b, WLR10, XWH^{+}16, XL23, YR01, ZHZ^{+}18, AS07b, BG98, BYH^{+}15, CAO11, CZCC14, DM14, FNQ00, GMSK09, HS08, HDM10, Kam96, LT02, LZL12, LMS04b, LG13b, MGK12, MCR10, NT00, PF95, SGSB^{+}15, SNSW12, TG09, TB10, WL10, WA11, WK13, XB07, YZ10, ZS04, ZNN^{+}10]. Modelling [ACZP21, YLF^{+}21, ZRK06]. Models [BPVRSP16, BBR19, CEC^{+}19, DME23, LXLC20, SA21, TT17, ZLY23, ALWD05, AS07b, BGK^{+}16, CFG08, FJ95, GLMM04, GS98, HL96a, IZC00, LJ09, LNR94, LTP10, MCS99, MA12, MBM09, NS03, Pax94, SD15a, SKV03, TMP07, ZCD97, ZL16, vRWZ09]. moderate [LMW16]. modern [SRS08]. modes [Tha04]. modification [WSMJ04]. Modular [PPS^{+}22, BYH^{+}15, IBM95, KR00, LY94]. Modularity [BMB19]. modulated [SRS03]. Modulation [CK10a, CZTX23, GSH^{+}22, LCW^{+}24, CGK10, EF08, LZZR12, YZBR14]. modulo [OdG96, SL95]. modulo- [OdG96, SL95]. Molecular [DGLM16]. Moment [PJ13]. Moment-based [PJ13]. Moments [XCL^{+}19, XCQ^{+}23]. Monetary [ZRH18]. Monetization [YCGH17]. Money [TH21]. mongering [DMC06]. Monitor [DGW^{+}17, HGM^{+}17, MHL^{+}14]. Monitoring [ABMT23, BRY^{+}19, CLK^{+}24, CFM^{+}19, DGNK21, FAS^{+}23, GSW^{+}23, KP23, LDW^{+}20, LZZ^{+}22a, PKK18, PBGMFM22, RAA^{+}24, SLD^{+}22, TCTP20, TXW^{+}21, XY09b, XCW^{+}20a, XCW^{+}20b, XWX^{+}24, ARK09, BTH11, BRISCSP11, BR06, CBSK07, JL12a, Kuc14, LCH^{+}06, RW93, RHC^{+}12, SLC^{+}07, SBDR10, TAG08, THRW12, WS05, XZB08, ZF96, ZGTG05].

Monitors [LLL^{+}22a]. Monocle [PKK18]. Monotone [DME23]. monotonicity [IK09]. Moral [RL23]. Morphing [WMCW22]. Morphism [FZQ^{+}22]. Motif [QZC^{+}22]. Motion [LLZ^{+}19]. motioncast [WHW^{+}11]. motivation [CSEZ93, WJLH06]. move [KM10]. Movement [AHL96, GCWC17, ZLL^{+}23a, SH12]. Movement-Based [ZLL^{+}23a, AHL96]. MP [CLL^{+}19]. MP-RDMA [CLL^{+}19]. MPEG [FNQ00, LS03b]. MPLS [CN10b, HM04, LBB08, SSFM08, WL08, dOSAU04, vDJJ^{+}22]. MPLS-based [HM04, LBB08]. MPR [BJY11]. MPR-aware [BJY11]. MPSoCs [FMCS20]. MPTCP [FKCA18, HGB^{+}19, KGPL13, OL16, XXZ^{+}23]. MRF [CLS07]. MSP [LS93a]. MST [CFM13]. MSXmin [KR00]. MTI [ZL15]. MTU [MG95]. MU [GHK18, XWJ22]. MU-MIMO [XWJ22]. Much [LL17a, LLY^{+}13, SSFM08]. MulTFRC [DW11]. Multi [AAAR19, ER20, AP17, AWM^{+}20, AM19, BGHS10, BVBV17, CBdV^{+}17, CBDCP19, CZGKB24, CE19, CHO^{+}19, CJLF16, CBZ16, CLM^{+}18, CLL^{+}19, CKZC19, DJK23, DZ18, DPTP24, DZ20, EGR^{+}16, EL24, FMPS20, FM23, GJCB18, GHK^{+}23, GZL^{+}17, GB18, GVGV17, GLS21, GCW21, HSH^{+}06, HVT18, HAB^{+}22, HRM22, JTL^{+}17, JTL^{+}18, KSNR20, KSRW22, KS19, KCH^{+}19, KLY^{+}23, LFZ^{+}22, LFC18, LPD^{+}18, LZL^{+}21, LLCJ22, LLL^{+}17, LCX^{+}19, LAJ20, LSC^{+}21, LZC22, LCZ^{+}23, LYKT21, LLM^{+}24, Med95, ML23, MAPZ18, NGK19, NLB19, PWL^{+}24, PG18, QHZC18, QCMY16, QLSW19, QLF23, QYC^{+}24, RTNS21, RZE^{+}21, SFM^{+}18, SLS^{+}23, SK21, SPM^{+}17, SFS^{+}22, TPW^{+}18, TWH24, TH97, TTM22, TZX^{+}22, WZH^{+}18, WZX^{+}22, WLY^{+}24, XSH^{+}15, XWL^{+}18, XOYL20, YZY^{+}20, YXAZ^{+}18, YXL18b, YXCH21, YYB^{+}22, YCZ^{+}23, YWHW24, YLL21, ZZ17, ZHZ^{+}18, ZK19, ZW22, ZQW^{+}23, ARS16, AAV09, BSH^{+}11, BESW08, CW16]. multi [CF94, CRS99, COS95, DV09, GKJ12, GSK08, HIM07, JS09, KN05, KS09b, KG16, LMS05a, LMS05b, LHB^{+}05, LRL08, LJ09, MHSC95, MRD08, Nee08, NL07, NSCR06, SKK07, SKE16, SCY15, TMH97, Voi07, XZTT08, YS07, ZL16, ZGS10, FLZ^{+}23]. multi- [ZGS10]. Multi-access [TH97]. Multi-Agent [PWL^{+}24, QYC^{+}24, WLY^{+}24]. Multi-AP [GB18]. Multi-Armed [AM19, HVT18, GKJ12]. Multi-AS-Key [FLZ^{+}23]. Multi-Associated [YYB^{+}22]. Multi-Attribute [ER20, LZL^{+}21]. Multi-Attributes-Based [WZH^{+}18]. Multi-AUV [QLF23]. multi-band [SKK07]. Multi-Beam [ZW22]. Multi-Bit [ZZ17]. multi-bit-rate [BSH^{+}11]. Multi-Bottleneck [YLL21]. Multi-Carrier [LPD^{+}18]. Multi-Category [LLL^{+}17, LCX^{+}19]. Multi-Cell [AP17, HRM22, MAPZ18]. Multi-Channel [CE19, CBZ16, DZ18, MHSC95]. multi-class [KG16, LMS05a, LMS05b]. multi-constrained [XZTT08]. Multi-Constraint [TTM22]. Multi-Core [CHO^{+}19, YYB^{+}22]. Multi-CQF [YCZ^{+}23]. Multi-Criteria [AWM^{+}20]. Multi-Dimensional [TPW^{+}18]. Multi-Domain [YZY^{+}20]. Multi-Edge [GHK^{+}23]. Multi-Factor [LCZ^{+}23]. Multi-Failure [LSC^{+}21]. Multi-Feature [GCW21]. Multi-Granularity [QHZC18]. Multi-Hop [BVBV17, GJCB18, GZL^{+}17, GVGV17, GLS21, KS19, LAJ20, LYKT21, SK21, SPM^{+}17, TWH24, YXAZ^{+}18, BESW08, CF94, DV09, GSK08, HIM07, JS09, KN05, KS09b, LHB^{+}05, LRL08, NL07, NSCR06, SKE16, TMH97, YS07, ZL16]. Multi-hour [Med95]. multi-lateral [SCY15]. Multi-Layer [KSNR20, LFC18, LJ09]. Multi-Link [CZGKB24, EGR^{+}16, YXL18b]. Multi-Mapping [YXCH21]. multi-match [CW16]. Multi-Modal [RZE^{+}21]. Multi-Mode [EL24]. Multi-node [XSH^{+}15]. Multi-Party [DZ20]. Multi-Path [CLM^{+}18, CLL^{+}19, CKZC19, LZC22, XOYL20, HSH^{+}06, CRS99, Voi07]. Multi-Platform [LLM^{+}24]. Multi-Point [RTNS21]. Multi-Population [QLSW19]. Multi-Protection [LLCJ22]. multi-radio [AAV09]. Multi-Rate [KCH^{+}19]. Multi-Resource [FMPS20, ML23, NLB19, PG18]. multi-ring [COS95]. Multi-Server [DPTP24, KSRW22, LFZ^{+}22]. Multi-Set [QCMY16]. Multi-Source [ZHZ^{+}18, ZQW^{+}23]. Multi-Stage [HAB^{+}22]. Multi-Stakeholder [SFS^{+}22]. multi-striding [ARS16]. multi-system [MRD08]. Multi-Tenant [CBdV^{+}17, CBDCP19, TZX^{+}22, WZX^{+}22]. Multi-Tier [AAAR19, AP17, JTL^{+}17, JTL^{+}18]. Multi-Timescale [MAPZ18]. Multi-Touch [XWL^{+}18]. Multi-Traffic [SFM^{+}18, Med95]. Multi-Transmitter [SLS^{+}23]. Multi-User [AM19, CJLF16, DJK23, KLY^{+}23, LLM^{+}24, NGK19, ZK19, Nee08]. Multi-VPN [BGHS10]. Multi-Wavelength [YWHW24]. Multiaccess [CEC^{+}19]. multiband [HG14]. multibeam [NMR03]. multibid [MT06]. multibit [SK03]. multibuffer [BBFG95]. multicarrier [AZ11, LCZC13, PWK^{+}13]. Multicast [AGKK03, BRY^{+}19, CGC^{+}17, DRW^{+}22, FFZ^{+}18, GMP13, GYLH17, GBG^{+}16, KPP93, Li09, LLLT10, LPW14, LHM02, LDHT02, MBG^{+}03, NKNK17, PLM^{+}16, QY04, QJZ^{+}16, QDD^{+}17, Ram96, SG96, SSR^{+}20, WFH12, ZLW^{+}17, ASW00, AC98, AK14, AADS05, ACKZ14, BCP13, BOY00, BO03, BLBS06, BV96, BAL10, BKTN03, BLS07, BKLM06, BL94, CBD02, CA03, CC95, CV12, CNS04, CH93, CHCH00, CGY00, CTG00, CGK10, CFD06, DS04, DEF^{+}96, DMS06, EAB02, FK07, FY07, FJL^{+}97, GLZC12, GLAMM11, GHK02, GJVZ06, GLSB08, HPR06, HGE04, HSE97, HL05, HL00, Jia98, KR00, KHTK00, KD00, KLS03, Kok10, KHW12, KK12, LNB00, LNB01, LLL06, LWL^{+}11, LLW^{+}12, LZZR12, Lia06, LO02b, LORS06, LG13b, LRM^{+}06, MP08, Mod99, MJ15, NBT98, OS05, PPV04, PSA96, QTWW16, RPGE04, RMM99, RGG11, RK06, RG98]. multicast [RKT02b, SA04, ST05, Ses97, SLS10, SG05, SM00, SV11, STL04, SL07b, SR14, THMK12, VHvdH01, VAS00, WZR08, WCY04, WQC06, WCAB15, XY10a, XFS06, XL11b, YFB02, YZBR14, YJH05, Zap04, ZSSK02, ZS03, ZS04, ZJS^{+}12, ZKO93]. Multicast-based [LDHT02]. Multicasting [ATE21, BAB20, AKS^{+}13, FMMLH06, HLL13, KEW06, LE13, LCZC13, Pan99, PZGLA98, SSM06]. multicasts [WL99]. Multichannel [CLW19, CSL21, GIKK11, NSY20, ZMMG22, AK14, BSYS12, CLSC15, CL16b, HL15, JGLS14, JGS^{+}15, JMI95, KV09, LZ09, LR09, MSH95, MS15, OY13, SKS16, SX10, WXR13, WLR10, WLZ11]. Multichannel-Spatial [ZMMG22]. multiclass [CN10a, JK96, KWC93, KL09]. multiCode [KCB03]. multiCode-CDMA [KCB03]. multicolumn [LSV99]. Multicommodity [GS98]. Multiconfiguration [JM00]. multiconstrained [Yua02]. multicore [GBL12]. multicost [CKV11]. multicriteria [SS10]. multidimensional [CW16, LH03, LS07, Sha94, ST13]. multidomain [DMB94, EST93]. multifiber [BPPP12, LS01]. multifractal [VR13]. Multifunctional [MFT^{+}20]. multigigabit [VS97]. multigranular [CAQ07]. Multigroup [XCL^{+}19, LQCC16]. Multihomed [KGdV^{+}21]. multihoming [AMS^{+}08, AMSS08, IAS06]. Multihop [BSS19, CAZG20, DZH19, DCZG19, QDD^{+}17, RL23, SCL^{+}23, SPLM17, TZL23, TTM23, URZ^{+}14, YK23, YZY^{+}18, AZLB16, BE08, BD07, Bej04, BB95, CFC01, CFZ97, CJZS14, EL11, EOSM10, EML12, GW94, GS97, GPM03, GGM10, GS11, HLW13, HK11, IBM95, JR14, JJS13a, JJS13b, JP09, JP13, JLS09, JL98, JM00, KWE^{+}10, Lab97, LDFK12, LSLL14, LK02, LE12b, LS06c, LHM02, LSS07, LLS10, LB04, LEY14, LG13b, MKS16, NT00, PSK^{+}15, QZZ^{+}13, RL93, RJJ^{+}11, SLS10, SPB16, SH14, TSR14, WB11, WSW12, WWT05, XWWC16, XW11, XLWT12, XE13, YSRL11, ZA95]. multihour [APSKPMGM12]. Multilane [KGdV^{+}21]. multilateral [AJF11]. Multilayer [ANTR17, LZ23, VLZL16, FDG^{+}10, SSV13]. multilayered [AEB02, VAS00]. multilevel [NR98]. multimatch [XLZC14]. multimedia [ALJ99, AW04, ACC^{+}94, CNS04, CCL99, CJJ09, CHH06, FqL98, GZT03, HL05, Jia98, KPP93, cLqL97, LAN97, LS97c, LMS99, RR93, RVR93, SL94, Wan04, WD05, YL97, ZLS96]. multimesh [TH97]. Multimodal [LYH^{+}23]. Multinet [Kim94]. multinetwork [FHSZ13]. Multiobjective [SBDR10, ZCW^{+}22]. multipacket [QAZ12, ZT03]. multiparented [GKT93]. multiparty [CSS06, LZL11]. multipass [KKSS12]. Multipath [BO07a, CZK^{+}21, FMH^{+}21a, JPS^{+}17, PWHL16, PPV17, PCW23, RRS^{+}14, TKXP20, WXH^{+}20, WCW^{+}17, ZLW^{+}20, AFT11, BD07, CER12, CWW^{+}15, GR16, GLSB08, HMM11, IAS06, JRY09, LMR07, NCK15, PM09, RDO^{+}07, SRP^{+}11, SKRK12, VWT^{+}14, ZPCS11, CKS17, KLVL19]. Multipaths [WXJ^{+}17, WSC^{+}23]. multipattern [BBK12]. multiperiod [BWS10]. Multiple [BBD^{+}18, BP19, CCW^{+}17, CLCL23, CZTX23, CCG20, DPTP24, GFW^{+}18, GDWX23, GDJX24, HKC^{+}20, HR14, KP23, KHČ^{+}09, LGS^{+}23, LS17, LYL^{+}22b, LJSB22, LSL^{+}21, MLX18, MVCS16, PPTP21, QLQ^{+}22, RMDJ16, SF23, TJHL21, VKO20, VN20, XZC^{+}19, XCL^{+}18, XZL20, XHY^{+}22, ZND^{+}16, ZCZ^{+}20, ZYY^{+}21, BRISCSP11, BB06, BKTN03, BH06, CU95a, CU95b, CT04b, CFZ97, CY14, DMC06, FUDA03, FP14, FMMLH06, GKT97, HC02, HKLS12, HL03, JVY06, JF04, JL12b, KHTK00, KA03, KK03a, LS94, LS06a, LE06, MSB97, MSSZ12, NMH99, PG94a, QGCL11, Ram08, RCOC03, SCN12, SDV06, SS06, SAKS13, SSM06, SPR08a, SKUB12, TNRP11, Tha04, WS93, WC08, ZBXH13, ZNZT16, ZWYY10]. multiple-access [CFZ97, SKUB12]. multiple-copy [SPR08a]. Multiple-Description [MVCS16]. Multiple-Hop [BP19]. Multiple-Message [ZYY^{+}21]. multiple-path [TNRP11]. multiple-plane [RCOC03]. multiple-primary-user [JL12b]. multiple-set [HKLS12]. Multiple-Unicast [HR14]. multiplexed [GV93, QM99]. multiplexer [BKH^{+}93, BBM93, CDS02, LMS99, SL94, SS98, SSD93, WM96]. multiplexer/scrambler [BKH^{+}93]. multiplexers [BGVC00, HLG94, KS01a]. Multiplexing [CBdV^{+}17, SJ95, SWH19, ZCdV^{+}18, BRM^{+}13, BS15, CP95, CJW11, CW10, FT06, cLqL97, LM95, Lee96, RRG10, Ros96, SD00, SR14]. multiplexors [PS98, SJ95]. Multipoint [CFM^{+}19, MGR02, ZRLD05]. multiprocessor [BG98, OKM94, SKT96]. Multiprotocol [YWZG23]. multiqueue [ZT03]. multiradio [CLSC15, LCG^{+}14, XWWC16]. Multirate [LE13, LWC^{+}14, PLM^{+}16, BD97, CH04, CSN06, FT07, GS97, KBV^{+}13, LDFK12, LY94, LNC98, LC96, LB04, MGR02, MG97b, MMR96, ST05]. Multiresource [JWSLC13]. Multiscale [FAF^{+}17, RRB06, YD07]. multiservice [Guo04, IZC00, MG97b, PL02, RG98, SD00]. Multiset [LGW^{+}17]. multisink [YYZ06]. Multisource [DYX12, TZL23, YYZ06]. multistability [RKA08]. multistage [CHA95, Kim94, SMSM06, YD07, YZ10]. Multistar [TYL94]. multistation [BBL95]. Multistring [WZL^{+}23b]. multiswarm [LZL11]. Multitenant [DXX^{+}23, LZW^{+}15]. Multitier [WWYY18, XRL^{+}22, CJH^{+}11]. Multiuser [DJK22, GB18, TW10, BRM^{+}13, BNS11, GNP^{+}13, LOP97, ORS93a]. Multivariate [PBGMFM22, SJSB22]. multiview [RCFC15]. Multiwavelength [RS98, RIM98]. multiway [LSV99]. mutation [YBX^{+}12]. Mutual [LLM^{+}24, FHH10, RC08, RCS14]. MVNO [LZL^{+}20]. My [WML^{+}18, ZZH^{+}10].

N [BKH^{+}93, RW95]. Nakamoto [SARM24]. Name [GYSZ19, JR21, LLY^{+}22, LNC93, TR98]. name-based [TR98]. Named [JR21, LLZ^{+}17, PRH17, YLF^{+}21]. Names [ABC^{+}16]. Namespaces [JCR21]. Naming [JR22]. nanoscale [LG13b]. Nash [BS09, IW08, KG05, SAM10]. Nation [HS19]. Nation-Wide [HS19]. Nationwide [DYJ^{+}23]. native [AKS96]. native-mode [AKS96]. nature [KL13, LTWW94, RSH^{+}11]. Navi [ZSL^{+}17]. Navigation [LJJ^{+}19, WCW19, ZCF^{+}24, ZSL^{+}17]. NB [PLS^{+}21, YHH^{+}21]. NB-Cache [PLS^{+}21]. NB-IoT [YHH^{+}21]. NCScale [HZLZ22]. nD [HBH93]. nD/D/1 [HBH93]. NDN [DLW^{+}17, LLY^{+}22, QLQ^{+}22]. Near [MBI^{+}17, Nee16b, PPV12, SS10, SUS20, WLL^{+}24, XRL^{+}22, HMNK13, JGS^{+}15, LLY^{+}16, SGD05, XAST12, YGKX10]. Near-Optimal [MBI^{+}17, Nee16b, SUS20, WLL^{+}24, PPV12, SS10, HMNK13, JGS^{+}15, LLY^{+}16, SGD05, YGKX10]. near-zero [XAST12]. Nearly [CCLL17]. need [TMH97]. Needed [LL17a]. Negative [CES22, ZSS^{+}20]. Neighbor [CZC^{+}22, CBZ16, CS17, CLV17, WML^{+}18, ZWZM18, CK11, MWC16, VAGT13, YWLL09].

Neighborhood [CWZY21, RJJ^{+}11, TAB^{+}15, GLG04, LS99, YDS10]. Neighborhood-centric [RJJ^{+}11]. neighboring [Kop96]. neighboring-queue [Kop96]. Neighbors [CBZ16]. nested [FHH10, LNC93]. Nesting [CXW^{+}18]. NET [DGLM16]. NetEgg [YLA^{+}18]. Netfind [SP94]. netflow [LDK12]. NetInventory} [BGJ^{+}04]. NetKernel [NSC^{+}22]. NetQuest [SQZ09]. NetVision [LCU^{+}20]. Network [AZLB16, AAR18, AVS04, ABMT23, ABS^{+}16, AMS22a, AMSB^{+}24, ACA16, BLC21, BCLS17, BHA^{+}20, BCL12, BBEF^{+}21, BDR22, BWES22, BM22, CBdV^{+}17, CBDCP19, CLTM22, CFC^{+}24, CPS17, CJS^{+}20, CCK16, CWHW18, CZL^{+}19, CWZ^{+}23, CGC^{+}24, CGL16, CH21, CHFH20, CL19, CLP^{+}17, CCCC17, CN19, CCG20, CEFS21, CBLVW06, CMY^{+}17, CXW^{+}18, CMY^{+}18, CJL^{+}19, DRMP18, DSYN24, DBW^{+}20, DMDM17, DMT^{+}19, DZL^{+}20, DKM^{+}17, DZL^{+}18, DZ20, DT15, DGLM16, DLLL16, DL04, DLPT06, EFFK18, ER23, EBJM18, EMAL17, EL24, ES05, FGRQ18, FLG^{+}20, FLM^{+}22, FZX^{+}23, FR07, FLTM18, FP14, FMPS20, FX17, FBRL18, FM22, FZQ^{+}22, FAS^{+}23, GCWC17, GXW^{+}19, GTU19, GJWZ16, GG94, GTC^{+}24, GCS06a, HWF^{+}20, HGM^{+}17, HWLL21, HSS^{+}21, HZLZ22, HCL18, HS18, HJG18, ISS22, JR21, JPM^{+}19, JJJ^{+}23, JD22, KRRR17, KSNR20, KSRW22, KSAK18, KAT^{+}22, KS19, KHH^{+}18, KJG18]. Network [KLR^{+}20, KW17, KLLT18, LCH^{+}06, LCK^{+}18, LGY16, LYSZ16, LWL17, LPD^{+}18, LGDC18, LCDW21, LZL^{+}21, LZS^{+}22, LXL^{+}22a, LGDC23, LZB^{+}23, LDS^{+}24, LTN^{+}19, LSCT17, LHW^{+}20, LJHB18, LN19, LL20, LZC20, LDZC20, LYL^{+}22b, LZC22, LDL^{+}22, LAL^{+}24, LSL^{+}18, LW17, LDRS18, LFL^{+}23, LKMK20, Ma16b, MHS^{+}17, MGLH18, MGZ^{+}23, MCMdlO23, MVCS16, MPMC^{+}22, MNZ23, MG97b, MFT^{+}20, MSM16, MRM17, MSTL17, MKG^{+}17, MHR^{+}20, NRB22, NJK^{+}19, NEH^{+}22, NLS19, NSC^{+}22, PDV^{+}24, PLS^{+}21, PPK15, PP17, PLR^{+}19, PLM^{+}16, PBGMFM22, PHC20, QYZX22, QL16a, QCMY16, QLF23, QDD^{+}17, QLQ^{+}22, REM17, RR19a, RRS^{+}14, RGY^{+}22, RAPP22, RAA^{+}24, RS19, RS20, RKPP16, SCN^{+}22, SRI^{+}18, SQ16, SDSY19, SWKA01, SAC^{+}18, SCPB19, SM14, SRB10, SGH^{+}19, SL17, SG17b, SM18, SGVO18, Sob17, SBM^{+}18, SRCDL19, SWL^{+}18, TL22, TY18, TXHL23, TZPZ23, THRW12, TSS21, TWY^{+}20, TGD^{+}20, UN11, VKPI17]. Network [VPC17, VLM16, VLDM17, WBWV16, WSXL16, WQY^{+}17, WMX17, WWC^{+}18, WLTJ19, WCW19, WWMZ20, WWW^{+}20a, WLC^{+}20, WRT^{+}21, WWMZ22, WDL^{+}23, WYZ^{+}24, WGvdS17, WBM^{+}18, WMT^{+}22, WZL^{+}23c, XYA^{+}21, XWH^{+}16, XWW^{+}19, XCW^{+}20a, XCW^{+}20b, XOW^{+}23, XLD^{+}24, XSM22, XL11b, YO17, YSC16, YDY^{+}24, YXL^{+}19, YSJL14, YLF^{+}21, YXZ17, YBQZ18, YXZ19, ZLG^{+}17, ZMH17, ZHZ^{+}18, ZWJ^{+}20, ZXW^{+}21, ZZX^{+}21a, ZSZN21, ZZX^{+}21b, ZLZ21a, ZCPP22, ZLY23, ZEV07a, ZLX^{+}23, ZCdV^{+}18, ZBdV23, ZZT^{+}17, ZLHM22, ZLN^{+}17, ZMWX18, AIN^{+}15, AP93a, Ada98, ACVS10, AS09, AM16, AD14, AD96, AVPG14, AZ09, ACKZ14, AC09, BMVB09, BSSLB95, BM09, BIV01, BGVC00, BSF16, BS97, BPS99, BE06, BLC11, CHML15, CFP^{+}09, CHM^{+}05, CC06, Cha10, CL07, CFS06, CBSK07, CTH10, CJH^{+}11, CLC12, CZM14, CBL15, CHLS07, CMN12, CDH^{+}10, CEFS99, CRB12, CCKK16, CBL06a, CK09, CN09, CM05c, CBL06b, DM95]. network [DMC06, DFMR15, DYH13, DBDJ14, DXT^{+}12, DK98, DLH^{+}14, DFZ06, DLT^{+}15, ESG11, EDBN12, EDM16, ES03, FWL08, FAB12, FK13, FSM14, FSH^{+}13, GKJ12, GLMM04, GGPS96, GCZ98, GLH95, GS98, GR14, GB99, GLLJ16, GCS06b, HAGL16, HBS96, HFC^{+}13, HC07, HSS08, Hou15, HKB14, HK11, IBM95, ILS97, JK15, JMI95, JAW11, JKJ13, JWSH15, JLM15, JS14, Kam10, KRLL11, KL07, KRH^{+}08, KL08, KKSS12, KLZ12, KHG^{+}14, KWS10, KBV^{+}13, KL03, KLS03, KM03, KSV07, KCB03, Kop96, KLO97, KWH11, Kuc14, Kum98, KHČ^{+}09, KCCM16, LE13, LSB06, LRJ08, LBFE09, LLW^{+}09, LK95, LL95, LZSS10, LMMN07, LS06b, LD95, LCH95, LC04a, LBL07, Lia06, LDGL13, LO02a, LZC09, Lin97, LS05a, LÜ14, LJC05, LJ09, LNL^{+}16, LDHT02, LMS04b, MJ01, MM13, MG97a, MMH^{+}15, MA12, MG16, MIB^{+}08, Med95, Mil95, Mil98]. network [MMR96, MW05, ME96, MRHWS14, MBRM96, Nee13, NT00, NS98, OF11, OMA^{+}10, OJRCC02, OR11, OWKS16, PPPW05, PYL99, PT00, PS09, PHL15, PRR06, PFC96, PS93, LZKT99, QL16b, QY12, QS04, RCW15, RGKR10, RJCE06, RW93, RS97a, RZC11, RS12, RVV^{+}15, Ros05, RKT02b, RW96, SKT96, SGR13, SKE16, SYDM09, SJGH10, SLG^{+}16, SJL^{+}16, SS06, ST04, SNXT13, SDW14, SLL15, SL07a, SSM06, SLL^{+}11, SC95, Sob05, SZM08, SQZ09, SV11, SV15, SK97, SKZ03, SCKB09, SZL^{+}14, SAS^{+}16c, TPC09, TK12, THP94, Tas96, Tas99, THDD05, TNML93, Tod94, TMP07, TKI^{+}15, THMK12, Tre11, VW09, VV09, VVP^{+}13, WBEGS05, WS06, WC08, WLC^{+}10, WDCL15, WM16, WSMJ04, WCAB15, WNV13, Wu94, WJK06, XY10a, XB07, XZB08, XL11a, YYZ06, YD04, YWA08, YW11, YSZL15, YL16, YASS15, YKKY08]. network [YR01, YS07, YGC10, YMKC08, YGKX10, YTL12, YCM11, YWZZ16, ZH08b, ZNK^{+}13, ZCX^{+}15, ZBA16, ZCB09, ZWYY10, ZGS10, ZKO93, Hu93]. Network-Aware [SGVO18, WRT^{+}21]. Network-Based [QYZX22]. network-coded [ACKZ14, THMK12]. Network-coding [XL11b].

Network-Coding-Based [SQ16, KWH11]. network-distributed [BM09]. network-edge [WBEGS05]. network-failure [LJ09]. Network-Flow [SM18]. network-internal [LDHT02]. Network-Layer [GTU19, LTN^{+}19, AZLB16, AC09]. Network-Level [DZL^{+}18, BCL12, WLC^{+}10]. network-on-chip [AIN^{+}15]. network-state [SZM08]. Network-Wide [BBEF^{+}21, TXHL23, WQY^{+}17, ZZX^{+}21b, FR07, THRW12, BSF16, GCS06b, LLW^{+}09, Tas96]. Network-Wise [TZPZ23]. Networked [AHEK24, CCZZ17, GSKN18, JL12a, VLDM17, ZWC^{+}24, CT01, DPR06]. Networking [ANTR17, ACDP17, BWG^{+}20, BBCD14, CPKL17, CGYZ16, CYH^{+}18, GTU19, GSM^{+}17, KSAK18, LLZ^{+}17, LLS^{+}23, LCH22, PGMR18, PRH17, QSW24, SM17, SM19, SDM20, SS16, SBC^{+}17, WBWV16, WBY^{+}17, XHC^{+}18, XYT^{+}21, XHY^{+}22, YLF^{+}21, ZXW^{+}20a, CCE^{+}06a, CCE^{+}06b, CPGZ15, HS06a, IGE^{+}03, LCL^{+}12b, LCG^{+}14, MHRR12, SRR08, TLS^{+}12, VT12, YL98]. Networks [AB21, ACC^{+}14, AMCD19, AdSD16, AGCFV18, ASKL18, AGBS23, ADT22, ER20, AY20, ALPK21, AdVS20, APSG14, AP17, AHP21, AGM^{+}17, AAF^{+}16, AMG^{+}17, AM19, BCO17, BTP^{+}17, BJK20, BVL^{+}19, BSS19, BTD^{+}17, BAB20, BK17, BTK^{+}17, BP19, BPST18, BBR19, BCD19, BBZ^{+}18, BMY^{+}17, BVBV17, CBDCP19, ÇTD22, CLK^{+}24, CLGSS17, CLWZ17, CPKL17, CCE^{+}17, CWL^{+}21, CE19, CP17, CAZG20, CLW16, CCLL17, CLS^{+}18, CLL^{+}18, CLW19, CH20, CMP16, CHO^{+}19, CWAO21, CWH^{+}16, CGC^{+}17, CS17, CLV17, CZX^{+}17, CZX18, CGC^{+}18, CLM^{+}18, CZL^{+}19, CDGZ20, CGZL20, CLZ^{+}20, CBHS20, CLHY22, CZTX23, CXL^{+}24, CNG^{+}16, CHW^{+}20, CWZY21, CSSG23, CAD^{+}17, CRS18, CCMW19, CEC^{+}19, CSN^{+}23, CMW^{+}20, CMP^{+}14, CFM^{+}19, CDW19, DAFZ^{+}18, DHK16, DRCM^{+}17, DPSA21, DPG^{+}24, DJS^{+}17, DZH19, DF20, DZL^{+}18, DYJ20, DZ20, DTN^{+}21, DYW^{+}16, DGW^{+}17, DCZG19, DGC^{+}20, DME23, DQYG23]. Networks [ENT^{+}24, FZ16, FXHY21, FMH^{+}21a, FKCA18, FSGH17, FLMS18, FFX^{+}17, FWK17, FMK^{+}18, FFZ^{+}18, FZW^{+}20, GHRH18, GJCB18, Gan20, GDC^{+}17, GZL^{+}17, GJD18, GKB^{+}16, GYLH17, GB18, GVGV17, GCX^{+}17, GLL^{+}18, GKCR21, GLS21, GSM16, GZJ^{+}18, GXL^{+}21, GCW21, GSL^{+}24, HKS16, HLZ^{+}21, HW22, HLZY23, HNW17, HAG19, HGM^{+}17, HVT18, HAB^{+}22, HWM^{+}24, HCL^{+}17, HZC^{+}19, HZB^{+}22, HR14, HHA17, HSM^{+}20, HMM^{+}20, HWJZ21, HLL^{+}21, HSM^{+}21, HSL20, HK14, HRM22, IYYI18, IKS17, JV17, JWW^{+}23, JLS^{+}17, JTL^{+}17, JTL^{+}18, JM17, KSUB^{+}18, KSM19, KYM22, KK16b, KE21, KPK^{+}16, KWH^{+}17, KIW^{+}17, KKH^{+}22, KIL24, KW19, KKS19, KSK17, KLKT16, KLP16, KK21, KLE16, LFC18, Le 18, LCK^{+}18, LMSR19, LMSS24, LWL17, LBP^{+}17, LXW^{+}17, LLX^{+}17, LWQ^{+}18, LWP^{+}19, LGDC19, LSDT19, LBZ^{+}20, LI21, LS22, LLY^{+}22, LLCJ22, LLS^{+}23, LXX^{+}17, LHW19, LBGL20, LWC^{+}23, LHY^{+}23, LJL^{+}16, LLL^{+}16, LDY^{+}16, LCZH17, LL17a, LWK^{+}18]. Networks [LFY^{+}19, LHZ^{+}19, LAJ20, LXZ^{+}21, LLM23, LAL^{+}24, LNLM24, LW22, LSHZ16, LGCG^{+}21, LSSK17, LYKT21, LSL17, LCU^{+}20, LFXY23, LZ23, LFF^{+}19, MLX18, MLS^{+}23, MYMY17, MMIY20, MLB21, MCC^{+}19, MMT14, MFR^{+}20, MBL19, MWW^{+}21, MAPZ18, MSM16, MRJ20, ML22a, ML22b, MGK20, MKS17, MJ17, MMG22, MMP17, NJK^{+}19, NSY20, NSY23, NK20, NLRS21, NDN^{+}18, NGRF19, NZW24, NSP^{+}16, OJSY16, ODJ23, PD16a, PC19, PWL^{+}24, PBV17, PP17, PL17, PDI20, PLM19, PJM^{+}19, PIST19, PLT^{+}20, QFH^{+}18, QJZ^{+}16, QZX^{+}17, QLSW19, RCR^{+}18, RRS23, RBPS21, RZS14, RTNS21, EGKM16, RS21, RZE^{+}21, RL23, SARM24, SNZ^{+}23, SRCT23, SK11, SCN^{+}24, SRMB^{+}23, SPQZ20, SS17, SQS20, SFM^{+}18, SLS^{+}23, SMD20, SKE19, SG17a, SdVK16, SRC^{+}20, SYZP19, SLD^{+}23, SCC^{+}17, SSY19, SE21, SCL^{+}23, SH23, SMC^{+}20, SLWW19, SSK^{+}17, SK21, SPLM17, SLH^{+}19, SSM20]. Networks [SBTH19, SRB^{+}20, SGL^{+}22, TE16, TKM20a, TKM20b, TJL^{+}19, TZCB23, TML22, TZL23, TZL^{+}24, TWTD17, TTM22, TTM23, TG21, TG23, TTCT19, TS14, URZ^{+}14, Van17, VVC^{+}17, VRR24, VPC17, WG16, WXN^{+}17, WLX^{+}17, WVZ17, WT17, WMP^{+}18, WDR^{+}20, WZLM22, WCL^{+}22, WMO^{+}23, WSZL20, WLC16, WCC14, WZZC17, WCZZ17, WFY^{+}18, WML^{+}18, WGZC21, XCS^{+}18, XWJ22, XTHL21, XS21, XXZ^{+}23, XXZ^{+}22b, XRL^{+}22, XXZ^{+}22a, XSZ^{+}22, XWW^{+}23, XHZ^{+}19, XCV^{+}20, YM16, YXC^{+}18, YSC18, YZY^{+}20, YCC^{+}21a, YLYL17, YXAZ^{+}18, YZL^{+}18, YY20, YXCH21, YYB^{+}22, YCZ^{+}23, YLD^{+}23, YK23, YDLM20, YLL21, YZGC23, YLK^{+}17, YXY^{+}18, YZY^{+}18, YZY^{+}21, YZZ^{+}21, YYL23, YLWH20, YKB^{+}23, ZBC^{+}22, ZZL^{+}24, ZCH^{+}24, ZFW14, ZWL^{+}16, ZV16, ZND^{+}16, ZYZ16, ZZ17, ZFW^{+}17a, ZWGC17, ZYL^{+}17, ZCZC17, ZM18, ZZH19, ZLRC20, ZGYB20, ZLX^{+}21, ZZX^{+}21b, ZSL^{+}21, ZY21, ZRP^{+}22, ZLL^{+}23a, ZMLR23, ZZW^{+}23, ZLL24, ZHCC24, ZZZ^{+}24, ZDB^{+}17, Zha17, ZJL^{+}19, ZMD^{+}20]. Networks [ZW22, ZMMG22, ZQ23, ZQW^{+}23, ZWZ^{+}24, ZDZ^{+}24, ZHT^{+}19, ZGLC20, ZJWY17, ZLTX17, ZWZM18, ZXW^{+}20b, ZLW^{+}16b, ZFW^{+}17b, ZLW^{+}17, ZYY^{+}21, ZHWH21, vDJJ^{+}22, AHK08, AS94, AC16, AS14, AK01, AA93, AACD^{+}96, AK00, AKA10, AEG^{+}13, AC98, AJV06, AK09, ARK09, ARK11, AA04, AA05, AHL96, ALJ99, AJDH01, AMP01, AEB02, AW04, AAM05, AMKY99, AA99, AGLM10, AGL16, Ali06, AK14, AK15, AS07a, AS07b, AB05, AWKN16, AD11, AABD13, ANSX13, ACCF12, AADS05, AZR97, ATB^{+}10, And04, AZ03, AL98, APSKPMGM12, AS96, AWFT15, ABJ^{+}13, ALMR14, AAV09, AJ06, AST11, BTH11, BCP13, BBG11, BE08, BD07, BO00, BCP00, BPPP12, BCGC15, BTC05, BM00, BO07b, BBFG95, BY06, BJ15, BBLV06a, BBLV06b, BCB99, BSH^{+}11, BV10, BV96, BF01, Bej04, BB06, BR06, BS97, BM93]. networks [BLCT97, BTC01, BK00, Ber00, BB95, BNJR12, BNJ16, BT93, BM97, BV05b, BI00, BLT02, BSS14, BSYS12, BD97, BP96, BC01b, Bor05, BMS14a, BCC07, BLB10, BS15, BL04, BLRC05, BDWS12, BDHR10, BGJ^{+}04, BM08, BZM08, BESW08, BWS10, BCMR04, CE09, CKS16, CLP12, CN10a, ÇY07, CAO11, CAQ07, CKL16, CZF^{+}16, CFM13, CDM13, CFPP96, CGM04, ÇM15, CFG08, CM05a, CT04a, CH04, CV12, CRL96, CB11, CCL11, CL15, CHA95, CT04b, CM05b, CRD08, CLC12, CGW^{+}12, CL12, CZM14, CSS^{+}14, CSSJ14, CS14, CC96, CCL99, CZ06, CYK09, CZC^{+}13, CLSC15, CTG00, CH11, CFC01, CE08, CF94, CFZ97, CZFF98, CPR99, CS98, CCA96, CSC04, CJZS14, CLW95, CMV10, CL05, CL09b, CW10, CKR93, CAL09, CGS93, CMGL11, CD96, CGEN98, CSEZ93, COS95, CJ97, CR98]. networks [CK00, CGK10, CN10b, CK11, CNP13, CG15a, CG15b, CL16b, Con11, CLG^{+}00a, CLG00b, DM03, DPBT11, DS99, DS04, DT93, DHSS14, DJ14, DSTM12, DMB94, DYX12, DZNT14, DV09, DLL^{+}11, DBT05, DFT06, DZH03, DRR98, DGG^{+}02, DM96, DJM97, EAB01, EAB02, EMPS06, EL11, ES96, EH11, ECN09, EM93, EFK07, ES07, EOSM10, EM09, EML12, FK07, FCL97, Fan05, FGK10, FRC98, FGL^{+}01, FC99, cFKSS99, FMMLH06, FEC13, FMSM^{+}11, FJ95, FT07, FM06, FML09, FqL98, FCT03, GP96a, GMP13, GMP08, GTS^{+}09, GW94, GLZC12, GDC^{+}16, GSK08, GGL09b, GGL09a, GV93, GH04, GNP^{+}13, GM03, GGC93, GGFS02, GT06, GCZ96, GSKR99, GS10a, GS13, GM00, GIKK11, GGH11, GP94, GKT93, GL93, GB10, GP96b, GZCX16, GRB09, GGK99, GEHM02, GVC97, GSA15, GT00, GS97]. networks [GT10, GT02, Gro99, GMYP16, GMS16, GO99, GPM03, GAA08, Guo04, GL10, GZDG06, GLS09, GS10b, GGM10, GS11, GMSK09, HS06a, HIM07, HLL13, HA16, HRCW08, HTAZ16, HKL06, HH10a, HBU95, HA96, HA97, HM04, HTC04, HSS08, HSPH09, HH10b, HKCL13, HS06b, HY08, HL98b, HL05, HMvdLM07, HMM11, HW12, HLW13, HN13, HK96, HKT95, HL00, HLL06, HK11, IKDD15, IW08, IK07, IMG98, IK09, JR14, JDSZ97, JS11, JMS07, JJS13a, JC13, JJS13b, JGLS14, JGS^{+}15, Jia98, JZC11, JSuRKH03, JPH08, JW10, JYT^{+}15, JLX^{+}16, JK05, JJ08, JP09, JL12a, JP13, JS09, JLS09, JS14, JBR16, JL98, JM00, JF04, KV96, KL12, KV98, KJF^{+}00, Kam10, KK16a, KWJY16, KWCR10, KKJ06, KKEE13, KA98, KK07, KIR08, KGL03, KS10, KE16, KT11, KS95, KA03, KCTI08, KNP05, KK93, KDHK15, KEW06]. networks [KSA12, KL95, Kim98, KqL99, KK00, KD00, KS09a, KLSS10, KWS^{+}11, KLS11a, KS11, KDYV12, KRKH10, KR05, KG10, KN05, KMZR12, KES13, KMHS09, KWE^{+}10, KNK^{+}14, KG05, KEY99, KHW12, KT06, KT07, Kri14, KS01b, Kuc14, KWZ08, KIR06, KRSY02, KLT15, KL09, KS98, KS09b, KT08, KGGZ11, KV09, LH07, Lab97, LBRA05, LM97, LMR99, LTS10, LSV01, LBB08, cLqL97, LDFK12, LSLL14, LK16a, LPKF10, LOP97, LMS12, LNS11, LS93c, LH95, LLD96, LT02, LS03a, LMS06, LMR07, LBHO07, LML10, LML11, LKC11, LLM11a, LKC^{+}13, LLM14, LCS12, LGC16, LMP08, LYRL07, LYWL08, LYS93, LA95c, LAN97, LMSKZ99, LZ06, LCH95, LSC99a, LSC99b, LKL00, LM01, LS01, LWL04, LHB^{+}05, LH05, LLL06, LLS07, LQXX07, LTZ08, LZF09, Li09, LGS09, LLLT10, LL10, LLM11b, LBX11, LYC11, LEYS11, LPF12, LE12a, LZL12, LE12b, LG13a, LSZW13, LZL^{+}14, LPW14]. networks [LZES14, LH14, LXY^{+}14, LZW^{+}15, LM15, LLE15b, LHZ^{+}16, LNA07, LLNC09, LH03, LS06c, Lia06, LWT^{+}15, LHM02, LO99, LWF96, Lie97, LNC98, LJA14, Lin93, LSMS06, LS06d, LS06e, LSS07, LR09, LLS10, LW13, LC96, LB04, LLL10, LFZS11, LK13, LCZC13, LEY14, LWR15, LNL^{+}16, LHC^{+}16, LWR^{+}16, LPCVC13, LO98, LG13b, LLS09, LV93, LBS99, LXC05, LLW^{+}14, LFL14, LTP10, LC94b, LNC04, LRM^{+}06, LKZ^{+}04, LYL07, LP07, LLY09, LRG10, LH10, MLLY06, MVRZ09, MCLG07, MBL10, MGCK15, MOR13, MRM99, MPS01, MSS^{+}12, MD04, MWQ^{+}10, MKS16, MQ05, MR98, MBLN93, MFL^{+}04, MGG^{+}05, MPF^{+}15, MSB97, MOZ05, MOY00, MBF^{+}02, MGR02, Med95, MG97b, MSP^{+}07, MWC16, MDMM09, MS95, MSSZ12, MJ13, MHSC95, MV14, MRD08, MHXT10, MEWP13, Mod99, MR96, MMS01, MS15, MA98, MK96, MK98, MAS09, NOF14, NSS96, NM06, NS03, NML08]. networks [Nee09, NL07, NPY07, NCT14, NSW11, NTS12, NLB15, NSCR06, ODC^{+}16, ORS93a, ORS93b, Ord99, OSZ^{+}06, OY13, OB03, OÇ10, PWDL05, PG95, Pan99, PM96, PSK^{+}15, PG93, PG94a, PNRMC13, PS05, PLR15, PL02, PLS07, PEA09, PA12, PCV08, PG94b, PT94, PPSV13, PBKG11, PRR06, PPV12, PS94, PGV16, PJ13, PCL15, PK01, PES^{+}12, QZZ^{+}13, QM99, QY04, QGCL11, QSS^{+}15, RrBG94, RGKR10, RP13, RDO^{+}07, RSM09, RGG11, RRG10, RM02, RCGS09, RR93, RL93, Ram96, RS04, RM08, RS08, Ram08, RS95a, RS97a, RS98, RRK96, RLKT98, RVS^{+}02, RJJ^{+}11, RSS09, RK06, RG98, RLP06, RB95, RD11a, RSR10, RWA^{+}08, RS07, Ros96, RZVZ06, RZZ06, RL94, RCGT06, RA95, RS97b, RKNS10, RYS12, SMG05a, SMG06, SMT98, SSV13, SKY10, SK10a, SK12a, SLP07]. networks [SPH04, SAS16a, SMGP15, SEK15, ST05, SZG09, SG13, SKR^{+}09, SAS^{+}16b, SJ12, SM14, SM16, SW04, SRB10, SLS10, SMS07, SM08, ST09, SSHK11, SKRK12, Sha94, SYP01, SYR05, She95, SH12, SCR08, SCY98, SS09, SS10, SL12, SK10b, SK12b, SBP03, SM00, SLH^{+}06, dSeSGM95, SMM11, SSAK12, SAKS13, SKS16, SM05, SMSM06, SR94, SEMO09, SR01, SKCW10, SSFM08, SH07, SZM08, STC12, SPR08b, SPR08a, SS04a, SGD05, SKUB12, SPB16, SX10, SB07, Ste08, SS04b, SV98a, SSZ03, SRD^{+}09, SDW00, SD00, SA05, SAS96, SAS99, SSA08, STL04, SD15b, SS07, SR14, SSR^{+}11, SN15, Świ96, TW10, TKN06, TXL^{+}12, TK12, TCS13, Tan16, TWL06, TX08, TAG08, TWHR11, THBR14, TSR14, TJ95, TSGR08, TKZ94, Tia05, TS08, TH97, TWL05, TEML09, TCPV13]. networks [TMH97, TdWC^{+}94, TAH99, TYLH09, fTL06, TLP^{+}16, TS09, VWT^{+}14, VJV14, VOK09, VG04, VS97, VAGT13, VRK09, VCM04, VA06, VA09, WZR08, WCY04, WY06, WXBZ04, Wan04, WLLD05, WQC06, WSC08, WYL09, WLL^{+}11, WB11, WA11, WVG12, WSW12, WKA^{+}13, WHM^{+}13, WLWL13, WCH95, WLS97, WCY00, WLL01, WK13, WKZL96, WM95, WWL02, WS05, WWT05, WS08, WFS09, WMFS10, WTS^{+}13, WFGZ13, WHTC15, XY10b, XTMM11, XL99, XXBC14, XK06a, XSHS12, XSH^{+}15, XWWC16, Xin07, XC08, XM99, XW11, XLWT12, XL11b, XK06b, XE13, XGF^{+}14, YMR00, YBG^{+}12, YD04, YD07, YLL10, YXF^{+}13, YKZ^{+}13, YJZW15, YJ15, YWLL09, YCV15, YS93, YHE04, YAA09, YLH15, YOY97, YGC10, YSRL11, YZBR14, YKGF08, YG10, YZ10, YNDM09, YM05, YBX^{+}10, YC12, ZOM03, ZA95, ZWDS00, ZSSK02, ZZZ^{+}07, ZY07b, ZH08a, ZKL11, ZSFZ11, ZA11, ZNK^{+}13, ZCJ^{+}13, ZYL^{+}14, ZZZ^{+}14, ZCW15, ZHC16, ZNZT16]. networks [ZT03, ZG08, ZR09, ZTS11, ZLC12, ZXH^{+}13, ZW14, ZL16, ZWTC16, ZF96, ZW10, ZPCS11, ZRP00, ZZZM03, ZRK06, ZJ12, ZZHZ13, ZM04, dFV02, dOSAU04, DKL01, vRWZ09]. Neumann [CLY06, YZLH17]. Neural [LBZ^{+}20, LLY^{+}22, RRS23, SRMB^{+}23, YXL^{+}19, YKB^{+}23, ZCH^{+}24, ZXW^{+}21, ZZZ^{+}24, CCL99]. Neutral [LSSK17, Ma16a]. Neutrality [LSK20, MM13]. Neutralizing [SKG^{+}18]. Neutrino [AAT^{+}23]. Never [CBZ16]. NewReno [PMW10]. Newton [SBNRS14]. Next [DRW^{+}22, SQS20, AMI^{+}07, ALMR14, DDPP00, DHSS14, MD04, THDD05, VA07]. Next-Generation [DRW^{+}22, SQS20, AMI^{+}07, ALMR14, DDPP00, DHSS14, MD04, THDD05]. NFA [ARS16]. NFA-based [ARS16]. NFV [BSM21, BVL^{+}19, CLX^{+}22, JWL^{+}18, KZH^{+}20, LHW^{+}20, LXZ^{+}21, MRJ20, WTJR22, XRL^{+}22]. NFV-Based [CLX^{+}22]. NFV-Enabled [LXZ^{+}21]. NFVnice [KZH^{+}20]. ng [MLJ^{+}22]. NIRA [YCB07]. No [CW19, CN16, QCMY16, SPGM13, VKO17, KS01b, MSS02, RK06, TT09]. NoC [EPS21, FMCS20]. NoC-Based [FMCS20]. Node [CS17, EE18, GJWZ16, KA20, MHS^{+}17, NTR18, PWLC23, SRB^{+}20, TT17, TXW^{+}19, YSC16, YWLL09, YY98, ZW14, ZM04, AGLM10, BM93, BKLS08, CDM13, CRB12, DT15, FM06, GGPS96, Ili00, JK15, JRY09, KKJ06, KRKH10, LYRL07, LG13a, MHXT10, NSS96, PM09, PG93, PG94a, LZKT99, SSHK11, TT09, THP94, WL07, XSH^{+}15, ZSCJ14, ZWYY10, YRB^{+}18]. Node- [YRB^{+}18]. Node-Based [EE18, PM09]. Node-Constrained [TXW^{+}19]. Nodes [NK20, SJ21, SGL^{+}22, VTBK21, CR14, GGL09b, GGL09a, GV06, IW08, KDHK15, LC03, MSB97, MEWP13, OWKS16, QY12, RPZ^{+}09, SNXT13, SK13, VJV14]. Noise [JHLW24, XCS^{+}18]. Noisy [GHZ20a, QJCR20, RFGL17, AC16, CLM^{+}16]. NOMA [GSL^{+}24, HRM22]. NOMA-Based [GSL^{+}24]. Non [APSKPMGM12, BT23, CG21, CW19, DME23, GA24, HKV23, HKB14, HSM^{+}21, LMS05b, LMSS24, LYW^{+}21, LSSK17, LSK20, ML22a, PLS^{+}21, SF23, TWL22, Van19, XZL20, ZRH18, BB96, CS00, KG16, LC03, MLT12, SYP01, YLH15]. Non- [APSKPMGM12, LYW^{+}21]. Non-Benign [LMSS24]. Non-blind [HKB14]. Non-Blocking [PLS^{+}21, YLH15]. non-bus-oriented [BB96]. Non-convex [LMS05b]. Non-Duplicate [HSM^{+}21]. Non-Exponential [HKV23, Van19]. non-FIFO [LC03]. non-homogeneous [KG16]. Non-Ideal [ML22a]. Non-Intrusive [CW19]. Non-Linear [DME23]. Non-Monetary [ZRH18]. Non-Neutral [LSSK17]. Non-Neutrality [LSK20]. non-optical [SYP01]. Non-Optimal [SF23]. Non-Orthogonal [XZL20]. non-prefix [MLT12]. non-real-time [CS00]. Non-Saturated [BT23]. Non-Stationary [GA24, TWL22]. Non-Uniform [CG21]. Nonblocking [KA20, MHSC95, YDLM20, CTH10, HL00, JPH08, LA95b, LNC98, LC96, MSH95, NPQ06, NWP09, NMH99, PB93, ZGS10]. Nonconcave [BMS14a]. Nonconvex [VL16]. noncooperative [BPPP12, KAEAS14, LO99, WHTC15, ZWTC16, ZW10]. nonequivalent [WXC16]. nonexclusive [SL14]. noninterruptive [HLL06]. Nonlinear [RAL04, CGMS13, PILR05, ZEV07b]. Nonnegative [CLY^{+}17]. nonovertaking [CCL09]. nonreal [HLG94]. nonreal-time [HLG94]. nonregulatory [MM13]. NonResilience [CJL^{+}19]. nonresponsive [ZDR04]. nonsaturated [MDL07]. nonstarvation [LZC09]. nonstationary [AZ06a, KZDM07, VR13]. Nonuniform [DD24, BBFG95, LA95b, NT00, WH97]. nonuniformly [MPL09]. nonzero [ZA11]. Norm [WGvdS17]. Normal [ACZP21, PPTP21, AM16]. Normalized [CFM^{+}09, Kuc14]. North [MHRR12]. note [ZCW15]. Notification [EPB14, GKPS06, JRL15, LAJS07, SCN12]. Novel [ÇTD22, CHFH20, GZL^{+}17, GJWZ16, HLHL22, LYH^{+}23, MFR^{+}20, TT17, TWH24, WWT05, WFY^{+}18, ZLTX17, AEB02, BO07b, BSS11b, CLC^{+}01, GIL^{+}15, HXLZ11, JMS08, KCB03, LSC99a, LMS04b, RSS09, TWHR11, WY95, YJ15, ZY07a, ZZZM03]. NP [CAP15, CBLVW06]. NP-completeness [CBLVW06]. NP-hard [CAP15]. NReee [HLHL22]. NSB [SCPB19]. NUM [FM22]. number [CL04, CTG00, GR14, GO02, JJSS04, LPIH11, NPY07, SZ07]. Numerical [SXEZ21, ZMW^{+}22]. NVS [KMZR12]. NWL [THRW12]. NWL-UFL [THRW12].

O [GNK^{+}21, qLP97]. OBEX [BH06].

Obfuscation [CAS^{+}20, HWLL21, RVV^{+}15]. Object [SCW^{+}21, YWW^{+}23, HQW^{+}16]. Objective [GXW^{+}19]. Objective-Driven [GXW^{+}19]. Objectives [CSN^{+}23, NLNL16, MR02, WC08]. Objects [ZZL^{+}21]. Oblivious [BBEF^{+}21, CLV17, CRS18, CSN^{+}23, KLOS09, LBZ^{+}20, TW23, YNDM09, BDS07, KLOS11, KLS11b, LSV01]. OBS [BV10, RZVZ06]. Observable [LLM23, LNLM24, REM17, VW09]. observation [DG01]. Observations [FMK^{+}18, HLZY23, KP23, LHW^{+}20, GKJ12, SMC02]. Observed [KLPS06, OPW^{+}10]. observers [WS93]. Observing [SRC^{+}20, ZLB17]. OBSS PD [LR22]. OBSS PD-Based [LR22]. Obstacle [DF20]. Obtaining [HFKC12].

OBUs [SMEH20]. OC [BKH^{+}93]. OC-48 [BKH^{+}93]. Occupancy [GYSPR14, SSG18]. OCO [WDL^{+}23]. Octopus [AZP^{+}23]. OFDM [KT06, LCW^{+}24, ODJ23, PWK^{+}13, WHZJ20]. OFDM-based [KT06]. OFDMA [AYS^{+}13, ASKR16, CJJ09, CLL^{+}14, CK10b, CG15b, EF08, GGM10, SR14, ZZW^{+}24]. Off [BT23, CQW^{+}18, FLS^{+}22, GSH^{+}22, QM99, Van17, BGK^{+}16, LCY^{+}19, SLD^{+}22, BBM93, MH02]. off-duty [BGK^{+}16]. Off-line [QM99]. Off-Path [CQW^{+}18, FLS^{+}22]. offered [GP94, PG94b]. Offering [JWSH15, KA03]. Offline [CMV10, NST^{+}16]. Offload [LSL17, PDV^{+}24]. Offloading [BSRdA16, BLM^{+}17, CJLF16, CZX18, CSR^{+}17, DPTP24, GZJ^{+}18, HFF^{+}24, JD20, MS17, QLY23, TJD23, WCL^{+}22, WYL24, XWW^{+}23, YCC^{+}21a, YZL^{+}18, ZHGF19, DRJ^{+}14, IGHT15, JWSH15, LLY^{+}13]. offs [FLC09, LA95b, SMS07, WKZL96]. Offset [HPP^{+}23, GCS06a]. Offsets [GPBL24]. OLAA [GSPV^{+}18]. Old [DLY^{+}22]. oligopoly [GS16]. omega [SYP01]. On-call [HKT95]. On-Demand [GXW^{+}19, HH18, KLK^{+}20, LCU^{+}20, NST^{+}16, ZZLW16, AF99, DYX12, MEVSS03, PWMC12, ZEV07a, ZEV07b]. on-duty [BGK^{+}16]. On-Line [XLW^{+}18, SMG06, ZY07b, BCN02, cFCcFW05, YKKY08, YF05]. On-Off [GSH^{+}22, BBM93, MH02]. On-Site [CZP18]. On-the-Fly [ZBZ^{+}19]. OnDisc [HTL^{+}19]. One [BHC^{+}21, FAS^{+}23, GCS06b, KIL24, LCZ^{+}23, NK20, OBS17, XSSK08, XWY^{+}18, YJL^{+}19, AS07a, CR99, FHH10, HLHD^{+}04, IW08, JK15, KM10, PEA09, XWG14, ZBXH13]. one- [CR99]. One-Bit [KIL24]. One-Dimensional [NK20, AS07a]. One-Hop [OBS17, PEA09]. one-mode [XWG14]. One-Pass [FAS^{+}23]. one-sender-multiple-receiver [ZBXH13]. one-shot [IW08, JK15]. one-time [FHH10]. one-to-many [HLHD^{+}04]. One-way [GCS06b]. Online [AP17, BSSU18, BBZ^{+}18, CKA16, CN19, DAFZ^{+}18, DBL^{+}19, DHHD18, DZH19, DZ20, FSSC18, GA24, GBMV21, GLS21, HTL^{+}19, HTW^{+}22, HKLM17, HWF^{+}20, HNP23, JWL^{+}18, JTL^{+}17, JTL^{+}18, KTvdSK18, KLS03, KLMW11, LCH22, LGS^{+}23, LL17b, LWW^{+}19b, LYW^{+}21, LW20, MHB^{+}21, MSS16, NLRS21, PDI20, PMAN16, RTLC17, SAMB18, SZ22, SZW^{+}16, SLF21, SZWW22, SKA^{+}18, SUS20, TJL^{+}19, TJHL21, TW23, TJD23, WLX^{+}17, WDR^{+}20, WHYC23, WMO^{+}23, WLD^{+}24, WCZZ17, XXZ^{+}23, XLL^{+}20, ZHW^{+}17, ZXW^{+}21, ZLZ21a, ZTH^{+}23, ZXW^{+}19, ZFLC18, BBMELH08, BLEM^{+}12, CFS^{+}10, CKV11, HZL16, JLX^{+}16, LZL12, MGK12, MKS16, PES^{+}12, XL11a, YKR11, ZLM16]. Only [SACH21]. Onto [BSRdA16]. ONU [NM06]. Open [KPK^{+}16, WLL^{+}16a, KSG11, TEML09, PJMM22, RRS23]. OpenFlow [CMFA14, KLC^{+}18, MLJ^{+}22]. OpenFlow-Compliant [KLC^{+}18]. OpenFunction [TML^{+}18]. Operation [CZGKB24, DYJ^{+}23, HHA17, ODJ23, BBL95, LC96]. Operational [CMP^{+}14, LHL^{+}23b, WXC^{+}24, FGL^{+}01, MIB^{+}08, NBTD07]. Operator [NJK^{+}19]. Operator-Defined [NJK^{+}19]. Opinions [KKS19]. opportunism [PD07]. Opportunistic [BCL^{+}09, BNJ16, BDR22, CS17, CW10, CPS13, HW22, JL12b, KW17, LDK12, LMODF18, LL18, LSL17, Nee19, Nee22, SKK07, SS16, WMS09, WSZL20, XWW^{+}23, ZMMG22, BGSSW13, BNJR12, CL09a, CB11, GSRS^{+}15, KYY^{+}12, KWH11, LS06b, LHZ^{+}16, LYS11, LHC^{+}16, Nee08, RGKR10, RHQZ13, SBD11, SK12b, TZP^{+}10]. opportunities [CKS16, GMLP10]. Opportunity [ZKL11, ZLSK15]. Optical [AdSD16, AAF^{+}16, BCO17, BBG^{+}10, CCE^{+}17, CWAO21, CWM^{+}17, CZTX23, DRW^{+}22, Dat17, ENT^{+}24, KW19, LS22, LZ23, NEH^{+}22, NPY07, RS21, SRCT23, TZX^{+}21, TWN^{+}20, WJ17, WZZC17, YZY^{+}20, YDY^{+}24, YYB^{+}22, YWHW24, ZYZ16, ZWZ^{+}24, ZHT^{+}19, ZLW^{+}17, ARK09, ARK11, AA99, Ali06, AZ09, APSKPMGM12, AJ06, BTH11, BPPP12, BM00, BSH^{+}11, BV10, BC01b, BL04, BLRC05, BM08, CAQ07, CJ14, CCL06, CV12, CCL09, CTH10, CSS^{+}14, CFC01, CCA96, CSC04, CJ07, CCF04, CL05, CLG00b, DS99, DMK05, DBDJ14, DHSS14, FJ07, FMSM^{+}11, GSKR99, HD07, JSuRKH03, JM00, KA98, KT11, KS01b, LBRA05, LSV01, LA95c, LQXX07, LYC11, LS06c, LHM02, LXC05, MBLN93, MBF^{+}02, MSSZ12, MMS01, MA98, MBRM96, NM06, NS03, OSZ^{+}06, OB03, PG95, Pan99, PEA09, QM99, RSM09, RIM98, RM02, RS04, RS08]. optical [Ram08, RS95a, RS97a, RS98, RZZ06, SMG05a, SK12a, SYDM09, SJ12, SYP01, SYR05, SEMO09, SKCW10, SS04a, SAS96, TWHR11, THBR14, TCPV13, TS09, WQC06, WS05, WYHL09, XTMM11, XL99, Xin07, XGF^{+}14, ZA11, ZJ12]. Optically [SS17]. optima [KLO97]. Optimal [ATE23, AAG^{+}16, AS96, AMS22b, AHEK24, BCP13, BPA20, BJK20, BFMF01, BPST18, CZF^{+}16, CE19, CL09a, CCLL17, CMP16, CH18, CAD^{+}17, CSN^{+}23, CL09b, CDM93, DEP17, DAFZ^{+}18, DS99, DJS^{+}17, DHHD18, DGW^{+}17, EMPS06, EKSV16, FLTM18, FMT03, FBFB17, FWK17, FM23, FCT03, GT06, GZCX16, GLS21, HNW17, HS14, HS16, HLHD^{+}04, HZLZ22, HY08, HJG18, Ili00, IMG98, JS11, JV17, JW23, JPS^{+}17, JBR16, KK16b, KKEE13, KE16, KA03, KLS11a, KA95, KGH^{+}20, KLKT16, KW17, LHL15, LFZ^{+}22, LMS12, LV00, LMMN07, LKL00, LCDW21, LZL11, LK16b, LO02a, LYL21, LPWP22, LLM23, LO02b, MAE19, MLB21, MKS16, MP08, MGK20, MBI^{+}17, MK98, NBV17, Nee16b, Nee19, PBSS23, PDSK04, PDE08, PS05, QLF23, QTE20, RBGK03, RGY^{+}22, RKH^{+}16, RT17, SRK22, SAKMB21, SA21, SV99, SLD^{+}23, SMC^{+}20, SAKS13]. Optimal [SSM06, SPLM17, SPM^{+}17, SM18, SGS20, SF23, SAM12, SUS20, TE16, TM13, Tan16, TJHL21, TWWG19, THP94, TW22, TS14, UZ93, VLM16, WMX17, WWMZ20, WLL^{+}24, WFS09, XYA^{+}21, XLL21, XRL^{+}22, YK23, YAA09, YN20, YLWH20, YBX^{+}10, YLY^{+}16, ZSCJ14, ZSL^{+}21, ZRH18, ZXW^{+}19, ZMWX18, dAF04, AS94, AABD13, BB94, BBLV06a, BBLV06b, CSSJ14, Coh94, CK09, DMC06, EOSM10, Geo08, GGFS02, Gro99, GMY13, HRCW08, HMNK13, HLW13, HN13, HL15, JAS10, JJS13b, JGS^{+}15, JJSS04, JL98, KK16a, KK07, KIR08, KGPL13, KDYV12, KNSV13, KWE^{+}10, KT07, LCM04, LLY^{+}16, LCL^{+}13b, LLE15a, LLE15b, LSS07, LTS05, LYS11, MAN15, MBG^{+}03, MRD08, MLC07, NDGL06, NM09, NML08, Nee08, PT96, PLS07, PPV12, LZKT99, SBD11, SZKT98, SL15b, ST09, SSHK11, SHZ16, SS10, SGD05, SX10, SAS99, TAH99, Val07, WB11]. optimal [XY10b, XCR11, XCR15, YWK07, YGKX10, ZB95, ZY07b]. Optimality [CGMS13, HH18, IYYI18, SCN^{+}24, SF23, XPL^{+}17, YK23, YN18, AWKN16, AEJV13, GS11, HN10, JGLS14, JW11, OY13, PL02, TWLC10, WZY^{+}16]. Optimally [BLC21, PBV17, WCC14]. Optimistic [AIL23]. Optimization [APSG14, AN20, BBCD14, BBR19, CPS17, CDKZ21, CLHY22, CMY^{+}18, DMK05, DMT^{+}19, FFZ^{+}18, GHRH18, GSKN18, HSO19, HCL18, Kar03, KW19, LPS19, LCSS17, LCS^{+}18, LDZC20, LL99, LSL^{+}21, LFF^{+}19, MLS^{+}23, MHS95, MCC^{+}19, MS17, NLS19, PDI20, QLSW19, QYC^{+}24, SYZP19, SLF21, SE21, WDR^{+}20, WRT^{+}21, WLD^{+}24, WCW^{+}17, XLAC16, YN19, YWW^{+}24, YWRK19, ZCW^{+}22, ZLZ^{+}23, ZXC^{+}18, ZGLC20, AZ09, BE08, BGHS10, BH06, BLRC05, CNS04, CBL13, CL16b, DT93, GKJ12, GCS06a, HIM07, HK11, JLM15, KK12, LMS05b, LS06e, LSXS16, MCLG07, MMR96, Nee16a, NLB15, PLR15, RS07, RA95, RHQZ13, SLG^{+}16, SDW14, SK10b, WLLD05, WD05, WLL01, YY98, YC12, ZHC16]. Optimization-Based [CMY^{+}18, LS06e]. Optimization-Enhanced [MCC^{+}19]. Optimizations [VL16]. Optimize [DNCK20, MGZ^{+}23, RWL^{+}22]. Optimized [ACC^{+}14, BT23, FML23, GHK^{+}23, SZWW22, CC06]. Optimizing [ASKL18, AWFT15, CCE^{+}17, CFZ94, CP20, GKCR21, HVT18, HHA17, JLX^{+}16, KSM19, KRS^{+}17, KLKP16, LYZ^{+}23a, MVRZ09, NCK15, NLT^{+}18, PIST19, RIM98, SHHP00, TKM20b, TX08, XYA^{+}21, ZT12, ZSLZ21, GSRS^{+}15, LO96, LEYS11, LLE16, SJL^{+}16, YMO97]. optimum [CD96]. option [MM13]. options [RS95b]. Orchestration [CHW^{+}20, DQW^{+}23, MNZ23, TZPZ23]. Order [GZY23, GLA19, HZG^{+}18, KLE16, MSS^{+}12, Nee08, SRCDL19, ACC^{+}94, FqL98, HLW13, KNR^{+}16, LSXS16, MAN15, Tia05]. order-optimal [HLW13, MAN15]. ordered [FP97]. ordering [QCLC16]. organization [GZDG06, KK07]. Organizing [GCMP20, QZC^{+}22, FLMM10, LPCVC13]. Orientation [TAH17]. Oriented [CZM^{+}24, HLZY23, NZW24, YSC16, BB96, CZ06, CZFF98, GS10a, GP96b, LWL04, WPL06, ZVN99]. Orienteering [XLX^{+}21]. origin [LTY06]. origin-destination [LTY06]. originators [FMMLH06]. origins [GMSK09]. ORLA [GSPV^{+}18]. ORLA/OLAA [GSPV^{+}18]. Ornstein [OS21]. Orthogonal [CYK09, GSPV^{+}18, XZL20, KN05]. OSA [CSS^{+}14]. Oscillator [FSGH17]. OSN [ZGY^{+}16]. OSPF [RBGK03, SDV06, SGD05]. OSPF/IS [SGD05]. OSPF/IS-IS [SGD05]. Othello [YBQZ18]. other [ACC^{+}94, KWC93]. OTN [LZ23]. Out-of-Band [XLZ^{+}19]. out-of-sequence [JID^{+}07]. Outage [GGH11]. outages [DSA^{+}14]. Outband [AMG^{+}17]. Outdated [YN19]. outer [AJV06, YYZ06]. Outlier [WZL^{+}23a]. Outlier-Concerned [WZL^{+}23a]. outlook [FEC13]. Output [CWZ^{+}23, CC95, CM93, GSD09, LS06a, MSS02, Nai97, OWMM97, PB93, PDT09]. output-queued [GSD09, LS06a]. output/input [PDT09]. Outsourced [YZHZ21, YDW18]. Outsourcing [WLW^{+}20]. Over-Provisioning [SC18b]. Over-the-Air [SRMB^{+}23]. Over-the-Top [AAAR19]. Overbooking [LW22, SMD20]. Overbooking-Empowered [LW22]. Overcoming [PRR06]. overflow [PV04, TG97, VL10]. Overhead [FST^{+}09, GSW^{+}23, GKB^{+}16, JLZJ19, KKH^{+}22, LYSZ16, BSS09, CB99, JJL15, SHN16, TD03]. Overheads [LPR17, KP96, YDS10]. overlaid [YGC10]. Overlapping [CWZY21, DMDM17, FZW^{+}20]. Overlay [FLM^{+}22, FBFB17, JW23, JPS^{+}17, KRLL11, LT16, TTM22, AADS05, BCMR04, CBSK07, CJV16, CR14, DLT^{+}15, DZH03, FK07, FY07, ILS97, KCTI08, KEAAH08, OR11, PGV16, RPZ^{+}09, SHHA09, ST08, SLL^{+}11, SRS08, TAB^{+}15, WZR08, XB14]. Overlay-Based [FBFB17]. overlays [BLBS06, KLOS09, MJ15]. Overload [CLZ^{+}23, GT06, LM15, NS98, Pil01, Rum93, Smi95]. Own [ZGY^{+}16, ZZH^{+}10].

P [LP24, CJ07, ZL15]. p-cycles [CJ07]. P-MTI [ZL15]. P-Sketch [LP24]. P2P [ANSX13, BQ08, FLMM10, LDH^{+}12, LYWL08, LZL11, MLLY06, MRR^{+}14, OAN15, PLD16, RBPS21, RS05, SQ16, STM^{+}12, SdVK16, SRS08, TAB^{+}15, WYL09, WLR10, WLZ11, YWLL09, YCL15, ZSCJ14, ZLC12, ZZLW16, ZLW16a, ZCL11, ZFC13, ZFC15]. P2P-TV [TAB^{+}15]. P3 [ZJL^{+}18]. P3-LOC [ZJL^{+}18]. P4DB [ZBZ^{+}19]. P4xos [DBW^{+}20]. Pacifier [KHW12]. Pacing [HCFC20, ZHWH21, EL11, SEMO09]. PACK [ZCM14]. Packet [AD96, BSF16, BPS99, BBCD14, BP19, CAS^{+}20, CFM^{+}19, DBL^{+}19, DKN21, DRM04, FZ16, FBQ^{+}23, FGR^{+}17, FGRQ18, FLM^{+}22, FLH^{+}17, FC17, GDC^{+}17, GJD18, GT00, GSKN18, HPP^{+}23, HKS16, HCFC20, Hu93, HHL^{+}19, HLL^{+}21, HSKY23, JFM^{+}22, KLC^{+}18, LTDM17, LC03, LQZ^{+}24, LMT16, LYZ^{+}17, MBG^{+}02, MLJ^{+}22, ML22b, NSP^{+}16, NLT^{+}18, PKVI17, RRS22, RZZ06, RS97b, SNLL16, SEMO09, SSZ05, TML22, TWL22, VKO17, VLZL16, VKPI17, WLD^{+}16, WQY^{+}17, WWC^{+}18, WXH^{+}20, XLT^{+}22, XXZ^{+}23, YLYL17, YDLT18, ZXW^{+}21, ZGS^{+}24, ZWZC23, AK01, AK00, ACP05, ABJ^{+}13, ARS16, BV05a, BO00, BAC12, BIV01, BBG^{+}10, BM93, BZ97, BBC^{+}02, BTC01, BB95, BLT02, BHL^{+}06, CLP12, CT95, CGM04, CL03, CV96, CSLH13, CW16, CRV13, CH93, CM93, CT04b, CCL09, CF94, CZFF98, CKKK09, CH98, CCKK16, CF98, CT96]. packet [CAH08, DM03, DLH^{+}14, DSR02, ENW96, EST93, EW08, FGK10, FK99, FMMR10, FJ95, GYB^{+}04, GKS05, GV93, Goo08, GVC97, Guo04, HM06, IM03, IKM08, JDSZ97, Jia06, JL98, JM00, JL12b, Kam96, KMR95, KR00, KGL03, KqL99, KK00, KK03a, KR08, KNR^{+}16, LS94, Le 02, LLLS07, LRC15, LZ06, LSC99b, LLJ^{+}14, LMT10, LBS99, LS07, LS09, LCB^{+}10, LRM^{+}06, MEVSS03, MFL^{+}04, MLT11, MLT12, MDMM09, MV16, ME96, NMC07, Pax99, QSS^{+}15, RCOC03, RSR11, RCGT06, RB09a, SL94, SM00, Smi02, Smi08, SC95, SPS^{+}02, SBDR08, ST13, SV98a, TT07, TC06, UBPE02, WLCC07, WH97, WY95, WKZL96, WXW11, XL05, XLZC14, YMKC08, ZKVM14]. Packet-Based [HKS16]. packet-by-packet [ABJ^{+}13].

Packet-dispersion [DRM04].

packet-forwarding [CLP12]. Packet-Level [FGRQ18, BSF16]. packet-loss [KK00]. Packet-mode [MBG^{+}02]. Packet-Scale [LYZ^{+}17]. Packet-Switched [FZ16, GT00, BO00, BTC01, JM00, MDMM09, SV98a]. packet-switches [RCGT06]. packet-switching [WH97, WKZL96]. Packets [CNDK18, HLH^{+}18, KK21, TSS14, BM09, CK07, JID^{+}07]. Packing [GH93, PG21, RTLC17, XLL21, CGY00, WJK06]. Padded [JMS08]. page [BMS14b]. pages [Bar95, SP94]. Paging [BPVRSP16, AHL96, SZ08]. Paid [WXM21]. Pair [XCC^{+}17, LL09]. pairs [XGF^{+}14]. Pairwise [LZL^{+}21, YM16, HMvdLM07, KWS10]. Pairwise-Based [LZL^{+}21]. PALS [LYSZ16]. PAM [LLM^{+}24]. PANDAS [YN20, XPL^{+}17]. Pando [DLZL17]. Paradigm [BCS^{+}19, LLS^{+}23, LYZ^{+}17, PPTP21, TWH24, ZJL^{+}18, AAV09, CPSWL96, LS97c, LMS99, MR96, WQZ^{+}13]. Paradigm-Driven [ZJL^{+}18]. paradox [RK15]. Parallel [DAA19, FM23, GLH95, GVM23, HWM^{+}24, JHM^{+}19, JHJL21, JHM^{+}21, LS22, LZZ^{+}22b, LYZ^{+}23a, OLZ17, VN22, XZG20, ZWZ20, BBHHR10, DW11, HW99, IM03, KG16, LZ09, MSS02, RB02, SMG05b, WF93b, ZHLL06, ZGS10, Kai93]. Parallel-Server [DAA19]. Parallelism [EBJM18, LXX^{+}24]. Parallelization [LYL21, ZYZ16]. parallelized [GBL12]. parallelizing [LO96]. Parameter [LYZ^{+}23a, ODT09, YKKY08]. Parameterization [LMSR19]. Parameters [YYB^{+}22, DT93, HR95, LO98, MR98, RVA00, VG05]. Parametric [TMH11]. parametrization [LZL^{+}14]. Pareto [BJK20, BNS11, KGPL13, RSS09]. Pareto-efficient [RSS09]. Pareto-optimal [KGPL13]. Parity [NBT98]. Parity-based [NBT98]. Parsing [HWZ^{+}23, LBZ^{+}20]. Part [NSC^{+}22, WMX17, DTM15, EMPS06, GP94, Kim98, PG94b, VW09]. Partial [ACC^{+}94, CN16, HLZY23, HZG^{+}18, JPM^{+}19, LYL21, MZZ^{+}23, NJM^{+}19, ODJ23, HS08, Kam96, KE16, Lab97, LTY06, MG97a, MGG^{+}05]. partial-express [MG97a]. Partial-order [ACC^{+}94]. Partially [LLM23, LNLM24, REM17, TWY^{+}20, Kim94, LC94b]. Participant [HW22]. partition [LO02b, LORS06, OS05, WM95]. partitioned [AN05]. Partitioning [ADR18, SCN^{+}22, SA21, SLSC20, WBWV16, YDLT18, ZCZ^{+}21, BZM08, CKKK09, GF95, LYWL08, YJH05]. Party [DZ20]. PASE [MBI^{+}17]. Pass [FAS^{+}23, WBEGS05]. Passenger [BSRdA16]. passing [Hon94, PHL15, dSeSGM95]. Passive [CFM^{+}19, HDQ^{+}16, LLL10, RDZ^{+}19, DHSS14, HQW^{+}16, LM13, LCH^{+}06, NM06, RW07, WJK^{+}12, Wu94, ZA11]. past [PP02]. PASTA [BMVB09]. patches [VG08]. patching [EKSV16]. Path [BCO17, CQW^{+}18, CP17, CWHW18, CLM^{+}18, CLL^{+}19, CLHY22, CKZC19, CFS09, DCZG19, DGC^{+}20, FGK10, FAWW22, FLS^{+}22, FLZ^{+}23, GR20a, GDWX23, GDJX24, HNW17, HS14, HS16, HCW^{+}16, JF04, KHYA20, LFC18, LCL16, LGDC23, LLL^{+}16, LZC22, MHS^{+}17, OL16, RRG10, RRS23, WWMZ20, XOYL20, YZY^{+}20, ZOM03, ZML^{+}19, ZXW^{+}19, AM16, AL98, AZ06b, BC01a, BV96, BL04, CL03, CZ06, CRS99, CN08, CFS11, Con11, CTVD14, GZS15, GDC^{+}16, GLAM97, Gro99, HSH^{+}06, HAGL16, HBB09, Ili00, IMG98, KLS09a, KMHS09, KK03b, KS09b, LH07, LOP97, LMG04, LWKD03, LL10, LJC05, MHL^{+}14, Med95, MJ13, MK96, NST00, OZPZ09, PCV08, RGKS10, RBC07, SHJ10, SYR05, SCY98, Sob02, TNRP11, VC14, Voi07, WLL01, XK06a, XCX^{+}06, XSZ^{+}07, YSRL11, Zap04, ZRP00, ZY16]. Path-Based [CP17, Med95]. path-finding [GLAM97]. path-loss [XK06a]. path-oriented [CZ06]. path-protecting [MJ13]. Path-protection [ZOM03]. pathological [BPS99]. Paths [BCO17, GLA19, KLVL19, YRB^{+}18, ZXC^{+}18, ARK09, BBO^{+}05, CSS08, CFZ94, DLT^{+}15, GCZ96, GR12, GR14, GSW02, GO02, HLHD^{+}04, IAS06, LO02b, SG05, TKI^{+}15, ZWYY10]. patrol [AVS04]. Pattern [WZL^{+}23b, YDW18, YBX^{+}12, BBHK14, LH13]. pattern-matching [LH13]. Patterns [DWCZ17, JYC^{+}16, XLW^{+}17b, ACVS10, CG04, VG04, YDS10, YBX^{+}10]. PAVI [HRLY21]. Payment [ZY21]. Payoff [XZC^{+}20, CY14]. PCA [PBGMFM22]. PCM [CP95]. PCN [BGK97, ML12]. PCN-based [ML12]. PCS [RB09a, AHL96, FCL97, HA97, IPG97, LVB96, LKL00, LH03, Lin97, MS95, VG04]. Peach [AMP01]. Peak [CGC^{+}24, LJJ^{+}19, LS97a]. PEDS [BBHHR10]. Peer [CZX18, GNK^{+}21, LCDW21, AB09, AJF11, BLL07, CJW11, CPS^{+}12, CZCC14, CE08, CY14, HS08, KT08, LLY06, LYRL07, LTZ08, Liu10, LCW05, MR09, NSW11, OAN15, SW04, SLL15, SNS12, SENB09, SMLN^{+}03, SRD^{+}09, TM13, WYL09, WXR13, WTS^{+}13]. peer-assisted [AJF11, CY14]. peer-division [CJW11]. Peer-to-Peer [LCDW21, AB09, BLL07, CPS^{+}12, CZCC14, CE08, HS08, KT08, LYRL07, LTZ08, Liu10, LCW05, MR09, SW04, SLL15, SNS12, SMLN^{+}03, SRD^{+}09, TM13, WXR13, WTS^{+}13]. Peering [GNK^{+}21, PD16a, SRP^{+}11, WXM21, BFF07]. PeerProbe [CZC^{+}22]. Penalty [LNLM24, TW23]. Pending [SGS20]. Per-Connection-Consistency [BWK^{+}22]. Per-domain [Jia06]. Per-Flow [CCC17, DHS^{+}23, GTC^{+}24, NS16, SL16a, LCL12a, CM12, GSK08, HLW13, JJS13b, LDK13]. Per-frame [SGSB^{+}15]. Per-Packet [GDC^{+}17]. Per-stream [PS98]. Perceiving [XWH^{+}16]. perception [VNS02]. perception-driven [VNS02]. Perceptions [NL16]. Perfect [TKM20a, LV06]. Perfectly [RDR17]. Performance [ACOR99, AEG^{+}17, ANTR17, AZP^{+}23, BE08, BIV01, BFS21, BTK^{+}17, BG98, BD96, CWGT14, CG21, CH04, CZCC14, CWM^{+}17, CLZ^{+}20, CCCC17, DAA19, EPS21, EF08, FM20, GP96a, Gan20, GP94, HKV23, HVT18, IM08, JSW^{+}20, JS09, Kam96, KK05, KGdV^{+}21, KqL99, KD00, KK03a, KTvdSK18, KEY99, KqL98, KSM05, LS93a, Lab97, LNB00, LR22, LXW^{+}17, LCP^{+}20, LS03b, MLS^{+}23, MKAE20, MGZ^{+}23, MCMdlO23, MS17, ML12, MZZ^{+}23, MKS17, NBK02, NT00, OWMM97, OKAS23, PLS^{+}21, PG94b, RMPG16, RLKT98, RPP^{+}19, RHX^{+}20, RW96, SPLP20, SQ16, SDM20, SS16, SPB16, SBLS19, SGPH98, SZT01, TXHL23, TJ95, TdWC^{+}94, TS09, VB94, VBHT17, VCM04, WLCC07, WWW^{+}20a, WLC^{+}20, WGL22, WZL^{+}23c, YDY^{+}24, YS93, ZZX^{+}21a, ZLY23, ZZZ^{+}24, ZMD^{+}20, ZWZC23, ZRK06, vRDHSP17, AKS96, AMS^{+}08, AMSS08, AZLB16, AK96, AW97, ACP05, BCL^{+}09]. performance [BPSK97, Ban99, BBFG95, BLPS10, BJ15, BV05b, BCR^{+}12, Bor05, BH06, CT95, CM12, CL03, CHA95, CMM95, CBAT06, CMGL11, CR98, CDM93, CYL16, DM14, DLH^{+}14, Fan05, FGK10, cFKSS99, FML09, FST^{+}09, GMP13, GYB^{+}04, GS13, GMD15, GS97, HP01, HKV^{+}13, HOT97, HGE04, JK96, JCJ95, JGS^{+}15, JIN^{+}12, JS14, JSBM02, KVR02, KWJY16, KKSS12, KGPL13, Kim94, KK00, KLS09a, Kum98, KG16, LBRA05, LM97, LMS00, LAJS07, LKC11, LH13, LLY01, LD95, LC04a, LK05, LBX11, LEYS11, LNA07, LK14, LMS99, LMS04b, LLS09, LLW^{+}14, LNR94, MMH^{+}15, MH02, MBC^{+}94, MG97b, OSW97, PFTK00, PWDL05, PPPW05, PYL99, PS15, RLZ10, SJL^{+}16, SD15a, SKKA01, SNSW12, SS96, SR02, SML04, SHHP00, SPGM13, SK13, Świ96, TCS13, Tas96, TB10, Tur09, VSR11, WEK97, WL07, WSKV08, WZLX12, WFH12, WDCL15]. performance [WJLH06, WNV13, WM96, WYHL09, XG05, YD04, YZ10, ZKL07, ZR09, ZHLL06, ZTS94, DKL01]. Performance-Aware [ZMD^{+}20, SPB16]. Performant [LZW^{+}21]. performing [ME96]. Period [LKC11, LHY^{+}23, YLL^{+}17]. Period-Area [LHY^{+}23]. Period-controlled [LKC11]. Periodic [FLG^{+}20, JD20, RDR17, CG15a, FJ94, OdG97, XLWT12]. periodically [KZDM07]. Permissioned [WWL^{+}24a]. Permissionless [HJL^{+}20]. permutation [MCR10, QM99, SYP01]. permutation-scanning [MCR10]. Perpetual [LFZS11]. Persistent [DSL^{+}18, DLL^{+}20, FLM^{+}22, HSM^{+}20, LP24, BHL^{+}06, DGK05, JS06]. Personal [DZL^{+}20, NST^{+}16, YWZG23, ZGZ22, ZLN^{+}17, BSNI06, BLDF09, HA96, MHS95]. Personalized [GCX^{+}17, PWL^{+}22, WZW^{+}20, ZQ99]. Perspective [CGC^{+}24, CKS17, CLP^{+}17, DZ20, ISS22, LBP^{+}17, LW17, RRS^{+}14, XLL^{+}20, YXZ19, YYFC24, DJ12, EKD12, GYJ^{+}16, GRB09, KH15, KK12, cLqL97, LO99, NOF14, SMS07, WL10, XB07]. Pervasive [LLZ^{+}23a, RMDJ16, SCY15]. PFC [HZC^{+}19]. PGPS [YTJQ05]. Phase [JRL15, SYL^{+}17, WHLL23, ANSX13, RKZG10, YZ10]. phase-type [YZ10]. Phased [CDKZ21]. Phenomenon [LYLW22]. Phone [ZZL^{+}21]. photonic [CEFS99, HM06, JPH08, ZGS10]. PHY [HZHZ18, HXZ23, SMC^{+}20]. PHY-Tree [HZHZ18]. Physarum [SRB^{+}20, CAP15]. Physarum-Inspired [SRB^{+}20]. Physical [ALY^{+}20, BMY^{+}17, CWLH20, DLR^{+}18, GSKN18, HGZJ21, HOZL16, HZHZ18, LDS^{+}24, LZC^{+}24, PPS^{+}22, WGL22, XZL20, XTHL21, YNZ^{+}17, YWH21, YYL23, HQY^{+}16, JC13, LTS10, MVRZ09, PDE08, SAS16a, SBNRS14, SHZ16, ZL15]. Physical-Layer [CWLH20, LZC^{+}24, XZL20, XTHL21, HQY^{+}16, SAS16a, ZL15]. Physical-Level [HGZJ21]. Physics [LYLW22]. Physics-Constraint [LYLW22]. PIAS [BCC^{+}17]. piecewise [FKT98]. PIM [DEF^{+}96]. Pipeline [BM09, ZZL^{+}22, WY95]. Pipelined [IK07, AMKY99, BN05, OKM94, XLZC14]. pipelines [AS09]. pipelining [Tas99]. PJ [ZZLM23]. place [GMZR13, HOZL16]. Placement [AMCD19, AAG^{+}16, BPW23, CLZ^{+}23, DWL^{+}18, DLY^{+}21, DQYG23, FMH^{+}21b, HTJ^{+}21, HGM^{+}17, JM17, LYS^{+}18, LZC^{+}17, MHR^{+}20, PLT^{+}20, RLZ^{+}18, SJ21, SRS21, SFS^{+}22, SRB^{+}20, WJH^{+}21, AKSS12, CN09, FMSM^{+}11, GZCX16, IMG98, KWS^{+}11, KR05, MHL^{+}14, MHXT10, NSS96, NSCR06, RPZ^{+}09, SAS99, TM13, YY98]. Placement-Sensitive [HTJ^{+}21]. placements [RIM98]. Placing [MSSZ12]. Plane [AAT^{+}23, ACDP17, BFK^{+}18, BZS23, CLX^{+}22, CLZ^{+}23, GWQ^{+}23, LLZ^{+}23b, LCL^{+}18, LZS^{+}22, LLY^{+}22, PKK18, SBC^{+}17, TML^{+}18, XGQ^{+}19, ZZH19, ZLX^{+}21, JRL15, NCK15, RCOC03, TLP^{+}16]. Planes [JR21, NHLB21, ZBZ^{+}19]. plaNET [GG94]. Planning [CZL^{+}19, DF20, DKM^{+}17, GHRH18, JLS^{+}17, LZ23, SRCT23, WWMZ20, BSNI06, BCC07, LGC16, SYDM09]. Platform [CHW^{+}20, LLM^{+}24, PDV^{+}24, SDM20, TML^{+}18, DYH13, YBG^{+}12]. Platform-Independent [TML^{+}18]. Platforms [CVV17, KNE^{+}17, YWZ^{+}23, TH21]. Platoon [SQS20]. Plausibility [ZJL^{+}19]. Play [JHM^{+}21]. playout [BLL07]. Plexus [AB09]. plugins [DDPP00]. Plus [LNLM24]. PMC [CH20]. PNNI [Ili00]. POEM [LS16]. Point [GHK^{+}23, LWL17, NLNL16, RTNS21, CHH06, DGG^{+}02, HGE04, KT07, KAMG07, KK06b, LB04, MGR02, MK10, MW06, NSW11, NS98, RKA08, SV06, ZRLD05]. point-process [SV06]. point-to-cloud [DGG^{+}02]. point-to-multipoint [MGR02, ZRLD05]. point-to-point [ZRLD05]. Points [LJSB22, BB06]. Poisson [BVBV17, CFG08, CLCL23, PF95, RCFC15, SH14, YHCL21]. PolarScout [SPR^{+}20]. PolarTracker [WZLM22]. Policer [SJZ^{+}24]. Policies [BVL^{+}19, CMR17, KSUB^{+}18, KRRR17, LSL^{+}18, MMT16, NCM18, RTNS21, WJ17, XNHM22, YLA^{+}18, AGGT16, BL15, BFMF01, CGMS13, CGK94, DM96, ESP05, FRC98, GGC93, GRHA15, GS11, JGLS14, LNB01, MCS99, NAA^{+}16, PLS07, RD11b, RV00, SV99, SM00, TGT01, TJ95, VCD15, YAA09, dOSAU04, dAF04]. policing [CFPP96, RL94]. Policy [ABS^{+}16, BCE^{+}19, DNCK20, FM20, JYC^{+}16, LCL^{+}18, LDRS18, LFXY23, SVG16, TSS21, VBC^{+}17, WSX^{+}21, WSXL16, XXZ^{+}22a, BCL12, BI00, BSP07, CSS06, CDRV11, FJB07, GBC^{+}95, JGS^{+}15, KV98, LS93b, LBX11, LCL^{+}12b, LCG^{+}14, RVS09, SCP99, SN15, TG96, WWL02, YW07]. Policy-Aware [ABS^{+}16]. Policy-Based [LFXY23, WSX^{+}21, LCL^{+}12b, LCG^{+}14]. Policy-Compliant [LDRS18, RVS09]. policy-free [GBC^{+}95]. Polling [KAZ01, LXL^{+}19, dSeSGM95, QCLC16, SA01a]. Polling-Based [LXL^{+}19, KAZ01]. pollution [OF11]. Polya [HAG19]. polymorphic [WLC^{+}10]. Polynomial [BB94, Dat17, LDFK12, RV01, SG17b, XZTT08, KLNS93, XGF^{+}14]. Polynomial-Size [Dat17]. Polynomial-time [LDFK12, XGF^{+}14]. PON [ALMR14]. PONs [FS17]. Pool [MKOY24, OPGT16, ZY07b]. Pooling [NJM^{+}19, RAPP22, WW16, BCR^{+}12, WRS^{+}15]. Pop [ML18]. Pop-Routing [ML18]. POPI [LCB^{+}10]. Popular [XCL^{+}19, CKC^{+}13, cFCcFW05, XY09b].

Popularity [DD24, SLS^{+}23, SS16, ZLL^{+}23b, CKR^{+}09]. Population [LXL^{+}17b, QLSW19]. Port [WGZC21]. portability [KCA97]. portals [CKC^{+}13]. portfolio [TNRP11]. ports [LGW^{+}11]. Pose [LCC^{+}20]. position [KDHK15, SC10]. Positioning [JLS^{+}17, YYT23, SK06, WWT05]. positive [SWL06, XK06a]. Possibility [SG17b]. possible [CB97, KGPL13]. Post [SBTH19, WCL^{+}22]. Post-Disaster [WCL^{+}22]. Post-Processing [SBTH19]. potato [TSGR08]. Potential [RRS^{+}14]. Power [BSSS21, BGMB^{+}20, CCE^{+}17, CLS^{+}18, CHO^{+}19, CGR^{+}18, DEP17, DLC^{+}17, DLC^{+}18a, DRW^{+}22, GCMP20, GCZY18, GXL^{+}21, HIM07, HHA17, KLC^{+}18, LYSZ16, LWAL17, LBGL20, LCZH17, NMR03, PYL^{+}17, PMN19, PPTP21, SRI^{+}18, SDW14, SFFF03, STC12, TSS14, TG21, VBHT17, WCWZ17, WHZJ20, WZLM22, WN16, WCC14, ZSS^{+}20, ZRP^{+}22, ZDB^{+}17, ZMD^{+}20, ZMMG22, AAZZ12, BBG11, BCP00, BO07b, BS08, BLEM^{+}12, CE09, CHH06, CPS13, CMFA14, DPBT11, HLS^{+}14b, HRCW08, KKEE13, KM10, KG05, LMS05a, LS06b, LSC99b, LSZW13, LWAT13, LS10, LRG10, PZS^{+}16, PT96, PLS07, QCS07, RKZG10, RSS09, SRR08, ST09, SK10b, SLH^{+}06, SKS16, TPC09, Tan16, VGP14, WCY04, Wan04, XY10b, XSC01, XSC03, XSHS12, XC08, ZKH10, ZH08a, dAF04]. Power-Aware [WN16, PZS^{+}16]. power-balancing [SK10b]. power-control [XSC03]. power-controlled [XSC01]. power-efficient [HLS^{+}14b, SLH^{+}06]. Power-Law [TSS14, CE09]. Power-Line [VBHT17]. power-proportional [LWAT13]. Power-Saving [CLS^{+}18, WCC14]. Power-Weight [LWAL17]. Powered [CGC^{+}24, XTHL21, HA16, RIM98]. powerful [CNP13]. Powering [ACC^{+}14]. Powerline [ALPK21]. PPQC [AWH^{+}22].

PPVC [SZWW22]. Practical

[AHX19, BCC^{+}17, CZW^{+}21, CZZ^{+}21, FLZ^{+}23, GLLL17, HZC^{+}19, HRM22, JHM^{+}19, KHH^{+}18, LW11, LWK^{+}18, LJHB18, MZK^{+}17, RD11b, SPQZ20, WB11, WQY^{+}17, YYT23, ZZW^{+}15, CFC01, EL11, JGS^{+}15, KRH^{+}08, LXY^{+}14, RGKR10, SPC10, SXLL08, WKZL96, YKGK13, ZZZ^{+}14]. Practicality [KHAWC17]. Practically [MGS^{+}21]. Practice [JLSB16, ES05]. Pre [CKZC19, AB07, BZM08, CCF04]. pre-cross-connected [CCF04]. Pre-Defined [CKZC19]. pre-partitioning [BZM08]. pre-provisioning [AB07]. Preacher [TSS21]. Precedence [CBV^{+}18, VN22]. Precise [FBQ^{+}23]. precision [KMH12, TX08, WWL02]. Precomputation [OS03]. precomputing [SG05]. Predicates [YLYL17, YL16]. predict [CJH^{+}11, CTVD14]. Predictable [BFS21, GDWX23, LGDC18, LLX^{+}19b, ZLSK15]. Predicting [ANSX13, JBDF07, SNZ^{+}23]. Prediction [ACZP21, CH18, CJ18, FX17, HCL18, JWZ^{+}21, LMODF18, WHC^{+}22, XWW^{+}23, ZGYB20, ZHZ^{+}24, ZCM14, Ada98, DFMR15, FR07, GMZR13, JHR05, LM01, LDGL13, MSBZ10, PPPW05]. Prediction-Based [LMODF18, ZCM14, JHR05]. Predictive [BRISCSP11, HZCL16, HBSX20, LH03, OOM^{+}18, AW04, HP00, QS04, SK06]. Predistribution [YM16, Zha17, HMvdLM07]. preemption [dOSAU04]. Preference [EFA19, LMSR19, XL23]. Preference-Aware [EFA19]. Preferential [DGW^{+}17, LGDC19, CHM^{+}05, GDW^{+}16]. Prefetching [GR20b, WCZZ17]. Prefix [LQZ^{+}24, RT17, SBLS19, BLC12, BBHK14, DKT06, LS05b, MLT12, PT10, PT12, RW07, ZZH^{+}10]. prefix-compressed [BLC12]. prefix-preserving [RW07]. Prefixes [BGMB^{+}20, DKN96, DKN97]. preplanned [MFB99]. prerecorded [AS02]. Presence [MMT16, QJCR20, CL05, JMMT12, JS12, KAEAS14, KKP15, KEAAH08, LGKV14, LYS11, SSM03]. presentation [Hos98]. Preservation [JWZ^{+}21, WZ16, WHTC15]. Preserving [AWH^{+}22, Cob02, JZW^{+}18, JZ18, JJJ^{+}23, LLX^{+}17, LZZ^{+}22a, LZC^{+}24, LCH20a, LCH20b, LLX^{+}19b, NT24, PWL^{+}22, WPZM16, WHC^{+}19, XGW^{+}20, ZYH^{+}21, ZGS^{+}24, ZJL^{+}18, ZJL^{+}19, CL12, CBL13, CBL15, DJ14, HGW^{+}16, RW07, SEK15]. Pressure [MMT16, ABJ^{+}13, BSS11b, JJS13a, LEY14, MS15, OWMM97, YSTL11, YSRL11]. PRESTO [LGS09]. preventing [AVS04, BHN11]. Prevention [DPG^{+}24, HZC^{+}19, KGL03]. Preventive [LLX19a]. Price [LH14, XXN^{+}19, YM05, GS16, KAS16, TC06, ZSFZ11]. Price-Anticipating [XXN^{+}19]. Price-based [YM05]. priced [JK05]. prices [HN10, VHNPM96]. Pricing [AAS10, CBHS20, CSEZ93, HHL18, JHS^{+}19, KS20, LCDW21, Ma16b, MHB^{+}21, MT06, PL02, TEE16, VNM22, WS06, WT17, WMX17, WD22, WYL23, WZD24, WYL24, YKZ^{+}13, ZGHH19, CN10a, CSMW02, CDFG06, JVJ05, JJ08, KA03, LSM^{+}14, Mar04, MW06, MAS09, PT00, RSS09, RS12, SC09, SS06, YMR00]. Pricing-Aware [WT17]. Pricing-based [YKZ^{+}13]. Primary [BCO17, BPL20, CAO11, CP20, GPM03, JL12b, YGC10]. primary-segmented [GPM03]. PRIME [GLAMM11, MR09]. primitive [YTL12]. Primitives [LYDA19]. principle [HLG94]. principles [ALWD05, MBRM96, OY95, ZS05]. Prior [WZH^{+}18]. priorities [BW98, CU95a, HC02, HLG94, YMO97]. Prioritization [BGMB^{+}20]. Prioritized [BF01, CP95, FBM^{+}21, JR96, GGM10].

Priority [CWAO21, CNDK18, Dat17,

Mar03, MKOY24, TWN^{+}20, BOY00, CSC94, CLG^{+}00a, Hon94, ITSO01, IK07, KK06b, LX97, LS93b, LS93c, LCB^{+}10, Mar04, McM95, RRK96, SZN00, WXBZ04]. Priority-Aware [MKOY24]. Priority-Based [CWAO21]. Privacy [AWH^{+}22, CL12, CBL15, CP18, CP20, DZ20, FGR^{+}17, GCX^{+}17, HW22, HRLY21, JZW^{+}18, JWZ^{+}21, JSXN18, JZ18, LLWB16, LLX^{+}17, LZZ^{+}22a, LJZ^{+}23, LZC^{+}24, LCH20a, LCH20b, LCH22, LCH23, LLX^{+}19b, MYYR13, NT24, PWL^{+}22, SS21, TXL^{+}18, WZ16, WPZM16, WHC^{+}19, WMYR16, WWX^{+}19, XGW^{+}20, XSM22, ZYH^{+}21, ZGS^{+}24, ZJL^{+}18, ZJL^{+}19, CBL13, HGW^{+}16, SCY15, WHTC15]. Privacy- [CL12]. Privacy-assured [WMYR16]. Privacy-Aware [HW22, JSXN18, LCH22]. Privacy-Preserving [AWH^{+}22, JZ18, LZZ^{+}22a, LZC^{+}24, LCH20a, LCH20b, LLX^{+}19b, PWL^{+}22, WPZM16, WHC^{+}19, XGW^{+}20, ZYH^{+}21, ZGS^{+}24, ZJL^{+}18, CBL15, CBL13, HGW^{+}16]. Private [GS19, ZZG^{+}16, CK00, DGG^{+}02, KAS16, KRSY02]. Proactive [CLSS09, DLR^{+}18, HWLL21, HZB^{+}22, LPS19, LJHB18, LW17, TEE16, TE16, ZHCL17, BD07, FY07, WMYR16]. Probabilistic [CLL^{+}18, Goo08, KK21, LX24, OKAS23, PBGMFM22, SL15b, SB07, SS04b, YHCL21, AEG^{+}13, BL04, LML10, LJ09, WLLZ16]. probabilities [CLW95, CKR93, FT07, GS13, KL09, PV04, ZRP00, vDP93]. Probability [LMSR19, LMODF18, GGH11, KS01a, LXC05, TCPV13, TG97, VL10, WF93a, Zeg95]. probes [DLPT06]. Probing [SL16a, SYW^{+}22, CL09a, GKPS06, LHZ^{+}16, TZP^{+}10, WMS09]. Problem [BFG^{+}14, CCE^{+}17, CSL21, CMY^{+}17, GCWC17, GZL^{+}17, GFW^{+}18, HNW17, KHYA20, LWK^{+}16, SR18, WN16, XLX^{+}21, XXW^{+}23, YXZ17, BRS06, CAP15, CGY00, FMSM^{+}11, GZS15, GSW02, KKP15, KWS^{+}11, KRS00, LGC16, LS01, LWCY12, LÜ14, SH12, wTjCjC97, WC08]. problematic [TLP^{+}16]. Problems [BSP21, GR20a, JD17, KSNR20, KW19, LAV16, MVCS16, SM18, CD96, GL10, HSS08]. procedures [AA96]. Process [OS21, SC17, WUZ^{+}19, ODT09, SV06]. processes [CLC^{+}01, NSW11, SSV13, VR13, YTJQ05]. Processing [ALR^{+}24, BBCD14, DBL^{+}19, FML23, KLC^{+}18, LLWB16, LLX^{+}17, LTZ^{+}22, LCU^{+}20, NLS19, PKVI17, SE21, SBTH19, VLZL16, VKPI17, WYL^{+}22, YZHZ21, ZZL^{+}22, CV96, GLH95, HKT95, KP96, PD16b, SKT96, SCR08, ZS05]. processing-constrained [SCR08]. Processing-While-Transmitting [LTZ^{+}22]. Processor [AGMY21, KCCM16, BMvU03, HW99, Kar10, PG93, PG94a, RPF^{+}14]. Processor-network [KCCM16]. processor-sharing [RPF^{+}14]. processors [KL08, KKSS12, THDD05]. Producers [SGS20]. product [LZL12]. Production [CZX^{+}17, NZW24]. products [LM97]. profile [AW04]. Profiles [SSK^{+}17]. Profiling [JWZ23, KP96, OPGT16, SYL^{+}17, YWRK19, FGM^{+}13, HFC^{+}13, LY10, TRKN10, XZB08]. Profiling-Based [SYL^{+}17]. Profit [SL14, ZHW^{+}17, CL13, LWLL16, SK12b, SSAK12]. profit-driven [SK12b]. profitability [STM^{+}12, XB07]. Profitable [LSK20]. ProgLIMI [WMP^{+}18]. ProgME [YCM11]. Programmability [DQYG23, GDL^{+}22, GDWX23, GDJX24]. Programmable [BWK^{+}22, BCER20, CSA^{+}21, ENT^{+}24, GWQ^{+}23, LZW^{+}21, LTN^{+}19, MSTL17, NHLB21, SHV^{+}23, WMP^{+}18, WZL^{+}23b, XLD^{+}24, ZZX^{+}21a, ZBZ^{+}19, YCM11]. Programming [CWHW18, CGAC20, LDS^{+}24, MKG^{+}17, SYZP19, YLA^{+}18, WC08]. Programs [LZS^{+}22]. ProGraph [LZB^{+}23]. progress [PWMC12]. Progressive [ABMT23, HHSS16, LLM^{+}24, ZAW^{+}22]. prohibition [SKZ03]. Project [NKL^{+}23]. projection [TAH99]. Projections [FAF^{+}17, XWG14]. Projective [RB09a]. Promoting [ACA16, FF99, AVS04]. Promotion [NZW24, WFY^{+}18]. proof [BLL07, PPV12, Sha97, WHTC15]. Proofs [WPZM16, Geo08]. Propagation [LZB^{+}23, WWW20b, ZHCC24, CKS16, GS98, KL12, MCR10, MH97, WH97, XW11]. Properties [OKAS23, RKPP16, YSC16, Zha17, CBL06a, GGC93, IK09, JBDF07, Le 02, LT95, LR03, QS05, YL16]. Property [FZQ^{+}22, NHLB21, Sob17, qLP97, SMH95]. Prophet [ZGYB20, ZHZ^{+}24]. proportion [ZDR04]. Proportional [DSR02, HLHL22, LWC^{+}14, PCL15, BS09, HS08, LLY01, LWAT13, MSA^{+}16, MS08, NZTD02, SV98c]. Proportionally [HG14]. Proposal [LSHZ16]. Prospect [LCH20a]. protect [NS98]. Protected [BCO17, Wu94]. Protecting [BPL20, SCY15, ZLTX17, MJ13]. Protection [CLG00b, LLWB16, LLCJ22, LCH23, OL16, PC19, RS21, VBC^{+}17, YY20, ABK15, BCP00, CLSS09, CCF04, hCgKsYwT96, FAB12, HTC04, Kam10, KRLL11, KGGZ11, LYL07, MJ13, RRG10, Ram08, SHJ10, ZOM03, ZZZ^{+}07]. Protective [ZZLM23, CGK94]. Protocol [CKZC19, CZK^{+}21, GSL^{+}24, HWZ^{+}23, Kai93, LYZ^{+}23b, NDS19, NMD^{+}17, PYL^{+}17, PLM19, SRBBG17, TML^{+}18, WLL^{+}24, WSMJ04, XCC^{+}17, XCL^{+}19, XGW^{+}20, YCC21b, YYFC24, ZLLY03, ZRD^{+}23, ZZLM23, AP93a, AP93b, AK00, AB09, ALJ99, BFM^{+}96, BD96, BWH^{+}07, CCG00, CDO97, CDM13, CT04b, CLM^{+}16, CYK09, CFC01, CLG^{+}00a, CFD06, CWW^{+}15, EH11, EPD94, EST93, FCAB00, FST^{+}09, GMP13, GYB^{+}04, GP98, GAA08, GCS06a, HP01, HR95, IZC00, JCJ95, KV96, KH15, KCA97, KIR06, KT08, KV09, LS93a, LHL15, LCH^{+}06, LSW15, LTB04, LA95c, LJA14, LS97b, LT94b, LQCC16, MWQ^{+}10, MP94, Mil98, NLB15, OdG97, PP93a, PFC96, RW04, RCS14, RSSZ13, SKK07, SKT96, SKRK12, SL07a, SMLN^{+}03, SA05, TNF97, TMMS01, TYLH09, TLP^{+}16, VS97, VL99, WBP^{+}11, WCH95, WMYR16, WF93b, YCV15, YWZZ16, ZB95, ZT03]. protocol [ZL13b, RBS02]. Protocols [AGM^{+}17, CCF17, FSGH17, HKC^{+}20, LDD21, LCX^{+}16, MRJ20, Sob17, SF23, WCC14, ZQ23, AACD^{+}96, AA96, ACOR99, BGH^{+}95, BG98, BS02, CFG08, CFZ97, CPR99, DC13, FTV^{+}10, FLC09, FB07, GLH95, GJVZ06, HOT97, JGKT07, JM00, KS06, KAZ01, LM13, LH95, LLY06, LO96, LLS07, LCL13a, LM96, LBS05b, MMR09, MWC16, MP93, OAN15, OdG96, ODC^{+}16, PDE08, PV10, PWMC12, PSA96, PS15, QCLC16, RW93, RS05, SL95, SMV93, SQ12, Spi97, Świ96, TNML93, fTL06, ZLC12, ZCY16]. PROTON [LA95c]. Provably [FHH10, ZLG^{+}20, HFKC12, LR09]. provided [AG16, Smi08]. Provider [DSYN24, SSA11, VNM22]. Provider-customer [SSA11]. Providers [DCN^{+}19, GSM^{+}17, KS20, LS17, XHY^{+}22, CY14, GHR14, MCL^{+}11]. Providing [CLY06, KKSS12, KS98, SRBBG17, WXBZ04, BCGM07, JR14, KZ97, WCH95]. Provision [WN17]. Provisioning [AA99, ATB^{+}10, CTG^{+}20, CNM20, CHO^{+}19, DHHD18, HJG18, KAK19, LW22, LKMK20, MFR^{+}20, SK11, SMD20, SC18b, SZW^{+}16, YZY^{+}20, YXZ19, ZWX^{+}24, ZHT^{+}19, ZLW^{+}17, AB07, CJ14, DZH03, GGPS96, HMM11, KZDM07, KRSY02, LC04b, LV93, RDO^{+}07, RSM09, RRG10, SK10a, SYR05, SL07a, TGT01, VWT^{+}14, WLZ11, XTMM11, ZZZ^{+}07, CCL99]. proxies [MPFK02]. Proximity [ZLG^{+}17, LLW^{+}14]. Proxy [GZT03, CC06, RV00, ZWDS00].

Proxy-assisted [GZT03]. proxy-driven [CC06]. proxy-server-based [ZWDS00]. pseudoserving [KG99]. PT [LQZ^{+}24]. PT-Tree [LQZ^{+}24]. PTA [BZS23]. Public [LYZ^{+}23b, SSR^{+}20, WQL^{+}21, WWYY18, YCGH17, MM13]. Public-Key [WQL^{+}21]. Publish [EPB14, CJV16, MJ14, OR11, BTK^{+}17]. publish-subscribe [OR11]. Publish/ [BTK^{+}17]. Publish/Subscribe [EPB14, CJV16, MJ14]. published [MYYR13]. publishing [CKC^{+}13]. PUE [DAFZ^{+}18]. pull [MV14]. Pulse [FSGH17, HS06b]. pulse-connected [HS06b]. pulsed [ZL13a]. PulseSync [LSW15]. PUMA [LCG^{+}14]. Purchasing [RLZ^{+}18]. purpose [GBC^{+}95]. pursuit [ZXH^{+}13]. Push [AMS22b, CW23, NDS19, Tas96, MV14]. Push-Down [AMS22b]. pushing [LK14]. Pyramid [LYY^{+}22].

Q [BLC21, CXL^{+}24, GS13, NTS12].

Q-CSMA [NTS12]. Q-DDCA [CXL^{+}24]. QFlow [BBR^{+}22]. QFQ [CRV13]. QoE [BBR^{+}22, CCY^{+}14, HH18, LMW16, QWL21, VC12, VC14, WZW^{+}20, YWW^{+}24]. QoE-aware [LMW16]. QoP [ZXTT08]. QoS [BCP13, BV10, Bej04, BBO^{+}05, BB06, CS99a, CM05a, CLW19, CNS04, CCL99, CKZC19, CS99b, CHH06, DNS23, DZH03, ES03, Gan20, GGPS96, GP98, GO99, KZ97, KV05, LNG^{+}21, LS06e, LO98, LO02b, LORS06, MPS01, Mar96, MS08, MLC07, NZTD02, Ord99, OS03, OS05, QS04, RRR02, SSW10, Sob02, SL08, TGT01, VK04, WWL02, WFS09, XG05, XL11b, XSZ^{+}07, XZTT08, YO17, YFB02, YXZ19, ZXTT08, ZWX^{+}24]. QoS-aware [YFB02, YO17, YXZ19]. QoS-based [BV10]. QoS-Provisioning [CCL99]. QoSMIC [YFB02]. QoT [SRCT23]. QRCode [XLP^{+}23]. QSPNET [BIV01]. quadratic [SN15]. Quality [AWH^{+}22, BBR19, GS16, GS19, GXL^{+}21, HHL18, KCM16, KW17, LOFH21, LL17b, LWK^{+}18, LSSK17, MHH20, MFR^{+}20, PGMR18, RCR^{+}18, RMDJ16, SARM24, SYG^{+}22, SN15, WCW^{+}17, AL98, Cob02, KA03, KS09a, KS13, MTK03, PD07, PD16b, SCP99, SJ12, SRS01, TAG08, WKA^{+}13, YBG^{+}12, YL98, Yua02, ZM09, ZXH^{+}13, ZF96]. Quality- [RMDJ16]. Quality-Aware [HHL18, KCM16, SYG^{+}22, WCW^{+}17]. Quality-of-Information [RCR^{+}18]. quality-of-recovery [SJ12]. Quality-of-Service [MFR^{+}20, KA03, SCP99, SRS01, Yua02, ZM09]. Quality-sensitive [GS16]. Quantifiable [JWZ^{+}21]. quantification [CBL15]. Quantifying [GK16, LXLC20, LK13, OZPZ09, VC12]. Quantitative [KP21, RSLL23, CK07, LC04b, MOZ05, MV16, ZCD97]. quantization [Kok10, KK12, LA95a]. quantized [JRL15]. Quantum [AGBS23, CXL^{+}24, LYLW22, MFR^{+}20, ZZL^{+}24, ZQ23, ZWZ^{+}24, ZLHM22, VLMN09]. Quasi [TZCB23, BIV01, KS13, PCV08]. quasi-experimental [KS13]. quasi-path [PCV08]. Quasi-Static [TZCB23]. quasi-synchronous [BIV01]. Queries [FLG^{+}23, JZW^{+}18, LLG^{+}17, LXW^{+}19, SdVK16, YLS^{+}17, CL12, SG13, XLWT12]. Query [FML23, LLWB16, LLX^{+}17, MZZ^{+}23, RAA^{+}24, YZHZ21, ZZ17, GZDG06, HP01]. Querying [CMW^{+}20, AK09]. Queue [BLPS10, CMR17, DXX^{+}23, HS14, HS16, HKV23, JMMT12, qLH93b, qLH93a, RMPG16, TAJ^{+}10, WFC18, YN18, CU95a, CS98, CH98, ES07, cFSKS02, HC02, HH98, HGG06, IK07, KV96, Kop96, KS04, LBS05a, LAJS07, LBX11, LT95, Low03, NTS12, RrBG94, RW95, SM14, SL07a, VL10, WSW12]. queue-based [LBS05a]. Queue-length [JMMT12]. queue-length-based [ES07, NTS12]. queue-overflow [VL10]. Queued [HYZH16, AZ03, GKS05, GSD09, KKLS05, KK03a, LS06a, LLLS07, LMNM01, MBG^{+}02, MBG^{+}03, McK99, MSS02, MS03, Mne08]. Queueing [CNDK18, FM22, LS93b, LS93c, MMT14, QS05, SM00, YTJQ05, BBLV06a, BBLV06b, BZ97, BTC01, BT93, BSS11b, CSC94, CM93, CMM95, CFM^{+}09, CJ97, ENW96, GLMM04, GP94, GVC97, GMS16, HS03, JBDF07, LS06a, LYS93, qLlH97, qLP97, LRL07, McM95, PB93, PG94b, RRB06, SV96, SV98a, SV98c, SSZ03, TGT01, TS08]. queueing-theoretic [LRL07]. Queues [AGMY21, Dat17, FM23, Hua17, LI21, LY22, OS21, RL18, TWN^{+}20, CCL06, HBH93, KG16, LS94, NMH99, SV06, TG97]. Queuing [GPBL24, JK96]. QUIC [MCM^{+}23]. Quick [LXW^{+}20]. Quorum [KSSK18, WCC14, CSS06, HL99]. Quorum-Based [KSSK18].

R [DXX^{+}23]. R-AQM [DXX^{+}23]. R2 [QW23]. RaaS [CYG^{+}14]. Race [KCTI08, VG08]. RACKs [AB21]. radar [GZCX16]. Radiation [DMLC18, DLY^{+}21]. Radio [BDR22, BP19, BCC07, BMY^{+}17, CBdV^{+}17, CLW16, CCLL17, DRMP18, DAFZ^{+}18, DZL^{+}18, GJCB18, Hu93, HZ20, KAHKB17, PLM19, RZS14, SPQZ20, AD14, AK14, AK15, AAV09, BIV01, BB95, CAO11, CFG08, CSC94, CZM14, CF94, CFZ97, FEC13, GSA15, HA16, JL98, KKEE13, KS10, LZES14, LWT^{+}15, ODC^{+}16, RL93, SKY10, STC12, SK97, SAS^{+}16c, WSW12, YKZ^{+}13, YGC10, ZYL^{+}14, CC06]. Radio-Based [PLM19]. Radios [RFGL17, YWZG23, PRR06, SX10]. Rails [LXW^{+}17]. Rain [HS19]. Ramp [SLF21]. RAN [PD16b]. Random [CLGSS17, CLD10, Dai22, EBJM18, FJ93, FAF^{+}17, FWK17, HLL^{+}21, LZ13, Mod99, MH97, PBSS23, PP17, PG21, TW22, URZ^{+}14, WW16, WMT^{+}22, YM16, YHCL21, AS07a, AS07b, AAB05, AEJV13, FM06, FML09, GP94, HLS14a, HSM^{+}13, HMvdLM07, IW08, JLM15, JS09, KDHK15, KS03, LM97, LMS00, LV06, LLM11a, LWR15, LFL14, LE06, OAN15, OWKS16, TS08, WL07, XK06b, YM05, ZGG05, dAF04]. random-access [IW08]. random-walk [HLS14a]. Randomization [TG21]. Randomized [BES22, BCE^{+}19, BGPS06, DAFZ^{+}18, HJL^{+}20, JD19, PG18, RS19, STQ13, TG23, Van19, WMO^{+}23, IKDD15, LE12a, LCL12a, LLS09, PP02]. randomizing [BV05b]. randomly [WY06]. Randomness [JHM^{+}21]. Range [AYL21, HCL^{+}17, HMM^{+}20, LLWB16, LXW^{+}19, LYL^{+}22a, TAH17, ZMW^{+}22, BSH^{+}11, CSLH13, CL12, ENW96, GB99, HL96b, LL10, NPY07, RVA00]. Range-Based [HMM^{+}20]. Range-Difference [ZMW^{+}22]. range-free [LL10]. Ranges [BBHH^{+}18, LXL^{+}22b, MRM17, SLH^{+}19, RKH^{+}16]. Ranging [RFGL17, ZXH^{+}13]. Rank [CHS^{+}20]. Ranking [GBMV21, KMT05]. Rapid [CHO^{+}19, CZX^{+}17, fTL06]. Raptor [Sho06]. Rate [CQW^{+}18, CGZL20, CDKZ21, DRW^{+}22, DD24, DZ18, DPTP24, EPS21, EAH^{+}18, GLL^{+}18, GSM16, HCFC20, HSS08, KWS10, Kok10, KCH^{+}19, KW17, MZK^{+}17, ML06, PL17, RUH^{+}18, Smi08, SV98c, TJD23, VLDM17, WD05, XWJ22, XPW^{+}18, YN18, ZGL^{+}19, ZRH18, AK01, AA04, AAM05, AZ06a, AAV09, AOM04, BSH^{+}11, BBC^{+}02, BGK97, BKTN03, BLT02, CK10a, CC06, CR99, CLY06, CRL96, CCY^{+}14, CTG00, CLK01, CLA07, DRR98, FGK10, cFKSS99, FNQ00, FSM14, Geo08, GM00, GV97, GMY13, HZC07, HKLM07, HL03, HP00, HDM13, JR14, Jia06, JP09, JBR16, KV98, KVR98, KWCR10, KK05, KR99, KMHS09, KqL98, KK06a, LA02, LMR99, LS97a, LC03, LMS05b, LCH95, LT95, LR03, LS03b, LRG10, MAN15, MKT96, PA12, PD16b, PLL13, RKZG10, RLA06, RT99, RYS12, SZKT98, SMGP15, SKKA01]. rate [SL94, SBP03, SV98b, SDW00, SA01b, SS05, Szy16, TCS13, Tha01, VWT^{+}14, VL05, Wan04, WH11, YL97, YDS06a, YJH05, YM05]. rate- [Wan04]. Rate-Adaptive [TJD23, ML06]. rate-based [KqL98, LR03, MKT96, YDS06a]. rate-control [LT95]. rate-controlled [BKTN03, KV98, ML06, YL97]. rate-distortion [CC06]. Rate-proportional [SV98c]. Rate-Splitting [DPTP24]. Rateless [APC21, DLLL16, LDZ^{+}17, SCY08, XAST12, YS15]. Rates [AGMY21, DME23, Van17, WHZJ20, ZP18, ATB^{+}10, BTC05, CG04, CLW95, HH10b, KN05, LMSKZ99, Rum93, TR98]. Rating [DLT^{+}15, MHH20, PMN19]. Ratio [AEG^{+}17, DHHD18, BLCT97, GMWD13, KCB03, PDT09]. Rational [KE21, JKJ13]. rationality [CY14]. Rayleigh [Tan16, YYC^{+}21, YYFC24]. Rayleigh-fading [Tan16]. Razor [LMT10]. RCA [HDM13, YBG^{+}12]. RCBR [Ada98, GKT97]. RCC [ZZW^{+}23]. RCS [RLZ10]. RDCD [ZWY^{+}18]. RDMA [CLL^{+}19, SLD^{+}23, XCV^{+}20, ZZW^{+}23]. RDMA-Based [XCV^{+}20]. re [BLRC05, KCA97, TG96, ZA95]. re-optimization [BLRC05]. re-usability [KCA97]. re-use [TG96, ZA95]. Reachability [SVG16, CBL15, LM96, LK13]. Reactive [LLX19a, ZGYB20, RSSZ13]. Read [ZLZL16]. Readers [AYL21]. Reading [LYDA19, LYC^{+}19]. Reading-Based [LYC^{+}19]. Readings [XCL^{+}18]. Reads [KAT^{+}22]. Ready [ZLW^{+}17, VS97]. ready-to-go [VS97]. Real [CDHM17, CM16, FXHY21, FDM^{+}17, LTDM17, LSSC22, LXL^{+}22a, LCZH17, MR98, NS16, OPGT16, RVA00, SXEZ21, TAG08, TG21, TG23, TZX^{+}22, XL98, YL16, ZWY^{+}18, Ada98, AA04, AAM05, BO03, BFM^{+}96, BCGM07, BC01b, BBM^{+}10, CNS04, CS00, FK03, GQ16, GV93, GP98, GPM03, GAA08, GF95, Hou14, HLG94, HGG06, IS00, KMR95, KWJY16, LBS05a, LLD96, MRM99, PSA96, SZN00, SGPH98, SA01b, Świ96, VAS00, VSR11, WXBZ04, YSZL15, ZLS96]. real- [HLG94]. Real-Time [CDHM17, FXHY21, FDM^{+}17, LTDM17, LSSC22, LXL^{+}22a, LCZH17, NS16, TG21, TG23, TZX^{+}22, ZWY^{+}18, MR98, RVA00, TAG08, XL98, YL16, Ada98, AA04, AAM05, BO03, BFM^{+}96, BCGM07, BC01b, BBM^{+}10, CNS04, CS00, FK03, GV93, GP98, GPM03, GAA08, GF95, Hou14, HGG06, IS00, KWJY16, LLD96, MRM99, PSA96, SZN00, SA01b, Świ96, VAS00, VSR11, WXBZ04, YSZL15, ZLS96]. Real-Trace-Based [CM16]. Real-World [SXEZ21, GQ16]. realistic [VV09]. realizable [LPP11]. realization [BSF16, HLS^{+}14b]. Realizing [KBV^{+}13]. Realtime [LBP^{+}17, RDZ^{+}19]. Rearrangeable [CTH10, NWP09, HLL06, RMM99, ZGS10]. rearrangeably [LC96]. reassembly [HW99, SC95]. Reassessment [YYFC24]. rebate [LSM^{+}14]. Receive [CDKZ21]. Receiver [AK15, LM15, ZZW^{+}23, CJW11, MR09, PM96, ZBXH13]. Receiver-based [AK15, LM15, CJW11]. Receiver-Driven [ZZW^{+}23, MR09]. Receivers [CLCL23, LYL^{+}22b, YHCL21, GHK02]. Reception [XZG21, ZT03]. Rechargeable [JWW^{+}23, LXX^{+}17, LHY^{+}23, MLX18, YLD^{+}23, CSSJ14, KKJ06, ZHC16]. Recognition [WLL^{+}24, XWL^{+}18]. Recommendation [CGYZ17]. Recommender [WLC16]. Reconcilable [FBRL18]. Reconciliation [HJL^{+}20]. Reconciling [XB14]. reconfigurability [LS03a, TS09]. Reconfigurable [JWH^{+}24, MKOY24, NJK^{+}19, PWL^{+}24, SRMB^{+}23, TZCB23, ZLM^{+}23, BM08, CM05b, KS11, Med95, APSKPMGM12]. Reconfiguration [DPM^{+}18, HM04, LWP^{+}19, LDL^{+}22, WJ17, WLTJ19, ZWZ20, BM00, ÇM15, Lab97]. Reconfigurations [ZYZ16, CVM^{+}15, VVP^{+}13]. reconfiguring [OMA^{+}10]. Reconnaissance [YXH^{+}21]. Reconstruction [DYW^{+}16, DCZG19, LLL^{+}16]. Reconvergence [BGMB^{+}20]. Recorded [WML^{+}18]. Recovery [AB21, BCLS17, CZX^{+}17, CLM^{+}18, GDWX23, GDJX24, LTDM17, ODJ23, TXL^{+}18, XWW^{+}18, XPW^{+}18, XWW^{+}19, XCW^{+}20a, XCW^{+}20b, ZAW^{+}22, ZXZ^{+}19, AA96, Ban99, BFF07, CSC04, FY07, HM04, KL95, KRKH10, KHČ^{+}09, LNB00, LESZ98, MEVSS03, MFB99, MBF^{+}02, MLC07, NBT98, QSS^{+}15, SJ12, SA01b, XFS06, ZXTT08]. Recruitment [HW22, LYW^{+}21, WZD24, XGW^{+}20]. rectification [FCA^{+}06]. Recurrent [LBZ^{+}20]. Recursive [HKS16, Ses97, YDY^{+}24, GYJ^{+}16, Val01]. Recyclable [NS16]. RED [CJOS01, LBS05b, RAL04, TL06]. Redirecting [WXH^{+}18]. Redirection [LYS^{+}18]. redirections [SCKB09]. Redistribution [GCZY18, ZWTC16]. Reduce [GKB^{+}16, HHL^{+}19, LZL^{+}18, SSG18, ZGZ22, CSG14, MMC05, WXC16]. Reduced [MGK20, LSC99b]. reduced-power [LSC99b]. Reduced-State [MGK20]. Reducing [FZ16, Lin97, MCES19, SCHG22, SC18b, ZRD^{+}23, BIS00, CMFA14, HA96, KP96, SZKT98]. Reduction [ZCM14, BSS11b, IM08, KBS12, LA95a, LT95, SSFM08]. Redundancy [DZH19, FGRQ18, GHBSWV17, JK21, WTJR22, WSL^{+}24, AKK13, GMP08, LCW^{+}15, SPGM13]. Redundant [DRCM^{+}17, LPR17, MFB99]. redux [YCL15]. Reed [WSL^{+}24]. reel [CDRV11].

Reexamining [GYJ^{+}16]. Refactorization [ZWJ^{+}20]. Reference [WCK^{+}20, BM09, LDK13]. references [ABA^{+}16]. refined [LBX11]. Reflection [ZLX^{+}21]. REFWA [TKN06]. regeneration [KT11]. regenerator [FMSM^{+}11]. regenerators [MSSZ12]. Regime [LY22, GGL09b, GGL09a, GGH11, LV06, XK06a]. regimes [LLE16]. Region [DWCZ17, AJV06, JP09, JLS09, LLS09, TKI^{+}15, UN11]. region-disjoint [TKI^{+}15]. Regional [HCW^{+}23, SBGJ18, TRVG20, WWMZ22]. Regions [CLCL23, Dai22, DZ18, LE06, TKI^{+}15]. Register [XCZ^{+}17]. registration [VG04]. Regret [BRK^{+}22]. Regular [LT16, LN19, MPN^{+}14, BAC12, FDG^{+}11, IBM95, KH07, LLE15b, PLT14, QM99]. regularity [LLE16]. Regularization [LZX^{+}24, TZX^{+}21]. Regularization-Based [TZX^{+}21]. regulate [KA95]. Regulated [CVV17, LZKT99]. regulation [AS94, CCLT02, IS00, LYS93]. regulations [SSW10]. Regulator [BM22, VG05]. Regulators [Le 18]. REINFORCE [KLR^{+}20]. Reinforcement [CCZL23, FAWW22, HLZY23, HFF^{+}24, HNP23, HGZ^{+}23, LGCG^{+}21, MNZ23, MMG22, SCPB19, SGL^{+}22, TWY^{+}20, WZW^{+}20, WLY^{+}24, XWYL23, YZGC23]. Reinforcing [LMSS24]. rekeying [ZLLY03]. Relation [KLLT18, QJZ^{+}16, JD03]. Relations [CGL16]. relationships [DEH^{+}07, Gao01]. Relative [LL20, SYL^{+}17]. relaxation [SYDM09]. Relaxed [ZGZC20]. Relay [AMG^{+}17, CCK16, FBFB17, MFR^{+}20, NK20, SRB^{+}20, ZGLC20, CFG08, CR14, DK98, DFT06, GMY13, LJNK12, MHXT10, MS15, RK06, SSHK11, SR14, XWWC16]. Relaying [KS06, TJHL21, BGHS10, KE16]. Relays [YXAZ^{+}18, BJ15, GSRS^{+}15, GMYP16, RP13, WSC08]. release [RVV^{+}15, ZVN99]. Releasing [CHO^{+}19]. relevance [GB99]. Reliability [ALR^{+}24, CM05a, CZL^{+}19, CJ07, DF20, LLM11a, LLL^{+}22a, LT94a, MBL19, MCM^{+}23, WYZ^{+}24, ZSZN21, CLP12, CZ12, FT06, GGH11, HLX^{+}15, LLM14, LE12b, LLY09, WK13]. Reliability-Aware [DF20]. Reliable [BLM^{+}17, CLTM22, CWLH20, CZZ^{+}21, CNG^{+}16, CZK^{+}21, EPB14, LMODF18, LCZ^{+}23, RDO^{+}07, SL16b, Ste08, XAST12, ZLL^{+}23a, ZWY^{+}18, ZJ12, AA05, AADS05, BSP07, CGK10, FJL^{+}97, GS98, GAA08, HPR06, KHTK00, KHW12, LNB00, NBT98, PPV04, PNRMC13, RPGE04, SHJ10, WCH95, WXW15, XFS06, ZLLY03]. REM [RMPG16]. remainder [Su15]. Remaining [WYL^{+}22]. Remedy [CJL^{+}19]. Remote [GNK^{+}21, OS21, QW23, TW22, YEMJ24, WQZ^{+}13]. Removal [WMO^{+}23]. Rendered [LL10]. Rendezvous [CCLL17, CSL21, CYK09, ZYL^{+}14]. reneging [CSC94]. reneging/dropping [CSC94]. renegotiation [MR98]. Renewable [CZP18, LSS07]. Renewal [WN17, XSHS12]. Reno [CBAT06, PFTK00, SKV03]. rent [KKP15]. rental [KKP15]. Reoptimization [DJCA21]. Reordering [HLL^{+}21, ML22b, WLD^{+}16, BPS99, BHL^{+}06, LGKV14, MSS^{+}12]. repair [HK94]. Repairs [SGVO18]. repeat [QY12]. Repeatable [ZMD^{+}20]. Repeated [MRHWS14]. repeater [VLMN09]. repeaters [BCL^{+}09]. rephrasing [WLCC07]. Replacement [RV00, XNHM22, PP02]. Replacements [VCVC17]. Replay [WCQ^{+}20]. Replay-Resilient [WCQ^{+}20]. replaying [PPK15]. replicated [BSSS01, KR05, RB02, ZAFB00]. Replication [BLV10, BFS21, JK21, LCL16, LSC^{+}21, MHL19, TML22, MV08, WS08, ZFC13, ZFC15]. report [SC10]. Reporting [HV22, DG08, YG10]. Representation [LJJ^{+}19, LS09, RBC07]. Representations [DMDM17, KNE^{+}17, Ili00]. reprovisioning [LTP10, SZM08]. Reputation [FS17, MHB^{+}21, NL16, SL15c]. Request [DJS^{+}17, FMH^{+}21b, GSW^{+}23, LYS^{+}18, PLT^{+}20, XYA^{+}21, ZWX^{+}24, GSW99, QY12]. Requests [LPR17, SCL^{+}23, XYQ^{+}17, ATB^{+}10, JPS04, MSSZ12]. required [Kok10]. Requirement [DHHD18, KGH^{+}20, LH13]. Requirements [LNG^{+}21, LZC20, LZC22, SAC^{+}18, CS99a, LE12b, LO02b, LLY09, OS05, PG95, Pan99, SZKT98, XM99, ZEV07a, ZM04, vDP93]. reroute [ABK15, WL08, CSA^{+}21]. Rerouting [BES22, BBD^{+}18, BPST18, EGR^{+}16, JLZJ19, YXL18b, BCN02, GR16, NLY^{+}07, RLKT98, WCY00, dOSAU04]. Rescue [WCL^{+}22, WSX^{+}23]. Resequencing [LZ09]. Reservation [HSS^{+}21, SK97, XYA^{+}21, CV12, CFS09, CFS11, DM03, HSM^{+}13, SK06, YCL09]. Reservation-based [SK97]. Reservations [FSSC18, TCPV13]. residence [FCL97]. Residency [AQK^{+}19]. residential [GQ16]. Residual [WYL09, WLL13]. Residual-based [WYL09]. residual-geometric [WLL13]. Resilience [MDMM09, NTR18, XNHM22, AEG^{+}13, LYRL07, LJ09, LCW05]. Resiliency [CNM^{+}17, GDL^{+}22, KLR^{+}20, SACH21, YM16, RC08]. Resilient [ABBH^{+}16, BLBS06, HMM^{+}20, LSC^{+}21, NEH^{+}22, PSST21, SPR^{+}20, WCQ^{+}20, YSC18, ZZS^{+}16, ZYY^{+}21, CER12, DBDJ14, FY07, QL16b, RSU^{+}09, YSJL14, YKR11]. Resistant [DHK16, LMR07, PSK^{+}15, YKGK13, ZW10]. Resolution [CZM^{+}24, CBL06b, DLT16, KHLC13, LS05b, SG96]. resolutions [LL09]. Resolvable [KR20]. Resolver [NKL^{+}23, GYJ^{+}16]. Resolvers [PJMM22]. Resolving [GS09, SHHA09]. Resonant [LGZ^{+}23]. Resource [AMCD19, AMSB^{+}24, BSSS01, BDR22, BWES22, CLK^{+}24, CNS04, CWAO21, CYH^{+}18, CJJ09, CL19, DGK05, DTN^{+}21, DQW^{+}23, DJB^{+}22, DGG^{+}02, ENT^{+}24, FMSM24, FHMS18, FMPS20, FSSC18, GCD23, GSW99, GF95, HSO19, HNP23, JTL^{+}17, JTL^{+}18, JD22, KK16b, KKH^{+}22, KSSK18, LFZ^{+}22, LAV16, LWW^{+}19b, LW22, LKMK20, MV09, MZK^{+}17, MNZ23, ML23, MHR^{+}20, MKOY24, NJM^{+}19, NS21, NLB19, NLNL16, PWL^{+}24, PG18, PG21, RAPP22, SMEH20, SC17, SC18b, SLD^{+}23, SZW^{+}16, SC22b, TWL05, WT17, WDL^{+}23, XYA^{+}21, XLL^{+}20, XLL21, ZLRC20, ZLG^{+}20, ZWX^{+}24, AYS^{+}13, BCP13, BLV10, BM00, BRISCSP11, BF01, BS08, CqLL98, CZ06, CR14, DS04, ES07, FP14, FJ95, GMYP16, HTAZ16, HSM^{+}13, HG14, JS11, JK05, KL03, LRJ08, LMS06, LAN97, LLE15a, LCW^{+}15, LPCVC13, LMW16, MBL10, MCS99, NDGL06, NM09, RSR10, SZKT98, SCR08, SSAK12, WBEGS05, WSW12]. resource [YJ15, YZBR14]. Resource-Allocation [LAV16]. Resource-aware [TWL05]. resource-constrained [LCW^{+}15]. Resources [ALR^{+}24, BLC21, DCN^{+}19, SMD20, SBM^{+}18, YYB^{+}22, KR05, KMZR12, LMG04, LO02a, MHS95, MM94, NCK15, PD07, WS06, WRS^{+}15, ZS05]. Response [CZP18, JR22, WXH^{+}18, XCV^{+}20, GT06, HH98, qLH93b, qLH93a, NJW16]. Responsive [CL17, VV09]. Restless [LAV16, WN16, XWYL23]. Restorable [CN16, CN10b, KKL03, KL03, KLS09b, KLOS11, LRJ08]. Restoration [XM99, AB07, BBO^{+}05, BKLS08, Con11, IMG98, KLS09a, LWKD03, MK98, PCV08, QGCL11, THBR14]. Restricted [AC98, ASW00, KK03a]. restrictions [WM16]. restrictive [Ili00]. Result [Nee22]. resulting [CJ97]. Results [DRMP18, FSGH17, SH12, SWL06]. Resynchronization [JPS04]. Rethinking [CFP^{+}09, TB10, SMM11]. Retirement [DLY^{+}22]. Retransmission [TSS14, LNM^{+}09, LWR^{+}16, MBA06, PSA96, SV11, dAF04]. retrial [LO02a]. retrials [VCM04]. Retrieval [FBQ^{+}23, HK14, LJL^{+}16, LZC^{+}17, XTW^{+}22, BM97, RR93, YJZW15]. retrodiction [LWR^{+}16]. Reuse [LR22, CGGS97, HL98b, LSC99b, RW96, SS93, SS94a, SS94b, Sha97]. Revealing [CSSG23]. Revelations [NBV17]. Revenue [CZTX23, DSYN24, JSW^{+}20, RRK96, YK23, ZWR^{+}23, AAG14, HN10, MW05]. Revenue-Based [ZWR^{+}23]. Reversal [RpLP^{+}17]. Reverse [DXX^{+}23, IPG97]. Reversible [SLC^{+}07]. Revisited [LNG^{+}21, VL16, VIT21, Geo08, MBM09]. Revisiting [CL16a, LZXF14, LZX^{+}21, SARM24, ZMLR23, LTCS22]. Reward [LXX^{+}17, RL18, XWX^{+}24]. rewards [GKJ12]. RF [DJK22, DJK23, GGM11, WSZL20, XWL^{+}18]. RF-based [GGM11]. RFID [CLM^{+}16, CCF17, GLM^{+}16, GLC^{+}16, GSN^{+}16, GLY17, GLLL17, HQY^{+}16, HOZL16, HZHZ18, LL09, LHL15, LWCY12, LCL13a, LWL^{+}23, LCX^{+}16, LXL^{+}17b, LLL^{+}17, LXL^{+}17a, LLG^{+}17, LCX^{+}19, LXL^{+}19, LXW^{+}19, LCY^{+}19, LCL^{+}20, LCC^{+}20, LYL^{+}22a, LCQL14, LQCC16, OLZ17, QZL^{+}16, QCLC16, SL15a, SL15b, SL16b, SYL^{+}17, SCW^{+}21, SLSC20, SSH^{+}23, WXJ^{+}17, WCQ^{+}20, WSC^{+}23, WLL^{+}24, XCC^{+}17, XXCC17, XWY^{+}18, XZC^{+}19, XCL^{+}19, YW11, YLL^{+}17, YGL^{+}19, YZP^{+}14, ZL13b, ZL14, ZCY16, ZSZ^{+}17]. RFID-Based [ZSZ^{+}17]. RFID-Enabled [QZL^{+}16]. RFIDs [ALY^{+}20, AYL21, LYDA19, ZLZL16]. Rhythmic [LWL^{+}23]. rich [LS93a]. Ride [ZLL^{+}23b]. Riders [WWW^{+}18]. Right [CKLS22, FZ16, LWAT13]. right-sizing [LWAT13]. Rigid [WWL24b]. Rigorous [GLLL17, NR13]. Ring [TS14, BO03, CM05b, CDRV11, Coh94, COS95, GGC93, Gro99, KK93, LS03a, LS01, LT94a, RW96, SMG06, TJ95, TG96, TMMS01]. ring-based [Gro99]. Rings [CXW^{+}18, RS21, YM16, AK96, BBMELH08, CGGS97, FCT03, FT06, GYB^{+}04, GRS00, HLHD^{+}04, RW95, SZ07, SF95, ZVN99, ZQ99, ZQ00]. Risk [GSKN18, SRS21, VTBK21, XTMM11, MW05, SYR05]. Risk-aware [XTMM11, MW05]. Risk-Sensitive [GSKN18]. Risks [FS17]. RL [SCPB19]. RL-NSB [SCPB19]. road [HLP11, SK06]. roadmap [FGM^{+}13]. roaming [MD04]. Robin [PK01, QFH^{+}18, TL22, CM03, LS94, LMS04a, OJRCC02, SMT98, SV96, RP06]. RobLoP [JZW^{+}18]. Robotic [HWJZ21, ZXW^{+}20b]. Robust [BR06, BLT02, BCD19, CGZL20, CLY^{+}17, DGNK21, DYW^{+}16, ESG11, EAH^{+}18, GJD18, HGM^{+}17, JZW^{+}18, KO13, KW17, LMSS24, LSZW13, LZB^{+}23, LPS19, LWC^{+}23, LDY^{+}16, LAL^{+}24, QJCR20, RPPA22, SHZ16, SYJ09, THDD05, TZCB23, VRK09, VCVC17, WZX^{+}22, WWL^{+}24a, WML^{+}18, XPL^{+}17, YLD^{+}23, YXZ19, ZCX^{+}15, ZY21, ZZLW16, AC06, CDM13, GJVZ06, GT99, HZL16, JLM15, KLC15, LMP08, LKZ^{+}04, RrBG94, RSSZ13, Smi95, XXBC14, YS93, YC12]. Robustness [CQLW22, LBS05a, LGCG^{+}21, QZX^{+}17, QLSW19, QZC^{+}22, QYC^{+}24, TZX^{+}22, XXN^{+}19, ZMH17, DSTM12, TPC09]. Robustness-Aware [TZX^{+}22]. RobustPay [ZY21]. ROC [YKR11]. rocketfuel [SMWA04]. ROHC [THDD05]. Role [WMX17, BMVB09, BM97, JS06, PDT09, SJGH10, SSA08]. Roles [LLX^{+}19b]. room [ZT03]. Root [FLM^{+}22, MRMR17, WWYY18, AST11, YBG^{+}12]. ROSE [QZX^{+}17]. rotating [LT94a]. Roulette [FMH^{+}21a]. Round [AEG^{+}17, PK01, QFH^{+}18, RP06, TL22, AAM05, CM03, LS94, LMS04a, OJRCC02, SMT98, SV96]. Round-Robin [TL22, CM03, LS94, LMS04a, SV96]. round-robin-based [OJRCC02]. Rounding [RS19]. Route [AMS22a, FLMS18, SVL^{+}16, XYL^{+}17, ZWCL17, AMS^{+}08, AMSS08, BLC12, CYG^{+}14, CDRV11, EST93, GCH^{+}15, KKL03, LWT^{+}15, LXX^{+}14, MRM99, YG10, ABC^{+}16]. Routed [SSR^{+}20, AM16, BM00, CV12, GL93, KS01b, RM02, SYR05, SAS99, ZKL11]. Router [DDPP00, KLSV12, MFT^{+}20, PDT09, CVM^{+}15, HPR06, HPV09, IKM08, LLW^{+}09, LS05b, LCB^{+}10, PPV04, PCB^{+}98, RPGE04, YLLY05, ZDR04]. router-assisted [HPR06, PPV04, RPGE04]. router-specific [LLW^{+}09]. router-wide [CVM^{+}15]. Routers [EPS21, HLH^{+}18, PLS^{+}21, VWNT17, BBG^{+}10, DDPP00, LBS11, NKS08, PZS^{+}16, PT12, SDV06, SKHL12, VSR11]. Routes [KSSD24, FR07, GV06, LP07, SK12b]. Routh [AOM04]. Routing [ACC^{+}14, ABBHP01, AdSD16, AGCFV18, ABC^{+}16, ASKL18, AGBS23, ER20, AAZZ12, AAF^{+}16, BSSLB95, BJK20, BBEF^{+}21, BO16, CCE^{+}17, CAZG20, CYG^{+}14, CZX^{+}17, CLQ^{+}19, CXL^{+}24, CRS18, CSN^{+}23, CLP^{+}17, DJS^{+}17, DZH19, DPT^{+}18, DPM^{+}18, DDP^{+}19, DMB94, DKN96, DKN97, EMAL17, FXHY21, FZX^{+}23, FMCS20, FM23, GR20a, GYLH17, GLNP01, GVGV17, HHSS16, HLP^{+}16, JV17, JFM^{+}22, JPS^{+}17, KKL03, KSSK18, KGH^{+}20, LMSS24, LMODF18, LLCJ22, LWC^{+}23, LYKT21, LNC04, LLY09, ML18, NGRF19, Ord99, OB03, PC19, PBT^{+}20, PLT^{+}20, QL16a, QFH^{+}18, RpLP^{+}17, RS95a, RS07, RZE^{+}21, SAC^{+}18, SACH21, SdVK16, SSY19, SCL^{+}23, SVL^{+}16, Sob17, SF23, SLH^{+}19, SGL^{+}22, TJL^{+}19, TWY^{+}20, TTCT19, TZX^{+}22, URZ^{+}14, WBWV16, WJYL16, XLH^{+}17, YSC18, YYB^{+}22, YN18, YLWH20, ZZL^{+}24, ZXW^{+}20a, ZY21, ZLL24, ZMMG22, ZLTX17, ZXW^{+}19, AKA10, AN05, AEB02, AS08, AC06]. routing [AZ06b, ABJ^{+}13, AFT11, BD07, BLV10, BCGC15, BO07a, BBFG95, BBLV06a, BBLV06b, BSH^{+}11, BGHS10, BKLS08, BNJR12, BNJ16, BP96, BWS10, CM05a, CJ14, CT04a, CV12, CNS04, CM05b, CSSJ14, CS14, CS15, CYK09, CTG00, CFC01, CEFS99, CER12, CMV10, CL05, CKR93, CRS99, CDRV11, CO94, COS95, CN08, CR14, CFD06, DEF^{+}96, DLT^{+}15, DFGV11, EAB01, EAB02, EFK07, FJB07, FC99, FEC13, FJ94, FB07, FML09, FSH^{+}13, GR01, GLAM97, GLAMM11, GLA93, GKT93, GLG04, GR16, GS03, GSW02, GO99, GRHA15, GLSB08, HP01, HHL06, HSH^{+}06, HY08, HL05, JRY09, Jia98, JHR05, KA98, KD00, KL03, KLS03, KLS09a, KLS09b, KLOS09, KLOS11, KLS11b, KPP93, KLO97, KWH11, KWZ08, KV05, KIR06, KHČ^{+}09, KS09b, KGGZ11, LMJ98, LABJ01, LAPS08]. routing [LQ13, LSV01, LRJ08, LDFK12, LS14, LK95, LT02, LMR07, LML10, LLM14, LMP08, LMMN07, LS99, LLW^{+}12, LXY^{+}14, LS06c, LGGZ10, LHM02, LS06e, LSS07, LR09, LB04, LTS05, LW11, LSXS16, LCW05, LS97c, LO98, LORS06, LRG10, LH10, LBP^{+}16, MG97a, MWQ^{+}10, MBLN93, MG97b, MHRR12, MMS01, MS15, MA98, MK96, NM09, NMR03, NZTD02, NXTY10, ORS93a, ORS93b, OS03, PM09, PSK^{+}15, PNRMC13, PYL99, PS05, PA12, Pax97, PMH95, PT10, QYZS06, RGKR10, RVS09, RM02, RM08, RHC^{+}12, RS00, RHQZ13, RYS12, SMG06, SPH04, SRP^{+}11, SHHA09, SM16, SRB10, SRS01, SLH^{+}06, SKCW10, Sob02, Sob05, SQ12, SNC^{+}07, SL08, SPR08b, SPR08a, SD00, TNRP11, TK12, TYJ16, TSGR08, TAH99, TYLH09, VWT^{+}14, VK04, VB94, VCD15, WCY04, WQGW09, WTXT11, WJZ^{+}12, WJK06, XWWC16, XCR11, XCR15, XZTT08, YFB02]. routing [YCB07, YXF^{+}13, YMO97, YSTL11, YSRL11, Yua02, YNDM09, ZOM03, Zap04, ZA95, ZHC16, ZZG^{+}16, ZGS10, ZW10, ZRP00]. Routing-as-a-service [CYG^{+}14]. Routing-Aware [LYKT21]. Routing-Oblivious [BBEF^{+}21]. routings [Ste08]. RSSA [KW19]. RSVP [Kar06]. RT [GAA08, BDS07]. Rule [RT17, VCVC17, YXC^{+}18, ZWZC23, CW16, KK03a, SBD11]. Rule-Set [ZWZC23]. rulebase [CKKK09]. RuleChecker [ZZH19]. RuleOut [XBM^{+}23]. rules [BDWS12, NS98]. RuleScope [WBY^{+}17]. rulings [SZG09]. Rumor [GCW21, DMC06]. run [VL05]. runs [HKLM07]. Runtime [SHV^{+}23, ZWL^{+}22]. Rural [AD18]. RWA [CKV11, JD17, ZOM03].

s [PES^{+}12, LLM^{+}24, WZL^{+}13, BLC12, PCB^{+}98]. S-ALOHA [WZL^{+}13]. s-based [HM06]. S4 [MWQ^{+}10]. SACK [SKV03]. SAF [PRH17]. Safe [DLC^{+}17, DLC^{+}18a, LXY^{+}14, LGGZ10, VCVC17, VVC^{+}17, AZR97, WJZ^{+}12, XDZ^{+}23, XXZ^{+}22a]. SAFE-ME [XXZ^{+}22a]. Safeguarding [FGR^{+}17]. Safety [SARM24, ZV16, SR02]. Safety-Awareness [ZV16]. Same [DKSC18, HH98]. Sample [HS14, HS16, LCL16, ZY16]. Sample-Path [HS14, HS16, LCL16]. sample-path-based [ZY16]. sampled [DLT05, HV06]. Sampleless [WCWZ17]. samples [PP02]. Sampling [CM18, DZL^{+}20, DHS^{+}23, HSM^{+}21, JHL22, LCL17, LCX^{+}19, OS21, PBSS23, VGKG10, BTC05, DT93, DG01, DG08, HLS14a, LQCC16, MV09, OAN15, PV04, SRD^{+}09, WLL13, ZGG05]. Sampling-Based [JHL22]. Sampling-on-Demand [CM18]. SAT [BS97]. SAT-based [BS97]. satellite [AMP01, AEB02, CDFG06, EAB01, EAB02, RLZ10, TKN06, Tha04, WCH95, ZRK06]. satellite-switched [Tha04]. Satellites [LLZ^{+}23a, FMT03, NMR03]. Satisfaction [CMY^{+}17, PPTP21, DBL13]. Satisfiability [RCR^{+}18]. satisfy [MSSZ12]. Saturated [BT23]. Saturation [ACDP17, JS12]. SAVE [DRR98]. Saving [BHC^{+}21, CLS^{+}18, LYSZ16, WCC14, CLP12]. Scalability [CKLS22, JMS07, LJ09, LL18, RCR^{+}18, XHC^{+}18, ZFW^{+}17a, ZR09, ZJWY17, AIN^{+}15, CRL96, GRHA15, HS06b, LJC05, LR03, TYJ16]. Scalable [AKK13, AC09, ARS16, BV05a, BAC12, BBHK14, CCC17, CWM^{+}17, CLH^{+}24, CHFH20, CEFS99, CKKK09, DPT^{+}18, EFA19, HYK^{+}23, KHTK00, LGW^{+}11, LZZR12, LLWB16, LYMA^{+}17, LWC^{+}23, LT16, MLS^{+}23, MEVSS03, MVB^{+}21, NKNK17, NB99, OWKS16, QZL^{+}16, RV21, RBPS21, SFAS05, SIYL09, SBLS19, SGL^{+}22, WLC^{+}20, WZL^{+}23b, XXZ^{+}22a, YLYL17, YDLT18, ZSSK02, ZEV07b, AC98, AB09, ASCG08, BGHS10, CBSK07, CLK01, EFK07, FCAB00, FHSZ13, GDW^{+}16, GSN^{+}16, IBM95, KL07, KNR^{+}16, KSV07, LSW15, LTB04, LLW^{+}12, OS05, PT12, QL16b, SA04, SLO^{+}14, SKHL12, SSZ03, SMLN^{+}03, STL04, WHM^{+}13, YF05, ZLLY03, ZEV07a, ZLSK15]. Scale [AS19, AAG^{+}16, BRY^{+}19, BFK^{+}18, CZW^{+}21, CXL^{+}24, CWZY21, CGL16, DGC^{+}20, GCD23, GLM^{+}16, GLY17, GLLL17, HV22, HAB^{+}22, HOZL16, HGZ^{+}23, JD17, LSDT19, LXL^{+}17b, LXW^{+}19, LYZ^{+}17, NTD17, PJMM22, QLY23, QZX^{+}17, QLSW19, RWL^{+}22, SLD^{+}23, Van19, VKO20, WWW^{+}20a, XXCC17, XLW^{+}17b, YHH^{+}21, YGL^{+}19, ZFW14, ZHZ^{+}18, AKA10, AF99, BBC^{+}02, BS00, CZF^{+}16, CKR^{+}09, CL03, CC95, CRL96, CCL11, CLM^{+}16, DZNT14, DLH^{+}14, ES03, FCA^{+}06, GSN^{+}16, Goo08, GKT97, GT03, HMvdLM07, JC13, Jia06, JYT^{+}15, KHLC13, LC03, LYWL08, LTB04, LTZ08, LZL12, LGD^{+}10, LCQL14, MA12, PYL99, PS05, PLS07, PJ13, LZKT99, SJL^{+}13, SQZ09, SXLL08, TK12, WDCL15, XY09a, XW11, YKKY08, YDS06a, ZSFZ11, ZW14, ZL13b, ZL14, ZKO93]. Scale-Free [BFK^{+}18, CGL16, QZX^{+}17, QLSW19]. scaleable [PPPW05]. ScaleTrust [HYK^{+}23]. Scaling [AK09, CFC^{+}24, CBL06a, DJK22, DJK23, FAF^{+}17, FDM^{+}17, HZLZ22, JWL^{+}18, LL17a, LY22, LW20, MYMY17, RRS23, SVL^{+}16, TWL22, VN20, VN22, WWL^{+}15, XDZ^{+}23, YGC10, AGLM10, AAZZ12, BSF16, BLC11, DFT06, EMPS06, GGL09b, GGL09a, HW12, KEW06, KCCM16, PES^{+}12, XK06a]. Scalpel [GDW^{+}16]. scan [DKC^{+}15, Tre11]. Scanner [HCW^{+}23]. Scanning [GLM^{+}16, QML^{+}24, MCR10]. SCAPE [DLC^{+}18a]. SCD2 [XHY^{+}22]. SCDP [APC21]. SCED [SCP99]. Scenario [MNZ23, YLA^{+}18]. Scenario-Based [YLA^{+}18]. Scenarios [LQZ^{+}24, SRBBG17]. Schedulability [LK05, FP97]. Schedule [BLC21, MRM17, CT04b, CD96]. schedule-sensing [CT04b]. Scheduled [CLGSS17, JP09]. Scheduler [HLHL22, LS22, TES19, WYL^{+}22, XXZ^{+}23, ASKR16, Guo04, PDSK04, RP06, SPC10, SKUB12, Tur09, WTS^{+}13]. schedulers [FKT98, GK16, KKV16, LS94, LMS04a, LK05, LE12a, MFL^{+}04]. schedules [CF94, DS99, RCGT06, RA95, WB11]. Scheduling [ER20, AdVS20, APSG14, AZ06a, AZ11, AWM^{+}20, AHEK24, AEJV13, BCC^{+}17, BSSS21, BWES22, BC01b, ÇM15, CSD22, CMP16, CGC^{+}17, CH18, CDGZ20, CJ18, DEP17, DMLC18, DWCZ17, DZH19, FLG^{+}20, FMH^{+}21b, GLA19, GB18, GHW22, GLS21, GGM10, HTL^{+}19, HTJ^{+}21, HTW^{+}22, HLZY23, HS14, HS16, HDF19, Hou14, HYZH16, HZCL16, HGB^{+}19, JHL22, JMI95, JJJ^{+}23, JE18, JD20, KSUB^{+}18, KSM19, KYM22, KSSD24, KCM16, KAA^{+}18, KZH^{+}20, KW17, KLE16, LPR17, LE12b, LLL^{+}22b, LWAL17, LBGL20, LEY14, LWW^{+}19b, LX21, MS14, MMT14, MEWP13, MGK20, MKS17, MHR^{+}20, NV21, Nee16b, Nee19, Nee22, PWL^{+}24, PS94, PK01, PG18, PG21, QLF23, RL93, RWL^{+}22, RDR17, SS17, SG17b, SSY19, SC22a, SC22b, SK21, SdVS22, TES19, TJL^{+}19, TZX^{+}21, TSN^{+}21, Tha04, THMK12, TG21, TG23, VIT21, WJ17, WT17, WZH^{+}18, WCW19, WYL23]. Scheduling [WLS23, WSZL20, WXH^{+}20, WH97, WW16, WLL^{+}16b, XPL^{+}17, XYL^{+}17, XLL21, YPA19, YCW^{+}19, YCZ^{+}23, YN20, YYC^{+}21, YYFC24, YLWH20, ZA11, ZWYD18, ZSS^{+}20, ZTH^{+}23, ZLWM18, ZLW18, ZGZC20, AS14, AD14, AF99, ALJ99, AS96, BGSSW13, BTC01, BHN11, BCR^{+}12, BRS10, BSYS12, Bor05, BESW08, BSS09, CKL16, CM12, CL09a, CM03, CRV13, CHCH00, CLSC15, CCA96, CJZS14, CGEN98, CK07, CK09, CK10b, CG15b, CAH08, DV09, DSR02, ESP05, ES07, GIKK11, GV97, GVC97, GSA15, GLS09, GS11, HH10a, HKV^{+}13, HY10, HLW13, HN13, HK96, IS00, ITSO01, IM08, IK07, JK96, JMMT12, JR14, JMS08, JS11, JAS10, JJS13a, JJS13b, JGLS14, JGS^{+}15, JW11, JJL15, JP13, JS09, JLS09, JLRS16, JL98, KWJY16, KWCR10, KKEE13, KAEAS14, KKLS05, KLMW11, KWE^{+}10, KCB03, LX97, LNS11, LLLS07, LMMN07, LK02, LLE15a, LLE15b, LHZ^{+}16, LLE16]. scheduling [LS06d, LR09, LW13, LYS11, LNL^{+}16, LLS09, LBS99, LRG10, MSWL06, MKS16, MSA^{+}16, MBG^{+}02, MBG^{+}03, McK99, MV16, Mod99, MS15, NJW16, NM06, Nee08, Nee09, Nee16a, OES16, PHL15, QZZ^{+}13, QM99, RSU^{+}09, RB09a, RS97b, SBD11, SMT98, SKT96, SAS16a, SCP99, SM16, SM00, SV98b, Su15, SS05, SR14, SCY08, TT09, TJ95, Tas96, Tas99, TZP^{+}10, TD03, Val07, VL10, WXBZ04, Wan04, WZY^{+}16, WFS09, WLLZ16, XL05, XLWT12, XE13, XME15, YSZL15, YL97, YDS10, ZQ99, ZJS^{+}12, ZCW15, ZL16, ZCL11, ZFC15]. scheduling-latency [IM08]. schema [Tre11]. Scheme [AGCFV18, BCO17, CJS^{+}20, CHO^{+}19, GGZC19, GZJ^{+}18, HFF^{+}24, HGZ^{+}23, JLZJ19, LWW^{+}19b, LSC^{+}21, LLM^{+}24, MAE19, QLSW19, QZC^{+}22, QLQ^{+}22, SPQZ20, SFM^{+}18, SJWH^{+}17, TZX^{+}22, WZX^{+}22, WSX^{+}23, WMT^{+}22, XZC^{+}20, YM16, YWH21, ZTH^{+}23, ZHCC24, Zha17, AA04, AJDH01, AMP01, AAM05, AB07, AB05, ABK15, AS02, ACP05, Bej09, BS97, BAL10, BBHHR10, CLC^{+}01, CSSJ14, CH97, CLG^{+}00a, EL11, GP96b, GPM03, HSH^{+}06, HA96, Hon94, IS00, IM08, KMR95, KCB03, KEY99, KqL98, LS93c, LH13, LPIH11, LSC99a, LSC99b, LBS05b, Mar04, ML06, NL99, PPV04, QS04, RSS09, RPV13, SS93, SG94, SK06, SV11, SC10, STL04, TKN06, TCS04, WM95, XSC03, XHN04, YG10, YZ10, ZB95, ZTS11, ZHLL06, ZW10]. Schemes [CLW16, CVV17, CLZ^{+}20, HRM22, KS95, LWL17, LSSC22, SS94b, VPC17, AS94, BCGC15, CSLH13, HP01, HL98a, JS09, KM10, KA95, KS03, LBS05a, LK95, MCLG07, MRD08, OJRCC02, OS03, PSA96, PP02, RPGE04, RLKT98, Rum93, TNF97, VB94]. SCI [PFC96]. science [XB07]. Scientific [NR98]. SCOQ [CM93]. SCORE [LTB04, NST^{+}16]. SCP [Smi95]. scrambler [BKH^{+}93]. Screen [XLP^{+}23]. ScreenID [XLP^{+}23]. scrubbing [WSMJ04]. SCTP [IAS06]. SDH [OSZ^{+}06, RRG10]. SDH/SONET [RRG10]. SDL [HBS96]. SDL-92 [HBS96]. SDMA [STKL01]. SDMA/TDMA [STKL01]. SDN [BBD^{+}18, CLX^{+}22, CM18, CSR^{+}20, DPM^{+}18, DJB^{+}22, HBSX20, KSC^{+}23, LLZ^{+}23b, LCL^{+}18, LTZ^{+}22, NRB22, PIST19, RZE^{+}21, SSG18, VCVC17, VVC^{+}17, WLX^{+}17, XBM^{+}23, XGQ^{+}19, XCL^{+}22, XLH^{+}17, YY20, YXH^{+}21, YLA^{+}18, ZML^{+}19, ZMLL21, ZWCL17, ZWY^{+}18, ZFW^{+}17b]. SDN-Based [DJB^{+}22, LTZ^{+}22, ZFW^{+}17b]. SDN-Enabled [RZE^{+}21]. SDN-Managed [NRB22]. SDNs [WXH^{+}18, XYL^{+}17, XYQ^{+}17, ZXC^{+}18, ZGZC20]. SDNShield [CLX^{+}22]. SDPA [SBC^{+}17]. SEAD [LXL^{+}22b]. Seamless [TCS04, ZWCL17, VVP^{+}13]. Seamlessly [PDV^{+}24]. Search [FBFB17, YGL^{+}19, AB09, CL07, CLM^{+}16, GH04, LSV99, LGC16]. Searching [YSC16, ZL13b]. Second [TL22, FqL98, LSXS16, Tia05, VFBD11]. second-order [FqL98, LSXS16, Tia05]. Secondary [CL13, AAS10, GS16, HGW^{+}16, MAS09, SL14]. Secrecy [CCMW19, ZFW14, CZF^{+}16, KES13, RCW15]. Secret [HK24, WFZ^{+}23, FHH10, ZAS12]. Sector [LCK^{+}18]. Secure [AGBS23, CDW19, FMCS20, HKC^{+}20, HYK^{+}23, HK24, HFF^{+}24, HZ20, HK14, KHH^{+}18, LWW^{+}19a, LDD21, RVS09, SVG16, WLW^{+}20, WFZ^{+}23, WSX^{+}23, WCM^{+}21, WGL00, WMT^{+}22, XHZ^{+}19, XHY^{+}22, YCC21b, ZHCC24, Zha17, FHH10, LMR07, SL07b, ZZZ^{+}14]. Securing [LAPS08, SBNRS14]. Security [BVL^{+}19, La17, LLZ^{+}17, LCL^{+}18, LCP^{+}20, LN19, NT24, PLR^{+}19, PBGMFM22, WSXL16, WXC^{+}24, XLP^{+}23, YXH^{+}21, YYL23, ZLHM22, JAW11, La16, LTS10, SKCW10, WSMJ04, XZB08, ZSFZ11]. seek [WL16]. Seeking [FXQ^{+}21]. Segment [CLP^{+}17, DPT^{+}18, HTC04, SAC^{+}18, SACH21, TWY^{+}20, LYL07]. segmentation [JYT^{+}15, MMC05]. Segmented [KLC^{+}18, ZWZ^{+}24, GPM03]. segments [HBB09, RS12]. Seismograph [ZLB17]. Sel [TZPZ23]. Sel-INT [TZPZ23]. select [BC01b, Mod99, PM96, PS94]. Selecting [XLT^{+}22, ZRD^{+}23]. Selection [BPA20, BZM^{+}22, BDR22, GHK18, HR95, JD22, KHAWC17, KCH^{+}19, LLM^{+}24, XYL^{+}17, YY20, ZWJ^{+}22, ZLZ^{+}23, ZDB^{+}17, ZQL^{+}23, BSP07, CN09, CG15b, GCH^{+}15, GMY13, JF04, KA98, KMHS09, KK03b, KT07, LH07, LWKD03, MRM99, RPV13, TNRP11, TGRR07, VC14, WS08, YWLL09, ZAFB00]. Selective [CSSG23, LYDA19, LFXY23, AHL96, GT00, KVF^{+}12, SR02]. Self [AACD^{+}96, AMS22b, CQLW22, CO94, CB97, DHS^{+}23, EF17, EL24, FLMM10, FX17, GCMP20, KS11, KLKP16, LXL^{+}22b, LYZ^{+}23b, LFF^{+}19, QZC^{+}22, QYC^{+}24, Spi97, WMCW22, WTSW97, ZSL^{+}17, ZLM^{+}23, BCP13, FCT03, GSRS^{+}15, GZDG06, HP00, KK07, KR05, LHK^{+}12, LTWW94, LGD^{+}10, LPCVC13, MK98, PYL99, SAS^{+}16b, SF95, TG97, Wu94, WWT05, XM99, ZGS10]. Self-Adaptive [DHS^{+}23, LXL^{+}22b, QYC^{+}24, BCP13]. Self-Adjusting [AMS22b, SAS^{+}16b]. Self-Certified [LYZ^{+}23b]. Self-chord [FLMM10]. Self-Competition [CQLW22]. self-configurable [WWT05]. Self-Deployable [ZSL^{+}17]. self-healing [FCT03, MK98, SF95, Wu94, XM99]. Self-Learning [EL24]. Self-Morphing [WMCW22]. Self-Optimization [LFF^{+}19]. Self-Optimizing [KLKP16, GSRS^{+}15]. self-organization [GZDG06, KK07]. Self-Organizing [GCMP20, QZC^{+}22, FLMM10, LPCVC13]. Self-Reconfigurable [ZLM^{+}23, KS11]. self-routing [PYL99, ZGS10]. self-similar [LHK^{+}12, LTWW94, TG97]. Self-similarity [CB97, WTSW97, LGD^{+}10]. Self-Stabilized [FX17]. Self-stabilizing [AACD^{+}96, Spi97, KR05]. Self-termination [CO94]. self-tuning [HP00]. Selfish [MYH21, PD16a, SLL^{+}11, BOGS^{+}16, IW08, JAW11, QYZS06]. semantics [YWZZ16]. semantics-aware [YWZZ16]. Semi [HSM^{+}13, LC96, XY09a, XL11a]. semi-Markov [XY09a]. Semi-random [HSM^{+}13]. Semi-rearrangeably [LC96]. semi-truthful [XL11a]. semiautonomous [DJ12]. semisoft [AS02]. Sender [ZDB^{+}17, ZBXH13]. Sense [KA20, SCN12]. Sensing [CZC^{+}22, CBZ16, CCX^{+}23, JYC^{+}16, JWH^{+}24, JSXN18, JZ18, JHS^{+}19, LXW^{+}20, LZY20, LLL^{+}16, LL18, LCC^{+}20, LYC^{+}19, LLX^{+}19b, SLD^{+}22, WZ16, WSC^{+}23, WLW^{+}17, YLL^{+}17, YXL^{+}19, YZZ^{+}21, ZGZ22, ZLWM18, CT04b, KNSV13, LZES14, MVRZ09, RZWQ12, ZG14, ZHC16, ZL15]. Sensitive [FKCA18, GSKN18, HTL^{+}19, HTJ^{+}21, LJL^{+}16, ML22a, ML22b, QSW24, QHZC18, TML22, XXZ^{+}22b, YCZ^{+}23, GS16, KLS11a, LL98, LNC04, RVV^{+}15]. Sensitivity [DKM^{+}17]. Sensor [ADT22, AGM^{+}17, BBR19, CWH^{+}16, CGC^{+}17, CGC^{+}18, CNG^{+}16, DPG^{+}24, DYW^{+}16, DGC^{+}20, DLLL16, GCWC17, HMM^{+}20, JWW^{+}23, JLS^{+}17, LLX^{+}17, LLZ^{+}19, LXX^{+}17, LHY^{+}23, LJL^{+}16, LLL^{+}16, LDY^{+}16, LCZH17, LXW^{+}19, MLX18, PBV17, PLM19, QZX^{+}17, QLSW19, SSY19, SLH^{+}19, SRB^{+}20, TT17, TW22, XCS^{+}18, XWJ22, YM16, YLD^{+}23, ZWL^{+}16, Zha17, ZZT^{+}17, AC16, AK09, AA05, ACCF12, BTC05, BDHR10, CLP12, ÇY07, CHML15, CT04a, CL12, CSSJ14, CZC^{+}13, CDH^{+}10, CK09, CK11, CNP13, DJ14, DLL^{+}11, DLH^{+}14, GTS^{+}09, GDC^{+}16, GT06, GIKK11, GZCX16, GAA08, GZDG06, HS06a, HLL13, HSS08, HKCL13, HY08, HMvdLM07, IKDD15, IGE^{+}03, JC13, JYT^{+}15, JL12a, JS14, KK07, KBS11, KLZ12, KLSS10, KWS^{+}11, KLS11a, KG10, KWZ08, KIR06, LGS09, LL10, LG13a, LZL^{+}14, LLNC09, LÜ14, LJW^{+}07, LLL10, LFZS11, LWR15, LHC^{+}16, LWR^{+}16, LP07, LH10, MCLG07, MHXT10, MEWP13, NLB15, ODC^{+}16]. sensor [OÇ10, PLS07, RLP06, RWA^{+}08, RKNS10, SMGP15, SGR13, SZG09, SM08, SH12, SK10b, SH07, SK13, SX10, SA05, SSA08, TXL^{+}12, TK12, TX08, TYLH09, VA06, VA09, WY06, WSC08, WA11, WVG12, WDCL15, WFS09, WMFS10, XXBC14, XSHS12, XSH^{+}15, XLWT12, YJZW15, YCV15, YHE04, YAA09, YG10, YZP^{+}14, YBX^{+}10, YBX^{+}12, ZCJ^{+}13, ZHC16, ZG08, ZXH^{+}13, ZPCS11, ZZHZ13, vRWZ09]. Sensor-Actuator [SSY19]. Sensor-Assisted [XWJ22]. Sensor-Augmented [LXW^{+}19]. sensor-enabled [YZP^{+}14]. Sensor/Controller [TW22]. Sensors [BZM^{+}22, CCG20, GFW^{+}18, KP23, MLX18, YWZG23, ZWYD18, AKK13, KKJ06]. sensors-to-sink [AKK13]. Sensory [LCY^{+}19]. Sentinel [ZXZ^{+}19]. separable [SN15]. Separating [RJCE06]. Separation [HLG94, SM16]. sequence [JID^{+}07, UZ93]. Sequences [CSL21, VL16, CU95b, CVM^{+}15, MP94, Nai97, UZ93]. Sequential [CXL18, CCK16, LWL17, XWW^{+}18, YS21, ZLWM18]. Serial [YPA19]. Served [OLZ17]. Server [BPA20, DPTP24, DAA19, GHK^{+}23, GHBSWV17, GVM23, KSRW22, KLLT18, LFZ^{+}22, LXZ^{+}21, RTLC17, VN20, WN17, ZHW^{+}17, ZWR^{+}23, ZLW^{+}19, BSP07, CG04, CJ97, DBDJ14, GCZ98, JIN^{+}12, KG99, LGW^{+}11, OKM94, RPF^{+}14, SNSW12, dSeSGM95, SLO^{+}14, SZT01, WS08, WLZ11, XL95, YLLY05, ZAFB00, ZWDS00]. Server-Assisted [GHK^{+}23]. server-centric [YLLY05]. server-side [KG99]. Serverless [TJD23, ZZL^{+}22]. Servers [AAR18, FM23, HCFC20, LLCJ22, LZZ^{+}22b, ML23, VN22, XLL21, ZCH^{+}24, AW97, CT01, GBL12, LGW^{+}11, NBK02, SV98b, SV98c]. Service [ACLX17, BSM21, BCLS17, BFG^{+}14, CTG^{+}20, CBHS20, CZD^{+}22, CWZ^{+}17, CHW^{+}20, CH21, DBW^{+}20, DKM^{+}17, DZH03, EMAL17, FMH^{+}21b, FBM^{+}21, GR20a, GXL^{+}21, GXS^{+}21, HV22, HJG18, JWL^{+}18, KS20, KZH^{+}20, LWT^{+}21, LS16, LYL21, LPM23, LS17, LW20, Ma16b, MCES19, MHH20, MFR^{+}20, NS98, PLT^{+}20, QSW24, RL18, RGY^{+}22, SMD20, SFS^{+}22, TTM22, WUZ^{+}19, WJH^{+}21, WZX^{+}22, WHYC23, WXC^{+}24, XDZ^{+}23, XS21, YZY^{+}20, ZHCL17, ZLG^{+}20, ZLZ^{+}21b, ZBdV23, ZJWY17, ZMWX18, AHK08, Ada98, ACC^{+}94, AL98, AAS14, Bar95, BTC01, BBL95, CCLT02, CLS07, CYG^{+}14, CLA07, CAL09, CF98, Cob02, Con11, CFD06, CAH08, DCGN03, DJ16, FP95, FP97, FJJ^{+}01, GS10a, GRB09, GKT97, Hon94, JDSZ97, JPS04, JF04, KA03, Kim98, KLOS09, KR99, KK06a, KK06b, LS93b, LV00, LL13, LLE15b, LLE16, LWF96, LMS04b, LV93, LFL14, MLLY06, MCL^{+}11, Mar03]. service [PD07, RRG10, RB09b, SCP99, SC09, SRS01, SYR05, She95, SG94, SLO^{+}14, SZN00, VWT^{+}14, WCH95, XB07, YBG^{+}12, YL98, YLLY05, YTL12, Yua02, ZM09, ZAFB00, ZT03, ZF96, vDP93]. service-curve [CAH08]. Service-Driven [DKM^{+}17]. service-guaranteed [JF04]. service-scheduling [BTC01]. Services [AMCD19, AEG^{+}17, CZD^{+}22, ECL^{+}20, EPB14, FLBR^{+}19, KLR^{+}20, SMEH20, SRS21, SCL^{+}23, TEE16, WFC18, WWW^{+}18, WWYY18, ZCH^{+}24, ZWC^{+}24, ZLW18, BM97, BLT02, BCGM07, CT01, CLY06, CZ06, CY14, CS00, CJJ09, CN09, DTM15, DSR02, DGG^{+}02, FK99, FT07, GV93, GM00, GVC97, GGM10, GS04, JJ08, KA03, KL95, LC97, LMS05b, LLY01, LK02, LC04b, Mar04, NS98, PPV04, PG93, PG94a, PT00, PILR05, SL94, SV11, SIYL09, SZN00, SDW00, Szy16, WXBZ04, YR01, ZSSK02, ZZZ^{+}07, DKL01]. Serving [HZCB17, LLS^{+}23, ZCH^{+}24, ZHCL17, CDI^{+}04, LEYS11]. Session [BSM21, DZL^{+}18, LWL17, BMM^{+}09, BSP07, RGKR10, Coh94]. Session-Based [DZL^{+}18]. sessions [AK01, FJL^{+}97, JVY06]. Set [FLG^{+}23, LWK^{+}16, QCMY16, SLH^{+}19, WLK^{+}17, YLS^{+}17, ZWZC23, HKLS12, JLRS16, KLT15, Lin93]. Sets [Fuk20, SCC^{+}17, XCC^{+}17, XZC^{+}19, ZCZ^{+}20, AZ06a, BNS11, MSSZ12]. Setting [PJM^{+}19, VG05]. settings [KBV^{+}13]. Settlement [WXM21, MCL^{+}10, MCL^{+}11]. Settlement-Free [WXM21]. settlements [SRP^{+}11]. setup [BV96, IPG97, Pil01]. several [HOT97]. SFC [FXHY21, TZX^{+}22]. SFCs [LGS^{+}23]. SFT [LGS^{+}23]. SFT-Box [LGS^{+}23]. SGX [KHH^{+}18]. SGX-Tor [KHH^{+}18]. Shadow [VHNPM96, LAN97]. Shallow [BHC^{+}21]. Shallow-Buffered [BHC^{+}21]. shaper [KL95]. shapers [Le 02]. Shaping [LZL^{+}18, SPR^{+}20, XSM22, ZCdV^{+}18, GGPS96]. Shapley [MCL^{+}10, XL23]. Sharded [SPLP20]. share [KCB03]. Shared [CP18, HWC22, SYR05, VTBK21, WXH^{+}20, XCL^{+}22, BT93, BL04, CM93, CH97, CH98, CK07, CW12, FJ07, GP94, GBC^{+}95, HTC04, KKV16, Kim94, KKS^{+}08, LWKD03, LYL07, MJ13, MM94, PG94b, RKT02a, SV99, SS03, ZY07b, ZY07a, ZKO93]. shared-buffer [FJ07, SV99]. shared-memory [CH98]. Sharing [AGMY21, ALPK21, AN20, ACA16, BLM^{+}17, BPL20, CP20, DCN^{+}19, FHMS18, GKCR21, HK24, HWF^{+}20, HSE97, HSS^{+}21, JCR21, LBP^{+}17, LCDW21, LSHZ16, MCES19, NJM^{+}19, NLNL16, RBPS21, SAMB18, SPQZ20, SGH^{+}19, SLD^{+}23, SC22a, SSM20, Van19, WFZ^{+}23, WSX^{+}23, XCZ^{+}17, XQG^{+}22, ZLRC20, Ali06, AdE07, BBG11, BSSS01, BMvU03, CL04, CZ06, CL13, Coh94, FCAB00, FLC09, FJ95, GSW99, GT10, HTAZ16, JR96, Kar10, KAS16, KL08, cLqL97, LCL12a, LCL^{+}13b, LZL11, LMW16, MR02, PLD16, PG93, PG94a, RPV13, RSR10, RPF^{+}14, SKY10, SSAK12, SMP^{+}14, SZN00, SRS08, TMH97, WM95, XL98, ZWYY10]. Sheep [HHD22]. Shell [SPR^{+}20]. Shell-Shaping [SPR^{+}20]. shield [RSU^{+}09]. Shielding [ZMH17]. Shift [DLR^{+}18, TWH24, CGEN98]. Shifting [FLG^{+}23, YLS^{+}17]. SHIP [SBLS19]. Shopping [ZSZ^{+}17]. Short [BK17, BBHH^{+}18, ZHT^{+}19, ZZLM23, KH15]. Short-Term [ZHT^{+}19, KH15]. Shortest [KHYA20, WYL^{+}22, ZXW^{+}19, AM16, AZ06b, CSS08, CN08, GO02, KS09b, NST00, RBC07, XCX^{+}06, YSRL11]. shortest-path [CN08, YSRL11]. shot [IW08, JK15]. Shoulders [DQW^{+}23]. SHR [hCgKsYwT96]. SHRiNK [PPPW05].

Shuffle [FLG^{+}20, YDLM20, IBM95, Lie97]. shuffle-exchange [Lie97]. shuffle-exchange-based [IBM95]. Shufflecast [DRW^{+}22]. shufflenet [GLNP01]. shufflenets [TYL94, YY98]. shutdown [SDV06]. SI [KK16a]. SIC [LYL^{+}22b]. SICS [WLW^{+}20]. Side [FLS^{+}22, GWYS19, KIW^{+}17, LLL^{+}22a, SLS^{+}23, ZWC^{+}24, ZLW^{+}19, GK16, KKV16, KG99, LP07]. Side-Channel [GWYS19, KIW^{+}17]. Sight [LJSB22]. Signal [TWL22, CH15]. Signaling [FST^{+}09, GLH95, HA96, JGKT07, LVB96, LC97, RW93, THBR14, ZS03, ZS04, ZS13]. signaling-free [THBR14]. signalized [HLP11]. signalling [IZC00]. Signals [KLY^{+}23, QYZX22, XWL^{+}18, BSH^{+}11, GH93, TZZ^{+}14]. Signature [ABBF19, CJS^{+}20, WLC^{+}10]. Signatures [HS18, WL99]. significant [CM05c]. SILK [CCY^{+}14]. similar [LHK^{+}12, LTWW94, TG97]. Similarity [NGL22, SSN^{+}23, SNC23, CB97, LGD^{+}10, WTSW97]. Simple [AB07, KM08, PK01, RBPS21, SG17a, XYA^{+}21, ZZLW16, Bej09, BTC01, CLP12, CHL16, CBAT06, CSSJ14, CLK01, FK03, GKT97, LDH^{+}12, PFTK00, SMT98, SS93, SCY98, ZTS11, ZCL11]. SimpleMAC [CHL16]. Simplification [BSRdA16, LS05a]. SIMPS [BLDF09]. simulating [FP01]. Simulation [ZMD^{+}20, AD96, And04, Con11, DT93, HAGL16, LV06, LY10, PPPW05, ST04, Val07, YKKY08]. simulations [Geo08, PV04]. simulcast [KK12]. Simultaneous [CLS^{+}21, WLY^{+}23, ZZ17]. Simultaneously [CMFA14, MLX18, XCL^{+}19]. Single [ATE22, ARK09, BPA21, CBZ16, CLS^{+}21, DJK22, DJK23, DZ18, KHYA20, LCC^{+}20, LCZ^{+}23, SNLL16, SPS^{+}02, YY20, YWHW24, ZZL^{+}21, BM93, BHN11, BB96, BBL95, CFG08, CTG00, CJ97, GS16, GS10b, Hon94, JMI95, JK05, KNP05, Kim98, KRKH10, KAMG07, LL09, LC94a, LS03a, LRL07, PDSK04, PG93, RKA08, RA95, SG96, SSFM08, SV11, SPR08b, SX10, TMMS01, YWK07]. Single- [CBZ16]. single-and [BHN11]. single-cell [YWK07]. Single-Channel [DZ18]. Single-Constrained [KHYA20]. single-copy [SPR08b]. single-cycle [SG96]. Single-Forking [BPA21]. single-hop [BB96, JMI95, LRL07, RA95, SV11]. single-hub [Kim98, LS03a]. Single-link [ARK09]. single-medium [BBL95]. single-node [KRKH10, PG93]. Single-Packet [SNLL16, SPS^{+}02]. single-relay [CFG08]. single-ring [TMMS01]. single-server [CJ97]. single-service [Hon94]. single-source [CFG08]. Single-Wavelength [YWHW24]. Sink [GCWC17, AA05, AKK13, CPSWL96, KWS^{+}11, LH10]. SINR [AKSS12, BRS10, CMP16, CJZS14, KWE^{+}10, Kuc14, QZZ^{+}13, SGJ17, TZL23, YZY^{+}18, ZYL^{+}17]. SINR-Based [TZL23, BRS10, KWE^{+}10]. SINR-constraint [Kuc14]. SIP [JIN^{+}12, SZ08, SNSW12]. SIR [HRCW08, KG05, ZY16]. SIR-based [KG05]. SIS [MLB21]. Site [CZP18]. sites [CDI^{+}04]. Situation [CWZ^{+}17]. Situation-Aware [CWZ^{+}17]. situations [RS95b]. Size [Dat17, DD24, GHBSWV17, LYY^{+}22, QJZ^{+}16, SGS20, TWN^{+}20, TRVG20, CFS06, DMS06, HLZ^{+}14]. Sizes [HKV23, Van19]. Sizing [LMSKZ99, SC95, LBS11, LLM11b, Lin93, LWAT13, PDT09]. Skeleton [LDY^{+}16]. skeletons [Bej09]. Sketch [FAS^{+}23, GTC^{+}24, HZL^{+}23, LP24, LX24, LFL^{+}23, MZZ^{+}23, TXHL23, XCQ^{+}23, YJL^{+}19, YXY^{+}18, ZXW^{+}20a]. Sketch-Based [MZZ^{+}23]. Sketches [VRR24, SLC^{+}07]. skew [LMS99]. Skewless [MMH^{+}15]. Skewness [FLBR^{+}19]. Skewness-Aware [FLBR^{+}19]. ski [KKP15]. ski-rental [KKP15]. Skype [CCY^{+}14, XYLL14]. Skype/SILK [CCY^{+}14]. SLA [CZ06, SBDR10]. SLAs [DZH03]. SLAW [LHK^{+}12]. Sleep [BSSS21, ZWYD18, WFS09]. Sleep-Wake [BSSS21]. Sleep-Wakeup [ZWYD18]. sleep/wake [WFS09]. sleeping [YHE04]. Slice [DSYN24, DQW^{+}23, JD22, SCPB19, ZLZ21a, ZAW^{+}22, WJYL16]. Slices [DSYN24, SMD20]. Slicing [AMSB^{+}24, CBdV^{+}17, CBDCP19, DRMP18, EL24, FMPS20, LKMK20, MPMC^{+}22, MNZ23, ZCdV^{+}18, ZBdV23]. Sliding [BEK^{+}22, Spi97]. SlimFL [YKB^{+}23]. Slimmable [YKB^{+}23]. slot [BB94, CEFS99, LHL15, STKL01, SS93, SS94a, SS94b, Sha97]. slots [ZVN99]. Slotted [BBF18, FZ16, HYLS21, ALJ99, CFG08, MMR09, NSS96, IZC00]. Slotted-Aloha [BBF18, HYLS21, MMR09]. Slow [GSM^{+}17]. Slowdown [GHBSWV17]. SMAC [GKB^{+}16]. Small [BAB20, CZX18, GZJ^{+}18, MPN^{+}14, NLRS21, RTNS21, WSX^{+}21, YM16, YZL^{+}18, YWW^{+}23, ASKR16, EW08, JAS10, Kuc14, MWQ^{+}10, SEMO09, SSZ05, SAS^{+}16c, VSR11, WH97, YLCP11]. Small-Cell [CZX18, YZL^{+}18, Kuc14]. SMAQ [qLlH97]. Smart [CSSG23, DLC^{+}18b, FMCS20, HH17, HHA17, LLY^{+}22, TEE16, ZCH^{+}24, KAZ01, LTS10, MMC05, STKL01, SS07, WMYR16, CS14]. Smarter [BGMB^{+}20]. Smartly [QML^{+}24]. Smartphone [LZY20, KCCM16, WZ16]. smartphones [YXFT16, DSM^{+}17, GND17, LPD^{+}18, LYC^{+}19, XLZ^{+}19]. SMDS [Lin93]. Smoking [ZWS^{+}17]. Smooth [TL16, HSG^{+}08, KKLS05]. Smoothed [JTL^{+}17, JTL^{+}18, DRR98]. Smoothing [RT99, LCY96, LV00, SZKT98]. SMS [TEML09]. SMS-capable [TEML09]. SMT [SCHG22]. SNAG [NRB22]. Snapshots [CXL18]. Sniffing [AHX19]. snoop [ML06]. SNR [CW23, LT94b]. Social [BBZ^{+}18, CGYZ16, CGYZ17, CS17, CGL16, DZ20, FZW^{+}20, FZQ^{+}22, GLZC12, GBMV21, GCX^{+}17, GCW21, HCL^{+}17, KK16b, KJG18, KKS19, KSK17, LCH20b, LCH22, LCH23, NZW24, OJSY16, QJZ^{+}16, SNZ^{+}23, SYG^{+}22, SH23, TWTD17, WLC16, WCZZ17, WGvdS17, WFY^{+}18, YSC16, YZHZ21, ZND^{+}16, ZHGF19, ZJL^{+}19, ZQW^{+}23, AAG14, CS14, CPGZ15, DZNT14, JLX^{+}16, KK16a, LZL12, LWLL16, LLW^{+}14, PES^{+}12, SLL15, WWL^{+}15, YKGF08, YGKX10, ZNZT16]. Social-Aware [LCH20b, ZHGF19, GLZC12]. social-network-aided [SLL15]. social-proximity [LLW^{+}14]. social-welfare [AAG14]. Socially [WCZZ17]. Socially-Driven [WCZZ17]. sociology [BLDF09]. sockets [YL98]. Soft [AZR97, GKB^{+}16, LXL^{+}22a, ZLWH17, JGKT07]. soft-state [JGKT07]. Software [AAR18, ACDP17, BTK^{+}17, CLK^{+}24, CPKL17, CYH^{+}18, CSR^{+}17, DBL^{+}19, DQYG23, FXHY21, FS17, FLMS18, GJD18, GSM^{+}17, GDL^{+}22, GDWX23, GDJX24, HNW17, HLH^{+}18, HSL20, KLKT16, LXZ^{+}21, MGZ^{+}23, MSM16, NJK^{+}19, PDV^{+}24, PKVI17, SM17, SM19, SBC^{+}17, SWH19, SGL^{+}22, TCTP20, TML^{+}18, TTM22, TTCT19, WMP^{+}18, WDR^{+}20, WBY^{+}17, WGZC21, XHC^{+}18, YXC^{+}18, YXCH21, YZGC23, YLK^{+}17, YXY^{+}18, ZXW^{+}20a, ZZH19, ZLX^{+}21, ZZX^{+}21b, DDPP00, Fel95, HA16, LNL^{+}16, WF93b]. Software-Defined [AAR18, ACDP17, BTK^{+}17, CLK^{+}24, CYH^{+}18, DQYG23, FLMS18, GJD18, GDL^{+}22, GDWX23, GDJX24, HNW17, HSL20, LXZ^{+}21, MGZ^{+}23, MSM16, NJK^{+}19, SM17, SM19, SBC^{+}17, SWH19, SGL^{+}22, TML^{+}18, TTM22, TTCT19, WMP^{+}18, WDR^{+}20, WBY^{+}17, WGZC21, XHC^{+}18, YXC^{+}18, YZGC23, YLK^{+}17, YXY^{+}18, ZXW^{+}20a, ZZH19, ZLX^{+}21, HA16]. Sold [BMBK21]. Sold-Out [BMBK21]. Solomon [WSL^{+}24]. SOLOR [GSRS^{+}15]. Solution [WJ17, XCZ^{+}17, CAP15, CLP12, KGPL13, MRHWS14, SRR08, XC08]. Solutions [CLH^{+}24, CAD^{+}17, FFX^{+}17, LSDT19, YYC^{+}21, BBHK14, CMN12, KHG^{+}14, MK10, SGD05]. Solvability [BSP21]. Solving [KW19, VL16]. Some [AS94, Le 02, MBRM96, PC19, SH12, ZSZN21, JK96]. Somewhat [YRB^{+}18]. SONET [OSZ^{+}06, RRG10, SZ07, ZQ00]. SONET/SDH [OSZ^{+}06]. SONET/WDM [ZQ00]. Sorted [YDLT18]. Sorted-Partitioning [YDLT18]. sorters [LC94a]. Source [CMW^{+}20, FFX^{+}17, FWK17, FLZ^{+}23, HL96a, HR14, MBM09, NKL^{+}23, LZKT99, SSR^{+}20, VAS00, ZHZ^{+}18, ZQW^{+}23, BK06, CFG08, CLS07, COS95, GV93, Hey97, KV98, KL95, LP07, RVS09, RJCE06, RL94, SAM12, WTSW97, ZY16]. Source-adaptive [VAS00]. source-based [SAM12]. source/channel [GV93]. Sources [CKA16, KP23, YEMJ24, BBM93, CP95, EM93, FNQ00, HA16, HS03, JJSS04, KWC93, LM95, LSS07, MH02, MR98, TSR14]. Sourcing [LL17b, SH23, NL16]. Southbound [ZLX^{+}23]. SPABox [FGR^{+}17]. Space [CGYZ17, CXW^{+}18, FLH^{+}17, JR21, KA20, LH95, LFL^{+}23, MBL19, SXEZ21, WSXL16, WLW^{+}17, AIN^{+}15, GP98, LTS10, PLT14, ST04, SM00, SSFM08, WXR13, WXC16, WXW15, ZNN^{+}10]. space-based [SM00]. Space-Efficient [LFL^{+}23]. Space-time [LH95]. Spaces [KGH^{+}20, SRI^{+}18, LQ13]. SPAF [RSR11]. Spam [ZGY^{+}16]. Span [CHO^{+}19]. Spanner [YNZ^{+}17, ZYL^{+}17, SS10]. Spanning [ZLTX17, GIKK11, GR16, QGCL11, YRO16]. spare [HBU95, HM04, KD10, LTS05, XM99]. spare-capacity [HBU95]. Sparse [DLLL16, HWZ^{+}23, SWL^{+}18, VIT21, WZL^{+}23a, ZSZN21, ZSK12, DPMK11, SSM06, SAS96, WLL^{+}11]. sparsely [ZLW16a]. Sparseness [YNZ^{+}17]. Spatial [AKSS12, BD07, CBdV^{+}17, GHRH18, LR22, MQL^{+}22, SYL^{+}17, VA06, WA11, XCL^{+}18, ZCZ^{+}20, ZMMG22, BRM^{+}13, CW10, CGGS97, HSPH09, HKCL13, NSW11, RW96, TWL04, TG96]. Spatial-Temporal [MQL^{+}22, SYL^{+}17, HKCL13]. Spatially [CWAO21, KW19, ZKH10]. Spatio [BTC05, PS09, RZWQ12]. Spatio-temporal [BTC05, PS09, RZWQ12]. Spatiotemporal [CEC^{+}19, HZ20, KWH^{+}17]. special [CCE^{+}06a, CCE^{+}06b, Tow06a]. Specialize [XCV^{+}20]. Specialized [CBV^{+}18]. specific [LLW^{+}09, WEK97]. Specification [HBS96, LT94b, CDO97, OdG97, SR02, TNF97]. Specifications [CWZ^{+}23, KLNS93, MP94]. Spectra [OKAS23]. Spectral [BMB19, SL94, FHT^{+}10, qLH93b, qLH93a, PJ13, SKK07]. Spectrally [CWAO21, KW19]. Spectrally-Spatially [CWAO21]. Spectrum [ALPK21, AAF^{+}16, BPL20, CGYZ17, CP18, CP20, DLC^{+}18b, DRQ^{+}16, GSPV^{+}18, GKCR21, GT10, JD17, JZ18, KS10, LHL^{+}21, LZY20, LSL^{+}18, MMG22, NBV17, QDD^{+}17, SAMB18, SAKMB21, SA21, SPQZ20, SGH^{+}19, SC22b, SSM20, WHC^{+}19, WZZC17, ZCPG^{+}23, ZYZ16, ZLZ^{+}23, ZLWM18, AAG14, AAS10, CZ12, CL09a, CL13, GS16, HGW^{+}16, JGMB03, JL12b, KYY^{+}12, KS12, MGCK15, MAS09, PWK^{+}13, RPV13, SKY10, SC09, SL14, SK12b, WHTC15, YKZ^{+}13, ZWTC16]. Spectrum-Aware [DLC^{+}18b]. speculative [IM08]. speech [MBM09]. Speed [BHC^{+}21, BSM21, BWK^{+}22, DLW^{+}17, EBJM18, HSM^{+}20, HSM^{+}21, LXW^{+}17, OJSY16, PJM^{+}19, RAA^{+}24, VIT21, VN20, VN22, WMO^{+}23, AACD^{+}96, AAZZ12, BS97, BK00, CCL99, CS98, CGS93, CGEN98, CT96, EM93, EVF06, FqL98, GLH95, GP96b, GGK99, HM06, HKT95, IK07, ILS97, KV96, KL13, KCCM16, LS93a, cLqL97, LH95, LMNM01, LYS93, LCH95, LLS07, LNM^{+}09, LLE15a, LBS05b, LT94b, LXX^{+}14, PLT14, RW07, SFAS05, SLC^{+}07, SS03, SSZ03, SXLL08, XW11, YLCP11]. speed-up [LMNM01]. Speeding [KR20]. Speedup [HYZH16, AD96, Kok10, MSS02, TT09, WYHL09]. SPIN [HTJ^{+}21]. spite [Cob02]. SplayNet [SAS^{+}16b]. Splicing [BSM21]. Split [HWHW18, SRCDL19, KD00, PGV16, XHN04]. Split-Central-Buffered [SRCDL19]. split-connection [XHN04]. split-incapable [PGV16]. Splitting [ATE23, DPTP24, ZLW^{+}17, BIS00, LL09, SSM06, WQC06, WTXT11]. Splitting-Aware [ZLW^{+}17]. Sponsored [LSSK17, VNM22]. Sponsoring [JWSH18]. spoofed [WJS07]. Spoofing [HHD22]. Spot [MAS09]. Spraying [HHL^{+}19, HLL^{+}21, BWS10]. Spread [HSM^{+}20, HSM^{+}21, WMO^{+}23, CFZ97, VOK09, YLCP11]. Spreader [LCL^{+}13b, MCZ^{+}22]. Spreading [CXL18, CP17, SSV13, fTL06, VNS02]. Sprout [ACLX17]. SPSA [BFMF01]. SPT [NST01]. Squeezing [WWW^{+}20a]. SQUID [SPC10]. SRLG [SYR05, ARK11]. SRLGs [ZJ12]. SRM [LESZ98]. SRPT [VN20]. SRR [Guo04]. SRv6 [RGY^{+}22]. SS7 [Rum93, RS95b]. SSED [AAR18]. Stability [CLCL23, CMR17, Dai22, HYLS21, JSZ14, LLCL11, LJA14, MMT16, MJ13, RMPG16, Tia05, TTCT19, VHT21, Voi07, ZHGF19, DKL01, AZ03, AOM04, AEJV13, BLPS10, CDRV11, FP14, GPLT15, JT01, LV06, LMNM01, Lie97, LLS09, LE06, PWDL05, RLA06, SLD14, TWLC10, YS93, YDS06b, ZKL07]. stabilization [AZ06b]. Stabilized [FX17]. stabilizes [TG96]. Stabilizing [GCH^{+}15, AACD^{+}96, KR05, LBS05b, Spi97]. Stable [AGGT16, ESP05, GR01, LWT^{+}21, OAN15, RBPS21, SdVK16, TZL23, XRL^{+}22, YZY^{+}18, AB05, CLK01, GSW02, JMS08, KNK^{+}14, KG16, YXF^{+}13]. Stack [LDS^{+}24, NSC^{+}22, SL17]. Stacked [BS19, SSFM08]. Stackelberg [DJB^{+}22, KLO97]. stacking [JSuRKH03].

Stadium [BMBK21]. Stage [CWGT14, HAB^{+}22, LLM^{+}24, SRCT23, XZC^{+}20, ZQW^{+}23, BHN11, HY10, HL00, KD00, LHZ^{+}16, SYP01]. staging [ZWDS00]. STAIR [BKLM06]. staircase [TCS04]. Stakeholder [SFS^{+}22]. Stakeholders [JWSH18, LSK20]. stale [SRS01]. Staleness [LCL16]. Stall [AAA18]. Stalls [ZLW^{+}19]. stamp [SA01b, WPZM16]. Stamping [SL16a]. Standing [DQW^{+}23]. Star [LYC11, DS99, LA95c, LS01, PM96]. Star-block [LYC11]. StarFront [LLZ^{+}23a]. Stars [LLZ^{+}19]. Start [ALYX22, DQW^{+}23, GVC97]. Start-time [GVC97]. Start-Up [ALYX22]. Starvation [VKO17, GSK08, GMSK09, Sha97]. State [AGMY21, CFC^{+}24, CCZZ17, CL19, CMP^{+}14, GHZ^{+}20b, HCL18, LHW^{+}20, MGK20, SXEZ21, TKM20a, TTM22, YN19, ZK19, AKA10, CLW95, CKR93, DW11, FB07, JGKT07, KK03b, LRC15, LB04, LWR15, MWQ^{+}10, Nai97, OdG97, QY12, RZC11, SRS01, SKV03, SZM08, VVP^{+}12, XHN04, XCR11, XCR15]. State-Dependent [AGMY21, TTM22, CLW95, CKR93, LB04]. State-Free [CCZZ17]. Stateful [BSM21, CKLS22, SBC^{+}17, VKPI17]. Stateless [CKLS22, LNG^{+}21, CB11, RSR11, SSZ03]. states [Kop96, LA95a]. Static [CAZG20, CV12, CNM^{+}17, LT02, TZCB23, CKL16, EM09, ITSO01, LYWL08, LS09, MWQ^{+}10, Mar04, PL02, WCY04, WXBZ04]. static-priority [ITSO01, WXBZ04]. station [AKSS12, GT00, LMS06, PT96, SH12, SKS16]. Stationary [GA24, TWL22, AAB05, LV06]. Statistical [CBdV^{+}17, CL03, DT93, GJCB18, KR08, cLqL97, MBA06, NMD^{+}17, PBGMFM22, RLP06, SD00, ZCdV^{+}18, CP95, CBL06a, FqL98, KKP15, LM95, Lee96, qLlH97, LMS04b, NR13, RRR02, SMH95, SGR13, SL94, WTSW97, WM96]. statistical-matching [qLlH97]. statistically [GV93]. Statistics [XYQ^{+}17, BCGC15, DLT05, HLZ^{+}14, HXLZ11, SHN16, WZLX12]. Status [BSSS21, FLG^{+}20, MKAE20]. Stay [SZ22]. Stay-or-Switch [SZ22]. Steady [QY12, XHN04, DW11, SKV03]. Steady-state [QY12, XHN04, DW11, SKV03]. Stealing [Van19]. stealth [DKC^{+}15]. Stealthy [CH21]. Steering [GHK18]. Steganography [ZHCC24]. Steganography-Based [ZHCC24]. Stein [FM06]. Steiner [AC98, CAP15]. steps [Geo08]. Still [LLX19a]. STINs [LTZ^{+}22]. Stitching [SWL^{+}18]. STM [IMG98]. Stochastic [ADR18, BM22, CCMW19, FK13, HLP11, HCL18, KJG18, LFL14, MW05, MMP17, PRH17, SAKMB21, SA21, SRS21, SE21, SLD^{+}22, WWC^{+}18, WLL01, XPL^{+}17, XC08, AAB05, BBM93, CE08, FMMR10, HN10, LYRL07, LRL08, NML08, Nee16a, NCT14, ORS93b, SKKA01, SR01, VG05, WWL02, XY10b, YAA09]. stop [LZ09, QY12]. stop-and-wait [LZ09, QY12]. Storage [ACLX17, AAA18, ECL^{+}20, GGZC19, HK24, HZLZ22, IKS17, LMD16, LK16b, LJL^{+}16, LS17, LSC^{+}21, LFC^{+}22, LPWP22, PLT^{+}20, RV21, EGKM16, SGVO18, WFZ^{+}23, XLAC16, YCC21b, AK09, BM97, DPR06, LK14, MPFK02, PT10, SK13, YJZW15]. storage-efficiency [PT10]. Store [ZLG^{+}17, CD96]. stored [SZKT98]. Stores [WSL^{+}24]. Storm [LWW^{+}19b]. Story [DLY^{+}22]. STPP [SYL^{+}17]. Straggler [AS19, BPA20, BPA21]. Strategic [GS19, HSS^{+}21, OJSY16, La16]. Strategies [CEC^{+}19, CEFS21, JK21, KLKP16, LW17, MBI^{+}17, SSK^{+}17, AC16, AAS10, HPR06, JK96, KLO97, KK06a, LS93b, LO02a, LS97c, MV14, Ram08, TAB^{+}15, VGKG10, XM99, ZZZ^{+}07, ZCL11, ZM04]. Strategy [QZX^{+}17, YZL^{+}18, AVPG14, JR96, LMP08, MHRR12, QSS^{+}15, SCY98, WHTC15]. strategy-proof [WHTC15]. stratified [Kar10, RP06]. Stream [CLS^{+}21, FDM^{+}17, XCZL20, KS13, PS98, SJ95]. Streaming [AAA18, AAAR19, BTD^{+}17, BBR^{+}22, CZM^{+}24, EAH^{+}18, EFA19, GWYS19, JSZ14, KHG^{+}14, KCM16, LKS^{+}16, LBP^{+}17, LLT^{+}16, MRR^{+}14, QWL21, SQ16, TPW^{+}18, TH21, TL16, WZH^{+}24, YWW^{+}24, ZZLW16, ACKZ14, CC06, CJW11, CZCC14, DM14, DYX12, FHSZ13, GMY13, JBR16, KL07, Liu10, MR09, MEVSS03, OWKS16, PWMC12, SLL15, SHN16, VNS02, VAM^{+}06, WXR13, WLCW16, WCAB15, WLZ11, ZSCJ14, ZEV07a, ZEV07b, ZLW16a, ZCL11]. Streams [DSL^{+}18, HH18, LDW^{+}20, LX24, PWL^{+}22, RWL^{+}22, RDR17, XCQ^{+}23, BD97, BS02, CM05c, GZT03, HL03, HH10b, SLC^{+}07, WD05]. street [LK95]. strength [CH15]. STRESS [HGE04]. stressed [BF01]. Stretch [YNZ^{+}17, LQ13, MWQ^{+}10]. Strict [LJZ^{+}23]. Strictly [JPH08]. striding [ARS16]. String [YDW18, NST01]. striped [DLPT06]. Strong [LLWB16, Tur09, ZHGF19]. Structural [CLL^{+}18, JLSB16, MP94, JL12a, PJ13, SMH95]. Structure [CGYZ16, FBFB17, FZW^{+}20, LHW^{+}20, SXQ^{+}23, WYZ^{+}24, BSS11b, DPBT11, DMS06, KLPS06, OPW^{+}10, OGLK14]. Structure-Aware [FBFB17]. Structured [CSN^{+}23, FLBR^{+}19, BFMF01, BQ08, KEAAH08, LCW05]. Structures [GYSZ19, VNM22, FDG^{+}10, MJ13, SJ12, VL97]. Structuring [BS02]. STS [BKH^{+}93]. STS-N [BKH^{+}93]. Student [MGS^{+}21]. Study [ACZP21, BMBK21, CWGT14, CZGKB24, FAF^{+}17, LS97a, LXW^{+}17, LSSC22, TH21, WWW^{+}20a, XNHM22, ZML^{+}19, ZMLL21, AT03, BM00, CLSC15, DYH13, ESG11, FST^{+}09, HJL^{+}12, HL98b, IZC00, KYY^{+}12, Kon06, KEAAH08, LS93b, OSW97, RrBG94, SML04, SZM08, SENB09, WLS97, XG05, XYLL14]. stuffing [CB99]. style [AB05, VGKG10]. sub [BFF07]. sub-50 [BFF07]. subcritical [GGL09a]. Subexponential [TWN^{+}20]. Subexponential-Size [TWN^{+}20]. Subject [CGAC20, QYZX22, NT00, XSZ^{+}07, ZWYY10]. Sublinear [YEMJ24]. Submarine [WWMZ20, WWMZ22]. Submodular [LYW^{+}21, XXW^{+}23, KLT15]. suboptimal [LLCL11]. Subscribe [BTK^{+}17, EPB14, CJV16, MJ14, OR11]. subscribers [GMZR13]. subscription [GJVZ06]. Subscriptions [JFM^{+}22]. subset [AB09]. Subsidization [Ma16a]. substitution [CDS02, PL02]. substrate [KMZR12]. Subtasks [BPA21]. Successive [HWM^{+}24, LTS05]. Succinct [LS09]. suffice [SX10]. Suffix [HWHW18]. suitability [LZSS10]. suite [BFM^{+}96]. Sum [HS14, HS16, Far95, McA94, TCS13]. Sum-Queue [HS14, HS16]. summaries [KM08]. Summarizing [LX24]. Summary [FCAB00]. SUNOS [PP93a]. Super [CZM^{+}24, LY22, MCZ^{+}22, GGL09b]. super-critical [GGL09b]. Super-Halfin [LY22]. Super-Resolution [CZM^{+}24]. superimposed [WM16]. superior [PT10]. superlinear [BLC11]. superlinearly [BS08]. Superposition [YKB^{+}23]. Superspreader [TXHL23]. supervised [HFC^{+}13]. supplemental [BK06]. supplementary [JWSH15]. Supply [QZL^{+}16]. Support [AMCD19, BVL^{+}19, LAJ20, MCMdlO23, Ada98, CPSWL96, GCZ98, KLSV12, SWKA01, YW11, YL98, ZM04]. Supporting [FKCA18, HGG06, Ram08, SZKT98, ZWZC23, ACR12, BM97, CJJ09, CL09b, FT07, Lin93, PGV16, RVS^{+}02, WM96, YD04, DKL01]. Suppression [RBPS21, HGE04]. Surface [LDY^{+}16, YLK^{+}17]. Surfaces [SRMB^{+}23]. surge [CLSS09]. surjective [FJ07]. surjective-mapping [FJ07]. surrounding [LLNC09]. surveillance [LJW^{+}07, YKR11]. Survivability [EM09, YO17, YXZ17, LML11]. Survivable [ACA16, HMM11, OSZ^{+}06, PBT^{+}20, ZLTX17, AM16, Ali06, BO07b, FCT03, HBU95, HC07, IMG98, KNP05, LGC16, LYC11, LTS05, MK96, SJ12, SZM08, YRO16]. survive [RS05]. SUSE [PT10]. Sustainable [LFC^{+}22]. Sustaining [ZLY23, AWKN16]. SVC [EAH^{+}18]. swapping [CO94, Coh94]. Swarm [DNS23, DC13, DPBT11]. swarming [MDL^{+}13]. Swarms [LDL^{+}22]. Sweep [GFW^{+}18]. SWEET [HZCB17]. Swing [VV09]. Swipe [LCZ^{+}23]. Swiss [NLRS21]. Switch [CWGT14, HBSX20, SZ22, SRCDL19, YDY^{+}24, AMI^{+}07, AMKY99, BL94, BS00, CL03, CC95, CM93, CAH08, GSD09, IM03, JK96, KJF^{+}00, KR00, Kim94, KK03a, LS06a, LK10, MS03, Mne08, OWMM97, ODC^{+}16, PYL99, RCOC03, She95, WY95, WLL01, YCL09, YZ10, Zal09]. Switch-Based [YDY^{+}24]. Switch-Controller [HBSX20]. Switched [FZ16, QYZX22, ZP18, BO00, BV10, BTC01, CHA95, Coh94, FGK10, FRC98, FCT03, GT00, JM00, LT02, MDMM09, RZZ06, SEMO09, SV98a, Tha04, WCY00, ZJS^{+}12]. Switched-Beam [ZP18]. Switches [BCER20, CSA^{+}21, CCCC17, CSR^{+}20, Dat17, HYZH16, HWC22, LZW^{+}21, SSG18, SHV^{+}23, TZX^{+}21, TWN^{+}20, VIT21, WZL^{+}23b, XLD^{+}24, YZLH17, ZZX^{+}21a, AZ03, ACP05, BHN11, BS00, CT95, CLY06, CH97, CH98, CMFA14, CDM93, GKS05, Geo08, HM06, HSG^{+}08, HY10, JMS08, JAS10, KKLS05, Kok10, LS94, LA95b, LLLS07, LMNM01, MBG^{+}02, MBG^{+}03, McK99, MS95, MSS02, NMC07, NPQ06, NMH99, OJRCC02, Pad95, PB93, RCGT06, RB09a, SV99, SPC10, SM00, Smi02, Smi08, TGT01, TT09, TD03, WYHL09, ZY07a]. Switching [AB23, FMH^{+}21a, KA20, KAA^{+}18, LHL^{+}23a, MSS02, QYZX22, QFH^{+}18, SZ22, SLF21, XHC^{+}18, ZWZ^{+}24, ZDZ^{+}24, BM93, BT93, CAQ07, CqLL98, CH93, CHCH00, CCL09, CSS^{+}14, CFS09, CT96, GKS05, GVC97, HSG^{+}08, IKDD15, LL95, LQXX07, Lia06, LWT^{+}15, LNC98, LC94b, MSH95, MHSC95, Mne08, NML98, NPY07, PMH95, QY04, RrBG94, Ses97, Sha94, Tas99, Tha01, Tha04, WH97, WKZL96, ZGS10, ZKO93]. sybil [YKGF08, HWJZ21, PSST21, WWW^{+}18, YGKX10, ZZS^{+}16]. Sybil-Resilient [PSST21, ZZS^{+}16]. SybilGuard [YKGF08]. SybilLimit [YGKX10]. Symbol [ODJ23]. Symmetric [XZL^{+}21, ZVN99]. Symphony [RKZG10]. Synchronizable [CU95b]. Synchronization [HKS16, LGW^{+}17, EGKM16, WGL22, Ber00, EPD94, FJ94, HS06b, LSW15, MMH^{+}15, RVB12, SKR^{+}09, SA05, VRK09, ZLS96, ZS03]. Synchronizations [ZMLL21]. Synchronize [XCL^{+}18, Lev95]. Synchronized [ASSK13, SLWW19, RR93, WFS09]. Synchronizing [TKZ94, Mil95]. synchronous [BIV01, BSSS01, BD97, CHA95, CK07, OSW97, RKZG10, WF93b, WTS^{+}13, ZB95]. Synchrony [JE18]. Synergy [JK21]. Synoptic [HFC^{+}13]. Synthesis [TR17, ZNN^{+}10]. Synthesizing [CWZ^{+}23, MBI^{+}17, ZJL^{+}19]. Synthetic [BMB19]. System [AHX19, APSG14, AAG^{+}16, BMBK21, BWG^{+}20, CXZ^{+}22, CLY^{+}17, CW19, DPG^{+}24, DLY^{+}22, DYJ^{+}23, GGZC19, GND17, HDQ^{+}16, KSRW22, LLT^{+}16, SVL^{+}16, VLMN09, WCC14, XCC^{+}17, XZC^{+}19, XCL^{+}19, YYT23, YC12, ZZS^{+}16, ZWS^{+}17, ZSL^{+}17, ZSZ^{+}17, ZCM14, AS09, AYS^{+}13, AKS^{+}13, BAC12, BLCT97, BGJ^{+}04, CSC94, CCLT02, CFZ94, CS99b, CTVD14, DM14, FGM^{+}13, Gao01, GBC^{+}95, HLSG04, HN10, JBDF07, LC97, LCH^{+}06, LY94, LCL13a, LZES14, LFV10, LHC05, McM95, MRD08, PBKG11, RD11a, SZG^{+}13, SL15c, SLL15, VGP14, WH97, YL98, YW07, YNDM09, vDP93]. System-level [YC12, RD11a]. system-theoretic [LFV10]. Systematic [APC21, SX16, CLSC15, LMT10, SSR^{+}11]. Systemizing [YLK^{+}17]. Systems [AAA18, AAAR19, ALYX22, AHEK24, BRY^{+}19, BSP21, BCD19, CCF17, CP18, CP20, CES22, DAA19, DME23, DLR^{+}18, DJB^{+}22, EBJM18, FHMS18, GLM^{+}16, GLC^{+}16, GLY17, GLLL17, GCZY18, GSKN18, HPP^{+}23, HV22, HOZL16, HJL^{+}20, HH17, HBSX20, HGZ^{+}23, JSXN18, JHS^{+}19, JD19, LFZ^{+}22, LMD16, LXL^{+}17b, LLG^{+}17, LCX^{+}19, LXL^{+}19, LXW^{+}19, LCL^{+}20, LX21, LPWP22, LYL^{+}22a, LHL^{+}23b, MZK^{+}17, MFT^{+}20, MHL19, NJM^{+}19, NLT^{+}18, OBS17, PPS^{+}22, QYZX22, SPLP20, SQ16, SC18a, Van19, VKO20, WLL^{+}24, WN17, XXCC17, XOYL20, XZL20, XQG^{+}22, XSM22, YLL^{+}17, YGL^{+}19, YWH21, ZCPG^{+}23, ZCPP22, AZ11, BB94, BSNI06, BS09, BNS11, BBL95, BMS14a, BSP07, BQ08, CKR^{+}09, CqLL98, CHCH00, CPS^{+}12, CZCC14, CLM^{+}16, CHLS07, CJJ09, DM15, DEH^{+}07, EF08, FUDA03, FLMM10, GSN^{+}16, HL99, HK94, HS03, HS08, HLP11, Hon94, HG14, KAEAS14, KD10, LVB96, LMS05a, LBHO07, LZSS10, LDH^{+}12]. systems [qLP97, LZZR12, LZL11, LPP11, LS97b, LS05a, LJW^{+}07, LCW05, LCQL14, MBC^{+}94, MV08, MDL^{+}13, PLD16, PD07, QCLC16, QS05, RW07, RD11b, SNS12, SHZ16, SWL06, SKG12, dSeSGM95, SS96, SS04b, SRS08, TM13, TAB^{+}15, WF93a, WXR13, WLR10, XSC01, XSC03, YZP^{+}14, ZGG05, ZLW16a, ZL13b, ZL14, dAF04].

T [AB21, PYL^{+}17, SJWH^{+}17, WSX^{+}21]. T-Cache [WSX^{+}21]. T-Chain [SJWH^{+}17]. T-RACKs [AB21]. Table [SSG18, SGS20, YXH^{+}21, ZXC^{+}18, AN05, ZGG05]. Tables [CNM^{+}17, LS05b, LS10, PT10, PT12, RTK^{+}16, XLZC14]. Tackling [ACDP17, AST11]. Tactile [PPS^{+}22]. Tag [GLY17, LCC^{+}20, QCLC16, SYL^{+}17, WLL^{+}24, XCC^{+}17, XWY^{+}18, XZC^{+}19, YGL^{+}19, ZCZ^{+}20, ZCZF20, CLM^{+}16, LL09, LHL15, LCQL14, ZL13b, ZL15]. Tag-ordering [QCLC16]. TagAttention [SCW^{+}21]. Tagbeat [YLL^{+}17]. Tagcaster [ALYX22]. Tagger [HZC^{+}19]. Tags [CCZZ17, HDQ^{+}16, JWH^{+}24, LXL^{+}17b, LXL^{+}17a, LCL^{+}20, OLZ17, SL16b, SSH^{+}23, ZCZF20, HQY^{+}16, HQW^{+}16, LCL13a, SL15a]. Tahoe [SKV03]. Tail [CDL^{+}19, RMPG16, TSS14, NJW16]. TailCutter [CDL^{+}19]. Tailed [LWAL17, MMT14, MMT16, ZHCL17, BMvU03, JMMT12, LGD^{+}10, NAA^{+}16, NJW16]. Tailored [MCM^{+}23, SMD20]. Tailoring [SSK^{+}17]. Takes [RR19b]. Taking [Bej09]. Tale [CVHM22, LDZC20, LLX^{+}19b]. Talk [ZWGC17, WS05]. Taming [CLWZ17, HZL16, LGDC18, TRKN12]. Tandem [RR19b]. Tango [RR19b]. Tapping [TWWG19]. Target [BMB19, GCWC17, Van17, YSC16, ACCF12, CDH^{+}10, SG13, SH07, YZP^{+}14, ZG08]. Target-Oriented [YSC16]. Targeted [LCH22, HBB09, KLMW11, KK06a]. Targeting [TMGB19]. Targets [CCW^{+}17]. Tash [LYDA19]. Task [CLZ^{+}23, JD19, LHZ^{+}19, LDL^{+}22, MHR^{+}20, SC22b, TJD23, WCL^{+}22, YCC^{+}21a]. Task-Aware [LHZ^{+}19]. Task-Based [LDL^{+}22]. Tasks [ALR^{+}24, CBV^{+}18, DMLC18, JD20, SE21, YPA19, ZLM16]. TCAM [BBHHR10, BBHH^{+}18, CSLH13, CW16, HZG^{+}18, HWHW18, LMT10, MLT11, MPN^{+}14, MRM17, NLT^{+}18, RKH^{+}16, WSX^{+}21, WXC16, ZHLL06]. TCAM-Based [HWHW18, NLT^{+}18, CW16, MLT11, ZHLL06]. TCAMs [LMT10, LS10, MLT12, SRK22]. TCP [CBAT06, AB21, AB23, AEG^{+}17, AMP01, AAB05, AT03, BH05, BHC^{+}21, BPSK97, BLPS10, BC01a, BV05b, BHL^{+}06, BBM^{+}10, BSS^{+}11a, CQW^{+}18, CM12, CDFG06, CBD02, CLM99, CMR17, CL16a, CR98, DW11, EL11, EW08, cFKSS99, FLS^{+}22, GLMM04, HSH^{+}06, JD03, JGMB03, KV98, KVR98, KVR02, KK05, KP96, KLVL19, Kum98, KK06a, KK06b, LM97, LMS00, LLS07, LXW^{+}17, Li24, LBS05b, LHZ^{+}19, Low03, MBA06, MGG^{+}05, MNR03, MMC05, MSBZ10, MG95, ML06, PFTK00, PP93a, PMW10, Pax94, PWHL16, PPV17, PWWP18, PCW23, PDT09, RUH^{+}18, RMPG16, RAL04, RLZ10, RKPP16, SM14, SKKA01, SCR08, SWL06, SKV03, SR02, SHHP00, SXEZ21, TKXP20, TL06, VSR11, WLLD05, WJLH06, WXH^{+}20, WFGZ13, WCW^{+}17, YSL^{+}14, YR01, ZWH^{+}17, ZLW^{+}20, ZLW^{+}19, ZRK06, ZHWH21]. TCP-compliant [BLPS10]. TCP-friendly [JGMB03]. TCP-like [CBD02, SWL06]. TCP-LP [KK06b]. TCP-Peach [AMP01]. TCP-RED [RAL04]. TCP-targeted [KK06a]. TCP/AQM [EW08, SCR08]. TCP/IP [AAB05, KP96, LM97, LMS00, PP93a, WLLD05]. TCPSbed [PPS^{+}22]. TD [Wan04]. TD-CDMA [Wan04]. TDM [BD97, Tha01, ZA11]. TDMA [CS99a, DHSS14, DV09, STKL01]. TDMA-based [DHSS14]. Teacher [MGS^{+}21]. Teacher-Student [MGS^{+}21]. Team [XLX^{+}21]. technique [CHLS07, FUDA03, KLS11a, WWT05, ZBXH13]. Techniques [JHL22, SBTH19, BMM^{+}09, BP96, CSS08, DRM04, GZT03, GS97, KR08, KT06, RR93, SXLL08, TBV^{+}13]. technologies [ALMR14, JKJ13, JWSH15]. Technology [CWLH20, CLS^{+}19, GHZ20a, GHZ^{+}20b, HLZ^{+}21, JYL^{+}19, KIW^{+}17, XZL^{+}21, SJGH10]. TEE [ZWC^{+}24]. TEE-Assisted [ZWC^{+}24]. telecom [HMM11, SZM08]. telecommunication [LC97]. telecommunications [KA03, MOZ05, ZWO^{+}96, dFV02]. teleconferencing [RB95]. Telemetry [JJJ^{+}23, TZPZ23]. Telephony [CHW^{+}20, GS04, XYLL14]. teletraffic [Lee96]. Templates [ZGY^{+}16]. Temporal [LCL17, MQL^{+}22, RZS14, SL17, SYL^{+}17, TT17, ZCZ^{+}20, BTC05, HSPH09, HKCL13, PS09, RZWQ12]. Temporal-Spatial [ZCZ^{+}20]. Temporally [NDN^{+}18]. Tenant [CBdV^{+}17, CBDCP19, TZX^{+}22, WZX^{+}22, CYG^{+}14]. tenant-directed [CYG^{+}14]. Tenants [ISS22]. Tenet [BFM^{+}96]. Tensor [DZL^{+}20, TXW^{+}21, XLW^{+}17a, XWW^{+}18, XWW^{+}19, XCW^{+}20b, XOW^{+}23]. Term [BK17, DSYN24, WDL^{+}23, ZHT^{+}19, KH15, SENB09]. Terminal [HR14, QYZX22, ZSZN21, BB95, KD10, XHN04]. terminals [JS12, VA07]. Terminating [GS04]. termination [CO94]. Ternary [KLC^{+}18, NLT^{+}18]. Terrain [CXK^{+}23]. terrestrial [ZRK06]. terrestrial-satellite [ZRK06]. test [CU95b, MP94, UZ93, ZKVM14]. Testbed [PPS^{+}22, KKM^{+}97]. Testing [CLH^{+}24, HLP^{+}16, ZZH19, ZZX^{+}21a, HLSG04, HKLS12, HBS96, LM13, LCH^{+}06, SMV93]. tests [FUDA03, MP93]. Tethering [LS16, HLS^{+}14b]. Their [GBMV21, FK07, Far95, LMP96, MCS99, McA94, SKG12, dSeSGM95, TLS^{+}12]. theorem [CS98, Su15]. theorems [HBH93, WJK06]. Theoretic [LCH20a, LCSS17, LCS^{+}18, MGLH18, WBM^{+}18, vDJJ^{+}22, BGSSW13, CL16b, DJ12, DM96, EML12, GSA15, KR99, Kon06, KK12, LFV10, LyT98, LRL07, LCW05, MLLY06, NOF14, RSS09, She95, SBP03, SM05, SXLL08, VT12, YMR00, YXF^{+}13, ZRLD05]. Theoretical [AB23, CSL21, CLV17, Dai22, DNS23, SWH19, CCG00, CSMW02, CGM04, KL13, LWT^{+}15]. Theoretically [LZZ^{+}22b]. Theory [ALR^{+}24, AWM^{+}20, AN20, HZG^{+}18, JLSB16, Le 18, ML18, NGL22, TXW^{+}19, YLL21, BCR^{+}12, CCE^{+}06a, CCE^{+}06b, CRB09, CCLT02, CL04, CG97, ES05, FHT^{+}10, GO99, KLT15, MRD08, RV01, SL05, SRP^{+}11, Sob05, SQ12, Tow06a].

Things [LCP^{+}20, QML^{+}24, BCS^{+}19, CQLW22, FWN^{+}22, LWW^{+}19a, QZC^{+}22, SDM20, YXZ19, YYC^{+}21]. thinnest [GZS15]. Thou [GNK^{+}21]. Thread [LFZ^{+}22]. Three [CWGT14, CVHM22, KL95, SAMB18, GR16, HL00, KD00, LPF12]. three-dimensional [LPF12]. Three-level [KL95]. Three-Stage [CWGT14, HL00, KD00]. Three-Tier [SAMB18]. Threshold [MSRG18, QLY23, WLY^{+}24, LS93c, LQCC16, NL99]. Threshold-Based [QLY23, LQCC16]. Thresholded [LDW^{+}20]. thresholds [CH98, HC02, RrBG94]. throttles [LT95, YLLY05]. Throughput [BLC21, BBF18, BVBV17, CCE^{+}17, CZGKB24, CHM^{+}05, CLS^{+}18, CGR^{+}18, CH18, CGZL20, CFS11, CCCC17, GB18, GGC93, GKCR21, JJS13b, JPS^{+}17, KSM19, KIR08, KNSV13, LLE15b, LK16b, LYS11, LZC20, LZC22, LYKT21, MAE19, MYMY17, MSRG18, NL07, PG21, SL12, SPLM17, SPM^{+}17, SSM20, SGJ17, TWH24, XY10b, XLH^{+}17, XSZ^{+}22, YS15, YN20, ZGYB20, AP93a, AWKN16, BM08, BZM08, CCG00, CBD02, CSS06, CFS06, CS15, CMGL11, CN10b, DW11, DFT06, EMPS06, EW08, FK99, FSM14, GSK08, GIKK11, HL15, JD03, JS12, JGLS14, JGS^{+}15, JW10, JW11, JJL15, KLC15, KH15, Kok10, KNK^{+}14, KHW12, KT06, LNS11, LH13, LMMN07, LLL06, LQXX07, LE12a, LZES14, LZC09, LE06, MBG^{+}03, MSBZ10, MS03, NTS12, OY13, PMW10, PPSV13, QY12, RPF^{+}14, RB09a, SGR13, SSM03, SPC10, SKV03, Smi95, TWL04, TWL06, VGP14, WZY^{+}16]. throughput [XLZC14]. Throughput-competitive [CFS11]. Throughput-Delay [LK16b, EMPS06, GIKK11]. Throughput-Optimal [SPLM17, SPM^{+}17, YN20, JJS13b, KIR08, KNSV13, LLE15b, LYS11]. throughput-optimality [JW11]. Throughputs [Van17]. throwboxes [BCL10]. thwart [KVF^{+}12]. Thwarting [BOGS^{+}16, WLC^{+}10]. THz [MMG22]. THz-Enabled [MMG22]. TIDE [DSM^{+}17]. Tie [CGYZ16, TGRR07]. Tier [AAAR19, AP17, HTW^{+}22, JTL^{+}17, JTL^{+}18, KPK^{+}16, SAMB18, DJ16, JID^{+}07, NBTD07]. Tier-1 [JID^{+}07, NBTD07]. Tiered [LLX^{+}17, SA21, RB09b]. ties [CPGZ15]. Tight [CLW16, KGH^{+}20, ZGZC20, CRV13]. Tiling [YWW^{+}23]. tilt [PLR15]. Time [AEG^{+}17, BJK20, Ber00, CDHM17, CCLL17, CRL96, CWH^{+}16, CZC^{+}13, CFZ97, CGEN98, DPSA21, DYW^{+}16, ER23, FZ16, FMSM24, FXHY21, FDM^{+}17, FHQ^{+}17, GMZR13, GSKN18, HV22, HHL^{+}19, KK16b, KP21, KG16, LTDM17, LWP^{+}19, LSSC22, LXL^{+}22a, LCZH17, LYL^{+}22a, ML22a, ML22b, Nee16b, Nee22, NS16, QSW24, RFGL17, SG18, SL16a, SLJJ16, SLWW19, SA05, TML22, TG21, TG23, TZX^{+}22, VMCB22, WXH^{+}18, WWW20b, WYL^{+}22, YCZ^{+}23, YDS06a, YSY16, ZHCL17, ZWY^{+}18, Ada98, AA04, AAM05, BOY00, BO03, BM09, BFM^{+}96, BB94, BCGM07, BC01b, BBM^{+}10, CE09, ÇM15, CNS04, CE08, CS00, DZNT14, ES03, FCA^{+}06, FHH10, FCL97, FK03, GV93, GP98, GVC97, GKT97, GT03, GPM03, GAA08, Guo04, GF95, GCS06a, HS03, Hou14, HS06b, HLG94, HGG06, IS00, Ili00, KMR95, KWJY16, KMH12, LDFK12, LH95, LLD96, Lev95, LSM^{+}14, LMS99]. time [LLS09, LCQL14, MRM99, MR98, Mil98, NJW16, NMR03, ODC^{+}16, PZS^{+}16, PSA96, LZKT99, RVA00, SKR^{+}09, SYP01, SK10b, SBP03, SZN00, SA01b, Świ96, TAG08, Tha04, TC06, VAS00, VSR11, WXBZ04, WFS09, XL98, XZTT08, XGF^{+}14, YSZL15, YL16, ZVN99, ZLS96, ZA11, ZPCS11]. Time-bounded [CZC^{+}13]. Time-clustering-based [GMZR13]. time-complexity [Guo04]. Time-Constrained [CWH^{+}16]. time-critical [DZNT14, ZPCS11]. Time-diffusion [SA05]. time-division [SYP01, Tha04]. time-driven [BOY00]. Time-Efficient [LYL^{+}22a]. time-of-day [LSM^{+}14]. Time-of-Flight [RFGL17]. Time-scale [YDS06a, GKT97, GT03]. Time-Sensitive [ML22a, ML22b, TML22, YCZ^{+}23]. Time-Sensitive-Service [QSW24]. Time-shift [CGEN98]. Time-Slotted [FZ16]. Time-spread [CFZ97]. Time-stable [KG16]. time-stamp [SA01b]. time-synchronized [WFS09]. Time-to-Live [ER23]. Time-To-Rendezvous [CCLL17]. Time-Triggered [KP21, LWP^{+}19, YCZ^{+}23]. time-variant [SBP03]. Time-Varying [BJK20, DPSA21, YSY16, ÇM15, KMH12, LLS09, NMR03, TC06]. Timed [MSM16, HR95, RW95, Świ96, ZB95]. timed-token [RW95, ZB95]. TimeFlip [MRM17]. Timely [CH18, DWCZ17, DZH19, EPB14, KP23, NABZ12]. Timely-Throughput [CH18]. timeout [LO02a, MBA06]. timeouts [dSeSGM95]. timer [HGE04, Hon94, Kar10, VL97]. timer-controlled [Hon94]. timer-suppression [HGE04]. timers [FUDA03]. times [AAM05, GPLT15, HK96, NAA^{+}16, PP02, SR01]. Timescale [MAPZ18, RYS12, BFMF01]. Timestamp [FBRL18, MRM17]. Timestamp-Based [MRM17]. Timing [SS21, AD96, GK16, KKV16, VL97]. tiny [LMSKZ99]. TinySet [EF17]. TipTop [LSDT19]. TLS [SNSW12, YWZ^{+}23]. Today [MYC^{+}19]. TOFU [XL11a]. token [AK96, GQ16, HR95, Hon94, RW95, dSeSGM95, Świ96, Tod94, ZB95]. token-passing [Hon94]. Tolerance [KSSK18, WYZ^{+}24, AA96, BDHR10, PT94]. tolerances [CS99a]. Tolerant [CWM^{+}17, HK14, LWK^{+}18, MKG^{+}17, PJM^{+}19, SZMD17, WLK^{+}17, WLC^{+}20, WDL^{+}23, XCS^{+}18, YSC18, ZCZC17, ZGYB20, ZZT^{+}17, AD11, AABD13, BWS10, CS14, GLZC12, HIM07, LSS^{+}13, Pad95, SS09, SAKS13, UN11, WMS09, WKA^{+}13, WMYR16, ZNK^{+}13]. Toll [ALYX22]. Tomography [BHA^{+}20, CH21, DGW^{+}17, FAWW22, GDC^{+}17, HGM^{+}17, LGDC18, LGDC19, LGDC23, REM17, DL04, DLPT06, EDBN12, GDW^{+}16, MG16]. tomorrow [CWSB05]. Tool [DSM^{+}17, qLlH97, LCB^{+}10, SP94]. Toolkit [YLA^{+}18, LBP^{+}16, WJZ^{+}12]. Tools [YHH^{+}21]. Top [AAAR19, BCE^{+}19, LLX^{+}17, LLG^{+}17, SXQ^{+}23, TXW^{+}21, YZL^{+}19]. Top- [BCE^{+}19, LLX^{+}17, LLG^{+}17, SXQ^{+}23, TXW^{+}21, YZL^{+}19]. topic [CJV16]. topic-based [CJV16]. Topological [DLL^{+}11, ES96, MLT11, Zha17, Ros05]. Topologies [MBLN93, VKO17, WJYL16, FMMLH06, HLHD^{+}04, HFKC12, KS01b, OMA^{+}10, PEA09, QM99, SA04, SMWA04, SKZ03, SRS08, YJZW15]. Topology [AS01, BSRdA16, BCD19, BGJ^{+}04, ÇTD22, CHFH20, CN16, CYX^{+}17, DJ14, GNP^{+}13, HWLL21, JPM^{+}19, KSC^{+}23, KLKT16, NOF14, QYC^{+}24, SJSB22, Su15, TZCB23, TTM22, WWMZ20, WGL22, YZY^{+}20, YXL18b, YWH21, YLY^{+}16, ZWGC17, ZWJ^{+}20, ZLX^{+}23, ZLTX17, AA93, AACD^{+}96, AM16, ALWD05, APSKPMGM12, Bej09, CA03, CF94, EDBN12, FHT^{+}10, GW94, GM03, GB10, HIM07, HSFK09, JL98, KH07, LA95c, LHB^{+}05, LH05, LNC04, MOZ05, MOY00, NXTY10, OY95, SLG^{+}16, SFFF03, SK06, SCY08, WC08, WL10, ZCD97]. Topology-Adaptive [CYX^{+}17]. topology-control [LHB^{+}05]. Topology-transparent [Su15, JL98, SCY08]. TopoX [ZWJ^{+}20].

Tor [AYM14, KHH^{+}18, LLY^{+}12, TWS^{+}22]. torus [SMG06]. Total [SG18, ZH08a]. totem [TMMS01]. Touch [XWL^{+}18]. Tour [JLS^{+}17]. Touring [KSG11]. Towers [XLW^{+}17b]. Trace [CM16, FWN^{+}22, PV04]. Trace-Back [FWN^{+}22]. trace-driven [PV04]. traceback [Goo08, SWKA01, SPS^{+}02, SXLL08]. traceroute [GS09]. traceroute-based [GS09]. Traceroutes [DDP^{+}19]. Traces [ZJL^{+}19, MYYR13]. Tracetree [SA04]. Tracing [DGC^{+}20, GJD18, SCW^{+}21]. track [CTVD14]. tracker [DC13]. tracker-based [DC13]. Tracking [FBRL18, KMH12, LXL^{+}17b, LLM23, LNLM24, XWY^{+}18, GSD09, HQW^{+}16, LHL15, MHS95, NL99, SZ08, SG13, SH07, TGT01]. Trade [LCY^{+}19, FLC09, LA95b, SMS07, WKZL96]. Trade-off [LCY^{+}19]. trade-offs [FLC09, LA95b, SMS07, WKZL96]. Tradeoff [DD24, JV17, PYL^{+}17, CZF^{+}16, GIKK11, HMNK13, KCCM16, LCQL14, MAN15, MV16, LZKT99, SMP^{+}14, SD15b, TPC09, WHW^{+}11]. Tradeoffs [LMS04a, Nee16b, ZWL^{+}22, BM00, JWSLC13, KNK^{+}14, LSMS06, PS05, SK13, XL05]. Trading [AMSB^{+}24, CV96, CP18, GCD23, LPM23, LSL^{+}18, NS21, SLD^{+}22, CL13, LWLL16, SL14, SML04]. TrafAda [WZH^{+}24]. Traffic [ACZP21, AHX19, AdVS20, BSRdA16, BM22, CCC17, CLH^{+}24, CKS17, CLP^{+}17, CN16, CGAC20, CDS02, CYX^{+}17, DGNK21, DWCZ17, DJ12, DK98, DHS^{+}23, FGRQ18, FAF^{+}17, GWYS19, HS08, HH18, HSM^{+}20, HWC22, HS18, JLZJ19, JE18, KLS11b, KAHKB17, Le 18, LZL^{+}18, LZZ^{+}22a, LXL^{+}22a, LZB^{+}23, LWAL17, LYH^{+}23, LGHL17, LJHB18, LAJ20, MQL^{+}22, MMT14, MMT16, NGK19, NDN^{+}18, OOM^{+}18, PCW^{+}16, RB95, RZS14, SK11, SSNS17, SAC^{+}18, SACH21, SFM^{+}18, SR18, SRC^{+}20, SJZ^{+}24, TWWG19, TCTP20, TWY^{+}20, TXW^{+}19, TG21, TG23, TR17, WDR^{+}20, WHLL23, WZH^{+}24, WHC^{+}22, XWW^{+}18, XPW^{+}18, XOW^{+}23, XSM22, XLW^{+}17b, YZGC23, YWZ^{+}23, ZSL^{+}21, ZYS^{+}23, ZLY23, ZHZ^{+}24, ZCdV^{+}18, ZXZ^{+}19, ZHT^{+}19, AS94, AS14, AA04, AJDH01, APSKPMGM12, AC06, BBFG95, BGVC00, BBMELH08, BK00, BGK97, BB96, BI00, BBK12, BBM^{+}10, BL94, CAQ07, CKL16, CS99a, CLC^{+}01]. traffic [CCLT02, CRD08, CC96, CS99b, CJOS01, CPSWL96, CN09, CB97, DM95, DTM15, DG01, EAB01, EM93, ENW96, EM09, FTV^{+}10, FRC98, FGL^{+}01, cFCcFW05, FMMR10, FKT98, FqL98, GP96a, GM03, GGPS96, GRS00, GP94, GKT97, GB99, GS10b, GLSB08, HA16, HL96a, HL96b, HL03, HFC^{+}13, HV06, Hou14, Hou15, HLG94, HGG06, IS00, ITSO01, JK96, JMMT12, JS06, JS11, JBDF07, KVR98, KJF^{+}00, KWJY16, KHG^{+}14, KL95, KR08, KZ97, KLS09a, KLS09b, KLOS09, KLOS11, KLPS06, KA95, KZDM07, LA02, LCM04, LBFE09, LA95b, LL98, LTWW94, LYS93, qLlH97, qLP97, LCL12a, LE12b, LLE16, LTY06, LS03b, LMS04b, LNR94, MJ01, MCS99, MG16, MR98, MBG^{+}03, MGR02, Med95, MBM09, MJ13, MW05, Mod99, MLC07, NS03, Nee09, NABZ12, NT00, OSW97, OMA^{+}10, PLD16, PSK^{+}15, PG94b]. traffic [PF95, PDT09, LZKT99, QK01, RHC^{+}12, RD11a, RCFC15, RZWQ12, SMG06, SK10a, SK12a, SHHA09, STM^{+}12, SW04, SJL^{+}13, SMC02, SAM10, SHN16, SM05, SSD93, SAM12, SNC^{+}07, SGD05, SV98b, SA01b, SS05, Świ96, TNRP11, TG09, TMH11, TG97, VV09, VSR11, WJS07, WH11, WA11, WZY^{+}16, WJK^{+}12, WH97, WTSW97, WM96, Xin07, XZB08, XCR11, XWG14, XCR15, YRRR12, YD04, YWK07, YSZL15, YTJQ05, YZ10, YNDM09, ZQ00, ZRLD05, ZCX^{+}15, ZBA16, ZDR04, ZZZM03, dOSAU04]. Traffic-Aware [CYX^{+}17, HWC22, RD11a, WH11].

Traffic-Based [GWYS19]. traffic-feature [FTV^{+}10]. Traffic-oblivious [KLS11b]. traffics [Low00]. TrafficShaper [LZL^{+}18]. trails [BTH11, CCF04, THRW12]. Train [LTCS22]. Training [FZX^{+}23, HWM^{+}24, WLD^{+}24, YFM^{+}22]. Training-Free [YFM^{+}22]. Trajectory [DG01, DG08, GJD18, LSL^{+}21, TXL^{+}18, XWW^{+}23, YSC18]. Trajectory-Based [YSC18]. transactions [BC01a, Tow06a]. transceiver [RS97b]. Transfer [DLC^{+}17, WHC^{+}22, XYT^{+}21, YWRK19, BKTN03, IAS06, LS97b, RW04, XL98, XSHS12]. Transfers [CDK^{+}17, NRB22, XWW^{+}23, YCW^{+}19, ZCB^{+}17, LSS^{+}13, MG95]. transformation [BCL12, MLT11, PT10]. Transformer [WLY^{+}24]. Transformers [YLYL17]. Transient [CAZG20, VWNT17, ZCZC17, AQJRS16, ANSX13, DGK05, EJ14, FB07, HBH93, NLY^{+}07, WQGW09]. Transiently [LDRS18]. Transit [ASKL18, PGMR18, CSG14, MCL^{+}11]. Transition [CSR^{+}20, ANSX13, TCS04]. Transitioning [WSL^{+}24]. translation [LSV01]. Transmission [CLZ^{+}20, Dai22, LZL^{+}18, LWP^{+}19, LTZ^{+}22, LLCJ22, PPTP21, SC22a, SSK^{+}17, SLH^{+}19, VPC17, WG16, WMT^{+}22, AABD13, ATB^{+}10, BSH^{+}11, CL09a, CF94, CPS13, CWW^{+}15, GMLP10, HH10a, HLG94, IM08, KWCR10, LZ13, MCLG07, MGK12, MSS^{+}12, NBT98, OÇ10, PLS07, RA95, SL07a, SH14, TSR14, UBPE02, WBP^{+}11, WQZ^{+}13, ZM09]. transmission-range [BSH^{+}11]. Transmissions [CLW19, DJK23, PBSS23, RTNS21, WCK^{+}20, XZG20, BB96, CCA96, PS94]. Transmit [CDKZ21, ZKH10, GMS16, QCS07]. Transmitter [BZM^{+}22, SLS^{+}23]. Transmitter-Side [SLS^{+}23]. Transmitting [CGC^{+}24, LTZ^{+}22]. Transparency [LLL^{+}22a, PLR^{+}19, GG94]. Transparent [AdSD16, BMB^{+}11, BCB99, CMV10, JL98, Su15, SCY08, WSMJ04, ZTS94]. Transport [APC21, CLL^{+}19, FST^{+}09, HZB^{+}22, LZX^{+}21, MBI^{+}17, RBS02, VHT21, ZRD^{+}23, ZQ23, AKS96, AA05, ACC^{+}94, AS02, BWH^{+}07, GAA08, HOT97, KMR95, LS93a, LyT98, LT94b, MG97a, MEWP13, OdG96, OSZ^{+}06, PDE08, PSA96, RG98, SL95, SKRK12, SS96, XK06a]. Transportation [DLC^{+}18b]. Transporting [LMR99, ZH08b]. Trap [TYJ16]. traveling [BRS06]. Travi [ZSL^{+}17]. Travi-Navi [ZSL^{+}17]. treatment [BY06]. Tree [CZX^{+}17, HZHZ18, LQZ^{+}24, QFH^{+}18, WWMZ20, XLD^{+}24, XLT^{+}22, YGL^{+}19, BO03, BGVC00, CAP15, CPSWL96, FY07, GL10, IKDD15, LHL15, MSWL06, NST00, Ram96, SMG05a, SA04, SL15b, WJK06, YNDM09, ÇTD22, CCC17, GGZC19, HZHZ18, LQZ^{+}24]. Tree-Based [HZHZ18, YGL^{+}19, IKDD15]. tree-packing [WJK06]. Trees [AMS22b, HS16, ZLTX17, AC98, BLS07, CA03, DMS06, GIKK11, GR16, HSE97, JRY09, LO02b, MFB99, QGCL11, RMM99, SG05, SSM06, YRO16, ZXTT08]. Trends [ZLL^{+}23b, KSG11]. Trie [GYSZ19, BLC12, SKHL12]. Trie-Based [GYSZ19]. tries [SK03]. Triggered [KP21, LWP^{+}19, QYZX22, YCZ^{+}23]. trilateration [YLL10]. trimming [GDW^{+}16]. TRINITY [SSK^{+}17]. Trip [AEG^{+}17, AAM05, LV06]. TrueTop [ZZS^{+}16]. Trunk [WWMZ20]. Trunk-and-Branch [WWMZ20]. Trust [GTU19, PSST21]. Trust-Graph [PSST21]. Trusted [LSL^{+}18, MFR^{+}20]. Trustworthiness [WWX^{+}19]. Truth [FXQ^{+}21, PWL^{+}22, NL16]. Truthful [AAG14, GS19, NBV17, WHC^{+}19, ZFLC18, MPF^{+}15, SK12b, XL11a]. truthfully [ZLM16]. TSearch [YSC16]. TSMA [CFZ97]. TSS [XLT^{+}22]. TSS-Combined [XLT^{+}22]. TTL [BSG^{+}18, CNM20, GMD15, XNHM22]. TTL-Based [BSG^{+}18, GMD15]. tuangou [CSG14]. Tunable [YRO16, YO17, CM03, TGRR07]. Tuning [CJOS01, TSN^{+}21, WZL^{+}23c, ZWH^{+}17, BO07b, CCG00, HP00, RS97b, ZA11]. tunneling [KRKH10]. Tunnels [HZCB17, SACH21, HTC04, KL03, LRJ08, LYL07]. Tuple [LQZ^{+}24]. TupleMerge [DBL^{+}19]. TupleTree [ZWZC23]. turn [SKZ03]. turn-prohibition [SKZ03]. Tussle [CWSB05, XB14]. TV [HH10a, HH10b, TAB^{+}15]. TVA [YWA08]. TVWS [BTD^{+}17]. Tweaking [ECL^{+}20]. Twelve [DD11]. Twins [HQW^{+}16]. Twitter [ZZS^{+}16]. Two [AS07b, BTP^{+}17, CSS08, CNDK18, GGZC19, KVR98, KPK^{+}16, KW17, LS94, LL09, LLX^{+}17, LWT^{+}15, LDZC20, LLX^{+}19b, LLM^{+}24, PBSS23, RR19b, TW22, WMX17, WHLL23, XZC^{+}20, YCC^{+}21a, ZQW^{+}23, BFMF01, BHN11, CR99, CLL^{+}14, FCA^{+}06, GCM^{+}16, HY10, HN10, KS06, LYC11, LHZ^{+}16, LESZ98, LJNK12, LS05b, RKZG10, SHJ10, TZP^{+}10, TdWC^{+}94, WLCC07]. Two-Connected [BTP^{+}17]. Two-Dimensional [YCC^{+}21a, AS07b, LS94, LWT^{+}15, CLL^{+}14, LS05b, WLCC07]. Two-Flow [KW17]. two-hop [LJNK12]. Two-Layer [GGZC19, CR99]. two-level [LYC11, TZP^{+}10]. Two-Part [WMX17]. two-path [SHJ10]. Two-Phase [WHLL23, RKZG10]. Two-Stage [LLM^{+}24, XZC^{+}20, ZQW^{+}23, BHN11, HY10, LHZ^{+}16]. Two-Tier [KPK^{+}16]. Two-Tiered [LLX^{+}17]. two-time-scale [FCA^{+}06]. two-timescale [BFMF01]. Two-Way [PBSS23, TW22, KVR98]. Type [BK17, ML23, VN22, Kam96, OWMM97, YZ10]. types [DEH^{+}07].

U2 [GGZC19]. U2-Tree [GGZC19]. UAV [HNP23, KSSD24, LHY^{+}23, LDL^{+}22, WSX^{+}23, WLY^{+}24, XWJ22, XSZ^{+}22].

UAV-Assisted [WLY^{+}24, HNP23, WSX^{+}23]. UAVs [LSL^{+}21, XXZ^{+}22b, XWX^{+}24, ZGLC20]. UAVs-Aided [ZGLC20]. Ubiquitous [ZWS^{+}17, LKZ^{+}04]. UDNs [LPS19]. UDP [FMMR10, PP93b]. UFinAKA [WCH^{+}24]. UFL [THRW12]. UHF [AYL21, HQY^{+}16]. Uhlenbeck [OS21]. UIO [CU95b]. Ultra [CLTM22, CGR^{+}18, CLZ^{+}20, NSY20, NSY23, YBQZ18, Szy16]. Ultra-Dense [CLZ^{+}20, NSY20, NSY23]. Ultra-Fast [YBQZ18]. ultra-low-latency [Szy16]. Ultra-Low-Power [CGR^{+}18]. Ultra-Reliable [CLTM22]. Ultrafast [HLHL22]. Ultrasonic [GSM16, SM17, SM19, SDM20, SMGP15]. Ultrasound [LGZ^{+}23]. Unachievability [DFZ06]. unambiguous [THRW12]. unbalanced [PG94b]. unbiased [SRD^{+}09, ZCB09]. Unbounded [YS21]. unbuffered [MM94]. Uncertain [FFX^{+}17, Fuk20, LZY^{+}22, NBV17, QDD^{+}17, XGQ^{+}19, LO98, SBP03, YNDM09]. Uncertainties [TE16]. Uncertainty [LHL^{+}21, ZDCW18, GTS^{+}09, HZL16, HKCL13, KLC15, MW05, YC12, dFV02]. Uncoded [GR20b]. Uncooperative [SSM20, FCA^{+}06]. Uncoordinated [CEC^{+}19]. Uncoupled [RR19a]. Uncovering [GNK^{+}21]. Underlay [LHL^{+}21, KNK^{+}14, XB14]. Underload [MFL^{+}04]. Understand [PBGMFM22]. Understanding [ALWD05, ALW09, AST11, DCZG19, ECN09, cFKSS99, GGM11, JLS09, LLZ^{+}19, LZL^{+}20, PWLC23, RPP^{+}19, RDZ^{+}19, TWL04, WL10, XLW^{+}17b, ZCY16, MA12, WQGW09]. Underwater [DPG^{+}24, GSL^{+}24, RZE^{+}21, HKCL13, ZPCS11]. undirected [JVY06, LLL06]. unequally [RIM98]. Unfairness [BK17]. Unicast [HR14, AADS05, DLPT06, ESG11, FML09, GLAMM11, GLSB08, JVY06, LNB01, LO02b, LORS06, OS05, QTWW16, RS00, SL05, ZNK^{+}13]. Unidirectional [KSSK18, hCgKsYwT96]. Unification [NLT^{+}18, WJK06]. Unified [CLZ^{+}20, LLX19a, AA96, CS00, GLSB08, LEYS11, LCG^{+}14, NCT14, PM09, RL07, SS07, TYJ16, Tha01]. Uniform [CG21, BB96, HL99, MM94, NT00]. uniform-traffic [BB96]. Unifying [TW22, JWSLC13, ZFC15]. unilateral [BSS14]. UniMIN [BS00]. unique [AM16, Nai97]. uniqueness [RKA08, TWLC07]. UniROPE [GJD18]. Unit [LWK^{+}16, SZMD17, WLK^{+}17]. Units [VLZL16]. Universal [DGC^{+}20, GJD18, GGZC19, HK24, LY22, XCQ^{+}23, Lev95]. universality [Sha94]. Unknown [BCD19, FM22, FM23, GLY17, LCL^{+}20, YZP^{+}14, ZCZF20, GKJ12, MS14, SZT01, ZWTC16]. Unknown-target [YZP^{+}14]. Unlicensed [CLGSS17, GSPV^{+}18]. Unmanned [WCL^{+}22]. Unmodified [HLP^{+}16]. Unmodulated [LJJ^{+}19]. unprecedented [CSS^{+}14]. Unpredictable [LGDC18, KLS09a]. unpunctual [Lee96]. Unreliable [GLY17, PBSS23, QTE20, ZCZ^{+}20, Zha17, DG08, Hou15, LCQL14, ZW14]. unsaturated [TS08]. Unslotted [CFC01]. Unstable [WWL^{+}24a]. Unstructured [SdVK16, XQG^{+}22, YCL15, RS05, SRS08, SRD^{+}09, WZR08, YWLL09]. Unsupervised [SL17, HFC^{+}13]. untuned [PRR06]. Unveiling [CKC^{+}13]. UPC [MR98]. upcalls [GP98]. UPCF [CHH06]. Updatable [KLC^{+}18, WCH^{+}24, XLT^{+}22]. Update [FXHY21, LCL17, MGZ^{+}23, TZX^{+}22, VVC^{+}17, XYL^{+}17, YWH21, AHL96, CVM^{+}15, Lin97]. Updates [BU21, BSSS21, FLMS18, GYSZ19, HZG^{+}18, LDRS18, MKAE20, MSM16, MRM17, VCVC17, WGZC21, ZWZC23, ZWCL17, ZGZC20, BN05, LXX^{+}14, NM09, SZM08]. Updating [ZWX^{+}24]. Upgrades [PIST19]. Upgrading [MK10]. Uplink [CLW19, DJK23, HTW^{+}22, ZCPG^{+}23, ASKR16, CS99b, CK07, DM15, HRCW08, SEK15]. uplinks [Nee08]. Upon [LLZ^{+}23a, BFF07]. upper [FP95]. Urban [BMBK21, CLQ^{+}19, PJM^{+}19, XLW^{+}17b, ZHZ^{+}18, ACVS10, CK10a, CAK12]. Urban-Scale [ZHZ^{+}18]. URC [MBL19]. URLLC [AdVS20]. Urn [GYSPR14]. URSA [LKZ^{+}04]. usability [KCA97]. Usage [ACVS10, KLLT18, Ma16b, CSN06, JK05, KL03, LRJ08, SKK07, WXC16]. Usage-Based [Ma16b]. usage-priced [JK05]. Use [TZPZ23, BCL^{+}09, BBL95, FF99, KAZ01, MCL^{+}10, MCS99, RK15, TNF97, TG96, ZAFB00, ZA95]. used [ZVN99]. User [AP17, AM19, BDR22, Bor05, CCLL17, CJLF16, CW19, CGL16, CJ18, CP20, DSM^{+}17, DJK23, GHK18, KLY^{+}23, LZZ^{+}22a, LPS19, LSCT17, LYW^{+}21, LCZ^{+}23, LYC^{+}19, LLM^{+}24, MQL^{+}22, NGK19, SSNS17, SGH^{+}19, SSK^{+}17, SS21, TXL^{+}18, WWX^{+}19, XGW^{+}20, ZZS^{+}16, ZHCL17, ZK19, AG16, AW04, Bar95, BMM^{+}09, CAO11, CKR^{+}09, DFMR15, GP98, HSPH09, JBR16, JL12b, KDYV12, KLC15, LAPS08, LCB^{+}10, Nee08, RD11b, TNML93, VG04, VCM04, XY09a, YD04]. User-Centric [DSM^{+}17, LSCT17, SGH^{+}19]. user-controlled [LAPS08]. User-Defined [LZZ^{+}22a]. User-level [Bor05, LCB^{+}10]. user-provided [AG16]. user-session [BMM^{+}09]. user-space [GP98]. Users [BPL20, CLCL23, DNS23, GBMV21, GS19, MS17, OJSY16, WPZM16, DJ12, FP14, GHR14, GH04, HLS14a, JKJ13, KS06, LPIH11, NL99]. Using [Ada98, AWM^{+}20, BLC21, BPVRSP16, CAS^{+}20, CSG14, CSD22, CJH^{+}11, CLS^{+}21, CN19, Dat17, FML23, FMCS20, FLH^{+}17, FBRL18, GSM^{+}17, HDQ^{+}16, HAG19, HV22, HWHW18, HWJZ21, JLS^{+}17, KE21, KLY^{+}23, LKS^{+}16, LBZ^{+}20, LLL^{+}16, MCZ^{+}22, MLB21, MGG^{+}05, MNZ23, MPN^{+}14, MRM17, NRB22, REM17, RUH^{+}18, RpLP^{+}17, RPP^{+}19, SAC^{+}18, SKE19, SC17, SHN16, SBGJ18, TR98, TWL22, WSX^{+}21, WCC14, WLL^{+}16a, XWL^{+}18, XLT^{+}22, XYQ^{+}17, YLYL17, YJL^{+}19, YBQZ18, ZLG^{+}20, ZHCC24, ZGY^{+}16, ASW00, ARK09, AN05, ABA^{+}16, AOM04, BLC12, BLBS06, BHL07, Ber00, BFMF01, BKTN03, BLDF09, BL04, BDWS12, BBHK14, CLP12, CAO11, CHM^{+}05, CLC^{+}01, CW16, CKKK09, CCF04, CFD06, DKT06, DLPT06, ES96, ES07, EM09, FWL08, GLA93, GMWD13, GLG04, GP98, GR12, GSD09, GT00, GCS06b, HQW^{+}16, HKLM07, HKLS12, HBS96, HK96, IPG97, IAS06]. using [Kam96, Kam10, KRLL11, KKL03, KBS12, KHTK00, KMH12, KRKH10, KL03, KLS03, Kop96, KLO97, KS13, LSV99, LRJ08, LBFE09, LDK12, LAN97, LS99, LZF09, LGW^{+}11, LTY06, LJ09, LMT16, MSS02, OWKS16, PYL99, PWMC12, PD16b, PSA96, PJ13, PWK^{+}13, PP02, RRK96, SGR13, SEK15, SRS03, SYDM09, SG96, SJ12, STKL01, SV96, SAM10, SK06, SNC^{+}07, SKZ03, SS04b, TNRP11, UZ93, VWT^{+}14, VS97, WJS07, WJK^{+}12, Wil96, WGL00, WWL02, WZL^{+}13, YD07, YL16, YKKY08, ZKH10, ZAS12, ZLS96, ZGG05]. Utilities [FM23]. Utility [CSD22, CPS^{+}12, CGYZ16, CP18, DTM^{+}17, DCN^{+}19, DMT^{+}19, FM22, GCX^{+}17, HN13, JWZ^{+}21, KSNR20, KSRW22, LA02, LFZ^{+}22, LZC20, LYW^{+}21, LNLM24, MLX18, Nee19, NCM18, NZW24, PLR15, PL17, PGMR18, RR19a, SGJ17, WWC^{+}18, YLF^{+}21, YN18, ZTH^{+}23, BNS11, BMS14a, EML12, HMNK13, HL15, JW10, KS03, LLCL11, LCZC13, Nee13, XSC03]. Utility-Aware [JWZ^{+}21]. Utility-Based [DTM^{+}17, DCN^{+}19, LA02, XSC03]. Utility-Centric [PGMR18]. utility-delay [HMNK13]. Utilization [JD17, KSSK18, LCLC18, ZFLC18, CZ12, QS04, SCY98].

Utilizing [PSST21, XLP^{+}23, CFM^{+}09, CS14, RS07, ZZHZ13].

V2X [NV21]. vacations [RW95].

Validation [LFY^{+}19, NKL^{+}23, WQL^{+}21, XGQ^{+}19, vRDHSP17, ALWD05, CBAT06, DM14, PFTK00]. validity [HDM10]. Valuable [DFGV11]. Value [Hua17, WSL^{+}24, XL23, ZLG^{+}17, MCL^{+}10]. Value-of-Information [Hua17]. VANET [DKSC18, LNL^{+}16]. VANETs [CLQ^{+}19, FGM^{+}13, HLP11, KGdV^{+}21, LMODF18]. Vanishing [YN18]. Variability [LGHL17, XLL^{+}20, LBFE09, SZKT98, WTSW97]. Variable [XPW^{+}18, BB94, BGK97, CR99, FNQ00, JMI95, KLS09b, KLOS09, Le 02, RT99, RKK14, SA01b, Tha01]. variable-bit-rate [RT99]. variable-increment [RKK14]. variable-length [JMI95]. variable-rate [FNQ00, Tha01]. variables [GKJ12, NM09]. variant [SBP03]. variation [JJSS04]. variations [HH98]. Various [CCW^{+}17, SWH19, AT03]. Varying [BJK20, DPSA21, KW17, YSY16, BLEM^{+}12, ÇM15, KMH12, LLS09, NMR03, SR01, TC06]. VBR [Ada98, BI00, CLC^{+}01, HL96a, HL96b, Hey97, KL95, KZ97, LyT98, LNR94, MCS99, RB95, WKZL96]. VBR-video [HL96b]. VCP [LY10]. Vector [Sob17, ZCJ^{+}13]. Vehicle [CXK^{+}23, SQS20]. Vehicles [BSRdA16, LDD21, WCL^{+}22]. Vehicular [ALR^{+}24, CZC^{+}22, CDW19, DLR^{+}18, HSL20, LAJ20, PJM^{+}19, SKA^{+}18, XWW^{+}23, YSC18, ZV16, CK10a, LNL^{+}16, LLW^{+}14, ZSK12, ZLSK15, SMEH20]. Vehicular-OBUs-As-On-Demand-Fogs [SMEH20]. Verifiable [HK24, YZHZ21, ZZG^{+}16]. Verification [FLZ^{+}23, JR21, KLKT16, SCHG22, SHV^{+}23, YLYL17, ZLZ^{+}21b, ZGS^{+}24, vDJJ^{+}22, KBS12, KVF^{+}12, OdG97, SR02, TYP^{+}15, YL16]. verifying [LK13]. version [AKS96, LTWW94, SKT96]. versions [AT03, Kum98]. Versus [LPM23, Van19, AD96, CFPP96, GKS05, LNB00, LSM^{+}14, MS15, RrBG94, SR14, XG05, YGC10]. Vertex [ZSLZ21, MFB99]. vertex-redundant [MFB99]. Vertical [AMCD19, JSuRKH03]. Very [GBG^{+}16, LNM^{+}09, LS10, VSR11]. Via [ADR18, AB23, AD18, BGHS10, BSSS21, BCR^{+}12, BZM08, CXL18, CAP15, CDFG06, CLZ^{+}23, CGAC20, DSYN24, DMDM17, DRJ^{+}14, FLS^{+}22, FM06, GA24, GLLJ16, HGZJ21, HZLZ22, JK21, KSAK18, KK06b, LXLC20, LLY^{+}22, LK16b, LXX^{+}17, LZY20, LGZ^{+}23, LL18, LJHB18, LZZ^{+}22b, LW17, MHS^{+}17, MQL^{+}22, MG16, MHRR12, Nee16a, NGRF19, PV04, RV21, RKT02a, SRMB^{+}23, SPR^{+}20, SYZP19, SCS^{+}22, SLWW19, SPB16, SFS^{+}22, Su15, SV06, SN15, THRW12, TRKN10, WZL^{+}23c, XDZ^{+}23, XCL^{+}22, XWG14, YS93, YKGF08, YCGH17, YZZ^{+}21, YC12, ZG14, ZMH17, ZP18, ZCW^{+}22, ZL15]. viable [SNC^{+}07]. Vibration [YLL^{+}17]. ViChaser [CZM^{+}24]. Video [AD14, AAA18, AAAR19, BTD^{+}17, BBR^{+}22, CZM^{+}24, EAH^{+}18, EFA19, FML23, GWYS19, HH18, JSZ14, KCM16, KS13, LKS^{+}16, LMSR19, LCU^{+}20, MLS^{+}23, MJ17, QWL21, RWL^{+}22, SZWW22, TPW^{+}18, TL16, WCW^{+}17, XYLL14, YZBR14, YWW^{+}24, ZWDS00, ZWJ^{+}22, ZZL^{+}22, Ada98, ABA^{+}16, BM97, CKR^{+}09, CT01, CR99, CPS^{+}12, DM14, DYX12, DRR98, FHSZ13, FNQ00, GH93, GMY13, HL96a, HL96b, HH10b, JBR16, KMR95, KL07, KMHS09, LMR99, LCY96, LY94, LZZR12, LZL11, LyT98, Liu10, LNR94, MCS99, OWKS16, PSK^{+}15, PD16b, RB95, RT99, RCFC15, SZKT98, SRS03, SHN16, SSD93, TM13, TAG08, TCS04, VC12, VAS00, WXR13, WKZL96, WLR10, VC14]. Video-aware [AD14]. video-conferencing [LZL11]. video-on-demand [TM13]. video-QoE [VC12, VC14]. videoconferences [Hey97]. Videoconferencing [CVHM22, TH96, BO00]. Videos [SUS20]. VidQ [FML23]. View [BS19, HJL^{+}20]. viewer [KS13]. viewing [SHN16]. Viewpoint [CZM^{+}24]. views [SAS16a]. ViNEYard [CRB12]. Violations [LCL^{+}18, ZF96]. Viral [NTD17, DZNT14]. Virtual [AL98, ACA16, BFG^{+}14, BCD19, CMR17, CL16a, CYX^{+}17, DYJ^{+}23, EMAL17, FMMLH06, GM03, GJWZ16, HTW^{+}19, HKLM17, JPM^{+}19, KRS^{+}17, KLLT18, LOP97, LWLL16, LLNC09, LWK^{+}18, LJJ^{+}19, MK96, NGRF19, PHC20, RS19, RS20, SC17, SZMD17, TYLH09, XCZ^{+}17, YLH17, YLY^{+}16, ZG14, ZLL24, ZLX^{+}23, ZZT^{+}17, AS09, APSKPMGM12, CFZ94, CRB12, CK00, DJ14, DGG^{+}02, EDM16, GW94, GCZ96, HLHD^{+}04, HL15, HK96, IPG97, JK15, KH07, KRSY02, KS04, LBS05a, LTB04, LMG04, Med95, OMA^{+}10, OSZ^{+}06, SBNRS14, SCY98, SKHL12, SZL^{+}14, VS97, VL99, WKA^{+}13, WRS^{+}15, WM95, XL95]. Virtual-coordinate-based [TYLH09]. Virtual-topology [GM03]. VirtualClock [FP95]. Virtualization [EMAL17, HYK^{+}23, KLR^{+}20, LXL^{+}22a, NTR18, WZL^{+}23c, CL15, FK13, FSH^{+}13]. Virtualization-Aware [LXL^{+}22a]. Virtualization-Based [KLR^{+}20]. Virtualized [CN19, GWQ^{+}23, LDS^{+}24, NSC^{+}22]. virtualizing [KMZR12]. Virus [DME23, VOK09]. Visibility [HKC^{+}20, LBP^{+}16]. Visible [BWG^{+}20, WG16]. Vision [SCW^{+}21, YWW^{+}23]. Vision-RFID [SCW^{+}21]. Visual [FML23, TZZ^{+}14]. Vitalizing [Ma16a]. VM [SC18a]. VNE [GJWZ16]. VNF [AMCD19, LW20, MCES19, SJ21, XDZ^{+}23, ZLZ21a]. VoD [AAG^{+}16, ZFC13, ZFC15]. Voice [ALYX22, WML^{+}18, LZ06, MTK03]. VoIP [CCY^{+}14, HLSG04, SZ08]. Volatile [ZWL^{+}22]. Volatility [SL17]. Volume [ABBF19, BCE^{+}19, HLZ^{+}14]. Volumes [ACZP21]. von-Neumann [YZLH17]. vote [HWZ^{+}23]. voting [WKWV16]. voting-based [WKWV16]. Voyager [LTCS22]. VP [SD00]. VPN [BGHS10]. VPNs [CL08, CL09b, RRK07]. vs [ATE22]. vSDN [ZWZ20]. vSFC [ZLZ^{+}21b]. vSwitch [RRS23]. Vu [SPGM13]. Vulnerabilities [HHD22, LHL^{+}23b]. Vulnerability [GCZY18, TWS^{+}22, DXT^{+}12, DT15, MYYR13, NZCM11, SNXT13].

WAIPO [GND17]. wait [LZ09, QY12]. Waiting [ZVN99]. Wake [BSSS21, PLM19, CK09, ODC^{+}16, WFS09]. Wake-Up [PLM19, CK09, ODC^{+}16]. Wakeup [ZWYD18, PZS^{+}16]. walk [FML09, HLS14a, LHK^{+}12, RSH^{+}11]. walks [LKC^{+}13, LZ13]. Walls [CW19]. WAN [DCGN03, WRS^{+}15, ZYS^{+}23]. WANs [GDL^{+}22, GDWX23, GDJX24, YCW^{+}19]. war [KAS16]. warning [FGM^{+}13]. Wars [YLK^{+}17]. wasted [BB96]. Watch [CH21, WXJ^{+}17]. Watching [TGD^{+}20]. watermarking [HKB14]. Wave [DF20, SKE19, YXAZ^{+}18, YLWH20, ZW22, ZWZM18, ZXW^{+}20b, AWFT15, DMK05]. wave-mixing [DMK05]. Waveband [CAQ07]. Wavebanding [TS14]. Waveforms [WCK^{+}20]. waveguide [NPQ06]. Wavelength [AdSD16, BM00, GYLH17, PG95, Pan99, WQC06, YWHW24, AM16, And04, AZ09, BPPP12, CV12, CM05b, CMV10, CL05, FT06, GSKR99, GLG04, GT00, KA98, KS01b, LSV01, LS99, LS01, LHM02, MBLN93, MA98, NPQ06, NPY07, OB03, QY04, RM02, RS95a, RS98, RZVZ06, SMG05a, SMG06, SYR05, SKCW10, SAS96, SAS99, XL99, ZOM03, ZA95, ZQ00, ZZZ^{+}07, ZY07b, ZKL11, ZRP00]. wavelength-convertible [ZZZ^{+}07]. Wavelength-routed [BM00, AM16, CV12, KS01b, RM02, SYR05, SAS99]. wavelength-routing [MBLN93, ZRP00]. wavelength-selective [GT00]. wavelengths [RIM98, SML04]. wavelet [KKS^{+}08, MJ01]. wavelet-based [KKS^{+}08]. Way [BPVRSP16, PBSS23, QZC^{+}22, TW22, BIS00, GCS06b, KVR98]. Waze [LZY20]. Waze-Inspired [LZY20]. WCS [SWL^{+}18]. WDM [SK11, AEG^{+}13, ANTR17, AA99, ATB^{+}10, And04, BSH^{+}11, BBMELH08, CV12, CM05b, CRD08, CEFS99, CMV10, CLG00b, DS99, DSTM12, EM09, FMMLH06, FCT03, GM03, GRS00, HD07, HLHD^{+}04, IBM95, JM00, JF04, KA98, KT11, KL09, LS03a, LML11, LLM14, LS01, LS06c, LXC05, LTP10, LLY09, MJ13, Mod99, MMS01, MBRM96, NPQ06, PM96, PS94, QY04, RA95, RS97b, SMG05a, SMG06, SK10a, SK12a, SSM06, SM05, TMP07, TCPV13, TMH97, TS14, VWT^{+}14, WQC06, XTMM11, Xin07, XGF^{+}14, ZOM03, ZA95, ZQ99, ZQ00, ZZZ^{+}07, ZY07b, ZA11, ZLTX17, ZZZM03]. WDM-based [LML11]. WDM/TDM [ZA11]. Weak [AKA10]. weakens [QTWW16]. Wearable [MFT^{+}20, SM17, SM19]. weaving [MLT12]. web [PP02, AW97, BMM^{+}09, BMS14b, CDHM17, CDI^{+}04, CJOS01, CB97, FCAB00, FRC98, HZCB17, LAJS07, MPFK02, RW04, RSB01, TRKN10, ZAFB00, ZRD^{+}23]. web-cache [PP02]. Web-conscious [MPFK02]. Webcam [LWT^{+}21]. Webex [CVHM22]. websites [XY09b, BS19]. Weight [GBG^{+}16, KAA^{+}18, LWAL17, MMT14, SZMD17, VL16, YRB^{+}18, BGH^{+}95, FJL^{+}97, JMMT12, LJA14, MS15, NJW16, NBTD07, PPV04, WBEGS05]. Weighted [Far95, FHT^{+}10, McA94, SRK22, SG18, SWL^{+}18, ZWL^{+}16, AS08, JLRS16, LWL^{+}11, SPB16, ZJ12]. weights [CL09b].

Welfare [ZHW^{+}17, AAG14, LWLL16]. Wheel [CDRV11, FMH^{+}21a]. wheels [Kar10, VL97]. Where [GNK^{+}21]. Which [RCS14]. While [LTZ^{+}22, AWKN16, CK09, KCB03]. Whispers [WXW15]. White [CGYZ17, SRI^{+}18, WLW^{+}17, Bar95, SP94]. whitespaces [MGCK15]. Whitt [LY22]. Wi [BMBK21, BTD^{+}17, CZGKB24, HLS^{+}14b, JYC^{+}16, MGLH18, MSRG18, SPR^{+}20, WCWZ17, XLZ^{+}19, XS21, YCGH17, ZZW^{+}24]. Wi-Fi [BMBK21, BTD^{+}17, CZGKB24, HLS^{+}14b, JYC^{+}16, MGLH18, MSRG18, SPR^{+}20, WCWZ17, XLZ^{+}19, XS21, YCGH17, ZZW^{+}24]. Wide [BBEF^{+}21, BFG^{+}14, CCX^{+}23, CB97, DQYG23, HS19, HCW^{+}23, KA20, PF95, SRI^{+}18, TXHL23, TRKN10, Wan04, WQY^{+}17, WZLM22, ZZX^{+}21b, ZRP^{+}22, ZLWM18, BSF16, CVM^{+}15, DSA^{+}14, DEF^{+}96, FCAB00, FR07, GCS06b, HL05, HK96, Jia98, KKM^{+}97, LLW^{+}09, LL13, LM01, Med95, MBRM96, Pax94, RVS^{+}02, STM^{+}12, THRW12, Tas96, ZWDS00, HGZJ21]. Wide-Area [BFG^{+}14, CCX^{+}23, SRI^{+}18, DEF^{+}96, FCAB00, HK96, KKM^{+}97, LM01, Med95, MBRM96, Pax94, RVS^{+}02, ZWDS00]. Wide-band [Wan04]. Wide-Sense [KA20]. Wideband [CDKZ21]. widest [SG05]. WiFED [MBN^{+}21]. WiFi [ACVS10, AYL21, BLM^{+}17, CLGSS17, CW19, GSPV^{+}18, GBG^{+}16, KLY^{+}23, LLY^{+}13, LS16, MBN^{+}21, MW06, PWWP18, RFGL17, WLY^{+}23, WLL^{+}16a, ZSK12]. WiFi-Based [CW19]. Wild [CJL^{+}19, DLY^{+}22, DYJ^{+}23, JHM^{+}21, LLL^{+}22a, SL16b, YHH^{+}21, ZXH^{+}13, ZZW^{+}15]. Wildcard [XYQ^{+}17, YXC^{+}18]. Wildcard-Based [XYQ^{+}17]. WiMax [EF08]. window [BLPS10, GBC^{+}95, JGMB03, KVR02, MW00, SL05, Spi97, TAJ^{+}10, YWK07]. window-based [JGMB03, MW00, SL05]. windowing [SG96]. Windows [BEK^{+}22]. wired [Bej04, BV05b]. wired-wireless [BV05b]. Wireless [ATE21, ATEY22, ATE23, ALYX22, AdVS20, APSG14, AGM^{+}17, AHEK24, BJK20, BBR^{+}22, BBR19, ÇTD22, CLGSS17, CAZG20, CLS^{+}18, CLW19, CMP16, CCK16, CGYZ16, CWH^{+}16, CGC^{+}17, CLV17, CGC^{+}18, CZL^{+}19, CDGZ20, CBHS20, CCX^{+}23, CGC^{+}24, CNG^{+}16, CEC^{+}19, CCG20, CXW^{+}18, CJL^{+}19, DLC^{+}17, DMLC18, DWL^{+}18, DLY^{+}21, DHK16, DRCM^{+}17, DWCZ17, DLY^{+}22, DTN^{+}21, DCZG19, FFZ^{+}18, FWN^{+}22, Gan20, GDC^{+}17, GZL^{+}17, GB18, GV17, HLZ^{+}21, HLZY23, HLP^{+}16, HWM^{+}24, HCL^{+}17, JWW^{+}23, JLS^{+}17, JM17, JE18, JD22, KSUB^{+}18, KSM19, KP23, KS19, KWH^{+}17, KIW^{+}17, KAA^{+}18, LTDM17, LCK^{+}18, LMSR19, LLE16, LYMA^{+}17, LZL^{+}21, LTN^{+}19, LHY^{+}23, LLL^{+}16, LDY^{+}16, LCZH17, LL17a, LWK^{+}18, LJSB22, LSHZ16, LGCG^{+}21, LYKT21, LCU^{+}20, LFF^{+}19, MLX18, MYMY17, MGK20, NSY20, NSY23, NK20, Nee16b, Nee22, NSP^{+}16, OBS17, ODJ23, PBV17, PLM19, PPV17, PPTP21, QZX^{+}17, QLSW19, RCW15, RCR^{+}18]. Wireless [RRS^{+}14, RHX^{+}20, SAMB18, SSY19, SMC^{+}20, SSK^{+}17, SPLM17, SLH^{+}19, SdVS22, SRB^{+}20, TKM20a, TKM20b, TT17, TZL23, TTM23, TG21, TG23, URZ^{+}14, WVZ17, WWW^{+}20a, WLD^{+}24, WN16, WCW^{+}17, WML^{+}18, YM16, YFM^{+}22, YLD^{+}23, YZY^{+}18, ZJS^{+}12, ZFW14, ZWL^{+}16, ZZ17, ZFW^{+}17a, ZWGC17, ZYL^{+}17, ZWYD18, ZLL24, ZDB^{+}17, Zha17, ZMD^{+}20, ZMMG22, ZZT^{+}17, ZWZM18, ZXW^{+}20b, ZYY^{+}21, AIN^{+}15, AS14, AK00, AJV06, AK09, AA04, AA05, ALJ99, AJDH01, AW04, AGLM10, AGL16, AK15, AS07a, AS07b, AVPG14, ACCF12, AZ06a, AZ11, AT03, AKS^{+}13, AWFT15, AAV09, AST11, BCP13, BBG11, BD07, BPSK97, BCP00, BCGC15, BTC05, Bej04, BHL07, BNJR12, BNJ16, BV05b, BRS10, BS09, BNS11, BSYS12, BE06, Bor05, BMS14a, BCC07, BLB10, BZM08, BESW08, CKS16, CLP12, ÇY07, CHML15, CFM13, CCV03, CT04a, CHL16, CLM99, CCL11]. wireless [CT04b, CLC12, CYK09, CZC^{+}13, CLSC15, CH11, CS00, CJJ09, CJZS14, CSN06, CHH06, CK07, CGK10, CG15a, CG15b, DPBT11, DYH13, DJ14, DYX12, DV09, DLL^{+}11, DFT06, EMPS06, EL11, ESP05, ES07, EOSM10, EML12, Fan05, FTZ^{+}13, FML09, FK13, FSM14, GHR14, GMP13, GDC^{+}16, GSK08, GMZR13, GS13, GHW14, GT10, GT02, GMY13, GMS16, GAA08, GLS09, GS10b, GS11, GMSK09, HIM07, HLL13, HSM^{+}13, HKV^{+}13, HSS08, HG14, Hou14, Hou15, HSPH09, HY08, HLW13, HK11, IKDD15, IW08, IGE^{+}03, IK09, JR14, JMS07, JCJ95, JJS13a, JC13, JJS13b, JGLS14, JGS^{+}15, JW10, JL12a, JP13, JS09, JLS09, JS14, JBR16, KL12, KWJY16, KWCR10, KK07, KIR08, KE16, KRH^{+}08, KDHK15, KEW06, KSA12, KK00, KS09a, KD10, KLSS10, KWS^{+}11, KS11, KDYV12, KBV^{+}13, KG10, KN05, KMZR12, KES13, KMHS09, KWE^{+}10]. wireless [KG05, KHW12, KT06, KT07, KIR06, KS12, KS09b, KV09, LTS10, LBB08, LDFK12, LSLL14, LK16a, LMS12, LNS11, LMS05a, LKC11, LGC16, LMP08, LAN97, LSC99b, LHB^{+}05, LH05, LZF09, Li09, LY10, LLLT10, LBX11, LPF12, LE12a, LZZR12, LG13a, LZ13, LLE15a, LLE15b, LHZ^{+}16, LLNC09, LJA14, LS06d, LSS07, LR09, LLS10, LÜ14, LLL10, LCZC13, LEY14, LHC^{+}16, LBS99, LRG10, LH10, MVRZ09, MCLG07, MBL10, MHS95, MSS^{+}12, MWQ^{+}10, MQ05, MPF^{+}15, MWC16, MRD08, MHXT10, MAS09, Nee08, Nee09, NSW11, NTS12, NSCR06, ODC^{+}16, OY13, OÇ10, PSK^{+}15, PT96, PNRMC13, PLS07, PHL15, PA12, PD07, PPSV13, PRR06, PPV12, PCL15, QCS07, QSS^{+}15, RGG11, RCGS09, RLKT98, RVS^{+}02, RJJ^{+}11, RSS09, RD11a, RD11b, RSR10, RWA^{+}08, RKNS10, SLP07, SPH04, SGR13, SZ08, SEK15, SYDM09, SRR08, SM14, SM16, SZG^{+}13]. wireless [SRB10, SLS10, ST09, SKRK12, SA01a, SH12, SSWK13, SS09, SS10, SL12, SK10b, SLH^{+}06, SSAK12, SMSM06, SH07, SV11, SKUB12, SPB16, SX10, SA05, STL04, SN15, TXL^{+}12, TCS13, Tan16, TX08, TS08, TYLH09, UBPE02, VJV14, VAGT13, VL10, VCM04, VA06, VA07, VA09, WCY04, WY06, Wan04, WSC08, WLL^{+}11, WB11, WA11, WVG12, WKA^{+}13, WHM^{+}13, WLWL13, WDCL15, WLL01, WK13, WWL02, WWT05, WHTC15, XY10b, XSC01, XSC03, XAST12, XXBC14, XHN04, XK06a, XSHS12, XSH^{+}15, XWWC16, XC08, XW11, XL11a, XE13, YWK07, YLL10, YJ15, YCV15, YASS15, YHE04, YAA09, YS07, YGC10, YSRL11, YG10, YBX^{+}10, YBX^{+}12, YC12, ZKH10, ZAS12, ZH08a, ZH08b, ZSFZ11, ZBXH13, ZCJ^{+}13, ZG14, ZCW15, ZT03, ZG08, ZR09, ZXH^{+}13, ZW14, ZL16, ZWTC16, ZW10, ZZHZ13, vRWZ09]. wireless-optical [SYDM09]. Wireless-Powered [CGC^{+}24]. WirelessHART [CSSG23]. Wirelessly [XTHL21]. wireline [BMS14a]. Wise [LZX^{+}24, TZPZ23, WXX^{+}24]. Wisely [CDL^{+}19, TZPZ23]. Within [NJM^{+}19, SKG^{+}18, ZGZC20, DM95]. Without [BBHH^{+}18, CBZ16, SL16a, WZH^{+}18, YXL18b, GR01, JJS13b, JR96, KBS12, KHW12, LS97a, LEY14, MGK14, MMH^{+}15, SS93, Sha97, VB94, WM16, WYH10]. WiZig [GHZ20a]. WLAN [ALMR14, BPK^{+}10, KLK^{+}20, KAHKB17, RP13, ZS13]. WLAN/WPAN [RP13]. WLANs [APB^{+}13, BCL^{+}09, BOGS^{+}16, BJY11, CZW^{+}21, DJK22, DJK23, GSRS^{+}15, GHK18, KAMG07, LNM^{+}09, LWC^{+}14, NKNK17, NGK19, RKZG10, RKA08, SRBBG17, ZK19]. WMGR [AGCFV18]. WOBAN [SYDM09]. Work [CVV17, Van19, GK16, She95, TG96]. work-conserving [GK16, TG96]. Workers [LZY^{+}22]. Workload [AHP21, MCS99, ZCZ^{+}21, AW97, LD95, VAM^{+}06]. Workloads [BPW23, CG04, KG16]. World [SXEZ21, GQ16, CB97, TRKN10]. Wormhole [KMS^{+}01]. wormholes [DLL^{+}11]. Worms [ZGTG05, MCR10, VG08, WLC^{+}10]. Worst [BGVC00, GSKR99, Lee96, TML22, BS15, Val01]. Worst-Case [TML22, BGVC00, GSKR99, Lee96, BS15, Val01]. WPAN [RP13, ZS13]. WRSNs [SLD^{+}22]. WSNs [HLX^{+}15, LSL^{+}21, RPP^{+}19, SCS^{+}22, TAH17, WJYL16]. WWW [NBK02].

X.500 [Bar95]. XCP [ALW09]. Xen [LXL^{+}22a]. XOR [AGGT16]. XOR-based [AGGT16]. XORs [KRH^{+}08]. XsVarSWITCH [KA20]. Xunet [KKM^{+}97].

Yao [ZYL^{+}17]. Year [ACZP21, DLY^{+}22, FAF^{+}17]. Years [TGD^{+}20, DD11, GCM^{+}16].

Z [RWA^{+}08]. Z-MAC [RWA^{+}08]. ZC-NET [ZRP^{+}22]. Zero [ABBF19, ALYX22, MHH20, PMN19, SCW^{+}21, PZS^{+}16, WLC^{+}10, XAST12]. Zero-Day [ABBF19, WLC^{+}10]. Zero-Rating [MHH20]. zerotime [PZS^{+}16]. ZigBee [CLS^{+}21, KCH^{+}19, SPR^{+}20, WCK^{+}20, WLY^{+}23]. ZigFi [GHZ^{+}20b]. Zone [XWX^{+}24, HP01, SPH04]. Zones [BPL20, TRKN12]. Zoom [CVHM22].

References

Abu-Amara:1993:FTM

[AA93] Hosame Abu-Amara. A fast topology maintenance algorithm for high-bandwidth networks. IEEE/ACM Transactions on Networking, 1(3):386–394, June 1993. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL

http://www.acm.org/pubs/
citations/journals/ton/
1993-1-3/p386-abu-amara/

.

Agarwal:1996:UAF

[AA96] Anjali Agarwal and J. William Atwood. A unified approach to fault-tolerance in communication protocols based on recovery procedures. IEEE/ACM Transactions on Networking, 4(5):785–795, October 1996. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1996-4-5/
p785-agarwal/.

Alanyali:1999:PAW

[AA99] Murat Alanyali and Ender Ayanoglu. Provisioning algorithms for WDM optical networks. IEEE/ ACM Transactions on Networking, 7(5):767–778, October 1999. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1999-7-5/
p767-alanyali/.

Akan:2004:AAR

[AA04] Özgür B. Akan and Ian F.

Akyildiz. ARC: the analytical rate control scheme for real-time traffic in wireless networks. IEEE/ACM Transactions on Networking, 12(4):634–644, August 2004. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Akan:2005:ESR

[AA05] Özgür B. Akan and Ian F.

Akyildiz. Event-to-sink reliable transport in wireless sensor networks. IEEE/ ACM Transactions on Networking, 13(5):1003–1016, October 2005. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Al-Abbasi:2018:VSD

[AAA18] Abubakr O. Al-Abbasi and Vaneet Aggarwal. Video streaming in distributed erasure-coded storage systems: Stall duration analysis. IEEE/ACM Transactions on Networking, 26(4): 1921–1932, August 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Al-Abbasi:2019:MTC

[AAAR19] Abubakr O. Al-Abbasi, Vaneet Aggarwal, and Moo-Ryong Ra. Multi-tier caching analysis in CDNbased over-the-top video streaming systems. IEEE/ ACM Transactions on Networking, 27(2):835–847, April 2019. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Altman:2005:SMT

[AAB05] Eitan Altman, Konstantin Avrachenkov, and Chadi Barakat. A stochastic model of TCP/IP with stationary random losses. IEEE/ACM Transactions on Networking, 13(2):356– 369, April 2005. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Altman:2013:COC

[AABD13] Eitan Altman, Amar Prakash Azad, Tamer Basar, and Francesco De Pellegrini. Combined optimal control of activation and transmission in delay-tolerant networks. IEEE/ACM Transactions on Networking, 21 (2):482–494, April 2013. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Abu-Amara:1996:STM

[AACD^{+}96] Hosame Abu-Amara, Brian A. Coan, Shlomi Dolev, Arkady Kanevsky, and Jennifer L. Welch. Self-stabilizing topology maintenance protocols for high-speed networks. IEEE/ACM Transactions on Networking, 4(6):902–912, December 1996. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL

http://www.acm.org/pubs/
citations/journals/ton/
1996-4-6/p902-abu-amara/

.

Amir:2005:CBF

[AADS05] Yair Amir, Baruch Awerbuch, Claudiu Danilov, and Jonathan Stanton. A costbenefit flow control for reliable multicast and unicast in overlay networks. IEEE/ ACM Transactions on Networking, 13(5):1094–1106, October 2005. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Archambault:2016:RSA

[AAF^{+}16] Emile Archambault, Nabih Alloune, Marija Furdek, Zhenyu Xu, Christine Tremblay, Ajmal Muhammad, Jiajia Chen, Lena Wosinska, Paul Littlewood, and Michel P. Belanger. Routing and spectrum assignment in elastic filterless optical networks. IEEE/ ACM Transactions on Networking, 24(6):3578–3592, December 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Al-Ayyoub:2014:TSA

[AAG14] Mahmoud Al-Ayyoub and Himanshu Gupta. Truthful spectrum auctions with approximate social-welfare or revenue. IEEE/ACM Transactions on Networking, 22(6):1873–1885, December 2014. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Applegate:2016:OCP

[AAG^{+}16] David Applegate, Aaron Archer, Vijay Gopalakrishnan, Seungjoon Lee, and K. K. Ramakrishnan. Optimal content placement for a large-scale VoD system. IEEE/ACM Transactions on Networking, 24(4): 2114–2127, August 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Akyildiz:2005:RCS

[AAM05] Ian F. Akyildiz, Özgür B. Akan, and Giacomo Morabito. A rate control scheme for adaptive realtime applications in IP networks with lossy links and long round trip times. IEEE/ACM Transactions on Networking, 13(3):554– 567, June 2005. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Alasadi:2018:SSU

[AAR18] Emad Alasadi and Hamed S.

Al-Raweshidy. SSED: Servers under softwaredefined network architectures to eliminate discovery messages. IEEE/ ACM Transactions on Networking, 26(1):104–117, February 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

AlDaoud:2010:PSS

[AAS10] Ashraf Al Daoud, Murat Alanyali, and David Starobinski. Pricing strategies for spectrum lease in secondary markets. IEEE/ ACM Transactions on Networking, 18(2):462–475, April 2010. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Antikainen:2014:DSA

[AAS14] Markku Antikainen, Tuomas Aura, and Mikko Särelä. Denial-of-service attacks in Bloom-filter-based forwarding. IEEE/ ACM Transactions on Networking, 22(5):1463–1476, October 2014. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Ahmad:2023:NFC

[AAT^{+}23] Mukhtiar Ahmad, Syed Muhammad Nawazish Ali, Muhammad Taimoor Tariq, Syed Usman Jafri, Adnan Abbas, Syeda Mashal Abbas Zaidi, Muhammad Basit Iqbal Awan, Zartash Afzal Uzmi, and Zafar Ayyub Qazi. Neutrino: a fast and consistent edge-based cellular control plane. IEEE/ ACM Transactions on Networking, 31(2):754–769, April 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2022.3202496.

Avallone:2009:CRA

[AAV09] Stefano Avallone, Ian F. Akyildiz, and Giorgio Ventre. A channel and rate assignment algorithm and a layer-2.5 forwarding paradigm for multi-radio wireless mesh networks. IEEE/ACM Transactions on Networking, 17(1):267– 280, February 2009. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Andrews:2012:RPM

[AAZZ12] Matthew Andrews, Antonio Fernández Anta, Lisa Zhang, and Wenbo Zhao. Routing for power minimization in the speed scaling model. IEEE/ ACM Transactions on Networking, 20(1):285–294, February 2012. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Alpcan:2005:GSA

[AB05] Tansu Alpcan and Tamer Basar. A globally stable adaptive congestion control scheme for Internetstyle networks with delay. IEEE/ACM Transactions on Networking, 13(6): 1261–1274, December 2005. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Alicherry:2007:SPP

[AB07] Mansoor Alicherry and Randeep Bhatia. Simple pre-provisioning scheme to enable fast restoration. IEEE/ACM Transactions on Networking, 15(2):400– 412, April 2007. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Ahmed:2009:PSP

[AB09] Reaz Ahmed and Raouf Boutaba. Plexus: a scalable peer-to-peer protocol enabling efficient subset search. IEEE/ACM Transactions on Networking, 17(1):130–143, February 2009. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Abdelmoniem:2021:RFR

[AB21] Ahmed M. Abdelmoniem and Brahim Bensaou. T-RACKs: a faster recovery mechanism for TCP in data center networks. IEEE/ACM Transactions on Networking, 29(3):1074– 1087, June 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2021.3059913.

Abdelmoniem:2023:ETH

[AB23] Ahmed M. Abdelmoniem and Brahim Bensaou. Enhancing TCP via hysteresis switching: Theoretical analysis and empirical evaluation. IEEE/ ACM Transactions on Networking, 31(6):2614– 2623, 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2023.3262564.

Anand:2016:EVA

[ABA^{+}16] Ashok Anand, Athula Balachandran, Aditya Akella, Vyas Sekar, and Srinivasan Seshan. Enhancing video accessibility and availability using informationbound references. IEEE/ ACM Transactions on Networking, 24(2):1223–1236, April 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Afek:2019:ZDS

[ABBF19] Yehuda Afek, Anat Bremler-Barr, and Shir Landau Feibish. Zero-day signature extraction for highvolume attacks. IEEE/ ACM Transactions on Networking, 27(2):691–706, April 2019. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Afek:2016:MDE

[ABBH^{+}16] Yehuda Afek, Anat Bremler-Barr, Yotam Harchol, David Hay, and Yaron Koral. Making DPI engines resilient to algorithmic complexity attacks. IEEE/ ACM Transactions on Networking, 24(6):3262–3275, December 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Afek:2001:RC

[ABBHP01] Yehuda Afek, Anat Bremler-Barr, and Sariel Har-Peled. Routing with a clue. IEEE/ ACM Transactions on Networking, 9(6):693–705, December 2001. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Ahmed:2016:RN

[ABC^{+}16] Reaz Ahmed, Faizul Bari, Shihabur Rahman Chowdhury, Golam Rabbani, Raouf Boutaba, Bertrand Mathieu, Reaz Ahmed, Faizul Bari, Shihabur Rahman Chowdhury, Golam Rabbani, Raouf Boutaba, and Bertrand Mathieu. αRoute: Routing on names. IEEE/ACM Transactions on Networking, 24(5):3070–3083, October 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Athanasopoulou:2013:BPB

[ABJ^{+}13] Eleftheria Athanasopoulou, Loc X. Bui, Tianxiong Ji, R. Srikant, and Alexander Stolyar. Back-pressure-based packet-by-packet adaptive routing in communication networks. IEEE/ ACM Transactions on Networking, 21(1):244–257, February 2013. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Antonakopoulos:2015:FPM

[ABK15] Spyridon Antonakopoulos, Yigal Bejerano, and Pramod Koppol. Full protection made easy: the DisPath IP fast reroute scheme. IEEE/ACM Transactions on Networking, 23(4):1229–1242, August 2015. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Arrigoni:2023:BAN

[ABMT23] Viviana Arrigoni, Novella Bartolini, Annalisa Massini, and Federico Trombetti. A Bayesian approach to network monitoring for progressive failure localization. IEEE/ACM Transactions on Networking, 31(2):770– 783, April 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2022.3200249.

Atya:2016:PAE

[ABS^{+}16] Ahmed Osama Fathy Atya, Ioannis Broustis, Shailendra Singh, Dimitris Syrivelis, Srikanth V. Krishnamurthy, and Thomas F. La Porta. A policyaware enforcement logic for appropriately invoking network coding. IEEE/ ACM Transactions on Networking, 24(4):2005–2018, August 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Aharoni:1998:RDS

[AC98] Ehud Aharoni and Reuven Cohen. Restricted dynamic Steiner trees for scalable multicast in datagram networks. IEEE/ ACM Transactions on Networking, 6(3):286–297, June 1998. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/
                                        journals/ton/1998-6-3/
                                        p286-aharoni/.

Applegate:2006:MRR

[AC06] David Applegate and Edith Cohen. Making routing robust to changing traffic demands: algorithms and evaluation. IEEE/ACM Transactions on Networking, 14(6):1193–1206, December 2006. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Argyraki:2009:SNL

[AC09] Katerina Argyraki and David R. Cheriton. Scalable network-layer defense against Internet bandwidth-flooding attacks. IEEE/ACM Transactions on Networking, 17(4):1284–1297, August 2009. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Abdolee:2016:DLS

[AC16] Reza Abdolee and Benoit Champagne. Diffusion LMS strategies in sensor networks with noisy input data. IEEE/ACM Transactions on Networking, 24 (1):3–14, February 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Ayoubi:2016:TPB

[ACA16] Sara Ayoubi, Yiheng Chen, and Chadi Assi. Towards promoting backupsharing in survivable virtual network design. IEEE/ ACM Transactions on Networking, 24(5):3218–3231, October 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Amer:1994:PTS

[ACC^{+}94] Paul D. Amer, Christophe Chassot, Thomas J. Connolly, Michel Diaz, and Phillip Conrad. Partialorder transport service for multimedia and other applications. IEEE/ACM Transactions on Networking, 2(5):440–456, October 1994. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1994-2-5/
p440-amer/.

Addis:2014:EMT

[ACC^{+}14] Bernardetta Addis, Antonio Capone, Giuliana Carello, Luca G. Gianoli, and Brunilde Sanso. Energy management through optimized routing and device powering for greener communication networks. IEEE/ACM Transactions on Networking, 22(1):313– 325, February 2014. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Amaldi:2012:DWS

[ACCF12] Edoardo Amaldi, Antonio Capone, Matteo Cesana, and Ilario Filippini. Design of wireless sensor networks for mobile target detection. IEEE/ACM Transactions on Networking, 20(3): 784–797, June 2012. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Ambrosin:2017:LTC

[ACDP17] Moreno Ambrosin, Mauro Conti, Fabio De Gaspari, and Radha Poovendran. LineSwitch: Tackling control plane saturation attacks in software-defined networking. IEEE/ACM Transactions on Networking, 25(2):1206–1219, April 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Anshelevich:2014:CAG

[ACKZ14] Elliot Anshelevich, Bugra Caskurlu, Koushik Kar, and Hang Zhang. Capacity allocation games for network-coded multicast streaming. IEEE/ACM Transactions on Networking, 22(2):595–607, April 2014. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Aggarwal:2017:SFC

[ACLX17] Vaneet Aggarwal, Yih-Farn Robin Chen, Tian Lan, and Yu Xiang. Sprout: a functional caching approach to minimize service latency in erasurecoded storage. IEEE/ACM Transactions on Networking, 25(6):3683–3694, December 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Agrawal:1999:PBF

[ACOR99] Rajeev Agrawal, Rene L.

Cruz, Clayton Okino, and Rajendran Rajan. Performance bonds for flow control protocols. IEEE/ ACM Transactions on Networking, 7(3):310–323, June 1999. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1999-7-3/
p310-agrawal/.

Ascia:2005:EMS

[ACP05] Giuseppe Ascia, Vincenzo Catania, and Daniela Panno. An evolutionary management scheme in high-performance packet switches. IEEE/ACM Transactions on Networking, 13(2):262–275, April 2005. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Aezladen:2012:ELB

[ACR12] Mhameed Aezladen, Reuven Cohen, and Danny Raz. Efficient location-based decision-supporting content distribution to mobile groups. IEEE/ACM Transactions on Networking, 20(5):1514–1526, October 2012. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Afanasyev:2010:UPU

[ACVS10] Mikhail Afanasyev, Tsuwei Chen, Geoffrey M. Voelker, and Alex C. Snoeren. Usage patterns in an urban WiFi network. IEEE/ ACM Transactions on Networking, 18(5):1359–1372, October 2010. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Alasmar:2021:ITV

[ACZP21] Mohammed Alasmar, Richard Clegg, Nickolay Zakhleniuk, and George Parisis. Internet traffic volumes are not Gaussian — they are log-normal: an 18-year longitudinal study with implications for modelling and prediction. IEEE/ACM Transactions on Networking, 29(3):1266– 1279, June 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.

acm.org/doi/10.1109/TNET.
2021.3059542.

Ahn:1996:PNS

[AD96] Jong Suk Ahn and Peter B. Danzig. Packet network simulation: speedup and accuracy versus timing granularity. IEEE/ ACM Transactions on Networking, 4(5):743–757, October 1996. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1996-4-5/
p743-ahn/.

Altman:2011:FCF

[AD11] Eitan Altman and Francesco De Pellegrini. Forward correction and fountain codes in delay-tolerant networks. IEEE/ACM Transactions on Networking, 19(1):1–13, February 2011. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Ahlehagh:2014:VAS

[AD14] Hasti Ahlehagh and Sujit Dey. Video-aware scheduling and caching in the radio access network. IEEE/ ACM Transactions on Networking, 22(5):1444–1462, October 2014. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Agarwal:2018:RBA

[AD18] Satyam Agarwal and Swades De. Rural broadband access via clustered collaborative communication. IEEE/ACM Transactions on Networking, 26(5):2160– 2173, October 2018. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Adas:1998:UAL

[Ada98] Abdelnaser Mohammad Adas. Using adaptive linear prediction to support real-time VBR video under RCBR network service model. IEEE/ACM Transactions on Networking, 6(5):635–644, October 1998. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1998-6-5/
p635-adas/.

Arifler:2007:FAA

[AdE07] Dogu Arifler, Gustavo de Veciana, and Brian L. Evans. A factor analytic approach to inferring congestion sharing based on flow level measurements. IEEE/ACM Transactions on Networking, 15(1):67– 79, February 2007. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Araldo:2018:CEC

[ADR18] Andrea Araldo, Gyorgy Dan, and Dario Rossi. Caching encrypted content via stochastic cache partitioning. IEEE/ACM Transactions on Networking, 26(1):548–561, February 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Agra:2016:MCD

[AdSD16] Agostinho Agra, Amaro de Sousa, and Mahdi Doostmohammadi. The minimum cost design of transparent optical networks combining grooming, routing, and wavelength assignment. IEEE/ ACM Transactions on Networking, 24(6):3702–3713, December 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Akram:2022:CAD

[ADT22] Vahid Khalilpour Akram, Orhan Dagdeviren, and Bulent Tavli. A coverageaware distributed k-connectivity maintenance algorithm for arbitrarily large k in mobile sensor networks. IEEE/ ACM Transactions on Networking, 30(1):62–75, February 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.

                                        acm.org/doi/10.1109/TNET.
                                        2021.3104356.

Anand:2020:JSU

[AdVS20] Arjun Anand, Gustavo de Veciana, and Sanjay Shakkottai. Joint scheduling of URLLC and eMBB traffic in 5G wireless networks. IEEE/ ACM Transactions on Networking, 28(2):477–490, April 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL https://

                                        dl.acm.org/doi/abs/10.
                                        1109/TNET.2020.2968373.

Akyildiz:2002:MNR

[AEB02] Ian F. Akyildiz, Eylem Ekici, and Michael D. Bender. MLSR: a novel routing algorithm for multilayered satellite IP networks. IEEE/ACM Transactions on Networking, 10(3):411– 424, June 2002. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Agarwal:2013:RWN

[AEG^{+}13] Pankaj K. Agarwal, Alon Efrat, Shashidhara K. Ganjugunte, David Hay, Swaminathan Sankararaman, and Gil Zussman. The resilience of WDM networks to probabilistic geographical failures. IEEE/ ACM Transactions on Networking, 21(5):1525–1538, October 2013. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Ahmed:2017:DLE

[AEG^{+}17] Faraz Ahmed, Jeffrey Erman, Zihui Ge, Alex X. Liu, Jia Wang, and He Yan. Detecting and localizing endto-end performance degradation for cellular data services based on TCP loss ratio and round trip time. IEEE/ACM Transactions on Networking, 25(6): 3709–3722, December 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Ayesta:2013:SRE

[AEJV13] Urtzi Ayesta, Martin Erausquin, Matthieu Jonckheere, and Ina Maria Verloop. Scheduling in a random environment: stability and asymptotic optimality. IEEE/ACM Transactions on Networking, 21(1): 258–271, February 2013. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Aksoy:1999:SAL

[AF99] Demet Aksoy and Michael Franklin. R × W: a scheduling approach for large-scale on-demand data broadcast. IEEE/ACM Transactions on Networking, 7(6):846–860, December 1999. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/
                                        journals/ton/1999-7-6/
                                        p846-aksoy/.

Augustin:2011:MMR

[AFT11] Brice Augustin, Timur Friedman, and Renata Teixeira. Measuring multipath routing in the Internet. IEEE/ACM Transactions on Networking, 19 (3):830–840, June 2011. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Afrasiabi:2016:EUP

[AG16] Mohammad Hadi Afrasiabi and Roch Guérin. Exploring user-provided connectivity. IEEE/ACM Transactions on Networking, 24(1):542–554, February 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Akhtar:2023:FSR

[AGBS23] Md Shahbaz Akhtar, Krishnakumar G., Vishnu B., and Abhishek Sinha. Fast and secure routing algorithms for quantum key distribution networks. IEEE/ACM Transactions on Networking, 31(5):2281– 2296, 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.

acm.org/doi/10.1109/TNET.
2023.3246114.

Aguirre-Guerrero:2018:WGC

[AGCFV18] Daniela Aguirre-Guerrero, Miguel Camelo, Lluis Fabrega, and Pere Vila. WMGR: a generic and compact routing scheme for data center networks. IEEE/ACM Transactions on Networking, 26(1):356– 369, February 2018. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Athanasiadou:2016:SXB

[AGGT16] Sophia Athanasiadou, Marios Gatzianas, Leonidas Georgiadis, and Leandros Tassiulas. Stable XORbased policies for the broadcast erasure channel with feedback. IEEE/ ACM Transactions on Networking, 24(1):476–491, February 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Auerbach:2003:MGM

[AGKK03] Joshua Auerbach, Madan Gopal, Marc Kaplan, and Shay Kutten. Multicast group membership management. IEEE/ ACM Transactions on Networking, 11(1):166–175, February 2003. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Alfano:2016:CCW

[AGL16] Giusi Alfano, Michele Garetto, and Emilio Leonardi. Content-centric wireless networks with limited buffers: when mobility hurts. IEEE/ACM Transactions on Networking, 24(1):299–311, February 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Alfano:2010:CSW

[AGLM10] Giusi Alfano, Michele Garetto, Emilio Leonardi, and Valentina Martina. Capacity scaling of wireless networks with inhomogeneous node density: lower bounds. IEEE/ ACM Transactions on Networking, 18(5):1624–1636, October 2010. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Aoudia:2017:GFM

[AGM^{+}17] Faycal Ait Aoudia, Matthieu Gautier, Michele Magno, Olivier Berder, and Luca Benini. A generic framework for modeling MAC protocols in wireless sensor networks. IEEE/ACM Transactions on Networking, 25(3):1489–1500, June 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Ahmadi:2021:PSQ

[AGMY21] Mahdieh Ahmadi, Morteza Golkarifard, Ali Movaghar, and Hamed Yousefi. Processor sharing queues with impatient customers and state-dependent rates. IEEE/ ACM Transactions on Networking, 29(6):2467–2477, December 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2021.3091189.

Ayan:2024:OFH

[AHEK24] Onur Ayan, Sandra Hirche, Anthony Ephremides, and Wolfgang Kellerer. Optimal finite horizon scheduling of wireless networked control systems. IEEE/ ACM Transactions on Networking, 32(2):927–942, April 2024. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2023.3300949.

Aad:2008:IDS

[AHK08] Imad Aad, Jean-Pierre Hubaux, and Edward W. Knightly. Impact of denial of service attacks on ad hoc networks. IEEE/ACM Transactions on Networking, 16(4):791–802, August 2008. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Akyildiz:1996:MLU

[AHL96] Ian F. Akyildiz, Joseph S. M. Ho, and Yi-Bing Lin. Movement-based location update and selective paging for PCS networks. IEEE/ACM Transactions on Networking, 4(4):629– 638, August 1996. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/
                                        journals/ton/1996-4-4/
                                        p629-akyildiz/.

Ao:2021:JWD

[AHP21] Weng Chon Ao, Po-Han Huang, and Konstantinos Psounis. Joint workload distribution and capacity augmentation in hybrid datacenter networks. IEEE/ACM Transactions on Networking, 29(1):120– 133, February 2021. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1109/
                                        TNET.2020.3027607.

Albazrqaoe:2019:PBT

[AHX19] Wahhab Albazrqaoe, Jun Huang, and Guoliang Xing. A practical Bluetooth traffic sniffing system: Design, implementation, and countermeasure. IEEE/ACM Transactions on Networking, 27(1):71–84, February 2019. CODEN IEANEP.

ISSN 1063-6692 (print), 1558-2566 (electronic).

Anderson:2023:LLO

[AIL23] Daron Anderson, George Iosifidis, and Douglas J. Leith. Lazy Lagrangians for optimistic learning with budget constraints. IEEE/ACM Transactions on Networking, 31(5):1935– 1949, 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2022.3222404.

Abadal:2015:AES

[AIN^{+}15] Sergi Abadal, Mario Iannazzo, Mario Nemirovsky, Albert Cabellos-Aparicio, Heekwan Lee, and Eduard Alarcón. On the area and energy scalability of wireless network-on-chip: a model-based benchmarked design space exploration. IEEE/ACM Transactions on Networking, 23(5):1501– 1513, October 2015. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Awduche:2006:DAC

[AJ06] Daniel O. Awduche and Bijan Jabbari. Demand assigned capacity management (DACM) in IP over optical (IPO) networks. IEEE/ACM Transactions on Networking, 14(4):900– 913, August 2006. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Akyildiz:2001:AFS

[AJDH01] Ian F. Akyildiz, Inwhee Joe, Henry Driver, and Yung-Lung Ho. An adaptive FEC scheme for data traffic in wireless ATM networks. IEEE/ACM Transactions on Networking, 9 (4):419–426, August 2001. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Aperjis:2011:BME

[AJF11] Christina Aperjis, Ramesh Johari, and Michael J. Freedman. Bilateral and multilateral exchanges for peer-assisted content distribution. IEEE/ACM Transactions on Networking, 19(5):1290–1303, October 2011. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Ahmad:2006:OBC

[AJV06] Sahand Haji Ali Ahmad, Aleksandar Jovičić, and Pramod Viswanath. On outer bounds to the capacity region of wireless networks. IEEE/ACM Transactions on Networking, 14 (SI):2770–2776, June 2006. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Altman:1996:BPM

[AK96] Eitan Altman and Daniel Kofman. Bounds for performance measures of token rings. IEEE/ ACM Transactions on Networking, 4(2):292–299, April 1996. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1996-4-2/
p292-altman/.

Acampora:2000:NAM

[AK00] Anthony S. Acampora and Srikanth V. Krishnamurthy. A new adaptive MAC layer protocol for broadband packet wireless networks in harsh fading and interference environments. IEEE/ACM Transactions on Networking, 8 (3):328–336, 2000. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/2000-8-3/
p328-acampora/.

Abraham:2001:NAA

[AK01] Santosh Paul Abraham and Anurag Kumar. A new approach for asynchronous distributed rate control of elastic sessions in integrated packet networks. IEEE/ACM Transactions on Networking, 9 (1):15–30, 2001. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/articles/
                                        journals/ton/2001-9-1/
                                        p15-abraham/p15-abraham.
                                        pdf;   http://www.acm.
                                        org/pubs/citations/journals/
                                        ton/2001-9-1/p15-abraham/

.

Ahn:2009:SLD

[AK09] Joon Ahn and Bhaskar Krishnamachari. Scaling laws for data-centric storage and querying in wireless sensor networks. IEEE/ ACM Transactions on Networking, 17(4):1242–1255, August 2009. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Almasaeid:2014:EMD

[AK14] Hisham M. Almasaeid and Ahmed E. Kamal. Exploiting multichannel diversity for cooperative multicast in cognitive radio mesh networks. IEEE/ACM Transactions on Networking, 22 (3):770–783, June 2014. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Almasaeid:2015:RBC

[AK15] Hisham M. Almasaeid and Ahmed E. Kamal. Receiver-based channel allocation in cognitive radio wireless mesh networks.

IEEE/ACM Transactions on Networking, 23(4):1286– 1299, August 2015. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Acer:2010:WSR

[AKA10] Utku Günay Acer, Shivkumar Kalyanaraman, and Alhussein A. Abouzeid. Weak state routing for large-scale dynamic networks. IEEE/ACM Transactions on Networking, 18(5):1450–1463, October 2010. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Al-Kofahi:2013:SRS

[AKK13] Osameh M. Al-Kofahi and Ahmed E. Kamal. Scalable redundancy for sensors-to-sink communication. IEEE/ACM Transactions on Networking, 21(6): 1774–1784, December 2013. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Ahuja:1996:DIP

[AKS96] R. Ahuja, S. Keshav, and H. Saran. Design, implementation, and performance measurement of a native-mode ATM transport layer (extended version). IEEE/ACM Transactions on Networking, 4(4):502–515, August 1996. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/
                                        journals/ton/1996-4-4/
                                        p502-ahuja/.

Aryafar:2013:AAB

[AKS^{+}13] Ehsan Aryafar, Mohammad Ali Khojastepour, Karthik Sundaresan, Sampath Rangarajan, and Edward Knightly. ADAM: an adaptive beamforming system for multicasting in wireless LANs. IEEE/ ACM Transactions on Networking, 21(5):1595–1608, October 2013. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Altman:2012:SSG

[AKSS12] Eitan Altman, Anurag Kumar, Chandramani Singh, and Rajesh Sundaresan. Spatial SINR games of base station placement and mobile association. IEEE/ ACM Transactions on Networking, 20(6):1856–1869, December 2012. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Anerousis:1998:VPC

[AL98] Nikolaos Anerousis and Aurel A. Lazar. Virtual path control for ATM networks with call level quality of service guarantees.

IEEE/ACM Transactions on Networking, 6(2):222– 236, April 1998. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1998-6-2/
p222-anerousis/.

Ali:2006:GSS

[Ali06] Maher Ali. Generalized sharing in survivable optical networks. IEEE/ACM Transactions on Networking, 14(6):1388–1399, December 2006. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Akyildiz:1999:SCP

[ALJ99] Ian F. Akyildiz, David A.

Levine, and Inwhee Joe. A slotted CDMA protocol with BER scheduling for wireless multimedia networks. IEEE/ ACM Transactions on Networking, 7(2):146–158, April 1999. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1999-7-2/
p146-akyildiz/.

Aurzada:2014:FAN

[ALMR14] Frank Aurzada, Martin Lévesque, Martin Maier, and Martin Reisslein. FiWi access networks based on next-generation PON and gigabit-class WLAN technologies: a capacity and delay analysis. IEEE/ ACM Transactions on Networking, 22(4):1176–1189, August 2014. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Ali:2021:DSS

[ALPK21] Kamran Ali, Alex X. Liu, Ioannis Pefkianakis, and Kyu-Han Kim. Distributed spectrum sharing for enterprise powerline communication networks. IEEE/ ACM Transactions on Networking, 29(3):1032–1045, June 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2021.3056512.

Abdisarabshali:2024:DTM

[ALR^{+}24]  Payam  Abdisarabshali,

Minghui Liwang, Amir Rajabzadeh, Mahmood Ahmadi, and Seyyedali Hosseinalipour. Decomposition theory meets reliability analysis: Processing of computation-intensive dependent tasks over vehicular clouds with dynamic resources. IEEE/ ACM Transactions on Networking, 32(1):475– 490, 2024. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2023.3286709.

Andrew:2009:UXE

[ALW09] Lachlan L. H. Andrew, Steven H. Low, and Bartek P. Wydrowski. Understanding XCP: equilibrium and fairness. IEEE/ ACM Transactions on Networking, 17(6):1697–1710, December 2009. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Alderson:2005:UIT

[ALWD05] David Alderson, Lun Li, Walter Willinger, and John C. Doyle. Understanding Internet topology: principles, models, and validation. IEEE/ ACM Transactions on Networking, 13(6):1205–1218, December 2005. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

An:2020:ABF

[ALY^{+}20] Zhenlin An, Qiongzheng Lin, Lei Yang, Wei Lou, and Lei Xie. Acquiring Bloom filters across commercial RFIDs in physical layer. IEEE/ACM Transactions on Networking, 28(4):1804–1817, August 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2020.2992851.

An:2022:TAW

[ALYX22] Zhenlin An, Qiongzheng Lin, Lei Yang, and Lei Xie. Tagcaster: Activating wireless voice of electronic toll collection systems with zero start-up cost. IEEE/ACM Transactions on Networking, 30(5): 2328–2342, 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2022.3169914.

Agrawal:2016:EIU

[AM16] Gaurav Agrawal and Deep Medhi. Embedding IP unique shortest path topology on a wavelength-routed network: normal and survivable design. IEEE/ACM Transactions on Networking, 24(2):1109–1124, April 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Avner:2019:MUC

[AM19] Orly Avner and Shie Mannor. Multi-user communication networks: a coordinated multi-armed bandit approach. IEEE/ACM Transactions on Networking, 27(6):2192–2207, December 2019. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL https://

dl.acm.org/doi/abs/10.
1109/TNET.2019.2935043.

Agarwal:2019:VPR

[AMCD19] Satyam Agarwal, Francesco Malandrino, Carla Fabiana Chiasserini, and Swades De. VNF placement and resource allocation for the support of vertical services in 5G networks. IEEE/ ACM Transactions on Networking, 27(1):433–446, February 2019. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Asadi:2017:DEF

[AMG^{+}17] Arash Asadi, Vincenzo Mancuso, Rohit Gupta, Arash Asadi, Vincenzo Mancuso, and Rohit Gupta. DORE: an experimental framework to enable outband D2D relay in cellular networks. IEEE/ ACM Transactions on Networking, 25(5):2930–2943, October 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Abel:2007:DIN

[AMI^{+}07] François Abel, Cyriel Minkenberg, Ilias Iliadis, Ton Engbersen, Mitchell Gusat, Ferdinand Gramsamer, and Ronald P. Luijten. Design issues in nextgeneration merchant switch fabrics. IEEE/ACM Transactions on Networking, 15 (6):1603–1615, December 2007. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Al-Mouhamed:1999:EPD

[AMKY99] Mayez A. Al-Mouhamed, Mohammed Kaleemuddin, and Habib Yousef. Evaluation of pipelined dilated banyan switch architectures for ATM networks. IEEE/ACM Transactions on Networking, 7(5):724–740, October 1999. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/
                                        journals/ton/1999-7-5/
                                        p724-al-mouhamed/.

Ammar:2002:E

[Amm02] Mostafa H. Ammar. Editorial. IEEE/ACM Transactions on Networking, 10(1): 1, February 2002. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Ammar:2003:E

[Amm03] Mostafa Ammar. Editorial. IEEE/ACM Transactions on Networking, 11(1): 1, February 2003. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Akyildiz:2001:TPN

[AMP01] Ian F. Akyildiz, Giacomo Morabito, and Sergio Palazzo. TCP-Peach: a new congestion control scheme for satellite IP networks. IEEE/ACM Transactions on Networking, 9 (3):307–321, June 2001. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Akella:2008:CPB

[AMS^{+}08] Aditya Akella, Bruce Maggs, Srinivasan Seshan, Anees Shaikh, and Ramesh K. Sitaraman. Corrections to ‘On the performance benefits of multihoming route control’. IEEE/ACM Transactions on Networking, 16(2):496, April 2008. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). See [AMSS08].

Avin:2022:DAN

[AMS22a] Chen Avin, Kaushik Mondal, and Stefan Schmid. Demand-aware network design with minimal congestion and route lengths. IEEE/ACM Transactions on Networking, 30(4):1838– 1848, 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2022.3153586.

Avin:2022:PTO

[AMS22b] Chen Avin, Kaushik Mondal, and Stefan Schmid. Push-down trees: Optimal self-adjusting complete trees. IEEE/ ACM Transactions on Networking, 30(6):2419– 2432, 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2022.3174118.

Ayepah-Mensah:2024:BEF

[AMSB^{+}24] Daniel Ayepah-Mensah, Guolin Sun, Gordon Owusu Boateng, Stephen Anokye, and Guisong Liu. Blockchainenabled federated learningbased resource allocation and trading for network slicing in 5G. IEEE/ ACM Transactions on Networking, 32(1):654– 669, 2024. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2023.3297390.

Akella:2008:PBM

[AMSS08] Aditya Akella, Bruce Maggs, Srinivasan Seshan, and Anees Shaikh. On the performance benefits of multihoming route control. IEEE/ACM Transactions on Networking, 16(1):91– 104, February 2008. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). See corrections [AMS^{+}08].

Akhbarizadeh:2005:HBI

[AN05] Mohammad J. Akhbarizadeh and Mehrdad Nourani. Hardware-based IP routing using partitioned lookup table. IEEE/ACM Transactions on Networking, 13 (4):769–781, August 2005. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Asheralieva:2020:CCT

[AN20] Alia Asheralieva and Dusit Niyato. Combining contract theory and Lyapunov optimization for content sharing with edge caching and device-to-device communications. IEEE/ACM Transactions on Networking, 28(3):1213– 1226, June 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2020.2978117.

Andrew:2004:FSW

[And04] Lachlan L. H. Andrew.

Fast simulation of wavelength continuous WDM networks. IEEE/ACM Transactions on Networking, 12(4):759–765, August 2004. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Altman:2013:PIM

[ANSX13] Eitan Altman, Philippe Nain, Adam Shwartz, and Yuedong Xu. Predicting the impact of measures against P2P networks: transient behavior and phase transition. IEEE/ACM Transactions on Networking, 21(3):935– 949, June 2013. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Al-Naday:2017:ICM

[ANTR17] Mays F. Al-Naday, Nikolaos Thomos, and Martin J. Reed. Informationcentric multilayer networking: Improving performance through an ICN/ WDM architecture. IEEE/ ACM Transactions on Networking, 25(1):83–97, February 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Aweya:2004:DSA

[AOM04] James Aweya, Michel Ouellette, and Delfin Y. Montuno. Design and stability analysis of a rate control algorithm using the Routh– Hurwitz stability criterion. IEEE/ACM Transactions on Networking, 12(4):719– 732, August 2004. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Abbott:1993:INT

[AP93a] Mark B. Abbott and Larry L. Peterson. Increasing network throughput by integrating protocol layers. IEEE/ACM Transactions on Networking, 1(5):600–610, October 1993. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1993-1-5/
p600-abbott/.

Abbott:1993:LAP

[AP93b] Mark B. Abbott and Larry L. Peterson. A language-based approach to protocol implementation. IEEE/ACM Transactions on Networking, 1(1): 4–19, February 1993. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL http://

www.acm.org/pubs/citations/
journals/ton/1993-1-1/
p4-abbott/.

Ao:2017:AAO

[AP17] Weng Chon Ao and Konstantinos Psounis. Approximation algorithms for online user association in multi-tier multi-cell mobile networks. IEEE/ ACM Transactions on Networking, 25(4):2361–2374, August 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Arslan:2013:AAC

[APB^{+}13] Mustafa Y. Arslan, Konstantinos Pelechrinis, Ioannis Broustis, Shailendra Singh, Srikanth V. Krishnamurthy, Sateesh Addepalli, and Konstantina Papagiannaki. ACORN: an auto-configuration framework for 802.11n WLANs. IEEE/ACM Transactions on Networking, 21(3):896– 909, June 2013. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Alasmar:2021:SSR

[APC21] Mohammed Alasmar, George Parisis, and Jon Crowcroft. SCDP: Systematic rateless coding for efficient data transport in data centers. IEEE/ACM Transactions on Networking, 29(6):2723–2736, December 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2021.3098386.

Anderson:2014:ODS

[APSG14] Eric Anderson, Caleb Phillips, Douglas Sicker, and Dirk Grunwald. Optimization decomposition for scheduling and system configuration in wireless networks. IEEE/ACM Transactions on Networking, 22(1):271–284, February 2014. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Aparicio-Pardo:2012:NRV

[APSKPMGM12] Ramon Aparicio-Pardo, Nina Skorin-Kapov, Pablo Pavon-Marino, and Belen Garcia-Manrubia. (Non- )reconfigurable virtual topology design under multihour traffic in optical networks. IEEE/ACM Transactions on Networking, 20(5):1567– 1580, October 2012. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Al-Qudah:2016:ITD

[AQJRS16] Zakaria Al-Qudah, Eamon Johnson, Michael Rabinovich, and Oliver Spatscheck. Internet with transient destinationcontrolled addressing. IEEE/ ACM Transactions on Networking, 24(2):731–744, April 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Atya:2019:CMI

[AQK^{+}19] Ahmed Osama Fathy Atya, Zhiyun Qian, Srikanth V. Krishnamurthy, Thomas La Porta, Patrick Mc-Daniel, and Lisa M. Marvel. Catch me if you can: a closer look at malicious co-residency on the cloud. IEEE/ACM Transactions on Networking, 27(2):560– 576, April 2019. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Ahuja:2009:SLF

[ARK09] Satyajeet S. Ahuja, Srinivasan Ramasubramanian, and Marwan M. Krunz. Single-link failure detection in all-optical networks using monitoring cycles and paths. IEEE/ ACM Transactions on Networking, 17(4):1080–1093, August 2009. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Ahuja:2011:SFL

[ARK11] Satyajeet S. Ahuja, Srinivasan Ramasubramanian, and Marwan Krunz. SRLG failure localization in optical networks. IEEE/ACM Transactions on Networking, 19(4):989–999, August 2011. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Avalle:2016:SAN

[ARS16] Matteo Avalle, Fulvio Risso, and Riccardo Sisto. Scalable algorithms for NFA multi-striding and NFA-based deep packet inspection on GPUs. IEEE/ ACM Transactions on Networking, 24(3):1704–1717, June 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Abdelaziz:1994:SOT

[AS94] Mohamed Abdelaziz and Ioannis Stavrakakis. Some optimal traffic regulation schemes for ATM networks: a Markov decision approach. IEEE/ ACM Transactions on Networking, 2(5):508–519, October 1994. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1994-2-5/
p508-abdelaziz/.

Asawa:1996:OSH

[AS96] Manjari Asawa and Wayne E.

Stark. Optimal scheduling of handoffs in cellular networks. IEEE/ ACM Transactions on Networking, 4(3):428–441, June 1996. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1996-4-3/
p428-asawa/.

Awerbuch:2001:TAD

[AS01] Baruch Awerbuch and Yuval Shavitt. Topology aggregation for directed graphs. IEEE/ACM Transactions on Networking, 9 (1):82–90, 2001. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/articles/
journals/ton/2001-9-1/
p82-awerbuch/p82-awerbuch.
pdf;   http://www.acm.
org/pubs/citations/journals/
ton/2001-9-1/p82-awerbuch/

.

Antoniou:2002:EDC

[AS02] Zoe Antoniou and Ioannis Stavrakakis. An efficient deadline-credit-based transport scheme for prerecorded semisoft continuous media applications. IEEE/ACM Transactions on Networking, 10(5):630– 643, October 2002. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Alparslan:2007:GRM

[AS07a] Denizhan N. Alparslan and Khosrow Sohraby. A generalized random mobility model for wireless ad hoc networks and its analysis: one-dimensional case. IEEE/ACM Transactions on Networking, 15(3):602– 615, June 2007. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Alparslan:2007:TDM

[AS07b] Denizhan N. Alparslan and Khosrow Sohraby. Twodimensional modeling and analysis of generalized random mobility models for wireless ad hoc networks. IEEE/ACM Transactions on Networking, 15(3):616– 629, June 2007. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Allalouf:2008:CDA

[AS08] Miriam Allalouf and Yuval Shavitt. Centralized and distributed algorithms for routing and weighted maxmin fair bandwidth allocation. IEEE/ACM Transactions on Networking, 16(5): 1015–1024, October 2008. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Agrawal:2009:HBN

[AS09] Banit Agrawal and Timothy Sherwood. Highbandwidth network memory system through virtual pipelines. IEEE/ ACM Transactions on Networking, 17(4):1029–1041, August 2009. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Abedini:2014:CCS

[AS14] Navid Abedini and Srinivas Shakkottai. Content caching and scheduling in wireless networks with elastic and inelastic traffic. IEEE/ACM Transactions on Networking, 22(3):864– 874, June 2014. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Aktas:2019:SMS

[AS19] Mehmet Fatih Aktaş and Emina Soljanin. Straggler mitigation at scale. IEEE/ ACM Transactions on Networking, 27(6):2266–2279, December 2019. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL https://

                                        dl.acm.org/doi/abs/10.
                                        1109/TNET.2019.2946464.

Ali:2008:SCA

[ASCG08] Zafar Ali, Waseem Sheikh, Edwin K. P. Chong, and Arif Ghafoor. A scalable call admission control algorithm. IEEE/ACM Transactions on Networking, 16 (2):424–434, April 2008. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Ahmed:2018:OIT

[ASKL18] Faraz Ahmed, M. Zubair Shafiq, Amir R. Khakpour, and Alex X. Liu. Optimizing Internet transit routing for content delivery networks. IEEE/ACM Transactions on Networking, 26 (1):76–89, February 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Arslan:2016:IIB

[ASKR16] Mustafa Y. Arslan, Karthikeyan Sundaresan, Srikanth V. Krishnamurthy, and Sampath Rangarajan. iBUS: an integrated beamformer and uplink scheduler for OFDMA small cells. IEEE/ ACM Transactions on Networking, 24(2):901–914, April 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Aryafar:2013:SCC

[ASSK13] Ehsan Aryafar, Theodoros Salonidis, Jingpu Shi, and Edward Knightly. Synchronized CSMA contention: model, implementation, and evaluation. IEEE/ ACM Transactions on Networking, 21(5):1349–1362, October 2013. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Aziz:2011:UTR

[AST11] Adel Aziz, David Starobinski, and Patrick Thiran. Understanding and tackling the root causes of instability in wireless mesh networks. IEEE/ ACM Transactions on Networking, 19(4):1178–1193, August 2011. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Abdalla:2000:KMR

[ASW00] Michel Abdalla, Yuval Shavitt, and Avishai Wool. Key management for restricted multicast using broadcast encryption. IEEE/ ACM Transactions on Networking, 8(4):443–454, 2000. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL

                                        http://www.acm.org/pubs/
                                        citations/journals/ton/
                                        2000-8-4/p443-abdalla/.

Anjum:2003:CSV

[AT03] Farooq Anjum and Leandros Tassiulas. Comparative study of various TCP versions over a wireless link with correlated losses. IEEE/ACM Transactions on Networking, 11(3):370– 383, June 2003. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Andrei:2010:PDD

[ATB^{+}10] Dragos Andrei, Massimo Tornatore, Marwan Batayneh, Charles U. Martel, and Biswanath Mukherjee. Provisioning of deadline-driven requests with flexible transmission rates in WDM mesh networks. IEEE/ ACM Transactions on Networking, 18(2):353–366, April 2010. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Abolhassani:2021:DGA

[ATE21] Bahman Abolhassani, John Tadrous, and Atilla Eryilmaz. Delay gain analysis of wireless multicasting for content distribution. IEEE/ACM Transactions on Networking, 29(2):529– 542, April 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2020.3039634.

Abolhassani:2022:SVD

[ATE22] Bahman Abolhassani, John Tadrous, and Atilla Eryilmaz. Single vs distributed edge caching for dynamic content. IEEE/ ACM Transactions on Networking, 30(2):669–682, April 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2021.3121098.

Abolhassani:2023:OLS

[ATE23] Bahman Abolhassani, John Tadrous, and Atilla Eryilmaz. Optimal loadsplitting and distributedcaching for dynamic content over the wireless edge. IEEE/ACM Transactions on Networking, 31(5):2178– 2190, 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2023.3244039.

Abolhassani:2022:FCD

[ATEY22] Bahman Abolhassani, John Tadrous, Atilla Eryilmaz, and Edmund Yeh. Fresh caching of dynamic content over the wireless edge. IEEE/ACM Transactions on Networking, 30(5):2315– 2327, 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2022.3170245.

Altieri:2014:ACS

[AVPG14] Andrés Altieri, Leonardo Rey Vega, Pablo Piantanida, and Cecilia G. Galarza. Analysis of a cooperative strategy for a large decentralized wireless network. IEEE/ACM Transactions on Networking, 22(4):1039– 1051, August 2014. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Albuquerque:2004:NBP

[AVS04] Célio Albuquerque, Brett J. Vickers, and Tatsuya Suda. Network border patrol: preventing congestion collapse and promoting fairness in the Internet. IEEE/ ACM Transactions on Networking, 12(1):173–186, February 2004. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Arlitt:1997:IWS

[AW97] Martin F. Arlitt and Carey L. Williamson. Internet Web servers: workload characterization and performance implications. IEEE/ACM Transactions on Networking, 5(5):631– 645, October 1997. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL http://

www.acm.org/pubs/citations/
journals/ton/1997-5-5/
p631-arlitt/.

Akyildiz:2004:PUM

[AW04] Ian F. Akyildiz and Wenye Wang. The predictive user mobility profile framework for wireless multimedia networks. IEEE/ACM Transactions on Networking, 12(6):1021–1035, December 2004. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Athanasiou:2015:OCA

[AWFT15] George Athanasiou, Pradeep Chathuranga Weeraddana, Carlo Fischione, and Leandros Tassiulas. Optimizing client association for load balancing and fairness in millimeter-wave wireless networks. IEEE/ACM Transactions on Networking, 23(3):836–850, June 2015. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

An:2022:PBB

[AWH^{+}22] Jian An, Zhenxing Wang, Xin He, Xiaolin Gui, Jindong Cheng, and Ruowei Gui. PPQC: a blockchainbased privacy-preserving quality control mechanism in crowdsensing applications. IEEE/ACM Transactions on Networking, 30 (3):1352–1367, June 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1109/TNET.2022.3141582.

Alresaini:2016:BDE

[AWKN16] Majed Alresaini, Kwame-Lante Wright, Bhaskar Krishnamachari, and Michael J. Neely. Backpressure delay enhancement for encounter-based mobile networks while sustaining throughput optimality. IEEE/ACM Transactions on Networking, 24(2):1196– 1208, April 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Arisdakessian:2020:FIM

[AWM^{+}20] Sarhad Arisdakessian, Omar Abdel Wahab, Azzam Mourad, Hadi Otrok, and Nadjia Kara. FoGMatch: an intelligent multi-criteria IoT-Fog scheduling approach using game theory. IEEE/ ACM Transactions on Networking, 28(4):1779–1789, August 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2020.2994015.

Albanna:2020:CML

[AY20] Amr Albanna and Homayoun Yousefi’Zadeh. Congestion minimization of LTE networks: a deep learning approach. IEEE/ ACM Transactions on Networking, 28(1):347–359, February 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL https://

dl.acm.org/doi/abs/10.
1109/TNET.2019.2960266.

An:2021:IUR

[AYL21] Zhenlin An, Lei Yang, and Qiongzheng Lin. Identifying UHF RFIDs in range of readers with WiFi. IEEE/ACM Transactions on Networking, 29(3):1252– 1265, June 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2021.3057392.

Akhoondi:2014:LLL

[AYM14] Masoud Akhoondi, Curtis Yu, and Harsha V. Madhyastha. LASTor: a low-latency AS-aware Tor client. IEEE/ACM Transactions on Networking, 22 (6):1742–1755, December 2014. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Arslan:2013:RMS

[AYS^{+}13] Mustafa Y. Arslan, Jongwon Yoon, Karthikeyan Sundaresan, Srikanth V. Krishnamurthy, and Suman Banerjee. A resource management system for interference mitigation in enterprise OFDMA femtocells. IEEE/ACM Transactions on Networking, 21(5): 1447–1460, October 2013. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Andrews:2003:ASN

[AZ03] Matthew Andrews and Lisa Zhang. Achieving stability in networks of inputqueued switches. IEEE/ ACM Transactions on Networking, 11(5):848–857, October 2003. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Andrews:2006:SNW

[AZ06a] Matthew Andrews and Lisa Zhang. Scheduling over nonstationary wireless channels with finite rate sets. IEEE/ACM Transactions on Networking, 14(5): 1067–1077, October 2006. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Arora:2006:LLS

[AZ06b] Anish Arora and Hongwei Zhang. LSRP: local stabilization in shortest path routing. IEEE/ACM Transactions on Networking, 14(3):520–531, 2006. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Andrews:2009:CWA

[AZ09] Matthew Andrews and Lisa Zhang. Complexity of wavelength assignment in optical network optimization. IEEE/ACM Transactions on Networking, 17 (2):646–657, April 2009. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Andrews:2011:SAM

[AZ11] Matthew Andrews and Lisa Zhang. Scheduling algorithms for multicarrier wireless data systems. IEEE/ACM Transactions on Networking, 19(2):447– 455, April 2011. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Al-Zubaidy:2016:NLP

[AZLB16] Hussein Al-Zubaidy, Jörg Liebeherr, and Almut Burchard. Network-layer performance analysis of multihop fading channels. IEEE/ACM Transactions on Networking, 24(1):204– 217, February 2016. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

An:2023:OEE

[AZP^{+}23] Congkai An, Anfu Zhou, Jialiang Pei, Xi Liu, Dongzhu Xu, Liang Liu, and Huadong Ma. Octopus: Exploiting the edge intelligence for accessible 5G mobile performance enhancement. IEEE/ACM Transactions on Networking, 31(6):2454– 2469, 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2022.3224369.

Andersin:1997:SSA

[AZR97] Michael Andersin, Jens Zander, and Zvi Rosberg. Soft and safe admission control in cellular networks. IEEE/ACM Transactions on Networking, 5(2):255– 265, April 1997. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/

journals/ton/1997-5-2/
p255-andersin/.

Bhattacharjee:2020:EEM

[BAB20] Sangeeta Bhattacharjee, Tamaghna Acharya, and Uma Bhattacharya. Energyefficient multicasting in hybrid cognitive small cell networks: a crosslayer approach. IEEE/ ACM Transactions on Networking, 28(1):262–274, February 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL https://

dl.acm.org/doi/abs/10.
1109/TNET.2019.2962309.

Bando:2012:SLR

[BAC12] Masanori Bando, N. Sertac Artan, and H. Jonathan Chao. Scalable lookahead regular expression detection system for deep packet inspection. IEEE/ACM Transactions on Networking, 20(3):699–714, June 2012. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Berbecaru:2010:FSM

[BAL10] Diana Berbecaru, Luca Albertalli, and Antonio Lioy. The ForwardDiffsig scheme for multicast authentication. IEEE/ACM Transactions on Networking, 18(6): 1855–1868, December 2010. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Banerjea:1999:FRG

[Ban99] Anindo Banerjea. Fault recovery for guaranteed performance communications connections. IEEE/ ACM Transactions on Networking, 7(5):653–668, October 1999. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1999-7-5/
p653-banerjea/.

Barker:1995:AUI

[Bar95] Paul Barker. An analysis of user input to an X.500 white pages directory service. IEEE/ ACM Transactions on Networking, 3(2):112–125, April 1995. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1995-3-2/
p112-barker/.

Barcaccia:1994:PTO

[BB94] Piera Barcaccia and Maurizio A. Bonuccelli. Polynomial time optimal algorithms for time slot assignment of variable bandwidth systems. IEEE/ ACM Transactions on Networking, 2(3):247–251, June 1994. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1994-2-3/
p247-barcaccia/.

Bertossi:1995:CAH

[BB95] Alan A. Bertossi and Maurizio A. Bonuccelli. Code assignment for hidden terminal interference avoidance in multihop packet radio networks. IEEE/ ACM Transactions on Networking, 3(4):441–449, August 1995. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1995-3-4/
p441-bertossi/.

Birk:1996:EDI

[BB96] Yitzhak Birk and Noam Bloch. The effects of destructive interference and wasted transmissions on the uniform-traffic capacity of non-bus-oriented singlehop interconnections. IEEE/ ACM Transactions on Networking, 4(3):442–448, June 1996. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1996-4-3/
p442-birk/.

Bejerano:2006:MFF

[BB06] Yigal Bejerano and Randeep S. Bhatia. MiFi: a framework for fairness and QoS assurance for current IEEE 802.11 networks with multiple access points. IEEE/ACM Transactions on Networking, 14(4):849– 862, August 2006. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Bhaskar:2016:LCB

[BB16] Sonia A. Bhaskar and Sonia A. Bhaskar. Localization from connectivity: a 1-bit maximum likelihood approach. IEEE/ ACM Transactions on Networking, 24(5):2939–2953, October 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Bennett:2002:DJB

[BBC^{+}02] Jon C. R. Bennett, Kent Benson, Anna Charny, William F. Courtney, and Jean-Yves Le Boudec. Delay jitter bounds and packet scale rate guarantee for expedited forwarding. IEEE/ACM Transactions on Networking, 10 (4):529–540, August 2002. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Bolla:2014:GNP

[BBCD14] Raffaele Bolla, Roberto Bruschi, Alessandro Carrega, and Franco Davoli. Green networking with packet processing engines: Modeling and optimization. IEEE/ACM Transactions on Networking, 22(1):110– 123, February 2014. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Basta:2018:ELF

[BBD^{+}18] Arsany Basta, Andreas Blenk, Szymon Dudycz, Arne Ludwig, Stefan Schmid, Stefan Schmid, Szymon Dudycz, Andreas Blenk, Arne Ludwig, and Arsany Basta. Efficient loop-free rerouting of multiple SDN flows. IEEE/ACM Transactions on Networking, 26 (2):948–961, April 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Ben-Basat:2021:RON

[BBEF^{+}21] Ran Ben-Basat, Gil Einziger, Shir Landau Feibish, Jalil Moraney, Bilal Tayh, and Danny Raz. Routingoblivious network-wide measurements. IEEE/ACM Transactions on Networking, 29(6):2386–2398, December 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.

                                        acm.org/doi/10.1109/TNET.
                                        2021.3061737.

Barletta:2018:TAD

[BBF18] Luca Barletta, Flaminio Borgonovo, and Ilario Filippini. The throughput and access delay of slotted-aloha with exponential backoff. IEEE/ ACM Transactions on Networking, 26(1):451–464, February 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Bannister:1995:PMD

[BBFG95] Joseph Bannister, Flaminio Borgonovo, Luigi Fratta, and Mario Gerla. A performance model of deflection routing in multibuffer networks with nonuniform traffic. IEEE/ACM Transactions on Networking, 3(5):509–520, October 1995. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/
                                        journals/ton/1995-3-5/
                                        p509-bannister/.

Beheshti:2010:OPB

[BBG^{+}10] Neda Beheshti, Emily Burmeister, Yashar Ganjali, John E. Bowers, Daniel J. Blumenthal, and Nick McKeown. Optical packet buffers for backbone Internet routers. IEEE/ ACM Transactions on Networking, 18(5):1599–1609, October 2010. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Baccelli:2011:DDP

[BBG11] François Baccelli, Nicholas Bambos, and Nicolas Gast. Distributed delay-power control algorithms for bandwidth sharing in wireless networks. IEEE/ ACM Transactions on Networking, 19(5):1458–1471, October 2011. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Bremler-Barr:2018:ESR

[BBHH^{+}18] Anat Bremler-Barr, Yotam Harchol, David Hay, Yacov Hel-Or, Yotam Harchol, Anat Bremler-Barr, David Hay, and Yacov Hel-Or. Encoding short ranges in TCAM without expansion: Efficient algorithm and applications. IEEE/ACM Transactions on Networking, 26(2):835–850, April 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Bremler-Barr:2010:PPE

[BBHHR10] Anat Bremler-Barr, David Hay, Danny Hendler, and Ron M. Roth. PEDS: a parallel error detection scheme for TCAM devices. IEEE/ ACM Transactions on Networking, 18(5):1665–1675, October 2010. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Bremler-Barr:2014:CSP

[BBHK14] Anat Bremler-Barr, David Hay, and Yaron Koral. CompactDFA: Scalable pattern matching using longest prefix match solutions. IEEE/ACM Transactions on Networking, 22 (2):415–428, April 2014. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Bremler-Barr:2012:AMM

[BBK12] Anat Bremler-Barr and Yaron Koral. Accelerating multipattern matching on compressed HTTP traffic. IEEE/ACM Transactions on Networking, 20(3):970– 983, June 2012. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Borst:1995:USL

[BBL95] Sem C. Borst, Onno J. Boxma, and Hanoch Levy. The use of service limits for efficient operation of multistation singlemedium communication systems. IEEE/ACM Transactions on Networking, 3(5):602–612, October 1995. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1995-3-5/
p602-borst/.

Barrenetxea:2006:CLN

[BBLV06a] Guillermo Barrenetxea, Baltasar Beferull-Lozano, and Martin Vetterli. Correction to ‘Lattice networks: Capacity limits, optimal routing, and queueing behavior’. IEEE/ACM Transactions on Networking, 14(5):1150, October 2006. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). See [BBLV06b].

Barrenetxea:2006:LNC

[BBLV06b] Guillermo Barrenetxea, Baltasar Berefull-Lozano, and Martin Vetterli. Lattice networks: capacity limits, optimal routing, and queueing behavior. IEEE/ ACM Transactions on Networking, 14(3):492–505, 2006. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). See correction [BBLV06a].

Baiocchi:1993:EAA

[BBM93] Andrea Baiocchi and Nicola Bléfari-Melazzi. An errorcontrolled approximate analysis of a stochastic fluid flow model applied to an ATM multiplexer with heterogeneous On-Off sources. IEEE/ACM Transactions on Networking, 1(6):628–637, December 1993. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/
                                        journals/ton/1993-1-6/
                                        p628-baiocchi/.

Brosh:2010:DFT

[BBM^{+}10] Eli Brosh, Salman Abdul Baset, Vishal Misra, Dan Rubenstein, and Henning Schulzrinne. The delayfriendliness of TCP for real-time traffic. IEEE/ ACM Transactions on Networking, 18(5):1478–1491, October 2010. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Benson:2008:CAO

[BBMELH08] Karyn Benson, Benjamin Birnbaum, Esteban Molina-Estolano, and Ran Libeskind-Hadas. Competitive analysis of online traffic grooming in WDM rings. IEEE/ ACM Transactions on Networking, 16(4):984–997, August 2008. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Bejerano:2005:ACQ

[BBO^{+}05] Yigal Bejerano, Yuri Breitbart, Ariel Orda, Rajeev Rastogi, and Alexander Sprintson. Algorithms for computing QoS paths with restoration. IEEE/ACM Transactions on Networking, 13(3):648–661, June 2005. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Boubrima:2019:DWS

[BBR19] Ahmed Boubrima, Walid Bechkit, and Herve Rivano. On the deployment of wireless sensor networks for air quality mapping: Optimization models and algorithms. IEEE/ ACM Transactions on Networking, 27(4):1629–1642, August 2019. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Bhattacharyya:2022:QLA

[BBR^{+}22] Rajarshi Bhattacharyya, Archana Bura, Desik Rengarajan, Mason Rumuly, Bainan Xia, Srinivas Shakkottai, Dileep Kalathil, Ricky K. P. Mok, and Amogh Dhamdhere. QFlow: a learning approach to high QoE video streaming at the wireless edge. IEEE/ ACM Transactions on Networking, 30(1):32–46, February 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.

                                        acm.org/doi/10.1109/TNET.
                                        2021.3106675.

Brinton:2018:EOS

[BBZ^{+}18] Christopher G. Brinton, Swapna Buccapatnam, Liang Zheng, Da Cao, Andrew S. Lan, Felix M. F. Wong, Sangtae Ha, Mung Chiang, and H. Vincent Poor. On the efficiency of online social learning networks. IEEE/ACM Transactions on Networking, 26(5):2076– 2089, October 2018. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Barford:2001:CPA

[BC01a] Paul Barford and Mark Crovella. Critical path analysis of TCP transactions. IEEE/ACM Transactions on Networking, 9 (3):238–248, June 2001. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Bonuccelli:2001:SRT

[BC01b] Maurizio A. Bonuccelli and M. Claudia Clò. Scheduling of real-time messages in optical broadcast-and-select networks. IEEE/ACM Transactions on Networking, 9(5):541–552, October 2001. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Basagni:1999:MTD

[BCB99] Stefano Basagni, Imrich Chlamtac, and Danilo Bruschi. A mobilitytransparent deterministic broadcast mechanism for ad hoc networks. IEEE/ ACM Transactions on Networking, 7(6):799–807, December 1999. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/articles/
journals/ton/1999-7-6/
p799-basagni/p799-basagni.
pdf;   http://www.acm.
org/pubs/citations/journals/
ton/1999-7-6/p799-basagni/

.

Bosio:2007:RPW

[BCC07] Sandro Bosio, Antonio Capone, and Matteo Cesana. Radio planning of wireless local area networks. IEEE/ACM Transactions on Networking, 15 (6):1414–1427, December 2007. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Bai:2017:PPI

[BCC^{+}17] Wei Bai, Li Chen, Kai Chen, Dongsu Han, Chen Tian, and Hao Wang. PIAS: Practical informationagnostic flow scheduling for commodity data centers. IEEE/ACM Transactions on Networking, 25(4): 1954–1967, August 2017.

CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Bouchoucha:2019:TIU

[BCD19] Taha Bouchoucha, Chen-Nee Chuah, and Zhi Ding. Topology inference of unknown networks based on robust virtual coordinate systems. IEEE/ ACM Transactions on Networking, 27(1):405–418, February 2019. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

BenBasat:2019:RAP

[BCE^{+}19] Ran Ben Basat, Xiaoqi Chen, Gil Einziger, Roy Friedman, and Yaron Kassner. Randomized admission policy for efficient top-k, frequency, and volume estimation. IEEE/ ACM Transactions on Networking, 27(4):1432–1445, August 2019. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Basat:2020:DHH

[BCER20] Ran Ben Basat, Xiaoqi Chen, Gil Einziger, and Ori Rottenstreich. Designing heavy-hitter detection algorithms for programmable switches. IEEE/ACM Transactions on Networking, 28(3):1172– 1185, June 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2020.2982739.

Banaei:2015:ASG

[BCGC15] Armin Banaei, Daren B. H.

Cline, Costas N. Georghiades, and Shuguang Cui. On asymptotic statistics for geometric routing schemes in wireless ad hoc networks. IEEE/ACM Transactions on Networking, 23 (2):559–573, April 2015. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Boggia:2007:FBC

[BCGM07] Gennaro Boggia, Pietro Camarda, Luigi Alfredo Grieco, and Saverio Mascolo. Feedback-based control for providing real-time services with the 802.11e MAC. IEEE/ACM Transactions on Networking, 15 (2):323–333, April 2007. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Bahl:2009:OUC

[BCL^{+}09] Paramvir Bahl, Ranveer Chandra, Patrick P. C. Lee, Vishal Misra, Jitendra Padhye, Dan Rubenstein, and Yan Yu. Opportunistic use of client repeaters to improve performance of WLANs. IEEE/ ACM Transactions on Networking, 17(4):1160–1171, August 2009. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Banerjee:2010:DFE

[BCL10] Nilanjan Banerjee, Mark D. Corner, and Brian Neil Levine. Design and field experimentation of an energyefficient architecture for DTN throwboxes. IEEE/ ACM Transactions on Networking, 18(2):554–567, April 2010. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Basile:2012:NLA

[BCL12] Cataldo Basile, Alberto Cappadonia, and Antonio Lioy. Network-level access control policy analysis and transformation. IEEE/ACM Transactions on Networking, 20(4):985– 998, August 2012. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Bartolini:2017:CSR

[BCLS17] Novella Bartolini, Stefano Ciavarella, Thomas F. La Porta, and Simone Silvestri. On critical service recovery after massive network failures. IEEE/ ACM Transactions on Networking, 25(4):2235–2249, August 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Byers:2004:ICD

[BCMR04] John W. Byers, Jeffrey Considine, Michael Mitzenmacher, and Stanislav Rost. Informed content delivery across adaptive overlay networks. IEEE/ACM Transactions on Networking, 12(5):767–780, October 2004. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Bejerano:2002:EHR

[BCN02] Yigal Bejerano, Israel Cidon, and Joseph (Seffi) Naor. Efficient handoff rerouting algorithms: a competitive on-line algorithmic approach. IEEE/ ACM Transactions on Networking, 10(6):749–760, December 2002. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Ba:2017:DSB

[BCO17] Seydou Ba, Bijoy Chand Chatterjee, and Eiji Oki. Defragmentation scheme based on exchanging primary and backup paths in 1 + 1 path protected elastic optical networks. IEEE/ACM Transactions on Networking, 25(3):1717– 1731, June 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Bambos:2000:CAA

[BCP00] Nicholas Bambos, Shou C. Chen, and Gregory J. Pottie. Channel access algorithms with active link protection for wireless communication networks with power control. IEEE/ACM Transactions on Networking, 8 (5):583–597, 2000. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/
                                        journals/ton/2000-8-5/
                                        p583-bambos/.

Baccarelli:2013:OSA

[BCP13] Enzo Baccarelli, Nicola Cordeschi, and Valentina Polli. Optimal self-adaptive QoS resource management in interference-affected multicast wireless networks. IEEE/ACM Transactions on Networking, 21(6):1750– 1759, December 2013. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Birand:2012:APG

[BCR^{+}12] Berk Birand, Maria Chudnovsky, Bernard Ries, Paul Seymour, Gil Zussman, and Yori Zwols. Analyzing the performance of greedy maximal scheduling via local pooling and graph theory. IEEE/ACM Transactions on Networking, 20(1): 163–176, February 2012. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Bakshi:2019:EEM

[BCS^{+}19] Arjun Bakshi, Lu Chen, Kannan Srinivasan, C. Emre Koksal, and Atilla Eryilmaz. EMIT: an efficient MAC paradigm for the Internet of Things. IEEE/ ACM Transactions on Networking, 27(4):1572–1583, August 2019. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Braun:1996:PEC

[BD96] Torsten Braun and Christophe Diot. Performance evaluation and cache analysis of an ILP protocol implementation. IEEE/ ACM Transactions on Networking, 4(3):318–330, June 1996. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1996-4-3/
p318-braun/.

Bolla:1997:CMS

[BD97] Raffaele Bolla and Franco Davoli. Control of multirate synchronous streams in hybrid TDM access networks. IEEE/ACM Transactions on Networking, 5 (2):291–304, April 1997. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL

                                        http://www.acm.org/pubs/
                                        citations/journals/ton/
                                        1997-5-2/p291-bolla/.

Baek:2007:SEB

[BD07] Seung Jun Baek and Gustavo De Veciana. Spatial energy balancing through proactive multipath routing in wireless multihop networks. IEEE/ACM Transactions on Networking, 15(1):93–104, February 2007. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Bredin:2010:DSN

[BDHR10] Jonathan L. Bredin, Erik D. Demaine, Mohammad Taghi Hajiaghayi, and Daniela Rus. Deploying sensor networks with guaranteed fault tolerance. IEEE/ ACM Transactions on Networking, 18(1):216–228, February 2010. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Biswas:2022:BAU

[BDR22] Nilanjan Biswas, Goutam Das, and Priyadip Ray. Buffer-aware user selection and resource allocation for an opportunistic cognitive radio network: a cross-layer approach. IEEE/ACM Transactions on Networking, 30(5):1940– 1954, 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2022.3159819.

Beimel:2007:ROE

[BDS07] Amos Beimel, Shlomi Dolev, and Noam Singer. RT oblivious erasure correcting. IEEE/ACM Transactions on Networking, 15(6):1321–1332, December 2007. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Brauckhoff:2012:AEB

[BDWS12] Daniela Brauckhoff, Xenofontas Dimitropoulos, Arno Wagner, and Kavé Salamatian. Anomaly extraction in backbone networks using association rules. IEEE/ ACM Transactions on Networking, 20(6):1788–1799, December 2012. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Borbash:2006:FMW

[BE06] Steven A. Borbash and Anthony Ephremides. The feasibility of matchings in a wireless network. IEEE/ ACM Transactions on Networking, 14(SI):2749–2755, June 2006. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Bader:2008:POI

[BE08] Ahmed Bader and Eylem Ekici. Performance optimization of interferencelimited multihop networks. IEEE/ACM Transactions on Networking, 16(5):1147– 1160, October 2008. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Bejerano:2004:EIM

[Bej04]  Yigal Bejerano.  Effi-

cient integration of multihop wireless and wired networks with QoS constraints. IEEE/ACM Transactions on Networking, 12(6):1064–1078, December 2004. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Bejerano:2009:TSC

[Bej09] Yigal Bejerano. Taking the skeletons out of the closets: a simple and efficient topology discovery scheme for large Ethernet LANs. IEEE/ACM Transactions on Networking, 17(5):1385– 1398, October 2009. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

BenBasat:2022:MMS

[BEK^{+}22] Ran Ben Basat, Gil Einziger, Isaac Keslassy, Ariel Orda, Shay Vargaftik, and Erez Waisbard. Memento: Making sliding windows efficient for heavy hitters. IEEE/ACM Transactions on Networking, 30(4): 1440–1453, 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2021.3132385.

Berthaud:2000:TSN

[Ber00] Jean-Marc Berthaud. Time synchronization over networks using convex closures. IEEE/ACM Transactions on Networking, 8 (2):265–277, 2000. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/2000-8-2/
p265-berthaud/.

Bankhamer:2022:LFR

[BES22] Gregor Bankhamer, Robert Elsässer, and Stefan Schmid. Local fast rerouting with low congestion: a randomized approach. IEEE/ ACM Transactions on Networking, 30(6):2403– 2418, 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2022.3174731.

Bui:2008:ACC

[BESW08] Loc Bui, Atilla Eryilmaz, R. Srikant, and Xinzhou Wu. Asynchronous congestion control in multihop wireless networks with maximal matching-based scheduling. IEEE/ACM Transactions on Networking, 16(4):826–839, August 2008. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Beard:2001:PRA

[BF01] Cory C. Beard and Victor S. Frost. Prioritized resource allocation for stressed networks. IEEE/ ACM Transactions on Networking, 9(5):618–633, October 2001. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Bonaventure:2007:ASM

[BFF07] Olivier Bonaventure, Clarence Filsfils, and Pierre François. Achieving sub-50 milliseconds recovery upon BGP peering link failures. IEEE/ACM Transactions on Networking, 15(5):1123– 1135, October 2007. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Bienkowski:2014:WAV

[BFG^{+}14] Marcin Bienkowski, Anja Feldmann, Johannes Grassler, Gregor Schaffrath, and Stefan Schmid. The widearea virtual service migration problem: a competitive analysis approach. IEEE/ACM Transactions on Networking, 22(1):165– 178, February 2014. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Blasius:2018:EES

[BFK^{+}18] Thomas Blasius, Tobias Friedrich, Anton Krohmer, Soren Laue, Anton Krohmer, Soren Laue, Tobias Friedrich, and Thomas Blasius. Efficient embedding of scalefree graphs in the hyperbolic plane. IEEE/ACM Transactions on Networking, 26(2):920–933, April 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Banerjea:1996:TRP

[BFM^{+}96] Anindo Banerjea, Domenico Ferrari, Bruce A. Mah, Mark Moran, Dinesh C. Verma, and Hui Zhang. The Tenet real-time protocol suite: design, implementation, and experiences. IEEE/ACM Transactions on Networking, 4(1):1–10, February 1996. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/
                                        journals/ton/1996-4-1/
                                        p1-banerjea/.

Bhatnagar:2001:OSF

[BFMF01] Shalabh Bhatnagar, Michael C. Fu, Steven I. Marcus, and Pedram J. Fard. Optimal structured feedback policies for ABR flow control using two-timescale SPSA. IEEE/ACM Transactions on Networking, 9(4):479– 491, August 2001. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Behrouzi-Far:2021:ERF

[BFS21] Amir Behrouzi-Far and Emina Soljanin. Efficient replication for fast and predictable performance in distributed computing. IEEE/ACM Transactions on Networking, 29(4):1467–1476, August 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2021.3062215.

Bjorkman:1998:PMM

[BG98] Mats Björkman and Per Gunningberg. Performance modeling of multiprocessor implementations of protocols. IEEE/ ACM Transactions on Networking, 6(3):262–273, June 1998. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1998-6-3/
p262-bjorkman/.

Bird:1995:KFL

[BGH^{+}95] Ray Bird, Inder Gopal, Amir Herzberg, Phil Janson, Shay Kutten, Refik Molva, and Moti Yung. The KryptoKnight family of light-weight protocols for authentication and key distribution. IEEE/ ACM Transactions on Networking, 3(1):31–41, February 1995. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1995-3-1/
p31-bird/.

Bateni:2010:MVO

[BGHS10] MohammadHossein Bateni, Alexandre Gerber, MohammadTaghi Hajiaghayi, and Subhabrata Sen. Multi-VPN optimization for scalable routing via relaying. IEEE/ACM Transactions on Networking, 18(5):1544– 1556, October 2010. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Breitbart:2004:TDH

[BGJ^{+}04] Yuri Breitbart, Minos Garofalakis, Ben Jai, Cliff Martin, Rajeev Rastogi, and Avi Silberschatz. Topology discovery in heterogeneous IP networks: the NetInventory system. IEEE/ACM Transactions on Networking, 12(3):401– 414, June 2004. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Bhattacharya:1997:DCA

[BGK97] Partha P. Bhattacharya, Leonidas Georgiadis, and Arvind Krishna. Distributed channel allocation for PCN with variable rate traffic. IEEE/ ACM Transactions on Networking, 5(6):907–923, December 1997. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/
                                        journals/ton/1997-5-6/
                                        p907-bhattacharya/.

Bienkowski:2016:DAD

[BGK^{+}16] Marcin Bienkowski, Leszek Gasieniec, Marek Klonowski, Miroslaw Korzeniowski, Bernard Mans, Stefan Schmid, and Roger Wattenhofer. Distributed alarming in the on-duty and off-duty models. IEEE/ ACM Transactions on Networking, 24(1):218–230, February 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Brenes:2020:PPP

[BGMB^{+}20] Juan Brenes, Alberto García-Martínez, Marcelo Bagnulo, Andra Lutu, and Cristel Pelsser. Power prefixes prioritization for smarter BGP reconvergence. IEEE/ACM Transactions on Networking, 28 (3):1074–1087, June 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL

https://dl.acm.org/doi/
10.1109/TNET.2020.2979665.

Boyd:2006:RGA

[BGPS06] Stephen Boyd, Arpita Ghosh, Balaji Prabhakar, and Devavrat Shah. Randomized gossip algorithms. IEEE/ACM Transactions on Networking, 14(SI): 2508–2530, June 2006. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Banchs:2013:GTA

[BGSSW13] Albert Banchs, Andres Garcia-Saavedra, Pablo Serrano, and Joerg Widmer. A game-theoretic approach to distributed opportunistic scheduling. IEEE/ACM Transactions on Networking, 21(5):1553– 1566, October 2013. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Barcelo:2000:WCT

[BGVC00] José M. Barceló, Jorge García-Vidal, and Olga Casals. Worst-case traffic in a tree network of ATM multiplexers. IEEE/ACM Transactions on Networking, 8(4):507–516, 2000. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL

                                        http://www.acm.org/pubs/
                                        citations/journals/ton/
                                        2000-8-4/p507-barcelo/.

Baccelli:2005:ITF

[BH05] François Baccelli and Dohy Hong. Interaction of TCP flows as billiards. IEEE/ ACM Transactions on Networking, 13(4):841–853, August 2005. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Boucouvalas:2006:OIP

[BH06] Anthony C. Boucouvalas and Pi Huang. OBEX over IrDA: performance analysis and optimization by considering multiple applications. IEEE/ACM Transactions on Networking, 14 (6):1292–1301, December 2006. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Bartolini:2020:FBF

[BHA^{+}20] Novella Bartolini, Ting He, Viviana Arrigoni, Annalisa Massini, Federico Trombetti, and Hana Khamfroush. On fundamental bounds on failure identifiability by Boolean network tomography. IEEE/ ACM Transactions on Networking, 28(2):588–601, April 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL https://

dl.acm.org/doi/abs/10.
1109/TNET.2020.2969523.

Bai:2021:OMC

[BHC^{+}21] Wei Bai, Shuihai Hu, Kai Chen, Kun Tan, and Yongqiang Xiong. One more config is enough: Saving (DC)TCP for highspeed extremely shallowbuffered datacenters. IEEE/ ACM Transactions on Networking, 29(2):489–502, April 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2020.3032999.

Bohacek:2006:NTP

[BHL^{+}06] Stephan Bohacek, Joao P.

Hespanha, Junsoo Lee, Chansook Lim, and Katia Obraczka. A new TCP for persistent packet reordering. IEEE/ACM Transactions on Networking, 14 (2):369–382, April 2006. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Bejerano:2007:FLB

[BHL07] Yigal Bejerano, Seung-Jae Han, and Li Li. Fairness and load balancing in wireless LANs using association control. IEEE/ACM Transactions on Networking, 15(3):560–573, June 2007. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Bianco:2011:CPS

[BHN11] Andrea Bianco, David Hay, and Fabio Neri. Crosstalkpreventing scheduling in single-and two-stage AWGbased cell switches. IEEE/ ACM Transactions on Networking, 19(1):142–155, February 2011. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Biswas:2000:DFB

[BI00] Subir K. Biswas and Rauf Izmailov. Design of a fair bandwidth allocation policy for VBR traffic in ATM networks. IEEE/ACM Transactions on Networking, 8 (2):212, 2000. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/

journals/ton/2000-8-2/
p212-biswas/.

Biswas:2000:CSE

[BIS00] Subir K. Biswas, Rauf Izmailov, and Bhaskar Sengupta. Connection splitting: an efficient way of reducing call blocking in ATM. IEEE/ACM Transactions on Networking, 8 (5):655–666, 2000. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/2000-8-5/
p655-biswas/.

Banerjee:2001:PEQ

[BIV01] Ayan Banerjee, Ronald A.

Iltis, and Emmanouel A. Varvarigos. Performance evaluation for a quasisynchronous packet radio network (QSPNET). IEEE/ACM Transactions on Networking, 9(5):567– 577, October 2001. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Barghi:2015:EAA

[BJ15] Sanaz Barghi and Hamid Jafarkhani. Exploiting asynchronous amplify-and-forward relays to enhance the performance of IEEE 802.11 networks. IEEE/ ACM Transactions on Networking, 23(2):479–490, April 2015. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Banirazi:2020:HDP

[BJK20] Reza Banirazi, Edmond Jonckheere, and Bhaskar Krishnamachari. Heatdiffusion: Pareto optimal dynamic routing for time-varying wireless networks. IEEE/ACM Transactions on Networking, 28(4):1520–1533, August 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2020.2991745.

Barghi:2011:MAM

[BJY11] Sanaz Barghi, Hamid Jafarkhani, and Homayoun Yousefi’zadeh. MIMOassisted MPR-aware MAC design for asynchronous WLANs. IEEE/ACM Transactions on Networking, 19(6):1652–1665, December 2011. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Berger:2000:DBE

[BK00] Arthur W. Berger and Yaakov Kogan. Dimensioning bandwidth for elastic traffic in high-speed data networks. IEEE/ACM Transactions on Networking, 8(5):643–654, 2000.

CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL

http://www.acm.org/pubs/
citations/journals/ton/
2000-8-5/p643-berger/.

Birk:2006:CDI

[BK06] Yitzhak Birk and Tomer Kol. Coding on demand by an informed source (IS-COD) for efficient broadcast of different supplemental data to caching clients. IEEE/ACM Transactions on Networking, 14 (SI):2825–2830, June 2006. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Bhattacharya:2017:AMI

[BK17] Abhijit Bhattacharya and Anurag Kumar. Analytical modeling of IEEE 802.11type CSMA/CA networks with short term unfairness. IEEE/ACM Transactions on Networking, 25(6): 3455–3472, December 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Bagheri:1993:SBM

[BKH^{+}93] Mehran Bagheri, Dennis T.

Kong, Wayne S. Holden, Fernando C. Irizarry, and Derek D. Mahoney. An STS-N byte-interleaving multiplexer/scrambler and demultiplexer/descrambler architecture and its experimental OC-48 implementation. IEEE/ ACM Transactions on Networking, 1(3):282–285, June 1993. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/
                                        journals/ton/1993-1-3/
                                        p282-bagheri/.

Byers:2006:FGL

[BKLM06] John W. Byers, Gu-In Kwon, Michael Luby, and Michael Mitzenmacher. Fine-grained layered multicast with STAIR. IEEE/ ACM Transactions on Networking, 14(1):81–93, February 2006. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Bhatia:2008:BGR

[BKLS08] Randeep S. Bhatia, Murali Kodialam, T. V. Lakshman, and Sudipta Sengupta. Bandwidth guaranteed routing with fast restoration against link and node failures. IEEE/ACM Transactions on Networking, 16(6):1321–1330, December 2008. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Bhattacharyya:2003:ERC

[BKTN03] Supratik Bhattacharyya, James F. Kurose, Don Towsley, and Ramesh Nagarajan. Efficient ratecontrolled bulk data transfer using multiple multicast groups. IEEE/ACM Transactions on Networking, 11(6):895–907, December 2003. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Byun:1994:DAA

[BL94] Jae W. Byun and Tony T.

Lee. The design and analysis of an ATM multicast switch with adaptive traffic controller. IEEE/ ACM Transactions on Networking, 2(3):288–298, June 1994. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1994-2-3/
p288-byun/.

Bouillet:2004:DCS

[BL04] Eric Bouillet and Jean-François Labourdette. Distributed computation of shared backup path in mesh optical networks using probabilistic methods. IEEE/ACM Transactions on Networking, 12(5):920– 930, October 2004. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Basile:2015:AAL

[BL15] Cataldo Basile and Antonio Lioy. Analysis of application-layer filtering policies with application to HTTP. IEEE/ACM Transactions on Networking, 23 (1):28–41, February 2015. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Bouabdallah:2010:DAM

[BLB10] Nizar Bouabdallah, Rami Langar, and Raouf Boutaba. Design and analysis of mobility-aware clustering algorithms for wireless mesh networks. IEEE/ ACM Transactions on Networking, 18(6):1677–1690, December 2010. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Banerjee:2006:RMU

[BLBS06] Suman Banerjee, Seungjoon Lee, Bobby Bhattacharjee, and Aravind Srinivasan. Resilient multicast using overlays. IEEE/ACM Transactions on Networking, 14(2):237–248, April 2006. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Burchard:2011:SSN

[BLC11] Almut Burchard, Jörg Liebeherr, and Florin Ciucu. On superlinear scaling of network delays. IEEE/ACM Transactions on Networking, 19(4):1043– 1056, August 2011. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Bando:2012:FBG

[BLC12] Masanori Bando, Yi-Li Lin, and H. Jonathan Chao. FlashTrie: beyond 100-Gb/s IP route lookup using hash-based prefixcompressed trie. IEEE/ ACM Transactions on Networking, 20(4):1262–1275, August 2012. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Bae:2021:LSN

[BLC21] Jeongmin Bae, Joohyun Lee, and Song Chong. Learning to schedule network resources throughput and delay optimally using Q^{+}-learning. IEEE/ ACM Transactions on Networking, 29(2):750–763, April 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2021.3051663.

Bensaou:1997:ECL

[BLCT97] Brahim Bensaou, Shirley T. C. Lam, Hon-Wai Chu, and Danny H. K. Tsang. Estimation of the cell loss ratio in ATM networks with a fuzzy system and application to measurement-based call admission control. IEEE/ ACM Transactions on Networking, 5(4):572–584, August 1997. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1997-5-4/
p572-bensaou/.

Borrel:2009:SUS

[BLDF09] Vincent Borrel, Franck Legendre, Marcelo Dias De Amorim, and Serge Fdida. SIMPS: using sociology for personal mobility. IEEE/ACM Transactions on Networking, 17(3):831– 842, June 2009. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Buchbinder:2012:DPA

[BLEM^{+}12] Niv Buchbinder, Liane Lewin-Eytan, Ishai Menache, Joseph Naor, and Ariel Orda. Dynamic power allocation under arbitrary varying channels: an online approach. IEEE/ACM Transactions on Networking, 20(2):477–487, April 2012. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Baughman:2007:CPP

[BLL07] Nathaniel E. Baughman, Marc Liberatore, and Brian Neil Levine. Cheatproof playout for centralized and peer-to-peer gaming. IEEE/ACM Transactions on Networking, 15 (1):1–13, February 2007. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Beyranvand:2017:TFE

[BLM^{+}17] Hamzeh Beyranvand, Martin Levesque, Martin Maier, Jawad A. Salehi, Christos Verikoukis, and David Tipper. Toward 5G: FiWi enhanced LTE-A HetNets with reliable low-latency fiber backhaul sharing and WiFi offloading. IEEE/ ACM Transactions on Networking, 25(2):690–707, April 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Barbera:2010:QSA

[BLPS10] Mario Barbera, Alfio Lombardo, Carla Panarello, and Giovanni Schembra. Queue stability analysis and performance evaluation of a TCP-compliant window management mechanism. IEEE/ACM Transactions on Networking, 18(4):1275– 1288, August 2010. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Bouillet:2005:LRO

[BLRC05] Eric Bouillet, Jean-François Labourdette, Ramu Ramamurthy, and Sid Chaudhuri. Lightpath reoptimization in mesh optical networks. IEEE/ACM Transactions on Networking, 13(2):437–447, April 2005. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Brosh:2007:AHA

[BLS07] Eli Brosh, Asaf Levin, and Yuval Shavitt. Approximation and heuristic algorithms for minimum-delay application-layer multicast trees. IEEE/ACM Transactions on Networking, 15 (2):473–484, April 2007. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Blanchini:2002:RRC

[BLT02] Franco Blanchini, Renato Lo Cigno, and Roberto Tempo. Robust rate control for integrated services packet networks. IEEE/ ACM Transactions on Networking, 10(5):644–652, October 2002. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Balasubramanian:2010:RRD

[BLV10] Aruna Balasubramanian, Brian Neil Levine, and Arun Venkataramani. Replication routing in DTNs: a resource allocation approach. IEEE/ACM Transactions on Networking, 18 (2):596–609, April 2010. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Benmohamed:1993:FCC

[BM93] Lotfi Benmohamed and Semyon M. Meerkov. Feedback control of congestion in packet switching networks: the case of a single congested node. IEEE/ ACM Transactions on Networking, 1(6):693–708, December 1993. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1993-1-6/
p693-benmohamed/.

Bianchi:1997:RLS

[BM97] Giuseppe Bianchi and Riccardo Melen. The role of local storage in supporting video retrieval services on ATM networks. IEEE/ ACM Transactions on Networking, 5(6):882–892, December 1997. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/

                                        journals/ton/1997-5-6/
                                        p882-bianchi/.

Banerjee:2000:WRO

[BM00] Dhritiman Banerjee and Biswanath Mukherjee. Wavelengthrouted optical networks: linear formulation, resource budgeting tradeoffs, and a reconfiguration study. IEEE/ACM Transactions on Networking, 8 (5):598–607, 2000. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/
                                        journals/ton/2000-8-5/
                                        p598-banerjee/.

Brzezinski:2008:ATR

[BM08] Andrew Brzezinski and Eytan Modiano. Achieving 100% throughput in reconfigurable optical networks. IEEE/ACM Transactions on Networking, 16 (4):970–983, August 2008. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Baldi:2009:PFP

[BM09] Mario Baldi and Guido Marchetto. Pipeline forwarding of packets based on a low-accuracy networkdistributed common time reference. IEEE/ACM Transactions on Networking, 17(6):1936–1949, December 2009. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Boroujeny:2022:DST

[BM22] Massieh Kordi Boroujeny and Brian L. Mark. Design of a stochastic traffic regulator for end-to-end network delay guarantees. IEEE/ACM Transactions on Networking, 30(6): 2531–2543, 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2022.3181858.

Balakrishnan:2011:MTE

[BMB^{+}11] Mahesh Balakrishnan, Tudor Marian, Kenneth P. Birman, Hakim Weatherspoon, and Lakshmi Ganesh. Maelstrom: transparent error correction for communication between data centers. IEEE/ACM Transactions on Networking, 19(3):617–629, June 2011. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Baldesi:2019:SGF

[BMB19] Luca Baldesi, Athina Markopoulou, and Carter T. Buttsc. Spectral graph forge: a framework for generating synthetic graphs with a target modularity. IEEE/ ACM Transactions on Networking, 27(5):2125–2136, October 2019. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Barrachina-Munoz:2021:WFC

[BMBK21] Sergio Barrachina-Muñoz, Boris Bellalta, and Edward W. Knightly. Wi-Fi channel bonding: an all-channel system and experimental study from urban hotspots to a soldout stadium. IEEE/ACM Transactions on Networking, 29(5):2101–2114, October 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2021.3077770.

Bianco:2009:WUS

[BMM^{+}09] Andrea Bianco, Gianluca Mardente, Marco Mellia, Maurizio Munafò, and Luca Muscariello. Web user-session inference by means of clustering techniques. IEEE/ACM Transactions on Networking, 17 (2):405–416, April 2009. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Borst:2014:NUM

[BMS14a] Sem C. Borst, Mihalis G. Markakis, and Iraj Saniee. Nonconcave utility maximization in locally coupled systems, with applications to wireless and wireline networks. IEEE/ACM Transactions on Networking, 22 (2):674–687, April 2014. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Butkiewicz:2014:CWP

[BMS14b] Michael Butkiewicz, Harsha V. Madhyastha, and Vyas Sekar. Characterizing Web page complexity and its impact. IEEE/ACM Transactions on Networking, 22(3):943–956, June 2014. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Baccelli:2009:RPN

[BMVB09] Françis Baccelli, Sridhar Machiraju, Darryl Veitch, and Jean Bolot. The role of PASTA in network measurement. IEEE/ ACM Transactions on Networking, 17(4):1340–1353, August 2009. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Borst:2003:GPS

[BMvU03] Sem Borst, Michel Mandjes, and Miranda van Uitert. Generalized processor sharing with lighttailed and heavy-tailed input. IEEE/ACM Transactions on Networking, 11 (5):821–834, October 2003. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Brown:2017:MCC

[BMY^{+}17] Michael Brown, Colin Marshall, Dejun Yang, Ming Li, Jian Lin, Guoliang Xue, Michael Brown, Colin Marshall, Dejun Yang, Ming Li, Jian Lin, and Guoliang Xue. Maximizing capacity in cognitive radio networks under physical interference model. IEEE/ ACM Transactions on Networking, 25(5):3003–3015, October 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Basu:2005:FIU

[BN05] Anindya Basu and Girija Narlikar. Fast incremental updates for pipelined forwarding engines. IEEE/ ACM Transactions on Networking, 13(3):690–703, June 2005. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Bhorkar:2016:ORC

[BNJ16] Abhijeet Bhorkar, Mohammad Naghshvar, and Tara Javidi. Opportunistic routing with congestion diversity in wireless ad hoc networks. IEEE/ACM Transactions on Networking, 24 (2):1167–1180, April 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Bhorkar:2012:AOR

[BNJR12] Abhijeet A. Bhorkar, Mohammad Naghshvar, Tara Javidi, and Bhaskar D. Rao. Adaptive opportunistic routing for wireless ad hoc networks. IEEE/ ACM Transactions on Networking, 20(1):243–256, February 2012. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Boche:2011:PBU

[BNS11] Holger Boche, Siddharth Naik, and Martin Schubert. Pareto boundary of utility sets for multiuser wireless systems. IEEE/ACM Transactions on Networking, 19(2):589–601, April 2011. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Baldi:2000:EED

[BO00] Mario Baldi and Yoram Ofek. End-to-end delay analysis of videoconferencing over packet-switched networks. IEEE/ACM Transactions on Networking, 8(4):479–492, 2000. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL

http://www.acm.org/pubs/
citations/journals/ton/
2000-8-4/p479-baldi/.

Baldi:2003:CRT

[BO03] Mario Baldi and Yoram Ofek. A comparison of ring and tree embedding for real-time group multicast. IEEE/ACM Transactions on Networking, 11(3): 451–464, June 2003. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Banner:2007:MRA

[BO07a] Ron Banner and Ariel Orda. Multipath routing algorithms for congestion minimization. IEEE/ACM Transactions on Networking, 15(2):413–424, April 2007. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Banner:2007:PTN

[BO07b] Ron Banner and Ariel Orda. The power of tuning: a novel approach for the efficient design of survivable networks. IEEE/ACM Transactions on Networking, 15(4):737–749, August 2007. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Blocq:2016:HGB

[BO16] Gideon Blocq and Ariel Orda. How good is bargained routing? IEEE/ ACM Transactions on Networking, 24(6):3493–3507, December 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Banchs:2016:TSB

[BOGS^{+}16] Albert Banchs, Jorge Ortin, Andres Garcia-Saavedra, Douglas J. Leith, and Pablo Serrano. Thwarting selfish behavior in 802.11 WLANs. IEEE/ ACM Transactions on Networking, 24(1):492–505, February 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Borst:2005:ULP

[Bor05] Sem Borst. User-level performance of channel-aware scheduling algorithms in wireless data networks. IEEE/ACM Transactions on Networking, 13(3):636– 647, June 2005. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Baldi:2000:AGM

[BOY00] Mario Baldi, Yoram Ofek, and Bülent Yener. Adaptive group multicast with time-driven priority. IEEE/ ACM Transactions on Networking, 8(1):31–43, February 2000. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/

journals/ton/2000-8-1/
p31-baldi/.

Bononi:1996:AEI

[BP96] Alberto Bononi and Paul R. Prucnal. Analytical evaluation of improved access techniques in deflection routing networks. IEEE/ ACM Transactions on Networking, 4(5):726–730, October 1996. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1996-4-5/
p726-bononi/.

Borkotoky:2019:FCB

[BP19] Siddhartha S. Borkotoky and Michael B. Pursley. Fountain-coded broadcast distribution in multiplehop packet radio networks. IEEE/ACM Transactions on Networking, 27(1):29– 41, February 2019. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Badita:2020:OSS

[BPA20] Ajay Badita, Parimal Parag, and Vaneet Aggarwal. Optimal server selection for straggler mitigation. IEEE/ACM Transactions on Networking, 28(2): 709–721, April 2020. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL https:

//dl.acm.org/doi/abs/
10.1109/TNET.2020.2973224.

Badita:2021:SFC

[BPA21] Ajay Badita, Parimal Parag, and Vaneet Aggarwal. Single-forking of coded subtasks for straggler mitigation. IEEE/ ACM Transactions on Networking, 29(6):2413–2424, December 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2021.3075377.

Broustis:2010:MDG

[BPK^{+}10] Ioannis Broustis, Konstantina Papagiannaki, Srikanth V. Krishnamurthy, Michalis Faloutsos, and Vivek P. Mhatre. Measurementdriven guidelines for 802.11 WLAN design. IEEE/ ACM Transactions on Networking, 18(3):722–735, June 2010. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Bhattarai:2020:DEZ

[BPL20] Sudeep Bhattarai, Jung-Min Park, and William Lehr. Dynamic exclusion zones for protecting primary users in database-driven spectrum sharing. IEEE/ACM Transactions on Networking, 28(4):1506–1519, August 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2020.2986410.

Bampas:2012:NMW

[BPPP12] Evangelos Bampas, Aris Pagourtzis, George Pierrakos, and Katerina Potika. On a noncooperative model for wavelength assignment in multifiber optical networks. IEEE/ACM Transactions on Networking, 20(4):1125–1137, August 2012. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Bennett:1999:PRP

[BPS99] Jon C. R. Bennett, Craig Partridge, and Nicholas Shectman. Packet reordering is not pathological network behavior. IEEE/ ACM Transactions on Networking, 7(6):789–798, December 1999. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/articles/
                                        journals/ton/1999-7-6/
                                        p789-bennett/p789-bennett.
                                        pdf;   http://www.acm.
                                        org/pubs/citations/journals/
                                        ton/1999-7-6/p789-bennett/

.

Balakrishnan:1997:CMI

[BPSK97] Hari Balakrishnan, Venkata N. Padmanabhan, Srinivasan Seshan, and Randy H.

Katz. A comparison of mechanisms for improving TCP performance over wireless links. IEEE/ ACM Transactions on Networking, 5(6):756–769, December 1997. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1997-5-6/
p756-balakrishnan/.

Borokhovich:2018:LOL

[BPST18] Michael Borokhovich, Yvonne-Anne Pignolet, Stefan Schmid, and Gilles Tredan. Load-optimal local fast rerouting for dense networks. IEEE/ACM Transactions on Networking, 26 (6):2583–2597, December 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Berrocal-Plaza:2016:EWA

[BPVRSP16] Victor Berrocal-Plaza, Miguel A.

Vega-Rodriguez, and Juan M. Sanchez-Perez. An efficient way of assigning paging areas by using mobility models. IEEE/ACM Transactions on Networking, 24(6): 3726–3739, December 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Bao:2023:DLB

[BPW23] Yixin Bao, Yanghua Peng, and Chuan Wu. Deep learning-based job placement in distributed machine learning clusters with heterogeneous workloads. IEEE/ACM Transactions on Networking, 31(2):634– 647, April 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2022.3202529.

Bustamante:2008:DLS

[BQ08] Fabián E. Bustamante and Yi Qiao. Designing lessstructured P2P systems for the expected high churn. IEEE/ACM Transactions on Networking, 16(3):617– 627, June 2008. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Bejerano:2006:RML

[BR06] Yigal Bejerano and Rajeev Rastogi. Robust monitoring of link delays and faults in IP networks. IEEE/ ACM Transactions on Networking, 14(5):1092–1103, October 2006. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Barlet-Ros:2011:PRM

[BRISCSP11] Pere Barlet-Ros, Gianluca Iannaccone, Josep Sanjuàs-Cuxart, and Josep Solé-Pareta. Predictive resource management of multiple monitoring applications. IEEE/ACM Transactions on Networking, 19 (3):788–801, June 2011. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Bura:2022:LCC

[BRK^{+}22] Archana Bura, Desik Rengarajan, Dileep Kalathil, Srinivas Shakkottai, and Jean-Francois Chamberland. Learning to cache and caching to learn: Regret analysis of caching algorithms. IEEE/ACM Transactions on Networking, 30(1):18–31, February 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2021.3105880.

Balan:2013:AED

[BRM^{+}13] Horia Vlad Balan, Ryan Rogalin, Antonios Michaloliakos, Konstantinos Psounis, and Giuseppe Caire. AirSync: enabling distributed multiuser MIMO with full spatial multiplexing. IEEE/ACM Transactions on Networking, 21(6): 1681–1695, December 2013. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Breitgand:2006:TMP

[BRS06] David Breitgand, Danny Raz, and Yuval Shavitt. The traveling miser problem. IEEE/ACM Transactions on Networking, 14 (4):711–724, August 2006. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Blough:2010:AAW

[BRS10] Douglas M. Blough, G. Resta, and P. Santi. Approximation algorithms for wireless link scheduling with SINRbased interference. IEEE/ ACM Transactions on Networking, 18(6):1701–1712, December 2010. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Bejerano:2019:DDM

[BRY^{+}19] Yigal Bejerano, Chandrashekhar Raman, Chun-Nam Yu, Varun Gupta, Craig Gutterman, Tomas Young, Hugo A. Infante, Yousef M. Abdelmalek, and Gil Zussman. DyMo: Dynamic monitoring of largescale LTE-multicast systems. IEEE/ACM Transactions on Networking, 27(1):258–271, February 2019. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Bellur:1997:SNA

[BS97] Bhargav R. Bellur and Galen H. Sasaki. A SAT-based network access scheme for fairness in high speed networks. IEEE/ACM Transactions on Networking, 5(3):371– 381, June 1997. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1997-5-3/
p371-bellur/.

Byun:2000:USA

[BS00] Sung Hyuk Byun and Dan Keun Sung. The UniMIN switch architecture for large-scale ATM switches. IEEE/ACM Transactions on Networking, 8(1):109–120, February 2000. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/2000-8-1/
p109-byun/.

Brassil:2002:SIM

[BS02] Jack Brassil and Henning Schulzrinne. Structuring Internet media streams with cueing protocols. IEEE/ACM Transactions on Networking, 10(4):466– 476, August 2002. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Boche:2008:SGC

[BS08] Holger Boche and Martin Schubert. A superlinearly and globally convergent algorithm for power control and resource allocation with general interference functions. IEEE/ACM Transactions on Networking, 16(2):383–395, April 2008. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Boche:2009:NBP

[BS09] Holger Boche and Martin Schubert. Nash bargaining and proportional fairness for wireless systems. IEEE/ ACM Transactions on Networking, 17(5):1453–1466, October 2009. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Bouillard:2015:EWC

[BS15] Anne Bouillard and Giovanni Stea. Exact worstcase delay in FIFO-multiplexing feed-forward networks. IEEE/ ACM Transactions on Networking, 23(5):1387–1400, October 2015. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Bajpai:2019:LVD

[BS19] Vaibhav Bajpai and Jurgen Schonwalder. A longitudinal view of dual-stacked Websites — failures, latency and happy eyeballs. IEEE/ACM Transactions on Networking, 27(2):577– 590, April 2019. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Beirami:2016:PLN

[BSF16] Ahmad Beirami, Mohsen Sardari, and Faramarz Fekri. Packet-level network compression: realization and scaling of the networkwide benefits. IEEE/ACM Transactions on Networking, 24(3):1588–1604, June 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Basu:2018:ATB

[BSG^{+}18] Soumya Basu, Aditya Sundarrajan, Javad Ghaderi, Sanjay Shakkottai, and Ramesh Sitaraman. Adaptive TTL-based caching for content delivery. IEEE/ ACM Transactions on Networking, 26(3):1063–1077, June 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Batayneh:2011:RTR

[BSH^{+}11] Marwan Batayneh, Dominic A. Schupke, Marco Hoffmann, Andreas Kirstaedter, and Biswanath Mukherjee. On routing and transmission-range determination of multi-bit-rate signals over mixed-line-rate WDM optical networks for carrier Ethernet. IEEE/ ACM Transactions on Networking, 19(5):1304–1316, October 2011. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Barbette:2021:CSC

[BSM21] Tom Barbette, Cyril Soldani, and Laurent Mathy. Combined stateful classification and session splicing for high-speed NFV service chaining. IEEE/ ACM Transactions on Networking, 29(6):2560–2573, December 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2021.3099240.

Bejerano:2006:ELA

[BSNI06] Yigal Bejerano, Mark A. Smith, Joseph (Seffi) Naor, and Nicole Immorlica. Efficient location area planning for personal communication systems. IEEE/ACM Transactions on Networking, 14(2):438–450, April 2006. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Bozinovski:2007:MAS

[BSP07] Marjan Bozinovski, Hans P. Schwefel, and Ramjee Prasad. Maximum availability server selection policy for efficient and reliable session control systems. IEEE/ACM Transactions on Networking, 15 (2):387–399, April 2007. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Boche:2021:ASC

[BSP21] Holger Boche, Rafael F.

Schaefer, and H. Vincent Poor. On the algorithmic solvability of channel dependent classification problems in communication systems. IEEE/ ACM Transactions on Networking, 29(3):1155–1168, June 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2021.3059920.

Baron:2016:OMD

[BSRdA16] Benjamin Baron, Promethee Spathis, Herve Rivano, and Marcelo Dias de Amorim. Offloading massive data onto passenger vehicles: Topology simplification and traffic assignment. IEEE/ACM Transactions on Networking, 24(6):3248– 3261, December 2016. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Bui:2009:DLS

[BSS09] Loc X. Bui, Sujay Sanghavi, and R. Srikant. Distributed link scheduling with constant overhead. IEEE/ACM Transactions on Networking, 17(5):1467– 1480, October 2009. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Budzisz:2011:FCL

[BSS^{+}11a] Lukasz Budzisz, Rade Stanojević, Arieh Schlote, Fred Baker, and Robert Shorten. On the fair coexistence of loss- and delaybased TCP. IEEE/ACM Transactions on Networking, 19(6):1811–1824, December 2011. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Bui:2011:NAR

[BSS11b] Loc X. Bui, R. Srikant, and Alexander Stolyar. A novel architecture for reduction of delay and queueing structure complexity in the back-pressure algorithm. IEEE/ACM Transactions on Networking, 19 (6):1597–1609, December 2011. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Blough:2014:FUI

[BSS14] Douglas M. Blough, Paolo Santi, and Ramya Srinivasan. On the feasibility of unilateral interference cancellation in MIMO networks. IEEE/ACM Transactions on Networking, 22 (6):1831–1844, December 2014. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Bedewy:2019:AIM

[BSS19] Ahmed M. Bedewy, Yin Sun, and Ness B. Shroff. The age of information in multihop networks. IEEE/ ACM Transactions on Networking, 27(3):1248–1257, June 2019. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Bala:1995:RLL

[BSSLB95] Krishna Bala, Thomas E.

Stern, David Simchi-Levi, and Kavita Bala. Routing in a linear lightwave network. IEEE/ ACM Transactions on Networking, 3(4):459–469, August 1995. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1995-3-4/
p459-bala/.

Begole:2001:RSR

[BSSS01] James Begole, Randall B.

Smith, Craig A. Struble, and Clifford A. Shaffer. Resource sharing for replicated synchronous groupware. IEEE/ACM Transactions on Networking, 9(6):833–843, December 2001. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Bedewy:2021:LPS

[BSSS21] Ahmed M. Bedewy, Yin Sun, Rahul Singh, and Ness B. Shroff. Low-power status updates via sleepwake scheduling. IEEE/ ACM Transactions on Networking, 29(5):2129–2141, October 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2021.3081102.

Bienkowski:2018:OAF

[BSSU18] Marcin Bienkowski, Nadi Sarrar, Stefan Schmid, and Steve Uhlig. Online aggregation of the forwarding information base: Accounting for locality and churn. IEEE/ACM Transactions on Networking, 26(1):591– 604, February 2018. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Bodas:2012:LCS

[BSYS12] Shreeshankar Bodas, Sanjay Shakkottai, Lei Ying, and R. Srikant. Lowcomplexity scheduling algorithms for multichannel downlink wireless networks. IEEE/ACM Transactions on Networking, 20(5):1608–1621, October 2012. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Bianchi:1993:IQA

[BT93] Giuseppe Bianchi and Jonathan S. Turner. Improved queueing analysis of shared buffer switching networks. IEEE/ ACM Transactions on Networking, 1(4):482–490, August 1993. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1993-1-4/
p482-bianchi/.

Baiocchi:2023:FCO

[BT23] Andrea Baiocchi and Ion Turcanu. On flow control and optimized backoff in non-saturated CSMA. IEEE/ACM Transactions on Networking, 31(5):2191– 2206, 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2023.3239410.

Bensaou:2001:CBF

[BTC01] Brahim Bensaou, Danny H. K. Tsang, and King Tung Chan. Credit-based fair queueing (CBFQ): a simple service-scheduling algorithm for packet-switched networks. IEEE/ACM Transactions on Networking, 9(5):591–604, October 2001. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Bandyopadhyay:2005:STS

[BTC05] Seema Bandyopadhyay, Qingjiang Tian, and Edward J. Coyle. Spatiotemporal sampling rates and energy efficiency in wireless sensor networks. IEEE/ACM Transactions on Networking, 13(6):1339– 1352, December 2005. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Bedogni:2017:DAV

[BTD^{+}17] Luca Bedogni, Angelo Trotta, Marco Di Felice, Yue Gao, Xingjian Zhang, Qianyun Zhang, Fabio Malabocchia, and Luciano Bononi. Dynamic adaptive video streaming on heterogeneous TVWS and Wi-Fi networks. IEEE/ ACM Transactions on Networking, 25(6):3253–3266, December 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Babarczi:2011:ALF

[BTH11] Péter Babarczi, János Tapolcai, and Pin-Han Ho. Adjacent link failure localization with monitoring trails in all-optical mesh networks. IEEE/ACM Transactions on Networking, 19(3):907–920, June 2011. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Bhowmik:2017:HPP

[BTK^{+}17] Sukanya Bhowmik, Muhammad Adnan Tariq, Boris Koldehofe, Frank Durr, Thomas Kohler, and Kurt Rothermel. High performance Publish/ subscribe middleware in softwaredefined networks. IEEE/ ACM Transactions on Networking, 25(3):1501–1516, June 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Babarczi:2017:DCT

[BTP^{+}17] Peter Babarczi, Janos Tapolcai, Alija Pasic, Lajos Ronyai, Erika R. Berczi-Kovacs, and Muriel Medard. Diversity coding in two-connected networks. IEEE/ACM Transactions on Networking, 25(4):2308– 2319, August 2017. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Bastopcu:2021:AIU

[BU21] Melih Bastopcu and Sennur Ulukus. Age of information for updates with distortion: Constant and age-dependent distortion constraints. IEEE/ACM Transactions on Networking, 29(6):2425–2438, December 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2021.3091493.

Bauer:1996:DAM

[BV96] Fred Bauer and Anujan Varma. Distributed algorithms for multicast path setup in data networks. IEEE/ACM Transactions on Networking, 4(2):181– 191, April 1996. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/
                                        journals/ton/1996-4-2/
                                        p181-bauer/.

Baboescu:2005:SPC

[BV05a] Florin Baboescu and George Varghese. Scalable packet classification. IEEE/ACM Transactions on Networking, 13(1):2–14, February 2005. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Biaz:2005:RCL

[BV05b] Saâd Biaz and Nitin H. Vaidya. “De-randomizing” congestion losses to improve TCP performance over wired-wireless networks. IEEE/ACM Transactions on Networking, 13 (3):596–608, June 2005. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Bathula:2010:QBM

[BV10] Balagangadhar G. Bathula and Vinod M. Vokkarane. QoS-based manycasting over optical burst-switched (OBS) networks. IEEE/ ACM Transactions on Networking, 18(1):271–283, February 2010. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Buratti:2017:EET

[BVBV17] Chiara Buratti, Roberto Verdone, Chiara Buratti, and Roberto Verdone. Endto-end throughput of ad hoc multi-hop networks in a Poisson field of interferers. IEEE/ACM Transactions on Networking, 25(5): 3189–3202, October 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Basile:2019:ASA

[BVL^{+}19] Cataldo Basile, Fulvio Valenza, Antonio Lioy, Diego R. Lopez, and Antonio Pastor Perales. Adding support for automatic enforcement of security policies in NFV networks. IEEE/ACM Transactions on Networking, 27(2):707– 720, April 2019. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Berger:1998:EBP

[BW98] Arthur W. Berger and Ward Whitt. Effective bandwidths with priorities. IEEE/ACM Transactions on Networking, 6(4):447– 460, August 1998. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1998-6-4/
p447-berger/.

Blocher:2022:HRS

[BWES22] Marcel Blöcher, Lin Wang, Patrick Eugster, and Max Schmidt. Holistic resource scheduling for data center in-network computing. IEEE/ACM Transactions on Networking, 30(6): 2448–2463, 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2022.3174783.

Beysens:2020:CFN

[BWG^{+}20] Jona Beysens, Qing Wang, Ander Galisteo, Domenico Giustiniano, and Sofie Pollin. A cell-free networking system with visible light. IEEE/ACM Transactions on Networking, 28(2): 461–476, April 2020. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL https:

//dl.acm.org/doi/abs/
10.1109/TNET.2020.2966322.

Bridges:2007:CET

[BWH^{+}07] Patrick G. Bridges, Gary T.

Wong, Matti Hiltunen, Richard D. Schlichting, and Matthew J. Barrick. A configurable and extensible transport protocol. IEEE/ ACM Transactions on Networking, 15(6):1254–1265, December 2007. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Barbette:2022:CHS

[BWK^{+}22] Tom Barbette, Erfan Wu, Dejan Kostić, Gerald Q. Maguire, Panagiotis Papadimitratos, and Marco Chiesa. Cheetah: a highspeed programmable loadbalancer framework with guaranteed per-connection-consistency. IEEE/ACM Transactions on Networking, 30(1):354–367, February 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2021.3113370.

Bulut:2010:CEM

[BWS10] Eyuphan Bulut, Zijian Wang, and Boleslaw Karol Szymanski. Cost-effective multiperiod spraying for routing in delay-tolerant networks. IEEE/ACM Transactions on Networking, 18(5):1530–1543, October 2010. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Barbero:2006:CLT

[BY06] Ángela I. Barbero and Øyvind Ytrehus. Cyclelogical treatment for “Cyclopathic” networks. IEEE/ ACM Transactions on Networking, 14(SI):2795–2804, June 2006. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Briat:2015:CIT

[BYH^{+}15] Corentin Briat, Emre Altug Yavuz, Håkan Hjalmarsson, Karl Henrik Johansson, Ulf T. Jönsson, Gunnar Karlsson, and Henrik Sandberg. The conservation of information, towards an axiomatized modular modeling approach to congestion control. IEEE/ACM Transactions on Networking, 23(3):851–865, June 2015. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Bennett:1997:HPF

[BZ97] Jon C. R. Bennett and Hui Zhang. Hierarchical packet fair queueing algorithms. IEEE/ACM Transactions on Networking, 5(5):675–689, October 1997. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1997-5-5/
p675-bennett/.

Brzezinski:2008:DTM

[BZM08] Andrew Brzezinski, Gil Zussman, and Eytan Modiano. Distributed throughput maximization in wireless mesh networks via prepartitioning. IEEE/ACM Transactions on Networking, 16(6):1406–1419, December 2008. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Bhattacharya:2022:SSE

[BZM^{+}22] Arani Bhattacharya, Caitao Zhan, Abhishek Maji, Himanshu Gupta, Samir R. Das, and Petar M. Djurić. Selection of sensors for efficient transmitter localization. IEEE/ACM Transactions on Networking, 30(1):107–119, February 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2021.3104000.

Bressana:2023:PFH

[BZS23] Pietro Bressana, Noa Zilberman, and Robert Soulé.

PTA: Finding hard-to-find data plane bugs. IEEE/ ACM Transactions on Networking, 31(3):1324–1337, June 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2022.3214062.

Chalmers:2003:TMT

[CA03] Robert C. Chalmers and Kevin C. Almeroth. On the topology of multicast trees. IEEE/ACM Transactions on Networking, 11(1):153– 165, February 2003. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Chiaraviglio:2017:LAI

[CAD^{+}17] Luca Chiaraviglio, Lavinia Amorosi, Paolo DellOlmo, William Liu, Jairo A. Gutierrez, Antonio Cianfrani, Marco Polverini, Esther Le Rouzic, and Marco Listanti. Lifetime-aware ISP networks: Optimal formulation and solutions. IEEE/ACM Transactions on Networking, 25(3):1924– 1937, June 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Cruz:2008:SCF

[CAH08] R. L. Cruz and Saleh Al-Harthi. A servicecurve framework for packet scheduling with switch configuration delays. IEEE/ ACM Transactions on Networking, 16(1):196–205, February 2008. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Camp:2012:CFU

[CAK12] Joseph Camp, Ehsan Aryafar, and Edward Knightly. Coupled 802.11 flows in urban channels: model and experimental evaluation. IEEE/ACM Transactions on Networking, 20(5): 1452–1465, October 2012. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Cicconetti:2009:FBA

[CAL09] Claudio Cicconetti, Ian F.

Akyildiz, and Luciano Lenzini. FEBA: a bandwidth allocation algorithm for service differentiation in IEEE 802.16 mesh networks. IEEE/ACM Transactions on Networking, 17 (3):884–897, June 2009. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Canberk:2011:PUA

[CAO11] Berk Canberk, Ian F. Akyildiz, and Sema Oktug. Primary user activity modeling using first-difference filter clustering and correlation in cognitive radio networks. IEEE/ACM Transactions on Networking, 19(1):170–183, February 2011. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Caleffi:2015:SST

[CAP15] Marcello Caleffi, Ian F. Akyildiz, and Luigi Paura. On the solution of the Steiner tree NP-hard problem via Physarum BioNetwork. IEEE/ACM Transactions on Networking, 23(4):1092–1106, August 2015. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Cao:2007:WSD

[CAQ07] Xiaojun Cao, Vishal Anand, and Chunming Qiao. Waveband switching for dynamic traffic demands in multigranular optical networks. IEEE/ACM Transactions on Networking, 15(4):957– 968, August 2007. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Cao:2020:PHO

[CAS^{+}20] Yue Cao, Ahmed Osama Fathy Atya, Shailendra Singh, Zhiyun Qian, Srikanth V. Krishnamurthy, Thomas F. La Porta, Prashant Krishnamurthy, and Lisa Marvel. Packet header obfuscation using MIMO. IEEE/ ACM Transactions on Networking, 28(4):1712–1725, August 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2020.2998398.

Champati:2020:TAM

[CAZG20] Jaya Prakash Champati, Hussein Al-Zubaidy, and James Gross. Transient analysis for multihop wireless networks under static routing. IEEE/ ACM Transactions on Networking, 28(2):722–735, April 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL https://

dl.acm.org/doi/abs/10.
1109/TNET.2020.2975616.

Crovella:1997:SWW

[CB97] Mark E. Crovella and Azer Bestavros. Self-similarity in World Wide Web traffic: evidence and possible causes. IEEE/ACM Transactions on Networking, 5(6):835–846, December 1997. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1997-5-6/
p835-crovella/.

Cheshire:1999:COB

[CB99] Stuart Cheshire and Mary Baker. Consistent overhead Byte stuffing. IEEE/ ACM Transactions on Networking, 7(2):159–172, April 1999. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/
                                        journals/ton/1999-7-2/
                                        p159-cheshire/.

Chau:2011:ALS

[CB11] Chi-Kin Chau and Prithwish Basu. Analysis of latency of stateless opportunistic forwarding in intermittently connected networks. IEEE/ACM Transactions on Networking, 19(4):1111–1124, August 2011. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Chen:2006:CMT

[CBAT06] Zesheng Chen, Tian Bu, Mostafa Ammar, and Don Towsley. Comments on “Modeling TCP Reno performance: a simple model and its empirical validation”. IEEE/ACM Transactions on Networking, 14 (2):451–453, April 2006. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). See [PFTK00].

Chaintreau:2002:ITL

[CBD02] Augustin Chaintreau, François Baccelli, and Christophe Diot. Impact of TCPlike congestion control on the throughput of multicast groups. IEEE/ACM Transactions on Networking, 10 (4):500–512, August 2002. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Caballero:2019:NSG

[CBDCP19] Pablo Caballero, Albert Banchs, Gustavo De Veciana, and Xavier Costa-Perez. Network slicing games: Enabling customization in multi-tenant mobile networks. IEEE/ ACM Transactions on Networking, 27(2):662–675, April 2019. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Caballero:2017:MTR

[CBdV^{+}17] Pablo Caballero, Albert Banchs, Gustavo de Veciana, Xavier Costa-Perez, Pablo Caballero, Albert Banchs, Gustavo de Veciana, and Xavier Costa-Perez. Multi-tenant radio access network slicing: Statistical multiplexing of spatial loads. IEEE/ ACM Transactions on Networking, 25(5):3044–3058, October 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Chen:2020:PBA

[CBHS20] Cheng Chen, Randall A.

Berry, Michael L. Honig, and Vijay G. Subramanian. Pricing, bandwidth allocation, and service competition in heterogeneous wireless networks. IEEE/ ACM Transactions on Networking, 28(5):2299–2308, October 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2020.3008141.

Ciucu:2006:SPS

[CBL06a] Florin Ciucu, Almut Burchard, and Jörg Liebeherr. Scaling properties of statistical end-to-end bounds in the network calculus. IEEE/ACM Transactions on Networking, 14 (SI):2300–2312, June 2006. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Cristescu:2006:LNC

[CBL06b] Razvan Cristescu and Baltasar Beferull-Lozano. Lossy network correlated data gathering with highresolution coding. IEEE/ ACM Transactions on Networking, 14(SI):2817–2824, June 2006. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Chen:2013:CDP

[CBL13] Fei Chen, Bezawada Bruhadeshwar, and Alex X. Liu.

Cross-domain privacy-preserving cooperative firewall optimization. IEEE/ACM Transactions on Networking, 21(3):857–868, June 2013. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Chen:2015:PPQ

[CBL15] Fei Chen, Bruhadeshwar Bezawada, and Alex X. Liu. Privacy-preserving quantification of crossdomain network reachability. IEEE/ACM Transactions on Networking, 23 (3):946–958, June 2015. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Cristescu:2006:NCD

[CBLVW06] Razvan Cristescu, Baltasar Beferull-Lozano, Martin Vetterli, and Roger Wattenhofer. Network correlated data gathering with explicit communication: NP-completeness and algorithms. IEEE/ACM Transactions on Networking, 14 (1):41–54, February 2006. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Chen:2007:ABS

[CBSK07] Yan Chen, David Bindel, Han Hee Song, and Randy H. Katz. Algebra-based scalable overlay network monitoring: algorithms, evaluation, and applications. IEEE/ACM Transactions on Networking, 15(5):1084– 1097, October 2007. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Chatterjee:2018:EFC

[CBV^{+}18] Avhishek Chatterjee, Michael Borokhovich, Lav R. Varshney, Sriram Vishwanath, Lav R. Varshney, Avhishek Chatterjee, Sriram Vishwanath, and Michael Borokhovich. Efficient and flexible crowdsourcing of specialized tasks with precedence constraints. IEEE/ACM Transactions on Networking, 26(2):879–892, April 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Chen:2016:NLN

[CBZ16] Lin Chen, Kaigui Bian, and Meng Zheng. Never live without neighbors: From single- to multi-channel neighbor discovery for mobile sensing applications. IEEE/ACM Transactions on Networking, 24(5):3148– 3161, October 2016. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Chao:1995:DAL

[CC95] H. Jonathan Chao and Byeong-Seog Choe. Design and analysis of a large-scale multicast output buffered ATM switch. IEEE/ACM Transactions on Networking, 3(2):126– 138, April 1995. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1995-3-2/
p126-chao/.

Cheng:1996:DFT

[CC96] Ray-Guang Cheng and Chung-Ju Chang. Design of a fuzzy traffic controller for ATM networks. IEEE/ACM Transactions on Networking, 4(3):460– 469, June 1996. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1996-4-3/
p460-cheng/.

Chakareski:2006:RER

[CC06] Jacob Chakareski and Philip A. Chou. RaDiO edge: rate-distortion optimized proxy-driven streaming from the network edge. IEEE/ACM Transactions on Networking, 14(6):1302– 1312, December 2006. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Choi:1996:EST

[CCA96] Hongsik Choi, Hyeong-Ah Choi, and Murat Azizoğlu. Efficient scheduling of transmissions in optical broadcast networks. IEEE/ ACM Transactions on Networking, 4(6):913–920, December 1996. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/
                                        journals/ton/1996-4-6/
                                        p913-choi/.

Chen:2017:CTS

[CCC17] Min Chen, Shigang Chen, and Zhiping Cai. Counter Tree: a scalable counter architecture for per-flow traffic measurement. IEEE/ ACM Transactions on Networking, 25(2):1249–1262, April 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Cohen:2017:CIT

[CCCC17] Rami Cohen, Yuval Cassuto, Rami Cohen, and Yuval Cassuto. Coding for improved throughput performance in network switches. IEEE/ACM Transactions on Networking, 25(5):2802– 2814, October 2017. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Cai:2006:ISIa

[CCE^{+}06a] Ning Cai, Mung Chiang, Michelle Effros, Ralf Koetter, Muriel Médard, Balaji Prabhakar, R. Srikant, Don Towsley, and Raymond W. Yeung. Introduction to the special issue on networking and information theory. IEEE/ACM Transactions on Networking, 14 (SI):2285–2288, June 2006. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Cai:2006:ISIb

[CCE^{+}06b] Ning Cai, Mung Chiang, Michelle Effros, Ralf Koetter, Muriel Médard, Balaji Prabhakar, R. Srikant, Don Towsley, and Raymond W. Yeung. Introduction to the special issue on networking and information theory. IEEE/ACM Transactions on Networking, 14(4): 674, August 2006. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Cao:2017:OTO

[CCE^{+}17] Zizhong Cao, Paul Claisse, Rene-Jean Essiambre, Murali Kodialam, and T. V. Lakshman. Optimizing throughput in optical networks: The joint routing and power control problem. IEEE/ACM Transactions on Networking, 25(1): 199–209, February 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Chow:2004:FOL

[CCF04] Timothy Y. Chow, Fabian Chudak, and Anthony M. Ffrench. Fast optical layer mesh protection using pre-cross-connected trails. IEEE/ACM Transactions on Networking, 12(3):539– 548, June 2004. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Chen:2017:LAA

[CCF17] Min Chen, Shigang Chen, and Yuguang Fang. Lightweight anonymous authentication protocols for RFID systems. IEEE/ACM Transactions on Networking, 25 (3):1475–1488, June 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Cali:2000:DTI

[CCG00] Frederico Calì, Marco Conti, and Enrico Gregori. Dynamic tuning of the IEEE 802.11 protocol to achieve a theoretical throughput limit. IEEE/ACM Transactions on Networking, 8(6):785– 799, 2000. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/articles/
                                        journals/ton/2000-8-6/
                                        p785-cal/p785-cal.pdf;
                                         http://www.acm.org/pubs/

citations/journals/ton/
2000-8-6/p785-cal/.

Cohen:2020:EDC

[CCG20] Alejandro Cohen, Asaf Cohen, and Omer Gurewitz. Efficient data collection over multiple access wireless sensors network. IEEE/ACM Transactions on Networking, 28(2):491– 504, April 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL https://

dl.acm.org/doi/abs/10.
1109/TNET.2020.2964764.

Chattopadhyay:2016:SDA

[CCK16] Arpan Chattopadhyay, Marceau Coupechoux, and Anurag Kumar. Sequential decision algorithms for measurementbased impromptu deployment of a wireless relay network along a line. IEEE/ ACM Transactions on Networking, 24(5):2954–2968, October 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Chrysos:2016:DNF

[CCKK16] Nikolaos Chrysos, Lydia Chen, Christoforos Kachris, and Manolis Katevenis. Discharging the network from its flow control headaches: packet drops and HOL blocking. IEEE/ACM Transactions on Networking, 24(1):15– 28, February 2016. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Cheng:1999:QPN

[CCL99] Ray-Guang Cheng, Chung-Ju Chang, and Li-Fong Lin. A QoS-Provisioning neural fuzzy connection admission controller for multimedia high-speed networks. IEEE/ACM Transactions on Networking, 7(1):111– 121, February 1999. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL http://

                                        www.acm.org/pubs/citations/
                                        journals/ton/1999-7-1/
                                        p111-cheng/.

Chang:2006:COF

[CCL06] Cheng-Shang Chang, Yi-Ting Chen, and Duan-Shin Lee. Constructions of optical FIFO queues. IEEE/ ACM Transactions on Networking, 14(SI):2838–2843, June 2006. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Chen:2009:CLC

[CCL09] Yi-Ting Chen, Jay Cheng, and Duan-Shin Lee. Constructions of linear compressors, nonovertaking delay lines, and flexible delay lines for optical packet switching. IEEE/ACM Transactions on Networking, 17(6):2014–2027, December 2009. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Chau:2011:CLS

[CCL11] Chi-Kin Chau, Minghua Chen, and Soung Chang Liew. Capacity of largescale CSMA wireless networks. IEEE/ACM Transactions on Networking, 19 (3):893–906, June 2011. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Chang:2017:EEU

[CCLL17] Cheng-Shang Chang, Cheng-Yu Chen, Duan-Shin Lee, and Wanjiun Liao. Efficient encoding of user IDs for nearly optimal expected time-to-rendezvous in heterogeneous cognitive radio networks. IEEE/ACM Transactions on Networking, 25(6):3323–3337, December 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Chang:2002:MST

[CCLT02] Cheng Shang Chang, Rene L.

Cruz, Jean Yves Le Boudec, and Patrick Thiran. A min,+ system theory for constrained traffic regulation and dynamic service guarantees. IEEE/ACM Transactions on Networking, 10(6):805–817, December 2002. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Chisci:2019:ISI

[CCMW19] Giovanni Chisci, Andrea Conti, Lorenzo Mucchi, and Moe Z. Win. Intrinsic secrecy in inhomogeneous stochastic networks. IEEE/ ACM Transactions on Networking, 27(4):1291–1304, August 2019. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Cen:2003:EED

[CCV03] Song Cen, Pamela C. Cosman, and Geoffrey M. Voelker. End-to-end differentiation of congestion and wireless losses. IEEE/ACM Transactions on Networking, 11(5):703–717, October 2003. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Chang:2017:FFG

[CCW^{+}17] Liqiong Chang, Xiaojiang Chen, Yu Wang, Dingyi Fang, Ju Wang, Tianzhang Xing, and Zhanyong Tang. FitLoc: Fine-grained and low-cost device-free localization for multiple targets over various areas. IEEE/ ACM Transactions on Networking, 25(4):1994–2007, August 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Chen:2023:TWA

[CCX^{+}23] Lili Chen, Kai Chen, Jie Xiong, Ke Li, Sunghoon Ivan Lee, Fuwei Wang, Zhanyong Tang, Zheng Wang, Dingyi Fang, and Xiaojiang Chen. Toward wide-area contactless wireless sensing. IEEE/ ACM Transactions on Networking, 31(2):590–605, April 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2022.3196744.

Chen:2014:MQR

[CCY^{+}14] Shannon Chen, Cing-Yu Chu, Su-Ling Yeh, Hao-Hua Chu, and Polly Huang. Modeling the QoE of rate changes in Skype/SILK VoIP calls. IEEE/ACM Transactions on Networking, 22(6):1781–1793, December 2014. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Chang:2023:DDR

[CCZL23] Hao-Hsuan Chang, Hao Chen, Jianzhong Zhang, and Lingjia Liu. Decentralized deep reinforcement learning meets mobility load balancing. IEEE/ ACM Transactions on Networking, 31(2):473–484, April 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2022.3176528.

Chen:2017:ISF

[CCZZ17] Min Chen, Shigang Chen, You Zhou, and Youlin Zhang. Identifying statefree networked tags. IEEE/ ACM Transactions on Networking, 25(3):1607–1620, June 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Clementi:1996:HAO

[CD96] Andrea E. F. Clementi and Miriam Di Ianni. On the hardness of approximating optimum schedule problems in store and forward networks. IEEE/ ACM Transactions on Networking, 4(2):272–280, April 1996. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/
                                        journals/ton/1996-4-2/
                                        p272-clementi/.

Celandroni:2006:LLT

[CDFG06] Nedo Celandroni, Franco Davoli, Erina Ferro, and Alberto Gotta. Long-lived TCP connections via satellite: cross-layer bandwidth allocation, pricing, and adaptive control. IEEE/ ACM Transactions on Networking, 14(5):1019–1030, October 2006. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Chen:2020:HSH

[CDGZ20] Tingjun Chen, Jelena Diakonikolas, Javad Ghaderi, and Gil Zussman. Hybrid scheduling in heterogeneous half- and fullduplex wireless networks. IEEE/ACM Transactions on Networking, 28(2):764– 777, April 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL https://

dl.acm.org/doi/abs/10.
1109/TNET.2020.2973371.

Chin:2010:DIM

[CDH^{+}10] Jren-Chit Chin, Yu Dong, Wing-Kai Hon, Chris Yu-Tak Ma, and David K. Y. Yau. Detection of intelligent mobile target in a mobile sensor network. IEEE/ACM Transactions on Networking, 18(1):41– 52, February 2010. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Carlucci:2017:CCW

[CDHM17] Gaetano Carlucci, Luca De Cicco, Stefan Holmer, and Saverio Mascolo. Congestion control for Web real-time communication. IEEE/ACM Transactions on Networking, 25(5):2629– 2642, October 2017. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Challenger:2004:ESD

[CDI^{+}04] James R. Challenger, Paul Dantzig, Arun Iyengar, Mark S. Squillante, and Li Zhang. Efficiently serving dynamic data at highly accessed Web sites. IEEE/ACM Transactions on Networking, 12(2):233– 246, April 2004. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Courcoubetis:2017:CCB

[CDK^{+}17] Costas A. Courcoubetis, Antonis Dimakis, Michalis Kanakakis, Costas A. Courcoubetis, Antonis Dimakis, and Michalis Kanakakis. Congestion control for background data transfers with minimal delay impact. IEEE/ACM Transactions on Networking, 25(5): 2743–2758, October 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Chen:2021:WFD

[CDKZ21] Tingjun Chen, Mahmood Baraani Dastjerdi, Harish Krishnaswamy, and Gil Zussman. Wideband fullduplex phased array with joint transmit and receive beamforming: Optimization and rate gains. IEEE/ ACM Transactions on Networking, 29(4):1591–1604, August 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2021.3069125.

Cui:2019:TWC

[CDL^{+}19] Yong Cui, Ningwei Dai, Zeqi Lai, Minming Li, Zhenhua Li, Yuming Hu, Kui Ren, and Yuchi Chen. TailCutter: Wisely cutting tail latency in cloud CDNs under cost constraints. IEEE/ACM Transactions on Networking, 27(4):1612– 1628, August 2019. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Coppo:1993:OCD

[CDM93] Paolo Coppo, Matteo D’Ambrosio, and Riccardo Melen. Optimal cost/performance design of ATM switches. IEEE/ ACM Transactions on Networking, 1(5):566–575, October 1993. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/

                                        journals/ton/1993-1-5/
                                        p566-coppo/.

CastroFernandes:2013:ERA

[CDM13] Natalia Castro Fernandes, Marcelo Duffles Donato Moreira, and Otto Carlos Muniz Bandeira Duarte. An efficient and robust addressing protocol for node autoconfiguration in ad hoc networks. IEEE/ACM Transactions on Networking, 21(3):845–856, June 2013. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Castelluccia:1997:GEP

[CDO97] Claude Castelluccia, Walid Dabbous, and Sean O’Malley. Generating efficient protocol code from an abstract specification. IEEE/ ACM Transactions on Networking, 5(4):514–524, August 1997. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/
                                        journals/ton/1997-5-4/
                                        p514-castelluccia/.

Coras:2016:AML

[CDPLCA16] Florin Coras, Jordi Domingo-Pascual, Darrel Lewis, and Albert Cabellos-Aparicio. An analytical model for Loc/ID mappings caches. IEEE/ACM Transactions on Networking, 24(1):506– 516, February 2016. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Cittadini:2011:WRR

[CDRV11] Luca Cittadini, Giuseppe Di Battista, Massimo Rimondini, and Stefano Vissicchio. Wheel + ring = reel: the impact of route filtering on the stability of policy routing. IEEE/ ACM Transactions on Networking, 19(4):1085–1096, August 2011. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Courcoubetis:2002:TES

[CDS02] Costas A. Courcoubetis, Antonis Dimakis, and George D. Stamoulis. Traffic equivalence and substitution in a multiplexer with applications to dynamic available capacity estimation. IEEE/ACM Transactions on Networking, 10 (2):217–231, April 2002. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Cui:2019:ABF

[CDW19] Hui Cui, Robert H. Deng, and Guilin Wang. An attribute-based framework for secure communications in vehicular ad hoc networks. IEEE/ACM Transactions on Networking, 27 (2):721–733, April 2019.

CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Chiu:2008:MFD

[CE08] Yuh-Ming Chiu and Do Young Eun. Minimizing file download time in stochastic peer-to-peer networks. IEEE/ACM Transactions on Networking, 16(2):253– 266, April 2008. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Cai:2009:CBD

[CE09] Han Cai and Do Young Eun. Crossing over the bounded domain: from exponential to power-law intermeeting time in mobile ad hoc networks. IEEE/ ACM Transactions on Networking, 17(5):1578–1591, October 2009. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Cayci:2019:OLD

[CE19] Semih Cayci and Atilla Eryilmaz. Optimal learning for dynamic coding in deadline-constrained multichannel networks. IEEE/ ACM Transactions on Networking, 27(3):1043–1054, June 2019. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Chisci:2019:UMW

[CEC^{+}19] Giovanni Chisci, Hesham Elsawy, Andrea Conti, Mohamed-Slim Alouini, and Moe Z. Win. Uncoordinated massive wireless networks: Spatiotemporal models and multiaccess strategies. IEEE/ACM Transactions on Networking, 27(3):918–931, June 2019. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Chlamtac:1999:SWA

[CEFS99] Imrich Chlamtac, Viktória Elek, Andrea Fumagalli, and Csaba Szabó. Scalable WDM access network architecture based on photonic slot routing. IEEE/ACM Transactions on Networking, 7(1):1–9, February 1999. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1999-7-1/
p1-chlamtac/.

Cohen:2021:ASN

[CEFS21] Itamar Cohen, Gil Einziger, Roy Friedman, and Gabriel Scalosub. Access strategies for network caching. IEEE/ACM Transactions on Networking, 29(2):609– 622, April 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.

                                        acm.org/doi/10.1109/TNET.
                                        2020.3043280.

Cho:2012:IDA

[CER12] Sangman Cho, Theodore Elhourani, and Srinivasan Ramasubramanian. Independent directed acyclic graphs for resilient multipath routing. IEEE/ ACM Transactions on Networking, 20(1):153–162, February 2012. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Cohen:2022:FNA

[CES22] Itamar Cohen, Gil Einziger, and Gabriel Scalosub. False negative awareness in indicator-based caching systems. IEEE/ACM Transactions on Networking, 30(6):2674–2687, 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1109/TNET.2022.3177282.

Chlamtac:1994:MTS

[CF94] Imrich Chlamtac and András Faragó. Making transmission schedules immune to topology changes in multihop packet radio networks. IEEE/ACM Transactions on Networking, 2(1):23–29, February 1994. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/

journals/ton/1994-2-1/
p23-chlamtac/.

Clark:1998:EAB

[CF98] David D. Clark and Wenjia Fang. Explicit allocation of best-effort packet delivery service. IEEE/ ACM Transactions on Networking, 6(4):362–373, August 1998. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1998-6-4/
p362-clark/.

Chich:2001:UDR

[CFC01] Thierry Chich, Pierre Fraigniaud, and Johanne Cohen. Unslotted deflection routing: a practical and efficient protocol for multihop optical networks. IEEE/ACM Transactions on Networking, 9 (1):47–59, 2001. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/articles/
journals/ton/2001-9-1/
p47-chich/p47-chich.pdf;
http://www.acm.org/pubs/
citations/journals/ton/
2001-9-1/p47-chich/.

Carvalho:2024:SDD

[CFC^{+}24] Fabrício B. Carvalho, Ronaldo A.

Ferreira, Ítalo Cunha, Marcos A. M. Vieira, and Murali K. Ramanathan. State disaggregation for dynamic scaling of network functions. IEEE/ACM Transactions on Networking, 32 (1):81–95, 2024. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2023.3282562.

Feng:2005:TCP

[cFCcFW05] Wu chang Feng, Francis Chang, Wu chi Feng, and Jonathan Walpole. A traffic characterization of popular on-line games. IEEE/ACM Transactions on Networking, 13(3):488– 500, June 2005. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Costa:2006:ISD

[CFD06] Luís Henrique M. K. Costa, Serge Fdida, and Otto Carlos M. B. Duarte. Incremental service deployment using the hop-by-hop multicast routing protocol. IEEE/ACM Transactions on Networking, 14 (3):543–556, 2006. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Cerutti:2008:DMS

[CFG08] Isabella Cerutti, Andrea Fumagalli, and Puja Gupta. Delay models of singlesource single-relay cooperative ARQ protocols in slotted radio networks with Poisson frame arrivals. IEEE/ACM Transactions on Networking, 16(2):371– 382, April 2008. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Feng:1999:UIT

[cFKSS99] Wu chang Feng, Dilip D.

Kandlur, Debanjan Saha, and Kang G. Shin. Understanding and improving TCP performance over networks with minimum rate guarantees. IEEE/ ACM Transactions on Networking, 7(2):173–187, April 1999. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1999-7-2/
p173-feng/.

Chen:2009:NQD

[CFM^{+}09] Mingyu Chen, Xingzhe Fan, Manohar N. Murthi, T. Dilusha Wickramarathna, and Kamal Premaratne. Normalized queueing delay: congestion control jointly utilizing delay and marking. IEEE/ACM Transactions on Networking, 17(2):618– 631, April 2009. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Caragiannis:2013:EIM

[CFM13] Ioannis Caragiannis, Michele Flammini, and Luca Moscardelli. An exponential improvement on the MST heuristic for minimum energy broadcasting in ad hoc wireless networks. IEEE/ ACM Transactions on Networking, 21(4):1322–1331, August 2013. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Cociglio:2019:MPM

[CFM^{+}19] Mauro Cociglio, Giuseppe Fioccola, Guido Marchetto, Amedeo Sapio, and Riccardo Sisto. Multipoint passive monitoring in packet networks. IEEE/ ACM Transactions on Networking, 27(6):2377–2390, December 2019. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL https://

                                        dl.acm.org/doi/abs/10.
                                        1109/TNET.2019.2950157.

Casado:2009:REN

[CFP^{+}09] Martín Casado, Michael J. Freedman, Justin Pettit, Jianying Luo, Natasha Gude, Nick McKeown, and Scott Shenker. Rethinking enterprise network control. IEEE/ACM Transactions on Networking, 17(4): 1270–1283, August 2009. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Castiglione:2021:CIC

[CFP^{+}21] Luca Maria Castiglione, Paolo Falcone, Alberto Petrillo, Simon Pietro Romano, and Stefania Santini. Cooperative intersection crossing over 5G. IEEE/ ACM Transactions on Networking, 29(1):303–317, February 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2020.3032652.

Catania:1996:CAF

[CFPP96] Vincenzo Catania, Giuseppe Ficili, Sergio Palazzo, and Daniela Panno. A comparative analysis of fuzzy versus conventional policing mechanisms for ATM networks. IEEE/ ACM Transactions on Networking, 4(3):449–459, June 1996. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1996-4-3/
p449-catania/.

Chekuri:2006:ATA

[CFS06] Chandra Chekuri, Christina Fragouli, and Emina Soljanin. On average throughput and alphabet size in network coding. IEEE/ ACM Transactions on Networking, 14(SI):2410–2424, June 2006. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Cohen:2009:PSG

[CFS09] Reuven Cohen, Niloofar Fazlollahi, and David Starobinski. Path switching and grading algorithms for advance channel reservation architectures. IEEE/ ACM Transactions on Networking, 17(5):1684–1695, October 2009. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Chambers:2010:COG

[CFS^{+}10] Chris Chambers, Wu-Chang Feng, Sambit Sahu, Debanjan Saha, and David Brandt. Characterizing online games. IEEE/ACM Transactions on Networking, 18(3):899–910, June 2010. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Cohen:2011:TCA

[CFS11] Reuven Cohen, Niloofar Fazlollahi, and David Starobinski. Throughputcompetitive advance reservation with bounded path dispersion. IEEE/ACM Transactions on Networking, 19(5):1265–1275, October 2011. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Feng:2002:BAQ

[cFSKS02] Wu chang Feng, Kang G.

Shin, Dilip D. Kandlur, and Debanjan Saha. The BLUE active queue management algorithms. IEEE/ACM Transactions on Networking, 10(4):513– 528, August 2002. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Chlamtac:1994:OSV

[CFZ94] Imrich Chlamtac, András Faragó, and Tao Zhang. Optimizing the system of virtual paths. IEEE/ ACM Transactions on Networking, 2(6):581–587, December 1994. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1994-2-6/
p581-chlamtac/.

Chlamtac:1997:TMT

[CFZ97] Imrich Chlamtac, András Faragó, and Hongbiao Zhang. Time-spread multiple-access (TSMA) protocols for multihop mobile radio networks. IEEE/ ACM Transactions on Networking, 5(6):804–812, December 1997. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/
                                        journals/ton/1997-5-6/
                                        p804-chlamtac/.

Cobb:1997:FT

[CG97] Jorge A. Cobb and Mohamed G. Gouda. Flow theory. IEEE/ACM Transactions on Networking, 5 (5):661–674, October 1997. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL

                                        http://www.acm.org/pubs/
                                        citations/journals/ton/
                                        1997-5-5/p661-cobb/.

Cherkasova:2004:AEM

[CG04] Ludmila Cherkasova and Minaxi Gupta. Analysis of enterprise media server workloads: access patterns, locality, content evolution, and rates of change. IEEE/ACM Transactions on Networking, 12 (5):781–794, October 2004. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Cohen:2015:EAP

[CG15a] Reuven Cohen and Guy Grebla. Efficient allocation of periodic feedback channels in broadband wireless networks. IEEE/ACM Transactions on Networking, 23(2):426–436, April 2015. CODEN IEANEP.

ISSN 1063-6692 (print), 1558-2566 (electronic).

Cohen:2015:JSF

[CG15b] Reuven Cohen and Guy Grebla. Joint scheduling and fast cell selection in OFDMA wireless networks. IEEE/ACM Transactions on Networking, 23(1):114– 125, February 2015. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Casale:2021:PAM

[CG21] Giuliano Casale and Nicolas Gast. Performance analysis methods for listbased caches with nonuniform access. IEEE/ ACM Transactions on Networking, 29(2):651–664, April 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2020.3042869.

Coniglio:2020:ETE

[CGAC20] Stefano Coniglio, Luca Giovanni Gianoli, Edoardo Amaldi, and Antonio Capone. Elastic traffic engineering subject to a fair bandwidth allocation via bilevel programming. IEEE/ACM Transactions on Networking, 28(6):2407–2420, December 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2020.3007572.

Chen:2017:CDD

[CGC^{+}17] Quan Chen, Hong Gao, Siyao Cheng, Xiaolin Fang, Zhipeng Cai, and Jianzhong Li. Centralized and distributed delay-bounded scheduling algorithms for multicast in duty-cycled wireless sensor networks. IEEE/ACM Transactions on Networking, 25(6):3573– 3586, December 2017. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Chen:2018:DLL

[CGC^{+}18] Quan Chen, Hong Gao, Zhipeng Cai, Lianglun Cheng, and Jianzhong Li. Distributed low-latency data aggregation for dutycycle wireless sensor networks. IEEE/ACM Transactions on Networking, 26(5):2347–2360, October 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Chen:2024:PAM

[CGC^{+}24] Quan Chen, Song Guo, Zhipeng Cai, Jing Li, Tuo Shi, and Hong Gao. Peak AoI minimization at wireless-powered network edge: From the perspective of both charging and transmitting. IEEE/ ACM Transactions on Networking, 32(1):806– 821, 2024. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2023.3303266.

Cobb:1998:TSS

[CGEN98] Jorge A. Cobb, Mohamed G. Gouda, and Amal El-Nahas. Timeshift scheduling — fair scheduling of flows in highspeed networks. IEEE/ ACM Transactions on Networking, 6(3):274–285, June 1998. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1998-6-3/
p274-cobb/.

Cidon:1997:IFA

[CGGS97] Israel Cidon, Leonidas Georgiadis, Roch Guérin, and Yuval Shavitt. Improved fairness algorithms for rings with spatial reuse. IEEE/ACM Transactions on Networking, 5(2):190– 204, April 1997. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1997-5-2/
p190-cidon/.

Cidon:1994:PBP

[CGK94] Israel Cidon, Roch Guérin, and Asad Khamisy. On protective buffer policies. IEEE/ACM Transactions on Networking, 2(3):240– 246, June 1994. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/
                                        journals/ton/1994-2-3/
                                        p240-cidon/.

Cohen:2010:CLH

[CGK10] Reuven Cohen, Guy Grebla, and Liran Katzir. Cross-layer hybrid FEC/ ARQ reliable multicast with adaptive modulation and coding in broadband wireless networks. IEEE/ ACM Transactions on Networking, 18(6):1908–1920, December 2010. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Chiasserini:2016:SNA

[CGL16] Carla-Fabiana Chiasserini, Michele Garetto, and Emilio Leonardi. Social network De-anonymization under scale-free user relations. IEEE/ACM Transactions on Networking, 24 (6):3756–3769, December 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Cavendish:2004:CTA

[CGM04] Dirceu Cavendish, Mario Gerla, and Saverio Mascolo. A control theoretical approach to congestion control in packet networks. IEEE/ACM Transactions on Networking, 12 (5):893–906, October 2004. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Cello:2013:OCC

[CGMS13] Marco Cello, Giorgio Gnecco, Mario Marchese, and Marcello Sanguineti. Optimality conditions for coordinate-convex policies in CAC with nonlinear feasibility boundaries. IEEE/ ACM Transactions on Networking, 21(5):1363–1377, October 2013. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Chen:2018:MBT

[CGR^{+}18] Tingjun Chen, Javad Ghaderi, Dan Rubenstein, Gil Zussman, Dan Rubenstein, Tingjun Chen, Javad Ghaderi, and Gil Zussman. Maximizing broadcast throughput under ultra-low-power constraints. IEEE/ACM Transactions on Networking, 26(2):779–792, April 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Cidon:1993:CEH

[CGS93] Israel Cidon, Inder S. Gopal, and Adrian Segall. Connection establishment in high-speed networks. IEEE/ACM Transactions on Networking, 1(4):469– 481, August 1993. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1993-1-4/
p469-cidon/.

Chen:2012:DGA

[CGW^{+}12] Kai Chen, Chuanxiong Guo, Haitao Wu, Jing Yuan, Zhenqian Feng, Yan Chen, Songwu Lu, and Wenfei Wu. DAC: generic and automatic address configuration for data center networks. IEEE/ACM Transactions on Networking, 20(1):84–99, February 2012. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Chen:2000:MPP

[CGY00] Shiwen Chen, Oktay Günlük, and Bülent Yener. The multicast packing problem. IEEE/ACM Transactions on Networking, 8 (3):311–318, 2000. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/2000-8-3/
p311-chen/.

Chen:2016:EST

[CGYZ16] Xu Chen, Xiaowen Gong, Lei Yang, and Junshan Zhang. Exploiting social tie structure for cooperative wireless networking: a social group utility maximization framework. IEEE/ ACM Transactions on Networking, 24(6):3593–3606, December 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Chen:2017:AWS

[CGYZ17] Xu Chen, Xiaowen Gong, Lei Yang, and Junshan Zhang. Amazon in the white space: Social recommendation aided distributed spectrum access. IEEE/ACM Transactions on Networking, 25(1):536– 549, February 2017. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Chen:2020:HTR

[CGZL20] Si Chen, Wei Gong, Jia Zhao, and Jiangchuan Liu. High-throughput and robust rate adaptation for backscatter networks. IEEE/ACM Transactions on Networking, 28(5):2120–2131, October 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.

                                        acm.org/doi/10.1109/TNET.
                                        2020.3002876.

Chen:1993:ACM

[CH93] Xing Chen and Jeremiah F. Hayes. Access control in multicast packet switching. IEEE/ACM Transactions on Networking, 1(6):638–649, December 1993. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/
                                        journals/ton/1993-1-6/
                                        p638-chen/.

Choudhury:1997:NBM

[CH97] Abhijit K. Choudhury and Ellen L. Hahne. A new buffer management scheme for hierarchical shared memory switches. IEEE/ ACM Transactions on Networking, 5(5):728–738, October 1997. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/
                                        journals/ton/1997-5-5/
                                        p728-choudhury/.

Choudhury:1998:DQL

[CH98] Abhijit K. Choudhury and Ellen L. Hahne. Dynamic queue length thresholds for shared-memory packet switches. IEEE/ ACM Transactions on Networking, 6(2):130–140, April 1998. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1998-6-2/
p130-choudhury/.

Chang:2004:PAH

[CH04] Ben-Jye Chang and Ren-Hung Hwang. Performance analysis for hierarchical multirate loss networks. IEEE/ACM Transactions on Networking, 12(1):187– 199, February 2004. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Chiang:2011:CLJ

[CH11] Jerry T. Chiang and Yih-Chun Hu. Cross-layer jamming detection and mitigation in wireless broadcast networks. IEEE/ ACM Transactions on Networking, 19(1):286–298, February 2011. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Cheng:2015:EYJ

[CH15] Chih-Chuan Cheng and Pi-Cheng Hsiu. Extend your journey: considering signal strength and fluctuation in location-based applications. IEEE/ACM Transactions on Networking, 23 (2):451–464, April 2015. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Chen:2018:TTO

[CH18] Kun Chen and Longbo Huang. Timely-throughput optimal scheduling with prediction. IEEE/ACM Transactions on Networking, 26(6):2457–2470, December 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Chang:2020:CDA

[CH20] Nai-Wen Chang and Sun-Yuan Hsieh. Conditional diagnosability of alternating group networks under the PMC model. IEEE/ ACM Transactions on Networking, 28(5):1968–1980, October 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2020.3002093.

Chiu:2021:SDA

[CH21] Cho-Chun Chiu and Ting He. Stealthy DGoS attack: DeGrading of service under the watch of network tomography. IEEE/ ACM Transactions on Networking, 29(3):1294–1307, June 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2021.3058230.

Chen:1995:NMP

[CHA95] C. Y. Roger Chen, Shuo-Hsien Hsiao, and Abdulaziz S. Almazyad. A new model for the performance evaluation of synchronous circuit switched multistage interconnection networks. IEEE/ACM Transactions on Networking, 3(6):708–715, December 1995. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1995-3-6/
p708-chen/.

Chamberland:2010:GAN

[Cha10] Steven Chamberland. Global access network evolution. IEEE/ACM Transactions on Networking, 18(1):136– 149, February 2010. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Chen:2000:ECS

[CHCH00] Wen-Tsuen Chen, Chun-Fu Huang, Yi-Luang Chang, and Wu-Yuin Hwang. An efficient cell-scheduling algorithm for multicast ATM switching systems. IEEE/ACM Transactions on Networking, 8(4):517– 525, 2000. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/

                                        journals/ton/2000-8-4/
                                        p517-chen/.

Chkirbene:2020:LNS

[CHFH20] Zina Chkirbene, Rachid Hadjidj, Sebti Foufou, and Ridha Hamila. LaScaDa: a novel scalable topology for data center network. IEEE/ACM Transactions on Networking, 28(5):2051–2064, October 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2020.3008512.

Chou:2006:UNP

[CHH06] Zi-Tsan Chou, Ching-Chi Hsu, and Shin-Neng Hsu. UPCF: a new point coordination function with QoS and power management for multimedia over wireless LANs. IEEE/ACM Transactions on Networking, 14 (4):807–820, August 2006. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Chang:2016:SSW

[CHL16] Sang-Yoon Chang, Yih-Chun Hu, and Nicola Laurenti. SimpleMAC: a simple wireless MAC-layer countermeasure to intelligent and insider jammers. IEEE/ACM Transactions on Networking, 24(2):1095– 1108, April 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Chi:2007:LFN

[CHLS07] Caixia Chi, Dawei Huang, David Lee, and XiaoRong Sun. Lazy flooding: a new technique for information dissemination in distributed network systems. IEEE/ACM Transactions on Networking, 15(1):80– 92, February 2007. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Chait:2005:TDU

[CHM^{+}05] Yossi Chait, C. V. Hollot, Vishal Misra, Don Towsley, Honggang Zhang, and Yong Cui. Throughput differentiation using coloring at the network edge and preferential marking at the core. IEEE/ACM Transactions on Networking, 13 (4):743–754, August 2005. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Cao:2015:LFL

[CHML15] Zhichao Cao, Yuan He, Qiang Ma, and Yunhao Liu. L^{2}: lazy forwarding in low-duty-cycle wireless sensor network. IEEE/ACM Transactions on Networking, 23(3):922–930, June 2015. CODEN IEANEP.

ISSN 1063-6692 (print), 1558-2566 (electronic).

Chatterjee:2019:SPM

[CHO^{+}19] Bijoy Chand Chatterjee, Fujun He, Eiji Oki, Andrea Fumagalli, and Naoaki Yamanaka. A span power management scheme for rapid lightpath provisioning and releasing in multicore fiber networks. IEEE/ ACM Transactions on Networking, 27(2):734–747, April 2019. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Chen:2020:DBI

[CHS^{+}20] Yongce Chen, Yan Huang, Yi Shi, Y. Thomas Hou, Wenjing Lou, and Sastry Kompella. On DoFbased interference cancellation under general channel rank conditions. IEEE/ ACM Transactions on Networking, 28(3):1002–1016, June 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2020.2974989.

Cheng:2020:HHS

[CHW^{+}20] Bo Cheng, Shoulu Hou, Ming Wang, Shuai Zhao, and Junliang Chen. HSOP: a hybrid service orchestration platform for Internettelephony networks. IEEE/ ACM Transactions on Networking, 28(3):1102–1115, June 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2020.2981477.

Cohen:1997:CRI

[CJ97] Joel E. Cohen and Clark Jeffries. Congestion resulting from increased capacity in single-server queueing networks. IEEE/ ACM Transactions on Networking, 5(2):305–310, April 1997. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1997-5-2/
p305-cohen/.

Cholda:2007:RAO

[CJ07] Piotr Chołda and Andrzej Jajszczyk. Reliability assessment of optical pcycles. IEEE/ACM Transactions on Networking, 15 (6):1579–1592, December 2007. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Chamania:2014:CAE

[CJ14] Mohit Chamania and Admela Jukan. A comparative analysis of the effects of dynamic optical circuit provisioning on IP routing. IEEE/ACM Transactions on Networking, 22 (2):429–442, April 2014. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Chowdhery:2018:ACP

[CJ18] Aakanksha Chowdhery and Kyle Jamieson. Aerial channel prediction and user scheduling in mobile drone hotspots. IEEE/ACM Transactions on Networking, 26(6):2679–2692, December 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Chen:2011:ULG

[CJH^{+}11] Shuyi Chen, Kaustubh R. Joshi, Matti A. Hiltunen, Richard D. Schlichting, and William H. Sanders. Using link gradients to predict the impact of network latency on multitier applications. IEEE/ACM Transactions on Networking, 19(3):855– 868, June 2011. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Choi:2009:RAO

[CJJ09] Kae Won Choi, Wha Sook Jeon, and Dong Geun Jeong. Resource allocation in OFDMA wireless communications systems supporting multimedia services. IEEE/ACM Transactions on Networking, 17 (3):926–935, June 2009.

CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Cui:2019:WNI

[CJL^{+}19] Yong Cui, Yimin Jiang, Zeqi Lai, Xiaomeng Chen, Y. Charlie Hu, Kun Tan, Minglong Dai, Kai Zheng, and Yi Li. Wireless network instabilities in the wild: Measurement, applications NonResilience, and OS Remedy. IEEE/ ACM Transactions on Networking, 27(1):214–230, February 2019. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Chen:2016:EMU

[CJLF16] Xu Chen, Lei Jiao, Wenzhong Li, and Xiaoming Fu. Efficient multi-user computation offloading for mobile-edge cloud computing. IEEE/ACM Transactions on Networking, 24(5): 2795–2808, October 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Christiansen:2001:TRW

[CJOS01] Mikkel Christiansen, Kevin Jeffay, David Ott, and F. Donelson Smith. Tuning RED for Web traffic. IEEE/ACM Transactions on Networking, 9(3): 249–264, June 2001. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Chang:2020:CHS

[CJS^{+}20] Jinyong Chang, Yanyan Ji, Bilin Shao, Maozhi Xu, and Rui Xue. Certificateless homomorphic signature scheme for network coding. IEEE/ACM Transactions on Networking, 28(6):2615–2628, December 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2020.3013902.

Chen:2016:ABD

[CJV16] Chen Chen, Hans-Arno Jacobsen, and Roman Vitenberg. Algorithms based on divide and conquer for topic-based publish/ subscribe overlay design. IEEE/ACM Transactions on Networking, 24(1):422– 436, February 2016. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Chang:2011:LSR

[CJW11] Hyunseok Chang, Sugih Jamin, and Wenjie Wang. Live streaming with receiverbased peer-division multiplexing. IEEE/ACM Transactions on Networking, 19(1):55–68, February 2011. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Choi:2014:DLS

[CJZS14] Jin-Ghoo Choi, Changhee Joo, Junshan Zhang, and Ness B. Shroff. Distributed link scheduling under SINR model in multihop wireless networks. IEEE/ ACM Transactions on Networking, 22(4):1204–1217, August 2014. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Cohen:2000:CVP

[CK00] Reuven Cohen and Gideon Kaempfer. On the cost of virtual private networks. IEEE/ACM Transactions on Networking, 8 (6):775–784, 2000. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/articles/
journals/ton/2000-8-6/
p775-cohen/p775-cohen.
pdf;   http://www.acm.
org/pubs/citations/journals/
ton/2000-8-6/p775-cohen/

.

Cohen:2007:GQA

[CK07] Reuven Cohen and Liran Katzir. A generic quantitative approach to the scheduling of synchronous packets in a shared uplink wireless channel. IEEE/ ACM Transactions on Networking, 15(4):932–943, August 2007. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Cohen:2009:OWS

[CK09] Reuven Cohen and Boris Kapchits. An optimal wake-up scheduling algorithm for minimizing energy consumption while limiting maximum delay in a mesh sensor network. IEEE/ACM Transactions on Networking, 17(2):570– 581, April 2009. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Camp:2010:MRA

[CK10a] Joseph Camp and Edward Knightly. Modulation rate adaptation in urban and vehicular environments: cross-layer implementation and experimental evaluation. IEEE/ ACM Transactions on Networking, 18(6):1949–1962, December 2010. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Cohen:2010:CAE

[CK10b] Reuven Cohen and Liran Katzir. Computational analysis and efficient algorithms for micro and macro OFDMA downlink scheduling. IEEE/ACM Transactions on Networking, 18 (1):15–26, February 2010. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Cohen:2011:CND

[CK11] Reuven Cohen and Boris Kapchits. Continuous neighbor discovery in asynchronous sensor networks. IEEE/ACM Transactions on Networking, 19(1):69– 79, February 2011. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Chau:2016:OAI

[CKA16] Chi-Kin Chau, Majid Khonji, and Muhammad Aftab. Online algorithms for information aggregation from distributed and correlated sources. IEEE/ ACM Transactions on Networking, 24(6):3714–3725, December 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Cuevas:2013:UIC

[CKC^{+}13] Rubén Cuevas, Michal Kryczka, Angel Cuevas, Sebastian Kaune, Carmen Guerrero, and Reza Rejaie. Unveiling the incentives for content publishing in popular BitTorrent portals. IEEE/ACM Transactions on Networking, 21(5): 1421–1435, October 2013. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Choi:2009:SPC

[CKKK09] Lynn Choi, Hyogon Kim, Sunil Kim, and Moon Hae Kim. Scalable packet classification through rulebase partitioning using the maximum entropy hashing. IEEE/ACM Transactions on Networking, 17(6): 1926–1935, December 2009. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Cao:2016:JSD

[CKL16] Zizhong Cao, Murali Kodialam, and T. V. Lakshman. Joint static and dynamic traffic scheduling in data center networks. IEEE/ACM Transactions on Networking, 24(3):1908– 1918, June 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Cohen:2022:LSA

[CKLS22] Reuven Cohen, Matty Kadosh, Alan Lo, and Qasem Sayah. LB scalability: Achieving the right balance between being stateful and stateless. IEEE/ ACM Transactions on Networking, 30(1):382–393, February 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2021.3112517.

Chung:1993:CAB

[CKR93] Shun-Ping Chung, Arik Kashper, and Keith W. Ross. Computing approximate blocking probabilities for large loss networks with state-dependent routing. IEEE/ACM Transactions on Networking, 1(1):105–115, February 1993. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1993-1-1/
p105-chung/.

Cha:2009:AVP

[CKR^{+}09] Meeyoung Cha, Haewoon Kwak, Pablo Rodriguez, Yong-Yeol Ahn, and Sue Moon. Analyzing the video popularity characteristics of large-scale user generated content systems. IEEE/ACM Transactions on Networking, 17(5):1357– 1370, October 2009. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Cai:2016:EDO

[CKS16] Han Cai, Irem Koprulu, and Ness B. Shroff. Exploiting double opportunities for latency-constrained content propagation in wireless networks. IEEE/ ACM Transactions on Networking, 24(2):1025–1037, April 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Chiesa:2017:TEE

[CKS17] Marco Chiesa, Guy Kindler, and Michael Schapira. Traffic engineering with Equal-Cost-MultiPath: an algorithmic perspective. IEEE/ACM Transactions on Networking, 25(2):779– 792, April 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Christodoulopoulos:2011:IDM

[CKV11] Konstantinos Christodoulopoulos, Panagiotis Kokkinos, and Emmanouel Manos Varvarigos. Indirect and direct multicost algorithms for online impairmentaware RWA. IEEE/ACM Transactions on Networking, 19(6):1759–1772, December 2011. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Chiariotti:2019:ADL

[CKZC19] Federico Chiariotti, Stepan Kucera, Andrea Zanella, and Holger Claussen. Analysis and design of a latency control protocol for multi-path data delivery with pre-defined QoS guarantees. IEEE/ACM Transactions on Networking, 27 (3):1165–1178, June 2019.

CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Chan:2003:SPG

[CL03] Man Chi Chan and Tony T.

Lee. Statistical performance guarantees in largescale cross-path packet switch. IEEE/ACM Transactions on Networking, 11 (2):325–337, April 2003. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Chang:2004:BST

[CL04] Cheng-Shang Chang and Zhen Liu. A bandwidth sharing theory for a large number of HTTP-like connections. IEEE/ACM Transactions on Networking, 12(5):952–962, October 2004. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Chu:2005:DRW

[CL05] Xiaowen Chu and Bo Li.

Dynamic routing and wavelength assignment in the presence of wavelength conversion for all-optical networks. IEEE/ACM Transactions on Networking, 13 (3):704–715, June 2005. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Chang:2007:CFS

[CL07] Nicholas B. Chang and Mingyan Liu. Controlled flooding search in a large network. IEEE/ACM Transactions on Networking, 15(2):436–449, April 2007. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Chu:2008:NAA

[CL08] Jian Chu and Chin-Tau Lea. New architecture and algorithms for fast construction of hose-model VPNs. IEEE/ACM Transactions on Networking, 16 (3):670–679, June 2008. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Chang:2009:OCP

[CL09a] Nicholas B. Chang and Mingyan Liu. Optimal channel probing and transmission scheduling for opportunistic spectrum access. IEEE/ACM Transactions on Networking, 17(6): 1805–1818, December 2009. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Chu:2009:OLW

[CL09b] Jian Chu and Chin-Tau Lea. Optimal link weights for IP-based networks supporting hose-model VPNs. IEEE/ACM Transactions on Networking, 17(3):778– 788, June 2009. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Chen:2012:PIP

[CL12] Fei Chen and Alex X.

Liu. Privacy- and integritypreserving range queries in sensor networks. IEEE/ ACM Transactions on Networking, 20(6):1774–1787, December 2012. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Chun:2013:SST

[CL13] Sung Hyun Chun and Richard J. La. Secondary spectrum trading: auction-based framework for spectrum allocation and profit sharing. IEEE/ ACM Transactions on Networking, 21(1):176–189, February 2013. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Checco:2015:FVN

[CL15] Alessandro Checco and Douglas J. Leith. Fair virtualization of 802.11 networks. IEEE/ACM Transactions on Networking, 23(1):148–160, February 2015. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Cheng:2016:RTC

[CL16a] Luwei Cheng and Francis C. M. Lau. Revisiting TCP congestion control in a virtual cluster environment. IEEE/ACM Transactions on Networking, 24(4):2154– 2167, August 2016. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Cohen:2016:DGT

[CL16b] Kobi Cohen and Amir Leshem. Distributed gametheoretic optimization and management of multichannel ALOHA networks. IEEE/ACM Transactions on Networking, 24(3):1718– 1731, June 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Checco:2017:FRD

[CL17] Alessandro Checco and Doug J. Leith. Fast, responsive decentralized graph coloring. IEEE/ ACM Transactions on Networking, 25(6):3628–3640, December 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Chuai:2019:AFR

[CL19] Jie Chuai and Victor O. K. Li. An analytical framework for resource allocation between data and delayed network state information. IEEE/ACM Transactions on Networking, 27(4):1487– 1500, August 2019. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Christin:2007:ECB

[CLA07] Nicolas Christin, Jörg Liebeherr, and Tarek Abdelzaher. Enhancing classbased service architectures with adaptive rate allocation and dropping mechanisms. IEEE/ACM Transactions on Networking, 15 (3):669–682, June 2007. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Chang:2001:NSU

[CLC^{+}01] Zhi-Ren Chang, I-Chung Lee, Cheng-Shang Chang, Chien-Hsin Li, and Ben-Li Sui. A novel scheme using the information of departure processes for delay guarantees of distributed VBR traffic. IEEE/ACM Transactions on Networking, 9(4):452–463, August 2001. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Chen:2012:BAN

[CLC12] Wei Chen, Khaled B.

Letaief, and Zhigang Cao. Buffer-aware network coding for wireless networks. IEEE/ACM Transactions on Networking, 20(5):1389– 1401, October 2012. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Chang:2023:SRC

[CLCL23] Chia-Ming Chang, Yi-Jheng Lin, Cheng-Shang Chang, and Duan-Shin Lee. On the stability regions of coded Poisson receivers with multiple classes of users and receivers. IEEE/ACM Transactions on Networking, 31 (1):234–247, 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2022.3188757.

Chen:2010:RAG

[CLD10] Lijun Chen, Steven H. Low, and John C. Doyle. Random access game and medium access control design. IEEE/ACM Transactions on Networking, 18(4): 1303–1316, August 2010. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Corner:2000:PSI

[CLG^{+}00a] Mark D. Corner, Jörg Liebeherr, Nada Golmie, Chatschik Bisdikian, and David H. Su. A priority scheme for the IEEE 802.14 MAC protocol for hybrid fiber-coax networks.

IEEE/ACM Transactions on Networking, 8(2):200– 211, 2000. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/2000-8-2/
p200-corner/.

Crochat:2000:PIW

[CLG00b] Olivier Crochat, Jean-Yves Le Boudec, and Ornan Gerstel. Protection interoperability for WDM optical networks. IEEE/ACM Transactions on Networking, 8(3):384– 395, 2000. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/2000-8-3/
p384-crochat/.

Cano:2017:FCS

[CLGSS17] Cristina Cano, Douglas J.

Leith, Andres Garcia-Saavedra, and Pablo Serrano. Fair coexistence of scheduled and random access wireless networks: Unlicensed LTE/WiFi. IEEE/ ACM Transactions on Networking, 25(6):3267–3281, December 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Chen:2024:TSL

[CLH^{+}24] Xiang Chen, Hongyan Liu, Qun Huang, Dong Zhang, Haifeng Zhou, Chunming Wu, and Xuan Liu. Toward scalable and low-cost traffic testing for evaluating DDoS defense solutions. IEEE/ACM Transactions on Networking, 32 (1):191–206, 2024. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2023.3281449.

Chen:2022:CPO

[CLHY22] Lin Chen, Shan Lin, Hua Huang, and Weihua Yang. Charging path optimization in mobile networks. IEEE/ACM Transactions on Networking, 30(5):2262– 2273, 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2022.3167781.

Chong:2001:SSS

[CLK01] Song Chong, Sangho Lee, and Sungho Kang. A simple, scalable, and stable explicit rate allocation algorithm for MAX-MIN flow control with minimum rate guarantee. IEEE/ACM Transactions on Networking, 9(3):322–335, June 2001. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Cai:2024:RCF

[CLK^{+}24] Mingxin Cai, Yutong Liu, Linghe Kong, Guihai Chen, Liang Liu, Meikang Qiu, and Shahid Mumtaz. Resource critical flow monitoring in software-defined networks. IEEE/ACM Transactions on Networking, 32(1):396–410, 2024. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL

https://dl.acm.org/doi/
10.1109/TNET.2023.3286691.

Cicconetti:2014:ETD

[CLL^{+}14] Claudio Cicconetti, Luciano Lenzini, Andrea Lodi, Silvano Martello, Enzo Mingozzi, and Michele Monaci. Efficient twodimensional data allocation in IEEE 802.16 OFDMA. IEEE/ACM Transactions on Networking, 22(5):1645– 1658, October 2014. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Chang:2018:PFS

[CLL^{+}18] Cheng-Shang Chang, Duan-Shin Lee, Li-Heng Liou, Sheng-Min Lu, and Mu-Huan Wu. A probabilistic framework for structural analysis and community detection in directed networks. IEEE/ACM Transactions on Networking, 26 (1):31–46, February 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Chen:2019:MRE

[CLL^{+}19] Guo Chen, Yuanwei Lu, Bojie Li, Kun Tan, Yongqiang Xiong, Peng Cheng, Jiansong Zhang, and Thomas Moscibroda. MP-RDMA: Enabling RDMA with multi-path transport in datacenters. IEEE/ACM Transactions on Networking, 27(6):2308–2323, December 2019. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL https://

                                        dl.acm.org/doi/abs/10.
                                        1109/TNET.2019.2948917.

Chaskar:1999:TWL

[CLM99] Hemant M. Chaskar, T. V. Lakshman, and U. Madhow. TCP over wireless with link level error control: analysis and design methodology. IEEE/ ACM Transactions on Networking, 7(5):605–615, October 1999. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/
                                        journals/ton/1999-7-5/
                                        p605-chaskar/.

Chen:2016:ETS

[CLM^{+}16] Min Chen, Wen Luo, Zhen Mo, Shigang Chen, and Yuguang Fang. An efficient tag search protocol in large-scale RFID systems with noisy channel. IEEE/ACM Transactions on Networking, 24 (2):703–716, April 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Chen:2018:FFM

[CLM^{+}18] Guo Chen, Yuanwei Lu, Yuan Meng, Bojie Li, Kun Tan, Dan Pei, Peng Cheng, Layong Luo, Yongqiang Xiong, Xiaoliang Wang, and Youjian Zhao. FUSO: Fast multi-path loss recovery for data center networks. IEEE/ACM Transactions on Networking, 26 (3):1376–1389, June 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Campobello:2012:IES

[CLP12] Giuseppe Campobello, Alessandro Leonardi, and Sergio Palazzo. Improving energy saving and reliability in wireless sensor networks using a simple CRT-based packet-forwarding solution. IEEE/ACM Transactions on Networking, 20(1):191– 205, February 2012. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Cianfrani:2017:IDS

[CLP^{+}17] Antonio Cianfrani, Marco Listanti, Marco Polverini, Antonio Cianfrani, Marco Listanti, and Marco Polverini. Incremental deployment of segment routing into an ISP network: a traffic engineering perspective. IEEE/ ACM Transactions on Networking, 25(5):3146–3160, October 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Chen:2019:DAG

[CLQ^{+}19] Chen Chen, Lei Liu, Tie Qiu, Dapeng Oliver Wu, and Zhiyuan Ren. Delay-aware grid-based geographic routing in urban VANETs: a backbone approach. IEEE/ ACM Transactions on Networking, 27(6):2324–2337, December 2019. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL https://

dl.acm.org/doi/abs/10.
1109/TNET.2019.2944595.

Lau:1997:SMB

[cLqL97] Wing cheong Lau and San qi Li. Statistical multiplexing and buffer sharing in multimedia high-speed networks: a frequencydomain perspective. IEEE/ ACM Transactions on Networking, 5(3):382–396, June 1997. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.

acm.org/pubs/citations/
journals/ton/1997-5-3/
p382-lau/.

Chen:2007:MFS

[CLS07] Cheng Chen, Zheng Guo Li, and Yeng Chai Soh. MRF: a framework for source and destination based bandwidth differentiation service. IEEE/ACM Transactions on Networking, 15(3):697–708, June 2007. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Chang:2018:GPS

[CLS^{+}18] Cheng-Shang Chang, Duan-Shin Lee, Chia-Kai Su, Cheng-Shang Chang, Chia-Kai Su, and Duan-Shin Lee. Greenput: a power-saving algorithm that achieves maximum throughput in wireless networks. IEEE/ACM Transactions on Networking, 26 (2):906–919, April 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Chi:2019:CCT

[CLS^{+}19] Zicheng Chi, Yan Li, Hongyu Sun, Yao, and Ting Zhu. Concurrent crosstechnology communication among heterogeneous IoT devices. IEEE/ACM Transactions on Networking, 27(3):932–947, June 2019. CODEN IEANEP.

ISSN 1063-6692 (print), 1558-2566 (electronic).

Chi:2021:SBD

[CLS^{+}21] Zicheng Chi, Yan Li, Hongyu Sun, Zhichuan Huang, and Ting Zhu. Simultaneous bi-directional communications and data forwarding using a single ZigBee data stream. IEEE/ACM Transactions on Networking, 29(2):821– 833, April 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2021.3054339.

Cheng:2015:SSM

[CLSC15] Yu Cheng, Hongkun Li, Devu Manikantan Shila, and Xianghui Cao. A systematic study of maximal scheduling algorithms in multiradio multichannel wireless networks. IEEE/ ACM Transactions on Networking, 23(4):1342–1355, August 2015. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Chou:2009:PSP

[CLSS09] Jerry Chi-Yuan Chou, Bill Lin, Subhabrata Sen, and Oliver Spatscheck. Proactive surge protection: a defense mechanism for bandwidth-based attacks. IEEE/ACM Transactions on Networking, 17(6):1711– 1723, December 2009. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Cai:2022:URD

[CLTM22] Yang Cai, Jaime Llorca, Antonia M. Tulino, and Andreas F. Molisch. Ultrareliable distributed cloud network control with endto-end latency constraints. IEEE/ACM Transactions on Networking, 30(6):2505– 2520, 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2022.3179349.

Chen:2017:OND

[CLV17] Lin Chen, Yong Li, and Athanasios V. Vasilakos. On oblivious neighbor discovery in distributed wireless networks with directional antennas: Theoretical foundation and algorithm design. IEEE/ ACM Transactions on Networking, 25(4):1982–1993, August 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Choudhury:1995:IAC

[CLW95] Gagan L. Choudhury, Kin K. Leung, and Ward Whitt. An inversion algorithm to compute blocking probabilities in loss networks with statedependent rates. IEEE/ ACM Transactions on Networking, 3(5):585–601, October 1995. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/
                                        journals/ton/1995-3-5/
                                        p585-choudhury/.

Chang:2016:TLB

[CLW16] Cheng-Shang Chang, Wanjiun Liao, and Tsung-Ying Wu. Tight lower bounds for channel hopping schemes in cognitive radio networks. IEEE/ACM Transactions on Networking, 24(4):2343– 2356, August 2016. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Chang:2019:AGF

[CLW19] Cheng-Shang Chang, Duan-Shin Lee, and Chun Wang. Asynchronous grant-free uplink transmissions in multichannel wireless networks with heterogeneous QoS guarantees. IEEE/ ACM Transactions on Networking, 27(4):1584–1597, August 2019. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Cao:2017:CTC

[CLWZ17] Zhichao Cao, Daibo Liu, Jiliang Wang, and Xiaolong Zheng. Chase: Taming concurrent broadcast for flooding in asynchronous duty cycle networks. IEEE/ ACM Transactions on Networking, 25(5):2872–2885, October 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Chen:2022:SNB

[CLX^{+}22] Kuan-Yin Chen, Sen Liu, Yang Xu, Ishant Kumar Siddhrau, Siyu Zhou, Zehua Guo, and H. Jonathan Chao. SDNShield: NFVbased defense framework against DDoS attacks on SDN control plane. IEEE/ ACM Transactions on Networking, 30(1):1–17, February 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2021.3105187.

Chang:2006:PGR

[CLY06] Cheng-Shang Chang, Duan-Shin Lee, and Chi-Yao Yue. Providing guaranteed rate services in the load balanced Birkhoff–von Neumann switches. IEEE/ ACM Transactions on Networking, 14(3):644–656, 2006. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Cheng:2017:DDC

[CLY^{+}17] Jie Cheng, Yaning Liu, Qiang Ye, Hongwei Du, and Athanasios V. Vasilakos. DISCS: a distributed coordinate system based on robust nonnegative matrix completion. IEEE/ACM Transactions on Networking, 25(2):934–947, April 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Chen:2020:PAJ

[CLZ^{+}20] Shuyi Chen, Xiqing Liu, Tianyu Zhao, Hsiao-Hwa Chen, and Weixiao Meng. Performance analysis of joint transmission schemes in ultra-dense networks — a unified approach. IEEE/ ACM Transactions on Networking, 28(1):154–167, February 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL https://

                                        dl.acm.org/doi/abs/10.
                                        1109/TNET.2019.2957319.

Chen:2023:ECP

[CLZ^{+}23] Xiang Chen, Hongyan Liu, Dong Zhang, Qun Huang, Haifeng Zhou, Chunming Wu, and Qiang Yang. Eliminating control plane overload via measurement task placement. IEEE/ACM Transactions on Networking, 31(4):1717– 1731, 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2022.3223420.

Chen:1993:SFP

[CM93] David X. Chen and Jon W.

Mark. SCOQ: a fast packet switch with shared concentration and output queueing. IEEE/ ACM Transactions on Networking, 1(1):142–151, February 1993. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1993-1-1/
p142-chen/.

Chaskar:2003:FST

[CM03] Hemant M. Chaskar and Upamanyu Madhow. Fair scheduling with tunable latency: a round-robin approach. IEEE/ACM Transactions on Networking, 11 (4):592–601, August 2003. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Chakrabarti:2005:RCR

[CM05a] Anirban Chakrabarti and G. Manimaran. Reliability constrained routing in QoS networks. IEEE/ACM Transactions on Networking, 13(3):662–675, June 2005. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Chen:2005:ERW

[CM05b] Li-Wei Chen and Eytan Modiano. Efficient routing and wavelength assignment for reconfigurable WDM ring networks with wavelength converters. IEEE/ ACM Transactions on Networking, 13(1):173–186, February 2005. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Cormode:2005:WNF

[CM05c] Graham Cormode and S. Muthukrishnan. What’s new: Finding significant differences in network data streams. IEEE/ACM Transactions on Networking, 13(6):1219–1232, December 2005. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Carofiglio:2012:ITP

[CM12] Giovanna Carofiglio and Luca Muscariello. On the impact of TCP and perflow scheduling on Internet performance. IEEE/ACM Transactions on Networking, 20(2):620–633, April 2012. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Celik:2015:SNT

[ÇM15] Güner D. Çelik and Eytan Modiano. Scheduling in networks with time-varying channels and reconfiguration delay. IEEE/ACM Transactions on Networking, 23(1):99–113, February 2015. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Chatzipapas:2016:MGM

[CM16] Angelos Chatzipapas and Vincenzo Mancuso. An M/G/1 model for gigabit energy efficient Ethernet links with coalescing and real-trace-based evaluation. IEEE/ACM Transactions on Networking, 24(5):2663– 2675, October 2016. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Cohen:2018:SDS

[CM18] Reuven Cohen and Evgeny Moroshko. Sampling-on-demand in SDN. IEEE/ ACM Transactions on Networking, 26(6):2612–2622, December 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Congdon:2014:SRL

[CMFA14] Paul T. Congdon, Prasant Mohapatra, Matthew Farrens, and Venkatesh Akella. Simultaneously reducing latency and power consumption in OpenFlow switches. IEEE/ACM Transactions on Networking, 22(3):1007–1020, June 2014. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Ciullo:2011:ICM

[CMGL11] Delia Ciullo, Valentina Martina, Michele Garetto, and Emilio Leonardi. Impact of correlated mobility on delay-throughput performance in mobile ad hoc networks. IEEE/ACM Transactions on Networking, 19(6):1745–1758, December 2011. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Chen:1995:QAP

[CMM95] C. Y. Roger Chen, Georges A. Makhoul, and Dikran S. Meliksetian. A queueing analysis of the performance of DQDB. IEEE/ ACM Transactions on Networking, 3(6):872–881, December 1995. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/
                                        journals/ton/1995-3-6/
                                        p872-chen/.

Chiaraviglio:2012:MIN

[CMN12] Luca Chiaraviglio, Marco Mellia, and Fabio Neri. Minimizing ISP network energy cost: formulation and solutions. IEEE/ACM Transactions on Networking, 20(2):463–476, April 2012. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Clad:2014:GCL

[CMP^{+}14] Francois Clad, Pascal Merindol, Jean-Jacques Pansiot, Pierre Francois, and Olivier Bonaventure. Graceful convergence in link-state IP networks: a lightweight algorithm ensuring minimal operational impact. IEEE/ ACM Transactions on Networking, 22(1):300–312, February 2014. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Chaporkar:2016:ODS

[CMP16] Prasanna Chaporkar, Stefan Magureanu, and Alexandre Proutiere. Optimal distributed scheduling in wireless networks under the SINR interference model. IEEE/ACM Transactions on Networking, 24(4):2033– 2045, August 2016. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Chavan:2017:TVQ

[CMR17] Santosh Chavan, Nizar Malangadan, and Gaurav Raina. TCP with virtual queue management policies: Stability and bifurcation analysis. IEEE/ACM Transactions on Networking, 25(2):1020–1033, April 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Christodoulopoulos:2010:ORW

[CMV10] Konstantinos Christodoulopoulos, Konstantinos Manousakis, and Emmanouel Varvarigos. Offline routing and wavelength assignment in transparent WDM networks. IEEE/ACM Transactions on Networking, 18(5):1557–1570, October 2010. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Choi:2020:ISF

[CMW^{+}20] Jaeyoung Choi, Sangwoo Moon, Jiin Woo, Kyunghwan Son, Jinwoo Shin, and Yung Yi. Information source finding in networks: Querying with budgets. IEEE/ACM Transactions on Networking, 28(5):2271–2284, October 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2020.3009946.

Cui:2017:LNC

[CMY^{+}17] Ying Cui, Muriel Medard, Edmund Yeh, Douglas Leith, Fan Lai, and Ken R. Duffy. A linear network code construction for general integer connections based on the constraint satisfaction problem. IEEE/ ACM Transactions on Networking, 25(6):3441–3454, December 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Cui:2018:OBL

[CMY^{+}18] Ying Cui, Muriel Medard, Edmund Yeh, Douglas Leith, and Ken R. Duffy. Optimization-based linear network coding for general connections of continuous flows. IEEE/ ACM Transactions on Networking, 26(5):2033–2047, October 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Cohen:2008:CCE

[CN08] Reuven Cohen and Gabi Nakibly. On the computational complexity and effectiveness of N-hub shortestpath routing. IEEE/ACM Transactions on Networking, 16(3):691–704, June 2008. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Cohen:2009:TEA

[CN09] Reuven Cohen and Gabi Nakibly. A traffic engineering approach for placement and selection of network services. IEEE/ACM Transactions on Networking, 17(2):487–500, April 2009. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Campos-Nanez:2010:DAA

[CN10a] Enrique Campos-Náñez. Decentralized algorithms for adaptive pricing in multiclass loss networks. IEEE/ACM Transactions on Networking, 18(3):830– 843, June 2010. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Cohen:2010:MRT

[CN10b] Reuven Cohen and Gabi Nakibly. Maximizing restorable throughput in MPLS networks. IEEE/ ACM Transactions on Networking, 18(2):568–581, April 2010. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Cohen:2016:RLT

[CN16] Reuven Cohen and Gabi Nakibly. Restorable logical topology in the face of no or partial traffic demand knowledge. IEEE/ ACM Transactions on Networking, 24(4):2074–2085, August 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Cohen:2019:CEV

[CN19] Reuven Cohen and Yuval Nezri. Cardinality estimation in a virtualized network device using online machine learning. IEEE/ ACM Transactions on Networking, 27(5):2098–2110, October 2019. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Chuprikov:2018:PQP

[CNDK18] Pavel Chuprikov, Sergey I.

Nikolenko, Alex Davydow, and Kirill Kogan. Priority queueing for packets with two characteristics. IEEE/ACM Transactions on Networking, 26(1): 342–355, February 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Cheng:2016:AER

[CNG^{+}16] Long Cheng, Jianwei Niu, Yu Gu, Chengwen Luo, and Tian He. Achieving efficient reliable flooding in low-duty-cycle wireless sensor networks. IEEE/ACM Transactions on Networking, 24(6):3676–3689, December 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Chiesa:2017:RSF

[CNM^{+}17] Marco Chiesa, Ilya Nikolaevskiy, Slobodan Mitrovic, Andrei Gurtov, Aleksander Madry, Michael Schapira, and Scott Shenker. On the resiliency of static forwarding tables. IEEE/ ACM Transactions on Networking, 25(2):1133–1146, April 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Carra:2020:EPC

[CNM20] Damiano Carra, Giovanni Neglia, and Pietro Michiardi. Elastic provisioning of cloud caches: a cost-aware TTL approach. IEEE/ACM Transactions on Networking, 28(3):1283– 1296, June 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2020.2980105.

Cohen:2013:ADF

[CNP13] Reuven Cohen, Ilia Nudelman, and Gleb Polevoy. On the admission of dependent flows in powerful sensor networks. IEEE/ ACM Transactions on Networking, 21(5):1461–1471, October 2013. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Charikar:2004:ROQ

[CNS04] Moses Charikar, Joseph Naor, and Baruch Schieber.

Resource optimization in QoS multicast routing of real-time multimedia. IEEE/ACM Transactions on Networking, 12(2):340– 348, April 2004. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Cohen:1994:SML

[CO94] Reuven Cohen and Yoram Ofek. Self-termination mechanism for label swapping routing. IEEE/ ACM Transactions on Networking, 2(5):541–545, October 1994. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1994-2-5/
p541-cohen/.

Cobb:2002:PQS

[Cob02] Jorge Arturo Cobb. Preserving quality of service guarantees in spite of flow aggregation. IEEE/ACM Transactions on Networking, 10(1):43–53, February 2002. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Cohen:1994:SSN

[Coh94] Reuven Cohen. “Session swapping”: a new approach for optimal bandwidth sharing of ring circuit switched channels. IEEE/ACM Transactions on Networking, 2(3):263– 268, June 1994. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/
                                        journals/ton/1994-2-3/
                                        p263-cohen/.

Conway:2011:FSS

[Con11] Adrian E. Conway. Fast simulation of service availability in mesh networks with dynamic path restoration. IEEE/ACM Transactions on Networking, 19 (1):92–101, February 2011. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Cohen:1995:NLS

[COS95] Reuven Cohen, Yoram Ofek, and Adrian Segall. A new label-based source routing for multi-ring networks. IEEE/ACM Transactions on Networking, 3 (3):320–328, June 1995. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL

                                        http://www.acm.org/pubs/
                                        citations/journals/ton/
                                        1995-3-3/p320-cohen/.

Chan:1995:PSM

[CP95] Ming H. Chan and John P. Princen. Prioritized statistical multiplexing of PCM sources. IEEE/ ACM Transactions on Networking, 3(5):549–559, October 1995. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1995-3-5/
p549-chan/.

Chai:2017:PBE

[CP17] Wei Koong Chai and George Pavlou. Path-based epidemic spreading in networks. IEEE/ACM Transactions on Networking, 25(1):565–578, February 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Clark:2018:TUP

[CP18] Matthew A. Clark and Konstantinos Psounis. Trading utility for privacy in shared spectrum access systems. IEEE/ACM Transactions on Networking, 26(1):259–273, February 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Clark:2020:OPU

[CP20] Matthew Clark and Konstantinos Psounis. Optimizing primary user privacy in spectrum sharing systems. IEEE/ ACM Transactions on Networking, 28(2):533–546, April 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL https://

                                        dl.acm.org/doi/abs/10.
                                        1109/TNET.2020.2967776.

Chen:2015:EST

[CPGZ15] Xu Chen, Brian Proulx, Xiaowen Gong, and Junshan Zhang. Exploiting social ties for cooperative D2D communications: a mobile social networking case. IEEE/ACM Transactions on Networking, 23(5): 1471–1484, October 2015. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Cao:2017:EMN

[CPKL17] Zizhong Cao, Shivendra S. Panwar, Murali Kodialam, and T. V. Lakshman. Enhancing mobile networks with software defined networking and cloud computing. IEEE/ACM Transactions on Networking, 25 (3):1431–1444, June 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Chlamtac:1999:ECA

[CPR99] Imrich Chlamtac, Chiara Petrioli, and Jason Redi. Energy-conserving access protocols for identification networks. IEEE/ ACM Transactions on Networking, 7(1):51–59, February 1999. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL http://www.

acm.org/pubs/citations/
journals/ton/1999-7-1/
p51-chlamtac/.

Chen:2012:UMP

[CPS^{+}12] Minghua Chen, Miroslav Ponec, Sudipta Sengupta, Jin Li, and Philip A. Chou. Utility maximization in peer-to-peer systems with applications to video conferencing. IEEE/ ACM Transactions on Networking, 20(6):1681–1694, December 2012. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Clementi:2013:OMM

[CPS13] Andrea Clementi, Francesco Pasquale, and Riccardo Silvestri. Opportunistic MANETs: mobility can make up for low transmission power. IEEE/ACM Transactions on Networking, 21(2):610–620, April 2013. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Ceselli:2017:MEC

[CPS17] Alberto Ceselli, Marco Premoli, and Stefano Secci. Mobile edge cloud network design optimization. IEEE/ACM Transactions on Networking, 25(3):1818– 1831, June 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Cohen:1996:STP

[CPSWL96] Reuven Cohen, Baiju V. Patel, Frank Schaffa, and Marc Willebeek-LeMair. The sink tree paradigm: connectionless traffic support on ATM LAN’s. IEEE/ACM Transactions on Networking, 4(3):363– 374, June 1996. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1996-4-3/
p363-cohen/.

Che:1998:ARM

[CqLL98] Hao Che, San qi Li, and Arthur Lin. Adaptive resource management for flow-based IP/ATM hybrid switching systems. IEEE/ ACM Transactions on Networking, 6(5):544–557, October 1998. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1998-6-5/
p544-che/.

Chen:2022:ARE

[CQLW22] Ning Chen, Tie Qiu, Zilong Lu, and Dapeng Oliver Wu. An adaptive robustness evolution algorithm with self-competition and its 3D deployment for Internet of Things. IEEE/ ACM Transactions on Networking, 30(1):368–381, February 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2021.3113916.

Cao:2018:PTE

[CQW^{+}18] Yue Cao, Zhiyun Qian, Zhongjie Wang, Tuan Dao, Srikanth V. Krishnamurthy, Lisa M. Marvel, Yue Cao, Tuan Dao, Lisa M. Marvel, Zhongjie Wang, Zhiyun Qian, and Srikanth V. Krishnamurthy. Off-path TCP exploits of the challenge ACK global rate limit. IEEE/ACM Transactions on Networking, 26 (2):765–778, April 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Cohen:1998:THP

[CR98] Reuven Cohen and Srinivas Ramanathan. TCP for high performance in hybrid fiber coaxial broadband access networks. IEEE/ACM Transactions on Networking, 6(1):15–29, February 1998. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1998-6-1/
p15-cohen/.

Chandra:1999:MOT

[CR99] Kavitha Chandra and Amy R. Reibman. Modeling one- and two-layer variable bit rate video. IEEE/ACM Transactions on Networking, 7(3):398– 413, June 1999. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/
                                        journals/ton/1999-7-3/
                                        p398-chandra/.

Cohen:2014:CER

[CR14] Rami Cohen and Danny Raz. Cost-effective resource allocation of overlay routing relay nodes. IEEE/ACM Transactions on Networking, 22(2):636– 646, April 2014. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Cardenas:2009:EDA

[CRB09] Alvaro A. Cárdenas, Svetlana Radosavac, and John S. Baras. Evaluation of detection algorithms for MAC layer misbehavior: theory and experiments. IEEE/ACM Transactions on Networking, 17(2):605– 617, April 2009. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Chowdhury:2012:VVN

[CRB12] Mosharaf Chowdhury, Muntasir Raihan Rahman, and Raouf Boutaba. ViNEYard: virtual network embedding algorithms with coordinated node and link mapping. IEEE/ACM Transactions on Networking, 20(1):206– 219, February 2012. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Chen:2008:HTG

[CRD08] Bensong Chen, George N.

Rouskas, and Rudra Dutta. On hierarchical traffic grooming in WDM networks. IEEE/ACM Transactions on Networking, 16(5):1226–1238, October 2008. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Charny:1996:TSA

[CRL96] Anna Charny, K. K. Ramakrishnan, and Anthony Lauck. Time scale analysis scalability issues for explicit rate allocation in ATM networks. IEEE/ ACM Transactions on Networking, 4(4):569–581, August 1996. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1996-4-4/
p569-charny/.

Cidon:1999:AMP

[CRS99] Israel Cidon, Raphael Rom, and Yuval Shavitt. Analysis of multi-path routing. IEEE/ACM Transactions on Networking, 7(6):885–896, December 1999. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/
                                        journals/ton/1999-7-6/
                                        p885-cidon/.

Chiesa:2018:ORI

[CRS18] Marco Chiesa, Gabor Retvari, and Michael Schapira. Oblivious routing in IP networks. IEEE/ACM Transactions on Networking, 26 (3):1292–1305, June 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Checconi:2013:QEP

[CRV13] Fabio Checconi, Luigi Rizzo, and Paolo Valente. QFQ: efficient packet scheduling with tight guarantees. IEEE/ACM Transactions on Networking, 21 (3):802–816, June 2013. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Choe:1998:CAA

[CS98] Jinwoo Choe and Ness B. Shroff. A central-limit-theorem-based approach for analyzing queue behavior in high-speed networks. IEEE/ACM Transactions on Networking, 6(5):659– 671, October 1998. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL http://

www.acm.org/pubs/citations/
journals/ton/1998-6-5/
p659-choe/.

Capone:1999:DQR

[CS99a] Jeffrey M. Capone and Ioannis Stavrakakis. Delivering QoS requirements to traffic with diverse delay tolerances in a TDMA environment. IEEE/ACM Transactions on Networking, 7(1):75–87, February 1999. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1999-7-1/
p75-capone/.

Choi:1999:UCS

[CS99b] Sunghyun Choi and Kang G.

Shin. An uplink CDMA system architecture with diverse QoS guarantees for heterogeneous traffic. IEEE/ACM Transactions on Networking, 7(5):616– 628, October 1999. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL http://

www.acm.org/pubs/citations/
journals/ton/1999-7-5/
p616-choi/.

Choi:2000:UWL

[CS00] Sunghyun Choi and Kang G.

Shin. A unified wireless LAN architecture for real-time and non-real-time communication services. IEEE/ACM Transactions on Networking, 8(1):44–59, February 2000. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/
                                        journals/ton/2000-8-1/
                                        p44-choi/.

Chen:2014:SUD

[CS14] Kang Chen and Haiying Shen. SMART: utilizing distributed social map for lightweight routing in delay-tolerant networks. IEEE/ACM Transactions on Networking, 22(5):1545– 1558, October 2014. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Chen:2015:DFI

[CS15] Kang Chen and Haiying Shen. DTN-FLOW: interlandmark data flow for high-throughput routing in DTNs. IEEE/ACM Transactions on Networking, 23(1):212–226, February 2015. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Chen:2017:FAN

[CS17] Kang Chen and Haiying Shen. FaceChange: Attaining neighbor node anonymity in mobile opportunistic social networks with fine-grained control. IEEE/ACM Transactions on Networking, 25(2):1176– 1189, April 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Chiesa:2021:FRP

[CSA^{+}21] Marco Chiesa, Roshan Sedar, Gianni Antichi, Michael Borokhovich, Andrzej Kamisiński, Georgios Nikolaidis, and Stefan Schmid. Fast ReRoute on programmable switches. IEEE/ACM Transactions on Networking, 29(2):637– 650, April 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2020.3045293.

Chang:1994:ACP

[CSC94] Chung-Ju Chang, Tian-Tsair Su, and Yueh-Yiing Chiang. Analysis of a cutoff priority cellular radio system with finite queueing and reneging/dropping. IEEE/ACM Transactions on Networking, 2(2):166– 175, April 1994. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1994-2-2/
p166-chang/.

Choi:2004:LRD

[CSC04] Hongsik Choi, Suresh Subramaniam, and Hyeong-Ah Choi. Loopback recovery from double-link failures in optical mesh networks. IEEE/ACM Transactions on Networking, 12 (6):1119–1130, December 2004. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Chackochan:2022:AAU

[CSD22] Reena Chackochan, Albert Sunny, and Senthilkumar Dhanasekaran. Approximate aggregate utility maximization using greedy maximal scheduling. IEEE/ACM Transactions on Networking, 30(6): 2521–2530, 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2022.3179451.

Cocchi:1993:PCN

[CSEZ93] Ron Cocchi, Scott Shenker, Deborah Estrin, and Lixia Zhang. Pricing in computer networks: motivation, formulation, and example. IEEE/ACM Transactions on Networking, 1(6):614–627, December 1993. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/

journals/ton/1993-1-6/
p614-cocchi/.

Castro:2014:UTR

[CSG14] Ignacio Castro, Rade Stanojevic, and Sergey Gorinsky. Using tuangou to reduce IP transit costs. IEEE/ ACM Transactions on Networking, 22(5):1415–1428, October 2014. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Chang:2021:TGC

[CSL21] Cheng-Shang Chang, Jang-Ping Sheu, and Yi-Jheng Lin. On the theoretical gap of channel hopping sequences with maximum rendezvous diversity in the multichannel rendezvous problem. IEEE/ ACM Transactions on Networking, 29(4):1620–1633, August 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2021.3067643.

Chang:2013:EGC

[CSLH13] Yeim-Kuan Chang, Cheng-Chien Su, Yung-Chieh Lin, and Sun-Yuan Hsieh. Efficient gray-code-based range encoding schemes for packet classification in TCAM. IEEE/ACM Transactions on Networking, 21(4):1201–1214, August 2013. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Cao:2002:IPG

[CSMW02] Xi-Ren Cao, Hong-Xia Shen, Rodolfo Milito, and Patrica Wirth. Internet pricing with a game theoretical approach: concepts and examples. IEEE/ACM Transactions on Networking, 10(2):208–216, April 2002. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Chou:2006:CBA

[CSN06] Chun-Ting Chou, Kang G. Shin, and Sai Shankar N. Contention-based airtime usage control in multirate IEEE 802.11 wireless LANs. IEEE/ACM Transactions on Networking, 14 (6):1179–1192, December 2006. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Chitavisutthivong:2023:OOR

[CSN^{+}23] Kanatip Chitavisutthivong, Sucha Supittayapornpong, Pooria Namyar, Mingyang Zhang, Minlan Yu, and Ramesh Govindan. Optimal oblivious routing with concave objectives for structured networks. IEEE/ACM Transactions on Networking, 31(6):2669– 2681, 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2023.3264632.

Cui:2017:SDC

[CSR^{+}17] Yong Cui, Jian Song, Kui Ren, Minming Li, Zongpeng Li, Qingmei Ren, and Yangjun Zhang. Software defined cooperative offloading for mobile cloudlets. IEEE/ACM Transactions on Networking, 25(3):1746– 1760, June 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Csikor:2020:TSH

[CSR^{+}20] Levente Csikor, Márk Szalay, Gábor Rétvári, Gergely Pongrácz, Dimitrios P. Pezaros, and László Toka. Transition to SDN is HARMLESS: Hybrid architecture for migrating legacy Ethernet switches to SDN. IEEE/ ACM Transactions on Networking, 28(1):275–288, February 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL https://

dl.acm.org/doi/abs/10.
1109/TNET.2019.2958762.

Chaporkar:2006:DQP

[CSS06] Prasanna Chaporkar, Saswati Sarkar, and Rahul Shetty. Dynamic quorum policy for maximizing throughput in limited information multiparty MAC. IEEE/ACM Transactions on Networking, 14(4):835–848, August 2006. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Chen:2008:TTF

[CSS08] Shigang Chen, Meongchul Song, and Sartaj Sahni. Two techniques for fast computation of constrained shortest paths. IEEE/ ACM Transactions on Networking, 16(1):105–115, February 2008. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Chen:2014:OOS

[CSS^{+}14] Kai Chen, Ankit Singla, Atul Singh, Kishore Ramachandran, Lei Xu, Yueping Zhang, Xitao Wen, and Yan Chen. OSA: an optical switching architecture for data center networks with unprecedented flexibility. IEEE/ACM Transactions on Networking, 22(2):498–511, April 2014. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Cheng:2023:RSS

[CSSG23] Xia Cheng, Junyang Shi, Mo Sha, and Linke Guo. Revealing smart selective jamming attacks in WirelessHART networks.

IEEE/ACM Transactions on Networking, 31(4):1611– 1625, 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2022.3224358.

Chen:2014:SAO

[CSSJ14] Shengbo Chen, Prasun Sinha, Ness B. Shroff, and Changhee Joo. A simple asymptotically optimal joint energy allocation and routing scheme in rechargeable sensor networks. IEEE/ACM Transactions on Networking, 22(4):1325–1336, August 2014. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Cao:1995:PMA

[CT95] Xi-Ren Cao and Don Towsley. A performance model for ATM switches with general packet length distributions. IEEE/ ACM Transactions on Networking, 3(3):299–309, June 1995. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1995-3-3/
p299-cao/.

Cruz:1996:CAA

[CT96] R. L. Cruz and Jung-Tsung Tsai. COD: alternative architectures for high speed packet switching. IEEE/ACM Transactions on Networking, 4(1):11–21, February 1996. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/
                                        journals/ton/1996-4-1/
                                        p11-cruz/.

Chan:2001:DSA

[CT01] S.-H Gary Chan and Fouad Tobagi. Distributed servers architecture for networked video services. IEEE/ACM Transactions on Networking, 9(2):125– 136, 2001. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/articles/
                                        journals/ton/2001-9-2/
                                        p125-chan/p125-chan.pdf;
                                         http://www.acm.org/pubs/
                                        citations/journals/ton/
                                        2001-9-2/p125-chan/.

Chang:2004:MLR

[CT04a] Jae-Hwan Chang and Leandros Tassiulas. Maximum lifetime routing in wireless sensor networks. IEEE/ACM Transactions on Networking, 12(4):609– 619, August 2004. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Chen:2004:HSS

[CT04b] Hsiao-Hwa Chen and Wee-Teck Tea. Hierarchy schedule-sensing protocol for CDMA wireless datacentric networks with multiple packet collision and capture effect. IEEE/ACM Transactions on Networking, 12(6):1036–1048, December 2004. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Cabuk:2022:MTN

[ÇTD22] Umut Can Çabuk, Mustafa Tosun, and Orhan Dagdeviren. MAX-Tree: a novel topology formation for maximal area coverage in wireless ad-hoc networks. IEEE/ACM Transactions on Networking, 30(1):162–175, February 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2021.3110675.

Cheung:2000:DMR

[CTG00] Chi-Chung Cheung, Danny H. K. Tsang, and Sanjay Gupta. Dynamic multicast routing based on mean number of new calls accepted before blocking for single rate loss networks. IEEE/ACM Transactions on Networking, 8 (5):679–688, 2000. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.

                                        acm.org/pubs/citations/
                                        journals/ton/2000-8-5/
                                        p679-cheung/.

Cao:2020:EFT

[CTG^{+}20] Xiaofeng Cao, Guoming Tang, Deke Guo, Yan Li, and Weiming Zhang. Edge federation: Towards an integrated service provisioning model. IEEE/ ACM Transactions on Networking, 28(3):1116–1129, June 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2020.2979361.

Chen:2010:RNO

[CTH10] Irene Y. Chen, Li-Da Tong, and Yi-Ming Huang. Rearrangeable nonblocking optical interconnection network fabrics with crosstalk constraints. IEEE/ACM Transactions on Networking, 18(5):1413–1421, October 2010. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Cunha:2014:DSP

[CTVD14] Ítalo Cunha, Renata Teixeira, Darryl Veitch, and Christophe Diot. DTRACK: a system to predict and track Internet path changes. IEEE/ACM Transactions on Networking, 22(4):1025– 1038, August 2014. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Chao:1995:AQM

[CU95a] H. Jonathan Chao and Necdet Uzun. An ATM queue manager handling multiple delay and loss priorities. IEEE/ACM Transactions on Networking, 3(6):652–659, December 1995. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1995-3-6/
p652-chao/.

Chen:1995:STS

[CU95b] Wen-Huei Chen and Hasan Ural. Synchronizable test sequences based on multiple UIO sequences. IEEE/ACM Transactions on Networking, 3(2):152– 157, April 1995. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1995-3-2/
p152-chen/.

Chandranmenon:1996:TPH

[CV96] Girish P. Chandranmenon and George Varghese. Trading packet headers for packet processing. IEEE/ ACM Transactions on Networking, 4(2):141–152, April 1996. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1996-4-2/
p141-chandranmenon/.

Charbonneau:2012:SRW

[CV12] Neal Charbonneau and Vinod M. Vokkarane. Static routing and wavelength assignment for multicast advance reservation in alloptical wavelength-routed WDM networks. IEEE/ ACM Transactions on Networking, 20(1):1–14, February 2012. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Chang:2022:TTV

[CVHM22] Hyunseok Chang, Matteo Varvello, Fang Hao, and Sarit Mukherjee. A tale of three videoconferencing applications: Zoom, Webex, and Meet. IEEE/ ACM Transactions on Networking, 30(5):2343– 2358, 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2022.3171467.

Clad:2015:CMU

[CVM^{+}15] Francois Clad, Stefano Vissicchio, Pascal Mérindol, Pierre Francois, and Jean-Jacques Pansiot. Computing minimal update sequences for graceful router-wide reconfigurations. IEEE/ACM Transactions on Networking, 23(5):1373–1386, October 2015. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Chatterjee:2017:WCR

[CVV17] Avhishek Chatterjee, Lav R.

Varshney, and Sriram Vishwanath. Work capacity of regulated freelance platforms: Fundamental limits and decentralized schemes. IEEE/ACM Transactions on Networking, 25(6):3641– 3654, December 2017. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Chu:2010:OCS

[CW10] Shan Chu and Xin Wang.

Opportunistic and cooperative spatial multiplexing in MIMO ad hoc networks. IEEE/ACM Transactions on Networking, 18(5):1610– 1623, October 2010. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Courcoubetis:2012:EIS

[CW12] Costas Courcoubetis and Richard Weber. Economic issues in shared infrastructures. IEEE/ACM Transactions on Networking, 20 (2):594–608, April 2012. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Chang:2016:TBM

[CW16] Dao-Yuan Chang and Pi-Chung Wang. TCAMbased multi-match packet classification using multidimensional rule layering. IEEE/ACM Transactions on Networking, 24(2):1125– 1138, April 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Cheng:2019:WNE

[CW19] Linsong Cheng and Jiliang Wang. Walls have no ears: a non-intrusive WiFi-Based user identification system for mobile devices. IEEE/ACM Transactions on Networking, 27(1):245– 257, February 2019. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Chen:2023:APS

[CW23] Qian Chen and Jiliang Wang. AlignTrack: Push the SNR limit of LoRa collision decoding. IEEE/ ACM Transactions on Networking, 31(5):2070– 2085, 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2023.3235041.

Chatterjee:2021:PBI

[CWAO21] Bijoy Chand Chatterjee, Abdul Wadud, Imran Ahmed, and Eiji Oki. Priority-based inter-core and inter-mode crosstalkavoided resource allocation for spectrally-spatially elastic optical networks. IEEE/ ACM Transactions on Networking, 29(4):1634–1647, August 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2021.3068212.

Cai:2014:SPT

[CWGT14] Yan Cai, Xiaolin Wang, Weibo Gong, and Don Towsley. A study on the performance of a three-stage load-balancing switch. IEEE/ACM Transactions on Networking, 22 (1):52–65, February 2014. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Chen:2016:TCD

[CWH^{+}16] Lin Chen, Wei Wang, Hua Huang, Shan Lin, Lin Chen, Wei Wang, Hua Huang, and Shan Lin. On time-constrained data harvesting in wireless sensor networks: Approximation algorithm design. IEEE/ ACM Transactions on Networking, 24(5):3123–3135, October 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Chen:2018:EPP

[CWHW18] Fei Chen, Chunming Wu, Xiaoyan Hong, and Bin Wang. Easy path programming: Elevate abstraction level for network functions. IEEE/ACM Transactions on Networking, 26(1):189– 202, February 2018. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Cao:2021:CXF

[CWL^{+}21] Zhichao Cao, Jiliang Wang, Daibo Liu, Qiang Ma, Xin Miao, and Xufei Mao. Chase++: Fountainenabled fast flooding in asynchronous duty cycle networks. IEEE/ ACM Transactions on Networking, 29(1):410–422, February 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2020.3034251.

Chen:2020:RPL

[CWLH20] Yongrui Chen, Shuai Wang, Zhijun Li, and Tian He. Reliable physical-layer crosstechnology communication with emulation error correction. IEEE/ ACM Transactions on Networking, 28(2):612–624, April 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL https://

dl.acm.org/doi/abs/10.
1109/TNET.2020.2963985.

Chen:2017:TSF

[CWM^{+}17] Kai Chen, Xitao Wen, Xingyu Ma, Yan Chen, Yong Xia, Chengchen Hu, Qunfeng Dong, and Yongqiang Liu. Toward a scalable, fault-tolerant, high-performance optical data center architecture. IEEE/ACM Transactions on Networking, 25(4):2281– 2294, August 2017. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Clark:2005:TCD

[CWSB05] David D. Clark, John Wroclawski, Karen R. Sollins, and Robert Braden. Tussle in cyberspace: defining tomorrow’s Internet. IEEE/ACM Transactions on Networking, 13(3):462– 475, June 2005. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Cui:2015:FFC

[CWW^{+}15] Yong Cui, Lian Wang, Xin Wang, Hongyi Wang, and Yining Wang. FMTCP: a fountain code-based multipath transmission control protocol. IEEE/ACM Transactions on Networking, 23(2):465–478, April 2015. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Cheng:2017:SAD

[CWZ^{+}17] Bo Cheng, Ming Wang, Shuai Zhao, Zhongyi Zhai, Da Zhu, and Junliang Chen. Situation-aware dynamic service coordination in an IoT environment. IEEE/ACM Transactions on Networking, 25(4):2082– 2095, August 2017. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Chen:2023:SFN

[CWZ^{+}23] Haoxian Chen, Chenyuan Wu, Andrew Zhao, Mukund Raghothaman, Mayur Naik, and Boon Thau Loo. Synthesizing formal network specifications from input– output examples. IEEE/ ACM Transactions on Networking, 31(3):994–1009, June 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2022.3208551.

Cheng:2021:LNI

[CWZY21] Fan Cheng, Congtao Wang, Xingyi Zhang, and Yun Yang. A local-neighborhood information based overlapping community detection algorithm for largescale complex networks. IEEE/ACM Transactions on Networking, 29(2):543– 556, April 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2020.3038756.

Chen:2023:LFC

[CXK^{+}23] Xi Chen, Qiao Xiang, Linghe Kong, Huisan Xu, and Xue Liu. Learning from FM communications: Toward accurate, efficient, all-terrain vehicle localization. IEEE/ACM Transactions on Networking, 31 (1):42–57, 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2022.3187885.

Cai:2018:ISF

[CXL18] Kechao Cai, Hong Xie, and John C. S. Lui. Information spreading forensics via sequential dependent snapshots. IEEE/ ACM Transactions on Networking, 26(1):478–491, February 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Chen:2024:QDD

[CXL^{+}24] Lutong Chen, Kaiping Xue, Jian Li, Ruidong Li, Nenghai Yu, Qibin Sun, and Jun Lu. Q-DDCA: Decentralized dynamic congestion avoid routing in large-scale quantum networks. IEEE/ACM Transactions on Networking, 32 (1):368–381, 2024. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2023.3285093.

Cui:2018:DND

[CXW^{+}18] Yong Cui, Shihan Xiao, Xin Wang, Zhenjie Yang, Shenghui Yan, Chao Zhu, Xiang-Yang Li, and Ning Ge. Diamond: Nesting the data center network with wireless rings in 3-D space. IEEE/ACM Transactions on Networking, 26(1):145– 160, February 2018. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Chen:2022:IDI

[CXZ^{+}22] Yanjiao Chen, Meng Xue, Jian Zhang, Runmin Ou, Qian Zhang, and Peng Kuang. DetectDUI : an incar detection system for drink driving and BACs. IEEE/ACM Transactions on Networking, 30(2):896– 910, April 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2021.3125950.

Camtepe:2007:CDK

[ÇY07] Seyit A. Çamtepe and Bülent Yener. Combinatorial design of key distribution mechanisms for wireless sensor networks. IEEE/ACM Transactions on Networking, 15(2):346– 358, April 2007. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Cho:2014:PMP

[CY14] Jeong-Woo Cho and Yung Yi. On the payoff mechanisms in peer-assisted services with multiple content providers: rationality and fairness. IEEE/ACM Transactions on Networking, 22(3):731–744, June 2014. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Chen:2014:RSR

[CYG^{+}14] Chao-Chih Chen, Lihua Yuan, Albert Greenberg, Chen-Nee Chuah, and Prasant Mohapatra. Routingas-a-service (RaaS): a framework for tenantdirected route control in data center. IEEE/ACM Transactions on Networking, 22(5):1401–1414, October 2014. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Chen:2018:JRA

[CYH^{+}18] Qingxia Chen, F. Richard Yu, Tao Huang, Renchao Xie, Jiang Liu, and Yunjie Liu. Joint resource allocation for software-defined networking, caching, and computing. IEEE/ACM Transactions on Networking, 26(1):274–287, February 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Cheng:2009:ORR

[CYK09] Bow-Nan Cheng, Murat Yuksel, and Shivkumar Kalyanaraman. Orthogonal rendezvous routing protocol for wireless mesh networks. IEEE/ACM Transactions on Networking, 17 (2):542–555, April 2009. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Cui:2016:EDP

[CYL16] Ying Cui, Edmund M. Yeh, and Ran Liu. Enhancing the delay performance of dynamic backpressure algorithms. IEEE/ACM Transactions on Networking, 24 (2):954–967, April 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Cao:2022:AGC

[CYTH22] Hankun Cao, Qifa Yan, Xiaohu Tang, and Guojun Han. Adaptive gradient coding. IEEE/ ACM Transactions on Networking, 30(2):717–734, April 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2021.3122873.

Cui:2017:TAV

[CYX^{+}17] Yong Cui, Zhenjie Yang, Shihan Xiao, Xin Wang, and Shenghui Yan. Trafficaware virtual machine migration in topologyadaptive DCN. IEEE/ ACM Transactions on Networking, 25(6):3427–3440, December 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Cheng:2006:DIS

[CZ06] Yu Cheng and Weihua Zhuang. Dynamic inter-SLA resource sharing in path-oriented differentiated services networks. IEEE/ ACM Transactions on Networking, 14(3):657–670, 2006. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Cao:2012:BRU

[CZ12] Lili Cao and Haitao Zheng.

Balancing reliability and utilization in dynamic spectrum access. IEEE/ACM Transactions on Networking, 20(3):651–661, June 2012. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Cheng:2013:TBE

[CZC^{+}13] Wei Cheng, Nan Zhang, Xiuzhen Cheng, Min Song, and Dechang Chen. Timebounded essential localization for wireless sensor networks. IEEE/ACM Transactions on Networking, 21 (2):400–412, April 2013. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Cai:2022:PEV

[CZC^{+}22] Yunxiang Cai, Hongzi Zhu, Shan Chang, Xiao Wang, Jiangang Shen, and Minyi Guo. PeerProbe: Estimating vehicular neighbor distribution with adaptive compressive sensing. IEEE/ACM Transactions on Networking, 30(4):1703– 1716, 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2022.3149008.

Chen:2014:PME

[CZCC14] Yishuai Chen, Baoxian Zhang, Changjia Chen, and Dah Ming Chiu. Performance modeling and evaluation of peer-to-peer live streaming systems under flash crowds. IEEE/ ACM Transactions on Networking, 22(4):1106–1120, August 2014. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Chen:2022:EES

[CZD^{+}22] Xianhao Chen, Guangyu Zhu, Haichuan Ding, Lan Zhang, Haixia Zhang, and Yuguang Fang. End-to-end service auction: a general double auction mechanism for edge computing services. IEEE/ ACM Transactions on Networking, 30(6):2616– 2629, 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2022.3179239.

Cao:2016:OSC

[CZF^{+}16] Xuanyu Cao, Jinbei Zhang, Luoyi Fu, Weijie Wu, and Xinbing Wang. Optimal secrecy capacity-delay tradeoff in large-scale mobile ad hoc networks. IEEE/ACM Transactions on Networking, 24(2):1139–1152, April 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Chlamtac:1998:DAE

[CZFF98] Imrich Chlamtac, Hongbiao Zhang, András Faragó, and Andrea Fumagalli. A deterministic approach to the end-to-end analysis of packet flows in connectionoriented networks. IEEE/ ACM Transactions on Networking, 6(4):422–431, August 1998. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1998-6-4/
p422-chlamtac/.

Carrascosa-Zamacois:2024:WFM

[CZGKB24] Marc Carrascosa-Zamacois, Giovanni Geraci, Edward Knightly, and Boris Bellalta. Wi-Fi multilink operation: an experimental study of latency and throughput. IEEE/ACM Transactions on Networking, 32(1):308– 322, 2024. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2023.3283154.

Chiariotti:2021:HPR

[CZK^{+}21] Federico Chiariotti, Andrea Zanella, Stepan Kucera, Kariem Fahmi, and Holger Claussen. The HOP protocol: Reliable latencybounded end-to-end multipath communication. IEEE/ ACM Transactions on Networking, 29(5):2281–2295, October 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.

acm.org/doi/10.1109/TNET.
2021.3084450.

Chen:2019:HRL

[CZL^{+}19] Qian Chen, Xiao Juan Zhang, Wei Lih Lim, Yuen Sam Kwok, and Sumei Sun. High reliability, low latency and cost effective network planning for industrial wireless mesh networks. IEEE/ ACM Transactions on Networking, 27(6):2354–2362, December 2019. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL https://

dl.acm.org/doi/abs/10.
1109/TNET.2019.2947077.

Chen:2014:EDC

[CZM14] Shaxun Chen, Kai Zeng, and Prasant Mohapatra. Efficient data capturing for network forensics in cognitive radio networks. IEEE/ ACM Transactions on Networking, 22(6):1988–2000, December 2014. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Chen:2024:VCY

[CZM^{+}24] Ning Chen, Sheng Zhang, Zhi Ma, Yu Chen, Yibo Jin, Jie Wu, Zhuzhong Qian, Yu Liang, and Sanglu Lu. ViChaser: Chase your viewpoint for live video streaming with blockoriented super-resolution.

IEEE/ACM Transactions on Networking, 32(1):445– 459, 2024. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2023.3286108.

Cao:2018:DCD

[CZP18] Xuanyu Cao, Junshan Zhang, and H. Vincent Poor. Data center demand response with on-site renewable generation: a bargaining approach. IEEE/ ACM Transactions on Networking, 26(6):2707–2720, December 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Chen:2023:MRA

[CZTX23] Cao Chen, Fen Zhou, Massimo Tornatore, and Shilin Xiao. Maximizing revenue with adaptive modulation and multiple FECs in flexible optical networks. IEEE/ACM Transactions on Networking, 31 (1):220–233, 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2022.3194982.

Chen:2021:EPL

[CZW^{+}21] Zhe Chen, Xu Zhang, Sulei Wang, Yuedong Xu, Jie Xiong, and Xin Wang. Enabling practical large-scale MIMO in WLANs with hybrid beamforming. IEEE/ ACM Transactions on Networking, 29(4):1605–1619, August 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2021.3073160.

Chen:2017:TRF

[CZX^{+}17] Guo Chen, Youjian Zhao, Hailiang Xu, Dan Pei, and Dan Li. F^{2} tree: Rapid failure recovery for routing in production data center networks. IEEE/ ACM Transactions on Networking, 25(4):1940–1953, August 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Chen:2018:CPO

[CZX18] Lixing Chen, Sheng Zhou, and Jie Xu. Computation peer offloading for energy-constrained mobile edge computing in smallcell networks. IEEE/ ACM Transactions on Networking, 26(4):1619–1632, August 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Chen:2021:RPB

[CZZ^{+}21] Si Chen, Maolin Zhang, Jia Zhao, Wei Gong, and Jiangchuan Liu. Reliable and practical Bluetooth backscatter with commodity devices. IEEE/ ACM Transactions on Networking, 29(4):1717–1729, August 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2021.3068865.

Chen:2012:EEE

[CZZY12] Binbin Chen, Ziling Zhou, Yuda Zhao, and Haifeng Yu. Efficient error estimating coding: feasibility and applications. IEEE/ACM Transactions on Networking, 20(1):29–44, February 2012. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Doncel:2019:PDP

[DAA19] Josu Doncel, Samuli Aalto, and Urtzi Ayesta. Performance degradation in parallel-server systems. IEEE/ACM Transactions on Networking, 27(2):875– 888, April 2019. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

delAngel:2004:OPR

[dAF04] Guillermo del Angel and Terrence L. Fine. Optimal power and retransmission control policies for random access systems. IEEE/ ACM Transactions on Networking, 12(6):1156–1166, December 2004. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Dabaghchian:2018:OLR

[DAFZ^{+}18] Monireh Dabaghchian, Amir Alipour-Fanid, Kai Zeng, Qingsi Wang, and Peter Auer. Online learning with randomized feedback graphs for optimal PUE attacks in cognitive radio networks. IEEE/ ACM Transactions on Networking, 26(5):2268–2281, October 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Dai:2022:TFR

[Dai22]  Lin Dai.  A theoreti-

cal framework for random access: Stability regions and transmission control. IEEE/ACM Transactions on Networking, 30(5): 2173–2200, 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2022.3164458.

Datta:2017:CPS

[Dat17] Amitava Datta. Construction of polynomial-size optical priority queues using linear switches and fiber delay lines. IEEE/ACM Transactions on Networking, 25(2):974–987, April 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Develder:2014:JDS

[DBDJ14] Chris Develder, Jens Buysse, Bart Dhoedt, and Brigitte Jaumard. Joint dimensioning of server and network infrastructure for resilient optical grids/clouds. IEEE/ACM Transactions on Networking, 22(5):1591– 1606, October 2014. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Duffy:2013:DCS

[DBL13] Ken R. Duffy, Charles Bordenave, and Douglas J. Leith. Decentralized constraint satisfaction. IEEE/ ACM Transactions on Networking, 21(4):1298–1308, August 2013. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Daly:2019:TFS

[DBL^{+}19] James Daly, Valerio Bruschi, Leonardo Linguaglossa, Salvatore Pontarelli, Dario Rossi, Jerome Tollet, Eric Torng, and Andrew Yourtchenko. TupleMerge: Fast software packet processing for online packet classification. IEEE/ACM Transactions on Networking, 27(4):1417– 1431, August 2019. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Dousse:2005:IIC

[DBT05] Olivier Dousse, François Baccelli, and Patrick Thiran. Impact of interferences on connectivity in ad hoc networks. IEEE/ACM Transactions on Networking, 13(2):425–436, April 2005. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Dang:2020:PCN

[DBW^{+}20] Huynh Tu Dang, Pietro Bressana, Han Wang, Ki Suh Lee, Noa Zilberman, Hakim Weatherspoon, Marco Canini, Fernando Pedone, and Robert Soulé. P4xos: Consensus as a network service. IEEE/ ACM Transactions on Networking, 28(4):1726–1738, August 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2020.2992106.

Dan:2013:CDP

[DC13] György Dán and Niklas Carlsson. Centralized and distributed protocols for tracker-based dynamic swarm management. IEEE/ACM Transactions on Networking, 21(1):297– 310, February 2013. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Dahlin:2003:EEW

[DCGN03] Michael Dahlin, Bharat Baddepudi V. Chandra, Lei Gao, and Amol Nayate. End-to-end WAN service availability. IEEE/ACM Transactions on Networking, 11(2):300–313, April 2003. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Dehghan:2019:SCR

[DCN^{+}19] Mostafa Dehghan, Weibo Chu, Philippe Nain, Don Towsley, and Zhi-Li Zhang. Sharing cache resources among content providers: a utility-based approach. IEEE/ACM Transactions on Networking, 27(2):477– 490, April 2019. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Dong:2019:UPR

[DCZG19] Wei Dong, Chenhong Cao, Xiaoyu Zhang, and Yi Gao. Understanding path reconstruction algorithms in multihop wireless networks. IEEE/ACM Transactions on Networking, 27(1):1–14, February 2019. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Dhamdhere:2011:TYE

[DD11] Amogh Dhamdhere and Constantine Dovrolis. Twelve years in the evolution of the Internet ecosystem. IEEE/ ACM Transactions on Networking, 19(5):1420–1433, October 2011. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Deng:2024:DCU

[DD24] Yong Deng and Min Dong.

Decentralized caching under nonuniform file popularity and size: Memoryrate tradeoff characterization. IEEE/ACM Transactions on Networking, 32 (1):175–190, 2024. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2023.3284347.

DiBartolomeo:2019:ERE

[DDP^{+}19] Marco Di Bartolomeo, Valentino Di Donato, Maurizio Pizzonia, Claudio Squarcella, and Massimo Rimondini. Extracting routing events from traceroutes: a matter of empathy. IEEE/ACM Transactions on Networking, 27 (3):1000–1012, June 2019. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Decasper:2000:RPS

[DDPP00] Dan Decasper, Zubin Dittia, Guru Parulkar, and Bernhard Plattner. Router plugins: a software architecture for next-generation routers. IEEE/ACM Transactions on Networking, 8(1):2–15, February 2000. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/2000-8-1/
p2-decasper/.

Deering:1996:PAW

[DEF^{+}96] Stephen Deering, Deborah L. Estrin, Dino Farinacci, Van Jacobson, Ching-Gung Liu, and Liming Wei. The PIM architecture for wide-area multicast routing. IEEE/ ACM Transactions on Networking, 4(2):153–162, April 1996. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1996-4-2/
p153-deering/.

DiBattista:2007:CTR

[DEH^{+}07] Giuseppe Di Battista, Thomas Erlebach, Alexander Hall, Maurizio Patrignani, Maurizio Pizzonia, and Thomas Schank. Computing the types of the relationships between autonomous systems. IEEE/ ACM Transactions on Networking, 15(2):267–280, April 2007. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

DOro:2017:OPA

[DEP17] Salvatore D’Oro, Eylem Ekici, and Sergio Palazzo. Optimal power allocation and scheduling under jamming attacks. IEEE/ACM Transactions on Networking, 25(3):1310–1323, June 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Devoti:2020:PMW

[DF20] Francesco Devoti and Ilario Filippini. Planning mmwave access networks under obstacle blockages: a reliability-aware approach. IEEE/ACM Transactions on Networking, 28(5):2203–2214, October 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2020.3006926.

Dubois-Ferriere:2011:VDL

[DFGV11] Henri Dubois-Ferrière, Matthias Grossglauser, and Martin Vetterli. Valuable detours: least-cost anypath routing. IEEE/ACM Transactions on Networking, 19(2):333– 346, April 2011. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

DellAmico:2015:UAP

[DFMR15] Matteo Dell’Amico, Maurizio Filippone, Pietro Michiardi, and Yves Roudier. On user availability prediction and network applications. IEEE/ACM Transactions on Networking, 23(4):1300–1313, August 2015. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Dousse:2006:TSW

[DFT06] Olivier Dousse, Massimo Franceschetti, and Patrick Thiran. On the throughput scaling of wireless relay networks. IEEE/ACM Transactions on Networking, 14 (SI):2756–2761, June 2006. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

dHalluin:2002:MCT

[dFV02] Yann d’Halluin, Peter A. Forsyth, and Kenneth R. Vetzal. Managing capacity for telecommunications networks under uncertainty. IEEE/ACM Transactions on Networking, 10 (4):579–587, August 2002. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Dougherty:2006:UNC

[DFZ06] Randall Dougherty, Chris Freiling, and Kenneth Zeger. Unachievability of network coding capacity. IEEE/ACM Transactions on Networking, 14 (SI):2365–2372, June 2006. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Duffield:2001:TSD

[DG01] N. G. Duffield and Matthias Grossglauser. Trajectory sampling for direct traffic observation. IEEE/ACM Transactions on Networking, 9(3):280–292, June 2001. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Duffield:2008:TSU

[DG08] Nick Duffield and Matthias Grossglauser. Trajectory sampling with unreliable reporting. IEEE/ACM Transactions on Networking, 16(1):37–50, February 2008. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Dong:2020:UPT

[DGC^{+}20] Wei Dong, Yi Gao, Chenhong Cao, Xiaoyu Zhang, and Wenbin Wu. Universal path tracing for large-scale sensor networks. IEEE/ACM Transactions on Networking, 28(2):447– 460, April 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL https://

dl.acm.org/doi/abs/10.
1109/TNET.2020.2965587.

Duffield:2002:RMH

[DGG^{+}02] N. G. Duffield, Pawan Goyal, Albert Greenberg, Partho Mishra, K. K. Ramakrishnan, and Jacobus E. van der Merwe. Resource management with hoses: point-to-cloud services for virtual private networks. IEEE/ACM Transactions on Networking, 10 (5):679–692, October 2002. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Deb:2005:RAB

[DGK05] Supratim Deb, Ayalvadi Ganesh, and Peter Key. Resource allocation between persistent and transient flows. IEEE/ACM Transactions on Networking, 13 (2):302–315, April 2005. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Donvito:2016:NNM

[DGLM16] Lidia Donvito, Laura Galluccio, Alfio Lombardo, and Giacomo Morabito. μ-NET: a network for molecular biology applications in microfluidic chips. IEEE/ ACM Transactions on Networking, 24(4):2525–2538, August 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Demianiuk:2021:RDM

[DGNK21] Vitalii Demianiuk, Sergey Gorinsky, Sergey I. Nikolenko, and Kirill Kogan. Robust distributed monitoring of traffic flows. IEEE/ ACM Transactions on Networking, 29(1):275–288, February 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2020.3034890.

Dong:2017:OMA

[DGW^{+}17] Wei Dong, Yi Gao, Wenbin Wu, Jiajun Bu, Chun Chen, and Xiang-Yang Li. Optimal monitor assignment for preferential link tomography in communication networks. IEEE/ ACM Transactions on Networking, 25(1):210–223, February 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Deng:2018:OCP

[DHHD18] Han Deng, I-Hong Hou, I-Hong Hou, and Han Deng. Optimal capacity provisioning for online job allocation with hard allocation ratio requirement. IEEE/ACM Transactions on Networking, 26(2):724– 736, April 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Dams:2016:JRL

[DHK16] Johannes Dams, Martin Hoefer, and Thomas Kesselheim. Jammingresistant learning in wireless networks. IEEE/ ACM Transactions on Networking, 24(5):2809–2818, October 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Du:2023:SAS

[DHS^{+}23] Yang Du, He Huang, Yu-E Sun, Shigang Chen, Guoju Gao, and Xiaocan Wu. Self-adaptive sampling based per-flow traffic measurement. IEEE/ ACM Transactions on Networking, 31(3):1010–1025, June 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2022.3212066.

Dhaini:2014:EET

[DHSS14] Ahmad R. Dhaini, Pin-Han Ho, Gangxiang Shen, and Basem Shihada. Energy efficiency in TDMAbased next-generation passive optical access networks. IEEE/ACM Transactions on Networking, 22 (3):850–863, June 2014. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

DiPalantino:2012:TES

[DJ12] Dominic DiPalantino and Ramesh Johari. Traffic engineering with semiautonomous users: a game-theoretic perspective. IEEE/ACM Transactions on Networking, 20(6):1938– 1949, December 2012. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Dhanapala:2014:TPM

[DJ14] Dulanjalie C. Dhanapala and Anura P. Jayasumana. Topology preserving maps: extracting layout maps of wireless sensor networks from virtual coordinates. IEEE/ACM Transactions on Networking, 22(3):784– 797, June 2014. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Dai:2016:IST

[DJ16] Wei Dai and Scott Jordan. ISP service tier design. IEEE/ACM Transactions on Networking, 24 (3):1434–1447, June 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Du:2022:SBR

[DJB^{+}22] Jun Du, Chunxiao Jiang, Abderrahim Benslimane, Song Guo, and Yong Ren. SDN-based resource allocation in edge and cloud computing systems: an evolutionary Stackelberg differential game approach. IEEE/ACM Transactions on Networking, 30(4):1613– 1628, 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2022.3152150.

Duong:2021:EMB

[DJCA21] Huy Duong, Brigitte Jaumard, David Coudert, and Romualdas Armolavicius. Efficient makebefore-break layer 2 reoptimization. IEEE/ACM Transactions on Networking, 29(5):1910–1921, October 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2021.3078581.

Dasala:2022:SMW

[DJK22] Keerthi Priya Dasala, Josep M. Jornet, and Edward W. Knightly. Scaling mmWave WLANs with single RF chain multiuser beamforming. IEEE/ ACM Transactions on Networking, 30(6):2630– 2643, 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2022.3182976.

Dasala:2023:SMU

[DJK23] Keerthi Priya Dasala, Josep Miquel Jornet, and Edward W. Knightly. Scaling multi-user mmWave WLANs: The case for concurrent uplink transmissions on a single RF chain. IEEE/ACM Transactions on Networking, 31(5):2133– 2146, 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2023.3239438.

Dziong:1997:FBM

[DJM97] Zbigniew Dziong, Marek Juda, and Lorne G. Mason. A framework for bandwidth management in ATM networks — aggregate equivalent bandwidth estimation approach. IEEE/ ACM Transactions on Networking, 5(1):134–147, February 1997. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1997-5-1/
p134-dziong/.

Dehghan:2017:COR

[DJS^{+}17] Mostafa Dehghan, Bo Jiang, Anand Seetharam, Ting He, Theodoros Salonidis, Jim Kurose, Don Towsley, and Ramesh Sitaraman. On the complexity of optimal request routing and content caching in heterogeneous cache networks. IEEE/ACM Transactions on Networking, 25(3):1635– 1648, June 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Dixit:1998:TDM

[DK98] Sudhir S. Dixit and Sharad Kumar. Traffic descriptor mapping and traffic control for frame relay over ATM network. IEEE/ACM Transactions on Networking, 6(1):56–70, February 1998. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1998-6-1/
p56-dixit/.

Dainotti:2015:ASS

[DKC^{+}15] Alberto Dainotti, Alistair King, Kimberly Claffy, Ferdinando Papale, and Antonio Pescapé. Analysis of a “/0” stealth scan from a botnet. IEEE/ACM Transactions on Networking, 23 (2):341–354, April 2015. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

deVeciana:2001:SPA

[DKL01] Gustavo De Veciana, Takis Konstantopoulos, and Tae-Jin Lee. Stability and performance analysis of networks supporting elastic services. IEEE/ACM Transactions on Networking, 9(1):2–14, 2001. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL http://

www.acm.org/pubs/articles/
journals/ton/2001-9-1/
p2-de\_veciana/p2-de\_
veciana.pdf;   http://
www.acm.org/pubs/citations/
journals/ton/2001-9-1/
p2-de\_veciana/;   http:
//www.acm.org/pubs/citations/
journals/ton/2001-9-1/
p2-de_veciana/.

DiFrancesco:2017:SAS

[DKM^{+}17] Paolo Di Francesco, Jacek Kibilda, Francesco Malandrino, Nicholas J. Kaminski, and Luiz A. DaSilva. Sensitivity analysis on service-driven network planning. IEEE/ACM Transactions on Networking, 25 (3):1417–1430, June 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Doeringer:1996:RLP

[DKN96] Willibald Doeringer, Günter Karjoth, and Mehdi Nassehi. Routing on longest-matching prefixes. IEEE/ACM Transactions on Networking, 4(1):86–97, February 1996. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1996-4-1/
p86-doeringer/. See cor-

rections [DKN97].

Doeringer:1997:CRL

[DKN97] Willibald Doeringer, Günter Karjoth, and Mehdi Nassehi. Corrections to “Routing on longest-matching prefixes”. IEEE/ACM Transactions on Networking, 5(4):600, August 1997. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1997-5-4/
p600-doeringer/.  See

[DKN96].

Demianiuk:2021:APC

[DKN21] Vitalii Demianiuk, Kirill Kogan, and Sergey Nikolenko. Approximate packet classifiers with controlled accuracy. IEEE/ ACM Transactions on Networking, 29(3):1141–1154, June 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2021.3056948.

Darzanos:2019:CFE

[DKS19] George Darzanos, Iordanis Koutsopoulos, and George D. Stamoulis. Cloud federations: Economics, games and benefits. IEEE/ ACM Transactions on Networking, 27(5):2111–2124, October 2019. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Dressler:2018:AVB

[DKSC18] Falko Dressler, Florian Klingler, Christoph Sommer, and Reuven Cohen. Not all VANET broadcasts are the same: Contextaware class based broadcast. IEEE/ACM Transactions on Networking, 26 (1):17–30, February 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Dharmapurikar:2006:LPM

[DKT06] Sarang Dharmapurikar, Praveen Krishnamurthy, and David E. Taylor. Longest prefix matching using Bloom filters. IEEE/ ACM Transactions on Networking, 14(2):397–409, April 2006. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Duffield:2004:NTM

[DL04] N. G. Duffield and Francesco Lo Presti. Network tomography from measured end-to-end delay covariance. IEEE/ACM Transactions on Networking, 12(6): 978–992, December 2004. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Dai:2017:SCW

[DLC^{+}17] Haipeng Dai, Yunhuai Liu, Guihai Chen, Xiaobing Wu, Tian He, Alex X. Liu, and Huizhen Ma. Safe charging for wireless power transfer. IEEE/ ACM Transactions on Networking, 25(6):3531–3544, December 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Dai:2018:SSC

[DLC^{+}18a] Haipeng Dai, Yunhuai Liu, Guihai Chen, Xiaobing Wu, Tian He, Alex X. Liu, and Yang Zhao. SCAPE: Safe charging with adjustable power. IEEE/ ACM Transactions on Networking, 26(1):520–533, February 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Ding:2018:IDT

[DLC^{+}18b] Haichuan Ding, Xuanheng Li, Ying Cai, Beatriz Lorenzo, and Yuguang Fang. Intelligent data transportation in smart cities: a spectrum-aware approach. IEEE/ACM Transactions on Networking, 26(6):2598–2611, December 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Dong:2014:MAP

[DLH^{+}14] Wei Dong, Yunhao Liu, Yuan He, Tong Zhu, and Chun Chen. Measurement and analysis on the packet delivery performance in a large-scale sensor network. IEEE/ACM Transactions on Networking, 22(6):1952– 1963, December 2014. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Dong:2011:TDW

[DLL^{+}11] Dezun Dong, Mo Li, Yunhao Liu, Xiang-Yang Li, and Xiangke Liao. Topological detection on wormholes in wireless ad hoc and sensor networks. IEEE/ ACM Transactions on Networking, 19(6):1787–1796, December 2011. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Dai:2020:FPI

[DLL^{+}20] Haipeng Dai, Meng Li, Alex X. Liu, Jiaqi Zheng, and Guihai Chen. Finding persistent items in distributed datasets. IEEE/ ACM Transactions on Networking, 28(1):1–14, February 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL https://

                                        dl.acm.org/doi/abs/10.
                                        1109/TNET.2019.2946417.

Du:2016:RDE

[DLLL16] Wan Du, Zhenjiang Li, Jansen Christian Liando, and Mo Li. From rateless to distanceless: Enabling sparse sensor network deployment in large areas. IEEE/ACM Transactions on Networking, 24(4):2498– 2511, August 2016. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Duffield:2006:NLT

[DLPT06] Nick Duffield, Francesco Lo Presti, Vern Paxson, and Don Towsley. Network loss tomography using striped unicast probes. IEEE/ACM Transactions on Networking, 14(4):697– 710, August 2006. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Du:2018:PDS

[DLR^{+}18] Jian Du, Xue Liu, Lei Rao, Jian Du, Lei Rao, and Xue Liu. Proactive Doppler shift compensation in vehicular cyber-physical systems. IEEE/ACM Transactions on Networking, 26 (2):807–818, April 2018.

CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Duffield:2005:EFD

[DLT05] Nick Duffield, Carsten Lund, and Mikkel Thorup. Estimating flow distributions from sampled flow statistics. IEEE/ACM Transactions on Networking, 13(5):933–946, October 2005. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Du:2015:RNP

[DLT^{+}15] Wei Du, Yongjun Liao, Narisu Tao, Pierre Geurts, Xiaoming Fu, and Guy Leduc. Rating network paths for locality-aware overlay construction and routing. IEEE/ACM Transactions on Networking, 23(5):1661–1673, October 2015. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Daly:2016:DRA

[DLT16] James Daly, Alex X. Liu, and Eric Torng. A difference resolution approach to compressing access control lists. IEEE/ACM Transactions on Networking, 24(1): 610–623, February 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Dai:2017:BHS

[DLW^{+}17] Huichen Dai, Jianyuan Lu, Yi Wang, Tian Pan, and Bin Liu. BFAST: Highspeed and memory-efficient approach for NDN forwarding engine. IEEE/ACM Transactions on Networking, 25(2):1235–1248, April 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Dai:2021:RCW

[DLY^{+}21] Haipeng Dai, Yunhuai Liu, Nan Yu, Chaofeng Wu, Guihai Chen, Tian He, and Alex X. Liu. Radiation constrained wireless charger placement. IEEE/ ACM Transactions on Networking, 29(1):48–64, February 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2020.3028704.

Ding:2022:CRL

[DLY^{+}22] Yi Ding, Ling Liu, Yu Yang, Yunhuai Liu, Desheng Zhang, and Tian He. From conception to retirement: a lifetime story of a 3-year-old wireless beacon system in the wild. IEEE/ ACM Transactions on Networking, 30(1):47–61, February 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.

acm.org/doi/10.1109/TNET.
2021.3107043.

Du:2017:PFE

[DLZL17] Wan Du, Jansen Christian Liando, Huanle Zhang, and Mo Li. Pando: Fountainenabled fast data dissemination with constructive interference. IEEE/ACM Transactions on Networking, 25(2):820–833, April 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

DAmbrosio:1995:ELB

[DM95] Matteo D’Ambrosio and Riccardo Melen. Evaluating the limit behavior of the ATM traffic within a network. IEEE/ ACM Transactions on Networking, 3(6):832–841, December 1995. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1995-3-6/
p832-d_ambrosio/.

Dziong:1996:FCA

[DM96] Zbigniew Dziong and Lorne G.

Mason. Fair-efficient call admission control policies for broadband networks — a game theoretic framework. IEEE/ ACM Transactions on Networking, 4(1):123–136, February 1996. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/
                                        journals/ton/1996-4-1/
                                        p123-dziong/.

Daigle:2003:APN

[DM03] John N. Daigle and Marcos Nascimento Magalhães. Analysis of packet networks having contention-based reservation with application to GPRS. IEEE/ACM Transactions on Networking, 11(4):602–615, August 2003. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

DeCicco:2014:AVS

[DM14] Luca De Cicco and Saverio Mascolo. An adaptive video streaming control system: modeling, validation, and performance evaluation. IEEE/ACM Transactions on Networking, 22(2):526–539, April 2014. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Deb:2015:LBU

[DM15] Supratim Deb and Pantelis Monogioudis. Learningbased uplink interference management in 4G LTE cellular systems. IEEE/ ACM Transactions on Networking, 23(2):398–411, April 2015. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Dimitrijevic:1994:RMN

[DMB94] Dragomir D. Dimitrijevic, Basil Maglaris, and Robert R. Boorstyn. Routing in multidomain networks. IEEE/ACM Transactions on Networking, 2 (3):252–262, June 1994. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL

http://www.acm.org/pubs/
citations/journals/ton/
1994-2-3/p252-dimitrijevic/

.

Deb:2006:AGN

[DMC06] Supratim Deb, Muriel Médard, and Clifford Choute. Algebraic gossip: a network coding approach to optimal multiple rumor mongering. IEEE/ACM Transactions on Networking, 14 (SI):2486–2507, June 2006. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Dau:2017:LNF

[DMDM17] Hoang Dau, Olgica Milenkovic, Hoang Dau, and Olgica Milenkovic. Latent network features and overlapping community discovery via Boolean intersection representations. IEEE/ ACM Transactions on Networking, 25(5):3219–3234, October 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Doshi:2023:CBV

[DME23] Vishwaraj Doshi, Shailaja Mallick, and Do Young Eun. Convergence of bivirus epidemic models with non-linear rates on networks — a monotone dynamical systems approach. IEEE/ACM Transactions on Networking, 31(3):1187– 1201, June 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2022.3213015.

Dasylva:2005:OOC

[DMK05] Abel Dasylva, Delfin Y. Montuno, and Prasad Kodaypak. Optimization of optical cross-connects with wave-mixing conversion. IEEE/ACM Transactions on Networking, 13(2):448– 458, April 2005. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Dai:2018:RCS

[DMLC18] Haipeng Dai, Huizhen Ma, Alex X. Liu, and Guihai Chen. Radiation constrained scheduling of wireless charging tasks. IEEE/ ACM Transactions on Networking, 26(1):314–327, February 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Deb:2014:AEI

[DMMS14] Supratim Deb, Pantelis Monogioudis, Jerzy Miernik, and James P. Seymour. Algorithms for enhanced inter-Cell interference coordination (eICIC) in LTE HetNets. IEEE/ACM Transactions on Networking, 22(1):137–150, February 2014. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Dolev:2006:MTS

[DMS06] Danny Dolev, Osnat Mokryn, and Yuval Shavitt. On multicast trees: structure and size estimation. IEEE/ ACM Transactions on Networking, 14(3):557–567, 2006. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Dehghan:2019:UOA

[DMT^{+}19] Mostafa Dehghan, Laurent Massoulie, Don Towsley, Daniel Sadoc Menasche, and Y. C. Tay. A utility optimization approach to network cache design. IEEE/ACM Transactions on Networking, 27(3):1013– 1027, June 2019. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Demianiuk:2020:NAO

[DNCK20] Vitalii Demianiuk, Sergey Nikolenko, Pavel Chuprikov, and Kirill Kogan. New alternatives to optimize policy classifiers. IEEE/ ACM Transactions on Networking, 28(3):1088–1101, June 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2020.2979400.

Danilchenko:2023:CUQ

[DNS23] Kiril Danilchenko, Zeev Nutov, and Michael Segal. Covering users with QoS by a connected swarm of drones: Graph theoretical approach and experiments. IEEE/ACM Transactions on Networking, 31(6):2483– 2498, 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2022.3231184.

deOliveira:2004:NPP

[dOSAU04] Jaudelice C. de Oliveira, Caterina Scoglio, Ian F. Akyildiz, and George Uhl. New preemption policies for DiffServ-aware traffic engineering to minimize rerouting in MPLS networks. IEEE/ACM Transactions on Networking, 12 (4):733–745, August 2004. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Das:2011:SPS

[DPBT11] Saumitra Das, Konstantina Papagiannaki, Suman Banerjee, and Y. C. Tay. SWARM: the power of structure in community wireless mesh networks. IEEE/ACM Transactions on Networking, 19(3):760– 773, June 2011. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Das:2024:AAI

[DPG^{+}24] Soumadeep Das, Aryan Mohammadi Pasikhani, Prosanta Gope, John A. Clark, Chintan Patel, and Biplab Sikdar. AIDPS: Adaptive intrusion detection and prevention system for underwater acoustic sensor networks. IEEE/ ACM Transactions on Networking, 32(2):1080–1095, April 2024. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2023.3313156.

Destounis:2018:MCS

[DPM^{+}18] Apostolos Destounis, Stefano Paris, Lorenzo Maggi, Georgios S. Paschos, and Jeremie Leguay. Minimum cost SDN routing with reconfiguration frequency constraints. IEEE/ ACM Transactions on Networking, 26(4):1577–1590, August 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Drugan:2011:DCS

[DPMK11] Ovidiu Valentin Drugan, Thomas Plagemann, and Ellen Munthe-Kaas. Detecting communities in sparse MANETs. IEEE/ ACM Transactions on Networking, 19(5):1434–1447, October 2011. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Dimakis:2006:DEC

[DPR06] Alexandros G. Dimakis, Vinod Prabhakaran, and Kannan Ramchandran. Decentralized erasure codes for distributed networked storage. IEEE/ACM Transactions on Networking, 14(SI):2809–2816, June 2006. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Das:2021:MLM

[DPSA21] Shrayan Das, Kirtan Gopal Panda, Debarati Sen, and Wasim Arif. Maximizing last-minute backup in endangered time-varying inter-datacenter networks. IEEE/ACM Transactions on Networking, 29(6):2646– 2663, December 2021. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL https:/

/dl.acm.org/doi/10.1109/
TNET.2021.3098766.

Desmouceaux:2018:SAA

[DPT^{+}18] Yoann Desmouceaux, Pierre Pfister, Jerome Tollet, Mark Townsley, Thomas Clausen, Jerome Tollet, Mark Townsley, Yoann Desmouceaux, Thomas Clausen, and Pierre Pfister. 6LB: Scalable and application-aware load balancing with segment routing. IEEE/ACM Transactions on Networking, 26 (2):819–834, April 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Diamanti:2024:DMR

[DPTP24] Maria Diamanti, Christos Pelekis, Eirini Eleni Tsiropoulou, and Symeon Papavassiliou. Delay minimization for ratesplitting multiple accessbased multi-server MEC offloading. IEEE/ACM Transactions on Networking, 32(2):1035–1047, April 2024. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL

https://dl.acm.org/doi/
10.1109/TNET.2023.3311131.

Dong:2023:SSG

[DQW^{+}23] Tianjian Dong, Qi Qi, Jingyu Wang, Zirui Zhuang, Haifeng Sun, Jianxin Liao, and Zhu Han. Standing on the shoulders of giants: Cross-slice federated meta learning for resource orchestration to cold-start slice. IEEE/ ACM Transactions on Networking, 31(2):828–845, April 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2022.3200853.

Dou:2023:EIC

[DQYG23] Songshi Dou, Li Qi, Chao Yao, and Zehua Guo. Exploring the impact of critical programmability on controller placement for software-defined wide area networks. IEEE/ ACM Transactions on Networking, 31(6):2575– 2588, 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2023.3252639.

Dao:2017:MRC

[DRCM^{+}17] Tuan Dao, Amit K. Roy-Chowdhury, Harsha V. Madhyastha, Srikanth V. Krishnamurthy, and Tom La Porta. Managing redundant content in bandwidth constrained wireless networks. IEEE/ACM Transactions on Networking, 25 (2):988–1003, April 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Dong:2014:IID

[DRJ^{+}14] Wei Dong, Swati Rallapalli, Rittwik Jana, Lili Qiu, K. K. Ramakrishnan, Leo Razoumov, Yin Zhang, and Tae Won Cho. iDEAL: incentivized dynamic cellular offloading via auctions. IEEE/ACM Transactions on Networking, 22(4):1271– 1284, August 2014. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Dovrolis:2004:PDT

[DRM04] Constantinos Dovrolis, Parameswaran Ramanathan, and David Moore. Packet-dispersion techniques and a capacityestimation methodology. IEEE/ACM Transactions on Networking, 12(6):963– 977, December 2004. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

DOro:2018:LCD

[DRMP18] Salvatore D’Oro, Francesco Restuccia, Tommaso Melodia, and Sergio Palazzo. Low-complexity distributed radio access network slicing: Algorithms and experimental results. IEEE/ ACM Transactions on Networking, 26(6):2815–2828, December 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Dong:2016:DAD

[DRQ^{+}16] Wei Dong, Swati Rallapalli, Lili Qiu, K. K. Ramakrishnan, and Yin Zhang. Double auctions for dynamic spectrum allocation. IEEE/ACM Transactions on Networking, 24(4): 2485–2497, August 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Duffield:1998:SAS

[DRR98] N. G. Duffield, K. K. Ramakrishnan, and Amy R. Reibman. SAVE: an algorithm for smoothed adaptive video over explicit rate networks. IEEE/ ACM Transactions on Networking, 6(6):717–728, December 1998. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/
                                        journals/ton/1998-6-6/
                                        p717-duffield/.

Das:2022:SOD

[DRW^{+}22] Sushovan Das, Afsaneh Rahbar, Xinyu Crystal Wu, Zhuang Wang, Weitao Wang, Ang Chen, and T. S. Eugene Ng. Shufflecast: an optical, datarate agnostic, and lowpower multicast architecture for next-generation compute clusters. IEEE/ ACM Transactions on Networking, 30(5):1970– 1985, 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2022.3158899.

Dasylva:1999:OWS

[DS99] Abel Dasylva and R. Srikant.

Optimal WDM schedules for optical star networks. IEEE/ACM Transactions on Networking, 7(3):446– 456, June 1999. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1999-7-3/
p446-dasylva/.

Deb:2004:CCF

[DS04] Supratim Deb and R. Srikant.

Congestion control for fair resource allocation in networks with multicast flows. IEEE/ACM Transactions on Networking, 12(2):274– 285, April 2004. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Dainotti:2014:ACW

[DSA^{+}14] Alberto Dainotti, Claudio Squarcella, Emile Aben, Kimberly C. Claffy, Marco Chiesa, Michele Russo, and Antonio Pescapé. Analysis of country-wide Internet outages caused by censorship. IEEE/ACM Transactions on Networking, 22(6): 1964–1977, December 2014. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Silva:1995:PSS

[dSeSGM95] Edmundo de Souza e. Silva, H. Richard Gail, and Richard R. Muntz. Polling systems with server timeouts and their application to token passing networks. IEEE/ ACM Transactions on Networking, 3(5):560–575, October 1995. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1995-3-5/
p560-de_souza_e_silva/.

Dai:2018:IEP

[DSL^{+}18] Haipeng Dai, Muhammad Shahzad, Alex X. Liu, Meng Li, Yuankun Zhong, and Guihai Chen. Identifying and estimating persistent items in data streams. IEEE/ACM Transactions on Networking, 26(6):2429– 2442, December 2018. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Dao:2017:TUC

[DSM^{+}17] Tuan A. Dao, Indrajeet Singh, Harsha V. Madhyastha, Srikanth V. Krishnamurthy, Guohong Cao, and Prasant Mohapatra. TIDE: a usercentric tool for identifying energy hungry applications on Smartphones. IEEE/ACM Transactions on Networking, 25(3):1459– 1474, June 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Dovrolis:2002:PDS

[DSR02] Constantinos Dovrolis, Dimitrios Stiliadis, and Parameswaran Ramanathan. Proportional differentiated services: delay differentiation and packet scheduling. IEEE/ACM Transactions on Networking, 10 (1):12–26, February 2002. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Dikbiyik:2012:EEC

[DSTM12] Ferhat Dikbiyik, Laxman Sahasrabuddhe, Massimo Tornatore, and Biswanath Mukherjee. Exploiting excess capacity to improve robustness of WDM mesh networks. IEEE/ ACM Transactions on Networking, 20(1):114–124, February 2012. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Dai:2024:MLT

[DSYN24] Miao Dai, Gang Sun, Hongfang Yu, and Dusit Niyato. Maximize the long-term average revenue of network slice provider via admission control among heterogeneous slices. IEEE/ACM Transactions on Networking, 32 (1):745–760, 2024. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2023.3297883.

Devetsikiotis:1993:SOD

[DT93] Michael Devetsikiotis and J. Keith Townsend. Statistical optimization of dynamic importance sampling parameters for efficient simulation of communication networks. IEEE/ ACM Transactions on Networking, 1(3):293–305, June 1993. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/
                                        journals/ton/1993-1-3/
                                        p293-devetsikiotis/.

Dinh:2015:NUJ

[DT15] Thang N. Dinh and My T. Thai. Network under joint node and link attacks: vulnerability assessment methods and analysis.

IEEE/ACM Transactions on Networking, 23(3):1001– 1011, June 2015. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Dikbiyik:2015:EEC

[DTM15] Ferhat Dikbiyik, Massimo Tornatore, and Biswanath Mukherjee. Exploiting excess capacity, part II: differentiated services under traffic growth. IEEE/ ACM Transactions on Networking, 23(5):1599–1609, October 2015. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

DAronco:2017:IUB

[DTM^{+}17] Stefano D’Aronco, Laura Toni, Sergio Mena, Xiaoqing Zhu, and Pascal Frossard. Improved utilitybased congestion control for delay-constrained communication. IEEE/ACM Transactions on Networking, 25(1):349–362, February 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Dinh:2021:FLW

[DTN^{+}21] Canh T. Dinh, Nguyen H.

Tran, Minh N. H. Nguyen, Choong Seon Hong, Wei Bao, Albert Y. Zomaya, and Vincent Gramoli. Federated learning over wireless networks: Convergence analysis and resource allocation. IEEE/ ACM Transactions on Networking, 29(1):398–409, February 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2020.3035770.

Djukic:2009:DAL

[DV09] Petar Djukic and Shahrokh Valaee. Delay aware link scheduling for multi-hop TDMA wireless networks. IEEE/ACM Transactions on Networking, 17(3):870– 883, June 2009. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Damjanovic:2011:ETS

[DW11] Dragana Damjanovic and Michael Welzl. An extension of the TCP steadystate throughput equation for parallel flows and its application in MulTFRC. IEEE/ACM Transactions on Networking, 19(6):1676– 1689, December 2011. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Deng:2017:TWF

[DWCZ17] Lei Deng, Chih-Chun Wang, Minghua Chen, and Shizhen Zhao. Timely wireless flows with general traffic patterns: Capacity region and scheduling algorithms. IEEE/ACM Transactions on Networking, 25 (6):3473–3486, December 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Dai:2018:WCP

[DWL^{+}18] Haipeng Dai, Xiaoyu Wang, Alex X. Liu, Huizhen Ma, Guihai Chen, and Wanchun Dou. Wireless charger placement for directional charging. IEEE/ ACM Transactions on Networking, 26(4):1865–1878, August 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Dinh:2012:NAA

[DXT^{+}12] Thang N. Dinh, Ying Xuan, My T. Thai, Panos M. Pardalos, and Taieb Znati. On new approaches of assessing network vulnerability: hardness and approximation. IEEE/ACM Transactions on Networking, 20(2):609–619, April 2012. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Du:2023:RAR

[DXX^{+}23] Xinle Du, Ke Xu, Lei Xu, Kai Zheng, Meng Shen, Bo Wu, and Tong Li. R-AQM: Reverse ACK active queue management in multitenant data centers. IEEE/ACM Transactions on Networking, 31(2):526– 541, April 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2022.3197973.

Deng:2013:FPH

[DYH13] Xi Deng, Yuanyuan Yang, and Sangjin Hong. A flexible platform for hardwareaware network experiments and a case study on wireless network coding. IEEE/ ACM Transactions on Networking, 21(1):149–161, February 2013. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Ding:2020:CMM

[DYJ20] Kai Ding, Homayoun Yousefi’zadeh, and Faryar Jabbari. Connectivity maintenance in mobile networks. IEEE/ ACM Transactions on Networking, 28(3):1269–1282, June 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2020.2979522.

Ding:2023:NDO

[DYJ^{+}23] Yi Ding, Yu Yang, Wenchao Jiang, Yunhuai Liu, Tian He, and Desheng Zhang. Nationwide deployment and operation of a virtual arrival detection system in the wild. IEEE/ACM Transactions on Networking, 31(2):574– 589, April 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2022.3196806.

Dong:2016:ART

[DYW^{+}16] Wei Dong, Jie Yu, Jiliang Wang, Xuefeng Zhang, Yi Gao, Chun Chen, and Jiajun Bu. Accurate and robust time reconstruction for deployed sensor networks. IEEE/ACM Transactions on Networking, 24(4):2372–2385, August 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Ding:2012:MVD

[DYX12] Yong Ding, Yang Yang, and Li Xiao. Multisource video on-demand streaming in wireless mesh networks. IEEE/ACM Transactions on Networking, 20 (6):1800–1813, December 2012. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Diakonikolas:2018:RRS

[DZ18] Jelena Diakonikolas and Gil Zussman. On the rate regions of single-channel and multi-channel fullduplex links. IEEE/ACM Transactions on Networking, 26(1):47–60, February 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Ding:2020:MPP

[DZ20] Kemi Ding and Junshan Zhang. Multi-party privacy conflict management in online social networks: a network game perspective. IEEE/ACM Transactions on Networking, 28(6):2685–2698, December 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2020.3016315.

Duan:2003:SON

[DZH03] Zhenhai Duan, Zhi-Li Zhang, and Yiwei Thomas Hou. Service overlay networks: SLAs, QoS, and bandwidth provisioning. IEEE/ACM Transactions on Networking, 11(6):870– 883, December 2003. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Deng:2019:ORS

[DZH19] Han Deng, Tao Zhao, and I-Hong Hou. Online routing and scheduling with capacity redundancy for timely delivery guarantees in multihop networks.

IEEE/ACM Transactions on Networking, 27(3):1258– 1271, June 2019. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Ding:2018:SBC

[DZL^{+}18] Haichuan Ding, Chi Zhang, Xuanheng Li, Jianqing Liu, Miao Pan, Yuguang Fang, Shigang Chen, Yuguang Fang, Chi Zhang, Miao Pan, Xuanheng Li, Jianqing Liu, Haichuan Ding, and Shigang Chen. Sessionbased cooperation in cognitive radio networks: a network-level approach. IEEE/ACM Transactions on Networking, 26(2):685– 698, April 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Deng:2020:NLE

[DZL^{+}20] Lei Deng, Haifeng Zheng, Xiao-Yang Liu, Xinxin Feng, and Zhizhang David Chen. Network latency estimation with leverage sampling for personal devices: an adaptive tensor completion approach. IEEE/ACM Transactions on Networking, 28(6):2797–2808, December 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.

                                        acm.org/doi/10.1109/TNET.
                                        2020.3022757.

Dinh:2014:CEV

[DZNT14] Thang N. Dinh, Huiyuan Zhang, Dzung T. Nguyen, and My T. Thai. Costeffective viral marketing for time-critical campaigns in large-scale social networks. IEEE/ACM Transactions on Networking, 22(6):2001– 2011, December 2014. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Ekici:2001:DRA

[EAB01] Eylem Ekici, Ian F. Akyildiz, and Michael D. Bender. A distributed routing algorithm for datagram traffic in LEO satellite networks. IEEE/ACM Transactions on Networking, 9 (2):137–147, 2001. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/articles/
                                        journals/ton/2001-9-2/
                                        p137-ekici/p137-ekici.
                                        pdf;   http://www.acm.
                                        org/pubs/citations/journals/
                                        ton/2001-9-2/p137-ekici/

.

Ekici:2002:MRA

[EAB02] Eylem Ekici, Ian F. Akyildiz, and Michael D. Bender. A multicast routing algorithm for LEO satellite IP networks. IEEE/ACM Transactions on Networking, 10(2):183–192, April 2002. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Elgabli:2018:LRR

[EAH^{+}18] Anis Elgabli, Vaneet Aggarwal, Shuai Hao, Feng Qian, and Subhabrata Sen. LBP: Robust rate adaptation algorithm for SVC video streaming. IEEE/ ACM Transactions on Networking, 26(4):1633–1645, August 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Engelmann:2018:EPR

[EBJM18] Anna Engelmann, Wolfgang Bziuk, Admela Jukan, and Muriel Medard. Exploiting parallelism with random linear network coding in high-speed Ethernet systems. IEEE/ACM Transactions on Networking, 26(6):2829–2842, December 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

E:2020:HTH

[ECL^{+}20] Jinlong E., Yong Cui, Zhenhua Li, Mingkang Ruan, and Ennan Zhai. HyCloud: Tweaking hybrid cloud storage services for cost-efficient filesystem hosting. IEEE/ACM Transactions on Networking, 28(6):2629–2642, December 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2020.3019571.

Elmeleegy:2009:UME

[ECN09] Khaled Elmeleegy, Alan L. Cox, and T. S. Eugene Ng. Understanding and mitigating the effects of count to infinity in Ethernet networks. IEEE/ ACM Transactions on Networking, 17(1):186–199, February 2009. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Eriksson:2012:ENT

[EDBN12] Brian Eriksson, Gautam Dasarathy, Paul Barford, and Robert Nowak. Efficient network tomography for Internet topology discovery. IEEE/ACM Transactions on Networking, 20 (3):931–943, June 2012. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Esposito:2016:DVN

[EDM16] Flavio Esposito, Donato Di Paola, and Ibrahim Matta. On distributed virtual network embedding with guarantees. IEEE/ACM Transactions on Networking, 24(1):569–582, February 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

ElAzzouni:2018:NBD

[EE18] Sherif ElAzzouni and Eylem Ekici. Node-based distributed channel access with enhanced delay characteristics. IEEE/ACM Transactions on Networking, 26(3):1474–1487, June 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Elayoubi:2008:PEA

[EF08] Salah-Eddine Elayoubi and Benoît Fourestié. Performance evaluation of admission control and adaptive modulation in OFDMA WiMax systems. IEEE/ ACM Transactions on Networking, 16(5):1200–1211, October 2008. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Einziger:2017:TAE

[EF17] Gil Einziger and Roy Friedman. TinySet — an access efficient self adjusting Bloom filter construction. IEEE/ACM Transactions on Networking, 25(4): 2295–2307, August 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Elgabli:2019:GPA

[EFA19] Anis Elgabli, Muhamad Felemban, and Vaneet Aggarwal. GroupCast: Preference-aware cooperative video streaming with scalable video coding. IEEE/ACM Transactions on Networking, 27(3):1138– 1150, June 2019. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Einziger:2018:IBI

[EFFK18] Gil Einziger, Benny Fellman, Roy Friedman, and Yaron Kassner. ICE buckets: Improved counter estimation for network measurement. IEEE/ACM Transactions on Networking, 26(3):1165–1178, June 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Eriksson:2007:DDA

[EFK07] Jakob Eriksson, Michalis Faloutsos, and Srikanth V. Krishnamurthy. DART: dynamic address routing for scalable ad hoc and mesh networks. IEEE/ ACM Transactions on Networking, 15(1):119–132, February 2007. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Rouayheb:2016:SDC

[EGKM16] Salim El Rouayheb, Sreechakra Goparaju, Han Mao Kiah, and Olgica Milenkovic. Synchronization and deduplication in coded distributed storage networks. IEEE/ACM Transactions on Networking, 24(5):3056– 3069, October 2016. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Elhourani:2016:IFR

[EGR^{+}16] Theodore Elhourani, Abishek Gopalan, Srinivasan Ramasubramanian, Theodore Elhourani, Abishek Gopalan, and Srinivasan Ramasubramanian. IP fast rerouting for multi-link failures. IEEE/ACM Transactions on Networking, 24(5):3014– 3025, October 2016. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Elhawary:2011:EEP

[EH11] Mohamed Elhawary and Zygmunt J. Haas. Energyefficient protocol for cooperative networks. IEEE/ ACM Transactions on Networking, 19(2):561–574, April 2011. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Eshete:2014:TBC

[EJ14] Addisu Tadesse Eshete and Yuming Jiang. On the transient behavior of CHOKe. IEEE/ACM Transactions on Networking, 22(3):875– 888, June 2014. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Elmokashfi:2012:BCE

[EKD12] Ahmed Elmokashfi, Amund Kvalbein, and Constantine Dovrolis. BGP churn evolution: a perspective from the core. IEEE/ACM Transactions on Networking, 20 (2):571–584, April 2012. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Eshghi:2016:OPC

[EKSV16] Soheil Eshghi, M. H. R. Khouzani, Saswati Sarkar, and Santosh S. Venkatesh. Optimal patching in clustered malware epidemics. IEEE/ACM Transactions on Networking, 24(1):283– 298, February 2016. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

ElRakabawy:2011:PAP

[EL11] Sherif M. ElRakabawy and Christoph Lindemann. A practical adaptive pacing scheme for TCP in multihop wireless networks.

IEEE/ACM Transactions on Networking, 19(4):975– 988, August 2011. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Esmat:2024:SLM

[EL24] Haitham H. Esmat and Beatriz Lorenzo. Selflearning multi-mode slicing mechanism for dynamic network architectures. IEEE/ACM Transactions on Networking, 32 (2):1048–1063, April 2024. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL

https://dl.acm.org/doi/
10.1109/TNET.2023.3305975.

Elwalid:1993:EBG

[EM93] Anwar I. Elwalid and Debasis Mitra. Effective bandwidth of general Markovian traffic sources and admission control of high speed networks. IEEE/ ACM Transactions on Networking, 1(3):329–343, June 1993. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1993-1-3/
p329-elwalid/.

Eshoul:2009:SAU

[EM09] Abdelhamid E. Eshoul and Hussein T. Mouftah. Survivability approaches using p-cycles in WDM mesh networks under static traffic. IEEE/ACM Transactions on Networking, 17 (2):671–683, April 2009. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Eramo:2017:ASF

[EMAL17] Vincenzo Eramo, Emanuele Miucci, Mostafa Ammar, and Francesco Giacinto Lavacca. An approach for service function chain routing and virtual function network instance migration in network function virtualization architectures. IEEE/ACM Transactions on Networking, 25(4):2008– 2025, August 2017. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Eswaran:2012:CTU

[EML12] Sharanya Eswaran, Archan Misra, and Thomas F. La Porta. Control-theoretic utility maximization in multihop wireless networks under mission dynamics. IEEE/ACM Transactions on Networking, 20(4):1082– 1095, August 2012. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

ElGamal:2006:OTD

[EMPS06] Abbas El Gamal, James Mammen, Balaji Prabhakar, and Devavrat Shah. Optimal throughput-delay scaling in wireless networks: part I: the fluid model. IEEE/ACM Transactions on Networking, 14 (SI):2568–2592, June 2006. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Etezadi:2024:PFO

[ENT^{+}24] Ehsan Etezadi, Carlos Natalino, Christine Tremblay, Lena Wosinska, and Marija Furdek. Programmable filterless optical networks: Architecture, design, and resource allocation. IEEE/ ACM Transactions on Networking, 32(2):1096–1109, April 2024. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2023.3319746.

Erramilli:1996:EQA

[ENW96] Ashok Erramilli, Onuttom Narayan, and Walter Willinger. Experimental queueing analysis with long-range dependent packet traffic. IEEE/ACM Transactions on Networking, 4(2):209– 223, April 1996. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1996-4-2/
p209-erramilli/.

Eryilmaz:2010:DCL

[EOSM10] Atilla Eryilmaz, Asuman Ozdaglar, Devavrat Shah, and Eytan Modiano. Distributed cross-layer algorithms for the optimal control of multihop wireless networks. IEEE/ACM Transactions on Networking, 18(2):638–651, April 2010. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Esposito:2014:RTE

[EPB14] Christian Esposito, Marco Platania, and Roberto Beraldi. Reliable and timely event notification for publish/subscribe services over the Internet. IEEE/ACM Transactions on Networking, 22(1):230– 243, February 2014. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Escobar:1994:FSP

[EPD94] Julio Escobar, Craig Partridge, and Debra Deutsch. Flow synchronization protocol. IEEE/ACM Transactions on Networking, 2 (2):111–121, April 1994. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL

                                        http://www.acm.org/pubs/
                                        citations/journals/ton/
                                        1994-2-2/p111-escobar/.

Ejaz:2021:HAI

[EPS21] Ahsen Ejaz, Vassilis Papaefstathiou, and Ioannis Sourdis. HighwayNoC: Approaching ideal NoC performance with dual data rate routers. IEEE/ ACM Transactions on Networking, 29(1):318–331, February 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2020.3034581.

Alaoui:2020:MMA

[ER20] Sara El Alaoui and Byrav Ramamurthy. MARS: a multi-attribute routing and scheduling algorithm for DTN interplanetary networks. IEEE/ ACM Transactions on Networking, 28(5):2065–2076, October 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2020.3008630.

Elsayed:2023:TLC

[ER23] Karim Elsayed and Amr Rizk. Time-to-live caching with network delays: Exact analysis and computable approximations. IEEE/ ACM Transactions on Networking, 31(3):1087–1100, June 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2022.3207914.

Elbaum:1996:TDL

[ES96] Reuven Elbaum and Moshe Sidi. Topological design of local-area networks using genetic algorithms. IEEE/ ACM Transactions on Networking, 4(5):766–778, October 1996. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/
                                        journals/ton/1996-4-5/
                                        p766-elbaum/.

Eun:2003:MAA

[ES03] Do Young Eun and Ness B. Shroff. A measurementanalytic approach for QoS estimation in a network based on the dominant time scale. IEEE/ACM Transactions on Networking, 11 (2):222–235, April 2003. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Eun:2005:NDT

[ES05] Do Young Eun and Ness B. Shroff. Network decomposition: theory and practice. IEEE/ACM Transactions on Networking, 13(3): 526–539, June 2005. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Eryilmaz:2007:FRA

[ES07] Atilla Eryilmaz and R. Srikant.

Fair resource allocation in wireless networks using queue-length-based scheduling and congestion control. IEEE/ACM Transactions on Networking, 15(6): 1333–1344, December 2007. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

ElRouayheb:2011:RNC

[ESG11] Salim El Rouayheb, Alex Sprintson, and Costas Georghiades. Robust network codes for unicast connections: a case study. IEEE/ACM Transactions on Networking, 19(3):644– 656, June 2011. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Eryilmaz:2005:SSP

[ESP05] Atilla Eryilmaz, R. Srikant, and James R. Perkins. Stable scheduling policies for fading wireless channels. IEEE/ACM Transactions on Networking, 13(2):411– 424, April 2005. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Estrin:1993:PRE

[EST93] Deborah Estrin, Martha Steenstrup, and Gene Tsudik. A protocol for route establishment and packet forwarding across multidomain Internets. IEEE/ACM Transactions on Networking, 1(1):56–70, February 1993. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1993-1-1/
p56-estrin/.

Estan:2006:BAC

[EVF06] Cristian Estan, George Varghese, and Michael Fisk. Bitmap algorithms for counting active flows on high-speed links. IEEE/ ACM Transactions on Networking, 14(5):925–937, October 2006. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Eun:2008:ATT

[EW08] Do Young Eun and Xinbing Wang. Achieving 100% throughput in TCP/AQM under aggressive packet marking with small buffer. IEEE/ACM Transactions on Networking, 16(4):945– 956, August 2008. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Francois:2012:FCP

[FAB12] Jérôme François, Issam Aib, and Raouf Boutaba.

FireCol: a collaborative protection network for the detection of flooding DDoS attacks. IEEE/ACM Transactions on Networking, 20(6):1828–1841, December 2012. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Fontugne:2017:SIT

[FAF^{+}17] Romain Fontugne, Patrice Abry, Kensuke Fukuda, Darryl Veitch, Kenjiro Cho, Pierre Borgnat, and Herwig Wendt. Scaling in Internet traffic: a 14 year and 3 day longitudinal study, with multiscale analyses and random projections. IEEE/ACM Transactions on Networking, 25(4):2152– 2165, August 2017. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Fang:2005:MPA

[Fan05] Yuguang Fang. Modeling and performance analysis for wireless mobile networks: a new analytical approach. IEEE/ ACM Transactions on Networking, 13(5):989–1002, October 2005. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Farkas:1995:CWS

[Far95]  Peter Farkas̆.  Com-

ments on “Weighted sum codes for error detection and their comparison with existing codes”. IEEE/ACM Transactions on Networking, 3(2):222– 223, April 1995. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1995-3-2/
p222-farkas/. See [McA94].

Fu:2023:OPC

[FAS^{+}23] Yongquan Fu, Lun An, Siqi Shen, Kai Chen, and Pere Barlet-Ros. A one-pass clustering based sketch method for network monitoring. IEEE/ ACM Transactions on Networking, 31(6):2604– 2613, 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2023.3251981.

Feng:2022:BIR

[FAWW22] Cuiying Feng, Jianwei An, Kui Wu, and Jianping Wang. Bound inference and reinforcement learningbased path construction in bandwidth tomography. IEEE/ACM Transactions on Networking, 30(2):501– 514, April 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.

acm.org/doi/10.1109/TNET.
2021.3118006.

Francois:2007:ATL

[FB07] Pierre François and Olivier Bonaventure. Avoiding transient loops during the convergence of link-state routing protocols. IEEE/ ACM Transactions on Networking, 15(6):1280–1292, December 2007. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Fu:2017:MSA

[FBFB17] Yongquan Fu, Ernst Biersack, Yongquan Fu, and Ernst Biersack. MCR: Structure-aware overlaybased latency-optimal greedy relay search. IEEE/ACM Transactions on Networking, 25(5):3016–3029, October 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Farkiani:2021:PDD

[FBM^{+}21] Behrooz Farkiani, Bahador Bakhshi, S. Ali MirHassani, Tim Wauters, Bruno Volckaert, and Filip De Turck. Prioritized deployment of dynamic service function chains. IEEE/ ACM Transactions on Networking, 29(3):979–993, June 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2021.3055074.

Faltelli:2023:MAP

[FBQ^{+}23] Marco Faltelli, Giacomo Belocchi, Francesco Quaglia, Salvatore Pontarelli, and Giuseppe Bianchi. Metronome: Adaptive and precise intermittent packet retrieval in DPDK. IEEE/ACM Transactions on Networking, 31(3):979–993, June 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL

https://dl.acm.org/doi/
10.1109/TNET.2022.3208799.

Fu:2018:ETC

[FBRL18] Yongquan Fu, Pere Barlet-Ros, and Dongsheng Li. Every timestamp counts: Accurate tracking of network latencies using reconcilable difference aggregator. IEEE/ACM Transactions on Networking, 26 (1):90–103, February 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Felstaine:1999:DRC

[FC99] Eyal Felstaine and Reuven Cohen. On the distribution of routing computation in hierarchical ATM networks. IEEE/ ACM Transactions on Networking, 7(6):906–916, December 1999. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1999-7-6/
p906-felstaine/.

FischereSilva:2017:EEE

[FC17] Renan Fischer e Silva and Paul M. Carpenter. Energy efficient Ethernet on MapReduce clusters: Packet coalescing to improve 10GbE links. IEEE/ ACM Transactions on Networking, 25(5):2731–2742, October 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Fan:2006:TTS

[FCA^{+}06] Xingzhe Fan, Kartikeya Chandrayana, Murat Arcak, Shivkumar Kalyanaraman, and John Ting-Yung Wen. A two-time-scale design for edge-based detection and rectification of uncooperative flows. IEEE/ ACM Transactions on Networking, 14(6):1313–1322, December 2006. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Fan:2000:SCS

[FCAB00] Li Fan, Pei Cao, Jussara Almeida, and Andrei Z. Broder. Summary cache: a scalable wide-area Web cache sharing protocol. IEEE/ACM Transactions on Networking, 8(3): 281–293, 2000. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/
                                        journals/ton/2000-8-3/
                                        p281-fan/.

Fang:1997:MPN

[FCL97] Yuguang Fang, Imrich Chlamtac, and Yi-Bing Lin. Modeling PCS networks under general call holding time and cell residence time distributions. IEEE/ACM Transactions on Networking, 5(6):893–906, December 1997. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/
                                        journals/ton/1997-5-6/
                                        p893-fang/.

Fumagalli:2003:ODS

[FCT03] Andrea Fumagalli, Isabella Cerutti, and Marco Tacca. Optimal design of survivable mesh networks based on line switched WDM selfhealing rings. IEEE/ACM Transactions on Networking, 11(3):501–512, June 2003. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Ficara:2010:ECB

[FDG^{+}10] Domenico Ficara, Andrea Di Pietro, Stefano Giordano, Gregorio Procissi, and Fabio Vitucci. Enhancing counting bloom filters through Huffmancoded multilayer structures. IEEE/ACM Transactions on Networking, 18 (6):1977–1987, December 2010. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Ficara:2011:DED

[FDG^{+}11] Domenico Ficara, Andrea Di Pietro, Stefano Giordano, Gregorio Procissi, Fabio Vitucci, and Gianni Antichi. Differential encoding of DFAs for fast regular expression matching. IEEE/ACM Transactions on Networking, 19(3): 683–694, June 2011. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Fu:2017:DAS

[FDM^{+}17] Tom Z. J. Fu, Jianbing Ding, Richard T. B. Ma, Marianne Winslett, Yin Yang, and Zhenjie Zhang. DRS: Auto-scaling for real-time stream analytics. IEEE/ACM Transactions on Networking, 25(6): 3338–3352, December 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Filippini:2013:NOR

[FEC13] Ilario Filippini, Eylem Ekici, and Matteo Cesana. A new outlook on routing in cognitive radio networks: minimummaintenance-cost routing. IEEE/ACM Transactions on Networking, 21(5):1484– 1498, October 2013. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Feldmeier:1995:FSI

[Fel95] David C. Feldmeier. Fast software implementation of error detection codes. IEEE/ACM Transactions on Networking, 3(6):640– 651, December 1995. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL http://

                                        www.acm.org/pubs/citations/
                                        journals/ton/1995-3-6/
                                        p640-feldmeier/.

Floyd:1999:PUE

[FF99] Sally Floyd and Kevin Fall. Promoting the use of end-to-end congestion control in the Internet. IEEE/ACM Transactions on Networking, 7(4):458– 472, August 1999. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/
                                        journals/ton/1999-7-4/
                                        p458-floyd/.

Fu:2017:DSD

[FFX^{+}17] Luoyi Fu, Xinzhe Fu, Zhiying Xu, Qianyang Peng, Xinbing Wang, and Songwu Lu. Determining source–destination connectivity in uncertain networks: Modeling and solutions. IEEE/ACM Transactions on Networking, 25 (6):3237–3252, December 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Fu:2018:JOM

[FFZ^{+}18] Luoyi Fu, Xinzhe Fu, Zesen Zhang, Zhiying Xu, Xudong Wu, Xinbing Wang, and Songwu Lu. Joint optimization of multicast energy in delayconstrained mobile wireless networks. IEEE/ ACM Transactions on Networking, 26(1):633–646, February 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Fashandi:2010:PDP

[FGK10] Shervan Fashandi, Shahab Oveis Gharan, and Amir K. Khandani. Path diversity over packet switched networks: performance analysis and rate allocation. IEEE/ACM Transactions on Networking, 18(5): 1373–1386, October 2010. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Feldmann:2001:DTD

[FGL^{+}01] Anja Feldmann, Albert Greenberg, Carsten Lund, Nick Reingold, Jennifer Rexford, and Fred True. Deriving traffic demands for operational IP networks: methodology and experience. IEEE/ACM Transactions on Networking, 9(3):265–280, June 2001. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Fogue:2013:ASB

[FGM^{+}13] Manuel Fogue, Piedad Garrido, Francisco J. Martinez, Juan-Carlos Cano, Carlos T. Calafate, and Pietro Manzoni. An adaptive system based on roadmap profiling to enhance warning message dissemination in VANETs. IEEE/ACM Transactions on Networking, 21(3):883–895, June 2013. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Fan:2017:SSP

[FGR^{+}17] Jingyuan Fan, Chaowen Guan, Kui Ren, Yong Cui, and Chunming Qiao. SPABox: Safeguarding privacy during deep packet inspection at a Middle-Box. IEEE/ACM Transactions on Networking, 25(6): 3753–3766, December 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Fan:2018:MBP

[FGRQ18] Jingyuan Fan, Chaowen Guan, Kui Ren, and Chunming Qiao. Middleboxbased packet-level redundancy elimination over encrypted network traffic. IEEE/ACM Transactions on Networking, 26(4):1742– 1753, August 2018. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Fan:2010:PSN

[FHH10] Chun-I Fan, Pei-Hsiu Ho, and Ruei-Hau Hsu. Provably secure nested onetime secret mechanisms for fast mutual authentication and key exchange in mobile communications. IEEE/ACM Transactions on Networking, 18(3):996– 1009, June 2010. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Fossati:2018:FRA

[FHMS18] Francesca Fossati, Sahar Hoteit, Stefano Moretti, and Stefano Secci. Fair resource allocation in systems with complete information sharing. IEEE/ ACM Transactions on Networking, 26(6):2801–2814, December 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Fukuda:2017:DMA

[FHQ^{+}17] Kensuke Fukuda, John Heidemann, Abdul Qadeer, Kensuke Fukuda, John Heidemann, and Abdul Qadeer. Detecting malicious activity with DNS backscatter over time. IEEE/ACM Transactions on Networking, 25(5):3203– 3218, October 2017. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Freris:2013:DAS

[FHSZ13] Nikolaos M. Freris, Cheng-Hsin Hsu, Jatinder Pal Singh, and Xiaoqing Zhu. Distortion-aware scalable video streaming to multinetwork clients. IEEE/ ACM Transactions on Networking, 21(2):469–481, April 2013. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Fay:2010:WSD

[FHT^{+}10] Damien Fay, Hamed Haddadi, Andrew Thomason, Andrew W. Moore, Richard Mortier, Almerima Jamakovic, Steve Uhlig, and Miguel Rio. Weighted spectral distribution for Internet topology analysis: theory and applications. IEEE/ACM Transactions on Networking, 18(1):164– 176, February 2010. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Floyd:1993:RED

[FJ93] Sally Floyd and Van Jacobson. Random early detection gateways for congestion avoidance. IEEE/ ACM Transactions on Networking, 1(4):397–413, August 1993. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1993-1-4/
p397-floyd/.

Floyd:1994:SPR

[FJ94] Sally Floyd and Van Jacobson. The synchronization of periodic routing messages. IEEE/ACM Transactions on Networking, 2(2):122– 136, April 1994. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1994-2-2/
p122-floyd/.

Floyd:1995:LRM

[FJ95] Sally Floyd and Van Jacobson. Link-sharing and resource management models for packet networks. IEEE/ACM Transactions on Networking, 3(4):365– 386, August 1995. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/
                                        journals/ton/1995-3-4/
                                        p365-floyd/.

Fayoumi:2007:SMB

[FJ07] Ayman G. Fayoumi and Anura P. Jayasumana. A surjective-mapping based model for optical sharedbuffer cross-connect. IEEE/ ACM Transactions on Networking, 15(1):226–233, February 2007. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Feamster:2007:IAE

[FJB07] Nick Feamster, Ramesh Johari, and Hari Balakrishnan. Implications of autonomy for the expressiveness of policy routing. IEEE/ ACM Transactions on Networking, 15(6):1266–1279, December 2007. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Francis:2001:IGI

[FJJ^{+}01] Paul Francis, Sugih Jamin, Cheng Jin, Yixin Jin, Danny Raz, Yuval Shavitt, and Lixia Zhang. IDMaps: a global Internet host distance estimation service. IEEE/ACM Transactions on Networking, 9(5):525– 540, October 2001. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Floyd:1997:RMF

[FJL^{+}97] Sally Floyd, Van Jacobson, Ching-Gung Liu, Steven McCanne, and Lixia Zhang. A reliable multicast framework for light-weight sessions and application level framing. IEEE/ ACM Transactions on Networking, 5(6):784–803, December 1997. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1997-5-6/
p784-floyd/.

Feng:1999:APM

[FK99] Wu-Chang Feng and Dilip D.

Kandlur. Adaptive packet marking for maintaining end-to-end throughput in a differentiatedservices Internet. IEEE/ ACM Transactions on Networking, 7(5):685–697, October 1999. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1999-7-5/
p685-feng/.

Fu:2003:SMR

[FK03] Huirong Fu and Edward W.

Knightly. A simple model of real-time flow aggregation. IEEE/ACM Transactions on Networking, 11(3): 422–435, June 2003. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Fahmy:2007:COM

[FK07] Sonia Fahmy and Minseok Kwon. Characterizing overlay multicast networks and their costs. IEEE/ACM Transactions on Networking, 15(2):373–386, April 2007. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Fu:2013:SGW

[FK13] Fangwen Fu and Ulas C. Kozat. Stochastic game for wireless network virtualization. IEEE/ACM Transactions on Networking, 21 (1):84–97, February 2013. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Ferlin:2018:MMF

[FKCA18] Simone Ferlin, Stepan Kucera, Holger Claussen, and Ozgu Alay. MPTCP meets FEC: Supporting latency-sensitive applications over heterogeneous networks. IEEE/ACM Transactions on Networking, 26(5):2005–2018, October 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Firoiu:1998:EAC

[FKT98] Victor Firoiu, Jim Kurose, and Don Towsley. Efficient admission control of piecewise linear traffic envelopes at EDF schedulers. IEEE/ ACM Transactions on Networking, 6(5):558–570, October 1998. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1998-6-5/
p558-firoiu/.

Fu:2019:SAM

[FLBR^{+}19] Yongquan Fu, Dongsheng Li, Pere Barlet-Ros, Chun Huang, Zhen Huang, Siqi Shen, and Huayou Su. A skewness-aware matrix factorization approach for mesh-structured cloud services. IEEE/ACM Transactions on Networking, 27(4):1598–1611, August 2019. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Fan:2009:DTO

[FLC09] Bin Fan, John C. S.

Lui, and Dah-Ming Chiu. The design trade-offs of BitTorrent-like file sharing protocols. IEEE/ACM Transactions on Networking, 17(2):365–376, April 2009. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Fan:2020:SSM

[FLG^{+}20] Yuqi Fan, Wenlong Liu, Dan Guo, Weili Wu, and Dingzhu Du. Shuffle scheduling for MapReduce jobs based on periodic network status. IEEE/ ACM Transactions on Networking, 28(4):1832–1844, August 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2020.2993945.

Fu:2023:SFF

[FLG^{+}23] Pengtao Fu, Lailong Luo, Deke Guo, Shangsen Li, and Yun Zhou. A shifting filter framework for dynamic set queries. IEEE/ ACM Transactions on Networking, 31(5):2329– 2344, 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2023.3247628.

Fiessler:2017:HEH

[FLH^{+}17] Andreas Fiessler, Claas Lorenz, Sven Hager, Bjorn Scheuermann, and Andrew W. Moore. HyPaFilter+: Enhanced hybrid packet filtering using hardware assisted classification and header space analysis. IEEE/ACM Transactions on Networking, 25(6): 3655–3669, December 2017.

CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Fang:2022:TAR

[FLM^{+}22] Chongrong Fang, Haoyu Liu, Mao Miao, Jie Ye, Lei Wang, Wansheng Zhang, Daxiang Kang, Biao Lyu, Shunmin Zhu, Peng Cheng, and Jiming Chen. Towards automatic root cause diagnosis of persistent packet loss in cloud overlay network. IEEE/ ACM Transactions on Networking, 30(3):1178–1192, June 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2021.3137557.

Forestiero:2010:SCB

[FLMM10] Agostino Forestiero, Emilio Leonardi, Carlo Mastroianni, and Michela Meo. Self-chord: a bio-inspired P2P framework for selforganizing distributed systems. IEEE/ACM Transactions on Networking, 18(5):1651–1664, October 2010. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Foerster:2018:LFR

[FLMS18] Klaus-Tycho Foerster, Arne Ludwig, Jan Marcinkowski, and Stefan Schmid. Loopfree route updates for software-defined networks. IEEE/ACM Transactions on Networking, 26(1):328– 341, February 2018. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Feng:2022:PTH

[FLS^{+}22] Xuewei Feng, Qi Li, Kun Sun, Chuanpu Fu, and Ke Xu. Off-path TCP hijacking attacks via the side channel of downgraded IPID. IEEE/ ACM Transactions on Networking, 30(1):409–422, February 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2021.3115517.

Feng:2018:ODC

[FLTM18] Hao Feng, Jaime Llorca, Antonia M. Tulino, and Andreas F. Molisch. Optimal dynamic cloud network control. IEEE/ ACM Transactions on Networking, 26(5):2118–2131, October 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Fu:2023:MPS

[FLZ^{+}23] Songtao Fu, Qi Li, Min Zhu, Xiaoliang Wang, Su Yao, Yangfei Guo, Xinle Du, and Ke Xu. MASK: Practical source and path verification based on Multi-AS-Key. IEEE/ ACM Transactions on Networking, 31(4):1478– 1493, 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2022.3222610.

Franceschetti:2006:CNL

[FM06] Massimo Franceschetti and Ronald Meester. Critical node lifetimes in random networks via the Chen-Stein method. IEEE/ ACM Transactions on Networking, 14(SI):2831–2837, June 2006. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Fu:2020:EEJ

[FM20] Jing Fu and Bill Moran.

Energy-efficient job-assignment policy with asymptotically guaranteed performance deviation. IEEE/ ACM Transactions on Networking, 28(3):1325–1338, June 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2020.2983460.

Fu:2022:LNN

[FM22] Xinzhe Fu and Eytan Modiano. Learning-NUM: Network utility maximization with unknown utility functions and queueing delay. IEEE/ACM Transactions on Networking, 30(6): 2788–2803, 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2022.3182890.

Fu:2023:ORP

[FM23] Xinzhe Fu and Eytan Modiano. Optimal routing to parallel servers with unknown utilities — multi– armed bandit with queues. IEEE/ACM Transactions on Networking, 31(5):1997– 2012, 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2022.3227136.

Fernandes:2020:USR

[FMCS20] Ramon Fernandes, César Marcon, Rodrigo Cataldo, and Johanna Sepúlveda. Using smart routing for secure and dependable NoCbased MPSoCs. IEEE/ ACM Transactions on Networking, 28(3):1158–1171, June 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2020.2979372.

Fan:2021:RWB

[FMH^{+}21a] Fujie Fan, Hangyu Meng, Bing Hu, Kwan L. Yeung, and Zhifeng Zhao. Roulette wheel balancing algorithm with dynamic flowlet switching for multipath datacenter networks. IEEE/ACM Transactions on Networking, 29(2):834– 847, April 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2021.3051995.

Farhadi:2021:SPR

[FMH^{+}21b] Vajiheh Farhadi, Fidan Mehmeti, Ting He, Thomas F. La Porta, Hana Khamfroush, Shiqiang Wang, Kevin S. Chan, and Konstantinos Poularakis. Service placement and request scheduling for dataintensive applications in edge clouds. IEEE/ ACM Transactions on Networking, 29(2):779–792, April 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2020.3048613.

Fu:2018:FFD

[FMK^{+}18] Luoyi Fu, Songjun Ma, Lingkun Kong, Shiyu Liang, and Xinbing Wang. FINE: a framework for distributed learning on incomplete observations for heterogeneous crowdsensing networks. IEEE/ACM Transactions on Networking, 26(3):1092–1109, June 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Froc:2009:DPW

[FML09] Gwillerm Froc, Issam Mabrouki, and Xavier Lagrange. Design and performance of wireless data gathering networks based on unicast random walk routing. IEEE/ ACM Transactions on Networking, 17(4):1214–1227, August 2009. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Felemban:2023:VVQ

[FML23] Noor Felemban, Fidan Mehmeti, and Thomas F. La Porta. VidQ: Video query using optimized audio-visual processing. IEEE/ACM Transactions on Networking, 31(3):1338– 1352, June 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2022.3215601.

Ferrel:2006:VTM

[FMMLH06] Ian Ferrel, Adrian Mettler, Edward Miller, and Ran Libeskind-Hadas. Virtual topologies for multicasting with multiple originators in WDM networks. IEEE/ACM Transactions on Networking, 14(1):183– 190, February 2006. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Finamore:2010:KSP

[FMMR10] Alessandro Finamore, Marco Mellia, Michela Meo, and Dario Rossi. KISS: stochastic packet inspection classifier for UDP traffic. IEEE/ ACM Transactions on Networking, 18(5):1505–1515, October 2010. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Fossati:2020:MRA

[FMPS20] Francesca Fossati, Stefano Moretti, Patrice Perny, and Stefano Secci. Multiresource allocation for network slicing. IEEE/ ACM Transactions on Networking, 28(3):1311–1324, June 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2020.2979667.

Flammini:2011:CRP

[FMSM^{+}11] Michele Flammini, Alberto Marchetti-Spaccamela, Gianpiero Monaco, Luca Moscardelli, and Shmuel Zaks. On the complexity of the regenerator placement problem in optical networks. IEEE/ACM Transactions on Networking, 19 (2):498–511, April 2011. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Faisal:2024:JJT

[FMSM24] Tooba Faisal, Damiano Di Francesco Maesa, Nishanth Sastry, and Simone Mangiante. JITRA: Justin-time resource allocation through the distributed ledgers for 5G and beyond. IEEE/ACM Transactions on Networking, 32(2):1201– 1211, April 2024. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2023.3318239.

Fu:2003:OEA

[FMT03] Alvin C. Fu, Eytan Modiano, and John N. Tsitsiklis. Optimal energy allocation and admission control for communications satellites. IEEE/ACM Transactions on Networking, 11(3): 488–500, June 2003. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Frey:2000:GBF

[FNQ00] Michael Frey and Son Nguyen-Quang. A gammabased framework for modeling variable-rate MPEG video sources: the GOP GBAR model. IEEE/ACM Transactions on Networking, 8(6):710–719, 2000.

CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL

http://www.acm.org/pubs/
articles/journals/ton/
2000-8-6/p710-frey/p710-
frey.pdf;   http://www.
acm.org/pubs/citations/
journals/ton/2000-8-6/
p710-frey/.

Figueira:1995:UBD

[FP95] Norival R. Figueira and Joseph Pasquale. An upper bound on delay for the VirtualClock service discipline. IEEE/ACM Transactions on Networking, 3(4):399–408, August 1995. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1995-3-4/
p399-figueira/.

Figueira:1997:SCD

[FP97] Norival R. Figueira and Joseph Pasquale. A schedulability condition for deadline-ordered service disciplines. IEEE/ ACM Transactions on Networking, 5(2):232–244, April 1997. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1997-5-2/
p232-figueira/.

Floyd:2001:DSI

[FP01] Sally Floyd and Vern Paxson. Difficulties in simulating the Internet. IEEE/ACM Transactions on Networking, 9(4):392– 403, August 2001. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Ferragut:2014:NRA

[FP14] Andrés Ferragut and Fernando Paganini. Network resource allocation for users with multiple connections: fairness and stability. IEEE/ACM Transactions on Networking, 22 (2):349–362, April 2014. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Fulton:1998:DJF

[FqL98] Cathy A. Fulton and San qi Li. Delay jitter firstorder and second-order statistical functions of general traffic on high-speed multimedia networks. IEEE/ ACM Transactions on Networking, 6(2):150–163, April 1998. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/
                                        journals/ton/1998-6-2/
                                        p150-fulton/.

Feamster:2007:NWP

[FR07] Nick Feamster and Jennifer Rexford. Networkwide prediction of BGP routes. IEEE/ACM Transactions on Networking, 15 (2):253–266, April 2007. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Feldmann:1998:EPC

[FRC98] Anja Feldmann, Jennifer Rexford, and Ramón Cáceres. Efficient policies for carrying Web traffic over flowswitched networks. IEEE/ ACM Transactions on Networking, 6(6):673–685, December 1998. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1998-6-6/
p673-feldmann/.

Fernandez:2017:EDR

[FS17] Alvaro Fernandez and Norvald Stol. Economic, dissatisfaction, and reputation risks of hardware and software failures in PONs. IEEE/ACM Transactions on Networking, 25(2):1119– 1132, April 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Ferrari:2017:CRP

[FSGH17] Lorenzo Ferrari, Anna Scaglione, Reinhard Gentz, and Yao-Win Peter Hong. Convergence results on pulse coupled oscillator protocols in locally connected networks. IEEE/ ACM Transactions on Networking, 25(2):1004–1019, April 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Fukushima:2013:MDR

[FSH^{+}13] Masaki Fukushima, Kohei Sugiyama, Teruyuki Hasegawa, Toru Hasegawa, and Akihiro Nakao. Minimum disclosure routing for network virtualization and its experimental evaluation. IEEE/ACM Transactions on Networking, 21(6): 1839–1851, December 2013. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Fu:2014:DRA

[FSM14] Amy Fu, Parastoo Sadeghi, and Muriel Médard. Dynamic rate adaptation for improved throughput and delay in wireless network coded broadcast. IEEE/ ACM Transactions on Networking, 22(6):1715–1728, December 2014. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Fuerst:2018:KOE

[FSSC18] Carlo Fuerst, Stefan Schmid, Lalith Suresh, and Paolo Costa. Kraken: Online and elastic resource reservations for cloud datacenters. IEEE/ACM Transactions on Networking, 26(1): 422–435, February 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Fu:2009:OPS

[FST^{+}09] Xiaoming Fu, Henning Schulzrinne, Hannes Tschofenig, Christian Dickmann, and Dieter Hogrefe. Overhead and performance study of the General Internet Signaling Transport (GIST) protocol. IEEE/ ACM Transactions on Networking, 17(1):158–171, February 2009. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Fumagalli:2006:DRD

[FT06] Andrea Fumagalli and Marco Tacca. Differentiated reliability (DiR) in wavelength division multiplexing rings. IEEE/ ACM Transactions on Networking, 14(1):159–168, February 2006. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Fodor:2007:BBP

[FT07] Gábor Fodor and Miklós Telek. Bounding the blocking probabilities in multirate CDMA networks supporting elastic services. IEEE/ACM Transactions on Networking, 15(4):944– 956, August 2007. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Tsaur:2006:CLA

[fTL06] Lih feng Tsaur and Daniel C. Lee. Closed-loop architecture and protocols for rapid dynamic spreading gain adaptation in CDMA networks. IEEE/ACM Transactions on Networking, 14 (4):821–834, August 2006. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Fadlullah:2010:DCA

[FTV^{+}10] Zubair M. Fadlullah, Tarik Taleb, Athanasios V. Vasilakos, Mohsen Guizani, and Nei Kato. DTRAB: combating against attacks on encrypted protocols through traffic-feature analysis. IEEE/ACM Transactions on Networking, 18(4):1234–1247, August 2010. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Fang:2013:FGC

[FTZ^{+}13] Ji Fang, Kun Tan, Yuanyang Zhang, Shouyuan Chen, Lixin Shi, Jiansong Zhang, Yongguang Zhang, and Zhenhui Tan. Fine-grained channel access in wireless LAN. IEEE/ACM Transactions on Networking, 21 (3):772–787, June 2013. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Fecko:2003:TGF

[FUDA03] Mariusz A. Fecko, M. Ümit Uyar, Ali Y. Duale, and Paul D. Amer. A technique to generate feasible tests for communications systems with multiple timers. IEEE/ACM Transactions on Networking, 11 (5):796–809, October 2003. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Fukunaga:2020:AAF

[Fuk20] Takuro Fukunaga. Adaptive algorithm for finding connected dominating sets in uncertain graphs. IEEE/ ACM Transactions on Networking, 28(1):387–398, February 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL https://

dl.acm.org/doi/abs/10.
1109/TNET.2019.2963361.

Fu:2017:WCO

[FWK17] Luoyi Fu, Xinbing Wang, and P. R. Kumar. Are we connected? optimal determination of source– destination connectivity in random networks. IEEE/ ACM Transactions on Networking, 25(2):751–764, April 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Fragouli:2008:EBU

[FWL08] Christina Fragouli, Jörg Widmer, and Jean-Yves Le Boudec. Efficient broadcasting using network coding. IEEE/ACM Transactions on Networking, 16 (2):450–463, April 2008. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Fu:2022:DTB

[FWN^{+}22] Junsong Fu, Na Wang, Leyao Nie, Baojiang Cui, and Bharat K. Bhargava. Defending traceback attack in 3D wireless Internet of Things. IEEE/ACM Transactions on Networking, 30(4):1765– 1779, 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2022.3149293.

Fu:2017:SSD

[FX17] Yongquan Fu and Xu Xiaoping. Self-stabilized distributed network distance prediction. IEEE/ ACM Transactions on Networking, 25(1):451–464, February 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Fan:2021:RTU

[FXHY21] Xingpeng Fan, Hongli Xu, He Huang, and Xuwei Yang. Real-time update of joint SFC and routing in software defined networks. IEEE/ACM Transactions on Networking, 29(6):2664–2677, December 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2021.3095935.

Fu:2021:STD

[FXQ^{+}21] Luoyi Fu, Jiasheng Xu, Shan Qu, Zhiying Xu, Xinbing Wang, and Guihai Chen. Seeking the truth in a decentralized manner. IEEE/ACM Transactions on Networking, 29(5):2296–2312, October 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2021.3085000.

Fei:2007:PTR

[FY07] Zongming Fei and Mengkun Yang. A proactive tree recovery mechanism for resilient overlay multicast. IEEE/ACM Transactions on Networking, 15(1):173– 186, February 2007. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Fabini:2016:RTR

[FZ16] Joachim Fabini and Tanja Zseby. The right time: Reducing effective end-to-end delay in time-slotted packet-switched networks. IEEE/ACM Transactions on Networking, 24(4):2251– 2263, August 2016. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Fu:2022:MSN

[FZQ^{+}22] Luoyi Fu, Jiapeng Zhang, Shan Qu, Huquan Kang, Xinbing Wang, and Guihai Chen. Measuring social network Deanonymizability by means of morphism property. IEEE/ACM Transactions on Networking, 30(6):2744– 2759, 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2022.3180158.

Fu:2020:ASN

[FZW^{+}20] Luoyi Fu, Jiapeng Zhang, Shuaiqi Wang, Xinyu Wu, Xinbing Wang, and Guihai Chen. De-anonymizing social networks with overlapping community structure.

IEEE/ACM Transactions on Networking, 28(1):360– 375, February 2020. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL https:

//dl.acm.org/doi/abs/
10.1109/TNET.2019.2962731.

Fang:2023:GGR

[FZX^{+}23] Jin Fang, Gongming Zhao, Hongli Xu, Changbo Wu, and Zhuolong Yu. GRID: Gradient routing with innetwork aggregation for distributed training. IEEE/ ACM Transactions on Networking, 31(5):2267– 2280, 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2023.3244794.

Ganguly:2024:OFL

[GA24] Bhargav Ganguly and Vaneet Aggarwal. Online federated learning via nonstationary detection and adaptation amidst concept drift. IEEE/ACM Transactions on Networking, 32 (1):643–653, 2024. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2023.3294366.

Gungor:2008:RTR

[GAA08] Vehbi Cagri Gungor, Özgür B.

Akan, and Ian F. Akyildiz. A real-time and reliable transport (RT)^{2} protocol for wireless sensor and actor networks. IEEE/ACM Transactions on Networking, 16(2):359–370, April 2008. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Ganesan:2020:PGD

[Gan20] Ashwin Ganesan. Performance guarantees of distributed algorithms for QoS in wireless ad hoc networks. IEEE/ACM Transactions on Networking, 28(1):182–195, February 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL https://

                                        dl.acm.org/doi/abs/10.
                                        1109/TNET.2019.2959797.

Gao:2001:IAS

[Gao01]  Lixin Gao.  On infer-

ring autonomous system relationships in the Internet. IEEE/ACM Transactions on Networking, 9(6): 733–745, December 2001. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Grossglauser:1999:RLR

[GB99] Matthias Grossglauser and Jean-Chrysostome Bolot. On the relevance of longrange dependence in network traffic. IEEE/ACM Transactions on Networking, 7(5):629–640, October 1999. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1999-7-5/
p629-grossglauser/.

Gobjuka:2010:ETD

[GB10] Hassan Gobjuka and Yuri J.

Breitbart. Ethernet topology discovery for networks with incomplete information. IEEE/ACM Transactions on Networking, 18(4): 1220–1233, August 2010. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Ge:2018:HTF

[GB18] Mengyao Ge and Douglas M. Blough. High throughput and fair scheduling for multi-AP multiuser MIMO in dense wireless networks. IEEE/ ACM Transactions on Networking, 26(5):2414–2427, October 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Gutekunst:1995:DPG

[GBC^{+}95] Thomas Gutekunst, Daniel Bauer, Germano Caronni, Bernhard Plattner, and Hasan. A distributed and policy-free general-purpose shared window system. IEEE/ACM Transactions on Networking, 3(1):51–62, February 1995. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/
                                        journals/ton/1995-3-1/
                                        p51-gutekunst/.

Gupta:2016:LWF

[GBG^{+}16] Varun Gupta, Yigal Bejerano, Craig Gutterman, Jaime Ferragut, Katherine Guo, Thyaga Nandagopal, and Gil Zussman. Lightweight feedback mechanism for WiFi multicast to very large groups — experimental evaluation. IEEE/ACM Transactions on Networking, 24(6):3826–3840, December 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Guo:2012:EPL

[GBL12] Danhua Guo, Laxmi Narayan Bhuyan, and Bin Liu. An efficient parallelized L7filter design for multicore servers. IEEE/ACM Transactions on Networking, 20(5):1426–1439, October 2012. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Giovanidis:2021:ROS

[GBMV21] Anastasios Giovanidis, Bruno Baynat, Clémence Magnien, and Antoine Vendeville. Ranking online social users by their influence. IEEE/ACM Transactions on Networking, 29(5):2198–2214, October 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2021.3085201.

Gao:2023:DDA

[GCD23] Shuqin Gao, Costas A.

Courcoubetis, and Lingjie Duan. Distributed double auction mechanisms for large-scale device-to-device resource trading. IEEE/ACM Transactions on Networking, 31(3):1308– 1323, June 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2022.3218552.

Godfrey:2015:SRS

[GCH^{+}15] P. Brighten Godfrey, Matthew Caesar, Ian Haken, Yaron Singer, Scott Shenker, and Ion Stoica. Stabilizing route selection in BGP. IEEE/ACM Transactions on Networking, 23(1):282– 299, February 2015. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Gonzalez:2016:AEG

[GCM^{+}16] Roberto Gonzalez, Ruben Cuevas, Reza Motamedi, Reza Rejaie, and Angel Cuevas. Assessing the evolution of Google+ in its first two years. IEEE/ACM Transactions on Networking, 24(3):1813–1826, June 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Guan:2020:DJP

[GCMP20] Zhangyu Guan, Nan Cen, Tommaso Melodia, and Scott M. Pudlewski. Distributed joint power, association and flight control for massive-MIMO selforganizing flying drones. IEEE/ACM Transactions on Networking, 28(4):1491– 1505, August 2020. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1109/
                                        TNET.2020.2985972.

Gurewitz:2006:NCT

[GCS06a] Omer Gurewitz, Israel Cidon, and Moshe Sidi. Network classless time protocol based on clock offset optimization. IEEE/ACM Transactions on Networking, 14(4):876–888, August 2006. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Gurewitz:2006:OWD

[GCS06b] Omer Gurewitz, Israel Cidon, and Moshe Sidi. One-way delay estimation using network-wide measurements. IEEE/ ACM Transactions on Networking, 14(SI):2710–2724, June 2006. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Guo:2021:MFD

[GCW21] Jianxiong Guo, Tiantian Chen, and Weili Wu. A multi-feature diffusion model: Rumor blocking in social networks. IEEE/ ACM Transactions on Networking, 29(1):386–397, February 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2020.3032893.

Gao:2017:EEA

[GCWC17] Xiaofeng Gao, Zhiyin Chen, Fan Wu, and Guihai Chen. Energy efficient algorithms for k-sink minimum movement target coverage problem in mobile sensor network. IEEE/ ACM Transactions on Networking, 25(6):3616–3627, December 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Gong:2017:SGU

[GCX^{+}17] Xiaowen Gong, Xu Chen, Kai Xing, Dong-Hoon Shin, Mengyuan Zhang, and Junshan Zhang. From social group utility maximization to personalized location privacy in mobile networks. IEEE/ACM Transactions on Networking, 25 (3):1703–1716, June 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Gerstel:1996:LVP

[GCZ96] Ornan Gerstel, Israel Cidon, and Shmuel Zaks. The layout of virtual paths in ATM networks. IEEE/ ACM Transactions on Networking, 4(6):873–884, December 1996. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1996-4-6/
p873-gerstel/.

Gerstel:1998:ESC

[GCZ98] Ornan Ori Gerstel, Israel Cidon, and Shmuel Zaks. Efficient support for client/server applications over heterogeneous ATM network. IEEE/ ACM Transactions on Networking, 6(4):432–446, August 1998. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1998-6-4/
p432-gerstel/.

Gulcu:2018:AVP

[GCZY18] Talha Cihad Gulcu, Vaggos Chatziafratis, Yingrui Zhang, and Osman Yagan. Attack vulnerability of power systems under an equal load redistribution model. IEEE/ACM Transactions on Networking, 26 (3):1306–1319, June 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Gao:2016:IPI

[GDC^{+}16] Yi Gao, Wei Dong, Chun Chen, Jiajun Bu, Wenbin Wu, and Xue Liu. iPath: path inference in wireless sensor networks. IEEE/ACM Transactions on Networking, 24(1):517– 528, February 2016. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Gao:2017:APP

[GDC^{+}17] Yi Gao, Wei Dong, Chun Chen, Xiaoyu Zhang, Jiajun Bu, and Xue Liu. Accurate per-packet delay tomography in wireless ad hoc networks. IEEE/ ACM Transactions on Networking, 25(1):480–491, February 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Guo:2024:TIP

[GDJX24] Zehua Guo, Songshi Dou, Wenchao Jiang, and Yuanqing Xia. Toward improved path programmability recovery for softwaredefined WANs under multiple controller failures. IEEE/ACM Transactions on Networking, 32(1):143– 158, 2024. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2023.3286456.

Guo:2022:MCR

[GDL^{+}22] Zehua Guo, Songshi Dou, Sen Liu, Wendi Feng, Wenchao Jiang, Yang Xu, and Zhi-Li Zhang. Maintaining control resiliency and flow programmability in software-defined WANs during controller failures. IEEE/ACM Transactions on Networking, 30(3):969– 984, June 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2021.3128771.

Gao:2016:SSP

[GDW^{+}16] Yi Gao, Wei Dong, Wenbin Wu, Chun Chen, Xiang-Yang Li, and Jiajun Bu. Scalpel: scalable preferential link tomography based on graph trimming. IEEE/ACM Transactions on Networking, 24(3):1392– 1403, June 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Guo:2023:TFP

[GDWX23] Zehua Guo, Songshi Dou, Wenfei Wu, and Yuanqing Xia. Toward flexible and predictable path programmability recovery under multiple controller failures in softwaredefined WANs. IEEE/ ACM Transactions on Networking, 31(5):1965– 1980, 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2022.3227423.

Gouda:2002:HIC

[GEHM02] Mohamed G. Gouda, E. N. (Mootaz)

Elnozahy, Chin-Tser Huang, and Tommy M. McGuire. Hop integrity in computer networks. IEEE/ACM Transactions on Networking, 10(3):308–319, June 2002. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Georgakopoulos:2008:BCB

[Geo08] George F. Georgakopoulos. Buffered cross-bar switches, revisited: design steps, proofs and simulations towards optimal rate and minimum buffer memory. IEEE/ACM Transactions on Networking, 16(6): 1340–1351, December 2008. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Gupta:1995:RPR

[GF95] Amit Gupta and Domenico Ferrari. Resource partitioning for real-time communication. IEEE/ACM Transactions on Networking, 3(5):501–508, October 1995. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/
                                        journals/ton/1995-3-5/
                                        p501-gupta/.

Gao:2018:AAS

[GFW^{+}18] Xiaofeng Gao, Jiahao Fan, Fan Wu, Guihai Chen, Jiahao Fan, Guihai Chen, Xiaofeng Gao, and Fan Wu. Approximation algorithms for sweep coverage problem with multiple mobile sensors. IEEE/ACM Transactions on Networking, 26 (2):990–1003, April 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Gopal:1994:NTP

[GG94] Inder Gopal and Roch Guérin. Network transparency: the plaNET approach. IEEE/ACM Transactions on Networking, 2 (3):226–239, June 1994. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL

http://www.acm.org/pubs/
citations/journals/ton/
1994-2-3/p226-gopal/.

Georgiadis:1993:TPF

[GGC93] Leonidas Georgiadis, Roch Guérin, and Israel Cidon. Throughput properties of fair policies in ring networks. IEEE/ACM Transactions on Networking, 1(6):718–728, December 1993. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL

http://www.acm.org/pubs/
citations/journals/ton/
1993-1-6/p718-georgiadis/

.

Georgiadis:2002:LOB

[GGFS02] Leonidas Georgiadis, Panos Georgatsos, Konstantinos Floros, and Stelios Sartzetakis. Lexicographically optimal balanced networks. IEEE/ACM Transactions on Networking, 10(6):818– 829, December 2002. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Giacomelli:2011:OPG

[GGH11] Riccardo Giacomelli, Radha Krishna Ganti, and Martin Haenggi. Outage probability of general ad hoc networks in the high-reliability regime. IEEE/ACM Transactions on Networking, 19(4):1151–1163, August 2011. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Gopal:1999:FBH

[GGK99] Ajei Gopal, Inder Gopal, and Shay Kutten. Fast broadcast in high-speed networks. IEEE/ACM Transactions on Networking, 7(2):262–275, April 1999. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL

                                        http://www.acm.org/pubs/
                                        citations/journals/ton/
                                        1999-7-2/p262-gopal/.

Garetto:2009:CSAb

[GGL09a] Michele Garetto, Paolo Giaccone, and Emilio Leonardi. Capacity scaling in ad hoc networks with heterogeneous mobile nodes: the subcritical regime. IEEE/ACM Transactions on Networking, 17(6):1888–1901, December 2009. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Garetto:2009:CSAa

[GGL09b] Michele Garetto, Paolo Giaccone, and Emilio Leonardi. Capacity scaling in ad hoc networks with heterogeneous mobile nodes: the supercritical regime. IEEE/ ACM Transactions on Networking, 17(5):1522–1535, October 2009. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Gupta:2010:SPS

[GGM10] Ashima Gupta, Debalina Ghosh, and Prasant Mohapatra. Scheduling prioritized services in multihop OFDMA networks. IEEE/ ACM Transactions on Networking, 18(6):1780–1792, December 2010. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Giorgetti:2011:ULR

[GGM11] Gianni Giorgetti, Sandeep Kumar S. Gupta, and Gianfranco Manes. Understanding the limits of RFbased collaborative localization. IEEE/ACM Transactions on Networking, 19 (6):1638–1651, December 2011. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Georgiadis:1996:ENQ

[GGPS96] Leonidas Georgiadis, Roch Guérin, Vinod Peris, and Kumar N. Sivarajan. Efficient network QoS provisioning based on per node traffic shaping. IEEE/ ACM Transactions on Networking, 4(4):482–501, August 1996. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/
                                        journals/ton/1996-4-4/
                                        p482-georgiadis/.

Gao:2019:UTU

[GGZC19] Xiaofeng Gao, Yuanning Gao, Yichen Zhu, and Guihai Chen. U2-Tree: a universal two-layer distributed indexing scheme for cloud storage system. IEEE/ACM Transactions on Networking, 27(1):201– 213, February 2019. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Ghanbari:1993:PCV

[GH93] Mohammad Ghanbari and Charles J. Hughes. Packing coded video signals into ATM cells. IEEE/ ACM Transactions on Networking, 1(5):505–509, October 1993. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/
                                        journals/ton/1993-1-5/
                                        p505-ghanbari/.

Gau:2004:CSM

[GH04] Rung-Hung Gau and Zygmunt J. Haas. Concurrent search of mobile users in cellular networks. IEEE/ACM Transactions on Networking, 12(1):117– 130, February 2004. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Gardner:2017:BMJ

[GHBSWV17] Kristen Gardner, Mor Harchol-Balter, Alan Scheller-Wolf, and Benny Van Houdt. A better model for job redundancy: Decoupling server slowdown and job size. IEEE/ACM Transactions on Networking, 25(6):3353–3367, December 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Gau:2002:MFC

[GHK02] Rung-Hung Gau, Zygmunt J. Haas, and Bhaskar Krishnamachari. On multicast flow control for heterogeneous receivers. IEEE/ACM Transactions on Networking, 10(1):86– 101, February 2002. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Ghasempour:2018:DBS

[GHK18]  Yasaman  Ghasempour,

Muhammad Kumail Haider, and Edward W. Knightly. Decoupling beam steering and user selection for MU– MIMO 60-GHz WLANs. IEEE/ACM Transactions on Networking, 26(5):2390– 2403, October 2018. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Ganguly:2023:MES

[GHK^{+}23] Bhargav Ganguly, Seyyedali Hosseinalipour, Kwang Taik Kim, Christopher G. Brinton, Vaneet Aggarwal, David J. Love, and Mung Chiang. Multi-edge serverassisted dynamic federated learning with an optimized floating aggregation point. IEEE/ACM Transactions on Networking, 31(6):2682– 2697, 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2023.3262482.

Gajic:2014:CWP

[GHR14] Vojislav Gajić, Jianwei Huang, and Bixio Rimoldi. Competition of wireless providers for atomic users. IEEE/ACM Transactions on Networking, 22(2):512– 525, April 2014. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

G:2018:SMP

[GHRH18] David Gonzalez G., Harri Hakula, Antti Rasila, and Jyri Hamalainen. Spatial mappings for planning and optimization of cellular networks. IEEE/ ACM Transactions on Networking, 26(1):175–188, February 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Goussevskaia:2014:AWC

[GHW14] Olga Goussevskaia, Magnús M.

Halldórsson, and Roger Wattenhofer. Algorithms for wireless capacity. IEEE/ ACM Transactions on Networking, 22(3):745–755, June 2014. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Gopalam:2022:DLS

[GHW22] Swaroop Gopalam, Stephen V.

Hanly, and Philip Whiting. Distributed and local scheduling algorithms for mmWave integrated access and backhaul. IEEE/ ACM Transactions on Networking, 30(4):1749– 1764, 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2022.3154367.

Guo:2020:WCT

[GHZ20a] Xiuzhen Guo, Yuan He, and Xiaolong Zheng. WiZig: Cross-technology energy communication over a noisy channel. IEEE/ACM Transactions on Networking, 28(6):2449–2460, December 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2020.3013921.

Guo:2020:ZHC

[GHZ^{+}20b] Xiuzhen Guo, Yuan He, Xiaolong Zheng, Liangcheng Yu, and Omprakash Gnawali. ZigFi: Harnessing channel state information for cross-technology communication. IEEE/ACM Transactions on Networking, 28(1):301–311, February 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL https://

                                        dl.acm.org/doi/abs/10.
                                        1109/TNET.2019.2962707.

Ghosh:2011:MSS

[GIKK11] Amitabha Ghosh, Özlem Durmaz Incel, V. S. Anil Kumar, and Bhaskar Krishnamachari. Multichannel scheduling and spanning trees: throughputdelay tradeoff for fast data collection in sensor networks. IEEE/ACM Transactions on Networking, 19 (6):1731–1744, December 2011. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Gregori:2015:NMA

[GIL^{+}15] Enrico Gregori, Alessandro Improta, Luciano Lenzini, Lorenzo Rossi, and Luca Sani. A novel methodology to address the Internet AS-level data incompleteness. IEEE/ACM Transactions on Networking, 23(4): 1314–1327, August 2015. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Gallardo:2018:SAC

[GJCB18] Guillaume Artero Gallardo, Gentian Jakllari, Lucile Canourgues, and Andre-Luc Beylot. Statistical admission control in multi-hop cognitive radio networks. IEEE/ACM Transactions on Networking, 26(3):1390–1403, June 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Gao:2018:UUR

[GJD18] Yi Gao, Yuan Jing, and Wei Dong. UniROPE: Universal and robust packet trajectory tracing for softwaredefined networks. IEEE/ ACM Transactions on Networking, 26(6):2515–2527, December 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Gorinsky:2006:DMP

[GJVZ06] Sergey Gorinsky, Sugat Jain, Harrick Vin, and Yongguang Zhang. Design of multicast protocols robust against inflated subscription. IEEE/ACM Transactions on Networking, 14(2):249–262, April 2006. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Gong:2016:NLC

[GJWZ16] Long Gong, Huihui Jiang, Yixiang Wang, and Zuqing Zhu. Novel locationconstrained virtual network embedding LC-VNE algorithms towards integrated node and link mapping. IEEE/ACM Transactions on Networking, 24(6):3648– 3661, December 2016. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Gong:2016:QIL

[GK16] Xun Gong and Negar Kiyavash. Quantifying the information leakage in timing side channels in deterministic work-conserving schedulers. IEEE/ACM Transactions on Networking, 24(3):1841–1852, June 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Garlapati:2016:SSM

[GKB^{+}16] Shravan Garlapati, Teja Kuruganti, Michael R. Buehrer, Jeffrey H. Reed, Shravan Garlapati, Teja Kuruganti, Michael R. Buehrer, and Jeffrey H. Reed. SMAC: a soft MAC to reduce control overhead and latency in CDMA-based AMI networks. IEEE/ACM Transactions on Networking, 24(5):2648–2662, October 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Gopal:2021:CAT

[GKCR21] Sneihil Gopal, Sanjit K.

Kaul, Rakesh Chaturvedi, and Sumit Roy. Coexistence of age and throughput optimizing networks: a spectrum sharing game. IEEE/ACM Transactions on Networking, 29(4):1494–1508, August 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2021.3067900.

Gai:2012:CNO

[GKJ12] Yi Gai, Bhaskar Krishnamachari, and Rahul Jain. Combinatorial network optimization with unknown variables: multiarmed bandits with linear rewards and individual observations. IEEE/ ACM Transactions on Networking, 20(5):1466–1478, October 2012. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Ganesh:2006:CNP

[GKPS06] Ayalvadi J. Ganesh, Peter B. Key, Damien Polis, and R. Srikant. Congestion notification and probing mechanisms for endpoint admission control. IEEE/ACM Transactions on Networking, 14 (3):568–578, 2006. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Ganjali:2005:CSV

[GKS05] Yashar Ganjali, Abtin Keshavarzian, and Devavrat Shah. Cell switching versus packet switching in inputqueued switches. IEEE/ ACM Transactions on Networking, 13(4):782–789, August 2005. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Gibbens:1993:DRM

[GKT93] Richard J. Gibbens, Frank P. Kelly, and Stephen R. E. Turner. Dynamic routing in multiparented networks. IEEE/ACM Transactions on Networking, 1(2):261– 270, April 1993. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/
                                        journals/ton/1993-1-2/
                                        p261-gibbens/.

Grossglauser:1997:RSE

[GKT97]  Matthias  Grossglauser,

Srinivasan Keshav, and David N. C. Tse. RCBR: a simple and efficient service for multiple timescale traffic. IEEE/ACM Transactions on Networking, 5(6):741–755, December 1997. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1997-5-6/
p741-grossglauser/.

Girard:1993:DAR

[GL93] André Girard and Bernard Liau. Dimensioning of adaptively routed networks. IEEE/ACM Transactions on Networking, 1 (4):460–468, August 1993. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL

http://www.acm.org/pubs/
citations/journals/ton/
1993-1-4/p460-girard/.

Guo:2010:DAM

[GL10] Song Guo and Victor C. M.

Leung. A distributed algorithm for min-max tree and max-min cut problems in communication networks. IEEE/ACM Transactions on Networking, 18(4):1067– 1076, August 2010. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Garcia-Lunes-Aceves:1993:LRU

[GLA93] J. J. Garcia-Lunes-Aceves. Loop-free routing using diffusing computations. IEEE/ACM Transactions on Networking, 1(1):130– 141, February 1993. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL http://

                                        www.acm.org/pubs/citations/
                                        journals/ton/1993-1-1/
                                        p130-garcia-lunes-aceves/

.

Garrido:2019:JSC

[GLA19] Pablo Garrido, Douglas J. Leith, and Ramon Aguero. Joint scheduling and coding for low in-order delivery delay over lossy paths with delayed feedback. IEEE/ ACM Transactions on Networking, 27(5):1987–2000, October 2019. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Garcia-Luna-Aceves:1997:PAL

[GLAM97] J. J. Garcia-Luna-Aceves and Shree Murthy. A path-finding algorithm for loop-free routing. IEEE/ ACM Transactions on Networking, 5(1):148–160, February 1997. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/
                                        journals/ton/1997-5-1/

p148-garcia-luna-aceves/

.

Garcia-Luna-Aceves:2011:PID

[GLAMM11] J. J. Garcia-Luna-Aceves and Rolando Menchaca-Mendez. PRIME: an interest-driven approach to integrated unicast and multicast routing in MANETs. IEEE/ACM Transactions on Networking, 19(6):1573– 1586, December 2011. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Gong:2016:FCC

[GLC^{+}16] Wei Gong, Haoxiang Liu, Lei Chen, Kebin Liu, and Yunhao Liu. Fast composite counting in RFID systems. IEEE/ACM Transactions on Networking, 24(5):2756–2767, October 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Gong:2004:CDW

[GLG04] Yongtao Gong, Peiyuan Lee, and Wanyi Gu. Comments on “Dynamic wavelength routing using congestion and neighborhood information”. IEEE/ACM Transactions on Networking, 12(3):572, June 2004. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). See [LS99].

Ghosal:1995:PAP

[GLH95] Dipak Ghosal, T. V. Lakshman, and Yennun Huang. Parallel architectures for processing high speed network signaling protocols. IEEE/ACM Transactions on Networking, 3(6):716–728, December 1995. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1995-3-6/
p716-ghosal/.

Gong:2018:CAR

[GLL^{+}18] Wei Gong, Haoxiang Liu, Jiangchuan Liu, Xiaoyi Fan, Kebin Liu, Qiang Ma, Xiaoyu Ji, Kebin Liu, Jiangchuan Liu, Xiaoyi Fan, Qiang Ma, Haoxiang Liu, Wei Gong, and Xiaoyu Ji. Channel-aware rate adaptation for backscatter networks. IEEE/ACM Transactions on Networking, 26(2):751–764, April 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Guo:2016:FNB

[GLLJ16] Jian Guo, Fangming Liu, John C. S. Lui, and Hai Jin. Fair network bandwidth allocation in IaaS datacenters via a cooperative game approach. IEEE/ACM Transactions on Networking, 24 (2):873–886, April 2016.

CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Gong:2017:TMR

[GLLL17] Wei Gong, Jiangchuan Liu, Kebin Liu, and Yunhao Liu. Toward more rigorous and practical cardinality estimation for largescale RFID systems. IEEE/ ACM Transactions on Networking, 25(3):1347–1358, June 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Gong:2016:FAC

[GLM^{+}16] Wei Gong, Haoxiang Liu, Xin Miao, Kebin Liu, Wenbo He, Lan Zhang, and Yunhao Liu. Fast and adaptive continuous scanning in large-scale RFID systems. IEEE/ACM Transactions on Networking, 24(6):3314–3325, December 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Garetto:2004:CQN

[GLMM04] Michele Garetto, Renato Lo Cigno, Michela Meo, and Marco Ajmone Marsan. Closed queueing network models of interacting longlived TCP flows. IEEE/ ACM Transactions on Networking, 12(2):300–311, April 2004. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Gerla:2001:RBS

[GLNP01] Mario Gerla, Emilio Leonardi, Fabio Neri, and Prasasth Palnati. Routing in the bidirectional shufflenet. IEEE/ACM Transactions on Networking, 9(1):91– 103, 2001. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/articles/
                                        journals/ton/2001-9-1/
                                        p91-gerla/p91-gerla.pdf;
                                         http://www.acm.org/pubs/
                                        citations/journals/ton/
                                        2001-9-1/p91-gerla/.

Gupta:2009:LCD

[GLS09] Abhinav Gupta, Xiaojun Lin, and R. Srikant. Lowcomplexity distributed scheduling algorithms for wireless networks. IEEE/ACM Transactions on Networking, 17(6):1846–1859, December 2009. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). See comments and corrections [ZCW15].

Gu:2021:AOO

[GLS21] Yan Gu, Bo Liu, and Xiaojun Shen. Asymptotically optimal online scheduling with arbitrary hard deadlines in multi-hop communication networks. IEEE/ ACM Transactions on Networking, 29(4):1452–1466, August 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2021.3065703.

Guven:2008:UFM

[GLSB08] Tuna Güven, Richard J.

La, Mark A. Shayman, and Bobby Bhattacharjee. A unified framework for multipath routing for unicast and multicast traffic. IEEE/ACM Transactions on Networking, 16(5): 1038–1051, October 2008. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Gong:2017:EUT

[GLY17] Wei Gong, Jiangchuan Liu, and Zhe Yang. Efficient unknown tag detection in large-scale RFID systems with unreliable channels. IEEE/ACM Transactions on Networking, 25(4):2528– 2539, August 2017. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Gao:2012:SAM

[GLZC12] Wei Gao, Qinghua Li, Bo Zhao, and Guohong Cao. Social-aware multicast in disruption-tolerant networks. IEEE/ACM Transactions on Networking, 20(5):1553–1566, October 2012. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Ghani:2000:EDE

[GM00] Nasir Ghani and Jon W. Mark. Enhanced distributed explicit rate allocation for ABR services in ATM networks. IEEE/ACM Transactions on Networking, 8(1):71–86, February 2000. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/
                                        journals/ton/2000-8-1/
                                        p71-ghani/.

Gencata:2003:VTA

[GM03] Aysegül Gençata and Biswanath Mukherjee. Virtualtopology adaptation for WDM mesh networks under dynamic traffic. IEEE/ ACM Transactions on Networking, 11(2):236–247, April 2003. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Gomaa:2015:HCP

[GMD15] Hazem Gomaa, Geoffrey G. Messier, and Robert Davies. Hierarchical cache performance analysis under TTL-based consistency. IEEE/ACM Transactions on Networking, 23(4):1190– 1201, August 2015. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Giustiniano:2010:MTO

[GMLP10] Domenico Giustiniano, David Malone, Douglas J. Leith, and Konstantina Papagiannaki. Measuring transmission opportunities in 802.11 links. IEEE/ACM Transactions on Networking, 18(5):1516–1529, October 2010. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Gandhi:2008:MBL

[GMP08] Rajiv Gandhi, Arunesh Mishra, and Srinivasan Parthasarathy. Minimizing broadcast latency and redundancy in ad hoc networks. IEEE/ACM Transactions on Networking, 16 (4):840–851, August 2008. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Galluccio:2013:GMG

[GMP13] Laura Galluccio, Giacomo Morabito, and Sergio Palazzo. GEographic Multicast (GEM) for dense wireless networks: protocol design and performance analysis. IEEE/ ACM Transactions on Networking, 21(4):1332–1346, August 2013. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Guan:2016:TTD

[GMS16] Zhangyu Guan, Tommaso Melodia, and Gesualdo Scutari. To transmit or not to transmit?: distributed queueing games in infrastructureless wireless networks. IEEE/ACM Transactions on Networking, 24 (2):1153–1166, April 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Gurewitz:2009:MMO

[GMSK09] Omer Gurewitz, Vincenzo Mancuso, Jingpu Shi, and Edward W. Knightly. Measurement and modeling of the origins of starvation of congestion-controlled flows in wireless mesh networks. IEEE/ACM Transactions on Networking, 17(6):1832– 1845, December 2009. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Gomaa:2013:EIC

[GMWD13] Hazem Gomaa, Geoffrey G. Messier, Carey Williamson, and Robert Davies. Estimating instantaneous cache hit ratio using Markov chain analysis. IEEE/ ACM Transactions on Networking, 21(5):1472–1483, October 2013. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Guan:2013:JOR

[GMY13] Zhangyu Guan, Tommaso Melodia, and Dongfeng Yuan. Jointly optimal rate control and relay selection for cooperative wireless video streaming. IEEE/ ACM Transactions on Networking, 21(4):1173–1186, August 2013. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Guan:2016:DRM

[GMYP16] Zhangyu Guan, Tommaso Melodia, Dongfeng Yuan, and Dimitris A. Pados. Distributed resource management for cognitive ad hoc networks with cooperative relays. IEEE/ACM Transactions on Networking, 24 (3):1675–1689, June 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Gatmir-Motahari:2013:TCB

[GMZR13]  Sara  Gatmir-Motahari,

Hui Zang, and Phyllis Reuther. Time-clustering-based place prediction for wireless subscribers. IEEE/ ACM Transactions on Networking, 21(5):1436–1446, October 2013. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Gu:2017:WFB

[GND17] Fei Gu, Jianwei Niu, and Lingjie Duan. WAIPO: a fusion-based collaborative indoor localization system on Smartphones. IEEE/ ACM Transactions on Networking, 25(4):2267–2280, August 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Giotsas:2021:PWA

[GNK^{+}21] Vasileios Giotsas, George Nomikos, Vasileios Kotronis, Pavlos Sermpezis, Petros Gigis, Lefteris Manassakis, Christoph Dietzel, Stavros Konstantaras, and Xenofontas Dimitropoulos. O peer, where art thou? Uncovering remote peering interconnections at IXPs. IEEE/ACM Transactions on Networking, 29(1):1–16, February 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2020.3025945.

Gelal:2013:TCE

[GNP^{+}13] Ece Gelal, Jianxia Ning, Konstantinos Pelechrinis, Tae-Suk Kim, Ioannis Broustis, Srikanth V. Krishnamurthy, and Bhaskar D. Rao. Topology control for effective interference cancellation in multiuser MIMO networks.

IEEE/ACM Transactions on Networking, 21(2):455– 468, April 2013. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Guerin:1999:QRN

[GO99] Roche A. Guérin and Ariel Orda. QoS routing in networks with inaccurate information: theory and algorithms. IEEE/ ACM Transactions on Networking, 7(3):350–364, June 1999. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1999-7-3/
p350-guerin/.

Guerin:2002:CSP

[GO02] Roch Guérin and Ariel Orda. Computing shortest paths for any number of hops. IEEE/ACM Transactions on Networking, 10 (5):613–620, October 2002. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Goodrich:2008:PPM

[Goo08] Michael T. Goodrich. Probabilistic packet marking for large-scale IP traceback. IEEE/ACM Transactions on Networking, 16 (1):15–24, February 2008. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Gianatti:1994:PAA

[GP94] Stefano Gianatti and Achille Pattavina. Performance analysis of ATM Banyan networks with shared queueing — part I: random offered traffic. IEEE/ ACM Transactions on Networking, 2(4):398–410, August 1994. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1994-2-4/
p398-gianatti/.

Gaiti:1996:PMI

[GP96a] Dominique Gaïti and Guy Pujolle. Performance management issues in ATM networks: traffic and congestion control. IEEE/ ACM Transactions on Networking, 4(2):249–257, April 1996. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1996-4-2/
p249-gaiti/.

Gong:1996:AEC

[GP96b] Fengmin Gong and Gurudatta M. Parulkar. An application-oriented error control scheme for highspeed networks. IEEE/ ACM Transactions on Networking, 4(5):669–683, October 1996. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1996-4-5/
p669-gong/.

Gopalakrishnan:1998:EUP

[GP98] R. Gopalakrishnan and Gurudatta M. Parulkar. Efficient user-space protocol implementations with QoS guarantees using realtime upcalls. IEEE/ACM Transactions on Networking, 6(4):374–388, August 1998. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1998-6-4/
p374-gopalakrishnan/.

Guidolin–Pina:2024:CGB

[GPBL24]  Damien  Guidolin-Pina,

Marc Boyer, and Jean-Yves Le Boudec. Configuration of guard band and offsets in cyclic queuing and forwarding. IEEE/ACM Transactions on Networking, 32(1): 598–612, 2024. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2023.3293050.

Georgiadis:2015:MET

[GPLT15] Leonidas Georgiadis, Georgios S. Paschos, Lavy Libman, and Leandros Tassiulas. Minimal evacuation times and stability. IEEE/ACM Transactions on Networking, 23(3): 931–945, June 2015. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Gummadi:2003:EPS

[GPM03] Krishna Phani Gummadi, Madhavarapu Jnana Pradeep, and C. Siva Ram Murthy. An efficient primary-segmented backup scheme for dependable real-time communication in multihop networks. IEEE/ACM Transactions on Networking, 11 (1):81–94, February 2003. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Gao:2016:ERW

[GQ16] Fengyu Gao and Hongyan Qian. Efficient, real-world token bucket configuration for residential gateways. IEEE/ACM Transactions on Networking, 24(1):462– 475, February 2016. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Gao:2001:SIR

[GR01] Lixin Gao and Jennifer Rexford. Stable Internet routing without global coordination. IEEE/ACM Transactions on Networking, 9(6):681–692, December 2001. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Gopalan:2012:IAL

[GR12] Abishek Gopalan and Srinivasan Ramasubramanian. On identifying additive link metrics using linearly independent cycles and paths. IEEE/ACM Transactions on Networking, 20(3):906– 916, June 2012. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Gopalan:2014:MNL

[GR14] Abishek Gopalan and Srinivasan Ramasubramanian. On the maximum number of linearly independent cycles and paths in a network. IEEE/ACM Transactions on Networking, 22(5):1373– 1388, October 2014. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Gopalan:2016:IFR

[GR16] Abishek Gopalan and Srinivasan Ramasubramanian. IP fast rerouting and disjoint multipath routing with three edge-independent spanning trees. IEEE/ACM Transactions on Networking, 24 (3):1336–1349, June 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Gao:2020:CMS

[GR20a] Lingnan Gao and George N. Rouskas. Congestion minimization for service chain routing problems with path length considerations. IEEE/ACM Transactions on Networking, 28(6):2643–2656, December 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2020.3017792.

Ghasemi:2020:ACC

[GR20b] Hooshang Ghasemi and Aditya Ramamoorthy. Asynchronous coded caching with uncoded prefetching. IEEE/ACM Transactions on Networking, 28(5):2146–2159, October 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2020.3003907.

Goodell:2009:DSP

[GRB09] Geoffrey Goodell, Mema Roussopoulos, and Scott Bradner. A directory service for perspective access networks. IEEE/ACM Transactions on Networking, 17(2):501–514, April 2009. CODEN IEANEP.

ISSN 1063-6692 (print), 1558-2566 (electronic).

Gulyas:2015:SRP

[GRHA15] András Gulyás, Gábor Rétvári, Zalán Heszberger, and Rachit Agarwal. On the scalability of routing with policies. IEEE/ ACM Transactions on Networking, 23(5):1610–1618, October 2015. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Grover:1999:HAP

[Gro99] Wayne D. Grover. High availability path design in ring-based optimal networks. IEEE/ACM Transactions on Networking, 7 (4):558–574, August 1999. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL

http://www.acm.org/pubs/
citations/journals/ton/
1999-7-4/p558-grover/.

Gerstel:2000:CET

[GRS00] Ornan Gerstel, Rajiv Ramaswami, and Galen H. Sasaki. Cost-effective traffic grooming in WDM rings. IEEE/ACM Transactions on Networking, 8 (5):618–630, 2000. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/

journals/ton/2000-8-5/
p618-gerstel/.

Greenberg:1997:CTA

[GS97] Albert G. Greenberg and R. Srikant. Computational techniques for accurate performance evaluation of multirate, multihop communication networks. IEEE/ACM Transactions on Networking, 5(2):266– 277, April 1997. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1997-5-2/
p266-greenberg/.

Girard:1998:MFM

[GS98] Andre Girard and Brunilde Sansó. Multicommodity flow models, failure propagation, and reliable loss network design. IEEE/ACM Transactions on Networking, 6(1):82–93, February 1998. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1998-6-1/
p82-girard/.

Gouda:2003:MRM

[GS03] Mohamed G. Gouda and Marco Schneider. Maximizable routing metrics. IEEE/ACM Transactions on Networking, 11(4):663– 675, August 2003. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Gurbani:2004:TTS

[GS04] Vijay K. Gurbani and Xian-He Sun. Terminating telephony services on the Internet. IEEE/ACM Transactions on Networking, 12(4):571–581, August 2004. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Gunes:2009:RIA

[GS09] Mehmet H. Gunes and Kamil Sarac. Resolving IP aliases in building traceroute-based Internet maps. IEEE/ACM Transactions on Networking, 17(6):1738–1751, December 2009. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Gerstel:2010:GFS

[GS10a] Ori Gerstel and G. Sasaki.

A general framework for service availability for bandwidth-efficient connectionoriented networks. IEEE/ ACM Transactions on Networking, 18(3):985–995, June 2010. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Gupta:2010:DAW

[GS10b] Gagan Raj Gupta and Ness B. Shroff. Delay analysis for wireless networks with single hop traffic and general interference constraints. IEEE/ACM Transactions on Networking, 18(2):393–405, April 2010. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Gupta:2011:DAO

[GS11] Gagan Raj Gupta and Ness B. Shroff. Delay analysis and optimality of scheduling policies for multihop wireless networks. IEEE/ACM Transactions on Networking, 19(1):129– 141, February 2011. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Ghaderi:2013:IAP

[GS13] Javad Ghaderi and R. Srikant. The impact of access probabilities on the delay performance of Q–CSMA algorithms in wireless networks. IEEE/ACM Transactions on Networking, 21(4):1063– 1075, August 2013. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Ghosh:2016:QSP

[GS16] Arnob Ghosh and Saswati Sarkar. Quality-sensitive price competition in secondary market spectrum oligopoly: single location game. IEEE/ACM Transactions on Networking, 24 (3):1894–1907, June 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Gong:2019:TMC

[GS19] Xiaowen Gong and Ness B.

Shroff. Truthful mobile crowdsensing for strategic users with private data quality. IEEE/ACM Transactions on Networking, 27(5):1959–1972, October 2019. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Gozupek:2015:GTA

[GSA15] Didem Gözüpek, Mordechai Shalom, and Fatih Alagöz. A graph-theoretic approach to scheduling in cognitive radio networks. IEEE/ ACM Transactions on Networking, 23(1):317–328, February 2015. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Gourgy:2009:TBO

[GSD09] Amir Gourgy, Ted H. Szymanski, and Douglas G. Down. On tracking the behavior of an output-queued switch using an inputqueued switch. IEEE/ACM Transactions on Networking, 17(6):1978–1988, December 2009. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Guo:2022:EAL

[GSH^{+}22] Xiuzhen Guo, Longfei Shangguan, Yuan He, Jia Zhang, Haotian Jiang, Awais Ahmad Siddiqi, and Yunhao Liu. Efficient ambient LoRa backscatter with on-off keying modulation. IEEE/ACM Transactions on Networking, 30(2):641– 654, April 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2021.3121787.

Garetto:2008:MPF

[GSK08] Michele Garetto, Theodoros Salonidis, and Edward W. Knightly. Modeling perflow throughput and capturing starvation in CSMA multi-hop wireless networks. IEEE/ACM Transactions on Networking, 16 (4):864–877, August 2008. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Guo:2018:RSA

[GSKN18] Xueying Guo, Rahul Singh, P. R. Kumar, and Zhisheng Niu. A risk-sensitive approach for packet interdelivery time optimization in networked cyberphysical systems. IEEE/ ACM Transactions on Networking, 26(4):1976–1989, August 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Gerstel:1999:WCA

[GSKR99] Ori Gerstel, Galen Sasaki, Shay Kutten, and Rajiv Ramaswami. Worst-case analysis of dynamic wavelength allocation in optical networks. IEEE/ ACM Transactions on Networking, 7(6):833–846, December 1999. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1999-7-6/
p833-gerstel/.

Guo:2024:HNB

[GSL^{+}24] Jiani Guo, Shanshan Song, Jun Liu, Hao Chen, Jun-Hong Cui, and Guangjie Han. A hybrid NOMA-Based MAC protocol for underwater acoustic networks. IEEE/ACM Transactions on Networking, 32 (2):1187–1200, April 2024. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL

https://dl.acm.org/doi/
10.1109/TNET.2023.3311682.

Guan:2016:DAJ

[GSM16] Zhangyu Guan, G. Enrico Santagati, and Tommaso Melodia. Distributed algorithms for joint channel access and rate control in ultrasonic intrabody networks. IEEE/ ACM Transactions on Networking, 24(5):3109–3122, October 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Gharakheili:2017:EFS

[GSM^{+}17] Hassan Habibi Gharakheili, Vijay Sivaraman, Tim Moors, Arun Vishwanath, John Matthews, and Craig Russell. Enabling fast and slow lanes for content providers using software defined networking. IEEE/ACM Transactions on Networking, 25(3):1373– 1385, June 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Gong:2016:FSC

[GSN^{+}16] Wei Gong, Ivan Stojmenovic, Amiya Nayak, Kebin Liu, and Haoxiang Liu. Fast and scalable counterfeits estimation for large-scale RFID systems. IEEE/ACM Transactions on Networking, 24(2):1052– 1064, April 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Garcia-Saavedra:2018:OOO

[GSPV^{+}18] Andres Garcia-Saavedra, Paul Patras, Victor Valls, Xavier Costa-Perez, and Douglas J. Leith. ORLA/ OLAA: Orthogonal coexistence of LAA and WiFi in unlicensed spectrum. IEEE/ACM Transactions on Networking, 26(6):2665– 2678, December 2018. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Garcia-Saavedra:2015:SSO

[GSRS^{+}15] Andres Garcia-Saavedra, Balaji Rengarajan, Pablo Serrano, Daniel Camps-Mur, and Xavier Costa-Pérez. SOLOR: selfoptimizing WLANs with legacy-compatible opportunistic relays. IEEE/ ACM Transactions on Networking, 23(4):1202–1215, August 2015. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Greenberg:1999:RSB

[GSW99] Albert G. Greenberg, R. Srikant, and Ward Whitt. Resource sharing for book-ahead and instantaneous-request calls. IEEE/ACM Transactions on Networking, 7(1):10–22, February 1999. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/
                                        journals/ton/1999-7-1/
                                        p10-greenberg/.

Griffin:2002:SPP

[GSW02] Timothy G. Griffin, F. Bruce Shepherd, and Gordon Wilfong. The stable paths problem and interdomain routing. IEEE/ACM Transactions on Networking, 10(2):232–243, April 2002. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Gao:2023:BBH

[GSW^{+}23] Kaihui Gao, Chen Sun, Shuai Wang, Dan Li, Yu Zhou, Hongqiang Harry Liu, Lingjun Zhu, Ming Zhang, Xiang Deng, Cheng Zhou, and Lu Lu. Bufferbased high-coverage and low-overhead request event monitoring in the cloud. IEEE/ACM Transactions on Networking, 31(4):1732– 1747, 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2022.3224610.

Grossglauser:1999:FRM

[GT99] Matthias Grossglauser and David N. C. Tse. A framework for robust measurementbased admission control. IEEE/ACM Transactions on Networking, 7(3):293– 309, June 1999. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1999-7-3/
p293-grossglauser/.

Grah:2000:PSL

[GT00] Adrian Grah and Terence D. Todd. Packetswitched local area networks using wavelengthselective station couplers. IEEE/ACM Transactions on Networking, 8(2):251– 264, 2000. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/2000-8-2/
p251-grah/.

Grossglauser:2002:MIC

[GT02] Matthias Grossglauser and David N. C. Tse. Mobility increases the capacity of ad hoc wireless networks. IEEE/ACM Transactions on Networking, 10 (4):477–486, August 2002. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Grossglauser:2003:TSD

[GT03] Matthias Grossglauser and David N. C. Tse. A timescale decomposition approach to measurementbased admission control.

IEEE/ACM Transactions on Networking, 11(4):550– 563, August 2003. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Georgiadis:2006:OOR

[GT06] Leonidas Georgiadis and Leandros Tassiulas. Optimal overload response in sensor networks. IEEE/ ACM Transactions on Networking, 14(SI):2684–2696, June 2006. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Grokop:2010:SSB

[GT10] Leonard H. Grokop and David N. C. Tse. Spectrum sharing between wireless networks. IEEE/ ACM Transactions on Networking, 18(5):1401–1412, October 2010. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Gu:2024:PFN

[GTC^{+}24] Liyuan Gu, Ye Tian, Wei Chen, Zhongxiang Wei, Cenman Wang, and Xinming Zhang. Perflow network measurement with distributed sketch. IEEE/ACM Transactions on Networking, 32(1):411– 426, 2024. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2023.3286879.

Ganeriwal:2009:ECU

[GTS^{+}09] Saurabh Ganeriwal, Ilias Tsigkogiannis, Hohyun Shim, Vlassios Tsiatsis, Mani B. Srivastava, and Deepak Ganesan. Estimating clock uncertainty for efficient duty-cycling in sensor networks. IEEE/ACM Transactions on Networking, 17(3):843–856, June 2009. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Ghali:2019:CWT

[GTU19] Cesar Ghali, Gene Tsudik, and Ersin Uzun. In content we trust: Networklayer trust in contentcentric networking. IEEE/ ACM Transactions on Networking, 27(5):1787–1800, October 2019. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Guo:2004:STC

[Guo04] Chuanxiong Guo. SRR:

an O(1) time-complexity packet scheduler for flows in multiservice packet networks. IEEE/ACM Transactions on Networking, 12 (6):1144–1155, December 2004. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Garrett:1993:JSC

[GV93] Mark W. Garrett and Martin Vetterli. Joint source/ channel coding of statistically multiplexed realtime services on packet networks. IEEE/ACM Transactions on Networking, 1(1):71–80, February 1993. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/
                                        journals/ton/1993-1-1/
                                        p71-garrett/.

Goyal:1997:GGR

[GV97] Pawan Goyal and Harrick M. Vin. Generalized guaranteed rate scheduling algorithms: a framework. IEEE/ACM Transactions on Networking, 5(4):561– 571, August 1997. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/
                                        journals/ton/1997-5-4/
                                        p561-goyal/.

Grossglauser:2006:LMN

[GV06] Matthias Grossglauser and Martin Vetterli. Locating mobile nodes with EASE: learning efficient routes from encounter histories alone. IEEE/ACM Transactions on Networking, 14 (3):457–469, 2006. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Gomez-Vilardebo:2017:CDE

[GV17] Jesus Gomez-Vilardebo.

Competitive design of energy harvesting communications in wireless fading channels. IEEE/ACM Transactions on Networking, 25(6):3863–3872, December 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Goyal:1997:SFQ

[GVC97] Pawan Goyal, Harrick M.

Vin, and Haichen Cheng. Start-time fair queueing: a scheduling algorithm for integrated services packet switching networks. IEEE/ ACM Transactions on Networking, 5(5):690–704, October 1997. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1997-5-5/
p690-goyal/.

Gomez-Vilardebo:2017:RAM

[GVGV17] Jesus Gomez-Vilardebo and Jesus Gomez-Vilardebo. Routing in accumulative multi-hop networks. IEEE/ ACM Transactions on Networking, 25(5):2815–2828, October 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Goren:2023:DDP

[GVM23] Guy Goren, Shay Vargaftik, and Yoram Moses. Distributed dispatching in the parallel server model. IEEE/ACM Transactions on Networking, 31(4):1521– 1534, 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2022.3220931.

Ganz:1994:EAV

[GW94] Aura Ganz and Xudong Wang. Efficient algorithm for virtual topology design in multihop lightwave networks. IEEE/ ACM Transactions on Networking, 2(3):217–225, June 1994. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1994-2-3/
p217-ganz/.

Gao:2023:DVF

[GWQ^{+}23] Kaihui Gao, Shuai Wang, Kun Qian, Dan Li, Rui Miao, Bo Li, Yu Zhou, Ennan Zhai, Chen Sun, Jiaqi Gao, Dai Zhang, Binzhang Fu, Frank Kelly, Dennis Cai, Hongqiang Harry Liu, Yan Li, Hongwei Yang, and Tao Sun. Dependable virtualized fabric on programmable data plane. IEEE/ACM Transactions on Networking, 31(4):1748– 1764, 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2022.3224617.

Gu:2019:TBS

[GWYS19] Jiaxi Gu, Jiliang Wang, Zhiwen Yu, and Kele Shen. Traffic-based side-channel attack in video streaming. IEEE/ACM Transactions on Networking, 27(3): 972–985, June 2019. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Guo:2021:AAP

[GXL^{+}21] Zehua Guo, Yang Xu, Ya-Feng Liu, Sen Liu, H. Jonathan Chao, Zhi-Li Zhang, and Yuanqing Xia. AggreFlow: Achieving power efficiency, load balancing, and quality of service in data center networks. IEEE/ ACM Transactions on Networking, 29(1):17–33, February 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2020.3026015.

Guo:2021:HFH

[GXS^{+}21] Deke Guo, Junjie Xie, Xiaofeng Shi, Haofan Cai, Chen Qian, and Honghui Chen. HDS: a fast hybrid data location service for hierarchical mobile edge computing. IEEE/ ACM Transactions on Networking, 29(3):1308–1320, June 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2021.3058401.

Gao:2019:ODD

[GXW^{+}19] Kai Gao, Qiao Xiang, Xin Wang, Yang Richard Yang, and Jun Bi. An objectivedriven on-demand network abstraction for adaptive applications. IEEE/ACM Transactions on Networking, 27(2):805–818, April 2019. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Gianvecchio:2011:HBI

[GXWW11] Steven Gianvecchio, Mengjun Xie, Zhenyu Wu, and Haining Wang. Humans and bots in Internet chat: measurement, analysis, and automated classification. IEEE/ACM Transactions on Networking, 19(5):1557– 1571, October 2011. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Gambiroza:2004:DAI

[GYB^{+}04] Violeta Gambiroza, Ping Yuan, Laura Balzano, Yonghe Liu, Steve Sheafor, and Edward Knightly. Design, analysis, and implementation of DVSR: a fair high-performance protocol for packet rings. IEEE/ACM Transactions on Networking, 12(1):85– 102, February 2004. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Gao:2016:RDG

[GYJ^{+}16] Hongyu Gao, Vinod Yegneswaran, Jian Jiang, Yan Chen, Phillip Porras, Shalini Ghosh, and Haixin Duan. Reexamining DNS from a global recursive resolver perspective. IEEE/ACM Transactions on Networking, 24 (1):43–57, February 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Ge:2017:MRW

[GYLH17] Mao Ge, Tong Ye, Tony T.

Lee, and Weisheng Hu. Multicast routing and wavelength assignment in AWG-based Clos networks. IEEE/ACM Transactions on Networking, 25(3):1892– 1909, June 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Goratti:2014:UOA

[GYSPR14] Leonardo Goratti, Ece Yaprak, Stefano Savazzi, and Carlos Pomalaza-Raez. An urn occupancy approach for modeling the energy consumption of distributed beaconing. IEEE/ ACM Transactions on Networking, 22(1):203–216, February 2014. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Ghasemi:2019:GTB

[GYSZ19] Chavoosh Ghasemi, Hamed Yousefi, Kang G. Shin, and Beichuan Zhang. On the granularity of triebased data structures for name lookups and updates. IEEE/ACM Transactions on Networking, 27(2):777– 789, April 2019. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Gueye:2006:CBG

[GZCF06] Bamba Gueye, Artur Ziviani, Mark Crovella, and Serge Fdida. Constraintbased geolocation of Internet hosts. IEEE/ACM Transactions on Networking, 14(6):1219–1232, December 2006. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Gong:2016:OPB

[GZCX16] Xiaowen Gong, Junshan Zhang, Douglas Cochran, and Kai Xing. Optimal placement for barrier coverage in bistatic radar sensor networks. IEEE/ ACM Transactions on Networking, 24(1):259–271, February 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Gupta:2006:CSC

[GZDG06] Himanshu Gupta, Zongheng Zhou, Samir R. Das, and Quinyi Gu. Connected sensor cover: self-organization of sensor networks for efficient query execution. IEEE/ACM Transactions on Networking, 14(1):55– 67, February 2006. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Guo:2018:ECO

[GZJ^{+}18] Fengxian Guo, Heli Zhang, Hong Ji, Xi Li, and Victor C. M. Leung. An efficient computation offloading management scheme in the densely deployed small cell networks with mobile edge computing. IEEE/ ACM Transactions on Networking, 26(6):2651–2664, December 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Gao:2017:NAM

[GZL^{+}17] Xiaofeng Gao, Xudong Zhu, Jun Li, Fan Wu, Guihai Chen, Ding-Zhu Du, and Shaojie Tang. A novel approximation for multi-hop connected clustering problem in wireless networks. IEEE/ ACM Transactions on Networking, 25(4):2223–2234, August 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Gao:2015:TPP

[GZS15] Jianhang Gao, Qing Zhao, and Ananthram Swami. The thinnest path problem. IEEE/ACM Transactions on Networking, 23(4): 1176–1189, August 2015. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Gao:2003:PAT

[GZT03] Lixin Gao, Zhi-Li Zhang, and Don Towsley. Proxyassisted techniques for delivering continuous multimedia streams. IEEE/ ACM Transactions on Networking, 11(6):884–894, December 2003. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Gao:2023:HOC

[GZY23] Yang Gao, Hongli Zhang, and Xiangzhan Yu. Higherorder community detection: On information degeneration and its elimination. IEEE/ACM Transactions on Networking, 31(2):891– 903, April 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2022.3201668.

Ho:1996:LAS

[HA96] Joseph S. M. Ho and Ian F. Akyildiz. Local anchor scheme for reducing signaling costs in personal communications networks. IEEE/ACM Transactions on Networking, 4(5):709–725, October 1996. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1996-4-5/
p709-ho/.

Ho:1997:DHD

[HA97] Joseph S. M. Ho and Ian F.

Akyildiz. Dynamic hierarchical database architecture for location management in PCS networks. IEEE/ACM Transactions on Networking, 5(5):646– 660, October 1997. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL http://

www.acm.org/pubs/citations/
journals/ton/1997-5-5/
p646-ho/.

Han:2016:TLB

[HA16] Tao Han and Nirwan Ansari. A traffic load balancing framework for software-defined radio access networks powered by hybrid energy sources. IEEE/ACM Transactions on Networking, 24(2):1038– 1051, April 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Hosseinalipour:2022:MSH

[HAB^{+}22] Seyyedali Hosseinalipour, Sheikh Shams Azam, Christopher G. Brinton, Nicolò Michelusi, Vaneet Aggarwal, David J. Love, and Huaiyu Dai. Multistage hybrid federated learning over large-scale D2D-enabled fog networks. IEEE/ACM Transactions on Networking, 30(4):1569– 1584, 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2022.3143495.

Hayhoe:2019:CEN

[HAG19] Mikhail Hayhoe, Fady Alajaji, and Bahman Gharesifard. Curing epidemics on networks using a Polya contagion model. IEEE/ ACM Transactions on Networking, 27(5):2085–2097, October 2019. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Han:2016:ECP

[HAGL16] Seon Yeong Han, Nael B.

Abu-Ghazaleh, and Dongman Lee. Efficient and consistent path loss model for mobile network simulation. IEEE/ACM Transactions on Networking, 24 (3):1774–1786, June 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Harfoush:2009:MCB

[HBB09] Khaled Harfoush, Azer Bestavros, and John Byers. Measuring capacity bandwidth of targeted path segments. IEEE/ACM Transactions on Networking, 17 (1):80–92, February 2009. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Humblet:1993:BTA

[HBH93] Pierre Humblet, Amit Bhargava, and Michael G. Hluchyj. Ballot theorems applied to the transient analysis of nD/D/1 queues. IEEE/ACM Transactions on Networking, 1(1):81–95, February 1993. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1993-1-1/
p81-humblet/.

Henniger:1996:STB

[HBS96] Olaf Henniger, Michel Barbeau, and Behçet Sarikaya. Specification and testing of the behavior of network management agents using SDL-92. IEEE/ ACM Transactions on Networking, 4(6):951–962, December 1996. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1996-4-6/
p951-henniger/.

Huang:2020:PSC

[HBSX20] Xi Huang, Simeng Bian, Ziyu Shao, and Hong Xu. Predictive switchcontroller association and control devolution for SDN systems. IEEE/ACM Transactions on Networking, 28(6):2783–2796, December 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2020.3021787.

Herzberg:1995:HAS

[HBU95] Meir Herzberg, Stephen J. Bye, and Anthony Utano. The hop-limit approach for spare-capacity assignment in survivable networks. IEEE/ACM Transactions on Networking, 3(6):775– 784, December 1995. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL http://

www.acm.org/pubs/citations/
journals/ton/1995-3-6/
p775-herzberg/.

Hahne:2002:DQL

[HC02] Ellen L. Hahne and Abhijit K. Choudhury. Dynamic queue length thresholds for multiple loss priorities. IEEE/ACM Transactions on Networking, 10(3): 368–380, June 2002. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Ho:2007:GSN

[HC07] Kwok Shing Ho and Kwok Wai Cheung. Generalized survivable network. IEEE/ACM Transactions on Networking, 15(4):750– 760, August 2007. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Hawari:2020:HAP

[HCFC20] Mohammed Hawari, Juan-Antonio Cordero-Fuertes, and Thomas Clausen. High-accuracy packet pacing on commodity servers for constant-rate flows. IEEE/ACM Transactions on Networking, 28(5):1953– 1967, October 2020. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL https:/

/dl.acm.org/doi/10.1109/
TNET.2020.3001672.

Chung:1996:CCU

[hCgKsYwT96] Sung hark Chung, Hu gon Kim, Yong seok Yoon, and Dong wan Tcha. Costminimizing construction of a unidirectional SHR with diverse protection. IEEE/ ACM Transactions on Networking, 4(6):921–928, December 1996. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1996-4-6/
p921-chung/.

Hu:2009:EBA

[HCL09] Yan Hu, Dah-Ming Chiu, and John C. S. Lui. Entropy based adaptive flow aggregation. IEEE/ACM Transactions on Networking, 17(3):698–711, June 2009. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Hou:2017:CHW

[HCL^{+}17] Ronghui Hou, Yu Cheng, Jiandong Li, Min Sheng, and King-Shan Lui. Capacity of hybrid wireless networks with long-range social contacts behavior. IEEE/ACM Transactions on Networking, 25(2):834– 848, April 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Huang:2018:LAS

[HCL18] Longbo Huang, Minghua Chen, and Yunxin Liu. Learning-aided stochastic network optimization with state prediction. IEEE/ ACM Transactions on Networking, 26(4):1810–1820, August 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Hu:2016:EPC

[HCW^{+}16] Shuihai Hu, Kai Chen, Haitao Wu, Wei Bai, Chang Lan, Hao Wang, Hongze Zhao, and Chuanxiong Guo. Explicit path control in commodity data centers: Design and applications. IEEE/ACM Transactions on Networking, 24(5):2768–2781, October 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Hou:2023:HED

[HCW^{+}23] Bingnan Hou, Zhiping Cai, Kui Wu, Tao Yang, and Tongqing Zhou. 6Scan: a high-efficiency dynamic Internet-wide IPv6 scanner with regional encoding. IEEE/ACM Transactions on Networking, 31(4): 1870–1885, 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.

                                        acm.org/doi/10.1109/TNET.
                                        2023.3233953.

Hamza:2007:WOI

[HD07] Haitham S. Hamza and Jitender S. Deogun. WDM optical interconnects: a balanced design approach. IEEE/ACM Transactions on Networking, 15(6):1565– 1578, December 2007. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

He:2019:JAS

[HDF19] Qing He, Gyorgy Dan, and Viktoria Fodor. Joint assignment and scheduling for minimizing age of correlated information. IEEE/ ACM Transactions on Networking, 27(5):1887–1900, October 2019. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Huang:2010:VIM

[HDM10] Kaidi Huang, Ken R. Duffy, and David Malone. On the validity of IEEE 802.11 MAC modeling hypotheses. IEEE/ACM Transactions on Networking, 18(6): 1935–1948, December 2010. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Huang:2013:HRC

[HDM13] Kaidi D. Huang, Ken R. Duffy, and David Malone.

H-RCA: 802.11 collisionaware rate control. IEEE/ ACM Transactions on Networking, 21(4):1021–1034, August 2013. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Han:2016:CCB

[HDQ^{+}16] Jinsong Han, Han Ding, Chen Qian, Wei Xi, Zhi Wang, Zhiping Jiang, Longfei Shangguan, Jizhong Zhao, Jinsong Han, Han Ding, Chen Qian, Wei Xi, Zhi Wang, Zhiping Jiang, Longfei Shangguan, and Jizhong Zhao. CBID: a customer behavior identification system using passive tags. IEEE/ACM Transactions on Networking, 24(5): 2885–2898, October 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Heyman:1997:GSM

[Hey97] Daniel P. Heyman. The GBAR source model for VBR videoconferences. IEEE/ACM Transactions on Networking, 5(4):554– 560, August 1997. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1997-5-4/
p554-heyman/.

Himura:2013:SGB

[HFC^{+}13] Yosuke Himura, Kensuke Fukuda, Kenjiro Cho, Pierre Borgnat, Patrice Abry, and Hiroshi Esaki. Synoptic graphlet: bridging the gap between supervised and unsupervised profiling of host-level network traffic. IEEE/ ACM Transactions on Networking, 21(4):1284–1297, August 2013. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

He:2024:BBS

[HFF^{+}24] Qiang He, Zheng Feng, Hui Fang, Xingwei Wang, Liang Zhao, Yudong Yao, and Keping Yu. A blockchain-based scheme for secure data offloading in healthcare with deep reinforcement learning. IEEE/ACM Transactions on Networking, 32 (1):65–80, 2024. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2023.3274631.

He:2012:OPL

[HFKC12] Yihua He, Michalis Faloutsos, Srikanth V. Krishnamurthy, and Marek Chrobak. Obtaining provably legitimate Internet topologies. IEEE/ACM Transactions on Networking, 20(1):271–284, February 2012. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Hou:2014:PFD

[HG14] I-Hong Hou and Piyush Gupta. Proportionally fair distributed resource allocation in multiband wireless systems. IEEE/ACM Transactions on Networking, 22(6):1819–1830, December 2014. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Hurtig:2019:LLS

[HGB^{+}19] Per Hurtig, Karl-Johan Grinnemo, Anna Brunstrom, Simone Ferlin, Ozgu Alay, and Nicolas Kuhn. Low-latency scheduling in MPTCP. IEEE/ACM Transactions on Networking, 27(1):302–315, February 2019. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Helmy:2004:SMB

[HGE04] Ahmed Helmy, Sandeep Gupta, and Deborah Estrin. The STRESS method for boundary-point performance analysis of endto-end multicast timersuppression mechanisms. IEEE/ACM Transactions on Networking, 12(1):44– 58, February 2004. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Huang:2006:SER

[HGG06] Yaqing Huang, Roch Guérin, and Pranav Gupta. Supporting excess real-time traffic with active drop queue. IEEE/ACM Transactions on Networking, 14 (5):965–977, October 2006. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

He:2017:REM

[HGM^{+}17] Ting He, Athanasios Gkelias, Liang Ma, Kin K. Leung, Ananthram Swami, and Don Towsley. Robust and efficient monitor placement for network tomography in dynamic networks. IEEE/ACM Transactions on Networking, 25 (3):1732–1745, June 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Huang:2016:GPP

[HGW^{+}16] Qianyi Huang, Yang Gui, Fan Wu, Guihai Chen, and Qian Zhang. A general privacy-preserving auction mechanism for secondary spectrum markets. IEEE/ACM Transactions on Networking, 24(3):1881– 1893, June 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Huang:2023:AMM

[HGZ^{+}23] Xiuqi Huang, Yuanning Gao, Xinyi Zhou, Xiaofeng Gao, and Guihai Chen. An adaptive metadata management scheme based on deep reinforcement learning for large-scale distributed file systems. IEEE/ACM Transactions on Networking, 31(6):2840– 2853, 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2023.3266400.

He:2021:WPL

[HGZJ21] Yuan He, Xiuzhen Guo, Jia Zhang, and Haotian Jiang. WIDE: Physicallevel CTC via digital emulation. IEEE/ACM Transactions on Networking, 29(4):1567–1579, August 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2021.3071782.

Hajek:1998:VQR

[HH98] Bruce Hajek and Linhai He. On variations of queue response for inputs with the same mean and autocorrelation function. IEEE/ ACM Transactions on Networking, 6(5):588–598, October 1998. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/
                                        journals/ton/1998-6-5/
                                        p588-hajek/.

Hefeeda:2010:BTS

[HH10a] Mohamed Hefeeda and Cheng-Hsin Hsu. On burst transmission scheduling in mobile TV broadcast networks. IEEE/ACM Transactions on Networking, 18 (2):610–623, April 2010. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Hsu:2010:BVS

[HH10b] Cheng-Hsin Hsu and Mohamed M. Hefeeda. Broadcasting video streams encoded with arbitrary bit rates in energy-constrained mobile TV networks. IEEE/ ACM Transactions on Networking, 18(3):681–694, June 2010. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Huang:2017:ISS

[HH17] Longbo Huang and Longbo Huang. Intelligence of smart systems: Model, bounds, and algorithms. IEEE/ACM Transactions on Networking, 25(5):2960– 2973, October 2017. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Hsieh:2018:HTA

[HH18] Ping-Chun Hsieh and I-Hong Hou. Heavy-traffic analysis of QoE optimality for on-demand video streams over fading channels. IEEE/ACM Transactions on Networking, 26(4): 1768–1781, August 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Huang:2017:SGE

[HHA17] Xueqing Huang, Tao Han, and Nirwan Ansari. Smart grid enabled mobile networks: Jointly optimizing BS operation and power distribution. IEEE/ACM Transactions on Networking, 25(3):1832–1845, June 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Hilton:2022:BIS

[HHD22] Alden Hilton, Joel Hirschmann, and Casey Deccio. Beware of IPs in sheep’s clothing: Measurement and disclosure of IP spoofing vulnerabilities. IEEE/ ACM Transactions on Networking, 30(4):1659– 1673, 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.

                                        acm.org/doi/10.1109/TNET.
                                        2022.3149011.

Haas:2006:GBA

[HHL06] Zygmunt J. Haas, Joseph Y. Halpern, and Li Li. Gossipbased ad hoc routing. IEEE/ACM Transactions on Networking, 14(3):479– 491, 2006. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Han:2018:QAP

[HHL18] Kai Han, He Huang, and Jun Luo. Qualityaware pricing for mobile crowdsensing. IEEE/ ACM Transactions on Networking, 26(4):1728–1741, August 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Hu:2019:CCB

[HHL^{+}19] Jinbin Hu, Jiawei Huang, Wenjun Lv, Yutao Zhou, Jianxin Wang, and Tian He. CAPS: Coding-based adaptive packet spraying to reduce flow completion time in data center. IEEE/ ACM Transactions on Networking, 27(6):2338–2353, December 2019. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL https://

                                        dl.acm.org/doi/abs/10.
                                        1109/TNET.2019.2945863.

Harks:2016:RGP

[HHSS16] Tobias Harks, Martin Hoefer, Kevin Schewior, and Alexander Skopalik. Routing games with progressive filling. IEEE/ACM Transactions on Networking, 24(4):2553–2562, August 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Hajiaghayi:2007:POF

[HIM07] Mohammad Taghi Hajiaghayi, Nicole Immorlica, and Vahab S. Mirrokni. Power optimization in fault-tolerant topology control algorithms for wireless multi-hop networks. IEEE/ACM Transactions on Networking, 15(6):1345– 1358, December 2007. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Huin:2018:ONS

[HJG18] Nicolas Huin, Brigitte Jaumard, and Frederic Giroire. Optimal network service chain provisioning. IEEE/ ACM Transactions on Networking, 26(3):1320–1333, June 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Han:2012:ABE

[HJL^{+}12] Bo Han, Lusheng Ji, Seungjoon Lee, Bobby Bhattacharjee, and Robert R.

Miller. Are all bits equal?: experimental study of IEEE 802.11 communication bit errors. IEEE/ ACM Transactions on Networking, 20(6):1695–1706, December 2012. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Hou:2020:RVR

[HJL^{+}20] Ruomu Hou, Irvan Jahja, Loi Luu, Prateek Saxena, and Haifeng Yu. Randomized view reconciliation in permissionless distributed systems. IEEE/ ACM Transactions on Networking, 28(5):1925–1938, October 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2020.3004834.

Hart:1994:ARC

[HK94] George W. Hart and Samir G. Kelekar. Automated repair of complex systems by fault compensation. IEEE/ ACM Transactions on Networking, 2(2):193–205, April 1994. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1994-2-2/
p193-hart/.

Hung:1996:BSW

[HK96] Anthony Hung and George Kesidis. Bandwidth scheduling for wide-area ATM networks using virtual finishing times. IEEE/ACM Transactions on Networking, 4(1):49–54, February 1996. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1996-4-1/
p49-hung/.

Hwang:2011:CLO

[HK11] June Hwang and Seong-Lyun Kim. Cross-layer optimization and network coding in CSMA/CA-based wireless multihop networks. IEEE/ACM Transactions on Networking, 19(4):1028– 1042, August 2011. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Hur:2014:SDR

[HK14] Junbeom Hur and Kyungtae Kang. Secure data retrieval for decentralized disruption-tolerant military networks. IEEE/ACM Transactions on Networking, 22(1):16–26, February 2014. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Hayashi:2024:UAC

[HK24] Masahito Hayashi and Takeshi Koshiba. Universal adaptive construction of verifiable secret sharing and its application to verifiable secure distributed data storage. IEEE/ACM Transactions on Networking, 32 (1):253–267, 2024. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2023.3283577.

Houmansadr:2014:NBW

[HKB14] Amir Houmansadr, Negar Kiyavash, and Nikita Borisov. Non-blind watermarking of network flows. IEEE/ACM Transactions on Networking, 22(4):1232– 1244, August 2014. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Han:2020:SMF

[HKC^{+}20] Juhyeng Han, Seongmin Kim, Daeyang Cho, Byungkwon Choi, Jaehyeong Ha, and Dongsu Han. A secure middlebox framework for enabling visibility over multiple encryption protocols. IEEE/ACM Transactions on Networking, 28(6):2727–2740, December 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2020.3016785.

Hsu:2013:EST

[HKCL13] Chih-Cheng Hsu, Ming-Shing Kuo, Cheng-Fu Chou, and Kate Ching-Ju Lin. The elimination of spatial-temporal uncertainty in underwater sensor networks. IEEE/ ACM Transactions on Networking, 21(4):1229–1242, August 2013. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Harvey:2006:CIN

[HKL06] Nicholas J. A. Harvey, Robert Kleinberg, and April Rasala Lehman. On the capacity of information networks. IEEE/ ACM Transactions on Networking, 14(SI):2345–2364, June 2006. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Hao:2007:FME

[HKLM07] Fang Hao, Murali Kodialam, T. V. Lakshman, and Shantidev Mohanty. Fast, memory efficient flow rate estimation using runs. IEEE/ACM Transactions on Networking, 15(6):1467– 1477, December 2007. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Hao:2017:OAV

[HKLM17] Fang Hao, Murali Kodialam, T. V. Lakshman, and Sarit Mukherjee. Online allocation of virtual machines in a distributed cloud. IEEE/ACM Transactions on Networking, 25(1):238– 249, February 2017. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Hao:2012:FDM

[HKLS12] Fang Hao, Murali Kodialam, T. V. Lakshman, and Haoyu Song. Fast dynamic multiple-set membership testing using combinatorial Bloom filters. IEEE/ACM Transactions on Networking, 20(1):295– 304, February 2012. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Hajikhani:2016:RMC

[HKS16] Mohammad Javad Hajikhani, Thomas Kunz, and Howard Schwartz. A recursive method for clock synchronization in asymmetric packet-based networks. IEEE/ACM Transactions on Networking, 24(4):2332– 2342, August 2016. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Hwang:1995:OPD

[HKT95] Ren-Hung Hwang, James F.

Kurose, and Don Towsley. On-call processing delay in high speed networks. IEEE/ACM Transactions on Networking, 3(6):628– 639, December 1995. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL http://

www.acm.org/pubs/citations/
journals/ton/1995-3-6/
p628-hwang/.

Hegde:2013:ECS

[HKV^{+}13] Malati Hegde, Pavan Kumar, K. R. Vasudev, N. N. Sowmya, S. V. R. Anand, Anurag Kumar, and Joy Kuri. Experiences with a centralized scheduling approach for performance management of IEEE 802.11 wireless LANs. IEEE/ACM Transactions on Networking, 21 (2):648–662, April 2013. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Hellemans:2023:PLB

[HKV23] Tim Hellemans, Grzegorz Kielanski, and Benny Van Houdt. Performance of load balancers with bounded maximum queue length in case of non-exponential job sizes. IEEE/ACM Transactions on Networking, 31(4): 1626–1641, 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2022.3221283.

Heyman:1996:SMV

[HL96a] Daniel P. Heyman and T. V. Lakshman. Source models for VBR broadcastvideo traffic. IEEE/ACM Transactions on Networking, 4(1):40–48, February 1996. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/
                                        journals/ton/1996-4-1/
                                        p40-heyman/.

Heyman:1996:WIL

[HL96b] Daniel P. Heyman and T. V. Lakshman. What are the implications of longrange dependence for VBRvideo traffic engineering? IEEE/ACM Transactions on Networking, 4(3):301– 317, June 1996. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/
                                        journals/ton/1996-4-3/
                                        p301-heyman/.

Hac:1998:DLM

[HL98a] Anna Hać and Bo Liu. Database and location management schemes for mobile communications. IEEE/ACM Transactions on Networking, 6(6):851– 865, December 1998. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL http://

www.acm.org/pubs/citations/
journals/ton/1998-6-6/
p851-hac/.

Huang:1998:SIC

[HL98b] Nen-Fu Huang and Huey-Ing Liu. A study of isochronous channel reuse in DQDB metropolitan area networks. IEEE/ ACM Transactions on Networking, 6(4):475–484, August 1998. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1998-6-4/
p475-huang/.

Haas:1999:AH

[HL99] Zygmunt J. Haas and Ben Liang. Ad hoc mobility management with uniform quorum systems. IEEE/ACM Transactions on Networking, 7(2):228– 240, April 1999. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1999-7-2/
p228-haas/.

Hwang:2000:NMT

[HL00] Frank K. Hwang and Sheng-Chyang Liaw. On nonblocking multicast threestage Clos networks. IEEE/ ACM Transactions on Networking, 8(4):535–539, 2000. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL

                                        http://www.acm.org/pubs/
                                        citations/journals/ton/
                                        2000-8-4/p535-hwang/.

Heyman:2003:MMI

[HL03] Daniel P. Heyman and David Lucantoni. Modeling multiple IP traffic streams with rate limits. IEEE/ACM Transactions on Networking, 11(6):948– 958, December 2003. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Huang:2005:CID

[HL05] Tzu-Lun Huang and D. T. Lee. Comments and an improvement on “A distributed algorithm of delaybounded multicast routing for multimedia applications in wide area networks”. IEEE/ACM Transactions on Networking, 13(6):1410– 1411, December 2005. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). See [Jia98].

Huang:2015:AOT

[HL15] Po-Kai Huang and Xiaojun Lin. Achieving optimal throughput utility and low delay with CSMA-like algorithms: a virtual multichannel approach. IEEE/ACM Transactions on Networking, 23 (2):505–518, April 2015. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Huang:1994:SPD

[HLG94] Chun-Chong Huang and Alberto Leon-Garcia. Separation principle of dynamic transmission and enqueueing priorities for realand nonreal-time traffic in ATM multiplexers. IEEE/ ACM Transactions on Networking, 2(6):588–601, December 1994. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1994-2-6/
p588-huang/.

Hong:2018:KCK

[HLH^{+}18] Cheol-Ho Hong, Kyungwoon Lee, Jaehyun Hwang, Hyunchan Park, and Chuck Yoo. Kafe: Can OS kernels forward packets fast enough for software routers? IEEE/ACM Transactions on Networking, 26(6):2734–2747, December 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Hartline:2004:OVT

[HLHD^{+}04] Jeff R. K. Hartline, Ran Libeskind-Hadas, Kurt M. Dresner, Ethan W. Drucker, and Katrina J. Ray. Optimal virtual topologies for one-to-many communication in WDM paths and rings. IEEE/ACM Transactions on Networking, 12 (2):375–383, April 2004. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Huang:2022:GNU

[HLHL22] Yan Huang, Shaoran Li, Y. Thomas Hou, and Wenjing Lou. GPF+ : a novel ultrafast GPU-based proportional fair scheduler for 5G NReee. IEEE/ ACM Transactions on Networking, 30(2):601–615, April 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2021.3118005.

Hwang:2006:NRN

[HLL06] Frank K. Hwang, Wen-Dar Lin, and Vadim Lioubimov. On noninterruptive rearrangeable networks. IEEE/ ACM Transactions on Networking, 14(5):1141–1149, October 2006. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Han:2013:DCA

[HLL13] Kai Han, Yang Liu, and Jun Luo. Duty-cycle-aware minimum-energy multicasting in wireless sensor networks. IEEE/ACM Transactions on Networking, 21 (3):910–923, June 2013. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Huang:2021:MPR

[HLL^{+}21] Jiawei Huang, Wenjun Lyu, Weihe Li, Jianxin Wang, and Tian He. Mitigating packet reordering for random packet spraying in data center networks. IEEE/ACM Transactions on Networking, 29(3):1183– 1196, June 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2021.3056601.

Ho:2011:SMC

[HLP11] Ivan Wang-Hei Ho, Kin K.

Leung, and John W. Polak. Stochastic model and connectivity dynamics for VANETs in signalized road systems. IEEE/ ACM Transactions on Networking, 19(1):195–208, February 2011. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Hoque:2016:AAT

[HLP^{+}16] Endadul Hoque, Hyojeong Lee, Rahul Potharaju, Charles Killian, and Cristina Nita-Rotaru. Automated adversarial testing of unmodified wireless routing implementations. IEEE/ ACM Transactions on Networking, 24(6):3369–3382, December 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Han:2014:YFM

[HLS14a] Bo Han, Jian Li, and Aravind Srinivasan. Your friends have more friends than you do: identifying influential mobile users through random-walk sampling. IEEE/ACM Transactions on Networking, 22(5):1389–1400, October 2014. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Han:2014:DRE

[HLS^{+}14b] Hao Han, Yunxin Liu, Guobin Shen, Yongguang Zhang, Qun Li, and Chiu C. Tan. Design, realization, and evaluation of DozyAP for power-efficient Wi-Fi tethering. IEEE/ ACM Transactions on Networking, 22(5):1672–1685, October 2014. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Hao:2004:ISI

[HLSG04] Ruibing Hao, David Lee, Rakesh K. Sinha, and Nancy Griffeth. Integrated system interoperability testing with applications to VoIP. IEEE/ACM Transactions on Networking, 12(5):823–836, October 2004. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Huang:2013:LCC

[HLW13] Po-Kai Huang, Xiaojun Lin, and Chih-Chun Wang. A low-complexity congestion control and scheduling algorithm for multihop wireless networks with order-optimal per-flow delay. IEEE/ACM Transactions on Networking, 21 (2):495–508, April 2013. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Han:2015:AEE

[HLX^{+}15] Kai Han, Jun Luo, Liu Xiang, Mingjun Xiao, and Liusheng Huang. Achieving energy efficiency and reliability for data dissemination in duty-cycled WSNs. IEEE/ACM Transactions on Networking, 23(4):1041– 1052, August 2015. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Hu:2014:DCF

[HLZ^{+}14] Chengchen Hu, Bin Liu, Hongbo Zhao, Kai Chen, Yan Chen, Yu Cheng, and Hao Wu. Discount counting for fast flow statistics on flow size and flow volume. IEEE/ACM Transactions on Networking, 22(3): 970–981, June 2014. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Han:2021:DLG

[HLZ^{+}21] Dianqi Han, Ang Li, Lili Zhang, Yan Zhang, Jiawei Li, Tao Li, Ting Zhu, and Yanchao Zhang. Deep learning-guided jamming for cross-technology wireless networks: Attack and defense. IEEE/ ACM Transactions on Networking, 29(5):1922–1932, October 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2021.3082839.

Hao:2023:DOS

[HLZY23] Yijun Hao, Fang Li, Cong Zhao, and Shusen Yang. Delay-oriented scheduling in 5G downlink wireless networks based on reinforcement learning with partial observations. IEEE/ACM Transactions on Networking, 31 (1):380–394, 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.

acm.org/doi/10.1109/TNET.
2022.3194953.

Ho:2004:RSC

[HM04] Pin-Han Ho and Hussein T.

Mouftah. Reconfiguration of spare capacity for MPLS-based recovery in the Internet backbone networks. IEEE/ACM Transactions on Networking, 12 (1):73–84, February 2004. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Harai:2006:HSB

[HM06] Hiroaki Harai and Masayuki Murata. High-speed buffer management for 40 Gb/s-based photonic packet switches. IEEE/ ACM Transactions on Networking, 14(1):191–204, February 2006. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Huang:2011:SMP

[HMM11] Sheng Huang, Charles U.

Martel, and Biswanath Mukherjee. Survivable multipath provisioning with differential delay constraint in telecom mesh networks. IEEE/ACM Transactions on Networking, 19(3):657– 669, June 2011. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Huang:2020:RRB

[HMM^{+}20] Haojun Huang, Wang Miao, Geyong Min, Chengqiang Huang, Xu Zhang, and Chen Wang. Resilient range-based d-dimensional localization for mobile sensor networks. IEEE/ ACM Transactions on Networking, 28(5):2037–2050, October 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2020.3002946.

Huang:2013:LBA

[HMNK13] Longbo Huang, Scott Moeller, Michael J. Neely, and Bhaskar Krishnamachari. LIFO-backpressure achieves near-optimal utilitydelay tradeoff. IEEE/ACM Transactions on Networking, 21(3):831–844, June 2013. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Huang:2007:MPK

[HMvdLM07] Dijiang Huang, Manish Mehta, Appie van de Liefvoort, and Deep Medhi. Modeling pairwise key establishment for random key predistribution in largescale sensor networks. IEEE/ACM Transactions on Networking, 15(5):1204– 1215, October 2007. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Huang:2010:OTP

[HN10] Longbo Huang and Michael J.

Neely. The optimality of two prices: maximizing revenue in a stochastic communication system. IEEE/ACM Transactions on Networking, 18(2):406– 419, April 2010. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Huang:2013:UOS

[HN13] Longbo Huang and Michael J.

Neely. Utility optimal scheduling in energyharvesting networks. IEEE/ ACM Transactions on Networking, 21(4):1117–1130, August 2013. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Hoang:2023:DRL

[HNP23] Linh T. Hoang, Chuyen T.

Nguyen, and Anh T. Pham. Deep reinforcement learning-based online resource management for UAV-assisted edge computing with dual connectivity. IEEE/ACM Transactions on Networking, 31(6): 2761–2776, 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.

acm.org/doi/10.1109/TNET.
2023.3263538.

Hari:2017:POP

[HNW17] Adiseshu Hari, Urs Niesen, and Gordon Wilfong. On the problem of optimal path encoding for softwaredefined networks. IEEE/ ACM Transactions on Networking, 25(1):189–198, February 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Hong:1994:AAT

[Hon94] Seung Ho Hong. Approximate analysis of timercontrolled priority scheme in the single-service tokenpassing systems. IEEE/ ACM Transactions on Networking, 2(2):206–215, April 1994. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1994-2-2/
p206-hong/.

Hoschka:1998:CEP

[Hos98] Philipp Hoschka. Compact and efficient presentation conversion code. IEEE/ACM Transactions on Networking, 6(4):389– 396, August 1998. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/

journals/ton/1998-6-4/
p389-hoschka/.

Heidemann:1997:MPH

[HOT97] John Heidemann, Katia Obraczka, and Joe Touch. Modeling the performance of HTTP over several transport protocols. IEEE/ ACM Transactions on Networking, 5(5):616–630, October 1997. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1997-5-5/
p616-heidemann/.

Hou:2014:SHR

[Hou14]  I-Hong Hou.  Schedul-

ing heterogeneous real-time traffic over fading wireless channels. IEEE/ ACM Transactions on Networking, 22(5):1631–1644, October 2014. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Hou:2015:BDC

[Hou15] I-Hong Hou. Broadcasting delay-constrained traffic over unreliable wireless links with network coding. IEEE/ACM Transactions on Networking, 23(3): 728–740, June 2015. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Hou:2016:PPL

[HOZL16] Yuxiao Hou, Jiajue Ou, Yuanqing Zheng, and Mo Li. PLACE: Physical layer cardinality estimation for large-scale RFID systems. IEEE/ ACM Transactions on Networking, 24(5):2702–2714, October 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Hu:2000:PST

[HP00] Rose Qingyang Hu and David W. Petr. A predictive self-tuning fuzzylogic feedback rate controller. IEEE/ACM Transactions on Networking, 8 (6):697–709, 2000. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/articles/
journals/ton/2000-8-6/
p697-hu/p697-hu.pdf;
http://www.acm.org/pubs/
citations/journals/ton/
2000-8-6/p697-hu/.

Haas:2001:PQC

[HP01] Zygmunt J. Haas and Marc R. Pearlman. The performance of query control schemes for the zone routing protocol. IEEE/ ACM Transactions on Networking, 9(4):427–438, August 2001. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Ha:2023:COE

[HPP^{+}23] Youngmok Ha, Eunji Pak, Jongkil Park, Taeho Kim, and Ji Won Yoon. Clock offset estimation for systems with asymmetric packet delays. IEEE/ ACM Transactions on Networking, 31(4):1838– 1853, 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2022.3229407.

He:2006:IDS

[HPR06] Xinming He, Christos Papadopoulos, and Pavlin Radoslavov. Incremental deployment strategies for router-assisted reliable multicast. IEEE/ACM Transactions on Networking, 14(4):779–792, August 2006. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Hohn:2009:CRC

[HPV09] Nicolas Hohn, Konstantina Papagiannaki, and Darryl Veitch. Capturing router congestion and delay. IEEE/ACM Transactions on Networking, 17(3): 789–802, June 2009. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Han:2016:TDF

[HQW^{+}16] Jinsong Han, Chen Qian, Xing Wang, Dan Ma, Jizhong Zhao, Wei Xi, Zhiping Jiang, and Zhi Wang. Twins: devicefree object tracking using passive tags. IEEE/ACM Transactions on Networking, 24(3):1605–1617, June 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Han:2016:GGA

[HQY^{+}16] Jinsong Han, Chen Qian, Panlong Yang, Dan Ma, Zhiping Jiang, Wei Xi, and Jizhong Zhao. GenePrint: generic and accurate physicallayer identification for UHF RFID tags. IEEE/ACM Transactions on Networking, 24(2):846–858, April 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Hamdaoui:1995:STT

[HR95] Moncef Hamdaoui and Parameswaran Ramanathan. Selection of timed token protocol parameters to guarantee message deadlines. IEEE/ACM Transactions on Networking, 3(3): 340–351, June 1995. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL http://

www.acm.org/pubs/citations/
journals/ton/1995-3-3/
p340-hamdaoui/.

Huang:2014:MUC

[HR14] Shurui Huang and Aditya Ramamoorthy. On the multiple-unicast capacity of 3-source, 3-terminal directed acyclic networks. IEEE/ACM Transactions on Networking, 22(1):285– 299, February 2014. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Hande:2008:DUP

[HRCW08] Prashanth Hande, Sundeep Rangan, Mung Chiang, and Xinzhou Wu. Distributed uplink power control for optimal SIR assignment in cellular data networks. IEEE/ACM Transactions on Networking, 16 (6):1420–1433, December 2008. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

He:2021:PBA

[HRLY21] Lin He, Gang Ren, Ying Liu, and Jiahai Yang. PAVI: Bootstrapping accountability and privacy to IPv6 Internet. IEEE/ ACM Transactions on Networking, 29(2):695–708, April 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2020.3047667.

Hussein:2022:HNM

[HRM22] Abdalla Hussein, Catherine Rosenberg, and Patrick Mitran. Hybrid NOMA in multi-cell networks: From a centralized analysis to practical schemes. IEEE/ ACM Transactions on Networking, 30(3):1268–1282, June 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2021.3135599.

He:2003:ECA

[HS03] Jiafu He and Khosrow Sohraby. An extended combinatorial analysis framework for discrete-time queueing systems with general sources. IEEE/ACM Transactions on Networking, 11(1):95–110, February 2003. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Haas:2006:NNM

[HS06a] Zygmunt J. Haas and Tara Small. A new networking model for biological applications of ad hoc sensor networks. IEEE/ACM Transactions on Networking, 14(1):27–40, February 2006. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Hu:2006:SCT

[HS06b] An-Swol Hu and Sergio D. Servetto. On the scalability of cooperative time synchronization in pulse-connected networks. IEEE/ACM Transactions on Networking, 14 (SI):2725–2748, June 2006. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Hefeeda:2008:TMP

[HS08] Mohamed Hefeeda and Osama Saleh. Traffic modeling and proportional partial caching for peer-to-peer systems. IEEE/ACM Transactions on Networking, 16(6):1447–1460, December 2008. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Hariharan:2014:SPO

[HS14] Srikanth Hariharan and Ness B. Shroff. On samplepath optimal dynamic scheduling for sum-queue minimization in forests. IEEE/ACM Transactions on Networking, 22(1):151– 164, February 2014. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Hariharan:2016:SPO

[HS16] Srikanth Hariharan and Ness B. Shroff. On sample-path optimal dynamic scheduling for sumqueue minimization in trees under the K-hop interference model. IEEE/ ACM Transactions on Networking, 24(4):2458–2471, August 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Hubballi:2018:NTC

[HS18] Neminath Hubballi and Mayank Swarnkar. BitCoding: Network traffic classification through encoded bit level signatures. IEEE/ ACM Transactions on Networking, 26(5):2334–2346, October 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Honda:2019:NWD

[HS19] Hirotada Honda and Hiroshi Saito. Nation-wide disaster avoidance control against heavy rain. IEEE/ACM Transactions on Networking, 27(3):1084– 1097, June 2019. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Herzog:1997:SCM

[HSE97] Shai Herzog, Scott Shenker, and Deborah Estrin. Sharing the “cost” of multicast trees: an axiomatic analysis. IEEE/ACM Transactions on Networking, 5(6):847–860, December 1997. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1997-5-6/
p847-herzog/.

He:2009:LLF

[HSFK09] Yihua He, Georgos Siganos, Michalis Faloutsos, and Srikanth Krishnamurthy. Lord of the links: a framework for discovering missing links in the Internet topology. IEEE/ACM Transactions on Networking, 17(2):391–404, April 2009. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

He:2008:GSS

[HSG^{+}08] Si-Min He, Shu-Tao Sun, Hong-Tao Guan, Qiang Zheng, You-Jian Zhao, and Wen Gao. On guaranteed smooth switching for buffered crossbar switches. IEEE/ACM Transactions on Networking, 16(3):718– 731, June 2008. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Han:2006:MPT

[HSH^{+}06] Huaizhong Han, Srinivas Shakkottai, C. V. Hollot, R. Srikant, and Don Towsley. Multi-path TCP: a joint congestion control and routing scheme to exploit path diversity in the Internet. IEEE/ ACM Transactions on Networking, 14(6):1260–1271, December 2006. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Hur:2023:PCF

[HSKY23] Junnyung Hur, Hyeon Gy Shon, Young Jae Kim, and Myungkeun Yoon. Packet chunking for file detection. IEEE/ACM Transactions on Networking, 31(3):1202– 1215, June 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2022.3215549.

Hui:2020:CCD

[HSL20] Yilong Hui, Zhou Su, and Tom H. Luan. Collaborative content delivery in software-defined heterogeneous vehicular networks. IEEE/ACM Transactions on Networking, 28(2):575– 587, April 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL https://

                                        dl.acm.org/doi/abs/10.
                                        1109/TNET.2020.2968746.

He:2013:SRB

[HSM^{+}13] Yong He, Jie Sun, Xiaojun Ma, Athanasios V. Vasilakos, Ruixi Yuan, and Weibo Gong. Semi-random backoff: towards resource reservation for channel access in wireless LANs. IEEE/ACM Transactions on Networking, 21(1):204– 217, February 2013. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Huang:2020:EPS

[HSM^{+}20] He Huang, Yu-E Sun, Chaoyi Ma, Shigang Chen, You Zhou, Wenjian Yang, Shaojie Tang, Hongli Xu, and Yan Qiao. An efficient K-persistent spread estimator for traffic measurement in high-speed networks. IEEE/ACM Transactions on Networking, 28(4):1463–1476, August 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2020.2982003.

Huang:2021:SEN

[HSM^{+}21] He Huang, Yu-E Sun, Chaoyi Ma, Shigang Chen, Yang Du, Haibo Wang, and Qingjun Xiao. Spread estimation with non-duplicate sampling in high-speed networks. IEEE/ACM Transactions on Networking, 29(5):2073–2086, October 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2021.3078725.

He:2019:OMB

[HSO19] Fujun He, Takehiro Sato, and Eiji Oki. Optimization model for backup resource allocation in middleboxes with importance. IEEE/ ACM Transactions on Networking, 27(4):1742–1755, August 2019. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Hsu:2009:MST

[HSPH09] Wei-Jen Hsu, Thrasyvoulos Spyropoulos, Konstantinos Psounis, and Ahmed Helmy. Modeling spatial and temporal dependencies of user mobility in wireless mobile networks. IEEE/ ACM Transactions on Networking, 17(5):1564–1577, October 2009. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Hou:2008:RAN

[HSS08] Y. Thomas Hou, Yi Shi, and Hanif D. Sherali. Rate allocation and network lifetime problems for wireless sensor networks. IEEE/ACM Transactions on Networking, 16(2):321– 334, April 2008. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Hou:2021:ESN

[HSS^{+}21] Jing Hou, Li Sun, Tao Shu, Yong Xiao, and Marwan Krunz. Economics of strategic network infrastructure sharing: a backup reservation approach. IEEE/ ACM Transactions on Networking, 29(2):665–680, April 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2020.3044875.

Hanif:2016:MFG

[HTAZ16] Ahmed Farhan Hanif, Hamidou Tembine, Mohamad Assaad, and Djamal Zeghlache. Mean-field games for resource sharing in cloud-based networks. IEEE/ACM Transactions on Networking, 24(1):624– 637, February 2016. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Ho:2004:SSP

[HTC04] Pin-Han Ho, János Tapolcai, and Tibor Cinkler. Segment shared protection in mesh communications networks with bandwidth guaranteed tunnels. IEEE/ ACM Transactions on Networking, 12(6):1105–1118, December 2004. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). See comments [LYL07].

Han:2021:SBJ

[HTJ^{+}21] Zhenhua Han, Haisheng Tan, Shaofeng H.-C. Jiang, Wanli Cao, Xiaoming Fu, Lan Zhang, and Francis C. M. Lau. SPIN: BSP job scheduling with placementsensitive execution. IEEE/ ACM Transactions on Networking, 29(5):2267–2280, October 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2021.3087221.

Han:2019:OOL

[HTL^{+}19] Zhenhua Han, Haisheng Tan, Xiang-Yang Li, Shaofeng H.-C. Jiang, Yupeng Li, and Francis C. M. Lau. OnDisc: Online latencysensitive job dispatching and scheduling in heterogeneous edge-clouds. IEEE/ ACM Transactions on Networking, 27(6):2472–2485, December 2019. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL https://

                                        dl.acm.org/doi/abs/10.
                                        1109/TNET.2019.2953806.

Han:2019:EED

[HTW^{+}19] Zhenhua Han, Haisheng Tan, Rui Wang, Guihai Chen, Yupeng Li, and Francis Chi Moon Lau.

Energy-efficient dynamic virtual machine management in data centers. IEEE/ACM Transactions on Networking, 27(1):344– 360, February 2019. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Han:2022:EOL

[HTW^{+}22] Zhenhua Han, Haisheng Tan, Rui Wang, Yuncong Hong, and Francis C. M. Lau. Efficient online learning based cross-tier uplink scheduling in Het-Nets. IEEE/ACM Transactions on Networking, 30(6): 2389–2402, 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2022.3173432.

Hu:1993:DCA

[Hu93] Limin Hu. Distributed code assignments for CDMA Packet Radio Network. IEEE/ACM Transactions on Networking, 1(6):668– 677, December 1993. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL http://

www.acm.org/pubs/citations/
journals/ton/1993-1-6/
p668-hu/.

Huang:2017:VIM

[Hua17] Longbo Huang. The valueof-information in matching with queues. IEEE/ACM Transactions on Networking, 25(1):29–42, February 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Hohn:2006:IST

[HV06] Nicolas Hohn and Darryl Veitch. Inverting sampled traffic. IEEE/ACM Transactions on Networking, 14 (1):68–80, February 2006. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Hellemans:2022:ILB

[HV22] Tim Hellemans and Benny Van Houdt. Improved load balancing in large scale systems using attained service time reporting. IEEE/ ACM Transactions on Networking, 30(1):341–353, February 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2021.3110186.

Henri:2018:MAB

[HVT18] Sebastien Henri, Christina Vlachou, and Patrick Thiran. Multi-armed bandit in action: Optimizing performance in dynamic hybrid networks. IEEE/ ACM Transactions on Networking, 26(4):1879–1892, August 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Hobson:1999:PEP

[HW99] Richard F. Hobson and P. S. Wong. A parallel embedded-processor architecture for ATM reassembly. IEEE/ACM Transactions on Networking, 7(1):23–37, February 1999. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1999-7-1/
p23-hobson/.

Huang:2012:CSG

[HW12] Wentao Huang and Xinbing Wang. Capacity scaling of general cognitive networks. IEEE/ ACM Transactions on Networking, 20(5):1501–1513, October 2012. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Han:2022:PAP

[HW22] Yanyan Han and Hongyi Wu. Privacy-aware participant recruitment in opportunistic device to device networks. IEEE/ ACM Transactions on Networking, 30(3):1340–1351, June 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.

acm.org/doi/10.1109/TNET.
2022.3141069.

Huang:2022:TAB

[HWC22] Sijiang Huang, Mowei Wang, and Yong Cui. Traffic-aware buffer management in shared memory switches. IEEE/ ACM Transactions on Networking, 30(6):2559– 2573, 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2022.3173930.

Hayes:2020:OIG

[HWF^{+}20] David A. Hayes, Michael Welzl, Simone Ferlin, David Ros, and Safiqul Islam. Online identification of groups of flows sharing a network bottleneck. IEEE/ ACM Transactions on Networking, 28(5):2229–2242, October 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2020.3007346.

Huang:2018:TBI

[HWHW18] Jhih-Yu Huang, Pi-Chung Wang, Jhih-Yu Huang, and Pi-Chung Wang. TCAMbased IP address lookup using longest suffix split. IEEE/ACM Transactions on Networking, 26(2):976– 989, April 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Huang:2021:DCS

[HWJZ21] Yong Huang, Wei Wang, Tao Jiang, and Qian Zhang. Detecting colluding Sybil attackers in robotic networks using backscatters. IEEE/ ACM Transactions on Networking, 29(2):793–804, April 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2020.3048126.

Hou:2021:CAN

[HWLL21] Tao Hou, Tao Wang, Zhuo Lu, and Yao Liu. Combating adversarial network topology inference by proactive topology obfuscation. IEEE/ACM Transactions on Networking, 29(6):2779–2792, December 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2021.3101692.

Hosseinalipour:2024:PSL

[HWM^{+}24] Seyyedali Hosseinalipour, Su Wang, Nicolò Michelusi, Vaneet Aggarwal, Christopher G. Brinton, David J. Love, and Mung Chiang. Parallel successive learning for dynamic distributed model training over heterogeneous wireless networks. IEEE/ACM Transactions on Networking, 32 (1):222–237, 2024. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2023.3286987.

He:2023:SPP

[HWZ^{+}23] Yunhua He, Yueting Wu, Cui Zhang, Jialong Shen, Ke Xiao, Keshav Sood, and Limin Sun. A sparse protocol parsing method for IIoT based on BPSOvote-HMM hybrid model. IEEE/ACM Transactions on Networking, 31(2):485– 496, April 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2022.3184751.

Hua:2011:BNE

[HXLZ11] Nan Hua, Jun Xu, Bill Lin, and Haiquan Zhao. BRICK: a novel exact active statistics counter architecture. IEEE/ACM Transactions on Networking, 19(3):670–682, June 2011. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Hou:2023:CCC

[HXZ23] Ningning Hou, Xianjin Xia, and Yuanqing Zheng.

CloakLoRa: a covert channel over LoRa PHY. IEEE/ ACM Transactions on Networking, 31(3):1159–1172, June 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2022.3209255.

Hua:2008:ORD

[HY08] Cunqing Hua and Tak-Shing Peter Yum. Optimal routing and data aggregation for maximizing lifetime of wireless sensor networks. IEEE/ACM Transactions on Networking, 16 (4):892–903, August 2008. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Hu:2010:FBS

[HY10] Bing Hu and Kwan L.

Yeung. Feedback-based scheduling for load-balanced two-stage switches. IEEE/ ACM Transactions on Networking, 18(4):1077–1090, August 2010. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Han:2023:SSV

[HYK^{+}23] Juhyeng Han, Insu Yun, Seongmin Kim, Taesoo Kim, Sooel Son, and Dongsu Han. Scalable and secure virtualization of HSM with ScaleTrust.

IEEE/ACM Transactions on Networking, 31(4):1595– 1610, 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2022.3220427.

Huang:2021:DSA

[HYLS21] Huanhuan Huang, Tong Ye, Tony T. Lee, and Weiqiang Sun. Delay and stability analysis of connection-based slotted-aloha. IEEE/ ACM Transactions on Networking, 29(1):203–219, February 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2020.3029774.

Hu:2016:ISI

[HYZH16] Bing Hu, Kwan L. Yeung, Qian Zhou, and Chunzhi He. On iterative scheduling for input-queued switches with a speedup of 2 − 1/N. IEEE/ACM Transactions on Networking, 24(6): 3565–3577, December 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Hu:2020:SAS

[HZ20] Yidan Hu and Rui Zhang. A spatiotemporal approach for secure crowdsourced radio environment map construction. IEEE/ACM Transactions on Networking, 28(4):1790–1803, August 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2020.2992939.

Hu:2022:ABB

[HZB^{+}22] Shuihai Hu, Gaoxiong Zeng, Wei Bai, Zilong Wang, Baochen Qiao, Kai Chen, Kun Tan, and Yi Wang. Aeolus: a building block for proactive transport in datacenter networks. IEEE/ ACM Transactions on Networking, 30(2):542–556, April 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2021.3119986.

Hande:2007:DRA

[HZC07] Prashanth Hande, Shengyu Zhang, and Mung Chiang. Distributed rate allocation for inelastic flows. IEEE/ ACM Transactions on Networking, 15(6):1240–1253, December 2007. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Hu:2019:TPP

[HZC^{+}19] Shuihai Hu, Yibo Zhu, Peng Cheng, Chuanxiong Guo, Kun Tan, Jitendra Padhye, and Kai Chen. Tagger: Practical PFC deadlock prevention in data center networks. IEEE/ACM Transactions on Networking, 27(2):889– 902, April 2019. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Houmansadr:2017:SSW

[HZCB17] Amir Houmansadr, Wenxuan Zhou, Matthew Caesar, and Nikita Borisov. SWEET: Serving the Web by exploiting email tunnels. IEEE/ACM Transactions on Networking, 25 (3):1517–1527, June 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Huang:2016:WBM

[HZCL16] Longbo Huang, Shaoquan Zhang, Minghua Chen, and Xin Liu. When backpressure meets predictive scheduling. IEEE/ ACM Transactions on Networking, 24(4):2237–2250, August 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

He:2018:POT

[HZG^{+}18] Peng He, Wenyuan Zhang, Hongtao Guan, Kave Salamatian, and Gaogang Xie. Partial order theory for fast TCAM updates. IEEE/ ACM Transactions on Networking, 26(1):217–230, February 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Hou:2018:PTP

[HZHZ18] Yuxiao Hou, Yuanqing Zheng, Yuxiao Hou, and Yuanqing Zheng. PHY-Tree: Physical layer treebased RFID identification. IEEE/ACM Transactions on Networking, 26(2):711– 723, April 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Han:2016:TUB

[HZL16] Kai Han, Chi Zhang, and Jun Luo. Taming the uncertainty: budget limited robust crowdsensing through online learning. IEEE/ACM Transactions on Networking, 24(3):1462– 1475, June 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Huang:2023:CEA

[HZL^{+}23] Jiawei Huang, Wenlu Zhang, Yijun Li, Lin Li, Zhaoyi Li, Jin Ye, and Jianxin Wang. ChainSketch: an efficient and accurate sketch for heavy flow detection. IEEE/ ACM Transactions on Networking, 31(2):738–753, April 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2022.3199506.

Hu:2022:NTO

[HZLZ22] Yuchong Hu, Xiaoyang Zhang, Patrick P. C. Lee, and Pan Zhou. NC-Scale: Toward optimal storage scaling via network coding. IEEE/ ACM Transactions on Networking, 30(1):271–284, February 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2021.3106394.

Iyengar:2006:CMT

[IAS06] Janardhan R. Iyengar, Paul D. Amer, and Randall Stewart. Concurrent multipath transfer using SCTP multihoming over independent end-to-end paths. IEEE/ACM Transactions on Networking, 14(5):951– 964, October 2006. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Iness:1995:GGS

[IBM95] Jason Iness, Subrata Banerjee, and Biswanath Mukherjee. GEMNET: a generalized, shuffle-exchange-based, regular, scalable, modular, multihop, WDM lightwave network. IEEE/ ACM Transactions on Networking, 3(4):470–476, August 1995. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1995-3-4/
p470-iness/.

Intanagonwiwat:2003:DDW

[IGE^{+}03] Chalermek Intanagonwiwat, Ramesh Govindan, Deborah Estrin, John Heidemann, and Fabio Silva. Directed diffusion for wireless sensor networking. IEEE/ACM Transactions on Networking, 11(1):2–16, February 2003. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Iosifidis:2015:DAM

[IGHT15] George Iosifidis, Lin Gao, Jianwei Huang, and Leandros Tassiulas. A doubleauction mechanism for mobile data-offloading markets. IEEE/ACM Transactions on Networking, 23(5): 1634–1647, October 2015. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Iosifidis:2017:EFC

[IGHT17] George Iosifidis, Lin Gao, Jianwei Huang, and Leandros Tassiulas. Efficient and fair collaborative mobile Internet access. IEEE/ACM Transactions on Networking, 25 (3):1386–1400, June 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Ioannou:2007:PHP

[IK07] Aggelos Ioannou and Manolis G. H. Katevenis. Pipelined heap (priority queue) management for advanced scheduling in high-speed networks. IEEE/ACM Transactions on Networking, 15(2):450–461, April 2007. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Issariyakul:2009:AFC

[IK09] Teerawat Issariyakul and Vikram Krishnamurthy. Amplify-and-forward cooperative diversity wireless networks: model, analysis, and monotonicity properties. IEEE/ACM Transactions on Networking, 17(1): 225–238, February 2009. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Imon:2015:EER

[IKDD15] Sk Kajal Arefin Imon, Adnan Khan, Mario Di Francesco, and Sajal K. Das. Energy-efficient randomized switching for maximizing lifetime in treebased wireless sensor networks. IEEE/ACM Transactions on Networking, 23(5):1401–1415, October 2015. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Iyer:2008:DPB

[IKM08] Sundar Iyer, Ramana Rao Kompella, and Nick McKeown. Designing packet buffers for router linecards. IEEE/ACM Transactions on Networking, 16(3):705– 717, June 2008. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Iosifidis:2017:DSC

[IKS17] George Iosifidis, Iordanis Koutsopoulos, and Georgios Smaragdakis. Distributed storage control algorithms for dynamic networks. IEEE/ACM Transactions on Networking, 25 (3):1359–1372, June 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Iliadis:2000:OPC

[Ili00]  Ilias Iliadis.  Optimal

PNNI complex node representations for restrictive costs and minimal path computation time. IEEE/ACM Transactions on Networking, 8(4):493– 506, 2000. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.

                                        acm.org/pubs/citations/
                                        journals/ton/2000-8-4/
                                        p493-iliadis/.

Izmailov:1997:DAC

[ILS97] Rauf Izmailov, Duan-Shin Lee, and Bhaskar Sengupta. Design and analysis of a congestion-free overlay on a high-speed network. IEEE/ACM Transactions on Networking, 5(6):970–980, December 1997. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/
                                        journals/ton/1997-5-6/
                                        p970-izmailov/.

Iyer:2003:APP

[IM03] Sundar Iyer and Nick W. McKeown. Analysis of the parallel packet switch architecture. IEEE/ACM Transactions on Networking, 11(2):314–324, April 2003. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Iliadis:2008:PST

[IM08] Ilias Iliadis and Cyriel Minkenberg. Performance of a speculative transmission scheme for schedulinglatency reduction. IEEE/ ACM Transactions on Networking, 16(1):182–195, February 2008. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Iraschko:1998:OCP

[IMG98] Rainer R. Iraschko, M. H.

MacGregor, and Wayne D. Grover. Optimal capacity placement for path restoration in STM or ATM mesh-survivable networks. IEEE/ACM Transactions on Networking, 6(3):325– 336, June 1998. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1998-6-3/
p325-iraschko/.

I:1997:PMM

[IPG97] Chih-Lin I., Gregory P.

Pollini, and Richard D. Gitlin. PCS mobility management using the reverse virtual call setup algorithm. IEEE/ACM Transactions on Networking, 5(1):13–24, February 1997. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1997-5-1/
p13-i/.

Iatrou:2000:DRS

[IS00] Steve Iatrou and Ioannis Stavrakakis. A dynamic regulation and scheduling scheme for real-time traffic management. IEEE/ ACM Transactions on Networking, 8(1):60–70, February 2000. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/2000-8-1/
p60-iatrou/.

Iqbal:2022:CAL

[ISS22] Hassan Iqbal, Anand Singh, and Muhammad Shahzad. Characterizing the availability and latency in AWS network from the perspective of tenants. IEEE/ACM Transactions on Networking, 30(4):1554– 1568, 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2022.3148701.

Iida:2001:DAC

[ITSO01] Katsuyoshi Iida, Tetsuya Takine, Hideki Sunahara, and Yuji Oie. Delay analysis for CBR traffic under static-priority scheduling. IEEE/ACM Transactions on Networking, 9(2): 177–185, 2001. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/articles/
journals/ton/2001-9-2/
p177-iida/p177-iida.pdf;
 http://www.acm.org/pubs/
citations/journals/ton/
2001-9-2/p177-iida/.

Inaltekin:2008:ANE

[IW08] Hazer Inaltekin and Stephen B.

Wicker. The analysis of Nash equilibria of the oneshot random-access game for wireless networks and the behavior of selfish nodes. IEEE/ACM Transactions on Networking, 16(5):1094–1107, October 2008. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Ioannidis:2018:ACN

[IYYI18] Stratis Ioannidis, Edmund Yeh, Edmund Yeh, and Stratis Ioannidis. Adaptive caching networks with optimality guarantees. IEEE/ ACM Transactions on Networking, 26(2):737–750, April 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Ivanovich:2000:SDM

[IZC00] Milosh Ivanovich, Moshe Zukerman, and Fraser Cameron. A study of deadlock models for a multiservice medium access protocol employing a Slotted Aloha signalling channel. IEEE/ACM Transactions on Networking, 8(6): 800–811, 2000. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/articles/
journals/ton/2000-8-6/

                                        p800-ivanovich/p800-ivanovich.
                                        pdf;   http://www.acm.
                                        org/pubs/citations/journals/
                                        ton/2000-8-6/p800-ivanovich/

.

Ji:2010:OSA

[JAS10] Tianxiong Ji, Eleftheria Athanasopoulou, and R. Srikant. On optimal scheduling algorithms for small generalized switches. IEEE/ACM Transactions on Networking, 18(5):1585– 1598, October 2010. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Jiang:2011:HBS

[JAW11] Libin Jiang, Venkat Anantharam, and Jean Walrand. How bad are selfish investments in network security? IEEE/ACM Transactions on Networking, 19 (2):549–560, April 2011. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Jin:2007:PPC

[JBDF07] Yasong Jin, Soshant Bali, Tyrone E. Duncan, and Victor S. Frost. Predicting properties of congestion events for a queueing system with fBm traffic. IEEE/ACM Transactions on Networking, 15(5): 1098–1108, October 2007. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Joseph:2016:ORA

[JBR16] Vinay Joseph, Sem Borst, and Martin I. Reiman. Optimal rate allocation for video streaming in wireless networks with user dynamics. IEEE/ACM Transactions on Networking, 24 (2):820–835, April 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Ji:2013:DDC

[JC13] Shouling Ji and Zhipeng Cai. Distributed data collection in large-scale asynchronous wireless sensor networks under the generalized physical interference model. IEEE/ACM Transactions on Networking, 21(4):1270–1283, August 2013. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Jeong:1995:DPE

[JCJ95] Dong Guen Jeong, Chong-Ho Choi, and Wha Sook Jeon. Design and performance evaluation of a new medium access control protocol for local wireless data communications. IEEE/ ACM Transactions on Networking, 3(6):742–752, December 1995. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/

journals/ton/1995-3-6/
p742-jeong/.

Jahanian:2021:GBN

[JCR21] Mohammad Jahanian, Jiachen Chen, and K. K. Ramakrishnan. Graph-based namespaces and load sharing for efficient information dissemination. IEEE/ ACM Transactions on Networking, 29(6):2439–2452, December 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2021.3094839.

Jain:2003:EEA

[JD03] Manish Jain and Constantinos Dovrolis. End-to-end available bandwidth: measurement methodology, dynamics, and relation with TCP throughput. IEEE/ACM Transactions on Networking, 11(4):537– 549, August 2003. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Jaumard:2017:ESU

[JD17] Brigitte Jaumard and Maryam Daryalal. Efficient spectrum utilization in large scale RWA problems. IEEE/ACM Transactions on Networking, 25 (2):1263–1278, April 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Josilo:2019:DAR

[JD19] Sladana Josilo and Gyorgy Dan. Decentralized algorithm for randomized task allocation in fog computing systems. IEEE/ACM Transactions on Networking, 27(1):85–97, February 2019. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Josilo:2020:COS

[JD20] Slaana Jošilo and György Dán. Computation offloading scheduling for periodic tasks in mobile edge computing. IEEE/ ACM Transactions on Networking, 28(2):667–680, April 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL https://

dl.acm.org/doi/abs/10.
1109/TNET.2020.2968209.

Josilo:2022:JWE

[JD22] Sladana Jošilo and György Dán. Joint wireless and edge computing resource management with dynamic network slice selection. IEEE/ACM Transactions on Networking, 30(4):1865– 1878, 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2022.3156178.

Jamin:1997:MAC

[JDSZ97] Sugih Jamin, Peter B. Danzig, Scott J. Shenker, and Lixia Zhang. A measurement-based admission control algorithm for integrated service packet networks. IEEE/ACM Transactions on Networking, 5(1):56–70, February 1997. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1997-5-1/
p56-jamin/.

Joo:2018:WSI

[JE18] Changhee Joo and Atilla Eryilmaz. Wireless scheduling for information freshness and synchrony: Driftbased design and heavytraffic analysis. IEEE/ ACM Transactions on Networking, 26(6):2556–2568, December 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Jukan:2004:PSM

[JF04] Admela Jukan and Gerald Franzl. Path selection methods with multiple constraints in serviceguaranteed WDM networks. IEEE/ACM Transactions on Networking, 12 (1):59–72, February 2004. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Jepsen:2022:FRP

[JFM^{+}22] Theo Jepsen, Ali Fattaholmanan, Masoud Moshref, Nate Foster, Antonio Carzaniga, and Robert Soulé. Forwarding and routing with packet subscriptions. IEEE/ ACM Transactions on Networking, 30(6):2464– 2479, 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2022.3172066.

Ji:2007:CHS

[JGKT07] Ping Ji, Zihui Ge, Jim Kurose, and Don Towsley. A comparison of hardstate and soft-state signaling protocols. IEEE/ACM Transactions on Networking, 15(2):281–294, April 2007. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Ji:2014:LCS

[JGLS14] Bo Ji, Gagan R. Gupta, Xiaojun Lin, and Ness B. Shroff. Low-complexity scheduling policies for achieving throughput and asymptotic delay optimality in multichannel wireless networks. IEEE/ACM Transactions on Networking, 22(6):1911–1924, December 2014. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Jin:2003:STF

[JGMB03] Shudong Jin, Liang Guo, Ibrahim Matta, and Azer Bestavros. A spectrum of TCP-friendly windowbased congestion control algorithms. IEEE/ACM Transactions on Networking, 11(3):341–355, June 2003. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Ji:2015:AOT

[JGS^{+}15] Bo Ji, Gagan R. Gupta, Manu Sharma, Xiaojun Lin, and Ness B. Shroff. Achieving optimal throughput and near-optimal asymptotic delay performance in multichannel wireless networks with low complexity: a practical greedy scheduling policy. IEEE/ACM Transactions on Networking, 23(3):880–893, June 2015. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Jin:2021:PBC

[JHJL21] Meng Jin, Yuan He, Chengkun Jiang, and Yunhao Liu. Parallel backscatter: Channel estimation and beyond. IEEE/ ACM Transactions on Networking, 29(3):1128–1140, June 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2021.3058977.

Jajoo:2022:CSB

[JHL22] Akshay Jajoo, Y. Charlie Hu, and Xiaojun Lin. A case for sampling-based learning techniques in coflow scheduling. IEEE/ ACM Transactions on Networking, 30(4):1494– 1508, 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2021.3138923.

Jin:2024:CCA

[JHLW24] Meng Jin, Yuan He, Yunhao Liu, and Xinbing Wang. Covert communication with acoustic noise. IEEE/ACM Transactions on Networking, 32(1):207– 221, 2024. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2023.3286692.

Jin:2019:FPP

[JHM^{+}19] Meng Jin, Yuan He, Xin Meng, Yilun Zheng, Dingyi Fang, and Xiaojiang Chen. FlipTracer: Practical parallel decoding for backscatter communication. IEEE/ ACM Transactions on Networking, 27(1):330–343, February 2019. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Jin:2021:PBW

[JHM^{+}21] Meng Jin, Yuan He, Xin Meng, Dingyi Fang, and Xiaojiang Chen. Parallel backscatter in the wild: When burstiness and randomness play with you. IEEE/ACM Transactions on Networking, 29(1):65– 77, February 2021. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1109/
                                        TNET.2020.3027735.

Jiang:2005:PBL

[JHR05] Shengming Jiang, Dajiang He, and Jianqiang Rao. A prediction-based link availability estimation for routing metrics in MANETs. IEEE/ACM Transactions on Networking, 13(6):1302– 1312, December 2005. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Jin:2019:DDP

[JHS^{+}19] Haiming Jin, Baoxiang He, Lu Su, Klara Nahrstedt, and Xinbing Wang. Data-driven pricing for sensing effort elicitation in mobile crowd sensing systems. IEEE/ACM Transactions on Networking, 27(6):2208–2221, December 2019. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL https://

dl.acm.org/doi/abs/10.
1109/TNET.2019.2938453.

Jia:1998:DAD

[Jia98] Xiaohua Jia. A distributed algorithm of delay-bounded multicast routing for multimedia applications in wide area networks. IEEE/ ACM Transactions on Networking, 6(6):828–837, December 1998. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1998-6-6/

p828-jia/. See comments [HL05].

Jiang:2006:PDP

[Jia06] Yuming Jiang. Per-domain packet scale rate guarantee for expedited forwarding. IEEE/ACM Transactions on Networking, 14(3): 630–643, 2006. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Jaiswal:2007:MCS

[JID^{+}07] Sharad Jaiswal, Gianluca Iannaccone, Christophe Diot, Jim Kurose, and Don Towsley. Measurement and classification of out-of-sequence packets in a Tier-1 IP backbone. IEEE/ACM Transactions on Networking, 15(1):54–66, February 2007. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Jiang:2012:DIP

[JIN^{+}12] Hongbo Jiang, Arun Iyengar, Erich Nahum, Wolfgang Segmuller, Asser N. Tantawi, and Charles P. Wright. Design, implementation, and performance of a load balancer for SIP server clusters. IEEE/ ACM Transactions on Networking, 20(4):1190–1202, August 2012. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Jin:2008:FDC

[JJ08] Nan Jin and Scott Jordan. On the feasibility of dynamic congestion-based pricing in differentiated services networks. IEEE/ ACM Transactions on Networking, 16(5):1001–1014, October 2008. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Jin:2023:SBN

[JJJ^{+}23] Yibo Jin, Lei Jiao, Mingtao Ji, Zhuzhong Qian, Sheng Zhang, Ning Chen, and Sanglu Lu. Scheduling in-band network telemetry with convergencepreserving federated learning. IEEE/ACM Transactions on Networking, 31(5): 2313–2328, 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2023.3253302.

Jin:2015:FLC

[JJL15] Hu Jin, Bang Chul Jung, and Victor C. M. Leung. Fundamental limits of CDF-based scheduling: throughput, fairness, and feedback overhead. IEEE/ACM Transactions on Networking, 23(3):894– 907, June 2015. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Ji:2013:DBB

[JJS13a] Bo Ji, Changhee Joo, and Ness B. Shroff. Delay-based back-pressure scheduling in multihop wireless networks. IEEE/ACM Transactions on Networking, 21(5):1539– 1552, October 2013. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Ji:2013:TOS

[JJS13b] Bo Ji, Changhee Joo, and Ness B. Shroff. Throughputoptimal scheduling in multihop wireless networks without per-flow information. IEEE/ACM Transactions on Networking, 21 (2):634–647, April 2013. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Jordan:2004:VOB

[JJSS04] Scott Jordan, Kalpana Jogi, Chunlin Shi, and Ikhlaq Sidhu. The variation of optimal bandwidth and buffer allocation with the number of sources. IEEE/ ACM Transactions on Networking, 12(6):1093–1104, December 2004. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Jacob:1996:DPS

[JK96] Lillykutty Jacob and Anurag Kumar. Delay performance of some scheduling strategies in an input queuing ATM switch with multiclass bursty traffic. IEEE/ACM Transactions on Networking, 4(2):258– 271, April 1996. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/
                                        journals/ton/1996-4-2/
                                        p258-jacob/.

Jin:2005:DUP

[JK05] Youngmi Jin and George Kesidis. Dynamics of usage-priced communication networks: the case of a single bottleneck resource. IEEE/ACM Transactions on Networking, 13(5):1041– 1053, October 2005. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Jarray:2015:DAV

[JK15] Abdallah Jarray and Ahmed Karmouch. Decomposition approaches for virtual network embedding with oneshot node and link mapping. IEEE/ACM Transactions on Networking, 23 (3):1012–1025, June 2015. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Joshi:2021:SRA

[JK21] Gauri Joshi and Dhruva Kaushal. Synergy via redundancy: Adaptive replication strategies and fundamental limits. IEEE/ ACM Transactions on Networking, 29(2):737–749, April 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2020.3047513.

Jin:2013:DDN

[JKJ13] Youngmi Jin, George Kesidis, and Ju Wook Jang. Diffusion dynamics of network technologies with bounded rational users: aspiration-based learning. IEEE/ACM Transactions on Networking, 21(1):28– 40, February 2013. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Ju:1998:OTS

[JL98] Ji-Her Ju and Victor O. K. Li. An optimal topology-transparent scheduling method in multihop packet radio networks. IEEE/ACM Transactions on Networking, 6 (3):298–306, June 1998. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL

                                        http://www.acm.org/pubs/
                                        citations/journals/ton/
                                        1998-6-3/p298-ju/.

Jindal:2012:NCW

[JL12a] Apoorva Jindal and Mingyan Liu. Networked computing in wireless sensor networks for structural health monitoring. IEEE/ ACM Transactions on Networking, 20(4):1203–1216, August 2012. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Jung:2012:OSA

[JL12b] Eric Jung and Xin Liu. Opportunistic spectrum access in multiple-primary-user environments under the packet collision constraint. IEEE/ACM Transactions on Networking, 20 (2):501–514, April 2012.

CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Johnston:2015:ROA

[JLM15] Matthew Johnston, Hyang-Won Lee, and Eytan Modiano. A robust optimization approach to backup network design with random failures. IEEE/ ACM Transactions on Networking, 23(4):1216–1228, August 2015. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Joo:2016:DGA

[JLRS16] Changhee Joo, Xiaojun Lin, Jiho Ryu, and Ness B. Shroff. Distributed greedy approximation to maximum weighted independent set for scheduling with fading channels. IEEE/ACM Transactions on Networking, 24(3):1476–1488, June 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Joo:2009:UCR

[JLS09] Changhee Joo, Xiaojun Lin, and Ness B. Shroff. Understanding the capacity region of the Greedy maximal scheduling algorithm in multihop wireless networks. IEEE/ACM Transactions on Networking, 17(4):1132– 1145, August 2009. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Jiang:2017:JCT

[JLS^{+}17] Guiyuan Jiang, Siew-Kei Lam, Yidan Sun, Lijia Tu, and Jigang Wu. Joint charging tour planning and depot positioning for wireless sensor networks using mobile chargers. IEEE/ ACM Transactions on Networking, 25(4):2250–2266, August 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Ji:2016:SDA

[JLSB16] Shouling Ji, Weiqing Li, Mudhakar Srivatsa, and Raheem Beyah. Structural data de-anonymization: Theory and practice. IEEE/ ACM Transactions on Networking, 24(6):3523–3536, December 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Jiang:2024:ICB

[JLW^{+}24] Wanchun Jiang, Haoyang Li, Jia Wu, Zheyuan Liu, Jiawei Huang, Danfeng Shan, and Jianxin Wang. Improvement of Copa: Behaviors and friendliness of delay-based congestion control algorithm. IEEE/ACM Transactions on Networking, 32 (1):127–142, 2024. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2023.3278677.

Jiao:2016:OCO

[JLX^{+}16] Lei Jiao, Jun Li, Tianyin Xu, Wei Du, and Xiaoming Fu. Optimizing cost for online social networks on geo-distributed clouds. IEEE/ACM Transactions on Networking, 24 (1):99–112, February 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Jia:2019:MET

[JLZJ19] Xuya Jia, Dan Li, Jing Zhu, and Yong Jiang. Metro: an efficient traffic fast rerouting scheme with low overhead. IEEE/ ACM Transactions on Networking, 27(5):2015–2027, October 2019. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Jue:2000:MMP

[JM00] Jason P. Jue and Biswanath Mukherjee. Multiconfiguration multihop protocols: a new class of protocols for packet-switched WDM optical networks. IEEE/ACM Transactions on Networking, 8(5):631– 642, 2000. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/
                                        journals/ton/2000-8-5/
                                        p631-jue/.

Johnston:2017:CPW

[JM17] Matthew Johnston and Eytan Modiano. Controller placement in wireless networks with delayed CSI. IEEE/ACM Transactions on Networking, 25(3):1775– 1788, June 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Jia:1995:SVM

[JMI95] Feiling Jia, Biswanath Mukherjee, and Jason Iness. Scheduling variablelength messages in a singlehop multichannel local lightwave network. IEEE/ ACM Transactions on Networking, 3(4):477–488, August 1995. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/
                                        journals/ton/1995-3-4/
                                        p477-jia/.

Jagannathan:2012:QLA

[JMMT12] Krishna Jagannathan, Mihalis Markakis, Eytan Modiano, and John N. Tsitsiklis. Queue-length asymptotics for generalized max-weight scheduling in the presence of heavy-tailed traffic. IEEE/ ACM Transactions on Networking, 20(4):1096–1111, August 2012. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Jelenkovic:2007:SWN

[JMS07] Predrag R. Jelenković, Petar Momčilović, and Mark S. Squillante. Scalability of wireless networks. IEEE/ACM Transactions on Networking, 15(2):295– 308, April 2007. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Jaramillo:2008:PFN

[JMS08] Juan José Jaramillo, Fabio Milan, and R. Srikant. Padded frames: a novel algorithm for stable scheduling in load-balanced switches. IEEE/ACM Transactions on Networking, 16(5):1212– 1225, October 2008. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Jindal:2009:ARR

[JP09] Apoorva Jindal and Konstantinos Psounis. The achievable rate region of 802.11-scheduled multihop networks. IEEE/ACM Transactions on Networking, 17(4):1118–1131, August 2009. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Jindal:2013:ECC

[JP13] Apoorva Jindal and Konstantinos Psounis. On the efficiency of CSMA-CA scheduling in wireless multihop networks. IEEE/ ACM Transactions on Networking, 21(5):1392–1406, October 2013. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Jiang:2008:SNC

[JPH08] Xiaohong Jiang, Achille Pattavina, and Susumu Horiguchi. Strictly nonblocking f-cast photonic networks. IEEE/ACM Transactions on Networking, 16(3):732–745, June 2008. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Jayasumana:2019:NTM

[JPM^{+}19] Anura P. Jayasumana, Randy Paffenroth, Gunjan Mahindre, Sridhar Ramasamy, and Kelum Gajamannage. Network topology mapping from partial virtual coordinates and graph geodesics. IEEE/ ACM Transactions on Networking, 27(6):2405–2417, December 2019. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL https://

dl.acm.org/doi/abs/10.
1109/TNET.2019.2953921.

Jamjoom:2004:RCB

[JPS04] Hani Jamjoom, Padmanabhan Pillai, and Kang G. Shin. Resynchronization and controllability of bursty service requests. IEEE/ACM Transactions on Networking, 12(4):582– 594, August 2004. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Jones:2017:OAT

[JPS^{+}17] Nathaniel M. Jones, Georgios S. Paschos, Brooke Shrader, Eytan Modiano, Nathaniel M. Jones, Georgios S. Paschos, Brooke Shrader, and Eytan Modiano. An overlay architecture for throughput optimal multipath routing. IEEE/ACM Transactions on Networking, 25(5):2615– 2628, October 2017. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Jiang:1996:PCB

[JR96] Hua Jiang and Stephen S.

Rappaport. Prioritized channel borrowing without locking: a channel sharing strategy for cellular communications. IEEE/ACM Transactions on Networking, 4(2):163– 172, April 1996. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/
                                        journals/ton/1996-4-2/
                                        p163-jiang/.

Jahromizadeh:2014:JRC

[JR14] Soroush Jahromizadeh and Veselin Rakocevic. Joint rate control and scheduling for providing bounded delay with high efficiency in multihop wireless networks. IEEE/ACM Transactions on Networking, 22(5):1686– 1698, October 2014. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Jahanian:2021:NSA

[JR21] Mohammad Jahanian and K. K. Ramakrishnan. Name space analysis: Verification of named data network data planes. IEEE/ ACM Transactions on Networking, 29(2):848–861, April 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2021.3050769.

Jahanian:2022:CCI

[JR22] Mohammad Jahanian and K. K. Ramakrishnan. CoN-ICE: Consensus in intermittentlyconnected environments by exploiting naming with application to emergency response. IEEE/ACM Transactions on Networking, 30 (5):1926–1939, 2022. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL https:/

/dl.acm.org/doi/10.1109/
TNET.2022.3156101.

Jiang:2015:PPA

[JRL15] Wanchun Jiang, Fengyuan Ren, and Chuang Lin. Phase plane analysis of quantized congestion notification for data center Ethernet. IEEE/ACM Transactions on Networking, 23 (1):1–14, February 2015. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Jayavelu:2009:MCT

[JRY09] Giridhar Jayavelu, Srinivasan Ramasubramanian, and Ossama Younis. Maintaining colored trees for disjoint multipath routing under node failures. IEEE/ACM Transactions on Networking, 17(1):346– 359, February 2009. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Jamjoom:2006:RCP

[JS06] Hani Jamjoom and Kang G.

Shin. On the role and controllability of persistent clients in traffic aggregates. IEEE/ACM Transactions on Networking, 14(2):410– 423, April 2006. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Joo:2009:PRA

[JS09] Changhee Joo and Ness B. Shroff. Performance of random access scheduling schemes in multi-hop wireless networks. IEEE/ ACM Transactions on Networking, 17(5):1481–1493, October 2009. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Jaramillo:2011:OSF

[JS11] Juan José Jaramillo and R. Srikant. Optimal scheduling for fair resource allocation in ad hoc networks with elastic and inelastic traffic. IEEE/ ACM Transactions on Networking, 19(4):1125–1136, August 2011. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Jang:2012:IST

[JS12] Beakcheol Jang and Mihail L. Sichitiu. IEEE 802.11 saturation throughput analysis in the presence of hidden terminals. IEEE/ACM Transactions on Networking, 20(2):557– 570, April 2012. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Joo:2014:DPN

[JS14] Changhee Joo and Ness B.

Shroff. On the delay performance of in-network aggregation in lossy wireless sensor networks. IEEE/ACM Transactions on Networking, 22(2):662–673, April 2014. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Jung:2002:DPE

[JSBM02] Jaeyeon Jung, Emil Sit, Hari Balakrishnan, and Robert Morris. DNS performance and the effectiveness of caching. IEEE/ ACM Transactions on Networking, 10(5):589–603, October 2002. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Jiang:2003:BBC

[JSuRKH03] Xiaohong Jiang, Hong Shen, Md. Mamun ur Rashid Khandker, and Susumu Horiguchi. Blocking behaviors of crosstalk-free optical Banyan networks on vertical stacking. IEEE/ACM Transactions on Networking, 11(6): 982–993, December 2003. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Jiang:2020:BIC

[JSW^{+}20] Yuxuan Jiang, Mohammad Shahrad, David Wentzlaff, Danny H. K. Tsang, and Carlee Joe-Wong. Burstable instances for clouds: Performance modeling, equilibrium analysis, and revenue maximization. IEEE/ACM Transactions on Networking, 28(6):2489–2502, December 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2020.3015523.

Jin:2018:IMP

[JSXN18] Haiming Jin, Lu Su, Houping Xiao, and Klara Nahrstedt. Incentive mechanism for privacy-aware data aggregation in mobile crowd sensing systems. IEEE/ACM Transactions on Networking, 26(5):2019– 2032, October 2018. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Jiang:2014:IFE

[JSZ14] Junchen Jiang, Vyas Sekar, and Hui Zhang. Improving fairness, efficiency, and stability in HTTP-based adaptive video streaming with Festive. IEEE/ ACM Transactions on Networking, 22(1):326–340, February 2014. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Johari:2001:EEC

[JT01] Ramesh Johari and David Kim Hong Tan. End-to-end congestion control for the Internet: delays and stability. IEEE/ACM Transactions on Networking, 9(6): 818–832, December 2001. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Jiao:2017:SOR

[JTL^{+}17] Lei Jiao, Antonia Maria Tulino, Jaime Llorca, Yue Jin, and Alessandra Sala. Smoothed online resource allocation in multi-tier distributed cloud networks. IEEE/ACM Transactions on Networking, 25(4):2556– 2570, August 2017. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). See correction [JTL^{+}18].

Jiao:2018:CSO

[JTL^{+}18] Lei Jiao, Antonia Maria Tulino, Jaime Llorca, Yue Jin, and Alessandra Sala. Corrections to “Smoothed Online Resource Allocation in Multi-Tier Distributed Cloud Networks”. IEEE/ ACM Transactions on Networking, 26(3):1531, June 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). See [JTL^{+}17].

Javidbakht:2017:DAT

[JV17] Omid Javidbakht and Parv Venkitasubramaniam. Delay anonymity tradeoff in mix networks: Optimal routing. IEEE/ACM Transactions on Networking, 25(2):1162–1175, April 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Jin:2005:DCB

[JVJ05] Nan Jin, Gayathri Venkitachalam, and Scott Jordan. Dynamic congestionbased pricing of bandwidth and buffer. IEEE/ACM Transactions on Networking, 13(6):1233–1246, December 2005. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Jain:2006:CMU

[JVY06] Kamal Jain, Vijay V. Vazirani, and Gideon Yuval. On the capacity of multiple unicast sessions in undirected graphs. IEEE/ ACM Transactions on Networking, 14(SI):2805–2809, June 2006. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Jiang:2010:DCA

[JW10] Libin Jiang and Jean Walrand. A distributed CSMA algorithm for throughput and utility maximization in wireless networks. IEEE/ACM Transactions on Networking, 18(3):960– 972, June 2010. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Jiang:2011:ATO

[JW11] Libin Jiang and Jean Walrand. Approaching throughput-optimality in distributed CSMA scheduling algorithms with collisions. IEEE/ACM Transactions on Networking, 19 (3):816–829, June 2011. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Jiang:2023:AOB

[JW23] Suhan Jiang and Jie Wu.

Approaching an optimal bitcoin mining overlay. IEEE/ACM Transactions on Networking, 31(5):2013– 2026, 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2023.3235307.

Jiao:2024:EDE

[JWH^{+}24] Wenli Jiao, Ju Wang, Yelu He, Xiangdong Xi, Fuwei Wang, Dingyi Fang, and Xiaojiang Chen. Eliminating design effort: a reconfigurable sensing framework for chipless, backscatter tags. IEEE/ACM Transactions on Networking, 32(2):1155–1170, April 2024. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1109/TNET.2023.3320263.

Jia:2018:OSN

[JWL^{+}18] Yongzheng Jia, Chuan Wu, Zongpeng Li, Franck Le, Alex Liu, Zongpeng Li, Yongzheng Jia, Chuan Wu, Franck Le, and Alex Liu. Online scaling of NFV service chains across geo-distributed datacenters. IEEE/ACM Transactions on Networking, 26 (2):699–710, April 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Joe-Wong:2015:OSN

[JWSH15] Carlee Joe-Wong, Soumya Sen, and Sangtae Ha. Offering supplementary network technologies: adoption behavior and offloading benefits. IEEE/ACM Transactions on Networking, 23(2):355–368, April 2015. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Joe-Wong:2018:SMD

[JWSH18] Carlee Joe-Wong, Soumya Sen, and Sangtae Ha. Sponsoring mobile data: Analyzing the impact on Internet stakeholders.

IEEE/ACM Transactions on Networking, 26(3):1179– 1192, June 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Joe-Wong:2013:MAF

[JWSLC13] Carlee Joe-Wong, Soumya Sen, Tian Lan, and Mung Chiang. Multiresource allocation: fairness-efficiency tradeoffs in a unifying framework. IEEE/ACM Transactions on Networking, 21(6):1785–1798, December 2013. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Jia:2023:ECM

[JWW^{+}23] Riheng Jia, Jinhao Wu, Xiong Wang, Jianfeng Lu, Feilong Lin, Zhonglong Zheng, and Minglu Li. Energy cost minimization in wireless rechargeable sensor networks. IEEE/ ACM Transactions on Networking, 31(5):2345– 2360, 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2023.3248088.

Jiang:2021:UAG

[JWZ^{+}21] Hongbo Jiang, Mengyuan Wang, Ping Zhao, Zhu Xiao, and Schahram Dustdar. A utility-aware general framework with quantifiable privacy preservation for destination prediction in LBSs. IEEE/ ACM Transactions on Networking, 29(5):2228–2241, October 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2021.3084251.

Jin:2023:KAI

[JWZ23] Meng Jin, Xinbing Wang, and Chenghu Zhou. Key agreement on IoT devices with echo profiling. IEEE/ACM Transactions on Networking, 31(4):1795– 1808, 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2022.3230642.

Jeong:2016:EEW

[JYC^{+}16] Jaeseong Jeong, Yung Yi, Jeong-Woo Cho, Do Young Eun, and Song Chong. Energy-efficient Wi-Fi sensing policy under generalized mobility patterns with aging. IEEE/ACM Transactions on Networking, 24(4):2416–2428, August 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Jiang:2019:BBC

[JYL^{+}19] Wenchao Jiang, Zhimeng Yin, Ruofeng Liu, Zhijun Li, Song Min Kim, and Tian He. Boosting the bitrate of crosstechnology communication on commodity IoT devices. IEEE/ACM Transactions on Networking, 27(3):1069– 1083, June 2019. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Jiang:2015:CBS

[JYT^{+}15] Hongbo Jiang, Tianlong Yu, Chen Tian, Guang Tan, and Chonggang Wang. Connectivity-based segmentation in large-scale 2-D/3-D sensor networks: algorithm and applications. IEEE/ACM Transactions on Networking, 23(1):15– 27, February 2015. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Jin:2018:PPC

[JZ18] Xiaocong Jin and Yanchao Zhang. Privacypreserving crowdsourced spectrum sensing. IEEE/ ACM Transactions on Networking, 26(3):1236–1249, June 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Jian:2011:AML

[JZC11] Ying Jian, Ming Zhang, and Shigang Chen. Achieving MAC-layer fairness in CSMA/CA networks. IEEE/ACM Transactions on Networking, 19(5):1472– 1484, October 2011. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Jiang:2018:RTR

[JZW^{+}18] Hongbo Jiang, Ping Zhao, Chen Wang, Chen Wang, Hongbo Jiang, and Ping Zhao. RobLoP: Towards robust privacy preserving against location dependent attacks in continuous LBS queries. IEEE/ACM Transactions on Networking, 26(2):1018–1032, April 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Konstantopoulos:1995:OFC

[KA95] Takis Konstantopoulos and Venkat Anantharam. Optimal flow control schemes that regulate the burstiness of traffic. IEEE/ ACM Transactions on Networking, 3(4):423–432, August 1995. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/
                                        journals/ton/1995-3-4/
                                        p423-konstantopoulos/.

Karasan:1998:EWR

[KA98] Ezhan Karasan and Ender Ayanoglu. Effects of wavelength routing and selection algorithms on wavelength conversion gain in WDM optical networks. IEEE/ACM Transactions on Networking, 6(2):186– 196, April 1998. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1998-6-2/
p186-karasan/.

Keon:2003:OPM

[KA03] Neil J. Keon and G. Anandalingam. Optimal pricing for multiple services in telecommunications networks offering quality-of-service guarantees. IEEE/ ACM Transactions on Networking, 11(1):66–80, February 2003. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Kabacinski:2020:WSN

[KA20] Wojciech Kabaciński and Mustafa Abdulsahib. Widesense nonblocking convertingspace-converting switching node architecture under XsVarSWITCH control algorithm. IEEE/ ACM Transactions on Networking, 28(4):1550–1561, August 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2020.2989639.

Krishnasamy:2018:AMW

[KAA^{+}18] Subhashini Krishnasamy, P. T. Akhil, Ari Arapostathis, Rajesh Sundaresan, and Sanjay Shakkottai. Augmenting maxweight with explicit learning for wireless scheduling with switching costs. IEEE/ACM Transactions on Networking, 26(6):2501– 2514, December 2018. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Kavitha:2014:FSC

[KAEAS14] Veeraruna Kavitha, Eitan Altman, R. El-Azouzi, and Rajesh Sundaresan. Fair scheduling in cellular systems in the presence of noncooperative mobiles. IEEE/ACM Transactions on Networking, 22(2):580– 594, April 2014. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Kornycky:2017:RFT

[KAHKB17] Joe Kornycky, Omar Abdul-Hameed, Ahmet Kondoz, and Brian C. Barber. Radio frequency traffic classification over WLAN. IEEE/ACM Transactions on Networking, 25(1):56– 68, February 2017. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Kaiserswerth:1993:PPE

[Kai93]  Matthias  Kaiserswerth.

The Parallel Protocol Engine. IEEE/ACM Transactions on Networking, 1(6):650–663, December 1993. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1993-1-6/
p650-kaiserswerth/.

Kiani:2019:HCP

[KAK19] Abbas Kiani, Nirwan Ansari, and Abdallah Khreishah. Hierarchical capacity provisioning for fog computing. IEEE/ACM Transactions on Networking, 27(3):962–971, June 2019. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Kamal:1996:PMP

[Kam96] Ahmed E. Kamal. Performance modeling of partial packet discarding using the end-of-packet indicator in AAL type 5. IEEE/ ACM Transactions on Networking, 4(6):929–940, December 1996. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1996-4-6/
p929-kamal/.

Kamal:2010:NPM

[Kam10] Ahmed E. Kamal. 1 + N network protection for mesh networks: network coding-based protection using p-cycles. IEEE/ACM Transactions on Networking, 18(1):67–80, February 2010. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Kumar:2007:NIF

[KAMG07] Anurag Kumar, Eitan Altman, Daniele Miorandi, and Munish Goyal. New insights from a fixedpoint analysis of single cell IEEE 802.11 WLANs. IEEE/ACM Transactions on Networking, 15(3):588– 601, June 2007. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Karbowski:2003:CSF

[Kar03] Andrzej Karbowski. Comments on “Optimization flow control, I: Basic algorithm and convergence”. IEEE/ACM Transactions on Networking, 11(2):338– 339, April 2003. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). See [LL99].

Karsten:2006:CEI

[Kar06]  Martin Karsten.  Col-

lected experience from implementing RSVP. IEEE/ ACM Transactions on Networking, 14(4):767–778, August 2006. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Karsten:2010:AGP

[Kar10] Martin Karsten. Approximation of generalized processor sharing with interleaved stratified timer wheels. IEEE/ACM Transactions on Networking, 18 (3):708–721, June 2010. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Kavurmacioglu:2016:CPC

[KAS16] Emir Kavurmacioglu, Murat Alanyali, and David Starobinski. Competition in private commons: price war or market sharing? IEEE/ACM Transactions on Networking, 24(1):29– 42, February 2016. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Kettaneh:2022:ARN

[KAT^{+}22] Ibrahim Kettaneh, Ahmed Alquraan, Hatem Takruri, Ali José Mashtizadeh, and Samer Al-Kiswany. Accelerating reads with innetwork consistency-aware load balancing. IEEE/ ACM Transactions on Networking, 30(3):954–968, June 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2021.3126203.

Krishnamurthy:2001:PBM

[KAZ01] Srikanth V. Krishnamurthy, Anthony S. Acampora, and Michele Zorzi. Pollingbased media access protocols for use with smart adaptive array antennas. IEEE/ACM Transactions on Networking, 9(2):148– 161, 2001. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/articles/
journals/ton/2001-9-2/
p148-krishnamurthy/p148-
krishnamurthy.pdf;   http:
//www.acm.org/pubs/citations/
journals/ton/2001-9-2/
p148-krishnamurthy/.

Kasbekar:2011:LCG

[KBS11] Gaurav S. Kasbekar, Yigal Bejerano, and Saswati Sarkar. Lifetime and coverage guarantees through distributed coordinate-free sensor activation. IEEE/ ACM Transactions on Networking, 19(2):470–483, April 2011. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Kasbekar:2012:GCV

[KBS12] Gaurav S. Kasbekar, Yigal Bejerano, and Saswati Sarkar. Generic coverage verification without location information using dimension reduction. IEEE/ ACM Transactions on Networking, 20(6):1991–2004, December 2012. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Kim:2013:RBW

[KBV^{+}13] Tae-Suk Kim, Ioannis Broustis, Serdar Vural, Dimitris Syrivelis, Shailendra Singh, Srikanth V. Krishnamurthy, and Thomas F. La Porta. Realizing the benefits of wireless network coding in multirate settings. IEEE/ACM Transactions on Networking, 21 (3):950–962, June 2013. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Krupczak:1997:IPR

[KCA97] Bobby Krupczak, Kenneth L. Calvert, and Mostafa H. Ammar. Increasing the portability and re-usability of protocol code. IEEE/ACM Transactions on Networking, 5(4):445–459, August 1997. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1997-5-4/
p445-krupczak/.

Kong:2003:NSS

[KCB03] Peng-Yong Kong, Kee-Chaing Chua, and Brahim Bensaou. A novel scheduling scheme to share dropping ratio while guaranteeing a delay bound in a multiCode-CDMA network. IEEE/ACM Transactions on Networking, 11 (6):994–1006, December 2003. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Kwak:2016:PNS

[KCCM16] Jeongho Kwak, Okyoung Choi, Song Chong, and Prasant Mohapatra. Processornetwork speed scaling for energy: delay tradeoff in smartphone applications. IEEE/ACM Transactions on Networking, 24(3):1647– 1660, June 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Kong:2019:MRS

[KCH^{+}19] Linghe Kong, Yifeng Cao, Liang He, Guihai Chen, Min-You Wu, and Tian He. Multi-rate selection in Zig-Bee. IEEE/ACM Transactions on Networking, 27 (3):1055–1068, June 2019. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Kim:2016:QAS

[KCM16] Joongheon Kim, Giuseppe Caire, and Andreas F. Molisch. Quality-aware streaming and scheduling for device-to-device video delivery. IEEE/ ACM Transactions on Networking, 24(4):2319–2331, August 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Keralapura:2008:RCC

[KCTI08] Ram Keralapura, Chen-Nee Chuah, Nina Taft, and Gianluca Iannaccone. Race conditions in coexisting overlay networks. IEEE/ACM Transactions on Networking, 16(1):1–14, February 2008. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Kim:2000:PSR

[KD00] Dongsoo S. Kim and Ding-Zhu Du. Performance of split routing algorithm for three-stage multicast networks. IEEE/ACM Transactions on Networking, 8 (4):526–534, 2000. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/2000-8-4/
p526-kim/.

Kim:2010:LDS

[KD10] Hongseok Kim and Gustavo De Veciana. Leveraging dynamic spare capacity in wireless systems to conserve mobile terminals’ energy. IEEE/ACM Transactions on Networking, 18(3): 802–815, June 2010. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Khabbazian:2015:BIW

[KDHK15] Majid Khabbazian, Stephane Durocher, Alireza Haghnegahdar, and Fabian Kuhn. Bounding interference in wireless ad hoc networks with nodes in random position. IEEE/ACM Transactions on Networking, 23(4): 1078–1091, August 2015. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Kim:2012:DOU

[KDYV12] Hongseok Kim, Gustavo De Veciana, Xiangying Yang, and Muthaiah Venkatachalam. Distributed α-optimal user association and cell load balancing in wireless networks. IEEE/ ACM Transactions on Networking, 20(1):177–190, February 2012. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Kashef:2016:OPR

[KE16] Mohamed Kashef and Anthony Ephremides. Optimal partial relaying for energy-harvesting wireless networks. IEEE/ ACM Transactions on Networking, 24(1):113–122, February 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Kesavareddigari:2021:CIC

[KE21] Himaja Kesavareddigari and Atilla Eryilmaz. Counterintuitive characteristics of rational decision-making using biased inputs in information networks. IEEE/ ACM Transactions on Networking, 29(4):1774–1785, August 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2021.3075430.

Krishnamurthy:2008:ASS

[KEAAH08] Supriya Krishnamurthy, Sameh El-Ansary, Erik Aurell, and Seif Haridi. An analytical study of a structured overlay in the presence of dynamic membership. IEEE/ACM Transactions on Networking, 16 (4):814–825, August 2008. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Koksal:2013:CWN

[KES13] C. Emre Koksal, Ozgur Ercetin, and Yunus Sarikaya. Control of wireless networks with secrecy. IEEE/ ACM Transactions on Networking, 21(1):324–337, February 2013. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Khisti:2006:FLS

[KEW06] Ashish Khisti, Uri Erez, and Gregory W. Wornell. Fundamental limits and scaling behavior of cooperative multicasting in wireless networks. IEEE/ACM Transactions on Networking, 14 (SI):2762–2770, June 2006. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Kousa:1999:PAN

[KEY99] Maan A. Kousa, Ahmed K. Elhakeem, and Hui Yang. Performance of ATM networks under hybrid ARQ/ FEC error control scheme. IEEE/ACM Transactions on Networking, 7(6):917– 925, December 1999. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL http://

                                        www.acm.org/pubs/citations/
                                        journals/ton/1999-7-6/
                                        p917-kousa/.

Kong:1999:MSS

[KG99] Keith Kong and Dipak Ghosal. Mitigating serverside congestion in the Internet through pseudoserving. IEEE/ACM Transactions on Networking, 7(4):530– 544, August 1999. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1999-7-4/
p530-kong/.

Koskie:2005:NGA

[KG05] Sarah Koskie and Zoran Gajic. A Nash game algorithm for SIR-based power control in 3G wireless CDMA networks. IEEE/ ACM Transactions on Networking, 13(5):1017–1026, October 2005. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Ko:2010:EBI

[KG10] Young Myoung Ko and Natarajan Gautam. Epidemicbased information dissemination in wireless mobile sensor networks. IEEE/ ACM Transactions on Networking, 18(6):1738–1751, December 2010. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Kwon:2016:TSP

[KG16] Soongeol Kwon and Natarajan Gautam. Timestable performance in parallel queues with nonhomogeneous and multiclass workloads. IEEE/ ACM Transactions on Networking, 24(3):1322–1335, June 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Kassir:2021:AMP

[KGdV^{+}21] Saadallah Kassir, Pablo Caballero Garces, Gustavo de Veciana, Nannan Wang, Xi Wang, and Paparao Palacharla. An analytical model and performance evaluation of multihomed multilane VANETs. IEEE/ ACM Transactions on Networking, 29(1):346–359, February 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2020.3032324.

Kwong:2011:FEP

[KGGZ11] Kin-Wah Kwong, Lixin Gao, Roch Guérin, and Zhi-Li Zhang. On the feasibility and efficacy of protection routing in IP networks. IEEE/ACM Transactions on Networking, 19(5):1543– 1556, October 2011. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Korosi:2020:MRH

[KGH^{+}20] Attila Kőrösi, András Gulyás, Zalan Heszberger, József Bíró, and Gábor Rétvári. On the memory requirement of hop-by-hop routing: Tight bounds and optimal address spaces. IEEE/ACM Transactions on Networking, 28(3):1353– 1363, June 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2020.2984761.

Karol:2003:PDL

[KGL03] Mark Karol, S. Jamaloddin Golestani, and David Lee. Prevention of deadlocks and livelocks in lossless backpressured packet networks. IEEE/ACM Transactions on Networking, 11(6):923–934, December 2003. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Khalili:2013:MPO

[KGPL13] Ramin Khalili, Nicolas Gast, Miroslav Popovic, and Jean-Yves Le Boudec. MPTCP is not Paretooptimal: performance issues and a possible solution. IEEE/ACM Transactions on Networking, 21(5):

1651–1665, October 2013. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Komolafe:2007:HFR

[KH07] Olufemi Komolafe and David Harle. An holistic framework for regular virtual topology design. IEEE/ACM Transactions on Networking, 15(6): 1555–1564, December 2007. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Kim:2015:DAM

[KH15] Yunbae Kim and Ganguk Hwang. Design and analysis of medium access protocol: throughput and shortterm fairness perspective. IEEE/ACM Transactions on Networking, 23(3):959– 972, June 2015. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Keshavarz-Haddad:2017:HSC

[KHAWC17] Alireza Keshavarz-Haddad, Ehsan Aryafar, Michael Wang, and Mung Chiang. HetNets selection by clients: Convergence, efficiency, and practicality. IEEE/ACM Transactions on Networking, 25(1): 406–419, February 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Kvalbein:2009:MRC

[KHČ^{+}09] Amund Kvalbein, Audun Fosselie Hansen, Tarik Čičic, Stein Gjessing, and Olav Lysne. Multiple routing configurations for fast IP network recovery. IEEE/ ACM Transactions on Networking, 17(2):473–486, April 2009. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Khan:2014:SSF

[KHG^{+}14] Faisal Khan, Nicholas Hosein, Soheil Ghiasi, Chen-Nee Chuah, and Puneet Sharma. Streaming solutions for fine-grained network traffic measurements and analysis. IEEE/ACM Transactions on Networking, 22(2):377–390, April 2014. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Kim:2018:STS

[KHH^{+}18] Seongmin Kim, Juhyeng Han, Jaehyeong Ha, Taesoo Kim, and Dongsu Han. SGX-Tor: a secure and practical Tor anonymity network with SGX enclaves. IEEE/ACM Transactions on Networking, 26(5):2174–2187, October 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Keys:2013:ISI

[KHLC13] Ken Keys, Young Hyun, Matthew Luckie, and Kim Claffy. Internet-scale IPv4 alias resolution with MI-DAR. IEEE/ACM Transactions on Networking, 21 (2):383–399, April 2013. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Kasera:2000:SRM

[KHTK00] Sneha Kumar Kasera, Gísli Hjálmtýsson, Donald F. Towsley, and James F. Kurose. Scalable reliable multicast using multiple multicast channels. IEEE/ACM Transactions on Networking, 8(3):294– 310, 2000. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/2000-8-3/
p294-kasera/.

Koutsonikolas:2012:PHT

[KHW12] Dimitrios Koutsonikolas, Y. Charlie Hu, and Chih-Chun Wang. Pacifier: highthroughput, reliable multicast without “Crying babies” in wireless mesh networks. IEEE/ACM Transactions on Networking, 20(5):1375–1388, October 2012. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Kou:2020:BEA

[KHYA20] Caixia Kou, Dedong Hu, Jianhua Yuan, and Wenbao Ai. Bisection and exact algorithms based on the Lagrangian dual for a single-constrained shortest path problem. IEEE/ ACM Transactions on Networking, 28(1):224–233, February 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL https://

dl.acm.org/doi/abs/10.
1109/TNET.2019.2955451.

Kim:2024:EDG

[KIL24] Junseon Kim, Youngbin Im, and Kyunghan Lee. Enabling delayguaranteed congestion control with one-bit feedback in cellular networks. IEEE/ACM Transactions on Networking, 32(1): 3–16, 2024. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2023.3268721.

Kim:1994:DPM

[Kim94]  Hyong S. Kim.  De-

sign and performance of Multinet switch: a multistage ATM switch architecture with partially shared buffers. IEEE/ ACM Transactions on Networking, 2(6):571–580, December 1994. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/
                                        journals/ton/1994-2-6/
                                        p571-kim/.

Kim:1998:DSG

[Kim98] Peter Kim. Deterministic service guarantees in IEEE 802.12 networks — part I: the single-hub case. IEEE/ ACM Transactions on Networking, 6(5):645–658, October 1998. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/
                                        journals/ton/1998-6-5/
                                        p645-kim/.

Kulkarni:2006:ALI

[KIR06] Sunil Kulkarni, Aravind Iyer, and Catherine Rosenberg. An address-light, integrated MAC and routing protocol for wireless sensor networks. IEEE/ACM Transactions on Networking, 14(4):793–806, August 2006. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Karnik:2008:TOC

[KIR08] Aditya Karnik, Aravind Iyer, and Catherine Rosenberg. Throughput-optimal configuration of fixed wireless networks. IEEE/ ACM Transactions on Networking, 16(5):1161–1174, October 2008. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Kim:2017:FSC

[KIW^{+}17] Song Min Kim, Shigemi Ishida, Shuai Wang, Tian He, Song Min Kim, Shigemi Ishida, Shuai Wang, and Tian He. Free side-channel cross-technology communication in wireless networks. IEEE/ACM Transactions on Networking, 25(5):2974– 2987, October 2017. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Kalyanaraman:2000:ESA

[KJF^{+}00] Shivkumar Kalyanaraman, Raj Jain, Sonia Fahmy, Rohit Goyal, and Bobby Vandalore. The ER-ICA switch algorithm for ABR traffic management in ATM networks. IEEE/ ACM Transactions on Networking, 8(1):87–98, February 2000. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/2000-8-1/
p87-kalyanaraman/.

Kolla:2018:CLS

[KJG18] Ravi Kumar Kolla, Krishna Jagannathan, and Aditya Gopalan. Collaborative learning of stochastic bandits over a social network. IEEE/ACM Transactions on Networking, 26(4):1782– 1795, August 2018. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Kessler:1993:CFR

[KK93] Ilan Kessler and Arvind Krishna. On the cost of fairness in ring networks. IEEE/ACM Transactions on Networking, 1(3):306– 313, June 1993. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/
                                        journals/ton/1993-1-3/
                                        p306-kessler/.

Kim:2000:BAW

[KK00] Jeong Geun Kim and Marwan M. Krunz. Bandwidth allocation in wireless networks with guaranteed packet-loss performance. IEEE/ACM Transactions on Networking, 8 (3):337–349, 2000. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/
                                        journals/ton/2000-8-3/
                                        p337-kim/.

Kim:2003:PAM

[KK03a] Hakyong Kim and Kiseon Kim. Performance analysis of the multiple inputqueued packet switch with the restricted rule. IEEE/ ACM Transactions on Networking, 11(3):478–487, June 2003. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Korkmaz:2003:BDC

[KK03b] Turgay Korkmaz and Marwan Krunz. Bandwidthdelay constrained path selection under inaccurate state information. IEEE/ ACM Transactions on Networking, 11(3):384–398, June 2003. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Karnik:2005:PTC

[KK05] Aditya Karnik and Anurag Kumar. Performance of TCP congestion control with explicit rate feedback. IEEE/ACM Transactions on Networking, 13(1):108– 120, February 2005. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Kuzmanovic:2006:LRT

[KK06a] Aleksandar Kuzmanovic and Edward W. Knightly. Low-rate TCP-targeted denial of service attacks and counter strategies. IEEE/ ACM Transactions on Networking, 14(4):683–696, August 2006. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Kuzmanovic:2006:TLL

[KK06b] Aleksandar Kuzmanovic and Edward W. Knightly. TCP-LP: low-priority service via end-point congestion control. IEEE/ACM Transactions on Networking, 14(4):739–752, August 2006. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Karnik:2007:DOS

[KK07] Aditya Karnik and Anurag Kumar. Distributed optimal self-organization in ad hoc wireless sensor networks. IEEE/ACM Transactions on Networking, 15(5):1035–1045, October 2007. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Kozica:2012:QTP

[KK12] Ermin Kozica and W. Bastiaan Kleijn. A quantization theoretic perspective on simulcast and layered multicast optimization. IEEE/ACM Transactions on Networking, 20 (2):585–593, April 2012. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Kandhway:2016:CHS

[KK16a] Kundan Kandhway and Joy Kuri. Campaigning in heterogeneous social networks: optimal control of SI information epidemics. IEEE/ACM Transactions on Networking, 24(1):383– 396, February 2016. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Kandhway:2016:ORA

[KK16b] Kundan Kandhway and Joy Kuri. Optimal resource allocation over time and degree classes for maximizing information dissemination in social networks. IEEE/ ACM Transactions on Networking, 24(5):3204–3217, October 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Kumar:2021:PFC

[KK21] B. R. Vinay Kumar and Navin Kashyap. Probabilistic forwarding of coded packets on networks. IEEE/ACM Transactions on Networking, 29(1):234– 247, February 2021. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL https:/

/dl.acm.org/doi/10.1109/
TNET.2020.3031467.

Karaca:2013:OSP

[KKEE13] Mehmet Karaca, Karim Khalil, Eylem Ekici, and Ozgur Ercetin. Optimal scheduling and power allocation in cooperate-to-join cognitive radio networks. IEEE/ACM Transactions on Networking, 21 (6):1708–1721, December 2013. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Kim:2022:MOB

[KKH^{+}22] Junghoon Kim, Taejoon Kim, Morteza Hashemi, David J. Love, and Christopher G. Brinton. Minimum overhead beamforming and resource allocation in D2D edge networks. IEEE/ACM Transactions on Networking, 30(4):1454– 1468, 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2021.3133022.

Kar:2006:DNA

[KKJ06] Koushik Kar, Ananth Krishnamurthy, and Neeraj Jaggi. Dynamic node activation in networks of rechargeable sensors. IEEE/ACM Transactions on Networking, 14(1):15– 26, February 2006. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Kar:2003:RRB

[KKL03] Koushik Kar, Murali Kodialam, and T. V. Lakshman. Routing restorable bandwidth guaranteed connections using maximum 2route flows. IEEE/ACM Transactions on Networking, 11(5):772–781, October 2003. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Keslassy:2005:GSS

[KKLS05] Isaac Keslassy, Murali Kodialam, T. V. Lakshman, and Dimitrios Stiliadis. On guaranteed smooth scheduling for input-queued switches. IEEE/ACM Transactions on Networking, 13(6):1364– 1375, December 2005. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Kalmanek:1997:XLE

[KKM^{+}97] Charles R. Kalmanek, Srinivasan Keshav, William T. Marshall, Samuel P. Morgan, and Robert C. Restrick III. Xunet 2: lessons from an early wide-area ATM testbed. IEEE/ACM Transactions on Networking, 5(1):40–55, February 1997. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1997-5-1/
p40-kalmanek/.

Khanafer:2015:RBP

[KKP15] Ali Khanafer, Murali Kodialam, and Krishna P. N. Puttaswamy. To rent or to buy in the presence of statistical information: the constrained ski-rental problem. IEEE/ACM Transactions on Networking, 23(4): 1067–1077, August 2015. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Kim:2008:WBA

[KKS^{+}08] Min Sik Kim, Taekhyun Kim, Yong-June Shin, Simon S. Lam, and Edward J. Powers. A wavelet-based approach to detect shared congestion. IEEE/ACM Transactions on Networking, 16(4):763–776, August 2008. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Koprulu:2019:BOE

[KKS19] Irem Koprulu, Yoora Kim, and Ness B. Shroff. Battle of opinions over evolving social networks. IEEE/ ACM Transactions on Networking, 27(2):532–545, April 2019. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Keslassy:2012:PPG

[KKSS12] Isaac Keslassy, Kirill Kogan, Gabriel Scalosub, and Michael Segal. Providing performance guarantees in multipass network processors. IEEE/ACM Transactions on Networking, 20(6): 1895–1909, December 2012. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Kadloor:2016:MTS

[KKV16] Sachin Kadloor, Negar Kiyavash, and Parv Venkitasubramaniam. Mitigating timing side channel in shared schedulers. IEEE/ACM Transactions on Networking, 24(3):1562– 1573, June 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Kim:1995:TTS

[KL95] Kyeong Soo Kim and Byeong Gi Lee. Three-level traffic shaper and its application to source clock frequency recovery for VBR services in ATM networks. IEEE/ACM Transactions on Networking, 3(4):450– 458, August 1995. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1995-3-4/
p450-kim/.

Kodialam:2003:DRR

[KL03] Murali Kodialam and T. V. Lakshman. Dynamic routing of restorable bandwidth-guaranteed tunnels using aggregated network resource usage information. IEEE/ACM Transactions on Networking, 11(3):399–410, June 2003. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). See comments [LRJ08].

Kang:2007:MBE

[KL07] Seong-Ryong Kang and Dmitri Loguinov. Modeling best-effort and FEC streaming of scalable video in lossy network channels. IEEE/ACM Transactions on Networking, 15(1):187– 200, February 2007. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Kencl:2008:ALS

[KL08] Lukas Kencl and Jean-Yves Le Boudec. Adaptive load sharing for network processors. IEEE/ACM Transactions on Networking, 16 (2):293–306, April 2008. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Kuppuswamy:2009:AAE

[KL09] Kalyan Kuppuswamy and Daniel C. Lee. An analytic approach to efficiently computing call blocking probabilities for multiclass WDM networks. IEEE/ACM Transactions on Networking, 17(2):658–670, April 2009. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Kai:2012:ABP

[KL12] Cai Hong Kai and Soung Chang Liew. Applications of belief propagation in CSMA wireless networks. IEEE/ ACM Transactions on Networking, 20(4):1276–1289, August 2012. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Khakpour:2013:ITA

[KL13] Amir R. Khakpour and Alex X. Liu. An informationtheoretical approach to high-speed flow nature identification. IEEE/ ACM Transactions on Networking, 21(4):1076–1089, August 2013. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Kim:2015:CBR

[KLC15] Jihwan Kim, Hyang-Won Lee, and Song Chong. CSMA-based robust AP throughput guarantee under user distribution uncertainty. IEEE/ACM Transactions on Networking, 23 (3):782–795, June 2015. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Kao:2018:DUT

[KLC^{+}18] Sheng-Chun Kao, Ding-Yuan Lee, Ting-Sheng Chen, An-Yeu Wu, Sheng-Chun Kao, Ding-Yuan Lee, An-Yeu Wu, and Ting-Sheng Chen. Dynamically updatable ternary segmented aging Bloom filter for OpenFlow-compliant low-power packet processing. IEEE/ACM Transactions on Networking, 26 (2):1004–1017, April 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Kwak:2016:HOM

[KLE16] Jaewook Kwak, Chul-Ho Lee, and Do Young Eun. A high-order Markov-chain-based scheduling algorithm for low delay in CSMA networks. IEEE/ACM Transactions on Networking, 24(4):2278–2290, August 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Kim:2020:DCD

[KLK^{+}20] Seokhyun Kim, Kimin Lee, Yeonkeun Kim, Jinwoo Shin, Seungwon Shin, and Song Chong. Dynamic control for on-demand interference-managed WLAN infrastructures. IEEE/ ACM Transactions on Networking, 28(1):84–97, February 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL https://

dl.acm.org/doi/abs/10.
1109/TNET.2019.2953597.

Kvaternik:2016:MDS

[KLKP16] Karla Kvaternik, Jaime Llorca, Daniel Kilper, and Lacra Pavel. A methodology for the design of self-optimizing, decentralized content-caching strategies. IEEE/ACM Transactions on Networking, 24(5): 2634–2647, October 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Kozat:2016:OTV

[KLKT16] Ulas C. Kozat, Guanfeng Liang, Koray Kokten, and Janos Tapolcai. On optimal topology verification and failure localization for software defined networks. IEEE/ACM Transactions on Networking, 24(5):2899– 2912, October 2016. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Kuo:2018:DCV

[KLLT18] Tung-Wei Kuo, Bang-Heng Liou, Kate Ching-Ju Lin, and Ming-Jer Tsai. Deploying chains of virtual network functions: On the relation between link and server usage. IEEE/ ACM Transactions on Networking, 26(4):1562–1576, August 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Kodialam:2011:OST

[KLMW11] Murali Kodialam, T. V. Lakshman, Sarit Mukherjee, and Limin Wang. Online scheduling of targeted advertisements for IPTV. IEEE/ACM Transactions on Networking, 19(6):1825– 1834, December 2011. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Kristol:1993:PAG

[KLNS93] David M. Kristol, David Lee, Arun N. Netravali, and Krishan Sabnani. A polynomial algorithm for gateway generation from formal specifications. IEEE/ ACM Transactions on Networking, 1(2):217–229, April 1993. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/
                                        journals/ton/1993-1-2/
                                        p217-kristol/.

Korilis:1997:ANO

[KLO97] Yannis A. Korilis, Aurel A. Lazar, and Ariel Orda. Achieving network optima using Stackelberg routing strategies. IEEE/ ACM Transactions on Networking, 5(1):161–173, February 1997. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1997-5-1/
p161-korilis/.

Kodialam:2009:ORH

[KLOS09] Murali Kodialam, T. V.

Lakshman, James B. Orlin, and Sudipta Sengupta. Oblivious routing of highly variable traffic in service overlays and IP backbones. IEEE/ACM Transactions on Networking, 17(2):459– 472, April 2009. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Kodialam:2011:EER

[KLOS11] M. Kodialam, T. V. Lakshman, James B. Orlin, and Sudipta Sengupta. Endto-end restorable oblivious routing of hose model traffic. IEEE/ACM Transactions on Networking, 19(4): 1223–1236, August 2011. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Kucera:2016:ICI

[KLP16] Stepan Kucera and David Lopez-Perez. Inter-cell interference coordination for control channels in LTE heterogeneous networks. IEEE/ACM Transactions on Networking, 24(5):2872– 2884, October 2016. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Kohler:2006:OSA

[KLPS06] Eddie Kohler, Jinyang Li, Vern Paxson, and Scott Shenker. Observed structure of addresses in IP traffic. IEEE/ACM Transactions on Networking, 14(6): 1207–1218, December 2006. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Kulkarni:2020:RAE

[KLR^{+}20] Sameer G. Kulkarni, Guyue Liu, K. K. Ramakrishnan, Mayutan Arumaithurai, Timothy Wood, and Xiaoming Fu. REINFORCE: Achieving efficient failure resiliency for network function virtualizationbased services. IEEE/ ACM Transactions on Networking, 28(2):695–708, April 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL https://

                                        dl.acm.org/doi/abs/10.
                                        1109/TNET.2020.2969961.

Kodialam:2003:OMR

[KLS03] Murali Kodialam, T. V. Lakshman, and Sudipta Sengupta. Online multicast routing with bandwidth guarantees: a new approach using multicast network flow. IEEE/ACM Transactions on Networking, 11(4):676–686, August 2003. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Kodialam:2009:GPR

[KLS09a] Murali Kodialam, T. V.

Lakshman, and Sudipta Sengupta. Guaranteed performance routing of unpredictable traffic with fast path restoration. IEEE/ ACM Transactions on Networking, 17(5):1427–1438, October 2009. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Kodialam:2009:LRR

[KLS09b] Murali Kodialam, T. V.

Lakshman, and Sudipta Sengupta. Locally restorable routing of highly variable traffic. IEEE/ACM Transactions on Networking, 17 (3):752–763, June 2009. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Kim:2011:OAT

[KLS11a] Joohwan Kim, Xiaojun Lin, and Ness B. Shroff. Optimal anycast technique for delay-sensitive energyconstrained asynchronous sensor networks. IEEE/ ACM Transactions on Networking, 19(2):484–497, April 2011. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Kodialam:2011:TOR

[KLS11b] Murali Kodialam, T. V. Lakshman, and Sudipta Sengupta. Traffic-oblivious routing in the hose model. IEEE/ACM Transactions on Networking, 19(3):774– 787, June 2011. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Kim:2010:MDM

[KLSS10] Joohwan Kim, Xiaojun Lin, Ness B. Shroff, and Prasun Sinha. Minimizing delay and maximizing lifetime for wireless sensor networks with anycast. IEEE/ACM Transactions on Networking, 18(2):515– 528, April 2010. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Kompella:2012:RSF

[KLSV12] Ramana Rao Kompella, Kirill Levchenko, Alex C. Snoeren, and George Varghese. Router support for fine-grained latency measurements. IEEE/ACM Transactions on Networking, 20(3):811–824, June 2012. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Kuo:2015:MSS

[KLT15] Tung-Wei Kuo, Kate Ching-Ju Lin, and Ming-Jer Tsai. Maximizing submodular set function with connectivity constraint: theory and application to networks. IEEE/ACM Transactions on Networking, 23 (2):533–546, April 2015. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Kimura:2019:ICM

[KLVL19] Bruno Y. L. Kimura, Demetrius C. S. F. Lima, Leandro A. Villas, and Antonio A. F. Loureiro. Interpath contention in MultiPath TCP disjoint paths. IEEE/ACM Transactions on Networking, 27(4):1387– 1400, August 2019. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Kong:2023:TMU

[KLY^{+}23] Hao Kong, Li Lu, Jiadi Yu, Yingying Chen, Xiangyu Xu, and Feng Lyu. Toward multi-user authentication using WiFi signals. IEEE/ACM Transactions on Networking, 31(5): 2117–2132, 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2023.3237686.

Keung:2012:IDM

[KLZ12] Gabriel Y. Keung, Bo Li, and Qian Zhang. The intrusion detection in mobile sensor network. IEEE/ ACM Transactions on Networking, 20(4):1152–1161, August 2012. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Koetter:2003:AAN

[KM03] Ralf Koetter and Muriel Médard. An algebraic approach to network coding. IEEE/ACM Transactions on Networking, 11 (5):782–795, October 2003. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Kirsch:2008:SSH

[KM08] Adam Kirsch and Michael Mitzenmacher. Simple summaries for hashing with choices. IEEE/ ACM Transactions on Networking, 16(1):218–231, February 2008. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Kirsch:2010:POM

[KM10] Adam Kirsch and Michael Mitzenmacher. The power of one move: hashing schemes for hardware. IEEE/ACM Transactions on Networking, 18(6):1752– 1765, December 2010. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Kim:2012:TLP

[KMH12] Hayang Kim, Xiaoli Ma, and Benjamin Russell Hamilton. Tracking lowprecision clocks with timevarying drifts using Kalman filtering. IEEE/ACM Transactions on Networking, 20(1):257–270, February 2012. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Kompella:2009:PSR

[KMHS09] Sastry Kompella, Shiwen Mao, Y. Thomas Hou, and Hanif D. Sherali. On path selection and rate allocation for video in wireless mesh networks. IEEE/ ACM Transactions on Networking, 17(1):212–224, February 2009. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Kanakia:1995:ACC

[KMR95] Hemant Kanakia, Partho P.

Mishra, and Amy R. Reibman. An adaptive congestion control scheme for real time packet video transport. IEEE/ACM Transactions on Networking, 3(6):671–682, December 1995. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/
                                        journals/ton/1995-3-6/
                                        p671-kanakia/.

Katevenis:2001:WIC

[KMS^{+}01] Manolis G. H. Katevenis, Iakovos Mavroidis, Georgios Sapountzis, Eva Kalyvianaki, Ioannis Mavroidis, and Georgios Glykopoulos. Wormhole IP over (connectionless) ATM. IEEE/ ACM Transactions on Networking, 9(5):650–661, October 2001. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Kabamba:2005:RAR

[KMT05] Pierre T. Kabamba, Semyon M. Meerkov, and Choon Yik Tang. Ranking and adaptive ranking CDMA. IEEE/ACM Transactions on Networking, 13(3):622–635, June 2005. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Kokku:2012:NSV

[KMZR12] Ravi Kokku, Rajesh Mahindra, Honghai Zhang, and Sampath Rangarajan. NVS: a substrate for virtualizing wireless resources in cellular networks. IEEE/ ACM Transactions on Networking, 20(5):1333–1346, October 2012. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Kodialam:2005:CAR

[KN05] Murali Kodialam and Thyaga Nandagopal. Characterizing achievable rates in multi-hop wireless mesh networks with orthogonal channels. IEEE/ACM Transactions on Networking, 13(4):868–880, August 2005. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Kogan:2017:EFR

[KNE^{+}17] Kirill Kogan, Sergey I.

Nikolenko, Patrick Eugster, Alexander Shalimov, and Ori Rottenstreich. Efficient FIB representations on distributed platforms. IEEE/ ACM Transactions on Networking, 25(6):3309–3322, December 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Kompella:2014:CCU

[KNK^{+}14] Sastry Kompella, Gam D.

Nguyen, Clement Kam, Jeffrey E. Wieselthier, and Anthony Ephremides. Cooperation in cognitive underlay networks: stable throughput tradeoffs. IEEE/ACM Transactions on Networking, 22(6):1756– 1768, December 2014. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Kerivin:2005:DCS

[KNP05] Hervé Kerivin, Dritan Nace, and Thi-Tuyet-Loan Pham. Design of capacitated survivable networks with a single facility. IEEE/ACM Transactions on Networking, 13(2):248– 261, April 2005. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Kogan:2016:EOI

[KNR^{+}16] Kirill Kogan, Sergey I. Nikolenko, Ori Rottenstreich, William Culhane, and Patrick Eugster. Exploiting order independence for scalable and expressive packet classification. IEEE/ACM Transactions on Networking, 24(2):1251– 1264, April 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Kim:2013:TOC

[KNSV13] Tae Hyun Kim, Jian Ni, R. Srikant, and Nitin H. Vaidya. Throughputoptimal CSMA with imperfect carrier sensing. IEEE/ ACM Transactions on Networking, 21(5):1636–1650, October 2013. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Karbasi:2013:RLI

[KO13] Amin Karbasi and Sewoong Oh. Robust localization from incomplete local information. IEEE/ ACM Transactions on Networking, 21(4):1131–1144, August 2013. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Koksal:2010:RQS

[Kok10] Can Emre Koksal. Rate quantization and the speedup required to achieve 100 throughput for multicast over crossbar switches. IEEE/ACM Transactions on Networking, 18(4):1207– 1219, August 2010. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Konorski:2006:GTS

[Kon06] Jerzy Konorski. A gametheoretic study of CSMA/ CA under a backoff attack. IEEE/ACM Transactions on Networking, 14(6): 1167–1178, December 2006. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Koppelman:1996:CBN

[Kop96] David M. Koppelman.

Congested Banyan network analysis using congestedqueue states and neighboringqueue effects. IEEE/ ACM Transactions on Networking, 4(1):106–111, February 1996. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1996-4-1/
p106-koppelman/.

Kay:1996:PRP

[KP96] Jonathan Kay and Joseph Pasquale. Profiling and reducing processing overheads in TCP/IP. IEEE/ ACM Transactions on Networking, 4(6):817–828, December 1996. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1996-4-6/
p817-kay/.

Kirner:2021:QAI

[KP21] Raimund Kirner and Peter Puschner. A quantitative analysis of interfaces to time-triggered communication buses. IEEE/ ACM Transactions on Networking, 29(4):1786–1797, August 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2021.3073460.

Kalor:2023:TMD

[KP23] Anders E. Kalør and Petar Popovski. Timely monitoring of dynamic sources with observations from multiple wireless sensors. IEEE/ACM Transactions on Networking, 31(3):1263– 1276, June 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2022.3212794.

Kim:2016:LBT

[KPK^{+}16] Dongmyoung Kim, Taejun Park, Seongwon Kim, Hyoil Kim, and Sunghyun Choi. Load balancing in two-tier cellular networks with open and hybrid access femtocells. IEEE/ACM Transactions on Networking, 24(6): 3397–3411, December 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Kompella:1993:MRM

[KPP93] Vachaspathi P. Kompella, Joseph C. Pasquale, and George C. Polyzos. Multicast routing for multimedia communication. IEEE/ACM Transactions on Networking, 1(3):286– 292, June 1993. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/

                                        journals/ton/1993-1-3/
                                        p286-kompella/.

Kulkarni:1998:PAR

[KqL98] Lalita A. Kulkarni and San qi Li. Performance analysis of a rate-based feedback control scheme. IEEE/ ACM Transactions on Networking, 6(6):797–810, December 1998. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/
                                        journals/ton/1998-6-6/
                                        p797-kulkarni/.

Kim:1999:PAD

[KqL99] Yonghwan Kim and San qi Li. Performance analysis of data packet discarding in ATM networks. IEEE/ACM Transactions on Networking, 7(2):216– 227, April 1999. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/
                                        journals/ton/1999-7-2/
                                        p216-kim/.

Kolarov:1999:CTA

[KR99] Aleksandar Kolarov and G. Ramamurthy. A control-theoretic approach to the design of an explicit rate controller for ABR service. IEEE/ ACM Transactions on Networking, 7(5):741–753, October 1999. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1999-7-5/
p741-kolarov/.

Kannan:2000:MMM

[KR00] Rajgopal Kannan and Sibabrata Ray. MSXmin: a modular multicast ATM packet switch with low delay and hardware complexity. IEEE/ACM Transactions on Networking, 8(3): 407–418, 2000. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/2000-8-3/
p407-kannan/.

Ko:2005:DSS

[KR05] Bong-Jun Ko and Dan Rubenstein. Distributed self-stabilizing placement of replicated resources in emerging networks. IEEE/ ACM Transactions on Networking, 13(3):476–487, June 2005. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Kim:2008:STD

[KR08] Seong Soo Kim and A. L. Narasimha Reddy. Statistical techniques for detecting traffic anomalies through packet header data. IEEE/ACM Transactions on Networking, 16(3):562–575, June 2008. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Konstantinidis:2020:RDS

[KR20] Konstantinos Konstantinidis and Aditya Ramamoorthy. Resolvable designs for speeding up distributed computing. IEEE/ACM Transactions on Networking, 28(4):1657–1670, August 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2020.2992989.

Katti:2008:XAP

[KRH^{+}08] Sachin Katti, Hariharan Rahul, Wenjun Hu, Dina Katabi, Muriel Médard, and Jon Crowcroft. XORs in the air: practical wireless network coding. IEEE/ACM Transactions on Networking, 16(3):497– 510, June 2008. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Kriegleder:2014:CAA

[Kri14] Maximilian Kriegleder. A correction to algorithm A2 in “Asynchronous distributed averaging on communication networks”. IEEE/ACM Transactions on Networking, 22(6):2026– 2027, December 2014. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). See [MSP^{+}07].

Kini:2010:FRD

[KRKH10] Shrinivasa Kini, Srinivasan Ramasubramanian, Amund Kvalbein, and Audun Fosselie Hansen. Fast recovery from duallink or single-node failures in IP networks using tunneling. IEEE/ACM Transactions on Networking, 18 (6):1988–1999, December 2010. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Kamal:2011:OPA

[KRLL11] Ahmed E. Kamal, Aditya Ramamoorthy, Long Long, and Shizheng Li. Overlay protection against link failures using network coding. IEEE/ACM Transactions on Networking, 19(4): 1071–1084, August 2011. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Kang:2017:ACN

[KRRR17] Nanxi Kang, Ori Rottenstreich, Sanjay G. Rao, and Jennifer Rexford. Alpaca: Compact network policies with attribute-encoded addresses. IEEE/ACM Transactions on Networking, 25(3):1846–1860, June 2017. CODEN IEANEP.

ISSN 1063-6692 (print), 1558-2566 (electronic).

Krishnan:2000:CLP

[KRS00] P. Krishnan, Danny Raz, and Yuval Shavitt. The cache location problem. IEEE/ACM Transactions on Networking, 8(5):568– 582, 2000. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/
                                        journals/ton/2000-8-5/
                                        p568-krishnan/.

Kanizo:2017:OVB

[KRS^{+}17] Yossi Kanizo, Ori Rottenstreich, Itai Segall, Jose Yallouz, Yossi Kanizo, Ori Rottenstreich, Itai Segall, and Jose Yallouz. Optimizing virtual backup allocation for middleboxes. IEEE/ACM Transactions on Networking, 25(5):2759– 2772, October 2017. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Kumar:2002:APV

[KRSY02] Amit Kumar, Rajeev Rastogi, Avi Silberschatz, and Bulent Yener. Algorithms for provisioning virtual private networks in the hose model. IEEE/ACM Transactions on Networking, 10 (4):565–578, August 2002. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Katzela:1995:SFI

[KS95] Irene Katzela and Mischa Schwartz. Schemes for fault identification in communication networks. IEEE/ ACM Transactions on Networking, 3(6):753–764, December 1995. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1995-3-6/
p753-katzela/.

Kweon:1998:PDD

[KS98] Seok-Kyu Kweon and Kang G. Shin. Providing deterministic delay guarantees in ATM networks. IEEE/ACM Transactions on Networking, 6(6):838–850, December 1998. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1998-6-6/
p838-kweon/.

Kim:2001:LPC

[KS01a] Han S. Kim and Ness B.

Shroff. Loss probability calculations and asymptotic analysis for finite buffer multiplexers. IEEE/ ACM Transactions on Networking, 9(6):755–768, December 2001. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Krishnaswamy:2001:DLT

[KS01b] Rajesh M. Krishnaswamy and Kumar N. Sivarajan. Design of logical topologies: a linear formulation for wavelengthrouted optical networks with no wavelength changers. IEEE/ACM Transactions on Networking, 9(2): 186–198, 2001. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/articles/
                                        journals/ton/2001-9-2/
                                        p186-krishnaswamy/p186-
                                        krishnaswamy.pdf;   http:
                                        //www.acm.org/pubs/citations/
                                        journals/ton/2001-9-2/
                                        p186-krishnaswamy/.

Kunniyur:2003:EEC

[KS03] Srisankar Kunniyur and R. Srikant. End-to-end congestion control schemes: utility functions, random losses and ECN marks. IEEE/ACM Transactions on Networking, 11(5):689– 702, October 2003. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Kunniyur:2004:AVQ

[KS04] Srisankar S. Kunniyur and R. Srikant. An adaptive virtual queue (AVQ) algorithm for active queue management. IEEE/ACM Transactions on Networking, 12(2):286–299, April 2004. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Katz:2006:RPT

[KS06] Michael Katz and Shlomo Shamai. Relaying protocols for two colocated users. IEEE/ACM Transactions on Networking, 14 (SI):2329–2344, June 2006. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Kim:2009:AAA

[KS09a] Kyu-Han Kim and Kang G.

Shin. On accurate and asymmetry-aware measurement of link quality in wireless mesh networks. IEEE/ ACM Transactions on Networking, 17(4):1172–1185, August 2009. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Kwon:2009:ASP

[KS09b] Sungoh Kwon and Ness B.

Shroff. Analysis of shortest path routing for large multi-hop wireless networks. IEEE/ACM Transactions on Networking, 17 (3):857–869, June 2009. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Kasbekar:2010:SAF

[KS10] Gaurav S. Kasbekar and Saswati Sarkar. Spectrum auction framework for access allocation in cognitive radio networks. IEEE/ ACM Transactions on Networking, 18(6):1841–1854, December 2010. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Kim:2011:SRW

[KS11] Kyu-Han Kim and Kang G. Shin. Self-reconfigurable wireless mesh networks. IEEE/ACM Transactions on Networking, 19(2):393– 404, April 2011. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Kumar:2012:DME

[KS12] Ashwini Kumar and Kang G. Shin. DSASync: managing end-to-end connections in dynamic spectrum access wireless LANs. IEEE/ ACM Transactions on Networking, 20(4):1068–1081, August 2012. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Krishnan:2013:VSQ

[KS13] S. Shunmuga Krishnan and Ramesh K. Sitaraman. Video stream quality impacts viewer behavior: inferring causality using quasi-experimental designs. IEEE/ACM Transactions on Networking, 21 (6):2001–2014, December 2013. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Khabbazian:2019:GEA

[KS19] Majid Khabbazian and Keyvan Gharouni Saffar. The gain of energy accumulation in multi-hop wireless network broadcast. IEEE/ ACM Transactions on Networking, 27(5):1830–1844, October 2019. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Key:2020:PCC

[KS20] Peter Key and Richard Steinberg. Pricing, competition and content for Internet service providers. IEEE/ACM Transactions on Networking, 28(5):2285– 2298, October 2020. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL https:/

/dl.acm.org/doi/10.1109/
TNET.2020.3010550.

Khouzani:2012:MDM

[KSA12] M. H. R. Khouzani, Saswati Sarkar, and Eitan Altman. Maximum damage malware attack in mobile wireless networks. IEEE/ ACM Transactions on Networking, 20(5):1347–1360, October 2012. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Keshtkarjahromi:2018:DDN

[KSAK18] Yasaman Keshtkarjahromi, Hulya Seferoglu, Rashid Ansari, and Ashfaq Khokhar. Device-to-device networking meets cellular via network coding. IEEE/ ACM Transactions on Networking, 26(1):370–383, February 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Kong:2023:CAD

[KSC^{+}23] Dezhang Kong, Yi Shen, Xiang Chen, Qiumei Cheng, Hongyan Liu, Dong Zhang, Xuan Liu, Shuangxi Chen, and Chunming Wu. Combination attacks and defenses on SDN topology discovery. IEEE/ACM Transactions on Networking, 31(2):904– 919, April 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2022.3203561.

Kalafut:2011:TDO

[KSG11] Andrew J. Kalafut, Craig A. Shue, and Minaxi Gupta. Touring DNS open houses for trends and configurations. IEEE/ACM Transactions on Networking, 19 (6):1666–1675, December 2011. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Kotnis:2017:ICS

[KSK17] Bhushan Kotnis, Albert Sunny, and Joy Kuri. Incentivized campaigning in social networks. IEEE/ ACM Transactions on Networking, 25(3):1621–1634, June 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Kwak:2005:PAE

[KSM05] Byung-Jae Kwak, Nah-Oak Song, and Leonard E. Miller. Performance analysis of exponential backoff. IEEE/ACM Transactions on Networking, 13(2):343– 355, April 2005. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Kadota:2019:SAO

[KSM19] Igor Kadota, Abhishek Sinha, and Eytan Modiano. Scheduling algorithms for optimizing age of information in wireless networks with throughput constraints. IEEE/ ACM Transactions on Networking, 27(4):1359–1372, August 2019. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Karakoc:2020:MLD

[KSNR20] Nurullah Karakoç, Anna Scaglione, Angelia Nedić, and Martin Reisslein. Multilayer decomposition of network utility maximization problems. IEEE/ACM Transactions on Networking, 28(5):2077–2091, October 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2020.3003925.

Karakoc:2022:FEN

[KSRW22] Nurullah Karakoç, Anna Scaglione, Martin Reisslein, and Ruiyuan Wu. Federated edge network utility maximization for a multi-server system: Algorithm and convergence. IEEE/ACM Transactions on Networking, 30(5):2002– 2017, 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2022.3156530.

Khochare:2024:IAC

[KSSD24] Aakash Khochare, Francesco Betti Sorbelli, Yogesh Simmhan, and Sajal K. Das. Improved algorithms for coscheduling of edge analytics and routes for UAV fleet missions. IEEE/ACM Transactions on Networking, 32(1):17–33, 2024. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL https:/

/dl.acm.org/doi/10.1109/
TNET.2023.3277810.

Kleinheksel:2018:EFT

[KSSK18] Cory J. Kleinheksel, Arun K.

Somani, Arun K. Somani, and Cory J. Kleinheksel. Enhancing fault tolerance and resource utilization in unidirectional quorum-based cycle routing. IEEE/ACM Transactions on Networking, 26 (2):934–947, April 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Kadota:2018:SPM

[KSUB^{+}18] Igor Kadota, Abhishek Sinha, Elif Uysal-Biyikoglu, Rahul Singh, and Eytan Modiano. Scheduling policies for minimizing age of information in broadcast wireless networks. IEEE/ ACM Transactions on Networking, 26(6):2637–2650, December 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Kompella:2007:SAD

[KSV07] Ramana Rao Kompella, Sumeet Singh, and George Varghese. On scalable attack detection in the network. IEEE/ACM Transactions on Networking, 15 (1):14–25, February 2007. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Koutsopoulos:2006:CLA

[KT06] Iordanis Koutsopoulos and Leandros Tassiulas. Crosslayer adaptive techniques for throughput enhancement in wireless OFDMbased networks. IEEE/ ACM Transactions on Networking, 14(5):1056–1066, October 2006. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Koutsopoulos:2007:JOA

[KT07] Iordanis Koutsopoulos and Leandros Tassiulas. Joint optimal access point selection and channel assignment in wireless networks. IEEE/ACM Transactions on Networking, 15(3):521– 532, June 2007. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Kwong:2008:BHP

[KT08] Kin-Wah Kwong and H. K. Tsang. Building heterogeneous peer-to-peer networks: protocol and analysis. IEEE/ACM Transactions on Networking, 16 (2):281–292, April 2008. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Katrinis:2011:DWO

[KT11] Kostas M. Katrinis and Anna Tzanakaki. On the dimensioning of WDM optical networks with impairment-aware regeneration. IEEE/ACM Transactions on Networking, 19 (3):735–746, June 2011. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

KlosneeMuller:2018:CAH

[KTvdSK18] Sabrina Klos nee Muller, Cem Tekin, Mihaela van der Schaar, and Anja Klein. Context-aware hierarchical online learning for performance maximization in mobile crowdsourcing. IEEE/ACM Transactions on Networking, 26(3):1334– 1347, June 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Kucera:2014:ECC

[Kuc14]  Stepan Kucera.  En-

abling co-channel smallcell deployments in SINRconstraint networks by distributed monitoring of normalized network capacity. IEEE/ACM Transactions on Networking, 22(5):1577– 1590, October 2014. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Kumar:1998:CPA

[Kum98]  Anurag Kumar.  Com-

parative performance analysis of versions of TCP in a local network with a lossy link. IEEE/ ACM Transactions on Networking, 6(4):485–498, August 1998. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1998-6-4/
p485-kumar/.

Kabatepe:1996:FDQ

[KV96] Mete Kabatepe and Kenneth S. Vastola. The fair distributed queue (FDQ) protocol for highspeed metropolitan-area networks. IEEE/ACM Transactions on Networking, 4(3):331–339, June 1996. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL

http://www.acm.org/pubs/
citations/journals/ton/
1996-4-3/p331-kabatepe/

.

Kalampoukas:1998:ASP

[KV98] Lampros Kalampoukas and Anujan Varma. Analysis of source policy and its effects on TCP in rate-controlled ATM networks. IEEE/ ACM Transactions on Networking, 6(5):599–610, October 1998. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1998-6-5/
p599-kalampoukas/.

Kuipers:2005:CIC

[KV05] Fernando A. Kuipers and Piet F. A. Van Mieghem. Conditions that impact the complexity of QoS routing. IEEE/ACM Transactions on Networking, 13 (4):717–730, August 2005. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Kyasanur:2009:CMW

[KV09] Pradeep Kyasanur and Nitin H. Vaidya. Capacity of multichannel wireless networks under the protocol model. IEEE/ACM Transactions on Networking, 17(2):515–527, April 2009. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Khanna:2012:ASV

[KVF^{+}12] Sanjeev Khanna, Santosh S. Venkatesh, Omid Fatemieh, Fariba Khan, and Carl A. Gunter. Adaptive selective verification: an efficient adaptive countermeasure to thwart DoS attacks. IEEE/ACM Transactions on Networking, 20(3):715–728, June 2012. CODEN IEANEP.

ISSN 1063-6692 (print), 1558-2566 (electronic).

Kalampoukas:1998:TTT

[KVR98] Lampros Kalampoukas, Anujan Varma, and K. K. Ramakrishnan. Two-way TCP traffic over rate controlled channels: effects and analysis. IEEE/ ACM Transactions on Networking, 6(6):729–743, December 1998. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1998-6-6/
p729-kalampoukas/.

Kalampoukas:2002:EWA

[KVR02] Lampros Kalampoukas, Anujan Varma, and K. K. Ramakrishnan. Explicit window adaptation: a method to enhance TCP performance. IEEE/ACM Transactions on Networking, 10(3):338–350, June 2002. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Kuo:2017:ROO

[KW17] Wei-Cheng Kuo and Chih-Chun Wang. Robust and optimal opportunistic scheduling for downlink two-flow network coding with varying channel quality and rate adaptation. IEEE/ACM Transactions on Networking, 25(1):

465–479, February 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Klinkowski:2019:EOF

[KW19] Miroslaw Klinkowski and Krzysztof Walkowiak. An efficient optimization framework for solving RSSA problems in spectrally and spatially flexible optical networks. IEEE/ ACM Transactions on Networking, 27(4):1474–1486, August 2019. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Kesidis:1993:EBM

[KWC93] George Kesidis, Jean Walrand, and Cheng-Shang Chang. Effective bandwidths for multiclass Markov fluids and other ATM sources. IEEE/ACM Transactions on Networking, 1(4):424–428, August 1993. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1993-1-4/
p424-kesidis/.

Kannan:2010:AAM

[KWCR10] Rajgopal Kannan, Shuangqing Wei, Vasu Chakravarthy, and Muralidhar Rangaswamy. Approximation algorithms for minimum energy transmission scheduling in rate and duty-cycle constrained wireless networks. IEEE/ ACM Transactions on Networking, 18(1):296–306, February 2010. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Kompella:2010:OSB

[KWE^{+}10] Sastry Kompella, Jeffrey E. Wieselthier, Anthony Ephremides, Hanif D. Sherali, and Gam D. Nguyen. On optimal SINRbased scheduling in multihop wireless networks. IEEE/ACM Transactions on Networking, 18(6):1713– 1724, December 2010. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Koutsonikolas:2011:ENC

[KWH11] Dimitrios Koutsonikolas, Chih-Chun Wang, and Y. Charlie Hu. Efficient network-coding-based opportunistic routing through cumulative coded acknowledgments. IEEE/ACM Transactions on Networking, 19(5):1368–1381, October 2011. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Kim:2017:ESC

[KWH^{+}17] Song Min Kim, Shuai Wang, Tian He, Song Min Kim, Shuai Wang, and Tian He. Exploiting spatiotemporal correlation for wireless networks under interference. IEEE/ACM Transactions on Networking, 25(5):3132–3145, October 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Kang:2016:PLD

[KWJY16] Xiaohan Kang, Weina Wang, Juan José Jaramillo, and Lei Ying. On the performance of largest-deficit-first for scheduling realtime traffic in wireless networks. IEEE/ACM Transactions on Networking, 24 (1):72–84, February 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Khreishah:2010:RCP

[KWS10] Abdallah Khreishah, Chih-Chun Wang, and Ness B. Shroff. Rate control with pairwise intersession network coding. IEEE/ACM Transactions on Networking, 18(3):816–829, June 2010. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Kim:2011:MDL

[KWS^{+}11] Donghyun Kim, Wei Wang, Nassim Sohaee, Changcun Ma, Weili Wu, Wonjun Lee, and Ding-Zhu Du. Minimum data-latency-bound k-sink placement problem in wireless sensor networks. IEEE/ACM Transactions on Networking, 19(5):1344– 1353, October 2011. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Kuhn:2008:AAG

[KWZ08] Fabian Kuhn, Roger Wattenhofer, and Aaron Zollinger. An algorithmic approach to geographic routing in ad hoc and sensor networks. IEEE/ACM Transactions on Networking, 16 (1):51–62, February 2008. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Kamran:2022:DJC

[KYM22] Khashayar Kamran, Edmund Yeh, and Qian Ma. DECO: Joint computation scheduling, caching, and communication in dataintensive computing networks. IEEE/ACM Transactions on Networking, 30 (3):1058–1072, June 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1109/TNET.2021.3136157.

Kone:2012:EOS

[KYY^{+}12] Vinod Kone, Lei Yang, Xue Yang, Ben Y. Zhao, and Haitao Zheng. The effectiveness of opportunistic spectrum access: a measurement study. IEEE/ ACM Transactions on Networking, 20(6):2005–2016, December 2012. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Knightly:1997:DAT

[KZ97] Edward W. Knightly and Hui Zhang. D-BIND: an accurate traffic model for providing QoS guarantees to VBR traffic. IEEE/ACM Transactions on Networking, 5(2):219– 231, April 1997. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1997-5-2/
p219-knightly/.

Krithikaivasan:2007:ABT

[KZDM07] Balaji Krithikaivasan, Yong Zeng, Kaushik Deka, and Deep Medhi. ARCHbased traffic forecasting and dynamic bandwidth provisioning for periodically measured nonstationary traffic. IEEE/ACM Transactions on Networking, 15(3):683–696, June 2007. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Kulkarni:2020:NDB

[KZH^{+}20] Sameer G. Kulkarni, Wei Zhang, Jinho Hwang, Shriram Rajagopalan, K. K. Ramakrishnan, Timothy Wood, Mayutan Arumaithurai, and Xiaoming Fu. NFVnice: Dynamic backpressure and scheduling for NFV service chains. IEEE/ACM Transactions on Networking, 28(2):639– 652, April 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL https://

dl.acm.org/doi/abs/10.
1109/TNET.2020.2969971.

Lea:1995:BQS

[LA95a] Chin-Tau Lea and Anwar Alyatama. Bandwidth quantization and states reduction in the broadband ISDN. IEEE/ ACM Transactions on Networking, 3(3):352–360, June 1995. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1995-3-3/
p352-lea/.

Lee:1995:CLA

[LA95b] Myung J. Lee and David S. Ahn. Cell loss analysis and design trade-offs of nonblocking ATM switches with nonuniform traffic. IEEE/ACM Transactions on Networking, 3(2):199– 210, April 1995. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1995-3-2/
p199-lee/.

Levine:1995:PMA

[LA95c] David A. Levine and Ian F. Akyildiz. PROTON: a media access control protocol for optical networks with star topology. IEEE/ACM Transactions on Networking, 3(2):158– 168, April 1995. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1995-3-2/
p158-levine/.

La:2002:UBR

[LA02] Richard J. La and Venkat Anantharam. Utility-based rate control in the Internet for elastic traffic. IEEE/ACM Transactions on Networking, 10(2):272– 286, April 2002. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

La:2016:ISS

[La16] Richard J. La. Interdependent security with strategic agents and cascades of infection. IEEE/ACM Transactions on Networking, 24 (3):1378–1391, June 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

La:2017:EDC

[La17] Richard J. La. Effects of degree correlations in interdependent security: Good or bad? IEEE/ACM Transactions on Networking, 25(4): 2484–2497, August 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Labourdette:1997:PIP

[Lab97] Jean-François P. Labourdette. Performance impact of partial reconfiguration on multihop lightwave networks. IEEE/ ACM Transactions on Networking, 5(3):351–358, June 1997. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/
                                        journals/ton/1997-5-3/
                                        p351-labourdette/.

Labovitz:2001:DIR

[LABJ01] Craig Labovitz, Abha Ahuja, Abhijit Bose, and Farnam Jahanian. Delayed Internet routing convergence. IEEE/ACM Transactions on Networking, 9 (3):293–306, June 2001. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Liu:2020:TFC

[LAJ20] Teng Liu, Alhussein A.

Abouzeid, and A. Agung Julius. Traffic flow control in vehicular multihop networks with data caching and infrastructure support. IEEE/ ACM Transactions on Networking, 28(1):376–386, February 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL https://

dl.acm.org/doi/abs/10.
1109/TNET.2020.2963930.

Le:2007:EAQ

[LAJS07] Long Le, Jay Aikat, Kevin Jeffay, and F. Donelson Smith. The effects of active queue management and explicit congestion notification on Web performance. IEEE/ACM Transactions on Networking, 15(6):1217– 1230, December 2007. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Liu:2024:FLG

[LAL^{+}24] Chenyi Liu, Vaneet Aggarwal, Tian Lan, Nan Geng, Yuan Yang, Mingwei Xu, and Qing Li. FERN: Leveraging graph attention networks for failure evaluation and robust network design. IEEE/ACM Transactions on Networking, 32(2):1003– 1018, April 2024. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2023.3311678.

Levine:1997:REC

[LAN97] David A. Levine, Ian F. Akyildiz, and Mahmoud Naghshineh. A resource estimation and call admission algorithm for wireless multimedia networks using the shadow cluster concept. IEEE/ACM Transactions on Networking, 5(1):1–12, February 1997. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/
                                        journals/ton/1997-5-1/
                                        p1-levine/.

Lakshminarayanan:2008:SUC

[LAPS08] Karthik Lakshminarayanan, Daniel Adkins, Adrian Perrig, and Ion Stoica. Securing user-controlled routing infrastructures. IEEE/ ACM Transactions on Networking, 16(3):549–561, June 2008. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Larrnaaga:2016:DCB

[LAV16] Maialen Larrnaaga, Urtzi Ayesta, and Ina Maria Verloop. Dynamic control of birth-and-death restless bandits: Application to resource-allocation problems. IEEE/ACM Transactions on Networking, 24(6): 3812–3825, December 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Liu:2004:FPA

[LB04] Mingyan Liu and John S.

Baras. Fixed point approximation for multirate multihop loss networks with state-dependent routing. IEEE/ACM Transactions on Networking, 12(2):361– 374, April 2004. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Langar:2008:CAM

[LBB08] Rami Langar, Nizar Bouabdallah, and Raouf Boutaba. A comprehensive analysis of mobility management in MPLS-based wireless access networks. IEEE/ACM Transactions on Networking, 16(4):918–931, August 2008. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Lazarou:2009:DNT

[LBFE09] Georgios Y. Lazarou, Julie Baca, Victor S. Frost, and Joseph B. Evans. Describing network traffic using the index of variability. IEEE/ACM Transactions on Networking, 17(5): 1672–1683, October 2009. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Lin:2020:IAP

[LBGL20] Michael Lin, Novella Bartolini, Michael Giallorenzo, and Thomas F. La Porta. On interference aware power adjustment and scheduling in femtocell networks. IEEE/ACM Transactions on Networking, 28 (2):736–749, April 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL

https://dl.acm.org/doi/
abs/10.1109/TNET.2020.
2973833.

Lee:2007:MCN

[LBHO07] Junsoo Lee, Stephan Bohacek, João P. Hespanha, and Katia Obraczka. Modeling communication networks with hybrid systems. IEEE/ACM Transactions on Networking, 15(3):630– 643, June 2007. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Li:2007:NCE

[LBL07] Chengzhi Li, Almut Burchard, and Jörg Liebeherr. A network calculus with effective bandwidth. IEEE/ ACM Transactions on Networking, 15(6):1442–1453, December 2007. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Lutu:2016:BVT

[LBP^{+}16] Andra Lutu, Marcelo Bagnulo, Cristel Pelsser, Olaf Maennel, and Jesus Cid-Sueiro. The BGP visibility toolkit: detecting anomalous Internet routing behavior. IEEE/ACM Transactions on Networking, 24 (2):1237–1250, April 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Li:2017:ISR

[LBP^{+}17] Jian Li, Rajarshi Bhattacharyya, Suman Paul, Srinivas Shakkottai, and Vijay Subramanian. Incentivizing sharing in realtime D2D streaming networks: a mean field game perspective. IEEE/ACM Transactions on Networking, 25 (1):3–17, February 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Labourdette:2005:FAD

[LBRA05] Jean-François Labourdette, Eric Bouillet, Ramu Ramamurthy, and Ahmet A. Akyamaç. Fast approximate dimensioning and performance analysis of mesh optical networks. IEEE/ACM Transactions on Networking, 13(4):906– 917, August 2005. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Lu:1999:FSW

[LBS99] Songwu Lu, Vaduvur Bharghavan, and R. Srikant. Fair scheduling in wireless packet networks. IEEE/ ACM Transactions on Networking, 7(4):473–489, August 1999. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/
                                        journals/ton/1999-7-4/
                                        p473-lu/.

Lakshmikantha:2005:RRV

[LBS05a] Ashvin Lakshmikantha, Carolyn L. Beck, and R. Srikant. Robustness of real and virtual queuebased active queue management schemes. IEEE/ACM Transactions on Networking, 13(1):81–93, February 2005. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Liu:2005:ERS

[LBS05b] Shao Liu, Tamer Başar, and R. Srikant. Exponential-RED: a stabilizing AQM scheme for low- and highspeed TCP protocols. IEEE/ACM Transactions on Networking, 13(5):1068– 1081, October 2005. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Lakshmikantha:2011:IFA

[LBS11] Ashvin Lakshmikantha, Carolyn Beck, and R. Srikant. Impact of file arrivals and departures on buffer sizing in core routers. IEEE/ ACM Transactions on Networking, 19(2):347–358, April 2011. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Li:2011:RPC

[LBX11] Bo Li, Cem Boyaci, and Ye Xia. A refined performance characterization of longest-queue-first policy in wireless networks. IEEE/ ACM Transactions on Networking, 19(5):1382–1395, October 2011. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Li:2020:AOL

[LBZ^{+}20] Hao Li, Zhengda Bian, Peng Zhang, Zhun Sun, Chengchen Hu, Qiang Fu, Tian Pan, and Jia Lv. Application-oblivious L7 parsing using recurrent neural networks. IEEE/ ACM Transactions on Networking, 28(5):2009–2022, October 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2020.3000430.

Lee:1994:DSF

[LC94a] Tsern-Huei Lee and Jin-Jye Chou. Diagnosis of single faults in bitonic sorters. IEEE/ ACM Transactions on Networking, 2(5):497–507, October 1994. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/
                                        journals/ton/1994-2-5/
                                        p497-lee/.

Luciani:1994:AMP

[LC94b] James V. Luciani and C. Y. Roger Chen. An analytical model for partially blocking finite-buffered switching networks. IEEE/ ACM Transactions on Networking, 2(5):533–540, October 1994. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/
                                        journals/ton/1994-2-5/
                                        p533-luciani/.

Liotopoulos:1996:SNO

[LC96] Fotios K. Liotopoulos and Suresh Chalasani. Semirearrangeably nonblocking operation of Clos networks in the multirate environment. IEEE/ ACM Transactions on Networking, 4(2):281–291, April 1996. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1996-4-2/
p281-liotopoulos/.

LaPorta:1997:DSS

[LC97] Thomas F. La Porta and Kuo-Wei Herman Chen. A direct signaling system for flexible access and deployment of telecommunication services. IEEE/ ACM Transactions on Networking, 5(4):489–501, August 1997. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1997-5-4/
p489-la_porta/.

LeBoudec:2003:PSR

[LC03] Jean-Yves Le Boudec and Anna Charny. Packet scale rate guarantee for non-FIFO nodes. IEEE/ACM Transactions on Networking, 11(5):810–820, October 2003. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Li:2004:MPE

[LC04a] Wei Li and Xiuli Chao.

Modeling and performance evaluation of a cellular mobile network. IEEE/ ACM Transactions on Networking, 12(1):131–145, February 2004. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Liao:2004:DCP

[LC04b] Raymond R.-F. Liao and Andrew T. Campbell. Dynamic core provisioning for quantitative differentiated services. IEEE/ACM Transactions on Networking, 12(3):429–442, June 2004. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Lu:2010:PUL

[LCB^{+}10] Guohan Lu, Yan Chen, Stefan Birrer, Fabián E. Bustamante, and Xing Li. POPI: a user-level tool for inferring router packet forwarding priority. IEEE/ACM Transactions on Networking, 18(1):1–14, February 2010. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Liu:2020:PSS

[LCC^{+}20] Jia Liu, Shigang Chen, Min Chen, Qingjun Xiao, and Lijun Chen. Pose sensing with a single RFID tag. IEEE/ACM Transactions on Networking, 28(5):2023–2036, October 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.

acm.org/doi/10.1109/TNET.
2020.3007830.

Li:2021:OPP

[LCDW21] Yunpeng Li, Costas A.

Courcoubetis, Lingjie Duan, and Richard Weber. Optimal pricing for peer-to-peer sharing with network externalities. IEEE/ ACM Transactions on Networking, 29(1):148–161, February 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2020.3029398.

Liu:2014:PPB

[LCG^{+}14] Changbin Liu, Ricardo Correa, Harjot Gill, Tanveer Gill, Xiaozhou Li, Shivkumar Muthukumar, Taher Saeed, Boon Thau Loo, and Prithwish Basu. PUMA: policy-based unified multiradio architecture for agile mesh networking. IEEE/ACM Transactions on Networking, 22(6): 1897–1910, December 2014. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Li:1995:LCA

[LCH95] San-Qi Li, Song Chong, and Chia-Lin Hwang. Link capacity allocation and network control by filtered input rate in highspeed networks. IEEE/ ACM Transactions on Networking, 3(1):10–25, February 1995. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/
                                        journals/ton/1995-3-1/
                                        p10-li/.

Lee:2006:NPS

[LCH^{+}06] David Lee, Dongluo Chen, Ruibing Hao, Raymond E. Miller, Jianping Wu, and Xia Yin. Network protocol system monitoring: a formal approach with passive testing. IEEE/ACM Transactions on Networking, 14 (2):424–437, April 2006. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Liao:2020:PTA

[LCH20a] Guocheng Liao, Xu Chen, and Jianwei Huang. Prospect theoretic analysis of privacypreserving mechanism. IEEE/ ACM Transactions on Networking, 28(1):71–83, February 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL https://

                                        dl.acm.org/doi/abs/10.
                                        1109/TNET.2019.2951713.

Liao:2020:SAP

[LCH20b] Guocheng Liao, Xu Chen, and Jianwei Huang. Socialaware privacy-preserving mechanism for correlated data. IEEE/ACM Transactions on Networking, 28(4): 1671–1683, August 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL

https://dl.acm.org/doi/
10.1109/TNET.2020.2994213.

Liao:2022:PAO

[LCH22] Guocheng Liao, Xu Chen, and Jianwei Huang. Privacyaware online social networking with targeted advertisement. IEEE/ ACM Transactions on Networking, 30(3):1312–1327, June 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2021.3137513.

Liao:2023:PPU

[LCH23] Guocheng Liao, Xu Chen, and Jianwei Huang. Privacy protection under incomplete social and data correlation information. IEEE/ACM Transactions on Networking, 31(6):2515– 2528, 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2023.3254549.

Lee:2018:ASC

[LCK^{+}18] Woong-Hee Lee, Jeong-Sik Choi, Yong-Hwa Kim, Jong-Ho Lee, and Seong-Cheol Kim. Adaptive sector coloring game for geometric network information-based intercell interference coordination in wireless cellular networks. IEEE/ACM Transactions on Networking, 26(1):288–301, February 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Li:2012:PFT

[LCL12a] Tao Li, Shigang Chen, and Yibei Ling. Per-flow traffic measurement through randomized counter sharing. IEEE/ACM Transactions on Networking, 20(5): 1622–1634, October 2012. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Liu:2012:DPB

[LCL^{+}12b] Changbin Liu, Ricardo Correa, Xiaozhou Li, Prithwish Basu, Boon Thau Loo, and Yun Mao. Declarative policy-based adaptive mobile ad hoc networking. IEEE/ACM Transactions on Networking, 20(3):770– 783, June 2012. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Li:2013:EPI

[LCL13a] Tao Li, Shigang Chen, and Yibei Ling. Efficient protocols for identifying the missing tags in a large RFID system. IEEE/ACM Transactions on Networking, 21(6):1974–1987, December 2013. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Li:2013:SCB

[LCL^{+}13b] Tao Li, Shigang Chen, Wen Luo, Ming Zhang, and Yan Qiao. Spreader classification based on optimal dynamic bit sharing. IEEE/ACM Transactions on Networking, 21(3): 817–830, June 2013. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Li:2016:SPS

[LCL16] Xiaoyong Li, Daren B. H.

Cline, and Dmitri Loguinov. On sample-path staleness in lazy data replication. IEEE/ACM Transactions on Networking, 24(5): 2858–2871, October 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Li:2017:TUD

[LCL17] Xiaoyong Li, Daren B. H.

Cline, and Dmitri Loguinov. Temporal update dynamics under blind sampling. IEEE/ACM Transactions on Networking, 25(1):363–376, February 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Li:2018:SPV

[LCL^{+}18] Qi Li, Yanyu Chen, Patrick P. C. Lee, Mingwei Xu, and Kui Ren. Security policy violations in SDN data plane. IEEE/ ACM Transactions on Networking, 26(4):1715–1727, August 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Liu:2020:FAD

[LCL^{+}20] Xiulong Liu, Sheng Chen, Jia Liu, Wenyu Qu, Fengjun Xiao, Alex X. Liu, Jiannong Cao, and Jiangchuan Liu. Fast and accurate detection of unknown tags for RFID systems — hash collisions are desirable. IEEE/ ACM Transactions on Networking, 28(1):126–139, February 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL https://

                                        dl.acm.org/doi/abs/10.
                                        1109/TNET.2019.2957239.

Lin:2018:FEC

[LCLC18] Chi-Han Lin, Yi-Ting Chen, Kate Ching-Ju Lin, and Wen-Tsuen Chen. FDoF: Enhancing channel utilization for 802.11ac. IEEE/ACM Transactions on Networking, 26(1):465– 477, February 2018. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Lagoa:2004:ACA

[LCM04] Constantino M. Lagoa, Hao Che, and Bernardo A. Movsichoff. Adaptive control algorithms for decentralized optimal traffic engineering in the Internet. IEEE/ACM Transactions on Networking, 12(3):415– 428, June 2004. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Li:2020:DAG

[LCP^{+}20] Yixin Li, Bin Cao, Mugen Peng, Long Zhang, Lei Zhang, Daquan Feng, and Jihong Yu. Direct acyclic graph-based ledger for Internet of things: Performance and security analysis. IEEE/ ACM Transactions on Networking, 28(4):1643–1656, August 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2020.2991994.

Luo:2014:MTD

[LCQL14] Wen Luo, Shigang Chen, Yan Qiao, and Tao Li. Missing-tag detection and energy-time tradeoff in large-scale RFID systems with unreliable channels. IEEE/ACM Transactions on Networking, 22(4):1079– 1091, August 2014. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Leith:2012:MMF

[LCS12] Douglas J. Leith, Qizhi Cao, and Vijay G. Subramanian. Max-min fairness in 802.11 mesh networks. IEEE/ACM Transactions on Networking, 20(3):756– 769, June 2012. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Liu:2018:CGT

[LCS^{+}18] Ye Liu, Chung Shue Chen, Chi Wan Sung, Chandramani Singh, Ye Liu, Chi Wan Sung, Chung Shue Chen, and Chandramani Singh. Corrections to “A Game Theoretic Distributed Algorithm for Fe-ICIC Optimization in LTE-A HetNets”. IEEE/ACM Transactions on Networking, 26(2):1033, April 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). See [LCSS17].

Liu:2017:GTD

[LCSS17] Ye Liu, Chung Shue Chen, Chi Wan Sung, and Chandramani Singh. A game theoretic distributed algorithm for FeICIC optimization in LTE-A Het-Nets. IEEE/ACM Transactions on Networking, 25(6): 3500–3513, December 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). See corrections [LCS^{+}18].

Lu:2020:NDV

[LCU^{+}20] Zongqing Lu, Kevin Chan, Rahul Urgaonkar, Shiliang Pu, and Thomas La Porta. NetVision: Ondemand video processing in wireless networks. IEEE/ ACM Transactions on Networking, 28(1):196–209, February 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL https://

dl.acm.org/doi/abs/10.
1109/TNET.2019.2954909.

Loguinov:2005:GTA

[LCW05] Dmitri Loguinov, Juan Casas, and Xiaoming Wang. Graph-theoretic analysis of structured peerto-peer systems: routing distances and fault resilience. IEEE/ACM Transactions on Networking, 13(5):1107–1120, October 2005. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Li:2015:RCA

[LCW^{+}15] Yu-Sian Li, Trang Minh Cao, Shu-Ting Wang, Xin Huang, Cheng-Hsin Hsu, and Po-Ching Lin. A resource-constrained asymmetric redundancy elimination algorithm. IEEE/ ACM Transactions on Networking, 23(4):1135–1148, August 2015. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Liu:2024:HGM

[LCW^{+}24] Xin Liu, Zicheng Chi, Wei Wang, Yao Yao, Pei Hao, and Ting Zhu. High-granularity modulation for OFDM backscatter. IEEE/ACM Transactions on Networking, 32(1): 338–351, 2024. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2023.3286880.

Liu:2016:ERG

[LCX^{+}16] Jia Liu, Min Chen, Bin Xiao, Feng Zhu, Shigang Chen, and Lijun Chen. Efficient RFID grouping protocols. IEEE/ACM Transactions on Networking, 24(5): 3177–3190, October 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Liu:2019:EIS

[LCX^{+}19] Jia Liu, Shigang Chen, Qingjun Xiao, Min Chen, Bin Xiao, and Lijun Chen. Efficient information sampling in multi-category RFID systems. IEEE/ ACM Transactions on Networking, 27(1):159–172, February 2019. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Lam:1996:LSA

[LCY96] Simon S. Lam, Simon Chow, and David K. Y. Yau. A lossless smoothing algorithm for compressed video. IEEE/ ACM Transactions on Networking, 4(5):697–708, October 1996. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1996-4-5/
p697-lam/.

Liu:2019:FRS

[LCY^{+}19] Xiulong Liu, Jiannong Cao, Yanni Yang, Wenyu Qu, Xibin Zhao, Keqiu Li, and Didi Yao. Fast RFID sensory data collection: Tradeoff between computation and communication costs. IEEE/ACM Transactions on Networking, 27(3):1179– 1191, June 2019. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Liu:2023:RMF

[LCZ^{+}23] Jianwei Liu, Kaiyan Cui, Xiang Zou, Jinsong Han, Feng Lin, and Kui Ren. Reliable multi-factor user authentication with one single finger swipe. IEEE/ ACM Transactions on Networking, 31(3):1117–1131, June 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2022.3208002.

Liu:2013:UMF

[LCZC13] Juan Liu, Wei Chen, Ying Jun Zhang, and Zhigang Cao. A utility maximization framework for fair and efficient multicasting in multicarrier wireless cellular networks. IEEE/ ACM Transactions on Networking, 21(1):110–120, February 2013. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Liu:2017:AAR

[LCZH17] Daibo Liu, Zhichao Cao, Yi Zhang, and Mengshu Hou. Achieving accurate and real-time link estimation for low power wireless sensor networks. IEEE/ ACM Transactions on Networking, 25(4):2096–2109, August 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Li:1995:ANP

[LD95] Guang-Liang Li and Patrick W.

Dowd. An analysis of network performance degradation induced by workload fluctuations. IEEE/ ACM Transactions on Networking, 3(4):433–440, August 1995. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1995-3-4/
p433-li/.

Limbasiya:2021:MSE

[LDD21] Trupil Limbasiya, Debasis Das, and Sajal K. Das. MComIoV: Secure and energy-efficient message communication protocols for Internet of vehicles. IEEE/ACM Transactions on Networking, 29(3):1349– 1361, June 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2021.3062766.

Laufer:2012:PTA

[LDFK12] Rafael Laufer, Henri Dubois-Ferrière, and Leonard Kleinrock. Polynomialtime algorithms for multirate anypath routing in wireless multihop networks. IEEE/ACM Transactions on Networking, 20(3):742– 755, June 2012. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Liao:2013:DDM

[LDGL13] Yongjun Liao, Wei Du, Pierre Geurts, and Guy Leduc. DMFSGD: a decentralized matrix factorization algorithm for network distance prediction. IEEE/ ACM Transactions on Networking, 21(5):1511–1524, October 2013. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Lehrieder:2012:CBL

[LDH^{+}12] Frank Lehrieder, György Dán, Tobias Hoßfeld, Simon Oechsner, and Vlad Singeorzan. Caching for BitTorrent-like P2P systems: a simple fluid model and its implications. IEEE/ ACM Transactions on Networking, 20(4):1176–1189, August 2012. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

LoPresti:2002:MBI

[LDHT02] Francesco Lo Presti, N. G. Duffield, Joe Horowitz, and Don Towsley. Multicastbased inference of networkinternal delay distributions. IEEE/ACM Transactions on Networking, 10(6):761– 775, December 2002. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Lee:2012:OFL

[LDK12] Myungjin Lee, Nick Duffield, and Ramana Rao Kompella. Opportunistic flowlevel latency estimation using consistent netflow. IEEE/ACM Transactions on Networking, 20(1):139– 152, February 2012. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Lee:2013:HFP

[LDK13] Myungjin Lee, Nick Duffield, and Ramana Rao Kompella. High-fidelity per-flow delay measurements with reference latency interpolation. IEEE/ACM Transactions on Networking, 21(5): 1567–1580, October 2013. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Liu:2022:TBN

[LDL^{+}22] Dianxiong Liu, Zhiyong Du, Xiaodu Liu, Heyu Luan, Yitao Xu, and Yifan Xu. Task-based network reconfiguration in distributed UAV swarms: a bilateral matching approach. IEEE/ACM Transactions on Networking, 30(6):2688– 2700, 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2022.3181036.

Ludwig:2018:TPC

[LDRS18] Arne Ludwig, Szymon Dudycz, Matthias Rost, and Stefan Schmid. Transiently policy-compliant network updates. IEEE/ ACM Transactions on Networking, 26(6):2569–2582, December 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Li:2024:PNS

[LDS^{+}24] Hao Li, Yihan Dang, Guangda Sun, Changhao Wu, Peng Zhang, Danfeng Shan, Tian Pan, and Chengchen Hu. Programming network stack for physical middleboxes and virtualized network functions. IEEE/ACM Transactions on Networking, 32 (2):971–986, April 2024. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1109/TNET.2023.3307641.

Li:2020:TMD

[LDW^{+}20] Meng Li, Haipeng Dai, Xiaoyu Wang, Rui Xia, Alex X. Liu, and Guihai Chen. Thresholded monitoring in distributed data streams. IEEE/ ACM Transactions on Networking, 28(3):1033–1046, June 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2020.2979654.

Liu:2016:TRS

[LDY^{+}16] Wenping Liu, Tianping Deng, Yang, Hongbo Jiang, Xiaofei Liao, Jiangchuan Liu, Bo Li, and Guoyin Jiang. Towards robust surface skeleton extraction and its applications in 3D wireless sensor networks. IEEE/ACM Transactions on Networking, 24(6):3300– 3313, December 2016. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Li:2017:RH

[LDZ^{+}17] Zhenjiang Li, Wan Du, Yuanqing Zheng, Mo Li, and Dapeng Wu. From rateless to hopless. IEEE/ ACM Transactions on Networking, 25(1):69–82, February 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Liu:2020:TTM

[LDZC20] Qingyu Liu, Lei Deng, Haibo Zeng, and Minghua Chen. A tale of two metrics in network delay optimization. IEEE/ ACM Transactions on Networking, 28(3):1241–1254, June 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2020.2983867.

LeBoudec:2002:SPV

[Le 02] Jean-Yves Le Boudec. Some properties of variable length packet shapers. IEEE/ACM Transactions on Networking, 10(3):329– 337, June 2002. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

LeBoudec:2018:TTR

[Le 18] Jean-Yves Le Boudec. A theory of traffic regulators for deterministic networks with application to interleaved regulators. IEEE/ ACM Transactions on Networking, 26(6):2721–2733, December 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Luo:2006:TCS

[LE06] Jie Luo and Anthony Ephremides. On the throughput, capacity, and stability regions of random multiple access. IEEE/ ACM Transactions on Networking, 14(SI):2593–2607, June 2006. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Li:2012:ETB

[LE12a] Bin Li and Atilla Eryilmaz. Exploring the throughput boundaries of randomized schedulers in wireless networks. IEEE/ ACM Transactions on Networking, 20(4):1112–1124, August 2012. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Li:2012:SEE

[LE12b] Ruogu Li and Atilla Eryilmaz. Scheduling for end-to-end deadline-constrained traffic with reliability requirements in multihop networks. IEEE/ACM Transactions on Networking, 20(5):1649–1662, October 2012. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Lakshminarayana:2013:MMI

[LE13] Subhash Lakshminarayana and Atilla Eryilmaz. Multirate multicasting with intralayer network coding. IEEE/ACM Transactions on Networking, 21(4):1256– 1269, August 2013. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Lee:1996:WFC

[Lee96] Daniel Chonghwan Lee.

Worst-case fraction of CBR teletraffic unpunctual due to statistical multiplexing. IEEE/ACM Transactions on Networking, 4(1):98–105, February 1996. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1996-4-1/
p98-lee/.

Liu:1998:LER

[LESZ98] Ching-Gung Liu, Deborah Estrin, Scott Shenker, and Lixia Zhang. Local error recovery in SRM: comparison of two approaches. IEEE/ ACM Transactions on Networking, 6(6):686–699, December 1998. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1998-6-6/
p686-liu/.

Levine:1995:AST

[Lev95] Judah Levine. An algorithm to synchronize the time of a computer to universal time. IEEE/ ACM Transactions on Networking, 3(1):42–50, February 1995. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1995-3-1/
p42-levine/.

Liu:2014:SMW

[LEY14] Shihuan Liu, Eylem Ekici, and Lei Ying. Scheduling in multihop wireless networks without backpressure. IEEE/ACM Transactions on Networking, 22(5):1477–1488, October 2014. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Li:2011:UAO

[LEYS11] Ruogu Li, Atilla Eryilmaz, Lei Ying, and Ness B. Shroff. A unified approach to optimizing performance in networks serving heterogeneous flows. IEEE/ACM Transactions on Networking, 19(1):223– 236, February 2011. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Lamali:2018:ACA

[LFC18] Mohamed Lamine Lamali, Nasreddine Fergani, and Johanne Cohen. Algorithmic and complexity aspects of path computation in multi-layer networks. IEEE/ACM Transactions on Networking, 26(6):2787– 2800, December 2018. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Liu:2022:IMS

[LFC^{+}22] Yunshu Liu, Zhixuan Fang, Man Hon Cheung, Wei Cai, and Jianwei Huang. An incentive mechanism for sustainable blockchain storage. IEEE/ACM Transactions on Networking, 30(5): 2131–2144, 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2022.3166459.

Lynch:2019:ASO

[LFF^{+}19] David Lynch, Michael Fenton, David Fagan, Stepan Kucera, Holger Claussen, and Michael O’Neill. Automated self-optimization in heterogeneous wireless communications networks. IEEE/ACM Transactions on Networking, 27(1):419– 432, February 2019. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Lubben:2014:SBE

[LFL14] Ralf Lübben, Markus Fidler, and Jörg Liebeherr. Stochastic bandwidth estimation in networks with random service. IEEE/ACM Transactions on Networking, 22(2):484– 497, April 2014. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Luo:2023:AFG

[LFL^{+}23] Lailong Luo, Pengtao Fu, Shangsen Li, Deke Guo, Qianzhen Zhang, and Huaimin Wang. Ark Filter: a general and spaceefficient sketch for network flow analysis. IEEE/ ACM Transactions on Networking, 31(6):2825– 2839, 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2023.3263839.

Liebeherr:2010:STA

[LFV10] Jörg Liebeherr, Markus Fidler, and Shahrokh Valaee. A system-theoretic approach to bandwidth estimation. IEEE/ACM Transactions on Networking, 18(4):1040–1053, August 2010. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Luo:2023:MCD

[LFXY23] Shouxi Luo, Pingzhi Fan, Huanlai Xing, and Hongfang Yu. Meeting coflow deadlines in data center networks with policybased selective completion. IEEE/ACM Transactions on Networking, 31 (1):178–191, 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.

                                        acm.org/doi/10.1109/TNET.
                                        2022.3187821.

Liu:2019:MAV

[LFY^{+}19] Jiaqi Liu, Luoyi Fu, Yuhang Yao, Xinzhe Fu, Xinbing Wang, and Guihai Chen. Modeling, analysis and validation of evolving networks with hybrid interactions. IEEE/ ACM Transactions on Networking, 27(1):126–142, February 2019. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Lai:2022:UOT

[LFZ^{+}22] Pan Lai, Rui Fan, Xiao Zhang, Wei Zhang, Fang Liu, and Joey Tianyi Zhou. Utility optimal thread assignment and resource allocation in multiserver systems. IEEE/ ACM Transactions on Networking, 30(2):735–748, April 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2021.3123817.

Liu:2011:PFD

[LFZS11] Ren-Shiou Liu, Kai-Wei Fan, Zizhan Zheng, and Prasun Sinha. Perpetual and fair data collection for environmental energy harvesting sensor networks. IEEE/ACM Transactions on Networking, 19 (4):947–960, August 2011. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Li:2013:NCD

[LG13a] Zhijun Li and Guang Gong.

On the node clone detection in wireless sensor networks. IEEE/ACM Transactions on Networking, 21 (6):1799–1811, December 2013. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Lorenzo:2013:CAN

[LG13b] Beatriz Lorenzo and Savo Glisic. Context-aware nanoscale modeling of multicast multihop cellular networks. IEEE/ACM Transactions on Networking, 21 (2):359–372, April 2013. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Lemamou:2016:HIL

[LGC16] Eunice Adjarath Lemamou, Philippe Galinier, and Steven Chamberland. A hybrid iterated local search algorithm for the global planning problem of survivable 4G wireless networks. IEEE/ACM Transactions on Networking, 24(1):137– 148, February 2016. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Lorenzo:2021:ARC

[LGCG^{+}21] Beatriz Lorenzo, Francisco Javier González-Castaño, Linke Guo, Felipe Gil-Castiñeira, and Yuguang Fang. Autonomous robustness control for fog reinforcement in dynamic wireless networks. IEEE/ACM Transactions on Networking, 29(6):2522–2535, December 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2021.3091332.

Loiseau:2010:ISS

[LGD^{+}10] Patrick Loiseau, Paulo Gonçalves, Guillaume Dewaele, Pierre Borgnat, Patrice Abry, and Pascale Vicat-Blanc Primet. Investigating self-similarity and heavy-tailed distributions on a large-scale experimental facility. IEEE/ ACM Transactions on Networking, 18(4):1261–1274, August 2010. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Li:2018:TBP

[LGDC18] Huikang Li, Yi Gao, Wei Dong, and Chun Chen. Taming both predictable and unpredictable link failures for network tomography. IEEE/ACM Transactions on Networking, 26 (3):1460–1473, June 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Li:2019:PLT

[LGDC19] Huikang Li, Yi Gao, Wei Dong, and Chun Chen. Preferential link tomography in dynamic networks. IEEE/ACM Transactions on Networking, 27(5):1801– 1814, October 2019. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Li:2023:BBN

[LGDC23] Huikang Li, Yi Gao, Wei Dong, and Chun Chen. Bound-based network tomography for inferring interesting path metrics. IEEE/ACM Transactions on Networking, 31 (1):1–14, 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2022.3180631.

Liao:2010:SIR

[LGGZ10] Yong Liao, Lixin Gao, Roch Guérin, and Zhi-Li Zhang. Safe interdomain routing under diverse commercial agreements. IEEE/ ACM Transactions on Networking, 18(6):1829–1840, December 2010. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Liu:2017:EEB

[LGHL17] Fangming Liu, Jian Guo, Xiaomeng Huang, and John C. S. Lui. eBA: Efficient bandwidth guarantee under traffic variability in datacenters. IEEE/ ACM Transactions on Networking, 25(1):506–519, February 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Lee:2014:FMM

[LGKV14] Myungjin Lee, Sharon Goldberg, Ramana Rao Kompella, and George Varghese. FineComb: measuring microscopic latency and loss in the presence of reordering. IEEE/ ACM Transactions on Networking, 22(4):1136–1149, August 2014. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Li:2009:PFD

[LGS09] Ming Li, Deepak Ganesan, and Prashant Shenoy. PRESTO: feedback-driven data management in sensor networks. IEEE/ ACM Transactions on Networking, 17(4):1256–1269, August 2009. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Lin:2023:SBO

[LGS^{+}23] Xu Lin, Deke Guo, Yulong Shen, Guoming Tang, Bangbang Ren, and Ming Xu. SFT-Box: an online approach for minimizing the embedding cost of multiple hybrid SFCs. IEEE/ACM Transactions on Networking, 31(4):1463– 1477, 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2022.3221868.

Li:2011:SCE

[LGW^{+}11] Dan Li, Chuanxiong Guo, Haitao Wu, Kun Tan, Yongguang Zhang, Songwu Lu, and Jianping Wu. Scalable and cost-effective interconnection of datacenter servers using dual server ports. IEEE/ ACM Transactions on Networking, 19(1):102–114, February 2011. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Luo:2017:EMS

[LGW^{+}17] Lailong Luo, Deke Guo, Jie Wu, Ori Rottenstreich, Qian He, Yudong Qin, and Xueshan Luo. Efficient multiset synchronization. IEEE/ACM Transactions on Networking, 25 (2):1190–1205, April 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Li:2016:BEN

[LGY16] Zhenhua Li, Zhiyang Guo, and Yuanyuan Yang. BCCC: an expandable network for data centers. IEEE/ACM Transactions on Networking, 24(6):3740–3755, December 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Lin:2023:MCA

[LGZ^{+}23] Feng Lin, Ming Gao, Lingfeng Zhang, Yimin Li, Weiye Xu, Jinsong Han, Xian Xu, Wenyao Xu, and Kui Ren. Mobile communication among COTS IoT devices via a resonant gyroscope with ultrasound. IEEE/ACM Transactions on Networking, 31(3):1026– 1041, June 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2022.3205151.

Lee:1995:SCA

[LH95] Whay Chiou Lee and Pierre A. Humblet. Spacetime characteristics of ALOHA protocols in highspeed bidirectional bus networks. IEEE/ACM Transactions on Networking, 3(5):613–622, October 1995. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1995-3-5/
p613-lee/.

Liang:2003:PDB

[LH03] Ben Liang and Zygmunt J.

Haas. Predictive distancebased mobility management for multidimensional PCS networks. IEEE/ACM Transactions on Networking, 11(5):718–732, October 2003. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Li:2005:LTC

[LH05] Ning Li and Jennifer C.

Hou. Localized topology control algorithms for heterogeneous wireless networks. IEEE/ACM Transactions on Networking, 13 (6):1313–1324, December 2005. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

La:2007:DPD

[LH07] Richard J. La and Yijie Han. Distribution of path durations in mobile ad hoc networks and path selection. IEEE/ACM Transactions on Networking, 15(5): 993–1006, October 2007.

CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Luo:2010:JSM

[LH10] Jun Luo and Jean-Pierre Hubaux. Joint sink mobility and routing to maximize the lifetime of wireless sensor networks: the case of constrained mobility. IEEE/ACM Transactions on Networking, 18(3): 871–884, June 2010. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Lee:2013:PMS

[LH13] Tsern-Huei Lee and Nai-Lun Huang. A patternmatching scheme with high throughput performance and low memory requirement. IEEE/ACM Transactions on Networking, 21(4):1104–1116, August 2013. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Li:2014:PDC

[LH14] Shuqin Li and Jianwei Huang. Price differentiation for communication networks. IEEE/ACM Transactions on Networking, 22(3):703–716, June 2014. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Li:2005:CBD

[LHB^{+}05] Li Li, Joseph Y. Halpern, Paramvir Bahl, Yi-Min Wang, and Roger Wattenhofer. A cone-based distributed topology-control algorithm for wireless multi-hop networks. IEEE/ ACM Transactions on Networking, 13(1):147–159, February 2005. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Lim:2005:CIC

[LHC05] Hyuk Lim, Jennifer C. Hou, and Chong-Ho Choi. Constructing Internet coordinate system based on delay measurement. IEEE/ACM Transactions on Networking, 13(3):513–525, June 2005. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Liu:2016:DDO

[LHC^{+}16] Daibo Liu, Mengshu Hou, Zhichao Cao, Jiliang Wang, Yuan He, and Yunhao Liu. Duplicate detectable opportunistic forwarding in duty-cycled wireless sensor networks. IEEE/ACM Transactions on Networking, 24(2):662–673, April 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Lee:2012:SSS

[LHK^{+}12] Kyunghan Lee, Seongik Hong, Seong Joon Kim, Injong Rhee, and Song Chong. SLAW: self-similar least-action human walk. IEEE/ACM Transactions on Networking, 20(2):515– 529, April 2012. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Lai:2015:OSA

[LHL15] Yuan-Cheng Lai, Ling-Yen Hsiao, and Bor-Shen Lin. Optimal slot assignment for binary tracking tree protocol in RFID tag identification. IEEE/ACM Transactions on Networking, 23(1): 255–268, February 2015. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Li:2021:MSE

[LHL^{+}21] Shaoran Li, Yan Huang, Chengzhang Li, Brian A. Jalaian, Y. Thomas Hou, Wenjing Lou, and Stephen Russell. Maximize spectrum efficiency in underlay coexistence with channel uncertainty. IEEE/ ACM Transactions on Networking, 29(2):764–778, April 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2020.3047760.

Liu:2023:AAL

[LHL^{+}23a] Jingling Liu, Jiawei Huang, Weihe Li, Jianxin Wang, and Tian He. Asymmetryaware load balancing with adaptive switching granularity in data center. IEEE/ACM Transactions on Networking, 31(3):1145– 1158, June 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2022.3208201.

Lu:2023:IOI

[LHL^{+}23b] Yu-Han Lu, Sandy Hsin-Yu Hsiao, Chi-Yu Li, Yi-Chen Hsieh, Po-Yi Chou, Yao-Yu Li, Tian Xie, and Guan-Hua Tu. Insecurity of operational IMS call systems: Vulnerabilities, attacks, and countermeasures. IEEE/ ACM Transactions on Networking, 31(2):800–815, April 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2022.3205183.

Libeskind-Hadas:2002:MRW

[LHM02] Ran Libeskind-Hadas and Rami Melhem. Multicast routing and wavelength assignment in multihop optical networks. IEEE/ACM Transactions on Networking, 10(5):621–629, October 2002. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Lin:2019:ICI

[LHW19] Kate Ching-Ju Lin, Kai-Cheng Hsu, and Hung-Yu Wei. Inter-client interference cancellation for fullduplex networks with halfduplex clients. IEEE/ ACM Transactions on Networking, 27(5):2150–2163, October 2019. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Lin:2020:LGN

[LHW^{+}20] Yilei Lin, Ting He, Shiqiang Wang, Kevin Chan, and Stephen Pasteris. Looking glass of NFV: Inferring the structure and state of NFV network from external observations. IEEE/ ACM Transactions on Networking, 28(4):1477–1490, August 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2020.2985908.

Lin:2023:MEE

[LHY^{+}23] Chi Lin, Shibo Hao, Wei Yang, Pengfei Wang, Lei Wang, Guowei Wu, and Qiang Zhang. Maximizing energy efficiency of periodarea coverage with a UAV for wireless rechargeable sensor networks. IEEE/ ACM Transactions on Networking, 31(4):1657– 1673, 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2022.3220927.

Li:2016:DOS

[LHZ^{+}16] Hang Li, Chuan Huang, Ping Zhang, Shuguang Cui, and Junshan Zhang. Distributed opportunistic scheduling for energy harvesting based wireless networks: a two-stage probing approach. IEEE/ACM Transactions on Networking, 24(3):1618–1631, June 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Liu:2019:TAT

[LHZ^{+}19] Sen Liu, Jiawei Huang, Yutao Zhou, Jianxin Wang, and Tian He. Task-aware TCP in data center networks. IEEE/ACM Transactions on Networking, 27(1):389–404, February 2019. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Li:2009:MCW

[Li09] Xiang-Yang Li. Multicast capacity of wireless ad hoc networks. IEEE/ACM Transactions on Networking, 17(3):950–961, June 2009. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Li:2021:CNC

[LI21] Yuanyuan Li and Stratis Ioannidis. Cache networks of counting queues. IEEE/ ACM Transactions on Networking, 29(6):2751–2764, December 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2021.3102518.

Li:2024:TFN

[Li24] Qian Li. TCP FlexiS: a new approach to incipient congestion detection and control. IEEE/ ACM Transactions on Networking, 32(2):1245–1260, April 2024. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2023.3319441.

Liang:2006:MGM

[Lia06] Xue-Bin Liang. Matrix games in the multicast networks: maximum information flows with network switching. IEEE/ ACM Transactions on Networking, 14(SI):2433–2466, June 2006. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Liew:1997:SSB

[Lie97]  Soung C. Liew.  On

the stability of shuffleexchange and bidirectional shuffle-exchange deflection networks. IEEE/ACM Transactions on Networking, 5(1):87–94, February 1997. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1997-5-1/
p87-liew/.

Lin:1993:LSS

[Lin93] Frank Y. S. Lin. Link set sizing for networks supporting SMDS. IEEE/ ACM Transactions on Networking, 1(6):729–739, December 1993. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1993-1-6/
p729-lin/.

Lin:1997:RLU

[Lin97] Yi-Bing Lin. Reducing location update cost in a PCS network. IEEE/ ACM Transactions on Networking, 5(1):25–33, February 1997. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1997-5-1/
p25-lin/.

Liu:2010:DBC

[Liu10] Yong Liu. Delay bounds of chunk-based peer-to-peer video streaming. IEEE/ ACM Transactions on Networking, 18(4):1195–1206, August 2010. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Liu:2009:SNF

[LJ09] Guanglei Liu and Chuanyi Ji. Scalability of networkfailure resilience: analysis using multi-layer probabilistic graphical models. IEEE/ACM Transactions on Networking, 17(1):319– 331, February 2009. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Lim:2014:SMW

[LJA14] Sungsu Lim, Kyomin Jung, and Matthew Andrews. Stability of the max-weight protocol in adversarial wireless networks. IEEE/ ACM Transactions on Networking, 22(6):1859–1872, December 2014. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Liu:2005:SNM

[LJC05] Guanglei Liu, Chuanyi Ji, and Vincent W. S. Chan. On the scalability of network management information for inter-domain lightpath assessment. IEEE/ ACM Transactions on Networking, 13(1):160–172, February 2005. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Liu:2018:PPD

[LJHB18] Zhuotao Liu, Hao Jin, Yih-Chun Hu, and Michael Bailey. Practical proactive DDoS-attack mitigation via endpoint-driven in-network traffic control. IEEE/ACM Transactions on Networking, 26(4):1948– 1961, August 2018. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Liu:2019:INV

[LJJ^{+}19] Wenping Liu, Hongbo Jiang, Guoyin Jiang, Jiangchuan Liu, Xiaoqiang Ma, Yufu Jia, and Fu Xiao. Indoor navigation with virtual graph representation: Exploiting peak intensities of unmodulated luminaries. IEEE/ACM Transactions on Networking, 27(1): 187–200, February 2019. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Liu:2016:DSL

[LJL^{+}16] Wenping Liu, Hongbo Jiang, Jiangchuan Liu, Xiaofei Liao, Hongzhi Lin, and Tianping Deng. On the distance-sensitive and load-balanced information storage and retrieval for 3D sensor networks. IEEE/ ACM Transactions on Networking, 24(6):3439–3449, December 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Liu:2012:GTH

[LJNK12] Jiajia Liu, Xiaohong Jiang, Hiroki Nishiyama, and Nei Kato. Generalized two-hop relay for flexible delay control in MANETs. IEEE/ ACM Transactions on Networking, 20(6):1950–1963, December 2012. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Liu:2022:MLS

[LJSB22] Yuchen Liu, Yubing Jian, Raghupathy Sivakumar, and Douglas M. Blough. Maximizing line-of-sight coverage for mmWave wireless LANs with multiple access points. IEEE/ ACM Transactions on Networking, 30(2):698–716, April 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2021.3122378.

Liu:2007:MLS

[LJW^{+}07] Hai Liu, Xiaohua Jia, Peng-Jun Wan, Chih-Wei Yi, S. Kami Makki, and Niki Pissinou. Maximizing lifetime of sensor surveillance systems. IEEE/ACM Transactions on Networking, 15(2):334–345, April 2007. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Li:2023:LCM

[LJZ^{+}23] Ruixuan Li, Xiaofeng Jia, Zhenyong Zhang, Jun Shao, Rongxing Lu, Jingqiang Lin, Xiaoqi Jia, and Guiyi Wei. A longitudinal and comprehensive measurement of DNS strict privacy. IEEE/ ACM Transactions on Networking, 31(6):2793– 2808, 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2023.3262651.

Lee:1995:BAR

[LK95] Wei-Tsong Lee and Ling-Yang Kung. Binary addressing and routing schemes in the Manhattan street network. IEEE/ ACM Transactions on Networking, 3(1):26–30, February 1995. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1995-3-1/
p26-lee/.

Li:2002:CMS

[LK02] Chengzhi Li and Edward W. Knightly. Coordinated multihop scheduling: a framework for end-to-end services. IEEE/ACM Transactions on Networking, 10(6):776–789, December 2002. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Li:2005:SCP

[LK05] Chengzhi Li and Edward W. Knightly. Schedulability criterion and performance analysis of coordinated schedulers. IEEE/ ACM Transactions on Networking, 13(2):276–287, April 2005. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Lin:2010:CMS

[LK10] Bill Lin and Isaac Keslassy. The concurrent matching switch architecture. IEEE/ ACM Transactions on Networking, 18(4):1330–1343, August 2010. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Liu:2013:QVR

[LK13] Alex X. Liu and Amir R. Khakpour. Quantifying and verifying reachability for access controlled networks. IEEE/ACM Transactions on Networking, 21 (2):551–565, April 2013. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Liang:2014:FCP

[LK14] Guanfeng Liang and Ulas C.

Kozat. Fast cloud: pushing the envelope on delay performance of cloud storage with coding. IEEE/ ACM Transactions on Networking, 22(6):2012–2025, December 2014. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Laufer:2016:CWC

[LK16a] Rafael Laufer and Leonard Kleinrock. The capacity of wireless CSMA/CA networks. IEEE/ACM Transactions on Networking, 24 (3):1518–1532, June 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Liang:2016:TDO

[LK16b] Guanfeng Liang and Ulas C.

Kozat. On throughputdelay optimal access to storage clouds via load adaptive coding and chunking. IEEE/ACM Transactions on Networking, 24(4): 2168–2181, August 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Lee:2011:PCM

[LKC11] Minsik Lee, Youngjip Kim, and Chong-Ho Choi. Period-controlled MAC for high performance in wireless networks. IEEE/ ACM Transactions on Networking, 19(4):1237–1250, August 2011. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Lee:2013:CDM

[LKC^{+}13] Kyunghan Lee, Yoora Kim, Song Chong, Injong Rhee, Yung Yi, and Ness B. Shroff. On the critical delays of mobile networks under Lévy walks and Lévy flights. IEEE/ ACM Transactions on Networking, 21(5):1621–1635, October 2013. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Li:2000:ODM

[LKL00] Jie Li, Hisao Kameda, and Keqin Li. Optimal dynamic mobility management for PCS networks. IEEE/ACM Transactions on Networking, 8 (3):319–327, 2000. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/
                                        journals/ton/2000-8-3/
                                        p319-li/.

Luu:2020:CAR

[LKMK20] Quang-Trung Luu, Sylvaine Kerboeuf, Alexandre Mouradian, and Michel Kieffer. A coverage-aware resource provisioning method for network slicing. IEEE/ ACM Transactions on Networking, 28(6):2393–2406, December 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2020.3019098.

Le:2016:MCV

[LKS^{+}16] Anh Le, Lorenzo Keller, Hulya Seferoglu, Blerim Cici, Christina Fragouli, and Athina Markopoulou. MicroCast: Cooperative video streaming using cellular and local connections. IEEE/ACM Transactions on Networking, 24(5):2983– 2999, October 2016. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Luo:2004:UUR

[LKZ^{+}04] Haiyun Luo, Jiejun Kong, Petros Zerfos, Songwu Lu, and Lixia Zhang. URSA: ubiquitous and robust access control for mobile ad hoc networks. IEEE/ACM Transactions on Networking, 12(6):1049–1063, December 2004. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Lee:1995:NDN

[LL95] Jeong Gyu Lee and Byeong Gi Lee. A new distribution network based on controlled switching elements and its applications. IEEE/ACM Transactions on Networking, 3(1):70–81, February 1995. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/
                                        journals/ton/1995-3-1/
                                        p70-lee/.

Lee:1998:CDT

[LL98] Tsern-Huei Lee and Kuen-Chu Lai. Characterization of delay-sensitive traffic. IEEE/ACM Transactions on Networking, 6(4):499– 504, August 1998. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/
                                        journals/ton/1998-6-4/
                                        p499-lee/.

Low:1999:OFC

[LL99] Steven H. Low and David E. Lapsley. Optimization flow control, I: basic algorithm and convergence. IEEE/ ACM Transactions on Networking, 7(6):861–874, December 1999. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1999-7-6/

p861-low/. See comments [Kar03].

Lai:2009:TBA

[LL09] Yuan-Cheng Lai and Chih-Chung Lin. Two blocking algorithms on adaptive binary splitting: single and pair resolutions for RFID tag identification. IEEE/ACM Transactions on Networking, 17(3): 962–975, June 2009. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Li:2010:RPR

[LL10] Mo Li and Yunhao Liu.

Rendered path: range-free localization in anisotropic sensor networks with holes. IEEE/ACM Transactions on Networking, 18(1):320– 332, February 2010. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Leonard:2013:DIW

[LL13] Derek Leonard and Dmitri Loguinov. Demystifying Internet-wide service discovery. IEEE/ACM Transactions on Networking, 21 (6):1760–1773, December 2013. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Liu:2017:HMC

[LL17a] An Liu and Vincent K. N. Lau. How much cache is needed to achieve linear capacity scaling in backhaullimited dense wireless networks? IEEE/ACM Transactions on Networking, 25(1):179–188, February 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Liu:2017:OLA

[LL17b] Yang Liu and Mingyan Liu. An online learning approach to improving the quality of crowd-sourcing. IEEE/ACM Transactions on Networking, 25(4):2166– 2179, August 2017. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Liu:2018:ELS

[LL18] Kaikai Liu and Xiaolin Li. Enhancing localization scalability and accuracy via opportunistic sensing. IEEE/ACM Transactions on Networking, 26 (3):1517–1530, June 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Liu:2020:FFE

[LL20] Tingwei Liu and John C. S. Lui. FAVE: a fast and efficient network flow AVailability estimation method with bounded relative error. IEEE/ ACM Transactions on Networking, 28(2):505–518, April 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL https://

dl.acm.org/doi/abs/10.
1109/TNET.2020.2965161.

Liu:2019:CG

[LLAS19] Yang Liu, Bo Li, Brian D. O. Anderson, and Guodong Shi. Clique gossiping. IEEE/ACM Transactions on Networking, 27(6):2418–2431, December 2019. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL https://

dl.acm.org/doi/abs/10.
1109/TNET.2019.2952082.

Li:2022:TFA

[LLCJ22] Xiao-Yan Li, Wanling Lin, Jou-Ming Chang, and Xiaohua Jia. Transmission failure analysis of multi-protection routing in data center networks with heterogeneous edgecore servers. IEEE/ ACM Transactions on Networking, 30(4):1689– 1702, 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2022.3147320.

Lan:2011:SBS

[LLCL11] Tian Lan, Xiaojun Lin, Mung Chiang, and Ruby B. Lee. Stability and benefits of suboptimal utility maximization. IEEE/ ACM Transactions on Networking, 19(4):1194–1207, August 2011. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Lee:1996:DRC

[LLD96] Tsern-Huei Lee, Kuen-Chu Lai, and Shii-Tyng Duann. Design of a real-time call admission controller for ATM networks. IEEE/ ACM Transactions on Networking, 4(5):758–765, October 1996. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/
                                        journals/ton/1996-4-5/
                                        p758-lee/.

Li:2015:OCS

[LLE15a] Bin Li, Ruogu Li, and Atilla Eryilmaz. On the optimal convergence speed of wireless scheduling for fair resource allocation. IEEE/ACM Transactions on Networking, 23(2):631– 643, April 2015. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Li:2015:TOS

[LLE15b] Bin Li, Ruogu Li, and Atilla Eryilmaz. Throughputoptimal scheduling design with regular service guarantees in wireless networks. IEEE/ACM Transactions on Networking, 23(5):1542– 1552, October 2015. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Li:2016:WSD

[LLE16] Bin Li, Ruogu Li, and Atilla Eryilmaz. Wireless scheduling design for optimizing both service regularity and mean delay in heavy-traffic regimes. IEEE/ACM Transactions on Networking, 24(3):1867– 1880, June 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Liu:2017:TQC

[LLG^{+}17] Xiulong Liu, Keqiu Li, Song Guo, Alex X. Liu, Peng Li, Kun Wang, Jie Wu, Xiulong Liu, Keqiu Li, Song Guo, Alex X. Liu, Peng Li, Kun Wang, and Jie Wu. Top-k queries for categorized RFID systems. IEEE/ACM Transactions on Networking, 25(5):2587– 2600, October 2017. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Lim:2014:BCP

[LLJ^{+}14] Hyesook Lim, Nara Lee, Geumdan Jin, Jungwon Lee, Youngju Choi, and Changhoon Yim. Boundary cutting for packet classification. IEEE/ACM Transactions on Networking, 22(2):443–456, April 2014. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Li:2006:AMM

[LLL06] Zongpeng Li, Baochun Li, and Lap Chi Lau. On achieving maximum multicast throughput in undirected networks. IEEE/ ACM Transactions on Networking, 14(SI):2467–2485, June 2006. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Liu:2010:PDW

[LLL10] Yunhao Liu, Kebin Liu, and Mo Li. Passive diagnosis for wireless sensor networks. IEEE/ ACM Transactions on Networking, 18(4):1132–1144, August 2010. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Liu:2016:PRD

[LLL^{+}16] Zhidan Liu, Zhenjiang Li, Mo Li, Wei Xing, and Dongming Lu. Path reconstruction in dynamic wireless sensor networks using compressive sensing. IEEE/ACM Transactions on Networking, 24(4):1948– 1960, August 2016. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Liu:2017:MCR

[LLL^{+}17] Xiulong Liu, Keqiu Li, Alex X. Liu, Song Guo, Muhammad Shahzad, Ann L. Wang, and Jie Wu. Multicategory RFID estimation. IEEE/ACM Transactions on Networking, 25(1):264– 277, February 2017. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Li:2022:ISC

[LLL^{+}22a] Bingyu Li, Jingqiang Lin, Fengjun Li, Qiongxiao Wang, Wei Wang, Qi Li, Guangshen Cheng, Jiwu Jing, and Congli Wang. The invisible side of certificate transparency: Exploring the reliability of monitors in the wild. IEEE/ ACM Transactions on Networking, 30(2):749–765, April 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2021.3123507.

Li:2022:SAI

[LLL^{+}22b] Chengzhang Li, Qingyu Liu, Shaoran Li, Yongce Chen, Y. Thomas Hou, Wenjing Lou, and Sastry Kompella. Scheduling with age of information guarantee. IEEE/ACM Transactions on Networking, 30(5): 2046–2059, 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2022.3156866.

Lee:2007:EPS

[LLLS07] Yong Lee, Jianyu Lou, Junzhou Luo, and Xiaojun Shen. An efficient packet scheduling algorithm with deadline guarantees for inputqueued switches. IEEE/ ACM Transactions on Networking, 15(1):212–225, February 2007. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Li:2010:MCW

[LLLT10] Xiang-Yang Li, Yunhao Liu, Shi Li, and Shao-Jie Tang. Multicast capacity of wireless ad hoc networks under Gaussian channel model. IEEE/ ACM Transactions on Networking, 18(4):1145–1157, August 2010. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Lee:2011:RLN

[LLM11a] Kayi Lee, Hyang-Won Lee, and Eytan Modiano. Reliability in layered networks with random link failures. IEEE/ACM Transactions on Networking, 19(6):1835– 1848, December 2011. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Li:2011:BSB

[LLM11b] Tianji Li, Douglas Leith, and David Malone. Buffer sizing for 802.11-based networks. IEEE/ACM Transactions on Networking, 19(1):156–169, February 2011. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Lee:2014:MRW

[LLM14] Hyang-Won Lee, Kayi Lee, and Eytan Modiano. Maximizing reliability in WDM networks through lightpath routing. IEEE/ ACM Transactions on Networking, 22(4):1052–1066, August 2014. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Liu:2023:TMO

[LLM23] Bai Liu, Qingkai Liang, and Eytan Modiano. Tracking MaxWeight: Optimal control for partially observable and controllable networks. IEEE/ACM Transactions on Networking, 31(4):1809– 1821, 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2022.3225752.

Luo:2024:PSS

[LLM^{+}24] Bin Luo, Xinghua Li, Yinbin Miao, Man Zhang, Ximeng Liu, Yanbing Ren, Xizhao Luo, and Robert H. Deng. PAM^{3}S: Progressive two-stage auction-based multi-platform multi-user mutual selection scheme in MCS. IEEE/ACM Transactions on Networking, 32 (1):729–744, 2024. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2023.3297258.

Lian:2009:VSF

[LLNC09] Jie Lian, Yunhao Liu, Kshirasagar Naik, and Lei Chen. Virtual surrounding face geocasting in wireless ad hoc and sensor networks. IEEE/ACM Transactions on Networking, 17(1):200–211, February 2009. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Li:2007:EET

[LLS07] Yee-Ting Li, Douglas Leith, and Robert N. Shorten. Experimental evaluation of TCP protocols for highspeed networks. IEEE/ ACM Transactions on Networking, 15(5):1109–1122, October 2007. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Lotfinezhad:2009:SRD

[LLS09] Mahdi Lotfinezhad, Ben Liang, and Elvino S. Sousa. On stability region and delay performance of linear-memory randomized scheduling for timevarying networks. IEEE/ ACM Transactions on Networking, 17(6):1860–1873, December 2009. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Lin:2010:LCD

[LLS10] Longbi Lin, Xiaojun Lin, and Ness B. Shroff. Lowcomplexity and distributed energy minimization in multihop wireless networks. IEEE/ACM Transactions on Networking, 18(2):501– 514, April 2010. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Li:2023:EDN

[LLS^{+}23] Yuanyuan Li, Yuezhou Liu, Lili Su, Edmund Yeh, and Stratis Ioannidis. Experimental design networks: a paradigm for serving heterogeneous learners under networking constraints. IEEE/ACM Transactions on Networking, 31(5):2236– 2250, 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2023.3243534.

Liao:2016:LCG

[LLT^{+}16] Xiaofei Liao, Li Lin, Guang Tan, Hai Jin, Xiaobin Yang, Wei Zhang, and Bo Li. LiveRender: a cloud gaming system based on compressed graphics streaming. IEEE/ ACM Transactions on Networking, 24(4):2128–2139, August 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Le:2009:DNW

[LLW^{+}09] Franck Le, Sihyung Lee, Tina Wong, Hyong S. Kim, and Darrell Newcomb. Detecting network-wide and router-specific misconfigurations through data mining. IEEE/ACM Transactions on Networking, 17 (1):66–79, February 2009. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Li:2012:EES

[LLW^{+}12] Dan Li, Yuanjie Li, Jianping Wu, Sen Su, and Jiangwei Yu. ESM: efficient and scalable data center multicast routing. IEEE/ACM Transactions on Networking, 20(3):944– 955, June 2012. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Lu:2014:BAP

[LLW^{+}14] Ning Lu, Tom H. Luan, Miao Wang, Xuemin Shen, and Fan Bai. Bounds of asymptotic performance limits of social-proximity vehicular networks. IEEE/ ACM Transactions on Networking, 22(3):812–825, June 2014. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Liu:2015:GGL

[LLW^{+}15] Zhenhua Liu, Minghong Lin, Adam Wierman, Steven Low, and Lachlan L. H. Andrew. Greening geographical load balancing. IEEE/ACM Transactions on Networking, 23(2):657– 671, April 2015. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Li:2016:FSR

[LLWB16] Rui Li, Alex X. Liu, Ann L. Wang, and Bezawada Bruhadeshwar. Fast and scalable range query processing with strong privacy protection for cloud computing. IEEE/ACM Transactions on Networking, 24(4):2305–2318, August 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Li:2017:PIP

[LLX^{+}17] Rui Li, Alex X. Liu, Sheng Xiao, Hongyue Xu, Bezawada Bruhadeshwar, and Ann L. Wang. Privacy and integrity preserving top-k query processing for two-tiered sensor networks. IEEE/ACM Transactions on Networking, 25(4):2334– 2346, August 2017. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Lin:2019:UPR

[LLX19a] Zongzong Lin, Wenlian Lu, and Shouhuai Xu. Unified preventive and reactive cyber defense dynamics is still globally convergent. IEEE/ACM Transactions on Networking, 27 (3):1098–1111, June 2019. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Luo:2019:PPP

[LLX^{+}19b] Chengwen Luo, Xiao Liu, Wanli Xue, Yiran Shen, Jianqiang Li, Wen Hu, and Alex X. Liu. Predictable privacy-preserving mobile crowd sensing: a tale of two roles. IEEE/ ACM Transactions on Networking, 27(1):361–374, February 2019. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Leung:2001:APD

[LLY01] Matthew K. H. Leung, John C. S. Lui, and David K. Y. Yau. Adaptive proportional delay differentiated services: characterization and performance evaluation. IEEE/ ACM Transactions on Networking, 9(6):801–817, December 2001. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Lee:2006:DCK

[LLY06] Patrick P. C. Lee, John C. S. Lui, and David K. Y. Yau. Distributed collaborative key agreement and authentication protocols for dynamic peer groups. IEEE/ACM Transactions on Networking, 14(2):263– 276, April 2006. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Luo:2009:RCD

[LLY09] Hongbin Luo, Lemin Li, and Hongfang Yu. Routing connections with differentiated reliability requirements in WDM mesh networks. IEEE/ACM Transactions on Networking, 17(1):253–266, February 2009. CODEN IEANEP.

ISSN 1063-6692 (print), 1558-2566 (electronic).

Ling:2012:NCC

[LLY^{+}12] Zhen Ling, Junzhou Luo, Wei Yu, Xinwen Fu, Dong Xuan, and Weijia Jia. A new cell-counting-based attack against Tor. IEEE/ ACM Transactions on Networking, 20(4):1245–1261, August 2012. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Lee:2013:MDO

[LLY^{+}13] Kyunghan Lee, Joohyun Lee, Yung Yi, Injong Rhee, and Song Chong. Mobile data offloading: how much can WiFi deliver? IEEE/ACM Transactions on Networking, 21(2):536– 550, April 2013. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Lee:2016:MDN

[LLY^{+}16] Jinsung Lee, Hojin Lee, Yung Yi, Song Chong, Edward W. Knightly, and Mung Chiang. Making 802.11 DCF nearoptimal: design, implementation, and evaluation. IEEE/ACM Transactions on Networking, 24(3):1745– 1758, June 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Li:2022:SNL

[LLY^{+}22] Zhuo Li, Jindian Liu, Liu Yan, Beichuan Zhang, Peng Luo, and Kaihua Liu. Smart name lookup for NDN forwarding plane via neural networks. IEEE/ ACM Transactions on Networking, 30(2):529–541, April 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2021.3119769.

Li:2017:CBS

[LLZ^{+}17] Qi Li, Patrick P. C. Lee, Peng Zhang, Purui Su, Liang He, Kui Ren, Qi Li, Patrick P. C. Lee, Peng Zhang, Purui Su, Liang He, and Kui Ren. Capabilitybased security enforcement in named data networking. IEEE/ACM Transactions on Networking, 25(5): 2719–2730, October 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Li:2019:FSF

[LLZ^{+}19] Xiang-Yang Li, Huiqi Liu, Lan Zhang, Zhenan Wu, Yaochen Xie, Ge Chen, Chunxiao Wan, and Zhongwei Liang. Finding the stars in the fireworks: Deep understanding of motion sensor fingerprint. IEEE/ ACM Transactions on Networking, 27(5):1945–1958, October 2019. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Lai:2023:SCC

[LLZ^{+}23a] Zeqi Lai, Hewu Li, Qi Zhang, Qian Wu, and Jianping Wu. StarFront: Cooperatively constructing pervasive and low-latency CDNs upon emerging LEO satellites and clouds. IEEE/ ACM Transactions on Networking, 31(6):2559– 2574, 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2023.3260166.

Lei:2023:MCB

[LLZ^{+}23b] Kai Lei, Guanjie Lin, Meimei Zhang, Keke Li, Qi Li, Xiaojun Jing, and Peng Wang. Measuring the consistency between data and control plane in SDN. IEEE/ACM Transactions on Networking, 31(2):511– 525, April 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2022.3193698.

Lee:1995:CAS

[LM95] Hyong W. Lee and Jon W. Mark. Capacity allocation in statistical multiplexing of ATM sources. IEEE/ACM Transactions on Networking, 3(2):139– 151, April 1995. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1995-3-2/
p139-lee/.

Liu:1996:GFR

[LM96] Hong Liu and Raymond E.

Miller. Generalized fair reachability analysis for cyclic protocols. IEEE/ ACM Transactions on Networking, 4(2):192–204, April 1996. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1996-4-2/
p192-liu/.

Lakshman:1997:PTN

[LM97] T. V. Lakshman and Upamanyu Madhow. The performance of TCP/IP for networks with high bandwidth-delay products and random loss. IEEE/ ACM Transactions on Networking, 5(3):336–350, June 1997. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1997-5-3/
p336-lakshman/.

Li:2001:JBD

[LM01] Qiong Li and David L.

Mills. Jitter-based delayboundary prediction of wide-area networks. IEEE/ ACM Transactions on Networking, 9(5):578–590, October 2001. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Lalanne:2013:FDC

[LM13] Felipe Lalanne and Stephane Maag. A formal datacentric approach for passive testing of communication protocols. IEEE/ACM Transactions on Networking, 21(3):788–801, June 2013. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Li:2015:RBF

[LM15] Chih-Ping Li and Eytan Modiano. Receiver-based flow control for networks in overload. IEEE/ACM Transactions on Networking, 23(2):616–630, April 2015. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Le:2016:ADS

[LMD16] Anh Le, Athina Markopoulou, and Alexandros G. Dimakis. Auditing for distributed storage systems. IEEE/ACM Transactions on Networking, 24(4):2182– 2195, August 2016. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Levy:2004:DAR

[LMG04] Hanoch Levy, Tsippy Mendelson, and Gilad Goren. Dynamic allocation of resources to virtual path agents. IEEE/ACM Transactions on Networking, 12 (4):746–758, August 2004. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Labovitz:1998:IRI

[LMJ98] Craig Labovitz, G. Robert Malan, and Farnam Jahanian. Internet routing instability. IEEE/ ACM Transactions on Networking, 6(5):515–528, October 1998. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1998-6-5/
p515-labovitz/.

Lee:2010:DRN

[LML10] Hyang-Won Lee, Eytan Modiano, and Kayi Lee. Diverse routing in networks with probabilistic failures. IEEE/ACM Transactions on Networking, 18(6):1895– 1907, December 2010. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Lee:2011:CLS

[LML11] Kayi Lee, Eytan Modiano, and Hyang-Won Lee. Cross-layer survivability in WDM-based networks. IEEE/ACM Transactions on Networking, 19(4):1000– 1013, August 2011. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Leonardi:2007:OSR

[LMMN07] Emilio Leonardi, Marco Mellia, Marco Ajmone Marsan, and Fabio Neri. Optimal scheduling and routing for maximum network throughput. IEEE/ ACM Transactions on Networking, 15(6):1541–1554, December 2007. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Leonardi:2001:SIQ

[LMNM01] Emilio Leonardi, Marco Mellia, Fabio Neri, and Marco Ajmone Marsan. On the stability of inputqueued switches with speedup. IEEE/ACM Transactions on Networking, 9(1): 104–118, 2001. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/articles/
journals/ton/2001-9-1/
p104-leonardi/p104-leonardi.
pdf;   http://www.acm.
org/pubs/citations/journals/

ton/2001-9-1/p104-leonardi/

.

Li:2018:PPB

[LMODF18] Ning Li, Jose-Fernan Martinez-Ortega, Vicente Hernandez Diaz, and Jose Antonio Sanchez Fernandez. Probability predictionbased reliable and efficient opportunistic routing algorithm for VANETs. IEEE/ ACM Transactions on Networking, 26(4):1933–1947, August 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Low:1996:ACC

[LMP96] Steven H. Low, Nicholas F.

Maxemchuk, and Sanjoy Paul. Anonymous credit cards and their collusion analysis. IEEE/ ACM Transactions on Networking, 4(6):809–816, December 1996. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1996-4-6/
p809-low/.

Lenders:2008:DBA

[LMP08] Vincent Lenders, Martin May, and Bernhard Plattner. Density-based anycast: a robust routing strategy for wireless ad hoc networks. IEEE/ACM Transactions on Networking, 16(4):852–863, August 2008. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Lakshman:1999:TCV

[LMR99] T. V. Lakshman, P. P. Mishra, and K. K. Ramakrishnan. Transporting compressed video over ATM networks with explicit-rate feedback control. IEEE/ ACM Transactions on Networking, 7(5):710–723, October 1999. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/
                                        journals/ton/1999-7-5/
                                        p710-lakshman/.

Lee:2007:DAS

[LMR07] Patrick P. C. Lee, Vishal Misra, and Dan Rubenstein. Distributed algorithms for secure multipath routing in attackresistant networks. IEEE/ ACM Transactions on Networking, 15(6):1490–1501, December 2007. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Lombardo:1999:DTP

[LMS99] Alifo Lombardo, Giacomo Morabito, and Giovanni Schembra. A discretetime paradigm to evaluate skew performance in a multimedia ATM multiplexer. IEEE/ACM Transactions on Networking, 7(1):122–139, February 1999. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1999-7-1/
p122-lombardo/.

Lakshman:2000:TIP

[LMS00] T. V. Lakshman, Upamanyu Madhow, and Bernhard Suter. TCP/IP performance with random loss and bidirectional congestion. IEEE/ACM Transactions on Networking, 8 (5):541–555, 2000. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/articles/
journals/ton/2000-8-5/
p541-lakshman/p541-lakshman.
pdf;   http://www.acm.
org/pubs/citations/journals/
ton/2000-8-5/p541-lakshman/

.

Lenzini:2004:TBL

[LMS04a] Luciano Lenzini, Enzo Mingozzi, and Giovanni Stea. Tradeoffs between low complexity, low latency, and fairness with deficit roundrobin schedulers. IEEE/ ACM Transactions on Networking, 12(4):681–693, August 2004. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Lombardo:2004:NAF

[LMS04b] Alfio Lombardo, Giacomo Morabito, and Giovanni Schembra. A novel analytical framework compounding statistical traffic modeling and aggregatelevel service curve disciplines: network performance and efficiency implications. IEEE/ACM Transactions on Networking, 12(3):443–455, June 2004. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Lee:2005:DPA

[LMS05a] Jang-Won Lee, Ravi R. Mazumdar, and Ness B. Shroff. Downlink power allocation for multi-class wireless systems. IEEE/ ACM Transactions on Networking, 13(4):854–867, August 2005. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Lee:2005:NCO

[LMS05b] Jang-Won Lee, Ravi R. Mazumdar, and Ness B. Shroff. Non-convex optimization and rate control for multi-class services in the Internet. IEEE/ACM Transactions on Networking, 13(4):827–840, August 2005. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Lee:2006:JRA

[LMS06] Jang-Won Lee, Ravi R.

Mazumdar, and Ness B. Shroff. Joint resource allocation and base-station assignment for the downlink in CDMA networks. IEEE/ACM Transactions on Networking, 14(1):1–14, February 2006. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Le:2012:OCW

[LMS12] Long Bao Le, Eytan Modiano, and Ness B. Shroff. Optimal control of wireless networks with finite buffers. IEEE/ACM Transactions on Networking, 20(4):1316–1329, August 2012. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Levy:1999:SEB

[LMSKZ99] Hanoch Levy, Tzippi Mendelson, Moshe Sidi, and Joseph Keren-Zvi. Sizing exit buffers in ATM networks: an intriguing coexistence of instability and tiny cell loss rates. IEEE/ ACM Transactions on Networking, 7(6):926–936, December 1999. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1999-7-6/
p926-levy/.

Lee:2019:IPP

[LMSR19] Ming-Chun Lee, Andreas F. Molisch, Nishanth Sastry, and Aravindh Raman. Individual preference probability modeling and parameterization for video content in wireless caching networks. IEEE/ACM Transactions on Networking, 27 (2):676–690, April 2019. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Lenzen:2024:RRM

[LMSS24] Christoph Lenzen, Moti Medina, Mehrdad Saberi, and Stefan Schmid. Robust routing made easy: Reinforcing networks against non-benign faults. IEEE/ ACM Transactions on Networking, 32(1):283– 297, 2024. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2023.3283184.

Liu:2010:TRS

[LMT10] Alex X. Liu, Chad R. Meiners, and Eric Torng. TCAM razor: a systematic approach towards minimizing packet classifiers in TCAMs. IEEE/ACM Transactions on Networking, 18(2):490–500, April 2010. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Liu:2016:PCU

[LMT16] Alex X. Liu, Chad R. Meiners, and Eric Torng. Packet classification using binary content addressable memory. IEEE/ACM Transactions on Networking, 24 (3):1295–1307, June 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Lu:2016:QAR

[LMW16] Yu Lu, Mehul Motani, and Wai-Choong Wong. A QoE-aware resource distribution framework incentivizing context sharing and moderate competition. IEEE/ACM Transactions on Networking, 24(3):1364– 1377, June 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Liu:2019:CAR

[LN19] Alex X. Liu and Eric Norige. A de-compositional approach to regular expression matching for network security. IEEE/ACM Transactions on Networking, 27(6):2179–2191, December 2019. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL https://

dl.acm.org/doi/abs/10.
1109/TNET.2019.2941920.

Lian:2007:FEP

[LNA07] Jie Lian, Kshirasagar Naik, and Gordon B. Agnew. A framework for evaluating the performance of cluster algorithms for hierarchical networks. IEEE/ ACM Transactions on Networking, 15(6):1478–1489, December 2007. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Lacher:2000:PCC

[LNB00] Martin S. Lacher, Jörg Nonnenmacher, and Ernst W. Biersack. Performance comparison of centralized versus distributed error recovery for reliable multicast. IEEE/ACM Transactions on Networking, 8 (2):224, 2000. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/2000-8-2/
p224-lacher/.

Legout:2001:BAP

[LNB01] Arnaud Legout, Jörg Nonnenmacher, and Ernst W. Biersack. Bandwidthallocation policies for unicast and multicast flows. IEEE/ACM Transactions on Networking, 9(4):464– 478, August 2001. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Liang:1993:NMN

[LNC93] Luping Liang, Gerald W.

Neufeld, and Samuel T. Chanson. A name model for nested group communication. IEEE/ACM Transactions on Networking, 1(4):414–423, August 1993. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1993-1-4/
p414-liang/.

Liew:1998:BNM

[LNC98] Soung C. Liew, Ming-Hung Ng, and Cathy W. Chan. Blocking and nonblocking multirate Clos switching networks. IEEE/ ACM Transactions on Networking, 6(3):307–318, June 1998. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1998-6-3/
p307-liew/.

Lui:2004:RTA

[LNC04] King-Shan Lui, Klara Nahrstedt, and Shigang Chen. Routing with topology aggregation in delaybandwidth sensitive networks. IEEE/ACM Transactions on Networking, 12 (1):17–29, February 2004. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Laki:2021:CSF

[LNG^{+}21] Sándor Laki, Szilveszter Nádas, Gergő Gombos, Ferenc Fejes, Péter Hudoba, Zoltán Turányi, Zoltán Kiss, and Csaba Keszei. Core-stateless forwarding with QoS revisited: Decoupling delay and bandwidth requirements. IEEE/ACM Transactions on Networking, 29(2):503– 516, April 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2020.3041235.

Liu:2016:CDS

[LNL^{+}16] Kai Liu, Joseph K. Y. Ng, Victor C. S. Lee, Sang H. Son, and Ivan Stojmenovic. Cooperative data scheduling in hybrid vehicular ad hoc networks: VANET as a software defined network. IEEE/ACM Transactions on Networking, 24 (3):1759–1773, June 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Liu:2024:TDP

[LNLM24] Bai Liu, Quang Minh Nguyen, Qingkai Liang, and Eytan Modiano. Tracking drift-plus-penalty: Utility maximization for partially observable and controllable networks. IEEE/ ACM Transactions on Networking, 32(2):1064–1079, April 2024. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2023.3307684.

Li:2009:AFR

[LNM^{+}09] Tianji Li, Qiang Ni, David Malone, Douglas Leith, Yang Xiao, and Thierry Turletti. Aggregation with fragment retransmission for very high-speed WLANs. IEEE/ACM Transactions on Networking, 17(2):591– 604, April 2009. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Lucantoni:1994:MPE

[LNR94] David M. Lucantoni, Marcel F. Neuts, and Amy R. Reibman. Methods for performance evaluation of VBR video traffic models. IEEE/ACM Transactions on Networking, 2(2):176– 180, April 1994. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1994-2-2/
p176-lucantoni/.

Leconte:2011:IBT

[LNS11] Mathieu Leconte, Jian Ni, and R. Srikant. Improved bounds on the throughput efficiency of greedy maximal scheduling in wireless networks. IEEE/ACM Transactions on Networking, 19(3):709–720, June 2011. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Leue:1996:POI

[LO96] Stefan Leue and Philippe A. Oechslin. On parallelizing and optimizing the implementation of communication protocols. IEEE/ ACM Transactions on Networking, 4(1):55–70, February 1996. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/
                                        journals/ton/1996-4-1/
                                        p55-leue/.

Lorenz:1998:QRN

[LO98] Dean H. Lorenz and Ariel Orda. QoS routing in networks with uncertain parameters. IEEE/ACM Transactions on Networking, 6(6):768–778, December 1998. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/
                                        journals/ton/1998-6-6/
                                        p768-lorenz/.

Libman:1999:DPA

[LO99] Lavy Libman and Ariel Orda. The designer’s perspective to atomic noncooperative networks. IEEE/ ACM Transactions on Networking, 7(6):875–884, December 1999. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1999-7-6/
p875-libman/.

Libman:2002:ORT

[LO02a] Lavy Libman and Ariel Orda. Optimal retrial and timeout strategies for accessing network resources. IEEE/ACM Transactions on Networking, 10(4):551– 564, August 2002. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Lorenz:2002:OPQ

[LO02b] Dean H. Lorenz and Ariel Orda. Optimal partition of QoS requirements on unicast paths and multicast trees. IEEE/ACM Transactions on Networking, 10(1):102–114, February 2002. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Li:2021:QEI

[LOFH21] Wenjie Li, Sharief M. A.

Oteafy, Marwan Fayed, and Hossam S. Hassanein. Quality of experience in ICN: Keep your lowbitrate close and highbitrate closer. IEEE/ ACM Transactions on Networking, 29(2):557–570, April 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2020.3044995.

Lazar:1997:VPB

[LOP97] Aurel A. Lazar, Ariel Orda, and Dimitrios E. Pendarakis. Virtual path bandwidth allocation in multiuser networks. IEEE/ ACM Transactions on Networking, 5(6):861–871, December 1997. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/
                                        journals/ton/1997-5-6/
                                        p861-lazar/.

Lorenz:2006:EQP

[LORS06] Dean H. Lorenz, Ariel Orda, Danny Raz, and Yuval Shavitt. Efficient QoS partition and routing of unicast and multicast. IEEE/ACM Transactions on Networking, 14(6): 1336–1347, December 2006. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Low:2000:EBB

[Low00] Steven H. Low. Equilibrium bandwidth and buffer allocations for elastic traffics. IEEE/ACM Transactions on Networking, 8(3): 373–383, 2000. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/2000-8-3/
p373-low/.

Low:2003:DMT

[Low03] Steven H. Low. A duality model of TCP and queue management algorithms. IEEE/ACM Transactions on Networking, 11 (4):525–536, August 2003. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Luo:2007:DRS

[LP07] Huiyu Luo and Gregory J.

Pottie. Designing routes for source coding with explicit side information in sensor networks. IEEE/ ACM Transactions on Networking, 15(6):1401–1413, December 2007. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Li:2024:PSF

[LP24] Weihe Li and Paul Patras. P-sketch: a fast and accurate sketch for persistent item lookup. IEEE/ACM Transactions on Networking, 32(2):987– 1002, April 2024. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2023.3306897.

Lopez-Perez:2013:DCR

[LPCVC13] David López-Pérez, Xiaoli Chu, Athanasios V. Vasilakos, and Holger Claussen. On distributed and coordinated resource allocation for interference mitigation in self-organizing LTE networks. IEEE/ ACM Transactions on Networking, 21(4):1145–1158, August 2013. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Li:2018:DCM

[LPD^{+}18] Yuanjie Li, Chunyi Peng, Haotian Deng, Zengwen Yuan, Guan-Hua Tu, Jiayao Li, Songwu Lu, and Xi Li. Device-customized multi-carrier network access on commodity Smartphones. IEEE/ACM Transactions on Networking, 26(6):2542–2555, December 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Li:2012:CBT

[LPF12] Pan Li, Miao Pan, and Yuguang Fang. Capacity bounds of threedimensional wireless ad hoc networks. IEEE/ ACM Transactions on Networking, 20(4):1304–1315, August 2012. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Lestas:2011:NES

[LPIH11] Marios Lestas, Andreas Pitsillides, Petros Ioannou, and George Hadjipollas. A new estimation scheme for the effective number of users in Internet congestion control. IEEE/ ACM Transactions on Networking, 19(5):1499–1512, October 2011. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Lee:2017:DLB

[LPJ^{+}17] Changhyun Lee, Chunjong Park, Keon Jang, Sue Moon, and Dongsu Han. DX: Latency-based congestion control for datacenters. IEEE/ACM Transactions on Networking, 25(1): 335–348, February 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Law:2010:DCH

[LPKF10] Lap Kong Law, Konstantinos Pelechrinis, Srikanth V. Krishnamurthy, and Michalis Faloutsos. Downlink capacity of hybrid cellular ad hoc networks. IEEE/ ACM Transactions on Networking, 18(1):243–256, February 2010. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Lindner:2023:ABE

[LPM23] Steffen Lindner, Gabriel Paradzik, and Michael Menth. Alternative best effort (ABE) for service differentiation: Trading loss versus delay. IEEE/ ACM Transactions on Networking, 31(4):1642– 1656, 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2022.3221553.

Liaskos:2011:TRL

[LPP11] Christos K. Liaskos, Sophia G. Petridou, and Georgios I. Papadimitriou. Towards realizable, low-cost broadcast systems for dynamic environments. IEEE/ACM Transactions on Networking, 19(2):383–392, April 2011. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Lee:2017:SRR

[LPR17] Kangwook Lee, Ramtin Pedarsani, and Kannan Ramchandran. On scheduling redundant requests with cancellation overheads. IEEE/ACM Transactions on Networking, 25 (2):1279–1290, April 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Liakopoulos:2019:ROF

[LPS19] Nikolaos Liakopoulos, Georgios S. Paschos, and Thrasyvoulos Spyropoulos. Robust optimization framework for proactive user association in UDNs: a data-driven approach. IEEE/ACM Transactions on Networking, 27(4):1683– 1695, August 2019. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Li:2014:MCM

[LPW14] Yixuan Li, Qiuyu Peng, and Xinbing Wang. Multicast capacity with max-min fairness for heterogeneous networks. IEEE/ACM Transactions on Networking, 22(2):622–635, April 2014. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Liu:2022:OCL

[LPWP22] Kaiyang Liu, Jun Peng, Jingrong Wang, and Jianping Pan. Optimal caching for low latency in distributed coded storage systems. IEEE/ ACM Transactions on Networking, 30(3):1132–1145, June 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2021.3133215.

Lam:2013:GRD

[LQ13] Simon S. Lam and Chen Qian. Geographic routing in d-dimensional spaces with guaranteed delivery and low stretch. IEEE/ ACM Transactions on Networking, 21(2):663–677, April 2013. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Luo:2016:EPR

[LQCC16] Wen Luo, Yan Qiao, Shigang Chen, and Min Chen. An efficient protocol for RFID multigroup threshold-based classification based on sampling and logical bitmap. IEEE/ ACM Transactions on Networking, 24(1):397–407, February 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Li:2007:MTO

[LQXX07] Jikai Li, Chunming Qiao, Jinhui Xu, and Dahai Xu. Maximizing throughput for optical burst switching networks. IEEE/ACM Transactions on Networking, 15(5):1163–1176, October 2007. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Liao:2024:PTC

[LQZ^{+}24] Zhengyu Liao, Shiyou Qian, Zhonglong Zheng, Jiange Zhang, Jian Cao, Guangtao Xue, and Minglu Li. PT-Tree: a cascading prefix tuple tree for packet classification in dynamic scenarios. IEEE/ ACM Transactions on Networking, 32(1):506– 519, 2024. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2023.3289029.

Loguinov:2003:EER

[LR03] Dmitri Loguinov and Hayder Radha. End-to-end rate-based congestion control: convergence properties and scalability analysis. IEEE/ACM Transactions on Networking, 11 (4):564–577, August 2003. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Lin:2009:DPE

[LR09] Xiaojun Lin and Shahzada B.

Rasool. Distributed and provably efficient algorithms for joint channelassignment, scheduling, and routing in multichannel ad hoc wireless networks. IEEE/ACM Transactions on Networking, 17 (6):1874–1887, December 2009. CODEN IEANEP.

ISSN 1063-6692 (print), 1558-2566 (electronic).

Lanante:2022:PAI

[LR22] Leonardo Lanante and Sumit Roy. Performance analysis of the IEEE 802.11ax OBSS PD-based spatial reuse. IEEE/ ACM Transactions on Networking, 30(2):616–628, April 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2021.3117816.

Leogrande:2015:MCP

[LRC15] Marco Leogrande, Fulvio Risso, and Luigi Ciminiera. Modeling complex packet filters with finite state automata. IEEE/ACM Transactions on Networking, 23(1):42–55, February 2015. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Luo:2010:EWM

[LRG10] Jun Luo, Catherine Rosenberg, and André Girard. Engineering wireless mesh networks: joint scheduling, routing, power control, and rate adaptation. IEEE/ ACM Transactions on Networking, 18(5):1387–1400, October 2010. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Lau:2008:CDR

[LRJ08] William Lau, Gustav Filip Rosenbaum, and Sanjay Jha. Comments on ‘Dynamic routing of restorable bandwidth-guaranteed tunnels using aggregated network resource usage information’. IEEE/ACM Transactions on Networking, 16(1):244–245, February 2008. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). See [KL03].

Liu:2007:QTF

[LRL07] Xiliang Liu, Kaliappa Ravindran, and Dmitri Loguinov. A queueingtheoretic foundation of available bandwidth estimation: single-hop analysis. IEEE/ACM Transactions on Networking, 15 (4):918–931, August 2007. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Liu:2008:SFA

[LRL08] Xiliang Liu, Kaliappa Ravindran, and Dmitri Loguinov. A stochastic foundation of available bandwidth estimation: multi-hop analysis. IEEE/ACM Transactions on Networking, 16(1):130– 143, February 2008. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Lun:2006:MCM

[LRM^{+}06] Desmond S. Lun, Niranjan Ratnakar, Muriel Médard, Ralf Koetter, David R. Karger, Tracey Ho, Ebad Ahmed, and Fang Zhao. Minimum-cost multicast over coded packet networks. IEEE/ACM Transactions on Networking, 14 (SI):2608–2623, June 2006. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

LaPorta:1993:PAM

[LS93a] Thomas F. La Porta and Mischa Schwartz. Performance analysis of MSP: feature-rich highspeed transport protocol. IEEE/ACM Transactions on Networking, 1(6):740– 753, December 1993. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL http://

                                        www.acm.org/pubs/citations/
                                        journals/ton/1993-1-6/
                                        p740-la_porta/.

Landry:1993:QSP

[LS93b] Randall Landry and Ioannis Stavrakakis. Queueing study of a 3-priority policy with distinct service strategies. IEEE/ ACM Transactions on Networking, 1(5):576–589, October 1993. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL http://www.

acm.org/pubs/citations/
journals/ton/1993-1-5/
p576-landry/.

Lee:1993:QAT

[LS93c] Duan-Shin Lee and Bhaskar Sengupta. Queueing analysis of a threshold based priority scheme for ATM networks. IEEE/ACM Transactions on Networking, 1(6):709–717, December 1993. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1993-1-6/
p709-lee/.

LaMaire:1994:TRS

[LS94] Richard O. LaMaire and Dimitrios N. Serpanos. Two-dimensional roundrobin schedulers for packet switches with multiple input queues. IEEE/ACM Transactions on Networking, 2(5):471–482, October 1994. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1994-2-5/
p471-lamaire/.

Landry:1997:SDJ

[LS97a] Randall Landry and Ioannis Stavrakakis. Study of delay jitter with and without peak rate enforcement. IEEE/ACM Transactions on Networking, 5(4):543– 553, August 1997. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1997-5-4/
p543-landry/.

Limb:1997:PET

[LS97b] John O. Limb and Dolors Sala. A protocol for efficient transfer of data over hybrid fiber/ coax systems. IEEE/ ACM Transactions on Networking, 5(6):872–881, December 1997. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1997-5-6/
p872-limb/.

Lombardo:1997:APC

[LS97c] Alfio Lombardo and Giovanni Schembra. An analytical paradigm to compare routing strategies in an ATM multimedia environment. IEEE/ACM Transactions on Networking, 5(6):958–969, December 1997. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1997-5-6/
p958-lombardo/.

Li:1999:DWR

[LS99] Ling Li and Arun K. Somani. Dynamic wavelength routing using congestion and neighborhood information. IEEE/ACM Transactions on Networking, 7(5):779–786, October 1999. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1999-7-5/

p779-li/. See comments [GLG04].

Li:2001:WAP

[LS01] Guangzhi Li and Rahul Simha. On the wavelength assignment problem in multifiber WDM star and ring networks. IEEE/ACM Transactions on Networking, 9(1):60– 68, 2001. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/articles/
journals/ton/2001-9-1/
p60-li/p60-li.pdf;   http:
//www.acm.org/pubs/citations/
journals/ton/2001-9-1/
p60-li/.

Lee:2003:RSH

[LS03a] Kayi Lee and Kai-Yeung Siu. On the reconfigurability of single-hub WDM ring networks. IEEE/ACM Transactions on Networking, 11(2):273–284, April 2003. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Lombardo:2003:PEA

[LS03b] Alfio Lombardo and Giovanni Schembra. Performance evaluation of an adaptive-rate MPEG encoder matching intserv traffic constraints. IEEE/ ACM Transactions on Networking, 11(1):47–65, February 2003. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Lin:2005:SND

[LS05a] Xiaojun Lin and Ness B. Shroff. Simplification of network dynamics in large systems. IEEE/ACM Transactions on Networking, 13(4):813–826, August 2005. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Lu:2005:CDR

[LS05b] Haibin Lu and Sartaj Sahni. Conflict detection and resolution in twodimensional prefix router tables. IEEE/ACM Transactions on Networking, 13 (6):1353–1363, December 2005. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Lee:2006:MOQ

[LS06a] Hyoung-Il Lee and Seung-Woo Seo. Matching output queueing with a multiple input/output-queued switch. IEEE/ACM Transactions on Networking, 14(1):121–132, February 2006. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Leung:2006:OPC

[LS06b] Kin-Kwong Leung and Chi Wan Sung. An opportunistic power control algorithm for cellular network. IEEE/ACM Transactions on Networking, 14 (3):470–478, 2006. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Liang:2006:GAA

[LS06c] Weifa Liang and Xiaojun Shen. A general approach for all-to-all routing in multihop WDM optical networks. IEEE/ACM Transactions on Networking, 14 (4):914–923, August 2006. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Lin:2006:IIS

[LS06d] Xiaojun Lin and Ness B.

Shroff. The impact of imperfect scheduling on cross-layer congestion control in wireless networks.

IEEE/ACM Transactions on Networking, 14(2):302– 315, April 2006. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Lin:2006:OBA

[LS06e] Xiaojun Lin and Ness B. Shroff. An optimizationbased approach for QoS routing in high-bandwidth networks. IEEE/ACM Transactions on Networking, 14(6):1348–1361, December 2006. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Lu:2007:MPC

[LS07] Haibin Lu and Sartaj Sahni. O(log W) multidimensional packet classification. IEEE/ACM Transactions on Networking, 15 (2):462–472, April 2007. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Lu:2009:SRS

[LS09] Wencheng Lu and Sartaj Sahni. Succinct representation of static packet classifiers. IEEE/ACM Transactions on Networking, 17(3): 803–816, June 2009. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Lu:2010:LPT

[LS10] Wencheng Lu and Sartaj Sahni. Low-power TCAMs for very large forwarding tables. IEEE/ACM Transactions on Networking, 18 (3):948–959, June 2010. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Le:2014:IRI

[LS14] Franck Le and João Luís Sobrinho. Interconnecting routing instances. IEEE/ ACM Transactions on Networking, 22(2):540–553, April 2014. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Lim:2016:PME

[LS16] Wan-Seon Lim and Kang G.

Shin. POEM: Minimizing energy consumption for WiFi tethering service. IEEE/ACM Transactions on Networking, 24(6):3785– 3797, December 2016. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Liu:2017:MCC

[LS17] Guoxin Liu and Haiying Shen. Minimum-cost cloud storage service across multiple cloud providers. IEEE/ACM Transactions on Networking, 25(4):2498– 2513, August 2017. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Li:2022:CSC

[LS22] Zhuozhao Li and Haiying Shen. Co-scheduler: a coflow-aware data-parallel job scheduler in hybrid electrical/optical datacenter networks. IEEE/ ACM Transactions on Networking, 30(4):1599– 1612, 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2022.3143232.

Langberg:2006:ECN

[LSB06] Michael Langberg, Alexander Sprintson, and Jehoshua Bruck. The encoding complexity of network coding. IEEE/ACM Transactions on Networking, 14 (SI):2386–2397, June 2006. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Li:1999:CCN

[LSC99a] Junyi Li, Ness B. Shroff, and Edwin K. P. Chong. Channel carrying: a novel handoff scheme for mobile cellular networks. IEEE/ACM Transactions on Networking, 7(1):38–50, February 1999. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1999-7-1/
p38-li/.

Li:1999:RPC

[LSC99b] Junyi Li, Ness B. Shroff, and K. P. Chong. A reduced-power channel reuse scheme for wireless packet cellular networks. IEEE/ ACM Transactions on Networking, 7(6):818–832, December 1999. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1999-7-6/
p818-li/.

Liu:2021:LCM

[LSC^{+}21] Jinwei Liu, Haiying Shen, Hongmei Chi, Husnu S. Narman, Yongyi Yang, Long Cheng, and Wingyan Chung. A low-cost multifailure resilient replication scheme for highdata availability in cloud storage. IEEE/ACM Transactions on Networking, 29(4):1436–1451, August 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2020.3027814.

Lin:2017:UCN

[LSCT17] Kate Ching-Ju Lin, Wei-Liang Shen, Ming-Syan Chen, and Kun Tan. Usercentric network MIMO with dynamic clustering. IEEE/ACM Transactions on Networking, 25(3):1910– 1923, June 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Li:2019:TAE

[LSDT19] Xiang Li, J. David Smith, Thang N. Dinh, and My T. Thai. TipTop: Almost exact solutions for influence maximization in billionscale networks. IEEE/ ACM Transactions on Networking, 27(2):649–661, April 2019. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Lopes:2016:PDF

[LSHZ16] Luis Amaral Lopes, Rute Sofia, Huseyin Haci, and Huiling Zhu. A proposal for dynamic frequency sharing in wireless networks. IEEE/ACM Transactions on Networking, 24(5):2621– 2633, October 2016. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Lotfi:2020:NNP

[LSK20] Mohammad Hassan Lotfi, Saswati Sarkar, and George Kesidis. Is non-neutrality profitable for the stakeholders of the Internet market? IEEE/ACM Transactions on Networking, 28(4):1435–1448, August 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2020.2981259.

Lu:2017:CDO

[LSL17] Zongqing Lu, Xiao Sun, and Thomas La Porta. Cooperative data offload in opportunistic networks: From mobile devices to infrastructure. IEEE/ACM Transactions on Networking, 25(6):3382–3395, December 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Lorenzo:2018:DST

[LSL^{+}18] Beatriz Lorenzo, Alireza Shams Shafigh, Jianqing Liu, Francisco J. Gonzalez-Castano, and Yuguang Fang. Data and spectrum trading policies in a trusted cognitive dynamic network architecture. IEEE/ACM Transactions on Networking, 26(3):1502–1516, June 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Luo:2021:FGT

[LSL^{+}21] Chuanwen Luo, Meghana N.

Satpute, Deying Li, Yongcai Wang, Wenping Chen, and Weili Wu. Finegrained trajectory optimization of multiple UAVs for efficient data gathering from WSNs. IEEE/ ACM Transactions on Networking, 29(1):162–175, February 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2020.3027555.

Laufer:2014:CLB

[LSLL14] Rafael Laufer, Theodoros Salonidis, Henrik Lundgren, and Pascal Le Guyadec. A cross-layer backpressure architecture for wireless multihop networks. IEEE/ACM Transactions on Networking, 22 (2):363–376, April 2014. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Loiseau:2014:IMI

[LSM^{+}14] Patrick Loiseau, Galina Schwartz, John Musacchio, Saurabh Amin, and S. Shankar Sastry. Incentive mechanisms for Internet congestion management: fixed-budget rebate versus time-of-day pricing. IEEE/ACM Transactions on Networking, 22(2):647– 661, April 2014. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Lin:2006:DDC

[LSMS06] Xiaojun Lin, Gaurav Sharma, Ravi R. Mazumdar, and Ness B. Shroff. Degenerate delay-capacity tradeoffs in ad-hoc networks with Brownian mobility. IEEE/ACM Transactions on Networking, 14(SI): 2777–2784, June 2006. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Lin:2007:AOE

[LSS07] Longbi Lin, Ness B. Shroff, and R. Srikant. Asymptotically optimal energy-aware routing for multihop wireless networks with renewable energy sources. IEEE/ ACM Transactions on Networking, 15(5):1021–1034, October 2007. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Laoutaris:2013:DTB

[LSS^{+}13] Nikolaos Laoutaris, Georgios Smaragdakis, Rade Stanojevic, Pablo Rodriguez, and Ravi Sundaram. Delay-tolerant bulk data transfers on the Internet. IEEE/ACM Transactions on Networking, 21(6): 1852–1865, December 2013. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Li:2022:SIM

[LSSC22] Zhuozhao Li, Tanmoy Sen, Haiying Shen, and Mooi Choo Chuah. A study on the impact of memory DoS attacks on cloud applications and exploring realtime detection schemes. IEEE/ACM Transactions on Networking, 30(4):1644– 1658, 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2022.3144895.

Lotfi:2017:EQS

[LSSK17] Mohammad Hassan Lotfi, Karthikeyan Sundaresan, Saswati Sarkar, and Mohammad Ali Khojastepour. Economics of quality sponsored data in non-neutral networks. IEEE/ACM Transactions on Networking, 25(4):2068–2081, August 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Lampson:1999:ILU

[LSV99] Butler Lampson, Venkatachary Srinivasan, and George Varghese. IP lookups using multiway and multicolumn search. IEEE/ACM Transactions on Networking, 7(3):324– 334, June 1999. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1999-7-3/
p324-lampson/.

Lang:2001:AOA

[LSV01] Jonathan P. Lang, Vishal Sharma, and Emmanouel A. Varvarigos. An analysis of oblivious and adaptive routing in optical networks with wavelength translation. IEEE/ACM Transactions on Networking, 9 (4):503–517, August 2001. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Lenzen:2015:PES

[LSW15] Christoph Lenzen, Philipp Sommer, and Roger Wattenhofer. PulseSync: an efficient and scalable clock synchronization protocol. IEEE/ACM Transactions on Networking, 23(3):717– 727, June 2015. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Liu:2016:JCC

[LSXS16] Jia Liu, Ness B. Shroff, Cathy H. Xia, and Hanif D. Sherali. Joint congestion control and routing optimization: an efficient second-order distributed approach. IEEE/ACM Transactions on Networking, 24(3):1404–1420, June 2016. CODEN IEANEP.

ISSN 1063-6692 (print), 1558-2566 (electronic).

Li:2013:RPA

[LSZW13] William Wei-Liang Li, Yuan Shen, Ying Jun Zhang, and Moe Z. Win. Robust power allocation for energy-efficient locationaware networks. IEEE/ ACM Transactions on Networking, 21(6):1918–1930, December 2013. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Logothetis:1994:RAD

[LT94a] Dimitris Logothetis and Kishor S. Trivedi. Reliability analysis of the double counter-rotating ring with concentrator attachments. IEEE/ACM Transactions on Networking, 2(5):520–532, October 1994. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/
                                        journals/ton/1994-2-5/
                                        p520-logothetis/.

Lundy:1994:SAS

[LT94b] Gilbert M. Lundy and H. Alphan Tipici. Specification and analysis of the SNR high-speed transport protocol. IEEE/ ACM Transactions on Networking, 2(5):483–496, October 1994. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1994-2-5/
p483-lundy/.

Liu:1995:BRP

[LT95] Zhen Liu and Don Towsley.

Burst reduction properties of rate-control throttles: downstream queue behavior. IEEE/ACM Transactions on Networking, 3(1):82–90, February 1995. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1995-3-1/
p82-liu/.

Lee:2002:SDA

[LT02] Young Lee and James M.

Tien. Static and dynamic approaches to modeling end-to-end routing in circuit-switched networks. IEEE/ACM Transactions on Networking, 10(5):693– 705, October 2002. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Liu:2016:OAA

[LT16] Alex X. Liu and Eric Torng. Overlay automata and algorithms for fast and scalable regular expression matching. IEEE/ ACM Transactions on Networking, 24(4):2400–2415, August 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Lety:2004:SSC

[LTB04] Emmanuel Léty, Thierry Turletti, and François Baccelli. SCORE: a scalable communication protocol for large-scale virtual environments. IEEE/ACM Transactions on Networking, 12 (2):247–260, April 2004. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Liu:2022:VRA

[LTCS22] Chang Liu, Jean Tourrilhes, Chen-Nee Chuah, and Puneet Sharma. Voyager: Revisiting available bandwidth estimation with a new class of methods — decreasing- chirptrain methods. IEEE/ ACM Transactions on Networking, 30(4):1717– 1732, 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2022.3152175.

Le:2017:RPL

[LTDM17] Anh Le, Arash Saber Tehrani, Alexandros Dimakis, and Athina Markopoulou. Recovery of packet losses in wireless broadcast for realtime applications. IEEE/ ACM Transactions on Networking, 25(2):676–689, April 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Liaskos:2019:NLM

[LTN^{+}19] Christos Liaskos, Ageliki Tsioliaridou, Shuai Nie, Andreas Pitsillides, Sotiris Ioannidis, and Ian F. Akyildiz. On the network-layer modeling and configuration of programmable wireless environments. IEEE/ ACM Transactions on Networking, 27(4):1696–1713, August 2019. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Lucerna:2010:AMB

[LTP10] Diego Lucerna, Massimo Tornatore, and Achille Pattavina. Algorithms and models for backup reprovisioning in WDM networks. IEEE/ACM Transactions on Networking, 18(6):1883– 1894, December 2010. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Liu:2005:AOS

[LTS05] Yu Liu, David Tipper, and Peerapon Siripongwutikorn. Approximating optimal spare capacity allocation by successive survivable routing. IEEE/ ACM Transactions on Networking, 13(1):198–211, February 2005. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Lakshmanan:2010:APS

[LTS10] Sriram Lakshmanan, Cheng-Lin Tsao, and Raghupathy Sivakumar. Aegis: physical space security for wireless networks with smart antennas. IEEE/ACM Transactions on Networking, 18(4): 1105–1118, August 2010. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Leland:1994:SNE

[LTWW94] Will E. Leland, Murad S. Taqqu, Walter Willinger, and Daniel V. Wilson. On the self-similar nature of Ethernet traffic (extended version). IEEE/ACM Transactions on Networking, 2(1):1–15, February 1994. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1994-2-1/
p1-leland/.

Liang:2006:FLA

[LTY06] Gang Liang, Nina Taft, and Bin Yu. A fast lightweight approach to origin-destination IP traffic estimation using partial measurements. IEEE/ ACM Transactions on Networking, 14(SI):2634–2648, June 2006. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Li:2008:ASE

[LTZ08] Yung-Ming Li, Yong Tan, and Yong-Pin Zhou. Analysis of scale effects in peerto-peer networks. IEEE/ ACM Transactions on Networking, 16(3):590–602, June 2008. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Li:2022:PWT

[LTZ^{+}22] Xu Li, Feilong Tang, Yanmin Zhu, Luoyi Fu, Jiadi Yu, Long Chen, and Jiacheng Liu. Processingwhile-transmitting: Costminimized transmission in SDN-based STINs. IEEE/ ACM Transactions on Networking, 30(1):243–256, February 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2021.3107413.

Lin:2014:EHA

[LÜ14] Hui Lin and Halit Üster.

Exact and heuristic algorithms for data-gathering cluster-based wireless sensor network design problem. IEEE/ACM Transactions on Networking, 22(3): 903–916, June 2014. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Low:1993:NAS

[LV93] Steven H. Low and Pravin P. Varaiya. A new approach to service provisioning in ATM networks. IEEE/ ACM Transactions on Networking, 1(5):547–553, October 1993. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1993-1-5/
p547-low/.

LeBoudec:2000:OSG

[LV00] Jean-Yves Le Boudec and Olivier Verscheure. Optimal smoothing for guaranteed service. IEEE/ACM Transactions on Networking, 8(6):689–696, 2000. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL

http://www.acm.org/pubs/
articles/journals/ton/
2000-8-6/p689-le\_boudec/
p689-le\_boudec.pdf;
http://www.acm.org/pubs/
citations/journals/ton/
2000-8-6/p689-le\_boudec/
;   http://www.acm.org/
pubs/citations/journals/
ton/2000-8-6/p689-le_
boudec/.

LeBoudec:2006:RTM

[LV06] Jean-Yves Le Boudec and Milan Vojnovic. The random trip model: stability, stationary regime, and perfect simulation. IEEE/ ACM Transactions on Networking, 14(6):1153–1166, December 2006. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

LaPorta:1996:CSL

[LVB96] Thomas F. La Porta, Malathi Veeraraghavan, and Richard W. Buskens. Comparison of signaling loads for PCS systems. IEEE/ACM Transactions on Networking, 4(6):840– 856, December 1996. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL http://

www.acm.org/pubs/citations/
journals/ton/1996-4-6/
p840-la_porta/.

Liu:2011:PRC

[LW11] Cong Liu and Jie Wu. Practical routing in a cyclic MobiSpace. IEEE/ACM Transactions on Networking, 19(2):369–382, April 2011. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Lin:2013:CAA

[LW13] Yunyue Lin and Qishi Wu.

Complexity analysis and algorithm design for advance bandwidth scheduling in dedicated networks. IEEE/ACM Transactions on Networking, 21(1):14– 27, February 2013. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Lu:2017:ENA

[LW17] Zhuo Lu and Cliff Wang. Enabling network antiinference via proactive strategies: a fundamental perspective. IEEE/ACM Transactions on Networking, 25(1):43–55, February 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Luo:2020:OAV

[LW20] Ziyue Luo and Chuan Wu. An online algorithm for VNF service chain scaling in datacenters. IEEE/ ACM Transactions on Networking, 28(3):1061–1073, June 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2020.2979263.

Liwang:2022:OEC

[LW22] Minghui Liwang and Xianbin Wang. Overbookingempowered computing resource provisioning in cloud-aided mobile edge networks. IEEE/ACM Transactions on Networking, 30(5):2289–2303, 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL

https://dl.acm.org/doi/
10.1109/TNET.2022.3167396.

Lin:2017:DBM

[LWAL17] Shih-Chun Lin, Pu Wang, Ian F. Akyildiz, and Min Luo. Delay-based maximum power-weight scheduling with heavytailed traffic. IEEE/ ACM Transactions on Networking, 25(4):2540–2555, August 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Lin:2013:DRS

[LWAT13] Minghong Lin, Adam Wierman, Lachlan L. H. Andrew, and Eno Thereska. Dynamic right-sizing for power-proportional data centers. IEEE/ACM Transactions on Networking, 21(5):1378–1391, October 2013. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Li:2014:AAP

[LWC^{+}14] Wei Li, Shengling Wang, Yong Cui, Xiuzhen Cheng, Ran Xin, Mznah A. Al-Rodhaan, and Abdullah Al-Dhelaan. AP association for proportional fairness in multirate WLANs. IEEE/ACM Transactions on Networking, 22(1):191– 202, February 2014. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Lin:2023:FSR

[LWC^{+}23] Fusheng Lin, Hongyu Wang, Guo Chen, Guihua Zhou, Tingting Xu, Dehui Wei, Li Chen, Yuanwei Lu, Andrew Qu, Hua Shao, and Hongbo Jiang. Fast, scalable and robust centralized routing for data center networks. IEEE/ACM Transactions on Networking, 31(6):2624– 2639, 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2023.3259541.

Li:2012:GEE

[LWCY12] Tao Li, Samuel S. Wu, Shigang Chen, and Mark C. K. Yang. Generalized energy-efficient algorithms for the RFID estimation problem. IEEE/ ACM Transactions on Networking, 20(6):1978–1990, December 2012. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Liebeherr:1996:EAC

[LWF96] Jörg Liebeherr, Dallas E. Wrege, and Domenico Ferrari. Exact admission control for networks with a bounded delay service. IEEE/ACM Transactions on Networking, 4(6):885–901, December 1996. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1996-4-6/
p885-liebeherr/.

Liu:2016:AMC

[LWK^{+}16] Bei Liu, Wei Wang, Donghyun Kim, Deying Li, Jingyi Wang, Alade O. Tokuta, and Yaolin Jiang. On approximating minimum 3-connected mdominating set problem in unit disk graph. IEEE/ ACM Transactions on Networking, 24(5):2690–2701, October 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Liu:2018:PCQ

[LWK^{+}18] Bei Liu, Wei Wang, Donghyun Kim, Yingshu Li, Sung-Sik Kwon, and Yaolin Jiang. On practical construction of quality fault-tolerant virtual backbone in homogeneous wireless networks. IEEE/ ACM Transactions on Networking, 26(1):412–421, February 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Li:2003:EDR

[LWKD03] Guangzhi Li, Dongmei Wang, Charles Kalmanek, and Robert Doverspike. Efficient distributed restoration path selection for shared mesh restoration. IEEE/ACM Transactions on Networking, 11(5):761– 771, October 2003. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Li:2004:DBM

[LWL04] C. Y. Li, P. K. A. Wai, and Victor O. K. Li. The decomposition of a blocking model for connectionoriented networks. IEEE/ ACM Transactions on Networking, 12(3):549–558, June 2004. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Li:2011:DAD

[LWL^{+}11] Dan Li, Jianping Wu, Jiangchuan Liu, Yong Cui, and Ke Xu. Defending against distance cheating in link-weighted applicationlayer multicast. IEEE/ ACM Transactions on Networking, 19(5):1448–1457, October 2011. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Li:2017:ISN

[LWL17] Xiaohang Li, Chih-Chun Wang, and Xiaojun Lin. Inter-session network coding schemes for 1-to-2 downlink access-point networks with sequential hard deadline constraints. IEEE/ACM Transactions on Networking, 25(1):624– 638, February 2017. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Li:2023:RRA

[LWL^{+}23] Jiawei Li, Chuyu Wang, Ang Li, Dianqi Han, Yan Zhang, Jinhang Zuo, Rui Zhang, Lei Xie, and Yanchao Zhang. Rhythmic RFID authentication. IEEE/ACM Transactions on Networking, 31(2):877– 890, April 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2022.3204204.

Li:2016:VMT

[LWLL16] Hongxing Li, Chuan Wu, Zongpeng Li, and Francis C. M. Lau. Virtual machine trading in a federation of clouds: individual profit and social welfare maximization. IEEE/ACM Transactions on Networking, 24(3):1827–1840, June 2016. CODEN IEANEP.

ISSN 1063-6692 (print), 1558-2566 (electronic).

Li:2019:ERD

[LWP^{+}19] Zonghui Li, Hai Wan, Zaiyu Pang, Qiubo Chen, Yangdong Deng, Xibin Zhao, Yue Gao, Xiaoyu Song, and Ming Gu. An enhanced reconfiguration for deterministic transmission in time-triggered networks. IEEE/ACM Transactions on Networking, 27(3):1124– 1137, June 2019. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Li:2018:BGN

[LWQ^{+}18] Bo Li, Junfeng Wu, Hongsheng Qi, Alexandre Proutiere, and Guodong Shi. Boolean gossip networks. IEEE/ACM Transactions on Networking, 26(1):118–130, February 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Liu:2015:FSE

[LWR15] Qiang Liu, Xin Wang, and Nageswara S. V. Rao. Fusion of state estimates over long-haul sensor networks with random loss and delay. IEEE/ACM Transactions on Networking, 23 (2):644–656, April 2015. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Liu:2016:ERR

[LWR^{+}16] Qiang Liu, Xin Wang, Nageswara S. V. Rao, Katharine Brigham, and B. V. K. Vijaya Kumar. Effect of retransmission and retrodiction on estimation and fusion in long-haul sensor networks. IEEE/ ACM Transactions on Networking, 24(1):449–461, February 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Liang:2015:TDR

[LWT^{+}15] Qingkai Liang, Xinbing Wang, Xiaohua Tian, Fan Wu, and Qian Zhang. Twodimensional route switching in cognitive radio networks: a game-theoretical framework. IEEE/ACM Transactions on Networking, 23(4):1053–1066, August 2015. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Li:2021:GIB

[LWT^{+}21] Qiang Li, Zhihao Wang, Dawei Tan, Jinke Song, Haining Wang, Limin Sun, and Jiqiang Liu. Geo-CAM: an IP-based geolocation service through finegrained and stable webcam landmarks. IEEE/ ACM Transactions on Networking, 29(4):1798–1812, August 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2021.3073926.

Li:2019:TSE

[LWW^{+}19a] Xin Li, Minmei Wang, Huazhe Wang, Ye Yu, and Chen Qian. Toward secure and efficient communication for the Internet of Things. IEEE/ACM Transactions on Networking, 27(2):621–634, April 2019. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Liu:2019:AOS

[LWW^{+}19b] Shengchao Liu, Jianping Weng, Jessie Hui Wang, Changqing An, Yipeng Zhou, and Jilong Wang. An adaptive online scheme for scheduling and resource enforcement in storm. IEEE/ ACM Transactions on Networking, 27(4):1373–1386, August 2019. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Lam:1997:GPS

[LX97] Simon S. Lam and Geoffrey G. Xie. Group priority scheduling. IEEE/ ACM Transactions on Networking, 5(2):205–218, April 1997. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1997-5-2/
p205-lam/.

Liu:2021:EEE

[LX21] Libin Liu and Hong Xu.

Elasecutor: Elastic executor scheduling in data analytics systems. IEEE/ ACM Transactions on Networking, 29(2):681–694, April 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2021.3050927.

Liu:2024:PSS

[LX24] Yongqiang Liu and Xike Xie. A probabilistic sketch for summarizing cold items of data streams. IEEE/ ACM Transactions on Networking, 32(2):1287–1302, April 2024. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2023.3316426.

Lu:2005:ABP

[LXC05] Kejie Lu, Gaoxi Xiao, and Imrich Chlamtac. Analysis of blocking probability for distributed lightpath establishment in WDM optical networks. IEEE/ ACM Transactions on Networking, 13(1):187–197, February 2005. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Liu:2017:REB

[LXL^{+}17a] Xiulong Liu, Bin Xiao, Keqiu Li, Alex X. Liu, Jie Wu, Xin Xie, and Heng Qi. RFID estimation with blocker tags. IEEE/ACM Transactions on Networking, 25(1):224– 237, February 2017. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Liu:2017:FTP

[LXL^{+}17b] Xiulong Liu, Xin Xie, Keqiu Li, Bin Xiao, Jie Wu, Heng Qi, and Dawei Lu. Fast tracking the population of key tags in largescale anonymous RFID systems. IEEE/ACM Transactions on Networking, 25(1):278–291, February 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Liu:2019:EPB

[LXL^{+}19] Jia Liu, Bin Xiao, Xuan Liu, Kai Bu, Lijun Chen, and Changhai Nie. Efficient polling-based information collection in RFID systems. IEEE/ACM Transactions on Networking, 27(3):948– 961, June 2019. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Li:2022:VAT

[LXL^{+}22a] Chong Li, Sisu Xi, Chenyang Lu, Roch Guérin, and Christopher D. Gill. Virtualizationaware traffic control for soft real-time network traffic on Xen. IEEE/ ACM Transactions on Networking, 30(1):257–270, February 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2021.3114055.

Liu:2022:SCS

[LXL^{+}22b] Xilai Liu, Yan Xu, Peng Liu, Tong Yang, Jiaqi Xu, Lun Wang, Gaogang Xie, Xiaoming Li, and Steve Uhlig. SEAD counter: Self-adaptive counters with different counting ranges. IEEE/ ACM Transactions on Networking, 30(1):90–106, February 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2021.3107418.

Li:2020:QDE

[LXLC20] Ye Li, Hong Xie, John C. S. Lui, and Kenneth L. Calvert. Quantifying deployability and evolvability of future Internet architectures via economic models. IEEE/ ACM Transactions on Networking, 28(5):1995–2008, October 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2020.3006207.

Li:2017:LMS

[LXW^{+}17] Li, Ke Xu, Dan Wang, Chunyi Peng, Kai Zheng, Rashid Mijumbi, and Qingyang Xiao. A longitudinal measurement study of TCP performance and behavior in 3G/4G networks over high speed rails. IEEE/ACM Transactions on Networking, 25(4):2195– 2208, August 2017. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Liu:2019:ERQ

[LXW^{+}19] Xiulong Liu, Xin Xie, Shangguang Wang, Jia Liu, Didi Yao, Jiannong Cao, and Keqiu Li. Efficient range queries for large-scale sensor-augmented RFID systems. IEEE/ACM Transactions on Networking, 27(5):1873–1886, October 2019. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Li:2020:QAF

[LXW^{+}20] Xiaocan Li, Kun Xie, Xin Wang, Gaogang Xie, Dongliang Xie, Zhenyu Li, Jigang Wen, Zulong Diao, and Tian Wang. Quick and accurate false data detection in mobile crowd sensing. IEEE/ ACM Transactions on Networking, 28(3):1339–1352, June 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2020.2982685.

Luo:2014:HHA

[LXX^{+}14] Layong Luo, Gaogang Xie, Yingke Xie, Laurent Mathy, and Kavé Salamatian. A hybrid hardware architecture for high-speed IP lookups and fast route updates. IEEE/ACM Transactions on Networking, 22 (3):957–969, June 2014. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Liang:2017:AAC

[LXX^{+}17] Weifa Liang, Zichuan Xu, Wenzheng Xu, Jiugen Shi, Guoqiang Mao, and Sajal K. Das. Approximation algorithms for charging reward maximization in rechargeable sensor networks via a mobile charger. IEEE/ACM Transactions on Networking, 25(5):3161– 3174, October 2017. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Liao:2024:AFL

[LXX^{+}24] Yunming Liao, Yang Xu, Hongli Xu, Zhiwei Yao, Lun Wang, and Chunming Qiao. Accelerating federated learning with data and model parallelism in edge computing. IEEE/ACM Transactions on Networking, 32(1): 904–918, 2024. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2023.3299851.

Li:2014:SPE

[LXY^{+}14] Qi Li, Mingwei Xu, Yuan Yang, Lixin Gao, Yong Cui, and Jianping Wu. Safe and practical energyefficient detour routing in IP networks. IEEE/ACM Transactions on Networking, 22(6):1925–1937, December 2014. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Liu:2021:ISD

[LXZ^{+}21] Jianchun Liu, Hongli Xu, Gongming Zhao, Chen Qian, Xingpeng Fan, Xuwei Yang, and He Huang. Incremental server deployment for software-defined NFV-enabled networks. IEEE/ACM Transactions on Networking, 29(1):248– 261, February 2021. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL https:/

/dl.acm.org/doi/10.1109/
TNET.2020.3030298.

Leung:1994:MMV

[LY94] Yiu-Wing Leung and Tak-Shing Yum. A modular multirate video distribution system: design and dimensioning. IEEE/ ACM Transactions on Networking, 2(6):549–557, December 1994. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1994-2-6/
p549-leung/.

Li:2010:ASI

[LY10] Xiaolong Li and Homayoun Yousefi’zadeh. Analysis, simulation, and implementation of VCP: a wireless profiling. IEEE/ ACM Transactions on Networking, 18(5):1345–1358, October 2010. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Liu:2022:USD

[LY22] Xin Liu and Lei Ying. Universal scaling of distributed queues under load balancing in the super-Halfin– Whitt regime. IEEE/ ACM Transactions on Networking, 30(1):190–201, February 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2021.3105480.

Li:2011:SBD

[LYC11] Jung-Shian Li, Ching-Fang Yang, and Jian-Hong Chen. Star-block design in twolevel survivable optical networks. IEEE/ACM Transactions on Networking, 19 (2):526–539, April 2011. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Lu:2019:LRB

[LYC^{+}19] Li Lu, Jiadi Yu, Yingying Chen, Hongbo Liu, Yanmin Zhu, Linghe Kong, and Minglu Li. Lip readingbased user authentication through acoustic sensing on Smartphones. IEEE/ ACM Transactions on Networking, 27(1):447–460, February 2019. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Lin:2021:MCD

[LYD^{+}21] Chi Lin, Ziwei Yang, Haipeng Dai, Liangxian Cui, Lei Wang, and Guowei Wu. Minimizing charging delay for directional charging. IEEE/ACM Transactions on Networking, 29(6):2478–2493, December 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2021.3095280.

Lin:2019:TTS

[LYDA19] Qiongzheng Lin, Lei Yang, Chunhui Duan, and Zhenlin An. Tash: Toward selective reading as hash primitives for Gen2 RFIDs. IEEE/ACM Transactions on Networking, 27(2):819– 834, April 2019. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Lin:2023:NMD

[LYH^{+}23] Peng Lin, Kejiang Ye, Yishen Hu, Yanying Lin, and Cheng-Zhong Xu. A novel multimodal deep learning framework for encrypted traffic classification. IEEE/ACM Transactions on Networking, 31(3): 1369–1384, June 2023. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL https:/

/dl.acm.org/doi/10.1109/
TNET.2022.3215507.

Lou:2021:BHA

[LYKT21] Jiadong Lou, Xu Yuan, Sastry Kompella, and Nian-Feng Tzeng. Boosting or hindering: AoI and throughput interrelation in routing-aware multi-hop wireless networks. IEEE/ ACM Transactions on Networking, 29(3):1008–1021, June 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2021.3059694.

Luo:2007:CSS

[LYL07] Hongbin Luo, Hongfang Yu, and Lemin Li. Comments on ‘Segment shared protection in mesh communication networks with bandwidth guaranteed tunnels’. IEEE/ACM Transactions on Networking, 15 (6):1616, December 2007. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). See [HTC04].

Lin:2021:TOP

[LYL21] I-Chieh Lin, Yu-Hsuan Yeh, and Kate Ching-Ju Lin. Toward optimal partial parallelization for service function chaining. IEEE/ACM Transactions on Networking, 29(5):2033–2044, October 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2021.3075709.

Liu:2022:TER

[LYL^{+}22a] Jia Liu, Xi Yu, Xuan Liu, Xingyu Chen, Haisong Liu, Yanyan Wang, and Lijun Chen. Time-efficient range detection in commodity RFID systems. IEEE/ACM Transactions on Networking, 30(3):1118– 1131, June 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2021.3138083.

Liu:2022:ARN

[LYL^{+}22b] Tzu-Hsuan Liu, Che-Hao Yu, Yi-Jheng Lin, Chia-Ming Chang, Cheng-Shang Chang, and Duan-Shin Lee. ALOHA receivers: a network calculus approach for analyzing coded multiple access with SIC. IEEE/ACM Transactions on Networking, 30(2):840– 854, April 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2021.3123685.

Li:2022:LQD

[LYLW22] Chun Li, Yunyun Yang, Hui Liang, and Boying Wu. Learning quantum drift-diffusion phenomenon by physics-constraint machine learning. IEEE/ ACM Transactions on Networking, 30(5):2090– 2101, 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.

                                        acm.org/doi/10.1109/TNET.
                                        2022.3158987.

Li:2017:SFW

[LYMA^{+}17] Songze Li, Qian Yu, Mohammad Ali Maddah-Ali, A. Salman Avestimehr, Songze Li, Qian Yu, Mohammad Ali Maddah-Ali, and A. Salman Avestimehr. A scalable framework for wireless distributed computing. IEEE/ACM Transactions on Networking, 25(5):2643–2654, October 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Leonard:2007:LBN

[LYRL07] Derek Leonard, Zhongmei Yao, Vivek Rai, and Dmitri Loguinov. On lifetime-based node failure and stochastic resilience of decentralized peer-to-peer networks. IEEE/ACM Transactions on Networking, 15(3):644–656, June 2007. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Leung:1993:CMT

[LYS93] Kin K. Leung, Raymond W. Yeung, and Bhaskar Sengupta. A credit manager for traffic regulation in high-speed networks: a queueing analysis. IEEE/ACM Transactions on Networking, 1(2):236– 245, April 1993. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1993-1-2/
p236-leung/.

Liu:2011:TOO

[LYS11] Shihuan Liu, Lei Ying, and R. Srikant. Throughputoptimal opportunistic scheduling in the presence of flowlevel dynamics. IEEE/ ACM Transactions on Networking, 19(4):1057–1070, August 2011. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Liu:2018:CAL

[LYS^{+}18] Jiayi Liu, Qinghai Yang, Gwendal Simon, Jiayi Liu, Gwendal Simon, and Qinghai Yang. Congestion avoidance and load balancing in content placement and request redirection for mobile CDN. IEEE/ACM Transactions on Networking, 26(2):851–863, April 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Li:2016:PSN

[LYSZ16] Dan Li, Yirong Yu, Junxiao Shi, and Beichuan Zhang. PALS: Saving network power with low overhead to ISPs and applications. IEEE/ACM Transactions on Networking, 24(5):2913–2925, October 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Liew:1998:CAA

[LyT98] Soung C. Liew and Derek Chi yin Tse. A control-theoretic approach to adapting VBR compressed video for transport over a CBR communications channel. IEEE/ ACM Transactions on Networking, 6(1):42–55, February 1998. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/
                                        journals/ton/1998-6-1/
                                        p42-liew/.

Lu:2018:LCC

[LYW^{+}18] Jianyuan Lu, Tong Yang, Yi Wang, Huichen Dai, Xi Chen, Linxiao Jin, Haoyu Song, and Bin Liu. Low computational cost Bloom filters. IEEE/ ACM Transactions on Networking, 26(5):2254–2267, October 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Liu:2021:DOU

[LYW^{+}21] Wenbin Liu, Yongjian Yang, En Wang, Hengzhi Wang, Zihe Wang, and Jie Wu. Dynamic online user recruitment with (non- ) submodular utility in mobile CrowdSensing. IEEE/ ACM Transactions on Networking, 29(5):2156–2169, October 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2021.3083955.

Leonard:2008:SDP

[LYWL08] Derek Leonard, Zhongmei Yao, Xiaoming Wang, and Dmitri Loguinov. On static and dynamic partitioning behavior of largescale P2P networks. IEEE/ ACM Transactions on Networking, 16(6):1475–1488, December 2008. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Li:2022:PFG

[LYY^{+}22] Yuanpeng Li, Xiang Yu, Yilong Yang, Yang Zhou, Tong Yang, Zhuo Ma, and Shigang Chen. Pyramid Family: Generic frameworks for accurate and fast flow size measurement. IEEE/ACM Transactions on Networking, 30(2):586– 600, April 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2021.3120085.

Lovewell:2017:PSC

[LYZ^{+}17] Rebecca Lovewell, Qianwen Yin, Tianrong Zhang, Jasleen Kaur, and Frank Donelson Smith. Packetscale congestion control paradigm. IEEE/ACM Transactions on Networking, 25(1):306–319, February 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Liu:2023:OPM

[LYZ^{+}23a] Xuezheng Liu, Zirui Yan, Yipeng Zhou, Di Wu, Xu Chen, and Jessie Hui Wang. Optimizing parameter mixing under constrained communications in parallel federated learning. IEEE/ACM Transactions on Networking, 31(6): 2640–2652, 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2023.3257236.

Lu:2023:LAK

[LYZ^{+}23b] Xiaofeng Lu, Fan Yang, Luwen Zou, Pietro Lio, and Pan Hui. An LTE authentication and key agreement protocol based on the ECC self-certified public key. IEEE/ACM Transactions on Networking, 31(3):1101– 1116, June 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2022.3207360.

Levy:2006:EPD

[LZ06] Hanoch Levy and Haim Zlatokrilov. The effect of packet dispersion on voice applications in IP networks. IEEE/ACM Transactions on Networking, 14 (2):277–288, April 2006. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Li:2009:RAS

[LZ09] Jun Li and Yiqiang Q.

Zhao. Resequencing analysis of stop-and-wait ARQ for parallel multichannel communications. IEEE/ ACM Transactions on Networking, 17(3):817–830, June 2009. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Li:2013:RWG

[LZ13] Yanhua Li and Zhi-Li Zhang. Random walks and Green’s function on digraphs: a framework for estimating wireless transmission costs. IEEE/ ACM Transactions on Networking, 21(1):135–148, February 2013. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Lv:2023:MPF

[LZ23] Qian Lv and Zuqing Zhu.

On the multilayer planning of filterless optical networks with OTN encryption. IEEE/ACM Transactions on Networking, 31(6): 2529–2544, 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2023.3256409.

Li:2023:PRN

[LZB^{+}23] Wenhao Li, Xiao-Yu Zhang, Huaifeng Bao, Haichao Shi, and Qiang Wang. ProGraph: Robust network traffic identification with graph propagation. IEEE/ACM Transactions on Networking, 31(3):1385– 1399, June 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2022.3216603.

Liew:2009:BMD

[LZC09] Soung Chang Liew, Ying Jun Zhang, and Da Rui Chen. Bounded-mean-delay throughput and nonstarvation conditions in Aloha network. IEEE/ACM Transactions on Networking, 17(5):1606– 1618, October 2009. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Luo:2017:AAC

[LZC^{+}17] Jingjing Luo, Jinbei Zhang, Ying Cui, Li Yu, and Xinbing Wang. Asymptotic analysis on content placement and retrieval in MANETs. IEEE/ACM Transactions on Networking, 25(2):1103–1118, April 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Liu:2020:NUM

[LZC20] Qingyu Liu, Haibo Zeng, and Minghua Chen. Network utility maximization under maximum delay constraints and throughput requirements. IEEE/ACM Transactions on Networking, 28(5):2132–2145, October 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2020.3007842.

Liu:2022:MAT

[LZC22] Qingyu Liu, Haibo Zeng, and Minghua Chen. Minimizing AoI with throughput requirements in multipath network communication. IEEE/ACM Transactions on Networking, 30(3): 1203–1216, June 2022. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL https:/

/dl.acm.org/doi/10.1109/
TNET.2021.3135494.

Li:2024:PPP

[LZC^{+}24] Yu Li, Jiaheng Zhang, Junjie Chen, Yicong Chen, Ning Xie, and Hongbin Li. Privacy-preserving physical-layer authentication under cooperative attacks. IEEE/ ACM Transactions on Networking, 32(2):1171–1186, April 2024. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2023.3311470.

Li:2014:MST

[LZES14] Shuang Li, Zizhan Zheng, Eylem Ekici, and Ness Shroff. Maximizing system throughput by cooperative sensing in cognitive radio networks. IEEE/ ACM Transactions on Networking, 22(4):1245–1256, August 2014. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Li:2009:ACW

[LZF09] Pan Li, Chi Zhang, and Yuguang Fang. Asymptotic connectivity in wireless ad hoc networks using directional antennas. IEEE/ ACM Transactions on Networking, 17(4):1106–1117, August 2009. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Presti:1999:STS

[LZKT99] Francesco Lo Presti, Zhi-Li Zhang, Jim Kurose, and Don Towsley. Source time scale and optimal buffer/ bandwidth tradeoff for heterogeneous regulated traffic in a network node. IEEE/ACM Transactions on Networking, 7(4):490– 501, August 1999. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1999-7-4/
p490-presti/.

Liang:2011:OBS

[LZL11] Chao Liang, Miao Zhao, and Yong Liu. Optimal bandwidth sharing in multiswarm multiparty P2P video-conferencing systems. IEEE/ACM Transactions on Networking, 19(6):1704–1716, December 2011. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Li:2012:MPA

[LZL12] Yongkun Li, Bridge Qiao Zhao, and John C. S. Lui. On modeling product advertisement in large-scale online social networks. IEEE/ACM Transactions on Networking, 20(5):1412– 1425, October 2012. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Li:2014:LLB

[LZL^{+}14] Feng Li, Chi Zhang, Jun Luo, Shi-Qing Xin, and Ying He. LBDP: localized boundary detection and parametrization for 3-D sensor networks. IEEE/ACM Transactions on Networking, 22(2):567– 579, April 2014. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Li:2018:TSI

[LZL^{+}18] Wenxin Li, Xiaobo Zhou, Keqiu Li, Heng Qi, and Deke Guo. TrafficShaper: Shaping inter-datacenter traffic to reduce the transmission cost. IEEE/ACM Transactions on Networking, 26(3):1193–1206, June 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Li:2020:UEA

[LZL^{+}20] Yang Li, Jianwei Zheng, Zhenhua Li, Yunhao Liu, Feng Qian, Sen Bai, Yao Liu, and Xianlong Xin. Understanding the ecosystem and addressing the fundamental concerns of commercial MVNO. IEEE/ ACM Transactions on Networking, 28(3):1364–1377, June 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.

acm.org/doi/10.1109/TNET.
2020.2981514.

Li:2021:PBM

[LZL^{+}21] Ning Li, Zhaoxin Zhang, Alex X. Liu, Xin Yuan, and Yexia Cheng. Pairwisebased multi-attribute decision making approach for wireless network. IEEE/ ACM Transactions on Networking, 29(4):1687–1702, August 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2021.3074002.

Li:2022:CCL

[LZS^{+}22] Hao Li, Peng Zhang, Guangda Sun, Wanyue Cao, Chengchen Hu, Danfeng Shan, Tian Pan, and Qiang Fu. Compiling crosslanguage network programs into hybrid data plane. IEEE/ACM Transactions on Networking, 30(3):1088– 1103, June 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2021.3132303.

Lee:2010:SEE

[LZSS10] Sanghwan Lee, Zhi-Li Zhang, Sambit Sahu, and Debanjan Saha. On suitability of Euclidean embedding for host-based network coordinate systems. IEEE/ACM Transactions on Networking, 18(1):27– 40, February 2010. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Li:2015:GHB

[LZW^{+}15] Dan Li, Jing Zhu, Jianping Wu, Junjie Guan, and Ying Zhang. Guaranteeing heterogeneous bandwidth demand in multitenant data center networks. IEEE/ ACM Transactions on Networking, 23(5):1648–1660, October 2015. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Li:2021:EPF

[LZW^{+}21] Guanyu Li, Menghao Zhang, Shicheng Wang, Chang Liu, Mingwei Xu, Ang Chen, Hongxin Hu, Guofei Gu, Qi Li, and Jianping Wu. Enabling performant, flexible and cost-efficient DDoS defense with programmable switches. IEEE/ACM Transactions on Networking, 29(4):1509–1526, August 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2021.3062621.

Li:2021:RAM

[LZX^{+}21] Tong Li, Kai Zheng, Ke Xu, Rahul Arvind Jadhav, Tao Xiong, Keith Winstein, and Kun Tan. Revisiting acknowledgment mechanism for transport control: Modeling, analysis, and implementation. IEEE/ACM Transactions on Networking, 29(6):2678–2692, December 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2021.3101011.

Liu:2024:ABW

[LZX^{+}24] Jianchun Liu, Qingmin Zeng, Hongli Xu, Yang Xu, Zhiyuan Wang, and He Huang. Adaptive blockwise regularization and knowledge distillation for enhancing federated learning. IEEE/ACM Transactions on Networking, 32(1): 791–805, 2024. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2023.3301972.

Li:2014:RDM

[LZXF14] Dan Li, Hongze Zhao, Mingwei Xu, and Xiaoming Fu. Revisiting the design of mega data centers: considering heterogeneity among containers. IEEE/ ACM Transactions on Networking, 22(5):1503–1515, October 2014. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Lin:2020:CSD

[LZY20] Sen Lin, Junshan Zhang, and Lei Ying. Crowdsensing for spectrum discovery: a waze-inspired design via smartphone sensing. IEEE/ACM Transactions on Networking, 28(2):750– 763, April 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL https://

                                        dl.acm.org/doi/abs/10.
                                        1109/TNET.2020.2976927.

Li:2022:HCB

[LZY^{+}22] Feng Li, Jichao Zhao, Dongxiao Yu, Xiuzhen Cheng, and Weifeng Lv. Harnessing context for budget-limited crowdsensing with massive uncertain workers. IEEE/ ACM Transactions on Networking, 30(5):2231– 2245, 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2022.3169180.

Li:2022:UDP

[LZZ^{+}22a] Meng Li, Liehuang Zhu, Zijian Zhang, Chhagan Lal, Mauro Conti, and Mamoun Alazab. Userdefined privacy-preserving traffic monitoring against n-by-1 jamming attack.

IEEE/ACM Transactions on Networking, 30(5):2060– 2073, 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2022.3157654.

Liu:2022:AFL

[LZZ^{+}22b] Xuezheng Liu, Zhicong Zhong, Yipeng Zhou, Di Wu, Xu Chen, Min Chen, and Quan Z. Sheng. Accelerating federated learning via parallel servers: a theoretically guaranteed approach. IEEE/ ACM Transactions on Networking, 30(5):2201– 2215, 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2022.3168939.

Li:2012:SVM

[LZZR12] Peilong Li, Honghai Zhang, Baohua Zhao, and Sampath Rangarajan. Scalable video multicast with adaptive modulation and coding in broadband wireless data systems. IEEE/ACM Transactions on Networking, 20(1):57–68, February 2012. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Mokhtar:1998:AWR

[MA98] Ahmed Mokhtar and Murat Azizoğlu. Adaptive wavelength routing in all-optical networks. IEEE/ACM Transactions on Networking, 6(2):197– 206, April 1998. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/
                                        journals/ton/1998-6-2/
                                        p197-mokhtar/.

Mao:2012:TBU

[MA12] Guoqiang Mao and Brian D. O. Anderson. Towards a better understanding of large-scale network models. IEEE/ACM Transactions on Networking, 20 (2):408–421, April 2012. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Ma:2016:SCV

[Ma16a] Richard T. B. Ma. Subsidization competition: Vitalizing the neutral Internet. IEEE/ACM Transactions on Networking, 24(4): 2563–2576, August 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Ma:2016:UBP

[Ma16b] Richard T. B. Ma. Usagebased pricing and competition in congestible network service markets. IEEE/ ACM Transactions on Networking, 24(5):3084–3097, October 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Maatouk:2019:EET

[MAE19] Ali Maatouk, Mohamad Assaad, and Anthony Ephremides. Energy efficient and throughput optimal CSMA scheme. IEEE/ ACM Transactions on Networking, 27(1):316–329, February 2019. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Maatouk:2020:AIC

[MAE20] Ali Maatouk, Mohamad Assaad, and Anthony Ephremides. On the age of information in a CSMA environment. IEEE/ ACM Transactions on Networking, 28(2):818–831, April 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL https://

dl.acm.org/doi/abs/10.
1109/TNET.2020.2971350.

Maddah-Ali:2015:DCC

[MAN15] Mohammad Ali Maddah-Ali and Urs Niesen. Decentralized coded caching attains order-optimal memoryrate tradeoff. IEEE/ ACM Transactions on Networking, 23(4):1029–1040, August 2015. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Michaloliakos:2018:ACM

[MAPZ18] Antonios Michaloliakos, Weng Chon Ao, Konstantinos Psounis, and Yonglong Zhang. Asynchronously coordinated multi-timescale beamforming architecture for multi-cell networks. IEEE/ACM Transactions on Networking, 26(1):61– 75, February 2018. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Marcus:1996:AQA

[Mar96] William S. Marcus. An architecture for QoS analysis and experimentation. IEEE/ACM Transactions on Networking, 4(4):597– 603, August 1996. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/
                                        journals/ton/1996-4-4/
                                        p597-marcus/.

Marbach:2003:PSM

[Mar03] Peter Marbach. Priority service and max-min fairness. IEEE/ACM Transactions on Networking, 11 (5):733–746, October 2003. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Marbach:2004:ASP

[Mar04]  Peter Marbach.  Anal-

ysis of a static pricing scheme for priority services. IEEE/ACM Transactions on Networking, 12 (2):312–325, April 2004. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Mutlu:2009:SPS

[MAS09] Huseyin Mutlu, Murat Alanyali, and David Starobinski. Spot pricing of secondary spectrum access in wireless cellular networks. IEEE/ACM Transactions on Networking, 17(6):1794– 1804, December 2009. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Ma:2006:SAT

[MBA06] Liangping Ma, Kenneth E.

Barner, and Gonzalo R. Arce. Statistical analysis of TCP’s retransmission timeout algorithm. IEEE/ACM Transactions on Networking, 14(2):383–396, April 2006. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Marsan:1994:LEP

[MBC^{+}94] Marco Ajmone Marsan, Andrea Bianco, Luigi Ciminiera, Riccardo Sisto, and Adriano Valenzano. A LOTOS extension for the performance analysis of distributed systems. IEEE/ACM Transactions on Networking, 2(2):151– 165, April 1994. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1994-2-2/
p151-marsan/.

Medard:2002:GLB

[MBF^{+}02] Muriel Médard, Richard A. Barry, Steven G. Finn, Wenbo He, and Steven S. Lumetta. Generalized loopback recovery in optical mesh networks. IEEE/ ACM Transactions on Networking, 10(1):153–164, February 2002. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Marsan:2002:PMS

[MBG^{+}02] Marco Ajmone Marsan, Andrea Bianco, Paolo Giaccone, Emilio Leonardi, and Fabio Neri. Packetmode scheduling in inputqueued cell-based switches. IEEE/ACM Transactions on Networking, 10(5):666– 678, October 2002. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Marsan:2003:MTI

[MBG^{+}03] Marco Ajmone Marsan, Andrea Bianco, Paolo Giaccone, Emilio Leonardi, and Fabio Neri. Multicast traffic in input-queued switches: optimal scheduling and maximum throughput. IEEE/ACM Transactions on Networking, 11 (3):465–477, June 2003. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Munir:2017:PSE

[MBI^{+}17] Ali Munir, Ghufran Baig, Syed Mohammad Irteza, Ihsan Ayyub Qazi, Alex X. Liu, and Fahad Rafique Dogar. PASE: Synthesizing existing transport strategies for near-optimal data center transport. IEEE/ ACM Transactions on Networking, 25(1):320–334, February 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Madan:2010:FAR

[MBL10] Ritesh Madan, Stephen P.

Boyd, and Sanjay Lall. Fast algorithms for resource allocation in wireless cellular networks. IEEE/ACM Transactions on Networking, 18(3):973–984, June 2010. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Mendis:2019:DBR

[MBL19] H. V. Kalpanie Mendis, Indika A. M. Balapuwaduge, and Frank Y. Li. Dependabilitybased reliability analysis in URC networks: Availability in the space domain. IEEE/ACM Transactions on Networking, 27(5):1915– 1930, October 2019. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Marsan:1993:TWA

[MBLN93] M. Ajmone Marsan, Andrea Bianco, Emilio Leonardi, and Fabio Neri. Topologies for wavelength-routing alloptical networks. IEEE/ ACM Transactions on Networking, 1(5):534–546, October 1993. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/
                                        journals/ton/1993-1-5/
                                        p534-marsan/.

Menth:2009:SMS

[MBM09] Michael Menth, Andreas Binzenhöfer, and Stefan Mühleck. Source models for speech traffic revisited. IEEE/ACM Transactions on Networking, 17(4): 1042–1051, August 2009. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Mohanti:2021:WMW

[MBN^{+}21] Subhramoy Mohanti, Elif Bozkaya, M. Yousof Naderi, Berk Canberk, Gokhan Secinti, and Kaushik R. Chowdhury. WiFED mobile: WiFi friendly energy delivery with mobile distributed beamforming. IEEE/ACM Transactions on Networking, 29(3):1362– 1375, June 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2021.3061082.

Mukherjee:1996:SPD

[MBRM96] Biswanath Mukherjee, Dhritiman Banerjee, S. Ramamurthy, and Amarnath Mukherjee. Some principles for designing a wide-area WDM optical network. IEEE/ ACM Transactions on Networking, 4(5):684–696, October 1996. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1996-4-5/
p684-mukherjee/.

McAuley:1994:WSC

[McA94] A. J. McAuley. Weighted sum codes for error detection and their comparison with existing codes. IEEE/ACM Transactions on Networking, 2(1):16–22, February 1994. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/

                                        journals/ton/1994-2-1/
                                        p16-mcauley/. See com-

ments [Far95].

Malandrino:2019:OEM

[MCC^{+}19] Francesco Malandrino, Carla Fabiana Chiasserini, Claudio Casetti, Giada Landi, and Marco Capitani. An optimization-enhanced MANO for energy-efficient 5G networks. IEEE/ACM Transactions on Networking, 27(4):1756–1769, August 2019. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Malandrino:2023:EDD

[MCdG23] Francesco Malandrino, Carla Fabiana Chiasserini, and Giuseppe di Giacomo. Efficient distributed DNNs in the mobile-edge-cloud continuum. IEEE/ACM Transactions on Networking, 31(4): 1702–1716, 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2022.3222640.

Malandrino:2019:RSD

[MCES19] Francesco Malandrino, Carla Fabiana Chiasserini, Gil Einziger, and Gabriel Scalosub. Reducing service deployment cost through VNF sharing. IEEE/ACM Transactions on Networking, 27(6):2363–2376, December 2019. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL https://

dl.acm.org/doi/abs/10.
1109/TNET.2019.2945127.

McKeown:1999:ISA

[McK99] Nick McKeown. The iS-LIP scheduling algorithm for input-queued switches. IEEE/ACM Transactions on Networking, 7(2):188– 201, April 1999. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1999-7-2/
p188-mckeown/.

Ma:2010:IEU

[MCL^{+}10] Richard T. B. Ma, Dah Ming Chiu, John C. S. Lui, Vishal Misra, and Dan Rubenstein. Internet economics: the use of Shapley value for ISP settlement. IEEE/ACM Transactions on Networking, 18(3):775– 787, June 2010. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Ma:2011:CSB

[MCL^{+}11] Richard T. B. Ma, Dah Ming Chiu, John C. S. Lui, Vishal Misra, and Dan Rubenstein. On cooperative settlement between content, transit, and eyeball Internet service providers. IEEE/ACM Transactions on Networking, 19(3):802–815, June 2011. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Madan:2007:MOT

[MCLG07] Ritesh Madan, Shuguang Cui, Sanjay Lall, and Andrea J. Goldsmith. Modeling and optimization of transmission schemes in energy-constrained wireless sensor networks. IEEE/ ACM Transactions on Networking, 15(6):1359–1372, December 2007. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

McMillan:1995:DAC

[McM95] David McMillan. Delay analysis of a cellular mobile priority queueing system. IEEE/ACM Transactions on Networking, 3(3):310– 319, June 1995. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/
                                        journals/ton/1995-3-3/
                                        p310-mcmillan/.

Michel:2023:FEQ

[MCM^{+}23] François Michel, Alejandro Cohen, Derya Malak, Quentin De Coninck, Muriel Médard, and Olivier Bonaventure. FlEC: Enhancing QUIC with applicationtailored reliability mechanisms. IEEE/ACM Transactions on Networking, 31 (2):606–619, April 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL

https://dl.acm.org/doi/
10.1109/TNET.2022.3195611.

Malandrino:2023:NSH

[MCMdlO23] Francesco Malandrino, Carla Fabiana Chiasserini, Nuria Molner, and Antonio de la Oliva. Network support for high-performance distributed machine learning. IEEE/ACM Transactions on Networking, 31(1): 264–278, 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2022.3189077.

Manna:2010:IPS

[MCR10] Parbati Kumar Manna, Shigang Chen, and Sanjay Ranka. Inside the permutation-scanning worms: propagation modeling and analysis. IEEE/ACM Transactions on Networking, 18(3):858–870, June 2010. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Manzoni:1999:WMV

[MCS99] Pietro Manzoni, Paolo Cremonesi, and Giuseppe Serazzi. Workload models of VBR video traffic and their use in resource allocation policies. IEEE/ACM Transactions on Networking, 7(3):387– 397, June 1999. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/
                                        journals/ton/1999-7-3/
                                        p387-manzoni/.

Ma:2022:SSI

[MCZ^{+}22] Chaoyi Ma, Shigang Chen, Youlin Zhang, Qingjun Xiao, and Olufemi O. Odegbile. Super spreader identification using geometricmin filter. IEEE/ACM Transactions on Networking, 30(1):299–312, February 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2021.3108033.

Mao:2004:DDA

[MD04] Zuji Mao and Christos Douligeris. A distributed database architecture for global roaming in nextgeneration mobile networks. IEEE/ACM Transactions on Networking, 12(1):146–160, February 2004. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Malone:2007:MDC

[MDL07] David Malone, Ken Duffy, and Doug Leith. Modeling the 802.11 distributed coordination function in nonsaturated heterogeneous conditions. IEEE/ACM Transactions on Networking, 15(1):159–172, February 2007. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Menasche:2013:CAB

[MDL^{+}13] Daniel S. Menasche, Antonio A. De A.Rocha, Bin Li, Don Towsley, and Arun Venkataramani. Content availability and bundling in swarming systems. IEEE/ ACM Transactions on Networking, 21(2):580–593, April 2013. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Menth:2009:RAP

[MDMM09] Michael Menth, Michael Duelli, Ruediger Martin, and Jens Milbrandt. Resilience analysis of packetswitched communication networks. IEEE/ACM Transactions on Networking, 17(6):1950–1963, December 2009. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Modiano:1996:EAP

[ME96] Eytan Modiano and Anthony Ephremides. Efficient algorithms for performing packet broadcasts in a mesh network. IEEE/ ACM Transactions on Networking, 4(4):639–648, August 1996. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1996-4-4/
p639-modiano/.

Medhi:1995:MMC

[Med95] D. Medhi. Multi-hour, multi-traffic class network design for virtual pathbased dynamically reconfigurable wide-area ATM networks. IEEE/ACM Transactions on Networking, 3(6):809–818, December 1995. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1995-3-6/
p809-medhi/.

Mahanti:2003:SDM

[MEVSS03] Anirban Mahanti, Derek L. Eager, Mary K. Vernon, and David J. Sundaram-Stukel. Scalable ondemand media streaming with packet loss recovery. IEEE/ACM Transactions on Networking, 11(2):195– 209, April 2003. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Moazzez-Estanjini:2013:SMN

[MEWP13] Reza Moazzez-Estanjini, Jing Wang, and Ioannis Ch. Paschalidis. Scheduling mobile nodes for cooperative data transport in sensor networks. IEEE/ACM Transactions on Networking, 21(3):974–989, June 2013. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Medard:1999:RTP

[MFB99] Muriel Médard, Steven G.

Finn, and Richard A. Barry. Redundant trees for preplanned recovery in arbitrary vertex-redundant or edge-redundant graphs. IEEE/ACM Transactions on Networking, 7(5):641– 652, October 1999. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL http://

www.acm.org/pubs/citations/
journals/ton/1999-7-5/
p641-medard/.

Marsan:2004:UIP

[MFL^{+}04] Marco Ajmone Marsan, Mirko Franceschinis, Emilio Leonardi, Fabio Neri, and Alessandro Tarello. Underload instabilities in packet networks with flow schedulers. IEEE/ACM Transactions on Networking, 12 (6):1131–1143, December 2004. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Mehic:2020:NAQ

[MFR^{+}20] Miralem Mehic, Peppino Fazio, Stefan Rass, Oliver Maurhart, Momtchil Peev, Andreas Poppe, Jan Rozhon, Marcin Niemiec, and Miroslav Voznak. A novel approach to quality-of-service provisioning in trusted relay quantum key distribution networks. IEEE/ ACM Transactions on Networking, 28(1):168–181, February 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL https://

                                        dl.acm.org/doi/abs/10.
                                        1109/TNET.2019.2956079.

Miyandoab:2020:MIC

[MFT^{+}20] Fardin Derogarian Miyandoab, João Canas Ferreira, Vítor M. Grade Tavares, José Machado da Silva, and Fernando J. Velez. A multifunctional integrated circuit router for body area network wearable systems. IEEE/ ACM Transactions on Networking, 28(5):1981–1994, October 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2020.3004550.

Moldeklev:1995:HLA

[MG95] Kjersti Moldeklev and Per Gunningberg. How a large ATM MTU causes deadlocks in TCP data transfers. IEEE/ACM Transactions on Networking, 3(4):409–422, August 1995. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1995-3-4/
p409-moldeklev/.

MacGregor:1997:DPR

[MG97a] M. H. MacGregor and Wayne D. Grover. Distributed partial-express routing of broad-band transport network demands. IEEE/ACM Transactions on Networking, 5(6):981–988, December 1997. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1997-5-6/
p981-macgregor/.

Medhi:1997:NDP

[MG97b] D. Medhi and Sujit Guptan. Network dimensioning and performance of multiservice, multirate loss networks with dynamic routing. IEEE/ ACM Transactions on Networking, 5(6):944–957, December 1997. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/

                                        journals/ton/1997-5-6/
                                        p944-medhi/.

Mardani:2016:ETA

[MG16] Morteza Mardani and Georgios B. Giannakis. Estimating traffic and anomaly maps via network tomography. IEEE/ACM Transactions on Networking, 24(3):1533–1547, June 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Madhavan:2015:ACN

[MGCK15] Mukundan Madhavan, Harish Ganapathy, Malolan Chetlur, and Shivkumar Kalyanaraman. Adapting cellular networks to whitespaces spectrum. IEEE/ ACM Transactions on Networking, 23(2):383–397, April 2015. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Marsan:2005:UPD

[MGG^{+}05] Marco Ajmone Marsan, Michele Garetto, Paolo Giaccone, Emilio Leonardi, Enrico Schiattarella, and Alessandro Tarello. Using partial differential equations to model TCP mice and elephants in large IP networks. IEEE/ACM Transactions on Networking, 13(6):1289–1301, December 2005. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Magistretti:2012:MDM

[MGK12] Eugenio Magistretti, Omer Gurewitz, and Edward W. Knightly. Measurementdriven modeling of transmission coordination for 802.11 online. IEEE/ ACM Transactions on Networking, 20(5):1635–1648, October 2012. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Magistretti:2014:CAC

[MGK14] Eugenio Magistretti, Omer Gurewitz, and Edward W. Knightly. 802.11ec: collision avoidance without control messages. IEEE/ACM Transactions on Networking, 22(6):1845–1858, December 2014. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Mohan:2020:RSO

[MGK20] Avinash Mohan, Aditya Gopalan, and Anurag Kumar. Reduced-state, optimal scheduling for decentralized medium access control of a class of wireless networks. IEEE/ ACM Transactions on Networking, 28(3):1017–1032, June 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2020.2976923.

Ma:2018:IWF

[MGLH18] Qian Ma, Lin Gao, Ya-Feng Liu, and Jianwei Huang. Incentivizing Wi-Fi network crowdsourcing: a contract theoretic approach. IEEE/ACM Transactions on Networking, 26(3):1035– 1048, June 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Meddeb:2002:IPM

[MGR02] Aref Meddeb, André Girard, and Catherine Rosenberg. The impact of point-to-multipoint traffic concentration on multirate networks design. IEEE/ ACM Transactions on Networking, 10(1):115–124, February 2002. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Meng:2021:PDH

[MGS^{+}21] Zili Meng, Yaning Guo, Yixin Shen, Jing Chen, Chao Zhou, Minhu Wang, Jia Zhang, Mingwei Xu, Chen Sun, and Hongxin Hu. Practically deploying heavyweight adaptive bitrate algorithms with teacher-student learning. IEEE/ACM Transactions on Networking, 29(2):723– 736, April 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2020.3048666.

Majidi:2023:MBU

[MGZ^{+}23] Akbar Majidi, Xiaofeng Gao, Shunjia Zhu, Nazila Jahanbakhsh, Jiaqi Zheng, and Guihai Chen. MiFi: Bounded update to optimize network performance in software-defined data centers. IEEE/ ACM Transactions on Networking, 31(1):322– 335, 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2022.3192167.

Murali:1997:RAL

[MH97] Ramaswamy Murali and Brian L. Hughes. Random access with large propagation delay. IEEE/ ACM Transactions on Networking, 5(6):924–935, December 1997. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1997-5-6/
p924-murali/.

Mao:2002:LPA

[MH02] Guoqiang Mao and Daryoush Habibi. Loss performance analysis for heterogeneous ON-OFF sources with application to connection admission control. IEEE/ACM Transactions on Networking, 10(1):125– 138, February 2002. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Ma:2021:RPD

[MHB^{+}21] Qian Ma, Jianwei Huang, Tamer Başar, Ji Liu, and Xudong Chen. Reputation and pricing dynamics in online markets. IEEE/ ACM Transactions on Networking, 29(4):1745–1759, August 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2021.3071506.

Malik:2020:ZRC

[MHH20] Fehmina Malik, Manjesh K. Hanawal, and Yezekael Hayel. Zero-rating of content and its effect on the quality of service in the Internet. IEEE/ ACM Transactions on Networking, 28(6):2671–2684, December 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2020.3022676.

Ma:2014:ILM

[MHL^{+}14] Liang Ma, Ting He, Kin K. Leung, Ananthram Swami, and Don Towsley. Inferring link metrics from end-to-end path measurements: identifiability and monitor placement. IEEE/ ACM Transactions on Networking, 22(4):1351–1368, August 2014. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Mukhopadhyay:2019:ARM

[MHL19] Arpan Mukhopadhyay, Nidhi Hegde, and Marc Lelarge. Asymptotics of replication and matching in large caching systems. IEEE/ ACM Transactions on Networking, 27(4):1657–1668, August 2019. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Munir:2020:NSC

[MHR^{+}20] Ali Munir, Ting He, Ramya Raghavendra, Franck Le, and Alex X. Liu. Network scheduling and compute resource aware task placement in datacenters. IEEE/ACM Transactions on Networking, 28(6):2435–2448, December 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2020.3013548.

Medina:2012:GRS

[MHRR12] Daniel Medina, Felix Hoffmann, Francesco Rossetto, and Carl-Herbert Rokitansky. A geographic routing strategy for North Atlantic in-flight Internet access via airborne mesh networking. IEEE/ACM Transactions on Networking, 20(4):1231– 1244, August 2012. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Madhow:1995:OWR

[MHS95] Upamanyu Madhow, Michael L. Honig, and Kenneth Steiglitz. Optimization of wireless resources for personal communications mobility tracking. IEEE/ ACM Transactions on Networking, 3(6):698–707, December 1995. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1995-3-6/
p698-madhow/.

Ma:2017:NCL

[MHS^{+}17] Liang Ma, Ting He, Ananthram Swami, Don Towsley, and Kin K. Leung. Network capability in localizing node failures via endto-end path measurements. IEEE/ACM Transactions on Networking, 25(1):434– 450, February 2017. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Min:1995:NCN

[MHSC95] Paul S. Min, Manjunath V.

Hegde, Hossein Saidi, and Alex Chandra. Nonblocking copy networks in multichannel switching. IEEE/ ACM Transactions on Networking, 3(6):857–871, December 1995. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1995-3-6/
p857-min/.

Misra:2010:CRN

[MHXT10] Satyajayant Misra, Seung Don Hong, Guoliang Xue, and Jian Tang. Constrained relay node placement in wireless sensor networks: formulation and approximations. IEEE/ACM Transactions on Networking, 18(2):434–447, April 2010. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Markopoulou:2008:CFO

[MIB^{+}08] Athina Markopoulou, Gianluca Iannaccone, Supratik Bhattacharyya, Chen-Nee Chuah, Yashar Ganjali, and Christophe Diot. Characterization of failures in an operational IP backbone network. IEEE/ACM Transactions on Networking, 16(4):749–762, August 2008. CODEN IEANEP.

ISSN 1063-6692 (print), 1558-2566 (electronic).

Mills:1995:IAS

[Mil95]  David L. Mills.  Im-

proved algorithms for synchronizing computer network clocks. IEEE/ ACM Transactions on Networking, 3(3):245–254, June 1995. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/
                                        journals/ton/1995-3-3/
                                        p245-mills/.

Mills:1998:AHC

[Mil98] David L. Mills. Adaptive hybrid clock discipline algorithm for the network time protocol. IEEE/ ACM Transactions on Networking, 6(5):505–514, October 1998. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/
                                        journals/ton/1998-6-5/
                                        p505-mills/.

Mitzenmacher:2002:CBF

[Mit02]  Michael  Mitzenmacher.

Compressed bloom filters. IEEE/ACM Transactions on Networking, 10(5):604– 612, October 2002. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Ma:2001:MHN

[MJ01] Sheng Ma and Chuanyi Ji. Modeling heterogeneous network traffic in wavelet domain. IEEE/ ACM Transactions on Networking, 9(5):634–649, October 2001. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Metnani:2013:SFC

[MJ13] Ammar Metnani and Brigitte Jaumard. Stability of FIPP p-cycles under dynamic traffic in WDM networks: dynamic traffic,failure-independent path-protecting (FIPP) pcycles,path protection,shared bandwidth protection,stability of protection structures. IEEE/ACM Transactions on Networking, 21(2):413– 425, April 2013. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Muthusamy:2014:IFC

[MJ14] Vinod Muthusamy and Hans-Arno Jacobsen. Infrastructurefree content-based publish/subscribe. IEEE/ ACM Transactions on Networking, 22(5):1516–1530, October 2014. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Mokhtarian:2015:MDM

[MJ15] Kianoosh Mokhtarian and Hans-Arno Jacobsen. Minimumdelay multicast algorithms for mesh overlays. IEEE/ ACM Transactions on Networking, 23(3):973–986, June 2015. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Mokhtarian:2017:FCA

[MJ17] Kianoosh Mokhtarian and Hans-Arno Jacobsen. Flexible caching algorithms for video content distribution networks. IEEE/ACM Transactions on Networking, 25(2):1062–1075, April 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Murakami:1996:VPR

[MK96] Kazutaka Murakami and Hyong S. Kim. Virtual path routing for survivable ATM networks. IEEE/ACM Transactions on Networking, 4(1):22–39, February 1996. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/
                                        journals/ton/1996-4-1/
                                        p22-murakami/.

Murakami:1998:OCF

[MK98] Kazutaka Murakami and Hyong S. Kim. Optimal capacity and flow assignment for self-healing ATM networks based on line and end-to-end restoration. IEEE/ACM Transactions on Networking, 6(2):207– 221, April 1998. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1998-6-2/
p207-murakami/.

Mondal:2010:UME

[MK10] Amit Mondal and Aleksandar Kuzmanovic. Upgrading mice to elephants: effects and end-point solutions. IEEE/ACM Transactions on Networking, 18 (2):367–378, April 2010. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Maatouk:2020:AII

[MKAE20] Ali Maatouk, Saad Kriouile, Mohamad Assaad, and Anthony Ephremides. The age of incorrect information: a new performance metric for status updates. IEEE/ACM Transactions on Networking, 28(5):2215–2228, October 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2020.3005549.

Moon:2017:CNA

[MKG^{+}17] YoungGyoun Moon, Donghwi Kim, Younghwan Go, Yeongjin Kim, Yung Yi, Song Chong, and Kyoung-Soo Park. Cedos: a network architecture and programming abstraction for delay-tolerant mobile apps. IEEE/ACM Transactions on Networking, 25(2):646– 661, April 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Murakami:2024:EEP

[MKOY24] Masaki Murakami, Takashi Kurimoto, Satoru Okamoto, and Naoaki Yamanaka. Experimental evaluation on priority-aware guaranteed resource allocation for resource pool based reconfigurable hardware. IEEE/ACM Transactions on Networking, 32(1):298– 307, 2024. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2023.3288021.

Mao:2016:OOS

[MKS16] Zhoujia Mao, Can Emre Koksal, and Ness B. Shroff. Optimal online scheduling with arbitrary hard deadlines in multihop communication networks. IEEE/ ACM Transactions on Networking, 24(1):177–189, February 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Moharir:2017:SDN

[MKS17] Sharayu Moharir, Subhashini Krishnasamy, and Sanjay Shakkottai. Scheduling in densified networks: Algorithms and performance. IEEE/ACM Transactions on Networking, 25(1):164–178, February 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Mishra:1996:HRC

[MKT96] Partho Pratim Mishra, Hemant Kanakia, and Satish K. Tripathi. On hop-by-hop rate-based congestion control. IEEE/ ACM Transactions on Networking, 4(2):224–239, April 1996. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1996-4-2/
p224-mishra/.

Moon:2006:RAS

[ML06] Ji-Cheol Moon and Byeong Gi Lee. Rate-adaptive snoop: a TCP enhancement scheme over rate-controlled lossy links. IEEE/ACM Transactions on Networking, 14 (3):603–615, 2006. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Menth:2012:PPB

[ML12] Michael Menth and Frank Lehrieder. Performance of PCN-based admission control under challenging conditions. IEEE/ACM Transactions on Networking, 20 (2):422–435, April 2012. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Maccari:2018:IRC

[ML18] Leonardo Maccari and Renato Lo Cigno. Improving routing convergence with centrality: Theory and implementation of pop-routing. IEEE/ ACM Transactions on Networking, 26(5):2216–2229, October 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Mohammadpour:2022:ADT

[ML22a] Ehsan Mohammadpour and Jean-Yves Le Boudec. Analysis of dampers in time-sensitive networks with non-ideal clocks. IEEE/ACM Transactions on Networking, 30(4):1780– 1794, 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2022.3152178.

Mohammadpour:2022:PRT

[ML22b] Ehsan Mohammadpour and Jean-Yves Le Boudec. On packet reordering in time-sensitive networks. IEEE/ACM Transactions on Networking, 30(3):1045– 1057, June 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2021.3129590.

Meskar:2023:FMR

[ML23] Erfan Meskar and Ben Liang. Fair multi-resource allocation in heterogeneous servers with an external resource type. IEEE/ ACM Transactions on Networking, 31(3):1244–1262, June 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2022.3213426.

Mai:2021:OCI

[MLB21] Van Sy Mai, Richard J.

La, and Abdella Battou. Optimal cybersecurity investments in large networks using SIS model: Algorithm design. IEEE/ACM Transactions on Networking, 29(6):2453–2466, December 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.

acm.org/doi/10.1109/TNET.
2021.3091856.

Movsichoff:2007:EEO

[MLC07] Bernardo A. Movsichoff, Constantino M. Lagoa, and Hao Che. End-to-end optimal algorithms for integrated QoS, traffic engineering, and failure recovery. IEEE/ACM Transactions on Networking, 15 (4):813–823, August 2007. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Matousek:2022:CNB

[MLJ^{+}22] Jiří Matoušek, Adam Lučanský, David Janeček, Jozef Sabo, Jan Kořenek, and Gianni Antichi. ClassBenchng: Benchmarking packet classification algorithms in the OpenFlow era. IEEE/ACM Transactions on Networking, 30(5):1912– 1925, 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2022.3155708.

Ma:2006:ISD

[MLLY06] Richard T. B. Ma, Sam C. M. Lee, John C. S. Lui, and David K. Y. Yau. Incentive and service differentiation in P2P networks: a game theoretic approach. IEEE/ACM Transactions on Networking, 14 (5):978–991, October 2006.

CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Ma:2015:EIE

[MLM15] Richard T. B. Ma, John C. S. Lui, and Vishal Misra. Evolution of the Internet economic ecosystem. IEEE/ACM Transactions on Networking, 23 (1):85–98, February 2015. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Miao:2012:CAD

[MLS12] Guowang Miao, Ye Li, and Ananthram Swami. Channel-aware distributed medium access control. IEEE/ACM Transactions on Networking, 20(4):1290– 1303, August 2012. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Ma:2023:PAO

[MLS^{+}23] Junchao Ma, Lingjia Liu, Bodong Shang, Shashank Jere, and Pingzhi Fan. Performance analysis and optimization for layer-based scalable video caching in 6G networks. IEEE/ ACM Transactions on Networking, 31(4):1494– 1506, 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.

                                        acm.org/doi/10.1109/TNET.
                                        2022.3222931.

Meiners:2011:TTA

[MLT11] Chad R. Meiners, Alex X. Liu, and Eric Torng. Topological transformation approaches to TCAMbased packet classification. IEEE/ACM Transactions on Networking, 19(1):237– 250, February 2011. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Meiners:2012:BWN

[MLT12] Chad R. Meiners, Alex X. Liu, and Eric Torng. Bit weaving: a non-prefix approach to compressing packet classifiers in TCAMs. IEEE/ACM Transactions on Networking, 20(2):488–500, April 2012. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Ma:2018:CUM

[MLX18] Yu Ma, Weifa Liang, and Wenzheng Xu. Charging utility maximization in wireless rechargeable sensor networks by charging multiple sensors simultaneously. IEEE/ACM Transactions on Networking, 26(4):1591–1604, August 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Mitra:1994:ECU

[MM94] Debasis Mitra and John A.

Morrison. Erlang capacity and uniform approximations for shared unbuffered resources. IEEE/ ACM Transactions on Networking, 2(6):558–570, December 1994. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1994-2-6/
p558-mitra/.

Ma:2013:PON

[MM13] Richard T. B. Ma and Vishal Misra. The public option: a nonregulatory alternative to network neutrality. IEEE/ ACM Transactions on Networking, 21(6):1866–1879, December 2013. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Mellia:2005:TSF

[MMC05] Marco Mellia, Michela Meo, and Claudio Casetti. TCP smart framing: a segmentation algorithm to reduce TCP latency. IEEE/ ACM Transactions on Networking, 13(2):316–329, April 2005. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Moorthy:2022:ERL

[MMG22] Sabarish Krishna Moorthy, Maxwell Mcmanus, and Zhangyu Guan. ESN reinforcement learning for spectrum and flight control in THz-Enabled drone networks. IEEE/ACM Transactions on Networking, 30 (2):782–795, April 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL

https://dl.acm.org/doi/
10.1109/TNET.2021.3128836.

Mallada:2015:SNC

[MMH^{+}15] Enrique Mallada, Xiaoqiao Meng, Michel Hack, Li Zhang, and Ao Tang. Skewless network clock synchronization without discontinuity: convergence and performance. IEEE/ ACM Transactions on Networking, 23(5):1619–1633, October 2015. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Mahdian:2020:KCN

[MMIY20] Milad Mahdian, Armin Moharrer, Stratis Ioannidis, and Edmund Yeh. Kelly cache networks. IEEE/ACM Transactions on Networking, 28(3):1130– 1143, June 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.

acm.org/doi/10.1109/TNET.
2020.2982863.

Munari:2017:SGA

[MMP17] Andrea Munari, Petri Mahonen, and Marina Petrova. A stochastic geometry approach to asynchronous Aloha full-duplex networks. IEEE/ACM Transactions on Networking, 25(6):3695–3708, December 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Mitra:1996:AND

[MMR96] Debasis Mitra, John A.

Morrison, and K. G. Ramakrishnan. ATM network design and optimization: a multirate loss network framework. IEEE/ ACM Transactions on Networking, 4(4):531–543, August 1996. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1996-4-4/
p531-mitra/.

Ma:2009:AGS

[MMR09] Richard T. B. Ma, Vishal Misra, and Dan Rubenstein. An analysis of generalized slotted-Aloha protocols. IEEE/ACM Transactions on Networking, 17(3): 936–949, June 2009. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Mohan:2001:EAR

[MMS01] G. Mohan, C. Siva Ram Murthy, and Arun K. Somani. Efficient algorithms for routing dependable connections in WDM optical networks. IEEE/ACM Transactions on Networking, 9(5):553–566, October 2001. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Markakis:2014:MWS

[MMT14] Mihalis G. Markakis, Eytan Modiano, and John N. Tsitsiklis. Max-weight scheduling in queueing networks with heavy-tailed traffic. IEEE/ACM Transactions on Networking, 22(1):257–270, February 2014. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Markakis:2016:DSB

[MMT16] Mihalis G. Markakis, Eytan Modiano, and John N. Tsitsiklis. Delay stability of back-pressure policies in the presence of heavy-tailed traffic. IEEE/ ACM Transactions on Networking, 24(4):2046–2059, August 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Mneimneh:2008:MFI

[Mne08] Saad Mneimneh. Matching from the first iteration: an iterative switching algorithm for an input queued switch. IEEE/ ACM Transactions on Networking, 16(1):206–217, February 2008. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Martin:2003:DBC

[MNR03] Jim Martin, Arne Nilsson, and Injong Rhee. Delaybased congestion avoidance for TCP. IEEE/ACM Transactions on Networking, 11(3):356–369, June 2003. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Mason:2023:UDR

[MNZ23] Federico Mason, Gianfranco Nencioni, and Andrea Zanella. Using distributed reinforcement learning for resource orchestration in a network slicing scenario. IEEE/ ACM Transactions on Networking, 31(1):88–102, 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL

https://dl.acm.org/doi/
10.1109/TNET.2022.3187310.

Modiano:1999:RAS

[Mod99] Eytan Modiano. Random algorithms for scheduling multicast traffic in WDM broadcast-and-select networks. IEEE/ACM Transactions on Networking, 7 (3):425–434, June 1999. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL

http://www.acm.org/pubs/
citations/journals/ton/
1999-7-3/p425-modiano/.

Manfredi:2013:DCL

[MOR13] Sabato Manfredi, Francesco Oliviero, and Simon Pietro Romano. A distributed control law for load balancing in content delivery networks. IEEE/ACM Transactions on Networking, 21 (1):55–68, February 2013. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Mayer:2000:LCD

[MOY00] Alain Mayer, Yoram Ofek, and Moti Yung. Local and congestion-driven fairness algorithm in arbitrary topology networks. IEEE/ACM Transactions on Networking, 8(3):362– 372, 2000. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/2000-8-3/
p362-mayer/.

Maxemchuk:2005:QMT

[MOZ05] Nicholas F. Maxemchuk, Iradj Ouveysi, and Moshe Zukerman. A quantitative measure for telecommunications networks topology design. IEEE/ACM Transactions on Networking, 13 (4):731–742, August 2005. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Miller:1993:GMC

[MP93] Raymond E. Miller and Sanjoy Paul. On the generation of minimallength conformance tests for communication protocols. IEEE/ACM Transactions on Networking, 1(1): 116–129, February 1993. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL

http://www.acm.org/pubs/
citations/journals/ton/
1993-1-1/p116-miller/.

Miller:1994:SAP

[MP94] Raymond E. Miller and Sanjoy Paul. Structural analysis of protocol specifications and generation of maximal fault coverage conformance test sequences. IEEE/ACM Transactions on Networking, 2(5):457–470, October 1994. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/

                                        journals/ton/1994-2-5/
                                        p457-miller/.

Micciancio:2008:OCC

[MP08] Daniele Micciancio and Saurabh Panjwani. Optimal communication complexity of generic multicast key distribution. IEEE/ ACM Transactions on Networking, 16(4):803–813, August 2008. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Martignon:2015:ETB

[MPF^{+}15] Fabio Martignon, Stefano Paris, Ilario Filippini, Lin Chen, and Antonio Capone. Efficient and truthful bandwidth allocation in wireless mesh community networks. IEEE/ ACM Transactions on Networking, 23(1):161–174, February 2015. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Markatos:2002:WCS

[MPFK02] Evangelos P. Markatos, Dionisios N. Pnevmatikatos, Michail D. Flouris, and Manolis G. H. Katevenis. Web-conscious storage management for Web proxies. IEEE/ACM Transactions on Networking, 10(6): 735–748, December 2002. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Martinez:2009:DFN

[MPL09] Christopher J. Martinez, Devang K. Pandya, and Wei-Ming Lin. On designing fast nonuniformly distributed IP address lookup hashing algorithms. IEEE/ ACM Transactions on Networking, 17(6):1916–1925, December 2009. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Martin-Perez:2022:KGN

[MPMC^{+}22] Jorge Martín-Pérez, Francesco Malandrino, Carla Fabiana Chiasserini, Milan Groshev, and Carlos J. Bernardos. KPI guarantees in network slicing. IEEE/ACM Transactions on Networking, 30(2):655– 668, April 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2021.3120318.

Meiners:2014:FRE

[MPN^{+}14] Chad R. Meiners, Jignesh Patel, Eric Norige, Alex X. Liu, and Eric Torng. Fast regular expression matching using small TCAM. IEEE/ACM Transactions on Networking, 22(1):94– 109, February 2014. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Mansour:2001:JCQ

[MPS01] Yishay Mansour and Boaz Patt-Shamir. Jitter control in QoS networks. IEEE/ACM Transactions on Networking, 9(4):492– 502, August 2001. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Marbach:2005:CWA

[MQ05] Peter Marbach and Ying Qiu. Cooperation in wireless ad hoc networks: a market-based approach. IEEE/ACM Transactions on Networking, 13(6):1325– 1338, December 2005. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Ma:2022:IHI

[MQL^{+}22] Xiaobo Ma, Jian Qu, Jianfeng Li, John C. S. Lui, Zhenhua Li, Wenmao Liu, and Xiaohong Guan. Inferring hidden IoT devices and user interactions via spatial-temporal traffic fingerprinting. IEEE/ ACM Transactions on Networking, 30(1):394–408, February 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2021.3112480.

Moghe:1996:ECP

[MR96] Pratyush Moghé and Izhak Rubin. Enhanced call: a paradigm for applications with dynamic clientmembership and clientlevel binding in ATM networks. IEEE/ACM Transactions on Networking, 4(4):615–628, August 1996. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1996-4-4/
p615-moghe/.

Mark:1998:RED

[MR98] Brian L. Mark and Gopalakrishnan Ramamurthy. Realtime estimation and dynamic renegotiation of UPC parameters for arbitrary traffic sources in ATM networks. IEEE/ ACM Transactions on Networking, 6(6):811–827, December 1998. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1998-6-6/
p811-mark/.

Massoulie:2002:BSO

[MR02] Laurent Massoulié and James Roberts. Bandwidth sharing: objectives and algorithms. IEEE/ACM Transactions on Networking, 10(3):320–328, June 2002. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Magharei:2009:PPP

[MR09] Nazanin Magharei and Reza Rejaie. PRIME: peerto-peer receiver-driven meshbased streaming. IEEE/ ACM Transactions on Networking, 17(4):1052–1065, August 2009. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Misra:2008:ITB

[MRD08] Archan Misra, Abhishek Roy, and Sajal K. Das. Information-theory based optimal location management schemes for integrated multi-system wireless networks. IEEE/ACM Transactions on Networking, 16(3):525–538, June 2008. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Mohsenian-Rad:2014:RIN

[MRHWS14] Hamed Mohsenian-Rad, Jianwei Huang, Vincent W. S. Wong, and Robert Schober. Repeated intersession network coding games: efficiency and minmax bargaining solution. IEEE/ACM Transactions on Networking, 22(4):1121– 1135, August 2014. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Mohammadkhan:2020:CIA

[MRJ20] Ali Mohammadkhan, K. K.

Ramakrishnan, and Vivek A. Jain. CleanG — improving the architecture and protocols for future cellular networks with NFV. IEEE/ ACM Transactions on Networking, 28(6):2559–2572, December 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2020.3015946.

Manimaran:1999:NDR

[MRM99] G. Manimaran, Hariharan Shankar Rahul, and C. Siva Ram Murthy. A new distributed route selection approach for channel establishment in realtime networks. IEEE/ ACM Transactions on Networking, 7(5):698–709, October 1999. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1999-7-5/
p698-manimaran/.

Mizrahi:2017:TUT

[MRM17] Tal Mizrahi, Ori Rottenstreich, and Yoram Moses. TimeFlip: Using timestamp-based TCAM ranges to accurately schedule network updates. IEEE/ ACM Transactions on Networking, 25(2):849–863, April 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Mohaisen:2017:LOD

[MRMR17] Aziz Mohaisen, Kui Ren, Aziz Mohaisen, and Kui Ren. Leakage of .onion at the DNS root: Measurements, causes, and countermeasures. IEEE/ ACM Transactions on Networking, 25(5):3059–3072, October 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Magharei:2014:IFL

[MRR^{+}14] Nazanin Magharei, Reza Rejaie, Ivica Rimac, Volker Hilt, and Markus Hofmann. ISP-friendly live P2P streaming. IEEE/ ACM Transactions on Networking, 22(1):244–256, February 2014. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Merchant:1995:ACS

[MS95] Arif Merchant and Bhaskar Sengupta. Assignment of cells to switches in PCS networks. IEEE/ ACM Transactions on Networking, 3(5):521–526, October 1995. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1995-3-5/
p521-merchant/.

Mneimneh:2003:ATI

[MS03] Saad Mneimneh and Kai-Yeung Siu. On achieving throughput in an inputqueued switch. IEEE/ACM Transactions on Networking, 11(5):858–867, October 2003. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Menache:2008:CME

[MS08] Ishai Menache and Nahum Shimkin. Capacity management and equilibrium for proportional QoS. IEEE/ ACM Transactions on Networking, 16(5):1025–1037, October 2008. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Maguluri:2014:SJU

[MS14] Siva Theja Maguluri and R. Srikant. Scheduling jobs with unknown duration in clouds. IEEE/ ACM Transactions on Networking, 22(6):1938–1951, December 2014. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Moharir:2015:MWV

[MS15] Sharayu Moharir and Sanjay Shakkottai. Max weight versus back pressure: routing and scheduling in multichannel relay networks. IEEE/ACM Transactions on Networking, 23(5):1584– 1598, October 2015. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Mehmeti:2017:PMA

[MS17] Fidan Mehmeti and Thrasyvoulos Spyropoulos. Performance modeling, analysis, and optimization of delayed mobile data offloading for mobile users. IEEE/ACM Transactions on Networking, 25(1):550– 564, February 2017. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Margolies:2016:EMP

[MSA^{+}16] Robert Margolies, Ashwin Sridharan, Vaneet Aggarwal, Rittwik Jana, N. K. Shankaranarayanan, Vinay A. Vaishampayan, and Gil Zussman. Exploiting mobility in proportional fair cellular scheduling: measurements and algorithms. IEEE/ACM Transactions on Networking, 24(1):355–367, February 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Matragi:1997:JCA

[MSB97] Wassim Matragi, Khosrow Sohraby, and Chatschik Bisdikian. Jitter calculus in ATM networks: multiple nodes. IEEE/ ACM Transactions on Networking, 5(1):122–133, February 1997. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1997-5-1/
p122-matragi/.

Mirza:2010:MLA

[MSBZ10] Mariyam Mirza, Joel Sommers, Paul Barford, and Xiaojin Zhu. A machine learning approach to TCP throughput prediction. IEEE/ACM Transactions on Networking, 18(4): 1026–1039, August 2010. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Min:1995:NAB

[MSH95] P. S. Min, H. Saidi, and M. V. Hegde. A nonblocking architecture for broadband multichannel switching. IEEE/ ACM Transactions on Networking, 3(2):181–198, April 1995. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/

                                        journals/ton/1995-3-2/
                                        p181-min/.

Mizrahi:2016:TCN

[MSM16] Tal Mizrahi, Efi Saat, and Yoram Moses. Timed consistent network updates in software-defined networks. IEEE/ACM Transactions on Networking, 24(6):3412– 3425, December 2016. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Mehyar:2007:ADA

[MSP^{+}07] Mortada Mehyar, Demetri Spanos, John Pongsajapan, Steven H. Low, and Richard M. Murray. Asynchronous distributed averaging on communication networks. IEEE/ACM Transactions on Networking, 15(3):512–520, June 2007. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). See correction [Kri14].

Mehrnoush:2018:AMW

[MSRG18] Morteza Mehrnoush, Vanlin Sathya, Sumit Roy, and Monisha Ghosh. Analytical modeling of Wi-Fi and LTE-LAA coexistence: Throughput and impact of energy detection threshold. IEEE/ACM Transactions on Networking, 26(4): 1990–2003, August 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Mneimneh:2002:SUP

[MSS02] Saad Mneimneh, Vishal Sharma, and Kai-Yeung Siu. Switching using parallel input-output queued switches with no speedup. IEEE/ACM Transactions on Networking, 10(5):653– 665, October 2002. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Manweiler:2012:OMT

[MSS^{+}12] Justin Manweiler, Naveen Santhapuri, Souvik Sen, Romit Roy Choudhury, Srihari Nelakuditi, and Kamesh Munagala. Order matters: transmission reordering in wireless networks. IEEE/ACM Transactions on Networking, 20 (2):353–366, April 2012. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Moharir:2016:OLB

[MSS16] Sharayu Moharir, Sujay Sanghavi, and Sanjay Shakkottai. Online load balancing under graph constraints. IEEE/ACM Transactions on Networking, 24(3):1690–1703, June 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Mertzios:2012:PRO

[MSSZ12] George B. Mertzios, Ignasi Sau, Mordechai Shalom, and Shmuel Zaks. Placing regenerators in optical networks to satisfy multiple sets of requests. IEEE/ ACM Transactions on Networking, 20(6):1870–1879, December 2012. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Monti:2017:HPN

[MSTL17] Massimo Monti, Manolis Sifalakis, Christian F. Tschudin, and Marco Luise. On hardware programmable network dynamics with a chemistryinspired abstraction. IEEE/ ACM Transactions on Networking, 25(4):2054–2067, August 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Mao:2006:JDJ

[MSWL06] Yinian Mao, Yan Sun, Min Wu, and K. J. Ray Liu. JET: dynamic joinexit-tree amortization and scheduling for contributory key management. IEEE/ ACM Transactions on Networking, 14(5):1128–1140, October 2006. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Maille:2006:PIM

[MT06] Patrick Maillé and Bruno Tuffin. Pricing the Internet with multibid auctions. IEEE/ACM Transactions on Networking, 14(5):992– 1004, October 2006. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Markopoulou:2003:AQV

[MTK03] Athina P. Markopoulou, Fouad A. Tobagi, and Mansour J. Karam. Assessing the quality of voice communications over Internet backbones. IEEE/ACM Transactions on Networking, 11(5):747–760, October 2003. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Massoulie:2008:CRS

[MV08] Laurent Massoulié and Milan Vojnovic. Coupon replication systems. IEEE/ ACM Transactions on Networking, 16(3):603–616, June 2008. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Mandjes:2009:RDT

[MV09] Michel Mandjes and Remco Van De Meent. Resource dimensioning through buffer sampling. IEEE/ACM Transactions on Networking, 17(5):1631–1644, October 2009. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Minnebo:2014:FCP

[MV14] Wouter Minnebo and Benny Van Houdt. A fair comparison of pull and push strategies in large distributed networks. IEEE/ACM Transactions on Networking, 22(3):996–1006, June 2014. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Mishra:2016:AFP

[MV16] Abhishek Mishra and Parv Venkitasubramaniam. Anonymity and fairness in packet scheduling: a quantitative tradeoff. IEEE/ACM Transactions on Networking, 24(2):688–702, April 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Mendelson:2021:ASC

[MVB^{+}21] Gal Mendelson, Shay Vargaftik, Katherine Barabash, Dean H. Lorenz, Isaac Keslassy, and Ariel Orda. AnchorHash: a scalable consistent hash. IEEE/ ACM Transactions on Networking, 29(2):517–528, April 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.

acm.org/doi/10.1109/TNET.
2020.3039547.

Malboubi:2016:DNI

[MVCS16] Mehdi Malboubi, Cuong Vu, Chen-Nee Chuah, and Puneet Sharma. Decentralizing network inference problems with multipledescription fusion estimation MDFE. IEEE/ ACM Transactions on Networking, 24(4):2539–2552, August 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Ma:2009:OWM

[MVRZ09] Hui Ma, Rajiv Vijayakumar, Sumit Roy, and Jing Zhu. Optimizing 802.11 wireless mesh networks based on physical carrier sensing. IEEE/ ACM Transactions on Networking, 17(5):1550–1563, October 2009. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Mishra:1998:EFC

[MW98] Shivakant Mishra and Lei Wu. An evaluation of flow control in group communication. IEEE/ACM Transactions on Networking, 6(5):571–587, October 1998. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.

                                        acm.org/pubs/citations/
                                        journals/ton/1998-6-5/
                                        p571-mishra/.

Mo:2000:FEE

[MW00] Jeonghoon Mo and Jean Walrand. Fair end-to-end window-based congestion control. IEEE/ACM Transactions on Networking, 8(5):556–567, 2000. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL

                                        http://www.acm.org/pubs/
                                        citations/journals/ton/
                                        2000-8-5/p556-mo/.

Mitra:2005:STE

[MW05] Debasis Mitra and Qiong Wang. Stochastic traffic engineering for demand uncertainty and risk-aware network revenue management. IEEE/ACM Transactions on Networking, 13 (2):221–233, April 2005. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Musacchio:2006:WAP

[MW06] John Musacchio and Jean Walrand. WiFi access point pricing as a dynamic game. IEEE/ACM Transactions on Networking, 14 (2):289–301, April 2006. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Meng:2016:CBN

[MWC16] Tong Meng, Fan Wu, and Guihai Chen. Code-based neighbor discovery protocols in mobile wireless networks. IEEE/ACM Transactions on Networking, 24 (2):806–819, April 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Mao:2010:SSS

[MWQ^{+}10] Yun Mao, Feng Wang, Lili Qiu, Simon Lam, and Jonathan Smith. S4: small state and small stretch compact routing protocol for large static wireless networks. IEEE/ACM Transactions on Networking, 18 (3):761–774, June 2010. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Meng:2021:IAD

[MWW^{+}21] Xuying Meng, Yequan Wang, Suhang Wang, Di Yao, and Yujun Zhang. Interactive anomaly detection in dynamic communication networks. IEEE/ ACM Transactions on Networking, 29(6):2602–2615, December 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2021.3097137.

Motamedi:2019:MIT

[MYC^{+}19] Reza Motamedi, Bahador Yeganeh, Balakrishnan Chandrasekaran, Reza Rejaie, Bruce M. Maggs, and Walter Willinger. On mapping the interconnections in today’s Internet. IEEE/ ACM Transactions on Networking, 27(5):2056–2070, October 2019. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Ma:2021:SCG

[MYH21] Qian Ma, Edmund Yeh, and Jianwei Huang. Selfish caching games on directed graphs. IEEE/ ACM Transactions on Networking, 29(2):709–722, April 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2020.3047940.

Mahdian:2017:TDS

[MYMY17] Milad Mahdian, Edmund M. Yeh, Milad Mahdian, and Edmund M. Yeh. Throughput and delay scaling of content-centric ad hoc and heterogeneous wireless networks. IEEE/ACM Transactions on Networking, 25(5):3030–3043, October 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Ma:2013:PVP

[MYYR13] Chris Y. T. Ma, David K. Y. Yau, Nung Kwan Yip, and Nageswara S. V. Rao. Privacy vulnerability of published anonymous mobility traces. IEEE/ ACM Transactions on Networking, 21(3):720–733, June 2013. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Marasevic:2017:RAR

[MZK^{+}17] Jelena Marasevic, Jin Zhou, Harish Krishnaswamy, Yuan Zhong, and Gil Zussman. Resource allocation and rate gains in practical full-duplex systems. IEEE/ACM Transactions on Networking, 25(1):292– 305, February 2017. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Miao:2023:CHP

[MZZ^{+}23] Ruijie Miao, Yinda Zhang, Zihao Zheng, Ruixin Wang, Ruwen Zhang, Tong Yang, Zaoxing Liu, and Junchen Jiang. CocoSketch: Highperformance sketch-based measurement over arbitrary partial key query. IEEE/ACM Transactions on Networking, 31(6):2653– 2668, 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.

                                        acm.org/doi/10.1109/TNET.
                                        2023.3257226.

Nair:2016:CFF

[NAA^{+}16] Jayakrishnan Nair, Martin Andreasson, Lachlan L. H. Andrew, Steven H. Low, and John C. Doyle. On channel failures, file fragmentation policies, and heavy-tailed completion times. IEEE/ACM Transactions on Networking, 24(1):529–541, February 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Nguyen:2012:TCM

[NABZ12] Thuy T. T. Nguyen, Grenville Armitage, Philip Branch, and Sebastian Zander. Timely and continuous machine-learning-based classification for interactive IP traffic. IEEE/ ACM Transactions on Networking, 20(6):1880–1894, December 2012. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Naik:1997:ECU

[Nai97] Kshirasagar Naik. Efficient computation of unique input/output sequences in finite-state machines. IEEE/ACM Transactions on Networking, 5(4):585– 599, August 1997. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1997-5-4/
p585-naik/.

Nonnenmacher:1999:SFL

[NB99] Jörg Nonnenmacher and Ernst W. Biersack. Scalable feedback for large groups. IEEE/ACM Transactions on Networking, 7 (3):375–386, June 1999. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL

http://www.acm.org/pubs/
citations/journals/ton/
1999-7-3/p375-nonnenmacher/

.

Nahum:2002:PIW

[NBK02] Erich Nahum, Tsipora Barzilai, and Dilip D. Kandlur. Performance issues in WWW servers. IEEE/ACM Transactions on Networking, 10(1):2–11, February 2002. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Nonnenmacher:1998:PLR

[NBT98] Jörg Nonnenmacher, Ernst W.

Biersack, and Don Towsley. Parity-based loss recovery for reliable multicast transmission. IEEE/ACM Transactions on Networking, 6(4):349–361, August 1998. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/
                                        journals/ton/1998-6-4/
                                        p349-nonnenmacher/.

Nucci:2007:ILW

[NBTD07] Antonio Nucci, Supratik Bhattacharyya, Nina Taft, and Christophe Diot. IGP link weight assignment for operational Tier-1 backbones. IEEE/ACM Transactions on Networking, 15 (4):789–802, August 2007. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Nadendla:2017:OSA

[NBV17] V. Sriram Siddhardh Nadendla, Swastik K. Brahma, and Pramod K. Varshney. Optimal spectrum auction design with 2-D truthful revelations under uncertain spectrum availability. IEEE/ACM Transactions on Networking, 25(1):420– 433, February 2017. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Nakibly:2015:ODP

[NCK15] Gabi Nakibly, Reuven Cohen, and Liran Katzir. Optimizing data plane resources for multipath flows. IEEE/ACM Transactions on Networking, 23(1):138– 147, February 2015. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Neglia:2018:CPL

[NCM18] Giovanni Neglia, Damiano Carra, and Pietro Michiardi. Cache policies for linear utility maximization. IEEE/ACM Transactions on Networking, 26(1): 302–313, February 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Nguyen:2014:USM

[NCT14] Van Minh Nguyen, Chung Shue Chen, and Laurent Thomas. A unified stochastic model of handover measurement in mobile networks. IEEE/ ACM Transactions on Networking, 22(5):1559–1576, October 2014. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Nace:2006:COM

[NDGL06] Dritan Nace, Nhat-Linh Doan, Eric Gourdin, and Bernard Liau. Computing optimal max-min fair resource allocation for elastic flows. IEEE/ACM Transactions on Networking, 14(6):1272–1281, December 2006. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Nevat:2018:ADA

[NDN^{+}18] Ido Nevat, Dinil Mon Divakaran, Sai Ganesh Nagarajan, Pengfei Zhang, Le Su, Li Ling Ko, and Vrizlynn L. L. Thing. Anomaly detection and attribution in networks with temporally correlated traffic. IEEE/ACM Transactions on Networking, 26(1): 131–144, February 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Naumann:2019:PBD

[NDS19] Roman Naumann, Stefan Dietzel, and Bjorn Scheuermann. Push the barrier: Discrete event protocol emulation. IEEE/ACM Transactions on Networking, 27(2):635–648, April 2019. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Neely:2008:OOD

[Nee08] Michael J. Neely. Order optimal delay for opportunistic scheduling in multi-user wireless uplinks and downlinks. IEEE/ ACM Transactions on Networking, 16(5):1188–1199, October 2008. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Neely:2009:DAM

[Nee09] Michael J. Neely. Delay analysis for maximal scheduling with flow control in wireless networks with bursty traffic. IEEE/ ACM Transactions on Networking, 17(4):1146–1159, August 2009. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Neely:2013:DBN

[Nee13] Michael J. Neely. Delaybased network utility maximization. IEEE/ACM Transactions on Networking, 21(1):41–54, February 2013. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Neely:2016:DSO

[Nee16a] Michael J. Neely. Distributed stochastic optimization via correlated scheduling. IEEE/ACM Transactions on Networking, 24(2):759–772, April 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Neely:2016:EAW

[Nee16b] Michael J. Neely. Energyaware wireless scheduling with near-optimal backlog and convergence time tradeoffs. IEEE/ACM Transactions on Networking, 24(4):2223–2236, August 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Neely:2019:CAU

[Nee19] Michael J. Neely. Convergence and adaptation for utility optimal opportunistic scheduling. IEEE/ACM Transactions on Networking, 27(3):904–917, June 2019. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Neely:2022:CRC

[Nee22]  Michael J. Neely.  A

converse result on convergence time for opportunistic wireless scheduling. IEEE/ACM Transactions on Networking, 30(4): 1540–1553, 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2022.3146126.

Nasralla:2022:BRO

[NEH^{+}22] Zaid H. Nasralla, Taisir E. H. Elgorashi, Ali Hammadi, Mohamed O. I. Musa, and Jaafar M. H. Elmirghani. Blackout resilient optical core network. IEEE/ACM Transactions on Networking, 30(4):1795– 1806, 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2022.3156529.

Nikkhah:2016:MII

[NG16] Mehdi Nikkhah and Roch Guerin. Migrating the Internet to IPv6: an exploration of the when and why. IEEE/ACM Transactions on Networking, 24(4): 2291–2304, August 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Nayak:2019:MMU

[NGK19] Peshal Nayak, Michele Garetto, and Edward W. Knightly. Modeling multiuser WLANs under closedloop traffic. IEEE/ACM Transactions on Networking, 27(2):763–776, April 2019. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Neglia:2022:SCT

[NGL22] Giovanni Neglia, Michele Garetto, and Emilio Leonardi. Similarity caching: Theory and algorithms. IEEE/ ACM Transactions on Networking, 30(2):475–486, April 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2021.3126368.

Nguyen:2019:RFV

[NGRF19] Thi-Minh Nguyen, Andre Girard, Catherine Rosenberg, and Serge Fdida. Routing via functions in virtual networks: The curse of choices. IEEE/ACM Transactions on Networking, 27(3):1192–1205, June 2019. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Neves:2021:DPE

[NHLB21] Miguel Neves, Bradley Huffaker, Kirill Levchenko, and Marinho Barcellos. Dynamic property enforcement in programmable data planes. IEEE/ ACM Transactions on Networking, 29(4):1540–1552, August 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2021.3068339.

Nam:2019:ODR

[NJK^{+}19] Jaehyun Nam, Hyeonseong Jo, Yeonkeun Kim, Phillip Porras, Vinod Yegneswaran, and Seungwon Shin. Operator-defined reconfigurable network OS for software-defined networks. IEEE/ACM Transactions on Networking, 27 (3):1206–1219, June 2019. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Nandigam:2019:SWL

[NJM^{+}19] Anvitha Nandigam, Suraj Jog, D. Manjunath, Jayakrishnan Nair, and Balakrishna J. Prabhu. Sharing within limits: Partial resource pooling in loss systems. IEEE/ACM Transactions on Networking, 27(4):1305–1318, August 2019. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Nair:2016:WHT

[NJW16] Jayakrishnan Nair, Krishna Jagannathan, and Adam Wierman. When heavytailed and light-tailed flows compete: the response time tail under generalized max-weight scheduling. IEEE/ACM Transactions on Networking, 24 (2):982–995, April 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Naveen:2020:COD

[NK20] K. P. Naveen and Anurag Kumar. Coverage in one-dimensional wireless networks with infrastructure nodes and relay extensions. IEEE/ACM Transactions on Networking, 28(1):140–153, February 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL https://

dl.acm.org/doi/abs/10.
1109/TNET.2019.2957752.

Nosyk:2023:CRP

[NKL^{+}23] Yevheniya Nosyk, Maciej Korczyński, Qasim Lone, Marcin Skwarek, Baptiste Jonglez, and Andrzej Duda. The Closed Resolver Project: Measuring the deployment of inbound source address validation. IEEE/ACM Transactions on Networking, 31(6):2589– 2603, 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2023.3257413.

Naribole:2017:SMH

[NKNK17] Sharan Naribole, Edward Knightly, Sharan Naribole, and Edward Knightly. Scalable multicast in highly-directional 60-GHz WLANs. IEEE/ ACM Transactions on Networking, 25(5):2844–2857, October 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Nam:2008:MBA

[NKS08] Seung Yeob Nam, Sunggon Kim, and Dan Keun Sung. Measurement-based admission control at edge routers. IEEE/ACM Transactions on Networking, 16(2):410– 423, April 2008. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Naor:1999:LLA

[NL99] Zohar Naor and Hanoch Levy. LATS: a loadadaptive threshold scheme for tracking mobile users. IEEE/ACM Transactions on Networking, 7(6):808– 817, December 1999. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL http://

www.acm.org/pubs/citations/
journals/ton/1999-7-6/
p808-naor/.

Ng:2007:TAI

[NL07] Ping Chung Ng and Soung Chang Liew. Throughput analysis of IEEE802.11 multihop ad hoc networks. IEEE/ACM Transactions on Networking, 15(2):309– 322, April 2007. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Naghizadeh:2016:PTM

[NL16] Parinaz Naghizadeh and Mingyan Liu. Perceptions and truth: a mechanism design approach to crowd-sourcing reputation. IEEE/ACM Transactions on Networking, 24(1):163– 176, February 2016. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Nistor:2015:HAP

[NLB15] Maricica Nistor, Daniel E.

Lucani, and João Barros. Hardware abstraction and protocol optimization for coded sensor networks. IEEE/ACM Transactions on Networking, 23(3):866– 879, June 2015. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Nguyen:2019:MBF

[NLB19] Duong Tung Nguyen, Long Bao Le, and Vijay K. Bhargava. A market-based framework for multi-resource allocation in fog computing. IEEE/ACM Transactions on Networking, 27(3):1151– 1164, June 2019. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Niu:2016:AFP

[NLNL16] Di Niu, Baochun Li, Di Niu, and Baochun Li. An asynchronous fixedpoint algorithm for resource sharing with coupled objectives. IEEE/ ACM Transactions on Networking, 24(5):2593–2606, October 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Neglia:2021:SAK

[NLRS21] Giovanni Neglia, Emilio Leonardi, Guilherme Iecker Ricardo, and Thrasyvoulos Spyropoulos. A Swiss Army knife for online caching in small cell networks. IEEE/ ACM Transactions on Networking, 29(6):2536–2547, December 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2021.3100757.

Nazemi:2019:DOF

[NLS19] Sepideh Nazemi, Kin K.

Leung, and Ananthram Swami. Distributed optimization framework for in-network data processing. IEEE/ACM Transactions on Networking, 27(6):2432–2443, December 2019. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL https://

dl.acm.org/doi/abs/10.
1109/TNET.2019.2953581.

Norige:2018:TUF

[NLT^{+}18] Eric Norige, Alex X. Liu, Eric Torng, Eric Torng, Eric Norige, and Alex X. Liu. A ternary unification framework for optimizing TCAM-based packet classification systems. IEEE/ ACM Transactions on Networking, 26(2):657–670, April 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Nelakuditi:2007:FLR

[NLY^{+}07] Srihari Nelakuditi, Sanghwan Lee, Yinzhe Yu, Zhi-Li Zhang, and Chen-Nee Chuah. Fast local rerouting for handling transient link failures. IEEE/ACM Transactions on Networking, 15(2):359–372, April 2007. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Naser:2006:JOI

[NM06] Hassan Naser and Hussein T. Mouftah. A joint-ONU interval-based dynamic scheduling algorithm for Ethernet passive optical networks. IEEE/ACM Transactions on Networking, 14(4):889–899, August 2006. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Nair:2009:DIO

[NM09] Jayakrishnan Nair and D. Manjunath. Distributed iterative optimal resource allocation with concurrent updates of routing and flow control variables. IEEE/ ACM Transactions on Networking, 17(4):1312–1325, August 2009. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Neely:2007:LDT

[NMC07] Michael J. Neely, Eytan Modiano, and Yuan-Sheng Cheng. Logarithmic delay for N × N packet switches under the crossbar constraint. IEEE/ACM Transactions on Networking, 15 (3):657–668, June 2007. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Nikkhah:2017:SEP

[NMD^{+}17] Mehdi Nikkhah, Aman Mangal, Constantine Dovrolis, Roch Guerin, Mehdi Nikkhah, Aman Mangal, Constantine Dovrolis, and Roch Guerin. A statistical exploration of protocol adoption. IEEE/ ACM Transactions on Networking, 25(5):2858–2871, October 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Nong:1999:ANA

[NMH99] Ge Nong, Jogesh K. Muppala, and Mounir Hamdi. Analysis of nonblocking ATM switches with multiple input queues. IEEE/ ACM Transactions on Networking, 7(1):60–74, February 1999. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1999-7-1/
p60-nong/.

Newman:1998:ISU

[NML98] Peter Newman, Greg Minshall, and Thomas L. Lyon. IP switching — ATM under IP. IEEE/ ACM Transactions on Networking, 6(2):117–129, April 1998. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/
                                        journals/ton/1998-6-2/
                                        p117-newman/.

Neely:2008:FOS

[NML08] Michael J. Neely, Eytan Modiano, and Chih-Ping Li. Fairness and optimal stochastic control for heterogeneous networks. IEEE/ACM Transactions on Networking, 16(2):396– 409, April 2008. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Neely:2003:PAR

[NMR03] Michael J. Neely, Eytan Modiano, and Charles E. Rohrs. Power allocation and routing in multibeam satellites with timevarying channels. IEEE/ ACM Transactions on Networking, 11(1):138–152, February 2003. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Nahir:2014:TDC

[NOF14] Amir Nahir, Ariel Orda, and Ari Freund. Topology design of communication networks: a game-theoretic perspective. IEEE/ACM Transactions on Networking, 22(2):405–414, April 2014. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Ngo:2006:CAN

[NPQ06] Hung Q. Ngo, Dazhen Pan, and Chunming Qiao. Constructions and analyses of nonblocking WDM switches based on arrayed waveguide grating and limited wavelength conversion. IEEE/ACM Transactions on Networking, 14(1):205– 217, February 2006. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Ngo:2007:OSN

[NPY07] Hung Q. Ngo, Dazhen Pan, and Yuanyuan Yang. Optical switching networks with minimum number of limited-range wavelength converters. IEEE/ACM Transactions on Networking, 15(4):969–979, August 2007. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Neelakantan:1998:SFM

[NR98] B. Neelakantan and S. V.

Raghavan. Scientific foundations to the multilevel method. IEEE/ ACM Transactions on Networking, 6(3):337–346, June 1998. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/
                                        journals/ton/1998-6-3/
                                        p337-neelakantan/.

Nguyen:2013:RSA

[NR13] Hung X. Nguyen and Matthew Roughan. Rigorous statistical analysis of Internet loss measurements. IEEE/ACM Transactions on Networking, 21 (3):734–745, June 2013. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Nadig:2022:SSM

[NRB22] Deepak Nadig, Byrav Ramamurthy, and Brian Bockelman. SNAG: SDNmanaged network architecture for GridFTP transfers using applicationawareness. IEEE/ACM Transactions on Networking, 30(4):1585–1598, 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1109/TNET.2022.3150000.

Northcote:1998:SCP

[NS98] Bruce S. Northcote and Donald E. Smith. Service control point overload rules to protect intelligent network services. IEEE/ACM Transactions on Networking, 6(1):71–81, February 1998. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1998-6-1/
p71-northcote/.

Nayak:2003:DON

[NS03] Tapan Kumar Nayak and Kumar N. Sivarajan. Dimensioning optical networks under traffic growth models. IEEE/ACM Transactions on Networking, 11(6):935–947, December 2003. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Nyang:2016:RCC

[NS16] DaeHun Nyang and DongOh Shin. Recyclable counter with confinement for real-time per-flow measurement. IEEE/ACM Transactions on Networking, 24(5):3191–3203, October 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Naveen:2021:DAM

[NS21] K. P. Naveen and Rajesh Sundaresan. Doubleauction mechanisms for resource trading markets. IEEE/ACM Transactions on Networking, 29(3):1210– 1223, June 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2021.3058251.

Niu:2022:NMN

[NSC^{+}22] Zhixiong Niu, Qiang Su, Peng Cheng, Yongqiang Xiong, Dongsu Han, Keith Winstein, Chun Jason Xue, and Hong Xu. NetKernel: Making network stack part of the virtualized infrastructure. IEEE/ ACM Transactions on Networking, 30(3):999–1013, June 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2021.3129806.

Nuggehalli:2006:ECP

[NSCR06] Pavan Nuggehalli, Vikram Srinivasan, Carla-Fabiana Chiasserini, and Ramesh R. Rao. Efficient cache placement in multi-hop wireless networks. IEEE/ ACM Transactions on Networking, 14(5):1045–1055, October 2006. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Ning:2016:FAP

[NSP^{+}16] Jianxia Ning, Shailendra Singh, Konstantinos Pelechrinis, Bin Liu, Srikanth V. Krishnamurthy, and Ramesh Govindan. Forensic analysis of packet losses in wireless networks.

IEEE/ACM Transactions on Networking, 24(4):1975– 1988, August 2016. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Narahari:1996:EAE

[NSS96] Bhagirath Narahari, Sunil Shende, and Rahul Simha. Efficient algorithms for erasure node placement on slotted dual bus networks. IEEE/ACM Transactions on Networking, 4(5):779– 784, October 1996. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL http://

www.acm.org/pubs/citations/
journals/ton/1996-4-5/
p779-narahari/.

Narvaez:2000:NDA

[NST00] Paolo Narváez, Kai-Yeung Siu, and Hong-Yi Tzeng. New dynamic algorithms for shortest path tree computation. IEEE/ACM Transactions on Networking, 8(6):734–746, 2000. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL

http://www.acm.org/pubs/
articles/journals/ton/
2000-8-6/p734-narvaez/
p734-narvaez.pdf;   http:
//www.acm.org/pubs/citations/
journals/ton/2000-8-6/
p734-narvaez/.

Narvaez:2001:NDS

[NST01] Paolo Narváez, Kai-Yeung Siu, and Hong-Yi Tzeng. New dynamic SPT algorithm based on a balland-string model. IEEE/ ACM Transactions on Networking, 9(6):706–718, December 2001. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Nencioni:2016:SEG

[NST^{+}16] Gianfranco Nencioni, Nishanth Sastry, Gareth Tyson, Vijay Badrinarayanan, Dmytro Karamshuk, Jigna Chandaria, and Jon Crowcroft. SCORE: Exploiting global broadcasts to create offline personal channels for ondemand access. IEEE/ ACM Transactions on Networking, 24(4):2429–2442, August 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Ni:2011:CSP

[NSW11] Jian Ni, R. Srikant, and Xinzhou Wu. Coloring spatial point processes with applications to peer discovery in large wireless networks. IEEE/ACM Transactions on Networking, 19 (2):575–588, April 2011. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Narasimha:2020:MFG

[NSY20] Dheeraj Narasimha, Srinivas Shakkottai, and Lei Ying. A mean field game analysis of distributed MAC in ultradense multichannel wireless networks. IEEE/ ACM Transactions on Networking, 28(5):1939–1952, October 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2020.3002912.

Narasimha:2023:ADD

[NSY23] Dheeraj Narasimha, Srinivas Shakkottai, and Lei Ying. Age-dependent distributed MAC for ultradense wireless networks. IEEE/ACM Transactions on Networking, 31(4):1674– 1687, 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2022.3228173.

Noel:2000:PMM

[NT00] Eric Noel and K. Wendy Tang. Performance modeling of multihop network subject to uniform and nonuniform geometric traffic. IEEE/ACM Transactions on Networking, 8(6): 763–774, 2000. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/articles/
                                        journals/ton/2000-8-6/
                                        p763-noel/p763-noel.pdf;
                                         http://www.acm.org/pubs/
                                        citations/journals/ton/
                                        2000-8-6/p763-noel/.

Nguyen:2024:PPS

[NT24] Truc Nguyen and My T. Thai. Preserving privacy and security in federated learning. IEEE/ ACM Transactions on Networking, 32(1):833– 843, 2024. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2023.3302016.

Nguyen:2017:BSA

[NTD17] Hung T. Nguyen, My T. Thai, and Thang N. Dinh. A billion-scale approximation algorithm for maximizing benefit in viral marketing. IEEE/ACM Transactions on Networking, 25(4): 2419–2429, August 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Nagy:2018:NVI

[NTR18] Mate Nagy, Janos Tapolcai, and Gabor Retvari. Node virtualization for IP level resilience. IEEE/ACM Transactions on Networking, 26(3):1250–1263, June 2018. CODEN IEANEP.

ISSN 1063-6692 (print), 1558-2566 (electronic).

Ni:2012:QCQ

[NTS12] Jian Ni, Bo Tan, and R. Srikant. Q-CSMA: queue-length-based CSMA/ CA algorithms for achieving maximum throughput and low delay in wireless networks. IEEE/ACM Transactions on Networking, 20(3):825–836, June 2012. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Naghsh:2021:CFS

[NV21] Zahra Naghsh and Shahrokh Valaee. Conflict-free scheduling in cellular V2X communications. IEEE/ ACM Transactions on Networking, 29(1):106–119, February 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2020.3030850.

Ngo:2009:RNW

[NWP09] Hung Q. Ngo, Yang Wang, and Dazhen Pan. Rearrangeable and nonblocking [w, f]-distributors. IEEE/ ACM Transactions on Networking, 17(3):990–1001, June 2009. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Ni:2010:EDR

[NXTY10] Jian Ni, Haiyong Xie, Sekhar Tatikonda, and Yang Richard Yang. Efficient and dynamic routing topology inference from end-to-end measurements. IEEE/ACM Transactions on Networking, 18(1):123– 135, February 2010. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Neumayer:2011:AVF

[NZCM11] Sebastian Neumayer, Gil Zussman, Reuven Cohen, and Eytan Modiano. Assessing the vulnerability of the fiber infrastructure to disasters. IEEE/ACM Transactions on Networking, 19(6):1610–1623, December 2011. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Nelakuditi:2002:APR

[NZTD02] Srihari Nelakuditi, Zhi Li Zhang, Rose P. Tsang, and David H. C. Du. Adaptive proportional routing: a localized QoS routing approach. IEEE/ACM Transactions on Networking, 10(6):790–804, December 2002. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Ni:2024:AOP

[NZW24] Peikun Ni, Jianming Zhu, and Guoqing Wang. Activityoriented production promotion utility maximization in metaverse social networks. IEEE/ACM Transactions on Networking, 32(2):1140– 1154, April 2024. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2023.3309624.

Oguz:2015:SDP

[OAN15] Barlas Oğuz, Venkat Anantharam, and Ilkka Norros. Stable distributed P2P protocols based on random peer sampling. IEEE/ ACM Transactions on Networking, 23(5):1444–1456, October 2015. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Ozdaglar:2003:RWA

[OB03] Asuman E. Ozdaglar and Dimitri P. Bertsekas. Routing and wavelength assignment in optical networks. IEEE/ACM Transactions on Networking, 11 (2):259–272, April 2003. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Ouyang:2017:LOH

[OBS17] Wenzhuo Ouyang, Jingwen Bai, and Ashutosh Sabharwal. Leveraging onehop information in massive MIMO full-duplex wireless systems. IEEE/ACM Transactions on Networking, 25(3):1528–1539, June 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Ozdemir:2010:IFD

[OÇ10] Suat Ozdemir and Hasan Çam. Integration of false data detection with data aggregation and confidential transmission in wireless sensor networks. IEEE/ACM Transactions on Networking, 18(3):736– 749, June 2010. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Oller:2016:TCS

[ODC^{+}16] Joaquim Oller, Ilker Demirkol, Jordi Casademont, Josep Paradells, Gerd Ulrich Gamm, and Leonhard Reindl. Has time come to switch from duty-cycled MAC protocols to wakeup radio for wireless sensor networks? IEEE/ACM Transactions on Networking, 24(2):674–687, April 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Olah:1996:CMT

[OdG96] András L. Oláh and Sonia M. Heemstra de Groot.

Comments on “Minimumlatency transport protocols with modulo-N”. IEEE/ ACM Transactions on Networking, 4(4):660–666, August 1996. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1996-4-4/

p660-olah/. See [SL95].

Olah:1997:ASV

[OdG97] András L. Oláh and Sonia M. Heemstra de Groot. Alternative specification and verification of a periodic state exchange protocol. IEEE/ACM Transactions on Networking, 5(4):525– 529, August 1997. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1997-5-4/
p525-olah/.

Ozan:2023:POS

[ODJ23] Waseem Ozan, Izzat Darwazeh, and Kyle Jamieson. Partial OFDM symbol recovery to improve interfering wireless networks operation in collision environments. IEEE/ ACM Transactions on Networking, 31(2):680–694, April 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.

                                        acm.org/doi/10.1109/TNET.
                                        2022.3202857.

Okamura:2009:MAP

[ODT09] Hiroyuki Okamura, Tadashi Dohi, and Kishor S. Trivedi. Markovian arrival process parameter estimation with group data. IEEE/ACM Transactions on Networking, 17(4):1326– 1339, August 2009. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Ouyang:2016:DSM

[OES16] Wenzhuo Ouyang, Atilla Eryilmaz, and Ness B. Shroff. Downlink scheduling over Markovian fading channels. IEEE/ACM Transactions on Networking, 24(3):1801–1812, June 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Oggier:2011:ACA

[OF11] Frédérique Oggier and Hanane Fathi. An authentication code against pollution attacks in network coding. IEEE/ACM Transactions on Networking, 19 (6):1587–1596, December 2011. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Orsini:2014:EIK

[OGLK14] Chiara Orsini, Enrico Gregori, Luciano Lenzini, and Dmitri Krioukov. Evolution of the Internet kdense structure. IEEE/ ACM Transactions on Networking, 22(6):1769–1780, December 2014. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Oki:2002:CRR

[OJRCC02] Eiji Oki, Zhigang Jing, Roberto Rojas-Cessa, and H. Jonathan Chao. Concurrent round-robin-based dispatching schemes for Closnetwork switches. IEEE/ ACM Transactions on Networking, 10(6):830–844, December 2002. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Ok:2016:MDS

[OJSY16] Jungseul Ok, Youngmi Jin, Jinwoo Shin, and Yung Yi. On maximizing diffusion speed over social networks with strategic users. IEEE/ ACM Transactions on Networking, 24(6):3798–3811, December 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Oikonomou:2023:PFP

[OKAS23] Konstantinos Oikonomou, George Koufoudakis, Sonia Aïssa, and Ioannis Stavrakakis. Probabilistic flooding performance analysis exploiting graph spectra properties. IEEE/ACM Transactions on Networking, 31 (1):133–146, 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2022.3192310.

Omundsen:1994:PMA

[OKM94] Daniel S. Omundsen, A. Roger Kaye, and Samy A. Mahmoud. A pipelined, multiprocessor architecture for a connectionless server for broadband ISDN. IEEE/ACM Transactions on Networking, 2(2):181– 192, April 1994. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1994-2-2/
p181-omundsen/.

Oh:2016:FBP

[OL16] Bong-Hwan Oh and Jaiyong Lee. Feedback-based path failure detection and buffer blocking protection for MPTCP. IEEE/ACM Transactions on Networking, 24(6):3450–3461, December 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Ou:2017:CSP

[OLZ17] Jiajue Ou, Mo Li, and Yuanqing Zheng. Come and be served: Parallel decoding for COTS RFID tags. IEEE/ACM Transactions on Networking, 25 (3):1569–1581, June 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Ohsita:2010:GRV

[OMA^{+}10] Yuichi Ohsita, Takashi Miyamura, Shin’ichi Arakawa, Shingo Ata, Eiji Oki, Kohei Shiomoto, and Masayuki Murata. Gradually reconfiguring virtual network topologies based on estimated traffic matrices. IEEE/ACM Transactions on Networking, 18(1):177– 189, February 2010. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Otoshi:2018:HMP

[OOM^{+}18] Tatsuya Otoshi, Yuichi Ohsita, Masayuki Murata, Yousuke Takahashi, Keisuke Ishibashi, Kohei Shiomoto, and Tomoaki Hashimoto. Hierarchical model predictive traffic engineering. IEEE/ ACM Transactions on Networking, 26(4):1754–1767, August 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Oya:2016:DPM

[OPGT16] Simon Oya, Fernando Perez-Gonzalez, and Carmela Troncoso. Design of pool mixes against profiling attacks in real conditions. IEEE/ACM Transactions on Networking, 24(6):3662– 3675, December 2016. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Oliveira:2010:COI

[OPW^{+}10] Ricardo Oliveira, Dan Pei, Walter Willinger, Beichuan Zhang, and Lixia Zhang. The (in)completeness of the observed Internet ASlevel structure. IEEE/ ACM Transactions on Networking, 18(1):109–122, February 2010. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Onus:2011:MMD

[OR11] Melih Onus and Andréa W. Richa. Minimum maximumdegree publish-subscribe overlay network design. IEEE/ACM Transactions on Networking, 19(5):1331– 1343, October 2011. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Orda:1999:REE

[Ord99] Ariel Orda. Routing with end-to-end QoS guarantees in broadband networks. IEEE/ACM Transactions on Networking, 7(3):365– 374, June 1999. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1999-7-3/
p365-orda/.

Orda:1993:CRM

[ORS93a] Ariel Orda, Raphael Rom, and Nahum Shimkin. Competitive routing in multiuser communication networks. IEEE/ACM Transactions on Networking, 1 (5):510–521, October 1993. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL

http://www.acm.org/pubs/
citations/journals/ton/
1993-1-5/p510-orda/.

Orda:1993:MDR

[ORS93b] Ariel Orda, Raphael Rom, and Moshe Sidi. Minimum delay routing in stochastic networks. IEEE/ ACM Transactions on Networking, 1(2):187–198, April 1993. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1993-1-2/
p187-orda/.

Orda:2003:PSQ

[OS03] Ariel Orda and Alexander Sprintson. Precomputation schemes for QoS routing. IEEE/ACM Transactions on Networking, 11 (4):578–591, August 2003. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Orda:2005:SAP

[OS05] Ariel Orda and Alexander Sprintson. A scalable approach to the partition of QoS requirements in unicast and multicast. IEEE/ ACM Transactions on Networking, 13(5):1146–1159, October 2005. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Ornee:2021:SRE

[OS21] Tasmeen Zaman Ornee and Yin Sun. Sampling and remote estimation for the Ornstein– Uhlenbeck process through queues: Age of information and beyond. IEEE/ ACM Transactions on Networking, 29(5):1962–1975, October 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2021.3078137.

Ofek:1997:ISA

[OSW97] Yoram Ofek, Khosrow Sohraby, and Ho-Ting Wu. Integration of synchronous and asynchronous traffic on the MetaRing and its performance study. IEEE/ ACM Transactions on Networking, 5(1):111–121, February 1997. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1997-5-1/
p111-ofek/.

Ou:2006:SVC

[OSZ^{+}06] Canhui Ou, Laxman H. Sahasrabuddhe, Keyao Zhu, Charles U. Martel, and Biswanath Mukherjee. Survivable virtual concatenation for data over SONET/ SDH in optical transport networks. IEEE/ ACM Transactions on Networking, 14(1):218–231, February 2006. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Ostovari:2016:SVS

[OWKS16] Pouya Ostovari, Jie Wu, Abdallah Khreishah, and Ness B. Shroff. Scalable video streaming with helper nodes using random linear network coding. IEEE/ACM Transactions on Networking, 24 (3):1574–1587, June 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Ohsaki:1997:PIB

[OWMM97] Hiroyuki Ohsaki, Naoki Wakamiya, Masayuki Murata, and Hideo Miyahara. Performance of an input/ output buffered-type ATM LAN switch with backpressure function. IEEE/ ACM Transactions on Networking, 5(2):278–290, April 1997. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1997-5-2/
p278-ohsaki/.

Ofek:1995:MPA

[OY95] Yoram Ofek and Moti Yung. METANET: principles of an arbitrary topology LAN. IEEE/ ACM Transactions on Networking, 3(2):169–180, April 1995. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1995-3-2/
p169-ofek/.

Ouyang:2013:ATO

[OY13] Ming Ouyang and Lei Ying. Approaching throughput optimality with limited feedback in multichannel wireless downlink networks. IEEE/ACM Transactions on Networking, 21 (6):1827–1838, December 2013. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Oliveira:2009:QPE

[OZPZ09] Ricardo Oliveira, Beichuan Zhang, Dan Pei, and Lixia Zhang. Quantifying path exploration in the Internet. IEEE/ACM Transactions on Networking, 17 (2):445–458, April 2009. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Passos:2012:JAR

[PA12] Diego Passos and Celio V. N. Albuquerque. A joint approach to routing metrics and rate adaptation in wireless mesh networks. IEEE/ACM Transactions on Networking, 20 (4):999–1009, August 2012. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Padmanabhan:1995:EAF

[Pad95] Krishnan Padmanabhan.

An efficient architecture for fault-tolerant ATM switches. IEEE/ACM Transactions on Networking, 3(5):527–537, October 1995. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1995-3-5/
p527-padmanabhan/.

Pankaj:1999:WRM

[Pan99] Rajesh K. Pankaj. Wavelength requirements for multicasting in all-optical networks. IEEE/ACM Transactions on Networking, 7(3):414–424, June 1999. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL

http://www.acm.org/pubs/
citations/journals/ton/
1999-7-3/p414-pankaj/.

Paxson:1994:EDA

[Pax94]  Vern Paxson.  Empiri-

cally derived analytic models of wide-area TCP connections. IEEE/ACM Transactions on Networking, 2(4):316–336, August 1994. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1994-2-4/
p316-paxson/.

Paxson:1997:ERB

[Pax97]  Vern Paxson.  End-to-

end routing behavior in the Internet. IEEE/ACM Transactions on Networking, 5(5):601–615, October 1997. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1997-5-5/
p601-paxson/.

Paxson:1999:EEI

[Pax99] Vern Paxson. End-to-end Internet packet dynamics.

IEEE/ACM Transactions on Networking, 7(3):277– 292, June 1999. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1999-7-3/
p277-paxson/.

Pattavina:1993:AIO

[PB93] Achille Pattavina and Giacomo Bruzzi. Analysis of input and output queueing for nonblocking ATM switches. IEEE/ACM Transactions on Networking, 1(3):314– 328, June 1993. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1993-1-3/
p314-pattavina/.

Perez-Bueno:2022:LPP

[PBGMFM22] Fernando Pérez-Bueno, Luz García, Gabriel Maciá-Fernández, and Rafael Molina. Leveraging a probabilistic PCA model to understand the multivariate statistical network monitoring framework for network security anomaly detection. IEEE/ACM Transactions on Networking, 30(3):1217– 1229, June 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.

                                        acm.org/doi/10.1109/TNET.
                                        2021.3138536.

Pelechrinis:2011:MDA

[PBKG11] Konstantinos Pelechrinis, Ioannis Broustis, Srikanth V. Krishnamurthy, and Christos Gkantsidis. A measurementdriven anti-jamming system for 802.11 networks. IEEE/ACM Transactions on Networking, 19(4):1208– 1222, August 2011. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Pan:2023:OSD

[PBSS23] Jiayu Pan, Ahmed M. Bedewy, Yin Sun, and Ness B. Shroff. Optimal sampling for data freshness: Unreliable transmissions with random two-way delay. IEEE/ACM Transactions on Networking, 31 (1):408–420, 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2022.3194417.

Pasic:2020:MCS

[PBT^{+}20] Alija Pašić, Péter Babarczi, János Tapolcai, Erika R. Bérczi-Kovács, Zoltán Király, and Lajos Rónyai. Minimum cost survivable routing algorithms for generalized diversity coding. IEEE/ACM Transactions on Networking, 28(1):289– 300, February 2020. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL https:

//dl.acm.org/doi/abs/
10.1109/TNET.2019.2963574.

Pananjady:2017:OAC

[PBV17] Ashwin Pananjady, Vivek Kumar Bagaria, and Rahul Vaze. Optimally approximating the coverage lifetime of wireless sensor networks. IEEE/ACM Transactions on Networking, 25 (1):98–111, February 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Pai:2019:DCC

[PC19] Kung-Jui Pai and Jou-Ming Chang. Dual-CISTs: Configuring a protection routing on some Cayley networks. IEEE/ACM Transactions on Networking, 27(3):1112–1123, June 2019. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Partridge:1998:IR

[PCB^{+}98] Craig Partridge, Philip P.

Carvey, Ed Burgess, Isidro Castineyra, Tom Clarke, Lise Graham, Michael Hathaway, Phil Herman, Allen King, Steve Kohalmi, Tracy Ma, John Mcallen, Trevor Mendez, Walter C. Milliken, Ronald Pettyjohn, John Rokosz, Joshua Seeger, Michael Sollins, Steve Storch, Benjamin Tober, and Gregory D. Troxel. A 50-Gb/s IP router. IEEE/ ACM Transactions on Networking, 6(3):237–248, June 1998. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/
                                        journals/ton/1998-6-3/
                                        p237-partridge/.

Premkumar:2015:PFC

[PCL15] Karumbu Premkumar, Xiaomin Chen, and Douglas J. Leith. Proportional fair coding for wireless mesh networks. IEEE/ ACM Transactions on Networking, 23(1):269–281, February 2015. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Patel:2008:IQP

[PCV08] Maulin Patel, R. Chandrasekaran, and S. Venkatesan. Improved quasi-path restoration in mesh networks. IEEE/ACM Transactions on Networking, 16(1):144–156, February 2008. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Peng:2016:TCT

[PCW^{+}16] Yang Peng, Kai Chen, Guohui Wang, Wei Bai, Yangming Zhao, Hao Wang, Yanhui Geng, Zhiqiang Ma, and Lin Gu. Towards comprehensive traffic forecasting in cloud computing: Design and application. IEEE/ACM Transactions on Networking, 24(4): 2210–2222, August 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Pokhrel:2023:FED

[PCW23] Shiva Raj Pokhrel, Jinho Choi, and Anwar Walid. Fair and efficient distributed edge learning with hybrid multipath TCP. IEEE/ACM Transactions on Networking, 31(4):1582– 1594, 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2022.3219924.

Patil:2007:MRQ

[PD07] Shailesh Patil and Gustavo De Veciana. Managing resources and quality of service in heterogeneous wireless systems exploiting opportunism. IEEE/ ACM Transactions on Networking, 15(5):1046–1058, October 2007. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Pacifici:2016:CSD

[PD16a] Valentino Pacifici and Gyorgy Dan. Coordinated selfish distributed caching for peering content-centric networks. IEEE/ACM Transactions on Networking, 24(6):3690–3701, December 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Pedersen:2016:EMV

[PD16b] Hasti A. Pedersen and Sujit Dey. Enhancing mobile video capacity and quality using rate adaptation, RAN caching and processing. IEEE/ACM Transactions on Networking, 24 (2):996–1010, April 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Papandriopoulos:2008:ODP

[PDE08] John Papandriopoulos, Subhrakanti Dey, and Jamie Evans. Optimal and distributed protocols for cross-layer design of physical and transport layers in MANETs. IEEE/ ACM Transactions on Networking, 16(6):1392–1405, December 2008. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Paschos:2020:OCO

[PDI20] Georgios S. Paschos, Apostolos Destounis, and George Iosifidis. Online convex optimization for caching networks. IEEE/ACM Transactions on Networking, 28 (2):625–638, April 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL

https://dl.acm.org/doi/
abs/10.1109/TNET.2020.
2968424.

Panagakis:2004:OCA

[PDSK04] Antonis Panagakis, Nandita Dukkipati, Ioannis Stavrakakis, and Joy Kuri. Optimal call admission control on a single link with a GPS scheduler. IEEE/ACM Transactions on Networking, 12(5):865– 878, October 2004. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Prasad:2009:RBS

[PDT09] Ravi S. Prasad, Constantine Dovrolis, and Marina Thottan. Router buffer sizing for TCP traffic and the role of the output/input capacity ratio. IEEE/ ACM Transactions on Networking, 17(5):1645–1658, October 2009. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Pacifico:2024:EHS

[PDV^{+}24] Racyus D. G. Pacífico, Lucas F. S. Duarte, Luiz F. M. Vieira, Barath Raghavan, José A. M. Nacif, and Marcos A. M. Vieira. eBPFlow: a hardware/software platform to seamlessly offload network functions leveraging eBPF. IEEE/ ACM Transactions on Networking, 32(2):1319–1332, April 2024. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2023.3318251.

Pascu:2009:CFA

[PEA09] Stefan Alexandru Pascu and Ahmed A. El-Amawy. On conflict-free all-to-all broadcast in one-hop optical networks of arbitrary topologies. IEEE/ ACM Transactions on Networking, 17(5):1619–1630, October 2009. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Pujol:2012:LEC

[PES^{+}12] Josep M. Pujol, Vijay Erramilli, Georgos Siganos, Xiaoyuan Yang, Nikolaos Laoutaris, Parminder Chhabra, and Pablo Rodriguez. The little engine(s) that could: scaling online social networks. IEEE/ ACM Transactions on Networking, 20(4):1162–1175, August 2012. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Paxson:1995:WAT

[PF95] Vern Paxson and Sally Floyd. Wide area traffic: the failure of Poisson modeling. IEEE/ ACM Transactions on Networking, 3(3):226–244, June 1995. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1995-3-3/
p226-paxson/.

Picker:1996:ESF

[PFC96] Dan Picker, Ronald D. Fellman, and Paul M. Chau. An extension to the SCI flow control protocol for increased network efficiency. IEEE/ACM Transactions on Networking, 4(1):71–85, February 1996. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1996-4-1/
p71-picker/.

Padhye:2000:MTR

[PFTK00] Jitendra Padhye, Victor Firoiu, Donald F. Towsley, and James F. Kurose. Modeling TCP Reno performance: a simple model and its empirical validation. IEEE/ACM Transactions on Networking, 8 (2):133–145, 2000. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/2000-8-2/
p133-padhye/. See com-

ments [CBAT06].

Parekh:1993:GPS

[PG93] Abhay K. Parekh and Robert G. Gallager. A generalized processor sharing approach to flow control in integrated services networks: the single-node case. IEEE/ACM Transactions on Networking, 1(3): 344–357, June 1993. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL http://

www.acm.org/pubs/citations/
journals/ton/1993-1-3/
p344-parekh/.

Parekh:1994:GPS

[PG94a] Abhay K. Parekh and Robert G. Gallagher. A generalized processor sharing approach to flow control in integrated services networks: the multiple node case. IEEE/ ACM Transactions on Networking, 2(2):137–150, April 1994. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1994-2-2/
p137-parekh/.

Pattavina:1994:PAA

[PG94b] Achille Pattavina and Stefano Gianatti. Performance analysis of ATM Banyan networks with shared queueing — part II: correlated/unbalanced offered traffic. IEEE/ ACM Transactions on Networking, 2(4):411–424, August 1994. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1994-2-4/
p411-pattavina/.

Pankaj:1995:WRA

[PG95] Rajesh K. Pankaj and Robert G. Gallager. Wavelength requirements of alloptical networks. IEEE/ ACM Transactions on Networking, 3(3):269–280, June 1995. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1995-3-3/
p269-pankaj/.

Psychas:2018:RAS

[PG18] Konstantinos Psychas and Javad Ghaderi. Randomized algorithms for scheduling multi-resource jobs in the cloud. IEEE/ACM Transactions on Networking, 26(5):2202–2215, October 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Psychasand:2021:HTP

[PG21] Konstantinos Psychasand and Javad Ghaderi. Highthroughput bin packing: Scheduling jobs with random resource demands in clusters. IEEE/ACM Transactions on Networking, 29(1):220–233, February 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2020.3034022.

Phan:2018:UCN

[PGMR18] Truong Khoa Phan, David Griffin, Elisa Maini, and Miguel Rio. Utility-centric networking: Balancing transit costs with quality of experience. IEEE/ ACM Transactions on Networking, 26(1):245–258, February 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Plante:2016:MOS

[PGV16] Jeremy M. Plante, Arush Gadkar, and Vinod M. Vokkarane. Manycast overlay in split-incapable networks for supporting bandwidth-intensive applications. IEEE/ACM Transactions on Networking, 24(1):342–354, February 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Pham:2020:CAE

[PHC20] Minh Pham, Doan B.

Hoang, and Zenon Chaczko. Congestion-aware and energyaware virtual network embedding. IEEE/ACM Transactions on Networking, 28(1):210–223, February 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL https://

dl.acm.org/doi/abs/10.
1109/TNET.2019.2958367.

Paschalidis:2015:MPA

[PHL15] Ioannis Ch. Paschalidis, Fuzhuo Huang, and Wei Lai. A message-passing algorithm for wireless network scheduling. IEEE/ ACM Transactions on Networking, 23(5):1528–1541, October 2015. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Pillai:2001:DOC

[Pil01] R. Radhakrishna Pillai. A distributed overload control algorithm for delaybounded call setup. IEEE/ ACM Transactions on Networking, 9(6):780–789, December 2001. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Pitsillides:2005:ANC

[PILR05] Andreas Pitsillides, Petros Ioannou, Marios Lestas, and Loukas Rossides. Adaptive nonlinear congestion controller for a differentiatedservices framework. IEEE/ ACM Transactions on Networking, 13(1):94–107, February 2005. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Poularakis:2019:OGS

[PIST19] Konstantinos Poularakis, George Iosifidis, Georgios Smaragdakis, and Leandros Tassiulas. Optimizing gradual SDN upgrades in ISP networks. IEEE/ACM Transactions on Networking, 27(1):288– 301, February 2019. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Preciado:2013:MBS

[PJ13] Victor M. Preciado and Ali Jadbabaie. Momentbased spectral analysis of large-scale networks using local structural information. IEEE/ACM Transactions on Networking, 21 (2):373–382, April 2013. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Pi:2018:AAH

[PJDS18] Yibo Pi, Sugih Jamin, Peter Danzig, and Jacob Shaha. AP-Atoms: a highaccuracy data-driven client aggregation for global load balancing. IEEE/ACM Transactions on Networking, 26(6):2748–2761, December 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Popescu:2019:IDS

[PJM^{+}19] Dalia Popescu, Philippe Jacquet, Bernard Mans, Robert Dumitru, Andra Pastrav, and Emanuel Puschita. Information dissemination speed in delay tolerant urban vehicular networks in a hyperfractal setting. IEEE/ACM Transactions on Networking, 27(5): 1901–1914, October 2019. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Park:2022:LSB

[PJMM22] Jeman Park, Rhongho Jang, Manar Mohaisen, and David Mohaisen. A large-scale behavioral analysis of the Open DNS resolvers on the Internet. IEEE/ACM Transactions on Networking, 30(1):76– 89, February 2022. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL https:/

/dl.acm.org/doi/10.1109/
TNET.2021.3105599.

Pronk:2001:CCR

[PK01] Verus Pronk and Jan Korst. Comments on “Carryover Round Robin: A Simple Cell Scheduling Mechanism for ATM networks”. IEEE/ACM Transactions on Networking, 9(3):373, June 2001. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). See [SMT98].

Peresini:2018:DFG

[PKK18] Peter Peresini, Maciej Kuzniar, and Dejan Kostic. Dynamic, fine-grained data plane monitoring with monocle. IEEE/ ACM Transactions on Networking, 26(1):534–547, February 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Papadogiannaki:2017:ESP

[PKVI17] Eva Papadogiannaki, Lazaros Koromilas, Giorgos Vasiliadis, and Sotiris Ioannidis. Efficient software packet processing on heterogeneous and asymmetric hardware architectures. IEEE/ACM Transactions on Networking, 25(3):1593– 1606, June 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Paschalidis:2002:PML

[PL02] Ioannis Ch. Paschalidis and Yong Liu. Pricing in multiservice loss networks: static pricing, asymptotic optimality, and demand substitution effects. IEEE/ACM Transactions on Networking, 10(3):425–438, June 2002. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Partov:2017:UFR

[PL17] Bahar Partov and Douglas J. Leith. Utility fair rate allocation in LTE/ 802.11 networks. IEEE/ ACM Transactions on Networking, 25(2):1076–1088, April 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Pacifici:2016:CBA

[PLD16] Valentino Pacifici, Frank Lehrieder, and György Dán. Cache bandwidth allocation for P2P filesharing systems to minimize inter-ISP traffic. IEEE/ACM Transactions on Networking, 24(1):437– 448, February 2016. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Pefkianakis:2013:TMA

[PLL13] Ioannis Pefkianakis, Suk-Bok Lee, and Songwu Lu. Towards MIMO-aware 802.11n rate adaptation. IEEE/ACM Transactions on Networking, 21(3):692– 705, June 2013. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Paschos:2016:NCC

[PLM^{+}16] Georgios S. Paschos, Chih-Ping Li, Eytan Modiano, Kostas Choumas, and Thanasis Korakis. Innetwork congestion control for multirate multicast. IEEE/ACM Transactions on Networking, 24(5): 3043–3055, October 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Pegatoquet:2019:WRB

[PLM19] Alain Pegatoquet, Trong Nhan Le, and Michele Magno. A wake-up radio-based MAC protocol for autonomous wireless sensor networks. IEEE/ACM Transactions on Networking, 27(1):56– 70, February 2019. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Partov:2015:UFO

[PLR15] Bahar Partov, Douglas J. Leith, and Rouzbeh Razavi. Utility fair optimization of antenna tilt angles in LTE networks. IEEE/ ACM Transactions on Networking, 23(1):175–185, February 2015. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Pappas:2019:NTB

[PLR^{+}19] Christos Pappas, Taeho Lee, Raphael M. Reischuk, Pawel Szalachowski, and Adrian Perrig. Network transparency for better Internet security. IEEE/ ACM Transactions on Networking, 27(5):2028–2042, October 2019. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Paschalidis:2007:AOT

[PLS07] Ioannis Ch. Paschalidis, Wei Lai, and David Starobinski. Asymptotically optimal transmission policies for large-scale low-power wireless sensor networks. IEEE/ACM Transactions on Networking, 15(1):105– 118, February 2007. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Pan:2021:NCN

[PLS^{+}21] Tian Pan, Xingchen Lin, Enge Song, Cheng Xu, Jiao Zhang, Hao Li, Jianhui Lv, Tao Huang, Bin Liu, and Beichuan Zhang. NB-Cache: Non-blocking in-network caching for high-performance content routers. IEEE/ACM Transactions on Networking, 29(5):1976–1989, October 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2021.3083599.

Patel:2014:BSE

[PLT14] Jignesh Patel, Alex X. Liu, and Eric Torng. Bypassing space explosion in high-speed regular expression matching. IEEE/ACM Transactions on Networking, 22(6):1701–1714, December 2014. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Poularakis:2020:SPR

[PLT^{+}20] Konstantinos Poularakis, Jaime Llorca, Antonia M. Tulino, Ian Taylor, and Leandros Tassiulas. Service placement and request routing in MEC networks with storage, computation, and communication constraints. IEEE/ ACM Transactions on Networking, 28(3):1047–1060, June 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2020.2980175.

Papadimitriou:1996:LAR

[PM96] Georgios I. Papadimitriou and Dimitris G. Maritsas. Learning automatabased receiver conflict avoidance algorithms for WDM broadcast-and-select star networks. IEEE/ ACM Transactions on Networking, 4(3):407–412, June 1996. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1996-4-3/
p407-papadimitriou/.

Paganini:2009:UAC

[PM09] Fernando Paganini and Enrique Mallada. A unified approach to congestion control and node-based multipath routing. IEEE/ ACM Transactions on Networking, 17(5):1413–1426, October 2009. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Pedarsani:2016:OCC

[PMAN16] Ramtin Pedarsani, Mohammad Ali Maddah-Ali, and Urs Niesen. Online coded caching. IEEE/ACM Transactions on Networking, 24(2):836–845, April 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Pollini:1995:ERI

[PMH95] Gregory P. Pollini and Kathleen S. Meier-Hellstern.

Efficient routing of information between interconnected cellular mobile switching centers. IEEE/ ACM Transactions on Networking, 3(6):765–774, December 1995. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/
                                        journals/ton/1995-3-6/
                                        p765-pollini/.

Phalak:2019:ZRP

[PMN19] Kunal Phalak, D. Manjunath, and Jayakrishnan Nair. Zero rating: The power in the middle. IEEE/ACM Transactions on Networking, 27(2):862– 874, April 2019. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Parvez:2010:ATM

[PMW10] Nadim Parvez, Anirban Mahanti, and Carey Williamson. An analytic throughput model for TCP NewReno. IEEE/ACM Transactions on Networking, 18(2):448– 461, April 2010. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Paris:2013:CLM

[PNRMC13] Stefano Paris, Cristina Nita-Rotaru, Fabio Martignon, and Antonio Capone. Cross-layer metrics for reliable routing in wireless mesh networks. IEEE/ ACM Transactions on Networking, 21(3):1003–1016, June 2013. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Papadopoulos:1993:EES

[PP93a] Christos Papadopoulos and Gurudatta M. Parulkar. Experimental evaluation of SUNOS IPC and TCP/IP protocol implementation. IEEE/ACM Transactions on Networking, 1(2):199– 216, April 1993. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1993-1-2/
p199-papadopoulos/.

Partridge:1993:FU

[PP93b] Craig Partridge and Stephen Pink. A faster UDP. IEEE/ACM Transactions on Networking, 1(4):429– 440, August 1993. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1993-1-4/
p429-partridge/.

Psounis:2002:ERW

[PP02] Konstantinos Psounis and Balaji Prabhakar. Efficient randomized web-cache replacement schemes using samples from past eviction times. IEEE/ACM Transactions on Networking, 10 (4):441–455, August 2002. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Papanikos:2017:DBR

[PP17] Nikolaos Papanikos and Evangelos Papapetrou. Deterministic broadcasting and random linear network coding in mobile ad hoc networks. IEEE/ACM Transactions on Networking, 25(3):1540–1554, June 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Papadopoulos:2015:NMR

[PPK15] Fragkiskos Papadopoulos, Constantinos Psomas, and Dmitri Krioukov. Network mapping by replaying hyperbolic growth. IEEE/ ACM Transactions on Networking, 23(1):198–211, February 2015. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Pan:2005:SME

[PPPW05] Rong Pan, Balaji Prabhakar, Konstantinos Psounis, and Damon Wischik. SHRiNK: a method for enabling scaleable performance prediction and efficient network simulation. IEEE/ACM Transactions on Networking, 13(5):975– 988, October 2005. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Polachan:2022:TMT

[PPS^{+}22] Kurian Polachan, Joydeep Pal, Chandramani Singh, T. V. Prabhakar, and Fernando A. Kuipers. TCPSbed: a modular testbed for tactile Internet-based cyber-physical systems. IEEE/ACM Transactions on Networking, 30(2):796– 811, April 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2021.3124767.

Pei:2013:AAT

[PPSV13] Guanhong Pei, Srinivasan Parthasarathy, Aravind Srinivasan, and Anil Kumar S. Vullikanti. Approximation algorithms for throughput maximization in wireless networks with delay constraints. IEEE/ ACM Transactions on Networking, 21(6):1988–2000, December 2013. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Promponas:2021:GNS

[PPTP21]  Panagiotis  Promponas,

Christos Pelekis, Eirini Eleni Tsiropoulou, and Symeon Papavassiliou. Games in normal and satisfaction form for efficient transmission power allocation under dual 5G wireless multiple access paradigm. IEEE/ ACM Transactions on Networking, 29(6):2574–2587, December 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2021.3095351.

Papadopoulos:2004:LWM

[PPV04] Christos Papadopoulos, Guru Parulkar, and George Varghese. Light-weight multicast services (LMS): a router-assisted scheme for reliable multicast. IEEE/ ACM Transactions on Networking, 12(3):456–468, June 2004. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Phan:2012:NOD

[PPV12] Khoa Tran Phan, Jaeok Park, and Mihaela Van Der Schaar. Near-optimal deviation-proof medium access control designs in wireless networks. IEEE/ ACM Transactions on Networking, 20(5):1581–1594, October 2012. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Pokhrel:2017:AMM

[PPV17] Shiva Raj Pokhrel, Manoj Panda, and Hai L. Vu. Analytical modeling of multipath TCP over lastmile wireless. IEEE/ACM Transactions on Networking, 25(3):1876–1891, June 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Posch:2017:SSA

[PRH17] Daniel Posch, Benjamin Rainer, and Hermann Hellwagner. SAF: Stochastic adaptive forwarding in named data networking. IEEE/ACM Transactions on Networking, 25(2):1089– 1102, April 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Petrovic:2006:OUR

[PRR06] Dragan Petrović, Kannan Ramchandran, and Jan Rabaey. Overcoming untuned radios in wireless networks with network coding. IEEE/ACM Transactions on Networking, 14 (SI):2649–2657, June 2006. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Pieris:1993:LLB

[PS93] Gerard R. Pieris and Galen H. Sasaki. A linear lightwave Benes network.

IEEE/ACM Transactions on Networking, 1(4):441– 445, August 1993. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1993-1-4/
p441-pieris/.

Pieris:1994:STW

[PS94] Gerard R. Pieris and Galen H. Sasaki. Scheduling transmissions in WDM broadcast-and-select networks. IEEE/ACM Transactions on Networking, 2 (2):105–110, April 1994. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL

http://www.acm.org/pubs/
citations/journals/ton/
1994-2-2/p105-pieris/.

Privalov:1998:PJA

[PS98] Aleksandr Privalov and Khosrow Sohraby. Perstream jitter analysis in CBR ATM multiplexors. IEEE/ACM Transactions on Networking, 6(2):141– 149, April 1998. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1998-6-2/
p141-privalov/.

Park:2005:OTL

[PS05] Taejoon Park and Kang G. Shin. Optimal tradeoffs for location-based routing in large-scale ad hoc networks. IEEE/ACM Transactions on Networking, 13 (2):398–410, April 2005. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Paschalidis:2009:STN

[PS09] Ioannis Ch. Paschalidis and Georgios Smaragdakis. Spatio-temporal network anomaly detection by assessing deviations of empirical measures. IEEE/ACM Transactions on Networking, 17(3):685–697, June 2009. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Picu:2015:DMF

[PS15] Andreea Picu and Thrasyvoulos Spyropoulos. DTNmeteo: forecasting the performance of DTN protocols under heterogeneous mobility. IEEE/ACM Transactions on Networking, 23 (2):587–602, April 2015. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Pejhan:1996:ECU

[PSA96] Sassan Pejhan, Mischa Schwartz, and Dimitris Anastassiou. Error control using retransmission schemes in multicast transport protocols for realtime media. IEEE/ ACM Transactions on Networking, 4(3):413–427, June 1996. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1996-4-3/
p413-pejhan/.

Papageorgiou:2015:DRR

[PSK^{+}15] George Papageorgiou, Shailendra Singh, Srikanth V. Krishnamurthy, Ramesh Govindan, and Tom La Porta. A distortionresistant routing framework for video traffic in wireless multihop networks. IEEE/ACM Transactions on Networking, 23(2):412– 425, April 2015. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Poupko:2021:BSR

[PSST21] Ouri Poupko, Gal Shahaf, Ehud Shapiro, and Nimrod Talmon. Building a Sybil-resilient digital community utilizing trustgraph connectivity. IEEE/ ACM Transactions on Networking, 29(5):2215–2227, October 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2021.3084303.

Peha:1994:AFT

[PT94] Jon M. Peha and Fouad A.

Tobagi. Analyzing the fault tolerance of doubleloop networks. IEEE/ ACM Transactions on Networking, 2(4):363–373, August 1994. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1994-2-4/
p363-peha/.

Papavassiliou:1996:JOC

[PT96] Symeon Papavassiliou and Leandros Tassiulas. Joint optimal channel base station and power assignment for wireless access. IEEE/ ACM Transactions on Networking, 4(6):857–872, December 1996. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1996-4-6/
p857-papavassiliou/.

Paschalidis:2000:CDP

[PT00] Ioannis Ch. Paschalidis and John N. Tsitsiklis. Congestion-dependent pricing of network services. IEEE/ACM Transactions on Networking, 8 (2):171–184, 2000. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/2000-8-2/
p171-paschalidis/.

Pong:2010:SSS

[PT10] Fong Pong and Nian-Feng Tzeng. SUSE: superior storage-efficiency for routing tables through prefix transformation and aggregation. IEEE/ACM Transactions on Networking, 18 (1):81–94, February 2010. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Pong:2012:CLT

[PT12] Fong Pong and Nian-Feng Tzeng. Concise lookup tables for IPv4 and IPv6 longest prefix matching in scalable routers. IEEE/ACM Transactions on Networking, 20(3):729– 741, June 2012. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Paschalidis:2004:ISE

[PV04] Ioannis Ch. Paschalidis and Spyridon Vassilaras. Importance sampling for the estimation of buffer overflow probabilities via tracedriven simulations. IEEE/ ACM Transactions on Networking, 12(5):907–919, October 2004. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Park:2010:MAC

[PV10] Jaeok Park and Mihaela Van Der Schaar. Medium access control protocols with memory. IEEE/ACM Transactions on Networking, 18(6):1921–1934, December 2010. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Paganini:2005:CCH

[PWDL05] Fernando Paganini, Zhikui Wang, John C. Doyle, and Steven H. Low. Congestion control for high performance, stability, and fairness in general networks. IEEE/ACM Transactions on Networking, 13 (1):43–56, February 2005. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Peng:2016:MTA

[PWHL16] Qiuyu Peng, Anwar Walid, Jaehyun Hwang, and Steven H. Low. Multipath TCP: analysis, design, and implementation. IEEE/ACM Transactions on Networking, 24(1):596–609, February 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Premnath:2013:BOB

[PWK^{+}13] Sriram N. Premnath, Daryl Wasden, Sneha K. Kasera, Neal Patwari, and Behrouz Farhang-Boroujeny. Beyond OFDM: best-effort dynamic spectrum access using filterbank multicarrier. IEEE/ACM Transactions on Networking, 21(3): 869–882, June 2013. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Pang:2022:TPP

[PWL^{+}22] Xiaoyi Pang, Zhibo Wang, Defang Liu, John C. S. Lui, Qian Wang, and Ju Ren. Towards personalized privacy-preserving truth discovery over crowdsourced data streams. IEEE/ACM Transactions on Networking, 30(1):327– 340, February 2022. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL https:/

/dl.acm.org/doi/10.1109/
TNET.2021.3110052.

Pan:2024:BMA

[PWL^{+}24] Qianqian Pan, Jun Wu, Jianhua Li, Wu Yang, and Mohsen Guizani. Blockchain and multi-agent learning empowered incentive IRS resource scheduling for intelligent reconfigurable networks. IEEE/ ACM Transactions on Networking, 32(2):943–958, April 2024. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2023.3309729.

Ping:2023:UNL

[PWLC23] Haodi Ping, Yongcai Wang, Deying Li, and Wenping Chen. Understanding node localizability in barycentric linear localization. IEEE/ ACM Transactions on Networking, 31(3):1353–1368, June 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2022.3216204.

Parvez:2012:IMS

[PWMC12] Khandoker Nadim Parvez, Carey Williamson, Anirban Mahanti, and Niklas Carlsson. Insights on media streaming progress using BitTorrent-like protocols for on-demand streaming. IEEE/ACM Transactions on Networking, 20(3): 637–650, June 2012. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Pokhrel:2018:MCT

[PWWP18] Shiva Raj Pokhrel, Carey Williamson, Carey Williamson, and Shiva Raj Pokhrel. Modeling compound TCP over WiFi for IoT. IEEE/ ACM Transactions on Networking, 26(2):864–878, April 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Park:1999:DSR

[PYL99] Jae-Hyun Park, Hyunsoo Yoon, and Heung-Kyu Lee. The deflection self-routing Banyan network: a large-scale ATM switch using the fully adaptive self-routing and its performance analyses. IEEE/ACM Transactions on Networking, 7(4):588– 604, August 1999. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1999-7-4/
p588-park/.

Pan:2017:PED

[PYL^{+}17] Xiaodan Pan, Tong Ye, Tony T. Lee, Weisheng Hu, Xiaodan Pan, Tong Ye, Tony T. Lee, and Weisheng Hu. Power efficiency and delay tradeoff of 10GBase-T energy efficient Ethernet protocol. IEEE/ ACM Transactions on Networking, 25(5):2773–2787, October 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Parsa:1998:IAD

[PZGLA98] Mehrdad Parsa, Qing Zhu, and J. J. Garcia-Luna-Aceves. An iterative algorithm for delayconstrained minimum-cost multicasting. IEEE/ACM Transactions on Networking, 6(4):461–474, August 1998. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1998-6-4/
p461-parsa/.

Pan:2016:TZT

[PZS^{+}16] Tian Pan, Ting Zhang, Junxiao Shi, Yang Li, Linxiao Jin, Fuliang Li, Jiahai Yang, Beichuan Zhang, Xueren Yang, Mingui Zhang, Huichen Dai, and Bin Liu. Towards zerotime wakeup of line cards in power-aware routers. IEEE/ACM Transactions on Networking, 24(3):1448– 1461, June 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Qazi:2012:CCM

[QAZ12] Ihsan Ayyub Qazi, Lachlan L. H. Andrew, and Taieb Znati. Congestion control with multipacket feedback. IEEE/ACM Transactions on Networking, 20(6): 1721–1733, December 2012. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Qiao:2016:TOP

[QCLC16] Yan Qiao, Shigang Chen, Tao Li, and Shiping Chen. Tag-ordering polling protocols in RFID systems. IEEE/ACM Transactions on Networking, 24(3):1548– 1561, June 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Qiao:2016:WBF

[QCMY16] Yan Qiao, Shigang Chen, Zhen Mo, and Myungkeun Yoon. When Bloom filters are no longer compact: Multi-set membership lookup for network applications. IEEE/ACM Transactions on Networking, 24(6):3326–3339, December 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Qiao:2007:IAT

[QCS07] Daji Qiao, Sunghyun Choi, and Kang G. Shin. Interference analysis and transmit power control in IEEE 802.11a/h wireless LANs. IEEE/ACM Transactions on Networking, 15(5):1007– 1020, October 2007. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Qu:2017:MMC

[QDD^{+}17] Yuben Qu, Chao Dong, Haipeng Dai, Fan Wu, Shaojie Tang, Hai Wang, and Chang Tian. Multicast in multihop CRNs under uncertain spectrum availability: a network coding approach. IEEE/ ACM Transactions on Networking, 25(4):2026–2039, August 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Qian:2018:GRR

[QFH^{+}18] Zhemin Qian, Fujie Fan, Bing Hu, Kwan L. Yeung, and Liyan Li. Global round robin: Efficient routing with cut-through switching in fat-tree data center networks. IEEE/ ACM Transactions on Networking, 26(5):2230–2241, October 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Qiu:2011:LRM

[QGCL11] Jian Qiu, Mohan Gurusamy, Kee Chaing Chua, and Yong Liu. Local restoration with multiple spanning trees in metro Ethernet networks. IEEE/ACM Transactions on Networking, 19(2):602– 614, April 2011. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Qian:2018:HMM

[QHZC18] Jiangbo Qian, Zhipeng Huang, Qiang Zhu, and Huahui Chen. Hamming metric multi-granularity locality-sensitive Bloom filter. IEEE/ACM Transactions on Networking, 26(4): 1660–1673, August 2018.

CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Qiao:2020:RLI

[QJCR20] Yan Qiao, Jun Jiao, Xinhong Cui, and Yuan Rao. Robust loss inference in the presence of noisy measurements and hidden fault diagnosis. IEEE/ACM Transactions on Networking, 28(1):43–56, February 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL https://

                                        dl.acm.org/doi/abs/10.
                                        1109/TNET.2019.2948818.

Qin:2016:ISR

[QJZ^{+}16] Yi Qin, Riheng Jia, Jinbei Zhang, Weijie Wu, and Xinbing Wang. Impact of social relation and group size in multicast ad hoc networks. IEEE/ ACM Transactions on Networking, 24(4):1989–2004, August 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Qiu:2001:MBA

[QK01] Jingyu Qiu and Edward W. Knightly. Measurementbased admission control with aggregate traffic envelopes. IEEE/ACM Transactions on Networking, 9(2):199–210, 2001. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL

http://www.acm.org/pubs/
articles/journals/ton/
2001-9-2/p199-qiu/p199-
qiu.pdf;   http://www.
acm.org/pubs/citations/
journals/ton/2001-9-2/
p199-qiu/.

Qian:2016:GRN

[QL16a] Chen Qian and Simon S.

Lam. Greedy routing by network distance embedding. IEEE/ACM Transactions on Networking, 24(4): 2100–2113, August 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Qian:2016:SRL

[QL16b] Chen Qian and Simon S.

Lam. A scalable and resilient layer-2 network with Ethernet compatibility. IEEE/ACM Transactions on Networking, 24(1): 231–244, February 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Qiu:2023:OBS

[QLF23] Tianyou Qiu, Yiping Li, and Xisheng Feng. Optimal broadcast scheduling algorithm for a multi-AUV acoustic communication network. IEEE/ ACM Transactions on Networking, 31(5):2058– 2069, 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2022.3232956.

Li:1993:QRIb

[qLH93a] San qi Li and Chia-Lin Hwang. Queue response to input correlation functions: continuous spectral analysis. IEEE/ ACM Transactions on Networking, 1(6):678–692, December 1993. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/
                                        journals/ton/1993-1-6/
                                        p678-li/.

Li:1993:QRIa

[qLH93b] San qi Li and Chia-Lin Hwang. Queue response to input correlation functions: discrete spectral analysis. IEEE/ ACM Transactions on Networking, 1(5):522–533, October 1993. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/
                                        journals/ton/1993-1-5/
                                        p522-li/.

Li:1997:CTM

[qLlH97] San qi Li and Chia lin Hwang. On the convergence of traffic measurement and queueing analysis: a statisticalmatching and queueing (SMAQ) tool. IEEE/ ACM Transactions on Networking, 5(1):95–110, February 1997. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1997-5-1/
p95-li/.

Li:1997:LLF

[qLP97] San qi Li and James D.

Pruneski. The linearity of low frequency traffic flow: an intrinsic I/O property in queueing systems. IEEE/ACM Transactions on Networking, 5(3):429– 443, June 1997. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1997-5-3/
p429-li/.

Qu:2022:ELC

[QLQ^{+}22] Dapeng Qu, Guoxin Lv, Shijun Qu, Haiying Shen, Yue Yang, and Zhaoyang Heng. An effective and lightweight countermeasure scheme to multiple network attacks in NDN. IEEE/ACM Transactions on Networking, 30(2):515– 528, April 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.

                                        acm.org/doi/10.1109/TNET.
                                        2021.3121001.

Qiu:2019:ROS

[QLSW19] Tie Qiu, Jie Liu, Weisheng Si, and Dapeng Oliver Wu. Robustness optimization scheme with multipopulation co-evolution for scale-free wireless sensor networks. IEEE/ACM Transactions on Networking, 27(3):1028–1042, June 2019. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Qin:2023:EDT

[QLY23] Xudong Qin, Bin Li, and Lei Ying. Efficient distributed threshold-based offloading for large-scale mobile cloud computing. IEEE/ACM Transactions on Networking, 31(1):308– 321, 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2022.3193073.

Qiao:1999:LPE

[QM99] Chunming Qiao and Yousong Mei. Off-line permutation embedding and scheduling in multiplexed optical networks with regular topologies. IEEE/ ACM Transactions on Networking, 7(2):241–250, April 1999. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1999-7-2/
p241-qiao/.

Qu:2024:SSI

[QML^{+}24] Jian Qu, Xiaobo Ma, Wenmao Liu, Hongqing Sang, Jianfeng Li, Lei Xue, Xiapu Luo, Zhenhua Li, Li Feng, and Xiaohong Guan. On smartly scanning of the Internet of things. IEEE/ACM Transactions on Networking, 32(2):1019– 1034, April 2024. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2023.3312162.

Qiu:2004:PFC

[QS04] Dongyu Qiu and Ness B.

Shroff. A predictive flow control scheme for efficient network utilization and QoS. IEEE/ACM Transactions on Networking, 12(1): 161–172, February 2004. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Qiu:2005:QPF

[QS05] Dongyu Qiu and Ness B.

Shroff. Queueing properties of feedback flow control systems. IEEE/ACM Transactions on Networking, 13 (1):57–68, February 2005. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Qiu:2015:CLL

[QSS^{+}15] Chenxi Qiu, Haiying Shen, Sohraab Soltani, Karan Sapra, Hao Jiang, and Jason O. Hallstrom. CEDAR: a low-latency and distributed strategy for packet recovery in wireless networks. IEEE/ACM Transactions on Networking, 23(5):1514–1527, October 2015. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Qi:2024:TDB

[QSW24] Jianpeng Qi, Xiao Su, and Rui Wang. Toward distributively build timesensitive-service coverage in compute first networking. IEEE/ACM Transactions on Networking, 32(1): 582–597, 2024. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2023.3289830.

Quan:2020:CCO

[QTE20] Guocong Quan, Jian Tan, and Atilla Eryilmaz. Counterintuitive characteristics of optimal distributed LRU caching over unreliable channels. IEEE/ACM Transactions on Networking, 28(6):2461–2474, December 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2020.3015474.

Qin:2016:MWD

[QTWW16] Yi Qin, Xiaohua Tian, Weijie Wu, and Xinbing Wang. Mobility weakens the distinction between multicast and unicast. IEEE/ACM Transactions on Networking, 24(3):1350–1363, June 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Qi:2023:RDR

[QW23] Jianpeng Qi and Rui Wang.

R2: a distributed remote function execution mechanism with builtin metadata. IEEE/ ACM Transactions on Networking, 31(2):710–723, April 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2022.3198467.

Qiao:2021:BQD

[QWL21] Chunyu Qiao, Jiliang Wang, and Yunhao Liu. Beyond QoE: Diversity adaptation in video streaming at the edge. IEEE/ ACM Transactions on Networking, 29(1):289–302, February 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2020.3032416.

Qin:2004:MCC

[QY04] Xiangdong Qin and Yuanyuan Yang. Multicast connection capacity of WDM switching networks with limited wavelength conversion. IEEE/ACM Transactions on Networking, 12(3):526– 538, June 2004. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Qin:2012:SST

[QY12] Yang Qin and Lie-Liang Yang. Steady-state throughput analysis of network coding nodes employing stop-and-wait automatic repeat request. IEEE/ ACM Transactions on Networking, 20(5):1402–1411, October 2012. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Qiu:2024:SAR

[QYC^{+}24] Tie Qiu, Xinwei Yang, Ning Chen, Songwei Zhang, Geyong Min, and Dapeng Oliver Wu. A self-adaptive robustness optimization method with evolutionary multiagent for IoT topology. IEEE/ACM Transactions on Networking, 32(2):1346– 1361, April 2024. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2023.3319499.

Qiu:2006:SRI

[QYZS06] Lili Qiu, Yang Richard Yang, Yin Zhang, and Scott Shenker. On selfish routing in Internet-like environments. IEEE/ACM Transactions on Networking, 14(4):725–738, August 2006. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Qi:2022:ETC

[QYZX22] Yiwen Qi, Wenke Yu, Xudong Zhao, and Xindi Xu. Event-triggered control for network-based switched systems with switching signals subject to dualterminal DoS attacks. IEEE/ACM Transactions on Networking, 30(3):1283– 1293, June 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2021.3135963.

Qiu:2022:BWS

[QZC^{+}22] Tie Qiu, Lidi Zhang, Ning Chen, Songwei Zhang, Wenyuan Liu, and Dapeng Oliver Wu. Born this way: a self-organizing evolution scheme with motif for Internet of Things robustness. IEEE/ACM Transactions on Networking, 30(6): 2644–2657, 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2022.3178408.

Qi:2016:SID

[QZL^{+}16] Saiyu Qi, Yuanqing Zheng, Mo Li, Yunhao Liu, and Jinli Qiu. Scalable industry data access control in RFID-enabled supply chain. IEEE/ACM Transactions on Networking, 24(6):3551–3564, December 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Qiu:2017:RRS

[QZX^{+}17] Tie Qiu, Aoyang Zhao, Feng Xia, Weisheng Si, Dapeng Oliver Wu, Tie Qiu, Aoyang Zhao, Feng Xia, Weisheng Si, and Dapeng Oliver Wu. ROSE: Robustness strategy for scale-free wireless sensor networks. IEEE/ACM Transactions on Networking, 25(5):2944–2959, October 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Qian:2013:DCA

[QZZ^{+}13] Dajun Qian, Dong Zheng, Junshan Zhang, Ness B. Shroff, and Changhee Joo.

Distributed CSMA algorithms for link scheduling in multihop MIMO networks under SINR model. IEEE/ACM Transactions on Networking, 21(3):746– 759, June 2013. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Rouskas:1995:AOT

[RA95] George N. Rouskas and Mostafa H. Ammar. Analysis and optimization of transmission schedules for single-hop WDM networks. IEEE/ACM Transactions on Networking, 3(2):211– 221, April 1995. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1995-3-2/
p211-rouskas/.

Reviriego:2024:CEA

[RAA^{+}24] Pedro Reviriego, Jim Apple, Alvaro Alonso, Otmar Ertl, and Niv Dayan. Cardinality estimation adaptive cuckoo filters (CE-ACF): Approximate membership check and distinct query count for highspeed network monitoring. IEEE/ACM Transactions on Networking, 32(2):959– 970, April 2024. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.

                                        acm.org/doi/10.1109/TNET.
                                        2023.3302306.

Ranjan:2004:NIT

[RAL04] Priya Ranjan, Eyad H. Abed, and Richard J. La. Nonlinear instabilities in TCP-RED. IEEE/ACM Transactions on Networking, 12(6):1079–1092, December 2004. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Ramanathan:1996:MTG

[Ram96] S. Ramanathan. Multicast tree generation in networks with asymmetric links. IEEE/ACM Transactions on Networking, 4(4):558– 568, August 1996. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/
                                        journals/ton/1996-4-4/
                                        p558-ramanathan/.

Ramasubramanian:2008:SMP

[Ram08] Srinivasan Ramasubramanian. Supporting multiple protection strategies in optical networks. IEEE/ ACM Transactions on Networking, 16(6):1352–1365, December 2008. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Rene:2022:CCF

[RAPP22] Sergi Rene, Onur Ascigil, Ioannis Psaras, and George Pavlou. A congestion control framework based on innetwork resource pooling. IEEE/ACM Transactions on Networking, 30(2):683– 697, April 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2021.3128384.

Reibman:1995:TDV

[RB95] Amy R. Reibman and Arthur W. Berger. Traffic descriptors for VBR video teleconferencing over ATM networks. IEEE/ ACM Transactions on Networking, 3(3):329–339, June 1995. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1995-3-3/
p329-reibman/.

Rodriguez:2002:DPA

[RB02] Pablo Rodriguez and Ernst W.

Biersack. Dynamic parallel access to replicated content in the Internet. IEEE/ACM Transactions on Networking, 10(4):455– 465, August 2002. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Ross:2009:PCS

[RB09a] Kevin Ross and Nicholas Bambos. Projective cone scheduling (PCS) algorithms for packet switches of maximal throughput. IEEE/ACM Transactions on Networking, 17(3):976– 989, June 2009. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Rouskas:2009:BTS

[RB09b] George N. Rouskas and Nikhil Baradwaj. On bandwidth tiered service. IEEE/ ACM Transactions on Networking, 17(6):1780–1793, December 2009. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Retvari:2007:SPR

[RBC07] Gábor Rétvári, József J. Bíró, and Tibor Cinkler. On shortest path representation. IEEE/ACM Transactions on Networking, 15(6):1293–1306, December 2007. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Rastogi:2003:OCO

[RBGK03] Rajeev Rastogi, Yuri Breitbart, Minos Garofalakis, and Amit Kumar. Optimal configuration of OSPF aggregates. IEEE/ACM Transactions on Networking, 11(2):181–194, April 2003. CODEN IEANEP.

ISSN 1063-6692 (print), 1558-2566 (electronic).

Reddyvari:2021:MSS

[RBPS21] Vamseedhar Reddyvari, Sarat Chandra Bobbili, Parimal Parag, and Srinivas Shakkottai. Modesuppression: a simple, stable and scalable chunksharing algorithm for P2P networks. IEEE/ACM Transactions on Networking, 29(6):2548–2559, December 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2021.3092008.

Raman:2002:IIT

[RBS02] Suchitra Raman, Hari Balakrishnan, and Murari Srinivasan. ITP: an Image Transport Protocol for the Internet. IEEE/ACM Transactions on Networking, 10(3):297–307, June 2002. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Ramasubramanian:2008:DLF

[RC08] Srinivasan Ramasubramanian and Amit Chandak. Dual-link failure resiliency through backup link mutual exclusion. IEEE/ ACM Transactions on Networking, 16(1):157–169, February 2008. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Rossi:2015:PHM

[RCFC15] Lorenzo Rossi, Jacob Chakareski, Pascal Frossard, and Stefania Colonnese. A Poisson hidden Markov model for multiview video traffic. IEEE/ACM Transactions on Networking, 23(2):547– 558, April 2015. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Raman:2009:FLA

[RCGS09] Bhaskaran Raman, Kameswari Chebrolu, Dattatraya Gokhale, and Sayandeep Sen. On the feasibility of the link abstraction in wireless mesh networks. IEEE/ACM Transactions on Networking, 17(2):528–541, April 2009. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Rosenblum:2006:AFS

[RCGT06] Michael Rosenblum, Constantine Caramanis, Michel X. Goemans, and Vahid Tarokh. Approximating fluid schedules in crossbar packetswitches and Banyan networks. IEEE/ACM Transactions on Networking, 14 (6):1374–1387, December 2006. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Rojas-Cessa:2003:CFD

[RCOC03] Roberto Rojas-Cessa, Eiji Oki, and H. Jonathan Chao. Concurrent fault detection for a multiple-plane packet switch. IEEE/ACM Transactions on Networking, 11(4):616–627, August 2003. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Rager:2018:SSQ

[RCR^{+}18] Scott T. Rager, Ertugrul N. Ciftcioglu, Ram Ramanathan, Thomas F. La Porta, and Ramesh Govindan. Scalability and satisfiability of quality-of-information in wireless networks. IEEE/ACM Transactions on Networking, 26(1):398–411, February 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Ramaswamy:2014:WPM

[RCS14] Vinod Ramaswamy, Diganto Choudhury, and Srinivas Shakkottai. Which protocol? Mutual interaction of heterogeneous congestion controllers. IEEE/ ACM Transactions on Networking, 22(2):457–469, April 2014. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Rabbachin:2015:WNI

[RCW15] Alberto Rabbachin, Andrea Conti, and Moe Z. Win. Wireless network intrinsic secrecy. IEEE/ACM Transactions on Networking, 23(1):56–69, February 2015. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Rengarajan:2011:AAE

[RD11a] Balaji Rengarajan and Gustavo De Veciana. Architecture and abstractions for environment and trafficaware system-level coordination of wireless networks. IEEE/ACM Transactions on Networking, 19(3):721– 734, June 2011. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Rengarajan:2011:PAU

[RD11b] Balaji Rengarajan and Gustavo De Veciana. Practical adaptive user association policies for wireless systems with dynamic interference. IEEE/ACM Transactions on Networking, 19(6):1690–1703, December 2011. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Rai:2007:RMP

[RDO^{+}07] Smita Rai, Omkar Deshpande, Canhui Ou, Charles U.

Martel, and Biswanath Mukherjee. Reliable multipath provisioning for highcapacity backbone mesh networks. IEEE/ACM Transactions on Networking, 15(4):803–812, August 2007. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Rottenstreich:2017:PPS

[RDR17] Ori Rottenstreich, Mario Di Francesco, and Yoram Revah. Perfectly periodic scheduling of collective data streams. IEEE/ACM Transactions on Networking, 25(3):1332–1346, June 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Roy:2019:ULP

[RDZ^{+}19] Arjun Roy, Rajdeep Das, Hongyi Zeng, Jasmeet Bagga, and Alex C. Snoeren. Understanding the limits of passive realtime datacenter fault detection and localization. IEEE/ ACM Transactions on Networking, 27(5):2001–2014, October 2019. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Rad:2017:DNT

[REM17] Neshat Etemadi Rad, Yariv Ephraim, and Brian L. Mark. Delay network tomography using a partially observable bivariate Markov chain. IEEE/ ACM Transactions on Networking, 25(1):126–138, February 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Rea:2017:FNT

[RFGL17] Maurizio Rea, Aymen Fakhreddine, Domenico Giustiniano, and Vincent Lenders. Filtering noisy 802.11 time-of-flight ranging measurements from commoditized WiFi radios. IEEE/ACM Transactions on Networking, 25(4):2514– 2527, August 2017. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Ravindran:1998:CAM

[RG98] K. Ravindran and Ting-Jian Gong. Cost analysis of multicast transport architectures in multiservice networks. IEEE/ ACM Transactions on Networking, 6(1):94–109, February 1998. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/
                                        journals/ton/1998-6-1/
                                        p94-ravindran/.

Rajawat:2011:CLD

[RGG11] Ketan Rajawat, Nikolaos Gatsis, and Georgios B. Giannakis. Cross-layer designs in coded wireless fading networks with multicast. IEEE/ACM Transactions on Networking, 19(5): 1276–1289, October 2011. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Radunovic:2010:TPO

[RGKR10] Božidar Radunović, Christos Gkantsidis, Peter Key, and Pablo Rodriguez. Toward practical opportunistic routing with intrasession network coding for mesh networks. IEEE/ ACM Transactions on Networking, 18(2):420–433, April 2010. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Ray:2010:AAD

[RGKS10] Saikat Ray, Roch Guérin, Kin-Wah Kwong, and Rute Sofia. Always acyclic distributed path computation. IEEE/ACM Transactions on Networking, 18(1):307– 319, February 2010. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Ren:2022:ODS

[RGY^{+}22] Bangbang Ren, Deke Guo, Yali Yuan, Guoming Tang, Weijun Wang, and Xiaoming Fu. Optimal deployment of SRv6 to enable network interconnection service. IEEE/ ACM Transactions on Networking, 30(1):120–133, February 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2021.3105959.

Raza:2012:MFR

[RHC^{+}12] Saqib Raza, Guanyao Huang, Chen-Nee Chuah, Srini Seetharaman, and Jatinder Pal Singh. MeasuRouting: a framework for routing assisted traffic monitoring. IEEE/ACM Transactions on Networking, 20(1):45–56, February 2012. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Rahman:2016:DMF

[RHMF16] Sazzadur Rahman, Ting-Kai Huang, Harsha V. Madhyastha, and Michalis Faloutsos. Detecting malicious Facebook applications. IEEE/ACM Transactions on Networking, 24 (2):773–787, April 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Rozner:2013:MDO

[RHQZ13] Eric Rozner, Mi Kyung Han, Lili Qiu, and Yin Zhang. Model-driven optimization of opportunistic routing. IEEE/ACM Transactions on Networking, 21(2):594–609, April 2013. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Renani:2020:HHD

[RHX^{+}20] Alireza Ameli Renani, Jun Huang, Guoliang Xing, Abdol-Hossein Esfahanian, and Weiguo Wu. Harnessing hardware defects for improving wireless link performance. IEEE/ACM Transactions on Networking, 28(5):1913–1924, October 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2020.3003338.

Ramamurthy:1998:OAP

[RIM98] Byrav Ramamurthy, Jason Iness, and Biswanath Mukherjee. Optimizing amplifier placements in a multiwavelength optical LAN/MAN: the unequally powered wavelengths case. IEEE/ ACM Transactions on Networking, 6(6):755–767, December 1998. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1998-6-6/
p755-ramamurthy/.

Ramamoorthy:2006:SDS

[RJCE06] Aditya Ramamoorthy, Kamal Jain, Philip A. Chou, and Michelle Effros. Separating distributed source coding from network coding. IEEE/ACM Transactions on Networking, 14 (SI):2785–2795, June 2006. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Rangwala:2011:NCC

[RJJ^{+}11] Sumit Rangwala, Apoorva Jindal, Ki-Young Jang, Konstantinos Psounis, and Ramesh Govindan. Neighborhoodcentric congestion control for multihop wireless mesh networks. IEEE/ACM Transactions on Networking, 19(6):1797–1810, December 2011. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Ratnakar:2006:MCD

[RK06] Niranjan Ratnakar and Gerhard Kramer. The multicast capacity of deterministic relay networks with no interference. IEEE/ ACM Transactions on Networking, 14(SI):2425–2432, June 2006. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Rottenstreich:2015:BPW

[RK15] Ori Rottenstreich and Isaac Keslassy. The Bloom paradox: when not to use a Bloom filter. IEEE/ACM Transactions on Networking, 23(3):703–716, June 2015. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Ramaiyan:2008:FPA

[RKA08] Venkatesh Ramaiyan, Anurag Kumar, and Eitan Altman. Fixed point analysis of single cell IEEE 802.11e WLANs: uniqueness and multistability. IEEE/ACM Transactions on Networking, 16(5):1080–1093, October 2008. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Rottenstreich:2016:OTE

[RKH^{+}16] Ori Rottenstreich, Isaac Keslassy, Avinatan Hassidim, Haim Kaplan, and Ely Porat. Optimal in/ out TCAM encodings of ranges. IEEE/ACM Transactions on Networking, 24(1):555–568, February 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Rottenstreich:2014:VIC

[RKK14] Ori Rottenstreich, Yossi Kanizo, and Isaac Keslassy. The variable-increment counting Bloom filter. IEEE/ACM Transactions on Networking, 22(4):1092– 1105, August 2014. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Ruhrup:2010:MEB

[RKNS10] Stefan Rührup, Hanna Kalosha, Amiya Nayak, and Ivan Stojmenović. Message-efficient beaconless georouting with guaranteed delivery in wireless sensor, ad hoc, and actuator networks. IEEE/ACM Transactions on Networking, 18(1):95–108, February 2010. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Ruiz:2016:TNC

[RKPP16] Hamlet Medina Ruiz, Michel Kieffer, and Beatrice Pesquet-Popescu. TCP and network coding: Equilibrium and dynamic properties. IEEE/ACM Transactions on Networking, 24(4):1935–1947, August 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Rubenstein:2002:DSC

[RKT02a] Dan Rubenstein, Jim Kurose, and Don Towsley. Detecting shared congestion of flows via end-to-end measurement. IEEE/ACM Transactions on Networking, 10(3):381–395, June 2002. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Rubenstein:2002:IML

[RKT02b] Dan Rubenstein, Jim Kurose, and Don Towsley. The impact of multicast layering on network fairness. IEEE/ACM Transactions on Networking, 10 (2):169–182, April 2002. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Ramachandran:2010:SST

[RKZG10] Kishore Ramachandran, Ravi Kokku, Honghai Zhang, and Marco Gruteser. Symphony: synchronous two-phase rate and power control in 802.11 WLANs. IEEE/ACM Transactions on Networking, 18(4):1289– 1302, August 2010. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Ramanathan:1993:SAM

[RL93] Subramanian Ramanathan and Errol L. Lloyd. Scheduling algorithms for multihop radio networks. IEEE/ ACM Transactions on Networking, 1(2):166–177, April 1993. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1993-1-2/
p166-ramanathan/.

Rosenberg:1994:HFS

[RL94] Catherine Rosenberg and Bruno Laguë. A heuristic framework for source policing in ATM networks. IEEE/ACM Transactions on Networking, 2(4):387– 397, August 1994. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/
                                        journals/ton/1994-2-4/
                                        p387-rosenberg/.

Radunovic:2007:UFM

[RL07] Bozidar Radunović and Jean-Yves Le Boudec. A unified framework for maxmin and min-max fairness with applications. IEEE/ ACM Transactions on Networking, 15(5):1073–1083, October 2007. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Raviv:2018:MSR

[RL18] Li-On Raviv and Amir Leshem. Maximizing service reward for queues with deadlines. IEEE/ ACM Transactions on Networking, 26(5):2296–2308, October 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Ruzomberkaand:2023:IMH

[RL23] Eric Ruzomberkaand and David J. Love. Interference moral hazard in large multihop networks. IEEE/ACM Transactions on Networking, 31 (1):15–29, 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2022.3186234.

Ranjan:2006:GSC

[RLA06] Priya Ranjan, Richard J.

La, and Eyad H. Abed. Global stability conditions for rate control with arbitrary communication delays. IEEE/ACM Transactions on Networking, 14 (1):94–107, February 2006. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Ramjee:1998:PEC

[RLKT98] Ramachandran Ramjee, Thomas F. La Porta, Jim Kurose, and Don Towsley. Performance evaluation of connection rerouting schemes for ATM-based wireless networks. IEEE/ ACM Transactions on Networking, 6(3):249–261, June 1998. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1998-6-3/
p249-ramjee/.

Ray:2006:SLD

[RLP06] Saikat Ray, Wei Lai, and Ioannis Ch. Paschalidis. Statistical location detection with sensor networks. IEEE/ACM Transactions on Networking, 14(SI): 2670–2683, June 2006. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Roseti:2010:APE

[RLZ10] Cesare Roseti, Michele Luglio, and Francesco Zampognaro. Analysis and performance evaluation of a burst-based TCP for satellite DVB RCS links. IEEE/ACM Transactions on Networking, 18(3):911– 921, June 2010. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Ren:2018:DMD

[RLZ^{+}18] Xiaoqi Ren, Palma London, Juba Ziani, Adam Wierman, Palma London, Juba Ziani, Adam Wierman, and Xiaoqi Ren. Datum: Managing data purchasing and data placement in a geo-distributed data market. IEEE/ACM Transactions on Networking, 26(2):893–905, April 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Ramamurthy:2002:FAR

[RM02] Ramu Ramamurthy and Biswanath Mukherjee. Fixedalternate routing and wavelength conversion in wavelengthrouted optical networks. IEEE/ACM Transactions on Networking, 10(3):351– 367, June 2002. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Ramasubramanian:2008:BBR

[RM08] Venugopalan Ramasubramanian and Daniel Mossé. BRA: a bidirectional routing abstraction for asymmetric mobile ad hoc networks. IEEE/ACM Transactions on Networking, 16(1):116–129, February 2008. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Roy:2016:DQE

[RMDJ16] Nirmalya Roy, Archan Misra, Sajal K. Das, and Christine Julien. Determining quality- and energyaware multiple contexts in pervasive computing environments. IEEE/ACM Transactions on Networking, 24(5):3026–3042, October 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Raghavan:1999:RAC

[RMM99] Sriram Raghavan, G. Manimaran, and C. Siva Ram Murthy. A rearrangeable algorithm for the construction delay-constrained dynamic multicast trees. IEEE/ACM Transactions on Networking, 7(4):514– 529, August 1999. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/
                                        journals/ton/1999-7-4/
                                        p514-raghavan/.

Raina:2016:SPA

[RMPG16] Gaurav Raina, Sreelakshmi Manjunath, Sai Prasad, and Krishnamurthy Giridhar. Stability and performance analysis of compound TCP with REM and drop-tail queue management. IEEE/ACM Transactions on Networking, 24(4):1961–1974, August 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Rosberg:1996:CMA

[Ros96] Zvi Rosberg. Cell multiplexing in ATM networks. IEEE/ACM Transactions on Networking, 4(1):112– 122, February 1996. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL http://

                                        www.acm.org/pubs/citations/
                                        journals/ton/1996-4-1/
                                        p112-rosberg/.

Rosenberg:2005:HTN

[Ros05] Eric Rosenberg. Hierarchical topological network design. IEEE/ACM Transactions on Networking, 13(6): 1402–1409, December 2005. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Ramabhadran:2006:SRR

[RP06] Sriram Ramabhadran and Joseph Pasquale. The Stratified Round Robin scheduler: design, analysis and implementation. IEEE/ACM Transactions on Networking, 14(6):1362– 1373, December 2006. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Radunovic:2013:DCC

[RP13] Bozidar Radunovic and Alexandre Proutiere. On downlink capacity of cellular data networks with WLAN/WPAN relays. IEEE/ ACM Transactions on Networking, 21(1):286–296, February 2013. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Rosberg:2014:IJA

[RPF^{+}14] Zvi Rosberg, Yu Peng, Jing Fu, Jun Guo, Eric W. M. Wong, and Moshe Zukerman. Insensitive job assignment with throughput and energy criteria for processor-sharing server farms. IEEE/ACM Transactions on Networking, 22(4):1257–1270, August 2014. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Radoslavov:2004:CAL

[RPGE04] Pavlin Radoslavov, Christos Papadopoulos, Ramesh Govindan, and Deborah Estrin. A comparison of application-level and router-assisted hierarchical schemes for reliable multicast. IEEE/ACM Transactions on Networking, 12 (3):469–482, June 2004. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Rai:2017:LFB

[RpLP^{+}17] Anurag Rai, Chih ping Li, Georgios Paschos, Eytan Modiano, Anurag Rai, Chih ping Li, Georgios Paschos, and Eytan Modiano. Loop-free backpressure routing using linkreversal algorithms. IEEE/ ACM Transactions on Networking, 25(5):2988–3002, October 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Rao:2019:UIP

[RPP^{+}19] Vijay S. Rao, R. Venkatesha Prasad, T. V. Prabhakar, Chayan Sarkar, Madhusudan Koppal, and Ignas Niemegeers. Understanding and improving the performance of constructive interference using destructive interference in WSNs. IEEE/ACM Transactions on Networking, 27 (2):505–517, April 2019. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Reisizadeh:2022:CFR

[RPPA22] Amirhossein Reisizadeh, Saurav Prakash, Ramtin Pedarsani, and Amir Salman Avestimehr. CodedReduce: a fast and robust framework for gradient aggregation in distributed learning. IEEE/ACM Transactions on Networking, 30(1):148–161, February 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2021.3109097.

Ren:2013:ESS

[RPV13] Shaolei Ren, Jaeok Park, and Mihaela Van Der Schaar. Entry and spectrum sharing scheme selection in femtocell communications markets. IEEE/ ACM Transactions on Networking, 21(1):218–232, February 2013. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Roy:2009:PIO

[RPZ^{+}09] Sabyasachi Roy, Himabindu Pucha, Zheng Zhang, Y. Charlie Hu, and Lili Qiu. On the placement of infrastructure overlay nodes. IEEE/ACM Transactions on Networking, 17(4):1298– 1311, August 2009. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Ramanathan:1993:AFT

[RR93] Srinivas Ramanathan and P. Venkat Rangan. Adaptive feedback techniques for synchronized multimedia retrieval over integrated networks. IEEE/ ACM Transactions on Networking, 1(2):246–260, April 1993. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/
                                        journals/ton/1993-1-2/
                                        p246-ramanathan/.

Ramakrishnan:2019:CUA

[RR19a] S. Ramakrishnan and Venkatesh Ramaiyan. Completely uncoupled algorithms for network utility maximization. IEEE/ACM Transactions on Networking, 27(2):607–620, April 2019. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Reviriego:2019:TCB

[RR19b] Pedro Reviriego and Ori Rottenstreich. The tandem counting Bloom filter — it takes two counters to tango. IEEE/ACM Transactions on Networking, 27(6):2252–2265, December 2019. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL https://

dl.acm.org/doi/abs/10.
1109/TNET.2019.2944954.

Ribeiro:2006:MQA

[RRB06] Vinay J. Ribeiro, Rudolf H.

Riedi, and Richard G. Baraniuk. Multiscale queueing analysis. IEEE/ ACM Transactions on Networking, 14(5):1005–1018, October 2006. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

R:1994:CSF

[RrBG94] Allen R., J. r. Bonde, and Sumit Ghosh. A comparative study of fuzzy versus “fixed” thresholds for robust queue management in cell-switching networks. IEEE/ACM Transactions on Networking, 2(4):337– 344, August 1994. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1994-2-4/
p337-bonde/.

Ramachandran:2010:PCA

[RRG10] Madanagopal Ramachandran, N. Usha Rani, and Timothy A. Gonsalves. Path computation algorithms for dynamic service provisioning with protection and inverse multiplexing in SDH/SONET networks. IEEE/ACM Transactions on Networking, 18(5):1492–1504, October 2010. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Ramesh:1996:RMA

[RRK96] Sridhar Ramesh, Catherine Rosenberg, and Anurag Kumar. Revenue maximization in ATM networks using the CLP capability and buffer priority management. IEEE/ ACM Transactions on Networking, 4(6):941–950, December 1996. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1996-4-6/
p941-ramesh/.

Raghunath:2007:MBC

[RRK07] Satish Raghunath, K. K. Ramakrishnan, and Shivkumar Kalyanaraman. Measurementbased characterization of IP VPNs. IEEE/ACM Transactions on Networking, 15(6):1428–1441, December 2007. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Reisslein:2002:FGS

[RRR02] Martin Reisslein, Keith W.

Ross, and Srinivas Rajagopal. A framework for guaranteeing statistical QoS. IEEE/ACM Transactions on Networking, 10 (1):27–42, February 2002. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Ramaswamy:2014:MWN

[RRS^{+}14] Vinod Ramaswamy, Vinith Reddy, Srinivas Shakkottai, Alex Sprintson, and Natarajan Gautam. Multipath wireless network coding: an augmented potential game perspective. IEEE/ACM Transactions on Networking, 22(1):217– 229, February 2014. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Rashelbach:2022:CAP

[RRS22] Alon Rashelbach, Ori Rottenstreich, and Mark Silberstein. A computational approach to packet classification. IEEE/ ACM Transactions on Networking, 30(3):1073–1087, June 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.

acm.org/doi/10.1109/TNET.
2021.3131879.

Rashelbach:2023:SLA

[RRS23] Alon Rashelbach, Ori Rottenstreich, and Mark Silberstein. Scaling by learning: Accelerating Open vSwitch data path with neural networks. IEEE/ ACM Transactions on Networking, 31(3):1230–1243, June 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2022.3215143.

Ramaswami:1995:RWA

[RS95a] Rajiv Ramaswami and Kumar N. Sivarajan. Routing and wavelength assignment in all-optical networks. IEEE/ACM Transactions on Networking, 3(5):489–500, October 1995. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1995-3-5/
p489-ramaswami/.

Rumsewicz:1995:CSC

[RS95b] Michael P. Rumsewicz and Donald E. Smith. A comparison of SS7 congestion control options during mass call-in situations. IEEE/ACM Transactions on Networking, 3(1):1–9, February 1995. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1995-3-1/
p1-rumsewicz/.

Ramaswami:1997:DNC

[RS97a] Rajiv Ramaswami and Adrian Segall. Distributed network control for optical networks. IEEE/ ACM Transactions on Networking, 5(6):936–943, December 1997. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1997-5-6/
p936-ramaswami/.

Rouskas:1997:PSB

[RS97b] George N. Rouskas and Vijay Sivaraman. Packet scheduling in broadcast WDM networks with arbitrary transceiver tuning latencies. IEEE/ ACM Transactions on Networking, 5(3):359–370, June 1997. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1997-5-3/
p359-rouskas/.

Ramaswami:1998:MON

[RS98] Rajiv Ramaswami and Galen Sasaki. Multiwavelength optical networks with limited wavelength conversion. IEEE/ ACM Transactions on Networking, 6(6):744–754, December 1998. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/
                                        journals/ton/1998-6-6/
                                        p744-ramaswami/.

Reeves:2000:DAD

[RS00] Douglas S. Reeves and Hussein F. Salama. A distributed algorithm for delay-constrained unicast routing. IEEE/ACM Transactions on Networking, 8(2):239–250, 2000. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL

                                        http://www.acm.org/pubs/
                                        citations/journals/ton/
                                        2000-8-2/p239-reeves/.

Ramasubramanian:2004:AON

[RS04] Srinivasan Ramasubramanian and Arun K. Somani. Analysis of optical networks with heterogeneous grooming architectures. IEEE/ACM Transactions on Networking, 12 (5):931–943, October 2004. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Rubenstein:2005:CUP

[RS05] Dan Rubenstein and Sambit Sahu. Can unstructured P2P protocols survive flash crowds? IEEE/ACM Transactions on Networking, 13(3):501–512, June 2005. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Riedl:2007:ROI

[RS07] Anton Riedl and Dominic A. Schupke. Routing optimization in IP networks utilizing additive and concave link metrics. IEEE/ACM Transactions on Networking, 15(5):1136– 1148, October 2007. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Ramasubramanian:2008:MFC

[RS08] Srinivasan Ramasubramanian and Arun K. Somani. MICRON: a framework for connection establishment in optical networks. IEEE/ACM Transactions on Networking, 16 (2):473–485, April 2008. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Reiter:2012:CDP

[RS12] Miklós Reiter and Richard Steinberg. Congestiondependent pricing and forward contracts for complementary segments of a communication network. IEEE/ACM Transactions on Networking, 20(2):436– 449, April 2012. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Rost:2019:VNE

[RS19] Matthias Rost and Stefan Schmid. Virtual network embedding approximations: Leveraging randomized rounding. IEEE/ ACM Transactions on Networking, 27(5):2071–2084, October 2019. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Rost:2020:HIV

[RS20] Matthias Rost and Stefan Schmid. On the hardness and inapproximability of virtual network embeddings. IEEE/ ACM Transactions on Networking, 28(2):791–803, April 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL https://

                                        dl.acm.org/doi/abs/10.
                                        1109/TNET.2020.2975646.

Rozic:2021:OPC

[RS21] Ćiril Rožić and Galen Sasaki. Optical protection cost of loop free alternates on completely connected IP networks over optical rings. IEEE/ACM Transactions on Networking, 29(3):1116– 1127, June 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2021.3061515.

Rodriguez:2001:AWC

[RSB01] Pablo Rodriguez, Christian Spanner, and Ernst W. Biersack. Analysis of Web caching architectures: hierarchical and distributed caching. IEEE/ACM Transactions on Networking, 9(4):404–418, August 2001. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Rhee:2011:LWN

[RSH^{+}11] Injong Rhee, Minsu Shin, Seongik Hong, Kyunghan Lee, Seong Joon Kim, and Song Chong. On the Levywalk nature of human mobility. IEEE/ACM Transactions on Networking, 19 (3):630–643, June 2011. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Ruan:2023:GQA

[RSLL23] Na Ruan, Hanyi Sun, Zenan Lou, and Jie Li. A general quantitative analysis framework for attacks in blockchain. IEEE/ ACM Transactions on Networking, 31(2):664–679, April 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.

                                        acm.org/doi/10.1109/TNET.
                                        2022.3201493.

Rai:2009:PAO

[RSM09] Smita Rai, Ching-Fong Su, and Biswanath Mukherjee. On provisioning in all-optical networks: an impairment-aware approach. IEEE/ACM Transactions on Networking, 17 (6):1989–2001, December 2009. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Rezaei:2010:DRS

[RSR10] Behnam A. Rezaei, Nima Sarshar, and Vwani P. Roychowdhury. Distributed resource sharing in lowlatency wireless ad hoc networks. IEEE/ACM Transactions on Networking, 18(1):190–201, February 2010. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Rolando:2011:SSF

[RSR11] Pierluigi Rolando, Riccardo Sisto, and Fulvio Risso. SPAF: stateless FSA-based packet filters. IEEE/ACM Transactions on Networking, 19(1):14– 27, February 2011. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Rasti:2009:PEG

[RSS09] Mehdi Rasti, Ahmad R. Sharafat, and Babak Seyfe.

Pareto-efficient and goaldriven power control in wireless networks: a gametheoretic approach with a novel pricing scheme. IEEE/ACM Transactions on Networking, 17(2):556– 569, April 2009. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Richa:2013:EFM

[RSSZ13] Andréa Richa, Christian Scheideler, Stefan Schmid, and Jin Zhang. An efficient and fair MAC protocol robust to reactive interference. IEEE/ACM Transactions on Networking, 21(3): 760–771, June 2013. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Ranjan:2009:DSD

[RSU^{+}09] Supranamaya Ranjan, Ram Swaminathan, Mustafa Uysal, Antonio Nucci, and Edward Knightly. DDoSshield: DDoS-resilient scheduling to counter application layer attacks. IEEE/ACM Transactions on Networking, 17(1):26– 39, February 2009. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Raz:2004:DCE

[RSZ04] Danny Raz, Yuval Shavitt, and Lixia Zhang. Distributed council election. IEEE/ACM Transactions on Networking, 12(3):483– 492, June 2004. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Rexford:1999:SVB

[RT99] Jennifer Rexford and Don Towsley. Smoothing variable-bit-rate video in an Internetwork. IEEE/ ACM Transactions on Networking, 7(2):202–215, April 1999. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/
                                        journals/ton/1999-7-2/
                                        p202-rexford/.

Rottenstreich:2017:ORC

[RT17] Ori Rottenstreich and Janos Tapolcai. Optimal rule caching and lossy compression for longest prefix matching. IEEE/ACM Transactions on Networking, 25(2):864–878, April 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Retvari:2016:CIF

[RTK^{+}16] Gábor Rétvári, János Tapolcai, Attila Kőrösi, András Majdán, and Zalán Heszberger. Compressing IP forwarding tables: towards entropy bounds and beyond. IEEE/ ACM Transactions on Networking, 24(1):149–162, February 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Ren:2017:CDP

[RTLC17] Runtian Ren, Xueyan Tang, Yusen Li, and Wentong Cai. Competitiveness of dynamic bin packing for online cloud server allocation. IEEE/ACM Transactions on Networking, 25 (3):1324–1331, June 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Ricardo:2021:CPD

[RTNS21] Guilherme Iecker Ricardo, Alina Tuholukova, Giovanni Neglia, and Thrasyvoulos Spyropoulos. Caching policies for delay minimization in small cell networks with coordinated multipoint joint transmissions. IEEE/ACM Transactions on Networking, 29(3):1105– 1115, June 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2021.3062269.

Rahman:2018:UAH

[RUH^{+}18] M. Saifur Rahman, Md.

Yusuf Sarwar Uddin, Tahmid Hasan, M. Sohel Rahman, and M. Kaykobad.

Using adaptive heartbeat rate on long-lived TCP connections. IEEE/ACM Transactions on Networking, 26(1):203–216, February 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Rumsewicz:1993:AES

[Rum93] Michael P. Rumsewicz. Analysis of the effects of SS7 message discard schemes on call completion rates during overload. IEEE/ACM Transactions on Networking, 1(4):491– 502, August 1993. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/
                                        journals/ton/1993-1-4/
                                        p491-rumsewicz/.

Rizzo:2000:RPP

[RV00] Luigi Rizzo and Lorenzo Vicisano. Replacement policies for a proxy cache. IEEE/ACM Transactions on Networking, 8(2):158– 170, 2000. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/
                                        journals/ton/2000-8-2/
                                        p158-rizzo/.

Rummukainen:2001:PCA

[RV01] Hannu Rummukainen and Jorma Virtamo. Polynomial cost approximations in Markov decision theory based call admission control. IEEE/ACM Transactions on Networking, 9(6): 769–779, December 2001. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Raman:2021:CSB

[RV21] Ravi Kiran Raman and Lav R. Varshney. Coding for scalable blockchains via dynamic distributed storage. IEEE/ACM Transactions on Networking, 29(6):2588–2601, December 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2021.3098613.

Roughan:2000:RTE

[RVA00] Matthew Roughan, Darryl Veitch, and Patrice Abry. Real-time estimation of the parameters of long-range dependence. IEEE/ACM Transactions on Networking, 8 (4):467–478, 2000. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/2000-8-4/
p467-roughan/.

Ridoux:2012:CFF

[RVB12] Julien Ridoux, Darryl Veitch, and Timothy Broomhead. The case for feedforward clock synchronization. IEEE/ACM Transactions on Networking, 20(1): 231–242, February 2012. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Rangan:1993:CAA

[RVR93] P. Venkat Rangan, Harrick M. Vin, and Srinivas Ramanathan. Communication architectures and algorithms for media mixing in multimedia conferences. IEEE/ACM Transactions on Networking, 1(1):20–30, February 1993. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/
                                        journals/ton/1993-1-1/
                                        p20-rangan/.

Ramjee:2002:HDB

[RVS^{+}02] Ramachandran Ramjee, Kannan Varadhan, Luca Salgarelli, Sandra R. Thuel, Shie-Yuan Wang, and Thomas La Porta. HAWAII: a domain-based approach for supporting mobility in wide-area wireless networks. IEEE/ACM Transactions on Networking, 10 (3):396–410, June 2002. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Raghavan:2009:SPC

[RVS09] Barath Raghavan, Patric Verkaik, and Alex C. Snoeren. Secure and policycompliant source routing. IEEE/ACM Transactions on Networking, 17(3):764– 777, June 2009. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Riboni:2015:OSD

[RVV^{+}15] Daniele Riboni, Antonio Villani, Domenico Vitali, Claudio Bettini, and Luigi V. Mancini. Obfuscation of sensitive data for incremental release of network flows. IEEE/ACM Transactions on Networking, 23(2):672–686, April 2015. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Ramaswami:1993:ALE

[RW93] V. Ramaswami and Jonathan L.

Wang. Analysis of the link error monitoring protocols in the common channel signaling network. IEEE/ACM Transactions on Networking, 1(1):31–47, February 1993. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1993-1-1/
p31-ramaswami/.

Rubin:1995:AMQ

[RW95] Izhak Rubin and James Chien-Hsing Wu. Analysis of an M/G/1/N queue with vacations and its iterative application to FDDI timed-token rings. IEEE/ ACM Transactions on Networking, 3(6):842–856, December 1995. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1995-3-6/
p842-rubin/.

Rubin:1996:PAD

[RW96] Izhak Rubin and Ho-Ting Wu. Performance analysis and design of CQBT algorithm for a ring network with spatial reuse. IEEE/ACM Transactions on Networking, 4(4):649– 659, August 1996. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1996-4-4/
p649-rubin/.

Rabinovich:2004:DEC

[RW04] Michael Rabinovich and Hua Wang. DHTTP: an efficient and cache-friendly transfer protocol for the Web. IEEE/ACM Transactions on Networking, 12(6): 1007–1020, December 2004. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Ramaswamy:2007:HSP

[RW07] Ramaswamy Ramaswamy and Tilman Wolf. Highspeed prefix-preserving IP address anonymization for passive measurement systems. IEEE/ACM Transactions on Networking, 15 (1):26–39, February 2007. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Rhee:2008:ZMH

[RWA^{+}08] Injong Rhee, Ajit Warrier, Mahesh Aia, Jeongki Min, and Mihail L. Sichitiu. Z-MAC: a hybrid MAC for wireless sensor networks. IEEE/ACM Transactions on Networking, 16(3):511– 524, June 2008. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Rong:2022:SMC

[RWL^{+}22] Chenghao Rong, Jessie Hui Wang, Juncai Liu, Jilong Wang, Fenghua Li, and Xiaolei Huang. Scheduling massive camera streams to optimize large-scale live video analytics. IEEE/ ACM Transactions on Networking, 30(2):867–880, April 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.

                                        acm.org/doi/10.1109/TNET.
                                        2021.3125359.

Rhee:2007:LEB

[RX07] Injong Rhee and Lisong Xu. Limitations of equation-based congestion control. IEEE/ACM Transactions on Networking, 15(4):852–865, August 2007. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Ryu:2012:TDR

[RYS12] Jung Ryu, Lei Ying, and Sanjay Shakkottai. Timescale decoupled routing and rate control in intermittently connected networks. IEEE/ACM Transactions on Networking, 20(4):1138–1151, August 2012. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Raza:2011:GNS

[RZC11] Saqib Raza, Yuanbo Zhu, and Chen-Nee Chuah. Graceful network state migrations. IEEE/ACM Transactions on Networking, 19(4):1097–1110, August 2011. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Ruby:2021:SEE

[RZE^{+}21] Rukhsana Ruby, Shuxin Zhong, Basem M. ElHalawany, Hanjiang Luo, and Kaishun Wu. SDN-Enabled energy-aware routing in underwater multi-modal communication networks. IEEE/ACM Transactions on Networking, 29(3):965– 978, June 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2021.3056772.

Ren:2014:TTD

[RZS14] Wei Ren, Qing Zhao, and Ananthram Swami. Temporal traffic dynamics improve the connectivity of ad hoc cognitive radio networks. IEEE/ACM Transactions on Networking, 22(1):124–136, February 2014. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Rosberg:2006:AON

[RZVZ06] Zvi Rosberg, Andrew Zalesky, Hai L. Vu, and Moshe Zukerman. Analysis of OBS networks with limited wavelength conversion. IEEE/ACM Transactions on Networking, 14(5): 1118–1127, October 2006. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Roughan:2012:STC

[RZWQ12] Matthew Roughan, Yin Zhang, Walter Willinger, and Lili Qiu. Spatiotemporal compressive sensing and Internet traffic matrices. IEEE/ACM Transactions on Networking, 20 (3):662–676, June 2012. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Rosberg:2006:PDO

[RZZ06] Zvi Rosberg, Andrew Zalesky, and Moshe Zukerman. Packet delay in optical circuit-switched networks. IEEE/ACM Transactions on Networking, 14 (2):341–354, April 2006. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Sharon:2001:EPM

[SA01a] Oran Sharon and Eitan Altman. An efficient polling MAC for wireless LANs. IEEE/ACM Transactions on Networking, 9 (4):439–451, August 2001. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Su:2001:JTS

[SA01b] Weilian Su and Ian F. Akyildiz. The jitter timestamp approach for clock recovery of real-time variable bit-rate traffic. IEEE/ ACM Transactions on Networking, 9(6):746–754, December 2001. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Sarac:2004:TSM

[SA04] Kamil Sarac and Kevin C.

Almeroth. Tracetree: a scalable mechanism to discover multicast tree topologies in the Internet. IEEE/ ACM Transactions on Networking, 12(5):795–808, October 2004. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Su:2005:TDS

[SA05] Weilian Su and Ian F. Akyildiz. Time-diffusion synchronization protocol for wireless sensor networks. IEEE/ACM Transactions on Networking, 13(2):384– 397, April 2005. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Saha:2021:OSP

[SA21] Gourav Saha and Alhussein A. Abouzeid. Optimal spectrum partitioning and licensing in tiered access under stochastic market models. IEEE/ ACM Transactions on Networking, 29(5):1948–1961, October 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2021.3077643.

Schuller:2018:TEU

[SAC^{+}18] Timmy Schuller, Nils Aschenbruck, Markus Chimani, Martin Horneffer, and Stefan Schnitter. Traffic engineering using segment routing and considering requirements of a carrier IP network. IEEE/ ACM Transactions on Networking, 26(4):1851–1864, August 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Schuller:2021:FRO

[SACH21] Timmy Schüller, Nils Aschenbruck, Markus Chimani, and Martin Horneffer. Failure resiliency with only a few tunnels — enabling segment routing for traffic engineering. IEEE/ ACM Transactions on Networking, 29(1):262–274, February 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2020.3030543.

Saha:2021:ODS

[SAKMB21] Gourav Saha, Alhussein A. Abouzeid, Zaheer Khan, and Marja Matinmikko-Blue. On the optimal duration of spectrum leases in exclusive license markets with stochastic demand. IEEE/ACM Transactions on Networking, 29(3):1060– 1073, June 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2021.3060088.

Singh:2013:OFD

[SAKS13] Chandramani Singh, Eitan Altman, Anurag Kumar, and Rajesh Sundaresan. Optimal forwarding in delay-tolerant networks with multiple destinations. IEEE/ACM Transactions on Networking, 21(6):1812– 1826, December 2013. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Shrimali:2010:CIT

[SAM10] Gireesh Shrimali, Aditya Akella, and Almir Mutapcic. Cooperative interdomain traffic engineering using Nash bargaining and decomposition. IEEE/ACM Transactions on Networking, 18(2):341– 352, April 2010. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Soldo:2012:OSB

[SAM12] Fabio Soldo, Katerina Argyraki, and Athina Markopoulou. Optimal source-based filtering of malicious traffic. IEEE/ ACM Transactions on Networking, 20(2):381–395, April 2012. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Saha:2018:OAL

[SAMB18] Gourav Saha, Alhussein A. Abouzeid, and Marja Matinmikko-Blue. Online algorithm for leasing wireless channels in a three-tier spectrum sharing framework. IEEE/ACM Transactions on Networking, 26 (6):2623–2636, December 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Saad:2024:RNC

[SARM24] Muhammad Saad, Afsah Anwar, Srivatsan Ravi, and David Mohaisen. Revisiting Nakamoto consensus in asynchronous networks: a comprehensive analysis of bitcoin safety and chain quality. IEEE/ACM Transactions on Networking, 32 (1):844–858, 2024. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2023.3302955.

Subramaniam:1996:ANS

[SAS96] Suresh Subramaniam, Murat Azizoğlu, and Arun K. Somani. All-optical networks with sparse wavelength conversion. IEEE/ ACM Transactions on Networking, 4(4):544–557, August 1996. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1996-4-4/
p544-subramaniam/.

Subramaniam:1999:OCP

[SAS99] Suresh Subramaniam, Murat Azizoğlu, and Arun K. Somani. On optimal converter placement in wavelength-routed networks. IEEE/ACM Transactions on Networking, 7 (5):754–766, October 1999. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL

http://www.acm.org/pubs/
citations/journals/ton/
1999-7-5/p754-subramaniam/

.

Santacruz:2016:LPL

[SAS16a] Pedro E. Santacruz, Vaneet Aggarwal, and Ashutosh Sabharwal. Leveraging physical-layer capabilities: distributed scheduling in interference networks with local views. IEEE/ACM Transactions on Networking, 24(1):368– 382, February 2016. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Schmid:2016:STL

[SAS^{+}16b] Stefan Schmid, Chen Avin, Christian Scheideler, Michael Borokhovich, Bernhard Haeupler, and Zvi Lotker. SplayNet: towards locally self-adjusting networks. IEEE/ACM Transactions on Networking, 24 (3):1421–1433, June 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Sundaresan:2016:FFC

[SAS^{+}16c] Karthikeyan Sundaresan, Mustafa Y. Arslan, Shailendra Singh, Sampath Rangarajan, and Srikanth V. Krishnamurthy. FluidNet: a flexible cloud-based radio access network for small cells. IEEE/ACM Transactions on Networking, 24 (2):915–928, April 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Stoica:2004:III

[SAZ^{+}04] Ion Stoica, Daniel Adkins, Shelley Zhuang, Scott Shenker, and Sonesh Surana. Internet indirection infrastructure. IEEE/ ACM Transactions on Networking, 12(2):205–218, April 2004. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Stauffer:2007:PHD

[SB07] Alexandre O. Stauffer and Valmir C. Barbosa. Probabilistic heuristics for disseminating information in networks. IEEE/ACM Transactions on Networking, 15(2):425–435, April 2007. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Sun:2017:STS

[SBC^{+}17] Chen Sun, Jun Bi, Haoxian Chen, Hongxin Hu, Zhilong Zheng, Shuyong Zhu, and Chenghui Wu. SDPA: Toward a stateful data plane in software-defined networking. IEEE/ACM Transactions on Networking, 25(6):3294–3308, December 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Sadiq:2011:DOO

[SBD11] Bilal Sadiq, Seung Jun Baek, and Gustavo De Veciana. Delay-optimal opportunistic scheduling and approximations: the log rule. IEEE/ACM Transactions on Networking, 19 (2):405–418, April 2011. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Sommers:2008:GAI

[SBDR08] Joel Sommers, Paul Barford, Nick Duffield, and Amos Ron. A geometric approach to improving active packet loss measurement. IEEE/ACM Transactions on Networking, 16 (2):307–320, April 2008. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Sommers:2010:MMS

[SBDR10] Joel Sommers, Paul Barford, Nick Duffield, and Amos Ron. Multiobjective monitoring for SLA compliance. IEEE/ACM Transactions on Networking, 18 (2):652–665, April 2010. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Swamy:2018:ECU

[SBGJ18] Peruru Subrahmanya Swamy, Venkata Pavan Kumar Bellam, Radha Krishna Ganti, and Krishna Jagannathan. Efficient CSMA using regional free energy approximations. IEEE/ ACM Transactions on Networking, 26(4):1796–1809, August 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Stimpfling:2019:SSH

[SBLS19] Thibaut Stimpfling, Normand Belanger, J. M. Pierre Langlois, and Yvon Savaria. SHIP: a scalable highperformance IPv6 lookup algorithm that exploits prefix characteristics. IEEE/ ACM Transactions on Networking, 27(4):1529–1542, August 2019. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Soule:2018:MLM

[SBM^{+}18] Robert Soule, Shrutarshi Basu, Parisa Jalili Marandi, Fernando Pedone, Robert Kleinberg, Emin Gun Sirer, and Nate Foster. Merlin: a language for managing network resources. IEEE/ ACM Transactions on Networking, 26(5):2188–2201, October 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Seibert:2014:NSV

[SBNRS14] Jeff Seibert, Sheila Becker, Cristina Nita-Rotaru, and Radu State. Newton: securing virtual coordinates by enforcing physical laws. IEEE/ACM Transactions on Networking, 22(3):798– 811, June 2014. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Sichitiu:2003:EUT

[SBP03] Mihail L. Sichitiu, Peter H.

Bauer, and Kamal Premaratne. The effect of uncertain time-variant delays in ATM networks with explicit rate feedback: a control theoretic approach. IEEE/ACM Transactions on Networking, 11(4):628– 637, August 2003. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Sun:2019:FHA

[SBTH19] Yahui Sun, Marcus Brazil, Doreen Thomas, and Saman Halgamuge. The fast heuristic algorithms and post-processing techniques to design large and low-cost communication networks. IEEE/ACM Transactions on Networking, 27(1):375– 388, February 2019. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Smith:1995:SPR

[SC95] Donald E. Smith and H. Jonathan Chao. Sizing a packet reassembly buffer at a host computer in an ATM network. IEEE/ ACM Transactions on Networking, 3(6):798–808, December 1995. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/
                                        journals/ton/1995-3-6/
                                        p798-smith/.

Sengupta:2009:EFD

[SC09] Shamik Sengupta and Mainak Chatterjee. An economic framework for dynamic spectrum access and service pricing. IEEE/ ACM Transactions on Networking, 17(4):1200–1213, August 2009. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Sue:2010:FRP

[SC10] Chuan-Ching Sue and Hsaing-Wen Cheng. A fitting report position scheme for the gated IPACT dynamic bandwidth algorithm in EPONs. IEEE/ ACM Transactions on Networking, 18(2):624–637, April 2010. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Shen:2017:DAV

[SC17] Haiying Shen and Liuhua Chen. Distributed autonomous virtual resource management in datacenters using finite-Markov decision process. IEEE/ACM Transactions on Networking, 25(6):3836–3849, December 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Shen:2018:CCV

[SC18a] Haiying Shen and Liuhua Chen. CompVM: a complementary VM allocation mechanism for cloud systems. IEEE/ACM Transactions on Networking, 26 (3):1348–1361, June 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Shen:2018:RDM

[SC18b] Haiying Shen and Liuhua Chen. Resource demand misalignment: an important factor to consider for reducing resource overprovisioning in cloud datacenters. IEEE/ACM Transactions on Networking, 26(3):1207–1221, June 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Shi:2022:DTS

[SC22a] Hehuan Shi and Lin Chen. Downlink transmission scheduling with data sharing. IEEE/ ACM Transactions on Networking, 30(3):1193–1202, June 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2021.3138940.

Shi:2022:SBC

[SC22b] Hehuan Shi and Lin Chen. From spectrum bonding to contiguous-resource batching task scheduling. IEEE/ ACM Transactions on Networking, 30(3):1245–1254, June 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2021.3138991.

Shi:2017:ECD

[SCC^{+}17] Tuo Shi, Siyao Cheng, Zhipeng Cai, Yingshu Li, and Jianzhong Li. Exploring connected dominating sets in energy harvest networks. IEEE/ACM Transactions on Networking, 25 (3):1803–1817, June 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Shao:2023:CCL

[SCH23] Qi Shao, Man Hon Cheung, and Jianwei Huang. Crowdfunding with cognitive limitations. IEEE/ ACM Transactions on Networking, 31(6):2714– 2729, 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2023.3274114.

Shao:2022:ABC

[SCHG22] Xiaozhe Shao, Zibin Chen, Daniel Holcomb, and Lixin Gao. Accelerating BGP configuration verification through reducing cycles in SMT constraints. IEEE/ ACM Transactions on Networking, 30(6):2493– 2504, 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2022.3176267.

Su:2009:DBA

[SCKB09] Ao-Jan Su, David R. Choffnes, Aleksandar Kuzmanovic, and Fabián E. Bustamante. Drafting behind Akamai: inferring network conditions based on CDN redirections. IEEE/ ACM Transactions on Networking, 17(6):1752–1765, December 2009. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Shi:2023:SMD

[SCL^{+}23] Tuo Shi, Zhipeng Cai, Jianzhong Li, Hong Gao, Jiancheng Chen, and Ming Yang. Services management and distributed multihop requests routing in mobile edge networks. IEEE/ACM Transactions on Networking, 31(2):497– 510, April 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2022.3196267.

Sen:2012:CCC

[SCN12] Souvik Sen, Romit Roy Choudhury, and Srihari Nelakuditi. CSMA/CN: carrier sense multiple access with collision notification. IEEE/ACM Transactions on Networking, 20 (2):544–556, April 2012. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Saad:2022:EPA

[SCN^{+}22] Muhammad Saad, Victor Cook, Lan Nguyen, My T. Thai, and David Mohaisen. Exploring partitioning attacks on the bitcoin network. IEEE/ ACM Transactions on Networking, 30(1):202–214, February 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2021.3105604.

Salem:2024:TID

[SCN^{+}24] Tareq Si Salem, Gabriele Castellano, Giovanni Neglia, Fabio Pianese, and Andrea Araldo. Toward inference delivery networks: Distributing machine learning with optimality guarantees. IEEE/ACM Transactions on Networking, 32 (1):859–873, 2024. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2023.3305922.

Sariowan:1999:SGS

[SCP99] Hanrijanto Sariowan, Rene L.

Cruz, and George C. Polyzos. SCED: a generalized scheduling policy for guaranteeing qualityof-service. IEEE/ACM Transactions on Networking, 7(5):669–684, October 1999. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/
                                        journals/ton/1999-7-5/
                                        p669-sariowan/.

Sciancalepore:2019:RNR

[SCPB19] Vincenzo Sciancalepore, Xavier Costa-Perez, and Albert Banchs. RL-NSB: Reinforcement learningbased 5G network slice broker. IEEE/ACM Transactions on Networking, 27(4): 1543–1557, August 2019. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Shin:2008:DRT

[SCR08] Minsu Shin, Song Chong, and Injong Rhee. Dualresource TCP/AQM for processing-constrained networks. IEEE/ACM Transactions on Networking, 16 (2):435–449, April 2008. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Shi:2022:LLF

[SCS^{+}22] Lei Shi, Yuhua Cheng, Jinliang Shao, Qingchen Liu, and Wei Xing Zheng. Locating link failures in WSNs via cluster consensus and graph decomposition. IEEE/ACM Transactions on Networking, 30(5): 2304–2314, 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2022.3171272.

Shi:2021:TMO

[SCW^{+}21] Xiaofeng Shi, Haofan Cai, Minmei Wang, Ge Wang, Baiwen Huang, Junjie Xie, and Chen Qian. TagAttention: Mobile object tracing with zero appearance knowledge by vision-RFID fusion. IEEE/ ACM Transactions on Networking, 29(2):890–903, April 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2021.3052805.

Shiomoto:1998:SBM

[SCY98] Kohei Shiomoto, Shinichiro Chaki, and Naoaki Yamanaka. A simple bandwidth management strategy based on measurements of instantaneous virtual path utilization in ATM networks. IEEE/ ACM Transactions on Networking, 6(5):625–634, October 1998. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1998-6-5/
p625-shiomoto/.

Syrotiuk:2008:RFE

[SCY08] Violet R. Syrotiuk, Charles J. Colbourn, and Sruthi Yellamraju. Rateless forward error correction for topology-transparent scheduling. IEEE/ACM Transactions on Networking, 16(2):464–472, April 2008. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Shu:2015:PML

[SCY15] Tao Shu, Yingying Chen, and Jie Yang. Protecting multi-lateral localization privacy in pervasive environments. IEEE/ACM Transactions on Networking, 23(5):1688–1701, October 2015. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Su:2000:SMM

[SD00] Ching-Fong Su and Gustavo De Veciana. Statistical multiplexing and mix-dependent alternative routing in multiservice VP networks. IEEE/ ACM Transactions on Networking, 8(1):99–108, February 2000. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/
                                        journals/ton/2000-8-1/
                                        p99-su/.

Shah:2015:HPC

[SD15a] Virag Shah and Gustavo De Veciana. Highperformance centralized content delivery infrastructure: models and asymptotics. IEEE/ACM Transactions on Networking, 23(5):1674–1687, October 2015. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Sun:2015:BDI

[SD15b] Xinghua Sun and Lin Dai.

Backoff design for IEEE 802.11 DCF networks: fundamental tradeoff and design criterion. IEEE/ ACM Transactions on Networking, 23(1):300–316, February 2015. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Santagati:2020:DPE

[SDM20] G. Enrico Santagati, Neil Dave, and Tommaso Melodia. Design and performance evaluation of an implantable ultrasonic networking platform for the Internet of Medical Things. IEEE/ACM Transactions on Networking, 28(1):29–42, February 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL https://

dl.acm.org/doi/abs/10.
1109/TNET.2019.2949805.

Satpathi:2019:LLE

[SDSY19] Siddhartha Satpathi, Supratim Deb, R. Srikant, and He Yan. Learning latent events from network message logs. IEEE/ ACM Transactions on Networking, 27(4):1728–1741, August 2019. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Shaikh:2006:AID

[SDV06] Aman Shaikh, Rohit Dube, and Anujan Varma. Avoiding instability during graceful shutdown of multiple OSPF routers. IEEE/ACM Transactions on Networking, 14(3):532–542, 2006. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Shah:2016:SAR

[SdVK16] Virag Shah, Gustavo de Veciana, and George Kesidis. A stable approach for routing queries in unstructured P2P networks. IEEE/ ACM Transactions on Networking, 24(5):3136–3147, October 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Song:2022:MSW

[SdVS22] Jianhan Song, Gustavo de Veciana, and Sanjay Shakkottai. Metascheduling for the wireless downlink through learning with bandit feedback. IEEE/ACM Transactions on Networking, 30(2):487– 500, April 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2021.3117783.

Su:2000:ERF

[SDW00] Ching-Fong Su, Gustavo De Veciana, and Jean Walrand. Explicit rate flow control for ABR services in ATM networks. IEEE/ACM Transactions on Networking, 8(3):350– 361, 2000. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/2000-8-3/
p350-su/.

Shen:2014:PON

[SDW14] Yuan Shen, Wenhan Dai, and Moe Z. Win. Power optimization for network localization. IEEE/ACM Transactions on Networking, 22(4):1337–1350, August 2014. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Shi:2021:FDS

[SE21] Zai Shi and Atilla Eryilmaz. A flexible distributed stochastic optimization framework for concurrent tasks in processing networks. IEEE/ ACM Transactions on Networking, 29(5):2045–2058, October 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2021.3078054.

Sarikaya:2015:CPC

[SEK15] Yunus Sarikaya, Ozgur Ercetin, and Can Emre Koksal. Confidentialitypreserving control of uplink cellular wireless networks using hybrid ARQ. IEEE/ ACM Transactions on Networking, 23(5):1457–1470, October 2015. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Sivaraman:2009:PPS

[SEMO09] Vijay Sivaraman, Hossam Elgindy, David Moreland, and Diethelm Ostry. Packet pacing in small buffer optical packet switched networks. IEEE/ACM Transactions on Networking, 17(4):1066–1079, August 2009. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Steiner:2009:LTS

[SENB09] Moritz Steiner, Taoufik En-Najjary, and Ernst W. Biersack. Long term study of peer behavior in the KAD DHT. IEEE/ACM Transactions on Networking, 17(5):1371–1384, October 2009. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Sestini:1997:RCG

[Ses97] Fabrizio Sestini. Recursive copy generation for multicast ATM switching. IEEE/ACM Transactions on Networking, 5(3):329– 335, June 1997. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1997-5-3/
p329-sestini/.

Shi:1995:HSR

[SF95] Jianxu Shi and John P.

Fonseka. Hierarchical selfhealing rings. IEEE/ ACM Transactions on Networking, 3(6):690–697, December 1995. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1995-3-6/
p690-shi/.

Sobrinho:2023:NOR

[SF23] João Luís Sobrinho and Miguel Alves Ferreira. From non-optimal routing protocols to routing on multiple optimality criteria. IEEE/ACM Transactions on Networking, 31(1): 294–307, 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2022.3191808.

Sangireddy:2005:SME

[SFAS05] Rama Sangireddy, Natsuhiko Futamura, Srinivas Aluru, and Arun K. Somani. Scalable, memory efficient, high-speed IP lookup algorithms. IEEE/ ACM Transactions on Networking, 13(4):802–812, August 2005. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Siganos:2003:PLL

[SFFF03] Georgos Siganos, Michalis Faloutsos, Petros Faloutsos, and Christos Faloutsos. Power laws and the AS-level Internet topology. IEEE/ACM Transactions on Networking, 11 (4):514–524, August 2003. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Sciancalepore:2018:MTI

[SFM^{+}18] Vincenzo Sciancalepore, Ilario Filippini, Vincenzo Mancuso, Antonio Capone, and Albert Banchs. A multi-traffic inter-cell interference coordination scheme in dense cellular networks. IEEE/ACM Transactions on Networking, 26(5):2361–2375, October 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Sterz:2022:MSS

[SFS^{+}22] Artur Sterz, Patrick Felka, Bernd Simon, Sabrina Klos, Anja Klein, Oliver Hinz, and Bernd Freisleben. Multi-stakeholder service placement via iterative bargaining with incomplete information. IEEE/ ACM Transactions on Networking, 30(4):1822– 1837, 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2022.3157040.

Simmons:1994:DED

[SG94] Jane M. Simmons and Robert G. Gallager. Design of error detection scheme for class C service in ATM. IEEE/ACM Transactions on Networking, 2(1):80–88, February 1994. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1994-2-1/
p80-simmons/.

Schultz:1996:MCR

[SG96] Kenneth J. Schultz and P. Glenn Gulak. Multicast contention resolution with single-cycle windowing using content addressable FIFO’s. IEEE/ ACM Transactions on Networking, 4(5):731–742, October 1996. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/
                                        journals/ton/1996-4-5/
                                        p731-schultz/.

Siachalou:2005:APC

[SG05] Stavroula Siachalou and Leonidas Georgiadis. Algorithms for precomputing constrained widest paths and multicast trees. IEEE/ ACM Transactions on Networking, 13(5):1174–1187, October 2005. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Sarkar:2013:DFT

[SG13] Rik Sarkar and Jie Gao. Differential forms for target tracking and aggregate queries in distributed networks. IEEE/ACM Transactions on Networking, 21(4):1159–1172, August 2013. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Shafiee:2017:SCA

[SG17a] Mehrnoosh Shafiee and Javad Ghaderi. A simple congestion-aware algorithm for load balancing in datacenter networks. IEEE/ ACM Transactions on Networking, 25(6):3670–3682, December 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Sharifnassab:2017:PNS

[SG17b] Arsalan Sharifnassab and S. Jamaloddin Golestani. On the possibility of network scheduling with polynomial complexity and delay. IEEE/ACM Transactions on Networking, 25(6): 3850–3862, December 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Shafiee:2018:IBM

[SG18] Mehrnoosh Shafiee and Javad Ghaderi. An improved bound for minimizing the total weighted completion time of coflows in datacenters. IEEE/ ACM Transactions on Networking, 26(4):1674–1687, August 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Sridharan:2005:ANO

[SGD05] Ashwin Sridharan, Roch Guérin, and Christophe Diot. Achieving nearoptimal traffic engineering solutions for current OSPF/IS-IS networks.

IEEE/ACM Transactions on Networking, 13(2):234– 247, April 2005. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Shafigh:2019:UCD

[SGH^{+}19] Alireza Shams Shafigh, Savo Glisic, Ekram Hossain, Beatriz Lorenzo, and Luiz A. DaSilva. User-centric distributed spectrum sharing in dynamic network architectures. IEEE/ACM Transactions on Networking, 27 (1):15–28, February 2019. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Swamy:2017:ACU

[SGJ17] Peruru Subrahmanya Swamy, Radha Krishna Ganti, and Krishna Jagannathan. Adaptive CSMA under the SINR model: Efficient approximation algorithms for throughput and utility maximization. IEEE/ ACM Transactions on Networking, 25(4):1968–1981, August 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Sun:2022:ESR

[SGL^{+}22] Penghao Sun, Zehua Guo, Junfei Li, Yang Xu, Julong Lan, and Yuxiang Hu. Enabling scalable routing in software-defined networks with deep reinforcement learning on critical nodes. IEEE/ACM Transactions on Networking, 30(2):629– 640, April 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2021.3126933.

Stone:1998:PCC

[SGPH98] Jonathan Stone, Michael Greenwald, Craig Partridge, and James Hughes. Performance of checksums and CRC’s over real data. IEEE/ACM Transactions on Networking, 6(5):529–543, October 1998. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1998-6-5/
p529-stone/.

Sappidi:2013:MAT

[SGR13] Rajasekhar Sappidi, André Girard, and Catherine Rosenberg. Maximum achievable throughput in a wireless sensor network using in-network computation for statistical functions. IEEE/ACM Transactions on Networking, 21(5):1581–1594, October 2013. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Sivaraman:2020:OPI

[SGS20] Vignesh Sivaraman, Dibyajyoti Guha, and Biplab Sikdar. Optimal pending interest table size for ICN with mobile producers. IEEE/ ACM Transactions on Networking, 28(4):1615–1628, August 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2020.2988713.

Serrano:2015:PFE

[SGSB^{+}15] Pablo Serrano, Andres Garcia-Saavedra, Giuseppe Bianchi, Albert Banchs, and Arturo Azcorra. Perframe energy consumption in 802.11 devices and its implication on modeling and design. IEEE/ ACM Transactions on Networking, 23(4):1243–1256, August 2015. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Sipos:2018:NAF

[SGVO18] Marton Sipos, Josh Gahm, Narayan Venkat, and Dave Oran. Network-aware feasible repairs for erasurecoded storage. IEEE/ACM Transactions on Networking, 26(3):1404–1417, June 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Song:2007:CLA

[SH07] Liang Song and Dimitrios Hatzinakos. A crosslayer architecture of wireless sensor networks for target tracking. IEEE/ ACM Transactions on Networking, 15(1):145–158, February 2007. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Shi:2012:SFR

[SH12] Yi Shi and Y. Thomas Hou. Some fundamental results on base station movement problem for wireless sensor networks. IEEE/ ACM Transactions on Networking, 20(4):1054–1067, August 2012. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Stamatiou:2014:DCM

[SH14] Kostas Stamatiou and Martin Haenggi. Delay characterization of multihop transmission in a Poisson field of interference. IEEE/ ACM Transactions on Networking, 22(6):1794–1807, December 2014. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Shi:2023:SSI

[SH23] Qi Shi and Dong Hao. Social sourcing: Incorporating social networks into crowdsourcing contest design. IEEE/ACM Transactions on Networking, 31(4): 1535–1549, 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2022.3223367.

Sharony:1994:UMS

[Sha94] Jacob Sharony. The universality of multidimensional switching networks. IEEE/ ACM Transactions on Networking, 2(6):602–612, December 1994. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1994-2-6/
p602-sharony/.

Sharon:1997:PLS

[Sha97] Oran Sharon. A proof for lack of starvation in DQDB with and without slot reuse. IEEE/ ACM Transactions on Networking, 5(3):410–420, June 1997. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1997-5-3/
p410-sharon/.

Shenker:1995:MGW

[She95] Scott J. Shenker. Making greed work in networks: a game-theoretic analysis of switch service disciplines. IEEE/ ACM Transactions on Networking, 3(6):819–831, December 1995. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1995-3-6/
p819-shenker/.

Seetharaman:2009:RCL

[SHHA09] Srini Seetharaman, Volker Hilt, Markus Hofmann, and Mostafa Ammar. Resolving cross-layer conflict between overlay routing and traffic engineering. IEEE/ ACM Transactions on Networking, 17(6):1964–1977, December 2009. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Spatscheck:2000:OTF

[SHHP00] Oliver Spatscheck, Jørgen S.

Hansen, John H. Hartman, and Larry L. Peterson. Optimizing TCP forwarder performance. IEEE/ACM Transactions on Networking, 8(2):146– 157, 2000. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/2000-8-2/
p146-spatscheck/.

She:2010:HRC

[SHJ10] Qingya She, Xiaodong Huang, and Jason P. Jue. How reliable can two-path protection be? IEEE/ACM Transactions on Networking, 18(3):922–933, June 2010. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Siekkinen:2016:UVS

[SHN16] Matti Siekkinen, Mohammad Ashraful Hoque, and Jukka K. Nurminen. Using viewing statistics to control energy and traffic overhead in mobile video streaming. IEEE/ACM Transactions on Networking, 24 (3):1489–1503, June 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Shokrollahi:2006:RC

[Sho06] Amin Shokrollahi. Raptor codes. IEEE/ACM Transactions on Networking, 14 (SI):2551–2567, June 2006. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Shukla:2023:RVP

[SHV^{+}23] Apoorv Shukla, Kevin Hudemann, Zsolt Vági, Lily Hügerich, Georgios Smaragdakis, Artur Hecker, Stefan Schmid, and Anja Feldmann. Runtime verification for programmable switches. IEEE/ACM Transactions on Networking, 31(4):1822–1837, 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL

https://dl.acm.org/doi/
10.1109/TNET.2023.3234931.

Shin:2016:RCE

[SHZ16] Dong-Hoon Shin, Shibo He, and Junshan Zhang. Robust and cost-effective design of cyber-physical systems: an optimal middleware deployment approach. IEEE/ACM Transactions on Networking, 24(2):1081– 1094, April 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Srivatsa:2009:SKM

[SIYL09] Mudhakar Srivatsa, Arun Iyengar, Jian Yin, and Ling Liu. Scalable key management algorithms for location-based services. IEEE/ACM Transactions on Networking, 17(5):1399– 1412, October 2009. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Sidhu:1995:MGB

[SJ95] Ikhlaq Sidhu and Scott Jordan. Multiplexing gains in bit stream multiplexors. IEEE/ACM Transactions on Networking, 3(6):785–797, December 1995. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/
                                        journals/ton/1995-3-6/
                                        p785-sidhu/.

Shakkottai:2010:DAC

[SJ10] Srinivas Shakkottai and Ramesh Johari. Demandaware content distribution on the Internet. IEEE/ ACM Transactions on Networking, 18(2):476–489, April 2010. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Sebbah:2012:DQR

[SJ12] Samir Sebbah and Brigitte Jaumard. Differentiated quality-of-recovery in survivable optical mesh networks using p-structures. IEEE/ACM Transactions on Networking, 20(3):798– 810, June 2012. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Sallam:2021:JPA

[SJ21] Gamal Sallam and Bo Ji. Joint placement and allocation of VNF nodes with budget and capacity constraints. IEEE/ ACM Transactions on Networking, 29(3):1238–1251, June 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2021.3058378.

Sen:2010:MDN

[SJGH10] Soumya Sen, Youngmi Jin, Roch Guérin, and Kartik Hosanagar. Modeling the dynamics of network technology adoption and the role of converters. IEEE/ ACM Transactions on Networking, 18(6):1793–1805, December 2010. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Shafiq:2013:LSM

[SJL^{+}13] M. Zubair Shafiq, Lusheng Ji, Alex X. Liu, Jeffrey Pang, and Jia Wang. Large-scale measurement and characterization of cellular machine-to-machine traffic. IEEE/ACM Transactions on Networking, 21 (6):1960–1973, December 2013. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Shafiq:2016:COC

[SJL^{+}16] M. Zubair Shafiq, Lusheng Ji, Alex X. Liu, Jeffrey Pang, Shobha Venkataraman, and Jia Wang. Characterizing and optimizing cellular network performance during crowded events. IEEE/ACM Transactions on Networking, 24 (3):1308–1321, June 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Smith:2022:TIM

[SJSB22] Kevin D. Smith, Saber Jafarpour, Ananthram Swami, and Francesco Bullo. Topology inference with multivariate cumulants: The Möbius inference algorithm. IEEE/ACM Transactions on Networking, 30 (5):2102–2116, 2022. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1109/
                                        TNET.2022.3164336.

Shin:2017:CGI

[SJWH^{+}17] Kyuyong Shin, Carlee Joe-Wong, Sangtae Ha, Yung Yi, Injong Rhee, and Douglas S. Reeves. T-Chain: a general incentive scheme for cooperative computing. IEEE/ACM Transactions on Networking, 25(4): 2122–2137, August 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Shan:2024:EFT

[SJZ^{+}24] Danfeng Shan, Linbing Jiang, Peng Zhang, Wanchun Jiang, Hao Li, Yazhe Tang, and Fengyuan Ren. Enforcing fairness in the traffic policer among heterogeneous congestion control algorithms. IEEE/ACM Transactions on Networking, 32(1):34–49, 2024. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL https:/

/dl.acm.org/doi/10.1109/
TNET.2023.3276410.

Stamatelos:1997:RBA

[SK97] George M. Stamatelos and Vassilios N. Koukoulidis. Reservation-based bandwidth allocation in a radio ATM network. IEEE/ACM Transactions on Networking, 5(3):420– 428, June 1997. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1997-5-3/
p420-stamatelos/.

Sahni:2003:ECM

[SK03] Sartaj Sahni and Kun Suk Kim. Efficient construction of multibit tries for IP lookup. IEEE/ACM Transactions on Networking, 11 (4):650–662, August 2003. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Soh:2006:PBR

[SK06] Wee-Seng Soh and Hyong S. Kim. A predictive bandwidth reservation scheme using mobile positioning and road topology information. IEEE/ ACM Transactions on Networking, 14(5):1078–1091, October 2006. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Saleh:2010:DPW

[SK10a] Mohammad A. Saleh and Ahmed E. Kamal. Design and provisioning of WDM networks with many-to-many traffic grooming. IEEE/ACM Transactions on Networking, 18(6):1869– 1882, December 2010. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). See erratum [SK11].

Shu:2010:CTO

[SK10b] Tao Shu and Marwan Krunz. Coverage-time optimization for clustered wireless sensor networks: a power-balancing approach. IEEE/ACM Transactions on Networking, 18(1):202– 215, February 2010. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Saleh:2011:EDP

[SK11] Mohammad A. Saleh and Ahmed E. Kamal. Erratum to Design and Provisioning of WDM Networks With Many-to-Many Traffic Grooming. IEEE/ACM Transactions on Networking, 19(1):299, February 2011. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). See [SK10a].

Saleh:2012:AAM

[SK12a] Mohammad A. Saleh and Ahmed E. Kamal. Approximation algorithms for many-to-many traffic grooming in optical WDM networks. IEEE/ ACM Transactions on Networking, 20(5):1527–1540, October 2012. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Shu:2012:FCR

[SK12b] Tao Shu and Marwan Krunz. Finding cheap routes in profit-driven opportunistic spectrum access networks: a truthful mechanism design approach. IEEE/ACM Transactions on Networking, 20 (2):530–543, April 2012. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Srivastava:2013:BPL

[SK13] Rahul Srivastava and Can Emre Koksal. Basic performance limits and tradeoffs in energy-harvesting sensor nodes with finite data and energy storage. IEEE/ ACM Transactions on Networking, 21(4):1049–1062, August 2013. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Singh:2021:ACD

[SK21] Rahul Singh and P. R. Kumar. Adaptive CSMA for decentralized scheduling of multi-hop networks with end-to-end deadline constraints. IEEE/ ACM Transactions on Networking, 29(3):1224–1237, June 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2021.3063626.

Sim:2018:OCA

[SKA^{+}18] Gek Hong Sim, Sabrina Klos, Arash Asadi, Anja Klein, and Matthias Hollick. An online contextaware machine learning algorithm for 5G mmWave vehicular communications. IEEE/ACM Transactions on Networking, 26(6):2487– 2500, December 2018. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Skorin-Kapov:2010:NAO

[SKCW10] Nina Skorin-Kapov, Jiajia Chen, and Lena Wosinska. A new approach to optical networks security: attack-aware routing and wavelength assignment. IEEE/ACM Transactions on Networking, 18 (3):750–760, June 2010. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Sarikaya:2016:DNC

[SKE16] Yunus Sarikaya, C. Emre Koksal, and Ozgur Ercetin. Dynamic network control for confidential multi-hop communications. IEEE/ ACM Transactions on Networking, 24(2):1181–1195, April 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Shabara:2019:BDU

[SKE19] Yahia Shabara, C. Emre Koksal, and Eylem Ekici. Beam discovery using linear block codes for millimeter wave communication networks. IEEE/ ACM Transactions on Networking, 27(4):1446–1459, August 2019. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Shue:2012:AMA

[SKG12] Craig A. Shue, Andrew J.

Kalafut, and Minaxi Gupta. Abnormally malicious autonomous systems and their Internet connectivity. IEEE/ACM Transactions on Networking, 20(1): 220–230, February 2012. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Sermpezis:2018:ANB

[SKG^{+}18] Pavlos Sermpezis, Vasileios Kotronis, Petros Gigis, Xenofontas Dimitropoulos, Danilo Cicalese, Alistair King, and Alberto Dainotti. ARTEMIS: Neutralizing BGP hijacking within a minute. IEEE/ACM Transactions on Networking, 26(6):2471–2486, December 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Song:2012:ETB

[SKHL12] Haoyu Song, Murali Kodialam, Fang Hao, and T. V. Lakshman. Efficient trie braiding in scalable virtual routers. IEEE/ ACM Transactions on Networking, 20(5):1489–1500, October 2012. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Sabharwal:2007:OSU

[SKK07] Ashutosh Sabharwal, Ahmad Khoshnevis, and Edward Knightly. Opportunistic spectral usage: bounds and a multiband CSMA/CA protocol. IEEE/ACM Transactions on Networking, 15(3):533– 545, June 2007. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Shakkottai:2001:TPE

[SKKA01] Sanjay Shakkottai, Anurag Kumar, Aditya Karnik, and Ajit Anvekar. TCP performance over endto-end rate control and stochastic available capacity. IEEE/ACM Transactions on Networking, 9 (4):377–391, August 2001. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Scheuermann:2009:TSD

[SKR^{+}09] Björn Scheuermann, Wolfgang Kiess, Magnus Roos, Florian Jarre, and Martin Mauve. On the time synchronization of distributed log files in networks with local broadcast media. IEEE/ACM Transactions on Networking, 17 (2):431–444, April 2009. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Sharma:2012:TPE

[SKRK12] Vicky Sharma, Koushik Kar, K. K. Ramakrishnan, and Shivkumar Kalyanaraman. A transport protocol to exploit multipath diversity in wireless networks. IEEE/ACM Transactions on Networking, 20(4):1024– 1039, August 2012. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Singh:2016:CBS

[SKS16] Chandramani Singh, Anurag Kumar, and Rajesh Sundaresan. Combined base station association and power control in multichannel cellular networks. IEEE/ACM Transactions on Networking, 24(2):1065– 1080, April 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Salehi:1996:EAS

[SKT96] James D. Salehi, James F. Kurose, and Don Towsley. The effectiveness of affinitybased scheduling in multiprocessor network protocol processing (extended version). IEEE/ACM Transactions on Networking, 4(4):516–530, August 1996. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/
                                        journals/ton/1996-4-4/
                                        p516-salehi/.

Sridharan:2012:GLS

[SKUB12] Arun Sridharan, C. Emre Koksal, and Elif Uysal-Biyikoglu. A greedy link scheduler for wireless networks with Gaussian multiple-access and broadcast channels. IEEE/ ACM Transactions on Networking, 20(1):100–113, February 2012. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Sikdar:2003:AML

[SKV03] Biplab Sikdar, Shivkumar Kalyanaraman, and Kenneth S. Vastola. Analytic models for the latency and steady-state throughput of TCP Tahoe, Reno, and SACK. IEEE/ACM Transactions on Networking, 11(6):959–971, December 2003. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Salameh:2010:CAS

[SKY10] Haythem A. Bany Salameh, Marwan Krunz, and Ossama Younis. Cooperative adaptive spectrum sharing in cognitive radio networks. IEEE/ACM Transactions on Networking, 18(4):1181– 1194, August 2010. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Starobinski:2003:ANC

[SKZ03] David Starobinski, Mark Karpovsky, and Lev A. Zakrevski. Application of network calculus to general topologies using turnprohibition. IEEE/ACM Transactions on Networking, 11(3):411–421, June 2003. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Sheng:1994:SAP

[SL94] Hong-Dah Sheng and San-Qi Li. Spectral analysis of packet loss rate at a statistical multiplexer for multimedia services. IEEE/ACM Transactions on Networking, 2(1):53–65, February 1994. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/
                                        journals/ton/1994-2-1/
                                        p53-sheng/.

Shankar:1995:MTP

[SL95] A. Udaya Shankar and David Lee. Minimumlatency transport protocols with modulo-N. IEEE/ACM Transactions on Networking, 3(3):255– 268, June 1995. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/
                                        journals/ton/1995-3-3/
                                        p255-shankar/. See com-

ments [OdG96].

Sastry:2005:CTW

[SL05] Nishanth R. Sastry and Simon S. Lam. CYRF: a theory of window-based unicast congestion control. IEEE/ACM Transactions on Networking, 13(2):330– 342, April 2005. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Shorten:2007:QPN

[SL07a] Robert N. Shorten and Douglas J. Leith. On queue provisioning, network efficiency and the transmission control protocol. IEEE/ACM Transactions on Networking, 15 (4):866–877, August 2007. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Sun:2007:HGA

[SL07b] Yan Sun and K. J. Ray Liu.

Hierarchical group access control for secure multicast communications. IEEE/ ACM Transactions on Networking, 15(6):1514–1526, December 2007. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Spitler:2008:IEE

[SL08] Stephen L. Spitler and Daniel C. Lee. Integration of explicit effectivebandwidth-based QoS routing with best-effort routing. IEEE/ACM Transactions on Networking, 16(4):957– 969, August 2008. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Shpungin:2012:TEE

[SL12] Hanan Shpungin and Zongpeng Li. Throughput and energy efficiency in wireless ad hoc networks with Gaussian channels. IEEE/ACM Transactions on Networking, 20(1):15–28, February 2012. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Sheng:2014:PIT

[SL14] Shang-Pin Sheng and Mingyan Liu. Profit incentive in trading nonexclusive access on a secondary spectrum market through contract design. IEEE/ ACM Transactions on Networking, 22(4):1190–1203, August 2014. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Shahzad:2015:FAE

[SL15a] Muhammad Shahzad and Alex X. Liu. Fast and accurate estimation of RFID tags. IEEE/ACM Transactions on Networking, 23(1): 241–254, February 2015. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Shahzad:2015:POT

[SL15b] Muhammad Shahzad and Alex X. Liu. Probabilistic optimal tree hopping for RFID identification. IEEE/ACM Transactions on Networking, 23(3): 796–809, June 2015. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Shen:2015:HAA

[SL15c] Haiying Shen and Ze Li.

A hierarchical accountaided reputation management system for MANETs. IEEE/ACM Transactions on Networking, 23(1):70– 84, February 2015. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Shahzad:2016:AEP

[SL16a] Muhammad Shahzad and Alex X. Liu. Accurate and efficient per-flow latency measurement without probing and time stamping. IEEE/ACM Transactions on Networking, 24(6):3477–3492, December 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Shahzad:2016:FRD

[SL16b] Muhammad Shahzad and Alex X. Liu. Fast and reliable detection and identification of missing RFID tags in the wild. IEEE/ ACM Transactions on Networking, 24(6):3770–3784, December 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Shamsi:2017:UCU

[SL17] Zain Shamsi and Dmitri Loguinov. Unsupervised clustering under temporal feature volatility in network stack fingerprinting. IEEE/ACM Transactions on Networking, 25(4):2430– 2443, August 2017. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Schweller:2007:RSE

[SLC^{+}07] Robert Schweller, Zhichun Li, Yan Chen, Yan Gao, Ashish Gupta, Yin Zhang, Peter A. Dinda, Ming-Yang Kao, and Gokhan Memik. Reversible sketches: enabling monitoring and analysis over high-speed data streams. IEEE/ ACM Transactions on Networking, 15(5):1059–1072, October 2007. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Shakya:2022:DCH

[SLC22] Nishant Shakya, Fan Li, and Jinyuan Chen. On distributed computing with heterogeneous communication constraints. IEEE/ ACM Transactions on Networking, 30(6):2776– 2787, 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2022.3181234.

Sojoudi:2014:BDS

[SLD14] Somayeh Sojoudi, Steven H.

Low, and John C. Doyle. Buffering dynamics and stability of Internet congestion controllers. IEEE/ ACM Transactions on Networking, 22(6):1808–1818, December 2014. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Sun:2022:TCS

[SLD^{+}22] Yu Sun, Chi Lin, Haipeng Dai, Pengfei Wang, Lei Wang, Guowei Wu, and Qiang Zhang. Trading off charging and sensing for stochastic events monitoring in WRSNs. IEEE/ACM Transactions on Networking, 30(2):557– 571, April 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2021.3122130.

Shen:2023:EDO

[SLD^{+}23] Dian Shen, Junzhou Luo, Fang Dong, Xiaolin Guo, Ciyuan Chen, Kai Wang, and John C. S. Lui. Enabling distributed and optimal RDMA resource sharing in large-scale data center networks: Modeling, analysis, and implementation. IEEE/ACM Transactions on Networking, 31(6): 2745–2760, 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2023.3263562.

Shi:2021:COC

[SLF21] Ming Shi, Xiaojun Lin, and Sonia Fahmy. Competitive online convex optimization with switching costs and ramp constraints. IEEE/ACM Transactions on Networking, 29(2):876– 889, April 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2021.3053910.

Shafigh:2016:FDN

[SLG^{+}16] Alireza Shams Shafigh, Beatriz Lorenzo, Savo Glisic, Jordi Pérez-Romero, Luiz A. DaSilva, Allen B. MacKenzie, and Juha Röning. A framework for dynamic network architecture and topology optimization. IEEE/ACM Transactions on Networking, 24(2):717–730, April 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Sikora:2006:BPE

[SLH^{+}06] Marcin Sikora, J. Nicholas Laneman, Martin Haenggi, Daniel J. Costello, Jr., and Thomas E. Fuja. Bandwidth- and powerefficient routing in linear wireless networks. IEEE/ ACM Transactions on Networking, 14(SI):2624–2633, June 2006. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Song:2019:MCD

[SLH^{+}19] Liang Song, Chunyan Liu, Hejiao Huang, Hongwei Du, and Xiaohua Jia. Minimum connected dominating set under routing cost constraint in wireless sensor networks with different transmission ranges. IEEE/ACM Transactions on Networking, 27(2):546– 559, April 2019. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Shi:2016:FTC

[SLJJ16] Guodong Shi, Bo Li, Mikael Johansson, and Karl Henrik Johansson. Finitetime convergent gossiping. IEEE/ACM Transactions on Networking, 24(5):2782– 2794, October 2016. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Smaragdakis:2011:SON

[SLL^{+}11] Georgios Smaragdakis, Nikolaos Laoutaris, Vassilis Lekakis, Azer Bestavros, John W. Byers, and Mema Roussopoulos. Selfish overlay network creation and maintenance. IEEE/ACM Transactions on Networking, 19(6):1624–1637, December 2011. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Shen:2015:SNA

[SLL15] Haiying Shen, Yuhua Lin, and Jin Li. A social-network-aided efficient peer-to-peer live streaming system. IEEE/ ACM Transactions on Networking, 23(3):987–1000, June 2015. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Smaragdakis:2014:DSM

[SLO^{+}14] Georgios Smaragdakis, Nikolaos Laoutaris, Konstantinos Oikonomou, Ioannis Stavrakakis, and Azer Bestavros. Distributed server migration for scalable Internet service deployment. IEEE/ACM Transactions on Networking, 22(3):917–930, June 2014. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Salido:2007:EBE

[SLP07] Javier Salido, Loukas Lazos, and Radha Poovendran. Energy and bandwidthefficient key distribution in wireless ad hoc networks: a cross-layer approach. IEEE/ACM Transactions on Networking, 15 (6):1527–1540, December 2007. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Shakkottai:2010:MCL

[SLS10] Srinivas Shakkottai, Xin Liu, and R. Srikant. The multicast capacity of large multihop wireless networks. IEEE/ACM Transactions on Networking, 18(6):1691– 1700, December 2010. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Serbetci:2023:MTC

[SLS^{+}23] Berksan Serbetci, Eleftherios Lampiris, Thrasyvoulos Spyropoulos, Giuseppe Caire, and Petros Elia. Multi-transmitter coded caching networks with transmitter-side knowledge of file popularity. IEEE/ ACM Transactions on Networking, 31(3):1277–1292, June 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2022.3219161.

Su:2020:PAR

[SLSC20] Jian Su, Alex X. Liu, Zhengguo Sheng, and Yongrui Chen. A partitioning approach to RFID identification. IEEE/ACM Transactions on Networking, 28(5):2160–2173, October 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2020.3004852.

Shrivastav:2019:GST

[SLWW19] Vishal Shrivastav, Ki Suh Lee, Han Wang, and Hakim Weatherspoon. Globally synchronized time via datacenter networks. IEEE/ ACM Transactions on Networking, 27(4):1401–1416, August 2019. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Sikdar:2000:QAS

[SM00] Biplab Sikdar and D. Manjunath. Queueing analysis of scheduling policies in copy networks of spacebased multicast packet switches. IEEE/ACM Transactions on Networking, 8(3):396–406, 2000. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL

                                        http://www.acm.org/pubs/
                                        citations/journals/ton/
                                        2000-8-3/p396-sikdar/.

Sinha:2005:ITA

[SM05] Sushant Sinha and C. Siva Ram Murthy. Information theoretic approach to traffic adaptive WDM networks. IEEE/ACM Transactions on Networking, 13(4):881– 894, August 2005. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Sharma:2008:CHS

[SM08] Gaurav Sharma and Ravi R.

Mazumdar. A case for hybrid sensor networks. IEEE/ACM Transactions on Networking, 16(5):1121– 1132, October 2008. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Seferoglu:2014:NCA

[SM14] Hulya Seferoglu and Athina Markopoulou. Network coding-aware queue management for TCP flows over coded wireless networks. IEEE/ACM Transactions on Networking, 22(4):1297– 1310, August 2014. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Seferoglu:2016:SRS

[SM16] Hulya Seferoglu and Eytan Modiano. Separation of routing and scheduling in backpressure: based wireless networks. IEEE/ACM Transactions on Networking, 24(3):1787–1800, June 2016. CODEN IEANEP.

ISSN 1063-6692 (print), 1558-2566 (electronic).

Santagati:2017:SDU

[SM17] G. Enrico Santagati and Tommaso Melodia. A software-defined ultrasonic networking framework for wearable devices. IEEE/ ACM Transactions on Networking, 25(2):960–973, April 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). See corrections [SM19].

Sinha:2018:OCG

[SM18] Abhishek Sinha and Eytan Modiano. Optimal control for generalized networkflow problems. IEEE/ ACM Transactions on Networking, 26(1):506–519, February 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Santagati:2019:CSD

[SM19] G. Enrico Santagati and Tommaso Melodia. Corrections to “A Software-Defined Ultrasonic Networking Framework for Wearable Devices”. IEEE/ ACM Transactions on Networking, 27(3):1289, June 2019. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). See [SM17].

Shannon:2002:BFO

[SMC02] Colleen Shannon, David Moore, and K. C. Claffy. Beyond folklore: observations on fragmented traffic. IEEE/ACM Transactions on Networking, 10(6): 709–720, December 2002. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Shifrin:2020:OPC

[SMC^{+}20] Mark Shifrin, Daniel S.

Menasché, Asaf Cohen, Dennis Goeckel, and Omer Gurewitz. Optimal PHY configuration in wireless networks. IEEE/ACM Transactions on Networking, 28(6):2601–2614, December 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2020.3015881.

Sexton:2020:PSO

[SMD20] Conor Sexton, Nicola Marchetti, and Luiz A. DaSilva. On provisioning slices and overbooking resources in service tailored networks of the future. IEEE/ACM Transactions on Networking, 28(5):2106–2119, October 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.

                                        acm.org/doi/10.1109/TNET.
                                        2020.3004443.

Sami:2020:VOD

[SMEH20] Hani Sami, Azzam Mourad, and Wassim El-Hajj. Vehicular-OBUs-As-On-Demand-Fogs: Resource and context aware deployment of containerized micro-services. IEEE/ACM Transactions on Networking, 28(2):778– 790, April 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL https://

                                        dl.acm.org/doi/abs/10.
                                        1109/TNET.2020.2973800.

Saengudomlert:2005:DWA

[SMG05a] Poompat Saengudomlert, Eytan H. Modiano, and Robert G. Gallager. Dynamic wavelength assignment for WDM all-optical tree networks. IEEE/ACM Transactions on Networking, 13(4):895–905, August 2005. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Shi:2005:LBP

[SMG05b] Weiguang Shi, M. H. MacGregor, and Pawel Gburzynski. Load balancing for parallel forwarding. IEEE/ACM Transactions on Networking, 13(4):790– 801, August 2005. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Saengudomlert:2006:LRW

[SMG06] Poompat Saengudomlert, Eytan Modiano, and Robert G. Gallager. On-line routing and wavelength assignment for dynamic traffic in WDM ring and torus networks. IEEE/ACM Transactions on Networking, 14(2):330– 340, April 2006. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Santagati:2015:MAC

[SMGP15] G. Enrico Santagati, Tommaso Melodia, Laura Galluccio, and Sergio Palazzo. Medium access control and rate adaptation for ultrasonic intrabody sensor networks. IEEE/ACM Transactions on Networking, 23(4):1121–1134, August 2015. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Saidi:1995:NSP

[SMH95] H. Saidi, P. S. Min, and M. V. Hegde. A new structural property of statistical data forks. IEEE/ACM Transactions on Networking, 3(3):289– 298, June 1995. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1995-3-3/
p289-saidi/.

Smith:1995:ERC

[Smi95] Donald E. Smith. Ensuring robust call throughput and fairness for SCP overload controls. IEEE/ ACM Transactions on Networking, 3(5):538–548, October 1995. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/
                                        journals/ton/1995-3-5/
                                        p538-smith/.

Smiljanic:2002:FBA

[Smi02] Aleksandra Smiljanić. Flexible bandwidth allocation in high-capacity packet switches. IEEE/ACM Transactions on Networking, 10(2):287–293, April 2002. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Smiljanic:2008:RDG

[Smi08] Aleksandra Smiljanic. Rate and delay guarantees provided by Clos packet switches with load balancing. IEEE/ACM Transactions on Networking, 16(1): 170–181, February 2008. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Somani:2004:TWF

[SML04] Arun K. Somani, Mani Mina, and Ling Li. On trading wavelengths with fibers: a cost-performance based study. IEEE/ACM Transactions on Networking, 12(5):944–951, October 2004. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Stoica:2003:CSP

[SMLN^{+}03] Ion Stoica, Robert Morris, David Liben-Nowell, David R. Karger, M. Frans Kaashoek, Frank Dabek, and Hari Balakrishnan. Chord: a scalable peerto-peer lookup protocol for Internet applications. IEEE/ACM Transactions on Networking, 11(1):17– 32, February 2003. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Singh:2011:IAH

[SMM11] Sumit Singh, Raghuraman Mudumbai, and Upamanyu Madhow. Interference analysis for highly directional 60-GHz mesh networks: the case for rethinking medium access control. IEEE/ACM Transactions on Networking, 19(5): 1513–1527, October 2011. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Skoberne:2014:IAS

[SMP^{+}14] Nejc Skoberne, Olaf Maennel, Iain Phillips, Randy Bush, Jan Zorz, and Mojca Ciglaric. IPv4 address sharing mechanism classification and tradeoff analysis. IEEE/ACM Transactions on Networking, 22 (2):391–404, April 2014. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Sharma:2007:DCT

[SMS07] Gaurav Sharma, Ravi Mazumdar, and Ness B. Shroff. Delay and capacity trade-offs in mobile ad hoc networks: a global perspective. IEEE/ACM Transactions on Networking, 15 (5):981–992, October 2007. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Sirkeci-Mergen:2006:AAM

[SMSM06] Birsen Sirkeci-Mergen, Anna Scaglione, and Gökhan Mergen. Asymptotic analysis of multistage cooperative broadcast in wireless networks. IEEE/ ACM Transactions on Networking, 14(SI):2531–2550, June 2006. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Saha:1998:CRR

[SMT98] Debanjan Saha, Sarit Mukherjee, and Satish K. Tripathi. Carry-over round robin: a simple cell scheduling mechanism for ATM networks. IEEE/ACM Transactions on Networking, 6(6):779–796, December 1998. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1998-6-6/

p779-saha/. See comments [PK01].

Sidhu:1993:THP

[SMV93] Deepinder P. Sidhu, Howard Motteler, and Raghu Vallurupalli. On testing hierarchies for protocols. IEEE/ ACM Transactions on Networking, 1(5):590–599, October 1993. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1993-1-5/
p590-sidhu/.

Spring:2004:MIT

[SMWA04] Neil Spring, Ratul Mahajan, David Wetherall, and Thomas Anderson. Measuring ISP topologies with rocketfuel. IEEE/ACM Transactions on Networking, 12(1):2–16, February 2004. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Supittayapornpong:2015:QIM

[SN15] Sucha Supittayapornpong and Michael J. Neely. Quality of information maximization for wireless networks via a fully separable quadratic policy. IEEE/ACM Transactions on Networking, 23(2):574– 586, April 2015. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Soule:2007:EDT

[SNC^{+}07] Augustin Soule, Antonio Nucci, Rene L. Cruz, Emilio Leonardi, and Nina Taft. Estimating dynamic traffic matrices by using viable routing changes. IEEE/ACM Transactions on Networking, 15(3):485– 498, June 2007. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Salem:2023:ASC

[SNC23] Tareq Si Salem, Giovanni Neglia, and Damiano Carra. Ascent similarity caching with approximate indexes. IEEE/ ACM Transactions on Networking, 31(3):1173–1186, June 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2022.3217012.

Shamsi:2016:HSP

[SNLL16] Zain Shamsi, Ankur Nandwani, Derek Leonard, and Dmitri Loguinov. Hershel: Single-packet OS fingerprinting. IEEE/ ACM Transactions on Networking, 24(4):2196–2209, August 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Sherman:2012:FDB

[SNS12] Alex Sherman, Jason Nieh, and Clifford Stein. Fair-Torrent: a deficit-based distributed algorithm to ensure fairness in peerto-peer systems. IEEE/ ACM Transactions on Networking, 20(5):1361–1374, October 2012. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Shen:2012:ITS

[SNSW12] Charles Shen, Erich Nahum, Henning Schulzrinne, and Charles P. Wright. The impact of TLS on SIP server performance: measurement and modeling. IEEE/ ACM Transactions on Networking, 20(4):1217–1230, August 2012. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Shen:2013:DCL

[SNXT13] Yilin Shen, Nam P. Nguyen, Ying Xuan, and My T. Thai. On the discovery of critical links and nodes for assessing network vulnerability. IEEE/ACM Transactions on Networking, 21 (3):963–973, June 2013. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Sahay:2023:PLI

[SNZ^{+}23] Rajeev Sahay, Serena Nicoll, Minjun Zhang, Tsung-Yen Yang, Carlee Joe-Wong, Kerrie A. Douglas, and Christopher G. Brinton. Predicting learning interactions in social learning networks: a deep learning enabled approach. IEEE/ACM Transactions on Networking, 31(5):2086– 2100, 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2023.3237978.

Sobrinho:2002:AAQ

[Sob02] João Luís Sobrinho. Algebra and algorithms for QoS path computation and hopby-hop routing in the Internet. IEEE/ACM Transactions on Networking, 10 (4):541–550, August 2002. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Sobrinho:2005:ATD

[Sob05] João Luís Sobrinho. An algebraic theory of dynamic network routing. IEEE/ ACM Transactions on Networking, 13(5):1160–1173, October 2005. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Sobrinho:2017:CRV

[Sob17] Joao Luis Sobrinho. Correctness of routing vector protocols as a property of network cycles. IEEE/ACM Transactions on Networking, 25(1):150– 163, February 2017. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Schwartz:1994:AIG

[SP94] Michael F. Schwartz and Calton Pu. Applying an information gathering architecture to Netfind: a white pages tool for a changing and growing Internet. IEEE/ACM Transactions on Networking, 2(5):426–439, October 1994. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1994-2-5/
p426-schwartz/.

Stai:2016:PAC

[SPB16] Eleni Stai, Symeon Papavassiliou, and John S. Baras. Performance-aware cross-layer design in wireless multihop networks via a weighted backpressure approach. IEEE/ ACM Transactions on Networking, 24(1):245–258, February 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Shen:2010:SPT

[SPC10] Yanming Shen, Shivendra S. Panwar, and H. Jonathan Chao. SQUID: a practical 100% throughput scheduler for crosspoint buffered switches. IEEE/ ACM Transactions on Networking, 18(4):1119–1131, August 2010. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Srebrny:2013:NMD

[SPGM13] Piotr H. Srebrny, Thomas Plagemann, Vera Goebel, and Andreas Mauthe. No more Déjà Vu: eliminating redundancy with cachecast: feasibility and performance gains. IEEE/ACM Transactions on Networking, 21 (6):1736–1749, December 2013. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Samar:2004:IZR

[SPH04] Prince Samar, Marc R. Pearlman, and Zygmunt J. Haas. Independent zone routing: an adaptive hybrid routing framework for ad hoc wireless networks. IEEE/ACM Transactions on Networking, 12(4):595– 608, August 2004. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Spinelli:1997:SSW

[Spi97] John M. Spinelli. Selfstabilizing sliding window ARQ protocols. IEEE/ ACM Transactions on Networking, 5(2):245–254, April 1997. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1997-5-2/
p245-spinelli/.

Sinha:2017:TOMa

[SPLM17] Abhishek Sinha, Georgios Paschos, Chih-Ping Li, and Eytan Modiano. Throughput-optimal multihop broadcast on directed acyclic wireless networks. IEEE/ACM Transactions on Networking, 25(1):377– 391, February 2017. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Saino:2020:LIC

[SPLP20] Lorenzo Saino, Ioannis Psaras, Emilio Leonardi, and George Pavlou. Load imbalance and caching performance of sharded systems. IEEE/ACM Transactions on Networking, 28(1):112–125, February 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL https://

dl.acm.org/doi/abs/10.
1109/TNET.2019.2957075.

Sinha:2017:TOMb

[SPM^{+}17] Abhishek Sinha, Georgios Paschos, Eytan Modiano, Abhishek Sinha, Georgios Paschos, and Eytan Modiano. Throughput-optimal multi-hop broadcast algorithms. IEEE/ACM Transactions on Networking, 25(5):3088–3101, October 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Sangdeh:2020:PSS

[SPQZ20] Pedram Kheirkhah Sangdeh, Hossein Pirayesh, Adnan Quadri, and Huacheng Zeng. A practical spectrum sharing scheme for cognitive radio networks: Design and experiments. IEEE/ ACM Transactions on Networking, 28(4):1818–1831, August 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2020.2994134.

Spyropoulos:2008:ERIb

[SPR08a] Thrasyvoulos Spyropoulos, Konstantinos Psounis, and Cauligi S. Raghavendra.

Efficient routing in intermittently connected mobile networks: the multiplecopy case. IEEE/ACM Transactions on Networking, 16(1):77–90, February 2008. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Spyropoulos:2008:ERIa

[SPR08b] Thrasyvoulos Spyropoulos, Konstantinos Psounis, and Cauligi S. Raghavendra. Efficient routing in intermittently connected mobile networks: the single-copy case. IEEE/ACM Transactions on Networking, 16 (1):63–76, February 2008. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Shao:2020:PWF

[SPR^{+}20] Chenglong Shao, Hoorin Park, Heejun Roh, Wonjun Lee, and Hyoil Kim. PolarScout: Wi-Fi interferenceresilient ZigBee communication via Shell-shaping. IEEE/ACM Transactions on Networking, 28(4):1587– 1600, August 2020. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL https:/

/dl.acm.org/doi/10.1109/
TNET.2020.2989387.

Snoeren:2002:SPI

[SPS^{+}02] Alex C. Snoeren, Craig Partridge, Luis A. Sanchez, Christine E. Jones, Fabrice Tchakountio, Beverly Schwartz, Stephen T. Kent, and W. Timothy Strayer. Single-packet IP traceback. IEEE/ACM Transactions on Networking, 10(6):721– 734, December 2002. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Sobrinho:2012:TCD

[SQ12] João Luís Sobrinho and Tiago Quelhas. A theory for the connectivity discovered by routing protocols. IEEE/ACM Transactions on Networking, 20(3): 677–689, June 2012. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Saleh:2016:PAN

[SQ16] Bassel Saleh and Dongyu Qiu. Performance analysis of network-coding-based P2P live streaming systems. IEEE/ACM Transactions on Networking, 24(4):2140–2153, August 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Sarker:2020:CMN

[SQS20] Ankur Sarker, Chenxi Qiu, and Haiying Shen. Connectivity maintenance for next-generation decentralized vehicle platoon networks. IEEE/ ACM Transactions on Networking, 28(4):1449–1462, August 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2020.2986252.

Song:2009:NFF

[SQZ09] Han Hee Song, Lili Qiu, and Yin Zhang. NetQuest: a flexible framework for large-scale network measurement. IEEE/ACM Transactions on Networking, 17(1):106–119, February 2009. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Sivarajan:1994:LNB

[SR94] Kumar N. Sivarajan and Rajiv Ramaswami. Lightwave networks based on de Bruijn graphs. IEEE/ ACM Transactions on Networking, 2(1):70–79, February 1994. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1994-2-1/
p70-sivarajan/.

Siwko:2001:CAC

[SR01] J. Siwko and I. Rubin.

Connection admission control for capacity-varying networks with stochastic capacity change times.

IEEE/ACM Transactions on Networking, 9(3):351– 360, June 2001. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Smith:2002:FSV

[SR02] Mark A. Smith and K. K. Ramakrishnan. Formal specification and verification of safety and performance of TCP selective acknowledgment. IEEE/ ACM Transactions on Networking, 10(2):193–207, April 2002. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Sundaresan:2014:CVM

[SR14] Karthikeyan Sundaresan and Sampath Rangarajan. Cooperation versus multiplexing: multicast scheduling algorithms for OFDMA relay networks. IEEE/ACM Transactions on Networking, 22(3):756– 769, June 2014. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Shan:2018:EMM

[SR18] Danfeng Shan and Fengyuan Ren. ECN marking with micro-burst traffic: Problem, analysis, and improvement. IEEE/ACM Transactions on Networking, 26(4):1533–1546, August 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Sengupta:2010:NCA

[SRB10] Sudipta Sengupta, Shravan Rayanchu, and Suman Banerjee. Network codingaware routing in wireless networks. IEEE/ ACM Transactions on Networking, 18(4):1158–1170, August 2010. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Sun:2020:PIA

[SRB^{+}20] Yahui Sun, Daniel Rehfeldt, Marcus Brazil, Doreen Thomas, and Saman Halgamuge. A physaruminspired algorithm for minimum-cost relay node placement in wireless sensor networks. IEEE/ ACM Transactions on Networking, 28(2):681–694, April 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL https://

dl.acm.org/doi/abs/10.
1109/TNET.2020.2971770.

Sanabria-Russo:2017:HEM

[SRBBG17] Luis Sanabria-Russo, Jaume Barcelo, Boris Bellalta, and Francesco Gringoli. A high efficiency MAC protocol for WLANs: Providing fairness in dense scenarios. IEEE/ACM Transactions on Networking, 25(1): 492–505, February 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Shan:2020:OMM

[SRC^{+}20] Danfeng Shan, Fengyuan Ren, Peng Cheng, Ran Shu, and Chuanxiong Guo. Observing and mitigating micro-burst traffic in data center networks. IEEE/ ACM Transactions on Networking, 28(1):98–111, February 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL https://

                                        dl.acm.org/doi/abs/10.
                                        1109/TNET.2019.2953793.

Sule:2019:SCB

[SRCDL19] Oladele Theophilus Sule, Roberto Rojas-Cessa, Ziqian Dong, and Chuan-Bi Lin. A split-central-buffered loadbalancing Clos-network switch with in-order forwarding. IEEE/ACM Transactions on Networking, 27(2):467–476, April 2019. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Salani:2023:DSP

[SRCT23] Matteo Salani, Cristina Rottondi, Leopoldo Ceré, and Massimo Tornatore. Dual-stage planning for elastic optical networks integrating machine-learning-assisted QoT estimation. IEEE/ACM Transactions on Networking, 31(3):1293– 1307, June 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2022.3213970.

Stutzbach:2009:USU

[SRD^{+}09] Daniel Stutzbach, Reza Rejaie, Nick Duffield, Subhabrata Sen, and Walter Willinger. On unbiased sampling for unstructured peer-to-peer networks. IEEE/ACM Transactions on Networking, 17 (2):377–390, April 2009. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Saifullah:2018:LPW

[SRI^{+}18] Abusayeed Saifullah, Mahbubur Rahman, Dali Ismail, Chenyang Lu, Jie Liu, and Ranveer Chandra. Low-power wide-area network over white spaces. IEEE/ACM Transactions on Networking, 26(4):1893– 1906, August 2018. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Sadeh:2022:OWL

[SRK22] Yaniv Sadeh, Ori Rottenstreich, and Haim Kaplan. Optimal weighted load balancing in TCAMs.

IEEE/ACM Transactions on Networking, 30(3):985– 998, June 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2021.3140124.

Sanchez:2023:AAC

[SRMB^{+}23] Sara Garcia Sanchez, Guillem Reus-Muns, Carlos Bocanegra, Yanyu Li, Ufuk Muncuk, Yousof Naderi, Yanzhi Wang, Stratis Ioannidis, and Kaushik Roy Chowdhury. AirNN: Overthe-air computation for neural networks via reconfigurable intelligent surfaces. IEEE/ACM Transactions on Networking, 31 (6):2470–2482, 2023. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL https:/

/dl.acm.org/doi/10.1109/
TNET.2022.3225883.

Secci:2011:PEM

[SRP^{+}11] Stefano Secci, Jean-Louis Rougier, Achille Pattavina, Fioravante Patrone, and Guido Maier. Peering equilibrium multipath routing: a game theory framework for Internet peering settlements. IEEE/ACM Transactions on Networking, 19 (2):419–432, April 2011. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Sarshar:2008:LLW

[SRR08] Nima Sarshar, Behnam A.

Rezaei, and Vwani P. Roychowdhury. Low latency wireless ad hoc networking: power and bandwidth challenges and a solution. IEEE/ACM Transactions on Networking, 16 (2):335–346, April 2008. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Shaikh:2001:EIS

[SRS01] Anees Shaikh, Jennifer Rexford, and Kang G. Shin. Evaluating the impact of stale link state on quality-of-service routing. IEEE/ACM Transactions on Networking, 9(2): 162–176, 2001. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/articles/
journals/ton/2001-9-2/
p162-shaikh/p162-shaikh.
pdf;   http://www.acm.
org/pubs/citations/journals/
ton/2001-9-2/p162-shaikh/

.

Sarkar:2003:MFL

[SRS03] Uttam K. Sarkar, Subramanian Ramakrishnan, and Dilip Sarkar. Modeling full-length video using Markov-modulated Gamma-based framework. IEEE/ACM Transactions on Networking, 11(4):638– 649, August 2003. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Stutzbach:2008:CUO

[SRS08] Daniel Stutzbach, Reza Rejaie, and Subhabrata Sen. Characterizing unstructured overlay topologies in modern P2P filesharing systems. IEEE/ ACM Transactions on Networking, 16(2):267–280, April 2008. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Shabtai:2021:RAS

[SRS21] Galia Shabtai, Danny Raz, and Yuval Shavitt. Risk aware stochastic placement of cloud services. IEEE/ACM Transactions on Networking, 29(2):805– 820, April 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2021.3052962.

Sharon:1993:SSS

[SS93] Oran Sharon and Adrian Segall. A simple scheme for slot reuse without latency for a dual bus configuration. IEEE/ACM Transactions on Networking, 1(1):96–104, February 1993. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1993-1-1/
p96-sharon/.

Sharon:1994:ESR

[SS94a] Oran Sharon and Adrian Segall. On the efficiency of slot reuse in the Dual Bus configuration. IEEE/ ACM Transactions on Networking, 2(1):89–100, February 1994. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1994-2-1/
p89-sharon/.

Sharon:1994:SSR

[SS94b] Oran Sharon and Adrian Segall. Schemes for slot reuse in CRMA. IEEE/ ACM Transactions on Networking, 2(3):269–278, June 1994. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1994-2-3/
p269-sharon/.

Smith:1996:IGP

[SS96] Peter J. Smith and Mansoor Shafi. The impact of G.826 on the performance of transport systems. IEEE/ACM Transactions on Networking, 4(4):604–614, August 1996. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/
                                        journals/ton/1996-4-4/
                                        p604-smith/.

Shroff:1998:ILC

[SS98] Ness B. Shroff and Mischa Schwartz. Improved loss calculations at an ATM multiplexer. IEEE/ ACM Transactions on Networking, 6(4):411–421, August 1998. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/
                                        journals/ton/1998-6-4/
                                        p411-shroff/.

Srinivasan:2003:AFE

[SS03] R. Srinivasan and Arun K. Somani. On achieving fairness and efficiency in highspeed shared medium access. IEEE/ACM Transactions on Networking, 11(1): 111–124, February 2003. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Sridharan:2004:BAO

[SS04a] Ashwin Sridharan and Kumar N. Sivarajan. Blocking in all-optical networks. IEEE/ACM Transactions on Networking, 12(2):384– 397, April 2004. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Steinder:2004:PFL

[SS04b] Małgorzata Steinder and Adarshpal S. Sethi. Probabilistic fault localization in communication systems using belief networks. IEEE/ ACM Transactions on Networking, 12(5):809–822, October 2004. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Sun:2005:EED

[SS05] Wei Sun and Kang G. Shin.

End-to-end delay bounds for traffic aggregates under guaranteed-rate scheduling algorithms. IEEE/ ACM Transactions on Networking, 13(5):1188–1201, October 2005. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Shakkottai:2006:ENP

[SS06] Srinivas Shakkottai and R. Srikant. Economics of network pricing with multiple ISPs. IEEE/ACM Transactions on Networking, 14(6):1233–1245, December 2006. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Sundaresan:2007:UML

[SS07] Karthikeyan Sundaresan and Raghupathy Sivakumar. A unified MAC layer framework for ad-hoc networks with smart antennas. IEEE/ACM Transactions on Networking, 15(3):546– 559, June 2007. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Shpungin:2009:LEF

[SS09] Hanan Shpungin and Michael Segal. Low-energy faulttolerant bounded-hop broadcast in wireless networks. IEEE/ACM Transactions on Networking, 17(2):582– 590, April 2009. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Shpungin:2010:NOM

[SS10] Hanan Shpungin and Michael Segal. Near-optimal multicriteria spanner constructions in wireless ad hoc networks. IEEE/ACM Transactions on Networking, 18 (6):1963–1976, December 2010. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Sermpezis:2016:ECP

[SS16] Pavlos Sermpezis and Thrasyvoulos Spyropoulos. Effects of content popularity on the performance of content-centric opportunistic networking: an analytical approach and applications. IEEE/ACM Transactions on Networking, 24 (6):3354–3368, December 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Sankaran:2017:DAS

[SS17] Ganesh C. Sankaran and Krishna M. Sivalingam. Design and analysis of scheduling algorithms for optically groomed data center networks. IEEE/ ACM Transactions on Networking, 25(6):3282–3293, December 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Sivaraman:2021:DMA

[SS21] Vignesh Sivaraman and Biplab Sikdar. A defense mechanism against timing attacks on user privacy in ICN. IEEE/ACM Transactions on Networking, 29(6):2709–2722, December 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2021.3097536.

Subramanian:2008:BSN

[SSA08] Sundar Subramanian, Sanjay Shakkottai, and Ari Arapostathis. Broadcasting in sensor networks: the role of local information. IEEE/ ACM Transactions on Networking, 16(5):1133–1146, October 2008. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Singh:2011:PCC

[SSA11] Chandramani Singh, Saswati Sarkar, and Alireza Aram. Provider-customer coalitional games. IEEE/ ACM Transactions on Networking, 19(5):1528–1542, October 2011. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Singh:2012:CPS

[SSAK12] Chandramani Singh, Saswati Sarkar, Alireza Aram, and Anurag Kumar. Cooperative profit sharing in coalition-based resource allocation in wireless networks. IEEE/ACM Transactions on Networking, 20 (1):69–83, February 2012. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Skelly:1993:HMV

[SSD93] Paul Skelly, Mischa Schwartz, and Sudhir Dixit. A histogram-based model for video traffic behavior in an ATM multiplexer. IEEE/ ACM Transactions on Networking, 1(4):446–459, August 1993. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1993-1-4/
p446-skelly/.

Solano:2008:LSR

[SSFM08] Fernando Solano, Thomas Stidsen, Ramon Fabregat, and Jose Luis Marzo. Label space reduction in MPLS networks: how much can a single stacked label do? IEEE/ACM Transactions on Networking, 16(6):1308– 1320, December 2008. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Shirali-Shahreza:2018:DIE

[SSG18] Sajad Shirali-Shahreza and Yashar Ganjali. Delayed installation and expedited eviction: an alternative approach to reduce flow table occupancy in SDN switches. IEEE/ ACM Transactions on Networking, 26(4):1547–1561, August 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Su:2023:IRT

[SSH^{+}23] Jian Su, Zhengguo Sheng, Chenxi Huang, Gang Li, Alex X. Liu, and Zhangjie Fu. Identifying RFID tags in collisions. IEEE/ ACM Transactions on Networking, 31(4):1507– 1520, 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2022.3219016.

Sharma:2011:OAR

[SSHK11] Sushant Sharma, Yi Shi, Y. Thomas Hou, and Sastry Kompella. An optimal algorithm for relay node assignment in cooperative ad hoc networks. IEEE/ACM Transactions on Networking, 19(3):879– 892, June 2011. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Singh:2017:TTW

[SSK^{+}17] Shailendra Singh, Karthikeyan Sundaresan, Srikanth V. Krishnamurthy, Xinyu Zhang, Mohammad Ali Amir Khojastepour, and Sampath Rangarajan. TRINITY: Tailoring wireless transmission strategies to user profiles in enterprise wireless networks. IEEE/ACM Transactions on Networking, 25(2):910–924, April 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Shakkottai:2003:BTC

[SSM03] Sanjay Shakkottai, R. Srikant, and Sean P. Meyn. Bounds on the throughput of congestion controllers in the presence of feedback delay. IEEE/ACM Transactions on Networking, 11(6): 972–981, December 2003. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Singhal:2006:OMM

[SSM06] Narendra K. Singhal, Laxman H. Sahasrabuddhe, and Biswanath Mukherjee. Optimal multicasting of multiple light-trees of different bandwidth granularities in a WDM mesh network with sparse splitting capabilities. IEEE/ ACM Transactions on Networking, 14(5):1104–1117, October 2006. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Stahlbuhk:2020:TMU

[SSM20] Thomas Stahlbuhk, Brooke Shrader, and Eytan Modiano. Throughput maximization in uncooperative spectrum sharing networks. IEEE/ACM Transactions on Networking, 28(6):2517–2530, December 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2020.3012273.

Sabnis:2023:GGD

[SSN^{+}23] Anirudh Sabnis, Tareq Si Salem, Giovanni Neglia, Michele Garetto, Emilio Leonardi, and Ramesh K. Sitaraman. GRADES: Gradient descent for similarity caching. IEEE/ACM Transactions on Networking, 31(1):30–41, 2023. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1109/
                                        TNET.2022.3187044.

Sapountzis:2017:UAH

[SSNS17] Nikolaos Sapountzis, Thrasyvoulos Spyropoulos, Navid Nikaein, and Umer Salim. User association in Het-Nets: Impact of traffic differentiation and backhaul limitations. IEEE/ACM Transactions on Networking, 25(6):3396–3410, December 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Sung:2011:TSD

[SSR^{+}11] Yu-Wei Eric Sung, Xin Sun, Sanjay G. Rao, Geoffrey G. Xie, and David A. Maltz. Towards systematic design of enterprise networks. IEEE/ACM Transactions on Networking, 19 (3):695–708, June 2011. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Shahbaz:2020:ESR

[SSR^{+}20] Muhammad Shahbaz, Lalith Suresh, Jennifer Rexford, Nick Feamster, Ori Rottenstreich, and Mukesh Hira. Elmo: Source routed multicast for public clouds. IEEE/ACM Transactions on Networking, 28(6):2587–2600, December 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2020.3020869.

Sahneh:2013:GEM

[SSV13] Faryad Darabi Sahneh, Caterina Scoglio, and Piet Van Mieghem. Generalized epidemic mean-field model for spreading processes over multilayer complex networks. IEEE/ ACM Transactions on Networking, 21(5):1609–1620, October 2013. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Shetty:2010:IQR

[SSW10] Nikhil Shetty, Galina Schwartz, and Jean Walrand. Internet QoS and regulations. IEEE/ACM Transactions on Networking, 18(6):1725–1737, December 2010. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Shin:2013:FCW

[SSWK13] Ji-Yong Shin, Emin Gn Sirer, Hakim Weatherspoon, and Darko Kirovski. On the feasibility of completely wireless datacenters. IEEE/ACM Transactions on Networking, 21(5): 1666–1679, October 2013. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Shi:2019:DGR

[SSY19] Junyang Shi, Mo Sha, and Zhicheng Yang. Distributed graph routing and scheduling for industrial wireless sensoractuator networks. IEEE/ ACM Transactions on Networking, 27(4):1669–1682, August 2019. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Stoica:2003:CSF

[SSZ03] Ion Stoica, Scott Shenker, and Hui Zhang. Corestateless fair queueing: a scalable architecture to approximate fair bandwidth allocations in high-speed networks. IEEE/ACM Transactions on Networking, 11(1):33–46, February 2003. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Sun:2005:PCC

[SSZ05] Xuehong Sun, Sartaj K.

Sahni, and Yiqiang Q. Zhao. Packet classification consuming small amount of memory. IEEE/ACM Transactions on Networking, 13(5):1135–1145, October 2005. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Shavitt:2004:BBS

[ST04] Yuval Shavitt and Tomer Tankel. Big-bang simulation for embedding network distances in Euclidean space. IEEE/ACM Transactions on Networking, 12 (6):993–1006, December 2004. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Sarkar:2005:FDC

[ST05] Saswati Sarkar and Leandros Tassiulas. Fair distributed congestion control in multirate multicast networks. IEEE/ACM Transactions on Networking, 13(1):121–133, February 2005. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Shavitt:2008:HEI

[ST08] Yuval Shavitt and Tomer Tankel. Hyperbolic embedding of Internet graph for distance estimation and overlay construction. IEEE/ACM Transactions on Networking, 16(1):25– 36, February 2008. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Sharma:2009:EBD

[ST09] Shrutivandana Sharma and Demosthenis Teneketzis. An externalities-based decentralized optimal power allocation algorithm for wireless networks. IEEE/ ACM Transactions on Networking, 17(6):1819–1831, December 2009. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Song:2013:AAB

[ST13] Haoyu Song and Jonathan S. Turner. ABC: adaptive binary cuttings for multidimensional packet classification. IEEE/ACM Transactions on Networking, 21 (1):98–109, February 2013. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Sorooshyari:2012:PCC

[STC12] Siamak Sorooshyari, Chee Wei Tan, and Mung Chiang. Power control for cognitive radio networks: axioms, algorithms, and analysis. IEEE/ACM Transactions on Networking, 20(3):878– 891, June 2012. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Stefanakos:2008:RRN

[Ste08] Stamatis Stefanakos. Reliable routings in networks with generalized link failure events. IEEE/ACM Transactions on Networking, 16(6):1331–1339, December 2008. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Shad:2001:DSA

[STKL01] Faisal Shad, Terence D.

Todd, Vytas Kezys, and John Litva. Dynamic slot allocation (DSA) in indoor SDMA/TDMA using smart antenna basestation. IEEE/ACM Transactions on Networking, 9 (1):69–81, 2001. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/articles/
journals/ton/2001-9-1/
p69-shad/p69-shad.pdf;
http://www.acm.org/pubs/
citations/journals/ton/
2001-9-1/p69-shad/.

Sun:2004:SMK

[STL04] Yan Sun, Wade Trappe, and K. J. Ray Liu. A scalable multicast key management scheme for heterogeneous wireless networks. IEEE/ACM Transactions on Networking, 12(4):653– 666, August 2004. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Seibert:2012:IWI

[STM^{+}12] Jeff Seibert, Ruben Torres, Marco Mellia, Maurizio M. Munafo, Cristina Nita-Rotaru, and Sanjay Rao. The Internet-wide impact of P2P traffic localization on ISP profitability. IEEE/ACM Transactions on Networking, 20(6): 1910–1923, December 2012. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Soh:2013:RID

[STQ13] De Wen Soh, Wee Peng Tay, and Tony Q. S. Quek. Randomized information dissemination in dynamic environments. IEEE/ACM Transactions on Networking, 21(3):681–691, June 2013. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Su:2015:TTS

[Su15] Yi-Sheng Su. Topologytransparent scheduling via the Chinese remainder theorem. IEEE/ACM Transactions on Networking, 23(5):1416–1429, October 2015. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Spiteri:2020:BNO

[SUS20] Kevin Spiteri, Rahul Urgaonkar, and Ramesh K. Sitaraman. BOLA: Nearoptimal bitrate adaptation for online videos. IEEE/ ACM Transactions on Networking, 28(4):1698–1711, August 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2020.2996964.

Shreedhar:1996:EFQ

[SV96] M. Shreedhar and George Varghese. Efficient fair queueing using deficit round-robin. IEEE/ ACM Transactions on Networking, 4(3):375–385, June 1996. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1996-4-3/
p375-shreedhar/.

Stiliadis:1998:EFQ

[SV98a] Dimitrios Stiliadis and Anujan Varma. Efficient fair queueing algorithms for packet-switched networks. IEEE/ACM Transactions on Networking, 6(2):175– 185, April 1998. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/

                                        journals/ton/1998-6-2/
                                        p175-stiliadis/.

Stiliadis:1998:LSG

[SV98b] Dimitrios Stiliadis and Anujan Varma. Latencyrate servers: a general model for analysis of traffic scheduling algorithms. IEEE/ACM Transactions on Networking, 6(5):611–624, October 1998. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/
                                        journals/ton/1998-6-5/
                                        p611-stiliadis/.

Stiliadis:1998:RSD

[SV98c] Dimitrios Stiliadis and Anujan Varma. Rateproportional servers: a design methodology for fair queueing algorithms. IEEE/ACM Transactions on Networking, 6(2):164– 174, April 1998. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/
                                        journals/ton/1998-6-2/
                                        p164-stiliadis/.

Sharma:1999:OBM

[SV99] Supriya Sharma and Yannis Viniotis. Optimal buffer management policies for shared-buffer ATM switches. IEEE/ACM Transactions on Networking, 7(4):575–587, August 1999. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1999-7-4/
p575-sharma/.

Sundaresan:2006:CQP

[SV06] Rajesh Sundaresan and Sergio Verdú. Capacity of queues via pointprocess channels. IEEE/ ACM Transactions on Networking, 14(SI):2697–2709, June 2006. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Sorour:2011:ANC

[SV11] Sameh Sorour and Shahrokh Valaee. An adaptive network coded retransmission scheme for single-hop wireless multicast broadcast services. IEEE/ACM Transactions on Networking, 19(3):869–878, June 2011. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Sorour:2015:CDM

[SV15] Sameh Sorour and Shahrokh Valaee. Completion delay minimization for instantly decodable network codes. IEEE/ACM Transactions on Networking, 23(5):1553– 1567, October 2015. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Song:2016:IAR

[SVG16] Yang Song, Arun Venkataramani, and Lixin Gao. Identifying and addressing reachability and policy attacks in “Secure” BGP. IEEE/ACM Transactions on Networking, 24(5):2969– 2982, October 2016. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Sobrinho:2016:SIR

[SVL^{+}16] Joao Luis Sobrinho, Laurent Vanbever, Franck Le, Andre Sousa, and Jennifer Rexford. Scaling the Internet routing system through distributed route aggregation. IEEE/ACM Transactions on Networking, 24(6): 3462–3476, December 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Sen:2004:APP

[SW04] Subhabrata Sen and Jia Wang. Analyzing peer-to-peer traffic across large networks. IEEE/ACM Transactions on Networking, 12 (2):219–232, April 2004. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Sun:2019:TAV

[SWH19] Elaine Y.-N. Sun, Hsiao-Chun Wu, and Scott C.- H. Huang. Theoretical analysis of various software-defined multiplexing codes. IEEE/ACM Transactions on Networking, 27(6):2444–2457, December 2019. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL https://

dl.acm.org/doi/abs/10.
1109/TNET.2019.2949823.

Swiderski:1996:APA

[Świ96] Jacek Świderski. Approximate performance analysis of real-time traffic over heavily loaded networks with timed token protocols. IEEE/ACM Transactions on Networking, 4(3):470– 478, June 1996. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1996-4-3/
p470-swiderski/.

Savage:2001:NSI

[SWKA01] Stefan Savage, David Wetherall, Anna Karlin, and Tom Anderson. Network support for IP traceback. IEEE/ACM Transactions on Networking, 9(3):226– 237, June 2001. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Shorten:2006:PSM

[SWL06] Robert Shorten, Fabian Wirth, and Douglas Leith. A positive systems model of TCP-like congestion control: asymptotic results. IEEE/ACM Transactions on Networking, 14 (3):616–629, 2006. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Sun:2018:WWC

[SWL^{+}18] Tianyuan Sun, Yongcai Wang, Deying Li, Zhaoquan Gu, and Jia Xu. WCS: Weighted component stitching for sparse network localization. IEEE/ ACM Transactions on Networking, 26(5):2242–2253, October 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Starobinski:2010:AOD

[SX10] David Starobinski and Weiyao Xiao. Asymptotically optimal data dissemination in multichannel wireless sensor networks: single radios suffice. IEEE/ACM Transactions on Networking, 18(3):695– 707, June 2010. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Sun:2016:ISA

[SX16] Xin Sun and Geoffrey G.

Xie. An integrated systematic approach to designing enterprise access control. IEEE/ACM Transactions on Networking, 24(6): 3508–3522, December 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Sun:2021:MAE

[SXEZ21] Wei Sun, Lisong Xu, Sebastian Elbaum, and Di Zhao. Model-agnostic and efficient exploration of numerical congestion control state space of real-world TCP implementations. IEEE/ ACM Transactions on Networking, 29(5):1990–2004, October 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2021.3078161.

Sung:2008:LSI

[SXLL08] Minho Sung, Jun Xu, Jun Li, and Li Li. Large-scale IP traceback in high-speed Internet: practical techniques and informationtheoretic foundation. IEEE/ ACM Transactions on Networking, 16(6):1253–1266, December 2008. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Shi:2023:CCA

[SXQ^{+}23] Qilong Shi, Yuchen Xu, Jiuhua Qi, Wenjun Li, Tong Yang, Yang Xu, and Yi Wang. Cuckoo counter: Adaptive structure of counters for accurate frequency and top-k estimation. IEEE/ACM Transactions on Networking, 31(4): 1854–1869, 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2022.3232098.

Sarkar:2009:HWO

[SYDM09] Suman Sarkar, Hong-Hsu Yen, Sudhir Dixit, and Biswanath Mukherjee. Hybrid wireless-optical broadband access network (WOBAN): network planning using Lagrangean relaxation. IEEE/ACM Transactions on Networking, 17(4):1094–1105, August 2009. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Shi:2022:QAI

[SYG^{+}22] Zhiguo Shi, Guang Yang, Xiaowen Gong, Shibo He, and Jiming Chen. Qualityaware incentive mechanisms under social influences in data crowdsourcing. IEEE/ACM Transactions on Networking, 30(1):176–189, February 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2021.3105427.

Sirivianos:2009:REI

[SYJ09] Michael Sirivianos, Xiaowei Yang, and Stanislaw Jarecki. Robust and efficient incentives for cooperative content distribution. IEEE/ACM Transactions on Networking, 17(6): 1766–1779, December 2009. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Shangguan:2017:SST

[SYL^{+}17] Longfei Shangguan, Zheng Yang, Alex X. Liu, Zimu Zhou, and Yunhao Liu. STPP: Spatial-temporal phase profiling-based method for relative RFID tag localization. IEEE/ACM Transactions on Networking, 25(1):596–609, February 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Shen:2001:EPC

[SYP01] Xiaojun Shen, Fan Yang, and Yi Pan. Equivalent permutation capabilities between time-division optical omega networks and non-optical extra-stage omega networks. IEEE/ ACM Transactions on Networking, 9(4):518–524, August 2001. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Shen:2005:SRL

[SYR05] Lu Shen, Xi Yang, and Byrav Ramamurthy. Shared risk link group (SRLG)- diverse path provisioning under hybrid service level agreements in wavelengthrouted optical mesh networks. IEEE/ACM Transactions on Networking, 13 (4):918–931, August 2005. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Song:2022:EEP

[SYW^{+}22] Guanglei Song, Jiahai Yang, Zhiliang Wang, Lin He, Jinlei Lin, Long Pan, Chenxin Duan, and Xiaowen Quan. DET: Enabling efficient probing of IPv6 active addresses. IEEE/ACM Transactions on Networking, 30(4):1629– 1643, 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2022.3145040.

Shen:2019:OMD

[SYZP19] Kaiming Shen, Wei Yu, Licheng Zhao, and Daniel P. Palomar. Optimization of MIMO device-to-device networks via matrix fractional programming: a minorization– maximization approach. IEEE/ACM Transactions on Networking, 27(5):2164– 2177, October 2019. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Shalom:2007:EAM

[SZ07] Mordechai Shalom and Shmuel Zaks. A 10/7 + ϵ approximation for minimizing the number of ADMs in SONET rings. IEEE/ ACM Transactions on Networking, 15(6):1593–1602, December 2007. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Sarikaya:2008:SPT

[SZ08] Behcet Sarikaya and Xiao Zheng. SIP paging and tracking of wireless LAN hosts for VoIP. IEEE/ ACM Transactions on Networking, 16(3):539–548, June 2008. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Shen:2022:COS

[SZ22] Zhirong Shen and Guanglin Zhang. Competitive online stay-or-switch algorithms with minimum commitment and switching cost. IEEE/ACM Transactions on Networking, 30(6): 2804–2817, 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2022.3183142.

Sarkar:2009:DRI

[SZG09] Rik Sarkar, Xianjin Zhu, and Jie Gao. Double rulings for information brokerage in sensor networks. IEEE/ACM Transactions on Networking, 17(6):1902– 1915, December 2009. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Sen:2013:DIA

[SZG^{+}13] Sayandeep Sen, Tan Zhang, Syed Gilani, Shreesha Srinath, Suman Banerjee, and Sateesh Addepalli. Design and implementation of an “approximate” communication system for wireless media applications. IEEE/ ACM Transactions on Networking, 21(4):1035–1048, August 2013. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Salehi:1998:SSV

[SZKT98] James D. Salehi, Shi-Li Zhang, Jim Kurose, and Don Towsley. Supporting stored video: reducing rate variability and end-to-end resource requirements through optimal smoothing. IEEE/ ACM Transactions on Networking, 6(4):397–410, August 1998. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1998-6-4/
p397-salehi/.

Su:2014:EAV

[SZL^{+}14] Sen Su, Zhongbao Zhang, Alex X. Liu, Xiang Cheng, Yiwen Wang, and Xinchao Zhao. Energy-aware virtual network embedding. IEEE/ACM Transactions on Networking, 22(5):1607– 1620, October 2014. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Song:2008:CSB

[SZM08] Lei Song, Jing Zhang, and Biswanath Mukherjee. A comprehensive study on backup-bandwidth reprovisioning after network-state updates in survivable telecom mesh networks. IEEE/ ACM Transactions on Networking, 16(6):1366–1377, December 2008. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Shi:2017:AAM

[SZMD17] Yishuo Shi, Zhao Zhang, Yuchang Mo, and Ding-Zhu Du. Approximation algorithm for minimum weight fault-tolerant virtual backbone in unit disk graphs. IEEE/ACM Transactions on Networking, 25(2):925– 933, April 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Stoica:2000:HFS

[SZN00] Ion Stoica, Hui Zhang, and T. S. Eugene Ng. A hierarchical fair service curve algorithm for link-sharing, real-time, and priority services. IEEE/ACM Transactions on Networking, 8 (2):185–199, 2000. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/
                                        journals/ton/2000-8-2/
                                        p185-stoica/.

Sun:2001:PBU

[SZT01] Hairong Sun, Xinyu Zang, and Kishor S. Trivedi. Performance of broadcast and unknown server (BUS) in ATM LAN emulation. IEEE/ACM Transactions on Networking, 9(3):361– 372, June 2001. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Shi:2016:OAF

[SZW^{+}16] Weijie Shi, Linquan Zhang, Chuan Wu, Zongpeng Li, and Francis C. M. Lau. An online auction framework for dynamic resource provisioning in cloud computing. IEEE/ACM Transactions on Networking, 24(4): 2060–2073, August 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Shi:2022:POL

[SZWW22] Zhengkai Shi, Yipeng Zhou, Di Wu, and Chen Wang. PPVC: Online learning toward optimized video content caching. IEEE/ACM Transactions on Networking, 30(3):1029– 1044, June 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2021.3132038.

Szymanski:2016:ULL

[Szy16] Ted H. Szymanski. An ultra-low-latency guaranteedrate Internet for cloud services. IEEE/ACM Transactions on Networking, 24(1):123–136, February 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Traverso:2015:NFS

[TAB^{+}15] Stefano Traverso, Luca Abeni, Robert Birke, Csaba Kiraly, Emilio Leonardi, Renato Lo Cigno, and Marco Mellia. Neighborhood filtering strategies for overlay construction in P2P-TV systems: design and experimental comparison. IEEE/ACM Transactions on Networking, 23 (3):741–754, June 2015. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Tao:2008:RTM

[TAG08] Shu Tao, John Apostolopoulos, and Roch Guérin. Real-time monitoring of video quality in IP networks. IEEE/ ACM Transactions on Networking, 16(5):1052–1065, October 2008. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Tsai:1999:CGP

[TAH99] Wei K. Tsai, John K. Antonio, and Garng M. Huang. Complexity of gradient projection method for optimal routing in data networks. IEEE/ ACM Transactions on Networking, 7(6):897–905, December 1999. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/
                                        journals/ton/1999-7-6/
                                        p897-tsai/.

Tran:2017:AOR

[TAH17] Tien Tran, Min Kyung An, and Dung T. Huynh. Antenna orientation and range assignment algorithms in directional WSNs. IEEE/ ACM Transactions on Networking, 25(6):3368–3381, December 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Tang:2010:QDW

[TAJ^{+}10] Ao Tang, Lachlan L. H.

Andrew, Krister Jacobsson, Karl H. Johansson, Håkan Hjalmarsson, and Steven H. Low. Queue dynamics with window flow control. IEEE/ACM Transactions on Networking, 18(5):1422–1435, October 2010. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Tan:2016:OPC

[Tan16] Chee Wei Tan. Optimal power control in Rayleighfading heterogeneous wireless networks. IEEE/ACM Transactions on Networking, 24(2):940–953, April 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Tassiulas:1996:PFL

[Tas96] Leandros Tassiulas. Push forward link-level scheduling for network-wide performance. IEEE/ACM Transactions on Networking, 4(3):398–406, June 1996. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL

                                        http://www.acm.org/pubs/
                                        citations/journals/ton/
                                        1996-4-3/p398-tassiulas/

.

Tassiulas:1999:CTS

[Tas99] Leandros Tassiulas. Cutthrough switching, pipelining, and scheduling for network evacuation. IEEE/ ACM Transactions on Networking, 7(1):88–97, February 1999. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/
                                        journals/ton/1999-7-1/
                                        p88-tassiulas/.

Tinnirello:2010:RIE

[TB10] Ilenia Tinnirello and Giuseppe Bianchi. Rethinking the IEEE 802.11e EDCA performance modeling methodology. IEEE/ACM Transactions on Networking, 18 (2):540–553, April 2010. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Tariq:2013:ATD

[TBV^{+}13] Mukarram Bin Tariq, Kaushik Bhandankar, Vytautas Valancius, Amgad Zeitoun, Nick Feamster, and Mostafa Ammar. Answering: techniques and deployment experience. IEEE/ACM Transactions on Networking, 21(1):1–13, February 2013. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Thommes:2006:DPM

[TC06] Richard W. Thommes and Mark J. Coates. Deterministic packet marking for time-varying congestion price estimation. IEEE/ ACM Transactions on Networking, 14(3):592–602, 2006. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Triay:2013:ABP

[TCPV13] Joan Triay, Cristina Cervelló-Pastor, and Vinod M. Vokkarane. Analytical blocking probability model for hybrid immediate and advance reservations in optical WDM networks. IEEE/ACM Transactions on Networking, 21(6):1890– 1903, December 2013. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Tseng:2004:SCT

[TCS04] Yu-Chee Tseng, Yu-Chi Chueh, and Jang-Ping Sheu. Seamless channel transition for the staircase video broadcasting scheme. IEEE/ACM Transactions on Networking, 12(3):559– 571, June 2004. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Tan:2013:FAP

[TCS13] Chee Wei Tan, Mung Chiang, and R. Srikant. Fast algorithms and performance bounds for sum rate maximization in wireless networks. IEEE/ACM Transactions on Networking, 21(3):706–719, June 2013. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Tangari:2020:AAA

[TCTP20] Gioacchino Tangari, Marinos Charalambides, Daphne Tuncer, and George Pavlou. Accuracy-aware adaptive traffic monitoring for software dataplanes. IEEE/ ACM Transactions on Networking, 28(3):986–1001, June 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2020.2976952.

Towles:2003:GSS

[TD03] Brian Towles and William J. Dally. Guaranteed scheduling for switches with configuration overhead. IEEE/ ACM Transactions on Networking, 11(5):835–847, October 2003. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Tsai:1994:PAT

[TdWC^{+}94] Zsehong Tsai, Wen der Wang, Chien-Hwa Chiou, Jin-Fu Chang, and Lung-Sing Liang. Performance analysis of two echo control designs in ATM networks. IEEE/ACM Transactions on Networking, 2(1):30–39, February 1994. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1994-2-1/
p30-tsai/.

Tadrous:2016:OPC

[TE16] John Tadrous and Atilla Eryilmaz. On optimal proactive caching for mobile networks with demand uncertainties. IEEE/ ACM Transactions on Networking, 24(5):2715–2727, October 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Tadrous:2016:JSP

[TEE16] John Tadrous, Atilla Eryilmaz, and Hesham El Gamal. Joint smart pricing and proactive content caching for mobile services. IEEE/ACM Transactions on Networking, 24(4):2357– 2371, August 2016. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Traynor:2009:MAO

[TEML09] Patrick Traynor, William Enck, Patrick McDaniel, and Thomas La Porta. Mitigating attacks on open functionality in SMScapable cellular networks. IEEE/ACM Transactions on Networking, 17(1):40– 53, February 2009. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Tadrous:2019:ABS

[TES19] John Tadrous, Atilla Eryilmaz, and Ashutosh Sabharwal. Action-based scheduling: Leveraging app interactivity for scheduler efficiency. IEEE/ACM Transactions on Networking, 27(1):112–125, February 2019. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Tassiulas:1996:WPS

[TG96] Leandros Tassiulas and Leonidas Georgiadis. Any work-conserving policy stabilizes the ring with spatial re-use. IEEE/ACM Transactions on Networking, 4 (2):205–208, April 1996. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL

                                        http://www.acm.org/pubs/
                                        citations/journals/ton/
                                        1996-4-2/p205-tassiulas/

.

Tsybakov:1997:STA

[TG97] Boris Tsybakov and Nicoals D.

Georganas. On selfsimilar traffic in ATM queues: definitions, overflow probability bound, and cell delay distribution. IEEE/ACM Transactions on Networking, 5(3):397– 409, June 1997. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1997-5-3/
p397-tsybakov/.

Terdik:2009:LFF

[TG09] György Terdik and Tibor Gyires. Lévy flights and fractal modeling of Internet traffic. IEEE/ACM Transactions on Networking, 17(1):120–129, February 2009. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Tsanikidis:2021:PRS

[TG21] Christos Tsanikidis and Javad Ghaderi. On the power of randomization for scheduling realtime traffic in wireless networks. IEEE/ACM Transactions on Networking, 29(4):1703–1716, August 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2021.3072279.

Tsanikidis:2023:RSR

[TG23] Christos Tsanikidis and Javad Ghaderi. Randomized scheduling of realtime traffic in wireless networks over fading channels. IEEE/ACM Transactions on Networking, 31(4): 1688–1701, 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2022.3223315.

Trevisan:2020:FYE

[TGD^{+}20] Martino Trevisan, Danilo Giordano, Idilio Drago, Maurizio Matteo Munafò, and Marco Mellia. Five years at the edge: Watching Internet from the ISP network. IEEE/ ACM Transactions on Networking, 28(2):561–574, April 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL https://

                                        dl.acm.org/doi/abs/10.
                                        1109/TNET.2020.2967588.

Teixeira:2007:TBT

[TGRR07] Renata Teixeira, Timothy G. Griffin, Mauricio G. C. Resende, and Jennifer Rexford. TIE breaking: tunable interdomain egress selection. IEEE/ACM Transactions on Networking, 15(4):761– 774, August 2007. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Tabatabaee:2001:QPT

[TGT01] Vahid Tabatabaee, Leonidas Georgiadis, and Leandros Tassiulas. QoS provisioning and tracking fluid policies in input queueing switches. IEEE/ACM Transactions on Networking, 9(5):605– 617, October 2001. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Turletti:1996:VI

[TH96] Thierry Turletti and Christian Huitema. Videoconferencing on the Internet. IEEE/ACM Transactions on Networking, 4(3):340– 351, June 1996. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1996-4-3/
p340-turletti/.

Todd:1997:MMM

[TH97] Terence D. Todd and Ellen L. Hahne. Multiaccess mesh (multimesh) networks. IEEE/ACM Transactions on Networking, 5(2):181–189, April 1997. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL

http://www.acm.org/pubs/
citations/journals/ton/
1997-5-2/p181-todd/.

Tang:2021:HDY

[TH21] Ming Tang and Jianwei Huang. How do you earn money on live streaming Platforms? — A study of donation-based markets. IEEE/ACM Transactions on Networking, 29(4):1813–1826, August 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2021.3071488.

Tham:2001:UAF

[Tha01] Yiu Kwok Tham. A unified algorithmic framework for variable-rate TDM switching assignments. IEEE/ ACM Transactions on Networking, 9(5):662–668, October 2001. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Tham:2004:SSS

[Tha04] Yiu Kwok Tham. Scheduling satellite-switched timedivision multiple access with general switching modes. IEEE/ACM Transactions on Networking, 12 (4):645–652, August 2004. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Tapolcai:2014:SFF

[THBR14] János Tapolcai, Pin-Han Ho, Péter Babarczi, and Lajos Rónyai. On signalingfree failure dependent restoration in all-optical mesh networks. IEEE/ ACM Transactions on Networking, 22(4):1067–1078, August 2014. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Taylor:2005:RHC

[THDD05] David E. Taylor, Andreas Herkersdorf, Andreas Döring, and Gero Dittmann. Robust header compression (ROHC) in next-generation network processors. IEEE/ACM Transactions on Networking, 13(4):755–768, August 2005. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Traskov:2012:SNC

[THMK12] Danail Traskov, Michael Heindlmaier, Muriel Médard, and Ralf Koetter. Scheduling for network-coded multicast. IEEE/ACM Transactions on Networking, 20(5):1479–1488, October 2012. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Tassiulas:1994:OBC

[THP94] Leandros Tassiulas, Yao Chung Hung, and Shivendra S. Panwar. Optimal buffer control during congestion in an ATM network node.

IEEE/ACM Transactions on Networking, 2(4):374– 386, August 1994. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/
                                        journals/ton/1994-2-4/
                                        p374-tassiulas/.

Tapolcai:2012:NWL

[THRW12] János Tapolcai, Pin-Han Ho, Lajos Rónyai, and Bin Wu. Network-wide local unambiguous failure localization (NWL-UFL) via monitoring trails. IEEE/ ACM Transactions on Networking, 20(6):1762–1773, December 2012. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Tian:2005:SAD

[Tia05]  Yu-Ping Tian.  Stabil-

ity analysis and design of the second-order congestion control for networks with heterogeneous delays. IEEE/ACM Transactions on Networking, 13(5):1082– 1093, October 2005. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Tassiulas:1995:PMS

[TJ95] Leandros Tassiulas and Jinoo Joung. Performance measures and scheduling policies in ring networks. IEEE/ACM Transactions on Networking, 3(5):576– 584, October 1995. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL http://

www.acm.org/pubs/citations/
journals/ton/1995-3-5/
p576-tassiulas/.

Tutuncuoglu:2023:OLR

[TJD23] Feridun Tütüncüoğlu, Slaana Jošilo, and György Dán. Online learning for rateadaptive task offloading under latency constraints in serverless edge computing. IEEE/ACM Transactions on Networking, 31(2):695– 709, April 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2022.3197669.

Tan:2021:AOO

[TJHL21] Haisheng Tan, Shaofeng H.-C. Jiang, Zhenhua Han, and Mingxia Li. Asymptotically optimal online caching on multiple caches with relaying and bypassing. IEEE/ ACM Transactions on Networking, 29(4):1841–1852, August 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2021.3077115.

Tan:2019:JOC

[TJL^{+}19] Haisheng Tan, Shaofeng H.-C. Jiang, Yupeng Li, Xiang-Yang Li, Chenzi Zhang, Zhenhua Han, and Francis Chi Moon Lau. Joint online coflow routing and scheduling in data center networks. IEEE/ ACM Transactions on Networking, 27(5):1771–1786, October 2019. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Tan:2012:GGR

[TK12] Guang Tan and Anne-Marie Kermarrec. Greedy geographic routing in largescale sensor networks: a minimum network decomposition approach. IEEE/ ACM Transactions on Networking, 20(3):864–877, June 2012. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Trajanovski:2015:FCR

[TKI^{+}15] Stojan Trajanovski, Fernando A. Kuipers, Aleksandar Ilić, Jon Crowcroft, and Piet Van Mieghem. Finding critical regions and regiondisjoint paths in a network. IEEE/ACM Transactions on Networking, 23(3):908– 921, June 2015. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Talak:2020:IAI

[TKM20a] Rajat Talak, Sertac Karaman, and Eytan Modiano. Improving age of information in wireless networks with perfect channel state information. IEEE/ ACM Transactions on Networking, 28(4):1765–1778, August 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2020.2996237.

Talak:2020:OIF

[TKM20b] Rajat Talak, Sertac Karaman, and Eytan Modiano. Optimizing information freshness in wireless networks under general interference constraints. IEEE/ ACM Transactions on Networking, 28(1):15–28, February 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL https://

dl.acm.org/doi/abs/10.
1109/TNET.2019.2946481.

Taleb:2006:REF

[TKN06] Tarik Taleb, Nei Kato, and Yoshiaki Nemoto. RE-FWA: an efficient and fair congestion control scheme for LEO satellite networks. IEEE/ACM Transactions on Networking, 14(5):1031– 1044, October 2006. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Thomas:2020:LLF

[TKXP20] Yannis Thomas, Merkourios Karaliopoulos, George Xylomenos, and George C. Polyzos. Low latency friendliness for multipath TCP. IEEE/ACM Transactions on Networking, 28(1):248–261, February 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL

                                        https://dl.acm.org/doi/
                                        abs/10.1109/TNET.2019.
                                        2961759.

Tel:1994:SAN

[TKZ94] Gerard Tel, Ephraim Korach, and Shmuel Zaks. Synchronizing ABD networks. IEEE/ACM Transactions on Networking, 2 (1):66–69, February 1994. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL

                                        http://www.acm.org/pubs/
                                        citations/journals/ton/
                                        1994-2-1/p66-tel/.

Tinnakornsrisuphap:2006:ABH

[TL06] Peerapol Tinnakornsrisuphap and Richard J. La. Asymptotic behavior of heterogeneous TCP flows and RED gateway. IEEE/ ACM Transactions on Networking, 14(1):108–120, February 2006. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Tian:2016:TAS

[TL16] Guibin Tian and Yong Liu.

Towards agile and smooth video adaptation in HTTP adaptive streaming. IEEE/ ACM Transactions on Networking, 24(4):2386–2399, August 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Tabatabaee:2022:DRR

[TL22] Seyed Mohammadhossein Tabatabaee and Jean-Yves Le Boudec. Deficit roundrobin: a second network calculus analysis. IEEE/ACM Transactions on Networking, 30(5):2216– 2230, 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2022.3164772.

Tu:2016:DPC

[TLP^{+}16] Guan-Hua Tu, Yuanjie Li, Chunyi Peng, Chi-Yu Li, and Songwu Lu. Detecting problematic control-plane protocol interactions in mobile networks. IEEE/ACM Transactions on Networking, 24(2):1209–1222, April 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Tekin:2012:ACG

[TLS^{+}12] Cem Tekin, Mingyan Liu, Richard Southwell, Jianwei Huang, and Sahand Haji Ali Ahmad. Atomic congestion games on graphs and their applications in networking. IEEE/ACM Transactions on Networking, 20(5):1541–1552, October 2012. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Tan:2013:OCP

[TM13] Bo Tan and Laurent Massoulié. Optimal content placement for peer-to-peer video-on-demand systems. IEEE/ACM Transactions on Networking, 21(2):566– 579, April 2013. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Tillman:2019:GCF

[TMGB19] Balint Tillman, Athina Markopoulou, Minas Gjoka, and Carter T. Buttsc. 2K+ graph construction framework: Targeting joint degree matrix and beyond. IEEE/ACM Transactions on Networking, 27(2):591– 606, April 2019. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Tridandapani:1997:CSM

[TMH97] Srini B. Tridandapani, Biswanath Mukherjee, and Geir Hallingstad. Channel sharing in multi-hop WDM lightwave networks: do we need more channels? IEEE/ACM Transactions on Networking, 5(5):719–727, October 1997. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1997-5-5/
p719-tridandapani/.

Thatte:2011:PMA

[TMH11] Gautam Thatte, Urbashi Mitra, and John Heidemann. Parametric methods for anomaly detection in aggregate traffic. IEEE/ACM Transactions on Networking, 19(2):512– 525, April 2011. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Tian:2018:OED

[TML^{+}18] Chen Tian, Ali Munir, Alex X. Liu, Jie Yang, and Yangming Zhao. Open-Function: an extensible data plane abstraction protocol for platformindependent software-defined middleboxes. IEEE/ACM Transactions on Networking, 26(3):1488–1501, June 2018. CODEN IEANEP.

ISSN 1063-6692 (print), 1558-2566 (electronic).

Thomas:2022:WCD

[TML22] Ludovic Thomas, Ahlem Mifdaoui, and Jean-Yves Le Boudec. Worst-case delay bounds in time-sensitive networks with packet replication and elimination. IEEE/ACM Transactions on Networking, 30(6):2701– 2715, 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2022.3180763.

Thomopoulos:2001:LAT

[TMMS01] Efstratios Thomopoulos, Louise E. Moser, and Peter M. Melliar-Smith. Latency analysis of the totem single-ring protocol. IEEE/ ACM Transactions on Networking, 9(5):669–680, October 2001. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Tornatore:2007:WND

[TMP07] Massimo Tornatore, Guido Maier, and Achille Pattavina. WDM network design by ILP models based on flow aggregation. IEEE/ACM Transactions on Networking, 15(3):709– 720, June 2007. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Tantiprasut:1997:APS

[TNF97] Duke Tantiprasut, John Neil, and Craig Farrell. ASN.1 protocol specification for use with arbitrary encoding schemes. IEEE/ACM Transactions on Networking, 5(4):502– 513, August 1997. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1997-5-4/
p502-tantiprasut/.

Thekkath:1993:INP

[TNML93] Chandramohan A. Thekkath, Thu D. Nguyen, Evelyn Moy, and Edward D. Lazowska. Implementing network protocols at user level. IEEE/ACM Transactions on Networking, 1(5):554–565, October 1993. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1993-1-5/
p554-thekkath/.

Tague:2011:JAT

[TNRP11] Patrick Tague, Sidharth Nabar, James A. Ritcey, and Radha Poovendran. Jamming-aware traffic allocation for multiple-path routing using portfolio selection. IEEE/ACM Transactions on Networking, 19(1):184–194, February 2011. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Todd:1994:TGN

[Tod94] Terence D. Todd. The token grid network. IEEE/ ACM Transactions on Networking, 2(3):279–287, June 1994. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/
                                        journals/ton/1994-2-3/
                                        p279-todd/.

Towsley:2006:AIT

[Tow06a] Don Towsley. Abstracts from the IEEE transactions on information theory, special issue, June 2006. IEEE/ACM Transactions on Networking, 14 (4):675–682, August 2006. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Towsley:2006:E

[Tow06b] Don Towsley. Editorial. IEEE/ACM Transactions on Networking, 14(4):673, August 2006. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Tan:2009:ERT

[TPC09] Chee Wei Tan, Daniel P. Palomar, and Mung Chiang. Energy-robustness tradeoff in cellular network power control. IEEE/ACM Transactions on Networking, 17(3):912–925, June 2009. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Tang:2018:MDA

[TPW^{+}18] Ming Tang, Haitian Pang, Shou Wang, Lin Gao, Jianwei Huang, and Lifeng Sun. Multi-dimensional auction mechanisms for crowdsourced mobile video streaming. IEEE/ACM Transactions on Networking, 26(5):2062–2075, October 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Thaler:1998:UNM

[TR98] David G. Thaler and Chinya V. Ravishankar. Using name-based mappings to increase hit rates. IEEE/ACM Transactions on Networking, 6(1):1–14, February 1998. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1998-6-1/
p1-thaler/.

Tune:2017:CST

[TR17] Paul Tune and Matthew Roughan. Controlled synthesis of traffic matrices. IEEE/ACM Transactions on Networking, 25(3):1582– 1592, June 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Treurniet:2011:NAC

[Tre11] Joanne Treurniet. A network activity classification schema and its application to scan detection. IEEE/ ACM Transactions on Networking, 19(5):1396–1404, October 2011. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Trestian:2010:GIP

[TRKN10] Ionut Trestian, Supranamaya Ranjan, Aleksandar Kuzmanovic, and Antonio Nucci. Googling the Internet: profiling Internet endpoints via the World Wide Web. IEEE/ACM Transactions on Networking, 18 (2):666–679, April 2010. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Trestian:2012:TMD

[TRKN12] Ionut Trestian, Supranamaya Ranjan, Aleksandar Kuzmanovic, and Antonio Nucci. Taming the mobile data deluge with drop zones. IEEE/ACM Transactions on Networking, 20(4):1010–1023, August 2012. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Tapolcai:2020:FER

[TRVG20] János Tapolcai, Lajos Rónyai, Balázs Vass, and László Gyimóthi. Fast enumeration of regional link failures caused by disasters with limited size. IEEE/ ACM Transactions on Networking, 28(6):2421–2434, December 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2020.3009297.

Tickoo:2008:MQC

[TS08] Omesh Tickoo and Biplab Sikdar. Modeling queueing and channel access delay in unsaturated IEEE 802.11 random access MAC based wireless networks. IEEE/ACM Transactions on Networking, 16(4):878– 891, August 2008. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Turkcu:2009:PON

[TS09] Onur Turkcu and Suresh Subramaniam. Performance of optical networks with limited reconfigurability. IEEE/ACM Transactions on Networking, 17(6): 2002–2013, December 2009. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Turkcu:2014:OWW

[TS14] Onur Turkcu and Suresh Subramaniam. Optimal wavebanding in WDM ring networks. IEEE/ ACM Transactions on Networking, 22(1):179–190, February 2014. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Teixeira:2008:IHP

[TSGR08] Renata Teixeira, Aman Shaikh, Timothy G. Griffin, and Jennifer Rexford. Impact of hot-potato routing changes in IP networks. IEEE/ACM Transactions on Networking, 16(6):1295– 1307, December 2008. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Tariq:2021:ATC

[TSN^{+}21] Isfar Tariq, Rajat Sen, Thomas Novlan, Salam Akoum, Milap Majmundar, Gustavo de Veciana, and Sanjay Shakkottai. Autotuning for cellular scheduling through bandit-learning and low-dimensional clustering. IEEE/ACM Transactions on Networking, 29(5):1933–1947, October 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2021.3077455.

Tapparello:2014:DCT

[TSR14] Cristiano Tapparello, Osvaldo Simeone, and Michele Rossi. Dynamic compressiontransmission for energyharvesting multihop networks with correlated sources. IEEE/ACM Transactions on Networking, 22(6):1729–1741, December 2014. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Tan:2014:RDB

[TSS14] Jian Tan, Swapna, and Shroff. Retransmission delays with bounded packets: Power-law body and exponential tail. IEEE/ACM Transactions on Networking, 22(1):27–38, February 2014. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Thimmaraju:2021:PNP

[TSS21] Kashyap Thimmaraju, Liron Schiff, and Stefan Schmid. Preacher: Network policy checker for adversarial environments. IEEE/ ACM Transactions on Networking, 29(5):2087–2100, October 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2021.3078143.

Taylor:2007:CPC

[TT07] David E. Taylor and Jonathan S. Turner. Class-Bench: a packet classification benchmark. IEEE/ ACM Transactions on Networking, 15(3):499–511, June 2007. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Tabatabaee:2009:MCN

[TT09] Vahid Tabatabaee and Leandros Tassiulas. MNCM: a critical node matching approach to scheduling for input buffered switches with no speedup. IEEE/ ACM Transactions on Networking, 17(1):294–304, February 2009. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Tan:2017:EHW

[TT17] Liansheng Tan and Shengda Tang. Energy harvesting wireless sensor node with temporal death: Novel models and analyses. IEEE/ACM Transactions on Networking, 25(2):896– 909, April 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Tseng:2019:RSH

[TTCT19] Shih-Hao Tseng, Ao Tang, Gagan L. Choudhury, and Simon Tse. Routing stability in hybrid softwaredefined networks. IEEE/ ACM Transactions on Networking, 27(2):790–804, April 2019. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Tran:2022:SDM

[TTM22] Hai-Anh Tran, Duc Tran, and Abdelhamid Mellouk. State-dependent multiconstraint topology configuration for software-defined service overlay networks. IEEE/ACM Transactions on Networking, 30(5):1986– 2001, 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2022.3155475.

Tripathi:2023:IFM

[TTM23] Vishrant Tripathi, Rajat Talak, and Eytan Modiano. Information freshness in multihop wireless networks. IEEE/ACM Transactions on Networking, 31(2):784– 799, April 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2022.3201751.

Turner:2009:SPG

[Tur09] Jonathan S. Turner. Strong performance guarantees for asynchronous buffered crossbar scheduler. IEEE/ ACM Transactions on Networking, 17(4):1017–1028, August 2009. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Tajer:2010:MDG

[TW10] Ali Tajer and Xiaodong Wang. Multiuser diversity gain in cognitive networks. IEEE/ACM Transactions on Networking, 18 (6):1766–1779, December 2010. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Tsai:2022:UAM

[TW22] Cho-Hsin Tsai and Chih-Chun Wang. Unifying AoI minimization and remote estimation — optimal sensor/controller coordination with random two-way delay. IEEE/ ACM Transactions on Networking, 30(1):229–242, February 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2021.3111495.

Tsai:2023:DOO

[TW23] Cho-Hsin Tsai and Chih-Chun Wang. Distributionoblivious online algorithms for age-of-information penalty minimization. IEEE/ ACM Transactions on Networking, 31(4):1779– 1794, 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2022.3230009.

Teymoori:2024:LNH

[TWH24] Peyman Teymoori, Michael Welzl, and David A. Hayes. LGCC: a novel high-throughput and low delay paradigm shift in multi-hop congestion control. IEEE/ACM Transactions on Networking, 32 (1):761–776, 2024. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2023.3301291.

Tapolcai:2011:NAF

[TWHR11] János Tapolcai, Bin Wu, Pin-Han Ho, and Lajos Rónyai. A novel approach for failure localization in all-optical mesh networks. IEEE/ACM Transactions on Networking, 19(1):275– 285, February 2011. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Tang:2004:UCT

[TWL04] Ao Tang, Jiantao Wang, and Steven H. Low. Understanding CHOKe: throughput and spatial characteristics. IEEE/ACM Transactions on Networking, 12 (4):694–707, August 2004. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Trappe:2005:RAC

[TWL05] Wade Trappe, Yuke Wang, and K. J. Ray Liu. Resource-aware conference key establishment for heterogeneous networks. IEEE/ACM Transactions on Networking, 13(1):134– 146, February 2005. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Tang:2006:CIT

[TWL06] Ao Tang, Jiantao Wang, and Steven H. Low. Counterintuitive throughput behaviors in networks under end-to-end control. IEEE/ACM Transactions on Networking, 14(2):355– 368, April 2006. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Tong:2021:CFL

[TWL^{+}21] Xinyu Tong, Yang Wan, Qianru Li, Xiaohua Tian, and Xinbing Wang. CSI fingerprinting localization with low human efforts. IEEE/ACM Transactions on Networking, 29(1):372– 385, February 2021. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL https:/

/dl.acm.org/doi/10.1109/
TNET.2020.3035210.

Tong:2022:CPC

[TWL22] Shuai Tong, Jiliang Wang, and Yunhao Liu. Combating packet collisions using non-stationary signal scaling in LPWANs. IEEE/ACM Transactions on Networking, 30(3):1104– 1117, June 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2021.3131704.

Tang:2007:EHC

[TWLC07] Ao Tang, Jiantao Wang, Steven H. Low, and Mung Chiang. Equilibrium of heterogeneous congestion control: existence and uniqueness. IEEE/ACM Transactions on Networking, 15 (4):824–837, August 2007. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Tang:2010:EHC

[TWLC10] Ao Tang, Xiaoliang Wei, Steven H. Low, and Mung Chiang. Equilibrium of heterogeneous congestion control: optimality and stability. IEEE/ACM Transactions on Networking, 18(3): 844–857, June 2010. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Tang:2020:CSS

[TWN^{+}20] Bin Tang, Xiaoliang Wang, Cam-Tu Nguyen, Baoliu Ye, and Sanglu Lu. Construction of subexponentialsize optical priority queues with switches and fiber delay lines. IEEE/ACM Transactions on Networking, 28(1):336–346, February 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL https://

dl.acm.org/doi/abs/10.
1109/TNET.2019.2960402.

Tan:2022:AVT

[TWS^{+}22] Qingfeng Tan, Xuebin Wang, Wei Shi, Jian Tang, and Zhihong Tian. An anonymity vulnerability in Tor. IEEE/ACM Transactions on Networking, 30(6): 2574–2587, 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2022.3174003.

Tong:2017:AIM

[TWTD17] Guangmo Tong, Weili Wu, Shaojie Tang, and Ding-Zhu Du. Adaptive influence maximization in dynamic social networks. IEEE/ACM Transactions on Networking, 25(1):112– 125, February 2017. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Tang:2019:TKD

[TWWG19] Guoming Tang, Huan Wang, Kui Wu, and Deke Guo. Tapping the knowledge of dynamic traffic demands for optimal CDN design. IEEE/ACM Transactions on Networking, 27 (1):98–111, February 2019. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Tian:2020:TEP

[TWY^{+}20] Ying Tian, Zhiliang Wang, Xia Yin, Xingang Shi, Yingya Guo, Haijun Geng, and Jiahai Yang. Traffic engineering in partially deployed segment routing over IPv6 network with deep reinforcement learning. IEEE/ ACM Transactions on Networking, 28(4):1573–1586, August 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2020.2987866.

Tang:2008:OLC

[TX08] Xueyan Tang and Jianliang Xu. Optimizing lifetime for continuous data aggregation with precision guarantees in wireless sensor networks. IEEE/ACM Transactions on Networking, 16 (4):904–917, August 2008. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Tang:2023:HPI

[TXHL23] Lu Tang, Yao Xiao, Qun Huang, and Patrick P. C. Lee. A high-performance invertible sketch for networkwide superspreader detection. IEEE/ACM Transactions on Networking, 31(2): 724–737, April 2023. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1109/
                                        TNET.2022.3198738.

Tan:2012:EDF

[TXL^{+}12] Rui Tan, Guoliang Xing, Benyuan Liu, Jianping Wang, and Xiaohua Jia. Exploiting data fusion to improve the coverage of wireless sensor networks. IEEE/ACM Transactions on Networking, 20(2):450– 462, April 2012. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Tu:2018:NPB

[TXL^{+}18] Zhen Tu, Fengli Xu, Yong Li, Pengyu Zhang, and Depeng Jin. A new privacy breach: User trajectory recovery from aggregated mobility data. IEEE/ACM Transactions on Networking, 26(3):1446–1459, June 2018. CODEN IEANEP.

ISSN 1063-6692 (print), 1558-2566 (electronic).

Trimponias:2019:NCT

[TXW^{+}19] George Trimponias, Yan Xiao, Xiaorui Wu, Hong Xu, and Yanhui Geng. Node-constrained traffic engineering: Theory and applications. IEEE/ACM Transactions on Networking, 27(4):1344–1358, August 2019. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Tian:2021:EIT

[TXW^{+}21] Jiazheng Tian, Kun Xie, Xin Wang, Gaogang Xie, Kenli Li, Jigang Wen, Dafang Zhang, and Jiannong Cao. Efficiently inferring top-k largest monitoring data entries based on discrete tensor completion. IEEE/ACM Transactions on Networking, 29(6):2737–2750, December 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2021.3103527.

Tang:2018:LAC

[TY18] Bin Tang and Shenghao Yang. An LDPC approach for chunked network codes. IEEE/ACM Transactions on Networking, 26(1):605– 617, February 2018. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Tan:2016:TAU

[TYJ16] Guang Tan, Zhimeng Yin, and Hongbo Jiang. Trap array: a unified model for scalability evaluation of geometric routing. IEEE/ ACM Transactions on Networking, 24(1):328–341, February 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

To:1994:MIE

[TYL94] Philip P. To, Tak-Shing P. Yum, and Yiu-Wing Leung. Multistar implementation of expandable shufflenets. IEEE/ACM Transactions on Networking, 2(4):345– 351, August 1994. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/
                                        journals/ton/1994-2-4/
                                        p345-to/.

Tsai:2009:VCB

[TYLH09] Ming-Jer Tsai, Hong-Yen Yang, Bing-Hong Liu, and Wen-Qian Huang. Virtualcoordinate-based deliveryguaranteed routing protocol in wireless sensor networks. IEEE/ACM Transactions on Networking, 17(4):1228–1241, August 2009. CODEN IEANEP.

ISSN 1063-6692 (print), 1558-2566 (electronic).

Tytgat:2015:AEV

[TYP^{+}15] Lieven Tytgat, Opher Yaron, Sofie Pollin, Ingrid Moerman, and Piet Demeester. Analysis and experimental verification of frequency-based interference avoidance mechanisms in IEEE 802.15.4. IEEE/ACM Transactions on Networking, 23(2):369– 382, April 2015. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Teh:2023:EQS

[TZCB23] Min Yee Teh, Shizhen Zhao, Peirui Cao, and Keren Bergman. Enabling quasi-static reconfigurable networks with robust topology engineering. IEEE/ACM Transactions on Networking, 31(3):1056– 1070, June 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2022.3210534.

Tian:2023:DSM

[TZL23] Xiang Tian, Baoxian Zhang, and Cheng Li. Distributed stable multisource global broadcast for SINR-based wireless multihop networks. IEEE/ACM Transactions on Networking, 31(2):620– 633, April 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2022.3198331.

Tian:2024:MEC

[TZL^{+}24] Xiaohua Tian, Fengyuan Zhu, Hao Li, Mingwei Ouyang, Luwei Feng, Xinyu Tong, and Xinbing Wang. MobiScatter: Enhancing capacity in drone-assisted high-concurrency backscatter networks. IEEE/ ACM Transactions on Networking, 32(1):535– 549, 2024. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2023.3290168.

Thejaswi:2010:DOS

[TZP^{+}10] P. S. Chandrashekhar Thejaswi, Junshan Zhang, Man-On Pun, H. Vincent Poor, and Dong Zheng. Distributed opportunistic scheduling with twolevel probing. IEEE/ ACM Transactions on Networking, 18(5):1464–1477, October 2010. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Tang:2023:HUB

[TZPZ23] Shaofei Tang, Sicheng Zhao, Xiaoqin Pan, and Zuqing Zhu. How to use in-band network telemetry wisely: Network-wise orchestration of Sel-INT. IEEE/ACM Transactions on Networking, 31(1):421– 435, 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2022.3194086.

Tan:2021:RBC

[TZX^{+}21] Haisheng Tan, Chi Zhang, Chao Xu, Yupeng Li, Zhenhua Han, and Xiang-Yang Li. Regularizationbased coflow scheduling in optical circuit switches. IEEE/ACM Transactions on Networking, 29(3):1280– 1293, June 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2021.3058164.

Tu:2022:RAR

[TZX^{+}22] Huaqing Tu, Gongming Zhao, Hongli Xu, Yangming Zhao, and Yutong Zhai. A robustness-aware real-time SFC routing update scheme in multitenant clouds. IEEE/ ACM Transactions on Networking, 30(3):1230–1244, June 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2021.3137418.

Teng:2014:ELI

[TZZ^{+}14] Jin Teng, Boying Zhang, Junda Zhu, Xinfeng Li, Dong Xuan, and Yuan F. Zheng. EV-Loc: integrating electronic and visual signals for accurate localization. IEEE/ACM Transactions on Networking, 22(4):1285–1296, August 2014. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Uysal-Biyikoglu:2002:EEP

[UBPE02] Elif Uysal-Biyikoglu, Balaji Prabhakar, and Abbas El Gamal. Energy-efficient packet transmission over a wireless link. IEEE/ACM Transactions on Networking, 10(4):487–499, August 2002. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Urgaonkar:2011:NCR

[UN11] Rahul Urgaonkar and Michael J. Neely. Network capacity region and minimum energy function for a delay-tolerant mobile ad hoc network. IEEE/ ACM Transactions on Networking, 19(4):1137–1150, August 2011. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Uddin:2014:JRM

[URZ^{+}14] Md. Forkan Uddin, Catherine Rosenberg, Weihua Zhuang, Patrick Mitran, and Andre Girard. Joint routing and medium access control in fixed random access wireless multihop networks. IEEE/ACM Transactions on Networking, 22 (1):80–93, February 2014. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Ural:1993:OLT

[UZ93] Hasan Ural and Keqin Zhu.

Optimal length test sequence generation using distinguishing sequences. IEEE/ACM Transactions on Networking, 1(3):358– 371, June 1993. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1993-1-3/
p358-ural/.

Vuran:2006:SCB

[VA06] Mehmet C. Vuran and Ian F. Akyildiz. Spatial correlation-based collaborative medium access control in wireless sensor networks. IEEE/ACM Transactions on Networking, 14 (2):316–329, April 2006. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Vuran:2007:MAM

[VA07] Mehmet C. Vuran and Ian F. Akyildiz. A-MAC: adaptive medium access control for next generation wireless terminals. IEEE/ACM Transactions on Networking, 15(3):574– 587, June 2007. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Vuran:2009:ECW

[VA09] Mehmet C. Vuran and Ian F. Akyildiz. Error control in wireless sensor networks: a cross layer analysis. IEEE/ACM Transactions on Networking, 17(4): 1186–1199, August 2009. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Vasudevan:2013:EAN

[VAGT13] Sudarshan Vasudevan, Micah Adler, Dennis Goeckel, and Don Towsley. Efficient algorithms for neighbor discovery in wireless networks. IEEE/ACM Transactions on Networking, 21(1):69– 83, February 2013. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Valaee:2001:REW

[Val01] Shahrokh Valaee. A recursive estimator of worst-case burstiness. IEEE/ACM Transactions on Networking, 9(2):211–222, 2001. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL

http://www.acm.org/pubs/
articles/journals/ton/
2001-9-2/p211-valaee/
p211-valaee.pdf;   http:
//www.acm.org/pubs/citations/
journals/ton/2001-9-2/
p211-valaee/.

Valente:2007:EGS

[Val07] Paolo Valente. Exact GPS simulation and optimal fair scheduling with logarithmic complexity. IEEE/ACM Transactions on Networking, 15(6):1454–1466, December 2007. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Veloso:2006:HCL

[VAM^{+}06] Eveline Veloso, Virgílio Almeida, Wagner Meira, Jr., Azer Bestavros, and Shudong Jin. A hierarchical characterization of a live streaming media workload. IEEE/ACM Transactions on Networking, 14(1): 133–146, February 2006. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

VanHoudt:2017:EBR

[Van17] Benny Van Houdt. Explicit back-off rates for achieving target throughputs in CSMA/CA networks. IEEE/ACM Transactions on Networking, 25 (2):765–778, April 2017.

CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

VanHoudt:2019:RWS

[Van19] Benny Van Houdt. Randomized work stealing versus sharing in largescale systems with nonexponential job sizes. IEEE/ ACM Transactions on Networking, 27(5):2137–2149, October 2019. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Vickers:2000:SAM

[VAS00] Brett J. Vickers, Célio Albuquerque, and Tatsuya Suda. Sourceadaptive multilayered multicast algorithms for realtime video distribution. IEEE/ACM Transactions on Networking, 8(6):720– 733, 2000. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/articles/
                                        journals/ton/2000-8-6/
                                        p720-vickers/p720-vickers.
                                        pdf;   http://www.acm.
                                        org/pubs/citations/journals/
                                        ton/2000-8-6/p720-vickers/

.

Varvarigos:1994:PHR

[VB94] Emmanouel A. Varvarigos and Dimitri P. Bertsekas. Performance of hypercube routing schemes with or without buffering. IEEE/ACM Transactions on Networking, 2(3):299– 311, June 1994. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1994-2-3/
p299-varvarigos/.

Valenza:2017:CAC

[VBC^{+}17] Fulvio Valenza, Cataldo Basile, Daniele Canavese, Antonio Lioy, Fulvio Valenza, Cataldo Basile, Daniele Canavese, and Antonio Lioy. Classification and analysis of communication protection policy anomalies. IEEE/ACM Transactions on Networking, 25(5): 2601–2614, October 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Vlachou:2017:HCC

[VBHT17] Christina Vlachou, Albert Banchs, Julien Herzen, and Patrick Thiran. How CSMA/CA with deferral affects performance and dynamics in power-line communications. IEEE/ ACM Transactions on Networking, 25(1):250–263, February 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Venkataraman:2012:QVQ

[VC12] Mukundan Venkataraman and Mainak Chatterjee. Quantifying video-QoE degradations of Internet links. IEEE/ACM Transactions on Networking, 20 (2):396–407, April 2012. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Venkataraman:2014:EIP

[VC14] Mukundan Venkataraman and Mainak Chatterjee. Effects of Internet path selection on video-QoE: analysis and improvements. IEEE/ACM Transactions on Networking, 22(3):689– 702, June 2014. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Vissicchio:2015:IRP

[VCD15] Stefano Vissicchio, Luca Cittadini, and Giuseppe Di Battista. On iBGP routing policies. IEEE/ ACM Transactions on Networking, 23(1):227–240, February 2015. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Vlajic:2004:PAD

[VCM04] Natalija Vlajic, Charalambos D. Charalambous, and Dimitrios Makrakis. Performance aspects of data broadcast in wireless networks with user retrials. IEEE/ACM Transactions on Networking, 12(4):620– 633, August 2004. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Vissicchio:2017:SER

[VCVC17] Stefano Vissicchio, Luca Cittadini, Stefano Vissicchio, and Luca Cittadini. Safe, efficient, and robust SDN updates by combining rule replacements and additions. IEEE/ ACM Transactions on Networking, 25(5):3102–3115, October 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

vanDuijn:2022:ATA

[vDJJ^{+}22] Ingo van Duijn, Peter Gjøl Jensen, Jesper Stenbjerg Jensen, Troels Beck Krøgh, Jonas Sand Madsen, Stefan Schmid, Jirí Srba, and Marc Tom Thorgersen. Automata-theoretic approach to verification of MPLS networks under link failures. IEEE/ ACM Transactions on Networking, 30(2):766–781, April 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2021.3126572.

vanDoorn:1993:BPL

[vDP93] Erik A. van Doorn and Frans J. M. Panken. Blocking probabilities in a loss system with arrivals in geometrically distributed batches and heterogeneous service requirements. IEEE/ACM Transactions on Networking, 1(6):664–667, December 1993. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/
                                        journals/ton/1993-1-6/
                                        p664-van_doorn/.

Varvello:2011:ESL

[VFBD11] Matteo Varvello, Stefano Ferrari, Ernst Biersack, and Christophe Diot. Exploring second life. IEEE/ ACM Transactions on Networking, 19(1):80–91, February 2011. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Varsamopoulos:2004:DAR

[VG04] Georgios Varsamopoulos and Sandeep K. S. Gupta. Dynamically adapting registration areas to user mobility and call patterns for efficient location management in PCS networks. IEEE/ACM Transactions on Networking, 12(5):837– 850, October 2004. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Valaee:2005:ERP

[VG05] Shahrokh Valaee and Jean-Charles Grégoire. An estimator of regulator parameters in a stochastic setting. IEEE/ACM Transactions on Networking, 13(6): 1376–1389, December 2005. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Vojnovic:2008:RWA

[VG08] Milan Vojnovic and Ayalvadi J. Ganesh. On the race of worms, alerts, and patches. IEEE/ACM Transactions on Networking, 16(5):1066–1079, October 2008. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Vojnovic:2010:SSE

[VGKG10] Milan Vojnović, Varun Gupta, Thomas Karagiannis, and Christos Gkantsidis. Sampling strategies for epidemic-style information dissemination. IEEE/ ACM Transactions on Networking, 18(4):1013–1025, August 2010. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Vaze:2014:DPA

[VGP14] Rahul Vaze, Rachit Garg, and Neetish Pathak. Dynamic power allocation for maximizing throughput in energy-harvesting communication system. IEEE/ ACM Transactions on Networking, 22(5):1621–1630, October 2014. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Vargas:1996:SPL

[VHNPM96] Cesar Vargas, Manjunath V. Hegde, Morteza Naraghi-Pour, and Paul S. Min. Shadow prices for LLR and ALBA. IEEE/ ACM Transactions on Networking, 4(5):796–807, October 1996. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1996-4-5/
p796-vargas/.

Vardoyan:2021:TSA

[VHT21] Gayane Vardoyan, C. V. Hollot, and Don Towsley. Towards stability analysis of data transport mechanisms: a fluid model and its applications. IEEE/ ACM Transactions on Networking, 29(4):1730–1744, August 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.

acm.org/doi/10.1109/TNET.
2021.3075837.

VanMieghem:2001:EM

[VHvdH01] Piet Van Mieghem, Gerard Hooghiemstra, and Remco van der Hofstad. On the efficiency of multicast. IEEE/ACM Transactions on Networking, 9(6): 719–732, December 2001. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Valls:2021:BDR

[VIT21] Víctor Valls, George Iosifidis, and Leandros Tassiulas. Birkhoff’s decomposition revisited: Sparse scheduling for high-speed circuit switches. IEEE/ ACM Transactions on Networking, 29(6):2399–2412, December 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2021.3088327.

VanDeVen:2014:BEH

[VJV14] P. M. Van De Ven, Augustus J. E. M. Janssen, and J. S. H. Van Leeuwaarden. Balancing exposed and hidden nodes in linear wireless networks. IEEE/ ACM Transactions on Networking, 22(5):1429–1443, October 2014. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

VanMieghem:2004:CEQ

[VK04] Piet Van Mieghem and Fernando A. Kuipers. Concepts of exact QoS routing algorithms. IEEE/ACM Transactions on Networking, 12(5):851–864, October 2004. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Vargaftik:2017:NPL

[VKO17] Shay Vargaftik, Isaac Keslassy, and Ariel Orda. No packet left behind: Avoiding starvation in dynamic topologies. IEEE/ ACM Transactions on Networking, 25(4):2571–2584, August 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Vargaftik:2020:LLB

[VKO20] Shay Vargaftik, Isaac Keslassy, and Ariel Orda. LSQ: Load balancing in large-scale heterogeneous systems with multiple dispatchers. IEEE/ACM Transactions on Networking, 28(3):1186–1198, June 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1109/TNET.2020.2980061.

Vasiliadis:2017:DIS

[VKPI17] Giorgos Vasiliadis, Lazaros Koromilas, Michalis Polychronakis, and Sotiris Ioannidis. Design and implementation of a stateful network packet processing framework for GPUs. IEEE/ACM Transactions on Networking, 25(1):610– 623, February 2017. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Varghese:1997:HHT

[VL97] George Varghese and Anthony Lauck. Hashed and hierarchical timing wheels: efficient data structures for implementing a timer facility. IEEE/ ACM Transactions on Networking, 5(6):824–834, December 1997. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1997-5-6/
p824-varghese/.

Varvarigos:1999:VCD

[VL99] Emmanouel A. Varvarigos and Jonathan P. Lang. A virtual circuit deflection protocol. IEEE/ACM Transactions on Networking, 7(3):335– 349, June 1999. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1999-7-3/
p335-varvarigos/.

Vojnovic:2005:LRB

[VL05] Milan Vojnović and Jean-Yves Le Boudec. On the long-run behavior of equation-based rate control. IEEE/ACM Transactions on Networking, 13(3): 568–581, June 2005. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Venkataramanan:2010:WSA

[VL10] V. J. Venkataramanan and Xiaojun Lin. On wireless scheduling algorithms for minimizing the queue-overflow probability. IEEE/ACM Transactions on Networking, 18(3):788– 801, June 2010. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Valls:2016:MWR

[VL16] Victor Valls and Douglas J. Leith. Max-weight revisited: Sequences of nonconvex optimizations solving convex optimizations. IEEE/ACM Transactions on Networking, 24(5):2676– 2689, October 2016. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Vyavahare:2017:MRN

[VLDM17] Pooja Vyavahare, Nutan Limaye, Ajit A. Diwan, and D. Manjunath.

On the maximum rate of networked computation in a capacitated network. IEEE/ACM Transactions on Networking, 25(4):2444– 2458, August 2017. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Vyavahare:2016:OEF

[VLM16] Pooja Vyavahare, Nutan Limaye, and D. Manjunath. Optimal embedding of functions for in-network computation: Complexity analysis and algorithms. IEEE/ACM Transactions on Networking, 24(4):2019– 2032, August 2016. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

VanMeter:2009:SDL

[VLMN09] Rodney Van Meter, Thaddeus D. Ladd, W. J. Munro, and Kae Nemoto. System design for a longline quantum repeater. IEEE/ACM Transactions on Networking, 17(3):1002– 1013, June 2009. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Varvello:2016:MPC

[VLZL16] Matteo Varvello, Rafael Laufer, Feixiong Zhang, and T. V. Lakshman. Multilayer packet classification with graphics processing units. IEEE/ACM Transactions on Networking, 24(5):2728–2741, October 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Veitch:2022:ILB

[VMCB22] Darryl Veitch, Sathiya Kumaran Mani, Yi Cao, and Paul Barford. iHorology: Lowering the barrier to microsecond-level Internet time. IEEE/ACM Transactions on Networking, 30(6): 2544–2558, 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2022.3174189.

Vaze:2020:MSS

[VN20] Rahul Vaze and Jayakrishnan Nair. Multiple server SRPT with speed scaling is competitive. IEEE/ ACM Transactions on Networking, 28(4):1739–1751, August 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2020.2993142.

Vaze:2022:SSP

[VN22] Rahul Vaze and Jayakrishnan Nair. Speed scaling on parallel servers with MapReduce type precedence constraints. IEEE/ ACM Transactions on Networking, 30(4):1509– 1524, 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2022.3142091.

Vyavahare:2022:SDE

[VNM22] Pooja Vyavahare, Jayakrishnan Nair, and D. Manjunath. Sponsored data: On the effect of ISP competition on pricing dynamics and content provider market structures. IEEE/ ACM Transactions on Networking, 30(5):2018– 2031, 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2022.3162856.

Varadarajan:2002:ESP

[VNS02]  Srivatsan  Varadarajan,

Hung Q. Ngo, and Jaideep Srivastava. Error spreading: a perception-driven approach to handling error in continuous media streaming. IEEE/ ACM Transactions on Networking, 10(1):139–152, February 2002. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Voice:2007:SMP

[Voi07] Thomas Voice. Stability of multi-path dual congestion control algorithms. IEEE/ ACM Transactions on Networking, 15(6):1231–1239, December 2007. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

VanMieghem:2009:VSN

[VOK09] Piet Van Mieghem, Jasmina Omic, and Robert Kooij. Virus spread in networks. IEEE/ACM Transactions on Networking, 17 (1):1–14, February 2009. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Voskoboynik:2017:NCS

[VPC17] Niv Voskoboynik, Haim H. Permuter, and Asaf Cohen. Network coding schemes for data exchange networks with arbitrary transmission delays. IEEE/ACM Transactions on Networking, 25 (3):1293–1309, June 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Vehel:2013:LDM

[VR13] Jacques Lévy Véhel and Michal Rams. Large deviation multifractal analysis of a class of additive processes with correlated nonstationary increments. IEEE/ ACM Transactions on Networking, 21(4):1309–1321, August 2013. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

vanRijswijk-Deij:2017:PIE

[vRDHSP17] Roland van Rijswijk-Deij, Kaspar Hageman, Anna Sperotto, and Aiko Pras. The performance impact of elliptic curve cryptography on DNSSEC validation. IEEE/ACM Transactions on Networking, 25 (2):738–750, April 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Veitch:2009:RSA

[VRK09] Darryl Veitch, Julien Ridoux, and Satish Babu Korada. Robust synchronization of absolute and difference clocks over networks. IEEE/ACM Transactions on Networking, 17 (2):417–430, April 2009. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Voronov:2024:FAD

[VRR24] Tomer Voronov, Danny Raz, and Ori Rottenstreich. A framework for anomaly detection in blockchain networks with sketches. IEEE/ACM Transactions on Networking, 32(1):686– 698, 2024. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.

acm.org/doi/10.1109/TNET.
2023.3298253.

vonRickenbach:2009:AMI

[vRWZ09] Pascal von Rickenbach, Roger Wattenhofer, and Aaron Zollinger. Algorithmic models of interference in wireless ad hoc and sensor networks. IEEE/ ACM Transactions on Networking, 17(1):172–185, February 2009. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Varvarigos:1997:RVC

[VS97] Emmanouel A. Varvarigos and Vishal Sharma. The ready-to-go virtual circuit protocol: a lossfree protocol for multigigabit networks using FIFO buffers. IEEE/ ACM Transactions on Networking, 5(5):705–718, October 1997. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1997-5-5/
p705-varvarigos/.

Vishwanath:2011:ALP

[VSR11] Arun Vishwanath, Vijay Sivaraman, and George N. Rouskas. Anomalous loss performance for mixed realtime and TCP traffic in routers with very small buffers. IEEE/ACM Transactions on Networking, 19 (4):933–946, August 2011. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Venkitasubramaniam:2012:GTA

[VT12] Parv Venkitasubramaniam and Lang Tong. A game-theoretic approach to anonymous networking. IEEE/ACM Transactions on Networking, 20(3):892– 905, June 2012. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Vass:2021:EMS

[VTBK21] Balázs Vass, János Tapolcai, and Erika R. Bérczi-Kovács. Enumerating maximal shared risk link groups of circular disk failures hitting k nodes. IEEE/ ACM Transactions on Networking, 29(4):1648–1661, August 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2021.3070100.

Vishwanath:2009:SRR

[VV09] Kashi Venkatesh Vishwanath and Amin Vahdat. Swing: realistic and responsive network traffic generation. IEEE/ACM Transactions on Networking, 17(3):712–725, June 2009. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Vissicchio:2017:SUH

[VVC^{+}17] Stefano Vissicchio, Laurent Vanbever, Luca Cittadini, Geoffrey G. Xie, and Olivier Bonaventure. Safe update of hybrid SDN networks. IEEE/ACM Transactions on Networking, 25 (3):1649–1662, June 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Vanbever:2012:LML

[VVP^{+}12] Laurent Vanbever, Stefano Vissicchio, Cristel Pelsser, Pierre Francois, and Olivier Bonaventure. Lossless migrations of link-state IGPs. IEEE/ACM Transactions on Networking, 20(6):1842– 1855, December 2012. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Vissicchio:2013:INA

[VVP^{+}13] Stefano Vissicchio, Laurent Vanbever, Cristel Pelsser, Luca Cittadini, Pierre Francois, and Olivier Bonaventure. Improving network agility with seamless BGP reconfigurations. IEEE/ACM Transactions on Networking, 21(3):990– 1002, June 2013. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

VanMieghem:2009:OPN

[VW09] Piet Van Mieghem and Huijuan Wang. The observable part of a network. IEEE/ACM Transactions on Networking, 17 (1):93–105, February 2009. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Vural:2017:CTD

[VWNT17] Serdar Vural, Ning Wang, Pirabakaran Navaratnam, and Rahim Tafazolli. Caching transient data in Internet content routers. IEEE/ ACM Transactions on Networking, 25(2):1048–1061, April 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Vadrevu:2014:DSP

[VWT^{+}14] Chaitanya S. K. Vadrevu, Rui Wang, Massimo Tornatore, Charles U. Martel, and Biswanath Mukherjee. Degraded service provisioning in mixed-line-rate WDM backbone networks using multipath routing. IEEE/ACM Transactions on Networking, 22(3):840– 849, June 2014. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Wang:2011:SCM

[WA11] Pu Wang and Ian F. Akyildiz. Spatial correlation and mobility-aware traffic modeling for wireless sensor networks. IEEE/ACM Transactions on Networking, 19(6):1860–1873, December 2011. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Wang:2004:WBT

[Wan04]  Xudong Wang.  Wide-

band TD-CDMA MAC with minimum-power allocation and rate- and BER-scheduling for wireless multimedia networks. IEEE/ACM Transactions on Networking, 12(1):103– 116, February 2004. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Wang:2011:PCO

[WB11] Peng Wang and Stephan Bohacek. Practical computation of optimal schedules in multihop wireless networks. IEEE/ACM Transactions on Networking, 19 (2):305–318, April 2011. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Wang:2005:IEP

[WBEGS05] Haining Wang, Abhijit Bose, Mohamed El-Gendy, and Kang G. Shin. IP Easy-pass: a light-weight network-edge resource access control. IEEE/ACM Transactions on Networking, 13(6):1247–1260, December 2005. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Wu:2018:CTA

[WBM^{+}18] Ning Wu, Yingjie Bi, Nithin Michael, Ao Tang, John C. Doyle, and Nikolai Matni. A controltheoretic approach to innetwork congestion management. IEEE/ACM Transactions on Networking, 26(6):2443–2456, December 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Wang:2011:LTP

[WBP^{+}11] Ruhai Wang, Scott C.

Burleigh, Paavan Parikh, Che-Jen Lin, and Bo Sun. Licklider transmission protocol (LTP)-based DTN for cislunar communications. IEEE/ACM Transactions on Networking, 19(2):359– 368, April 2011. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Wang:2016:CNC

[WBWV16] Sen Wang, Jun Bi, Jianping Wu, and Athanasios V. Vasilakos. CPHR: In-network caching for information-centric networking with partitioning and hash-routing. IEEE/ ACM Transactions on Networking, 24(5):2742–2755, October 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Wen:2017:RIF

[WBY^{+}17] Xitao Wen, Kai Bu, Bo Yang, Yan Chen, Li Erran Li, Xiaolin Chen, Jianfeng Yang, and Xue Leng. RuleScope: Inspecting forwarding faults for softwaredefined networking. IEEE/ ACM Transactions on Networking, 25(4):2347–2360, August 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Wang:2008:IGA

[WC08] Chen-Shu Wang and Ching-Ter Chang. Integrated genetic algorithm and goal programming for network topology design problem with multiple objectives and multiple criteria. IEEE/ACM Transactions on Networking, 16(3):680– 690, June 2008. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Widmer:2015:EIN

[WCAB15] Joerg Widmer, Andrea Capalbo, Antonio Fernández Anta, and Albert Banchs. Efficient interlayer network codes for fair layered multicast streaming. IEEE/ ACM Transactions on Networking, 23(4):1107–1120, August 2015. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Wu:2014:OAP

[WCC14] Shan-Hung Wu, Ming-Syan Chen, and Chung-Min Chen. Optimally adaptive power-saving protocols for ad hoc networks using the hyper quorum system. IEEE/ACM Transactions on Networking, 22(1):1–15, February 2014. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Wang:2018:DID

[WCCM18] An Wang, Wentao Chang, Songqing Chen, and Aziz Mohaisen. Delving into Internet DDoS attacks by botnets: Characterization and analysis. IEEE/ACM Transactions on Networking, 26(6):2843–2855, December 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Ward:1995:DLC

[WCH95] Christopher Ward, Cheong H.

Choi, and Thomas F. Hain. A data link control protocol for LEO satellite networks providing a reliable datagram service. IEEE/ ACM Transactions on Networking, 3(1):91–103, February 1995. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1995-3-1/
p91-ward/.

Wang:2024:UFB

[WCH^{+}24] Mei Wang, Jing Chen, Kun He, Ruozhou Yu, Ruiying Du, and Zhihao Qian. UFinAKA: Fingerprintbased authentication and key agreement with updatable blind credentials. IEEE/ACM Transactions on Networking, 32(2):1110– 1123, April 2024. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2023.3311130.

Wang:2020:RWF

[WCK^{+}20] Zhe Wang, Yifeng Cao, Linghe Kong, Guihai Chen, Jiadi Yu, Shaojie Tang, and Yingying Chen. Reference waveforms forward concurrent transmissions in ZigBee communications. IEEE/ACM Transactions on Networking, 28(4):1629–1642, August 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.

acm.org/doi/10.1109/TNET.
2020.2992271.

Wang:2022:TOP

[WCL^{+}22] Yuntao Wang, Weiwei Chen, Tom H. Luan, Zhou Su, Qichao Xu, Ruidong Li, and Nan Chen. Task offloading for post-disaster rescue in unmanned aerial vehicles networks. IEEE/ ACM Transactions on Networking, 30(4):1525– 1539, 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2022.3140796.

Wei:2021:CEF

[WCM^{+}21] Jianghong Wei, Xiaofeng Chen, Jianfeng Ma, Xuexian Hu, and Kui Ren. Communication-efficient and fine-grained forward-secure asynchronous messaging. IEEE/ACM Transactions on Networking, 29(5):2242– 2253, October 2021. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL https:/

/dl.acm.org/doi/10.1109/
TNET.2021.3084692.

Wang:2020:HFR

[WCQ^{+}20] Ge Wang, Haofan Cai, Chen Qian, Jinsong Han, Shouqian Shi, Xin Li, Han Ding, Wei Xi, and Jizhong Zhao. Hu-Fu: Replay-resilient RFID authentication. IEEE/ ACM Transactions on Networking, 28(2):547–560, April 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL https://

                                        dl.acm.org/doi/abs/10.
                                        1109/TNET.2020.2964290.

Wu:2017:QAE

[WCW^{+}17] Jiyan Wu, Bo Cheng, Ming Wang, Junliang Chen, Jiyan Wu, Bo Cheng, Ming Wang, and Junliang Chen. Quality-aware energy optimization in wireless video communication with multipath TCP. IEEE/ACM Transactions on Networking, 25(5):2701–2718, October 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Wang:2019:NNS

[WCW19] Tianheng Wang, Andrea Conti, and Moe Z. Win. Network navigation with scheduling: Distributed algorithms. IEEE/ACM Transactions on Networking, 27(4):1319–1329, August 2019. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Wang:2017:SWF

[WCWZ17] Wei Wang, Yingjie Chen, Lu Wang, and Qian Zhang. Sampleless Wi-Fi: Bringing low power to Wi-Fi communications. IEEE/ ACM Transactions on Networking, 25(3):1663–1672, June 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Wong:2000:ARC

[WCY00] Eric W. M. Wong, Andy K. M. Chan, and Tak-Shing Peter Yum. Analysis of rerouting in circuitswitched networks. IEEE/ ACM Transactions on Networking, 8(3):419–427, 2000. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL

http://www.acm.org/pubs/
citations/journals/ton/
2000-8-3/p419-wong/.

Wan:2004:MPM

[WCY04] Peng-Jun Wan, Gruia Cãlinescu, and Chih-Wei Yi. Minimum-power multicast routing in static ad hoc wireless networks. IEEE/ACM Transactions on Networking, 12(3):507– 514, June 2004. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Wu:2017:SDL

[WCZZ17] Chao Wu, Xu Chen, Wenwu Zhu, and Yaoxue Zhang. Socially-driven learning-based prefetching in mobile online social networks. IEEE/ACM Transactions on Networking, 25(4):2320–2333, August 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Weber:2005:RAM

[WD05] Steven Weber and Gustavo De Veciana. Rate adaptive multimedia streams: optimization and admission control. IEEE/ACM Transactions on Networking, 13(6):1275–1288, December 2005. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Wang:2022:DPM

[WD22] Xuehe Wang and Lingjie Duan. Dynamic pricing and mean field analysis for controlling age of information. IEEE/ACM Transactions on Networking, 30(6): 2588–2600, 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2022.3174114.

Wang:2015:DPL

[WDCL15] Jiliang Wang, Wei Dong, Zhichao Cao, and Yunhao Liu. On the delay performance in a largescale wireless sensor network: measurement, analysis, and implications. IEEE/ACM Transactions on Networking, 23(1):186– 197, February 2015. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Wang:2023:DTO

[WDL^{+}23] Juncheng Wang, Min Dong, Ben Liang, Gary Boudreau, and Hatem Abou-Zeid. Delay-tolerant OCO with long-term constraints: Algorithm and its application to network resource allocation. IEEE/ACM Transactions on Networking, 31(1):147– 163, 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2022.3188285.

Wang:2020:JOO

[WDR^{+}20] Xiong Wang, Qi Deng, Jing Ren, Mehdi Malboubi, Sheng Wang, Shizhong Xu, and Chen-Nee Chuah. The joint optimization of online traffic matrix measurement and traffic engineering for softwaredefined networks. IEEE/ ACM Transactions on Networking, 28(1):234–247, February 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL https://

dl.acm.org/doi/abs/10.
1109/TNET.2019.2957008.

Wallach:1997:AAH

[WEK97] Deborah A. Wallach, Dawson R. Engler, and M. Frans Kaashoek. ASHs: applicationspecific handlers for highperformance messaging. IEEE/ACM Transactions on Networking, 5(4):460– 474, August 1997. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1997-5-4/
p460-wallach/.

Wang:1993:EEC

[WF93a] Qinglin Wang and Victor S. Frost. Efficient estimation of cell blocking probability for ATM systems. IEEE/ACM Transactions on Networking, 1(2):230– 235, April 1993. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1993-1-2/
p230-wang/.

Woodside:1993:ASA

[WF93b] C. Murray Woodside and R. Greg Franks. Alternative software architectures for parallel protocol execution with synchronous IPC. IEEE/ACM Transactions on Networking, 1(2):178– 186, April 1993. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/

journals/ton/1993-1-2/
p178-woodside/.

Wang:2018:DJI

[WFC18] Chunpu Wang, Chen Feng, and Julian Cheng. Distributed join-the-idle-queue for low latency cloud services. IEEE/ACM Transactions on Networking, 26(5):2309–2319, October 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Wu:2013:IIC

[WFGZ13] Haitao Wu, Zhenqian Feng, Chuanxiong Guo, and Yongguang Zhang. ICTCP: incast congestion control for TCP in data-center networks. IEEE/ACM Transactions on Networking, 21 (2):345–358, April 2013. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Wang:2012:MPH

[WFH12] Xinbing Wang, Luoyi Fu, and Chenhui Hu. Multicast performance with hierarchical cooperation. IEEE/ACM Transactions on Networking, 20(3):917– 930, June 2012. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Wu:2009:OSW

[WFS09] Yan Wu, Sonia Fahmy, and Ness B. Shroff. Optimal sleep/wake scheduling for time-synchronized sensor networks with QoS guarantees. IEEE/ACM Transactions on Networking, 17(5): 1508–1521, October 2009. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Wu:2018:GNF

[WFY^{+}18] Xudong Wu, Luoyi Fu, Yuhang Yao, Xinzhe Fu, Xinbing Wang, and Guihai Chen. GLP: a novel framework for group-level location promotion in geosocial networks. IEEE/ ACM Transactions on Networking, 26(6):2870–2883, December 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Wang:2023:SDI

[WFZ^{+}23] Na Wang, Junsong Fu, Shancheng Zhang, Zheng Zhang, Jiawen Qiao, Jianwei Liu, and Bharat K. Bhargava. Secure and distributed IoT data storage in clouds based on secret sharing and collaborative blockchain. IEEE/ ACM Transactions on Networking, 31(4):1550– 1565, 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2022.3218933.

Wang:2016:IFB

[WG16] Qing Wang and Domenico Giustiniano. Intra-frame bidirectional transmission in networks of visible LEDs. IEEE/ACM Transactions on Networking, 24 (6):3607–3619, December 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Wong:2000:SGC

[WGL00] Chung Kei Wong, Mohamed Gouda, and Simon S. Lam. Secure group communications using key graphs. IEEE/ ACM Transactions on Networking, 8(1):16–30, February 2000. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/2000-8-1/
p16-wong/.

Wang:2022:IST

[WGL22] Shuai Wang, Jinkun Geng, and Dan Li. Impact of synchronization topology on DML performance: Both logical topology and physical topology. IEEE/ ACM Transactions on Networking, 30(2):572–585, April 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2021.3117042.

Wu:2017:SNI

[WGvdS17] Chuchu Wu, Mario Gerla, and Mihaela van der Schaar. Social norm incentives for network coding in Manets. IEEE/ACM Transactions on Networking, 25(3):1761–1774, June 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Wu:2021:AFL

[WGZC21] Guanhao Wu, Xiaofeng Gao, Jiaqi Zheng, and Guihai Chen. Achieving fast loop-free updates with ingress port in softwaredefined networks. IEEE/ ACM Transactions on Networking, 29(4):1527–1539, August 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2021.3068177.

Weller:1997:SNT

[WH97] Timothy Weller and Bruce Hajek. Scheduling nonuniform traffic in a packetswitching system with small propagation delay. IEEE/ACM Transactions on Networking, 5(6):813– 823, December 1997. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL http://

                                        www.acm.org/pubs/citations/
                                        journals/ton/1997-5-6/
                                        p813-weller/.

Wang:2011:BBT

[WH11] Shao-Cheng Wang and Ahmed Helmy. BEWARE: background traffic-aware rate adaptation for IEEE 802.11. IEEE/ACM Transactions on Networking, 19(4):1164–1177, August 2011. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Wang:2019:PPT

[WHC^{+}19] Qian Wang, Jing Huang, Yanjiao Chen, Xin Tian, and Qian Zhang. Privacypreserving and truthful double auction for heterogeneous spectrum. IEEE/ ACM Transactions on Networking, 27(2):848–861, April 2019. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Wu:2022:DTL

[WHC^{+}22] Qiong Wu, Kaiwen He, Xu Chen, Shuai Yu, and Junshan Zhang. Deep transfer learning across cities for mobile traffic prediction. IEEE/ ACM Transactions on Networking, 30(3):1255–1267, June 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2021.3136707.

Wang:2023:TPA

[WHLL23] Yipeng Wang, Huijie He, Yingxu Lai, and Alex X. Liu. A two-phase approach to fast and accurate classification of encrypted traffic. IEEE/ACM Transactions on Networking, 31(3):1071– 1086, June 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2022.3209979.

Wang:2013:ECI

[WHM^{+}13] Yin Wang, Yuan He, Xufei Mao, Yunhao Liu, and Xiang-Yang Li. Exploiting constructive interference for scalable flooding in wireless networks. IEEE/ ACM Transactions on Networking, 21(6):1880–1889, December 2013. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Wu:2015:TPP

[WHTC15] Fan Wu, Qianyi Huang, Yixin Tao, and Guihai Chen. Towards privacy preservation in strategyproof spectrum auction mechanisms for noncooperative wireless networks. IEEE/ACM Transactions on Networking, 23(4):1271– 1285, August 2015. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Wang:2011:DCT

[WHW^{+}11] Xinbing Wang, Wentao Huang, Shangxing Wang, Jinbei Zhang, and Chenhui Hu. Delay and capacity tradeoff analysis for motioncast. IEEE/ ACM Transactions on Networking, 19(5):1354–1367, October 2011. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Wang:2023:OLF

[WHYC23] Chen Wang, Qin Hu, Dongxiao Yu, and Xiuzhen Cheng. Online learning for failure-aware edge backup of service function chains with the minimum latency. IEEE/ACM Transactions on Networking, 31(6):2730– 2744, 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2023.3265127.

Wang:2020:ELP

[WHZJ20] Wei Wang, Shiyue He, Qian Zhang, and Tao Jiang. Enabling low-power OFDM for IoT by exploiting asymmetric clock rates. IEEE/ACM Transactions on Networking, 28(2):602– 611, April 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL https://

                                        dl.acm.org/doi/abs/10.
                                        1109/TNET.2020.2966112.

Williamson:1996:DBA

[Wil96] Carey L. Williamson. Dynamic bandwidth allocation using loss-load curves. IEEE/ACM Transactions on Networking, 4(6):829– 839, December 1996. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL http://

                                        www.acm.org/pubs/citations/
                                        journals/ton/1996-4-6/
                                        p829-williamson/.

Wang:2017:APS

[WJ17] Chang-Heng Wang and Tara Javidi. Adaptive policies for scheduling with reconfiguration delay: an end-to-end solution for all-optical data centers. IEEE/ACM Transactions on Networking, 25(3):1555– 1568, June 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Wang:2021:SPC

[WJH^{+}21] Lin Wang, Lei Jiao, Ting He, Jun Li, and Henri Bal. Service placement for collaborative edge applications. IEEE/ACM Transactions on Networking, 29(1):34–47, February 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.

acm.org/doi/10.1109/TNET.
2020.3025985.

Wu:2006:UNC

[WJK06] Yunnan Wu, Kamal Jain, and Sun-Yuan Kung. A unification of network coding and tree-packing (routing) theorems. IEEE/ ACM Transactions on Networking, 14(SI):2398–2409, June 2006. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Wei:2012:ITP

[WJK^{+}12] Wei Wei, Sharad Jaiswal, Jim Kurose, Don Towsley, Kyoungwon Suh, and Bing Wang. Identifying 802.11 traffic from passive measurements using iterative Bayesian inference. IEEE/ ACM Transactions on Networking, 20(2):325–338, April 2012. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Wei:2006:FTM

[WJLH06] David X. Wei, Cheng Jin, Steven H. Low, and Sanjay Hegde. FAST TCP: motivation, architecture, algorithms, performance. IEEE/ACM Transactions on Networking, 14(6):1246– 1259, December 2006. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Wang:2007:DAS

[WJS07] Haining Wang, Cheng Jin, and Kang G. Shin. Defense against spoofed IP traffic using hop-count filtering. IEEE/ACM Transactions on Networking, 15 (1):40–53, February 2007. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Wang:2016:SEG

[WJYL16] Chen Wang, Hongbo Jiang, Tianlong Yu, and John C. S. Lui. SLICE: Enabling greedy routing in high genus 3-D WSNs with general topologies. IEEE/ ACM Transactions on Networking, 24(4):2472–2484, August 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Wang:2012:FFA

[WJZ^{+}12] Anduo Wang, Limin Jia, Wenchao Zhou, Yiqing Ren, Boon Thau Loo, Jennifer Rexford, Vivek Nigam, Andre Scedrov, and Carolyn Talcott. FSR: formal analysis and implementation toolkit for safe interdomain routing. IEEE/ ACM Transactions on Networking, 20(6):1814–1827, December 2012. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Woo:2013:EIM

[WK13] Shinuk Woo and Hwangnam Kim. An empirical interference modeling for link reliability assessment in wireless networks. IEEE/ACM Transactions on Networking, 21(1):272– 285, February 2013. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Wang:2013:CQF

[WKA^{+}13] Wei Wang, Donghyun Kim, Min Kyung An, Wei Gao, Xianyue Li, Zhao Zhang, and Weili Wu. On construction of quality fault-tolerant virtual backbone in wireless networks. IEEE/ACM Transactions on Networking, 21(5):1499– 1510, October 2013. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Wang:2016:VBF

[WKWV16] Chih-Yu Wang, Chun-Han Ko, Hung-Yu Wei, and Athanasios V. Vasilakos. A voting-based femtocell downlink cellbreathing control mechanism. IEEE/ACM Transactions on Networking, 24 (1):85–98, February 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Wrege:1996:DDB

[WKZL96] Dallas E. Wrege, Edward W. Knightly, Hui Zhang, and Jörg Liebeherr. Deterministic delay bounds for VBR video in packet-switching networks: fundamental limits and practical trade-offs. IEEE/ACM Transactions on Networking, 4(3):352– 362, June 1996. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1996-4-3/
p352-wrege/.

Wong:1999:DSF

[WL99] Chung Kei Wong and Simon S. Lam. Digital signatures for flows and multicasts. IEEE/ACM Transactions on Networking, 7(4):502–513, August 1999. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1999-7-4/
p502-wong/.

Wang:2007:LBP

[WL07] Xiaoming Wang and Dmitri Loguinov. Load-balancing performance of consistent hashing: asymptotic analysis of random node join. IEEE/ACM Transactions on Networking, 15(4):892– 905, August 2007. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Wang:2008:EDB

[WL08] Dongmei Wang and Guangzhi Li. Efficient distributed bandwidth management for MPLS fast reroute. IEEE/ACM Transactions on Networking, 16(2):486– 495, April 2008. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Wang:2010:UMI

[WL10] Xiaoming Wang and Dmitri Loguinov. Understanding and modeling the Internet topology: economics and evolution perspective. IEEE/ACM Transactions on Networking, 18(1):257– 270, February 2010. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Wang:2016:LHS

[WL16] Qingsi Wang and Mingyan Liu. Learning in hide-and-seek. IEEE/ACM Transactions on Networking, 24 (2):1279–1292, April 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Wang:2010:TZD

[WLC^{+}10] Lanjia Wang, Zhichun Li, Yan Chen, Zhi Fu, and Xing Li. Thwarting zeroday polymorphic worms with network-level lengthbased signature generation. IEEE/ACM Transactions on Networking, 18(1):53– 66, February 2010. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Wong:2016:ESR

[WLC16] Felix Ming Fai Wong, Zhenming Liu, and Mung Chiang. On the efficiency of social recommender networks. IEEE/ ACM Transactions on Networking, 24(4):2512–2524, August 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Wang:2020:SHP

[WLC^{+}20] Songtao Wang, Dan Li, Yang Cheng, Jinkun Geng, Yanshu Wang, Shuai Wang, Shutao Xia, and Jianping Wu. A scalable, highperformance, and faulttolerant network architecture for distributed machine learning. IEEE/ ACM Transactions on Networking, 28(4):1752–1764, August 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2020.2999377.

Wang:2007:PIT

[WLCC07] Pi-Chung Wang, Chun-Liang Lee, Chia-Tai Chan, and Hung-Yi Chang. Performance improvement of two-dimensional packet classification by filter rephrasing. IEEE/ACM Transactions on Networking, 15 (4):906–917, August 2007. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Wang:2016:MTC

[WLCW16] Feng Wang, Jiangchuan Liu, Minghua Chen, and Haiyang Wang. Migration towards cloud-assisted live media streaming. IEEE/ ACM Transactions on Networking, 24(1):272–282, February 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Wang:2016:EPC

[WLD^{+}16] Jiliang Wang, Shuo Lian, Wei Dong, Xiang-Yang Li, and Yunhao Liu. Every packet counts: Loss and reordering identification and its application in delay measurement. IEEE/ ACM Transactions on Networking, 24(6):3426–3438, December 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Wang:2024:JOO

[WLD^{+}24] Juncheng Wang, Ben Liang, Min Dong, Gary Boudreau, and Hatem Abou-Zeid. Joint online optimization of model training and analog aggregation for wireless edge learning. IEEE/ACM Transactions on Networking, 32(2):1212– 1228, April 2024. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2023.3318474.

Wang:2017:NCF

[WLK^{+}17] Wei Wang, Bei Liu, Donghyun Kim, Deying Li, Jingyi Wang, and Wei Gao. A new constant factor approximation to construct highly faulttolerant connected dominating set in unit disk graph. IEEE/ACM Transactions on Networking, 25 (1):18–28, February 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Wong:2001:SCP

[WLL01] Vincent W. S. Wong, Mark E. Lewis, and Victor C. M. Leung. Stochastic control of path optimization for inter-switch handoffs in wireless ATM networks. IEEE/ACM Transactions on Networking, 9 (3):336–350, June 2001.

CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Wang:2011:CBL

[WLL^{+}11] Xiaoping Wang, Jun Luo, Yunhao Liu, Shanshan Li, and Dezun Dong. Component-based localization in sparse wireless networks. IEEE/ACM Transactions on Networking, 19 (2):540–548, April 2011. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Wang:2013:MRG

[WLL13] Xiaoming Wang, Xiaoyong Li, and Dmitri Loguinov. Modeling residualgeometric flow sampling. IEEE/ACM Transactions on Networking, 21(4):1090– 1103, August 2013. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Wu:2016:ALO

[WLL^{+}16a] Di Wu, Qiang Liu, Yong Li, Julie A. McCann, Amelia C. Regan, and Nalini Venkatasubramanian. Adaptive lookup of open WiFi using crowdsensing. IEEE/ACM Transactions on Networking, 24(6):3634–3647, December 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Wu:2016:CCS

[WLL^{+}16b] Huasen Wu, Xiaojun Lin, Xin Liu, Kun Tan, and Yongguang Zhang. CoSchd: Coordinated scheduling with channel and load awareness for alleviating cellular congestion. IEEE/ ACM Transactions on Networking, 24(5):2579–2592, October 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Wang:2024:NOP

[WLL^{+}24] Xiujun Wang, Zhi Liu, Alex X. Liu, Xiao Zheng, Hao Zhou, Ammar Hawbani, and Zhe Dang. A near-optimal protocol for continuous tag recognition in mobile RFID systems. IEEE/ACM Transactions on Networking, 32(2):1303– 1318, April 2024. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2023.3317875.

Wang:2005:CLO

[WLLD05] Jiantao Wang, Lun Li, Steven H. Low, and John C. Doyle. Crosslayer optimization in TCP/ IP networks. IEEE/ACM Transactions on Networking, 13(3):582–595, June 2005. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Wu:2016:ALS

[WLLZ16] Huasen Wu, Xiaojun Lin, Xin Liu, and Youguang Zhang. Application-level scheduling with probabilistic deadline constraints. IEEE/ACM Transactions on Networking, 24(3):1504– 1517, June 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Wu:2010:MAM

[WLR10] Di Wu, Yong Liu, and Keith W. Ross. Modeling and analysis of multichannel P2P live video systems. IEEE/ACM Transactions on Networking, 18(4):1248– 1260, August 2010. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Willebeek-LeMair:1997:ADM

[WLS97] Marc Willebeek-LeMair and Perwez Shahabuddin. Approximating dependability measures of computer networks: an FDDI case study. IEEE/ ACM Transactions on Networking, 5(2):311–327, April 1997. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1997-5-2/
p311-willebeek-lemair/.

Wang:2018:TCB

[WLS^{+}18] Haiyang Wang, Tong Li, Ryan Shea, Xiaoqiang Ma, Feng Wang, Jiangchuan Liu, and Ke Xu. Toward cloud-based distributed interactive applications: Measurement, modeling, and analysis. IEEE/ACM Transactions on Networking, 26(1):3–16, February 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Wang:2023:MLF

[WLS23] Haoyu Wang, Zetian Liu, and Haiying Shen. Machine learning feature based job scheduling for distributed machine learning clusters. IEEE/ACM Transactions on Networking, 31 (1):58–73, 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2022.3190797.

Wang:2019:DCN

[WLTJ19] Chang-Heng Wang, Jaime Llorca, Antonia M. Tulino, and Tara Javidi. Dynamic cloud network control under reconfiguration delay and cost. IEEE/ACM Transactions on Networking, 27(2):491–504, April 2019. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Wu:2017:CEI

[WLW^{+}17] Fan Wu, Dongxin Liu, Zhihao Wu, Yuan Zhang, and Guihai Chen. Cost-efficient indoor white space exploration through compressive sensing. IEEE/ACM Transactions on Networking, 25(3):1686–1702, June 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Wang:2020:SSD

[WLW^{+}20] Huazhe Wang, Xin Li, Yang Wang, Yu Zhao, Ye Yu, Hongkun Yang, and Chen Qian. SICS: Secure and dynamic middlebox outsourcing. IEEE/ ACM Transactions on Networking, 28(6):2713–2726, December 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2020.3023386.

Wang:2013:MIC

[WLWL13] Xinbing Wang, Xiaojun Lin, Qingsi Wang, and Wentao Luan. Mobility increases the connectivity of wireless networks. IEEE/ACM Transactions on Networking, 21(2):440– 454, April 2013. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Wang:2017:EOA

[WLX^{+}17] Tao Wang, Fangming Liu, Hong Xu, Tao Wang, Fangming Liu, and Hong Xu. An efficient online algorithm for dynamic SDN controller assignment in data center networks. IEEE/ ACM Transactions on Networking, 25(5):2788–2801, October 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Wang:2023:SDD

[WLY^{+}23] Wei Wang, Xin Liu, Yao Yao, Zicheng Chi, Stuart Ray, Ting Zhu, and Yanchao Zhang. Simultaneous data dissemination among WiFi and ZigBee devices. IEEE/ACM Transactions on Networking, 31(6):2545– 2558, 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2023.3243070.

Wang:2024:ETA

[WLY^{+}24] Hao Wang, Chi Harold Liu, Haoming Yang, Guoren Wang, and Kin K. Leung. Ensuring threshold AoI for UAV-Assisted mobile crowdsensing by multi-agent deep reinforcement learning with transformer. IEEE/ACM Transactions on Networking, 32 (1):566–581, 2024. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2023.3289172.

Wu:2011:DSP

[WLZ11] Chuan Wu, Baochun Li, and Shuqiao Zhao. On dynamic server provisioning in multichannel P2P live streaming. IEEE/ ACM Transactions on Networking, 19(5):1317–1330, October 2011. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Wu:1995:BAS

[WM95] Guo-Liang Wu and Jon W.

Mark. A buffer allocation scheme for ATM networks: complete sharing based on virtual partition. IEEE/ ACM Transactions on Networking, 3(6):660–670, December 1995. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1995-3-6/
p660-wu/.

Wu:1996:CMP

[WM96] Guo-Liang Wu and Jon W.

Mark. Computational methods for performance evaluation of a statistical multiplexer supporting bursty traffic. IEEE/ ACM Transactions on Networking, 4(3):386–397, June 1996. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/
                                        journals/ton/1996-4-3/
                                        p386-wu/.

Wang:2016:ANC

[WM16] Xudong Wang and Wenguang Mao. Analog network coding without restrictions on superimposed frames. IEEE/ACM Transactions on Networking, 24 (2):788–805, April 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Wang:2022:FAC

[WMCW22] Haibo Wang, Chaoyi Ma, Shigang Chen, and Yuanda Wang. Fast and accurate cardinality estimation by self-morphing bitmaps. IEEE/ACM Transactions on Networking, 30(4):1674– 1688, 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2022.3147204.

Wu:2010:CML

[WMFS10] Yan Wu, Zhoujia Mao, Sonia Fahmy, and Ness B. Shroff. Constructing maximum-lifetime data gathering forests in sensor networks. IEEE/ ACM Transactions on Networking, 18(5):1571–1584, October 2010. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Wu:2018:YRM

[WML^{+}18] Fan Wu, Tong Meng, Aijing Li, Guihai Chen, and Nitin H. Vaidya. Have you recorded my voice: Toward robust neighbor discovery in mobile wireless networks. IEEE/ACM Transactions on Networking, 26 (3):1432–1445, June 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Wang:2023:RER

[WMO^{+}23] Haibo Wang, Chaoyi Ma, Olufemi O. Odegbile, Shigang Chen, and Jih-Kwon Peir. Randomized error removal for online spread estimation in highspeed networks. IEEE/ ACM Transactions on Networking, 31(2):558–573, April 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2022.3197968.

Wang:2018:PPL

[WMP^{+}18] Xiong Wang, Mehdi Malboubi, Zhihao Pan, Jing Ren, Sheng Wang, Shizhong Xu, and Chen-Nee Chuah.

ProgLIMI: Programmable LInk Metric Identification in software-defined networks. IEEE/ACM Transactions on Networking, 26(5):2376–2389, October 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Wang:2009:OEE

[WMS09] Wei Wang, Mehul Motani, and Vikram Srinivasan. Opportunistic energy-efficient contact probing in delaytolerant applications. IEEE/ ACM Transactions on Networking, 17(5):1592–1605, October 2009. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Wu:2022:GST

[WMT^{+}22] Renyong Wu, Jieming Ma, Zhixiang Tang, Xiehua Li, and Kim-Kwang Raymond Choo. A generic secure transmission scheme based on random linear network coding. IEEE/ ACM Transactions on Networking, 30(2):855–866, April 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2021.3124890.

Wang:2017:RDC

[WMX17] Xin Wang, Richard T. B. Ma, and Yinlong Xu. The role of data cap in optimal two-part network pricing. IEEE/ACM Transactions on Networking, 25(6): 3602–3615, December 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Won:2016:PAA

[WMYR16] Jongho Won, Chris Y. T.

Ma, David K. Y. Yau, and Nageswara S. V. Rao. Privacy-assured aggregation protocol for smart metering: a proactive faulttolerant approach. IEEE/ ACM Transactions on Networking, 24(3):1661–1674, June 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Wei:2016:PAW

[WN16] Xiaohan Wei and Michael J.

Neely. Power-aware wireless file downloading: a Lyapunov indexing approach to a constrained restless bandit problem. IEEE/ACM Transactions on Networking, 24(4):2264– 2277, August 2016. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Wei:2017:DCS

[WN17] Xiaohan Wei and Michael J.

Neely. Data center server provision: Distributed asynchronous control for coupled renewal systems. IEEE/ACM Transactions on Networking, 25(4):2180– 2194, August 2017. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Wolf:2013:HPC

[WNV13] Tilman Wolf, Sriram Natarajan, and Kamlesh T. Vasudevan. High-performance capabilities for 1-hop containment of network attacks. IEEE/ACM Transactions on Networking, 21 (6):1931–1946, December 2013. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Wang:2006:AOF

[WPL06] Wei-Hua Wang, Marimuthu Palaniswami, and Steven H. Low. Application-oriented flow control: fundamentals, algorithms and fairness. IEEE/ACM Transactions on Networking, 14(6): 1282–1291, December 2006. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Wang:2016:SEP

[WPZM16] Xinlei Wang, Amit Pande, Jindan Zhu, and Prasant Mohapatra. STAMP: Enabling privacy-preserving location proofs for mobile users. IEEE/ACM Transactions on Networking, 24 (6):3276–3289, December 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Wang:2006:WAM

[WQC06] Jianping Wang, Xiangtong Qi, and Biao Chen. Wavelength assignment for multicast in all-optical WDM networks with splitting constraints. IEEE/ ACM Transactions on Networking, 14(1):169–182, February 2006. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Wang:2009:UTI

[WQGW09] Feng Wang, Jian Qiu, Lixin Gao, and Jia Wang. On understanding transient interdomain routing failures. IEEE/ACM Transactions on Networking, 17(3):740– 751, June 2009. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Wang:2021:CVP

[WQL^{+}21] Minmei Wang, Chen Qian, Xin Li, Shouqian Shi, and Shigang Chen. Collaborative validation of public-key certificates for IoT by distributed caching. IEEE/ ACM Transactions on Networking, 29(1):92–105, February 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.

                                        acm.org/doi/10.1109/TNET.
                                        2020.3029135.

Wang:2017:PNW

[WQY^{+}17] Huazhe Wang, Chen Qian, Ye Yu, Hongkun Yang, Simon S. Lam, Huazhe Wang, Chen Qian, Ye Yu, Hongkun Yang, and Simon S. Lam. Practical network-wide packet behavior identification by AP classifier. IEEE/ ACM Transactions on Networking, 25(5):2886–2899, October 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Wu:2013:FTR

[WQZ^{+}13] Qianhong Wu, Bo Qin, Lei Zhang, Josep Domingo-Ferrer, and Jesús A. Manjón. Fast transmission to remote cooperative groups: a new key management paradigm. IEEE/ACM Transactions on Networking, 21(2):621– 633, April 2013. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Wood:2015:CDP

[WRS^{+}15] Timothy Wood, K. K. Ramakrishnan, Prashant Shenoy, Jacobus Van Der Merwe, Jinho Hwang, Guyue Liu, and Lucas Chaufournier. CloudNet: dynamic pooling of cloud resources by live WAN migration of virtual machines. IEEE/ACM Transactions on Networking, 23(5):1568– 1583, October 2015. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Wang:2021:NAO

[WRT^{+}21] Su Wang, Yichen Ruan, Yuwei Tu, Satyavrat Wagle, Christopher G. Brinton, and Carlee Joe-Wong. Network-aware optimization of distributed learning for fog computing. IEEE/ ACM Transactions on Networking, 29(5):2019–2032, October 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2021.3075432.

Wang:1993:FDM

[WS93] Clark Wang and Mischa Schwartz. Fault detection with multiple observers. IEEE/ACM Transactions on Networking, 1(1):48–55, February 1993. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1993-1-1/
p48-wang/.

Wu:2005:CTA

[WS05] Tao Wu and Arun K. Somani. Cross-talk attack monitoring and localization in all-optical networks. IEEE/ACM Transactions on Networking, 13(6):1390– 1401, December 2005. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Wang:2006:PNR

[WS06] Xin Wang and Henning Schulzrinne. Pricing network resources for adaptive applications. IEEE/ACM Transactions on Networking, 14(3):506–519, 2006. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Wu:2008:CAS

[WS08] Tao Wu and David Starobinski. A comparative analysis of server selection in content replication networks. IEEE/ACM Transactions on Networking, 16(6):1461– 1474, December 2008. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Wang:2008:ELW

[WSC08] Wei Wang, Vikram Srinivasan, and Kee-Chaing Chua. Extending the lifetime of wireless sensor networks through mobile relays. IEEE/ACM Transactions on Networking, 16(5): 1108–1120, October 2008. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Wang:2023:GMC

[WSC^{+}23] Ge Wang, Xiaofeng Shi, Haofan Cai, Chen Qian, Han Ding, Wei Xi, Kun Zhao, Jizhong Zhao, and Jinsong Han. A generalized method to combat multipaths for RFID sensing. IEEE/ACM Transactions on Networking, 31(1): 336–351, 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2022.3190862.

Wang:2008:MAL

[WSKV08] Xin Wang, Henning Schulzrinne, Dilip Kandlur, and Dinesh Verma. Measurement and analysis of LDAP performance. IEEE/ACM Transactions on Networking, 16(1):232–243, February 2008. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Wu:2024:ERS

[WSL^{+}24] Si Wu, Zhirong Shen, Patrick P. C. Lee, Zhiwei Bai, and Yinlong Xu. Elastic Reed–Solomon codes for efficient redundancy transitioning in distributed key–value stores. IEEE/ACM Transactions on Networking, 32(1):670– 685, 2024. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2023.3303865.

Watson:2004:PSN

[WSMJ04] David Watson, Matthew Smart, G. Robert Malan, and Farnam Jahanian. Protocol scrubbing: network security through transparent flow modification. IEEE/ACM Transactions on Networking, 12(2):261– 273, April 2004. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Wang:2012:DRA

[WSW12] Wei Wang, Kang G. Shin, and Wenbo Wang. Distributed resource allocation based on queue balancing in multihop cognitive radio networks. IEEE/ACM Transactions on Networking, 20(3):837–850, June 2012. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Wan:2021:CEP

[WSX^{+}21] Ying Wan, Haoyu Song, Yang Xu, Yilun Wang, Tian Pan, Chuwen Zhang, Yi Wang, and Bin Liu. T-Cache: Efficient policybased forwarding using small TCAM. IEEE/ACM Transactions on Networking, 29(6):2693–2708, December 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2021.3098320.

Wang:2023:SID

[WSX^{+}23] Yuntao Wang, Zhou Su, Qichao Xu, Ruidong Li, Tom H. Luan, and Pinghui Wang. A secure and intelligent data sharing scheme for UAV-assisted disaster rescue. IEEE/ ACM Transactions on Networking, 31(6):2422– 2438, 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2022.3226458.

Wang:2016:ENS

[WSXL16] Xiang Wang, Weiqi Shi, Yang Xiang, and Jun Li. Efficient network security policy enforcement with policy space analysis. IEEE/ACM Transactions on Networking, 24(5): 2926–2938, October 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Wei:2020:DOS

[WSZL20] Ziling Wei, Jinshu Su, Baokang Zhao, and Xicheng Lu. Distributed opportunistic scheduling in cooperative networks with RF energy harvesting. IEEE/ACM Transactions on Networking, 28(5):2257–2270, October 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2020.3011839.

Wang:2017:PAR

[WT17] You-Chiun Wang and Tzung-Yu Tsai. A pricingaware resource scheduling framework for LTE networks. IEEE/ACM Transactions on Networking, 25 (3):1445–1458, June 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Tcha:1997:NLB

[wTjCjC97] Dong wan Tcha, Yong joo Chung, and Taek jin Choi. A new lower bound for the frequency assignment problem. IEEE/ ACM Transactions on Networking, 5(1):34–39, February 1997. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/
                                        journals/ton/1997-5-1/
                                        p34-tcha/.

Woldeyohannes:2022:CEA

[WTJR22] Yordanos Tibebu Woldeyohannes, Besmir Tola, Yuming Jiang, and K. K. Ramakrishnan. CoShare: an efficient approach for redundancy allocation in NFV. IEEE/ACM Transactions on Networking, 30 (3):1014–1028, June 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL

https://dl.acm.org/doi/
10.1109/TNET.2021.3132279.

Wang:2017:MCC

[WTK^{+}17] Liang Wang, Gareth Tyson, Jussi Kangasharju, Jon Crowcroft, Liang Wang, Gareth Tyson, Jussi Kangasharju, and Jon Crowcroft. Milking the cache cow with fairness in mind. IEEE/ ACM Transactions on Networking, 25(5):2686–2700, October 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Wu:2013:FSD

[WTS^{+}13] Xinzhou Wu, Saurabha Tavildar, Sanjay Shakkottai, Tom Richardson, Junyi Li, Rajiv Laroia, and Aleksandar Jovicic. FlashLinQ: a synchronous distributed scheduler for peer-to-peer ad hoc networks. IEEE/ ACM Transactions on Networking, 21(4):1215–1228, August 2013. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Willinger:1997:STH

[WTSW97] Walter Willinger, Murad S.

Taqqu, Robert Sherman, and Daniel V. Wilson. Selfsimilarity through highvariability: statistical analysis of Ethernet LAN traffic at the source level. IEEE/ACM Transactions on Networking, 5(1):71–86, February 1997. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1997-5-1/
p71-willinger/.

Wang:2011:CSR

[WTXT11] Meng Wang, Chee Wei Tan, Weiyu Xu, and Ao Tang. Cost of not splitting in routing: characterization and estimation. IEEE/ACM Transactions on Networking, 19(6): 1849–1859, December 2011. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Wu:1994:PPS

[Wu94] Tsong-Ho Wu. A passive protected self-healing mesh network architecture and applications. IEEE/ ACM Transactions on Networking, 2(1):40–52, February 1994. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1994-2-1/
p40-wu/.

Wang:2019:DSM

[WUZ^{+}19] Shiqiang Wang, Rahul Urgaonkar, Murtaza Zafer, Ting He, Kevin Chan, and Kin K. Leung. Dynamic service migration in mobile edge computing based on Markov decision process. IEEE/ACM Transactions on Networking, 27 (3):1272–1288, June 2019. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Wang:2012:CLA

[WVG12] Yunbo Wang, Mehmet C.

Vuran, and Steve Goddard. Cross-layer analysis of the end-to-end delay distribution in wireless sensor networks. IEEE/ ACM Transactions on Networking, 20(1):305–318, February 2012. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Wang:2017:FAF

[WVZ17] Shu Wang, Vignesh Venkateswaran, and Xinyu Zhang. Fundamental analysis of fullduplex gains in wireless networks. IEEE/ACM Transactions on Networking, 25(3):1401–1416, June 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Wildman:2016:CLP

[WW16] Jeffrey Wildman and Steven Weber. On characterizing the local pooling factor of greedy maximal scheduling in random graphs. IEEE/ ACM Transactions on Networking, 24(4):2086–2099, August 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Wang:2018:DPF

[WWC^{+}18] Yitu Wang, Wei Wang, Ying Cui, Kang G. Shin, and Zhaoyang Zhang. Distributed packet forwarding and caching based on stochastic network utility maximization. IEEE/ACM Transactions on Networking, 26(3):1264–1277, June 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Wu:2002:DCA

[WWL02] Si Wu, K. Y. Michael Wong, and Bo Li. A dynamic call admission policy with precision QoS guarantee using stochastic control for mobile wireless networks. IEEE/ACM Transactions on Networking, 10 (2):257–271, April 2002. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Wu:2015:SSM

[WWL^{+}15] Yu Wu, Chuan Wu, Bo Li, Linquan Zhang, Zongpeng Li, and Francis C. M. Lau. Scaling social media applications into geo-distributed clouds. IEEE/ACM Transactions on Networking, 23 (3):689–702, June 2015. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Wang:2024:RPB

[WWL^{+}24a] Zihao Wang, Hang Wang, Zhuowen Li, Xinghua Li, Yinbin Miao, Yanbing Ren, Yunwei Wang, Zhe Ren, and Robert H. Deng. Robust permissioned blockchain consensus for unstable communication in FANET. IEEE/ACM Transactions on Networking, 32(1):699– 712, 2024. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2023.3295378.

Wei:2024:EEA

[WWL24b] Qinhan Wei, Yongcai Wang, and Deying Li. EMI: an efficient algorithm for identifying maximal rigid clusters in 3D generic graphs. IEEE/ACM Transactions on Networking, 32 (1):460–474, 2024. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2023.3287822.

Wang:2020:OSC

[WWMZ20] Zengfu Wang, Qing Wang, Bill Moran, and Moshe Zukerman. Optimal submarine cable path planning and trunk-and-branch tree network topology design. IEEE/ACM Transactions on Networking, 28(4):1562– 1572, August 2020. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1109/
                                        TNET.2020.2988047.

Wang:2022:SCN

[WWMZ22] Tianjiao Wang, Zengfu Wang, Bill Moran, and Moshe Zukerman. Submarine cable network design for regional connectivity. IEEE/ACM Transactions on Networking, 30(6): 2480–2492, 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2022.3171832.

Wu:2005:NSC

[WWT05] Hongyi Wu, Chong Wang, and Nian-Feng Tzeng. Novel self-configurable positioning technique for multihop wireless networks. IEEE/ACM Transactions on Networking, 13(3):609– 621, June 2005. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Wei:2011:MBI

[WWTK11] Wei Wei, Bing Wang, Don Towsley, and Jim Kurose. Model-based identification of dominant congested links. IEEE/ACM Transactions on Networking, 19(2):456–469, April 2011. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Wang:2018:GRS

[WWW^{+}18] Gang Wang, Bolun Wang, Tianyi Wang, Ana Nika, Haitao Zheng, and Ben Y. Zhao. Ghost riders: Sybil attacks on crowdsourced mobile mapping services. IEEE/ACM Transactions on Networking, 26(3):1123– 1136, June 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Wang:2020:SGE

[WWW^{+}20a] Haibo Wang, Jessie Hui Wang, Jilong Wang, Weizhen Dang, Jing’an Xue, Fenghua Li, and Jinzhe Shan. Squeezing the gap: an empirical study on DHCP performance in a largescale wireless network. IEEE/ACM Transactions on Networking, 28(2):832– 845, April 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL https://

                                        dl.acm.org/doi/abs/10.
                                        1109/TNET.2020.2971551.

Wang:2020:MAC

[WWW20b] Jie Wang, Wenye Wang, and Cliff Wang. Modeling and analysis of conflicting information propagation in a finite time horizon. IEEE/ACM Transactions on Networking, 28(3):972– 985, June 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2020.2976972.

Wu:2019:EDT

[WWX^{+}19] Haiqin Wu, Liangmin Wang, Guoliang Xue, Jian Tang, and Dejun Yang. Enabling data trustworthiness and user privacy in mobile crowdsensing. IEEE/ ACM Transactions on Networking, 27(6):2294–2307, December 2019. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL https://

                                        dl.acm.org/doi/abs/10.
                                        1109/TNET.2019.2944984.

Weng:2018:RCA

[WWYY18] Jianping Weng, Jessie Hui Wang, Jiahai Yang, and Yang. Root cause analysis of anomalies of multitier services in public clouds. IEEE/ACM Transactions on Networking, 26(4):1646– 1659, August 2018. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Wang:2004:PAD

[WXBZ04] Shengquan Wang, Dong Xuan, Riccardo Bettati, and Wei Zhao. Providing absolute differentiated services for realtime applications in staticpriority scheduling networks. IEEE/ACM Transactions on Networking, 12 (2):326–339, April 2004. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Wei:2016:FNC

[WXC16] Rihua Wei, Yang Xu, and H. Jonathan Chao. Finding nonequivalent classifiers in Boolean space to reduce TCAM usage. IEEE/ACM Transactions on Networking, 24(2):968–981, April 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Wang:2024:DOC

[WXC^{+}24] Sihan Wang, Tian Xie, Min-Yue Chen, Guan-Hua Tu, Chi-Yu Li, Xinyu Lei, Po-Yi Chou, Fucheng Hsieh, Yiwen Hu, Li Xiao, and Chunyi Peng. Dissecting operational cellular IoT service security: Attacks and defenses. IEEE/ ACM Transactions on Networking, 32(2):1229–1244, April 2024. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2023.3313557.

Wang:2018:MCR

[WXH^{+}18] Pengzhan Wang, Hongli Xu, Liusheng Huang, Chen Qian, Shaowei Wang, and Yanjing Sun. Minimizing controller response time through flow redirecting in SDNs. IEEE/ACM Transactions on Networking, 26(1):562–575, February 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Wei:2020:SBB

[WXH^{+}20] Wenjia Wei, Kaiping Xue, Jiangping Han, David S. L. Wei, and Peilin Hong. Shared bottleneckbased congestion control and packet scheduling for multipath TCP. IEEE/ ACM Transactions on Networking, 28(2):653–666, April 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL https://

                                        dl.acm.org/doi/abs/10.
                                        1109/TNET.2020.2970032.

Wang:2017:DWE

[WXJ^{+}17] Ju Wang, Jie Xiong, Hongbo Jiang, Xiaojiang Chen, and Dingyi Fang. D-Watch: Embracing “Bad” multipaths for device-free localization with COTS RFID devices. IEEE/ACM Transactions on Networking, 25(6):3559–3572, December 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Wang:2021:PPS

[WXM21] Xin Wang, Yinlong Xu, and Richard T. B. Ma. Paid peering, settlementfree peering, or both? IEEE/ACM Transactions on Networking, 29(2):585– 594, April 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2020.3045220.

Wang:2017:ECA

[WXN^{+}17] Peng Wang, Hong Xu, Zhixiong Niu, Dongsu Han, Yongqiang Xiong, Peng Wang, Hong Xu, Zhixiong Niu, Dongsu Han, and Yongqiang Xiong. Expeditus: Congestion-aware load balancing in Clos data center networks. IEEE/ ACM Transactions on Networking, 25(5):3175–3188, October 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Wang:2013:EDS

[WXR13] Miao Wang, Lisong Xu, and Byrav Ramamurthy. Exploring the design space of multichannel peer-to-peer live video streaming systems. IEEE/ ACM Transactions on Networking, 21(1):162–175, February 2013. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Wu:2011:DIF

[WXW11] Zhenyu Wu, Mengjun Xie, and Haining Wang. Design and implementation of a fast dynamic packet filter. IEEE/ACM Transactions on Networking, 19(5): 1405–1419, October 2011. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Wu:2015:WHS

[WXW15] Zhenyu Wu, Zhang Xu, and Haining Wang. Whispers in the hyper-space: highbandwidth and reliable covert channel attacks inside the cloud. IEEE/ACM Transactions on Networking, 23(2):603–614, April 2015. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Wang:2024:BBW

[WXX^{+}24] Lun Wang, Yang Xu, Hongli Xu, Zhida Jiang, Min Chen, Wuyang Zhang, and Chen Qian. BOSE: Block-wise federated learning in heterogeneous edge computing. IEEE/ACM Transactions on Networking, 32(2):1362–1377, April 2024. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL

https://dl.acm.org/doi/
10.1109/TNET.2023.3316421.

Wong:1995:DAN

[WY95] P. C. Wong and M. S. Yeung. Design and analysis of a novel fast packet switch: pipeline Banyan. IEEE/ACM Transactions on Networking, 3(1):63–69, February 1995. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1995-3-1/
p63-wong/.

Wan:2006:CRD

[WY06] Peng-Jun Wan and Chih-Wei Yi. Coverage by randomly deployed wireless sensor networks. IEEE/ ACM Transactions on Networking, 14(SI):2658–2669, June 2006. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Wu:2010:IFC

[WYH10] Bin Wu, Kwan L. Yeung, and Pin-Han Ho.

ILP formulations for pcycle design without candidate cycle enumeration. IEEE/ACM Transactions on Networking, 18(1):284– 295, February 2010. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Wu:2009:MIS

[WYHL09] Bin Wu, Kwan L. Yeung, Mounir Hamdi, and Xin Li. Minimizing internal speedup for performance guaranteed switches with optical fabrics. IEEE/ACM Transactions on Networking, 17(2):632–645, April 2009. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Wang:2009:RBE

[WYL09] Xiaoming Wang, Zhongmei Yao, and Dmitri Loguinov. Residual-based estimation of peer and link lifetimes in P2P networks. IEEE/ACM Transactions on Networking, 17(3):726– 739, June 2009. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Wang:2022:ADS

[WYL^{+}22] Zhiyuan Wang, Jiancheng Ye, Dong Lin, Yipei Chen, and John C. S. Lui. Approximate and deployable shortest remaining processing time scheduler. IEEE/ ACM Transactions on Networking, 30(3):1368–1381, June 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2022.3142148.

Wang:2023:DSD

[WYL23] Xiong Wang, Jiancheng Ye, and John C. S. Lui. Decentralized scheduling and dynamic pricing for edge computing: a mean field game approach. IEEE/ ACM Transactions on Networking, 31(3):965–978, June 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2022.3204698.

Wang:2024:MFG

[WYL24] Xiong Wang, Jiancheng Ye, and John C. S. Lui. Mean field graph based D2D collaboration and offloading pricing in mobile edge computing. IEEE/ACM Transactions on Networking, 32 (1):491–505, 2024. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2023.3288558.

Wang:2024:NMF

[WYZ^{+}24] Guijuan Wang, Jiguo Yu, Yifei Zou, Jianxi Fan, and Wei Cheng. A new measure of fault-tolerance for network reliability: Double-structure connectivity. IEEE/ACM Transactions on Networking, 32 (1):874–889, 2024. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2023.3305611.

Wang:2016:PPC

[WZ16] Wei Wang and Qian Zhang. Privacy preservation for context sensing on Smartphone. IEEE/ACM Transactions on Networking, 24 (6):3235–3247, December 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Wang:2024:DPC

[WZD24] Xuehe Wang, Shensheng Zheng, and Lingjie Duan. Dynamic pricing for client recruitment in federated learning. IEEE/ACM Transactions on Networking, 32(2):1273–1286, April 2024. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1109/TNET.2023.3312208.

Wang:2018:MAB

[WZH^{+}18] Shuo Wang, Jiao Zhang, Tao Huang, Tian Pan, Jiang Liu, and Yunjie Liu. Multi-attributes-based coflow scheduling without prior knowledge. IEEE/ ACM Transactions on Networking, 26(4):1962–1975, August 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Wang:2024:TCA

[WZH^{+}24] Yizong Wang, Dong Zhao, Chenghao Huang, Fuyu Yang, Teng Gao, Anfu Zhou, Huanhuan Zhang, Huadong Ma, Yang Du, and Aiyun Chen. TrafAda: Cost-aware traffic adaptation for maximizing bitrates in live streaming. IEEE/ACM Transactions on Networking, 32(1):96– 109, 2024. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2023.3285812.

Wu:2013:FAA

[WZL^{+}13] Huasen Wu, Chenxi Zhu, Richard J. La, Xin Liu, and Youguang Zhang. FASA: accelerated S-ALOHA using access history for eventdriven M2M communications. IEEE/ACM Transactions on Networking, 21 (6):1904–1917, December 2013. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Wang:2023:OCD

[WZL^{+}23a] En Wang, Mijia Zhang, Wenbin Liu, Haoyi Xiong, Bo Yang, Yongjian Yang, and Jie Wu. Outlierconcerned data completion exploiting intra- and interdata correlations in sparse CrowdSensing. IEEE/ ACM Transactions on Networking, 31(2):648–663, April 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2022.3201545.

Wang:2023:BSC

[WZL^{+}23b] Shicheng Wang, Menghao Zhang, Guanyu Li, Chang Liu, Zhiliang Wang, Ying Liu, and Mingwei Xu. Bolt: Scalable and cost-efficient multistring pattern matching with programmable switches. IEEE/ACM Transactions on Networking, 31(2):846– 861, April 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2022.3202523.

Wu:2023:PTL

[WZL^{+}23c] Qiang Wu, Xiangping Bryce Zhai, Xi Liu, Chun-Ming Wu, Fangliang Lou, and Hongke Zhang. Performance tuning via lean measurements for acceleration of network functions virtualization. IEEE/ ACM Transactions on Networking, 31(1):366– 379, 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2022.3193686.

Wang:2022:PAA

[WZLM22] Yuting Wang, Xiaolong Zheng, Liang Liu, and Huadong Ma. Polar-Tracker: Attitude-aware channel access for floating low power wide area networks. IEEE/ACM Transactions on Networking, 30 (4):1807–1821, 2022. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL https:/

/dl.acm.org/doi/10.1109/
TNET.2022.3154937.

Wang:2012:DBS

[WZLX12] Hao Wang, Haiquan Zhao, Bill Lin, and Jun Xu. DRAM-based statistics counter array architecture with performance guarantee. IEEE/ACM Transactions on Networking, 20(4): 1040–1053, August 2012. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Walters:2008:FMA

[WZR08] Aaron Walters, David Zage, and Cristina Nita Rotaru. A framework for mitigating attacks against measurement-based adaptation mechanisms in unstructured multicast overlay networks. IEEE/ACM Transactions on Networking, 16(6):1434–1446, December 2008. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Wang:2020:DTP

[WZW^{+}20] Fangxin Wang, Cong Zhang, Feng Wang, Jiangchuan Liu, Yifei Zhu, Haitian Pang, and Lifeng Sun. DeepCast: Towards personalized QoE for edgeassisted crowdcast with deep reinforcement learning. IEEE/ACM Transactions on Networking, 28(3): 1255–1268, June 2020. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL https:/

/dl.acm.org/doi/10.1109/
TNET.2020.2979966.

Wang:2022:RSM

[WZX^{+}22] Jingzhou Wang, Gongming Zhao, Hongli Xu, Yutong Zhai, Qianyu Zhang, He Huang, and Yongqiang Yang. A robust service mapping scheme for multitenant clouds. IEEE/ ACM Transactions on Networking, 30(3):1146–1161, June 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2021.3133293.

Wang:2016:MSM

[WZY^{+}16] Weina Wang, Kai Zhu, Lei Ying, Jian Tan, and Li Zhang. MapTask scheduling in MapReduce with data locality: throughput and heavytraffic optimality. IEEE/ ACM Transactions on Networking, 24(1):190–203, February 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Wu:2017:DSA

[WZZC17] Haitao Wu, Fen Zhou, Zuqing Zhu, and Yaojun Chen. On the distance spectrum assignment in elastic optical networks. IEEE/ACM Transactions on Networking, 25(4):2391– 2404, August 2017. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Xiao:2012:RRW

[XAST12] Weiyao Xiao, Sachin Agarwal, David Starobinski, and Ari Trachtenberg. Reliable rateless wireless broadcasting with near-zero feedback. IEEE/ACM Transactions on Networking, 20(6): 1924–1937, December 2012. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Xiao:2007:ANS

[XB07] Jin Xiao and Raouf Boutaba. Assessing network service profitability: modeling from market science perspective. IEEE/ACM Transactions on Networking, 15(6):1307–1320, December 2007. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Xiao:2014:ROU

[XB14] Jin Xiao and Raouf Boutaba. Reconciling the overlay and underlay tussle. IEEE/ ACM Transactions on Networking, 22(5):1489–1502, October 2014. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Xi:2023:RFA

[XBM^{+}23] Shaoke Xi, Kai Bu, Wensen Mao, Xiaoyu Zhang, Kui Ren, and Xinxin Ren. Rule-Out forwarding anomalies for SDN. IEEE/ ACM Transactions on Networking, 31(1):395– 407, 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2022.3194970.

Xing:2008:SLS

[XC08] Yiping Xing and R. Chandramouli. Stochastic learning solution for distributed discrete power control game in wireless data networks. IEEE/ACM Transactions on Networking, 16 (4):932–944, August 2008. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Xiao:2017:AJE

[XCC^{+}17] Qingjun Xiao, Shigang Chen, Min Chen, Yian Zhou, Zhiping Cai, Junzhou Luo, Qingjun Xiao, Shigang Chen, Min Chen, Yian Zhou, Zhiping Cai, and Junzhou Luo. Adaptive joint estimation protocol for arbitrary pair of tag sets in a distributed RFID system. IEEE/ ACM Transactions on Networking, 25(5):2670–2685, October 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Xie:2018:SIR

[XCL^{+}18] Lei Xie, Qingliang Cai, Alex X. Liu, Wei Wang, Yafeng Yin, and Sanglu Lu. Synchronize inertial readings from multiple mobile devices in spatial dimension. IEEE/ACM Transactions on Networking, 26(5): 2146–2159, October 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Xiao:2019:PSE

[XCL^{+}19] Qingjun Xiao, Shigang Chen, Jia Liu, Guang Cheng, and Junzhou Luo. A protocol for simultaneously estimating moments and popular groups in a multigroup RFID system. IEEE/ACM Transactions on Networking, 27(1):143– 158, February 2019. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Xie:2022:DSC

[XCL^{+}22] Renjie Xie, Jiahao Cao, Qi Li, Kun Sun, Guofei Gu, Mingwei Xu, and Yuan Yang. Disrupting the SDN control channel via shared links: Attacks and countermeasures. IEEE/ACM Transactions on Networking, 30(5):2158– 2172, 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2022.3169136.

Xiao:2023:UAS

[XCQ^{+}23] Qingjun Xiao, Xuyuan Cai, Yifei Qin, Zhiying Tang, Shigang Chen, and Yu Liu. Universal and accurate sketch for estimating heavy hitters and moments in data streams. IEEE/ACM Transactions on Networking, 31(5):1919– 1934, 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2022.3216025.

Xu:2011:LSR

[XCR11] Dahai Xu, Mung Chiang, and Jennifer Rexford. Linkstate routing with hop-by-hop forwarding can achieve optimal traffic engineering. IEEE/ACM Transactions on Networking, 19(6): 1717–1730, December 2011. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). See corrections [XCR15].

Xu:2015:CLS

[XCR15] Dahai Xu, Mung Chiang, and Jennifer Rexford. Corrections to “Link-state routing with hop-by-hop forwarding can achieve optimal traffic engineering”. IEEE/ACM Transactions on Networking, 23(5):1702– 1703, October 2015. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). See [XCR11].

Xiao:2018:NTL

[XCS^{+}18] Fu Xiao, Lei Chen, Chaoheng Sha, Lijuan Sun, Ruchuan Wang, Alex X. Liu, and Faraz Ahmed. Noise tolerant localization for sensor networks. IEEE/ ACM Transactions on Networking, 26(4):1701–1714, August 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Xue:2020:DDS

[XCV^{+}20] Jiachen Xue, Muhammad Usama Chaudhry, Balajee Vamanan, T. N. Vijaykumar, and Mithuna Thottethodi. Dart: Divide and specialize for fast response to congestion in RDMA-based datacenter networks. IEEE/ ACM Transactions on Networking, 28(1):322–335, February 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL https://

dl.acm.org/doi/abs/10.
1109/TNET.2019.2961671.

Xie:2020:AFRa

[XCW^{+}20a] Kun Xie, Yuxiang Chen, Xin Wang, Gaogang Xie, Jiannong Cao, and Jigang Wen. Accurate and fast recovery of network monitoring data: a GPU accelerated matrix completion. IEEE/ACM Transactions on Networking, 28(3):958– 971, June 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2020.2976129.

Xie:2020:AFRb

[XCW^{+}20b] Kun Xie, Yuxiang Chen, Xin Wang, Gaogang Xie, Jiannong Cao, Jigang Wen, Guangming Yang, and Jiaqi Sun. Accurate and fast recovery of network monitoring data with GPU-accelerated tensor completion. IEEE/ ACM Transactions on Networking, 28(4):1601–1614, August 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2020.2987845.

Xu:2006:CAF

[XCX^{+}06] Dahai Xu, Yang Chen, Yizhi Xiong, Chunming Qiao, and Xin He. On the complexity of and algorithms for finding the shortest path with a disjoint counterpart. IEEE/ ACM Transactions on Networking, 14(1):147–158, February 2006. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Xiao:2017:CEE

[XCZ^{+}17] Qingjun Xiao, Shigang Chen, You Zhou, Min Chen, Junzhou Luo, Tengli Li, and Yibei Ling. Cardinality estimation for elephant flows: a compact solution based on virtual register sharing. IEEE/ ACM Transactions on Networking, 25(6):3738–3752, December 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Xiao:2020:ECA

[XCZL20] Qingjun Xiao, Shigang Chen, You Zhou, and Junzhou Luo. Estimating cardinality for arbitrarily large data stream with improved memory efficiency. IEEE/ACM Transactions on Networking, 28(2):433– 446, April 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL https://

                                        dl.acm.org/doi/abs/10.
                                        1109/TNET.2020.2970860.

Xia:2023:SSA

[XDZ^{+}23] Rui Xia, Haipeng Dai, Jiaqi Zheng, Rong Gu, Xiaoyu Wang, Weijun Wang, and Guihai Chen. SAFE: Service availability via failure elimination through VNF scaling. IEEE/ACM Transactions on Networking, 31(5): 2042–2057, 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2022.3233488.

Xue:2013:DGC

[XE13] Dongyue Xue and Eylem Ekici. Delay-guaranteed cross-layer scheduling in multihop wireless networks. IEEE/ACM Transactions on Networking, 21(6):1696– 1707, December 2013. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Xie:2018:GFA

[XFCW18] Pengjin Xie, Jingchao Feng, Zhichao Cao, and Jiliang Wang. GeneWave: Fast authentication and key agreement on commodity mobile devices. IEEE/ ACM Transactions on Networking, 26(4):1688–1700, August 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Xie:2006:ILR

[XFS06] Feng Xie, Gang Feng, and Chee Kheong Siew. The impact of loss recovery on congestion control for reliable multicast. IEEE/ ACM Transactions on Networking, 14(6):1323–1335, December 2006. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Xu:2005:IQV

[XG05] Ying Xu and Roch Guérin.

Individual QoS versus aggregate QoS: a loss performance study. IEEE/ACM Transactions on Networking, 13(2):370–383, April 2005. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Xue:2014:PTA

[XGF^{+}14] Guoliang Xue, Ravi Gottapu, Xi Fang, Dejun Yang, and Krishnaiyan Thulasiraman. A polynomial-time algorithm for computing disjoint lightpath pairs in minimum isolated-failure-immune WDM optical networks. IEEE/ACM Transactions on Networking, 22 (2):470–483, April 2014. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Xie:2019:VDS

[XGQ^{+}19] Junjie Xie, Deke Guo, Chen Qian, Lei Liu, Bangbang Ren, and Honghui Chen. Validation of distributed SDN control plane under uncertain failures. IEEE/ACM Transactions on Networking, 27(3):1234– 1247, June 2019. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Xiao:2020:PPU

[XGW^{+}20] Mingjun Xiao, Guoju Gao, Jie Wu, Sheng Zhang, and Liusheng Huang. Privacypreserving user recruitment protocol for mobile crowdsensing. IEEE/ ACM Transactions on Networking, 28(2):519–532, April 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL https://

dl.acm.org/doi/abs/10.
1109/TNET.2019.2962362.

Xu:2018:AHS

[XHC^{+}18] Hongli Xu, He Huang, Shigang Chen, Gongming Zhao, and Liusheng Huang. Achieving high scalability through hybrid switching in software-defined networking. IEEE/ACM Transactions on Networking, 26(1): 618–632, February 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Xia:2005:ABC

[XHK^{+}05] Yong Xia, David Harrison, Shivkumar Kalyanaraman, Kishore Ramachandran, and Arvind Venkatesan. Accumulationbased congestion control. IEEE/ACM Transactions on Networking, 13(1):69– 80, February 2005. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Xie:2004:SSA

[XHN04] Feng Xie, Joseph L. Hammond, and Daniel L. Noneaker. Steady-state analysis of a split-connection scheme for Internet access through a wireless terminal. IEEE/ACM Transactions on Networking, 12 (3):515–525, June 2004. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Xue:2022:SSC

[XHY^{+}22] Kaiping Xue, Peixuan He, Jiayu Yang, Qiudong Xia, and David S. L. Wei. SCD2: Secure content delivery and deduplication with multiple content providers in information centric networking. IEEE/ACM Transactions on Networking, 30(4): 1849–1864, 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2022.3155110.

Xue:2019:SEA

[XHZ^{+}19] Kaiping Xue, Peixuan He, Xiang Zhang, Qiudong Xia, David S. L. Wei, Hao Yue, and Feng Wu. A secure, efficient, and accountable edge-based access control framework for information centric networks. IEEE/ACM Transactions on Networking, 27(3):1220– 1233, June 2019. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Xin:2007:BAD

[Xin07] Chunsheng Xin. Blocking analysis of dynamic traffic grooming in mesh WDM optical networks. IEEE/ACM Transactions on Networking, 15(3):721– 733, June 2007. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Xie:2006:PLA

[XK06a] Liang-Liang Xie and P. R.

Kumar. On the pathloss attenuation regime for positive cost and linear scaling of transport capacity in wireless networks. IEEE/ACM Transactions on Networking, 14 (SI):2313–2328, June 2006. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Xue:2006:ICL

[XK06b] Feng Xue and P. R. Kumar. On the θ-coverage and connectivity of large random networks. IEEE/ ACM Transactions on Networking, 14(SI):2289–2299, June 2006. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Xie:1995:DGV

[XL95] Geoffrey G. Xie and Simon S. Lam. Delay guarantee of virtual clock server. IEEE/ACM Transactions on Networking, 3(6):683–689, December 1995. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/

                                        journals/ton/1995-3-6/
                                        p683-xie/.

Xie:1998:RBT

[XL98] Geoffrey G. Xie and Simon S. Lam. Realtime block transfer under a link-sharing hierarchy. IEEE/ACM Transactions on Networking, 6(1):30–41, February 1998. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/
                                        journals/ton/1998-6-1/
                                        p30-xie/.

Xiao:1999:AAW

[XL99] Gaoxi Xiao and Yiu-Wing Leung. Algorithms for allocating wavelength converters in all-optical networks. IEEE/ACM Transactions on Networking, 7(4):545– 557, August 1999. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/
                                        journals/ton/1999-7-4/
                                        p545-xiao/.

Xu:2005:FTB

[XL05] Jun Xu and Richard J. Lipton. On fundamental tradeoffs between delay bounds and computational complexity in packet scheduling algorithms. IEEE/ACM Transactions on Networking, 13(1):15–28, February 2005. CODEN IEANEP.

ISSN 1063-6692 (print), 1558-2566 (electronic).

Xu:2011:TST

[XL11a] Ping Xu and Xiang-Yang Li. TOFU: semi-truthful online frequency allocation mechanism for wireless network. IEEE/ACM Transactions on Networking, 19 (2):433–446, April 2011. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Xuan:2011:NCM

[XL11b] Yuanzhe Xuan and Chin-Tau Lea. Network-coding multicast networks with QoS guarantees. IEEE/ ACM Transactions on Networking, 19(1):265–274, February 2011. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Xie:2023:CPA

[XL23] Hong Xie and John C. S.

Lui. Cooperation preference aware Shapley value: Modeling, algorithms and applications. IEEE/ ACM Transactions on Networking, 31(6):2439– 2453, 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2022.3228933.

Xiang:2016:JLC

[XLAC16] Yu Xiang, Tian Lan, Vaneet Aggarwal, and Yih-Farn R. Chen. Joint latency and cost optimization for erasure-coded data center storage. IEEE/ ACM Transactions on Networking, 24(4):2443–2457, August 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Xie:2024:ENC

[XLD^{+}24] Guorui Xie, Qing Li, Guanglin Duan, Jiaye Lin, Yutao Dong, Yong Jiang, Dan Zhao, and Yuan Yang. Empowering in-network classification in programmable switches by binary decision tree and knowledge distillation. IEEE/ACM Transactions on Networking, 32 (1):382–395, 2024. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2023.3287091.

Xu:2017:IDT

[XLH^{+}17] Hongli Xu, Xiang-Yang Li, Liusheng Huang, Hou Deng, He Huang, and Haibo Wang. Incremental deployment and throughput maximization routing for a hybrid SDN. IEEE/ACM Transactions on Networking, 25(3):1861– 1875, June 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Xu:2020:ORA

[XLL^{+}20] Huanle Xu, Yang Liu, Wing Cheong Lau, Tiantong Zeng, Jun Guo, and Alex X. Liu. Online resource allocation with machine variability: a bandit perspective. IEEE/ ACM Transactions on Networking, 28(5):2243–2256, October 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2020.3006906.

Xu:2021:OJS

[XLL21] Huanle Xu, Yang Liu, and Wing Cheong Lau. Optimal job scheduling with resource packing for heterogeneous servers. IEEE/ ACM Transactions on Networking, 29(4):1553–1566, August 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2021.3068201.

Xue:2023:SEQ

[XLP^{+}23] Guangtao Xue, Yijie Li, Hao Pan, Lanqing Yang, Yi-Chao Chen, Xiaoyu Ji, and Jiadi Yu. ScreenID: Enhancing QRCode security by utilizing screen dimming feature. IEEE/ ACM Transactions on Networking, 31(2):862–876, April 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2022.3203044.

Xiang:2013:CDA

[XLR13] Liu Xiang, Jun Luo, and Catherine Rosenberg. Compressed data aggregation: energy-efficient and high-fidelity data collection. IEEE/ACM Transactions on Networking, 21(6): 1722–1735, December 2013. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Xin:2022:FBU

[XLT^{+}22] Yao Xin, Wenjun Li, Guoming Tang, Tong Yang, Xiaohe Hu, and Yi Wang. FPGA-based updatable packet classification using TSS-combined bit-selecting tree. IEEE/ACM Transactions on Networking, 30(6): 2760–2775, 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2022.3181295.

Xie:2017:FTF

[XLW^{+}17a] Kun Xie, Xiaocan Li, Xin Wang, Gaogang Xie, Jigang Wen, Jiannong Cao, and Dafang Zhang. Fast tensor factorization for accurate Internet anomaly detection. IEEE/ACM Transactions on Networking, 25(6):3794–3807, December 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Xu:2017:UMT

[XLW^{+}17b] Fengli Xu, Yong Li, Huandong Wang, Pengyu Zhang, and Depeng Jin. Understanding mobile traffic patterns of large scale cellular towers in urban environment. IEEE/ACM Transactions on Networking, 25 (2):1147–1161, April 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Xie:2018:LAD

[XLW^{+}18] Kun Xie, Xiaocan Li, Xin Wang, Jiannong Cao, Gaogang Xie, Jigang Wen, Dafang Zhang, and Zheng Qin. On-line anomaly detection with high accuracy. IEEE/ACM Transactions on Networking, 26 (3):1222–1235, June 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Xu:2012:ESP

[XLWT12] XiaoHua Xu, Xiang-Yang Li, Peng-Jun Wan, and ShaoJie Tang. Efficient scheduling for periodic aggregation queries in multihop sensor networks. IEEE/ACM Transactions on Networking, 20(3):690– 698, June 2012. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Xu:2021:AAG

[XLX^{+}21] Wenzheng Xu, Weifa Liang, Zichuan Xu, Jian Peng, Dezhong Peng, Tang Liu, Xiaohua Jia, and Sajal K. Das. Approximation algorithms for the generalized team orienteering problem and its applications. IEEE/ ACM Transactions on Networking, 29(1):176–189, February 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2020.3027434.

Xia:2019:EBC

[XLZ^{+}19] Xianjin Xia, Shining Li, Yu Zhang, Bingqi Li, Yuanqing Zheng, and Tao Gu. Enabling out-of-band coordination of Wi-Fi communications on Smartphones. IEEE/ACM Transactions on Networking, 27(2):518– 531, April 2019. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Xu:2014:HTM

[XLZC14] Yang Xu, Zhaobo Liu, Zhuoyuan Zhang, and H. Jonathan Chao. Highthroughput and memoryefficient multimatch packet classification based on distributed and pipelined hash tables. IEEE/ACM Transactions on Networking, 22 (3):982–995, June 2014. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Xiong:1999:RSS

[XM99] Yijun Xiong and Lorne G.

Mason. Restoration strategies and spare capacity requirements in self-healing ATM networks. IEEE/ ACM Transactions on Networking, 7(1):98–110, February 1999. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1999-7-1/
p98-xiong/.

Xue:2015:CAD

[XME15] Dongyue Xue, Robert Murawski, and Eylem Ekici. Capacity achieving distributed scheduling with finite buffers. IEEE/ACM Transactions on Networking, 23(2):519–532, April 2015. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Xie:2022:ARC

[XNHM22] Tian Xie, Namitha Nambiar, Ting He, and Patrick McDaniel. Attack resilience of cache replacement policies: a study based on TTL approximation. IEEE/ACM Transactions on Networking, 30(6): 2433–2447, 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2022.3171720.

Xie:2023:DAT

[XOW^{+}23] Kun Xie, Yudian Ouyang, Xin Wang, Gaogang Xie, Kenli Li, Wei Liang, Jiannong Cao, and Jigang Wen. Deep adversarial tensor completion for accurate network traffic measurement. IEEE/ACM Transactions on Networking, 31(5):2101– 2116, 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2022.3233908.

Xie:2020:MLA

[XOYL20] Ning Xie, Le Ou-Yang, and Alex X. Liu. A machine learning approach to blind multi-path classification for massive MIMO systems. IEEE/ACM Transactions on Networking, 28(5):2309–2322, October 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2020.3008287.

Xie:2017:PRL

[XPL^{+}17] Qiaomin Xie, Mayank Pundir, Yi Lu, Cristina L. Abad, and Roy H. Campbell. Pandas: Robust locality-aware scheduling with stochastic delay optimality. IEEE/ACM Transactions on Networking, 25 (2):662–675, April 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Xie:2018:ARIb

[XPW^{+}18] Kun Xie, Can Peng, Xin Wang, Gaogang Xie, Jigang Wen, Jiannong Cao, Dafang Zhang, and Zheng Qin. Accurate recovery of Internet traffic data under variable rate measurements. IEEE/ACM Transactions on Networking, 26 (3):1137–1150, June 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Xie:2022:CEI

[XQG^{+}22] Junjie Xie, Chen Qian, Deke Guo, Minmei Wang, Ge Wang, and Honghui Chen. COIN: an efficient indexing mechanism for unstructured data sharing systems. IEEE/ ACM Transactions on Networking, 30(1):313–326, February 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2021.3110782.

Xu:2022:NOL

[XRL^{+}22] Zichuan Xu, Haozhe Ren, Weifa Liang, Qiufen Xia, Wanlei Zhou, Pan Zhou, Wenzheng Xu, Guowei Wu, and Mingchu Li. Near optimal learningdriven mechanisms for stable NFV markets in multitier cloud networks. IEEE/ACM Transactions on Networking, 30(6):2601– 2615, 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2022.3179295.

Xin:2021:CCD

[XS21] Liangxiao Xin and David Starobinski. Countering cascading denial of service attacks on Wi-Fi networks. IEEE/ACM Transactions on Networking, 29(3):1335– 1348, June 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2021.3062363.

Xiao:2001:DAC

[XSC01] Mingbo Xiao, Ness B. Shroff, and Edwin K. P. Chong. Distributed admission control for powercontrolled cellular wireless systems. IEEE/ACM Transactions on Networking, 9(6):790–800, December 2001. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Xiao:2003:UBP

[XSC03] Mingbo Xiao, Ness B.

Shroff, and Edwin K. P. Chong. A utility-based power-control scheme in wireless cellular systems. IEEE/ACM Transactions on Networking, 11(2):210– 221, April 2003. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Xie:2015:MNW

[XSH^{+}15] Liguang Xie, Yi Shi, Y. Thomas Hou, Wenjing Lou, Hanif D. Sherali, and Scott F. Midkiff. Multinode wireless energy charging in sensor networks. IEEE/ACM Transactions on Networking, 23(2):437– 450, April 2015. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Xie:2012:MSN

[XSHS12] Liguang Xie, Yi Shi, Y. Thomas Hou, and Hanif D. Sherali. Making sensor networks immortal: an energy-renewal approach with wireless power transfer. IEEE/ ACM Transactions on Networking, 20(6):1748–1761, December 2012. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Xiong:2022:NTS

[XSM22] Sijie Xiong, Anand D. Sarwate, and Narayan B. Mandayam. Network traffic shaping for enhancing privacy in IoT systems. IEEE/ACM Transactions on Networking, 30(3):1162– 1177, June 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2021.3140174.

Xia:2008:OMB

[XSSK08] Yong Xia, Lakshminarayanan Subramanian, Ion Stoica, and Shivkumar Kalyanaraman. One more bit is enough. IEEE/ACM Transactions on Networking, 16(6):1281–1294, December 2008. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Xue:2007:FPS

[XSZ^{+}07] Guoliang Xue, Arunabha Sen, Weiyi Zhang, Jian Tang, and Krishnaiya Thulasiraman. Finding a path subject to many additive QoS constraints. IEEE/ ACM Transactions on Networking, 15(1):201–211, February 2007. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Xu:2022:TMU

[XSZ^{+}22] Wenzheng Xu, Yueying Sun, Rui Zou, Weifa Liang, Qiufen Xia, Feng Shan, Tian Wang, Xiaohua Jia, and Zheng Li. Throughput maximization of UAV networks. IEEE/ACM Transactions on Networking, 30(2):881– 895, April 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2021.3125982.

Xie:2021:PLA

[XTHL21] Ning Xie, Haijun Tan, Lei Huang, and Alex X. Liu. Physical-layer authentication in wirelessly powered communication networks. IEEE/ACM Transactions on Networking, 29(4):1827–1840, August 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2021.3071670.

Xia:2011:RAP

[XTMM11] Ming Xia, Massimo Tornatore, Charles U. Martel, and Biswanath Mukherjee. Risk-aware provisioning for optical WDM mesh networks. IEEE/ACM Transactions on Networking, 19 (3):921–931, June 2011. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Xie:2022:FRL

[XTW^{+}22] Kun Xie, Jiazheng Tian, Xin Wang, Gaogang Xie, Jiannong Cao, Hongbo Jiang, and Jigang Wen. Fast retrieval of large entries with incomplete measurement data. IEEE/ ACM Transactions on Networking, 30(5):1955– 1969, 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2022.3160233.

Xu:2011:LIP

[XW11] Yi Xu and Wenye Wang. The limit of information propagation speed in largescale multihop wireless networks. IEEE/ACM Transactions on Networking, 19(1):209–222, February 2011. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Xu:2014:BAI

[XWG14] Kuai Xu, Feng Wang, and Lin Gu. Behavior analysis of Internet traffic via bipartite graphs and one-mode projections. IEEE/ACM Transactions on Networking, 22(3):931–942, June 2014. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Xie:2016:GCF

[XWH^{+}16] Yi Xie, Yu Wang, Haitao He, Yang Xiang, Shunzheng Yu, and Xincheng Liu. A general collaborative framework for modeling and perceiving distributed network behavior. IEEE/ACM Transactions on Networking, 24(5): 3162–3176, October 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Xiao:2022:SAR

[XWJ22] Xuedou Xiao, Wei Wang, and Tao Jiang. Sensorassisted rate adaptation for UAV MU-MIMO networks. IEEE/ACM Transactions on Networking, 30(4):1481– 1493, 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2021.3136911.

Xie:2018:MTA

[XWL^{+}18] Lei Xie, Chuyu Wang, Alex X. Liu, Jianqiang Sun, and Sanglu Lu. Multitouch in the air: Concurrent micromovement recognition using RF signals.

IEEE/ACM Transactions on Networking, 26(1):231– 244, February 2018. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Xie:2018:ARIa

[XWW^{+}18] Kun Xie, Lele Wang, Xin Wang, Gaogang Xie, Jigang Wen, Guangxing Zhang, Jiannong Cao, Dafang Zhang, Kun Xie, Xin Wang, Dafang Zhang, Jiannong Cao, Lele Wang, Gaogang Xie, Jigang Wen, and Guangxing Zhang. Accurate recovery of Internet traffic data: a sequential tensor completion approach. IEEE/ACM Transactions on Networking, 26 (2):793–806, April 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Xie:2019:ARM

[XWW^{+}19] Kun Xie, Xiangge Wang, Xin Wang, Yuxiang Chen, Gaogang Xie, Yudian Ouyang, Jigang Wen, Jiannong Cao, and Dafang Zhang. Accurate recovery of missing network measurement data with localized tensor completion. IEEE/ACM Transactions on Networking, 27(6):2222–2235, December 2019. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL https://

dl.acm.org/doi/abs/10.
1109/TNET.2019.2940147.

Xu:2023:OET

[XWW^{+}23] Chao Xu, Jessie Hui Wang, Jilong Wang, Tao Yu, Yipeng Zhou, Yuedong Xu, Di Wu, and Changqing An. Offloading elastic transfers to opportunistic vehicular networks based on imperfect trajectory prediction. IEEE/ACM Transactions on Networking, 31 (1):279–293, 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2022.3189047.

Xie:2016:CRR

[XWWC16] Kun Xie, Xin Wang, Jigang Wen, and Jiannong Cao. Cooperative routing with relay assignment in multiradio multihop wireless networks. IEEE/ACM Transactions on Networking, 24 (2):859–872, April 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Xu:2024:RMD

[XWX^{+}24] Wenzheng Xu, Chengxi Wang, Hongbin Xie, Weifa Liang, Haipeng Dai, Zichuan Xu, Ziming Wang, Bing Guo, and Sajal K. Das. Reward maximization for disaster zone monitoring with heterogeneous UAVs. IEEE/ACM Transactions on Networking, 32(1):890– 903, 2024. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2023.3300174.

Xiao:2018:OMT

[XWY^{+}18] Fu Xiao, Zhongqin Wang, Ning Ye, Ruchuan Wang, and Xiang-Yang Li. One more tag enables finegrained RFID localization and tracking. IEEE/ ACM Transactions on Networking, 26(1):161–174, February 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Xiong:2023:RLD

[XWYL23] Guojun Xiong, Shufan Wang, Gang Yan, and Jian Li. Reinforcement learning for dynamic dimensioning of cloud caches: a restless bandit approach. IEEE/ACM Transactions on Networking, 31(5):2147– 2161, 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2023.3235480.

Xiao:2014:ILW

[XXBC14] Qingjun Xiao, Bin Xiao, Kai Bu, and Jiannong Cao. Iterative localization of wireless sensor networks: an accurate and robust approach. IEEE/ACM Transactions on Networking, 22 (2):608–621, April 2014. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Xiao:2017:CAC

[XXCC17] Qingjun Xiao, Bin Xiao, Shigang Chen, and Jiming Chen. Collision-aware churn estimation in largescale dynamic RFID systems. IEEE/ACM Transactions on Networking, 25(1):392–405, February 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Xu:2019:RPA

[XXN^{+}19] Yuedong Xu, Zhujun Xiao, Tianyu Ni, Jessie Hui Wang, Xin Wang, and Eitan Altman. On the robustness of price-anticipating Kelly mechanism. IEEE/ ACM Transactions on Networking, 27(4):1558–1571, August 2019. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Xu:2023:AAH

[XXW^{+}23] Wenzheng Xu, Hongbin Xie, Chenxi Wang, Weifa Liang, Xiaohua Jia, Zichuan Xu, Pan Zhou, Weigang Wu, and Xiang Chen. An approximation algorithm for the h-hop independently submodular maximization problem and its applications. IEEE/ACM Transactions on Networking, 31(3):1216– 1229, June 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2022.3210825.

Xu:2022:SMS

[XXZ^{+}22a] Hongli Xu, Peng Xi, Gongming Zhao, Jianchun Liu, Chen Qian, and Liusheng Huang. SAFE-ME: Scalable and flexible policy enforcement in middlebox networks. IEEE/ ACM Transactions on Networking, 30(5):2246– 2261, 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2022.3167169.

Xu:2022:MDC

[XXZ^{+}22b] Wenzheng Xu, Tao Xiao, Junqi Zhang, Weifa Liang, Zichuan Xu, Xuxun Liu, Xiaohua Jia, and Sajal K. Das. Minimizing the deployment cost of UAVs for delay-sensitive data collection in IoT networks. IEEE/ACM Transactions on Networking, 30(2):812– 825, April 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2021.3123606.

Xing:2023:OLA

[XXZ^{+}23] Yitao Xing, Kaiping Xue, Yuan Zhang, Jiangping Han, Jian Li, and David S. L. WeiMember. An online learning assisted packet scheduler for MPTCP in mobile networks. IEEE/ ACM Transactions on Networking, 31(5):2297– 2312, 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2023.3246168.

Xie:2009:LSH

[XY09a] Yi Xie and Shun-Zheng Yu.

A large-scale hidden semi-Markov model for anomaly detection on user browsing behaviors. IEEE/ACM Transactions on Networking, 17(1):54–65, February 2009. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Xie:2009:MAL

[XY09b] Yi Xie and Shun-Zheng Yu.

Monitoring the applicationlayer DDoS attacks for popular websites. IEEE/ACM Transactions on Networking, 17(1):15–25, February 2009. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Xi:2010:DAM

[XY10a] Yufang Xi and Edmund M. Yeh. Distributed algorithms for minimum cost multicast with network coding. IEEE/ACM Transactions on Networking, 18 (2):379–392, April 2010. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Xi:2010:TOD

[XY10b] Yufang Xi and Edmund M. Yeh. Throughput optimal distributed power control of stochastic wireless networks. IEEE/ ACM Transactions on Networking, 18(4):1054–1066, August 2010. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Xiang:2021:ODL

[XYA^{+}21] Qiao Xiang, Haitao Yu, James Aspnes, Franck Le, Chin Guok, Linghe Kong, and Y. Richard Yang. Optimizing in the dark: Learning optimal network resource reservation through a simple request interface. IEEE/ACM Transactions on Networking, 29(2):571– 584, April 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2020.3045595.

Xu:2017:JRS

[XYL^{+}17] Hongli Xu, Zhuolong Yu, Xiang-Yang Li, Liusheng Huang, Chen Qian, Taeho Jung, Hongli Xu, Zhuolong Yu, Xiang-Yang Li, Liusheng Huang, Chen Qian, and Taeho Jung. Joint route selection and update scheduling for lowlatency update in SDNs. IEEE/ACM Transactions on Networking, 25(5):3073– 3087, October 2017. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Xu:2014:VTE

[XYLL14] Yang Xu, Chenguang Yu, Jingjiang Li, and Yong Liu. Video telephony for endconsumers: measurement study of Google+, iChat and Skype. IEEE/ACM Transactions on Networking, 22(3):826–839, June 2014. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Xu:2017:MFS

[XYQ^{+}17] Hongli Xu, Zhuolong Yu, Chen Qian, Xiang-Yang Li, Zichun Liu, and Liusheng Huang. Minimizing flow statistics collection cost using wildcard-based requests in SDNs. IEEE/ACM Transactions on Networking, 25(6):3587–3601, December 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Xu:2021:ACB

[XYT^{+}21] Zhiyuan Xu, Dejun Yang, Jian Tang, Yinan Tang, Tongtong Yuan, Yanzhi Wang, and Guoliang Xue. An actor–critic-based transfer learning framework for experience-driven networking. IEEE/ACM Transactions on Networking, 29(1):360–371, February 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2020.3037231.

Xu:2008:ITB

[XZB08] Kuai Xu, Zhi-Li Zhang, and Supratik Bhattacharyya. Internet traffic behavior profiling for network security monitoring. IEEE/ ACM Transactions on Networking, 16(6):1241–1252, December 2008. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Xiao:2019:ECA

[XZC^{+}19] Qingjun Xiao, Youlin Zhang, Shigang Chen, Min Chen, Jia Liu, Guang Cheng, and Junzhou Luo. Estimating cardinality of arbitrary expression of multiple tag sets in a distributed RFID system.

IEEE/ACM Transactions on Networking, 27(2):748– 762, April 2019. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Xia:2020:TSG

[XZC^{+}20] Hui Xia, Rui Zhang, Xiangguo Cheng, Tie Qiu, and Dapeng Oliver Wu. Two-stage game design of payoff decision-making scheme for crowdsourcing dilemmas. IEEE/ACM Transactions on Networking, 28(6):2741–2754, December 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2020.3018448.

Xia:2020:FPD

[XZG20] Xianjin Xia, Yuanqing Zheng, and Tao Gu. FTrack: Parallel decoding for LoRa transmissions. IEEE/ACM Transactions on Networking, 28(6):2573–2586, December 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2020.3018020.

Xia:2021:LDL

[XZG21] Xianjin Xia, Yuanqing Zheng, and Tao Gu. Lite-Nap: Downclocking LoRa reception. IEEE/ACM Transactions on Networking, 29(6):2632–2645, December 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2021.3096990.

Xie:2020:PLA

[XZL20] Ning Xie, Shengli Zhang, and Alex X. Liu. Physicallayer authentication in nonorthogonal multiple access systems. IEEE/ ACM Transactions on Networking, 28(3):1144–1157, June 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2020.2979058.

Xia:2021:ICC

[XZL^{+}21] Dan Xia, Xiaolong Zheng, Liang Liu, Chaoyu Wang, and Huadong Ma. c–Chirp: Towards symmetric crosstechnology communication over asymmetric channels. IEEE/ACM Transactions on Networking, 29(3):1169– 1182, June 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2021.3061083.

Xue:2008:PTA

[XZTT08] Guoliang Xue, Weiyi Zhang, Jian Tang, and Krishnaiyan Thulasiraman. Polynomial time approximation algorithms for multiconstrained QoS routing. IEEE/ACM Transactions on Networking, 16(3):656– 669, June 2008. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Ye:2009:OSP

[YAA09] Zhenzhen Ye, Alhussein A.

Abouzeid, and Jing Ai. Optimal stochastic policies for distributed data aggregation in wireless sensor networks. IEEE/ ACM Transactions on Networking, 17(5):1494–1507, October 2009. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Yazdanpanah:2015:DNC

[YASS15] Mina Yazdanpanah, Chadi Assi, Samir Sebbah, and Yousef Shayan. Does network coding combined with interference cancellation bring any gain to a wireless network? IEEE/ ACM Transactions on Networking, 23(5):1485–1500, October 2015. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Yan:2012:GRG

[YBG^{+}12] He Yan, Lee Breslau, Zihui Ge, Dan Massey, Dan Pei, and Jennifer Yates. G-RCA: a generic root cause analysis platform for service quality management in large IP networks. IEEE/ ACM Transactions on Networking, 20(6):1734–1747, December 2012. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Yu:2018:MEU

[YBQZ18] Ye Yu, Djamal Belazzougui, Chen Qian, and Qin Zhang. Memoryefficient and ultra-fast network lookup and forwarding using Othello hashing. IEEE/ACM Transactions on Networking, 26(3):1151– 1164, June 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Yun:2010:ODP

[YBX^{+}10] Ziqiu Yun, Xiaole Bai, Dong Xuan, Ten H. Lai, and Weijia Jia. Optimal deployment patterns for full coverage and kconnectivity (k <= 6) wireless sensor networks. IEEE/ACM Transactions on Networking, 18(3):934– 947, June 2010. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Yun:2012:PMW

[YBX^{+}12] Ziqiu Yun, Xiaole Bai, Dong Xuan, Weijia Jia, and Wei Zhao. Pattern mutation in wireless sensor deployment. IEEE/ ACM Transactions on Networking, 20(6):1964–1977, December 2012. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Yun:2012:SLO

[YC12] Sungho Yun and Constantine Caramanis. Systemlevel optimization in wireless networks: managing interference and uncertainty via robust optimization. IEEE/ACM Transactions on Networking, 20(2):339– 352, April 2012. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Yang:2007:NNI

[YCB07] Xiaowei Yang, David Clark, and Arthur W. Berger. NIRA: a new inter-domain routing architecture. IEEE/ ACM Transactions on Networking, 15(4):775–788, August 2007. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Yan:2021:TDT

[YCC^{+}21a] Zun Yan, Peng Cheng, Zhuo Chen, Branka Vucetic, and Yonghui Li. Twodimensional task offloading for mobile networks: an imitation learning framework. IEEE/ACM Transactions on Networking, 29(6):2494–2507, December 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2021.3093452.

Yang:2021:CIA

[YCC21b] Yang Yang, Yanjiao Chen, and Fei Chen. A compressive integrity auditing protocol for secure cloud storage. IEEE/ ACM Transactions on Networking, 29(3):1197–1209, June 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2021.3058130.

Yu:2017:PWF

[YCGH17] Haoran Yu, Man Hon Cheung, Lin Gao, and Jianwei Huang. Public Wi-Fi monetization via advertising. IEEE/ACM Transactions on Networking, 25(4): 2110–2121, August 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Yu:2009:CSL

[YCL09] Chao-Lin Yu, Cheng-Shang Chang, and Duan-Shin Lee. CR switch: a load-balanced switch with contention and reservation. IEEE/ACM Transactions on Networking, 17(5):1659–1671, October 2009. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Yao:2015:UPL

[YCL15] Zhongmei Yao, Daren B. H. Cline, and Dmitri Loguinov. Unstructured P2P link lifetimes redux. IEEE/ACM Transactions on Networking, 23(3):755– 767, June 2015. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Yuan:2011:PTP

[YCM11] Lihua Yuan, Chen-Nee Chuah, and Prasant Mohapatra. ProgME: towards programmable network measurement. IEEE/ ACM Transactions on Networking, 19(1):115–128, February 2011. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Yao:2015:EEE

[YCV15] Yanjun Yao, Qing Cao, and Athanasios V. Vasilakos. EDAL: an energy-efficient, delay-aware, and lifetimebalancing data collection protocol for heterogeneous wireless sensor networks. IEEE/ACM Transactions on Networking, 23(3):810– 823, June 2015. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Yang:2019:CES

[YCW^{+}19] Zhenjie Yang, Yong Cui, Xin Wang, Yadong Liu, Minming Li, Shihan Xiao, and Chuming Li. Costefficient scheduling of bulk transfers in inter-datacenter WANs. IEEE/ACM Transactions on Networking, 27(5):1973–1986, October 2019. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Yang:2023:BAT

[YCZ^{+}23] Dong Yang, Zongrong Cheng, Weiting Zhang, Hongke Zhang, and Xuemin Shen. Burst-aware timetriggered flow scheduling with enhanced multi-CQF in time-sensitive networks. IEEE/ACM Transactions on Networking, 31(6):2809– 2824, 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2023.3264583.

Yang:2004:EBN

[YD04] Shanchieh Jay Yang and Gustavo De Veciana. Enhancing both network and user performance for networks supporting best effort traffic. IEEE/ACM Transactions on Networking, 12(2):349–360, April 2004. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Yang:2007:IMC

[YD07] Xiangying Yang and Gustavo De Veciana. Inducing multiscale clustering using multistage MAC contention in CDMA ad hoc networks. IEEE/ACM Transactions on Networking, 15(6):1387–1400, December 2007. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Yao:2022:HTL

[YDCF^{+}22] Zhiyuan Yao, Yoann Desmouceaux, Juan-Antonio Cordero-Fuertes, Mark Townsley, and Thomas Clausen. HLB: Toward load-aware load balancing. IEEE/ ACM Transactions on Networking, 30(6):2658– 2673, 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2022.3177163.

Ye:2020:ABN

[YDLM20] Tong Ye, Jingjie Ding, Tony T. Lee, and Guido Maier. AWG-based nonblocking shuffle–exchange networks. IEEE/ACM Transactions on Networking, 28(6):2699–2712, December 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2020.3017500.

Yingchareonthawornchai:2018:SPA

[YDLT18] Sorrachai Yingchareonthawornchai, James Daly, Alex X. Liu, and Eric Torng. A sorted-partitioning approach to fast and scalable dynamic packet classification. IEEE/ACM Transactions on Networking, 26(4):1907–1920, August 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Yi:2006:TSD

[YDS06a] Yung Yi, Supratim Deb, and Sanjay Shakkottai. Time-scale decomposition and equivalent rate-based marking. IEEE/ACM Transactions on Networking, 14(5):938–950, October 2006. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Ying:2006:GSI

[YDS06b] Lei Ying, Geir E. Dullerud, and R. Srikant. Global stability of Internet congestion controllers with heterogeneous delays. IEEE/ACM Transactions on Networking, 14(3):579–591, 2006. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Yi:2010:MSL

[YDS10] Yung Yi, Gustavo De Veciana, and Sanjay Shakkottai. MAC scheduling with low overheads by learning neighborhood contention patterns. IEEE/ ACM Transactions on Networking, 18(5):1637–1650, October 2010. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Yuan:2018:ASP

[YDW18] Xingliang Yuan, Huayi Duan, and Cong Wang. Assuring string pattern matching in outsourced middleboxes. IEEE/ACM Transactions on Networking, 26(3):1362–1375, June 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Yan:2024:PIR

[YDY^{+}24] Fulong Yan, Xiong Deng, Changshun Yuan, Boyuan Yan, and Chongjin Xie. On the performance investigation of a recursive fast optical switch-based high performance computing network architecture. IEEE/ACM Transactions on Networking, 32 (1):777–790, 2024. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2023.3302650.

Yun:2024:RED

[YEMJ24] Jihyeon Yun, Atilla Eryilmaz, Jun Moon, and Changhee Joo. Remote estimation for dynamic IoT sources under sublinear communication costs. IEEE/ACM Transactions on Networking, 32(2):1333– 1345, April 2024. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2023.3314506.

Younis:2005:FSL

[YF05] Ossama Younis and Sonia Fahmy. FlowMate: scalable on-line flow clustering. IEEE/ACM Transactions on Networking, 13 (2):288–301, April 2005. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Yan:2002:QAM

[YFB02] Shuqian Yan, Michalis Faloutsos, and Anindo Banerjea. QoS-aware multicast routing for the Internet: the design and evaluation of QoSMIC. IEEE/ACM Transactions on Networking, 10(1):54– 66, February 2002. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Yang:2022:WTF

[YFM^{+}22] Edwin Yang, Song Fang, Ian Markwood, Yao Liu, Shangqing Zhao, Zhuo Lu, and Haojin Zhu. Wireless training-free keystroke inference attack and defense. IEEE/ACM Transactions on Networking, 30(4):1733– 1748, 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2022.3147721.

Yu:2010:DRF

[YG10] Zhen Yu and Yong Guan. A dynamic en-route filtering scheme for data reporting in wireless sensor networks. IEEE/ACM Transactions on Networking, 18(1):150– 163, February 2010. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Yin:2010:SLO

[YGC10] Changchuan Yin, Long Gao, and Shuguang Cui. Scaling laws for overlaid wireless networks: a cognitive radio network versus a primary network. IEEE/ ACM Transactions on Networking, 18(4):1317–1329, August 2010. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Yu:2010:SNO

[YGKX10] Haifeng Yu, Phillip B. Gibbons, Michael Kaminsky, and Feng Xiao. SybilLimit: a near-optimal social network defense against Sybil attacks. IEEE/ACM Transactions on Networking, 18(3):885–898, June 2010. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Yu:2019:ETB

[YGL^{+}19] Jihong Yu, Wei Gong, Jiangchuan Liu, Lin Chen, and Kehao Wang. On efficient tree-based tag search in large-scale RFID systems. IEEE/ACM Transactions on Networking, 27 (1):42–55, February 2019. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Yu:2021:PRP

[YHCL21] Che-Hao Yu, Lin Huang, Cheng-Shang Chang, and Duan-Shin Lee. Poisson receivers: a probabilistic framework for analyzing coded random access. IEEE/ACM Transactions on Networking, 29(2):862– 875, April 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2021.3050485.

Ye:2004:MAC

[YHE04] Wei Ye, John Heidemann, and Deborah Estrin. Medium access control with coordinated adaptive sleeping for wireless sensor networks. IEEE/ACM Transactions on Networking, 12 (3):493–506, June 2004. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Yang:2021:FLE

[YHH^{+}21] Deliang Yang, Xuan Huang, Jun Huang, Xiangmao Chang, Guoliang Xing, and Yang Yang. A first look at energy consumption of NB-IoT in the wild: Tools and large-scale measurement. IEEE/ACM Transactions on Networking, 29(6):2616–2631, December 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2021.3096656.

Yang:2015:NCC

[YJ15] Chao Yang and Scott Jordan. A novel coordinated connection access control and resource allocation framework for 4G wireless networks. IEEE/ ACM Transactions on Networking, 23(4):1328–1341, August 2015. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Yousefizadeh:2005:LMM

[YJH05] Homayoun Yousefi’zadeh, Hamid Jafarkhani, and Amir Habibi. Layered media multicast control (LMMC): rate allocation and partitioning. IEEE/ ACM Transactions on Networking, 13(3):540–553, June 2005. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Yang:2019:AMU

[YJL^{+}19] Tong Yang, Jie Jiang, Peng Liu, Qun Huang, Junzhi Gong, Yang Zhou, Rui Miao, Xiaoming Li, and Steve Uhlig. Adaptive measurements using one elastic sketch. IEEE/ACM Transactions on Networking, 27(6):2236–2251, December 2019. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL https://

                                        dl.acm.org/doi/abs/10.
                                        1109/TNET.2019.2943939.

Yang:2015:DIS

[YJZW15] Yang Yang, Miao Jin, Yao Zhao, and Hongyi Wu. Distributed information storage and retrieval in 3-D sensor networks with general topologies. IEEE/ ACM Transactions on Networking, 23(4):1149–1162, August 2015. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Yasodharan:2023:ROB

[YK23] Sarath Yasodharan and Anurag Kumar. Revenue optimal bandwidth allocation in a class of multihop networks: Algorithms and asymptotic optimality. IEEE/ACM Transactions on Networking, 31(5): 2251–2266, 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2023.3249480.

Yun:2023:SFL

[YKB^{+}23] Won Joon Yun, Yunseok Kwak, Hankyul Baek, Soyi Jung, Mingyue Ji, Mehdi Bennis, Jihong Park, and Joongheon Kim. SlimFL: Federated learning with superposition coding over slimmable neural networks. IEEE/ACM Transactions on Networking, 31(6):2499– 2514, 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2022.3231864.

Yu:2008:SDA

[YKGF08] Haifeng Yu, Michael Kaminsky, Phillip B. Gibbons, and Abraham D. Flaxman. SybilGuard: defending against sybil attacks via social networks. IEEE/ACM Transactions on Networking, 16(3):576– 589, June 2008. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Young:2013:TPC

[YKGK13] Maxwell Young, Aniket Kate, Ian Goldberg, and Martin Karsten. Towards practical communication in Byzantine-resistant DHTs. IEEE/ACM Transactions on Networking, 21(1):190– 203, February 2013. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Ye:2008:LSN

[YKKY08] Tao Ye, Hema T. Kaur, Shivkumar Kalyanaraman, and Murat Yuksel. Largescale network parameter configuration using an online simulation framework. IEEE/ACM Transactions on Networking, 16(4):777– 790, August 2008. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Younis:2011:RRO

[YKR11] Ossama Mohamed Younis, Marwan M. Krunz, and Srinivasan Ramasubramanian. ROC: resilient online coverage for surveillance applications. IEEE/ ACM Transactions on Networking, 19(1):251–264, February 2011. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Yang:2013:PBD

[YKZ^{+}13] Lei Yang, Hongseok Kim, Junshan Zhang, Mung Chiang, and Chee Wei Tan. Pricing-based decentralized spectrum access control in cognitive radio networks. IEEE/ACM Transactions on Networking, 21(2):522– 535, April 2013. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Yau:1997:ARS

[YL97] David K. Y. Yau and Simon S. Lam. Adaptive rate-controlled scheduling for multimedia applications. IEEE/ACM Transactions on Networking, 5(4):475–488, August 1997. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1997-5-4/
p475-yau/.

Yau:1998:MSS

[YL98] David K. Y. Yau and Simon S. Lam. Migrating sockets — end system support for networking with quality of service guarantees. IEEE/ ACM Transactions on Networking, 6(6):700–716, December 1998. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/
                                        journals/ton/1998-6-6/
                                        p700-yau/.

Yang:2016:RTV

[YL16] Hongkun Yang and Simon S. Lam. Real-time verification of network properties using atomic predicates. IEEE/ACM Transactions on Networking, 24 (2):887–900, April 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Yuan:2018:NSB

[YLA^{+}18] Yifei Yuan, Dong Lin, Siri Anil, Harsh Verma, Anirudh Chelluri, Rajeev Alur, and Boon Thau Loo. NetEgg: a scenario-based programming toolkit for SDN policies. IEEE/ ACM Transactions on Networking, 26(5):2104–2117, October 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Yoon:2011:FCS

[YLCP11] MyungKeun Yoon, Tao Li, Shigang Chen, and Jih-Kwon Peir. Fit a compact spread estimator in small high-speed memory. IEEE/ ACM Transactions on Networking, 19(5):1253–1264, October 2011. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Yang:2023:RWR

[YLD^{+}23] Wei Yang, Chi Lin, Haipeng Dai, Pengfei Wang, Jiankang Ren, Lei Wang, Guowei Wu, and Qiang Zhang. Robust wireless rechargeable sensor networks. IEEE/ ACM Transactions on Networking, 31(3):949–964, June 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2022.3199389.

Ye:2021:DBN

[YLF^{+}21] Yuhang Ye, Brian Lee, Ronan Flynn, Jin Xu, Guiming Fang, and Yuansong Qiao. Delay-based network utility maximization modelling for congestion control in named data networking. IEEE/ ACM Transactions on Networking, 29(5):2184–2197, October 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2021.3090174.

Ye:2015:ABN

[YLH15] Tong Ye, Tony T. Lee, and Weisheng Hu. AWGbased non-blocking Clos networks. IEEE/ACM Transactions on Networking, 23(2):491–504, April 2015. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Yan:2017:CAE

[YLH17] Fangfang Yan, Tony T. Lee, and Weisheng Hu. Congestion-aware embedding of heterogeneous bandwidth virtual data centers with hose model abstraction. IEEE/ACM Transactions on Networking, 25(2):806–819, April 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Yoon:2017:FWS

[YLK^{+}17] Changhoon Yoon, Seungsoo Lee, Heedo Kang, Taejune Park, Seungwon Shin, Vinod Yegneswaran, Phillip Porras, and Guofei Gu. Flow wars: Systemizing the attack surface and defenses in softwaredefined networks. IEEE/ ACM Transactions on Networking, 25(6):3514–3530, December 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Yang:2010:BTL

[YLL10] Zheng Yang, Yunhao Liu, and Xiang-Yang Li. Beyond trilateration: on the localizability of wireless ad hoc networks. IEEE/ACM Transactions on Networking, 18(6):1806–1814, December 2010. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Yang:2017:TSM

[YLL^{+}17] Lei Yang, Yao Li, Qiongzheng Lin, Huanyu Jia, Xiang-Yang Li, and Yunhao Liu. Tagbeat: Sensing mechanical vibration period with COTS RFID systems. IEEE/ACM Transactions on Networking, 25(6):3823– 3835, December 2017. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Ye:2021:CBM

[YLL21] Jiancheng Ye, Ka-Cheong Leung, and Steven H. Low. Combating bufferbloat in multi-bottleneck networks: Theory and algorithms. IEEE/ACM Transactions on Networking, 29(4):1477– 1493, August 2021. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL https:/

/dl.acm.org/doi/10.1109/
TNET.2021.3066505.

Yau:2005:DAD

[YLLY05] David K. Y. Yau, John C. S. Lui, Feng Liang, and Yeung Yam. Defending against distributed denial-of-service attacks with max-min fair servercentric router throttles. IEEE/ACM Transactions on Networking, 13(1):29– 42, February 2005. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Yang:2017:SFS

[YLS^{+}17] Tong Yang, Alex X. Liu, Muhammad Shahzad, Dongsheng Yang, Qiaobin Fu, Gaogang Xie, Xiaoming Li, Tong Yang, Alex X. Liu, Muhammad Shahzad, Dongsheng Yang, Qiaobin Fu, Gaogang Xie, and Xiaoming Li. A shifting framework for set queries. IEEE/ACM Transactions on Networking, 25(5):3116– 3131, October 2017. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Yuan:2020:OAA

[YLWH20] Dingwen Yuan, Hsuan-Yin Lin, Jörg Widmer, and Matthias Hollick. Optimal and approximation algorithms for joint routing and scheduling in millimeter-wave cellular networks. IEEE/ACM Transactions on Networking, 28(5):2188–2202, October 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.

acm.org/doi/10.1109/TNET.
2020.3006312.

Yun:2016:DOC

[YLY^{+}16] Donggyu Yun, Dongmyung Lee, Se-Young Yun, Jinwoo Shin, and Yung Yi. Delay optimal CSMA with linear virtual channels under a general topology. IEEE/ ACM Transactions on Networking, 24(5):2847–2857, October 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Yang:2017:SVN

[YLYL17] Hongkun Yang, Simon S.

Lam, Hongkun Yang, and Simon S. Lam. Scalable verification of networks with packet transformers using atomic predicates. IEEE/ACM Transactions on Networking, 25(5):2900– 2915, October 2017. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Yuen:2005:PBR

[YM05] Clement Yuen and Peter Marbach. Price-based rate control in random access networks. IEEE/ ACM Transactions on Networking, 13(5):1027–1040, October 2005. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Yagan:2016:WSN

[YM16] Osman Yagan and Armand M. Makowski. Wireless sensor networks under the random pairwise key predistribution scheme: Can resiliency be achieved with small key rings? IEEE/ACM Transactions on Networking, 24(6):3383– 3396, December 2016. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Yu:2008:MBA

[YMKC08] Xunqi Yu, James W. Modestino, Ragip Kurceren, and Yee Sin Chan. A model-based approach to evaluation of the efficacy of FEC coding in combating network packet losses. IEEE/ACM Transactions on Networking, 16(3):628– 641, June 2008. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Yener:1997:IAO

[YMO97] Bülent Yener, Spyridon Matsoukas, and Yoram Ofek. Iterative approach to optimizing convergence routing priorities. IEEE/ ACM Transactions on Networking, 5(4):530–542, August 1997. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/

journals/ton/1997-5-4/
p530-yener/.

Yaiche:2000:GTF

[YMR00] Haïkel Yaïche, Ravi R.

Mazumdar, and Catherine Rosenberg. A game theoretic framework for bandwidth allocation and pricing in broadband networks. IEEE/ACM Transactions on Networking, 8 (5):667–678, 2000. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/2000-8-5/
p667-yaiche/.

Yu:2018:NBA

[YN18] Hao Yu and Michael J.

Neely. A new backpressure algorithm for joint rate control and routing with vanishing utility optimality gaps and finite queue lengths. IEEE/ ACM Transactions on Networking, 26(4):1605–1618, August 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Yu:2019:LAO

[YN19] Hao Yu and Michael J.

Neely. Learning-aided optimization for energyharvesting devices with outdated state information. IEEE/ACM Transactions on Networking, 27(4):1501– 1514, August 2019. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Yekkehkhany:2020:BGP

[YN20] Ali Yekkehkhany and Rakesh Nagi. Blind GB-PANDAS: a blind throughput-optimal load balancing algorithm for affinity scheduling. IEEE/ACM Transactions on Networking, 28(3):1199– 1212, June 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2020.2978195.

Yuan:2009:ORF

[YNDM09] Xin Yuan, Wickus Nienaber, Zhenhai Duan, and Rami Melhem. Oblivious routing in fat-tree based system area networks with uncertain traffic demands. IEEE/ACM Transactions on Networking, 17(5):1439– 1452, October 2009. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Yu:2017:DSC

[YNZ^{+}17] Dongxiao Yu, Li Ning, Yifei Zou, Jiguo Yu, Xiuzhen Cheng, and Francis C. M. Lau. Distributed spanner construction with physical interference: Constant stretch and linear sparseness. IEEE/ACM Transactions on Networking, 25(4):

2138–2151, August 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Yallouz:2017:TQA

[YO17] Jose Yallouz and Ariel Orda. Tunable QoSaware network survivability. IEEE/ACM Transactions on Networking, 25(1): 139–149, February 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Yener:1997:CDC

[YOY97] Bülent Yener, Yoram Ofek, and Moti Yung. Combinatorial design of congestionfree networks. IEEE/ACM Transactions on Networking, 5(6):989–1000, December 1997. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1997-5-6/
p989-yener/.

Yang:2019:CAS

[YPA19] Chien-Sheng Yang, Ramtin Pedarsani, and A. Salman Avestimehr. Communicationaware scheduling of serial tasks for dispersed computing. IEEE/ACM Transactions on Networking, 27(4): 1330–1343, August 2019. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Yang:2021:ECD

[YPA21] Chien-Sheng Yang, Ramtin Pedarsani, and A. Salman Avestimehr. Edge computing in the dark: Leveraging contextual-combinatorial bandit and coded computing. IEEE/ACM Transactions on Networking, 29(3): 1022–1031, June 2021. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1109/
                                        TNET.2021.3058685.

Yeom:2001:MTB

[YR01] Ikjun Yeom and A. L. Narasimha Reddy. Modeling TCP behavior in a differentiated services network. IEEE/ACM Transactions on Networking, 9(1):31– 46, 2001. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/articles/
                                        journals/ton/2001-9-1/
                                        p31-yeom/p31-yeom.pdf;
                                         http://www.acm.org/pubs/
                                        citations/journals/ton/
                                        2001-9-1/p31-yeom/.

Yallouz:2018:MWL

[YRB^{+}18] Jose Yallouz, Ori Rottenstreich, Peter Babarczi, Avi Mendelson, and Ariel Orda. Minimumweight link-disjoint Node- “Somewhat disjoint” paths. IEEE/ACM Transactions on Networking, 26(3):1110– 1122, June 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Yallouz:2016:TSS

[YRO16] Jose Yallouz, Ori Rottenstreich, and Ariel Orda. Tunable survivable spanning trees. IEEE/ACM Transactions on Networking, 24(3):1853–1866, June 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Yadav:2012:DAG

[YRRR12] Sandeep Yadav, Ashwath Kumar Krishna Reddy, A. L. Narasimha Reddy, and Supranamaya Ranjan. Detecting algorithmically generated domain-flux attacks with DNS traffic analysis. IEEE/ACM Transactions on Networking, 20(5): 1663–1677, October 2012. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Yaron:1993:PSC

[YS93] Opher Yaron and Moshe Sidi. Performance and stability of communication networks via robust exponential bounds. IEEE/ACM Transactions on Networking, 1(3):372– 385, June 1993. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.

                                        acm.org/pubs/citations/
                                        journals/ton/1993-1-3/
                                        p372-yaron/.

Yi:2007:HHC

[YS07] Yung Yi and Sanjay Shakkottai. Hop-by-hop congestion control over a wireless multi-hop network. IEEE/ACM Transactions on Networking, 15(1):133– 144, February 2007. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Yang:2015:TRC

[YS15] Yang Yang and Ness B. Shroff. Throughput of rateless codes over broadcast erasure channels. IEEE/ ACM Transactions on Networking, 23(1):126–137, February 2015. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Yang:2021:CIU

[YS21] Feihong Yang and Yuan Shen. Critical intensity for unbounded sequential localizability. IEEE/ ACM Transactions on Networking, 29(3):1321–1334, June 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2021.3059743.

Yan:2016:TTO

[YSC16] Li Yan, Haiying Shen, and Kang Chen. TSearch: Target-oriented low-delay node searching in DTNs with social network properties. IEEE/ACM Transactions on Networking, 24(6): 3841–3855, December 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Yan:2018:MDC

[YSC18] Li Yan, Haiying Shen, and Kang Chen. MobiT: Distributed and congestionresilient trajectory-based routing for vehicular delay tolerant networks. IEEE/ ACM Transactions on Networking, 26(3):1078–1091, June 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Yao:2014:NCR

[YSJL14] Hongyi Yao, Danilo Silva, Sidharth Jaggi, and Michael Langberg. Network codes resilient to jamming and eavesdropping. IEEE/ ACM Transactions on Networking, 22(6):1978–1987, December 2014. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Yang:2014:TCA

[YSL^{+}14] Peng Yang, Juan Shao, Wen Luo, Lisong Xu, Jitender Deogun, and Ying Lu. TCP congestion avoidance algorithm identification. IEEE/ACM Transactions on Networking, 22(4): 1311–1324, August 2014. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Ying:2011:CSP

[YSRL11] Lei Ying, Sanjay Shakkottai, Aneesh Reddy, and Shihuan Liu. On combining shortest-path and back-pressure routing over multihop wireless networks. IEEE/ACM Transactions on Networking, 19(3):841– 854, June 2011. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Ying:2011:CBB

[YSTL11] Lei Ying, R. Srikant, Don Towsley, and Shihuan Liu. Cluster-based backpressure routing algorithm. IEEE/ACM Transactions on Networking, 19(6):1773– 1786, December 2011. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Yun:2016:DMA

[YSY16] Se-Young Yun, Jinwoo Shin, and Yung Yi. Distributed medium access over time-varying channels. IEEE/ACM Transactions on Networking, 24(5):3000– 3013, October 2016. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Yang:2015:DAS

[YSZL15] Lei Yang, Yalin E. Sagduyu, Junshan Zhang, and Jason H. Li. Deadlineaware scheduling with adaptive network coding for real-time traffic. IEEE/ ACM Transactions on Networking, 23(5):1430–1443, October 2015. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Yu:2005:QPG

[YTJQ05] Xiang Yu, Ian Li-Jin Thng, Yuming Jiang, and Chunming Qiao. Queueing processes in GPS and PGPS with LRD traffic inputs. IEEE/ACM Transactions on Networking, 13(3):676– 689, June 2005. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Yu:2012:LEN

[YTL12] Minlan Yu, Marina Thottan, and Li Li. Latency equalization as a new network service primitive. IEEE/ACM Transactions on Networking, 20(1): 125–138, February 2012. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Yuan:2002:HAM

[Yua02] Xin Yuan. Heuristic algorithms for multiconstrained quality-of-service routing. IEEE/ACM Transactions on Networking, 10(2):244– 256, April 2002. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Yin:2007:BAA

[YW07] Heng Yin and Haining Wang. Building an application-aware IPsec policy system. IEEE/ACM Transactions on Networking, 15(6):1502–1513, December 2007. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Yang:2011:FFI

[YW11] Zhipeng Yang and Hongyi Wu. FINDERS: a featherlight information network with delay-endurable RFID support. IEEE/ACM Transactions on Networking, 19(4):961–974, August 2011. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Yang:2008:TLN

[YWA08] Xiaowei Yang, David Wetherall, and Thomas Anderson. TVA: a DoS-limiting network architecture. IEEE/ ACM Transactions on Networking, 16(6):1267–1280, December 2008. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Yu:2021:FHP

[YWH21] Tianqi Yu, Xianbin Wang, and Jianling Hu. A fast hierarchical physical topology update scheme for edge-cloud collaborative IoT systems. IEEE/ ACM Transactions on Networking, 29(5):2254–2266, October 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2021.3085031.

Yang:2024:AAB

[YWHW24] Xiaoxue Yang, Hao Wang, Bing Hu, and Chunming Wu. ABOI: AWGR-Based optical interconnects for single-wavelength and multi-wavelength. IEEE/ ACM Transactions on Networking, 32(2):1124–1139, April 2024. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2023.3314096.

Yang:2007:DOC

[YWK07] Yaling Yang, Jun Wang, and Robin Kravets. Distributed optimal contention window control for elastic traffic in single-cell wireless LANs. IEEE/ACM Transactions on Networking, 15(6):1373–1386, December 2007. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Yao:2009:NIM

[YWLL09] Zhongmei Yao, Xiaoming Wang, Derek Leonard, and Dmitri Loguinov. Node isolation model and age-based neighbor selection in unstructured P2P networks. IEEE/ACM Transactions on Networking, 17(1):144– 157, February 2009. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Yun:2019:ABD

[YWRK19] Daqing Yun, Chase Q. Wu, Nageswara S. V. Rao, and Rajkumar Kettimuthu. Advising big data transfer over dedicated connections based on profiling optimization. IEEE/ACM Transactions on Networking, 27(6):2280–2293, December 2019. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL https://

                                        dl.acm.org/doi/abs/10.
                                        1109/TNET.2019.2943884.

Yang:2023:ETS

[YWW^{+}23] Zheng Yang, Xu Wang, Jiahang Wu, Yi Zhao, Qiang Ma, Xin Miao, Li Zhang, and Zimu Zhou. Edge-Duet: Tiling small object detection for edge assisted autonomous mobile vision. IEEE/ACM Transactions on Networking, 31(4):1765– 1778, 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2022.3223412.

Yuan:2024:JOQ

[YWW^{+}24] Yali Yuan, Weijun Wang, Yuhan Wang, Sripriya Srikant Adhatarao, Bangbang Ren, Kai Zheng, and Xiaoming Fu. Joint optimization of QoE and fairness for adaptive video streaming in heterogeneous mobile environments. IEEE/ACM Transactions on Networking, 32 (1):50–64, 2024. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2023.3277729.

Yun:2023:ETT

[YWZ^{+}23] Xiaochun Yun, Yipeng Wang, Yongzheng Zhang, Chen Zhao, and Zijian Zhao. Encrypted TLS traffic classification on cloud platforms. IEEE/ ACM Transactions on Networking, 31(1):164– 177, 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.

                                        acm.org/doi/10.1109/TNET.
                                        2022.3191312.

Yuan:2023:MBC

[YWZG23] Longzhi Yuan, Qiwei Wang, Jia Zhao, and Wei Gong. Multiprotocol backscatter with commodity radios for personal IoT sensors. IEEE/ACM Transactions on Networking, 31(3):1132– 1144, June 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2022.3213913.

Yun:2016:SAA

[YWZZ16] Xiaochun Yun, Yipeng Wang, Yongzheng Zhang, and Yu Zhou. A semanticsaware approach to the automated network protocol identification. IEEE/ ACM Transactions on Networking, 24(1):583–595, February 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Yang:2018:AMW

[YXAZ^{+}18] Guang Yang, Ming Xiao, Hussein Al-Zubaidy, Yongming Huang, and James Gross. Analysis of millimeterwave multi-hop networks with full-duplex buffered relays. IEEE/ACM Transactions on Networking, 26(1):576–590, February 2018. CODEN IEANEP.

ISSN 1063-6692 (print), 1558-2566 (electronic).

Yan:2018:AWR

[YXC^{+}18] Bo Yan, Yang Xu, H. Jonathan Chao, H. Jonathan Chao, Yang Xu, and Bo Yan. Adaptive wildcard rule cache management for software-defined networks. IEEE/ACM Transactions on Networking, 26(2):962– 975, April 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Yang:2021:IMM

[YXCH21] Xuwei Yang, Hongli Xu, Shigang Chen, and He Huang. Indirect multi-mapping for burstiness management in software defined networks. IEEE/ACM Transactions on Networking, 29(5):2059–2072, October 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2021.3078132.

Yang:2013:GTA

[YXF^{+}13] Dejun Yang, Guoliang Xue, Xi Fang, Satyajayant Misra, and Jin Zhang. A game-theoretic approach to stable routing in max-min fair networks. IEEE/ACM Transactions on Networking, 21(6):1947–1959, December 2013. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Yang:2016:IMC

[YXFT16] Dejun Yang, Guoliang Xue, Xi Fang, and Jian Tang. Incentive mechanisms for crowdsensing: crowdsourcing with smartphones. IEEE/ACM Transactions on Networking, 24(3):1732– 1744, June 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Yu:2021:FTS

[YXH^{+}21] Mingli Yu, Tian Xie, Ting He, Patrick McDaniel, and Quinn K. Burke. Flow table security in SDN: Adversarial reconnaissance and intelligent attacks. IEEE/ ACM Transactions on Networking, 29(6):2793–2806, December 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2021.3099717.

Yang:2018:CIL

[YXL^{+}18a] Tong Yang, Gaogang Xie, Alex X. Liu, Qiaobin Fu, Yanbiao Li, Xiaoming Li, and Laurent Mathy. Constant IP lookup with FIB explosion. IEEE/ ACM Transactions on Networking, 26(4):1821–1836, August 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Yang:2018:FRA

[YXL18b] Yuan Yang, Mingwei Xu, and Qi Li. Fast rerouting against multi-link failures without topology constraint. IEEE/ACM Transactions on Networking, 26(1):384–397, February 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Yang:2019:CCD

[YXL^{+}19] Kang Yang, Tianzhang Xing, Yang Liu, Zhenjiang Li, Xiaoqing Gong, Xiaojiang Chen, and Dingyi Fang. cDeepArch: a compact deep neural network architecture for mobile sensing. IEEE/ACM Transactions on Networking, 27(5):2043–2055, October 2019. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Yu:2018:CLC

[YXY^{+}18] Xiwen Yu, Hongli Xu, Da Yao, Haibo Wang, and Liusheng Huang. Count-Max: a lightweight and cooperative sketch measurement for software-defined networks. IEEE/ACM Transactions on Networking, 26(6):2774–2786, December 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Yu:2017:CNF

[YXZ17] Ruozhou Yu, Guoliang Xue, and Xiang Zhang. The critical network flow problem: Migratability and survivability. IEEE/ACM Transactions on Networking, 25(6):3545–3558, December 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Yu:2019:PQA

[YXZ19] Ruozhou Yu, Guoliang Xue, and Xiang Zhang. Provisioning QoS-aware and robust applications in Internet of Things: a network perspective. IEEE/ ACM Transactions on Networking, 27(5):1931–1944, October 2019. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Yeung:1998:NPO

[YY98] Kwan Lawrence Yeung and Tak-Shing Peter Yum. Node placement optimization in ShuffleNets. IEEE/ ACM Transactions on Networking, 6(3):319–324, June 1998. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/

journals/ton/1998-6-3/
p319-yeung/.

Yang:2020:SCS

[YY20] Ze Yang and Kwan L. Yeung. SDN candidate selection in hybrid IP/SDN networks for single link failure protection. IEEE/ ACM Transactions on Networking, 28(1):312–321, February 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL https://

dl.acm.org/doi/abs/10.
1109/TNET.2019.2959588.

Yang:2022:MAP

[YYB^{+}22] Hui Yang, Qiuyan Yao, Bowen Bao, Ao Yu, Jie Zhang, and Athanasios V. Vasilakos. Multi-associated parameters aggregationbased routing and resources allocation in multi-core elastic optical networks. IEEE/ACM Transactions on Networking, 30(5):2145– 2157, 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2022.3164869.

Yu:2021:ELS

[YYC^{+}21] Kan Yu, Jiguo Yu, Xiuzhen Cheng, Dongxiao Yu, and Anming Dong. Efficient link scheduling solutions for the Internet of Things under Rayleigh fading. IEEE/ ACM Transactions on Networking, 29(6):2508–2521, December 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2021.3093306.

Yu:2024:RAP

[YYFC24] Kan Yu, Jiguo Yu, Zhiyong Feng, and Honglong Chen. A reassessment on applying protocol interference model under Rayleigh fading: From perspective of link scheduling. IEEE/ACM Transactions on Networking, 32(1): 238–252, 2024. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2023.3284433.

Yu:2023:IMP

[YYL23] Kan Yu, Jiguo Yu, and Chuanwen Luo. The impact of mobility on physical layer security of 5G IoT networks. IEEE/ ACM Transactions on Networking, 31(3):1042–1055, June 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2022.3208212.

Yan:2023:CAP

[YYT23] Shangyao Yan, Zhimeng Yin, and Guang Tan.

CurveLight: an accurate and practical light positioning system. IEEE/ ACM Transactions on Networking, 31(5):1950– 1964, 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2022.3224817.

Yan:2006:OBM

[YYZ06] Xijin Yan, Jun Yang, and Zhen Zhang. An outer bound for multisource multisink network coding with minimum cost consideration. IEEE/ACM Transactions on Networking, 14 (SI):2373–2385, June 2006. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Yu:2010:PMS

[YZ10] Ming Yu and Mengchu Zhou. A performance modeling scheme for multistage switch networks with phase-type and bursty traffic. IEEE/ACM Transactions on Networking, 18(4): 1091–1104, August 2010. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Yoon:2014:VMJ

[YZBR14] Jongwon Yoon, Honghai Zhang, Suman Banerjee, and Sampath Rangarajan. Video multicast with joint resource allocation and adaptive modulation and coding in 4G networks. IEEE/ACM Transactions on Networking, 22(5):1531– 1544, October 2014. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Ye:2023:FFD

[YZGC23] Minghao Ye, Junjie Zhang, Zehua Guo, and H. Jonathan Chao. FlexDATE: Flexible and disturbance-aware traffic engineering with reinforcement learning in software-defined networks. IEEE/ACM Transactions on Networking, 31(4):1433– 1448, 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2022.3217083.

Yao:2021:VQP

[YZHZ21] Xin Yao, Rui Zhang, Dingquan Huang, and Yanchao Zhang. Verifiable query processing over outsourced social graph. IEEE/ACM Transactions on Networking, 29(5):2313–2326, October 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2021.3085574.

Yang:2018:DCO

[YZL^{+}18] Lichao Yang, Heli Zhang, Xi Li, Hong Ji, and Victor C. M. Leung. A distributed computation offloading strategy in smallcell networks integrated with mobile edge computing. IEEE/ACM Transactions on Networking, 26(6): 2762–2773, December 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Yang:2019:HAA

[YZL^{+}19] Tong Yang, Haowei Zhang, Jinyang Li, Junzhi Gong, Steve Uhlig, Shigang Chen, and Xiaoming Li. Heavy-Keeper: an accurate algorithm for finding topk elephant flows. IEEE/ ACM Transactions on Networking, 27(5):1845–1858, October 2019. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Ye:2017:DCB

[YZLH17] Tong Ye, Jinghui Zhang, Tony T. Lee, and Weisheng Hu. Deflection-compensated Birkhoff–von-Neumann switches. IEEE/ACM Transactions on Networking, 25(2):879– 895, April 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Yue:2014:UTI

[YZP^{+}14] Hao Yue, Chi Zhang, Miao Pan, Yuguang Fang, and Shigang Chen. Unknowntarget information collection in sensor-enabled RFID systems. IEEE/ ACM Transactions on Networking, 22(4):1164–1175, August 2014. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Yu:2018:SLB

[YZY^{+}18] Dongxiao Yu, Yifei Zou, Jiguo Yu, Xiuzhen Cheng, Qiang-Sheng Hua, Hai Jin, and Francis C. M. Lau. Stable local broadcast in multihop wireless networks under SINR. IEEE/ACM Transactions on Networking, 26(3):1278–1291, June 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Yan:2020:SFP

[YZY^{+}20] Boyuan Yan, Yongli Zhao, Xiaosong Yu, Yajie Li, Sabidur Rahman, Yongqi He, Xiangjun Xin, and Jie Zhang. Service function path provisioning with topology aggregation in multi-domain optical networks. IEEE/ACM Transactions on Networking, 28(6):2755–2767, December 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2020.3019708.

Yu:2021:DBD

[YZY^{+}21] Dongxiao Yu, Yifei Zou, Jiguo Yu, Yu Wu, Weifeng Lv, Xiuzhen Cheng, Falko Dressler, and Francis C. M. Lau. Distributed broadcasting in dynamic networks. IEEE/ACM Transactions on Networking, 29(5):2142–2155, October 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2021.3087818.

Yu:2021:EIA

[YZZ^{+}21] Dongxiao Yu, Yifei Zou, Yong Zhang, Hao Sheng, Weifeng Lv, and Xiuzhen Cheng. An exact implementation of the abstract MAC layer via carrier sensing in dynamic networks. IEEE/ACM Transactions on Networking, 29(3):994– 1007, June 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2021.3057890.

Zhang:1995:HWA

[ZA95] Zhensheng Zhang and Anthony S. Acampora. A heuristic wavelength assignment algorithm for multihop WDM networks with wavelength routing and wavelength re-use. IEEE/ACM Transactions on Networking, 3(3):281– 288, June 1995. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1995-3-3/
p281-zhang/.

Zhang:2011:SHW

[ZA11] Jingjing Zhang and Nirwan Ansari. Scheduling hybrid WDM/TDM passive optical networks with nonzero laser tuning time. IEEE/ ACM Transactions on Networking, 19(4):1014–1027, August 2011. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Zegura:2000:ALA

[ZAFB00] Ellen W. Zegura, Mostafa H. Ammar, Zongming Fei, and Samrat Bhattacharjee. Application-layer anycasting: a server selection architecture and use in a replicated Web service. IEEE/ACM Transactions on Networking, 8 (4):455–466, 2000. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/2000-8-4/
p455-zegura/.

Zalesky:2009:BCS

[Zal09] Andrew Zalesky. To burst or circuit switch? IEEE/ ACM Transactions on Networking, 17(1):305–318, February 2009. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Zappala:2004:APR

[Zap04] Daniel Zappala. Alternate path routing for multicast. IEEE/ACM Transactions on Networking, 12 (1):30–43, February 2004. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Zafer:2012:LGS

[ZAS12] Murtaza Zafer, Dakshi Agrawal, and Mudhakar Srivatsa. Limitations of generating a secret key using wireless fading under active adversary. IEEE/ ACM Transactions on Networking, 20(5):1440–1451, October 2012. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Zhang:2022:PSR

[ZAW^{+}22] Qiaolun Zhang, Omran Ayoub, Jun Wu, Francesco Musumeci, Gaolei Li, and Massimo Tornatore. Progressive slice recovery with guaranteed slice connectivity after massive failures.

IEEE/ACM Transactions on Networking, 30(2):826– 839, April 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2021.3130576.

Zhang:1995:OSB

[ZB95] Sijing Zhang and Alan Burns. An optimal synchronous bandwidth allocation scheme for guaranteeing synchronous message deadlines with the timed-token MAC protocol. IEEE/ACM Transactions on Networking, 3(6): 729–741, December 1995. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL

http://www.acm.org/pubs/
citations/journals/ton/
1995-3-6/p729-zhang/.

Zhang:2016:EEI

[ZBA16] Zhi Zhang, Yigal Bejerano, and Spyridon Antonakopoulos. Energy-efficient IP core network configuration under general traffic demands. IEEE/ACM Transactions on Networking, 24(2):745–758, April 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Zeng:2022:CCC

[ZBC^{+}22] Gaoxiong Zeng, Wei Bai, Ge Chen, Kai Chen, Dongsu Han, Yibo Zhu, and Lei Cui. Congestion control for cross-datacenter networks. IEEE/ACM Transactions on Networking, 30(5):2074–2089, 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL

https://dl.acm.org/doi/
10.1109/TNET.2022.3161580.

Zheng:2023:CNS

[ZBdV23] Jiaxiao Zheng, Albert Banchs, and Gustavo de Veciana. Constrained network slicing games: Achieving service guarantees and network efficiency. IEEE/ACM Transactions on Networking, 31(6):2698– 2713, 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2023.3262810.

Zhang:2013:EOS

[ZBXH13] Zhenghao Zhang, Steven Bronson, Jin Xie, and Wei Hu. Employing the one-sender-multiple-receiver technique in wireless LANs. IEEE/ACM Transactions on Networking, 21(4):1243–1255, August 2013. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Zhou:2019:PBF

[ZBZ^{+}19] Yu Zhou, Jun Bi, Cheng Zhang, Bingyang Liu, Zhaogeng Li, Yangyang Wang, and Mingli Yu. P4DB: On-the-fly debugging for programmable data planes. IEEE/ ACM Transactions on Networking, 27(4):1714–1727, August 2019. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Zhao:2009:TUE

[ZCB09] Yao Zhao, Yan Chen, and David Bindel. Towards unbiased end-to-end network diagnosis. IEEE/ ACM Transactions on Networking, 17(6):1724–1737, December 2009. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Zhang:2017:GDI

[ZCB^{+}17] Hong Zhang, Kai Chen, Wei Bai, Dongsu Han, Chen Tian, Hao Wang, Haibing Guan, and Ming Zhang. Guaranteeing deadlines for inter-data center transfers. IEEE/ ACM Transactions on Networking, 25(1):579–595, February 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Zegura:1997:QCG

[ZCD97] Ellen W. Zegura, Kenneth L. Calvert, and Michael J. Donahoo. A quantitative comparison of graph-based models for Internet topology. IEEE/ ACM Transactions on Networking, 5(6):770–783, December 1997. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1997-5-6/
p770-zegura/.

Zheng:2018:SMT

[ZCdV^{+}18] Jiaxiao Zheng, Pablo Caballero, Gustavo de Veciana, Seung Jun Baek, and Albert Banchs. Statistical multiplexing and traffic shaping games for network slicing. IEEE/ACM Transactions on Networking, 26(6):2528–2541, December 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Zeng:2024:AAA

[ZCF^{+}24] Liekang Zeng, Haowei Chen, Daipeng Feng, Xiaoxi Zhang, and Xu Chen. A3D: Adaptive, accurate, and autonomous navigation for edge-assisted drones. IEEE/ACM Transactions on Networking, 32 (1):713–728, 2024. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.

                                        acm.org/doi/10.1109/TNET.
                                        2023.3297876.

Zeng:2024:SGN

[ZCH^{+}24] Liekang Zeng, Xu Chen, Peng Huang, Ke Luo, Xiaoxi Zhang, and Zhi Zhou. Serving graph neural networks with distributed fog servers for smart IoT services. IEEE/ACM Transactions on Networking, 32 (1):550–565, 2024. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2023.3293052.

Zhang:2013:MLV

[ZCJ^{+}13] Liang Zhang, Shigang Chen, Ying Jian, Yuguang Fang, and Zhen Mo. Maximizing lifetime vector in wireless sensor networks. IEEE/ACM Transactions on Networking, 21(4):1187– 1200, August 2013. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Zhou:2011:SMC

[ZCL11] Yipeng Zhou, Dah-Ming Chiu, and John C. S. Lui. A simple model for chunk-scheduling strategies in P2P streaming. IEEE/ACM Transactions on Networking, 19(1):42– 54, February 2011. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Zohar:2014:PPB

[ZCM14] Eyal Zohar, Israel Cidon, and Osnat Mokryn. PACK: Prediction-based cloud bandwidth and cost reduction system. IEEE/ACM Transactions on Networking, 22(1):39–51, February 2014. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Zeng:2023:AUD

[ZCPG^{+}23] Yijing Zeng, Roberto Calvo-Palomino, Domenico Giustiniano, Gerome Bovet, and Suman Banerjee. Adaptive uplink data compression in spectrum crowdsensing systems. IEEE/ ACM Transactions on Networking, 31(5):2207– 2221, 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2023.3239378.

Zhang:2022:AAA

[ZCPP22] Chaoyun Zhang, Xavier Costa-Pérez, and Paul Patras. Adversarial attacks against deep learningbased network intrusion detection systems and defense mechanisms. IEEE/ ACM Transactions on Networking, 30(3):1294–1311, June 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.

                                        acm.org/doi/10.1109/TNET.
                                        2021.3137084.

Zhang:2015:CCN

[ZCW15] Fan Zhang, Yewen Cao, and Deqiang Wang. Comments and corrections: a note on “Low-complexity distributed scheduling algorithms for wireless networks”. IEEE/ACM Transactions on Networking, 23(4):1367–1369, August 2015. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). See [GLS09].

Zhang:2022:DBG

[ZCW^{+}22] Lei Zhang, Yong Cui, Mowei Wang, Kewei Zhu, Yibo Zhu, and Yong Jiang. DeepCC: Bridging the gap between congestion control and applications via multiobjective optimization. IEEE/ACM Transactions on Networking, 30(5): 2274–2288, 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2022.3167713.

Zhang:2015:RNT

[ZCX^{+}15] Jun Zhang, Xiao Chen, Yang Xiang, Wanlei Zhou, and Jie Wu. Robust network traffic classification. IEEE/ACM Transactions on Networking, 23(4): 1257–1270, August 2015. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Zhou:2016:URC

[ZCY16] Ziling Zhou, Binbin Chen, and Haifeng Yu. Understanding RFID counting protocols. IEEE/ ACM Transactions on Networking, 24(1):312–327, February 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Zhang:2020:EAT

[ZCZ^{+}20] Youlin Zhang, Shigang Chen, You Zhou, Olufemi O. Odegbile, and Yuguang Fang. Efficient anonymous temporal-spatial joint estimation at category level over multiple tag sets with unreliable channels. IEEE/ ACM Transactions on Networking, 28(5):2174–2187, October 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2020.3011347.

Zeng:2021:CCD

[ZCZ^{+}21] Liekang Zeng, Xu Chen, Zhi Zhou, Lei Yang, and Junshan Zhang. CoEdge: Cooperative DNN inference with adaptive workload partitioning over heterogeneous edge devices. IEEE/ACM Transactions on Networking, 29(2):595– 608, April 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2020.3042320.

Zhang:2017:TCD

[ZCZC17] Xiaomei Zhang, Guohong Cao, Xiaomei Zhang, and Guohong Cao. Transient community detection and its application to data forwarding in delay tolerant networks. IEEE/ ACM Transactions on Networking, 25(5):2829–2843, October 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Zhang:2020:MTD

[ZCZF20] Youlin Zhang, Shigang Chen, You Zhou, and Yuguang Fang. Missingtag detection with unknown tags. IEEE/ ACM Transactions on Networking, 28(3):1297–1310, June 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2020.2984706.

Zhao:2017:AGS

[ZDB^{+}17] Zhiwei Zhao, Wei Dong, Jiajun Bu, Tao Gu, and Geyong Min. Accurate and generic sender selection for bulk data dissemination in low-power wireless networks. IEEE/ACM Transactions on Networking, 25 (2):948–959, April 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Zhang:2018:JBG

[ZDCW18] Ying Zhang, Lei Deng, Minghua Chen, and Peijian Wang. Joint bidding and geographical load balancing for datacenters: Is uncertainty a blessing or a curse? IEEE/ACM Transactions on Networking, 26 (3):1049–1062, June 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Zhao:2004:MEP

[ZDR04] Zhili Zhao, Swaroop Darbha, and A. L. Narasimha Reddy. A method for estimating the proportion of nonresponsive traffic at a router. IEEE/ACM Transactions on Networking, 12 (4):708–718, August 2004. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Zheng:2024:LCC

[ZDZ^{+}24] Jiaqi Zheng, Zhuoxuan Du, Zhenqing Zha, Zixuan Yang, Xiaofeng Gao, and Guihai Chen. Learning to configure converters in hybrid switching data center networks.

IEEE/ACM Transactions on Networking, 32(1):520– 534, 2024. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2023.3294803.

Zegura:1995:EBP

[Zeg95] Ellen Witte Zegura. Evaluating blocking probability in generalized connectors. IEEE/ACM Transactions on Networking, 3(4):387– 398, August 1995. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/
                                        journals/ton/1995-3-4/
                                        p387-zegura/.

Zegura:2003:Eb

[Zeg03a] Ellen Zegura. Editorial. IEEE/ACM Transactions on Networking, 11(6):869, December 2003. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Zegura:2003:Ea

[Zeg03b] Ellen W. Zegura. Editorial. IEEE/ACM Transactions on Networking, 11(4): 513, August 2003. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Zegura:2004:E

[Zeg04] Ellen Zegura. Editorial. IEEE/ACM Transactions on Networking, 12(1):1, February 2004. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Zegura:2005:Ea

[Zeg05a] Ellen W. Zegura. Editorial. IEEE/ACM Transactions on Networking, 13(1): 1, February 2005. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Zegura:2005:Eb

[Zeg05b] Ellen W. Zegura. Editorial. IEEE/ACM Transactions on Networking, 13(3): 461, June 2005. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Zhao:2007:NBR

[ZEV07a] Yanping Zhao, Derek L.

Eager, and Mary K. Vernon. Network bandwidth requirements for scalable on-demand streaming. IEEE/ACM Transactions on Networking, 15(4):878– 891, August 2007. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Zhao:2007:SDS

[ZEV07b] Yanping Zhao, Derek L.

Eager, and Mary K. Vernon. Scalable on-demand streaming of nonlinear media. IEEE/ACM Transactions on Networking, 15(5): 1149–1162, October 2007. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Zho:1996:IMC

[ZF96] Hongbo Zho and Victor S. Frost. In-service monitoring for cell loss quality of service violations in ATM networks. IEEE/ACM Transactions on Networking, 4(2):240– 248, April 1996. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1996-4-2/
p240-zho/.

Zhou:2013:RAP

[ZFC13] Yipeng Zhou, Tom Z. J. Fu, and Dah Ming Chiu. On replication algorithm in P2P VoD. IEEE/ ACM Transactions on Networking, 21(1):233–243, February 2013. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Zhou:2015:UMA

[ZFC15] Yipeng Zhou, Tom Z. J. Fu, and Dah Ming Chiu. A unifying model and analysis of P2P VoD replication and scheduling. IEEE/ ACM Transactions on Networking, 23(4):1163–1175, August 2015. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Zhu:2018:TOA

[ZFLC18] Yifei Zhu, Silvery D. Fu, Jiangchuan Liu, and Yong Cui. Truthful online auction toward maximized instance utilization in the cloud. IEEE/ACM Transactions on Networking, 26(5):2132–2145, October 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Zhang:2014:AAS

[ZFW14] Jinbei Zhang, Luoyi Fu, and Xinbing Wang. Asymptotic analysis on secrecy capacity in large-scale wireless networks. IEEE/ACM Transactions on Networking, 22(1):66–79, February 2014. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Zhang:2017:CAW

[ZFW^{+}17a] Jinbei Zhang, Luoyi Fu, Qi Wang, Liang Liu, Xinyu Wang, and Xinbing Wang. Connectivity analysis in wireless networks with correlated mobility and cluster scalability. IEEE/ ACM Transactions on Networking, 25(4):2375–2390, August 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Zhu:2017:EAC

[ZFW^{+}17b] Tingwei Zhu, Dan Feng, Fang Wang, Yu Hua, Qingyu Shi, Jiahao Liu, Yongli Cheng, and Yong Wan. Efficient anonymous communication in SDNbased data center networks. IEEE/ACM Transactions on Networking, 25(6):3767– 3780, December 2017. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Zhao:2008:LMC

[ZG08] Qun Zhao and Mohan Gurusamy. Lifetime maximization for connected target coverage in wireless sensor networks. IEEE/ACM Transactions on Networking, 16(6):1378–1391, December 2008. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Zhang:2014:VFD

[ZG14] Lei Zhang and Dongning Guo. Virtual full duplex wireless broadcasting via compressed sensing. IEEE/ ACM Transactions on Networking, 22(5):1659–1671, October 2014. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Zhang:2005:ILL

[ZGG05] Hui Zhang, Ashish Goel, and Ramesh Govindan. Improving lookup latency in distributed hash table systems using random sampling. IEEE/ACM Transactions on Networking, 13(5):1121–1134, October 2005. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Zhang:2019:HPM

[ZGHH19] Meng Zhang, Lin Gao, Jianwei Huang, and Michael L. Honig. Hybrid pricing for mobile collaborative Internet access. IEEE/ACM Transactions on Networking, 27(3):986–999, June 2019. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Zhang:2019:DFD

[ZGL^{+}19] Han Zhang, Haijun Geng, Yahui Li, Xia Yin, Xingang Shi, Zhiliang Wang, Qianhong Wu, and Jianwei Liu. DA&FD-deadline-aware and flow durationbased rate control for mixed flows in DCNs. IEEE/ACM Transactions on Networking, 27(6):2458– 2471, December 2019. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL https:

//dl.acm.org/doi/abs/
10.1109/TNET.2019.2951925.

Zhong:2020:JOR

[ZGLC20] Xijian Zhong, Yan Guo, Ning Li, and Yancheng Chen. Joint optimization of relay deployment, channel allocation, and relay assignment for UAVs-aided D2D networks. IEEE/ ACM Transactions on Networking, 28(2):804–817, April 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL https://

dl.acm.org/doi/abs/10.
1109/TNET.2020.2970744.

Zheng:2010:PSR

[ZGS10] Si Qing Zheng, Ashwin Gumaste, and Hong Shen. A parallel selfrouting rearrangeable nonblocking multi-log_{2} N photonic switching network. IEEE/ACM Transactions on Networking, 18(2):529– 539, April 2010. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Zhang:2024:PPL

[ZGS^{+}24] Xiaoli Zhang, Wei Geng, Yiqiao Song, Hongbing Cheng, Ke Xu, and Qi Li. Privacy-preserving and lightweight verification of deep packet inspection in clouds. IEEE/ACM Transactions on Networking, 32 (1):159–174, 2024. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2023.3282100.

Zou:2005:MED

[ZGTG05] Cliff C. Zou, Weibo Gong, Don Towsley, and Lixin Gao. The monitoring and early detection of Internet worms. IEEE/ACM Transactions on Networking, 13 (5):961–974, October 2005. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Zhu:2016:BAC

[ZGY^{+}16] Tiantian Zhu, Hongyu Gao, Yi Yang, Kai Bu, Yan Chen, Doug Downey, Kathy Lee, and Alok N. Choudhary. Beating the artificial chaos: Fighting OSN spam using its own templates. IEEE/ACM Transactions on Networking, 24(6):3856–3869, December 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Zhang:2020:PTF

[ZGYB20] Jingxuan Zhang, Kai Gao, Y. Richard Yang, and Jun Bi. Prophet: Toward fast, error-tolerant modelbased throughput prediction for reactive flows in DC networks. IEEE/ACM Transactions on Networking, 28(6):2475–2488, December 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.

acm.org/doi/10.1109/TNET.
2020.3016838.

Zhang:2022:MCR

[ZGZ22] Yu Zhang, Tao Gu, and Xi Zhang. MDLdroid: a ChainSGD-reduce approach to mobile deep learning for personal mobile sensing. IEEE/ ACM Transactions on Networking, 30(1):134–147, February 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2021.3103846.

Zhou:2020:SRL

[ZGZC20] Hao Zhou, Xiaofeng Gao, Jiaqi Zheng, and Guihai Chen. Scheduling relaxed loop-free updates within tight lower bounds in SDNs. IEEE/ACM Transactions on Networking, 28(6):2503–2516, December 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2020.3017771.

Zhang:2008:ACT

[ZH08a] Honghai Zhang and Jennifer C. Hou. Asymptotic critical total power for kconnectivity of wireless networks. IEEE/ACM Transactions on Networking, 16 (2):347–358, April 2008. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Zhang:2008:AMT

[ZH08b] Honghai Zhang and Jennifer C. Hou. On the asymptotic minimum transporting energy and its implication on the wireless network capacity. IEEE/ ACM Transactions on Networking, 16(5):1175–1187, October 2008. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Zhao:2017:TPS

[Zha17] Jun Zhao. Topological properties of secure wireless sensor networks under the q-composite key predistribution scheme with unreliable links. IEEE/ACM Transactions on Networking, 25(3):1789–1802, June 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Zhang:2016:DGO

[ZHC16] Yongmin Zhang, Shibo He, and Jiming Chen. Data gathering optimization by dynamic sensing and routing in rechargeable sensor networks. IEEE/ACM Transactions on Networking, 24(3):1632–1646, June 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Zhang:2024:ESH

[ZHCC24] Liping Zhang, Wenshuo Han, Shukai Chen, and Kim-Kwang Raymond Choo. An efficient and secure health data propagation scheme using steganographybased approach for electronic health networks. IEEE/ACM Transactions on Networking, 32(2):1261– 1272, April 2024. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2023.3313160.

Zhang:2017:PSD

[ZHCL17] Shaoquan Zhang, Longbo Huang, Minghua Chen, and Xin Liu. Proactive serving decreases user delay exponentially: The lighttailed service time case. IEEE/ACM Transactions on Networking, 25(2):708– 723, April 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Zhang:2019:SAE

[ZHGF19] Xiaonan Zhang, Pei Huang, Linke Guo, and Yuguang Fang. Social-aware energyefficient data offloading with strong stability. IEEE/ ACM Transactions on Networking, 27(4):1515–1528, August 2019. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Zheng:2006:TBD

[ZHLL06] Kai Zheng, Chengchen Hu, Hongbin Lu, and Bin Liu. A TCAM-based distributed parallel IP lookup scheme and performance analysis. IEEE/ACM Transactions on Networking, 14(4):863– 875, August 2006. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Zhong:2019:PST

[ZHT^{+}19] Zhizhen Zhong, Nan Hua, Massimo Tornatore, Jialong Li, Yanhe Li, Xiaoping Zheng, and Biswanath Mukherjee. Provisioning short-term traffic fluctuations in elastic optical networks. IEEE/ACM Transactions on Networking, 27(4):1460–1473, August 2019. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Zhang:2017:OAI

[ZHW^{+}17] Xiaoxi Zhang, Zhiyi Huang, Chuan Wu, Zongpeng Li, and Francis C. M. Lau. Online auctions in IaaS clouds: Welfare and profit maximization with server costs. IEEE/ACM Transactions on Networking, 25(2):1034– 1047, April 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Zou:2021:FAA

[ZHWH21] Shaojun Zou, Jiawei Huang, Jianxin Wang, and Tian He. Flow-aware adaptive pacing to mitigate TCP incast in data center networks. IEEE/ ACM Transactions on Networking, 29(1):134–147, February 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2020.3027749.

Zhang:2018:USH

[ZHZ^{+}18] Desheng Zhang, Tian He, Fan Zhang, Chengzhong Xu, Tian He, Fan Zhang, Desheng Zhang, and Chengzhong Xu. Urban-scale human mobility modeling with multi-source urban network data. IEEE/ACM Transactions on Networking, 26 (2):671–684, April 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Zhang:2024:PTE

[ZHZ^{+}24] Yuntian Zhang, Ning Han, Tengteng Zhu, Junjie Zhang, Minghao Ye, Songshi Dou, and Zehua Guo. Prophet: Traffic engineeringcentric traffic matrix prediction. IEEE/ACM Transactions on Networking, 32(1):822–832, 2024. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL

https://dl.acm.org/doi/
10.1109/TNET.2023.3293098.

Zhu:2012:RCC

[ZJ12] Yi Zhu and Jason P. Jue.

Reliable collective communications with weighted SRLGs in optical networks. IEEE/ACM Transactions on Networking, 20(3):851– 863, June 2012. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Zhao:2018:PLP

[ZJL^{+}18] Ping Zhao, Hongbo Jiang, John C. S. Lui, Chen Wang, Fanzi Zeng, Fu Xiao, and Zhetao Li. P3-LOC: a privacy-preserving paradigm-driven framework for indoor localization. IEEE/ACM Transactions on Networking, 26(6): 2856–2869, December 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Zhao:2019:SPP

[ZJL^{+}19] Ping Zhao, Hongbo Jiang, Jie Li, Fanzi Zeng, Xiao Zhu, Kun Xie, and Guanglin Zhang. Synthesizing privacy preserving traces: Enhancing plausibility with social networks. IEEE/ ACM Transactions on Networking, 27(6):2391–2404, December 2019. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL https://

                                        dl.acm.org/doi/abs/10.
                                        1109/TNET.2019.2947452.

Zhang:2012:WMS

[ZJS^{+}12] Honghai Zhang, Yuanxi Jiang, Karthik Sundaresan, Sampath Rangarajan, and Baohua Zhao. Wireless multicast scheduling with switched beamforming antennas. IEEE/ ACM Transactions on Networking, 20(5):1595–1607, October 2012. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Zhou:2017:ECA

[ZJWY17] Hao Zhou, Yusheng Ji, Xiaoyan Wang, and Shigeki Yamada. eICIC configuration algorithm with service scalability in heterogeneous cellular networks. IEEE/ACM Transactions on Networking, 25(1):520– 535, February 2017. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Zhang:2019:ICS

[ZK19] Xu Zhang and Edward W. Knightly. CSIsnoop: Inferring channel state information in multi-user MIMO WLANs. IEEE/ ACM Transactions on Networking, 27(1):231–244, February 2019. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Zhou:2023:EDM

[ZKEN23] Xujin Zhou, Irem Koprulu, Atilla Eryilmaz, and Michael J. Neely. Efficient distributed MAC for dynamic demands: Congestion and age based designs. IEEE/ACM Transactions on Networking, 31 (1):74–87, 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2022.3191607.

Zafer:2010:TPE

[ZKH10] Murtaza Zafer, Bong Jun Ko, and Ivan Wang-Hei Ho. Transmit power estimation using spatially diverse measurements under wireless fading. IEEE/ ACM Transactions on Networking, 18(4):1171–1180, August 2010. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Zhang:2007:DIS

[ZKL07] Yueping Zhang, Seong-Ryong Kang, and Dmitri Loguinov. Delay-independent stability and performance of distributed congestion control. IEEE/ACM Transactions on Networking, 15(4):838–851, August 2007. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Zhang:2011:OCA

[ZKL11] Xiaolan Joy Zhang, Sun-Il Kim, and Steven S. Lumetta. Opportunity cost analysis for dynamic wavelength routed mesh networks. IEEE/ACM Transactions on Networking, 19 (3):747–759, June 2011. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Zhong:1993:CNS

[ZKO93] Wen De Zhong, Jaidev Kaniyil, and Y. Onozato. A copy network with shared buffers for large-scale multicast ATM switching. IEEE/ACM Transactions on Networking, 1(2):157– 165, April 1993. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/
                                        journals/ton/1993-1-2/
                                        p157-de_zhong/.

Zeng:2014:ATP

[ZKVM14] Hongyi Zeng, Peyman Kazemian, George Varghese, and Nick McKeown. Automatic test packet generation. IEEE/ACM Transactions on Networking, 22(2):554–566, April 2014. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Zarikoff:2013:MPI

[ZL13a] Brad W. Zarikoff and Douglas J. Leith. Measuring pulsed interference in 802.11 links. IEEE/ACM Transactions on Networking, 21(2):509–521, April 2013. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Zheng:2013:FTS

[ZL13b] Yuanqing Zheng and Mo Li.

Fast tag searching protocol for large-scale RFID systems. IEEE/ACM Transactions on Networking, 21 (3):924–934, June 2013. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Zheng:2014:TME

[ZL14] Yuanqing Zheng and Mo Li.

Towards more efficient cardinality estimation for large-scale RFID systems. IEEE/ACM Transactions on Networking, 22(6):1886– 1896, December 2014. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Zheng:2015:PMP

[ZL15] Yuanqing Zheng and Mo Li.

P-MTI: physical-layer missing tag identification via compressive sensing. IEEE/ ACM Transactions on Networking, 23(4):1356–1366, August 2015. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Zhao:2016:DSA

[ZL16] Shizhen Zhao and Xiaojun Lin. Design of scheduling algorithms for end-to-end backlog minimization in wireless multi-hop networks under k-hop interference models. IEEE/ACM Transactions on Networking, 24(2):1265–1278, April 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Zhang:2017:SOM

[ZLB17] Mingwei Zhang, Jun Li, and Scott Brooks. I-Seismograph: Observing, measuring, and analyzing Internet earthquakes. IEEE/ACM Transactions on Networking, 25(6):3411– 3426, December 2017. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Zhao:2012:MFA

[ZLC12] Bridge Qiao Zhao, John C. S. Lui, and Dah-Ming Chiu. A mathematical framework for analyzing adaptive incentive protocols in P2P networks. IEEE/ACM Transactions on Networking, 20(2):367– 380, April 2012. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Zhang:2017:CEN

[ZLG^{+}17] Yiming Zhang, Dongsheng Li, Chuanxiong Guo, Haitao Wu, Yongqiang Xiong, and Xicheng Lu. CubicRing: Exploiting network proximity for distributed in-memory keyvalue store. IEEE/ACM Transactions on Networking, 25(4):2040–2053, August 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Zhang:2020:PER

[ZLG^{+}20] Sheng Zhang, Yu Liang, Jidong Ge, Mingjun Xiao, and Jie Wu. Provably efficient resource allocation for edge service entities using Hermes. IEEE/ ACM Transactions on Networking, 28(4):1684–1697, August 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2020.2989307.

Zhou:2022:QNS

[ZLHM22] Hongyi Zhou, Kefan Lv, Longbo Huang, and Xiongfeng Ma. Quantum network: Security assessment and key management. IEEE/ ACM Transactions on Networking, 30(3):1328–1339, June 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2021.3136943.

Zhang:2023:MBR

[ZLL^{+}23a] Zhehui Zhang, Yuanjie Li, Qianru Li, Jinghao Zhao, Ghufran Baig, Lili Qiu, and Songwu Lu. Movement-based reliable mobility management for beyond 5G cellular networks. IEEE/ACM Transactions on Networking, 31 (1):192–207, 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2022.3190788.

Zong:2023:CEC

[ZLL^{+}23b] Tongyu Zong, Chen Li, Yuanyuan Lei, Guangyu Li, Houwei Cao, and Yong Liu. Cocktail edge caching: Ride dynamic trends of content popularity with ensemble learning. IEEE/ACM Transactions on Networking, 31(1): 208–219, 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2022.3193680.

Zhang:2024:AAC

[ZLL24] Weiguang Zhang, Jiarong Liang, and Xinyu Liang. Approximation algorithms for computing virtual backbones considering routing costs in wireless networks. IEEE/ACM Transactions on Networking, 32 (1):323–337, 2024. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2023.3284051.

Zhang:2003:PDS

[ZLLY03] X. Brian Zhang, Simon S.

Lam, Dong-Young Lee, and Y. Richard Yang. Protocol design for scalable and reliable group rekeying. IEEE/ACM Transactions on Networking, 11(6): 908–922, December 2003. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Zhao:2016:BFO

[ZLM16] Dong Zhao, Xiang-Yang Li, and Huadong Ma. Budgetfeasible online incentive mechanisms for crowdsourcing tasks truthfully. IEEE/ACM Transactions on Networking, 24(2):647– 661, April 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Zou:2023:BBB

[ZLM^{+}23] Renpeng Zou, Xixiang Lyu, Jing Ma, Bowen Zhang, and Danfang Wu. BCMIX: a blockchain-based dynamic self-reconfigurable Mixnet. IEEE/ACM Transactions on Networking, 31(5):2222–2235, 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL

https://dl.acm.org/doi/
10.1109/TNET.2023.3244962.

Zhu:2017:NLE

[ZLN^{+}17] Rui Zhu, Bang Liu, Di Niu, Zongpeng Li, and Hong Vicky Zhao. Network latency estimation for personal devices: a matrix completion approach. IEEE/ACM Transactions on Networking, 25(2):724–737, April 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Zhang:2020:ECR

[ZLRC20] Yongmin Zhang, Xiaolong Lan, Ju Ren, and Lin Cai. Efficient computing resource sharing for mobile edge-cloud computing networks. IEEE/ ACM Transactions on Networking, 28(3):1227–1240, June 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2020.2979807.

Zarros:1996:ISR

[ZLS96] Panagiotis N. Zarros, Myung J.

Lee, and Tarek N. Saadawi. Interparticipant synchronization in real-time multimedia conferencing using feedback. IEEE/ ACM Transactions on Networking, 4(2):173–180, April 1996. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1996-4-2/
p173-zarros/.

Zheng:2015:EPC

[ZLSK15] Zizhan Zheng, Zhixue Lu, Prasun Sinha, and Santosh Kumar. Ensuring predictable contact opportunity for scalable vehicular Internet access on the go. IEEE/ACM Transactions on Networking, 23(3): 768–781, June 2015. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Zhou:2017:NSL

[ZLTX17] Zhili Zhou, Tachun Lin, Krishnaiyan Thulasiraman, and Guoliang Xue. Novel survivable logical topology routing by logical protecting spanning trees in IP-over-WDM networks. IEEE/ACM Transactions on Networking, 25(3):1673– 1685, June 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Zhao:2016:SCS

[ZLW16a] Can Zhao, Xiaojun Lin, and Chuan Wu. The streaming capacity of sparsely connected P2P systems with distributed control. IEEE/ACM Transactions on Networking, 24(1):58– 71, February 2016. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Zhu:2016:CCC

[ZLW^{+}16b] Ming Zhu, Dan Li, Fangxin Wang, Anke Li, K. K. Ramakrishnan, Ying Liu, Jianping Wu, Nan Zhu, and Xue Liu. CCDN: Content-centric data center networks. IEEE/ACM Transactions on Networking, 24(6):3537–3550, December 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Zhu:2017:ISA

[ZLW^{+}17] Zuqing Zhu, Xiahe Liu, Yixiang Wang, Wei Lu, Long Gong, Shui Yu, and Nirwan Ansari. Impairmentand splitting-aware cloudready multicast provisioning in elastic optical networks. IEEE/ACM Transactions on Networking, 25 (2):1220–1234, April 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Zhou:2018:SFC

[ZLW18] Ruiting Zhou, Zongpeng Li, and Chuan Wu. Scheduling frameworks for cloud container services. IEEE/ ACM Transactions on Networking, 26(1):436–450, February 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Zhou:2019:TSS

[ZLW^{+}19] Jianer Zhou, Zhenyu Li, Qinghua Wu, Peter Steenkiste, Steve Uhlig, Jun Li, and Gaogang Xie. TCP stalls at the server side: Measurement and mitigation. IEEE/ACM Transactions on Networking, 27(1):272– 287, February 2019. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Zhao:2020:MAE

[ZLW^{+}20] Jia Zhao, Jiangchuan Liu, Haiyang Wang, Chi Xu, Wei Gong, and Changqiao Xu. Measurement, analysis, and enhancement of multipath TCP energy efficiency for datacenters. IEEE/ACM Transactions on Networking, 28(1):57–70, February 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL https://

                                        dl.acm.org/doi/abs/10.
                                        1109/TNET.2019.2950908.

Zhou:2017:ECM

[ZLWH17] Ruiting Zhou, Zongpeng Li, Chuan Wu, and Zhiyi Huang. An efficient cloud market mechanism for computing jobs with soft deadlines. IEEE/ACM Transactions on Networking, 25 (2):793–805, April 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Zhao:2018:SCS

[ZLWM18] Jie Zhao, Qiang Liu, Xin Wang, and Shiwen Mao. Scheduling of collaborative sequential compressed sensing over wide spectrum band. IEEE/ACM Transactions on Networking, 26(1):492–505, February 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Zhang:2021:CPR

[ZLX^{+}21] Menghao Zhang, Guanyu Li, Lei Xu, Jiasong Bai, Mingwei Xu, Guofei Gu, and Jianping Wu. Control plane reflection attacks and defenses in software-defined networks. IEEE/ACM Transactions on Networking, 29(2):623– 636, April 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2020.3040773.

Zhao:2023:SMD

[ZLX^{+}23] Gongming Zhao, Luyao Luo, Hongli Xu, Chun-Jen Chung, and Liguang Xie. Southbound message delivery with virtual network topology awareness in clouds. IEEE/ACM Transactions on Networking, 31 (1):248–263, 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2022.3190730.

Zhang:2023:SHP

[ZLY23] Jielun Zhang, Fuhao Li, and Feng Ye. Sustaining the high performance of AI-based network traffic classification models. IEEE/ACM Transactions on Networking, 31(2):816– 827, April 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2022.3203227.

Zhang:2021:OAI

[ZLZ21a] Qixia Zhang, Fangming Liu, and Chaobing Zeng. Online adaptive interferenceaware VNF deployment and migration for 5G network slice. IEEE/ACM Transactions on Networking, 29(5):2115–2128, October 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2021.3080197.

Zhang:2021:VGA

[ZLZ^{+}21b] Xiaoli Zhang, Qi Li, Zeyu Zhang, Jianping Wu, and Jiahai Yang. vSFC: Generic and agile verification of service function chains in the cloud. IEEE/ ACM Transactions on Networking, 29(1):78–91, February 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2020.3028846.

Zhang:2023:EFL

[ZLZ^{+}23] Tinghao Zhang, Kwok-Yan Lam, Jun Zhao, Feng Li, Huimei Han, and Norziana Jamil. Enhancing federated learning with spectrum allocation optimization and device selection. IEEE/ACM Transactions on Networking, 31(5):1981– 1996, 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2022.3231986.

Zheng:2016:RBD

[ZLZL16] Yuanqing Zheng, Mo Li, Yuanqing Zheng, and Mo Li. Read bulk data from computational RFIDs. IEEE/ACM Transactions on Networking, 24(5):3098–3108, October 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Znati:2004:NDA

[ZM04] Taieb F. Znati and Rami Melhem. Node delay assignment strategies to support end-to-end delay requirements in heterogeneous networks. IEEE/ ACM Transactions on Networking, 12(5):879–892, October 2004. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Zafer:2009:CAE

[ZM09] Murtaza A. Zafer and Eytan Modiano. A calculus approach to energyefficient data transmission with quality-of-service constraints. IEEE/ACM Transactions on Networking, 17(3):898–911, June 2009. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Zhang:2018:CIN

[ZM18] Jianan Zhang and Eytan Modiano. Connectivity in interdependent networks. IEEE/ACM Transactions on Networking, 26(5):2090– 2103, October 2018. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Zhao:2020:ELD

[ZMD^{+}20] Zhiwei Zhao, Geyong Min, Wei Dong, Xue Liu, Weifeng Gao, Tao Gu, and Minghang Yang. Exploiting link diversity for performance-aware and repeatable simulation in low-power wireless networks. IEEE/ACM Transactions on Networking, 28(6):2545–2558, December 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2020.3016056.

Zhang:2017:ENR

[ZMH17] Jianan Zhang, Eytan Modiano, and David Hay. Enhancing network robustness via shielding. IEEE/ ACM Transactions on Networking, 25(4):2209–2222, August 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Zhang:2019:HAI

[ZML^{+}19] Ziyao Zhang, Liang Ma, Kin K. Leung, Franck Le, Sastry Kompella, and Leandros Tassiulas. How advantageous is it? An analytical study of controllerassisted path construction in distributed SDN. IEEE/ ACM Transactions on Networking, 27(4):1643–1656, August 2019. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Zhang:2021:MAB

[ZMLL21] Ziyao Zhang, Liang Ma, Kin K. Leung, and Franck Le. More is not always better: an analytical study of controller synchronizations in distributed SDN. IEEE/ ACM Transactions on Networking, 29(4):1580–1590, August 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2021.3066580.

Zhang:2023:RCD

[ZMLR23] Yiran Zhang, Qingkai Meng, Yifan Liu, and Fengyuan Ren. Revisiting congestion detection in lossless networks. IEEE/ACM Transactions on Networking, 31(5):2361– 2375, 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2023.3250484.

Zhao:2022:JMS

[ZMMG22] Zhiwei Zhao, Wenliang Mao, Geyong Min, and Weifeng Gao. Joint multichannel-spatial diversity for efficient opportunistic routing in lowpower wireless networks. IEEE/ACM Transactions on Networking, 30(6):2716– 2729, 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2022.3181581.

Zeng:2022:LRD

[ZMW^{+}22] Guangyang Zeng, Biqiang Mu, Jieqiang Wei, Wing Shing Wong, and Junfeng Wu. Localizability with rangedifference measurements: Numerical computation and error bound analysis. IEEE/ACM Transactions on Networking, 30(5): 2117–2130, 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2022.3162930.

Zou:2018:OSD

[ZMWX18] Mao Zou, Richard T. B. Ma, Xin Wang, and Yinlong Xu. On optimal service differentiation in congested network markets. IEEE/ACM Transactions on Networking, 26(6):2693– 2706, December 2018. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Zhang:2016:CTL

[ZND^{+}16] Huiyuan Zhang, Dung T. Nguyen, Soham Das, Huiling Zhang, and My T. Thai. Corrections to “Least Cost Influence Maximization Across Multiple Social Networks” [Apr 16 929– 939]. IEEE/ACM Transactions on Networking, 24(5): 3232, October 2016. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Zhang:2013:BNC

[ZNK^{+}13] Xiaolan Zhang, Giovanni Neglia, Jim Kurose, Don Towsley, and Haixiang Wang. Benefits of network coding for unicast application in disruptiontolerant networks. IEEE/ ACM Transactions on Networking, 21(5):1407–1420, October 2013. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Zhang:2010:MBA

[ZNN^{+}10] Bo Zhang, Tze Sing Eugene Ng, Animesh Nandi, Rudolf H. Riedi, Peter Druschel, and Guohui Wang. Measurement-based analysis, modeling, and synthesis of the Internet delay space. IEEE/ACM Transactions on Networking, 18(1):229– 242, February 2010. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Zhang:2016:LCI

[ZNZT16] Huiyuan Zhang, Dung T. Nguyen, Huiling Zhang, and My T. Thai. Least cost influence maximization across multiple social networks. IEEE/ACM Transactions on Networking, 24(2):929–939, April 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Zang:2003:PPR

[ZOM03] Hui Zang, Canhui Ou, and Biswanath Mukherjee. Path-protection routing and wavelength assignment (RWA) in WDM mesh networks under duct-layer constraints. IEEE/ACM Transactions on Networking, 11(2):248–258, April 2003. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Zhang:2018:CHM

[ZP18] Yonglong Zhang and Konstantinos Psounis. Consistently high MIMO rates via switched-beam antennas. IEEE/ACM Transactions on Networking, 26(5): 2320–2333, October 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Zhou:2011:EMC

[ZPCS11] Zhong Zhou, Zheng Peng, Jun-Hong Cui, and Zhijie Shi. Efficient multipath communication for time-critical applications in underwater acoustic sensor networks. IEEE/ACM Transactions on Networking, 19(1):28–41, February 2011. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Zhang:1999:SAA

[ZQ99] Xijun Zhang and Chunming Qiao. On scheduling all-to-all personalized connections and cost-effective designs in WDM rings. IEEE/ACM Transactions on Networking, 7(3):435– 445, June 1999. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1999-7-3/
p435-zhang/.

Zhang:2000:ECA

[ZQ00] Xijun Zhang and Chunming Qiao. An effective and comprehensive approach for traffic grooming and wavelength assignment in SONET/WDM rings. IEEE/ACM Transactions on Networking, 8 (5):608–617, 2000. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/2000-8-5/
p608-zhang/.

Zhao:2023:DTP

[ZQ23] Yangming Zhao and Chunming Qiao. Distributed transport protocols for quantum data networks. IEEE/ACM Transactions on Networking, 31(6):2777– 2792, 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2023.3262547.

Zhou:2023:MLB

[ZQL^{+}23] Jianer Zhou, Xinyi Qiu, Zhenyu Li, Qing Li, Gareth Tyson, Jingpu Duan, Yi Wang, Heng Pan, and Qinghua Wu. A machine learning-based framework for dynamic selection of congestion control algorithms. IEEE/ ACM Transactions on Networking, 31(4):1566– 1581, 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2022.3220225.

Zhao:2023:FIM

[ZQW^{+}23] Yi Zhao, Meina Qiao, Haiyang Wang, Rui Zhang, Dan Wang, and Ke Xu. Friendship inference in mobile social networks: Exploiting multi-source information with two-stage deep learning framework. IEEE/ACM Transactions on Networking, 31(2):542– 557, April 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2022.3198105.

Zhao:2009:SPE

[ZR09] Suli Zhao and Dipankar Raychaudhuri. Scalability and performance evaluation of hierarchical hybrid wireless networks. IEEE/ ACM Transactions on Networking, 17(5):1536–1549, October 2009. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Zhang:2023:RMW

[ZRD^{+}23] Jia Zhang, Shaorui Ren, Enhuan Dong, Zili Meng, Yuan Yang, Mingwei Xu, Sijie Yang, Miao Zhang, and Yang Yue. Reducing mobile Web latency through adaptively selecting transport protocol. IEEE/ACM Transactions on Networking, 31(5): 2162–2177, 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2023.3235907.

Zhao:2018:NMM

[ZRH18] Tao Zhao, Korok Ray, and I-Hong Hou. A non-monetary mechanism for optimal rate control through efficient cost allocation. IEEE/ACM Transactions on Networking, 26 (3):1418–1431, June 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Zhu:2006:PMT

[ZRK06] Jing Zhu, Sumit Roy, and Jae H. Kim. Performance modelling of TCP enhancements in terrestrialsatellite hybrid networks. IEEE/ACM Transactions on Networking, 14(4):753– 766, August 2006. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Zhang:2005:EPP

[ZRLD05] Yin Zhang, Matthew Roughan, Carsten Lund, and David L. Donoho. Estimating pointto-point and point-to-multipoint traffic matrices: an information-theoretic approach. IEEE/ACM Transactions on Networking, 13(5):947–960, October 2005. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Zhu:2000:PDA

[ZRP00] Yuhong Zhu, George N. Rouskas, and Harry G. Perros. A path decomposition approach for computing blocking probabilities in wavelength-routing networks. IEEE/ACM Transactions on Networking, 8 (6):747–762, 2000. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/articles/
journals/ton/2000-8-6/
p747-zhu/p747-zhu.pdf;
 http://www.acm.org/pubs/
citations/journals/ton/
2000-8-6/p747-zhu/.

Zhang:2022:ZAH

[ZRP^{+}22] Zhenghao Zhang, Raghav Rathi, Steven Perez, Jumana Bukhari, and Yaoguang Zhong. ZCNET: Achieving high capacity in low power wide area networks. IEEE/ACM Transactions on Networking, 30(5):2032– 2045, 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2022.3158482.

Zhang:2003:DAF

[ZS03] Xi Zhang and Kang G.

Shin. Delay analysis of feedback-synchronization signaling for multicast flow control. IEEE/ACM Transactions on Networking, 11(3):436–450, June 2003. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Zhang:2004:MCM

[ZS04] Xi Zhang and Kang G.

Shin. Markov-chain modeling for multicast signaling delay analysis. IEEE/ACM Transactions on Networking, 12(4):667–680, August 2004. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Zhou:2005:AFJ

[ZS05] Yunkai Zhou and Harish Sethu. On achieving fairness in the joint allocation of processing and bandwidth resources: principles and algorithms. IEEE/ ACM Transactions on Networking, 13(5):1054–1067, October 2005. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Zhang:2013:CCS

[ZS13] Xinyu Zhang and Kang G. Shin. Cooperative carrier signaling: harmonizing coexisting WPAN and WLAN devices. IEEE/ ACM Transactions on Networking, 21(2):426–439, April 2013. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Zhang:2014:ODP

[ZSCJ14] Shaoquan Zhang, Ziyu Shao, Minghua Chen, and Libin Jiang. Optimal distributed P2P streaming under node degree bounds. IEEE/ACM Transactions on Networking, 22(3):717– 730, June 2014. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Zhang:2011:PSL

[ZSFZ11] Chi Zhang, Yang Song, Yuguang Fang, and Yanchao Zhang. On the price of security in largescale wireless ad hoc networks. IEEE/ACM Transactions on Networking, 19 (2):319–332, April 2011. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Zhang:2016:FCD

[ZSH^{+}16] Huazi Zhang, Kairan Sun, Qiuyuan Huang, Yonggang Wen, and Dapeng Wu. FUN coding: Design and analysis. IEEE/ ACM Transactions on Networking, 24(6):3340–3353, December 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Zheng:2012:SWD

[ZSK12] Zizhan Zheng, Prasun Sinha, and Santosh Kumar. Sparse WiFi deployment for vehicular Internet access with bounded interconnection gap. IEEE/ ACM Transactions on Networking, 20(3):956–969, June 2012. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Zheng:2017:TNS

[ZSL^{+}17] Yuanqing Zheng, Guobin Shen, Liqun Li, Chunshui Zhao, Mo Li, Feng Zhao, Yuanqing Zheng, Guobin Shen, Liqun Li, Chunshui Zhao, Mo Li, and Feng Zhao. Travi-Navi: Self-deployable indoor navigation system. IEEE/ ACM Transactions on Networking, 25(5):2655–2669, October 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Zhang:2021:OCD

[ZSL^{+}21] Jianan Zhang, Abhishek Sinha, Jaime Llorca, Antonia M. Tulino, and Eytan Modiano. Optimal control of distributed computing networks with mixedcast traffic flows. IEEE/ ACM Transactions on Networking, 29(4):1760–1773, August 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2021.3070699.

Zhao:2021:OMV

[ZSLZ21] Tianming Zhao, Weisheng Si, Wei Li, and Albert Y. Zomaya. Optimizing the maximum vertex coverage attacks under knapsack constraint. IEEE/ ACM Transactions on Networking, 29(3):1088–1104, June 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2021.3056450.

Zhang:2020:SPA

[ZSS^{+}20] Zhanzhan Zhang, Yin Sun, Ashutosh Sabharwal, Zhiyong Chen, and Bin Xia. Scheduling and power allocation dampens the negative effect of channel misreporting in massive MIMO. IEEE/ACM Transactions on Networking, 28(6):2531–2544, December 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2020.3014630.

Zhang:2002:SFC

[ZSSK02] Xi Zhang, Kang G. Shin, Debanjan Saha, and Dilip D. Kandlur. Scalable flow control for multicast ABR services in ATM networks. IEEE/ACM Transactions on Networking, 10(1):67– 85, February 2002. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Zhou:2017:DIR

[ZSZ^{+}17] Zimu Zhou, Longfei Shangguan, Xiaolong Zheng, Lei Yang, and Yunhao Liu. Design and implementation of an RFID-based customer shopping behavior mining system. IEEE/ ACM Transactions on Networking, 25(4):2405–2418, August 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Zhang:2021:MTN

[ZSZN21] Zuyuan Zhang, Fangming Shao, Nan Zhang, and Yifeng Niu. Maximizing k-terminal network reliability in some sparse graphs. IEEE/ ACM Transactions on Networking, 29(1):190–202, February 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2020.3030819.

Zhao:2003:MSR

[ZT03] Qing Zhao and Lang Tong. A multiqueue service room MAC protocol for wireless networks with multipacket reception. IEEE/ ACM Transactions on Networking, 11(1):125–137, February 2003. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Zhang:2012:OCA

[ZT12] Lu Zhang and Xueyan Tang. Optimizing client assignment for enhancing interactivity in distributed interactive applications. IEEE/ACM Transactions on Networking, 20(6):1707– 1720, December 2012. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Zhang:2023:OAS

[ZTH^{+}23] Chi Zhang, Haisheng Tan, Haoqiang Huang, Zhenhua Han, Shaofeng H.-C. Jiang, Guopeng Li, and Xiang-Yang Li. Online approximation scheme for scheduling heterogeneous utility jobs in edge computing. IEEE/ACM Transactions on Networking, 31 (1):352–365, 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2022.3193381.

Zitterbart:1994:HPT

[ZTS94] Martina Zitterbart, Ahmed N.

Tantawy, and Dimitrios N. Serpanos. A high performance transparent bridge. IEEE/ACM Transactions on Networking, 2(4):352– 362, August 1994. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1994-2-4/
p352-zitterbart/.

Zhao:2011:SCL

[ZTS11] Qinglin Zhao, Danny H. K. Tsang, and Taka Sakurai. A simple criticalload-based CAC scheme for IEEE 802.11 DCF networks. IEEE/ACM Transactions on Networking, 19(5):1485–1498, October 2011. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Zhang:2016:CCV

[ZV16] Le Zhang and Shahrokh Valaee. Congestion control for vehicular networks with safety-awareness. IEEE/ ACM Transactions on Networking, 24(6):3290–3299, December 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Zafirovic-Vukotic:1999:WTE

[ZVN99] Mirjana Zafirovic-Vukotic and Ignatius G. M. M. Niemegeers. Waiting time estimates in symmetric ATM-oriented rings with the destination release of used slots. IEEE/ ACM Transactions on Networking, 7(2):251–261, April 1999. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/
                                        journals/ton/1999-7-2/
                                        p251-zafirovic-vukotic/

.

Zhong:2010:CRR

[ZW10] Sheng Zhong and Fan Wu.

A collusion-resistant routing scheme for noncooperative wireless ad hoc networks. IEEE/ACM Transactions on Networking, 18 (2):582–595, April 2010. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Zhao:2014:NDD

[ZW14] Shizhen Zhao and Xinbing Wang. Node density and delay in large-scale wireless networks with unreliable links. IEEE/ ACM Transactions on Networking, 22(4):1150–1163, August 2014. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Zhao:2022:EMB

[ZW22] Jie Zhao and Xin Wang.

On the efficiency of multibeam medium access for millimeter-wave networks. IEEE/ACM Transactions on Networking, 30(4):1469– 1480, 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2021.3137562.

Zhang:2024:IBS

[ZWC^{+}24] Xiaohan Zhang, Jinwen Wang, Yueqiang Cheng, Qi Li, Kun Sun, Yao Zheng, Ning Zhang, and Xinghua Li. Interfacebased side channel in TEEassisted networked services. IEEE/ACM Transactions on Networking, 32 (1):613–626, 2024. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2023.3294019.

Zhou:2017:SLL

[ZWCL17] Haifeng Zhou, Chunming Wu, Qiumei Cheng, and Qianjun Liu. SDN–LIRU: a lossless and seamless method for SDN interdomain route updates. IEEE/ACM Transactions on Networking, 25(4):2473– 2483, August 2017. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Zhang:2000:VSP

[ZWDS00] Zhi-Li Zhang, Yuewei Wang, David H. C. Du, and Dongli Shu. Video staging: a proxy-server-based approach to end-to-end video delivery over wide-area networks. IEEE/ACM Transactions on Networking, 8 (4):429–442, 2000. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/2000-8-4/
p429-zhang/.

Zhang:2017:FTA

[ZWGC17] Chaoli Zhang, Fan Wu, Xiaofeng Gao, and Guihai Chen. Free talk in the air: a hierarchical topology for 60 GHz wireless data center networks. IEEE/ ACM Transactions on Networking, 25(6):3723–3737, December 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Zhang:2017:TAT

[ZWH^{+}17] Tao Zhang, Jianxin Wang, Jiawei Huang, Jianer Chen, Yi Pan, and Geyong Min. Tuning the aggressive TCP behavior for highly concurrent HTTP connections in intra-datacenter. IEEE/ ACM Transactions on Networking, 25(6):3808–3822, December 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Zhang:2020:TTR

[ZWJ^{+}20] Yiming Zhang, Haonan Wang, Menghan Jia, Jinyan Wang, Dongsheng Li, Guangtao Xue, and Kian-Lee Tan. TopoX: Topology refactorization for minimizing network communication in graph computations. IEEE/ACM Transactions on Networking, 28(6):2768–2782, December 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2020.3020813.

Zhang:2022:ACS

[ZWJ^{+}22] Sheng Zhang, Can Wang, Yibo Jin, Jie Wu, Zhuzhong Qian, Mingjun Xiao, and Sanglu Lu. Adaptive configuration selection and bandwidth allocation for edge-based video analytics. IEEE/ACM Transactions on Networking, 30(1):285– 298, February 2022. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1109/
                                        TNET.2021.3106937.

Zhang:2016:AML

[ZWL^{+}16] Zhao Zhang, James Willson, Zaixin Lu, Weili Wu, Xuding Zhu, and Ding-Zhu Du. Approximating maximum lifetime kcoverage through minimizing weighted k-cover in homogeneous wireless sensor networks. IEEE/ACM Transactions on Networking, 24(6):3620–3633, December 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Zhang:2022:MLV

[ZWL^{+}22] Xiaoxi Zhang, Jianyu Wang, Li-Feng Lee, Tom Yang, Akansha Kalra, Gauri Joshi, and Carlee Joe-Wong. Machine learning on volatile instances: Convergence, runtime, and cost tradeoffs. IEEE/ ACM Transactions on Networking, 30(1):215–228, February 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2021.3112082.

Zibman:1996:AAM

[ZWO^{+}96] Israel Zibman, Carl Woolf, Peter O’Reilly, Larry Strickland, David Willis, and John Visser. An architectural approach to minimizing feature interactions in telecommunications. IEEE/ACM Transactions on Networking, 4(4):582–596, August 1996. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/ton/1996-4-4/
p582-zibman/.

Zhang:2023:ERB

[ZWR^{+}23] Yongmin Zhang, Wei Wang, Ju Ren, Jinge Huang, Shibo He, and Yaoxue Zhang. Efficient revenuebased MEC server deployment and management in mobile edge-cloud computing. IEEE/ACM Transactions on Networking, 31(4): 1449–1462, 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2022.3217280.

Zheng:2017:DIC

[ZWS^{+}17] Xiaolong Zheng, Jiliang Wang, Longfei Shangguan, Zimu Zhou, and Yunhao Liu. Design and implementation of a CSI-based ubiquitous smoking detection system. IEEE/ACM Transactions on Networking, 25(6):3781–3793, December 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Zheng:2016:AUC

[ZWTC16] Zhenzhe Zheng, Fan Wu, Shaojie Tang, and Guihai Chen. AEGIS: an unknown combinatorial auction mechanism framework for heterogeneous spectrum redistribution in noncooperative wireless networks. IEEE/ACM Transactions on Networking, 24(3):1919– 1932, June 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Zhao:2024:JRU

[ZWX^{+}24] Gongming Zhao, Jingzhou Wang, Hongli Xu, Yangming Zhao, Xuwei Yang, and He Huang. Joint request updating and elastic resource provisioning with QoS guarantee in clouds. IEEE/ACM Transactions on Networking, 32 (1):110–126, 2024. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2023.3276881.

Zhou:2018:SRR

[ZWY^{+}18] Haifeng Zhou, Chunming Wu, Chengyu Yang, Pengfei Wang, Qi Yang, Zhouhao Lu, and Qiumei Cheng. SDN–RDCD: a real-time and reliable method for detecting compromised SDN devices. IEEE/ACM Transactions on Networking, 26(5):2048–2061, October 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Zhang:2018:BFS

[ZWYD18] Zhao Zhang, Weili Wu, Jing Yuan, and Ding-Zhu Du. Breach-free sleepwakeup scheduling for barrier coverage with heterogeneous wireless sensors. IEEE/ACM Transactions on Networking, 26(5):2404– 2413, October 2018. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Zheng:2010:MCM

[ZWYY10] S. Q. Zheng, Jianping Wang, Bing Yang, and Mei Yang. Minimum-cost multiple paths subject to minimum link and node sharing in a network. IEEE/ ACM Transactions on Networking, 18(5):1436–1449, October 2010. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Zhao:2020:PHV

[ZWZ20] Sicheng Zhao, Xing Wu, and Zuqing Zhu. On parallel and hitless vSDN reconfiguration. IEEE/ACM Transactions on Networking, 28(6):2657–2670, December 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2020.3014655.

Zhao:2024:SEE

[ZWZ^{+}24] Gongming Zhao, Jingzhou Wang, Yangming Zhao, Hongli Xu, Liusheng Huang, and Chunming Qiao. Segmented entanglement establishment with all-optical switching in quantum networks. IEEE/ACM Transactions on Networking, 32 (1):268–282, 2024. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2023.3281901.

Zhong:2023:THP

[ZWZC23] Jincheng Zhong, Ziling Wei, Shuang Zhao, and Shuhui Chen. Tuple-Tree: a high-performance packet classification algorithm supporting fast rule-set updates. IEEE/ ACM Transactions on Networking, 31(5):2027– 2041, 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2022.3227206.

Zhou:2018:FAD

[ZWZM18] Anfu Zhou, Teng Wei, Xinyu Zhang, and Huadong Ma. FastND: Accelerating directional neighbor discovery for 60-GHz millimeterwave wireless networks. IEEE/ACM Transactions on Networking, 26(5):2282– 2295, October 2018. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Zhao:2018:JOF

[ZXC^{+}18] Gongming Zhao, Hongli Xu, Shigang Chen, Liusheng Huang, and Pengzhan Wang. Joint optimization of flow table and group table for default paths in SDNs. IEEE/ ACM Transactions on Networking, 26(4):1837–1850, August 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Zhao:2013:LWS

[ZXH^{+}13] Jizhong Zhao, Wei Xi, Yuan He, Yunhao Liu, Xiang-Yang Li, Lufeng Mo, and Zheng Yang. Localization of wireless sensor networks in the wild: pursuit of ranging quality. IEEE/ACM Transactions on Networking, 21(1): 311–323, February 2013. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Zhang:2008:FAC

[ZXTT08] Weiyi Zhang, Guoliang Xue, Jian Tang, and Krishnaiyan Thulasiraman. Faster algorithms for construction of recovery trees enhancing QoP and QoS. IEEE/ACM Transactions on Networking, 16(3):642– 655, June 2008. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Zhou:2019:TOA

[ZXW^{+}19] Pan Zhou, Jie Xu, Wei Wang, Yuchong Hu, Dapeng Oliver Wu, and Shouling Ji. Toward optimal adaptive online shortest path routing with acceleration under jamming attack. IEEE/ ACM Transactions on Networking, 27(5):1815–1829, October 2019. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Zhai:2020:JRS

[ZXW^{+}20a] Yutong Zhai, Hongli Xu, Haibo Wang, Zeyu Meng, and He Huang. Joint routing and sketch configuration in software-defined networking. IEEE/ACM Transactions on Networking, 28(5):2092–2105, October 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2020.3002783.

Zhou:2020:RMW

[ZXW^{+}20b] Anfu Zhou, Shaoqing Xu, Song Wang, Jingqi Huang, Shaoyuan Yang, Teng Wei, Xinyu Zhang, and Huadong Ma. Robotic millimeter-wave wireless networks. IEEE/ACM Transactions on Networking, 28(4):1534–1549, August 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2020.2990498.

Zhang:2021:FOP

[ZXW^{+}21] Xinyi Zhang, Gaogang Xie, Xin Wang, Penghao Zhang, Yanbiao Li, and Kavé Salamatian. Fast online packet classification with convolutional neural network. IEEE/ACM Transactions on Networking, 29(6):2765–2778, December 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2021.3100114.

Zheng:2019:SFR

[ZXZ^{+}19] Jiaqi Zheng, Hong Xu, Xiaojun Zhu, Guihai Chen, and Yanhui Geng. Sentinel: Failure recovery in centralized traffic engineering. IEEE/ACM Transactions on Networking, 27(5): 1859–1872, October 2019. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Zhang:2007:NAM

[ZY07a] Zhenghao Zhang and Yuanyuan Yang. A novel analytical model for switches with shared buffer. IEEE/ ACM Transactions on Networking, 15(5):1191–1203, October 2007. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Zhang:2007:LOW

[ZY07b] Zhenghao Zhang and Yuanyuan Yang. On-line optimal wavelength assignment in WDM networks with shared wavelength converter pool. IEEE/ ACM Transactions on Networking, 15(1):234–245, February 2007. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Zhu:2016:ISD

[ZY16] Kai Zhu and Lei Ying.

Information source detection in the SIR model: a sample-path-based approach. IEEE/ACM Transactions on Networking, 24(1):408–421, February 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Zhang:2021:RSX

[ZY21] Yuhui Zhang and Dejun Yang. RobustPay^{+}: Robust payment routing with approximation guarantee in blockchain-based payment channel networks. IEEE/ ACM Transactions on Networking, 29(4):1676–1686, August 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2021.3069725.

Zhang:2021:PPD

[ZYH^{+}21] Mengyuan Zhang, Lei Yang, Shibo He, Ming Li, and Junshan Zhang. Privacy-preserving data aggregation for mobile crowdsensing with externality: an auction approach. IEEE/ACM Transactions on Networking, 29(3):1046– 1059, June 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2021.3056490.

Zhang:2014:CHB

[ZYL^{+}14] Yifan Zhang, Gexin Yu, Qun Li, Haodong Wang, Xiaojun Zhu, and Baosheng Wang. Channel-hopping-based communication rendezvous in cognitive radio networks. IEEE/ACM Transactions on Networking, 22(3):889–902, June 2014. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Zhang:2017:LAY

[ZYL^{+}17] Xiujuan Zhang, Jiguo Yu, Wei Li, Xiuzhen Cheng, Dongxiao Yu, and Feng Zhao. Localized algorithms for Yao graph-based spanner construction in wireless networks under SINR. IEEE/ACM Transactions on Networking, 25(4):2459– 2472, August 2017. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Zhang:2023:AHA

[ZYS^{+}23] Han Zhang, Xia Yin, Xingang Shi, Jilong Wang, Zhiliang Wang, Yingya Guo, Tian Lan, Yahui Li, Yongqing Zhu, Ke Ruan, and Haijun Geng. Achieving high availability in inter-DC WAN traffic engineering. IEEE/ACM Transactions on Networking, 31(6):2406–2421, 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL

https://dl.acm.org/doi/
10.1109/TNET.2022.3216592.

Zou:2021:DBR

[ZYY^{+}21] Yifei Zou, Dongxiao Yu, Jiguo Yu, Yong Zhang, Falko Dressler, and Xiuzhen Cheng. Distributed Byzantine-resilient multiplemessage dissemination in wireless networks. IEEE/ ACM Transactions on Networking, 29(4):1662–1675, August 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2021.3069324.

Zhang:2016:PSD

[ZYZ16] Mingyang Zhang, Changsheng You, and Zuqing Zhu. On the parallelization of spectrum defragmentation reconfigurations in elastic optical networks. IEEE/ACM Transactions on Networking, 24(5):2819– 2833, October 2016. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Zhang:2017:ABF

[ZZ17] Zhenghao Zhang and Zhenghao Zhang. Analog Bloom filter and contention-free multi-bit simultaneous query for centralized wireless networks. IEEE/ACM Transactions on Networking, 25(5):2916–2929, October 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Zhao:2016:PVI

[ZZG^{+}16] Mingchen Zhao, Wenchao Zhou, Alexander J. T. Gurney, Andreas Haeberlen, Micah Sherr, and Boon Thau Loo. Private and verifiable interdomain routing decisions. IEEE/ACM Transactions on Networking, 24(2):1011– 1024, April 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Zhang:2010:IDI

[ZZH^{+}10] Zheng Zhang, Ying Zhang, Y. Charlie Hu, Z. Morley Mao, and Randy Bush. iSPY: detecting IP prefix hijacking on my own. IEEE/ACM Transactions on Networking, 18(6):1815– 1828, December 2010. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Zhang:2019:FDP

[ZZH19] Peng Zhang, Cheng Zhang, and Chengchen Hu. Fast data plane testing for software-defined networks with RuleChecker. IEEE/ ACM Transactions on Networking, 27(1):173–186, February 2019. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Zhu:2013:AEF

[ZZHZ13] Ting Zhu, Ziguo Zhong, Tian He, and Zhi-Li Zhang. Achieving efficient flooding by utilizing link correlation in wireless sensor networks. IEEE/ACM Transactions on Networking, 21(1):121– 134, February 2013. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Zhu:2021:LAO

[ZZL^{+}21] Hongzi Zhu, Yuxiao Zhang, Zifan Liu, Xiao Wang, Shan Chang, and Yingying Chen. Localizing acoustic objects on a single phone. IEEE/ACM Transactions on Networking, 29(5):2170–2183, October 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2021.3080820.

Zhang:2022:CAP

[ZZL^{+}22] Miao Zhang, Yifei Zhu, Jiangchuan Liu, Feng Wang, and Fangxin Wang. CharmSeeker: Automated pipeline configuration for serverless video processing. IEEE/ACM Transactions on Networking, 30(6): 2730–2743, 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2022.3183231.

Zeng:2024:ERD

[ZZL^{+}24] Yiming Zeng, Jiarui Zhang, Ji Liu, Zhenhua Liu, and Yuanyuan Yang. Entanglement routing design over quantum networks. IEEE/ACM Transactions on Networking, 32 (1):352–367, 2024. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2023.3282560.

Zhu:2023:CPP

[ZZLM23] Shaopeng Zhu, Xiaolong Zheng, Liang Liu, and Huadong Ma. CSMA/PJ: a protective jamming based MAC protocol to harmonize the long and short links. IEEE/ACM Transactions on Networking, 31 (1):118–132, 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2022.3193027.

Zhao:2016:CPD

[ZZLW16] Can Zhao, Jian Zhao, Xiaojun Lin, and Chuan Wu. Capacity of P2P on-demand streaming with simple, robust, and decentralized control. IEEE/ ACM Transactions on Networking, 24(5):2607–2620, October 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Zhang:2016:TSR

[ZZS^{+}16] Jinxue Zhang, Rui Zhang, Jingchao Sun, Yanchao Zhang, Chi Zhang, Jinxue Zhang, Rui Zhang, Jingchao Sun, Yanchao Zhang, and Chi Zhang. TrueTop: a Sybil-resilient system for user influence measurement on Twitter. IEEE/ ACM Transactions on Networking, 24(5):2834–2846, October 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Zhou:2017:FTV

[ZZT^{+}17] Jiao Zhou, Zhao Zhang, Shaojie Tang, Xiaohui Huang, Yuchang Mo, and Ding-Zhu Du. Faulttolerant virtual backbone in heterogeneous wireless sensor network. IEEE/ ACM Transactions on Networking, 25(6):3487–3499, December 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Zhou:2015:PCG

[ZZW^{+}15] Xia Zhou, Zengbin Zhang, Gang Wang, Xiaoxiao Yu, Ben Y. Zhao, and Haitao Zheng. Practical conflict graphs in the wild. IEEE/ACM Transactions on Networking, 23(3):824– 835, June 2015. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Zhang:2023:RER

[ZZW^{+}23] Jiao Zhang, Xiaolong Zhong, Zirui Wan, Yu Tian, Tian Pan, and Tao Huang. RCC: Enabling receiverdriven RDMA congestion control with congestion divide-and-conquer in datacenter networks. IEEE/ACM Transactions on Networking, 31(1):103– 117, 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/10.1109/TNET.
                                        2022.3185105.

Zhu:2024:EOW

[ZZW^{+}24] Fengyuan Zhu, Renjie Zhao, Bingbing Wang, Xinbing Wang, Xinping Guan, Chenghu Zhou, and Xiaohua Tian. Enabling OFDMA in Wi-Fi backscatter. IEEE/ ACM Transactions on Networking, 32(1):427– 444, 2024. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2023.3290370.

Zhang:2021:HHP

[ZZX^{+}21a] Dai Zhang, Yu Zhou, Zhaowei Xi, Yangyang Wang, Mingwei Xu, and Jianping Wu. Hyper-Tester: High-performance network testing driven by programmable switches. IEEE/ACM Transactions on Networking, 29(5):2005– 2018, October 2021. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL https:/

/dl.acm.org/doi/10.1109/
TNET.2021.3077652.

Zhang:2021:NWF

[ZZX^{+}21b] Peng Zhang, Fangzheng Zhang, Shimin Xu, Zuoru Yang, Hao Li, Qi Li, Huanzhao Wang, Chao Shen, and Chengchen Hu. Network-wide forwarding anomaly detection and localization in software defined networks. IEEE/ ACM Transactions on Networking, 29(1):332–345, February 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.
acm.org/doi/10.1109/TNET.
2020.3033588.

Zhang:2007:AAP

[ZZZ^{+}07] Jing Zhang, Keyao Zhu, Hui Zang, Norman S.

Matloff, and Biswanath Mukherjee. Availabilityaware provisioning strategies for differentiated protection services in wavelengthconvertible WDM mesh networks. IEEE/ACM Transactions on Networking, 15(5):1177–1190, October 2007. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Zhang:2014:DSP

[ZZZ^{+}14] Xin Zhang, Fanfu Zhou, Xinyu Zhu, Haiyang Sun, Adrian Perrig, Athanasios V. Vasilakos, and Haibing Guan. DFL: secure and practical fault localization for datacenter networks. IEEE/ACM Transactions on Networking, 22(4):1218– 1231, August 2014. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Zhang:2024:LTH

[ZZZ^{+}24] Junxue Zhang, Chaoliang Zeng, Hong Zhang, Shuihai Hu, and Kai Chen. LiteFlow: Toward highperformance adaptive neural networks for kernel datapath. IEEE/ACM Transactions on Networking, 32 (1):627–642, 2024. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.

acm.org/doi/10.1109/TNET.
2023.3293152.

Zhu:2003:NGG

[ZZZM03] Hongyue Zhu, Hui Zang, Keyao Zhu, and Biswanath Mukherjee. A novel generic graph model for traffic grooming in heterogeneous WDM mesh networks. IEEE/ACM Transactions on Networking, 11 (2):285–299, April 2003. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).